XEmacs Lisp Reference Manual

Version 3.4 (for XEmacs 21.1), May 1999

by Ben Wing

Based on the GNU Emacs Lisp Reference Manual by Bil Lewis, Dan LaLiberte, Richard Stallman and the GNU Manual Group

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1990, 1991, 1992, 1993, 1994, 1995 Free Software Foundation, Inc. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994, 1995 Sun Microsystems, Inc. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995, 1996 Ben Wing.

Version 3.3 Revised for XEmacs Versions 21.1, April 1998.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the section entitled “GNU General Public License” is included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that the section entitled “GNU General Public License” may be included in a translation approved by the Free Software Foundation instead of in the original English. Cover art by Etienne Suvasa.

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc.

675 Mass Ave, Cambridge, MA 02139, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating that you

changed the files and the date of any change. b. You must cause any work that you distribute or publish, that in whole or in part

contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it. Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable source code,

which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY AP-PLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IM-PLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE EN-TIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

Copyright (C) 19yy name of author
This program is free software; you can redistribute it and/or
modify it under the terms of the GNU General Public License
as published by the Free Software Foundation; either version 2
of the License, or (at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License
along with this program; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail. If the program is interactive, make it output a short notice like this when it starts in an

interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
type `show w'.  This is free software, and you are welcome
to redistribute it under certain conditions; type `show c'
for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright
interest in the program `Gnomovision'
(which makes passes at compilers) written
by James Hacker.

signature of Ty Coon, 1 April 1989

Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

8 XEmacs Lisp Reference Manual

1 Introduction

Most of the XEmacs text editor is written in the programming language called XEmacs Lisp. You can write new code in XEmacs Lisp and install it as an extension to the editor. However, XEmacs Lisp is more than a mere “extension language”; it is a full computer programming language in its own right. You can use it as you would any other programming language.

Because XEmacs Lisp is designed for use in an editor, it has special features for scanning and parsing text as well as features for handling files, buffers, displays, subprocesses, and so on. XEmacs Lisp is closely integrated with the editing facilities; thus, editing commands are functions that can also conveniently be called from Lisp programs, and parameters for customization are ordinary Lisp variables.

This manual describes XEmacs Lisp, presuming considerable familiarity with the use of XEmacs for editing. (See The XEmacs Reference Manual, for this basic information.) Generally speaking, the earlier chapters describe features of XEmacs Lisp that have counterparts in many programming languages, and later chapters describe features that are peculiar to XEmacs Lisp or relate specifically to editing.

This is edition 3.3.

1.1 Caveats

This manual has gone through numerous drafts. It is nearly complete but not flawless. There are a few topics that are not covered, either because we consider them secondary (such as most of the individual modes) or because they are yet to be written. Because we are not able to deal with them completely, we have left out several parts intentionally. This includes most information about usage on VMS.

The manual should be fully correct in what it does cover, and it is therefore open to criticism on anything it says—from specific examples and descriptive text, to the ordering of chapters and sections. If something is confusing, or you find that you have to look at the sources or experiment to learn something not covered in the manual, then perhaps the manual should be fixed. Please let us know.

As you use the manual, we ask that you mark pages with corrections so you can later look them up and send them in. If you think of a simple, real-life example for a function or group of functions, please make an effort to write it up and send it in. Please reference any comments to the chapter name, section name, and function name, as appropriate, since page numbers and chapter and section numbers will change and we may have trouble finding the text you are talking about. Also state the number of the edition you are criticizing.

This manual was originally written for FSF Emacs 19 and was updated by Ben Wing (wing@666.com) for Lucid Emacs 19.10 and later for XEmacs 19.12, 19.13, 19.14, and 20.0. It was further updated by the XEmacs Development Team for 19.15, version 20 and 21. Please send comments and corrections relating to XEmacs-specific portions of this manual to

     xemacs@xemacs.org

or post to the newsgroup

     comp.emacs.xemacs

–Ben Wing

1.2 Lisp History

Lisp (LISt Processing language) was first developed in the late 1950’s at the Massachusetts Institute of Technology for research in artificial intelligence. The great power of the Lisp language makes it superior for other purposes as well, such as writing editing commands.

Dozens of Lisp implementations have been built over the years, each with its own idiosyncrasies. Many of them were inspired by Maclisp, which was written in the 1960’s at MIT’s Project MAC. Eventually the implementors of the descendants of Maclisp came together and developed a standard for Lisp systems, called Common Lisp.

XEmacs Lisp is largely inspired by Maclisp, and a little by Common Lisp. If you know Common Lisp, you will notice many similarities. However, many of the features of Common Lisp have been omitted or simplified in order to reduce the memory requirements of XEmacs. Sometimes the simplifications are so drastic that a Common Lisp user might be very confused. We will occasionally point out how XEmacs Lisp differs from Common Lisp. If you don’t know Common Lisp, don’t worry about it; this manual is self-contained.

1.3 Conventions

This section explains the notational conventions that are used in this manual. You may want to skip this section and refer back to it later.

1.3.1 Some Terms

Throughout this manual, the phrases “the Lisp reader” and “the Lisp printer” are used to refer to those routines in Lisp that convert textual representations of Lisp objects into actual Lisp objects, and vice versa. See Section 2.1 [Printed Representation], page 15, for more details. You, the person reading this manual, are thought of as “the programmer” and are addressed as “you”. “The user” is the person who uses Lisp programs, including those you write.

Examples of Lisp code appear in this font or form: (list 1 2 3). Names that represent arguments or metasyntactic variables appear in this font or form: first-number.

1.3.2 nil and t

In Lisp, the symbol nil has three separate meanings: it is a symbol with the name ‘nil’; it is the logical truth value false; and it is the empty list—the list of zero elements. When used as a variable, nil always has the value nil.

As far as the Lisp reader is concerned, ‘()’ and ‘nil’ are identical: they stand for the same object, the symbol nil. The different ways of writing the symbol are intended entirely for human readers. After the Lisp reader has read either ‘()’ or ‘nil’, there is no way to determine which representation was actually written by the programmer.

In this manual, we use () when we wish to emphasize that it means the empty list, and we use nil when we wish to emphasize that it means the truth value false. That is a good convention to use in Lisp programs also.

(cons 'foo ()) ; Emphasize the empty list (not nil) ; Emphasize the truth value false

In contexts where a truth value is expected, any non-nil value is considered to be true. However, t is the preferred way to represent the truth value true. When you need to choose a value which represents true, and there is no other basis for choosing, use t. The symbol t always has value t.

In XEmacs Lisp, nil and t are special symbols that always evaluate to themselves. This is so that you do not need to quote them to use them as constants in a program. An attempt to change their values results in a setting-constant error. See Section 10.6 [Accessing Variables], page 151.

1.3.3 Evaluation Notation

A Lisp expression that you can evaluate is called a form. Evaluating a form always produces a result, which is a Lisp object. In the examples in this manual, this is indicated with ‘⇒’:

     (car '(1 2))
          ⇒ 1

You can read this as “(car '(1 2)) evaluates to 1”.

When a form is a macro call, it expands into a new form for Lisp to evaluate. We show the result of the expansion with ‘↦’. We may or may not show the actual result of the evaluation of the expanded form.

     (news-cadr '(a b c))
          ↦ (car (cdr '(a b c)))
          ⇒ b

Sometimes to help describe one form we show another form that produces identical results. The exact equivalence of two forms is indicated with ‘ ≡ ’.

     (cons 'a nil) ≡ (list 'a)

1.3.4 Printing Notation

Many of the examples in this manual print text when they are evaluated. If you execute example code in a Lisp Interaction buffer (such as the buffer ‘*scratch*’), the printed text is inserted into the buffer. If you execute the example by other means (such as by evaluating the function eval-region), the printed text is displayed in the echo area. You should be aware that text displayed in the echo area is truncated to a single line.

Examples in this manual indicate printed text with ‘ ⊣ ’, irrespective of where that text goes. The value returned by evaluating the form (here bar) follows on a separate line.

     (progn (print 'foo) (print 'bar))
          ⊣ foo

          ⊣ bar
          ⇒ bar

1.3.5 Error Messages

Some examples signal errors. This normally displays an error message in the echo area. We show the error message on a line starting with ‘ error ’. Note that ‘ error ’ itself does not appear in the echo area.

     (+ 23 'x)
      error Wrong type argument: integer-or-marker-p, x

1.3.6 Buffer Text Notation

Some examples show modifications to text in a buffer, with “before” and “after” versions of the text. These examples show the contents of the buffer in question between two lines of dashes containing the buffer name. In addition, ‘⋆’ indicates the location of point. (The symbol for point, of course, is not part of the text in the buffer; it indicates the place between two characters where point is located.)

     ---------- Buffer: foo ----------
     This is the ⋆contents of foo.
     ---------- Buffer: foo ----------
     (insert "changed ")
          ⇒ nil
     ---------- Buffer: foo ----------
     This is the changed ⋆contents of foo.
     ---------- Buffer: foo ----------

1.3.7 Format of Descriptions

Functions, variables, macros, commands, user options, and special forms are described in this manual in a uniform format. The first line of a description contains the name of the item followed by its arguments, if any. The category—function, variable, or whatever—is printed next to the right margin. The description follows on succeeding lines, sometimes with examples.

1.3.7.1 A Sample Function Description

In a function description, the name of the function being described appears first. It is followed on the same line by a list of parameters. The names used for the parameters are also used in the body of the description.

The appearance of the keyword &optional in the parameter list indicates that the arguments for subsequent parameters may be omitted (omitted parameters default to nil). Do not write &optional when you call the function.

The keyword &rest (which will always be followed by a single parameter) indicates that any number of arguments can follow. The value of the single following parameter will be a list of all these arguments. Do not write &rest when you call the function.

Here is a description of an imaginary function foo:

Function foo integer1 &optional integer2 &rest integers The function foo subtracts integer1 from integer2, then adds all the rest of the arguments to the result. If integer2 is not supplied, then the number 19 is used by default.

          (foo 1 5 3 9)
               ⇒ 16
          (foo 5)
               ⇒ 14

More generally,

          (foo w x y...)

≡

          (+ (- x w) y...)

Any parameter whose name contains the name of a type (e.g., integer, integer1 or buffer) is expected to be of that type. A plural of a type (such as buffers) often means a list of objects of that type. Parameters named object may be of any type. (See Chapter 2 [Lisp Data Types], page 15, for a list of XEmacs object types.) Parameters with other sorts of names (e.g., new-file) are discussed specifically in the description of the function. In some sections, features common to parameters of several functions are described at the beginning.

See Section 11.2 [Lambda Expressions], page 164, for a more complete description of optional and rest arguments.

Command, macro, and special form descriptions have the same format, but the word ‘Function’ is replaced by ‘Command’, ‘Macro’, or ‘Special Form’, respectively. Commands are simply functions that may be called interactively; macros process their arguments differently from functions (the arguments are not evaluated), but are presented the same way.

Special form descriptions use a more complex notation to specify optional and repeated parameters because they can break the argument list down into separate arguments in more complicated ways. ‘[optional-arg]’ means that optional-arg is optional and ‘repeatedargs...’ stands for zero or more arguments. Parentheses are used when several arguments are grouped into additional levels of list structure. Here is an example:

Special Form count-loop (var [from to [inc]]) body … This imaginary special form implements a loop that executes the body forms and then increments the variable var on each iteration. On the first iteration, the variable has the value from; on subsequent iterations, it is incremented by 1 (or by inc if that is given). The loop exits before executing body if var equals to. Here is an example:

          (count-loop (i 0 10)
            (prin1 i) (princ " ")
            (prin1 (aref vector i)) (terpri))

If from and to are omitted, then var is bound to nil before the loop begins, and the loop exits if var is non-nil at the beginning of an iteration. Here is an example:

          (count-loop (done)
            (if (pending)
                (fixit)

              (setq done t)))

In this special form, the arguments from and to are optional, but must both be present or both absent. If they are present, inc may optionally be specified as well. These arguments are grouped with the argument var into a list, to distinguish them from body, which includes all remaining elements of the form.

1.3.7.2 A Sample Variable Description

A variable is a name that can hold a value. Although any variable can be set by the user, certain variables that exist specifically so that users can change them are called user options. Ordinary variables and user options are described using a format like that for functions except that there are no arguments.

Here is a description of the imaginary electric-future-map variable.

Variable electric-future-map

The value of this variable is a full keymap used by Electric Command Future mode. The functions in this map allow you to edit commands you have not yet thought about executing.

User option descriptions have the same format, but ‘Variable’ is replaced by ‘User Option’.

1.4 Acknowledgements

This manual was based on the GNU Emacs Lisp Reference Manual, version 2.4, written by Robert Krawitz, Bil Lewis, Dan LaLiberte, Richard M. Stallman and Chris Welty, the volunteers of the GNU manual group, in an effort extending over several years. Robert J. Chassell helped to review and edit the manual, with the support of the Defense Advanced Research Projects Agency, ARPA Order 6082, arranged by Warren A. Hunt, Jr. of Computational Logic, Inc.

Ben Wing adapted this manual for XEmacs 19.14 and 20.0, and earlier for Lucid Emacs 19.10, XEmacs 19.12, and XEmacs 19.13. He is the sole author of many of the manual sections, in particular the XEmacs-specific sections: events, faces, extents, glyphs, specifiers, toolbar, menubars, scrollbars, dialog boxes, devices, consoles, hash tables, range tables, char tables, databases, and others. The section on annotations was originally written by Chuck Thompson. Corrections to v3.1 and later were done by Martin Buchholz, Steve Baur, and Hrvoje Niksic.

Corrections to the original GNU Emacs Lisp Reference Manual were supplied by Karl Berry, Jim Blandy, Bard Bloom, Stephane Boucher, David Boyes, Alan Carroll, Richard Davis, Lawrence R. Dodd, Peter Doornbosch, David A. Duff, Chris Eich, Beverly Erlebacher, David Eckelkamp, Ralf Fassel, Eirik Fuller, Stephen Gildea, Bob Glickstein, Eric Hanchrow, George Hartzell, Nathan Hess, Masayuki Ida, Dan Jacobson, Jak Kirman, Bob Knighten, Frederick M. Korz, Joe Lammens, Glenn M. Lewis, K. Richard Magill, Brian Marick, Roland McGrath, Skip Montanaro, John Gardiner Myers, Thomas A. Peterson, Francesco Potorti, Friedrich Pukelsheim, Arnold D. Robbins, Raul Rockwell, Per Starback, Shinichirou Sugou, Kimmo Suominen, Edward Tharp, Bill Trost, Rickard Westman, Jean White, Matthew Wilding, Carl Witty, Dale Worley, Rusty Wright, and David D. Zuhn.

2 Lisp Data Types

A Lisp object is a piece of data used and manipulated by Lisp programs. For our purposes, a type or data type is a set of possible objects.

Every object belongs to at least one type. Objects of the same type have similar structures and may usually be used in the same contexts. Types can overlap, and objects can belong to two or more types. Consequently, we can ask whether an object belongs to a particular type, but not for “the” type of an object.

A few fundamental object types are built into XEmacs. These, from which all other types are constructed, are called primitive types. Each object belongs to one and only one primitive type. These types include integer, character (starting with XEmacs 20.0), float, cons, symbol, string, vector, bit-vector, subr, compiled-function, hashtable, range-table, char-table, weak-list, and several special types, such as buffer, that are related to editing. (See Section 2.5 [Editing Types], page 29.)

Each primitive type has a corresponding Lisp function that checks whether an object is a member of that type.

Note that Lisp is unlike many other languages in that Lisp objects are self-typing: the primitive type of the object is implicit in the object itself. For example, if an object is a vector, nothing can treat it as a number; Lisp knows it is a vector, not a number.

In most languages, the programmer must declare the data type of each variable, and the type is known by the compiler but not represented in the data. Such type declarations do not exist in XEmacs Lisp. A Lisp variable can have any type of value, and it remembers whatever value you store in it, type and all.

This chapter describes the purpose, printed representation, and read syntax of each of the standard types in Emacs Lisp. Details on how to use these types can be found in later chapters.

2.1 Printed Representation and Read Syntax

The printed representation of an object is the format of the output generated by the Lisp printer (the function prin1) for that object. The read syntax of an object is the format of the input accepted by the Lisp reader (the function read) for that object. Most objects have more than one possible read syntax. Some types of object have no read syntax; except for these cases, the printed representation of an object is also a read syntax for it.

In other languages, an expression is text; it has no other form. In Lisp, an expression is primarily a Lisp object and only secondarily the text that is the object’s read syntax. Often there is no need to emphasize this distinction, but you must keep it in the back of your mind, or you will occasionally be very confused.

Every type has a printed representation. Some types have no read syntax, since it may not make sense to enter objects of these types directly in a Lisp program. For example, the buffer type does not have a read syntax. Objects of these types are printed in hash notation: the characters ‘#<’ followed by a descriptive string (typically the type name followed by the name of the object), and closed with a matching ‘>’. Hash notation cannot be read at all, so the Lisp reader signals the error invalid-read-syntax whenever it encounters ‘#<’.

     (current-buffer)
          ⇒ #<buffer "objects.texi">

When you evaluate an expression interactively, the Lisp interpreter first reads the textual representation of it, producing a Lisp object, and then evaluates that object (see Chapter 8 [Evaluation], page 119). However, evaluation and reading are separate activities. Reading returns the Lisp object represented by the text that is read; the object may or may not be evaluated later. See Section 17.3 [Input Functions], page 253, for a description of read, the basic function for reading objects.

2.2 Comments

A comment is text that is written in a program only for the sake of humans that read the program, and that has no effect on the meaning of the program. In Lisp, a semicolon (‘;’) starts a comment if it is not within a string or character constant. The comment continues to the end of line. The Lisp reader discards comments; they do not become part of the Lisp objects which represent the program within the Lisp system.

The ‘#@count’ construct, which skips the next count characters, is useful for programgenerated comments containing binary data. The XEmacs Lisp byte compiler uses this in its output files (see Chapter 15 [Byte Compilation], page 205). It isn’t meant for source files, however.

See Section A.4 [Comment Tips], page 762, for conventions for formatting comments.

2.3 Primitive Types

For reference, here is a list of all the primitive types that may exist in XEmacs. Note that some of these types may not exist in some XEmacs executables; that depends on the options that XEmacs was configured with.

• bit-vector

• buffer

• char-table

• character

• charset

• coding-system

• cons

• color-instance

• compiled-function

• console

• database

• device

• event

• extent

• face • float

• font-instance

• frame

• glyph

• hashtable

• image-instance

• integer

• keymap

• marker

• process

• range-table

• specifier

• string

• subr

• subwindow

• symbol

• toolbar-button

• tooltalk-message

• tooltalk-pattern

• vector

• weak-list

• window

• window-configuration

• x-resource In addition, the following special types are created internally but will never be seen by

Lisp code. You may encounter them, however, if you are debugging XEmacs. The printed representation of these objects begins ‘#<INTERNAL EMACS BUG’, which indicates to the Lisp programmer that he has found an internal bug in XEmacs if he ever encounters any of these objects.

• char-table-entry

• command-builder

• extent-auxiliary

• extent-info

• lcrecord-list

• lstream

• opaque

• opaque-list

• popup-data

• symbol-value-buffer-local • symbol-value-forward

• symbol-value-lisp-magic

• symbol-value-varalias

• toolbar-data

2.4 Programming Types

There are two general categories of types in XEmacs Lisp: those having to do with Lisp programming, and those having to do with editing. The former exist in many Lisp implementations, in one form or another. The latter are unique to XEmacs Lisp.

2.4.1 Integer Type

The range of values for integers in XEmacs Lisp is −134217728 to 134217727 (28 bits; i.e., −2^{27} to 2^{28} − 1) on most machines. (Some machines, in particular 64-bit machines such as the DEC Alpha, may provide a wider range.) It is important to note that the XEmacs Lisp arithmetic functions do not check for overflow. Thus (1+ 134217727) is −134217728 on most machines. (However, you will get an error if you attempt to read an out-of-range number using the Lisp reader.)

The read syntax for integers is a sequence of (base ten) digits with an optional sign at the beginning. (The printed representation produced by the Lisp interpreter never has a leading ‘+’.)

-1 ; The integer -1. 1 ; The integer 1. +1 ; Also the integer 1. 268435457 ; Causes an error on a 28-bit implementation.

See Chapter 3 [Numbers], page 45, for more information.

2.4.2 Floating Point Type

XEmacs supports floating point numbers. The precise range of floating point numbers is machine-specific.

The printed representation for floating point numbers requires either a decimal point (with at least one digit following), an exponent, or both. For example, ‘1500.0’, ‘15e2’, ‘15.0e2’, ‘1.5e3’, and ‘.15e4’ are five ways of writing a floating point number whose value is 1500. They are all equivalent.

See Chapter 3 [Numbers], page 45, for more information.

2.4.3 Character Type

In XEmacs version 19, and in all versions of FSF GNU Emacs, a character in XEmacs Lisp is nothing more than an integer. This is yet another holdover from XEmacs Lisp’s derivation from vintage-1980 Lisps; modern versions of Lisp consider this equivalence a bad idea, and have separate character types. In XEmacs version 20, the modern convention is followed, and characters are their own primitive types. (This change was necessary in order for MULE, i.e. Asian-language, support to be correctly implemented.)

Even in XEmacs version 20, remnants of the equivalence between characters and integers still exist; this is termed the char-int confoundance disease. In particular, many functions such as eq, equal, and memq have equivalent functions (old-eq, old-equal, old-memq, etc.) that pretend like characters are integers are the same. Byte code compiled under any version 19 Emacs will have all such functions mapped to their old- equivalents when the byte code is read into XEmacs 20. This is to preserve compatibility – Emacs 19 converts all constant characters to the equivalent integer during byte-compilation, and thus there is no other way to preserve byte-code compatibility even if the code has specifically been written with the distinction between characters and integers in mind.

Every character has an equivalent integer, called the character code. For example, the character A is represented as the integer 65, following the standard ASCII representation of characters. If XEmacs was not compiled with MULE support, the range of this integer will always be 0 to 255 – eight bits, or one byte. (Integers outside this range are accepted but silently truncated; however, you should most decidedly not rely on this, because it will not work under XEmacs with MULE support.) When MULE support is present, the range of character codes is much larger. (Currently, 19 bits are used.)

FSF GNU Emacs uses kludgy character codes above 255 to represent keyboard input of ASCII characters in combination with certain modifiers. XEmacs does not use this (a more general mechanism is used that does not distinguish between ASCII keys and other keys), so you will never find character codes above 255 in a non-MULE XEmacs.

Individual characters are not often used in programs. It is far more common to work with strings, which are sequences composed of characters. See Section 2.4.8 [String Type], page 25.

The read syntax for characters begins with a question mark, followed by the character (if it’s printable) or some symbolic representation of it. In XEmacs 20, where characters are their own type, this is also the print representation. In XEmacs 19, however, where characters are really integers, the printed representation of a character is a decimal number. This is also a possible read syntax for a character, but writing characters that way in Lisp programs is a very bad idea. You should always use the special read syntax formats that XEmacs Lisp provides for characters.

The usual read syntax for alphanumeric characters is a question mark followed by the character; thus, ‘?A’ for the character A, ‘?B’ for the character B, and ‘?a’ for the character

a.

For example:

;; Under XEmacs 20:

     ?Q ⇒ ?Q    ?q ⇒ ?q
     (char-int ?Q) ⇒ 81

;; Under XEmacs 19:

     ?Q ⇒ 81     ?q ⇒ 113

You can use the same syntax for punctuation characters, but it is often a good idea to add a ‘\’ so that the Emacs commands for editing Lisp code don’t get confused. For example, ‘?\ ’ is the way to write the space character. If the character is ‘\’, you must use a second ‘\’ to quote it: ‘?\\’. XEmacs 20 always prints punctuation characters with a ‘\’ in front of them, to avoid confusion.

You can express the characters Control-g, backspace, tab, newline, vertical tab, formfeed, return, and escape as ‘?\a’, ‘?\b’, ‘?\t’, ‘?\n’, ‘?\v’, ‘?\f’, ‘?\r’, ‘?\e’, respectively. Their character codes are 7, 8, 9, 10, 11, 12, 13, and 27 in decimal. Thus,

;; Under XEmacs 20:

?\a ⇒ ?\^G              ; C-g
(char-int ?\a) ⇒ 7

?\b ⇒ ?\^H ; backspace, ^{⟨}\underline{\overline{BS}}^{⟩}, C-h

(char-int ?\b) ⇒ 8

?\t ⇒ ?\t ; tab, ^{⟨}\underline{\overline{TAB}}^{⟩}, C-i

(char-int ?\t) ⇒ 9

?\n ⇒ ?\n ; newline, ^{⟨}\underline{\overline{LFD}}^{⟩}, C-j ?\v ⇒ ?\^K ; vertical tab, C-k ?\f ⇒ ?\^L ; formfeed character, C-l ?\r ⇒ ?\r ; carriage return, ^{⟨}\underline{\overline{RET}}^{⟩}, C-m ?\e ⇒ ?\^[ ; escape character, ^{⟨}\underline{\overline{ESC}}^{⟩}, C-[ ?\\ ⇒ ?\\ ; backslash character, \ ;; Under XEmacs 19:

?\a ⇒ 7                 ; C-g

?\b ⇒ 8 ; backspace, ^{⟨}\underline{\overline{BS}}^{⟩}, C-h ?\t ⇒ 9 ; tab, ^{⟨}\underline{\overline{TAB}}^{⟩}, C-i ?\n ⇒ 10 ; newline, ^{⟨}\underline{\overline{LFD}}^{⟩}, C-j ?\v ⇒ 11 ; vertical tab, C-k ?\f ⇒ 12 ; formfeed character, C-l ?\r ⇒ 13 ; carriage return, ^{⟨}\underline{\overline{RET}}^{⟩}, C-m ?\e ⇒ 27 ; escape character, ^{⟨}\underline{\overline{ESC}}^{⟩}, C-[ ?\\ ⇒ 92 ; backslash character, \

These sequences which start with backslash are also known as escape sequences, because backslash plays the role of an escape character; this usage has nothing to do with the character ^{⟨}\underline{\overline{ESC}}^{⟩}.

Control characters may be represented using yet another read syntax. This consists of a question mark followed by a backslash, caret, and the corresponding non-control character, in either upper or lower case. For example, both ‘?\^I’ and ‘?\^i’ are valid read syntax for the character C-i, the character whose value is 9.

Instead of the ‘^’, you can use ‘C-’; thus, ‘?\C-i’ is equivalent to ‘?\^I’ and to ‘?\^i’: ;; Under XEmacs 20:

?\^I ⇒ ?\t   ?\C-I ⇒ ?\t
(char-int ?\^I) ⇒ 9

;; Under XEmacs 19:

?\^I ⇒ 9     ?\C-I ⇒ 9

There is also a character read syntax beginning with ‘\M-’. This sets the high bit of the character code (same as adding 128 to the character code). For example, ‘?\M-A’ stands for the character with character code 193, or 128 plus 65. You should not use this syntax in your programs. It is a holdover of yet another confoundance disease from earlier Emacsen. (This was used to represent keyboard input with the ^{⟨}\underline{\overline{META}}^{⟩} key set, thus the ‘M’; however, it conflicts with the legitimate ISO-8859-1 interpretation of the character code. For example, character code 193 is a lowercase ‘a’ with an acute accent, in ISO-8859-1.)

Finally, the most general read syntax consists of a question mark followed by a backslash and the character code in octal (up to three octal digits); thus, ‘?\101’ for the character A, ‘?\001’ for the character C-a, and ?\002 for the character C-b. Although this syntax can represent any ASCII character, it is preferred only when the precise octal value is more important than the ASCII representation.

;; Under XEmacs 20:

     ?\012 ⇒ ?\n        ?\n ⇒ ?\n        ?\C-j ⇒ ?\n
     ?\101 ⇒ ?A         ?A ⇒ ?A

;; Under XEmacs 19:

     ?\012 ⇒ 10         ?\n ⇒ 10         ?\C-j ⇒ 10
     ?\101 ⇒ 65         ?A ⇒ 65

A backslash is allowed, and harmless, preceding any character without a special escape meaning; thus, ‘?\+’ is equivalent to ‘?+’. There is no reason to add a backslash before most characters. However, you should add a backslash before any of the characters ‘()\|;'`"#.,’ to avoid confusing the Emacs commands for editing Lisp code. Also add a backslash before whitespace characters such as space, tab, newline and formfeed. However, it is cleaner to use one of the easily readable escape sequences, such as ‘\t’, instead of an actual whitespace character such as a tab.

2.4.4 Symbol Type

A symbol in XEmacs Lisp is an object with a name. The symbol name serves as the printed representation of the symbol. In ordinary use, the name is unique—no two symbols have the same name.

A symbol can serve as a variable, as a function name, or to hold a property list. Or it may serve only to be distinct from all other Lisp objects, so that its presence in a data structure may be recognized reliably. In a given context, usually only one of these uses is intended. But you can use one symbol in all of these ways, independently.

A symbol name can contain any characters whatever. Most symbol names are written with letters, digits, and the punctuation characters ‘-+=*/’. Such names require no special punctuation; the characters of the name suffice as long as the name does not look like a number. (If it does, write a ‘\’ at the beginning of the name to force interpretation as a symbol.) The characters ‘_~!@$%^&:<>{}’ are less often used but also require no special punctuation. Any other characters may be included in a symbol’s name by escaping them with a backslash. In contrast to its use in strings, however, a backslash in the name of a symbol simply quotes the single character that follows the backslash. For example, in a string, ‘\t’ represents a tab character; in the name of a symbol, however, ‘\t’ merely quotes the letter t. To have a symbol with a tab character in its name, you must actually use a tab (preceded with a backslash). But it’s rare to do such a thing.

Common Lisp note: In Common Lisp, lower case letters are always “folded” to upper case, unless they are explicitly escaped. In Emacs Lisp, upper case and lower case letters are distinct.

Here are several examples of symbol names. Note that the ‘+’ in the fifth example is escaped to prevent it from being read as a number. This is not necessary in the sixth example because the rest of the name makes it invalid as a number. foo ; A symbol named ‘foo’. FOO ; A symbol named ‘FOO’, different from ‘foo’.

     char-to-string      ; A symbol named ‘char-to-string’.

1+ ; A symbol named ‘1+’

; (not ‘+1’, which is an integer).

\+1 ; A symbol named ‘+1’

; (not a very readable name).

\(*\ 1\ 2\) ; A symbol named ‘(* 1 2)’ (a worse name).

     +-*/_~!@$%^&=:<>{}  ; A symbol named ‘+-*/_~!@$%^&=:<>{}’.

; These characters need not be escaped.

2.4.5 Sequence Types

A sequence is a Lisp object that represents an ordered set of elements. There are two kinds of sequence in XEmacs Lisp, lists and arrays. Thus, an object of type list or of type array is also considered a sequence.

Arrays are further subdivided into strings, vectors, and bit vectors. Vectors can hold elements of any type, but string elements must be characters, and bit vector elements must be either 0 or 1. However, the characters in a string can have extents (see Chapter 40 [Extents], page 585) and text properties (see Section 36.18 [Text Properties], page 539) like characters in a buffer; vectors do not support extents or text properties even when their elements happen to be characters.

Lists, strings, vectors, and bit vectors are different, but they have important similarities. For example, all have a length l, and all have elements which can be indexed from zero to l minus one. Also, several functions, called sequence functions, accept any kind of sequence. For example, the function elt can be used to extract an element of a sequence, given its index. See Chapter 6 [Sequences Arrays Vectors], page 101.

It is impossible to read the same sequence twice, since sequences are always created anew upon reading. If you read the read syntax for a sequence twice, you get two sequences with equal contents. There is one exception: the empty list () always stands for the same object,

nil.

2.4.6 Cons Cell and List Types

A cons cell is an object comprising two pointers named the car and the cdr. Each of them can point to any Lisp object.

A list is a series of cons cells, linked together so that the cdr of each cons cell points either to another cons cell or to the empty list. See Chapter 5 [Lists], page 77, for functions that work on lists. Because most cons cells are used as part of lists, the phrase list structure has come to refer to any structure made out of cons cells.

The names car and cdr have only historical meaning now. The original Lisp implementation ran on an IBM 704 computer which divided words into two parts, called the “address” part and the “decrement”; car was an instruction to extract the contents of the address part of a register, and cdr an instruction to extract the contents of the decrement. By contrast, “cons cells” are named for the function cons that creates them, which in turn is named for its purpose, the construction of cells.

Because cons cells are so central to Lisp, we also have a word for “an object which is not a cons cell”. These objects are called atoms.

The read syntax and printed representation for lists are identical, and consist of a left parenthesis, an arbitrary number of elements, and a right parenthesis.

Upon reading, each object inside the parentheses becomes an element of the list. That is, a cons cell is made for each element. The car of the cons cell points to the element, and its cdr points to the next cons cell of the list, which holds the next element in the list. The cdr of the last cons cell is set to point to nil.

A list can be illustrated by a diagram in which the cons cells are shown as pairs of boxes. (The Lisp reader cannot read such an illustration; unlike the textual notation, which can be understood by both humans and computers, the box illustrations can be understood only by humans.) The following represents the three-element list (rose violet buttercup):

         __________________
        |___|___|--> |___|___|--> |___|___|--> nil
          |            |            |
          |            |            |
           --> rose     --> violet   --> buttercup

In this diagram, each box represents a slot that can refer to any Lisp object. Each pair of boxes represents a cons cell. Each arrow is a reference to a Lisp object, either an atom or another cons cell.

In this example, the first box, the car of the first cons cell, refers to or “contains” rose (a symbol). The second box, the cdr of the first cons cell, refers to the next pair of boxes, the second cons cell. The car of the second cons cell refers to violet and the cdr refers to the third cons cell. The cdr of the third (and last) cons cell refers to nil.

Here is another diagram of the same list, (rose violet buttercup), sketched in a different manner:

      ---------------       ----------------       -------------------
     | car   | cdr   |     | car    | cdr   |     | car       | cdr   |
     | rose  |   o-------->| violet |   o-------->| buttercup |  nil  |
     |       |       |     |        |       |     |           |       |
      ---------------       ----------------       -------------------

A list with no elements in it is the empty list; it is identical to the symbol nil. In other words, nil is both a symbol and a list.

Here are examples of lists written in Lisp syntax:

(A 2 "A") ; A list of three elements.

() ; A list of no elements (the empty list). nil ; A list of no elements (the empty list). ("A ()") ; A list of one element: the string "A ()". (A ()) ; A list of two elements: A and the empty list. (A nil) ; Equivalent to the previous. ((A B C)) ; A list of one element

; (which is a list of three elements).

Here is the list (A ()), or equivalently (A nil), depicted with boxes and arrows:

         ____________
        |___|___|--> |___|___|--> nil
          |            |
          |            |
           --> A        --> nil

2.4.6.1 Dotted Pair Notation

Dotted pair notation is an alternative syntax for cons cells that represents the car and cdr explicitly. In this syntax, (a . b) stands for a cons cell whose car is the object a, and whose cdr is the object b. Dotted pair notation is therefore more general than list syntax. In the dotted pair notation, the list ‘(1 2 3)’ is written as ‘(1 . (2 . (3 . nil)))’. For nil-terminated lists, the two notations produce the same result, but list notation is usually clearer and more convenient when it is applicable. When printing a list, the dotted pair notation is only used if the cdr of a cell is not a list.

Here’s how box notation can illustrate dotted pairs. This example shows the pair (rose

. violet):
         ______
        |___|___|--> violet
          |
          |
           --> rose

Dotted pair notation can be combined with list notation to represent a chain of cons cells with a non-nil final cdr. For example, (rose violet . buttercup) is equivalent to

(rose . (violet . buttercup)). The object looks like this:
         ____________
        |___|___|--> |___|___|--> buttercup
          |            |
          |            |
           --> rose     --> violet

These diagrams make it evident why (rose . violet . buttercup) is invalid syntax; it would require a cons cell that has three parts rather than two.

The list (rose violet) is equivalent to (rose . (violet)) and looks like this:

         ____________
        |___|___|--> |___|___|--> nil
          |            |
          |            |
           --> rose     --> violet

Similarly, the three-element list (rose violet buttercup) is equivalent to (rose .

(violet . (buttercup))).

2.4.6.2 Association List Type

An association list or alist is a specially-constructed list whose elements are cons cells. In each element, the car is considered a key, and the cdr is considered an associated value.

(In some cases, the associated value is stored in the car of the cdr.) Association lists are often used as stacks, since it is easy to add or remove associations at the front of the list.

For example,

     (setq alist-of-colors
           '((rose . red) (lily . white)  (buttercup . yellow)))

sets the variable alist-of-colors to an alist of three elements. In the first element, rose is the key and red is the value.

See Section 5.8 [Association Lists], page 92, for a further explanation of alists and for functions that work on alists.

2.4.7 Array Type

An array is composed of an arbitrary number of slots for referring to other Lisp objects, arranged in a contiguous block of memory. Accessing any element of an array takes the same amount of time. In contrast, accessing an element of a list requires time proportional to the position of the element in the list. (Elements at the end of a list take longer to access than elements at the beginning of a list.)

XEmacs defines three types of array, strings, vectors, and bit vectors. A string is an array of characters, a vector is an array of arbitrary objects, and a bit vector is an array of 1’s and 0’s. All are one-dimensional. (Most other programming languages support multidimensional arrays, but they are not essential; you can get the same effect with an array of arrays.) Each type of array has its own read syntax; see Section 2.4.8 [String Type], page 25, Section 2.4.9 [Vector Type], page 26, and Section 2.4.10 [Bit Vector Type], page 26.

An array may have any length up to the largest integer; but once created, it has a fixed size. The first element of an array has index zero, the second element has index 1, and so on. This is called zero-origin indexing. For example, an array of four elements has indices 0, 1, 2, and 3.

The array type is contained in the sequence type and contains the string type, the vector type, and the bit vector type.

2.4.8 String Type

A string is an array of characters. Strings are used for many purposes in XEmacs, as can be expected in a text editor; for example, as the names of Lisp symbols, as messages for the user, and to represent text extracted from buffers. Strings in Lisp are constants: evaluation of a string returns the same string.

The read syntax for strings is a double-quote, an arbitrary number of characters, and another double-quote, "like this". The Lisp reader accepts the same formats for reading the characters of a string as it does for reading single characters (without the question mark that begins a character literal). You can enter a nonprinting character such as tab or C-a using the convenient escape sequences, like this: "\t, \C-a". You can include a double-quote in a string by preceding it with a backslash; thus, "\"" is a string containing just a single double-quote character. (See Section 2.4.3 [Character Type], page 18, for a description of the read syntax for characters.)

The printed representation of a string consists of a double-quote, the characters it contains, and another double-quote. However, you must escape any backslash or double-quote characters in the string with a backslash, like this: "this \" is an embedded quote".

The newline character is not special in the read syntax for strings; if you write a new line between the double-quotes, it becomes a character in the string. But an escaped newline— one that is preceded by ‘\’—does not become part of the string; i.e., the Lisp reader ignores an escaped newline while reading a string.

     "It is useful to include newlines
     in documentation strings,
     but the newline is \
     ignored if escaped."
          ⇒ "It is useful to include newlines
     in documentation strings,
     but the newline is ignored if escaped."

A string can hold extents and properties of the text it contains, in addition to the characters themselves. This enables programs that copy text between strings and buffers to preserve the extents and properties with no special effort. See Chapter 40 [Extents], page 585, See Section 36.18 [Text Properties], page 539.

Note that FSF GNU Emacs has a special read and print syntax for strings with text properties, but XEmacs does not currently implement this. It was judged better not to include this in XEmacs because it entails that equal return nil when passed a string with text properties and the equivalent string without text properties, which is often counterintuitive.

See Chapter 4 [Strings and Characters], page 59, for functions that work on strings.

2.4.9 Vector Type

A vector is a one-dimensional array of elements of any type. It takes a constant amount of time to access any element of a vector. (In a list, the access time of an element is proportional to the distance of the element from the beginning of the list.)

The printed representation of a vector consists of a left square bracket, the elements, and a right square bracket. This is also the read syntax. Like numbers and strings, vectors are considered constants for evaluation.

[1 "two" (three)] ; A vector of three elements.

          ⇒ [1 "two" (three)]

See Section 6.4 [Vectors], page 106, for functions that work with vectors.

2.4.10 Bit Vector Type

A bit vector is a one-dimensional array of 1’s and 0’s. It takes a constant amount of time to access any element of a bit vector, as for vectors. Bit vectors have an extremely compact internal representation (one machine bit per element), which makes them ideal for keeping track of unordered sets, large collections of boolean values, etc.

The printed representation of a bit vector consists of ‘#*’ followed by the bits in the vector. This is also the read syntax. Like numbers, strings, and vectors, bit vectors are considered constants for evaluation.

#*00101000 ; A bit vector of eight elements.

          ⇒ #*00101000

See Section 6.6 [Bit Vectors], page 107, for functions that work with bit vectors.

2.4.11 Function Type

Just as functions in other programming languages are executable, Lisp function objects are pieces of executable code. However, functions in Lisp are primarily Lisp objects, and only secondarily the text which represents them. These Lisp objects are lambda expressions: lists whose first element is the symbol lambda (see Section 11.2 [Lambda Expressions], page 164).

In most programming languages, it is impossible to have a function without a name. In Lisp, a function has no intrinsic name. A lambda expression is also called an anonymous function (see Section 11.7 [Anonymous Functions], page 172). A named function in Lisp is actually a symbol with a valid function in its function cell (see Section 11.4 [Defining Functions], page 168).

Most of the time, functions are called when their names are written in Lisp expressions in Lisp programs. However, you can construct or obtain a function object at run time and then call it with the primitive functions funcall and apply. See Section 11.5 [Calling Functions], page 169.

2.4.12 Macro Type

A Lisp macro is a user-defined construct that extends the Lisp language. It is represented as an object much like a function, but with different parameter-passing semantics. A Lisp macro has the form of a list whose first element is the symbol macro and whose cdr is a Lisp function object, including the lambda symbol.

Lisp macro objects are usually defined with the built-in defmacro function, but any list that begins with macro is a macro as far as XEmacs is concerned. See Chapter 12 [Macros], page 177, for an explanation of how to write a macro.

2.4.13 Primitive Function Type

A primitive function is a function callable from Lisp but written in the C programming language. Primitive functions are also called subrs or built-in functions. (The word “subr” is derived from “subroutine”.) Most primitive functions evaluate all their arguments when they are called. A primitive function that does not evaluate all its arguments is called a special form (see Section 8.2.7 [Special Forms], page 125).

It does not matter to the caller of a function whether the function is primitive. However, this does matter if you try to substitute a function written in Lisp for a primitive of the same name. The reason is that the primitive function may be called directly from C code. Calls to the redefined function from Lisp will use the new definition, but calls from C code may still use the built-in definition.

The term function refers to all Emacs functions, whether written in Lisp or C. See Section 2.4.11 [Function Type], page 27, for information about the functions written in Lisp.

Primitive functions have no read syntax and print in hash notation with the name of the subroutine.

     (symbol-function 'car)          ; Access the function cell

; of the symbol.

          ⇒ #<subr car>
     (subrp (symbol-function 'car))  ; Is this a primitive function?

⇒ t ; Yes.

2.4.14 Compiled-Function Type

The byte compiler produces compiled-function objects. The evaluator handles this data type specially when it appears as a function to be called. See Chapter 15 [Byte Compilation], page 205, for information about the byte compiler.

The printed representation for a compiled-function object is normally ‘#<compiled-function...>’. If print-readably is true, however, it is ‘#[...]’.

2.4.15 Autoload Type

An autoload object is a list whose first element is the symbol autoload. It is stored as the function definition of a symbol as a placeholder for the real definition; it says that the real definition is found in a file of Lisp code that should be loaded when necessary. The autoload object contains the name of the file, plus some other information about the real definition.

After the file has been loaded, the symbol should have a new function definition that is not an autoload object. The new definition is then called as if it had been there to begin with. From the user’s point of view, the function call works as expected, using the function definition in the loaded file.

An autoload object is usually created with the function autoload, which stores the object in the function cell of a symbol. See Section 14.2 [Autoload], page 198, for more details.

2.4.16 Char Table Type

(not yet documented)

2.4.17 Hash Table Type

A hash table is a table providing an arbitrary mapping from one Lisp object to another, using an internal indexing method called hashing. Hash tables are very fast (much more efficient that using an association list, when there are a large number of elements in the table).

Hash tables have no read syntax. They print in hash notation (The “hash” in “hash notation” has nothing to do with the “hash” in “hash table”), giving the number of elements, total space allocated for elements, and a unique number assigned at the time the hash table was created. (Hash tables automatically resize as necessary so there is no danger of running out of space for elements.)

     (make-hashtable 50)
          ⇒ #<hashtable 0/71 0x313a>

See Chapter 46 [Hash Tables], page 665, for information on how to create and work with hash tables.

2.4.18 Range Table Type

A range table is a table that maps from ranges of integers to arbitrary Lisp objects. Range tables automatically combine overlapping ranges that map to the same Lisp object, and operations are provided for mapping over all of the ranges in a range table.

Range tables have a special read syntax beginning with ‘#s(range-table’ (this is an example of structure read syntax, which is also used for char tables and faces).

     (setq x (make-range-table))
     (put-range-table 20 50 'foo x)
     (put-range-table 100 200 "bar" x)
     x
          ⇒ #s(range-table data ((20 50) foo (100 200) "bar"))

See Chapter 47 [Range Tables], page 669, for information on how to create and work with range tables.

2.4.19 Weak List Type

(not yet documented)

2.5 Editing Types

The types in the previous section are common to many Lisp dialects. XEmacs Lisp provides several additional data types for purposes connected with editing.

2.5.1 Buffer Type

A buffer is an object that holds text that can be edited (see Chapter 30 [Buffers], page 431). Most buffers hold the contents of a disk file (see Chapter 28 [Files], page 391) so they can be edited, but some are used for other purposes. Most buffers are also meant to be seen by the user, and therefore displayed, at some time, in a window (see Chapter 31 [Windows], page 445). But a buffer need not be displayed in any window.

The contents of a buffer are much like a string, but buffers are not used like strings in XEmacs Lisp, and the available operations are different. For example, insertion of text into a buffer is very efficient, whereas “inserting” text into a string requires concatenating substrings, and the result is an entirely new string object.

Each buffer has a designated position called point (see Chapter 34 [Positions], page 487). At any time, one buffer is the current buffer. Most editing commands act on the contents of the current buffer in the neighborhood of point. Many of the standard Emacs functions manipulate or test the characters in the current buffer; a whole chapter in this manual is devoted to describing these functions (see Chapter 36 [Text], page 511).

Several other data structures are associated with each buffer:

• a local syntax table (see Chapter 38 [Syntax Tables], page 569);

• a local keymap (see Chapter 20 [Keymaps], page 315);

• a local variable binding list (see Section 10.9 [Buffer-Local Variables], page 157);

• a list of extents (see Chapter 40 [Extents], page 585);

• and various other related properties.

The local keymap and variable list contain entries that individually override global bindings or values. These are used to customize the behavior of programs in different buffers, without actually changing the programs.

A buffer may be indirect, which means it shares the text of another buffer. See Section 30.11 [Indirect Buffers], page 442.

Buffers have no read syntax. They print in hash notation, showing the buffer name.

     (current-buffer)
          ⇒ #<buffer "objects.texi">

2.5.2 Marker Type

A marker denotes a position in a specific buffer. Markers therefore have two components: one for the buffer, and one for the position. Changes in the buffer’s text automatically relocate the position value as necessary to ensure that the marker always points between the same two characters in the buffer.

Markers have no read syntax. They print in hash notation, giving the current character position and the name of the buffer.

     (point-marker)
          ⇒ #<marker at 50661 in objects.texi>

See Chapter 35 [Markers], page 499, for information on how to test, create, copy, and move markers.

2.5.3 Extent Type

An extent specifies temporary alteration of the display appearance of a part of a buffer (or string). It contains markers delimiting a range of the buffer, plus a property list (a list whose elements are alternating property names and values). Extents are used to present parts of the buffer temporarily in a different display style. They have no read syntax, and print in hash notation, giving the buffer name and range of positions.

Extents can exist over strings as well as buffers; the primary use of this is to preserve extent and text property information as text is copied from one buffer to another or between different parts of a buffer.

Extents have no read syntax. They print in hash notation, giving the range of text they cover, the name of the buffer or string they are in, the address in core, and a summary of some of the properties attached to the extent.

     (extent-at (point))
          ⇒ #<extent [51742, 51748) font-lock text-prop 0x90121e0 in buffer objects.texi>

See Chapter 40 [Extents], page 585, for how to create and use extents.

Extents are used to implement text properties. See Section 36.18 [Text Properties], page 539.

2.5.4 Window Type

A window describes the portion of the frame that XEmacs uses to display a buffer. (In standard window-system usage, a window is what XEmacs calls a frame; XEmacs confusingly uses the term “window” to refer to what is called a pane in standard windowsystem usage.) Every window has one associated buffer, whose contents appear in the window. By contrast, a given buffer may appear in one window, no window, or several windows.

Though many windows may exist simultaneously, at any time one window is designated the selected window. This is the window where the cursor is (usually) displayed when XEmacs is ready for a command. The selected window usually displays the current buffer, but this is not necessarily the case.

Windows are grouped on the screen into frames; each window belongs to one and only one frame. See Section 2.5.5 [Frame Type], page 31.

Windows have no read syntax. They print in hash notation, giving the name of the buffer being displayed and a unique number assigned at the time the window was created. (This number can be useful because the buffer displayed in any given window can change frequently.)

     (selected-window)
          ⇒ #<window on "objects.texi" 0x266c>

See Chapter 31 [Windows], page 445, for a description of the functions that work on windows.

2.5.5 Frame Type

A frame is a rectangle on the screen (a window in standard window-system terminology) that contains one or more non-overlapping Emacs windows (panes in standard windowsystem terminology). A frame initially contains a single main window (plus perhaps a minibuffer window) which you can subdivide vertically or horizontally into smaller windows.

Frames have no read syntax. They print in hash notation, giving the frame’s type, name as used for resourcing, and a unique number assigned at the time the frame was created.

     (selected-frame)
          ⇒ #<x-frame "emacs" 0x9db>

See Chapter 32 [Frames], page 471, for a description of the functions that work on frames.

2.5.6 Device Type

A device represents a single display on which frames exist. Normally, there is only one device object, but there may be more than one if XEmacs is being run on a multiheaded display (e.g. an X server with attached color and mono screens) or if XEmacs is simultaneously driving frames attached to different consoles, e.g. an X display and a TTY connection.

Devices do not have a read syntax. They print in hash notation, giving the device’s type, connection name, and a unique number assigned at the time the device was created.

     (selected-device)
          ⇒ #<x-device on ":0.0" 0x5b9>

See Chapter 33 [Consoles and Devices], page 483, for a description of several functions related to devices.

2.5.7 Console Type

A console represents a single keyboard to which devices (i.e. displays on which frames exist) are connected. Normally, there is only one console object, but there may be more than one if XEmacs is simultaneously driving frames attached to different X servers and/or TTY connections. (XEmacs is capable of driving multiple X and TTY connections at the same time, and provides a robust mechanism for handling the differing display capabilities of such heterogeneous environments. A buffer with embedded glyphs and multiple fonts and colors, for example, will display reasonably if it simultaneously appears on a frame on a color X display, a frame on a mono X display, and a frame on a TTY connection.)

Consoles do not have a read syntax. They print in hash notation, giving the console’s type, connection name, and a unique number assigned at the time the console was created.

     (selected-console)
          ⇒ #<x-console on "localhost:0" 0x5b7>

See Chapter 33 [Consoles and Devices], page 483, for a description of several functions related to consoles.

2.5.8 Window Configuration Type

A window configuration stores information about the positions, sizes, and contents of the windows in a frame, so you can recreate the same arrangement of windows later.

Window configurations do not have a read syntax. They print in hash notation, giving a unique number assigned at the time the window configuration was created.

     (current-window-configuration)
          ⇒ #<window-configuration 0x2db4>

See Section 31.16 [Window Configurations], page 468, for a description of several functions related to window configurations.

2.5.9 Event Type

(not yet documented)

2.5.10 Process Type

The word process usually means a running program. XEmacs itself runs in a process of this sort. However, in XEmacs Lisp, a process is a Lisp object that designates a subprocess created by the XEmacs process. Programs such as shells, GDB, ftp, and compilers, running in subprocesses of XEmacs, extend the capabilities of XEmacs.

An Emacs subprocess takes textual input from Emacs and returns textual output to Emacs for further manipulation. Emacs can also send signals to the subprocess.

Process objects have no read syntax. They print in hash notation, giving the name of the process, its associated process ID, and the current state of the process:

     (process-list)
          ⇒ (#<process "shell" pid 2909 state:run>)

See Chapter 49 [Processes], page 673, for information about functions that create, delete, return information about, send input or signals to, and receive output from processes.

2.5.11 Stream Type

A stream is an object that can be used as a source or sink for characters—either to supply characters for input or to accept them as output. Many different types can be used this way: markers, buffers, strings, and functions. Most often, input streams (character sources) obtain characters from the keyboard, a buffer, or a file, and output streams (character sinks) send characters to a buffer, such as a ‘*Help*’ buffer, or to the echo area.

The object nil, in addition to its other meanings, may be used as a stream. It stands for the value of the variable standard-input or standard-output. Also, the object t as a stream specifies input using the minibuffer (see Chapter 18 [Minibuffers], page 261) or output in the echo area (see Section 45.3 [The Echo Area], page 648).

Streams have no special printed representation or read syntax, and print as whatever primitive type they are.

See Chapter 17 [Read and Print], page 251, for a description of functions related to streams, including parsing and printing functions.

2.5.12 Keymap Type

A keymap maps keys typed by the user to commands. This mapping controls how the user’s command input is executed.

NOTE: In XEmacs, a keymap is a separate primitive type. In FSF GNU Emacs, a keymap is actually a list whose car is the symbol keymap.

See Chapter 20 [Keymaps], page 315, for information about creating keymaps, handling prefix keys, local as well as global keymaps, and changing key bindings.

2.5.13 Syntax Table Type

Under XEmacs 20, a syntax table is a particular type of char table. Under XEmacs 19, a syntax table a vector of 256 integers. In both cases, each element defines how one character is interpreted when it appears in a buffer. For example, in C mode (see Section 26.1 [Major Modes], page 361), the ‘+’ character is punctuation, but in Lisp mode it is a valid character in a symbol. These modes specify different interpretations by changing the syntax table entry for ‘+’.

Syntax tables are used only for scanning text in buffers, not for reading Lisp expressions. The table the Lisp interpreter uses to read expressions is built into the XEmacs source code and cannot be changed; thus, to change the list delimiters to be ‘{’ and ‘}’ instead of ‘(’ and ‘)’ would be impossible.

See Chapter 38 [Syntax Tables], page 569, for details about syntax classes and how to make and modify syntax tables.

2.5.14 Display Table Type

A display table specifies how to display each character code. Each buffer and each window can have its own display table. A display table is actually a vector of length 256, although in XEmacs 20 this may change to be a particular type of char table. See Section 45.11 [Display Tables], page 659.

2.5.15 Database Type

(not yet documented)

2.5.16 Charset Type

(not yet documented)

2.5.17 Coding System Type

(not yet documented)

2.5.18 ToolTalk Message Type

(not yet documented)

2.5.19 ToolTalk Pattern Type

(not yet documented)

2.6 Window-System Types

XEmacs also has some types that represent objects such as faces (collections of display characters), fonts, and pixmaps that are commonly found in windowing systems.

2.6.1 Face Type

(not yet documented)

2.6.2 Glyph Type

(not yet documented)

2.6.3 Specifier Type

(not yet documented)

2.6.4 Font Instance Type

(not yet documented)

2.6.5 Color Instance Type

(not yet documented)

2.6.6 Image Instance Type

(not yet documented)

2.6.7 Toolbar Button Type

(not yet documented)

2.6.8 Subwindow Type

(not yet documented)

2.6.9 X Resource Type

(not yet documented)

2.7 Type Predicates

The XEmacs Lisp interpreter itself does not perform type checking on the actual arguments passed to functions when they are called. It could not do so, since function arguments in Lisp do not have declared data types, as they do in other programming languages. It is therefore up to the individual function to test whether each actual argument belongs to a type that the function can use.

All built-in functions do check the types of their actual arguments when appropriate, and signal a wrong-type-argument error if an argument is of the wrong type. For example, here is what happens if you pass an argument to + that it cannot handle:

     (+ 2 'a)
           error Wrong type argument: integer-or-marker-p, a

If you want your program to handle different types differently, you must do explicit type checking. The most common way to check the type of an object is to call a type predicate function. Emacs has a type predicate for each type, as well as some predicates for combinations of types.

A type predicate function takes one argument; it returns t if the argument belongs to the appropriate type, and nil otherwise. Following a general Lisp convention for predicate functions, most type predicates’ names end with ‘p’.

Here is an example which uses the predicates listp to check for a list and symbolp to check for a symbol.

     (defun add-on (x)
       (cond ((symbolp x)
              ;; If X is a symbol, put it on LIST.
              (setq list (cons x list)))
             ((listp x)

              ;; If X is a list, add its elements to LIST.
              (setq list (append x list)))
             (t
              ;; We only handle symbols and lists.
              (error "Invalid argument %s in add-on" x))))

Here is a table of predefined type predicates, in alphabetical order, with references to further information.

annotationp

See Section 44.2 [Annotation Primitives], page 642.

arrayp See Section 6.3 [Array Functions], page 104. atom See Section 5.3 [List-related Predicates], page 78.

bit-vector-p

See Section 6.7 [Bit Vector Functions], page 108.

bitp See Section 6.7 [Bit Vector Functions], page 108.

boolean-specifier-p

See Section 41.4 [Specifier Types], page 602.

buffer-glyph-p

See Section 43.3 [Glyph Types], page 638.

buffer-live-p

See Section 30.10 [Killing Buffers], page 441.

bufferp See Section 30.1 [Buffer Basics], page 431.

button-event-p

See Section 19.5.3 [Event Predicates], page 294.

button-press-event-p

See Section 19.5.3 [Event Predicates], page 294.

button-release-event-p

See Section 19.5.3 [Event Predicates], page 294.

case-table-p

See Section 4.12 [Case Tables], page 71.

char-int-p

See Section 4.5 [Character Codes], page 62.

char-or-char-int-p

See Section 4.5 [Character Codes], page 62.

char-or-string-p

See Section 4.2 [Predicates for Strings], page 60.

char-table-p

See Section 4.13 [Char Tables], page 73.

characterp

See Section 4.4 [Predicates for Characters], page 62.

color-instance-p

See Section 42.3 [Colors], page 623.

color-pixmap-image-instance-p

See Section 43.2.3.1 [Image Instance Types], page 635.

color-specifier-p

See Section 41.4 [Specifier Types], page 602.

commandp See Section 19.3 [Interactive Call], page 286.

compiled-function-p

See Section 2.4.14 [Compiled-Function Type], page 28.

console-live-p

See Section 33.4 [Connecting to a Console or Device], page 485.

consolep See Chapter 33 [Consoles and Devices], page 483. consp See Section 5.3 [List-related Predicates], page 78.

database-live-p

See Section 48.1 [Connecting to a Database], page 671.

databasep

See Chapter 48 [Databases], page 671.

device-live-p

See Section 33.4 [Connecting to a Console or Device], page 485.

device-or-frame-p

See Section 33.2 [Basic Device Functions], page 484.

devicep See Chapter 33 [Consoles and Devices], page 483.

eval-event-p

See Section 19.5.3 [Event Predicates], page 294.

event-live-p

See Section 19.5.3 [Event Predicates], page 294.

eventp See Section 19.5 [Events], page 290.

extent-live-p

See Section 40.2 [Creating and Modifying Extents], page 586.

extentp See Chapter 40 [Extents], page 585.

face-boolean-specifier-p

See Section 41.4 [Specifier Types], page 602.

facep See Section 42.1.2 [Basic Face Functions], page 616. floatp See Section 3.3 [Predicates on Numbers], page 46.

font-instance-p

See Section 42.2 [Fonts], page 621.

font-specifier-p

See Section 41.4 [Specifier Types], page 602.

frame-live-p

See Section 32.4 [Deleting Frames], page 476.

framep See Chapter 32 [Frames], page 471.

functionp

(not yet documented)

generic-specifier-p

See Section 41.4 [Specifier Types], page 602.

glyphp See Chapter 43 [Glyphs], page 625.

hashtablep

See Chapter 46 [Hash Tables], page 665.

icon-glyph-p

See Section 43.3 [Glyph Types], page 638.

image-instance-p

See Section 43.2 [Images], page 630.

image-specifier-p

See Section 41.4 [Specifier Types], page 602.

integer-char-or-marker-p

See Section 35.2 [Predicates on Markers], page 500.

integer-or-char-p

See Section 4.4 [Predicates for Characters], page 62.

integer-or-marker-p

See Section 35.2 [Predicates on Markers], page 500.

integer-specifier-p

See Section 41.4 [Specifier Types], page 602.

integerp See Section 3.3 [Predicates on Numbers], page 46. itimerp (not yet documented)

key-press-event-p

See Section 19.5.3 [Event Predicates], page 294.

keymapp See Section 20.3 [Creating Keymaps], page 316. keywordp (not yet documented) listp See Section 5.3 [List-related Predicates], page 78. markerp See Section 35.2 [Predicates on Markers], page 500.

misc-user-event-p

See Section 19.5.3 [Event Predicates], page 294.

mono-pixmap-image-instance-p

See Section 43.2.3.1 [Image Instance Types], page 635.

motion-event-p

See Section 19.5.3 [Event Predicates], page 294.

mouse-event-p

See Section 19.5.3 [Event Predicates], page 294.

natnum-specifier-p

See Section 41.4 [Specifier Types], page 602.

natnump See Section 3.3 [Predicates on Numbers], page 46. nlistp See Section 5.3 [List-related Predicates], page 78.

nothing-image-instance-p

See Section 43.2.3.1 [Image Instance Types], page 635.

number-char-or-marker-p

See Section 35.2 [Predicates on Markers], page 500.

number-or-marker-p

See Section 35.2 [Predicates on Markers], page 500.

numberp See Section 3.3 [Predicates on Numbers], page 46.

pointer-glyph-p

See Section 43.3 [Glyph Types], page 638.

pointer-image-instance-p

See Section 43.2.3.1 [Image Instance Types], page 635.

process-event-p

See Section 19.5.3 [Event Predicates], page 294.

processp See Chapter 49 [Processes], page 673.

range-table-p

See Chapter 47 [Range Tables], page 669.

ringp (not yet documented)

sequencep

See Section 6.1 [Sequence Functions], page 101.

specifierp

See Chapter 41 [Specifiers], page 599.

stringp See Section 4.2 [Predicates for Strings], page 60. subrp See Section 11.8 [Function Cells], page 173.

subwindow-image-instance-p

See Section 43.2.3.1 [Image Instance Types], page 635.

subwindowp

See Section 43.6 [Subwindows], page 640.

symbolp See Chapter 7 [Symbols], page 111.

syntax-table-p

See Chapter 38 [Syntax Tables], page 569.

text-image-instance-p

See Section 43.2.3.1 [Image Instance Types], page 635.

timeout-event-p

See Section 19.5.3 [Event Predicates], page 294.

toolbar-button-p

See Chapter 23 [Toolbar], page 351.

toolbar-specifier-p

See Chapter 23 [Toolbar], page 351.

user-variable-p

See Section 10.5 [Defining Variables], page 149.

vectorp See Section 6.4 [Vectors], page 106.

weak-list-p

See Section 5.10 [Weak Lists], page 98.

window-configuration-p

See Section 31.16 [Window Configurations], page 468.

window-live-p

See Section 31.3 [Deleting Windows], page 449.

windowp See Section 31.1 [Basic Windows], page 445.

The most general way to check the type of an object is to call the function type-of. Recall that each object belongs to one and only one primitive type; type-of tells you which one (see Chapter 2 [Lisp Data Types], page 15). But type-of knows nothing about nonprimitive types. In most cases, it is more convenient to use type predicates than type-of.

Function type-of object

This function returns a symbol naming the primitive type of object. The value

     is one of bit-vector, buffer, char-table, character, charset, coding-
     system, cons, color-instance, compiled-function, console, database,
     device, event, extent, face, float, font-instance, frame, glyph, hashtable,
     image-instance, integer, keymap, marker, process, range-table, specifier,
     string, subr, subwindow, symbol, toolbar-button, tooltalk-message,
     tooltalk-pattern, vector, weak-list, window, window-configuration, or
     x-resource.
          (type-of 1)
               ⇒ integer
          (type-of 'nil)
               ⇒ symbol

     (type-of '())    ; () is nil.
          ⇒ symbol
     (type-of '(x))
          ⇒ cons

2.8 Equality Predicates

Here we describe two functions that test for equality between any two objects. Other functions test equality between objects of specific types, e.g., strings. For these predicates, see the appropriate chapter describing the data type.

Function eq object1 object2 This function returns t if object1 and object2 are the same object, nil otherwise. The “same object” means that a change in one will be reflected by the same change in the other. eq returns t if object1 and object2 are integers with the same value. Also, since symbol names are normally unique, if the arguments are symbols with the same name, they are eq. For other types (e.g., lists, vectors, strings), two arguments with the same contents or elements are not necessarily eq to each other: they are eq only if they are the same object. (The make-symbol function returns an uninterned symbol that is not interned in the standard obarray. When uninterned symbols are in use, symbol names are no longer unique. Distinct symbols with the same name are not eq. See Section 7.3 [Creating Symbols], page 113.) NOTE: Under XEmacs 19, characters are really just integers, and thus characters and integers are eq. Under XEmacs 20, it was necessary to preserve remnants of this in function such as old-eq in order to maintain byte-code compatibility. Byte code compiled under any Emacs 19 will automatically have calls to eq mapped to old-eq when executed under XEmacs 20.

     (eq 'foo 'foo)
          ⇒ t
     (eq 456 456) ⇒ t
     (eq "asdf" "asdf")
          ⇒ nil
     (eq '(1 (2 (3))) '(1 (2 (3)))) ⇒ nil
     (setq foo '(1 (2 (3))))
          ⇒ (1 (2 (3)))
     (eq foo foo)
          ⇒ t
     (eq foo '(1 (2 (3))))
          ⇒ nil
     (eq [(1 2) 3] [(1 2) 3]) ⇒ nil

(eq (point-marker) (point-marker))
     ⇒ nil

Function old-eq obj1 obj2 This function exists under XEmacs 20 and is exactly like eq except that it suffers from the char-int confoundance disease. In other words, it returns t if given a character and the equivalent integer, even though the objects are of different types! You should not ever call this function explicitly in your code. However, be aware that all calls to eq in byte code compiled under version 19 map to old-

eq in XEmacs 20. (Likewise for old-equal, old-memq, old-member, old-assq
and old-assoc.)

;; Remember, this does not apply under XEmacs 19.

?A

     ⇒ ?A

(char-int ?A)
     ⇒ 65
(old-eq ?A 65)

⇒ t ; Eek, we’ve been infected.

(eq ?A 65)

⇒ nil ; We are still healthy.

Function equal object1 object2 This function returns t if object1 and object2 have equal components, nil otherwise. Whereas eq tests if its arguments are the same object, equal looks inside nonidentical arguments to see if their elements are the same. So, if two objects are eq, they are equal, but the converse is not always true.

(equal 'foo 'foo)
     ⇒ t
(equal 456 456) ⇒ t
(equal "asdf" "asdf")
     ⇒ t
(eq "asdf" "asdf")
     ⇒ nil
(equal '(1 (2 (3))) '(1 (2 (3)))) ⇒ t
(eq '(1 (2 (3))) '(1 (2 (3)))) ⇒ nil
(equal [(1 2) 3] [(1 2) 3]) ⇒ t
(eq [(1 2) 3] [(1 2) 3]) ⇒ nil
(equal (point-marker) (point-marker))
     ⇒ t

     (eq (point-marker) (point-marker))
          ⇒ nil

Comparison of strings is case-sensitive. Note that in FSF GNU Emacs, comparison of strings takes into account their text properties, and you have to use string-equal if you want only the strings themselves compared. This difference does not exist in XEmacs; equal and string-equal always return the same value on the same strings.

     (equal "asdf" "ASDF")
          ⇒ nil

Two distinct buffers are never equal, even if their contents are the same.

The test for equality is implemented recursively, and circular lists may therefore cause infinite recursion (leading to an error).

44 XEmacs Lisp Reference Manual

3 Numbers

XEmacs supports two numeric data types: integers and floating point numbers. Integers are whole numbers such as −3, 0, #b0111, #xFEED, #o744. Their values are exact. The number prefixes ‘#b’, ‘#o’, and ‘#x’ are supported to represent numbers in binary, octal, and hexadecimal notation (or radix). Floating point numbers are numbers with fractional parts, such as −4.5, 0.0, or 2.71828. They can also be expressed in exponential notation: 1.5e2 equals 150; in this example, ‘e2’ stands for ten to the second power, and is multiplied by 1.5. Floating point values are not exact; they have a fixed, limited amount of precision.

3.1 Integer Basics

The range of values for an integer depends on the machine. The minimum range is −134217728 to 134217727 (28 bits; i.e., −2^{27} to 2^{27} − 1), but some machines may provide a wider range. Many examples in this chapter assume an integer has 28 bits.

The Lisp reader reads an integer as a sequence of digits with optional initial sign and optional final period.

1 ; The integer 1.

1. ; The integer 1. +1 ; Also the integer 1. -1 ; The integer −1. 268435457 ; Also the integer 1, due to overflow. 0 ; The integer 0. -0 ; The integer 0.

To understand how various functions work on integers, especially the bitwise operators (see Section 3.8 [Bitwise Operations], page 53), it is often helpful to view the numbers in their binary form.

In 28-bit binary, the decimal integer 5 looks like this:

     0000  0000 0000  0000 0000  0000 0101

(We have inserted spaces between groups of 4 bits, and two spaces between groups of 8 bits, to make the binary integer easier to read.)

The integer −1 looks like this:

     1111  1111 1111  1111 1111  1111 1111

−1 is represented as 28 ones. (This is called two’s complement notation.)

The negative integer, −5, is creating by subtracting 4 from −1. In binary, the decimal integer 4 is 100. Consequently, −5 looks like this:

     1111  1111 1111  1111 1111  1111 1011

In this implementation, the largest 28-bit binary integer is the decimal integer 134,217,727. In binary, it looks like this:

     0111  1111 1111  1111 1111  1111 1111

Since the arithmetic functions do not check whether integers go outside their range, when you add 1 to 134,217,727, the value is the negative integer −134,217,728:

     (+ 1 134217727)
          ⇒ -134217728

          ⇒ 1000  0000 0000  0000 0000  0000 0000

Many of the following functions accept markers for arguments as well as integers. (See Chapter 35 [Markers], page 499.) More precisely, the actual arguments to such functions may be either integers or markers, which is why we often give these arguments the name int-or-marker. When the argument value is a marker, its position value is used and its buffer is ignored.

3.2 Floating Point Basics

XEmacs supports floating point numbers. The precise range of floating point numbers is machine-specific; it is the same as the range of the C data type double on the machine in question.

The printed representation for floating point numbers requires either a decimal point (with at least one digit following), an exponent, or both. For example, ‘1500.0’, ‘15e2’, ‘15.0e2’, ‘1.5e3’, and ‘.15e4’ are five ways of writing a floating point number whose value is 1500. They are all equivalent. You can also use a minus sign to write negative floating point numbers, as in ‘-1.0’.

Most modern computers support the IEEE floating point standard, which provides for positive infinity and negative infinity as floating point values. It also provides for a class of values called NaN or “not-a-number”; numerical functions return such values in cases where there is no correct answer. For example, (sqrt -1.0) returns a NaN. For practical purposes, there’s no significant difference between different NaN values in XEmacs Lisp, and there’s no rule for precisely which NaN value should be used in a particular case, so this manual doesn’t try to distinguish them. XEmacs Lisp has no read syntax for NaNs or infinities; perhaps we should create a syntax in the future.

You can use logb to extract the binary exponent of a floating point number (or estimate the logarithm of an integer):

Function logb number This function returns the binary exponent of number. More precisely, the value is the logarithm of number base 2, rounded down to an integer.

3.3 Type Predicates for Numbers

The functions in this section test whether the argument is a number or whether it is a certain sort of number. The functions integerp and floatp can take any type of Lisp object as argument (the predicates would not be of much use otherwise); but the zerop predicate requires a number as its argument. See also integer-or-marker-p, integer-

char-or-marker-p, number-or-marker-p and number-char-or-marker-p, in Section 35.2

[Predicates on Markers], page 500.

Function floatp object This predicate tests whether its argument is a floating point number and returns t if so, nil otherwise. floatp does not exist in Emacs versions 18 and earlier.

Function integerp object

This predicate tests whether its argument is an integer, and returns t if so, nil otherwise.

Function numberp object

This predicate tests whether its argument is a number (either integer or floating point), and returns t if so, nil otherwise.

Function natnump object

The natnump predicate (whose name comes from the phrase “natural-number-p”) tests to see whether its argument is a nonnegative integer, and returns t if so, nil otherwise. 0 is considered non-negative.

Function zerop number This predicate tests whether its argument is zero, and returns t if so, nil otherwise. The argument must be a number. These two forms are equivalent: (zerop x) ≡ (= x 0).

3.4 Comparison of Numbers

To test numbers for numerical equality, you should normally use =, not eq. There can be many distinct floating point number objects with the same numeric value. If you use eq to compare them, then you test whether two values are the same object. By contrast, = compares only the numeric values of the objects.

At present, each integer value has a unique Lisp object in XEmacs Lisp. Therefore, eq is equivalent to = where integers are concerned. It is sometimes convenient to use eq for comparing an unknown value with an integer, because eq does not report an error if the unknown value is not a number—it accepts arguments of any type. By contrast, = signals an error if the arguments are not numbers or markers. However, it is a good idea to use = if you can, even for comparing integers, just in case we change the representation of integers in a future XEmacs version.

There is another wrinkle: because floating point arithmetic is not exact, it is often a bad idea to check for equality of two floating point values. Usually it is better to test for approximate equality. Here’s a function to do this:

(defconst fuzz-factor 1.0e-6)
(defun approx-equal (x y)
  (or (and (= x 0) (= y 0))
      (< (/ (abs (- x y))
            (max (abs x) (abs y)))
         fuzz-factor)))

Common Lisp note: Comparing numbers in Common Lisp always requires = because Common Lisp implements multi-word integers, and two distinct integer objects can have the same numeric value. XEmacs Lisp can have just one integer object for any given value because it has a limited range of integer values.

In addition to numbers, all of the following functions also accept characters and markers as arguments, and treat them as their number equivalents.

Function = number &rest more-numbers This function returns t if all of its arguments are numerically equal, nil otherwise.

(= 5)
     ⇒ t
(= 5 6)
     ⇒ nil
(= 5 5.0)
     ⇒ t
(= 5 5 6)
     ⇒ nil

Function /= number &rest more-numbers This function returns t if no two arguments are numerically equal, nil otherwise.

(/= 5 6)
     ⇒ t
(/= 5 5 6)
     ⇒ nil
(/= 5 6 1)
     ⇒ t

Function < number &rest more-numbers This function returns t if the sequence of its arguments is monotonically increasing, nil otherwise.

(< 5 6)
     ⇒ t
(< 5 6 6)
     ⇒ nil
(< 5 6 7)
     ⇒ t

Function <= number &rest more-numbers This function returns t if the sequence of its arguments is monotonically nondecreasing, nil otherwise.

(<= 5 6)
     ⇒ t
(<= 5 6 6)
     ⇒ t
(<= 5 6 5)
     ⇒ nil

Function > number &rest more-numbers This function returns t if the sequence of its arguments is monotonically decreasing, nil otherwise.

Function >= number &rest more-numbers This function returns t if the sequence of its arguments is monotonically nonincreasing, nil otherwise.

Function max number &rest more-numbers This function returns the largest of its arguments.

     (max 20)
          ⇒ 20
     (max 1 2.5)
          ⇒ 2.5
     (max 1 3 2.5)
          ⇒ 3

Function min number &rest more-numbers This function returns the smallest of its arguments.

     (min -4 1)
          ⇒ -4

3.5 Numeric Conversions

To convert an integer to floating point, use the function float.

Function float number This returns number converted to floating point. If number is already a floating point number, float returns it unchanged.

There are four functions to convert floating point numbers to integers; they differ in how they round. These functions accept integer arguments also, and return such arguments unchanged.

Function truncate number

This returns number, converted to an integer by rounding towards zero.

Function floor number &optional divisor This returns number, converted to an integer by rounding downward (towards negative infinity).

If divisor is specified, number is divided by divisor before the floor is taken; this is the division operation that corresponds to mod. An arith-error results if divisor is 0.

Function ceiling number

This returns number, converted to an integer by rounding upward (towards positive infinity).

Function round number This returns number, converted to an integer by rounding towards the nearest integer. Rounding a value equidistant between two integers may choose the integer closer to zero, or it may prefer an even integer, depending on your machine.

3.6 Arithmetic Operations

XEmacs Lisp provides the traditional four arithmetic operations: addition, subtraction, multiplication, and division. Remainder and modulus functions supplement the division functions. The functions to add or subtract 1 are provided because they are traditional in Lisp and commonly used.

All of these functions except % return a floating point value if any argument is floating. It is important to note that in XEmacs Lisp, arithmetic functions do not check for

overflow. Thus (1+ 134217727) may evaluate to −134217728, depending on your hardware.

Function 1+ number-or-marker

This function returns number-or-marker plus 1. For example,

     (setq foo 4)
          ⇒ 4
     (1+ foo)
          ⇒ 5

This function is not analogous to the C operator ++—it does not increment a variable. It just computes a sum. Thus, if we continue,

     foo

          ⇒ 4

If you want to increment the variable, you must use setq, like this:

     (setq foo (1+ foo))
          ⇒ 5

Now that the cl package is always available from lisp code, a more convenient and natural way to increment a variable is (incf foo).

Function 1- number-or-marker This function returns number-or-marker minus 1.

Function abs number This returns the absolute value of number.

Function + &rest numbers-or-markers This function adds its arguments together. When given no arguments, + returns 0.

     (+)
          ⇒ 0
     (+ 1)
          ⇒ 1
     (+ 1 2 3 4)
          ⇒ 10

Function - &optional number-or-marker &rest other-numbers-or-markers The - function serves two purposes: negation and subtraction. When - has a single argument, the value is the negative of the argument. When there are multiple arguments, - subtracts each of the other-numbers-or-markers from number-or-marker, cumulatively. If there are no arguments, the result is 0.

     (- 10 1 2 3 4)
          ⇒ 0
     (- 10)
          ⇒ -10
     (-)

          ⇒ 0

Function * &rest numbers-or-markers This function multiplies its arguments together, and returns the product. When given no arguments, * returns 1.

     (*)
          ⇒ 1
     (* 1)
          ⇒ 1
     (* 1 2 3 4)
          ⇒ 24

Function / dividend divisor &rest divisors This function divides dividend by divisor and returns the quotient. If there are additional arguments divisors, then it divides dividend by each divisor in turn. Each argument may be a number or a marker. If all the arguments are integers, then the result is an integer too. This means the result has to be rounded. On most machines, the result is rounded towards zero after each division, but some machines may round differently with negative arguments. This is because the Lisp function / is implemented using the C division operator, which also permits machine-dependent rounding. As a practical matter, all known machines round in the standard fashion. If you divide by 0, an arith-error error is signaled. (See Section 9.5.3 [Errors], page 136.)

     (/ 6 2) ⇒ 3
     (/ 5 2)
          ⇒ 2
     (/ 25 3 2)
          ⇒ 4
     (/ -17 6)
          ⇒ -2

The result of (/ -17 6) could in principle be -3 on some machines.

Function % dividend divisor This function returns the integer remainder after division of dividend by divisor. The arguments must be integers or markers. For negative arguments, the remainder is in principle machine-dependent since the quotient is; but in practice, all known machines behave alike. An arith-error results if divisor is 0.

     (% 9 4)
          ⇒ 1

     (% -9 4)
          ⇒ -1
     (% 9 -4)
          ⇒ 1
     (% -9 -4)
          ⇒ -1

For any two integers dividend and divisor,

(+ (% dividend divisor)

(* (/ dividend divisor) divisor))

always equals dividend.

Function mod dividend divisor This function returns the value of dividend modulo divisor; in other words, the remainder after division of dividend by divisor, but with the same sign as divisor. The arguments must be numbers or markers. Unlike %, mod returns a well-defined result for negative arguments. It also permits floating point arguments; it rounds the quotient downward (towards minus infinity) to an integer, and uses that quotient to compute the remainder. An arith-error results if divisor is 0.

     (mod 9 4) ⇒ 1

     (mod -9 4) ⇒ 3
     (mod 9 -4) ⇒ -3
     (mod -9 -4) ⇒ -1
     (mod 5.5 2.5) ⇒ .5

For any two numbers dividend and divisor,

(+ (mod dividend divisor)

(* (floor dividend divisor) divisor))

always equals dividend, subject to rounding error if either argument is floating point. For floor, see Section 3.5 [Numeric Conversions], page 49.

3.7 Rounding Operations

The functions ffloor, fceiling, fround and ftruncate take a floating point argument and return a floating point result whose value is a nearby integer. ffloor returns the nearest integer below; fceiling, the nearest integer above; ftruncate, the nearest integer in the direction towards zero; fround, the nearest integer.

Function ffloor float

This function rounds float to the next lower integral value, and returns that value as a floating point number.

Function fceiling float

This function rounds float to the next higher integral value, and returns that value as a floating point number.

Function ftruncate float

This function rounds float towards zero to an integral value, and returns that value as a floating point number.

Function fround float

This function rounds float to the nearest integral value, and returns that value as a floating point number.

3.8 Bitwise Operations on Integers

In a computer, an integer is represented as a binary number, a sequence of bits (digits which are either zero or one). A bitwise operation acts on the individual bits of such a sequence. For example, shifting moves the whole sequence left or right one or more places, reproducing the same pattern “moved over”.

The bitwise operations in XEmacs Lisp apply only to integers.

Function lsh integer1 count lsh, which is an abbreviation for logical shift, shifts the bits in integer1 to the left count places, or to the right if count is negative, bringing zeros into the vacated bits. If count is negative, lsh shifts zeros into the leftmost (mostsignificant) bit, producing a positive result even if integer1 is negative. Contrast this with ash, below. Here are two examples of lsh, shifting a pattern of bits one place to the left. We show only the low-order eight bits of the binary pattern; the rest are all zero.

     (lsh 5 1)
          ⇒ 10

;; Decimal 5 becomes decimal 10.

     00000101 ⇒ 00001010
     (lsh 7 1)
          ⇒ 14

;; Decimal 7 becomes decimal 14.

     00000111 ⇒ 00001110

As the examples illustrate, shifting the pattern of bits one place to the left produces a number that is twice the value of the previous number. Shifting a pattern of bits two places to the left produces results like this (with 8-bit binary numbers):

     (lsh 3 2)
          ⇒ 12

;; Decimal 3 becomes decimal 12.

     00000011 ⇒ 00001100

On the other hand, shifting one place to the right looks like this:

     (lsh 6 -1)
          ⇒ 3

;; Decimal 6 becomes decimal 3.

     00000110 ⇒ 00000011
     (lsh 5 -1)
          ⇒ 2

;; Decimal 5 becomes decimal 2.

     00000101 ⇒ 00000010

As the example illustrates, shifting one place to the right divides the value of a positive integer by two, rounding downward. The function lsh, like all XEmacs Lisp arithmetic functions, does not check for overflow, so shifting left can discard significant bits and change the sign of the number. For example, left shifting 134,217,727 produces −2 on a 28-bit machine:

     (lsh 134217727 1)          ; left shift
          ⇒ -2

In binary, in the 28-bit implementation, the argument looks like this:

     ;; Decimal 134,217,727 0111  1111 1111  1111 1111  1111 1111

which becomes the following when left shifted:

     ;; Decimal −2 1111  1111 1111  1111 1111  1111 1110

Function ash integer1 count ash (arithmetic shift) shifts the bits in integer1 to the left count places, or to the right if count is negative. ash gives the same results as lsh except when integer1 and count are both negative. In that case, ash puts ones in the empty bit positions on the left, while lsh puts zeros in those bit positions. Thus, with ash, shifting the pattern of bits one place to the right looks like this:

     (ash -6 -1) ⇒ -3

;; Decimal −6 becomes decimal −3.

     1111  1111 1111  1111 1111  1111 1010

⇒

     1111  1111 1111  1111 1111  1111 1101

In contrast, shifting the pattern of bits one place to the right with lsh looks like this:

     (lsh -6 -1) ⇒ 134217725

;; Decimal −6 becomes decimal 134,217,725.

     1111  1111 1111  1111 1111  1111 1010

⇒

     0111  1111 1111  1111 1111  1111 1101

Here are other examples:

; 28-bit binary values

(lsh 5 2) ; 5 = 0000 0000 0000 0000 0000 0000 0101

⇒ 20 ; = 0000 0000 0000 0000 0000 0001 0100

     (ash 5 2)
          ⇒ 20

(lsh -5 2) ; -5 = 1111 1111 1111 1111 1111 1111 1011

⇒ -20 ; = 1111 1111 1111 1111 1111 1110 1100

     (ash -5 2)
          ⇒ -20

(lsh 5 -2) ; 5 = 0000 0000 0000 0000 0000 0000 0101 ⇒ 1 ; = 0000 0000 0000 0000 0000 0000 0001

     (ash 5 -2) ⇒ 1

(lsh -5 -2) ; -5 = 1111 1111 1111 1111 1111 1111 1011 ⇒ 4194302 ; = 0011 1111 1111 1111 1111 1111 1110

(ash -5 -2) ; -5 = 1111 1111 1111 1111 1111 1111 1011 ⇒ -2 ; = 1111 1111 1111 1111 1111 1111 1110

Function logand &rest ints-or-markers This function returns the “logical and” of the arguments: the nth bit is set in the result if, and only if, the nth bit is set in all the arguments. (“Set” means that the value of the bit is 1 rather than 0.) For example, using 4-bit binary numbers, the “logical and” of 13 and 12 is 12: 1101 combined with 1100 produces 1100. In both the binary numbers, the leftmost two bits are set (i.e., they are 1’s), so the leftmost two bits of the returned value are set. However, for the rightmost two bits, each is zero in at least one of the arguments, so the rightmost two bits of the returned value are 0’s. Therefore,

     (logand 13 12) ⇒ 12

If logand is not passed any argument, it returns a value of −1. This number is an identity element for logand because its binary representation consists entirely of ones. If logand is passed just one argument, it returns that argument.

; 28-bit binary values

(logand 14 13) ; 14 = 0000 0000 0000 0000 0000 0000 1110

; 13 = 0000 0000 0000 0000 0000 0000 1101

⇒ 12 ; 12 = 0000 0000 0000 0000 0000 0000 1100 (logand 14 13 4) ; 14 = 0000 0000 0000 0000 0000 0000 1110

; 13 = 0000 0000 0000 0000 0000 0000 1101 ; 4 = 0000 0000 0000 0000 0000 0000 0100

⇒ 4 ; 4 = 0000 0000 0000 0000 0000 0000 0100 (logand) ⇒ -1 ; -1 = 1111 1111 1111 1111 1111 1111 1111 Function logior &rest ints-or-markers

This function returns the “inclusive or” of its arguments: the nth bit is set in the result if, and only if, the nth bit is set in at least one of the arguments. If there are no arguments, the result is zero, which is an identity element for this operation. If logior is passed just one argument, it returns that argument.

; 28-bit binary values

(logior 12 5) ; 12 = 0000 0000 0000 0000 0000 0000 1100

; 5 = 0000 0000 0000 0000 0000 0000 0101

⇒ 13 ; 13 = 0000 0000 0000 0000 0000 0000 1101 (logior 12 5 7) ; 12 = 0000 0000 0000 0000 0000 0000 1100

; 5 = 0000 0000 0000 0000 0000 0000 0101 ; 7 = 0000 0000 0000 0000 0000 0000 0111

⇒ 15 ; 15 = 0000 0000 0000 0000 0000 0000 1111

Function logxor &rest ints-or-markers This function returns the “exclusive or” of its arguments: the nth bit is set in the result if, and only if, the nth bit is set in an odd number of the arguments. If there are no arguments, the result is 0, which is an identity element for this operation. If logxor is passed just one argument, it returns that argument.

; 28-bit binary values

(logxor 12 5) ; 12 = 0000 0000 0000 0000 0000 0000 1100

; 5 = 0000 0000 0000 0000 0000 0000 0101

⇒ 9 ; 9 = 0000 0000 0000 0000 0000 0000 1001 (logxor 12 5 7) ; 12 = 0000 0000 0000 0000 0000 0000 1100

; 5 = 0000 0000 0000 0000 0000 0000 0101 ; 7 = 0000 0000 0000 0000 0000 0000 0111

⇒ 14 ; 14 = 0000 0000 0000 0000 0000 0000 1110

Function lognot integer This function returns the logical complement of its argument: the nth bit is one in the result if, and only if, the nth bit is zero in integer, and vice-versa.

     (lognot 5)
          ⇒ -6

;; 5 = 0000 0000 0000 0000 0000 0000 0101

;; becomes

;; -6 = 1111 1111 1111 1111 1111 1111 1010

3.9 Standard Mathematical Functions

These mathematical functions are available if floating point is supported (which is the normal state of affairs). They allow integers as well as floating point numbers as arguments.

Function sin arg Function cos arg Function tan arg

These are the ordinary trigonometric functions, with argument measured in radians.

Function asin arg

The value of (asin arg) is a number between −pi/2 and pi/2 (inclusive) whose sine is arg; if, however, arg is out of range (outside [-1, 1]), then the result is a NaN.

Function acos arg

The value of (acos arg) is a number between 0 and pi (inclusive) whose cosine is arg; if, however, arg is out of range (outside [-1, 1]), then the result is a NaN.

Function atan arg

The value of (atan arg) is a number between −pi/2 and pi/2 (exclusive) whose tangent is arg.

Function sinh arg Function cosh arg Function tanh arg

These are the ordinary hyperbolic trigonometric functions.

Function asinh arg Function acosh arg Function atanh arg

These are the inverse hyperbolic trigonometric functions.

Function exp arg

This is the exponential function; it returns e to the power arg. e is a fundamental mathematical constant also called the base of natural logarithms.

Function log arg &optional base This function returns the logarithm of arg, with base base. If you don’t specify base, the base e is used. If arg is negative, the result is a NaN.

Function log10 arg

This function returns the logarithm of arg, with base 10. If arg is negative, the result is a NaN. (log10 x) ≡ (log x 10), at least approximately.

Function expt x y

This function returns x raised to power y. If both arguments are integers and y is positive, the result is an integer; in this case, it is truncated to fit the range of possible integer values.

Function sqrt arg

This returns the square root of arg. If arg is negative, the value is a NaN.

Function cube-root arg

This returns the cube root of arg.

3.10 Random Numbers

A deterministic computer program cannot generate true random numbers. For most purposes, pseudo-random numbers suffice. A series of pseudo-random numbers is generated in a deterministic fashion. The numbers are not truly random, but they have certain properties that mimic a random series. For example, all possible values occur equally often in a pseudo-random series.

In XEmacs, pseudo-random numbers are generated from a “seed” number. Starting from any given seed, the random function always generates the same sequence of numbers. XEmacs always starts with the same seed value, so the sequence of values of random is actually the same in each XEmacs run! For example, in one operating system, the first call to (random) after you start XEmacs always returns -1457731, and the second one always returns -7692030. This repeatability is helpful for debugging.

If you want truly unpredictable random numbers, execute (random t). This chooses a new seed based on the current time of day and on XEmacs’s process id number.

Function random &optional limit This function returns a pseudo-random integer. Repeated calls return a series of pseudo-random integers. If limit is a positive integer, the value is chosen to be nonnegative and less than limit. If limit is t, it means to choose a new seed based on the current time of day and on XEmacs’s process id number. On some machines, any integer representable in Lisp may be the result of random. On other machines, the result can never be larger than a certain maximum or less than a certain (negative) minimum.

4 Strings and Characters

A string in XEmacs Lisp is an array that contains an ordered sequence of characters. Strings are used as names of symbols, buffers, and files, to send messages to users, to hold text being copied between buffers, and for many other purposes. Because strings are so important, XEmacs Lisp has many functions expressly for manipulating them. XEmacs Lisp programs use strings more often than individual characters.

4.1 String and Character Basics

Strings in XEmacs Lisp are arrays that contain an ordered sequence of characters. Characters are their own primitive object type in XEmacs 20. However, in XEmacs 19, characters are represented in XEmacs Lisp as integers; whether an integer was intended as a character or not is determined only by how it is used. See Section 2.4.3 [Character Type], page 18.

The length of a string (like any array) is fixed and independent of the string contents, and cannot be altered. Strings in Lisp are not terminated by a distinguished character code. (By contrast, strings in C are terminated by a character with ASCII code 0.) This means that any character, including the null character (ASCII code 0), is a valid element of a string.

Since strings are considered arrays, you can operate on them with the general array functions. (See Chapter 6 [Sequences Arrays Vectors], page 101.) For example, you can access or change individual characters in a string using the functions aref and aset (see Section 6.3 [Array Functions], page 104).

Strings use an efficient representation for storing the characters in them, and thus take up much less memory than a vector of the same length.

Sometimes you will see strings used to hold key sequences. This exists for backward compatibility with Emacs 18, but should not be used in new code, since many key chords can’t be represented at all and others (in particular meta key chords) are confused with accented characters.

Strings are useful for holding regular expressions. You can also match regular expressions against strings (see Section 37.3 [Regexp Search], page 557). The functions match-string (see Section 37.6.1 [Simple Match Data], page 562) and replace-match (see Section 37.6.2 [Replacing Match], page 563) are useful for decomposing and modifying strings based on regular expression matching.

Like a buffer, a string can contain extents in it. These extents are created when a function such as buffer-substring is called on a region with duplicable extents in it. When the string is inserted into a buffer, the extents are inserted along with it. See Section 40.9 [Duplicable Extents], page 596.

See Chapter 36 [Text], page 511, for information about functions that display strings or copy them into buffers. See Section 2.4.3 [Character Type], page 18, and Section 2.4.8 [String Type], page 25, for information about the syntax of characters and strings.

4.2 The Predicates for Strings

For more information about general sequence and array predicates, see Chapter 6 [Sequences Arrays Vectors], page 101, and Section 6.2 [Arrays], page 103.

Function stringp object

This function returns t if object is a string, nil otherwise.

Function char-or-string-p object

This function returns t if object is a string or a character, nil otherwise. In XEmacs addition, this function also returns t if object is an integer that can be represented as a character. This is because of compatibility with previous XEmacs and should not be depended on.

4.3 Creating Strings

The following functions create strings, either from scratch, or by putting strings together, or by taking them apart.

Function string &rest characters

This function returns a new string made up of characters.

     (string ?X ?E ?m ?a ?c ?s)
          ⇒ "XEmacs"
     (string)
          ⇒ ""

Analogous functions operating on other data types include list, cons (see Section 5.5 [Building Lists], page 82), vector (see Section 6.4 [Vectors], page 106) and bit-vector (see Section 6.6 [Bit Vectors], page 107). This function has not been available in XEmacs prior to 21.0 and FSF Emacs prior to 20.3.

Function make-string count character This function returns a string made up of count repetitions of character. If count is negative, an error is signaled.

     (make-string 5 ?x)
          ⇒ "xxxxx"
     (make-string 0 ?x)
          ⇒ ""

Other functions to compare with this one include char-to-string (see Section 4.7 [String Conversion], page 65), make-vector (see Section 6.4 [Vectors], page 106), and make-list (see Section 5.5 [Building Lists], page 82).

Function substring string start &optional end This function returns a new string which consists of those characters from string in the range from (and including) the character at the index start up to (but excluding) the character at the index end. The first character is at index zero.

     (substring "abcdefg" 0 3)
          ⇒ "abc"

Here the index for ‘a’ is 0, the index for ‘b’ is 1, and the index for ‘c’ is 2. Thus, three letters, ‘abc’, are copied from the string "abcdefg". The index 3 marks the character position up to which the substring is copied. The character whose index is 3 is actually the fourth character in the string. A negative number counts from the end of the string, so that −1 signifies the index of the last character of the string. For example:

     (substring "abcdefg" -3 -1)
          ⇒ "ef"

In this example, the index for ‘e’ is −3, the index for ‘f’ is −2, and the index for ‘g’ is −1. Therefore, ‘e’ and ‘f’ are included, and ‘g’ is excluded. When nil is used as an index, it stands for the length of the string. Thus,

     (substring "abcdefg" -3 nil)
          ⇒ "efg"

Omitting the argument end is equivalent to specifying nil. It follows that (substring string 0) returns a copy of all of string.

     (substring "abcdefg" 0)
          ⇒ "abcdefg"

But we recommend copy-sequence for this purpose (see Section 6.1 [Sequence Functions], page 101). If the characters copied from string have duplicable extents or text properties, those are copied into the new string also. See Section 40.9 [Duplicable Extents], page 596. A wrong-type-argument error is signaled if either start or end is not an integer or nil. An args-out-of-range error is signaled if start indicates a character following end, or if either integer is out of range for string. Contrast this function with buffer-substring (see Section 36.2 [Buffer Contents], page 512), which returns a string containing a portion of the text in the current buffer. The beginning of a string is at index 0, but the beginning of a buffer is at index 1.

Function concat &rest sequences This function returns a new string consisting of the characters in the arguments passed to it (along with their text properties, if any). The arguments may be strings, lists of numbers, or vectors of numbers; they are not themselves changed. If concat receives no arguments, it returns an empty string.

     (concat "abc" "-def")
          ⇒ "abc-def"
     (concat "abc" (list 120 (+ 256 121)) [122])
          ⇒ "abcxyz"

;; nil is an empty sequence.

     (concat "abc" nil "-def")
          ⇒ "abc-def"
     (concat "The " "quick brown " "fox.")
          ⇒ "The quick brown fox."
     (concat)

          ⇒ ""

The second example above shows how characters stored in strings are taken modulo 256. In other words, each character in the string is stored in one byte. The concat function always constructs a new string that is not eq to any existing string. When an argument is an integer (not a sequence of integers), it is converted to a string of digits making up the decimal printed representation of the integer. Don’t use this feature; we plan to eliminate it. If you already use this feature, change your programs now! The proper way to convert an integer to a decimal number in this way is with format (see Section 4.10 [Formatting Strings], page 67) or number-to-string (see Section 4.7 [String Conversion], page 65).

     (concat 137)
          ⇒ "137"
     (concat 54 321)
          ⇒ "54321"

For information about other concatenation functions, see the description of mapconcat in Section 11.6 [Mapping Functions], page 171, vconcat in Section 6.4 [Vectors], page 106, bvconcat in Section 6.6 [Bit Vectors], page 107, and append in Section 5.5 [Building Lists], page 82.

4.4 The Predicates for Characters

Function characterp object

This function returns t if object is a character. Some functions that work on integers (e.g. the comparison functions <, <=, =, /=, etc. and the arithmetic functions +, -, *, etc.) accept characters and implicitly convert them into integers. In general, functions that work on characters also accept char-ints and implicitly convert them into characters. WARNING: Neither of these behaviors is very desirable, and they are maintained for backward compatibility with old E-Lisp programs that confounded characters and integers willy-nilly. These behaviors may change in the future; therefore, do not rely on them. Instead, convert the characters explicitly using char-int.

Function integer-or-char-p object

This function returns t if object is an integer or character.

4.5 Character Codes

Function char-int ch

This function converts a character into an equivalent integer. The resulting integer will always be non-negative. The integers in the range 0 - 255 map to characters as follows:

0 - 31 Control set 0

32 - 127 ASCII 128 - 159 Control set 1

160 - 255 Right half of ISO-8859-1

If support for MULE does not exist, these are the only valid character values. When MULE support exists, the values assigned to other characters may vary depending on the particular version of XEmacs, the order in which character sets were loaded, etc., and you should not depend on them.

Function int-char integer

This function converts an integer into the equivalent character. Not all integers correspond to valid characters; use char-int-p to determine whether this is the case. If the integer cannot be converted, nil is returned.

Function char-int-p object

This function returns t if object is an integer that can be converted into a character.

Function char-or-char-int-p object

This function returns t if object is a character or an integer that can be converted into one.

4.6 Comparison of Characters and Strings

Function char-equal character1 character2 This function returns t if the arguments represent the same character, nil otherwise. This function ignores differences in case if case-fold-search is non-nil.

     (char-equal ?x ?x)
          ⇒ t
     (let ((case-fold-search t))
       (char-equal ?x ?X))
          ⇒ t
     (let ((case-fold-search nil))
       (char-equal ?x ?X))
          ⇒ nil

Function char= character1 character2 This function returns t if the arguments represent the same character, nil otherwise. Case is significant.

     (char= ?x ?x)
          ⇒ t
     (char= ?x ?X)
          ⇒ nil
     (let ((case-fold-search t))
       (char-equal ?x ?X))
          ⇒ nil
     (let ((case-fold-search nil))
       (char-equal ?x ?X))
          ⇒ nil

Function string= string1 string2 This function returns t if the characters of the two strings match exactly; case is significant.

     (string= "abc" "abc")
          ⇒ t
     (string= "abc" "ABC")
          ⇒ nil
     (string= "ab" "ABC")
          ⇒ nil

Function string-equal string1 string2

string-equal is another name for string=.

Function string< string1 string2 This function compares two strings a character at a time. First it scans both the strings at once to find the first pair of corresponding characters that do not match. If the lesser character of those two is the character from string1, then string1 is less, and this function returns t. If the lesser character is the one from string2, then string1 is greater, and this function returns nil. If the two strings match entirely, the value is nil.

Pairs of characters are compared by their ASCII codes. Keep in mind that lower case letters have higher numeric values in the ASCII character set than their upper case counterparts; numbers and many punctuation characters have a lower numeric value than upper case letters.

     (string< "abc" "abd")
          ⇒ t
     (string< "abd" "abc")
          ⇒ nil
     (string< "123" "abc")
          ⇒ t

When the strings have different lengths, and they match up to the length of string1, then the result is t. If they match up to the length of string2, the result is nil. A string of no characters is less than any other string.

     (string< "" "abc")
          ⇒ t
     (string< "ab" "abc")
          ⇒ t
     (string< "abc" "")
          ⇒ nil
     (string< "abc" "ab")
          ⇒ nil
     (string< "" "")
          ⇒ nil

Function string-lessp string1 string2

string-lessp is another name for string<.

See also compare-buffer-substrings in Section 36.3 [Comparing Text], page 513, for a way to compare text in buffers. The function string-match, which matches a regular expression against a string, can be used for a kind of string comparison; see Section 37.3 [Regexp Search], page 557.

4.7 Conversion of Characters and Strings

This section describes functions for conversions between characters, strings and integers. format and prin1-to-string (see Section 17.5 [Output Functions], page 256) can also convert Lisp objects into strings. read-from-string (see Section 17.3 [Input Functions], page 253) can “convert” a string representation of a Lisp object into an object.

See Chapter 27 [Documentation], page 381, for functions that produce textual descriptions of text characters and general input events (single-key-description and textchar-description). These functions are used primarily for making help messages.

Function char-to-string character

This function returns a new string with a length of one character. The value of character, modulo 256, is used to initialize the element of the string. This function is similar to make-string with an integer argument of 1. (See Section 4.3 [Creating Strings], page 60.) This conversion can also be done with format using the ‘%c’ format specification. (See Section 4.10 [Formatting Strings], page 67.)

     (char-to-string ?x)
          ⇒ "x"
     (char-to-string (+ 256 ?x))
          ⇒ "x"
     (make-string 1 ?x)
          ⇒ "x"

Function string-to-char string

This function returns the first character in string. If the string is empty, the function returns 0. (Under XEmacs 19, the value is also 0 when the first character of string is the null character, ASCII code 0.)

     (string-to-char "ABC")

⇒ ?A ;; Under XEmacs 20. ⇒ 65 ;; Under XEmacs 19.

     (string-to-char "xyz")

⇒ ?x ;; Under XEmacs 20. ⇒ 120 ;; Under XEmacs 19.

     (string-to-char "")
          ⇒ 0
     (string-to-char "\000")

⇒ ?\^;; Under XEmacs 20. ⇒ 0 ;; Under XEmacs 20.

This function may be eliminated in the future if it does not seem useful enough to retain.

Function number-to-string number

This function returns a string consisting of the printed representation of number, which may be an integer or a floating point number. The value starts with a sign if the argument is negative.

     (number-to-string 256)
          ⇒ "256"
     (number-to-string -23)
          ⇒ "-23"
     (number-to-string -23.5)
          ⇒ "-23.5"

int-to-string is a semi-obsolete alias for this function.

See also the function format in Section 4.10 [Formatting Strings], page 67.

Function string-to-number string &optional base This function returns the numeric value of the characters in string, read in base. It skips spaces and tabs at the beginning of string, then reads as much of string as it can interpret as a number. (On some systems it ignores other whitespace at the beginning, not just spaces and tabs.) If the first character after the ignored whitespace is not a digit or a minus sign, this function returns 0.

If base is not specified, it defaults to ten. With base other than ten, only integers can be read.

     (string-to-number "256")
          ⇒ 256
     (string-to-number "25 is a perfect square.")
          ⇒ 25
     (string-to-number "X256")
          ⇒ 0
     (string-to-number "-4.5")
          ⇒ -4.5
     (string-to-number "ffff" 16)
          ⇒ 65535

string-to-int is an obsolete alias for this function.

4.8 Modifying Strings

You can modify a string using the general array-modifying primitives. See Section 6.2 [Arrays], page 103. The function aset modifies a single character; the function fillarray sets all characters in the string to a specified character.

Each string has a tick counter that starts out at zero (when the string is created) and is incremented each time a change is made to that string.

Function string-modified-tick string

This function returns the tick counter for ‘string’.

4.9 String Properties

Similar to symbols, extents, faces, and glyphs, you can attach additional information to strings in the form of string properties. These differ from text properties, which are logically attached to particular characters in the string.

To attach a property to a string, use put. To retrieve a property from a string, use get. You can also use remprop to remove a property from a string and object-props to retrieve a list of all the properties in a string.

4.10 Formatting Strings

Formatting means constructing a string by substitution of computed values at various places in a constant string. This string controls how the other values are printed as well as where they appear; it is called a format string.

Formatting is often useful for computing messages to be displayed. In fact, the functions message and error provide the same formatting feature described here; they differ from format only in how they use the result of formatting.

Function format string &rest objects This function returns a new string that is made by copying string and then replacing any format specification in the copy with encodings of the corresponding objects. The arguments objects are the computed values to be formatted.

A format specification is a sequence of characters beginning with a ‘%’. Thus, if there is a ‘%d’ in string, the format function replaces it with the printed representation of one of the values to be formatted (one of the arguments objects). For example:

     (format "The value of fill-column is %d." fill-column)
          ⇒ "The value of fill-column is 72."

If string contains more than one format specification, the format specifications correspond with successive values from objects. Thus, the first format specification in string uses the first such value, the second format specification uses the second such value, and so on. Any extra format specifications (those for which there are no corresponding values) cause unpredictable behavior. Any extra values to be formatted are ignored.

Certain format specifications require values of particular types. However, no error is signaled if the value actually supplied fails to have the expected type. Instead, the output is likely to be meaningless.

Here is a table of valid format specifications:

‘%s’ Replace the specification with the printed representation of the object, made without quoting. Thus, strings are represented by their contents alone, with no ‘"’ characters, and symbols appear without ‘\’ characters. This is equivalent to printing the object with princ. If there is no corresponding object, the empty string is used.

‘%S’ Replace the specification with the printed representation of the object, made with quoting. Thus, strings are enclosed in ‘"’ characters, and ‘\’ characters appear where necessary before special characters. This is equivalent to printing the object with prin1. If there is no corresponding object, the empty string is used.

‘%o’ Replace the specification with the base-eight representation of an integer.

‘%d’

‘%i’ Replace the specification with the base-ten representation of an integer.

‘%x’ Replace the specification with the base-sixteen representation of an integer, using lowercase letters.

‘%X’ Replace the specification with the base-sixteen representation of an integer, using uppercase letters.

‘%c’ Replace the specification with the character which is the value given.

‘%e’ Replace the specification with the exponential notation for a floating point number (e.g. ‘7.85200e+03’).

‘%f’ Replace the specification with the decimal-point notation for a floating point number.

‘%g’ Replace the specification with notation for a floating point number, using a “pretty format”. Either exponential notation or decimal-point notation will be used (usually whichever is shorter), and trailing zeroes are removed from the fractional part.

‘%%’ A single ‘%’ is placed in the string. This format specification is unusual in that it does not use a value. For example, (format "%% %d" 30) returns "% 30".

Any other format character results in an ‘Invalid format operation’ error. Here are several examples:

(format "The name of this buffer is %s." (buffer-name))
⇒ "The name of this buffer is strings.texi."
(format "The buffer object prints as %s." (current-buffer))
⇒ "The buffer object prints as #<buffer strings.texi>."
(format "The octal value of %d is %o,
    and the hex value is %x." 18 18 18)
⇒ "The octal value of 18 is 22,
    and the hex value is 12."

There are many additional flags and specifications that can occur between the ‘%’ and the format character, in the following order:

1. An optional repositioning specification, which is a positive integer followed by a ‘$’.

2. Zero or more of the optional flag characters ‘-’, ‘+’, ‘ ’, ‘0’, and ‘#’.

3. An asterisk (‘*’, meaning that the field width is now assumed to have been specified as an argument.

4. An optional minimum field width.

5. An optional precision, preceded by a ‘.’ character.

A repositioning specification changes which argument to format is used by the current and all following format specifications. Normally the first specification uses the first argument, the second specification uses the second argument, etc. Using a repositioning specification, you can change this. By placing a number N followed by a ‘$’ between the ‘%’ and the format character, you cause the specification to use the Nth argument. The next specification will use the N+1’th argument, etc.

For example:

     (format "Can't find file `%s' in directory `%s'."
             "ignatius.c" "loyola/")
          ⇒ "Can't find file `ignatius.c' in directory `loyola/'."
     (format "In directory `%2$s', the file `%1$s' was not found."
             "ignatius.c" "loyola/")
          ⇒ "In directory `loyola/', the file `ignatius.c' was not found."
     (format
         "The numbers %d and %d are %1$x and %x in hex and %1$o and %o in octal."
         37 12)
     ⇒ "The numbers 37 and 12 are 25 and c in hex and 45 and 14 in octal."

As you can see, this lets you reprocess arguments more than once or reword a format specification (thereby moving the arguments around) without having to actually reorder the arguments. This is especially useful in translating messages from one language to another: Different languages use different word orders, and this sometimes entails changing the order of the arguments. By using repositioning specifications, this can be accomplished without having to embed knowledge of particular languages into the location in the program’s code where the message is displayed.

All the specification characters allow an optional numeric prefix between the ‘%’ and the character, and following any repositioning specification or flag. The optional numeric prefix defines the minimum width for the object. If the printed representation of the object contains fewer characters than this, then it is padded. The padding is normally on the left, but will be on the right if the ‘-’ flag character is given. The padding character is normally a space, but if the ‘0’ flag character is given, zeros are used for padding.

     (format "%06d is padded on the left with zeros" 123)
          ⇒ "000123 is padded on the left with zeros"
     (format "%-6d is padded on the right" 123)
          ⇒ "123    is padded on the right"

format never truncates an object’s printed representation, no matter what width you specify. Thus, you can use a numeric prefix to specify a minimum spacing between columns with no risk of losing information.

In the following three examples, ‘%7s’ specifies a minimum width of 7. In the first case, the string inserted in place of ‘%7s’ has only 3 letters, so 4 blank spaces are inserted for padding. In the second case, the string "specification" is 13 letters wide but is not truncated. In the third case, the padding is on the right.

     (format "The word `%7s' actually has %d letters in it."
             "foo" (length "foo"))
          ⇒ "The word `    foo' actually has 3 letters in it."
     (format "The word `%7s' actually has %d letters in it."
             "specification" (length "specification"))
          ⇒ "The word `specification' actually has 13 letters in it."
     (format "The word `%-7s' actually has %d letters in it."
             "foo" (length "foo"))
          ⇒ "The word `foo    ' actually has 3 letters in it."

After any minimum field width, a precision may be specified by preceding it with a ‘.’ character. The precision specifies the minimum number of digits to appear in ‘%d’, ‘%i’, ‘%o’, ‘%x’, and ‘%X’ conversions (the number is padded on the left with zeroes as necessary); the number of digits printed after the decimal point for ‘%f’, ‘%e’, and ‘%E’ conversions; the number of significant digits printed in ‘%g’ and ‘%G’ conversions; and the maximum number of non-padding characters printed in ‘%s’ and ‘%S’ conversions. The default precision for floating-point conversions is six.

The other flag characters have the following meanings:

• The ‘ ’ flag means prefix non-negative numbers with a space.

• The ‘+’ flag means prefix non-negative numbers with a plus sign.

• The ‘#’ flag means print numbers in an alternate, more verbose format: octal numbers begin with zero; hex numbers begin with a ‘0x’ or ‘0X’; a decimal point is printed in ‘%f’, ‘%e’, and ‘%E’ conversions even if no numbers are printed after it; and trailing zeroes are not omitted in ‘%g’ and ‘%G’ conversions.

4.11 Character Case

The character case functions change the case of single characters or of the contents of strings. The functions convert only alphabetic characters (the letters ‘A’ through ‘Z’ and ‘a’ through ‘z’); other characters are not altered. The functions do not modify the strings that are passed to them as arguments.

The examples below use the characters ‘X’ and ‘x’ which have ASCII codes 88 and 120 respectively.

Function downcase string-or-char This function converts a character or a string to lower case. When the argument to downcase is a string, the function creates and returns a new string in which each letter in the argument that is upper case is converted to lower case. When the argument to downcase is a character, downcase returns the corresponding lower case character. (This value is actually an integer under XEmacs 19.) If the original character is lower case, or is not a letter, then the value equals the original character.

          (downcase "The cat in the hat")
               ⇒ "the cat in the hat"
          (downcase ?X)

⇒ ?x ;; Under XEmacs 20. ⇒ 120 ;; Under XEmacs 19.

Function upcase string-or-char This function converts a character or a string to upper case. When the argument to upcase is a string, the function creates and returns a new string in which each letter in the argument that is lower case is converted to upper case. When the argument to upcase is a character, upcase returns the corresponding upper case character. (This value is actually an integer under XEmacs 19.) If the original character is upper case, or is not a letter, then the value equals the original character.

     (upcase "The cat in the hat")
          ⇒ "THE CAT IN THE HAT"
     (upcase ?x)

⇒ ?X ;; Under XEmacs 20. ⇒ 88 ;; Under XEmacs 19.

Function capitalize string-or-char This function capitalizes strings or characters. If string-or-char is a string, the function creates and returns a new string, whose contents are a copy of stringor-char in which each word has been capitalized. This means that the first character of each word is converted to upper case, and the rest are converted to lower case. The definition of a word is any sequence of consecutive characters that are assigned to the word constituent syntax class in the current syntax table (see Section 38.2.1 [Syntax Class Table], page 570). When the argument to capitalize is a character, capitalize has the same result as upcase.

     (capitalize "The cat in the hat")
          ⇒ "The Cat In The Hat"
     (capitalize "THE 77TH-HATTED CAT")
          ⇒ "The 77th-Hatted Cat"
     (capitalize ?x)

⇒ ?X ;; Under XEmacs 20. ⇒ 88 ;; Under XEmacs 19.

4.12 The Case Table

You can customize case conversion by installing a special case table. A case table specifies the mapping between upper case and lower case letters. It affects both the string and character case conversion functions (see the previous section) and those that apply to text in the buffer (see Section 36.17 [Case Changes], page 538). You need a case table if you are using a language which has letters other than the standard ASCII letters.

A case table is a list of this form:

(downcase upcase canonicalize equivalences) where each element is either nil or a string of length 256. The element downcase says how to map each character to its lower-case equivalent. The element upcase maps each character to its upper-case equivalent. If lower and upper case characters are in one-to-one correspondence, use nil for upcase; then XEmacs deduces the upcase table from downcase.

For some languages, upper and lower case letters are not in one-to-one correspondence. There may be two different lower case letters with the same upper case equivalent. In these cases, you need to specify the maps for both directions.

The element canonicalize maps each character to a canonical equivalent; any two characters that are related by case-conversion have the same canonical equivalent character.

The element equivalences is a map that cyclicly permutes each equivalence class (of characters with the same canonical equivalent). (For ordinary ASCII, this would map ‘a’ into ‘A’ and ‘A’ into ‘a’, and likewise for each set of equivalent characters.)

When you construct a case table, you can provide nil for canonicalize; then Emacs fills in this string from upcase and downcase. You can also provide nil for equivalences; then Emacs fills in this string from canonicalize. In a case table that is actually in use, those components are non-nil. Do not try to specify equivalences without also specifying canonicalize.

Each buffer has a case table. XEmacs also has a standard case table which is copied into each buffer when you create the buffer. Changing the standard case table doesn’t affect any existing buffers.

Here are the functions for working with case tables:

Function case-table-p object

This predicate returns non-nil if object is a valid case table.

Function set-standard-case-table table

This function makes table the standard case table, so that it will apply to any buffers created subsequently.

Function standard-case-table

This returns the standard case table.

Function current-case-table

This function returns the current buffer’s case table.

Function set-case-table table

This sets the current buffer’s case table to table.

The following three functions are convenient subroutines for packages that define non-ASCII character sets. They modify a string downcase-table provided as an argument; this should be a string to be used as the downcase part of a case table. They also modify the standard syntax table. See Chapter 38 [Syntax Tables], page 569.

Function set-case-syntax-pair uc lc downcase-table

This function specifies a pair of corresponding letters, one upper case and one lower case.

Function set-case-syntax-delims l r downcase-table

This function makes characters l and r a matching pair of case-invariant delimiters.

Function set-case-syntax char syntax downcase-table This function makes char case-invariant, with syntax syntax.

Command describe-buffer-case-table

This command displays a description of the contents of the current buffer’s case table.

You can load the library ‘iso-syntax’ to set up the standard syntax table and define a case table for the 8-bit ISO Latin 1 character set.

4.13 The Char Table

A char table is a table that maps characters (or ranges of characters) to values. Char tables are specialized for characters, only allowing particular sorts of ranges to be assigned values. Although this loses in generality, it makes for extremely fast (constant-time) lookups, and thus is feasible for applications that do an extremely large number of lookups (e.g. scanning a buffer for a character in a particular syntax, where a lookup in the syntax table must occur once per character).

Note that char tables as a primitive type, and all of the functions in this section, exist only in XEmacs 20. In XEmacs 19, char tables are generally implemented using a vector of 256 elements.

When MULE support exists, the types of ranges that can be assigned values are

• all characters

• an entire charset

• a single row in a two-octet charset

• a single character

When MULE support is not present, the types of ranges that can be assigned values are

• all characters

• a single character

Function char-table-p object

This function returns non-nil if object is a char table.

4.13.1 Char Table Types

Each char table type is used for a different purpose and allows different sorts of values. The different char table types are

category Used for category tables, which specify the regexp categories that a character is in. The valid values are nil or a bit vector of 95 elements. Higher-level Lisp functions are provided for working with category tables. Currently categories and category tables only exist when MULE support is present.

char A generalized char table, for mapping from one character to another. Used for case tables, syntax matching tables, keyboard-translate-table, etc. The valid values are characters.

generic An even more generalized char table, for mapping from a character to anything.

display Used for display tables, which specify how a particular character is to appear when displayed. #### Not yet implemented.

syntax Used for syntax tables, which specify the syntax of a particular character. Higher-level Lisp functions are provided for working with syntax tables. The valid values are integers.

Function char-table-type table

This function returns the type of char table table.

Function char-table-type-list

This function returns a list of the recognized char table types.

Function valid-char-table-type-p type

This function returns t if type if a recognized char table type.

4.13.2 Working With Char Tables

Function make-char-table type

This function makes a new, empty char table of type type. type should be a

symbol, one of char, category, display, generic, or syntax.

Function put-char-table range val table

This function sets the value for chars in range to be val in table. range specifies one or more characters to be affected and should be one of the following:

• t (all characters are affected)

• A charset (only allowed when MULE support is present)

• A vector of two elements: a two-octet charset and a row number (only allowed when MULE support is present)

• A single character

val must be a value appropriate for the type of table.

Function get-char-table ch table

This function finds the value for char ch in table.

Function get-range-char-table range table &optional multi This function finds the value for a range in table. If there is more than one value, multi is returned (defaults to nil).

Function reset-char-table table

This function resets a char table to its default state.

Function map-char-table function table &optional range This function maps function over entries in table, calling it with two args, each key and value in the table. range specifies a subrange to map over and is in the same format as the range argument to put-range-table. If omitted or t, it defaults to the entire table.

Function valid-char-table-value-p value char-table-type

This function returns non-nil if value is a valid value for char-table-type.

Function check-valid-char-table-value value char-table-type

This function signals an error if value is not a valid value for char-table-type.

76 XEmacs Lisp Reference Manual

5 Lists

A list represents a sequence of zero or more elements (which may be any Lisp objects). The important difference between lists and vectors is that two or more lists can share part of their structure; in addition, you can insert or delete elements in a list without copying the whole list.

5.1 Lists and Cons Cells

Lists in Lisp are not a primitive data type; they are built up from cons cells. A cons cell is a data object that represents an ordered pair. It records two Lisp objects, one labeled as the car, and the other labeled as the cdr. These names are traditional; see Section 2.4.6 [Cons Cell Type], page 22. cdr is pronounced “could-er.”

A list is a series of cons cells chained together, one cons cell per element of the list. By convention, the cars of the cons cells are the elements of the list, and the cdrs are used to chain the list: the cdr of each cons cell is the following cons cell. The cdr of the last cons cell is nil. This asymmetry between the car and the cdr is entirely a matter of convention; at the level of cons cells, the car and cdr slots have the same characteristics.

Because most cons cells are used as part of lists, the phrase list structure has come to mean any structure made out of cons cells.

The symbol nil is considered a list as well as a symbol; it is the list with no elements. For convenience, the symbol nil is considered to have nil as its cdr (and also as its car).

The cdr of any nonempty list l is a list containing all the elements of l except the first.

5.2 Lists as Linked Pairs of Boxes

A cons cell can be illustrated as a pair of boxes. The first box represents the car and the second box represents the cdr. Here is an illustration of the two-element list, (tulip lily), made from two cons cells:

      ---------------         ---------------
     | car   | cdr   |       | car   | cdr   |
     | tulip |   o---------->| lily  |  nil  |
     |       |       |       |       |       |
      ---------------         ---------------

Each pair of boxes represents a cons cell. Each box “refers to”, “points to” or “contains” a Lisp object. (These terms are synonymous.) The first box, which is the car of the first cons cell, contains the symbol tulip. The arrow from the cdr of the first cons cell to the second cons cell indicates that the cdr of the first cons cell points to the second cons cell.

The same list can be illustrated in a different sort of box notation like this:

         ____________
        |___|___|--> |___|___|--> nil
          |            |
          |            |
           --> tulip    --> lily

Here is a more complex illustration, showing the three-element list, ((pine needles) oak maple), the first element of which is a two-element list:

    __________________
   |___|___|--> |___|___|--> |___|___|--> nil
     |            |            |
     |            |            |
     |             --> oak      --> maple
     |
     |     ____________
      --> |___|___|--> |___|___|--> nil
            |            |
            |            |
             --> pine     --> needles

The same list represented in the first box notation looks like this:

 --------------       --------------       --------------
| car   | cdr  |     | car   | cdr  |     | car   | cdr  |
|   o   |   o------->| oak   |   o------->| maple |  nil |
|   |   |      |     |       |      |     |       |      |
 -- | ---------       --------------       --------------
    |
    |
    |        --------------       ----------------
    |       | car   | cdr  |     | car     | cdr  |
     ------>| pine  |   o------->| needles |  nil |
            |       |      |     |         |      |
             --------------       ----------------

See Section 2.4.6 [Cons Cell Type], page 22, for the read and print syntax of cons cells and lists, and for more “box and arrow” illustrations of lists.

5.3 Predicates on Lists

The following predicates test whether a Lisp object is an atom, is a cons cell or is a list, or whether it is the distinguished object nil. (Many of these predicates can be defined in terms of the others, but they are used so often that it is worth having all of them.)

Function consp object This function returns t if object is a cons cell, nil otherwise. nil is not a cons cell, although it is a list.

Function atom object This function returns t if object is an atom, nil otherwise. All objects except cons cells are atoms. The symbol nil is an atom and is also a list; it is the only Lisp object that is both.

     (atom object) ≡ (not (consp object))

Function listp object This function returns t if object is a cons cell or nil. Otherwise, it returns

nil.

     (listp '(1)) ⇒ t
     (listp '()) ⇒ t

Function nlistp object

This function is the opposite of listp: it returns t if object is not a list. Otherwise, it returns nil.

     (listp object) ≡ (not (nlistp object))

Function null object This function returns t if object is nil, and returns nil otherwise. This function is identical to not, but as a matter of clarity we use null when object is considered a list and not when it is considered a truth value (see not in Section 9.3 [Combining Conditions], page 132).

     (null '(1)) ⇒ nil
     (null '()) ⇒ t

5.4 Accessing Elements of Lists

Function car cons-cell This function returns the value pointed to by the first pointer of the cons cell cons-cell. Expressed another way, this function returns the car of cons-cell.

As a special case, if cons-cell is nil, then car is defined to return nil; therefore, any list is a valid argument for car. An error is signaled if the argument is not a cons cell or nil.

     (car '(a b c)) ⇒ a
     (car '()) ⇒ nil

Function cdr cons-cell This function returns the value pointed to by the second pointer of the cons cell cons-cell. Expressed another way, this function returns the cdr of cons-cell.

As a special case, if cons-cell is nil, then cdr is defined to return nil; therefore, any list is a valid argument for cdr. An error is signaled if the argument is not a cons cell or nil.

     (cdr '(a b c))
          ⇒ (b c)
     (cdr '()) ⇒ nil

Function car-safe object

This function lets you take the car of a cons cell while avoiding errors for other data types. It returns the car of object if object is a cons cell, nil otherwise. This is in contrast to car, which signals an error if object is not a list.

     (car-safe object)

≡

     (let ((x object))
       (if (consp x)
           (car x)
         nil))

Function cdr-safe object

This function lets you take the cdr of a cons cell while avoiding errors for other data types. It returns the cdr of object if object is a cons cell, nil otherwise. This is in contrast to cdr, which signals an error if object is not a list.

     (cdr-safe object)

≡

     (let ((x object))
       (if (consp x)
           (cdr x)
         nil))

Function nth n list This function returns the nth element of list. Elements are numbered starting with zero, so the car of list is element number zero. If the length of list is n or less, the value is nil. If n is negative, nth returns the first element of list.

     (nth 2 '(1 2 3 4)) ⇒ 3
     (nth 10 '(1 2 3 4)) ⇒ nil
     (nth -3 '(1 2 3 4))
          ⇒ 1
     (nth n x) ≡ (car (nthcdr n x))

Function nthcdr n list

This function returns the nth cdr of list. In other words, it removes the first n links of list and returns what follows. If n is zero or negative, nthcdr returns all of list. If the length of list is n or less, nthcdr returns nil.

     (nthcdr 1 '(1 2 3 4)) ⇒ (2 3 4)
     (nthcdr 10 '(1 2 3 4))
          ⇒ nil
     (nthcdr -3 '(1 2 3 4)) ⇒ (1 2 3 4)

Many convenience functions are provided to make it easier for you to access particular elements in a nested list. All of these can be rewritten in terms of the functions just described.

Function caar cons-cell Function cadr cons-cell Function cdar cons-cell Function cddr cons-cell Function caaar cons-cell Function caadr cons-cell Function cadar cons-cell Function caddr cons-cell Function cdaar cons-cell Function cdadr cons-cell Function cddar cons-cell Function cdddr cons-cell Function caaaar cons-cell Function caaadr cons-cell Function caadar cons-cell Function caaddr cons-cell Function cadaar cons-cell Function cadadr cons-cell Function caddar cons-cell Function cadddr cons-cell Function cdaaar cons-cell Function cdaadr cons-cell Function cdadar cons-cell Function cdaddr cons-cell Function cddaar cons-cell Function cddadr cons-cell Function cdddar cons-cell Function cddddr cons-cell

Each of these functions is equivalent to one or more applications of car and/or cdr. For example,

          (cadr x)

is equivalent to

          (car (cdr x))

and

          (cdaddr x)

is equivalent to

          (cdr (car (cdr (cdr x))))

That is to say, read the a’s and d’s from right to left and apply a car or cdr for each a or d found, respectively.

Function first list

This is equivalent to (nth 0 list), i.e. the first element of list. (Note that this is also equivalent to car.)

Function second list

This is equivalent to (nth 1 list), i.e. the second element of list.

Function third list Function fourth list Function fifth list Function sixth list Function seventh list Function eighth list Function ninth list Function tenth list

These are equivalent to (nth 2 list) through (nth 9 list) respectively, i.e. the third through tenth elements of list.

5.5 Building Cons Cells and Lists

Many functions build lists, as lists reside at the very heart of Lisp. cons is the fundamental list-building function; however, it is interesting to note that list is used more times in the source code for Emacs than cons.

Function cons object1 object2 This function is the fundamental function used to build new list structure. It creates a new cons cell, making object1 the car, and object2 the cdr. It then returns the new cons cell. The arguments object1 and object2 may be any Lisp objects, but most often object2 is a list.

     (cons 1 '(2)) ⇒ (1 2)

     (cons 1 '()) ⇒ (1)
     (cons 1 2) ⇒ (1 . 2)

cons is often used to add a single element to the front of a list. This is called consing the element onto the list. For example:

     (setq list (cons newelt list))

Note that there is no conflict between the variable named list used in this example and the function named list described below; any symbol can serve both purposes.

Function list &rest objects This function creates a list with objects as its elements. The resulting list is always nil-terminated. If no objects are given, the empty list is returned.

     (list 1 2 3 4 5) ⇒ (1 2 3 4 5)

     (list 1 2 '(3 4 5) 'foo)
          ⇒ (1 2 (3 4 5) foo)
     (list) ⇒ nil

Function make-list length object This function creates a list of length length, in which all the elements have the identical value object. Compare make-list with make-string (see Section 4.3 [Creating Strings], page 60).

     (make-list 3 'pigs)
          ⇒ (pigs pigs pigs)
     (make-list 0 'pigs)
          ⇒ nil

Function append &rest sequences This function returns a list containing all the elements of sequences. The sequences may be lists, vectors, or strings, but the last one should be a list. All arguments except the last one are copied, so none of them are altered. More generally, the final argument to append may be any Lisp object. The final argument is not copied or converted; it becomes the cdr of the last cons cell in the new list. If the final argument is itself a list, then its elements become in effect elements of the result list. If the final element is not a list, the result is a “dotted list” since its final cdr is not nil as required in a true list. See nconc in Section 5.6.3 [Rearrangement], page 87, for a way to join lists with no copying. Here is an example of using append:

     (setq trees '(pine oak))
          ⇒ (pine oak)
     (setq more-trees (append '(maple birch) trees))
          ⇒ (maple birch pine oak)
     trees

          ⇒ (pine oak)

     more-trees
          ⇒ (maple birch pine oak)
     (eq trees (cdr (cdr more-trees)))
          ⇒ t

You can see how append works by looking at a box diagram. The variable trees is set to the list (pine oak) and then the variable more-trees is set to the list (maple birch pine oak). However, the variable trees continues to refer to the original list:

     more-trees                trees
     |                           |
     |     ____________-> ____________
      --> |___|___|--> |___|___|--> |___|___|--> |___|___|--> nil
            |            |            |            |
            |            |            |            |
             --> maple    -->birch     --> pine     --> oak

An empty sequence contributes nothing to the value returned by append. As a consequence of this, a final nil argument forces a copy of the previous argument.

     trees

          ⇒ (pine oak)

     (setq wood (append trees ()))
          ⇒ (pine oak)
     wood

          ⇒ (pine oak)

     (eq wood trees)
          ⇒ nil

This once was the usual way to copy a list, before the function copy-sequence was invented. See Chapter 6 [Sequences Arrays Vectors], page 101. With the help of apply, we can append all the lists in a list of lists:

     (apply 'append '((a b c) nil (x y z) nil))
          ⇒ (a b c x y z)

If no sequences are given, nil is returned:

     (append) ⇒ nil

Here are some examples where the final argument is not a list:

     (append '(x y) 'z)
          ⇒ (x y . z)
     (append '(x y) [z])
          ⇒ (x y . [z])

The second example shows that when the final argument is a sequence but not a list, the sequence’s elements do not become elements of the resulting list. Instead, the sequence becomes the final cdr, like any other non-list final argument. The append function also allows integers as arguments. It converts them to strings of digits, making up the decimal print representation of the integer, and then uses the strings instead of the original integers. Don’t use this feature; we plan to eliminate it. If you already use this feature, change your programs now! The proper way to convert an integer to a decimal number in this way is with format (see Section 4.10 [Formatting Strings], page 67) or number-to-string (see Section 4.7 [String Conversion], page 65).

Function reverse list

This function creates a new list whose elements are the elements of list, but in reverse order. The original argument list is not altered.

     (setq x '(1 2 3 4)) ⇒ (1 2 3 4)
     (reverse x)
          ⇒ (4 3 2 1)
     x

          ⇒ (1 2 3 4)

5.6 Modifying Existing List Structure

You can modify the car and cdr contents of a cons cell with the primitives setcar and

setcdr.

Common Lisp note: Common Lisp uses functions rplaca and rplacd to alter list structure; they change structure the same way as setcar and setcdr, but the Common Lisp functions return the cons cell while setcar and setcdr return the new car or cdr.

5.6.1 Altering List Elements with setcar

Changing the car of a cons cell is done with setcar. When used on a list, setcar replaces one element of a list with a different element.

Function setcar cons object This function stores object as the new car of cons, replacing its previous car. It returns the value object. For example:

     (setq x '(1 2)) ⇒ (1 2)

     (setcar x 4) ⇒ 4
     x  ⇒ (4 2)

When a cons cell is part of the shared structure of several lists, storing a new car into the cons changes one element of each of these lists. Here is an example:

;; Create two lists that are partly shared.

(setq x1 '(a b c))
     ⇒ (a b c)
(setq x2 (cons 'z (cdr x1)))
     ⇒ (z b c)

;; Replace the car of a shared link.

(setcar (cdr x1) 'foo)
     ⇒ foo

x1 ; Both lists are changed.

     ⇒ (a foo c)
x2

     ⇒ (z foo c)

;; Replace the car of a link that is not shared.

(setcar x1 'baz)
     ⇒ baz

x1 ; Only one list is changed.

     ⇒ (baz foo c)
x2

     ⇒ (z foo c)

Here is a graphical depiction of the shared structure of the two lists in the variables x1 and x2, showing why replacing b changes them both:

        __________________
x1---> |___|___|----> |___|___|--> |___|___|--> nil
         |        -->   |            |
         |       |      |            |
          --> a  |       --> b        --> c
                 |
       ______|
x2--> |___|___|--
        |
        |
         --> z

Here is an alternative form of box diagram, showing the same relationship:

x1:
 --------------       --------------       --------------
| car   | cdr  |     | car   | cdr  |     | car   | cdr  |
|   a   |   o------->|   b   |   o------->|   c   |  nil |
|       |      |  -->|       |      |     |       |      |
 --------------  |    --------------       --------------
                 |
x2:              |
 --------------  |
| car   | cdr  | |
|   z   |   o----
|       |      |
 --------------

5.6.2 Altering the CDR of a List

The lowest-level primitive for modifying a cdr is setcdr:

Function setcdr cons object This function stores object as the new cdr of cons, replacing its previous cdr. It returns the value object.

Here is an example of replacing the cdr of a list with a different list. All but the first element of the list are removed in favor of a different sequence of elements. The first element is unchanged, because it resides in the car of the list, and is not reached via the cdr.

(setq x '(1 2 3)) ⇒ (1 2 3)
(setcdr x '(4)) ⇒ (4)
x  ⇒ (1 4)

You can delete elements from the middle of a list by altering the cdrs of the cons cells in the list. For example, here we delete the second element, b, from the list (a b c), by changing the cdr of the first cell:

(setq x1 '(a b c))
     ⇒ (a b c)
(setcdr x1 (cdr (cdr x1)))
     ⇒ (c)
x1

     ⇒ (a c)

Here is the result in box notation:

                   --------------------
                  |                    |
 --------------   |   --------------   |    --------------
| car   | cdr  |  |  | car   | cdr  |   -->| car   | cdr  |
|   a   |   o-----   |   b   |   o-------->|   c   |  nil |
|       |      |     |       |      |      |       |      |
 --------------       --------------        --------------

The second cons cell, which previously held the element b, still exists and its car is still b, but it no longer forms part of this list.

It is equally easy to insert a new element by changing cdrs:

(setq x1 '(a b c))
     ⇒ (a b c)
(setcdr x1 (cons 'd (cdr x1)))
     ⇒ (d b c)
x1

     ⇒ (a d b c)

Here is this result in box notation:

 --------------        -------------       -------------
| car  | cdr   |      | car  | cdr  |     | car  | cdr  |
|   a  |   o   |   -->|   b  |   o------->|   c  |  nil |
|      |   |   |  |   |      |      |     |      |      |
 --------- | --   |    -------------       -------------
           |      |
     -----         --------
    |                      |
    |    ---------------   |
    |   | car   | cdr   |  |
     -->|   d   |   o------
        |       |       |
         ---------------

5.6.3 Functions that Rearrange Lists

Here are some functions that rearrange lists “destructively” by modifying the cdrs of their component cons cells. We call these functions “destructive” because they chew up the original lists passed to them as arguments, to produce a new list that is the returned value.

The function delq in the following section is another example of destructive list manipulation.

Function nconc &rest lists This function returns a list containing all the elements of lists. Unlike append (see Section 5.5 [Building Lists], page 82), the lists are not copied. Instead, the last cdr of each of the lists is changed to refer to the following list. The last of the lists is not altered. For example:

     (setq x '(1 2 3)) ⇒ (1 2 3)

     (nconc x '(4 5)) ⇒ (1 2 3 4 5)
     x  ⇒ (1 2 3 4 5)

Since the last argument of nconc is not itself modified, it is reasonable to use a constant list, such as '(4 5), as in the above example. For the same reason, the last argument need not be a list:

     (setq x '(1 2 3)) ⇒ (1 2 3)
     (nconc x 'z)
          ⇒ (1 2 3 . z)
     x  ⇒ (1 2 3 . z)

A common pitfall is to use a quoted constant list as a non-last argument to nconc. If you do this, your program will change each time you run it! Here is what happens:

(defun add-foo (x) ; We want this function to add

       (nconc '(foo) x))           ;   foo to the front of its arg.
     (symbol-function 'add-foo)
          ⇒ (lambda (x) (nconc (quote (foo)) x))
     (setq xx (add-foo '(1 2)))    ; It seems to work.
          ⇒ (foo 1 2)
     (setq xy (add-foo '(3 4)))    ; What happened?
          ⇒ (foo 1 2 3 4)
     (eq xx xy) ⇒ t
     (symbol-function 'add-foo)
          ⇒ (lambda (x) (nconc (quote (foo 1 2 3 4) x)))

Function nreverse list

This function reverses the order of the elements of list. Unlike reverse, nreverse alters its argument by reversing the cdrs in the cons cells forming the list. The cons cell that used to be the last one in list becomes the first cell of the value.

For example:

     (setq x '(1 2 3 4)) ⇒ (1 2 3 4)

     x

          ⇒ (1 2 3 4)

     (nreverse x)
          ⇒ (4 3 2 1)

;; The cell that was first is now last.

     x

          ⇒ (1)

To avoid confusion, we usually store the result of nreverse back in the same variable which held the original list:

     (setq x (nreverse x))

Here is the nreverse of our favorite example, (a b c), presented graphically: Original list head: Reversed list:

      -------------        -------------        ------------
     | car  | cdr  |      | car  | cdr  |      | car | cdr  |
     |   a  |  nil |<--   |   b  |   o  |<--   |   c |   o  |
     |      |      |   |  |      |   |  |   |  |     |   |  |
      -------------    |   --------- | -    |   -------- | -
                       |             |      |            |
                        -------------        ------------

Function sort list predicate This function sorts list stably, though destructively, and returns the sorted list. It compares elements using predicate. A stable sort is one in which elements with equal sort keys maintain their relative order before and after the sort. Stability is important when successive sorts are used to order elements according to different criteria. The argument predicate must be a function that accepts two arguments. It is called with two elements of list. To get an increasing order sort, the predicate should return t if the first element is “less than” the second, or nil if not. The destructive aspect of sort is that it rearranges the cons cells forming list by changing cdrs. A nondestructive sort function would create new cons cells to store the elements in their sorted order. If you wish to make a sorted copy without destroying the original, copy it first with copy-sequence and then sort. Sorting does not change the cars of the cons cells in list; the cons cell that originally contained the element a in list still has a in its car after sorting, but it now appears in a different position in the list due to the change of cdrs. For example:

     (setq nums '(1 3 2 6 5 4 0))
          ⇒ (1 3 2 6 5 4 0)
     (sort nums '<)
          ⇒ (0 1 2 3 4 5 6)
     nums ⇒ (1 2 3 4 5 6)

Note that the list in nums no longer contains 0; this is the same cons cell that it was before, but it is no longer the first one in the list. Don’t assume a variable that formerly held the argument now holds the entire sorted list! Instead, save the result of sort and use that. Most often we store the result back into the variable that held the original list:

     (setq nums (sort nums '<))

See Section 36.14 [Sorting], page 529, for more functions that perform sorting. See documentation in Section 27.2 [Accessing Documentation], page 382, for a useful example of sort.

5.7 Using Lists as Sets

A list can represent an unordered mathematical set—simply consider a value an element of a set if it appears in the list, and ignore the order of the list. To form the union of two sets, use append (as long as you don’t mind having duplicate elements). Other useful functions for sets include memq and delq, and their equal versions, member and delete.

Common Lisp note: Common Lisp has functions union (which avoids duplicate elements) and intersection for set operations, but XEmacs Lisp does not have them. You can write them in Lisp if you wish.

Function memq object list This function tests to see whether object is a member of list. If it is, memq returns a list starting with the first occurrence of object. Otherwise, it returns nil. The letter ‘q’ in memq says that it uses eq to compare object against the elements of the list. For example:

     (memq 'b '(a b c b a))
          ⇒ (b c b a)
     (memq '(2) '((1) (2)))    ; (2) and (2) are not eq.
          ⇒ nil

Function delq object list This function destructively removes all elements eq to object from list. The letter ‘q’ in delq says that it uses eq to compare object against the elements of the list, like memq.

When delq deletes elements from the front of the list, it does so simply by advancing down the list and returning a sublist that starts after those elements:

(delq 'a '(a b c)) ≡ (cdr '(a b c))

When an element to be deleted appears in the middle of the list, removing it involves changing the cdrs (see Section 5.6.2 [Setcdr], page 86).

(setq sample-list '(a b c (4)))
     ⇒ (a b c (4))
(delq 'a sample-list)
     ⇒ (b c (4))
sample-list
     ⇒ (a b c (4))
(delq 'c sample-list)
     ⇒ (a b (4))
sample-list
     ⇒ (a b (4))

Note that (delq 'c sample-list) modifies sample-list to splice out the third element, but (delq 'a sample-list) does not splice anything—it just returns a shorter list. Don’t assume that a variable which formerly held the argument list now has fewer elements, or that it still holds the original list! Instead, save the result of delq and use that. Most often we store the result back into the variable that held the original list:

(setq flowers (delq 'rose flowers))

In the following example, the (4) that delq attempts to match and the (4) in the

sample-list are not eq:
(delq '(4) sample-list)
     ⇒ (a c (4))

The following two functions are like memq and delq but use equal rather than eq to compare elements. They are new in Emacs 19.

Function member object list The function member tests to see whether object is a member of list, comparing members with object using equal. If object is a member, member returns a list starting with its first occurrence in list. Otherwise, it returns nil.

Compare this with memq:

     (member '(2) '((1) (2)))  ; (2) and (2) are equal.
          ⇒ ((2))
     (memq '(2) '((1) (2)))    ; (2) and (2) are not eq.
          ⇒ nil

;; Two strings with the same contents are equal.

     (member "foo" '("foo" "bar"))
          ⇒ ("foo" "bar")

Function delete object list This function destructively removes all elements equal to object from list. It is to delq as member is to memq: it uses equal to compare elements with object, like member; when it finds an element that matches, it removes the element just as delq would. For example:

     (delete '(2) '((2) (1) (2))) ⇒ '((1))

Common Lisp note: The functions member and delete in XEmacs Lisp are derived from Maclisp, not Common Lisp. The Common Lisp versions do not use equal to compare elements.

See also the function add-to-list, in Section 10.7 [Setting Variables], page 152, for another way to add an element to a list stored in a variable.

5.8 Association Lists

An association list, or alist for short, records a mapping from keys to values. It is a list of cons cells called associations: the car of each cell is the key, and the cdr is the associated value.^{1}

Here is an example of an alist. The key pine is associated with the value cones; the key oak is associated with acorns; and the key maple is associated with seeds.

     '((pine . cones)
       (oak . acorns)
       (maple . seeds))

The associated values in an alist may be any Lisp objects; so may the keys. For example, in the following alist, the symbol a is associated with the number 1, and the string "b" is associated with the list (2 3), which is the cdr of the alist element:

     ((a . 1) ("b" 2 3))

Sometimes it is better to design an alist to store the associated value in the car of the cdr of the element. Here is an example:

     '((rose red) (lily white) (buttercup yellow))

Here we regard red as the value associated with rose. One advantage of this method is that you can store other related information—even a list of other items—in the cdr of the cdr. One disadvantage is that you cannot use rassq (see below) to find the element containing a given value. When neither of these considerations is important, the choice is a matter of taste, as long as you are consistent about it for any given alist.

Note that the same alist shown above could be regarded as having the associated value in the cdr of the element; the value associated with rose would be the list (red).

Association lists are often used to record information that you might otherwise keep on a stack, since new associations may be added easily to the front of the list. When searching an association list for an association with a given key, the first one found is returned, if there is more than one.

In XEmacs Lisp, it is not an error if an element of an association list is not a cons cell. The alist search functions simply ignore such elements. Many other versions of Lisp signal errors in such cases.

Note that property lists are similar to association lists in several respects. A property list behaves like an association list in which each key can occur only once. See Section 5.9 [Property Lists], page 96, for a comparison of property lists and association lists.

Function assoc key alist This function returns the first association for key in alist. It compares key against the alist elements using equal (see Section 2.8 [Equality Predicates], page 41). It returns nil if no association in alist has a car equal to key. For example:

1 This usage of “key” is not related to the term “key sequence”; it means a value used to look up an item in a table. In this case, the table is the alist, and the alist associations are the items.

     (setq trees '((pine . cones) (oak . acorns) (maple . seeds)))
          ⇒ ((pine . cones) (oak . acorns) (maple . seeds))
     (assoc 'oak trees)
          ⇒ (oak . acorns)
     (cdr (assoc 'oak trees))
          ⇒ acorns
     (assoc 'birch trees)
          ⇒ nil

Here is another example, in which the keys and values are not symbols:

     (setq needles-per-cluster
           '((2 "Austrian Pine" "Red Pine")
             (3 "Pitch Pine")
             (5 "White Pine")))
     (cdr (assoc 3 needles-per-cluster))
          ⇒ ("Pitch Pine")
     (cdr (assoc 2 needles-per-cluster))
          ⇒ ("Austrian Pine" "Red Pine")

Function rassoc value alist This function returns the first association with value value in alist. It returns nil if no association in alist has a cdr equal to value.

rassoc is like assoc except that it compares the cdr of each alist association instead of the car. You can think of this as “reverse assoc”, finding the key for a given value.

Function assq key alist This function is like assoc in that it returns the first association for key in alist, but it makes the comparison using eq instead of equal. assq returns nil if no association in alist has a car eq to key. This function is used more often than assoc, since eq is faster than equal and most alists use symbols as keys. See Section 2.8 [Equality Predicates], page 41.

     (setq trees '((pine . cones) (oak . acorns) (maple . seeds)))
          ⇒ ((pine . cones) (oak . acorns) (maple . seeds))
     (assq 'pine trees)
          ⇒ (pine . cones)

On the other hand, assq is not usually useful in alists where the keys may not be symbols:

     (setq leaves
           '(("simple leaves" . oak)
             ("compound leaves" . horsechestnut)))
     (assq "simple leaves" leaves)
          ⇒ nil
     (assoc "simple leaves" leaves)
          ⇒ ("simple leaves" . oak)

Function rassq value alist This function returns the first association with value value in alist. It returns nil if no association in alist has a cdr eq to value. rassq is like assq except that it compares the cdr of each alist association instead of the car. You can think of this as “reverse assq”, finding the key for a given value. For example:

     (setq trees '((pine . cones) (oak . acorns) (maple . seeds)))
     (rassq 'acorns trees)
          ⇒ (oak . acorns)
     (rassq 'spores trees)
          ⇒ nil

Note that rassq cannot search for a value stored in the car of the cdr of an element:

     (setq colors '((rose red) (lily white) (buttercup yellow)))
     (rassq 'white colors)
          ⇒ nil

In this case, the cdr of the association (lily white) is not the symbol white, but rather the list (white). This becomes clearer if the association is written in dotted pair notation:

     (lily white) ≡ (lily . (white))

Function remassoc key alist

This function deletes by side effect any associations with key key in alist – i.e. it removes any elements from alist whose car is equal to key. The modified alist is returned. If the first member of alist has a car that is equal to key, there is no way to remove it by side effect; therefore, write (setq foo (remassoc key foo)) to be sure of changing the value of foo.

Function remassq key alist This function deletes by side effect any associations with key key in alist – i.e. it removes any elements from alist whose car is eq to key. The modified alist is returned. This function is exactly like remassoc, but comparisons between key and keys in alist are done using eq instead of equal.

Function remrassoc value alist This function deletes by side effect any associations with value value in alist – i.e. it removes any elements from alist whose cdr is equal to value. The modified alist is returned. If the first member of alist has a car that is equal to value, there is no way to remove it by side effect; therefore, write (setq foo (remassoc value foo)) to be sure of changing the value of foo. remrassoc is like remassoc except that it compares the cdr of each alist association instead of the car. You can think of this as “reverse remassoc”, removing an association based on its value instead of its key.

Function remrassq value alist This function deletes by side effect any associations with value value in alist – i.e. it removes any elements from alist whose cdr is eq to value. The modified alist is returned.

This function is exactly like remrassoc, but comparisons between value and values in alist are done using eq instead of equal.

Function copy-alist alist

This function returns a two-level deep copy of alist: it creates a new copy of each association, so that you can alter the associations of the new alist without changing the old one.

     (setq needles-per-cluster
           '((2 . ("Austrian Pine" "Red Pine"))
             (3 . ("Pitch Pine"))
             (5 . ("White Pine"))))

⇒

     ((2 "Austrian Pine" "Red Pine")
      (3 "Pitch Pine")
      (5 "White Pine"))
     (setq copy (copy-alist needles-per-cluster))

⇒

     ((2 "Austrian Pine" "Red Pine")
      (3 "Pitch Pine")
      (5 "White Pine"))
     (eq needles-per-cluster copy)
          ⇒ nil
     (equal needles-per-cluster copy)
          ⇒ t
     (eq (car needles-per-cluster) (car copy))
          ⇒ nil
     (cdr (car (cdr needles-per-cluster)))
          ⇒ ("Pitch Pine")
     (eq (cdr (car (cdr needles-per-cluster)))
         (cdr (car (cdr copy))))
          ⇒ t

This example shows how copy-alist makes it possible to change the associations of one copy without affecting the other:

     (setcdr (assq 3 copy) '("Martian Vacuum Pine"))
     (cdr (assq 3 needles-per-cluster))
          ⇒ ("Pitch Pine")

5.9 Property Lists

A property list (or plist) is another way of representing a mapping from keys to values. Instead of the list consisting of conses of a key and a value, the keys and values alternate as successive entries in the list. Thus, the association list

((a . 1) (b . 2) (c . 3))

has the equivalent property list form

(a 1 b 2 c 3)

Property lists are used to represent the properties associated with various sorts of objects, such as symbols, strings, frames, etc. The convention is that property lists can be modified in-place, while association lists generally are not.

Plists come in two varieties: normal plists, whose keys are compared with eq, and lax plists, whose keys are compared with equal,

Function valid-plist-p plist

Given a plist, this function returns non-nil if its format is correct. If it returns nil, check-valid-plist will signal an error when given the plist; that means it’s a malformed or circular plist or has non-symbols as keywords.

Function check-valid-plist plist

Given a plist, this function signals an error if there is anything wrong with it. This means that it’s a malformed or circular plist.

5.9.1 Working With Normal Plists

Function plist-get plist prop &optional default This function extracts a value from a property list. The function returns the value corresponding to the given prop, or default if prop is not one of the properties on the list.

Function plist-put plist prop val This function changes the value in plist of prop to val. If prop is already a property on the list, its value is set to val, otherwise the new prop val pair is added. The new plist is returned; use (setq x (plist-put x prop val)) to be sure to use the new value. The plist is modified by side effects.

Function plist-remprop plist prop

This function removes from plist the property prop and its value. The new plist

is returned; use (setq x (plist-remprop x prop val)) to be sure to use the

new value. The plist is modified by side effects.

Function plist-member plist prop

This function returns t if prop has a value specified in plist.

In the following functions, if optional arg nil-means-not-present is non-nil, then a property with a nil value is ignored or removed. This feature is a virus that has infected old Lisp implementations (and thus E-Lisp, due to RMS’s enamorment with old Lisps), but should not be used except for backward compatibility.

Function plists-eq a b &optional nil-means-not-present This function returns non-nil if property lists A and B are eq (i.e. their values are eq).

Function plists-equal a b &optional nil-means-not-present This function returns non-nil if property lists A and B are equal (i.e. their values are equal; their keys are still compared using eq).

Function canonicalize-plist plist &optional nil-means-not-present This function destructively removes any duplicate entries from a plist. In such cases, the first entry applies. The new plist is returned. If nil-means-not-present is given, the return value may not be eq to the passed-in value, so make sure to setq the value back into where it came from.

5.9.2 Working With Lax Plists

Recall that a lax plist is a property list whose keys are compared using equal instead of

eq.

Function lax-plist-get lax-plist prop &optional default This function extracts a value from a lax property list. The function returns the value corresponding to the given prop, or default if prop is not one of the properties on the list.

Function lax-plist-put lax-plist prop val This function changes the value in lax-plist of prop to val.

Function lax-plist-remprop lax-plist prop

This function removes from lax-plist the property prop and its value. The new

plist is returned; use (setq x (lax-plist-remprop x prop val)) to be sure to

use the new value. The lax-plist is modified by side effects.

Function lax-plist-member lax-plist prop

This function returns t if prop has a value specified in lax-plist.

In the following functions, if optional arg nil-means-not-present is non-nil, then a property with a nil value is ignored or removed. This feature is a virus that has infected old Lisp implementations (and thus E-Lisp, due to RMS’s enamorment with old Lisps), but should not be used except for backward compatibility.

Function lax-plists-eq a b &optional nil-means-not-present This function returns non-nil if lax property lists A and B are eq (i.e. their values are eq; their keys are still compared using equal).

Function lax-plists-equal a b &optional nil-means-not-present This function returns non-nil if lax property lists A and B are equal (i.e. their values are equal).

Function canonicalize-lax-plist lax-plist &optional nil-means-not-present This function destructively removes any duplicate entries from a lax plist. In such cases, the first entry applies. The new plist is returned. If nil-means-not-present is given, the return value may not be eq to the passed-in value, so make sure to setq the value back into where it came from.

5.9.3 Converting Plists To/From Alists

Function alist-to-plist alist

This function converts association list alist into the equivalent property-list form. The plist is returned. This converts from

     ((a . 1) (b . 2) (c . 3))

into

     (a 1 b 2 c 3)

The original alist is not modified.

Function plist-to-alist plist

This function converts property list plist into the equivalent association-list form. The alist is returned. This converts from

     (a 1 b 2 c 3)

into

     ((a . 1) (b . 2) (c . 3))

The original plist is not modified.

The following two functions are equivalent to the preceding two except that they destructively modify their arguments, using cons cells from the original list to form the new list rather than allocating new cons cells.

Function destructive-alist-to-plist alist

This function destructively converts association list alist into the equivalent property-list form. The plist is returned.

Function destructive-plist-to-alist plist

This function destructively converts property list plist into the equivalent association-list form. The alist is returned.

5.10 Weak Lists

A weak list is a special sort of list whose members are not counted as references for the purpose of garbage collection. This means that, for any object in the list, if there are no references to the object anywhere outside of the list (or other weak list or weak hash table), that object will disappear the next time a garbage collection happens. Weak lists can be useful for keeping track of things such as unobtrusive lists of another function’s buffers or markers. When that function is done with the elements, they will automatically disappear from the list.

Weak lists are used internally, for example, to manage the list holding the children of an extent – an extent that is unused but has a parent will still be reclaimed, and will automatically be removed from its parent’s list of children.

Weak lists are similar to weak hash tables (see Section 46.3 [Weak Hash Tables], page 666).

Function weak-list-p object

This function returns non-nil if object is a weak list.

Weak lists come in one of four types:

simple Objects in the list disappear if not referenced outside of the list.

assoc Objects in the list disappear if they are conses and either the car or the cdr of the cons is not referenced outside of the list.

key-assoc

Objects in the list disappear if they are conses and the car is not referenced outside of the list.

value-assoc

Objects in the list disappear if they are conses and the cdr is not referenced outside of the list.

Function make-weak-list &optional type

This function creates a new weak list of type type. type is a symbol (one of simple, assoc, key-assoc, or value-assoc, as described above) and defaults

     to simple.

Function weak-list-type weak

This function returns the type of the given weak-list object.

Function weak-list-list weak

This function returns the list contained in a weak-list object.

Function set-weak-list-list weak new-list

This function changes the list contained in a weak-list object.

100 XEmacs Lisp Reference Manual

6 Sequences, Arrays, and Vectors

Recall that the sequence type is the union of four other Lisp types: lists, vectors, bit vectors, and strings. In other words, any list is a sequence, any vector is a sequence, any bit vector is a sequence, and any string is a sequence. The common property that all sequences have is that each is an ordered collection of elements.

An array is a single primitive object that has a slot for each elements. All the elements are accessible in constant time, but the length of an existing array cannot be changed. Strings, vectors, and bit vectors are the three types of arrays.

A list is a sequence of elements, but it is not a single primitive object; it is made of cons cells, one cell per element. Finding the nth element requires looking through n cons cells, so elements farther from the beginning of the list take longer to access. But it is possible to add elements to the list, or remove elements.

The following diagram shows the relationship between these types:

 ___________________________________
|                                   |
|          Sequence                 |
|  ____________________________|
| |      | |                      | |
| | List | |         Array        | |
| |      | |  _______________| |
| |______| | |        | |       | | |
|          | | Vector | | String| | |
|          | |________| |_______| | |
|          |  __________________| |
|          | |                  | | |
|          | |    Bit Vector    | | |
|          | |__________________| | |
|          |______________________| |
|___________________________________|

The elements of vectors and lists may be any Lisp objects. The elements of strings are all characters. The elements of bit vectors are the numbers 0 and 1.

6.1 Sequences

In XEmacs Lisp, a sequence is either a list, a vector, a bit vector, or a string. The common property that all sequences have is that each is an ordered collection of elements. This section describes functions that accept any kind of sequence.

Function sequencep object

Returns t if object is a list, vector, bit vector, or string, nil otherwise.

Function copy-sequence sequence

Returns a copy of sequence. The copy is the same type of object as the original sequence, and it has the same elements in the same order. Storing a new element into the copy does not affect the original sequence, and vice versa. However, the elements of the new sequence are not copies; they are identical (eq) to the elements of the original. Therefore, changes made within these elements, as found via the copied sequence, are also visible in the original sequence. If the sequence is a string with extents or text properties, the extents and text properties in the copy are also copied, not shared with the original. (This means that modifying the extents or text properties of the original will not affect the copy.) However, the actual values of the properties are shared. See Chapter 40 [Extents], page 585, See Section 36.18 [Text Properties], page 539. See also append in Section 5.5 [Building Lists], page 82, concat in Section 4.3 [Creating Strings], page 60, vconcat in Section 6.4 [Vectors], page 106, and bvconcat in Section 6.6 [Bit Vectors], page 107, for other ways to copy sequences.

(setq bar '(1 2)) ⇒ (1 2)
(setq x (vector 'foo bar))
     ⇒ [foo (1 2)]
(setq y (copy-sequence x))
     ⇒ [foo (1 2)]
(eq x y) ⇒ nil
(equal x y) ⇒ t
(eq (elt x 1) (elt y 1))
     ⇒ t

;; Replacing an element of one sequence.

(aset x 0 'quux)
x ⇒ [quux (1 2)]
y ⇒ [foo (1 2)]

;; Modifying the inside of a shared element.

(setcar (aref x 1) 69)
x ⇒ [quux (69 2)]
y ⇒ [foo (69 2)]

;; Creating a bit vector.

(bit-vector 1 0 1 1 0 1 0 0)
     ⇒ #*10110100

Function length sequence Returns the number of elements in sequence. If sequence is a cons cell that is not a list (because the final cdr is not nil), a wrong-type-argument error is signaled.

(length '(1 2 3)) ⇒ 3
(length ()) ⇒ 0
(length "foobar")
    ⇒ 6

          (length [1 2 3]) ⇒ 3
          (length #*01101) ⇒ 5

Function elt sequence index This function returns the element of sequence indexed by index. Legitimate values of index are integers ranging from 0 up to one less than the length of sequence. If sequence is a list, then out-of-range values of index return nil; otherwise, they trigger an args-out-of-range error.

          (elt [1 2 3 4] 2) ⇒ 3
          (elt '(1 2 3 4) 2) ⇒ 3
          (char-to-string (elt "1234" 2))
               ⇒ "3"
          (elt #*00010000 3) ⇒ 1
          (elt [1 2 3 4] 4)
                error Args out of range: [1 2 3 4], 4
          (elt [1 2 3 4] -1)
                error Args out of range: [1 2 3 4], -1

This function generalizes aref (see Section 6.3 [Array Functions], page 104)

and nth (see Section 5.4 [List Elements], page 79).

6.2 Arrays

An array object has slots that hold a number of other Lisp objects, called the elements of the array. Any element of an array may be accessed in constant time. In contrast, an element of a list requires access time that is proportional to the position of the element in the list.

When you create an array, you must specify how many elements it has. The amount of space allocated depends on the number of elements. Therefore, it is impossible to change the size of an array once it is created; you cannot add or remove elements. However, you can replace an element with a different value.

XEmacs defines three types of array, all of which are one-dimensional: strings, vectors, and bit vectors. A vector is a general array; its elements can be any Lisp objects. A string is a specialized array; its elements must be characters. A bit vector is another specialized array; its elements must be bits (an integer, either 0 or 1). Each type of array has its own read syntax. See Section 2.4.8 [String Type], page 25, Section 2.4.9 [Vector Type], page 26, and Section 2.4.10 [Bit Vector Type], page 26.

All kinds of array share these characteristics:

• The first element of an array has index zero, the second element has index 1, and so on. This is called zero-origin indexing. For example, an array of four elements has indices 0, 1, 2, and 3.

• The elements of an array may be referenced or changed with the functions aref and aset, respectively (see Section 6.3 [Array Functions], page 104).

In principle, if you wish to have an array of text characters, you could use either a string or a vector. In practice, we always choose strings for such applications, for four reasons:

• They usually occupy one-fourth the space of a vector of the same elements. (This is one-eighth the space for 64-bit machines such as the DEC Alpha, and may also be different when MULE support is compiled into XEmacs.)

• Strings are printed in a way that shows the contents more clearly as characters.

• Strings can hold extent and text properties. See Chapter 40 [Extents], page 585, See Section 36.18 [Text Properties], page 539.

• Many of the specialized editing and I/O facilities of XEmacs accept only strings. For example, you cannot insert a vector of characters into a buffer the way you can insert a string. See Chapter 4 [Strings and Characters], page 59.

By contrast, for an array of keyboard input characters (such as a key sequence), a vector may be necessary, because many keyboard input characters are non-printable and are represented with symbols rather than with characters. See Section 19.6.1 [Key Sequence Input], page 301.

Similarly, when representing an array of bits, a bit vector has the following advantages over a regular vector:

• They occupy 1/32nd the space of a vector of the same elements. (1/64th on 64-bit machines such as the DEC Alpha.)

• Bit vectors are printed in a way that shows the contents more clearly as bits.

6.3 Functions that Operate on Arrays

In this section, we describe the functions that accept strings, vectors, and bit vectors.

Function arrayp object

This function returns t if object is an array (i.e., a string, vector, or bit vector).

          (arrayp "asdf")
          ⇒ t
          (arrayp [a])
          ⇒ t
          (arrayp #*101)
          ⇒ t

Function aref array index This function returns the indexth element of array. The first element is at index zero.

          (setq primes [2 3 5 7 11 13])
               ⇒ [2 3 5 7 11 13]
          (aref primes 4)
               ⇒ 11
          (elt primes 4)
               ⇒ 11

(aref "abcdefg" 1)
     ⇒ ?b
(aref #*1101 2) ⇒ 0

See also the function elt, in Section 6.1 [Sequence Functions], page 101.

Function aset array index object

This function sets the indexth element of array to be object. It returns object.

(setq w [foo bar baz])
     ⇒ [foo bar baz]
(aset w 0 'fu)
     ⇒ fu
w
     ⇒ [fu bar baz]
(setq x "asdfasfd")
     ⇒ "asdfasfd"
(aset x 3 ?Z)
     ⇒ ?Z
x
     ⇒ "asdZasfd"
(setq bv #*1111)
     ⇒ #*1111
(aset bv 2 0)
     ⇒ 0
bv

     ⇒ #*1101

If array is a string and object is not a character, a wrong-type-argument error results.

Function fillarray array object This function fills the array array with object, so that each element of array is object. It returns array.

(setq a [a b c d e f g])
     ⇒ [a b c d e f g]
(fillarray a 0)
     ⇒ [0 0 0 0 0 0 0]
a

     ⇒ [0 0 0 0 0 0 0]

(setq s "When in the course")
     ⇒ "When in the course"
(fillarray s ?-)
     ⇒ "------------------"
(setq bv #*1101)
     ⇒ #*1101
(fillarray bv 0)
     ⇒ #*0000

If array is a string and object is not a character, a wrong-type-argument error results.

The general sequence functions copy-sequence and length are often useful for objects known to be arrays. See Section 6.1 [Sequence Functions], page 101.

6.4 Vectors

Arrays in Lisp, like arrays in most languages, are blocks of memory whose elements can be accessed in constant time. A vector is a general-purpose array; its elements can be any Lisp objects. (The other kind of array in XEmacs Lisp is the string, whose elements must be characters.) Vectors in XEmacs serve as obarrays (vectors of symbols), although this is a shortcoming that should be fixed. They are also used internally as part of the representation of a byte-compiled function; if you print such a function, you will see a vector in it.

In XEmacs Lisp, the indices of the elements of a vector start from zero and count up from there.

Vectors are printed with square brackets surrounding the elements. Thus, a vector whose elements are the symbols a, b and a is printed as [a b a]. You can write vectors in the same way in Lisp input.

A vector, like a string or a number, is considered a constant for evaluation: the result of evaluating it is the same vector. This does not evaluate or even examine the elements of the vector. See Section 8.2.1 [Self-Evaluating Forms], page 122.

Here are examples of these principles:

     (setq avector [1 two '(three) "four" [five]])
          ⇒ [1 two (quote (three)) "four" [five]]
     (eval avector)
          ⇒ [1 two (quote (three)) "four" [five]]
     (eq avector (eval avector))
          ⇒ t

6.5 Functions That Operate on Vectors

Here are some functions that relate to vectors:

Function vectorp object

This function returns t if object is a vector.

          (vectorp [a])
               ⇒ t
          (vectorp "asdf")
               ⇒ nil

Function vector &rest objects This function creates and returns a vector whose elements are the arguments, objects.

          (vector 'foo 23 [bar baz] "rats")
               ⇒ [foo 23 [bar baz] "rats"]
          (vector)
               ⇒ []

Function make-vector length object This function returns a new vector consisting of length elements, each initialized to object.

     (setq sleepy (make-vector 9 'Z))
          ⇒ [Z Z Z Z Z Z Z Z Z]

Function vconcat &rest sequences This function returns a new vector containing all the elements of the sequences. The arguments sequences may be lists, vectors, or strings. If no sequences are given, an empty vector is returned.

The value is a newly constructed vector that is not eq to any existing vector.

     (setq a (vconcat '(A B C) '(D E F)))
          ⇒ [A B C D E F]
     (eq a (vconcat a))
          ⇒ nil
     (vconcat)
          ⇒ []
     (vconcat [A B C] "aa" '(foo (6 7)))
          ⇒ [A B C 97 97 foo (6 7)]

The vconcat function also allows integers as arguments. It converts them to strings of digits, making up the decimal print representation of the integer, and then uses the strings instead of the original integers. Don’t use this feature; we plan to eliminate it. If you already use this feature, change your programs now! The proper way to convert an integer to a decimal number in this way is with format (see Section 4.10 [Formatting Strings], page 67) or number-to-string (see Section 4.7 [String Conversion], page 65).

For other concatenation functions, see mapconcat in Section 11.6 [Mapping Functions], page 171, concat in Section 4.3 [Creating Strings], page 60, append in Section 5.5 [Building Lists], page 82, and bvconcat in Section 6.7 [Bit Vector Functions], page 108.

The append function provides a way to convert a vector into a list with the same elements (see Section 5.5 [Building Lists], page 82):

(setq avector [1 two (quote (three)) "four" [five]])
     ⇒ [1 two (quote (three)) "four" [five]]
(append avector nil)
     ⇒ (1 two (quote (three)) "four" [five])

6.6 Bit Vectors

Bit vectors are specialized vectors that can only represent arrays of 1’s and 0’s. Bit vectors have a very efficient representation and are useful for representing sets of boolean (true or false) values.

There is no limit on the size of a bit vector. You could, for example, create a bit vector with 100,000 elements if you really wanted to.

Bit vectors have a special printed representation consisting of ‘#*’ followed by the bits of the vector. For example, a bit vector whose elements are 0, 1, 1, 0, and 1, respectively, is printed as

#*01101

Bit vectors are considered constants for evaluation, like vectors, strings, and numbers. See Section 8.2.1 [Self-Evaluating Forms], page 122.

6.7 Functions That Operate on Bit Vectors

Here are some functions that relate to bit vectors:

Function bit-vector-p object

This function returns t if object is a bit vector.

     (bit-vector-p #*01)
          ⇒ t
     (bit-vector-p [0 1])
          ⇒ nil
     (bit-vector-p "01")
          ⇒ nil

Function bitp object This function returns t if object is either 0 or 1.

Function bit-vector &rest objects This function creates and returns a bit vector whose elements are the arguments objects. The elements must be either of the two integers 0 or 1.

     (bit-vector 0 0 0 1 0 0 0 0 1 0)
          ⇒ #*0001000010
     (bit-vector)
          ⇒ #*

Function make-bit-vector length object

This function creates and returns a bit vector consisting of length elements, each initialized to object.

     (setq picket-fence (make-bit-vector 9 1))
          ⇒ #*111111111

Function bvconcat &rest sequences This function returns a new bit vector containing all the elements of the sequences. The arguments sequences may be lists, vectors, or bit vectors, all of whose elements are the integers 0 or 1. If no sequences are given, an empty bit vector is returned. The value is a newly constructed bit vector that is not eq to any existing bit vector.

     (setq a (bvconcat '(1 1 0) '(0 0 1)))
          ⇒ #*110001
     (eq a (bvconcat a))
          ⇒ nil

     (bvconcat)
          ⇒ #*
     (bvconcat [1 0 0 0 0] #*111 '(0 0 0 0 1))
          ⇒ #*1000011100001

For other concatenation functions, see mapconcat in Section 11.6 [Mapping Functions], page 171, concat in Section 4.3 [Creating Strings], page 60, vconcat in Section 6.5 [Vector Functions], page 106, and append in Section 5.5 [Building Lists], page 82.

The append function provides a way to convert a bit vector into a list with the same elements (see Section 5.5 [Building Lists], page 82):

(setq bv #*00001110)
     ⇒ #*00001110
(append bv nil)
     ⇒ (0 0 0 0 1 1 1 0)

110 XEmacs Lisp Reference Manual

7 Symbols

A symbol is an object with a unique name. This chapter describes symbols, their components, their property lists, and how they are created and interned. Separate chapters describe the use of symbols as variables and as function names; see Chapter 10 [Variables], page 145, and Chapter 11 [Functions], page 163. For the precise read syntax for symbols, see Section 2.4.4 [Symbol Type], page 21.

You can test whether an arbitrary Lisp object is a symbol with symbolp:

Function symbolp object

This function returns t if object is a symbol, nil otherwise.

7.1 Symbol Components

Each symbol has four components (or “cells”), each of which references another object:

Print name The print name cell holds a string that names the symbol for reading and printing. See symbol-name in Section 7.3 [Creating Symbols], page 113.

Value The value cell holds the current value of the symbol as a variable. When a symbol is used as a form, the value of the form is the contents of the symbol’s value cell. See symbol-value in Section 10.6 [Accessing Variables], page 151.

Function The function cell holds the function definition of the symbol. When a symbol is used as a function, its function definition is used in its place. This cell is also used to make a symbol stand for a keymap or a keyboard macro, for editor command execution. Because each symbol has separate value and function cells, variables and function names do not conflict. See symbol-function in Section 11.8 [Function Cells], page 173.

Property list The property list cell holds the property list of the symbol. See symbol-plist in Section 7.4 [Symbol Properties], page 116.

The print name cell always holds a string, and cannot be changed. The other three cells can be set individually to any specified Lisp object.

The print name cell holds the string that is the name of the symbol. Since symbols are represented textually by their names, it is important not to have two symbols with the same name. The Lisp reader ensures this: every time it reads a symbol, it looks for an existing symbol with the specified name before it creates a new one. (In XEmacs Lisp, this lookup uses a hashing algorithm and an obarray; see Section 7.3 [Creating Symbols], page 113.)

In normal usage, the function cell usually contains a function or macro, as that is what the Lisp interpreter expects to see there (see Chapter 8 [Evaluation], page 119). Keyboard macros (see Section 19.13 [Keyboard Macros], page 313), keymaps (see Chapter 20 [Keymaps], page 315) and autoload objects (see Section 8.2.8 [Autoloading], page 126) are also sometimes stored in the function cell of symbols. We often refer to “the function foo” when we really mean the function stored in the function cell of the symbol foo. We make the distinction only when necessary.

The property list cell normally should hold a correctly formatted property list (see Section 5.9 [Property Lists], page 96), as a number of functions expect to see a property list there.

The function cell or the value cell may be void, which means that the cell does not reference any object. (This is not the same thing as holding the symbol void, nor the same as holding the symbol nil.) Examining a cell that is void results in an error, such as

‘Symbol's value as variable is void’.
   The four functions symbol-name, symbol-value, symbol-plist, and symbol-function

return the contents of the four cells of a symbol. Here as an example we show the contents of the four cells of the symbol buffer-file-name:

     (symbol-name 'buffer-file-name)
          ⇒ "buffer-file-name"
     (symbol-value 'buffer-file-name)
          ⇒ "/gnu/elisp/symbols.texi"
     (symbol-plist 'buffer-file-name)
          ⇒ (variable-documentation 29529)
     (symbol-function 'buffer-file-name)
          ⇒ #<subr buffer-file-name>

Because this symbol is the variable which holds the name of the file being visited in the current buffer, the value cell contents we see are the name of the source file of this chapter of the XEmacs Lisp Manual. The property list cell contains the list (variable-documentation 29529) which tells the documentation functions where to find the documentation string for the variable buffer-file-name in the ‘DOC’ file. (29529 is the offset from the beginning of the ‘DOC’ file to where that documentation string begins.) The function cell contains the function for returning the name of the file. buffer-file-name names a primitive function, which has no read syntax and prints in hash notation (see Section 2.4.13 [Primitive Function Type], page 27). A symbol naming a function written in Lisp would have a lambda expression (or a byte-code object) in this cell.

7.2 Defining Symbols

A definition in Lisp is a special form that announces your intention to use a certain symbol in a particular way. In XEmacs Lisp, you can define a symbol as a variable, or define it as a function (or macro), or both independently.

A definition construct typically specifies a value or meaning for the symbol for one kind of use, plus documentation for its meaning when used in this way. Thus, when you define a symbol as a variable, you can supply an initial value for the variable, plus documentation for the variable.

defvar and defconst are special forms that define a symbol as a global variable. They are documented in detail in Section 10.5 [Defining Variables], page 149.

defun defines a symbol as a function, creating a lambda expression and storing it in the function cell of the symbol. This lambda expression thus becomes the function definition of the symbol. (The term “function definition”, meaning the contents of the function cell, is derived from the idea that defun gives the symbol its definition as a function.) defsubst, define-function and defalias are other ways of defining a function. See Chapter 11 [Functions], page 163.

defmacro defines a symbol as a macro. It creates a macro object and stores it in the function cell of the symbol. Note that a given symbol can be a macro or a function, but not both at once, because both macro and function definitions are kept in the function cell, and that cell can hold only one Lisp object at any given time. See Chapter 12 [Macros], page 177.

In XEmacs Lisp, a definition is not required in order to use a symbol as a variable or function. Thus, you can make a symbol a global variable with setq, whether you define it first or not. The real purpose of definitions is to guide programmers and programming tools. They inform programmers who read the code that certain symbols are intended to be used as variables, or as functions. In addition, utilities such as ‘etags’ and ‘make-docfile’ recognize definitions, and add appropriate information to tag tables and the ‘DOC’ file. See Section 27.2 [Accessing Documentation], page 382.

7.3 Creating and Interning Symbols

To understand how symbols are created in XEmacs Lisp, you must know how Lisp reads them. Lisp must ensure that it finds the same symbol every time it reads the same set of characters. Failure to do so would cause complete confusion.

When the Lisp reader encounters a symbol, it reads all the characters of the name. Then it “hashes” those characters to find an index in a table called an obarray. Hashing is an efficient method of looking something up. For example, instead of searching a telephone book cover to cover when looking up Jan Jones, you start with the J’s and go from there. That is a simple version of hashing. Each element of the obarray is a bucket which holds all the symbols with a given hash code; to look for a given name, it is sufficient to look through all the symbols in the bucket for that name’s hash code.

If a symbol with the desired name is found, the reader uses that symbol. If the obarray does not contain a symbol with that name, the reader makes a new symbol and adds it to the obarray. Finding or adding a symbol with a certain name is called interning it, and the symbol is then called an interned symbol.

Interning ensures that each obarray has just one symbol with any particular name. Other like-named symbols may exist, but not in the same obarray. Thus, the reader gets the same symbols for the same names, as long as you keep reading with the same obarray.

No obarray contains all symbols; in fact, some symbols are not in any obarray. They are called uninterned symbols. An uninterned symbol has the same four cells as other symbols; however, the only way to gain access to it is by finding it in some other object or as the value of a variable.

In XEmacs Lisp, an obarray is actually a vector. Each element of the vector is a bucket; its value is either an interned symbol whose name hashes to that bucket, or 0 if the bucket is empty. Each interned symbol has an internal link (invisible to the user) to the next symbol in the bucket. Because these links are invisible, there is no way to find all the symbols in an obarray except using mapatoms (below). The order of symbols in a bucket is not significant.

In an empty obarray, every element is 0, and you can create an obarray with (makevector length 0). This is the only valid way to create an obarray. Prime numbers as lengths tend to result in good hashing; lengths one less than a power of two are also good.

Do not try to put symbols in an obarray yourself. This does not work—only intern can enter a symbol in an obarray properly. Do not try to intern one symbol in two obarrays. This would garble both obarrays, because a symbol has just one slot to hold the following symbol in the obarray bucket. The results would be unpredictable.

It is possible for two different symbols to have the same name in different obarrays; these symbols are not eq or equal. However, this normally happens only as part of the abbrev mechanism (see Chapter 39 [Abbrevs], page 579).

Common Lisp note: In Common Lisp, a single symbol may be interned in several obarrays.

Most of the functions below take a name and sometimes an obarray as arguments. A wrong-type-argument error is signaled if the name is not a string, or if the obarray is not a vector.

Function symbol-name symbol

This function returns the string that is symbol’s name. For example:

     (symbol-name 'foo)
          ⇒ "foo"

Changing the string by substituting characters, etc, does change the name of the symbol, but fails to update the obarray, so don’t do it!

Function make-symbol name

This function returns a newly-allocated, uninterned symbol whose name is name (which must be a string). Its value and function definition are void, and its property list is nil. In the example below, the value of sym is not eq to foo because it is a distinct uninterned symbol whose name is also ‘foo’.

     (setq sym (make-symbol "foo"))
          ⇒ foo
     (eq sym 'foo)
          ⇒ nil

Function intern name &optional obarray This function returns the interned symbol whose name is name. If there is no such symbol in the obarray obarray, intern creates a new one, adds it to the obarray, and returns it. If obarray is omitted, the value of the global variable

obarray is used.
     (setq sym (intern "foo"))
          ⇒ foo
     (eq sym 'foo)
          ⇒ t
     (setq sym1 (intern "foo" other-obarray))
          ⇒ foo
     (eq sym 'foo)
          ⇒ nil

Function intern-soft name &optional obarray This function returns the symbol in obarray whose name is name, or nil if obarray has no symbol with that name. Therefore, you can use intern-soft to test whether a symbol with a given name is already interned. If obarray is omitted, the value of the global variable obarray is used.

     (intern-soft "frazzle")        ; No such symbol exists.
          ⇒ nil
     (make-symbol "frazzle")        ; Create an uninterned one.
          ⇒ frazzle
     (intern-soft "frazzle")        ; That one cannot be found.
          ⇒ nil
     (setq sym (intern "frazzle"))  ; Create an interned one.
          ⇒ frazzle
     (intern-soft "frazzle")        ; That one can be found!
          ⇒ frazzle

(eq sym 'frazzle) ; And it is the same one.

          ⇒ t

Variable obarray

This variable is the standard obarray for use by intern and read.

Function mapatoms function &optional obarray This function calls function for each symbol in the obarray obarray. It returns nil. If obarray is omitted, it defaults to the value of obarray, the standard obarray for ordinary symbols.

     (setq count 0)
          ⇒ 0
     (defun count-syms (s)
       (setq count (1+ count)))
          ⇒ count-syms
     (mapatoms 'count-syms)
          ⇒ nil
     count

          ⇒ 1871

See documentation in Section 27.2 [Accessing Documentation], page 382, for another example using mapatoms.

Function unintern symbol &optional obarray This function deletes symbol from the obarray obarray. If symbol is not actually in the obarray, unintern does nothing. If obarray is nil, the current obarray is used.

If you provide a string instead of a symbol as symbol, it stands for a symbol name. Then unintern deletes the symbol (if any) in the obarray which has that name. If there is no such symbol, unintern does nothing.

If unintern does delete a symbol, it returns t. Otherwise it returns nil.

7.4 Symbol Properties

A property list (plist for short) is a list of paired elements stored in the property list cell of a symbol. Each of the pairs associates a property name (usually a symbol) with a property or value. Property lists are generally used to record information about a symbol, such as its documentation as a variable, the name of the file where it was defined, or perhaps even the grammatical class of the symbol (representing a word) in a language-understanding system.

Many objects other than symbols can have property lists associated with them, and XEmacs provides a full complement of functions for working with property lists. See Section 5.9 [Property Lists], page 96.

The property names and values in a property list can be any Lisp objects, but the names are usually symbols. They are compared using eq. Here is an example of a property list, found on the symbol progn when the compiler is loaded:

     (lisp-indent-function 0 byte-compile byte-compile-progn)

Here lisp-indent-function and byte-compile are property names, and the other two elements are the corresponding values.

7.4.1 Property Lists and Association Lists

Association lists (see Section 5.8 [Association Lists], page 92) are very similar to property lists. In contrast to association lists, the order of the pairs in the property list is not significant since the property names must be distinct.

Property lists are better than association lists for attaching information to various Lisp function names or variables. If all the associations are recorded in one association list, the program will need to search that entire list each time a function or variable is to be operated on. By contrast, if the information is recorded in the property lists of the function names or variables themselves, each search will scan only the length of one property list, which is usually short. This is why the documentation for a variable is recorded in a property named variable-documentation. The byte compiler likewise uses properties to record those functions needing special treatment.

However, association lists have their own advantages. Depending on your application, it may be faster to add an association to the front of an association list than to update a property. All properties for a symbol are stored in the same property list, so there is a possibility of a conflict between different uses of a property name. (For this reason, it is a good idea to choose property names that are probably unique, such as by including the name of the library in the property name.) An association list may be used like a stack where associations are pushed on the front of the list and later discarded; this is not possible with a property list.

7.4.2 Property List Functions for Symbols

Function symbol-plist symbol

This function returns the property list of symbol.

Function setplist symbol plist This function sets symbol’s property list to plist. Normally, plist should be a well-formed property list, but this is not enforced.

     (setplist 'foo '(a 1 b (2 3) c nil))
          ⇒ (a 1 b (2 3) c nil)
     (symbol-plist 'foo)
          ⇒ (a 1 b (2 3) c nil)

For symbols in special obarrays, which are not used for ordinary purposes, it may make sense to use the property list cell in a nonstandard fashion; in fact, the abbrev mechanism does so (see Chapter 39 [Abbrevs], page 579).

Function get symbol property This function finds the value of the property named property in symbol’s property list. If there is no such property, nil is returned. Thus, there is no distinction between a value of nil and the absence of the property. The name property is compared with the existing property names using eq, so any object is a legitimate property. See put for an example.

Function put symbol property value This function puts value onto symbol’s property list under the property name property, replacing any previous property value. The put function returns value.

     (put 'fly 'verb 'transitive)
          ⇒'transitive
     (put 'fly 'noun '(a buzzing little bug))
          ⇒ (a buzzing little bug)
     (get 'fly 'verb)
          ⇒ transitive
     (symbol-plist 'fly)
          ⇒ (verb transitive noun (a buzzing little bug))

7.4.3 Property Lists Outside Symbols

These functions are useful for manipulating property lists that are stored in places other than symbols:

Function getf plist property &optional default This returns the value of the property property stored in the property list plist. For example,

     (getf '(foo 4) 'foo)
          ⇒ 4

Function putf plist property value This stores value as the value of the property property in the property list plist. It may modify plist destructively, or it may construct a new list structure without altering the old. The function returns the modified property list, so you can store that back in the place where you got plist. For example,

     (setq my-plist '(bar t foo 4))
          ⇒ (bar t foo 4)
     (setq my-plist (putf my-plist 'foo 69))
          ⇒ (bar t foo 69)
     (setq my-plist (putf my-plist 'quux '(a)))
          ⇒ (quux (a) bar t foo 5)

Function plists-eq a b

This function returns non-nil if property lists a and b are eq. This means that the property lists have the same values for all the same properties, where comparison between values is done using eq.

Function plists-equal a b

This function returns non-nil if property lists a and b are equal.

Both of the above functions do order-insensitive comparisons.

(plists-eq '(a 1 b 2 c nil) '(b 2 a 1))
     ⇒ t
(plists-eq '(foo "hello" bar "goodbye") '(bar "goodbye" foo "hello"))
     ⇒ nil
(plists-equal '(foo "hello" bar "goodbye") '(bar "goodbye" foo "hello"))
     ⇒ t

8 Evaluation

The evaluation of expressions in XEmacs Lisp is performed by the Lisp interpreter—a program that receives a Lisp object as input and computes its value as an expression. How it does this depends on the data type of the object, according to rules described in this chapter. The interpreter runs automatically to evaluate portions of your program, but can also be called explicitly via the Lisp primitive function eval.

A Lisp object that is intended for evaluation is called an expression or a form. The fact that expressions are data objects and not merely text is one of the fundamental differences between Lisp-like languages and typical programming languages. Any object can be evaluated, but in practice only numbers, symbols, lists and strings are evaluated very often.

It is very common to read a Lisp expression and then evaluate the expression, but reading and evaluation are separate activities, and either can be performed alone. Reading per se does not evaluate anything; it converts the printed representation of a Lisp object to the object itself. It is up to the caller of read whether this object is a form to be evaluated, or serves some entirely different purpose. See Section 17.3 [Input Functions], page 253.

Do not confuse evaluation with command key interpretation. The editor command loop translates keyboard input into a command (an interactively callable function) using the active keymaps, and then uses call-interactively to invoke the command. The execution of the command itself involves evaluation if the command is written in Lisp, but that is not a part of command key interpretation itself. See Chapter 19 [Command Loop], page 281.

Evaluation is a recursive process. That is, evaluation of a form may call eval to evaluate parts of the form. For example, evaluation of a function call first evaluates each argument of the function call, and then evaluates each form in the function body. Consider evaluation of the form (car x): the subform x must first be evaluated recursively, so that its value can be passed as an argument to the function car.

Evaluation of a function call ultimately calls the function specified in it. See Chapter 11 [Functions], page 163. The execution of the function may itself work by evaluating the function definition; or the function may be a Lisp primitive implemented in C, or it may be a byte-compiled function (see Chapter 15 [Byte Compilation], page 205).

The evaluation of forms takes place in a context called the environment, which consists of the current values and bindings of all Lisp variables.^{1} Whenever the form refers to a variable without creating a new binding for it, the value of the binding in the current environment is used. See Chapter 10 [Variables], page 145.

Evaluation of a form may create new environments for recursive evaluation by binding variables (see Section 10.3 [Local Variables], page 146). These environments are temporary and vanish by the time evaluation of the form is complete. The form may also make changes that persist; these changes are called side effects. An example of a form that produces side

effects is (setq foo 1).

The details of what evaluation means for each kind of form are described below (see Section 8.2 [Forms], page 121).

1 This definition of “environment” is specifically not intended to include all the data that can affect the result of a program.

8.1 Eval

Most often, forms are evaluated automatically, by virtue of their occurrence in a program being run. On rare occasions, you may need to write code that evaluates a form that is computed at run time, such as after reading a form from text being edited or getting one from a property list. On these occasions, use the eval function.

Please note: it is generally cleaner and more flexible to call functions that are stored in data structures, rather than to evaluate expressions stored in data structures. Using functions provides the ability to pass information to them as arguments.

The functions and variables described in this section evaluate forms, specify limits to the evaluation process, or record recently returned values. Loading a file also does evaluation (see Chapter 14 [Loading], page 195).

Function eval form

This is the basic function for performing evaluation. It evaluates form in the current environment and returns the result. How the evaluation proceeds depends on the type of the object (see Section 8.2 [Forms], page 121). Since eval is a function, the argument expression that appears in a call to eval is evaluated twice: once as preparation before eval is called, and again by the eval function itself. Here is an example:

     (setq foo 'bar)
          ⇒ bar
     (setq bar 'baz)
          ⇒ baz

;; eval receives argument bar, which is the value of foo

     (eval foo)
          ⇒ baz
     (eval 'foo)
          ⇒ bar

The number of currently active calls to eval is limited to max-lisp-eval-depth (see below).

Command eval-region start end &optional stream This function evaluates the forms in the current buffer in the region defined by the positions start and end. It reads forms from the region and calls eval on them until the end of the region is reached, or until an error is signaled and not handled. If stream is supplied, standard-output is bound to it during the evaluation. You can use the variable load-read-function to specify a function for evalregion to use instead of read for reading expressions. See Section 14.1 [How Programs Do Loading], page 195.

eval-region always returns nil.

Command eval-buffer buffer &optional stream This is like eval-region except that it operates on the whole contents of buffer.

Variable max-lisp-eval-depth

This variable defines the maximum depth allowed in calls to eval, apply, and funcall before an error is signaled (with error message "Lisp nesting exceeds max-lisp-eval-depth"). This counts internal uses of those functions, such as for calling the functions mentioned in Lisp expressions, and recursive evaluation of function call arguments and function body forms.

This limit, with the associated error when it is exceeded, is one way that Lisp avoids infinite recursion on an ill-defined function.

The default value of this variable is 500. If you set it to a value less than 100, Lisp will reset it to 100 if the given value is reached.

max-specpdl-size provides another limit on nesting. See Section 10.3 [Local Variables], page 146.

Variable values

The value of this variable is a list of the values returned by all the expressions that were read from buffers (including the minibuffer), evaluated, and printed. The elements are ordered most recent first.

     (setq x 1) ⇒ 1
     (list 'A (1+ 2) auto-save-default)
          ⇒ (A 3 t)
     values

          ⇒ ((A 3 t) 1 ...)

This variable is useful for referring back to values of forms recently evaluated. It is generally a bad idea to print the value of values itself, since this may be very long. Instead, examine particular elements, like this:

;; Refer to the most recent evaluation result.

     (nth 0 values)
          ⇒ (A 3 t)

;; That put a new element on, ;; so all elements move back one.

     (nth 1 values)
          ⇒ (A 3 t)

;; This gets the element that was next-to-most-recent ;; before this example.

     (nth 3 values)
          ⇒ 1

8.2 Kinds of Forms

A Lisp object that is intended to be evaluated is called a form. How XEmacs evaluates a form depends on its data type. XEmacs has three different kinds of form that are evaluated differently: symbols, lists, and “all other types”. This section describes all three kinds, starting with “all other types” which are self-evaluating forms.

8.2.1 Self-Evaluating Forms

A self-evaluating form is any form that is not a list or symbol. Self-evaluating forms evaluate to themselves: the result of evaluation is the same object that was evaluated. Thus, the number 25 evaluates to 25, and the string "foo" evaluates to the string "foo". Likewise, evaluation of a vector does not cause evaluation of the elements of the vector—it returns the same vector with its contents unchanged.

'123 ; An object, shown without evaluation.

          ⇒ 123

123 ; Evaluated as usual—result is the same.

          ⇒ 123

(eval '123) ; Evaluated “by hand”—result is the same.

          ⇒ 123

(eval (eval '123)) ; Evaluating twice changes nothing.

          ⇒ 123

It is common to write numbers, characters, strings, and even vectors in Lisp code, taking advantage of the fact that they self-evaluate. However, it is quite unusual to do this for types that lack a read syntax, because there’s no way to write them textually. It is possible to construct Lisp expressions containing these types by means of a Lisp program. Here is an example:

;; Build an expression containing a buffer object.

     (setq buffer (list 'print (current-buffer)))
          ⇒ (print #<buffer eval.texi>)

;; Evaluate it.

     (eval buffer)
          ⊣ #<buffer eval.texi>
          ⇒ #<buffer eval.texi>

8.2.2 Symbol Forms

When a symbol is evaluated, it is treated as a variable. The result is the variable’s value, if it has one. If it has none (if its value cell is void), an error is signaled. For more information on the use of variables, see Chapter 10 [Variables], page 145.

In the following example, we set the value of a symbol with setq. Then we evaluate the symbol, and get back the value that setq stored.

     (setq a 123) ⇒ 123
     (eval 'a) ⇒ 123
     a  ⇒ 123

The symbols nil and t are treated specially, so that the value of nil is always nil, and the value of t is always t; you cannot set or bind them to any other values. Thus, these two symbols act like self-evaluating forms, even though eval treats them like any other symbol.

8.2.3 Classification of List Forms

A form that is a nonempty list is either a function call, a macro call, or a special form, according to its first element. These three kinds of forms are evaluated in different ways, described below. The remaining list elements constitute the arguments for the function, macro, or special form.

The first step in evaluating a nonempty list is to examine its first element. This element alone determines what kind of form the list is and how the rest of the list is to be processed. The first element is not evaluated, as it would be in some Lisp dialects such as Scheme.

8.2.4 Symbol Function Indirection

If the first element of the list is a symbol then evaluation examines the symbol’s function cell, and uses its contents instead of the original symbol. If the contents are another symbol, this process, called symbol function indirection, is repeated until it obtains a non-symbol. See Section 11.3 [Function Names], page 167, for more information about using a symbol as a name for a function stored in the function cell of the symbol.

One possible consequence of this process is an infinite loop, in the event that a symbol’s function cell refers to the same symbol. Or a symbol may have a void function cell, in which case the subroutine symbol-function signals a void-function error. But if neither of these things happens, we eventually obtain a non-symbol, which ought to be a function or other suitable object.

More precisely, we should now have a Lisp function (a lambda expression), a byte-code function, a primitive function, a Lisp macro, a special form, or an autoload object. Each of these types is a case described in one of the following sections. If the object is not one of these types, the error invalid-function is signaled.

The following example illustrates the symbol indirection process. We use fset to set the function cell of a symbol and symbol-function to get the function cell contents (see Section 11.8 [Function Cells], page 173). Specifically, we store the symbol car into the function cell of first, and the symbol first into the function cell of erste.

;; Build this function cell linkage:

     ;;   -------------       -----        -------        -------
     ;;  | #<subr car> | <-- | car |  <-- | first |  <-- | erste |
     ;;   -------------       -----        -------        -------
     (symbol-function 'car)
          ⇒ #<subr car>
     (fset 'first 'car)
          ⇒ car
     (fset 'erste 'first)
          ⇒ first

(erste '(1 2 3)) ; Call the function referenced by erste.

          ⇒ 1

By contrast, the following example calls a function without any symbol function indirection, because the first element is an anonymous Lisp function, not a symbol.

     ((lambda (arg) (erste arg))
      '(1 2 3))
          ⇒ 1

Executing the function itself evaluates its body; this does involve symbol function indirection when calling erste.

The built-in function indirect-function provides an easy way to perform symbol function indirection explicitly.

Function indirect-function function

This function returns the meaning of function as a function. If function is a symbol, then it finds function’s function definition and starts over with that value. If function is not a symbol, then it returns function itself. Here is how you could define indirect-function in Lisp:

          (defun indirect-function (function)
            (if (symbolp function)
                (indirect-function (symbol-function function))
              function))

8.2.5 Evaluation of Function Forms

If the first element of a list being evaluated is a Lisp function object, byte-code object or primitive function object, then that list is a function call. For example, here is a call to the function +:

     (+ 1 x)

The first step in evaluating a function call is to evaluate the remaining elements of the list from left to right. The results are the actual argument values, one value for each list element. The next step is to call the function with this list of arguments, effectively using the function apply (see Section 11.5 [Calling Functions], page 169). If the function is written in Lisp, the arguments are used to bind the argument variables of the function (see Section 11.2 [Lambda Expressions], page 164); then the forms in the function body are evaluated in order, and the value of the last body form becomes the value of the function call.

8.2.6 Lisp Macro Evaluation

If the first element of a list being evaluated is a macro object, then the list is a macro call. When a macro call is evaluated, the elements of the rest of the list are not initially evaluated. Instead, these elements themselves are used as the arguments of the macro. The macro definition computes a replacement form, called the expansion of the macro, to be evaluated in place of the original form. The expansion may be any sort of form: a selfevaluating constant, a symbol, or a list. If the expansion is itself a macro call, this process of expansion repeats until some other sort of form results.

Ordinary evaluation of a macro call finishes by evaluating the expansion. However, the macro expansion is not necessarily evaluated right away, or at all, because other programs also expand macro calls, and they may or may not evaluate the expansions.

Normally, the argument expressions are not evaluated as part of computing the macro expansion, but instead appear as part of the expansion, so they are computed when the expansion is computed.

For example, given a macro defined as follows:

     (defmacro cadr (x)
       (list 'car (list 'cdr x)))

an expression such as (cadr (assq 'handler list)) is a macro call, and its expansion is:

     (car (cdr (assq 'handler list)))

Note that the argument (assq 'handler list) appears in the expansion.

See Chapter 12 [Macros], page 177, for a complete description of XEmacs Lisp macros.

8.2.7 Special Forms

A special form is a primitive function specially marked so that its arguments are not all evaluated. Most special forms define control structures or perform variable bindings—things which functions cannot do.

Each special form has its own rules for which arguments are evaluated and which are used without evaluation. Whether a particular argument is evaluated may depend on the results of evaluating other arguments.

Here is a list, in alphabetical order, of all of the special forms in XEmacs Lisp with a reference to where each is described. and see Section 9.3 [Combining Conditions], page 132 catch see Section 9.5.1 [Catch and Throw], page 134 cond see Section 9.2 [Conditionals], page 130

condition-case

see Section 9.5.3.3 [Handling Errors], page 138

defconst see Section 10.5 [Defining Variables], page 149 defmacro see Section 12.4 [Defining Macros], page 179 defun see Section 11.4 [Defining Functions], page 168 defvar see Section 10.5 [Defining Variables], page 149 function see Section 11.7 [Anonymous Functions], page 172 if see Section 9.2 [Conditionals], page 130

interactive

see Section 19.3 [Interactive Call], page 286

let

let* see Section 10.3 [Local Variables], page 146 or see Section 9.3 [Combining Conditions], page 132

prog1
prog2

progn see Section 9.1 [Sequencing], page 129 quote see Section 8.3 [Quoting], page 126

save-current-buffer

see Section 34.3 [Excursions], page 495

save-excursion

see Section 34.3 [Excursions], page 495

save-restriction

see Section 34.4 [Narrowing], page 496

save-selected-window

see Section 34.3 [Excursions], page 495

save-window-excursion

see Section 31.16 [Window Configurations], page 468

setq see Section 10.7 [Setting Variables], page 152

setq-default

see Section 10.9.2 [Creating Buffer-Local], page 157

unwind-protect

see Section 9.5 [Nonlocal Exits], page 134

while see Section 9.4 [Iteration], page 133

with-output-to-temp-buffer

see Section 45.8 [Temporary Displays], page 656

Common Lisp note: here are some comparisons of special forms in XEmacs Lisp and Common Lisp. setq, if, and catch are special forms in both XEmacs Lisp and Common Lisp. defun is a special form in XEmacs Lisp, but a macro in Common Lisp. save-excursion is a special form in XEmacs Lisp, but doesn’t exist in Common Lisp. throw is a special form in Common Lisp (because it must be able to throw multiple values), but it is a function in XEmacs Lisp (which doesn’t have multiple values).

8.2.8 Autoloading

The autoload feature allows you to call a function or macro whose function definition has not yet been loaded into XEmacs. It specifies which file contains the definition. When an autoload object appears as a symbol’s function definition, calling that symbol as a function automatically loads the specified file; then it calls the real definition loaded from that file. See Section 14.2 [Autoload], page 198.

8.3 Quoting

The special form quote returns its single argument, as written, without evaluating it. This provides a way to include constant symbols and lists, which are not self-evaluating objects, in a program. (It is not necessary to quote self-evaluating objects such as numbers, strings, and vectors.)

Special Form quote object This special form returns object, without evaluating it.

Because quote is used so often in programs, Lisp provides a convenient read syntax for it. An apostrophe character (‘'’) followed by a Lisp object (in read syntax) expands to a list whose first element is quote, and whose second element is the object. Thus, the read syntax 'x is an abbreviation for (quote x).

Here are some examples of expressions that use quote:

     (quote (+ 1 2)) ⇒ (+ 1 2)
     (quote foo) ⇒ foo
     'foo ⇒ foo
     ''foo
          ⇒ (quote foo)
     '(quote foo)
          ⇒ (quote foo)
     ['foo]
          ⇒ [(quote foo)]

Other quoting constructs include function (see Section 11.7 [Anonymous Functions], page 172), which causes an anonymous lambda expression written in Lisp to be compiled, and ‘`’ (see Section 12.5 [Backquote], page 179), which is used to quote only part of a list, while computing and substituting other parts.

128 XEmacs Lisp Reference Manual

9 Control Structures

A Lisp program consists of expressions or forms (see Section 8.2 [Forms], page 121). We control the order of execution of the forms by enclosing them in control structures. Control structures are special forms which control when, whether, or how many times to execute the forms they contain.

The simplest order of execution is sequential execution: first form a, then form b, and so on. This is what happens when you write several forms in succession in the body of a function, or at top level in a file of Lisp code—the forms are executed in the order written. We call this textual order. For example, if a function body consists of two forms a and b, evaluation of the function evaluates first a and then b, and the function’s value is the value of b.

Explicit control structures make possible an order of execution other than sequential.

XEmacs Lisp provides several kinds of control structure, including other varieties of sequencing, conditionals, iteration, and (controlled) jumps—all discussed below. The builtin control structures are special forms since their subforms are not necessarily evaluated or not evaluated sequentially. You can use macros to define your own control structure constructs (see Chapter 12 [Macros], page 177).

9.1 Sequencing

Evaluating forms in the order they appear is the most common way control passes from one form to another. In some contexts, such as in a function body, this happens automatically. Elsewhere you must use a control structure construct to do this: progn, the simplest control construct of Lisp.

A progn special form looks like this:

     (progn a b c ...)

and it says to execute the forms a, b, c and so on, in that order. These forms are called the body of the progn form. The value of the last form in the body becomes the value of the entire progn.

In the early days of Lisp, progn was the only way to execute two or more forms in succession and use the value of the last of them. But programmers found they often needed to use a progn in the body of a function, where (at that time) only one form was allowed. So the body of a function was made into an “implicit progn”: several forms are allowed just as in the body of an actual progn. Many other control structures likewise contain an implicit progn. As a result, progn is not used as often as it used to be. It is needed now most often inside an unwind-protect, and, or, or in the then-part of an if.

Special Form progn forms…

This special form evaluates all of the forms, in textual order, returning the result of the final form.

     (progn (print "The first form")
            (print "The second form")
            (print "The third form"))
          ⊣ "The first form"
          ⊣ "The second form"
          ⊣ "The third form"
     ⇒ "The third form"

Two other control constructs likewise evaluate a series of forms but return a different value:

Special Form prog1 form1 forms… This special form evaluates form1 and all of the forms, in textual order, returning the result of form1.

     (prog1 (print "The first form")
            (print "The second form")
            (print "The third form"))
          ⊣ "The first form"
          ⊣ "The second form"
          ⊣ "The third form"
     ⇒ "The first form"

Here is a way to remove the first element from a list in the variable x, then return the value of that former element:

     (prog1 (car x) (setq x (cdr x)))

Special Form prog2 form1 form2 forms… This special form evaluates form1, form2, and all of the following forms, in textual order, returning the result of form2.

     (prog2 (print "The first form")
            (print "The second form")
            (print "The third form"))
          ⊣ "The first form"
          ⊣ "The second form"
          ⊣ "The third form"
     ⇒ "The second form"

9.2 Conditionals

Conditional control structures choose among alternatives. XEmacs Lisp has two conditional forms: if, which is much the same as in other languages, and cond, which is a generalized case statement.

Special Form if condition then-form else-forms… if chooses between the then-form and the else-forms based on the value of condition. If the evaluated condition is non-nil, then-form is evaluated and the result returned. Otherwise, the else-forms are evaluated in textual order, and the value of the last one is returned. (The else part of if is an example of an implicit progn. See Section 9.1 [Sequencing], page 129.)

If condition has the value nil, and no else-forms are given, if returns nil. if is a special form because the branch that is not selected is never evaluated— it is ignored. Thus, in the example below, true is not printed because print is never called.

     (if nil
         (print 'true)
       'very-false)
     ⇒ very-false

Special Form cond clause… cond chooses among an arbitrary number of alternatives. Each clause in the cond must be a list. The car of this list is the condition; the remaining elements, if any, the body-forms. Thus, a clause looks like this:

(condition body-forms...) cond tries the clauses in textual order, by evaluating the condition of each clause. If the value of condition is non-nil, the clause “succeeds”; then cond evaluates its body-forms, and the value of the last of body-forms becomes the value of the cond. The remaining clauses are ignored. If the value of condition is nil, the clause “fails”, so the cond moves on to the following clause, trying its condition. If every condition evaluates to nil, so that every clause fails, cond returns nil. A clause may also look like this:

(condition) Then, if condition is non-nil when tested, the value of condition becomes the value of the cond form. The following example has four clauses, which test for the cases where the value of x is a number, string, buffer and symbol, respectively:

     (cond ((numberp x) x)
           ((stringp x) x)
           ((bufferp x)
            (setq temporary-hack x) ; multiple body-forms
            (buffer-name x))        ; in one clause
           ((symbolp x) (symbol-value x)))

Often we want to execute the last clause whenever none of the previous clauses was successful. To do this, we use t as the condition of the last clause, like this: (t body-forms). The form t evaluates to t, which is never nil, so this clause never fails, provided the cond gets to it at all. For example,

     (cond ((eq a 'hack) 'foo)
           (t "default"))
     ⇒ "default"

This expression is a cond which returns foo if the value of a is 1, and returns the string "default" otherwise.

Any conditional construct can be expressed with cond or with if. Therefore, the choice between them is a matter of style. For example:

(if a b c)

≡

(cond (a b) (t c))

9.3 Constructs for Combining Conditions

This section describes three constructs that are often used together with if and cond to express complicated conditions. The constructs and and or can also be used individually as kinds of multiple conditional constructs.

Function not condition This function tests for the falsehood of condition. It returns t if condition is nil, and nil otherwise. The function not is identical to null, and we recommend using the name null if you are testing for an empty list.

Special Form and conditions… The and special form tests whether all the conditions are true. It works by evaluating the conditions one by one in the order written.

If any of the conditions evaluates to nil, then the result of the and must be nil regardless of the remaining conditions; so and returns right away, ignoring the remaining conditions.

If all the conditions turn out non-nil, then the value of the last of them becomes the value of the and form.

Here is an example. The first condition returns the integer 1, which is not nil. Similarly, the second condition returns the integer 2, which is not nil. The third condition is nil, so the remaining condition is never evaluated.

     (and (print 1) (print 2) nil (print 3))
          ⊣ 1

          ⊣ 2

     ⇒ nil

Here is a more realistic example of using and:

     (if (and (consp foo) (eq (car foo) 'x))
         (message "foo is a list starting with x"))

Note that (car foo) is not executed if (consp foo) returns nil, thus avoiding an error.

and can be expressed in terms of either if or cond. For example:

(and arg1 arg2 arg3)

≡ (if arg1 (if arg2 arg3)) ≡

     (cond (arg1 (cond (arg2 arg3))))

Special Form or conditions… The or special form tests whether at least one of the conditions is true. It works by evaluating all the conditions one by one in the order written. If any of the conditions evaluates to a non-nil value, then the result of the or must be non-nil; so or returns right away, ignoring the remaining conditions. The value it returns is the non-nil value of the condition just evaluated. If all the conditions turn out nil, then the or expression returns nil. For example, this expression tests whether x is either 0 or nil:

     (or (eq x nil) (eq x 0))

Like the and construct, or can be written in terms of cond. For example: (or arg1 arg2 arg3) ≡

     (cond (arg1)

(arg2) (arg3))

You could almost write or in terms of if, but not quite:

(if arg1 arg1 (if arg2 arg2 arg3))

This is not completely equivalent because it can evaluate arg1 or arg2 twice. By contrast, (or arg1 arg2 arg3) never evaluates any argument more than once.

9.4 Iteration

Iteration means executing part of a program repetitively. For example, you might want to repeat some computation once for each element of a list, or once for each integer from 0 to n. You can do this in XEmacs Lisp with the special form while:

Special Form while condition forms… while first evaluates condition. If the result is non-nil, it evaluates forms in textual order. Then it reevaluates condition, and if the result is non-nil, it evaluates forms again. This process repeats until condition evaluates to nil. There is no limit on the number of iterations that may occur. The loop will continue until either condition evaluates to nil or until an error or throw jumps out of it (see Section 9.5 [Nonlocal Exits], page 134). The value of a while form is always nil.

     (setq num 0) ⇒ 0
     (while (< num 4)
       (princ (format "Iteration %d." num))
       (setq num (1+ num)))
          ⊣ Iteration 0.
          ⊣ Iteration 1.
          ⊣ Iteration 2.
          ⊣ Iteration 3.
          ⇒ nil

If you would like to execute something on each iteration before the end-test, put it together with the end-test in a progn as the first argument of while, as shown here:

          (while (progn
                   (forward-line 1)
                   (not (looking-at "^$"))))

This moves forward one line and continues moving by lines until it reaches an empty. It is unusual in that the while has no body, just the end test (which also does the real work of moving point).

9.5 Nonlocal Exits

A nonlocal exit is a transfer of control from one point in a program to another remote point. Nonlocal exits can occur in XEmacs Lisp as a result of errors; you can also use them under explicit control. Nonlocal exits unbind all variable bindings made by the constructs being exited.

9.5.1 Explicit Nonlocal Exits: catch and throw

Most control constructs affect only the flow of control within the construct itself. The function throw is the exception to this rule of normal program execution: it performs a nonlocal exit on request. (There are other exceptions, but they are for error handling only.) throw is used inside a catch, and jumps back to that catch. For example:

     (catch 'foo
       (progn
         ...
         (throw 'foo t)
         ...))

The throw transfers control straight back to the corresponding catch, which returns immediately. The code following the throw is not executed. The second argument of throw is used as the return value of the catch.

The throw and the catch are matched through the first argument: throw searches for a catch whose first argument is eq to the one specified. Thus, in the above example, the throw specifies foo, and the catch specifies the same symbol, so that catch is applicable. If there is more than one applicable catch, the innermost one takes precedence.

Executing throw exits all Lisp constructs up to the matching catch, including function calls. When binding constructs such as let or function calls are exited in this way, the bindings are unbound, just as they are when these constructs exit normally (see Section 10.3 [Local Variables], page 146). Likewise, throw restores the buffer and position saved by save-excursion (see Section 34.3 [Excursions], page 495), and the narrowing status saved

by save-restriction and the window selection saved by save-window-excursion (see

Section 31.16 [Window Configurations], page 468). It also runs any cleanups established with the unwind-protect special form when it exits that form (see Section 9.5.4 [Cleanups], page 142).

The throw need not appear lexically within the catch that it jumps to. It can equally well be called from another function called within the catch. As long as the throw takes place chronologically after entry to the catch, and chronologically before exit from it, it has access to that catch. This is why throw can be used in commands such as exitrecursive-edit that throw back to the editor command loop (see Section 19.10 [Recursive Editing], page 310).

Common Lisp note: Most other versions of Lisp, including Common Lisp, have several ways of transferring control nonsequentially: return, return-from, and go, for example. XEmacs Lisp has only throw.

Special Form catch tag body… catch establishes a return point for the throw function. The return point is distinguished from other such return points by tag, which may be any Lisp object. The argument tag is evaluated normally before the return point is established. With the return point in effect, catch evaluates the forms of the body in textual order. If the forms execute normally, without error or nonlocal exit, the value of the last body form is returned from the catch. If a throw is done within body specifying the same value tag, the catch exits immediately; the value it returns is whatever was specified as the second argument of throw.

Function throw tag value The purpose of throw is to return from a return point previously established with catch. The argument tag is used to choose among the various existing return points; it must be eq to the value specified in the catch. If multiple return points match tag, the innermost one is used. The argument value is used as the value to return from that catch. If no return point is in effect with tag tag, then a no-catch error is signaled with data (tag value).

9.5.2 Examples of catch and throw

One way to use catch and throw is to exit from a doubly nested loop. (In most languages, this would be done with a “go to”.) Here we compute (foo i j) for i and j varying from 0 to 9:

(defun search-foo ()
  (catch 'loop
    (let ((i 0))
      (while (< i 10)
        (let ((j 0))
          (while (< j 10)
            (if (foo i j)
                (throw 'loop (list i j)))
            (setq j (1+ j))))
        (setq i (1+ i))))))

If foo ever returns non-nil, we stop immediately and return a list of i and j. If foo always returns nil, the catch returns normally, and the value is nil, since that is the result of the

while.

Here are two tricky examples, slightly different, showing two return points at once. First, two return points with the same tag, hack:

     (defun catch2 (tag)
       (catch tag
         (throw 'hack 'yes)))
     ⇒ catch2
     (catch 'hack
       (print (catch2 'hack))
       'no)
     ⊣ yes
     ⇒ no

Since both return points have tags that match the throw, it goes to the inner one, the one established in catch2. Therefore, catch2 returns normally with value yes, and this value is printed. Finally the second body form in the outer catch, which is 'no, is evaluated and returned from the outer catch.

Now let’s change the argument given to catch2:

     (defun catch2 (tag)
       (catch tag
         (throw 'hack 'yes)))
     ⇒ catch2
     (catch 'hack
       (print (catch2 'quux))
       'no)
     ⇒ yes

We still have two return points, but this time only the outer one has the tag hack; the inner one has the tag quux instead. Therefore, throw makes the outer catch return the value yes. The function print is never called, and the body-form 'no is never evaluated.

9.5.3 Errors

When XEmacs Lisp attempts to evaluate a form that, for some reason, cannot be evaluated, it signals an error.

When an error is signaled, XEmacs’s default reaction is to print an error message and terminate execution of the current command. This is the right thing to do in most cases, such as if you type C-f at the end of the buffer.

In complicated programs, simple termination may not be what you want. For example, the program may have made temporary changes in data structures, or created temporary buffers that should be deleted before the program is finished. In such cases, you would use unwind-protect to establish cleanup expressions to be evaluated in case of error. (See Section 9.5.4 [Cleanups], page 142.) Occasionally, you may wish the program to continue execution despite an error in a subroutine. In these cases, you would use condition-case to establish error handlers to recover control in case of error.

Resist the temptation to use error handling to transfer control from one part of the program to another; use catch and throw instead. See Section 9.5.1 [Catch and Throw], page 134.

9.5.3.1 How to Signal an Error

Most errors are signaled “automatically” within Lisp primitives which you call for other purposes, such as if you try to take the car of an integer or move forward a character at the end of the buffer; you can also signal errors explicitly with the functions error and

signal.

Quitting, which happens when the user types C-g, is not considered an error, but it is handled almost like an error. See Section 19.8 [Quitting], page 306.

Function error format-string &rest args This function signals an error with an error message constructed by applying format (see Section 4.7 [String Conversion], page 65) to format-string and args. These examples show typical uses of error:

     (error "You have committed an error.
             Try something else.")
           error You have committed an error.
             Try something else.
     (error "You have committed %d errors." 10)
           error You have committed 10 errors.

error works by calling signal with two arguments: the error symbol error, and a list containing the string returned by format. If you want to use your own string as an error message verbatim, don’t just write (error string). If string contains ‘%’, it will be interpreted as a format specifier, with undesirable results. Instead, use (error "%s" string).

Function signal error-symbol data This function signals an error named by error-symbol. The argument data is a list of additional Lisp objects relevant to the circumstances of the error. The argument error-symbol must be an error symbol—a symbol bearing a property error-conditions whose value is a list of condition names. This is how XEmacs Lisp classifies different sorts of errors. The number and significance of the objects in data depends on error-symbol. For example, with a wrong-type-arg error, there are two objects in the list: a predicate that describes the type that was expected, and the object that failed to fit that type. See Section 9.5.3.4 [Error Symbols], page 140, for a description of error symbols. Both error-symbol and data are available to any error handlers that handle the error: condition-case binds a local variable to a list of the form (errorsymbol . data) (see Section 9.5.3.3 [Handling Errors], page 138). If the error is not handled, these two values are used in printing the error message. The function signal never returns (though in older Emacs versions it could sometimes return).

     (signal 'wrong-number-of-arguments '(x y))
           error Wrong number of arguments: x, y
     (signal 'no-such-error '("My unknown error condition."))
           error peculiar error: "My unknown error condition."

Common Lisp note: XEmacs Lisp has nothing like the Common Lisp concept of continuable errors.

9.5.3.2 How XEmacs Processes Errors

When an error is signaled, signal searches for an active handler for the error. A handler is a sequence of Lisp expressions designated to be executed if an error happens in part of the Lisp program. If the error has an applicable handler, the handler is executed, and control resumes following the handler. The handler executes in the environment of the condition-case that established it; all functions called within that condition-case have already been exited, and the handler cannot return to them.

If there is no applicable handler for the error, the current command is terminated and control returns to the editor command loop, because the command loop has an implicit handler for all kinds of errors. The command loop’s handler uses the error symbol and associated data to print an error message.

An error that has no explicit handler may call the Lisp debugger. The debugger is enabled if the variable debug-on-error (see Section 16.1.1 [Error Debugging], page 217) is non-nil. Unlike error handlers, the debugger runs in the environment of the error, so that you can examine values of variables precisely as they were at the time of the error.

9.5.3.3 Writing Code to Handle Errors

The usual effect of signaling an error is to terminate the command that is running and return immediately to the XEmacs editor command loop. You can arrange to trap errors occurring in a part of your program by establishing an error handler, with the special form condition-case. A simple example looks like this:

     (condition-case nil
         (delete-file filename)
       (error nil))

This deletes the file named filename, catching any error and returning nil if an error occurs.

The second argument of condition-case is called the protected form. (In the example above, the protected form is a call to delete-file.) The error handlers go into effect when this form begins execution and are deactivated when this form returns. They remain in effect for all the intervening time. In particular, they are in effect during the execution of functions called by this form, in their subroutines, and so on. This is a good thing, since, strictly speaking, errors can be signaled only by Lisp primitives (including signal and error) called by the protected form, not by the protected form itself.

The arguments after the protected form are handlers. Each handler lists one or more condition names (which are symbols) to specify which errors it will handle. The error symbol specified when an error is signaled also defines a list of condition names. A handler applies to an error if they have any condition names in common. In the example above, there is one handler, and it specifies one condition name, error, which covers all errors.

The search for an applicable handler checks all the established handlers starting with the most recently established one. Thus, if two nested condition-case forms offer to handle the same error, the inner of the two will actually handle it.

When an error is handled, control returns to the handler. Before this happens, XEmacs unbinds all variable bindings made by binding constructs that are being exited and executes the cleanups of all unwind-protect forms that are exited. Once control arrives at the handler, the body of the handler is executed.

After execution of the handler body, execution continues by returning from the conditioncase form. Because the protected form is exited completely before execution of the handler, the handler cannot resume execution at the point of the error, nor can it examine variable bindings that were made within the protected form. All it can do is clean up and proceed.

condition-case is often used to trap errors that are predictable, such as failure to open a file in a call to insert-file-contents. It is also used to trap errors that are totally unpredictable, such as when the program evaluates an expression read from the user.

Error signaling and handling have some resemblance to throw and catch, but they are entirely separate facilities. An error cannot be caught by a catch, and a throw cannot be handled by an error handler (though using throw when there is no suitable catch signals an error that can be handled).

Special Form condition-case var protected-form handlers… This special form establishes the error handlers handlers around the execution of protected-form. If protected-form executes without error, the value it returns becomes the value of the condition-case form; in this case, the conditioncase has no effect. The condition-case form makes a difference when an error occurs during protected-form. Each of the handlers is a list of the form (conditions body...). Here conditions is an error condition name to be handled, or a list of condition names; body is one or more Lisp expressions to be executed when this handler handles an error. Here are examples of handlers:

     (error nil)
     (arith-error (message "Division by zero"))
     ((arith-error file-error)
      (message
       "Either division by zero or failure to open a file"))

Each error that occurs has an error symbol that describes what kind of error it is. The error-conditions property of this symbol is a list of condition names (see Section 9.5.3.4 [Error Symbols], page 140). Emacs searches all the active condition-case forms for a handler that specifies one or more of these condition names; the innermost matching condition-case handles the error. Within this condition-case, the first applicable handler handles the error. After executing the body of the handler, the condition-case returns normally, using the value of the last form in the handler body as the overall value. The argument var is a variable. condition-case does not bind this variable when executing the protected-form, only when it handles an error. At that time, it binds var locally to a list of the form (error-symbol . data), giving the particulars of the error. The handler can refer to this list to decide what to do.

For example, if the error is for failure opening a file, the file name is the second element of data—the third element of var. If var is nil, that means no variable is bound. Then the error symbol and associated data are not available to the handler.

Here is an example of using condition-case to handle the error that results from dividing by zero. The handler prints out a warning message and returns a very large number.

     (defun safe-divide (dividend divisor)
       (condition-case err

;; Protected form.

           (/ dividend divisor)

;; The handler.

         (arith-error                        ; Condition.
          (princ (format "Arithmetic error: %s" err))
          1000000)))
     ⇒ safe-divide
     (safe-divide 5 0)
          ⊣ Arithmetic error: (arith-error)
     ⇒ 1000000

The handler specifies condition name arith-error so that it will handle only division-by-zero errors. Other kinds of errors will not be handled, at least not by this condition-case. Thus,

     (safe-divide nil 3)
           error Wrong type argument: integer-or-marker-p, nil

Here is a condition-case that catches all kinds of errors, including those signaled with

error:
     (setq baz 34) ⇒ 34
     (condition-case err
         (if (eq baz 35)
             t

;; This is a call to the function error.

           (error "Rats!  The variable %s was %s, not 35" 'baz baz))

;; This is the handler; it is not a form.

       (error (princ (format "The error was: %s" err))
              2))
     ⊣ The error was: (error "Rats!  The variable baz was 34, not 35")
     ⇒ 2

9.5.3.4 Error Symbols and Condition Names

When you signal an error, you specify an error symbol to specify the kind of error you have in mind. Each error has one and only one error symbol to categorize it. This is the finest classification of errors defined by the XEmacs Lisp language.

These narrow classifications are grouped into a hierarchy of wider classes called error conditions, identified by condition names. The narrowest such classes belong to the error symbols themselves: each error symbol is also a condition name. There are also condition names for more extensive classes, up to the condition name error which takes in all kinds of errors. Thus, each error has one or more condition names: error, the error symbol if that is distinct from error, and perhaps some intermediate classifications.

In order for a symbol to be an error symbol, it must have an error-conditions property which gives a list of condition names. This list defines the conditions that this kind of error belongs to. (The error symbol itself, and the symbol error, should always be members of this list.) Thus, the hierarchy of condition names is defined by the error-conditions properties of the error symbols.

In addition to the error-conditions list, the error symbol should have an errormessage property whose value is a string to be printed when that error is signaled but not handled. If the error-message property exists, but is not a string, the error message

‘peculiar error’ is used.

Here is how we define a new error symbol, new-error:

     (put 'new-error
          'error-conditions
          '(error my-own-errors new-error))
     ⇒ (error my-own-errors new-error)
     (put 'new-error 'error-message "A new error")
     ⇒ "A new error"

This error has three condition names: new-error, the narrowest classification; my-own-errors, which we imagine is a wider classification; and error, which is the widest of all.

The error string should start with a capital letter but it should not end with a period. This is for consistency with the rest of Emacs.

Naturally, XEmacs will never signal new-error on its own; only an explicit call to signal (see Section 9.5.3.1 [Signaling Errors], page 137) in your code can do this:

     (signal 'new-error '(x y))
           error A new error: x, y

This error can be handled through any of the three condition names. This example handles new-error and any other errors in the class my-own-errors:

     (condition-case foo
         (bar nil t)
       (my-own-errors nil))

The significant way that errors are classified is by their condition names—the names used to match errors with handlers. An error symbol serves only as a convenient way to specify the intended error message and list of condition names. It would be cumbersome to give signal a list of condition names rather than one error symbol.

By contrast, using only error symbols without condition names would seriously decrease the power of condition-case. Condition names make it possible to categorize errors at various levels of generality when you write an error handler. Using error symbols alone would eliminate all but the narrowest level of classification.

See Appendix C [Standard Errors], page 773, for a list of all the standard error symbols and their conditions.

9.5.4 Cleaning Up from Nonlocal Exits

The unwind-protect construct is essential whenever you temporarily put a data structure in an inconsistent state; it permits you to ensure the data are consistent in the event of an error or throw.

Special Form unwind-protect body cleanup-forms… unwind-protect executes the body with a guarantee that the cleanup-forms will be evaluated if control leaves body, no matter how that happens. The body may complete normally, or execute a throw out of the unwind-protect, or cause an error; in all cases, the cleanup-forms will be evaluated.

If the body forms finish normally, unwind-protect returns the value of the last body form, after it evaluates the cleanup-forms. If the body forms do not finish, unwind-protect does not return any value in the normal sense.

Only the body is actually protected by the unwind-protect. If any of the cleanup-forms themselves exits nonlocally (e.g., via a throw or an error), unwind-protect is not guaranteed to evaluate the rest of them. If the failure of one of the cleanup-forms has the potential to cause trouble, then protect it with another unwind-protect around that form.

The number of currently active unwind-protect forms counts, together with the number of local variable bindings, against the limit max-specpdl-size (see Section 10.3 [Local Variables], page 146).

For example, here we make an invisible buffer for temporary use, and make sure to kill it before finishing:

(save-excursion
  (let ((buffer (get-buffer-create " *temp*")))
    (set-buffer buffer)
    (unwind-protect

body

      (kill-buffer buffer))))

You might think that we could just as well write (kill-buffer (current-buffer)) and dispense with the variable buffer. However, the way shown above is safer, if body happens to get an error after switching to a different buffer! (Alternatively, you could write another save-excursion around the body, to ensure that the temporary buffer becomes current in time to kill it.)

Here is an actual example taken from the file ‘ftp.el’. It creates a process (see Chapter 49 [Processes], page 673) to try to establish a connection to a remote machine. As the function ftp-login is highly susceptible to numerous problems that the writer of the function cannot anticipate, it is protected with a form that guarantees deletion of the process in the event of failure. Otherwise, XEmacs might fill up with useless subprocesses.

     (let ((win nil))
       (unwind-protect
           (progn
             (setq process (ftp-setup-buffer host file))
             (if (setq win (ftp-login process host user password))
                 (message "Logged in")
               (error "Ftp login failed")))
         (or win (and process (delete-process process)))))

This example actually has a small bug: if the user types C-g to quit, and the quit happens immediately after the function ftp-setup-buffer returns but before the variable process is set, the process will not be killed. There is no easy way to fix this bug, but at least it is very unlikely.

Here is another example which uses unwind-protect to make sure to kill a temporary buffer. In this example, the value returned by unwind-protect is used.

     (defun shell-command-string (cmd)
       "Return the output of the shell command CMD, as a string."
       (save-excursion
         (set-buffer (generate-new-buffer " OS*cmd"))
         (shell-command cmd t)
         (unwind-protect
             (buffer-string)
           (kill-buffer (current-buffer)))))

144 XEmacs Lisp Reference Manual

10 Variables

A variable is a name used in a program to stand for a value. Nearly all programming languages have variables of some sort. In the text of a Lisp program, variables are written using the syntax for symbols.

In Lisp, unlike most programming languages, programs are represented primarily as Lisp objects and only secondarily as text. The Lisp objects used for variables are symbols: the symbol name is the variable name, and the variable’s value is stored in the value cell of the symbol. The use of a symbol as a variable is independent of its use as a function name. See Section 7.1 [Symbol Components], page 111.

The Lisp objects that constitute a Lisp program determine the textual form of the program—it is simply the read syntax for those Lisp objects. This is why, for example, a variable in a textual Lisp program is written using the read syntax for the symbol that represents the variable.

10.1 Global Variables

The simplest way to use a variable is globally. This means that the variable has just one value at a time, and this value is in effect (at least for the moment) throughout the Lisp system. The value remains in effect until you specify a new one. When a new value replaces the old one, no trace of the old value remains in the variable.

You specify a value for a symbol with setq. For example,

     (setq x '(a b))

gives the variable x the value (a b). Note that setq does not evaluate its first argument, the name of the variable, but it does evaluate the second argument, the new value.

Once the variable has a value, you can refer to it by using the symbol by itself as an expression. Thus,

     x ⇒ (a b)

assuming the setq form shown above has already been executed.

If you do another setq, the new value replaces the old one:

     x  ⇒ (a b)
     (setq x 4) ⇒ 4
     x  ⇒ 4

10.2 Variables That Never Change

In XEmacs Lisp, some symbols always evaluate to themselves: the two special symbols nil and t, as well as keyword symbols, that is, symbols whose name begins with the character ‘:’. These symbols cannot be rebound, nor can their value cells be changed. An attempt to change the value of nil or t signals a setting-constant error.

     nil ≡ 'nil ⇒ nil
     (setq nil 500)
      error Attempt to set constant symbol: nil

10.3 Local Variables

Global variables have values that last until explicitly superseded with new values. Sometimes it is useful to create variable values that exist temporarily—only while within a certain part of the program. These values are called local, and the variables so used are called local variables.

For example, when a function is called, its argument variables receive new local values that last until the function exits. The let special form explicitly establishes new local values for specified variables; these last until exit from the let form.

Establishing a local value saves away the previous value (or lack of one) of the variable. When the life span of the local value is over, the previous value is restored. In the mean time, we say that the previous value is shadowed and not visible. Both global and local values may be shadowed (see Section 10.8.1 [Scope], page 154).

If you set a variable (such as with setq) while it is local, this replaces the local value; it does not alter the global value, or previous local values that are shadowed. To model this behavior, we speak of a local binding of the variable as well as a local value.

The local binding is a conceptual place that holds a local value. Entry to a function, or a special form such as let, creates the local binding; exit from the function or from the let removes the local binding. As long as the local binding lasts, the variable’s value is stored within it. Use of setq or set while there is a local binding stores a different value into the local binding; it does not create a new binding.

We also speak of the global binding, which is where (conceptually) the global value is kept.

A variable can have more than one local binding at a time (for example, if there are nested let forms that bind it). In such a case, the most recently created local binding that still exists is the current binding of the variable. (This is called dynamic scoping; see Section 10.8 [Variable Scoping], page 154.) If there are no local bindings, the variable’s global binding is its current binding. We also call the current binding the most-local existing binding, for emphasis. Ordinary evaluation of a symbol always returns the value of its current binding.

The special forms let and let* exist to create local bindings.

Special Form let (bindings…) forms… This special form binds variables according to bindings and then evaluates all of the forms in textual order. The let-form returns the value of the last form in forms. Each of the bindings is either (i) a symbol, in which case that symbol is bound to nil; or (ii) a list of the form (symbol value-form), in which case symbol is bound to the result of evaluating value-form. If value-form is omitted, nil is used.

All of the value-forms in bindings are evaluated in the order they appear and before any of the symbols are bound. Here is an example of this: Z is bound to the old value of Y, which is 2, not the new value, 1.

     (setq Y 2) ⇒ 2
     (let ((Y 1) (Z Y)) (list Y Z)) ⇒ (1 2)

Special Form let* (bindings…) forms… This special form is like let, but it binds each variable right after computing its local value, before computing the local value for the next variable. Therefore, an expression in bindings can reasonably refer to the preceding symbols bound in this let* form. Compare the following example with the example above for

let.
     (setq Y 2) ⇒ 2
     (let* ((Y 1)

(Z Y)) ; Use the just-established value of Y.

       (list Y Z))
          ⇒ (1 1)

Here is a complete list of the other facilities that create local bindings:

• Function calls (see Chapter 11 [Functions], page 163).

• Macro calls (see Chapter 12 [Macros], page 177).

• condition-case (see Section 9.5.3 [Errors], page 136).

Variables can also have buffer-local bindings (see Section 10.9 [Buffer-Local Variables], page 157). These kinds of bindings work somewhat like ordinary local bindings, but they are localized depending on “where” you are in Emacs, rather than localized in time.

Variable max-specpdl-size

This variable defines the limit on the total number of local variable bindings and unwind-protect cleanups (see Section 9.5 [Nonlocal Exits], page 134) that are allowed before signaling an error (with data "Variable binding depth

exceeds max-specpdl-size").

This limit, with the associated error when it is exceeded, is one way that Lisp avoids infinite recursion on an ill-defined function.

The default value is 600.

max-lisp-eval-depth provides another limit on depth of nesting. See Section 8.1 [Eval], page 120.

10.4 When a Variable is “Void”

If you have never given a symbol any value as a global variable, we say that that symbol’s global value is void. In other words, the symbol’s value cell does not have any Lisp object in it. If you try to evaluate the symbol, you get a void-variable error rather than a value.

Note that a value of nil is not the same as void. The symbol nil is a Lisp object and can be the value of a variable just as any other object can be; but it is a value. A void variable does not have any value.

After you have given a variable a value, you can make it void once more using

makunbound.

Function makunbound symbol

This function makes the current binding of symbol void. Subsequent attempts to use this symbol’s value as a variable will signal the error void-variable, unless or until you set it again. makunbound returns symbol.

(makunbound 'x) ; Make the global value

; of x void.

          ⇒ x
     x
      error Symbol's value as variable is void: x

If symbol is locally bound, makunbound affects the most local existing binding. This is the only way a symbol can have a void local binding, since all the constructs that create local bindings create them with values. In this case, the voidness lasts at most as long as the binding does; when the binding is removed due to exit from the construct that made it, the previous or global binding is reexposed as usual, and the variable is no longer void unless the newly reexposed binding was void all along.

(setq x 1) ; Put a value in the global binding.

          ⇒ 1

(let ((x 2)) ; Locally bind it.

(makunbound 'x) ; Void the local binding.

       x)
      error Symbol's value as variable is void: x

x ; The global binding is unchanged.

          ⇒ 1

(let ((x 2)) ; Locally bind it.

       (let ((x 3))           ; And again.

(makunbound 'x) ; Void the innermost-local binding. x)) ; And refer: it’s void.

      error Symbol's value as variable is void: x
     (let ((x 2))
       (let ((x 3))

(makunbound 'x)) ; Void inner binding, then remove it.

x) ; Now outer let binding is visible.

          ⇒ 2

A variable that has been made void with makunbound is indistinguishable from one that has never received a value and has always been void.

You can use the function boundp to test whether a variable is currently void.

Function boundp variable boundp returns t if variable (a symbol) is not void; more precisely, if its current binding is not void. It returns nil otherwise.

          (boundp 'abracadabra)          ; Starts out void.
               ⇒ nil
          (let ((abracadabra 5))         ; Locally bind it.
            (boundp 'abracadabra))
               ⇒ t
          (boundp 'abracadabra)          ; Still globally void.
               ⇒ nil

(setq abracadabra 5) ; Make it globally nonvoid.

               ⇒ 5
          (boundp 'abracadabra)
               ⇒ t

10.5 Defining Global Variables

You may announce your intention to use a symbol as a global variable with a variable definition: a special form, either defconst or defvar.

In XEmacs Lisp, definitions serve three purposes. First, they inform people who read the code that certain symbols are intended to be used a certain way (as variables). Second, they inform the Lisp system of these things, supplying a value and documentation. Third, they provide information to utilities such as etags and make-docfile, which create data bases of the functions and variables in a program.

The difference between defconst and defvar is primarily a matter of intent, serving to inform human readers of whether programs will change the variable. XEmacs Lisp does not restrict the ways in which a variable can be used based on defconst or defvar declarations. However, it does make a difference for initialization: defconst unconditionally initializes the variable, while defvar initializes it only if it is void.

One would expect user option variables to be defined with defconst, since programs do not change them. Unfortunately, this has bad results if the definition is in a library that is not preloaded: defconst would override any prior value when the library is loaded. Users would like to be able to set user options in their init files, and override the default values given in the definitions. For this reason, user options must be defined with defvar.

Special Form defvar symbol [value [doc-string]] This special form defines symbol as a value and initializes it. The definition informs a person reading your code that symbol is used as a variable that programs are likely to set or change. It is also used for all user option variables except in the preloaded parts of XEmacs. Note that symbol is not evaluated; the symbol to be defined must appear explicitly in the defvar.

If symbol already has a value (i.e., it is not void), value is not even evaluated, and symbol’s value remains unchanged. If symbol is void and value is specified, defvar evaluates it and sets symbol to the result. (If value is omitted, the value of symbol is not changed in any case.) When you evaluate a top-level defvar form with C-M-x in Emacs Lisp mode (eval-defun), a special feature of eval-defun evaluates it as a defconst. The purpose of this is to make sure the variable’s value is reinitialized, when you ask for it specifically. If symbol has a buffer-local binding in the current buffer, defvar sets the default value, not the local value. See Section 10.9 [Buffer-Local Variables], page 157. If the doc-string argument appears, it specifies the documentation for the variable. (This opportunity to specify documentation is one of the main benefits of defining the variable.) The documentation is stored in the symbol’s variabledocumentation property. The XEmacs help functions (see Chapter 27 [Documentation], page 381) look for this property. If the first character of doc-string is ‘*’, it means that this variable is considered a user option. This lets users set the variable conveniently using the commands

set-variable and edit-options.

For example, this form defines foo but does not set its value:

     (defvar foo)
          ⇒ foo

The following example sets the value of bar to 23, and gives it a documentation string:

     (defvar bar 23
       "The normal weight of a bar.")
          ⇒ bar

The following form changes the documentation string for bar, making it a user option, but does not change the value, since bar already has a value. (The addition (1+ 23) is not even performed.)

     (defvar bar (1+ 23)
       "*The normal weight of a bar.")
          ⇒ bar
     bar ⇒ 23

Here is an equivalent expression for the defvar special form:

(defvar symbol value doc-string)

≡

     (progn
       (if (not (boundp 'symbol))

(setq symbol value))

       (put 'symbol 'variable-documentation 'doc-string)

'symbol) The defvar form returns symbol, but it is normally used at top level in a file where its value does not matter.

Special Form defconst symbol [value [doc-string]] This special form defines symbol as a value and initializes it. It informs a person reading your code that symbol has a global value, established here, that will not normally be changed or locally bound by the execution of the program. The user, however, may be welcome to change it. Note that symbol is not evaluated; the symbol to be defined must appear explicitly in the defconst. defconst always evaluates value and sets the global value of symbol to the result, provided value is given. If symbol has a buffer-local binding in the current buffer, defconst sets the default value, not the local value. Please note: Don’t use defconst for user option variables in libraries that are not standardly preloaded. The user should be able to specify a value for such a variable in the ‘.emacs’ file, so that it will be in effect if and when the library is loaded later. Here, pi is a constant that presumably ought not to be changed by anyone (attempts by the Indiana State Legislature notwithstanding). As the second form illustrates, however, this is only advisory.

     (defconst pi 3.1415 "Pi to five places.")
          ⇒ pi
     (setq pi 3) ⇒ pi
     pi ⇒ 3

Function user-variable-p variable

This function returns t if variable is a user option—a variable intended to be set by the user for customization—and nil otherwise. (Variables other than user options exist for the internal purposes of Lisp programs, and users need not know about them.) User option variables are distinguished from other variables by the first character of the variable-documentation property. If the property exists and is a string, and its first character is ‘*’, then the variable is a user option.

If a user option variable has a variable-interactive property, the set-variable command uses that value to control reading the new value for the variable. The property’s value is used as if it were the argument to interactive.

Warning: If the defconst and defvar special forms are used while the variable has a local binding, they set the local binding’s value; the global binding is not changed. This is not what we really want. To prevent it, use these special forms at top level in a file, where normally no local binding is in effect, and make sure to load the file before making a local binding for the variable.

10.6 Accessing Variable Values

The usual way to reference a variable is to write the symbol which names it (see Section 8.2.2 [Symbol Forms], page 122). This requires you to specify the variable name when you write the program. Usually that is exactly what you want to do. Occasionally you need to choose at run time which variable to reference; then you can use symbol-value.

Function symbol-value symbol

This function returns the value of symbol. This is the value in the innermost local binding of the symbol, or its global value if it has no local bindings.

(setq abracadabra 5)
     ⇒ 5
(setq foo 9) ⇒ 9
;; Here the symbol abracadabra

;; is the symbol whose value is examined.

(let ((abracadabra 'foo))
  (symbol-value 'abracadabra))
     ⇒ foo

;; Here the value of abracadabra, ;; which is foo, ;; is the symbol whose value is examined.

(let ((abracadabra 'foo))
  (symbol-value abracadabra))
     ⇒ 9
(symbol-value 'abracadabra)
     ⇒ 5

A void-variable error is signaled if symbol has neither a local binding nor a global value.

10.7 How to Alter a Variable Value

The usual way to change the value of a variable is with the special form setq. When you need to compute the choice of variable at run time, use the function set.

Special Form setq [symbol form]… This special form is the most common method of changing a variable’s value. Each symbol is given a new value, which is the result of evaluating the corresponding form. The most-local existing binding of the symbol is changed. setq does not evaluate symbol; it sets the symbol that you write. We say that this argument is automatically quoted. The ‘q’ in setq stands for “quoted.” The value of the setq form is the value of the last form.

(setq x (1+ 2)) ⇒ 3

x ; x now has a global value.

     ⇒ 3
(let ((x 5))

(setq x 6) ; The local binding of x is set.

  x)
     ⇒ 6

x ; The global value is unchanged.

     ⇒ 3

Note that the first form is evaluated, then the first symbol is set, then the second form is evaluated, then the second symbol is set, and so on:

(setq x 10 ; Notice that x is set before y (1+ x)) ; the value of y is computed.

     ⇒ 11

Function set symbol value This function sets symbol’s value to value, then returns value. Since set is a function, the expression written for symbol is evaluated to obtain the symbol to set. The most-local existing binding of the variable is the binding that is set; shadowed bindings are not affected.

(set one 1)
 error Symbol's value as variable is void: one
(set 'one 1) ⇒ 1
(set 'two 'one)
     ⇒ one

(set two 2) ; two evaluates to symbol one.

     ⇒ 2

one ; So it is one that was set.

     ⇒ 2

(let ((one 1)) ; This binding of one is set, (set 'one 3) ; not the global value.

  one)
     ⇒ 3
one ⇒ 2

If symbol is not actually a symbol, a wrong-type-argument error is signaled.

(set '(x y) 'z)
 error Wrong type argument: symbolp, (x y)

Logically speaking, set is a more fundamental primitive than setq. Any use of setq can be trivially rewritten to use set; setq could even be defined as a macro, given the availability of set. However, set itself is rarely used; beginners hardly need to know about it. It is useful only for choosing at run time which variable to set. For example, the command set-variable, which reads a variable name from the user and then sets the variable, needs to use set.

Common Lisp note: In Common Lisp, set always changes the symbol’s special value, ignoring any lexical bindings. In XEmacs Lisp, all variables and all bindings are (in effect) special, so set always affects the most local existing binding.

One other function for setting a variable is designed to add an element to a list if it is not already present in the list.

Function add-to-list symbol element This function sets the variable symbol by consing element onto the old value, if element is not already a member of that value. It returns the resulting list, whether updated or not. The value of symbol had better be a list already before the call. The argument symbol is not implicitly quoted; add-to-list is an ordinary function, like set and unlike setq. Quote the argument yourself if that is what you want. Here’s a scenario showing how to use add-to-list:

          (setq foo '(a b))
               ⇒ (a b)
          (add-to-list 'foo 'c)     ;; Add c.
               ⇒ (c a b)
          (add-to-list 'foo 'b)     ;; No effect.
               ⇒ (c a b)

foo ;; foo was changed.

               ⇒ (c a b)

An equivalent expression for (add-to-list 'var value) is this:

     (or (member value var)
         (setq var (cons value var)))

10.8 Scoping Rules for Variable Bindings

A given symbol foo may have several local variable bindings, established at different places in the Lisp program, as well as a global binding. The most recently established binding takes precedence over the others.

Local bindings in XEmacs Lisp have indefinite scope and dynamic extent. Scope refers to where textually in the source code the binding can be accessed. Indefinite scope means that any part of the program can potentially access the variable binding. Extent refers to when, as the program is executing, the binding exists. Dynamic extent means that the binding lasts as long as the activation of the construct that established it.

The combination of dynamic extent and indefinite scope is called dynamic scoping. By contrast, most programming languages use lexical scoping, in which references to a local variable must be located textually within the function or block that binds the variable.

Common Lisp note: Variables declared “special” in Common Lisp are dynamically scoped, like variables in XEmacs Lisp.

10.8.1 Scope

XEmacs Lisp uses indefinite scope for local variable bindings. This means that any function anywhere in the program text might access a given binding of a variable. Consider the following function definitions:

     (defun binder (x)   ; x is bound in binder.

(foo 5)) ; foo is some other function.

     (defun user ()      ; x is used in user.
       (list x))

In a lexically scoped language, the binding of x in binder would never be accessible in user, because user is not textually contained within the function binder. However, in dynamically scoped XEmacs Lisp, user may or may not refer to the binding of x established in binder, depending on circumstances:

• If we call user directly without calling binder at all, then whatever binding of x is found, it cannot come from binder.

• If we define foo as follows and call binder, then the binding made in binder will be

seen in user:

         (defun foo (lose)
           (user))

• If we define foo as follows and call binder, then the binding made in binder will not

be seen in user:

         (defun foo (x)
           (user))

Here, when foo is called by binder, it binds x. (The binding in foo is said to shadow the one made in binder.) Therefore, user will access the x bound by foo instead of the one bound by binder.

10.8.2 Extent

Extent refers to the time during program execution that a variable name is valid. In XEmacs Lisp, a variable is valid only while the form that bound it is executing. This is called dynamic extent. “Local” or “automatic” variables in most languages, including C and Pascal, have dynamic extent.

One alternative to dynamic extent is indefinite extent. This means that a variable binding can live on past the exit from the form that made the binding. Common Lisp and Scheme, for example, support this, but XEmacs Lisp does not.

To illustrate this, the function below, make-add, returns a function that purports to add n to its own argument m. This would work in Common Lisp, but it does not work as intended in XEmacs Lisp, because after the call to make-add exits, the variable n is no longer bound to the actual argument 2.

     (defun make-add (n)
         (function (lambda (m) (+ n m))))  ; Return a function.
          ⇒ make-add
     (fset 'add2 (make-add 2))  ; Define function add2
                                ; with (make-add 2).
          ⇒ (lambda (m) (+ n m))

(add2 4) ; Try to add 2 to 4.

      error Symbol's value as variable is void: n

Some Lisp dialects have “closures”, objects that are like functions but record additional variable bindings. XEmacs Lisp does not have closures.

10.8.3 Implementation of Dynamic Scoping

A simple sample implementation (which is not how XEmacs Lisp actually works) may help you understand dynamic binding. This technique is called deep binding and was used in early Lisp systems.

Suppose there is a stack of bindings: variable-value pairs. At entry to a function or to a let form, we can push bindings on the stack for the arguments or local variables created there. We can pop those bindings from the stack at exit from the binding construct.

We can find the value of a variable by searching the stack from top to bottom for a binding for that variable; the value from that binding is the value of the variable. To set the variable, we search for the current binding, then store the new value into that binding.

As you can see, a function’s bindings remain in effect as long as it continues execution, even during its calls to other functions. That is why we say the extent of the binding is dynamic. And any other function can refer to the bindings, if it uses the same variables while the bindings are in effect. That is why we say the scope is indefinite.

The actual implementation of variable scoping in XEmacs Lisp uses a technique called shallow binding. Each variable has a standard place in which its current value is always found—the value cell of the symbol.

In shallow binding, setting the variable works by storing a value in the value cell. Creating a new binding works by pushing the old value (belonging to a previous binding) on a stack, and storing the local value in the value cell. Eliminating a binding works by popping the old value off the stack, into the value cell.

We use shallow binding because it has the same results as deep binding, but runs faster, since there is never a need to search for a binding.

10.8.4 Proper Use of Dynamic Scoping

Binding a variable in one function and using it in another is a powerful technique, but if used without restraint, it can make programs hard to understand. There are two clean ways to use this technique:

• Use or bind the variable only in a few related functions, written close together in one file. Such a variable is used for communication within one program. You should write comments to inform other programmers that they can see all uses of the variable before them, and to advise them not to add uses elsewhere.

• Give the variable a well-defined, documented meaning, and make all appropriate functions refer to it (but not bind it or set it) wherever that meaning is relevant. For example, the variable case-fold-search is defined as “non-nil means ignore case when searching”; various search and replace functions refer to it directly or through their subroutines, but do not bind or set it. Then you can bind the variable in other programs, knowing reliably what the effect will be.

In either case, you should define the variable with defvar. This helps other people understand your program by telling them to look for inter-function usage. It also avoids a warning from the byte compiler. Choose the variable’s name to avoid name conflicts—don’t use short names like x.

10.9 Buffer-Local Variables

Global and local variable bindings are found in most programming languages in one form or another. XEmacs also supports another, unusual kind of variable binding: buffer-local bindings, which apply only to one buffer. XEmacs Lisp is meant for programming editing commands, and having different values for a variable in different buffers is an important customization method.

10.9.1 Introduction to Buffer-Local Variables

A buffer-local variable has a buffer-local binding associated with a particular buffer. The binding is in effect when that buffer is current; otherwise, it is not in effect. If you set the variable while a buffer-local binding is in effect, the new value goes in that binding, so the global binding is unchanged; this means that the change is visible in that buffer alone.

A variable may have buffer-local bindings in some buffers but not in others. The global binding is shared by all the buffers that don’t have their own bindings. Thus, if you set the variable in a buffer that does not have a buffer-local binding for it, the new value is visible in all buffers except those with buffer-local bindings. (Here we are assuming that there are no let-style local bindings to complicate the issue.)

The most common use of buffer-local bindings is for major modes to change variables that control the behavior of commands. For example, C mode and Lisp mode both set the variable paragraph-start to specify that only blank lines separate paragraphs. They do this by making the variable buffer-local in the buffer that is being put into C mode or Lisp mode, and then setting it to the new value for that mode.

The usual way to make a buffer-local binding is with make-local-variable, which is what major mode commands use. This affects just the current buffer; all other buffers (including those yet to be created) continue to share the global value.

A more powerful operation is to mark the variable as automatically buffer-local by calling make-variable-buffer-local. You can think of this as making the variable local in all buffers, even those yet to be created. More precisely, the effect is that setting the variable automatically makes the variable local to the current buffer if it is not already so. All buffers start out by sharing the global value of the variable as usual, but any setq creates a buffer-local binding for the current buffer. The new value is stored in the buffer-local binding, leaving the (default) global binding untouched. The global value can no longer be changed with setq; you need to use setq-default to do that.

Local variables in a file you edit are also represented by buffer-local bindings for the buffer that holds the file within XEmacs. See Section 26.1.3 [Auto Major Mode], page 367.

10.9.2 Creating and Deleting Buffer-Local Bindings

Command make-local-variable variable

This function creates a buffer-local binding in the current buffer for variable (a symbol). Other buffers are not affected. The value returned is variable. The buffer-local value of variable starts out as the same value variable previously had. If variable was void, it remains void.

;; In buffer ‘b1’:

(setq foo 5) ; Affects all buffers.

          ⇒ 5
     (make-local-variable 'foo)  ; Now it is local in ‘b1’.
          ⇒ foo

foo ; That did not change ⇒ 5 ; the value. (setq foo 6) ; Change the value ⇒ 6 ; in ‘b1’.

     foo ⇒ 6

;; In buffer ‘b2’, the value hasn’t changed.

     (save-excursion
       (set-buffer "b2")
       foo)
          ⇒ 5

Making a variable buffer-local within a let-binding for that variable does not work. This is because let does not distinguish between different kinds of bindings; it knows only which variable the binding was made for. Please note: do not use make-local-variable for a hook variable. Instead, use make-local-hook. See Section 26.4 [Hooks], page 378.

Command make-variable-buffer-local variable

This function marks variable (a symbol) automatically buffer-local, so that any subsequent attempt to set it will make it local to the current buffer at the time. The value returned is variable.

Function local-variable-p variable &optional buffer This returns t if variable is buffer-local in buffer buffer (which defaults to the current buffer); otherwise, nil.

Function buffer-local-variables &optional buffer

This function returns a list describing the buffer-local variables in buffer buffer. It returns an association list (see Section 5.8 [Association Lists], page 92) in which each association contains one buffer-local variable and its value. When a buffer-local variable is void in buffer, then it appears directly in the resulting list. If buffer is omitted, the current buffer is used.

     (make-local-variable 'foobar)
     (makunbound 'foobar)
     (make-local-variable 'bind-me)
     (setq bind-me 69)
     (setq lcl (buffer-local-variables))

;; First, built-in variables local in all buffers:

     ⇒ ((mark-active . nil)
         (buffer-undo-list nil)
         (mode-name . "Fundamental")
         ...

;; Next, non-built-in local variables. ;; This one is local and void:

         foobar

;; This one is local and nonvoid:

         (bind-me . 69))

Note that storing new values into the cdrs of cons cells in this list does not change the local values of the variables.

Command kill-local-variable variable

This function deletes the buffer-local binding (if any) for variable (a symbol) in the current buffer. As a result, the global (default) binding of variable becomes visible in this buffer. Usually this results in a change in the value of variable, since the global value is usually different from the buffer-local value just eliminated. If you kill the local binding of a variable that automatically becomes local when set, this makes the global value visible in the current buffer. However, if you set the variable again, that will once again create a local binding for it.

kill-local-variable returns variable.

This function is a command because it is sometimes useful to kill one bufferlocal variable interactively, just as it is useful to create buffer-local variables interactively.

Function kill-all-local-variables

This function eliminates all the buffer-local variable bindings of the current buffer except for variables marked as “permanent”. As a result, the buffer will see the default values of most variables. This function also resets certain other information pertaining to the buffer: it sets the local keymap to nil, the syntax table to the value of standard-syntax-

table, and the abbrev table to the value of fundamental-mode-abbrev-table.

Every major mode command begins by calling this function, which has the effect of switching to Fundamental mode and erasing most of the effects of the previous major mode. To ensure that this does its job, the variables that major modes set should not be marked permanent.

kill-all-local-variables returns nil.

A local variable is permanent if the variable name (a symbol) has a permanent-local property that is non-nil. Permanent locals are appropriate for data pertaining to where the file came from or how to save it, rather than with how to edit the contents.

10.9.3 The Default Value of a Buffer-Local Variable

The global value of a variable with buffer-local bindings is also called the default value, because it is the value that is in effect except when specifically overridden.

The functions default-value and setq-default access and change a variable’s default value regardless of whether the current buffer has a buffer-local binding. For example, you could use setq-default to change the default setting of paragraph-start for most buffers; and this would work even when you are in a C or Lisp mode buffer that has a buffer-local value for this variable.

The special forms defvar and defconst also set the default value (if they set the variable at all), rather than any local value.

Function default-value symbol

This function returns symbol’s default value. This is the value that is seen in buffers that do not have their own values for this variable. If symbol is not buffer-local, this is equivalent to symbol-value (see Section 10.6 [Accessing Variables], page 151).

Function default-boundp symbol

The function default-boundp tells you whether symbol’s default value is non-

void. If (default-boundp 'foo) returns nil, then (default-value 'foo)

would get an error.

default-boundp is to default-value as boundp is to symbol-value.

Special Form setq-default symbol value

This sets the default value of symbol to value. It does not evaluate symbol, but does evaluate value. The value of the setq-default form is value.

If a symbol is not buffer-local for the current buffer, and is not marked automatically buffer-local, setq-default has the same effect as setq. If symbol is buffer-local for the current buffer, then this changes the value that other buffers will see (as long as they don’t have a buffer-local value), but not the value that the current buffer sees.

;; In buffer ‘foo’:

     (make-local-variable 'local)
          ⇒ local
     (setq local 'value-in-foo)
          ⇒ value-in-foo
     (setq-default local 'new-default)
          ⇒ new-default
     local
          ⇒ value-in-foo
     (default-value 'local)
          ⇒ new-default

;; In (the new) buffer ‘bar’:

     local
          ⇒ new-default
     (default-value 'local)
          ⇒ new-default
     (setq local 'another-default)
          ⇒ another-default
     (default-value 'local)
          ⇒ another-default

;; Back in buffer ‘foo’:

     local
          ⇒ value-in-foo
     (default-value 'local)
          ⇒ another-default

Function set-default symbol value

This function is like setq-default, except that symbol is evaluated.

     (set-default (car '(a b c)) 23)
          ⇒ 23
     (default-value 'a)
          ⇒ 23

10.10 Variable Aliases

You can define a variable as an alias for another. Any time you reference the former variable, the current value of the latter is returned. Any time you change the value of the former variable, the value of the latter is actually changed. This is useful in cases where you want to rename a variable but still make old code work (see Section 27.6 [Obsoleteness], page 389).

Function defvaralias variable alias This function defines variable as an alias for alias. Thenceforth, any operations performed on variable will actually be performed on alias. Both variable and alias should be symbols. If alias is nil, remove any aliases for variable. alias can itself be aliased, and the chain of variable aliases will be followed appropriately. If variable already has a value, this value will be shadowed until the alias is removed, at which point it will be restored. Currently variable cannot be a built-in variable, a variable that has a buffer-local value in any buffer, or the symbols nil or t.

Function variable-alias variable

If variable is aliased to another variable, this function returns that variable. variable should be a symbol. If variable is not aliased, this function returns

nil.

Function indirect-variable object

This function returns the variable at the end of object’s variable-alias chain. If object is a symbol, follow all variable aliases and return the final (non-aliased) symbol. If object is not a symbol, just return it. Signal a cyclic-variable-indirection error if there is a loop in the variable chain of symbols.

162 XEmacs Lisp Reference Manual

11 Functions

A Lisp program is composed mainly of Lisp functions. This chapter explains what functions are, how they accept arguments, and how to define them.

11.1 What Is a Function?

In a general sense, a function is a rule for carrying on a computation given several values called arguments. The result of the computation is called the value of the function. The computation can also have side effects: lasting changes in the values of variables or the contents of data structures.

Here are important terms for functions in XEmacs Lisp and for other function-like objects.

function In XEmacs Lisp, a function is anything that can be applied to arguments in a Lisp program. In some cases, we use it more specifically to mean a function written in Lisp. Special forms and macros are not functions.

primitive A primitive is a function callable from Lisp that is written in C, such as car or append. These functions are also called built-in functions or subrs. (Special forms are also considered primitives.) Usually the reason that a function is a primitives is because it is fundamental, because it provides a low-level interface to operating system services, or because it needs to run fast. Primitives can be modified or added only by changing the C sources and recompiling the editor. See section “Writing Lisp Primitives” in XEmacs Internals Manual.

lambda expression A lambda expression is a function written in Lisp. These are described in the following section.

special form A special form is a primitive that is like a function but does not evaluate all of its arguments in the usual way. It may evaluate only some of the arguments, or may evaluate them in an unusual order, or several times. Many special forms are described in Chapter 9 [Control Structures], page 129.

macro A macro is a construct defined in Lisp by the programmer. It differs from a function in that it translates a Lisp expression that you write into an equivalent expression to be evaluated instead of the original expression. Macros enable Lisp programmers to do the sorts of things that special forms can do. See Chapter 12 [Macros], page 177, for how to define and use macros.

command A command is an object that command-execute can invoke; it is a possible definition for a key sequence. Some functions are commands; a function written in Lisp is a command if it contains an interactive declaration (see Section 19.2 [Defining Commands], page 282). Such a function can be called from Lisp expressions like other functions; in this case, the fact that the function is a command makes no difference.

Keyboard macros (strings and vectors) are commands also, even though they are not functions. A symbol is a command if its function definition is a command; such symbols can be invoked with M-x. The symbol is a function as well if the definition is a function. See Section 19.1 [Command Overview], page 281.

keystroke command A keystroke command is a command that is bound to a key sequence (typically one to three keystrokes). The distinction is made here merely to avoid confusion with the meaning of “command” in non-Emacs editors; for Lisp programs, the distinction is normally unimportant.

compiled function A compiled function is a function that has been compiled by the byte compiler. See Section 2.4.14 [Compiled-Function Type], page 28.

Function subrp object

This function returns t if object is a built-in function (i.e., a Lisp primitive).

          (subrp 'message)            ; message is a symbol,

⇒ nil ; not a subr object.

          (subrp (symbol-function 'message))
               ⇒ t

Function compiled-function-p object

This function returns t if object is a compiled function. For example:

          (compiled-function-p (symbol-function 'next-line))
               ⇒ t

11.2 Lambda Expressions

A function written in Lisp is a list that looks like this:

(lambda (arg-variables...) [documentation-string] [interactive-declaration] body-forms...)

Such a list is called a lambda expression. In XEmacs Lisp, it actually is valid as an expression—it evaluates to itself. In some other Lisp dialects, a lambda expression is not a valid expression at all. In either case, its main use is not to be evaluated as an expression, but to be called as a function.

11.2.1 Components of a Lambda Expression

The first element of a lambda expression is always the symbol lambda. This indicates that the list represents a function. The reason functions are defined to start with lambda is so that other lists, intended for other uses, will not accidentally be valid as functions.

The second element is a list of symbols–the argument variable names. This is called the lambda list. When a Lisp function is called, the argument values are matched up against the variables in the lambda list, which are given local bindings with the values provided. See Section 10.3 [Local Variables], page 146.

The documentation string is a Lisp string object placed within the function definition to describe the function for the XEmacs help facilities. See Section 11.2.4 [Function Documentation], page 167.

The interactive declaration is a list of the form (interactive code-string). This declares how to provide arguments if the function is used interactively. Functions with this declaration are called commands; they can be called using M-x or bound to a key. Functions not intended to be called in this way should not have interactive declarations. See Section 19.2 [Defining Commands], page 282, for how to write an interactive declaration.

The rest of the elements are the body of the function: the Lisp code to do the work of the function (or, as a Lisp programmer would say, “a list of Lisp forms to evaluate”). The value returned by the function is the value returned by the last element of the body.

11.2.2 A Simple Lambda-Expression Example

Consider for example the following function:

     (lambda (a b c) (+ a b c))

We can call this function by writing it as the car of an expression, like this:

     ((lambda (a b c) (+ a b c))
      1 2 3)

This call evaluates the body of the lambda expression with the variable a bound to 1, b bound to 2, and c bound to 3. Evaluation of the body adds these three numbers, producing the result 6; therefore, this call to the function returns the value 6.

Note that the arguments can be the results of other function calls, as in this example:

     ((lambda (a b c) (+ a b c))
      1 (* 2 3) (- 5 4))

This evaluates the arguments 1, (* 2 3), and (- 5 4) from left to right. Then it applies the lambda expression to the argument values 1, 6 and 1 to produce the value 8.

It is not often useful to write a lambda expression as the car of a form in this way. You can get the same result, of making local variables and giving them values, using the special form let (see Section 10.3 [Local Variables], page 146). And let is clearer and easier to use. In practice, lambda expressions are either stored as the function definitions of symbols, to produce named functions, or passed as arguments to other functions (see Section 11.7 [Anonymous Functions], page 172).

However, calls to explicit lambda expressions were very useful in the old days of Lisp, before the special form let was invented. At that time, they were the only way to bind and initialize local variables.

11.2.3 Advanced Features of Argument Lists

Our simple sample function, (lambda (a b c) (+ a b c)), specifies three argument variables, so it must be called with three arguments: if you try to call it with only two arguments or four arguments, you get a wrong-number-of-arguments error.

It is often convenient to write a function that allows certain arguments to be omitted. For example, the function substring accepts three arguments—a string, the start index and the end index—but the third argument defaults to the length of the string if you omit it. It is also convenient for certain functions to accept an indefinite number of arguments, as the functions list and + do.

To specify optional arguments that may be omitted when a function is called, simply include the keyword &optional before the optional arguments. To specify a list of zero or more extra arguments, include the keyword &rest before one final argument.

Thus, the complete syntax for an argument list is as follows:

(required-vars...

[&optional optional-vars...]

[&rest rest-var]) The square brackets indicate that the &optional and &rest clauses, and the variables that follow them, are optional.

A call to the function requires one actual argument for each of the required-vars. There may be actual arguments for zero or more of the optional-vars, and there cannot be any actual arguments beyond that unless the lambda list uses &rest. In that case, there may be any number of extra actual arguments.

If actual arguments for the optional and rest variables are omitted, then they always default to nil. There is no way for the function to distinguish between an explicit argument of nil and an omitted argument. However, the body of the function is free to consider nil an abbreviation for some other meaningful value. This is what substring does; nil as the third argument to substring means to use the length of the string supplied.

Common Lisp note: Common Lisp allows the function to specify what default value to use when an optional argument is omitted; XEmacs Lisp always uses

     nil.

For example, an argument list that looks like this:

     (a b &optional c d &rest e)

binds a and b to the first two actual arguments, which are required. If one or two more arguments are provided, c and d are bound to them respectively; any arguments after the first four are collected into a list and e is bound to that list. If there are only two arguments, c is nil; if two or three arguments, d is nil; if four arguments or fewer, e is nil.

There is no way to have required arguments following optional ones—it would not make sense. To see why this must be so, suppose that c in the example were optional and d were required. Suppose three actual arguments are given; which variable would the third argument be for? Similarly, it makes no sense to have any more arguments (either required or optional) after a &rest argument.

Here are some examples of argument lists and proper calls:

     ((lambda (n) (1+ n))                ; One required:

1) ; requires exactly one argument.

          ⇒ 2

((lambda (n &optional n1) ; One required and one optional:

              (if n1 (+ n n1) (1+ n)))   ; 1 or 2 arguments.
      1 2)

          ⇒ 3

((lambda (n &rest ns) ; One required and one rest:

              (+ n (apply '+ ns)))       ; 1 or more arguments.
      1 2 3 4 5)
          ⇒ 15

11.2.4 Documentation Strings of Functions

A lambda expression may optionally have a documentation string just after the lambda list. This string does not affect execution of the function; it is a kind of comment, but a systematized comment which actually appears inside the Lisp world and can be used by the XEmacs help facilities. See Chapter 27 [Documentation], page 381, for how the documentation-string is accessed.

It is a good idea to provide documentation strings for all the functions in your program, even those that are only called from within your program. Documentation strings are like comments, except that they are easier to access.

The first line of the documentation string should stand on its own, because apropos displays just this first line. It should consist of one or two complete sentences that summarize the function’s purpose.

The start of the documentation string is usually indented in the source file, but since these spaces come before the starting double-quote, they are not part of the string. Some people make a practice of indenting any additional lines of the string so that the text lines up in the program source. This is a mistake. The indentation of the following lines is inside the string; what looks nice in the source code will look ugly when displayed by the help commands.

You may wonder how the documentation string could be optional, since there are required components of the function that follow it (the body). Since evaluation of a string returns that string, without any side effects, it has no effect if it is not the last form in the body. Thus, in practice, there is no confusion between the first form of the body and the documentation string; if the only body form is a string then it serves both as the return value and as the documentation.

11.3 Naming a Function

In most computer languages, every function has a name; the idea of a function without a name is nonsensical. In Lisp, a function in the strictest sense has no name. It is simply a list whose first element is lambda, or a primitive subr-object.

However, a symbol can serve as the name of a function. This happens when you put the function in the symbol’s function cell (see Section 7.1 [Symbol Components], page 111). Then the symbol itself becomes a valid, callable function, equivalent to the list or subrobject that its function cell refers to. The contents of the function cell are also called the symbol’s function definition. The procedure of using a symbol’s function definition in place of the symbol is called symbol function indirection; see Section 8.2.4 [Function Indirection], page 123.

In practice, nearly all functions are given names in this way and referred to through their names. For example, the symbol car works as a function and does what it does because the primitive subr-object #<subr car> is stored in its function cell.

We give functions names because it is convenient to refer to them by their names in Lisp expressions. For primitive subr-objects such as #<subr car>, names are the only way you can refer to them: there is no read syntax for such objects. For functions written in Lisp, the name is more convenient to use in a call than an explicit lambda expression. Also, a function with a name can refer to itself—it can be recursive. Writing the function’s name in its own definition is much more convenient than making the function definition point to itself (something that is not impossible but that has various disadvantages in practice).

We often identify functions with the symbols used to name them. For example, we often speak of “the function car”, not distinguishing between the symbol car and the primitive subr-object that is its function definition. For most purposes, there is no need to distinguish.

Even so, keep in mind that a function need not have a unique name. While a given function object usually appears in the function cell of only one symbol, this is just a matter of convenience. It is easy to store it in several symbols using fset; then each of the symbols is equally well a name for the same function.

A symbol used as a function name may also be used as a variable; these two uses of a symbol are independent and do not conflict.

11.4 Defining Functions

We usually give a name to a function when it is first created. This is called defining a function, and it is done with the defun special form.

Special Form defun name argument-list body-forms defun is the usual way to define new Lisp functions. It defines the symbol name as a function that looks like this:

(lambda argument-list . body-forms) defun stores this lambda expression in the function cell of name. It returns the value name, but usually we ignore this value. As described previously (see Section 11.2 [Lambda Expressions], page 164), argument-list is a list of argument names and may include the keywords &optional and &rest. Also, the first two forms in body-forms may be a documentation string and an interactive declaration. There is no conflict if the same symbol name is also used as a variable, since the symbol’s value cell is independent of the function cell. See Section 7.1 [Symbol Components], page 111. Here are some examples:

          (defun foo () 5)
               ⇒ foo
          (foo)⇒ 5
          (defun bar (a &optional b &rest c)
              (list a b c))
               ⇒ bar
          (bar 1 2 3 4 5) ⇒ (1 2 (3 4 5))

     (bar 1)
          ⇒ (1 nil nil)
     (bar)
      error Wrong number of arguments.
     (defun capitalize-backwards ()
       "Upcase the last letter of a word."
       (interactive)
       (backward-word 1)
       (forward-word 1)
       (backward-char 1)
       (capitalize-word 1))
          ⇒ capitalize-backwards

Be careful not to redefine existing functions unintentionally. defun redefines even primitive functions such as car without any hesitation or notification. Redefining a function already defined is often done deliberately, and there is no way to distinguish deliberate redefinition from unintentional redefinition.

Function define-function name definition

Function defalias name definition These equivalent special forms define the symbol name as a function, with definition definition (which can be any valid Lisp function). The proper place to use define-function or defalias is where a specific function name is being defined—especially where that name appears explicitly in the source file being loaded. This is because define-function and defalias record which file defined the function, just like defun. (see Section 14.5 [Unloading], page 203). By contrast, in programs that manipulate function definitions for other purposes, it is better to use fset, which does not keep such records.

See also defsubst, which defines a function like defun and tells the Lisp compiler to open-code it. See Section 11.9 [Inline Functions], page 175.

11.5 Calling Functions

Defining functions is only half the battle. Functions don’t do anything until you call them, i.e., tell them to run. Calling a function is also known as invocation.

The most common way of invoking a function is by evaluating a list. For example, evaluating the list (concat "a" "b") calls the function concat with arguments "a" and "b". See Chapter 8 [Evaluation], page 119, for a description of evaluation.

When you write a list as an expression in your program, the function name is part of the program. This means that you choose which function to call, and how many arguments to give it, when you write the program. Usually that’s just what you want. Occasionally you need to decide at run time which function to call. To do that, use the functions funcall

and apply.

Function funcall function &rest arguments funcall calls function with arguments, and returns whatever function returns.

Since funcall is a function, all of its arguments, including function, are evaluated before funcall is called. This means that you can use any expression to obtain the function to be called. It also means that funcall does not see the expressions you write for the arguments, only their values. These values are not evaluated a second time in the act of calling function; funcall enters the normal procedure for calling a function at the place where the arguments have already been evaluated.

The argument function must be either a Lisp function or a primitive function. Special forms and macros are not allowed, because they make sense only when given the “unevaluated” argument expressions. funcall cannot provide these because, as we saw above, it never knows them in the first place.

     (setq f 'list)
          ⇒ list
     (funcall f 'x 'y 'z)
          ⇒ (x y z)
     (funcall f 'x 'y '(z))
          ⇒ (x y (z))
     (funcall 'and t nil)
      error Invalid function: #<subr and>

Compare these example with the examples of apply.

Function apply function &rest arguments apply calls function with arguments, just like funcall but with one difference: the last of arguments is a list of arguments to give to function, rather than a single argument. We also say that apply spreads this list so that each individual element becomes an argument.

apply returns the result of calling function. As with funcall, function must either be a Lisp function or a primitive function; special forms and macros do not make sense in apply.

     (setq f 'list)
          ⇒ list
     (apply f 'x 'y 'z)
      error Wrong type argument: listp, z
     (apply '+ 1 2 '(3 4)) ⇒ 10
     (apply '+ '(1 2 3 4)) ⇒ 10
     (apply 'append '((a b c) nil (x y z) nil))
          ⇒ (a b c x y z)

For an interesting example of using apply, see the description of mapcar, in Section 11.6 [Mapping Functions], page 171.

It is common for Lisp functions to accept functions as arguments or find them in data structures (especially in hook variables and property lists) and call them using funcall or apply. Functions that accept function arguments are often called functionals.

Sometimes, when you call a functional, it is useful to supply a no-op function as the argument. Here are two different kinds of no-op function:

Function identity arg

This function returns arg and has no side effects.

Function ignore &rest args This function ignores any arguments and returns nil.

11.6 Mapping Functions

A mapping function applies a given function to each element of a list or other collection. XEmacs Lisp has three such functions; mapcar and mapconcat, which scan a list, are described here. For the third mapping function, mapatoms, see Section 7.3 [Creating Symbols], page 113.

Function mapcar function sequence mapcar applies function to each element of sequence in turn, and returns a list of the results. The argument sequence may be a list, a vector, or a string. The result is always a list. The length of the result is the same as the length of sequence. For example:

     (mapcar 'car '((a b) (c d) (e f)))
          ⇒ (a c e)
     (mapcar '1+ [1 2 3])
          ⇒ (2 3 4)
     (mapcar 'char-to-string "abc")
          ⇒ ("a" "b" "c")

;; Call each function in my-hooks.

     (mapcar 'funcall my-hooks)
     (defun mapcar* (f &rest args)
       "Apply FUNCTION to successive cars of all ARGS.
     Return the list of results."

;; If no list is exhausted,

       (if (not (memq 'nil args))

;; apply function to CARs.

           (cons (apply f (mapcar 'car args))
                 (apply 'mapcar* f

;; Recurse for rest of elements.

                        (mapcar 'cdr args)))))
     (mapcar* 'cons '(a b c) '(1 2 3 4))
          ⇒ ((a . 1) (b . 2) (c . 3))

Function mapconcat function sequence separator mapconcat applies function to each element of sequence: the results, which must be strings, are concatenated. Between each pair of result strings, mapconcat inserts the string separator. Usually separator contains a space or comma or other suitable punctuation. The argument function must be a function that can take one argument and return a string.

          (mapconcat 'symbol-name
                     '(The cat in the hat)
                     " ")
               ⇒ "The cat in the hat"
          (mapconcat (function (lambda (x) (format "%c" (1+ x))))
                     "HAL-8000"
                     "")
               ⇒ "IBM.9111"

11.7 Anonymous Functions

In Lisp, a function is a list that starts with lambda, a byte-code function compiled from such a list, or alternatively a primitive subr-object; names are “extra”. Although usually functions are defined with defun and given names at the same time, it is occasionally more concise to use an explicit lambda expression—an anonymous function. Such a list is valid wherever a function name is.

Any method of creating such a list makes a valid function. Even this:

     (setq silly (append '(lambda (x)) (list (list '+ (* 3 4) 'x))))
     ⇒ (lambda (x) (+ 12 x))

This computes a list that looks like (lambda (x) (+ 12 x)) and makes it the value (not the function definition!) of silly.

Here is how we might call this function:

     (funcall silly 1)
     ⇒ 13

(It does not work to write (silly 1), because this function is not the function definition of silly. We have not given silly any function definition, just a value as a variable.)

Most of the time, anonymous functions are constants that appear in your program. For example, you might want to pass one as an argument to the function mapcar, which applies any given function to each element of a list. Here we pass an anonymous function that multiplies a number by two:

     (defun double-each (list)
       (mapcar '(lambda (x) (* 2 x)) list))
     ⇒ double-each
     (double-each '(2 11))
     ⇒ (4 22)

In such cases, we usually use the special form function instead of simple quotation to quote the anonymous function.

Special Form function function-object This special form returns function-object without evaluating it. In this, it is equivalent to quote. However, it serves as a note to the XEmacs Lisp compiler that function-object is intended to be used only as a function, and therefore can safely be compiled. Contrast this with quote, in Section 8.3 [Quoting], page 126.

Using function instead of quote makes a difference inside a function or macro that you are going to compile. For example:

     (defun double-each (list)
       (mapcar (function (lambda (x) (* 2 x))) list))
     ⇒ double-each
     (double-each '(2 11))
     ⇒ (4 22)

If this definition of double-each is compiled, the anonymous function is compiled as well. By contrast, in the previous definition where ordinary quote is used, the argument passed to mapcar is the precise list shown:

     (lambda (x) (* x 2))

The Lisp compiler cannot assume this list is a function, even though it looks like one, since it does not know what mapcar does with the list. Perhaps mapcar will check that the car of the third element is the symbol *! The advantage of function is that it tells the compiler to go ahead and compile the constant function.

We sometimes write function instead of quote when quoting the name of a function, but this usage is just a sort of comment.

(function symbol) ≡ (quote symbol) ≡ 'symbol See documentation in Section 27.2 [Accessing Documentation], page 382, for a realistic

example using function and an anonymous function.

11.8 Accessing Function Cell Contents

The function definition of a symbol is the object stored in the function cell of the symbol. The functions described here access, test, and set the function cell of symbols.

See also the function indirect-function in Section 8.2.4 [Function Indirection], page 123.

Function symbol-function symbol

This returns the object in the function cell of symbol. If the symbol’s function cell is void, a void-function error is signaled. This function does not check that the returned object is a legitimate function.

          (defun bar (n) (+ n 2))
               ⇒ bar
          (symbol-function 'bar)
               ⇒ (lambda (n) (+ n 2))
          (fset 'baz 'bar)
               ⇒ bar
          (symbol-function 'baz)
               ⇒ bar

If you have never given a symbol any function definition, we say that that symbol’s function cell is void. In other words, the function cell does not have any Lisp object in it. If you try to call such a symbol as a function, it signals a void-function error.

Note that void is not the same as nil or the symbol void. The symbols nil and void are Lisp objects, and can be stored into a function cell just as any other object can be (and they can be valid functions if you define them in turn with defun). A void function cell contains no object whatsoever.

You can test the voidness of a symbol’s function definition with fboundp. After you have given a symbol a function definition, you can make it void once more using fmakunbound.

Function fboundp symbol

This function returns t if the symbol has an object in its function cell, nil otherwise. It does not check that the object is a legitimate function.

Function fmakunbound symbol

This function makes symbol’s function cell void, so that a subsequent attempt to access this cell will cause a void-function error. (See also makunbound, in Section 10.3 [Local Variables], page 146.)

     (defun foo (x) x)
          ⇒ x
     (foo 1) ⇒1
     (fmakunbound 'foo)
          ⇒ x
     (foo 1)
      error Symbol's function definition is void: foo

Function fset symbol object This function stores object in the function cell of symbol. The result is object. Normally object should be a function or the name of a function, but this is not checked. There are three normal uses of this function:

• Copying one symbol’s function definition to another. (In other words, making an alternate name for a function.)

• Giving a symbol a function definition that is not a list and therefore cannot be made with defun. For example, you can use fset to give a symbol s1 a function definition which is another symbol s2; then s1 serves as an alias for whatever definition s2 presently has.

• In constructs for defining or altering functions. If defun were not a primitive, it could be written in Lisp (as a macro) using fset.

Here are examples of the first two uses:

;; Give first the same definition car has.

     (fset 'first (symbol-function 'car))
          ⇒ #<subr car>
     (first '(1 2 3)) ⇒ 1

;; Make the symbol car the function definition of xfirst.

          (fset 'xfirst 'car)
               ⇒ car
          (xfirst '(1 2 3)) ⇒ 1
          (symbol-function 'xfirst)
               ⇒ car
          (symbol-function (symbol-function 'xfirst))
               ⇒ #<subr car>

;; Define a named keyboard macro.

          (fset 'kill-two-lines "\^u2\^k")
               ⇒ "\^u2\^k"

See also the related functions define-function and defalias, in Section 11.4 [Defining Functions], page 168.

When writing a function that extends a previously defined function, the following idiom is sometimes used:

     (fset 'old-foo (symbol-function 'foo))
     (defun foo ()
       "Just like old-foo, except more so."
       (old-foo) (more-so))

This does not work properly if foo has been defined to autoload. In such a case, when foo calls old-foo, Lisp attempts to define old-foo by loading a file. Since this presumably defines foo rather than old-foo, it does not produce the proper results. The only way to avoid this problem is to make sure the file is loaded before moving aside the old definition

of foo.

But it is unmodular and unclean, in any case, for a Lisp file to redefine a function defined elsewhere.

11.9 Inline Functions

You can define an inline function by using defsubst instead of defun. An inline function works just like an ordinary function except for one thing: when you compile a call to the function, the function’s definition is open-coded into the caller.

Making a function inline makes explicit calls run faster. But it also has disadvantages. For one thing, it reduces flexibility; if you change the definition of the function, calls already inlined still use the old definition until you recompile them. Since the flexibility of redefining functions is an important feature of XEmacs, you should not make a function inline unless its speed is really crucial.

Another disadvantage is that making a large function inline can increase the size of compiled code both in files and in memory. Since the speed advantage of inline functions is greatest for small functions, you generally should not make large functions inline.

It’s possible to define a macro to expand into the same code that an inline function would execute. But the macro would have a limitation: you can use it only explicitly—a macro cannot be called with apply, mapcar and so on. Also, it takes some work to convert an ordinary function into a macro. (See Chapter 12 [Macros], page 177.) To convert it into an inline function is very easy; simply replace defun with defsubst. Since each argument of an inline function is evaluated exactly once, you needn’t worry about how many times the body uses the arguments, as you do for macros. (See Section 12.6.1 [Argument Evaluation], page 180.)

Inline functions can be used and open-coded later on in the same file, following the definition, just like macros.

11.10 Other Topics Related to Functions

Here is a table of several functions that do things related to function calling and function definitions. They are documented elsewhere, but we provide cross references here. apply See Section 11.5 [Calling Functions], page 169. autoload See Section 14.2 [Autoload], page 198.

call-interactively

See Section 19.3 [Interactive Call], page 286.

commandp See Section 19.3 [Interactive Call], page 286.

documentation

See Section 27.2 [Accessing Documentation], page 382.

eval See Section 8.1 [Eval], page 120. funcall See Section 11.5 [Calling Functions], page 169. ignore See Section 11.5 [Calling Functions], page 169.

indirect-function

See Section 8.2.4 [Function Indirection], page 123.

interactive

See Section 19.2.1 [Using Interactive], page 282.

interactive-p

See Section 19.3 [Interactive Call], page 286.

mapatoms See Section 7.3 [Creating Symbols], page 113. mapcar See Section 11.6 [Mapping Functions], page 171.

mapconcat

See Section 11.6 [Mapping Functions], page 171.

undefined

See Section 20.8 [Key Lookup], page 324.

12 Macros

Macros enable you to define new control constructs and other language features. A macro is defined much like a function, but instead of telling how to compute a value, it tells how to compute another Lisp expression which will in turn compute the value. We call this expression the expansion of the macro.

Macros can do this because they operate on the unevaluated expressions for the arguments, not on the argument values as functions do. They can therefore construct an expansion containing these argument expressions or parts of them.

If you are using a macro to do something an ordinary function could do, just for the sake of speed, consider using an inline function instead. See Section 11.9 [Inline Functions], page 175.

12.1 A Simple Example of a Macro

Suppose we would like to define a Lisp construct to increment a variable value, much like the ++ operator in C. We would like to write (inc x) and have the effect of (setq x (1+ x)). Here’s a macro definition that does the job:

     (defmacro inc (var)
        (list 'setq var (list '1+ var)))

When this is called with (inc x), the argument var has the value x—not the value of x. The body of the macro uses this to construct the expansion, which is (setq x (1+ x)). Once the macro definition returns this expansion, Lisp proceeds to evaluate it, thus incrementing x.

12.2 Expansion of a Macro Call

A macro call looks just like a function call in that it is a list which starts with the name of the macro. The rest of the elements of the list are the arguments of the macro.

Evaluation of the macro call begins like evaluation of a function call except for one crucial difference: the macro arguments are the actual expressions appearing in the macro call. They are not evaluated before they are given to the macro definition. By contrast, the arguments of a function are results of evaluating the elements of the function call list.

Having obtained the arguments, Lisp invokes the macro definition just as a function is invoked. The argument variables of the macro are bound to the argument values from the macro call, or to a list of them in the case of a &rest argument. And the macro body executes and returns its value just as a function body does.

The second crucial difference between macros and functions is that the value returned by the macro body is not the value of the macro call. Instead, it is an alternate expression for computing that value, also known as the expansion of the macro. The Lisp interpreter proceeds to evaluate the expansion as soon as it comes back from the macro.

Since the expansion is evaluated in the normal manner, it may contain calls to other macros. It may even be a call to the same macro, though this is unusual.

You can see the expansion of a given macro call by calling macroexpand.

Function macroexpand form &optional environment This function expands form, if it is a macro call. If the result is another macro call, it is expanded in turn, until something which is not a macro call results. That is the value returned by macroexpand. If form is not a macro call to begin with, it is returned as given. Note that macroexpand does not look at the subexpressions of form (although some macro definitions may do so). Even if they are macro calls themselves, macroexpand does not expand them. The function macroexpand does not expand calls to inline functions. Normally there is no need for that, since a call to an inline function is no harder to understand than a call to an ordinary function. If environment is provided, it specifies an alist of macro definitions that shadow the currently defined macros. Byte compilation uses this feature.

          (defmacro inc (var)
              (list 'setq var (list '1+ var)))
               ⇒ inc
          (macroexpand '(inc r))
               ⇒ (setq r (1+ r))
          (defmacro inc2 (var1 var2)
              (list 'progn (list 'inc var1) (list 'inc var2)))
               ⇒ inc2
          (macroexpand '(inc2 r s))
               ⇒ (progn (inc r) (inc s))  ; inc not expanded here.

12.3 Macros and Byte Compilation

You might ask why we take the trouble to compute an expansion for a macro and then evaluate the expansion. Why not have the macro body produce the desired results directly? The reason has to do with compilation.

When a macro call appears in a Lisp program being compiled, the Lisp compiler calls the macro definition just as the interpreter would, and receives an expansion. But instead of evaluating this expansion, it compiles the expansion as if it had appeared directly in the program. As a result, the compiled code produces the value and side effects intended for the macro, but executes at full compiled speed. This would not work if the macro body computed the value and side effects itself—they would be computed at compile time, which is not useful.

In order for compilation of macro calls to work, the macros must be defined in Lisp when the calls to them are compiled. The compiler has a special feature to help you do this: if a file being compiled contains a defmacro form, the macro is defined temporarily for the rest of the compilation of that file. To use this feature, you must define the macro in the same file where it is used and before its first use.

Byte-compiling a file executes any require calls at top-level in the file. This is in case the file needs the required packages for proper compilation. One way to ensure that necessary macro definitions are available during compilation is to require the files that define them (see Section 14.4 [Named Features], page 201). To avoid loading the macro definition files when someone runs the compiled program, write eval-when-compile around the require calls (see Section 15.5 [Eval During Compile], page 209).

12.4 Defining Macros

A Lisp macro is a list whose car is macro. Its cdr should be a function; expansion of the macro works by applying the function (with apply) to the list of unevaluated argumentexpressions from the macro call.

It is possible to use an anonymous Lisp macro just like an anonymous function, but this is never done, because it does not make sense to pass an anonymous macro to functionals such as mapcar. In practice, all Lisp macros have names, and they are usually defined with the special form defmacro.

Special Form defmacro name argument-list body-forms…

defmacro defines the symbol name as a macro that looks like this:

(macro lambda argument-list . body-forms) This macro object is stored in the function cell of name. The value returned by evaluating the defmacro form is name, but usually we ignore this value. The shape and meaning of argument-list is the same as in a function, and the keywords &rest and &optional may be used (see Section 11.2.3 [Argument List], page 165). Macros may have a documentation string, but any interactive declaration is ignored since macros cannot be called interactively.

12.5 Backquote

Macros often need to construct large list structures from a mixture of constants and nonconstant parts. To make this easier, use the macro ‘`’ (often called backquote).

Backquote allows you to quote a list, but selectively evaluate elements of that list. In the simplest case, it is identical to the special form quote (see Section 8.3 [Quoting], page 126). For example, these two forms yield identical results:

     `(a list of (+ 2 3) elements)
          ⇒ (a list of (+ 2 3) elements)
     '(a list of (+ 2 3) elements)
          ⇒ (a list of (+ 2 3) elements)

The special marker ‘,’ inside of the argument to backquote indicates a value that isn’t constant. Backquote evaluates the argument of ‘,’ and puts the value in the list structure:

     (list 'a 'list 'of (+ 2 3) 'elements)
          ⇒ (a list of 5 elements)
     `(a list of ,(+ 2 3) elements)
          ⇒ (a list of 5 elements)

You can also splice an evaluated value into the resulting list, using the special marker ‘,@’. The elements of the spliced list become elements at the same level as the other elements of the resulting list. The equivalent code without using ‘`’ is often unreadable. Here are some examples:

(setq some-list '(2 3))
     ⇒ (2 3)
(cons 1 (append some-list '(4) some-list))
     ⇒ (1 2 3 4 2 3)
`(1 ,@some-list 4 ,@some-list)
     ⇒ (1 2 3 4 2 3)
(setq list '(hack foo bar))
     ⇒ (hack foo bar)
(cons 'use
  (cons 'the
    (cons 'words (append (cdr list) '(as elements)))))
     ⇒ (use the words foo bar as elements)
`(use the words ,@(cdr list) as elements)
     ⇒ (use the words foo bar as elements)

Before Emacs version 19.29, ‘`’ used a different syntax which required an extra level of parentheses around the entire backquote construct. Likewise, each ‘,’ or ‘,@’ substitution required an extra level of parentheses surrounding both the ‘,’ or ‘,@’ and the following expression. The old syntax required whitespace between the ‘`’, ‘,’ or ‘,@’ and the following expression. This syntax is still accepted, but no longer recommended except for compatibility with old Emacs versions.

12.6 Common Problems Using Macros

The basic facts of macro expansion have counterintuitive consequences. This section describes some important consequences that can lead to trouble, and rules to follow to avoid trouble.

12.6.1 Evaluating Macro Arguments Repeatedly

When defining a macro you must pay attention to the number of times the arguments will be evaluated when the expansion is executed. The following macro (used to facilitate iteration) illustrates the problem. This macro allows us to write a simple “for” loop such as one might find in Pascal.

(defmacro for (var from init to final do &rest body)
  "Execute a simple \"for\" loop.
For example, (for i from 1 to 10 do (print i))."
  (list 'let (list (list var init))
        (cons 'while (cons (list '<= var final)
                           (append body (list (list 'inc var)))))))
⇒ for
(for i from 1 to 3 do
   (setq square (* i i))
   (princ (format "\n%d %d" i square)))

↦

     (let ((i 1))
       (while (<= i 3)
         (setq square (* i i))
         (princ (format "%d      %d" i square))
         (inc i)))
     ⊣ 1       1 ⊣ 2       4 ⊣ 3       9 ⇒ nil

(The arguments from, to, and do in this macro are “syntactic sugar”; they are entirely ignored. The idea is that you will write noise words (such as from, to, and do) in those positions in the macro call.)

Here’s an equivalent definition simplified through use of backquote:

     (defmacro for (var from init to final do &rest body)
       "Execute a simple \"for\" loop.
     For example, (for i from 1 to 10 do (print i))."
       `(let ((,var ,init))
          (while (<= ,var ,final)
            ,@body
            (inc ,var))))

Both forms of this definition (with backquote and without) suffer from the defect that final is evaluated on every iteration. If final is a constant, this is not a problem. If it is a more complex form, say (long-complex-calculation x), this can slow down the execution significantly. If final has side effects, executing it more than once is probably incorrect.

A well-designed macro definition takes steps to avoid this problem by producing an expansion that evaluates the argument expressions exactly once unless repeated evaluation is part of the intended purpose of the macro. Here is a correct expansion for the for macro:

     (let ((i 1)
           (max 3))
       (while (<= i max)
         (setq square (* i i))
         (princ (format "%d      %d" i square))
         (inc i)))

Here is a macro definition that creates this expansion:

     (defmacro for (var from init to final do &rest body)
       "Execute a simple for loop: (for i from 1 to 10 do (print i))."
       `(let ((,var ,init)
              (max ,final))
          (while (<= ,var max)
            ,@body
            (inc ,var))))

Unfortunately, this introduces another problem.

12.6.2 Local Variables in Macro Expansions

The new definition of for has a new problem: it introduces a local variable named max which the user does not expect. This causes trouble in examples such as the following:

     (let ((max 0))
       (for x from 0 to 10 do
         (let ((this (frob x)))
           (if (< max this)
               (setq max this)))))

The references to max inside the body of the for, which are supposed to refer to the user’s binding of max, really access the binding made by for.

The way to correct this is to use an uninterned symbol instead of max (see Section 7.3 [Creating Symbols], page 113). The uninterned symbol can be bound and referred to just like any other symbol, but since it is created by for, we know that it cannot already appear in the user’s program. Since it is not interned, there is no way the user can put it into the program later. It will never appear anywhere except where put by for. Here is a definition of for that works this way:

     (defmacro for (var from init to final do &rest body)
       "Execute a simple for loop: (for i from 1 to 10 do (print i))."
       (let ((tempvar (make-symbol "max")))
         `(let ((,var ,init)
                (,tempvar ,final))
            (while (<= ,var ,tempvar)
              ,@body
              (inc ,var)))))

This creates an uninterned symbol named max and puts it in the expansion instead of the usual interned symbol max that appears in expressions ordinarily.

12.6.3 Evaluating Macro Arguments in Expansion

Another problem can happen if you evaluate any of the macro argument expressions during the computation of the expansion, such as by calling eval (see Section 8.1 [Eval], page 120). If the argument is supposed to refer to the user’s variables, you may have trouble if the user happens to use a variable with the same name as one of the macro arguments. Inside the macro body, the macro argument binding is the most local binding of this variable, so any references inside the form being evaluated do refer to it. Here is an example:

     (defmacro foo (a)
       (list 'setq (eval a) t))
          ⇒ foo

     (setq x 'b)
     (foo x) ↦ (setq b t)

⇒ t ; and b has been set. ;; but

     (setq a 'c)
     (foo a) ↦ (setq a t)

⇒ t ; but this set a, not c.

It makes a difference whether the user’s variable is named a or x, because a conflicts with the macro argument variable a.

Another reason not to call eval in a macro definition is that it probably won’t do what you intend in a compiled program. The byte-compiler runs macro definitions while compiling the program, when the program’s own computations (which you might have wished to access with eval) don’t occur and its local variable bindings don’t exist.

The safe way to work with the run-time value of an expression is to put the expression into the macro expansion, so that its value is computed as part of executing the expansion.

12.6.4 How Many Times is the Macro Expanded?

Occasionally problems result from the fact that a macro call is expanded each time it is evaluated in an interpreted function, but is expanded only once (during compilation) for a compiled function. If the macro definition has side effects, they will work differently depending on how many times the macro is expanded.

In particular, constructing objects is a kind of side effect. If the macro is called once, then the objects are constructed only once. In other words, the same structure of objects is used each time the macro call is executed. In interpreted operation, the macro is reexpanded each time, producing a fresh collection of objects each time. Usually this does not matter— the objects have the same contents whether they are shared or not. But if the surrounding program does side effects on the objects, it makes a difference whether they are shared. Here is an example:

     (defmacro empty-object ()
       (list 'quote (cons nil nil)))
     (defun initialize (condition)
       (let ((object (empty-object)))
         (if condition
             (setcar object condition))
         object))

If initialize is interpreted, a new list (nil) is constructed each time initialize is called. Thus, no side effect survives between calls. If initialize is compiled, then the macro empty-object is expanded during compilation, producing a single “constant” (nil) that is reused and altered each time initialize is called.

One way to avoid pathological cases like this is to think of empty-object as a funny kind of constant, not as a memory allocation construct. You wouldn’t use setcar on a constant such as '(nil), so naturally you won’t use it on (empty-object) either.

184 XEmacs Lisp Reference Manual

13 Writing Customization Definitions

This chapter describes how to declare user options for customization, and also customization groups for classifying them. We use the term customization item to include both kinds of customization definitions—as well as face definitions.

13.1 Common Keywords for All Kinds of Items

All kinds of customization declarations (for variables and groups, and for faces) accept keyword arguments for specifying various information. This section describes some keywords that apply to all kinds.

All of these keywords, except :tag, can be used more than once in a given item. Each use of the keyword has an independent effect. The keyword :tag is an exception because any given item can only display one name.

:tag name

Use name, a string, instead of the item’s name, to label the item in customization menus and buffers.

:group group

Put this customization item in group group. When you use :group in a defgroup, it makes the new group a subgroup of group. If you use this keyword more than once, you can put a single item into more than one group. Displaying any of those groups will show this item. Be careful not to overdo this!

:link link-data Include an external link after the documentation string for this item. This is a sentence containing an active field which references some other documentation. There are three alternatives you can use for link-data:

(custom-manual info-node)

Link to an Info node; info-node is a string which specifies the node name, as in "(emacs)Top". The link appears as ‘[manual]’ in the customization buffer.

(info-link info-node)

Like custom-manual except that the link appears in the customization buffer with the Info node name.

(url-link url)

Link to a web page; url is a string which specifies the url. The link appears in the customization buffer as url.

You can specify the text to use in the customization buffer by adding :tag name after the first element of the link-data; for example, (info-link :tag "foo" "(emacs)Top") makes a link to the Emacs manual which appears in the buffer as ‘foo’. An item can have more than one external link; however, most items have none at all.

:load file Load file file (a string) before displaying this customization item. Loading is done with load-library, and only if the file is not already loaded.

:require feature

Require feature feature (a symbol) when installing a value for this item (an option or a face) that was saved using the customization feature. This is done by calling require. The most common reason to use :require is when a variable enables a feature such as a minor mode, and just setting the variable won’t have any effect unless the code which implements the mode is loaded.

13.2 Defining Custom Groups

Each Emacs Lisp package should have one main customization group which contains all the options, faces and other groups in the package. If the package has a small number of options and faces, use just one group and put everything in it. When there are more than twelve or so options and faces, then you should structure them into subgroups, and put the subgroups under the package’s main customization group. It is OK to put some of the options and faces in the package’s main group alongside the subgroups.

The package’s main or only group should be a member of one or more of the standard customization groups. (To display the full list of them, use M-x customize.) Choose one or more of them (but not too many), and add your group to each of them using the :group keyword.

The way to declare new customization groups is with defgroup.

Macro defgroup group members doc [keyword value]...

Declare group as a customization group containing members. Do not quote the symbol group. The argument doc specifies the documentation string for the group.

The argument members is a list specifying an initial set of customization items to be members of the group. However, most often members is nil, and you specify the group’s members by using the :group keyword when defining those members.

If you want to specify group members through members, each element should have the form (name widget). Here name is a symbol, and widget is a widget type for editing that symbol. Useful widgets are custom-variable for a variable, custom-face for a face, and custom-group for a group.

In addition to the common keywords (see Section 13.1 [Common Keywords], page 185), you can use this keyword in defgroup:

:prefix prefix

If the name of an item in the group starts with prefix, then the tag for that item is constructed (by default) by omitting prefix. One group can have any number of prefixes.

13.3 Defining Customization Variables

Use defcustom to declare user-editable variables.

Macro defcustom option default doc [keyword value]...

Declare option as a customizable user option variable. Do not quote option. The argument doc specifies the documentation string for the variable. If option is void, defcustom initializes it to default. default should be an expression to compute the value; be careful in writing it, because it can be evaluated on more than one occasion. The following additional keywords are defined:

:type type

Use type as the data type for this option. It specifies which values are legitimate, and how to display the value. See Section 13.4 [Customization Types], page 188, for more information.

:options list

Specify list as the list of reasonable values for use in this option. Currently this is meaningful only when the type is hook. In that case, the elements of list should be functions that are useful as elements of the hook value. The user is not restricted to using only these functions, but they are offered as convenient alternatives.

:version version

This option specifies that the variable was first introduced, or its default value was changed, in Emacs version version. The value version must be a string. For example,

     (defcustom foo-max 34
       "*Maximum number of foo's allowed."
       :type 'integer
       :group 'foo
       :version "20.3")

:set setfunction Specify setfunction as the way to change the value of this option. The function setfunction should take two arguments, a symbol and the new value, and should do whatever is necessary to update the value properly for this option (which may not mean simply setting the option as a Lisp variable). The default for setfunction is set-

default.

:get getfunction Specify getfunction as the way to extract the value of this option. The function getfunction should take one argument, a symbol, and should return the “current value” for that symbol (which need not be the symbol’s Lisp value). The default is default-value.

:initialize function

function should be a function used to initialize the variable when the defcustom is evaluated. It should take two arguments, the symbol and value. Here are some predefined functions meant for use in this way:

custom-initialize-set

Use the variable’s :set function to initialize the variable, but do not reinitialize it if it is already non-void. This is the default :initialize function.

custom-initialize-default

Like custom-initialize-set, but use the function set-default to set the variable, instead of the variable’s :set function. This is the usual choice for a variable whose :set function enables or disables a minor mode; with this choice, defining the variable will not call the minor mode function, but customizing the variable will do so.

custom-initialize-reset

Always use the :set function to initialize the variable. If the variable is already non-void, reset it by calling the :set function using the current value (returned by the :get method).

custom-initialize-changed

Use the :set function to initialize the variable, if it is already set or has been customized; otherwise, just use

set-default.

The :require option is useful for an option that turns on the operation of a certain feature. Assuming that the package is coded to check the value of the option, you still need to arrange for the package to be loaded. You can do that with :require. See Section 13.1 [Common Keywords], page 185. Here is an example, from the library ‘paren.el’:

(defcustom show-paren-mode nil
"Toggle Show Paren mode...."
:set (lambda (symbol value)
(show-paren-mode (or value 0)))
:initialize 'custom-initialize-default
:type 'boolean
:group 'paren-showing
:require 'paren)

Internally, defcustom uses the symbol property standard-value to record the expression for the default value, and saved-value to record the value saved by the user with the customization buffer. The saved-value property is actually a list whose car is an expression which evaluates to the value.

13.4 Customization Types

When you define a user option with defcustom, you must specify its customization type. That is a Lisp object which describes (1) which values are legitimate and (2) how to display the value in the customization buffer for editing.

You specify the customization type in defcustom with the :type keyword. The argument of :type is evaluated; since types that vary at run time are rarely useful, normally you use a quoted constant. For example:

     (defcustom diff-command "diff"
       "*The command to use to run diff."
       :type '(string)
       :group 'diff)

In general, a customization type is a list whose first element is a symbol, one of the customization type names defined in the following sections. After this symbol come a number of arguments, depending on the symbol. Between the type symbol and its arguments, you can optionally write keyword-value pairs (see Section 13.4.4 [Type Keywords], page 192).

Some of the type symbols do not use any arguments; those are called simple types. For a simple type, if you do not use any keyword-value pairs, you can omit the parentheses around the type symbol. For example just string as a customization type is equivalent to

(string).

13.4.1 Simple Types

This section describes all the simple customization types.

sexp The value may be any Lisp object that can be printed and read back. You can use sexp as a fall-back for any option, if you don’t want to take the time to work out a more specific type to use.

integer The value must be an integer, and is represented textually in the customization buffer.

number The value must be a number, and is represented textually in the customization buffer.

string The value must be a string, and the customization buffer shows just the contents, with no delimiting ‘"’ characters and no quoting with ‘\’.

regexp Like string except that the string must be a valid regular expression.

character

The value must be a character code. A character code is actually an integer, but this type shows the value by inserting the character in the buffer, rather than by showing the number.

file The value must be a file name, and you can do completion with M-^{⟨}\underline{\overline{TAB}}^{⟩}.

(file :must-match t)

The value must be a file name for an existing file, and you can do completion with M-^{⟨}\underline{\overline{TAB}}^{⟩}.

directory

The value must be a directory name, and you can do completion with M-^{⟨}\underline{\overline{TAB}}^{⟩}.

symbol The value must be a symbol. It appears in the customization buffer as the name of the symbol. function The value must be either a lambda expression or a function name. When it is

a function name, you can do completion with M-^{⟨}\underline{\overline{TAB}}^{⟩}.

variable The value must be a variable name, and you can do completion with M-^{⟨}\underline{\overline{TAB}}^{⟩}.

face The value must be a symbol which is a face name.

boolean The value is boolean—either nil or t. Note that by using choice and const together (see the next section), you can specify that the value must be nil or t, but also specify the text to describe each value in a way that fits the specific meaning of the alternative.

13.4.2 Composite Types

When none of the simple types is appropriate, you can use composite types, which build new types from other types. Here are several ways of doing that:

(restricted-sexp :match-alternatives criteria)

The value may be any Lisp object that satisfies one of criteria. criteria should be a list, and each elements should be one of these possibilities:

• A predicate—that is, a function of one argument that returns non-nil if the argument fits a certain type. This means that objects of that type are acceptable.

• A quoted constant—that is, 'object. This means that object itself is an acceptable value.

For example,

     (restricted-sexp :match-alternatives (integerp 't 'nil))

allows integers, t and nil as legitimate values. The customization buffer shows all legitimate values using their read syntax, and the user edits them textually.

(cons car-type cdr-type)

The value must be a cons cell, its car must fit car-type, and its cdr must fit cdr-type. For example, (cons string symbol) is a customization type which matches values such as ("foo" . foo). In the customization buffer, the car and the cdr are displayed and edited separately, each according to the type that you specify for it.

(list element-types...)

The value must be a list with exactly as many elements as the element-types you have specified; and each element must fit the corresponding element-type. For example, (list integer string function) describes a list of three elements; the first element must be an integer, the second a string, and the third a function. In the customization buffer, the each element is displayed and edited separately, according to the type specified for it.

(vector element-types...)

Like list except that the value must be a vector instead of a list. The elements work the same as in list.

(choice alternative-types...)

The value must fit at least one of alternative-types. For example, (choice integer string) allows either an integer or a string. In the customization buffer, the user selects one of the alternatives using a menu, and can then edit the value in the usual way for that alternative. Normally the strings in this menu are determined automatically from the choices; however, you can specify different strings for the menu by including the :tag keyword in the alternatives. For example, if an integer stands for a number of spaces, while a string is text to use verbatim, you might write the customization type this way,

     (choice (integer :tag "Number of spaces")
             (string :tag "Literal text"))

so that the menu offers ‘Number of spaces’ and ‘Literal Text’. In any alternative for which nil is not a valid value, other than a const, you should specify a valid default for that alternative using the :value keyword. See Section 13.4.4 [Type Keywords], page 192.

(const value)

The value must be value—nothing else is allowed. The main use of const is inside of choice. For example, (choice integer (const nil)) allows either an integer or nil. :tag is often used with const, inside of choice. For example,

     (choice (const :tag "Yes" t)
             (const :tag "No" nil)
             (const :tag "Ask" foo))
(function-item function)

Like const, but used for values which are functions. This displays the documentation string as well as the function name. The documentation string is either the one you specify with :doc, or function’s own documentation string.

(variable-item variable)

Like const, but used for values which are variable names. This displays the documentation string as well as the variable name. The documentation string is either the one you specify with :doc, or variable’s own documentation string.

(set elements...)

The value must be a list and each element of the list must be one of the elements specified. This appears in the customization buffer as a checklist.

(repeat element-type)

The value must be a list and each element of the list must fit the type elementtype. This appears in the customization buffer as a list of elements, with ‘[INS]’ and ‘[DEL]’ buttons for adding more elements or removing elements.

13.4.3 Splicing into Lists

The :inline feature lets you splice a variable number of elements into the middle of a list or vector. You use it in a set, choice or repeat type which appears among the element-types of a list or vector.

Normally, each of the element-types in a list or vector describes one and only one element of the list or vector. Thus, if an element-type is a repeat, that specifies a list of unspecified length which appears as one element.

But when the element-type uses :inline, the value it matches is merged directly into the containing sequence. For example, if it matches a list with three elements, those become three elements of the overall sequence. This is analogous to using ‘,@’ in the backquote construct.

For example, to specify a list whose first element must be t and whose remaining arguments should be zero or more of foo and bar, use this customization type:

     (list (const t) (set :inline t foo bar))

This matches values such as (t), (t foo), (t bar) and (t foo bar).

When the element-type is a choice, you use :inline not in the choice itself, but in (some of) the alternatives of the choice. For example, to match a list which must start with a file name, followed either by the symbol t or two strings, use this customization type:

     (list file
           (choice (const t)
                   (list :inline t string string)))

If the user chooses the first alternative in the choice, then the overall list has two elements and the second element is t. If the user chooses the second alternative, then the overall list has three elements and the second and third must be strings.

13.4.4 Type Keywords

You can specify keyword-argument pairs in a customization type after the type name symbol. Here are the keywords you can use, and their meanings:

:value default This is used for a type that appears as an alternative inside of choice; it specifies the default value to use, at first, if and when the user selects this alternative with the menu in the customization buffer. Of course, if the actual value of the option fits this alternative, it will appear showing the actual value, not default. If nil is not a valid value for the alternative, then it is essential to specify a valid default with :value.

:format format-string This string will be inserted in the buffer to represent the value corresponding to the type. The following ‘%’ escapes are available for use in format-string:

‘%[button%]’ Display the text button marked as a button. The :action attribute specifies what the button will do if the user invokes it; its value is a function which takes two arguments—the widget which the button appears in, and the event. There is no way to specify two different buttons with different actions.

‘%{sample%}’ Show sample in a special face specified by :sample-face.

‘%v’ Substitute the item’s value. How the value is represented depends on the kind of item, and (for variables) on the customization type.

‘%d’ Substitute the item’s documentation string.

‘%h’ Like ‘%d’, but if the documentation string is more than one line, add an active field to control whether to show all of it or just the first line.

‘%t’ Substitute the tag here. You specify the tag with the :tag keyword.

‘%%’ Display a literal ‘%’.

:action action

Perform action if the user clicks on a button.

:button-face face

Use the face face (a face name or a list of face names) for button text displayed

with ‘%[...%]’.
:button-prefix prefix
:button-suffix suffix

These specify the text to display before and after a button. Each can be:

nil No text is inserted.

a string The string is inserted literally.

a symbol The symbol’s value is used.

:tag tag Use tag (a string) as the tag for the value (or part of the value) that corresponds to this type.

:doc doc Use doc as the documentation string for this value (or part of the value) that corresponds to this type. In order for this to work, you must specify a value for :format, and use ‘%d’ or ‘%h’ in that value. The usual reason to specify a documentation string for a type is to provide more information about the meanings of alternatives inside a :choice type or the parts of some other composite type.

:help-echo motion-doc

When you move to this item with widget-forward or widget-backward, it will display the string motion-doc in the echo area.

:match function Specify how to decide whether a value matches the type. The corresponding value, function, should be a function that accepts two arguments, a widget and a value; it should return non-nil if the value is acceptable.

14 Loading

Loading a file of Lisp code means bringing its contents into the Lisp environment in the form of Lisp objects. XEmacs finds and opens the file, reads the text, evaluates each form, and then closes the file.

The load functions evaluate all the expressions in a file just as the eval-current-buffer function evaluates all the expressions in a buffer. The difference is that the load functions read and evaluate the text in the file as found on disk, not the text in an Emacs buffer.

The loaded file must contain Lisp expressions, either as source code or as byte-compiled code. Each form in the file is called a top-level form. There is no special format for the forms in a loadable file; any form in a file may equally well be typed directly into a buffer and evaluated there. (Indeed, most code is tested this way.) Most often, the forms are function definitions and variable definitions.

A file containing Lisp code is often called a library. Thus, the “Rmail library” is a file containing code for Rmail mode. Similarly, a “Lisp library directory” is a directory of files containing Lisp code.

14.1 How Programs Do Loading

XEmacs Lisp has several interfaces for loading. For example, autoload creates a placeholder object for a function in a file; trying to call the autoloading function loads the file to get the function’s real definition (see Section 14.2 [Autoload], page 198). require loads a file if it isn’t already loaded (see Section 14.4 [Named Features], page 201). Ultimately, all these facilities call the load function to do the work.

Function load filename &optional missing-ok nomessage nosuffix This function finds and opens a file of Lisp code, evaluates all the forms in it, and closes the file. To find the file, load first looks for a file named ‘filename.elc’, that is, for a file whose name is filename with ‘.elc’ appended. If such a file exists, it is loaded. If there is no file by that name, then load looks for a file named ‘filename.el’. If that file exists, it is loaded. Finally, if neither of those names is found, load looks for a file named filename with nothing appended, and loads it if it exists. (The load function is not clever about looking at filename. In the perverse case of a file named ‘foo.el.el’, evaluation of (load "foo.el") will indeed find it.) If the optional argument nosuffix is non-nil, then the suffixes ‘.elc’ and ‘.el’ are not tried. In this case, you must specify the precise file name you want. If filename is a relative file name, such as ‘foo’ or ‘baz/foo.bar’, load searches for the file using the variable load-path. It appends filename to each of the directories listed in load-path, and loads the first file it finds whose name matches. The current default directory is tried only if it is specified in loadpath, where nil stands for the default directory. load tries all three possible suffixes in the first directory in load-path, then all three suffixes in the second directory, and so on.

If you get a warning that ‘foo.elc’ is older than ‘foo.el’, it means you should consider recompiling ‘foo.el’. See Chapter 15 [Byte Compilation], page 205.

Messages like ‘Loading foo...’ and ‘Loading foo...done’ appear in the echo area during loading unless nomessage is non-nil.

Any unhandled errors while loading a file terminate loading. If the load was done for the sake of autoload, any function definitions made during the loading are undone.

If load can’t find the file to load, then normally it signals the error file-error (with ‘Cannot open load file filename’). But if missing-ok is non-nil, then load just returns nil.

You can use the variable load-read-function to specify a function for load to use instead of read for reading expressions. See below.

load returns t if the file loads successfully.

User Option load-path

The value of this variable is a list of directories to search when loading files with load. Each element is a string (which must be a directory name) or nil (which stands for the current working directory). The value of loadpath is initialized from the environment variable EMACSLOADPATH, if that exists; otherwise its default value is specified in ‘emacs/src/paths.h’ when XEmacs is built.

The syntax of EMACSLOADPATH is the same as used for PATH; ‘:’ (or ‘;’, according to the operating system) separates directory names, and ‘.’ is used for the current default directory. Here is an example of how to set your EMACSLOADPATH variable from a csh ‘.login’ file:

     setenv EMACSLOADPATH .:/user/bil/emacs:/usr/lib/emacs/lisp

Here is how to set it using sh:

     export EMACSLOADPATH
     EMACSLOADPATH=.:/user/bil/emacs:/usr/local/lib/emacs/lisp

Here is an example of code you can place in a ‘.emacs’ file to add several directories to the front of your default load-path:

     (setq load-path
           (append (list nil "/user/bil/emacs"
                         "/usr/local/lisplib"
                         "~/emacs")
                   load-path))

In this example, the path searches the current working directory first, followed then by the ‘/user/bil/emacs’ directory, the ‘/usr/local/lisplib’ directory, and the ‘~/emacs’ directory, which are then followed by the standard directories for Lisp code.

The command line options ‘-l’ or ‘-load’ specify a Lisp library to load as part of Emacs startup. Since this file might be in the current directory, Emacs 18 temporarily adds the current directory to the front of load-path so the file can be found there. Newer Emacs versions also find such files in the current directory, but without altering load-path. Dumping Emacs uses a special value of load-path. If the value of load-path at the end of dumping is unchanged (that is, still the same special value), the dumped Emacs switches to the ordinary load-path value when it starts up, as described above. But if load-path has any other value at the end of dumping, that value is used for execution of the dumped Emacs also. Therefore, if you want to change load-path temporarily for loading a few li-

braries in ‘site-init.el’ or ‘site-load.el’, you should bind load-path lo-

cally with let around the calls to load.

Function locate-file filename path-list &optional suffixes mode This function searches for a file in the same way that load does, and returns the file found (if any). (In fact, load uses this function to search through load-path.) It searches for filename through path-list, expanded by one of the optional suffixes (string of suffixes separated by ‘:’s), checking for access mode (0|1|2|4 = exists|executable|writeable|readable), default readable. locate-file keeps hash tables of the directories it searches through, in order to speed things up. It tries valiantly to not get confused in the face of a changing and unpredictable environment, but can occasionally get tripped up. In this case, you will have to call locate-file-clear-hashing to get it back on track. See that function for details.

Function locate-file-clear-hashing path

This function clears the hash records for the specified list of directories. locatefile uses a hashing scheme to speed lookup, and will correctly track the following environmental changes:

• changes of any sort to the list of directories to be searched.

• addition and deletion of non-shadowing files (see below) from the directories in the list.

• byte-compilation of a .el file into a .elc file.

locate-file will primarily get confused if you add a file that shadows (i.e. has the same name as) another file further down in the directory list. In this case,

you must call locate-file-clear-hashing.

Variable load-in-progress

This variable is non-nil if Emacs is in the process of loading a file, and it is nil otherwise.

Variable load-read-function

This variable specifies an alternate expression-reading function for load and eval-region to use instead of read. The function should accept one argument, just as read does. Normally, the variable’s value is nil, which means those functions should use

read.

User Option load-warn-when-source-newer

This variable specifies whether load should check whether the source is newer than the binary. If this variable is true, then when a ‘.elc’ file is being loaded and the corresponding ‘.el’ is newer, a warning message will be printed. The default is nil, but it is bound to t during the initial loadup.

User Option load-warn-when-source-only

This variable specifies whether load should warn when loading a ‘.el’ file instead of an ‘.elc’. If this variable is true, then when load is called with a filename without an extension, and the ‘.elc’ version doesn’t exist but the ‘.el’ version does, then a message will be printed. If an explicit extension is passed to load, no warning will be printed. The default is nil, but it is bound to t during the initial loadup.

User Option load-ignore-elc-files

This variable specifies whether load should ignore ‘.elc’ files when a suffix is not given. This is normally used only to bootstrap the ‘.elc’ files when building XEmacs, when you use the command ‘make all-elc’. (This forces the ‘.el’ versions to be loaded in the process of compiling those same files, so that existing out-of-date ‘.elc’ files do not make it mess things up.)

To learn how load is used to build XEmacs, see Section B.1 [Building XEmacs], page 765.

14.2 Autoload

The autoload facility allows you to make a function or macro known in Lisp, but put off loading the file that defines it. The first call to the function automatically reads the proper file to install the real definition and other associated code, then runs the real definition as if it had been loaded all along.

There are two ways to set up an autoloaded function: by calling autoload, and by writing a special “magic” comment in the source before the real definition. autoload is the low-level primitive for autoloading; any Lisp program can call autoload at any time. Magic comments do nothing on their own; they serve as a guide for the command updatefile-autoloads, which constructs calls to autoload and arranges to execute them when Emacs is built. Magic comments are the most convenient way to make a function autoload, but only for packages installed along with Emacs.

Function autoload function filename &optional docstring interactive type This function defines the function (or macro) named function so as to load automatically from filename. The string filename specifies the file to load to get the real definition of function. The argument docstring is the documentation string for the function. Normally, this is the identical to the documentation string in the function definition itself. Specifying the documentation string in the call to autoload makes it possible to look at the documentation without loading the function’s real definition. If interactive is non-nil, then the function can be called interactively. This lets completion in M-x work without loading the function’s real definition. The complete interactive specification need not be given here; it’s not needed unless the user actually calls function, and when that happens, it’s time to load the real definition. You can autoload macros and keymaps as well as ordinary functions. Specify type as macro if function is really a macro. Specify type as keymap if function is really a keymap. Various parts of Emacs need to know this information without loading the real definition. An autoloaded keymap loads automatically during key lookup when a prefix key’s binding is the symbol function. Autoloading does not occur for other kinds of access to the keymap. In particular, it does not happen when a Lisp program gets the keymap from the value of a variable and calls define-key; not even if the variable name is the same symbol function. If function already has a non-void function definition that is not an autoload object, autoload does nothing and returns nil. If the function cell of function is void, or is already an autoload object, then it is defined as an autoload object like this:

(autoload filename docstring interactive type) For example,

     (symbol-function 'run-prolog)
          ⇒ (autoload "prolog" 169681 t nil)

In this case, "prolog" is the name of the file to load, 169681 refers to the documentation string in the ‘DOC’ file (see Section 27.1 [Documentation Basics], page 381), t means the function is interactive, and nil that it is not a macro or a keymap.

The autoloaded file usually contains other definitions and may require or provide one or more features. If the file is not completely loaded (due to an error in the evaluation of its contents), any function definitions or provide calls that occurred during the load are undone. This is to ensure that the next attempt to call any function autoloading from this file will try again to load the file. If not for this, then some of the functions in the file might appear defined, but they might fail to work properly for the lack of certain subroutines defined later in the file and not loaded successfully.

XEmacs as distributed comes with many autoloaded functions. The calls to autoload are in the file ‘loaddefs.el’. There is a convenient way of updating them automatically.

If the autoloaded file fails to define the desired Lisp function or macro, then an error is

signaled with data "Autoloading failed to define function function-name".

A magic autoload comment looks like ‘;;;###autoload’, on a line by itself, just before the real definition of the function in its autoloadable source file. The command M-x updatefile-autoloads writes a corresponding autoload call into ‘loaddefs.el’. Building Emacs

loads ‘loaddefs.el’ and thus calls autoload. M-x update-directory-autoloads is even

more powerful; it updates autoloads for all files in the current directory.

The same magic comment can copy any kind of form into ‘loaddefs.el’. If the form following the magic comment is not a function definition, it is copied verbatim. You can also use a magic comment to execute a form at build time without executing it when the file itself is loaded. To do this, write the form on the same line as the magic comment. Since it is in a comment, it does nothing when you load the source file; but update-file-autoloads copies it to ‘loaddefs.el’, where it is executed while building Emacs.

The following example shows how doctor is prepared for autoloading with a magic comment:

     ;;;###autoload
     (defun doctor ()
       "Switch to *doctor* buffer and start giving psychotherapy."
       (interactive)
       (switch-to-buffer "*doctor*")
       (doctor-mode))

Here’s what that produces in ‘loaddefs.el’:

     (autoload 'doctor "doctor"
       "\
     Switch to *doctor* buffer and start giving psychotherapy."
       t)

The backslash and newline immediately following the double-quote are a convention used only in the preloaded Lisp files such as ‘loaddefs.el’; they tell make-docfile to put the documentation string in the ‘DOC’ file. See Section B.1 [Building XEmacs], page 765.

14.3 Repeated Loading

You may load one file more than once in an Emacs session. For example, after you have rewritten and reinstalled a function definition by editing it in a buffer, you may wish to return to the original version; you can do this by reloading the file it came from.

When you load or reload files, bear in mind that the load and load-library functions automatically load a byte-compiled file rather than a non-compiled file of similar name. If you rewrite a file that you intend to save and reinstall, remember to byte-compile it if necessary; otherwise you may find yourself inadvertently reloading the older, byte-compiled file instead of your newer, non-compiled file!

When writing the forms in a Lisp library file, keep in mind that the file might be loaded more than once. For example, the choice of defvar vs. defconst for defining a variable depends on whether it is desirable to reinitialize the variable if the library is reloaded: defconst does so, and defvar does not. (See Section 10.5 [Defining Variables], page 149.)

The simplest way to add an element to an alist is like this:

     (setq minor-mode-alist
           (cons '(leif-mode " Leif") minor-mode-alist))

But this would add multiple elements if the library is reloaded. To avoid the problem, write this:

     (or (assq 'leif-mode minor-mode-alist)
         (setq minor-mode-alist
               (cons '(leif-mode " Leif") minor-mode-alist)))

To add an element to a list just once, use add-to-list (see Section 10.7 [Setting Variables], page 152).

Occasionally you will want to test explicitly whether a library has already been loaded. Here’s one way to test, in a library, whether it has been loaded before:

     (defvar foo-was-loaded)
     (if (not (boundp 'foo-was-loaded))

execute-first-time-only)

     (setq foo-was-loaded t)

If the library uses provide to provide a named feature, you can use featurep to test whether the library has been loaded.

14.4 Features

provide and require are an alternative to autoload for loading files automatically. They work in terms of named features. Autoloading is triggered by calling a specific function, but a feature is loaded the first time another program asks for it by name.

A feature name is a symbol that stands for a collection of functions, variables, etc. The file that defines them should provide the feature. Another program that uses them may ensure they are defined by requiring the feature. This loads the file of definitions if it hasn’t been loaded already.

To require the presence of a feature, call require with the feature name as argument. require looks in the global variable features to see whether the desired feature has been provided already. If not, it loads the feature from the appropriate file. This file should call provide at the top level to add the feature to features; if it fails to do so, require signals an error.

Features are normally named after the files that provide them, so that require need not be given the file name.

For example, in ‘emacs/lisp/prolog.el’, the definition for run-prolog includes the following code:

     (defun run-prolog ()
       "Run an inferior Prolog process, input and output via buffer *prolog*."
       (interactive)
       (require 'comint)
       (switch-to-buffer (make-comint "prolog" prolog-program-name))
       (inferior-prolog-mode))

The expression (require 'comint) loads the file ‘comint.el’ if it has not yet been loaded. This ensures that make-comint is defined.

The ‘comint.el’ file contains the following top-level expression:

     (provide 'comint)

This adds comint to the global features list, so that (require 'comint) will henceforth know that nothing needs to be done.

When require is used at top level in a file, it takes effect when you byte-compile that file (see Chapter 15 [Byte Compilation], page 205) as well as when you load it. This is in case the required package contains macros that the byte compiler must know about.

Although top-level calls to require are evaluated during byte compilation, provide calls are not. Therefore, you can ensure that a file of definitions is loaded before it is bytecompiled by including a provide followed by a require for the same feature, as in the following example.

(provide 'my-feature) ; Ignored by byte compiler,

; evaluated by load.

(require 'my-feature) ; Evaluated by byte compiler. The compiler ignores the provide, then processes the require by loading the file in question. Loading the file does execute the provide call, so the subsequent require call does nothing while loading.

Function provide feature

This function announces that feature is now loaded, or being loaded, into the current XEmacs session. This means that the facilities associated with feature are or will be available for other Lisp programs. The direct effect of calling provide is to add feature to the front of the list features if it is not already in the list. The argument feature must be a symbol. provide returns feature.

     features
          ⇒ (bar bish)
     (provide 'foo)
          ⇒ foo
     features
          ⇒ (foo bar bish)

When a file is loaded to satisfy an autoload, and it stops due to an error in the evaluating its contents, any function definitions or provide calls that occurred during the load are undone. See Section 14.2 [Autoload], page 198.

Function require feature &optional filename This function checks whether feature is present in the current XEmacs session (using (featurep feature); see below). If it is not, then require loads filename with load. If filename is not supplied, then the name of the symbol feature is used as the file name to load. If loading the file fails to provide feature, require signals an error, ‘Required

feature feature was not provided’.

Function featurep fexp

This function returns t if feature fexp is present in this Emacs. Use this to conditionalize execution of lisp code based on the presence or absence of emacs or environment extensions. fexp can be a symbol, a number, or a list. If fexp is a symbol, it is looked up in the ‘features’ variable, and t is returned if it is found, nil otherwise. If fexp is a number, the function returns t if this Emacs has an equal or greater number than fexp, nil otherwise. Note that minor Emacs version is expected to be 2 decimal places wide, so (featurep 20.4) will return nil on XEmacs 20.4—you must write (featurep 20.04), unless you wish to match for XEmacs 20.40. If fexp is a list whose car is the symbol and, the function returns t if all the features in its cdr are present, nil otherwise. If fexp is a list whose car is the symbol or, the function returns t if any the features in its cdr are present, nil otherwise. If fexp is a list whose car is the symbol not, the function returns t if the feature is not present, nil otherwise. Examples:

     (featurep 'xemacs)

⇒ ; t on XEmacs.

     (featurep '(and xemacs gnus))

⇒ ; t on XEmacs with Gnus loaded.

     (featurep '(or tty-frames (and emacs 19.30)))

⇒ ; t if this Emacs supports TTY frames.

     (featurep '(or (and xemacs 19.15) (and emacs 19.34)))

⇒ ; t on XEmacs 19.15 and later, or on

; FSF Emacs 19.34 and later.

Please note: The advanced arguments of this function (anything other than a symbol) are not yet supported by FSF Emacs. If you feel they are useful for supporting multiple Emacs variants, lobby Richard Stallman at ‘<bug-gnu-emacs@prep.ai.mit.edu>’.

Variable features

The value of this variable is a list of symbols that are the features loaded in the current XEmacs session. Each symbol was put in this list with a call to provide. The order of the elements in the features list is not significant.

14.5 Unloading

You can discard the functions and variables loaded by a library to reclaim memory for other Lisp objects. To do this, use the function unload-feature:

Command unload-feature feature &optional force This command unloads the library that provided feature feature. It undefines all functions, macros, and variables defined in that library with defconst, defvar,

defun, defmacro, defsubst, definf-function and defalias. It then restores

any autoloads formerly associated with those symbols. (Loading saves these in the autoload property of the symbol.) Ordinarily, unload-feature refuses to unload a library on which other loaded libraries depend. (A library a depends on library b if a contains a require for b.) If the optional argument force is non-nil, dependencies are ignored and you can unload any library.

The unload-feature function is written in Lisp; its actions are based on the variable

load-history.

Variable load-history

This variable’s value is an alist connecting library names with the names of functions and variables they define, the features they provide, and the features they require. Each element is a list and describes one library. The car of the list is the name of the library, as a string. The rest of the list is composed of these kinds of objects:

• Symbols that were defined by this library.

• Lists of the form (require . feature) indicating features that were required.

• Lists of the form (provide . feature) indicating features that were provided.

The value of load-history may have one element whose car is nil. This element describes definitions made with eval-buffer on a buffer that is not visiting a file.

The command eval-region updates load-history, but does so by adding the symbols defined to the element for the file being visited, rather than replacing that element.

14.6 Hooks for Loading

Variable after-load-alist

An alist of expressions to evaluate if and when particular libraries are loaded. Each element looks like this:

(filename forms...) When load is run and the file-name argument is filename, the forms in the corresponding element are executed at the end of loading. filename must match exactly! Normally filename is the name of a library, with no directory specified, since that is how load is normally called. An error in forms does not undo the load, but does prevent execution of the rest of the forms.

15 Byte Compilation

XEmacs Lisp has a compiler that translates functions written in Lisp into a special representation called byte-code that can be executed more efficiently. The compiler replaces Lisp function definitions with byte-code. When a byte-coded function is called, its definition is evaluated by the byte-code interpreter.

Because the byte-compiled code is evaluated by the byte-code interpreter, instead of being executed directly by the machine’s hardware (as true compiled code is), byte-code is completely transportable from machine to machine without recompilation. It is not, however, as fast as true compiled code.

In general, any version of Emacs can run byte-compiled code produced by recent earlier versions of Emacs, but the reverse is not true. In particular, if you compile a program with XEmacs 20, the compiled code may not run in earlier versions. See Section 15.3 [Docs and Compilation], page 208.

See Section 16.3 [Compilation Errors], page 227, for how to investigate errors occurring in byte compilation.

15.1 Performance of Byte-Compiled Code

A byte-compiled function is not as efficient as a primitive function written in C, but runs much faster than the version written in Lisp. Here is an example:

     (defun silly-loop (n)
       "Return time before and after N iterations of a loop."
       (let ((t1 (current-time-string)))
         (while (> (setq n (1- n))
                   0))
         (list t1 (current-time-string))))
     ⇒ silly-loop
     (silly-loop 5000000)
     ⇒ ("Fri Nov 28 20:56:16 1997"
         "Fri Nov 28 20:56:39 1997")  ; 23 seconds
     (byte-compile 'silly-loop)
     ⇒ #<compiled-function
     (from "loadup.el")
     (n)
     "...(23)"
     [current-time-string t1 n 0]
     2
     "Return time before and after N iterations of a loop.">
     (silly-loop 5000000)
     ⇒ ("Fri Nov 28 20:57:49 1997"
         "Fri Nov 28 20:57:55 1997")  ; 6 seconds

In this example, the interpreted code required 23 seconds to run, whereas the bytecompiled code required 6 seconds. These results are representative, but actual results will vary greatly.

15.2 The Compilation Functions

You can byte-compile an individual function or macro definition with the byte-compile function. You can compile a whole file with byte-compile-file, or several files with

byte-recompile-directory or batch-byte-compile.

When you run the byte compiler, you may get warnings in a buffer called ‘*Compile-Log*’. These report things in your program that suggest a problem but are not necessarily erroneous.

Be careful when byte-compiling code that uses macros. Macro calls are expanded when they are compiled, so the macros must already be defined for proper compilation. For more details, see Section 12.3 [Compiling Macros], page 178.

Normally, compiling a file does not evaluate the file’s contents or load the file. But it does execute any require calls at top level in the file. One way to ensure that necessary macro definitions are available during compilation is to require the file that defines them (see Section 14.4 [Named Features], page 201). To avoid loading the macro definition files when someone runs the compiled program, write eval-when-compile around the require calls (see Section 15.5 [Eval During Compile], page 209).

Function byte-compile symbol

This function byte-compiles the function definition of symbol, replacing the previous definition with the compiled one. The function definition of symbol must be the actual code for the function; i.e., the compiler does not follow indirection to another symbol. byte-compile returns the new, compiled definition of symbol. If symbol’s definition is a compiled-function object, byte-compile does nothing and returns nil. Lisp records only one function definition for any symbol, and if that is already compiled, non-compiled code is not available anywhere. So there is no way to “compile the same definition again.”

     (defun factorial (integer)
       "Compute factorial of INTEGER."
       (if (= 1 integer) 1
         (* integer (factorial (1- integer)))))
     ⇒ factorial
     (byte-compile 'factorial)
     ⇒ #<compiled-function
     (from "loadup.el")
     (integer)
     "...(21)"
     [integer 1 factorial]
     3
     "Compute factorial of INTEGER.">

The result is a compiled-function object. The string it contains is the actual byte-code; each character in it is an instruction or an operand of an instruction. The vector contains all the constants, variable names and function names used by the function, except for certain primitives that are coded as special instructions.

Command compile-defun &optional arg

This command reads the defun containing point, compiles it, and evaluates the result. If you use this on a defun that is actually a function definition, the effect is to install a compiled version of that function. If arg is non-nil, the result is inserted in the current buffer after the form; otherwise, it is printed in the minibuffer.

Command byte-compile-file filename &optional load This function compiles a file of Lisp code named filename into a file of byte-code. The output file’s name is made by appending ‘c’ to the end of filename. If load is non-nil, the file is loaded after having been compiled. Compilation works by reading the input file one form at a time. If it is a definition of a function or macro, the compiled function or macro definition is written out. Other forms are batched together, then each batch is compiled, and written so that its compiled code will be executed when the file is read. All comments are discarded when the input file is read. This command returns t. When called interactively, it prompts for the file name.

     % ls -l push*
     -rw-r--r--  1 lewis     791 Oct  5 20:31 push.el
     (byte-compile-file "~/emacs/push.el")
          ⇒ t
     % ls -l push*
     -rw-r--r--  1 lewis     791 Oct  5 20:31 push.el
     -rw-rw-rw-  1 lewis     638 Oct  8 20:25 push.elc

Command byte-recompile-directory directory &optional flag

This function recompiles every ‘.el’ file in directory that needs recompilation. A file needs recompilation if a ‘.elc’ file exists but is older than the ‘.el’ file. When a ‘.el’ file has no corresponding ‘.elc’ file, then flag says what to do. If it is nil, these files are ignored. If it is non-nil, the user is asked whether to compile each such file. The returned value of this command is unpredictable.

Function batch-byte-compile

This function runs byte-compile-file on files specified on the command line. This function must be used only in a batch execution of Emacs, as it kills Emacs on completion. An error in one file does not prevent processing of subsequent files. (The file that gets the error will not, of course, produce any compiled code.)

     % emacs -batch -f batch-byte-compile *.el

Function batch-byte-recompile-directory

This function is similar to batch-byte-compile but runs the command byterecompile-directory on the files remaining on the command line.

Variable byte-recompile-directory-ignore-errors-p

If non-nil, this specifies that byte-recompile-directory will continue compiling even when an error occurs in a file. This is normally nil, but is bound

     to t by batch-byte-recompile-directory.

Function byte-code code-string data-vector max-stack This function actually interprets byte-code. A byte-compiled function is actually defined with a body that calls byte-code. Don’t call this function yourself. Only the byte compiler knows how to generate valid calls to this function. In newer Emacs versions (19 and up), byte-code is usually executed as part of a compiled-function object, and only rarely due to an explicit call to byte-code.

15.3 Documentation Strings and Compilation

Functions and variables loaded from a byte-compiled file access their documentation strings dynamically from the file whenever needed. This saves space within Emacs, and makes loading faster because the documentation strings themselves need not be processed while loading the file. Actual access to the documentation strings becomes slower as a result, but normally not enough to bother users.

Dynamic access to documentation strings does have drawbacks:

• If you delete or move the compiled file after loading it, Emacs can no longer access the documentation strings for the functions and variables in the file.

• If you alter the compiled file (such as by compiling a new version), then further access to documentation strings in this file will give nonsense results.

If your site installs Emacs following the usual procedures, these problems will never normally occur. Installing a new version uses a new directory with a different name; as long as the old version remains installed, its files will remain unmodified in the places where they are expected to be.

However, if you have built Emacs yourself and use it from the directory where you built it, you will experience this problem occasionally if you edit and recompile Lisp files. When it happens, you can cure the problem by reloading the file after recompiling it.

Byte-compiled files made with Emacs 19.29 will not load into older versions because the older versions don’t support this feature. You can turn off this feature by setting bytecompile-dynamic-docstrings to nil. Once this is done, you can compile files that will load into older Emacs versions. You can do this globally, or for one source file by specifying a file-local binding for the variable. Here’s one way to do that:

     -*-byte-compile-dynamic-docstrings: nil;-*-

Variable byte-compile-dynamic-docstrings

If this is non-nil, the byte compiler generates compiled files that are set up for dynamic loading of documentation strings.

The dynamic documentation string feature writes compiled files that use a special Lisp reader construct, ‘#@count’. This construct skips the next count characters. It also uses the ‘#$’ construct, which stands for “the name of this file, as a string.” It is best not to use these constructs in Lisp source files.

15.4 Dynamic Loading of Individual Functions

When you compile a file, you can optionally enable the dynamic function loading feature (also known as lazy loading). With dynamic function loading, loading the file doesn’t fully read the function definitions in the file. Instead, each function definition contains a placeholder which refers to the file. The first time each function is called, it reads the full definition from the file, to replace the place-holder.

The advantage of dynamic function loading is that loading the file becomes much faster. This is a good thing for a file which contains many separate commands, provided that using one of them does not imply you will soon (or ever) use the rest. A specialized mode which provides many keyboard commands often has that usage pattern: a user may invoke the mode, but use only a few of the commands it provides.

The dynamic loading feature has certain disadvantages:

• If you delete or move the compiled file after loading it, Emacs can no longer load the remaining function definitions not already loaded.

• If you alter the compiled file (such as by compiling a new version), then trying to load any function not already loaded will get nonsense results.

If you compile a new version of the file, the best thing to do is immediately load the new compiled file. That will prevent any future problems.

The byte compiler uses the dynamic function loading feature if the variable bytecompile-dynamic is non-nil at compilation time. Do not set this variable globally, since dynamic loading is desirable only for certain files. Instead, enable the feature for specific source files with file-local variable bindings, like this:

     -*-byte-compile-dynamic: t;-*-

Variable byte-compile-dynamic

If this is non-nil, the byte compiler generates compiled files that are set up for dynamic function loading.

Function fetch-bytecode function

This immediately finishes loading the definition of function from its bytecompiled file, if it is not fully loaded already. The argument function may be a compiled-function object or a function name.

15.5 Evaluation During Compilation

These features permit you to write code to be evaluated during compilation of a program.

Special Form eval-and-compile body

This form marks body to be evaluated both when you compile the containing code and when you run it (whether compiled or not).

You can get a similar result by putting body in a separate file and referring to that file with require. Using require is preferable if there is a substantial amount of code to be executed in this way.

Special Form eval-when-compile body

This form marks body to be evaluated at compile time and not when the compiled program is loaded. The result of evaluation by the compiler becomes a constant which appears in the compiled program. When the program is interpreted, not compiled at all, body is evaluated normally.

At top level, this is analogous to the Common Lisp idiom (eval-when (compile eval) ...). Elsewhere, the Common Lisp ‘#.’ reader macro (but not when interpreting) is closer to what eval-when-compile does.

15.6 Compiled-Function Objects

Byte-compiled functions have a special data type: they are compiled-function objects.

A compiled-function object is a bit like a vector; however, the evaluator handles this data type specially when it appears as a function to be called. The printed representation for a compiled-function object normally begins with ‘#<compiled-function’ and ends with ‘>’. However, if the variable print-readably is non-nil, the object is printed beginning with ‘#[’ and ending with ‘]’. This representation can be read directly by the Lisp reader, and is used in byte-compiled files (those ending in ‘.elc’).

In Emacs version 18, there was no compiled-function object data type; compiled functions used the function byte-code to run the byte code.

A compiled-function object has a number of different elements. They are:

arglist The list of argument symbols.

instructions The string containing the byte-code instructions.

constants The vector of Lisp objects referenced by the byte code. These include symbols used as function names and variable names.

stacksize The maximum stack size this function needs.

doc-string The documentation string (if any); otherwise, nil. The value may be a number or a list, in case the documentation string is stored in a file. Use the function documentation to get the real documentation string (see Section 27.2 [Accessing Documentation], page 382).

interactive The interactive spec (if any). This can be a string or a Lisp expression. It is nil for a function that isn’t interactive.

domain The domain (if any). This is only meaningful if I18N3 (message-translation) support was compiled into XEmacs. This is a string defining which domain to find the translation for the documentation string and interactive prompt. See Section 54.2.4 [Domain Specification], page 730.

Here’s an example of a compiled-function object, in printed representation. It is the definition of the command backward-sexp.

#<compiled-function
(from "lisp.elc")
(&optional arg)
"...(15)" [arg 1 forward-sexp] 2 854740 "_p">

The primitive way to create a compiled-function object is with make-byte-code:

Function make-byte-code &rest elements

This function constructs and returns a compiled-function object with elements as its elements. Please note: Unlike all other Emacs-lisp functions, calling this with five arguments is not the same as calling it with six arguments, the last of which is nil. If the interactive arg is specified as nil, then that means that this function was defined with (interactive). If the arg is not specified, then that means the function is not interactive. This is terrible behavior which is retained for compatibility with old ‘.elc’ files which expected these semantics.

You should not try to come up with the elements for a compiled-function object yourself, because if they are inconsistent, XEmacs may crash when you call the function. Always leave it to the byte compiler to create these objects; it makes the elements consistent (we hope).

The following primitives are provided for accessing the elements of a compiled-function object.

Function compiled-function-arglist function

This function returns the argument list of compiled-function object function.

Function compiled-function-instructions function

This function returns a string describing the byte-code instructions of compiledfunction object function.

Function compiled-function-constants function

This function returns the vector of Lisp objects referenced by compiled-function object function.

Function compiled-function-stack-size function

This function returns the maximum stack size needed by compiled-function object function.

Function compiled-function-doc-string function

This function returns the doc string of compiled-function object function, if available.

Function compiled-function-interactive function

This function returns the interactive spec of compiled-function object function, if any. The return value is nil or a two-element list, the first element of which is the symbol interactive and the second element is the interactive spec (a string or Lisp form).

Function compiled-function-domain function

This function returns the domain of compiled-function object function, if any. The result will be a string or nil. See Section 54.2.4 [Domain Specification], page 730.

15.7 Disassembled Byte-Code

People do not write byte-code; that job is left to the byte compiler. But we provide a disassembler to satisfy a cat-like curiosity. The disassembler converts the byte-compiled code into humanly readable form.

The byte-code interpreter is implemented as a simple stack machine. It pushes values onto a stack of its own, then pops them off to use them in calculations whose results are themselves pushed back on the stack. When a byte-code function returns, it pops a value off the stack and returns it as the value of the function.

In addition to the stack, byte-code functions can use, bind, and set ordinary Lisp variables, by transferring values between variables and the stack.

Command disassemble object &optional stream This function prints the disassembled code for object. If stream is supplied, then output goes there. Otherwise, the disassembled code is printed to the stream standard-output. The argument object can be a function name or a lambda expression.

As a special exception, if this function is used interactively, it outputs to a

     buffer named ‘*Disassemble*’.

Here are two examples of using the disassemble function. We have added explanatory comments to help you relate the byte-code to the Lisp source; these do not appear in the output of disassemble. These examples show unoptimized byte-code. Nowadays byte-code is usually optimized, but we did not want to rewrite these examples, since they still serve their purpose.

     (defun factorial (integer)
       "Compute factorial of an integer."
       (if (= 1 integer) 1
         (* integer (factorial (1- integer)))))
          ⇒ factorial
     (factorial 4)
          ⇒ 24
     (disassemble 'factorial)
          ⊣ byte-code for factorial:
      doc: Compute factorial of an integer.
      args: (integer)

0 constant 1 ; Push 1 onto stack.

1   varref   integer        ; Get value of integer

; from the environment ; and push the value ; onto the stack.

2 eqlsign ; Pop top two values off stack, ; compare them, ; and push result onto stack.

3 goto-if-nil 10 ; Pop and test top of stack; ; if nil, go to 10, ; else continue.

6 constant 1 ; Push 1 onto top of stack.

7 goto 17 ; Go to 17 (in this case, 1 will be ; returned by the function).

10 constant * ; Push symbol * onto stack.

11  varref   integer        ; Push value of integer onto stack.
12  constant factorial      ; Push factorial onto stack.
13  varref   integer        ; Push value of integer onto stack.

14 sub1 ; Pop integer, decrement value, ; push new value onto stack.

; Stack now contains: ; − decremented value of integer

; − factorial
; − value of integer
; − *
15  call     1              ; Call function factorial using

; the first (i.e., the top) element ; of the stack as the argument; ; push returned value onto stack.

; Stack now contains: ; − result of recursive

;  call to factorial
; − value of integer
; − *

16 call 2 ; Using the first two ; (i.e., the top two) ; elements of the stack ; as arguments, ; call the function *, ; pushing the result onto the stack.

17 return ; Return the top element

; of the stack.

⇒ nil

The silly-loop function is somewhat more complex:

(defun silly-loop (n)
"Return time before and after N iterations of a loop."
(let ((t1 (current-time-string)))
(while (> (setq n (1- n))
0))
(list t1 (current-time-string))))
⇒ silly-loop
(disassemble 'silly-loop)
⊣ byte-code for silly-loop:
doc: Return time before and after N iterations of a loop.
args: (n)
0   constant current-time-string  ; Push
      ;   current-time-string

; onto top of stack.

1   call     0              ; Call current-time-string

; with no argument, ; pushing result onto stack.

2 varbind t1 ; Pop stack and bind t1 ; to popped value.

3 varref n ; Get value of n from ; the environment and push ; the value onto the stack.

4 sub1 ; Subtract 1 from top of stack.

5 dup ; Duplicate the top of the stack; ; i.e., copy the top of ; the stack and push the ; copy onto the stack.

6 varset n ; Pop the top of the stack, ; and bind n to the value.

; In effect, the sequence dup varset ; copies the top of the stack ; into the value of n ; without popping it.

7 constant 0 ; Push 0 onto stack.

8 gtr ; Pop top two values off stack, ; test if n is greater than 0 ; and push result onto stack.

9   goto-if-nil-else-pop 17 ; Goto 17 if n <= 0

; (this exits the while loop). ; else pop top of stack ; and continue

12  constant nil            ; Push nil onto stack

; (this is the body of the loop).

13 discard ; Discard result of the body ; of the loop (a while loop ; is always evaluated for ; its side effects).

14 goto 3 ; Jump back to beginning ; of while loop.

17 discard ; Discard result of while loop ; by popping top of stack. ; This result is the value nil that ; was not popped by the goto at 9.

18 varref t1 ; Push value of t1 onto stack.

19  constant current-time-string  ; Push
      ;   current-time-string

; onto top of stack.

20  call     0              ; Call current-time-string again.

21 list2 ; Pop top two elements off stack, ; create a list of them, ; and push list onto stack.

22 unbind 1 ; Unbind t1 in local environment.

23 return ; Return value of the top of stack.

216 XEmacs Lisp Reference Manual

16 Debugging Lisp Programs

There are three ways to investigate a problem in an XEmacs Lisp program, depending on what you are doing with the program when the problem appears.

• If the problem occurs when you run the program, you can use a Lisp debugger (either the default debugger or Edebug) to investigate what is happening during execution.

• If the problem is syntactic, so that Lisp cannot even read the program, you can use the XEmacs facilities for editing Lisp to localize it.

• If the problem occurs when trying to compile the program with the byte compiler, you need to know how to examine the compiler’s input buffer.

Another useful debugging tool is the dribble file. When a dribble file is open, XEmacs copies all keyboard input characters to that file. Afterward, you can examine the file to find out what input was used. See Section 50.8 [Terminal Input], page 706.

For debugging problems in terminal descriptions, the open-termscript function can be useful. See Section 50.9 [Terminal Output], page 709.

16.1 The Lisp Debugger

The Lisp debugger provides the ability to suspend evaluation of a form. While evaluation is suspended (a state that is commonly known as a break), you may examine the run time stack, examine the values of local or global variables, or change those values. Since a break is a recursive edit, all the usual editing facilities of XEmacs are available; you can even run programs that will enter the debugger recursively. See Section 19.10 [Recursive Editing], page 310.

16.1.1 Entering the Debugger on an Error

The most important time to enter the debugger is when a Lisp error happens. This allows you to investigate the immediate causes of the error.

However, entry to the debugger is not a normal consequence of an error. Many commands frequently get Lisp errors when invoked in inappropriate contexts (such as C-f at the end of the buffer) and during ordinary editing it would be very unpleasant to enter the debugger each time this happens. If you want errors to enter the debugger, set the variable debug-

on-error to non-nil.

User Option debug-on-error

This variable determines whether the debugger is called when an error is signaled and not handled. If debug-on-error is t, all errors call the debugger. If it is nil, none call the debugger. The value can also be a list of error conditions that should call the debugger. For example, if you set it to the list (void-variable), then only errors about a variable that has no value invoke the debugger. When this variable is non-nil, Emacs does not catch errors that happen in process filter functions and sentinels. Therefore, these errors also can invoke the debugger. See Chapter 49 [Processes], page 673.

User Option debug-ignored-errors

This variable specifies certain kinds of errors that should not enter the debugger. Its value is a list of error condition symbols and/or regular expressions. If the error has any of those condition symbols, or if the error message matches any of the regular expressions, then that error does not enter the debugger, regardless

     of the value of debug-on-error.

The normal value of this variable lists several errors that happen often during editing but rarely result from bugs in Lisp programs.

To debug an error that happens during loading of the ‘.emacs’ file, use the option ‘-debug-init’, which binds debug-on-error to t while ‘.emacs’ is loaded and inhibits use of condition-case to catch init file errors.

If your ‘.emacs’ file sets debug-on-error, the effect may not last past the end of loading ‘.emacs’. (This is an undesirable byproduct of the code that implements the ‘-debug-init’ command line option.) The best way to make ‘.emacs’ set debug-on-error permanently is with after-init-hook, like this:

     (add-hook 'after-init-hook
               '(lambda () (setq debug-on-error t)))

User Option debug-on-signal

This variable is similar to debug-on-error but breaks whenever an error is signalled, regardless of whether it would be handled.

16.1.2 Debugging Infinite Loops

When a program loops infinitely and fails to return, your first problem is to stop the loop. On most operating systems, you can do this with C-g, which causes quit.

Ordinary quitting gives no information about why the program was looping. To get more information, you can set the variable debug-on-quit to non-nil. Quitting with C-g is not considered an error, and debug-on-error has no effect on the handling of C-g. Likewise, debug-on-quit has no effect on errors.

Once you have the debugger running in the middle of the infinite loop, you can proceed from the debugger using the stepping commands. If you step through the entire loop, you will probably get enough information to solve the problem.

User Option debug-on-quit

This variable determines whether the debugger is called when quit is signaled and not handled. If debug-on-quit is non-nil, then the debugger is called whenever you quit (that is, type C-g). If debug-on-quit is nil, then the debugger is not called when you quit. See Section 19.8 [Quitting], page 306.

16.1.3 Entering the Debugger on a Function Call

To investigate a problem that happens in the middle of a program, one useful technique is to enter the debugger whenever a certain function is called. You can do this to the function in which the problem occurs, and then step through the function, or you can do this to a function called shortly before the problem, step quickly over the call to that function, and then step through its caller.

Command debug-on-entry function-name

This function requests function-name to invoke the debugger each time it is called. It works by inserting the form (debug 'debug) into the function definition as the first form. Any function defined as Lisp code may be set to break on entry, regardless of whether it is interpreted code or compiled code. If the function is a command, it will enter the debugger when called from Lisp and when called interactively (after the reading of the arguments). You can’t debug primitive functions (i.e., those written in C) this way. When debug-on-entry is called interactively, it prompts for function-name in the minibuffer. If the function is already set up to invoke the debugger on entry, debug-on-entry does nothing. Please note: if you redefine a function after using debug-on-entry on it, the code to enter the debugger is lost. debug-on-entry returns function-name.

     (defun fact (n)
       (if (zerop n) 1
           (* n (fact (1- n)))))
          ⇒ fact
     (debug-on-entry 'fact)
          ⇒ fact
     (fact 3)
     ------ Buffer: *Backtrace* ------
     Entering:
     * fact(3)
       eval-region(4870 4878 t)
       byte-code("...")
       eval-last-sexp(nil)
       (let ...)
       eval-insert-last-sexp(nil)
     * call-interactively(eval-insert-last-sexp)
     ------ Buffer: *Backtrace* ------
     (symbol-function 'fact)
          ⇒ (lambda (n)
               (debug (quote debug))
               (if (zerop n) 1 (* n (fact (1- n)))))

Command cancel-debug-on-entry function-name

This function undoes the effect of debug-on-entry on function-name. When called interactively, it prompts for function-name in the minibuffer. If functionname is nil or the empty string, it cancels debugging for all functions. If cancel-debug-on-entry is called more than once on the same function, the second call does nothing. cancel-debug-on-entry returns function-name.

16.1.4 Explicit Entry to the Debugger

You can cause the debugger to be called at a certain point in your program by writing the expression (debug) at that point. To do this, visit the source file, insert the text ‘(debug)’ at the proper place, and type C-M-x. Be sure to undo this insertion before you save the file!

The place where you insert ‘(debug)’ must be a place where an additional form can be evaluated and its value ignored. (If the value of (debug) isn’t ignored, it will alter the execution of the program!) The most common suitable places are inside a progn or an implicit progn (see Section 9.1 [Sequencing], page 129).

16.1.5 Using the Debugger

When the debugger is entered, it displays the previously selected buffer in one window and a buffer named ‘*Backtrace*’ in another window. The backtrace buffer contains one line for each level of Lisp function execution currently going on. At the beginning of this buffer is a message describing the reason that the debugger was invoked (such as the error message and associated data, if it was invoked due to an error).

The backtrace buffer is read-only and uses a special major mode, Debugger mode, in which letters are defined as debugger commands. The usual XEmacs editing commands are available; thus, you can switch windows to examine the buffer that was being edited at the time of the error, switch buffers, visit files, or do any other sort of editing. However, the debugger is a recursive editing level (see Section 19.10 [Recursive Editing], page 310) and it is wise to go back to the backtrace buffer and exit the debugger (with the q command) when you are finished with it. Exiting the debugger gets out of the recursive edit and kills the backtrace buffer.

The backtrace buffer shows you the functions that are executing and their argument values. It also allows you to specify a stack frame by moving point to the line describing that frame. (A stack frame is the place where the Lisp interpreter records information about a particular invocation of a function.) The frame whose line point is on is considered the current frame. Some of the debugger commands operate on the current frame.

The debugger itself must be run byte-compiled, since it makes assumptions about how many stack frames are used for the debugger itself. These assumptions are false if the debugger is running interpreted.

16.1.6 Debugger Commands

Inside the debugger (in Debugger mode), these special commands are available in addition to the usual cursor motion commands. (Keep in mind that all the usual facilities of XEmacs, such as switching windows or buffers, are still available.)

The most important use of debugger commands is for stepping through code, so that you can see how control flows. The debugger can step through the control structures of an interpreted function, but cannot do so in a byte-compiled function. If you would like to step through a byte-compiled function, replace it with an interpreted definition of the same function. (To do this, visit the source file for the function and type C-M-x on its definition.)

Here is a list of Debugger mode commands:

c Exit the debugger and continue execution. This resumes execution of the program as if the debugger had never been entered (aside from the effect of any variables or data structures you may have changed while inside the debugger). Continuing when an error or quit was signalled will cause the normal action of the signalling to take place. If you do not want this to happen, but instead want the program execution to continue as if the call to signal did not occur, use the r command.

d Continue execution, but enter the debugger the next time any Lisp function is called. This allows you to step through the subexpressions of an expression, seeing what values the subexpressions compute, and what else they do. The stack frame made for the function call which enters the debugger in this way will be flagged automatically so that the debugger will be called again when the frame is exited. You can use the u command to cancel this flag.

b Flag the current frame so that the debugger will be entered when the frame is exited. Frames flagged in this way are marked with stars in the backtrace buffer.

u Don’t enter the debugger when the current frame is exited. This cancels a b command on that frame.

e Read a Lisp expression in the minibuffer, evaluate it, and print the value in the echo area. The debugger alters certain important variables, and the current buffer, as part of its operation; e temporarily restores their outside-the-debugger values so you can examine them. This makes the debugger more transparent. By contrast, M-: does nothing special in the debugger; it shows you the variable values within the debugger.

q Terminate the program being debugged; return to top-level XEmacs command execution. If the debugger was entered due to a C-g but you really want to quit, and not debug, use the q command.

r Return a value from the debugger. The value is computed by reading an expression with the minibuffer and evaluating it.

The r command is useful when the debugger was invoked due to exit from a Lisp call frame (as requested with b); then the value specified in the r command is used as the value of that frame. It is also useful if you call debug and use its return value. If the debugger was entered at the beginning of a function call, r has the same effect as c, and the specified return value does not matter. If the debugger was entered through a call to signal (i.e. as a result of an error or quit), then returning a value will cause the call to signal itself to return, rather than throwing to top-level or invoking a handler, as is normal. This allows you to correct an error (e.g. the type of an argument was wrong) or continue from a debug-on-quit as if it never happened. Note that some errors (e.g. any error signalled using the error function, and many errors signalled from a primitive function) are not continuable. If you return a value from them and continue execution, then the error will immediately be signalled again. Other errors (e.g. wrong-type-argument errors) will be continually resignalled until the problem is corrected.

16.1.7 Invoking the Debugger

Here we describe fully the function used to invoke the debugger.

Function debug &rest debugger-args This function enters the debugger. It switches buffers to a buffer named ‘*Backtrace*’ (or ‘*Backtrace*<2>’ if it is the second recursive entry to the debugger, etc.), and fills it with information about the stack of Lisp function calls. It then enters a recursive edit, showing the backtrace buffer in Debugger mode. The Debugger mode c and r commands exit the recursive edit; then debug switches back to the previous buffer and returns to whatever called debug. This is the only way the function debug can return to its caller. If the first of the debugger-args passed to debug is nil (or if it is not one of the special values in the table below), then debug displays the rest of its arguments at the top of the ‘*Backtrace*’ buffer. This mechanism is used to display a message to the user. However, if the first argument passed to debug is one of the following special values, then it has special significance. Normally, these values are passed to debug only by the internals of XEmacs and the debugger, and not by programmers calling debug. The special values are:

lambda A first argument of lambda means debug was called because of entry to a function when debug-on-next-call was non-nil. The debugger displays ‘Entering:’ as a line of text at the top of the buffer.

debug debug as first argument indicates a call to debug because of entry to a function that was set to debug on entry. The debugger displays ‘Entering:’, just as in the lambda case. It also marks the stack frame for that function so that it will invoke the debugger when exited.

t When the first argument is t, this indicates a call to debug due to evaluation of a list form when debug-on-next-call is non-nil. The debugger displays the following as the top line in the buffer:

     Beginning evaluation of function call form:

exit When the first argument is exit, it indicates the exit of a stack frame previously marked to invoke the debugger on exit. The second argument given to debug in this case is the value being returned from the frame. The debugger displays ‘Return value:’ on the top line of the buffer, followed by the value being returned.

error When the first argument is error, the debugger indicates that it is being entered because an error or quit was signaled and not handled, by displaying ‘Signaling:’ followed by the error signaled and any arguments to signal. For example,

     (let ((debug-on-error t))
       (/ 1 0))
     ------ Buffer: *Backtrace* ------
     Signaling: (arith-error)
       /(1 0)
     ...
     ------ Buffer: *Backtrace* ------

If an error was signaled, presumably the variable debug-on-error is non-nil. If quit was signaled, then presumably the variable

debug-on-quit is non-nil.

nil Use nil as the first of the debugger-args when you want to enter the debugger explicitly. The rest of the debugger-args are printed on the top line of the buffer. You can use this feature to display messages—for example, to remind yourself of the conditions under which debug is called.

16.1.8 Internals of the Debugger

This section describes functions and variables used internally by the debugger.

Variable debugger

The value of this variable is the function to call to invoke the debugger. Its value must be a function of any number of arguments (or, more typically, the name of a function). Presumably this function will enter some kind of debugger. The default value of the variable is debug.

The first argument that Lisp hands to the function indicates why it was called. The convention for arguments is detailed in the description of debug.

Command backtrace &optional stream detailed This function prints a trace of Lisp function calls currently active. This is the function used by debug to fill up the ‘*Backtrace*’ buffer. It is written in C, since it must have access to the stack to determine which function calls are active. The return value is always nil.

The backtrace is normally printed to standard-output, but this can be changed by specifying a value for stream. If detailed is non-nil, the backtrace also shows places where currently active variable bindings, catches, condition-cases, and unwind-protects were made as well as function calls.

In the following example, a Lisp expression calls backtrace explicitly. This prints the backtrace to the stream standard-output: in this case, to the buffer ‘backtrace-output’. Each line of the backtrace represents one function call. The line shows the values of the function’s arguments if they are all known. If they are still being computed, the line says so. The arguments of special forms are elided.

     (with-output-to-temp-buffer "backtrace-output"
       (let ((var 1))
         (save-excursion
           (setq var (eval '(progn
                              (1+ var)
                              (list 'testing (backtrace))))))))

     ----------- Buffer: backtrace-output ------------
       backtrace()
       (list ...computing arguments...)
       (progn ...)
       eval((progn (1+ var) (list (quote testing) (backtrace))))
       (setq ...)
       (save-excursion ...)
       (let ...)
       (with-output-to-temp-buffer ...)
       eval-region(1973 2142 #<buffer *scratch*>)
       byte-code("...  for eval-print-last-sexp ...")
       eval-print-last-sexp(nil)
     * call-interactively(eval-print-last-sexp)
     ----------- Buffer: backtrace-output ------------

The character ‘*’ indicates a frame whose debug-on-exit flag is set.

Variable debug-on-next-call

If this variable is non-nil, it says to call the debugger before the next eval,

apply or funcall. Entering the debugger sets debug-on-next-call to nil.

The d command in the debugger works by setting this variable.

Function backtrace-debug level flag

This function sets the debug-on-exit flag of the stack frame level levels down the stack, giving it the value flag. If flag is non-nil, this will cause the debugger to be entered when that frame later exits. Even a nonlocal exit through that frame will enter the debugger. This function is used only by the debugger.

Variable command-debug-status

This variable records the debugging status of the current interactive command. Each time a command is called interactively, this variable is bound to nil. The debugger can set this variable to leave information for future debugger invocations during the same command. The advantage, for the debugger, of using this variable rather than another global variable is that the data will never carry over to a subsequent command invocation.

Function backtrace-frame frame-number

The function backtrace-frame is intended for use in Lisp debuggers. It returns information about what computation is happening in the stack frame framenumber levels down. If that frame has not evaluated the arguments yet (or is a special form), the value is (nil function arg-forms...). If that frame has evaluated its arguments and called its function already, the value is (t function arg-values...). In the return value, function is whatever was supplied as the car of the evaluated list, or a lambda expression in the case of a macro call. If the function has a &rest argument, that is represented as the tail of the list arg-values.

If frame-number is out of range, backtrace-frame returns nil.

16.2 Debugging Invalid Lisp Syntax

The Lisp reader reports invalid syntax, but cannot say where the real problem is. For example, the error “End of file during parsing” in evaluating an expression indicates an excess of open parentheses (or square brackets). The reader detects this imbalance at the end of the file, but it cannot figure out where the close parenthesis should have been. Likewise, “Invalid read syntax: ")"” indicates an excess close parenthesis or missing open parenthesis, but does not say where the missing parenthesis belongs. How, then, to find what to change?

If the problem is not simply an imbalance of parentheses, a useful technique is to try C-M-e at the beginning of each defun, and see if it goes to the place where that defun appears to end. If it does not, there is a problem in that defun.

However, unmatched parentheses are the most common syntax errors in Lisp, and we can give further advice for those cases.

16.2.1 Excess Open Parentheses

The first step is to find the defun that is unbalanced. If there is an excess open parenthesis, the way to do this is to insert a close parenthesis at the end of the file and type C-M-b (backward-sexp). This will move you to the beginning of the defun that is unbalanced. (Then type C-^{⟨}\underline{\overline{SPC}}^{⟩} C-_C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} to set the mark there, undo the insertion of the close parenthesis, and finally return to the mark.)

The next step is to determine precisely what is wrong. There is no way to be sure of this except to study the program, but often the existing indentation is a clue to where the parentheses should have been. The easiest way to use this clue is to reindent with C-M-q and see what moves.

Before you do this, make sure the defun has enough close parentheses. Otherwise, C-M-q will get an error, or will reindent all the rest of the file until the end. So move to the end of the defun and insert a close parenthesis there. Don’t use C-M-e to move there, since that too will fail to work until the defun is balanced.

Now you can go to the beginning of the defun and type C-M-q. Usually all the lines from a certain point to the end of the function will shift to the right. There is probably a missing close parenthesis, or a superfluous open parenthesis, near that point. (However, don’t assume this is true; study the code to make sure.) Once you have found the discrepancy, undo the C-M-q with C-_, since the old indentation is probably appropriate to the intended parentheses.

After you think you have fixed the problem, use C-M-q again. If the old indentation actually fit the intended nesting of parentheses, and you have put back those parentheses, C-M-q should not change anything.

16.2.2 Excess Close Parentheses

To deal with an excess close parenthesis, first insert an open parenthesis at the beginning of the file, back up over it, and type C-M-f to find the end of the unbalanced defun.

(Then type C-^{⟨}\underline{\overline{SPC}}^{⟩} C-_C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} to set the mark there, undo the insertion of the open parenthesis, and finally return to the mark.)

Then find the actual matching close parenthesis by typing C-M-f at the beginning of the defun. This will leave you somewhere short of the place where the defun ought to end. It is possible that you will find a spurious close parenthesis in that vicinity.

If you don’t see a problem at that point, the next thing to do is to type C-M-q at the beginning of the defun. A range of lines will probably shift left; if so, the missing open parenthesis or spurious close parenthesis is probably near the first of those lines. (However, don’t assume this is true; study the code to make sure.) Once you have found the discrepancy, undo the C-M-q with C-_, since the old indentation is probably appropriate to the intended parentheses.

After you think you have fixed the problem, use C-M-q again. If the old indentation actually fit the intended nesting of parentheses, and you have put back those parentheses, C-M-q should not change anything.

16.3 Debugging Problems in Compilation

When an error happens during byte compilation, it is normally due to invalid syntax in the program you are compiling. The compiler prints a suitable error message in the ‘*Compile-Log*’ buffer, and then stops. The message may state a function name in which the error was found, or it may not. Either way, here is how to find out where in the file the error occurred.

What you should do is switch to the buffer ‘ *Compiler Input*’. (Note that the buffer name starts with a space, so it does not show up in M-x list-buffers.) This buffer contains the program being compiled, and point shows how far the byte compiler was able to read.

If the error was due to invalid Lisp syntax, point shows exactly where the invalid syntax was detected. The cause of the error is not necessarily near by! Use the techniques in the previous section to find the error.

If the error was detected while compiling a form that had been read successfully, then point is located at the end of the form. In this case, this technique can’t localize the error precisely, but can still show you which function to check.

16.4 Edebug

Edebug is a source-level debugger for XEmacs Lisp programs that provides the following features:

• Step through evaluation, stopping before and after each expression.

• Set conditional or unconditional breakpoints, install embedded breakpoints, or a global break event.

• Trace slow or fast stopping briefly at each stop point, or each breakpoint.

• Display expression results and evaluate expressions as if outside of Edebug. Interface with the custom printing package for printing circular structures.

• Automatically reevaluate a list of expressions and display their results each time Edebug updates the display.

• Output trace info on function enter and exit.

• Errors stop before the source causing the error.

• Display backtrace without Edebug calls.

• Allow specification of argument evaluation for macros and defining forms.

• Provide rudimentary coverage testing and display of frequency counts.

The first three sections should tell you enough about Edebug to enable you to use it.

16.4.1 Using Edebug

To debug an XEmacs Lisp program with Edebug, you must first instrument the Lisp code that you want to debug. If you want to just try it now, load ‘edebug.el’, move point into a definition and do C-u C-M-x (eval-defun with a prefix argument). See Section 16.4.2 [Instrumenting], page 229 for alternative ways to instrument code.

Once a function is instrumented, any call to the function activates Edebug. Activating Edebug may stop execution and let you step through the function, or it may update the display and continue execution while checking for debugging commands, depending on the selected Edebug execution mode. The initial execution mode is step, by default, which does stop execution. See Section 16.4.3 [Edebug Execution Modes], page 230.

Within Edebug, you normally view an XEmacs buffer showing the source of the Lisp function you are debugging. This is referred to as the source code buffer—but note that it is not always the same buffer depending on which function is currently being executed.

An arrow at the left margin indicates the line where the function is executing. Point initially shows where within the line the function is executing, but you can move point yourself.

If you instrument the definition of fac (shown below) and then execute (fac 3), here is what you normally see. Point is at the open-parenthesis before if.

     (defun fac (n)
     =>⋆(if (< 0 n)
           (* n (fac (1- n)))
         1))

The places within a function where Edebug can stop execution are called stop points. These occur both before and after each subexpression that is a list, and also after each variable reference. Here we show with periods the stop points found in the function fac:

     (defun fac (n)
       .(if .(< 0 n.).
           .(* n. .(fac (1- n.).).).
         1).)

While the source code buffer is selected, the special commands of Edebug are available in it, in addition to the commands of XEmacs Lisp mode. (The buffer is temporarily made read-only, however.) For example, you can type the Edebug command ^{⟨}\underline{\overline{SPC}}^{⟩} to execute until the next stop point. If you type ^{⟨}\underline{\overline{SPC}}^{⟩} once after entry to fac, here is the display you will see:

     (defun fac (n)
     =>(if ⋆(< 0 n)

           (* n (fac (1- n)))
         1))

When Edebug stops execution after an expression, it displays the expression’s value in the echo area.

Other frequently used commands are b to set a breakpoint at a stop point, g to execute until a breakpoint is reached, and q to exit to the top-level command loop. Type ? to display a list of all Edebug commands.

16.4.2 Instrumenting for Edebug

In order to use Edebug to debug Lisp code, you must first instrument the code. Instrumenting a form inserts additional code into it which invokes Edebug at the proper places. Furthermore, if Edebug detects a syntax error while instrumenting, point is left at the erroneous code and an invalid-read-syntax error is signaled.

Once you have loaded Edebug, the command C-M-x (eval-defun) is redefined so that when invoked with a prefix argument on a definition, it instruments the definition before evaluating it. (The source code itself is not modified.) If the variable edebug-all-defs is non-nil, that inverts the meaning of the prefix argument: then C-M-x instruments the definition unless it has a prefix argument. The default value of edebug-all-defs is nil. The command M-x edebug-all-defs toggles the value of the variable edebug-all-defs.

   If edebug-all-defs is non-nil, then the commands eval-region, eval-current-

buffer, and eval-buffer also instrument any definitions they evaluate. Similarly, edebugall-forms controls whether eval-region should instrument any form, even non-defining forms. This doesn’t apply to loading or evaluations in the minibuffer. The command M-x edebug-all-forms toggles this option.

Another command, M-x edebug-eval-top-level-form, is available to instrument any top-level form regardless of the value of edebug-all-defs or edebug-all-forms.

Just before Edebug instruments any code, it calls any functions in the variable edebugsetup-hook and resets its value to nil. You could use this to load up Edebug specifications associated with a package you are using but only when you also use Edebug. For example, ‘my-specs.el’ may be loaded automatically when you use my-package with Edebug by including the following code in ‘my-package.el’.

     (add-hook 'edebug-setup-hook
       (function (lambda () (require 'my-specs))))

While Edebug is active, the command I (edebug-instrument-callee) instruments the definition of the function or macro called by the list form after point, if is not already instrumented. If the location of the definition is not known to Edebug, this command cannot be used. After loading Edebug, eval-region records the position of every definition it evaluates, even if not instrumenting it. Also see the command i (Section 16.4.4 [Jumping], page 231) which steps into the callee.

Edebug knows how to instrument all the standard special forms, an interactive form with an expression argument, anonymous lambda expressions, and other defining forms. (Specifications for macros defined by ‘cl.el’ (version 2.03) are provided in ‘cl-specs.el’.) Edebug cannot know what a user-defined macro will do with the arguments of a macro call so you must tell it. See Section 16.4.16 [Instrumenting Macro Calls], page 240 for the details.

Note that a couple ways remain to evaluate expressions without instrumenting them. Loading a file via the load subroutine does not instrument expressions for Edebug. Evaluations in the minibuffer via eval-expression (M-ESC) are not instrumented.

To remove instrumentation from a definition, simply reevaluate it with one of the noninstrumenting commands, or reload the file.

See Section 16.4.9 [Edebug Eval], page 235 for other evaluation functions available inside of Edebug.

16.4.3 Edebug Execution Modes

Edebug supports several execution modes for running the program you are debugging. We call these alternatives Edebug execution modes; do not confuse them with major or minor modes. The current Edebug execution mode determines how Edebug displays the progress of the evaluation, whether it stops at each stop point, or continues to the next breakpoint, for example.

Normally, you specify the Edebug execution mode by typing a command to continue the program in a certain mode. Here is a table of these commands. All except for S resume execution of the program, at least for a certain distance.

S Stop: don’t execute any more of the program for now, just wait for more Edebug

          commands (edebug-stop).

⟨\underline{\overline{SPC}}^{⟩} Step: stop at the next stop point encountered (edebug-step-mode).

n Next: stop at the next stop point encountered after an expression (edebugnext-mode). Also see edebug-forward-sexp in Section 16.4.5 [Edebug Misc], page 232.

t Trace: pause one second at each Edebug stop point (edebug-trace-mode).

T Rapid trace: update at each stop point, but don’t actually pause (edebug-

          Trace-fast-mode).

g Go: run until the next breakpoint (edebug-go-mode). See Section 16.4.6 [Breakpoints], page 232.

c Continue: pause for one second at each breakpoint, but don’t stop (edebug-

          continue-mode).

C Rapid continue: update at each breakpoint, but don’t actually pause (edebug-

          Continue-fast-mode).

G Go non-stop: ignore breakpoints (edebug-Go-nonstop-mode). You can still stop the program by hitting any key.

In general, the execution modes earlier in the above list run the program more slowly or stop sooner.

When you enter a new Edebug level, the initial execution mode comes from the value of the variable edebug-initial-mode. By default, this specifies step mode. Note that you may reenter the same Edebug level several times if, for example, an instrumented function is called several times from one command.

While executing or tracing, you can interrupt the execution by typing any Edebug command. Edebug stops the program at the next stop point and then executes the command that you typed. For example, typing t during execution switches to trace mode at the next stop point. You can use S to stop execution without doing anything else.

If your function happens to read input, a character you hit intending to interrupt execution may be read by the function instead. You can avoid such unintended results by paying attention to when your program wants input.

Keyboard macros containing Edebug commands do not work; when you exit from Edebug, to resume the program, whether you are defining or executing a keyboard macro is forgotten. Also, defining or executing a keyboard macro outside of Edebug does not affect the command loop inside Edebug. This is usually an advantage. But see edebug-continue-

kbd-macro.

16.4.4 Jumping

Commands described here let you jump to a specified location. All, except i, use temporary breakpoints to establish the stop point and then switch to go mode. Any other breakpoint reached before the intended stop point will also stop execution. See Section 16.4.6 [Breakpoints], page 232 for the details on breakpoints.

f Run the program forward over one expression (edebug-forward-sexp). More precisely, set a temporary breakpoint at the position that C-M-f would reach, then execute in go mode so that the program will stop at breakpoints. With a prefix argument n, the temporary breakpoint is placed n sexps beyond point. If the containing list ends before n more elements, then the place to stop is after the containing expression. Be careful that the position C-M-f finds is a place that the program will really get to; this may not be true in a cond, for example. This command does forward-sexp starting at point rather than the stop point. If you want to execute one expression from the current stop point, type w first, to move point there.

o Continue “out of” an expression (edebug-step-out). It places a temporary breakpoint at the end of the sexp containing point. If the containing sexp is a function definition itself, it continues until just before the last sexp in the definition. If that is where you are now, it returns from the function and then stops. In other words, this command does not exit the currently executing function unless you are positioned after the last sexp.

I Step into the function or macro after point after first ensuring that it is instrumented. It does this by calling edebug-on-entry and then switching to go mode. Although the automatic instrumentation is convenient, it is not later automatically uninstrumented. h Proceed to the stop point near where point is using a temporary breakpoint

          (edebug-goto-here).

All the commands in this section may fail to work as expected in case of nonlocal exit, because a nonlocal exit can bypass the temporary breakpoint where you expected the program to stop.

16.4.5 Miscellaneous

Some miscellaneous commands are described here.

? Display the help message for Edebug (edebug-help).

C-] Abort one level back to the previous command level (abort-recursive-edit).

q Return to the top level editor command loop (top-level). This exits all recursive editing levels, including all levels of Edebug activity. However, instrumented code protected with unwind-protect or condition-case forms may resume debugging.

Q Like q but don’t stop even for protected code (top-level-nonstop).

r Redisplay the most recently known expression result in the echo area (edebug-

          previous-result).

d Display a backtrace, excluding Edebug’s own functions for clarity (edebug-

          backtrace).

You cannot use debugger commands in the backtrace buffer in Edebug as you would in the standard debugger. The backtrace buffer is killed automatically when you continue execution.

From the Edebug recursive edit, you may invoke commands that activate Edebug again recursively. Any time Edebug is active, you can quit to the top level with q or abort one recursive edit level with C-]. You can display a backtrace of all the pending evaluations with d.

16.4.6 Breakpoints

There are three more ways to stop execution once it has started: breakpoints, the global break condition, and embedded breakpoints.

While using Edebug, you can specify breakpoints in the program you are testing: points where execution should stop. You can set a breakpoint at any stop point, as defined in Section 16.4.1 [Using Edebug], page 228. For setting and unsetting breakpoints, the stop point that is affected is the first one at or after point in the source code buffer. Here are the Edebug commands for breakpoints:

b Set a breakpoint at the stop point at or after point (edebug-set-breakpoint). If you use a prefix argument, the breakpoint is temporary (it turns off the first time it stops the program).

u Unset the breakpoint (if any) at the stop point at or after the current point

          (edebug-unset-breakpoint).

x condition ^{⟨}\underline{\overline{RET}}^{⟩} Set a conditional breakpoint which stops the program only if condition evaluates

          to a non-nil value (edebug-set-conditional-breakpoint). If you use a prefix

argument, the breakpoint is temporary (it turns off the first time it stops the program).

B Move point to the next breakpoint in the definition (edebug-next-breakpoint).

While in Edebug, you can set a breakpoint with b and unset one with u. First you must move point to a position at or before the desired Edebug stop point, then hit the key to change the breakpoint. Unsetting a breakpoint that has not been set does nothing.

Reevaluating or reinstrumenting a definition clears all its breakpoints.

A conditional breakpoint tests a condition each time the program gets there. To set a conditional breakpoint, use x, and specify the condition expression in the minibuffer. Setting a conditional breakpoint at a stop point that already has a conditional breakpoint puts the current condition expression in the minibuffer so you can edit it.

You can make both conditional and unconditional breakpoints temporary by using a prefix arg to the command to set the breakpoint. After breaking at a temporary breakpoint, it is automatically cleared.

Edebug always stops or pauses at a breakpoint except when the Edebug mode is Gononstop. In that mode, it ignores breakpoints entirely.

To find out where your breakpoints are, use B, which moves point to the next breakpoint in the definition following point, or to the first breakpoint if there are no following breakpoints. This command does not continue execution—it just moves point in the buffer.

16.4.6.1 Global Break Condition

In contrast to breaking when execution reaches specified locations, you can also cause a break when a certain event occurs. The global break condition is a condition that is repeatedly evaluated at every stop point. If it evaluates to a non-nil value, then execution is stopped or paused depending on the execution mode, just like a breakpoint. Any errors that might occur as a result of evaluating the condition are ignored, as if the result were

nil.

You can set or edit the condition expression, stored in edebug-global-break-condition,

using X (edebug-set-global-break-condition).

Using the global break condition is perhaps the fastest way to find where in your code some event occurs, but since it is rather expensive you should reset the condition to nil when not in use.

16.4.6.2 Embedded Breakpoints

Since all breakpoints in a definition are cleared each time you reinstrument it, you might rather create an embedded breakpoint which is simply a call to the function edebug. You can, of course, make such a call conditional. For example, in the fac function, insert the first line as shown below to stop when the argument reaches zero:

     (defun fac (n)
       (if (= n 0) (edebug))
       (if (< 0 n)
           (* n (fac (1- n)))
         1))

When the fac definition is instrumented and the function is called, Edebug will stop before the call to edebug. Depending on the execution mode, Edebug will stop or pause.

However, if no instrumented code is being executed, calling edebug will instead invoke debug. Calling debug will always invoke the standard backtrace debugger.

16.4.7 Trapping Errors

An error may be signaled by subroutines or XEmacs Lisp code. If a signal is not handled by a condition-case, this indicates an unrecognized situation has occurred. If Edebug is not active when an unhandled error is signaled, debug is run normally (if debug-on-error is non-nil). But while Edebug is active, debug-on-error and debug-on-quit are bound to edebug-on-error and edebug-on-quit, which are both t by default. Actually, if debugon-error already has a non-nil value, that value is still used.

It is best to change the values of edebug-on-error or edebug-on-quit when Edebug is not active since their values won’t be used until the next time Edebug is invoked at a deeper command level. If you only change debug-on-error or debug-on-quit while Edebug is active, these changes will be forgotten when Edebug becomes inactive. Furthermore, during Edebug’s recursive edit, these variables are bound to the values they had outside of Edebug.

Edebug shows you the last stop point that it knew about before the error was signaled. This may be the location of a call to a function which was not instrumented, within which the error actually occurred. For an unbound variable error, the last known stop point might be quite distant from the offending variable. If the cause of the error is not obvious at first, note that you can also get a full backtrace inside of Edebug (see Section 16.4.5 [Edebug Misc], page 232).

Edebug can also trap signals even if they are handled. If debug-on-error is a list of signal names, Edebug will stop when any of these errors are signaled. Edebug shows you the last known stop point just as for unhandled errors. After you continue execution, the error is signaled again (but without being caught by Edebug). Edebug can only trap errors that are handled if they are signaled in Lisp code (not subroutines) since it does so by temporarily replacing the signal function.

16.4.8 Edebug Views

The following Edebug commands let you view aspects of the buffer and window status that obtained before entry to Edebug.

v View the outside window configuration (edebug-view-outside).

p Temporarily display the outside current buffer with point at its outside position (edebug-bounce-point). If prefix arg is supplied, sit for that many seconds instead. w Move point back to the current stop point (edebug-where) in the source code

buffer. Also, if you use this command in another window displaying the same buffer, this window will be used instead to display the buffer in the future.

W Toggle the edebug-save-windows variable which indicates whether the outside window configuration is saved and restored (edebug-toggle-save-windows). Also, each time it is toggled on, make the outside window configuration the same as the current window configuration. With a prefix argument, edebug-toggle-save-windows only toggles saving and restoring of the selected window. To specify a window that is not displaying the source code buffer, you must use C-xXW from the global keymap.

You can view the outside window configuration with v or just bounce to the current point in the current buffer with p, even if it is not normally displayed. After moving point, you may wish to pop back to the stop point with w from a source code buffer.

By using W twice, Edebug again saves and restores the outside window configuration, but to the current configuration. This is a convenient way to, for example, add another buffer to be displayed whenever Edebug is active. However, the automatic redisplay of ‘*edebug*’ and ‘*edebug-trace*’ may conflict with the buffers you wish to see unless you have enough windows open.

16.4.9 Evaluation

While within Edebug, you can evaluate expressions “as if” Edebug were not running. Edebug tries to be invisible to the expression’s evaluation and printing. Evaluation of expressions that cause side effects will work as expected except for things that Edebug explicitly saves and restores. See Section 16.4.15 [The Outside Context], page 239 for details on this process. Also see Section 16.4.11 [Reading in Edebug], page 237 and Section 16.4.12 [Printing in Edebug], page 237 for topics related to evaluation.

e exp ^{⟨}\underline{\overline{RET}}^{⟩} Evaluate expression exp in the context outside of Edebug (edebug-eval-expression). In other words, Edebug tries to avoid altering the effect of exp.

M-^{⟨}\underline{\overline{ESC}}^{⟩} exp ^{⟨}\underline{\overline{RET}}^{⟩} Evaluate expression exp in the context of Edebug itself.

C-x C-e Evaluate the expression before point, in the context outside of Edebug (edebug-

          eval-last-sexp).

Edebug supports evaluation of expressions containing references to lexically bound symbols created by the following constructs in ‘cl.el’ (version 2.03 or later): lexical-let,

macrolet, and symbol-macrolet.

16.4.10 Evaluation List Buffer

You can use the evaluation list buffer, called ‘*edebug*’, to evaluate expressions interactively. You can also set up the evaluation list of expressions to be evaluated automatically each time Edebug updates the display.

E Switch to the evaluation list buffer ‘*edebug*’ (edebug-visit-eval-list).

In the ‘*edebug*’ buffer you can use the commands of Lisp Interaction as well as these special commands:

LFD Evaluate the expression before point, in the outside context, and insert the

     value in the buffer (edebug-eval-print-last-sexp).

C-x C-e Evaluate the expression before point, in the context outside of Edebug (edebug-

     eval-last-sexp).

C-c C-u Build a new evaluation list from the first expression of each group, reevaluate and redisplay (edebug-update-eval-list). Groups are separated by comment lines.

C-c C-d Delete the evaluation list group that point is in (edebug-delete-eval-item).

C-c C-w Switch back to the source code buffer at the current stop point (edebug-where).

You can evaluate expressions in the evaluation list window with LFD or C-x C-e, just as you would in ‘*scratch*’; but they are evaluated in the context outside of Edebug.

The expressions you enter interactively (and their results) are lost when you continue execution unless you add them to the evaluation list with C-c C-u. This command builds a new list from the first expression of each evaluation list group. Groups are separated by comment lines. Be careful not to add expressions that execute instrumented code otherwise an infinite loop will result.

When the evaluation list is redisplayed, each expression is displayed followed by the result of evaluating it, and a comment line. If an error occurs during an evaluation, the error message is displayed in a string as if it were the result. Therefore expressions that, for example, use variables not currently valid do not interrupt your debugging.

Here is an example of what the evaluation list window looks like after several expressions have been added to it:

(current-buffer)
#<buffer *scratch*>
;---------------------------------------------------------------
(selected-window)
#<window 16 on *scratch*>
;---------------------------------------------------------------
(point)
196
;---------------------------------------------------------------
bad-var
"Symbol's value as variable is void: bad-var"
;---------------------------------------------------------------
(recursion-depth)
0
;---------------------------------------------------------------
this-command
eval-last-sexp
;---------------------------------------------------------------

To delete a group, move point into it and type C-c C-d, or simply delete the text for the group and update the evaluation list with C-c C-u. When you add a new group, be sure it is separated from its neighbors by a comment line.

After selecting ‘*edebug*’, you can return to the source code buffer with C-c C-w. The ‘*edebug*’ buffer is killed when you continue execution, and recreated next time it is needed.

16.4.11 Reading in Edebug

To instrument a form, Edebug first reads the whole form. Edebug replaces the standard Lisp Reader with its own reader that remembers the positions of expressions. This reader is

used by the Edebug replacements for eval-region, eval-defun, eval-buffer, and eval-
current-buffer.

Another package, ‘cl-read.el’, replaces the standard reader with one that understands Common Lisp reader macros. If you use that package, Edebug will automatically load ‘edebug-cl-read.el’ to provide corresponding reader macros that remember positions of expressions. If you define new reader macros, you will have to define similar reader macros for Edebug.

16.4.12 Printing in Edebug

If the result of an expression in your program contains a circular reference, you may get an error when Edebug attempts to print it. You can set print-length to a non-zero value to limit the print length of lists (the number of cdrs), and in Emacs 19, set print-level to a non-zero value to limit the print depth of lists. But you can print such circular structures and structures that share elements more informatively by using the ‘cust-print’ package.

To load ‘cust-print’ and activate custom printing only for Edebug, simply use the command M-x edebug-install-custom-print. To restore the standard print functions, use M-x edebug-uninstall-custom-print. You can also activate custom printing for printing in any Lisp code; see the package for details.

Here is an example of code that creates a circular structure:

     (progn
       (edebug-install-custom-print)
       (setq a '(x y))
       (setcar a a))

Edebug will print the result of the setcar as ‘Result: #1=(#1# y)’. The ‘#1=’ notation names the structure that follows it, and the ‘#1#’ notation references the previously named structure. This notation is used for any shared elements of lists or vectors.

Independent of whether ‘cust-print’ is active, while printing results Edebug binds

print-length, print-level, and print-circle to edebug-print-length (50), edebug-

print-level (50), and edebug-print-circle (t) respectively, if these values are nonnil. Also, print-readably is bound to nil since some objects simply cannot be printed readably.

16.4.13 Tracing

In addition to automatic stepping through source code, which is also called tracing (see Section 16.4.3 [Edebug Execution Modes], page 230), Edebug can produce a traditional trace listing of execution in a separate buffer, ‘*edebug-trace*’.

If the variable edebug-trace is non-nil, each function entry and exit adds lines to the trace buffer. On function entry, Edebug prints ‘::::{’ followed by the function name and argument values. On function exit, Edebug prints ‘::::}’ followed by the function name and result of the function. The number of ‘:’s is computed from the recursion depth. The balanced braces in the trace buffer can be used to find the matching beginning or end of function calls. These displays may be customized by replacing the functions edebug-print-trace-before and edebug-print-trace-after, which take an arbitrary message string to print.

The macro edebug-tracing provides tracing similar to function enter and exit tracing, but for arbitrary expressions. This macro should be explicitly inserted by you around expressions you wish to trace the execution of. The first argument is a message string (evaluated), and the rest are expressions to evaluate. The result of the last expression is returned.

Finally, you can insert arbitrary strings into the trace buffer with explicit calls to edebugtrace. The arguments of this function are the same as for message, but a newline is always inserted after each string printed in this way.

edebug-tracing and edebug-trace insert lines in the trace buffer even if Edebug is not active. Every time the trace buffer is added to, the window is scrolled to show the last lines inserted. (There may be some display problems if you use tracing along with the evaluation list.)

16.4.14 Coverage Testing

Edebug provides a rudimentary coverage tester and display of execution frequency. Frequency counts are always accumulated, both before and after evaluation of each instrumented expression, even if the execution mode is Go-nonstop. Coverage testing is only done if the option edebug-test-coverage is non-nil because this is relatively expensive.

Both data sets are displayed by M-x edebug-display-freq-count.

Command edebug-display-freq-count

Display the frequency count data for each line of the current definition. The frequency counts are inserted as comment lines after each line, and you can undo all insertions with one undo command. The counts are inserted starting under the ( before an expression or the ) after an expression, or on the last char of a symbol. The counts are only displayed when they differ from previous counts on the same line. If coverage is being tested, whenever all known results of an expression are eq, the char = will be appended after the count for that expression. Note that this is always the case for an expression only evaluated once. To clear the frequency count and coverage data for a definition, reinstrument it.

For example, after evaluating (fac 5) with an embedded breakpoint, and setting edebug-test-coverage to t, when the breakpoint is reached, the frequency data is looks like this:

     (defun fac (n)
       (if (= n 0) (edebug))
     ;#6           1      0 =5
       (if (< 0 n)
     ;#5         =
           (* n (fac (1- n)))
     ;#    5               0
         1))
     ;#   0

The comment lines show that fac has been called 6 times. The first if statement has returned 5 times with the same result each time, and the same is true for the condition on the second if. The recursive call of fac has not returned at all.

16.4.15 The Outside Context

Edebug tries to be transparent to the program you are debugging. In addition, most evaluations you do within Edebug (see Section 16.4.9 [Edebug Eval], page 235) occur in the same outside context which is temporarily restored for the evaluation. But Edebug is not completely successful and this section explains precisely how it fails. Edebug operation unavoidably alters some data in XEmacs, and this can interfere with debugging certain programs. Also notice that Edebug’s protection against change of outside data means that any side effects intended by the user in the course of debugging will be defeated.

16.4.15.1 Checking Whether to Stop

Whenever Edebug is entered just to think about whether to take some action, it needs to save and restore certain data.

  • max-lisp-eval-depth and max-specpdl-size are both incremented one time to re-

duce Edebug’s impact on the stack. You could, however, still run out of stack space when using Edebug.

• The state of keyboard macro execution is saved and restored. While Edebug is active,

    executing-macro is bound to edebug-continue-kbd-macro.

16.4.15.2 Edebug Display Update

When Edebug needs to display something (e.g., in trace mode), it saves the current window configuration from “outside” Edebug. When you exit Edebug (by continuing the program), it restores the previous window configuration.

XEmacs redisplays only when it pauses. Usually, when you continue execution, the program comes back into Edebug at a breakpoint or after stepping without pausing or reading input in between. In such cases, XEmacs never gets a chance to redisplay the “outside” configuration. What you see is the same window configuration as the last time Edebug was active, with no interruption.

Entry to Edebug for displaying something also saves and restores the following data, but some of these are deliberately not restored if an error or quit signal occurs.

• Which buffer is current, and where point and mark are in the current buffer are saved and restored.

• The Edebug Display Update, is saved and restored if edebug-save-windows is non-nil. It is not restored on error or quit, but the outside selected window is reselected even on error or quit in case a save-excursion is active. If the value of edebug-save-windows is a list, only the listed windows are saved and restored. The window start and horizontal scrolling of the source code buffer are not restored, however, so that the display remains coherent.

• The value of point in each displayed buffer is saved and restored if edebug-save-

displayed-buffer-points is non-nil.
• The variables overlay-arrow-position and overlay-arrow-string are saved and

restored. So you can safely invoke Edebug from the recursive edit elsewhere in the same buffer.

• cursor-in-echo-area is locally bound to nil so that the cursor shows up in the window.

16.4.15.3 Edebug Recursive Edit

When Edebug is entered and actually reads commands from the user, it saves (and later restores) these additional data:

• The current match data, for whichever buffer was current.

• last-command, this-command, last-command-char, last-input-char, last-input-
event, last-command-event, last-event-frame, last-nonmenu-event, and track-

mouse . Commands used within Edebug do not affect these variables outside of Edebug. The key sequence returned by this-command-keys is changed by executing commands within Edebug and there is no way to reset the key sequence from Lisp. For Emacs 18, Edebug cannot save and restore the value of unread-command-char. Entering Edebug while this variable has a nontrivial value can interfere with execution of the program you are debugging.

• Complex commands executed while in Edebug are added to the variable commandhistory. In rare cases this can alter execution.

• Within Edebug, the recursion depth appears one deeper than the recursion depth outside Edebug. This is not true of the automatically updated evaluation list window.

• standard-output and standard-input are bound to nil by the recursive-edit, but

Edebug temporarily restores them during evaluations.

• The state of keyboard macro definition is saved and restored. While Edebug is active,

defining-kbd-macro is bound to edebug-continue-kbd-macro.

16.4.16 Instrumenting Macro Calls

When Edebug instruments an expression that calls a Lisp macro, it needs additional advice to do the job properly. This is because there is no way to tell which subexpressions of the macro call may be evaluated. (Evaluation may occur explicitly in the macro body, or when the resulting expansion is evaluated, or any time later.) You must explain the format of macro call arguments by using def-edebug-spec to define an Edebug specification for each macro.

Macro def-edebug-spec macro specification Specify which expressions of a call to macro macro are forms to be evaluated. For simple macros, the specification often looks very similar to the formal argument list of the macro definition, but specifications are much more general than macro arguments. The macro argument may actually be any symbol, not just a macro name. Unless you are using Emacs 19 or XEmacs, this macro is only defined in Edebug, so you may want to use the following which is equivalent: (put 'macro

     'edebug-form-spec 'specification)

Here is a simple example that defines the specification for the for macro described in the XEmacs Lisp Reference Manual, followed by an alternative, equivalent specification.

     (def-edebug-spec for
       (symbolp "from" form "to" form "do" &rest form))
     (def-edebug-spec for
       (symbolp ['from form] ['to form] ['do body]))

Here is a table of the possibilities for specification and how each directs processing of arguments.

•t All arguments are instrumented for evaluation.

•0 None of the arguments is instrumented.

•a symbol The symbol must have an Edebug specification which is used instead. This indirection is repeated until another kind of specification is found. This allows you to inherit the specification for another macro.

•a list The elements of the list describe the types of the arguments of a calling form. The possible elements of a specification list are described in the following sections.

16.4.16.1 Specification List

A specification list is required for an Edebug specification if some arguments of a macro call are evaluated while others are not. Some elements in a specification list match one or more arguments, but others modify the processing of all following elements. The latter, called keyword specifications, are symbols beginning with ‘&’ (e.g. &optional).

A specification list may contain sublists which match arguments that are themselves lists, or it may contain vectors used for grouping. Sublists and groups thus subdivide the specification list into a hierarchy of levels. Keyword specifications only apply to the remainder of the sublist or group they are contained in and there is an implicit grouping around a keyword specification and all following elements in the sublist or group.

If a specification list fails at some level, then backtracking may be invoked to find some alternative at a higher level, or if no alternatives remain, an error will be signaled. See Section 16.4.16.2 [Backtracking], page 244 for more details.

Edebug specifications provide at least the power of regular expression matching. Some context-free constructs are also supported: the matching of sublists with balanced parentheses, recursive processing of forms, and recursion via indirect specifications.

Each element of a specification list may be one of the following, with the corresponding type of argument:

sexp A single unevaluated expression.

form A single evaluated expression, which is instrumented.

place A place as in the Common Lisp setf place argument. It will be instrumented just like a form, but the macro is expected to strip the instrumentation. Two functions, edebug-unwrap and edebug-unwrap*, are provided to strip the instrumentation one level or recursively at all levels.

body  Short for &rest form. See &rest below.
function-form

A function form: either a quoted function symbol, a quoted lambda expression, or a form (that should evaluate to a function symbol or lambda expression). This is useful when function arguments might be quoted with quote rather than function since the body of a lambda expression will be instrumented either way.

lambda-expr

An unquoted anonymous lambda expression.

&optional

All following elements in the specification list are optional; as soon as one does not match, Edebug stops matching at this level. To make just a few elements optional followed by non-optional elements, use [&optional specs...]. To specify that several elements should all succeed together, use &optional [specs...]. See the defun example below.

&rest All following elements in the specification list are repeated zero or more times. All the elements need not match in the last repetition, however. To repeat only a few elements, use [&rest specs...]. To specify all elements must match on every repetition, use &rest [specs...].

&or Each of the following elements in the specification list is an alternative, processed left to right until one matches. One of the alternatives must match otherwise the &or specification fails. Each list element following &or is a single alternative even if it is a keyword specification. (This breaks the implicit grouping rule.) To group two or more list elements as a single alternative, enclose them in [...].

&not Each of the following elements is matched as alternatives as if by using &or, but if any of them match, the specification fails. If none of them match, nothing is matched, but the &not specification succeeds.

&define Indicates that the specification is for a defining form. The defining form itself is not instrumented (i.e. Edebug does not stop before and after the defining form), but forms inside it typically will be instrumented. The &define keyword should be the first element in a list specification. Additional specifications that may only appear after &define are described here. See the defun example below.

name The argument, a symbol, is the name of the defining form. But a defining form need not be named at all, in which case a unique name will be created for it. The name specification may be used more than once in the specification and each subsequent use will append the corresponding symbol argument to the previous name with ‘@’ between them. This is useful for generating unique but meaningful names for definitions such

as defadvice and defmethod.

:name The element following :name should be a symbol; it is used as an additional name component for the definition. This is useful to add a unique, static component to the name of the definition. It may be used more than once. No argument is matched.

arg The argument, a symbol, is the name of an argument of the defining form. However, lambda list keywords (symbols starting with ‘&’) are not allowed. See lambda-list and the example below.

lambda-list

This matches the whole argument list of an XEmacs Lisp lambda expression, which is a list of symbols and the keywords &optional

and &rest

def-body The argument is the body of code in a definition. This is like body, described above, but a definition body must be instrumented with a different Edebug call that looks up information associated with the definition. Use def-body for the highest level list of forms within the definition.

def-form The argument is a single, highest-level form in a definition. This is like def-body, except use this to match a single form rather than a list of forms. As a special case, def-form also means that tracing information is not output when the form is executed. See the interactive example below.

nil This is successful when there are no more arguments to match at the current argument list level; otherwise it fails. See sublist specifications and the backquote example below.

gate No argument is matched but backtracking through the gate is disabled while matching the remainder of the specifications at this level. This is primarily used to generate more specific syntax error messages. See Section 16.4.16.2 [Backtracking], page 244 for more details. Also see the let example below. other-symbol

Any other symbol in a specification list may be a predicate or an indirect specification. If the symbol has an Edebug specification, this indirect specification should be either a list specification that is used in place of the symbol, or a function that is called to process the arguments. The specification may be defined with def-edebug-spec just as for macros. See the defun example below. Otherwise, the symbol should be a predicate. The predicate is called with the argument and the specification fails if the predicate fails. The argument is not instrumented. Predicates that may be used include: symbolp, integerp, stringp, vectorp, atom (which matches a number, string, symbol, or vector), keywordp, and lambda-list-keywordp. The last two, defined in ‘edebug.el’, test whether the argument is a symbol starting with ‘:’ and ‘&’ respectively.

[elements...] Rather than matching a vector argument, a vector treats the elements as a single group specification.

"string" The argument should be a symbol named string. This specification is equivalent to the quoted symbol, 'symbol, where the name of symbol is the string, but the string form is preferred.

'symbol or (quote symbol)

The argument should be the symbol symbol. But use a string specification instead.

(vector elements...)

The argument should be a vector whose elements must match the elements in the specification. See the backquote example below.

(elements...)

Any other list is a sublist specification and the argument must be a list whose elements match the specification elements. A sublist specification may be a dotted list and the corresponding list argument may then be a dotted list. Alternatively, the last cdr of a dotted list specification may be another sublist specification (via a grouping or an indirect specification, e.g. (spec . [(more specs...)])) whose elements match the non-dotted list arguments. This is useful in recursive specifications such as in the backquote example below. Also see the description of a nil specification above for terminating such recursion. Note that a sublist specification of the form (specs . nil) means the same as

(specs), and (specs . (sublist-elements...)) means the same as (specs
sublist-elements...).

16.4.16.2 Backtracking

If a specification fails to match at some point, this does not necessarily mean a syntax error will be signaled; instead, backtracking will take place until all alternatives have been exhausted. Eventually every element of the argument list must be matched by some element in the specification, and every required element in the specification must match some argument.

Backtracking is disabled for the remainder of a sublist or group when certain conditions occur, described below. Backtracking is reenabled when a new alternative is established by &optional, &rest, or &or. It is also reenabled initially when processing a sublist or group specification or an indirect specification.

You might want to disable backtracking to commit to some alternative so that Edebug can provide a more specific syntax error message. Normally, if no alternative matches, Edebug reports that none matched, but if one alternative is committed to, Edebug can report how it failed to match.

First, backtracking is disabled while matching any of the form specifications (i.e. form, body, def-form, and def-body). These specifications will match any form so any error must be in the form itself rather than at a higher level.

Second, backtracking is disabled after successfully matching a quoted symbol or string specification, since this usually indicates a recognized construct. If you have a set of alternative constructs that all begin with the same symbol, you can usually work around this constraint by factoring the symbol out of the alternatives, e.g., ["foo" &or [first case]

[second case] ...].

Third, backtracking may be explicitly disabled by using the gate specification. This is useful when you know that no higher alternatives may apply.

16.4.16.3 Debugging Backquote

Backquote (`) is a macro that results in an expression that may or may not be evaluated. It is often used to simplify the definition of a macro to return an expression that is evaluated, but Edebug does not know when this is the case. However, the forms inside unquotes (, and ,@) are evaluated and Edebug instruments them.

Nested backquotes are supported by Edebug, but there is a limit on the support of quotes inside of backquotes. Quoted forms (with ') are not normally evaluated, but if the quoted form appears immediately within , and ,@ forms, Edebug treats this as a backquoted form at the next higher level (even if there is not a next higher level - this is difficult to fix).

If the backquoted forms happen to be code intended to be evaluated, you can have Edebug instrument them by using edebug-` instead of the regular `. Unquoted forms can always appear inside edebug-` anywhere a form is normally allowed. But (, form) may be used in two other places specially recognized by Edebug: wherever a predicate specification would match, and at the head of a list form in place of a function name or lambda expression. The form inside a spliced unquote, (,@ form), will be wrapped, but the unquote form itself will not be wrapped since this would interfere with the splicing.

There is one other complication with using edebug-`. If the edebug-` call is in a macro and the macro may be called from code that is also instrumented, and if unquoted forms contain any macro arguments bound to instrumented forms, then you should modify the specification for the macro as follows: the specifications for those arguments must use defform instead of form. (This is to reestablish the Edebugging context for those external forms.)

For example, the for macro (see section “Problems with Macros” in XEmacs Lisp Reference Manual) is shown here but with edebug-` substituted for regular `.

     (defmacro inc (var)
       (list 'setq var (list '1+ var)))
     (defmacro for (var from init to final do &rest body)
       (let ((tempvar (make-symbol "max")))
         (edebug-` (let (((, var) (, init))
                         ((, tempvar) (, final)))
                     (while (<= (, var) (, tempvar))
                       (, body)
                       (inc (, var)))))))

Here is the corresponding modified Edebug specification and some code that calls the macro:

     (def-edebug-spec for
       (symbolp "from" def-form "to" def-form "do" &rest def-form))
     (let ((n 5))
       (for i from n to (* n (+ n 1)) do
         (message "%s" i)))

After instrumenting the for macro and the macro call, Edebug first steps to the beginning of the macro call, then into the macro body, then through each of the unquoted expressions in the backquote showing the expressions that will be embedded in the backquote form. Then when the macro expansion is evaluated, Edebug will step through the let form and each time it gets to an unquoted form, it will jump back to an argument of the macro call to step through that expression. Finally stepping will continue after the macro call. Even more convoluted execution paths may result when using anonymous functions.

When the result of an expression is an instrumented expression, it is difficult to see the expression inside the instrumentation. So you may want to set the option edebug-unwrap-results to a non-nil value while debugging such expressions, but it would slow Edebug down to always do this.

16.4.16.4 Specification Examples

Here we provide several examples of Edebug specifications to show many of its capabilities.

A let special form has a sequence of bindings and a body. Each of the bindings is either a symbol or a sublist with a symbol and optional value. In the specification below, notice the gate inside of the sublist to prevent backtracking.

     (def-edebug-spec let
       ((&rest
         &or symbolp (gate symbolp &optional form))
        body))

Edebug uses the following specifications for defun and defmacro and the associated argument list and interactive specifications. It is necessary to handle the expression argument of an interactive form specially since it is actually evaluated outside of the function body.

(def-edebug-spec defmacro defun)      ; Indirect ref to defun spec
(def-edebug-spec defun
  (&define name lambda-list

[&optional stringp] ; Match the doc string, if present.

           [&optional ("interactive" interactive)]
           def-body))
(def-edebug-spec lambda-list
  (([&rest arg]
    [&optional ["&optional" arg &rest arg]]
    &optional ["&rest" arg]
    )))
(def-edebug-spec interactive
  (&optional &or stringp def-form))    ; Notice: def-form

The specification for backquote below illustrates how to match dotted lists and use nil to terminate recursion. It also illustrates how components of a vector may be matched. (The actual specification provided by Edebug does not support dotted lists because doing so causes very deep recursion that could fail.)

(def-edebug-spec ` (backquote-form))  ;; alias just for clarity
(def-edebug-spec backquote-form
  (&or ([&or "," ",@"] &or ("quote" backquote-form) form)
       (backquote-form . [&or nil backquote-form])
       (vector &rest backquote-form)
       sexp))

16.4.17 Edebug Options

These options affect the behavior of Edebug:

User Option edebug-setup-hook

Functions to call before Edebug is used. Each time it is set to a new value, Edebug will call those functions once and then edebug-setup-hook is reset to nil. You could use this to load up Edebug specifications associated with a package you are using but only when you also use Edebug. See Section 16.4.2 [Instrumenting], page 229.

User Option edebug-all-defs

If non-nil, normal evaluation of any defining forms (e.g. defun and defmacro) will instrument them for Edebug. This applies to eval-defun, eval-region,

and eval-current-buffer.

Use the command M-x edebug-all-defs to toggle the value of this variable. You may want to make this variable local to each buffer by calling (make-

local-variable 'edebug-all-defs) in your emacs-lisp-mode-hook. See

Section 16.4.2 [Instrumenting], page 229.

User Option edebug-all-forms

If non-nil, normal evaluation of any forms by eval-defun, eval-region, and eval-current-buffer will instrument them for Edebug. Use the command M-x edebug-all-forms to toggle the value of this option. See Section 16.4.2 [Instrumenting], page 229.

User Option edebug-save-windows

If non-nil, save and restore window configuration on Edebug calls. It takes some time to do this, so if your program does not care what happens to data about windows, you may want to set this variable to nil. If the value is a list, only the listed windows are saved and restored. M-x edebug-toggle-save-windows may be used to change this variable. This command is bound to W in source code buffers. See Section 16.4.15.2 [Edebug Display Update], page 239.

User Option edebug-save-displayed-buffer-points

If non-nil, save and restore point in all displayed buffers. This is necessary if you are debugging code that changes the point of a buffer which is displayed in a non-selected window. If Edebug or the user then selects the window, the buffer’s point will be changed to the window’s point. This is an expensive operation since it visits each window and therefore each displayed buffer twice for each Edebug activation, so it is best to avoid it if you can. See Section 16.4.15.2 [Edebug Display Update], page 239.

User Option edebug-initial-mode

If this variable is non-nil, it specifies the initial execution mode for Edebug when it is first activated. Possible values are step, next, go, Go-nonstop,

trace, Trace-fast, continue, and Continue-fast.

The default value is step. See Section 16.4.3 [Edebug Execution Modes], page 230.

User Option edebug-trace

Non-nil means display a trace of function entry and exit. Tracing output is displayed in a buffer named ‘*edebug-trace*’, one function entry or exit per line, indented by the recursion level. The default value is nil. Also see edebug-tracing. See Section 16.4.13 [Tracing], page 238.

User Option edebug-test-coverage

If non-nil, Edebug tests coverage of all expressions debugged. This is done by comparing the result of each expression with the previous result. Coverage is considered OK if two different results are found. So to sufficiently test the coverage of your code, try to execute it under conditions that evaluate all expressions more than once, and produce different results for each expression. Use M-x edebug-display-freq-count to display the frequency count and coverage information for a definition. See Section 16.4.14 [Coverage Testing], page 238.

User Option edebug-continue-kbd-macro

If non-nil, continue defining or executing any keyboard macro that is executing outside of Edebug. Use this with caution since it is not debugged. See Section 16.4.3 [Edebug Execution Modes], page 230.

User Option edebug-print-length

If non-nil, bind print-length to this while printing results in Edebug. The default value is 50. See Section 16.4.12 [Printing in Edebug], page 237.

User Option edebug-print-level

If non-nil, bind print-level to this while printing results in Edebug. The default value is 50.

User Option edebug-print-circle

If non-nil, bind print-circle to this while printing results in Edebug. The default value is nil.

User Option edebug-on-error

debug-on-error is bound to this while Edebug is active. See Section 16.4.7 [Trapping Errors], page 234.

User Option edebug-on-quit

debug-on-quit is bound to this while Edebug is active. See Section 16.4.7 [Trapping Errors], page 234.

User Option edebug-unwrap-results

Non-nil if Edebug should unwrap results of expressions. This is useful when debugging macros where the results of expressions are instrumented expressions. But don’t do this when results might be circular or an infinite loop will result. See Section 16.4.16.3 [Debugging Backquote], page 245.

User Option edebug-global-break-condition

If non-nil, an expression to test for at every stop point. If the result is non-nil, then break. Errors are ignored. See Section 16.4.6.1 [Global Break Condition], page 233.

250 XEmacs Lisp Reference Manual

17 Reading and Printing Lisp Objects

Printing and reading are the operations of converting Lisp objects to textual form and vice versa. They use the printed representations and read syntax described in Chapter 2 [Lisp Data Types], page 15.

This chapter describes the Lisp functions for reading and printing. It also describes streams, which specify where to get the text (if reading) or where to put it (if printing).

17.1 Introduction to Reading and Printing

Reading a Lisp object means parsing a Lisp expression in textual form and producing a corresponding Lisp object. This is how Lisp programs get into Lisp from files of Lisp code. We call the text the read syntax of the object. For example, the text ‘(a . 5)’ is the read syntax for a cons cell whose car is a and whose cdr is the number 5.

Printing a Lisp object means producing text that represents that object—converting the object to its printed representation. Printing the cons cell described above produces the text ‘(a . 5)’.

Reading and printing are more or less inverse operations: printing the object that results from reading a given piece of text often produces the same text, and reading the text that results from printing an object usually produces a similar-looking object. For example, printing the symbol foo produces the text ‘foo’, and reading that text returns the symbol foo. Printing a list whose elements are a and b produces the text ‘(a b)’, and reading that text produces a list (but not the same list) with elements a and b.

However, these two operations are not precisely inverses. There are three kinds of exceptions:

• Printing can produce text that cannot be read. For example, buffers, windows, frames, subprocesses and markers print into text that starts with ‘#’; if you try to read this text, you get an error. There is no way to read those data types.

• One object can have multiple textual representations. For example, ‘1’ and ‘01’ represent the same integer, and ‘(a b)’ and ‘(a . (b))’ represent the same list. Reading will accept any of the alternatives, but printing must choose one of them.

• Comments can appear at certain points in the middle of an object’s read sequence without affecting the result of reading it.

17.2 Input Streams

Most of the Lisp functions for reading text take an input stream as an argument. The input stream specifies where or how to get the characters of the text to be read. Here are the possible types of input stream:

buffer The input characters are read from buffer, starting with the character directly after point. Point advances as characters are read.

marker The input characters are read from the buffer that marker is in, starting with the character directly after the marker. The marker position advances as characters are read. The value of point in the buffer has no effect when the stream is a marker.

string The input characters are taken from string, starting at the first character in the string and using as many characters as required.

function The input characters are generated by function, one character per call. Normally function is called with no arguments, and should return a character. Occasionally function is called with one argument (always a character). When that happens, function should save the argument and arrange to return it on the next call. This is called unreading the character; it happens when the Lisp reader reads one character too many and wants to “put it back where it came from”.

t t used as a stream means that the input is read from the minibuffer. In fact, the minibuffer is invoked once and the text given by the user is made into a string that is then used as the input stream.

nil nil supplied as an input stream means to use the value of standard-input instead; that value is the default input stream, and must be a non-nil input stream.

symbol A symbol as input stream is equivalent to the symbol’s function definition (if any).

Here is an example of reading from a stream that is a buffer, showing where point is located before and after:

---------- Buffer: foo ----------
This⋆ is the contents of foo.
---------- Buffer: foo ----------
(read (get-buffer "foo"))
⇒ is
(read (get-buffer "foo"))
⇒ the
---------- Buffer: foo ----------
This is the⋆ contents of foo.
---------- Buffer: foo ----------

Note that the first read skips a space. Reading skips any amount of whitespace preceding the significant text.

In Emacs 18, reading a symbol discarded the delimiter terminating the symbol. Thus, point would end up at the beginning of ‘contents’ rather than after ‘the’. The Emacs 19 behavior is superior because it correctly handles input such as ‘bar(foo)’, where the open-parenthesis that ends one object is needed as the beginning of another object.

Here is an example of reading from a stream that is a marker, initially positioned at the beginning of the buffer shown. The value read is the symbol This.

---------- Buffer: foo ----------
This is the contents of foo.
---------- Buffer: foo ----------

(setq m (set-marker (make-marker) 1 (get-buffer "foo")))
     ⇒ #<marker at 1 in foo>
(read m) ⇒ This
m

⇒ #<marker at 5 in foo> ;; Before the first space. Here we read from the contents of a string:

(read "(When in) the course")
     ⇒ (When in)

The following example reads from the minibuffer. The prompt is: ‘Lisp expression: ’. (That is always the prompt used when you read from the stream t.) The user’s input is shown following the prompt.

(read t)
     ⇒ 23
---------- Buffer: Minibuffer ----------
Lisp expression: 23 ^{⟨}\underline{\overline{RET}}^{⟩}
---------- Buffer: Minibuffer ----------

Finally, here is an example of a stream that is a function, named useless-stream. Before we use the stream, we initialize the variable useless-list to a list of characters. Then each call to the function useless-stream obtains the next character in the list or unreads a character by adding it to the front of the list.

(setq useless-list (append "XY()" nil))
     ⇒ (88 89 40 41)
(defun useless-stream (&optional unread)
  (if unread
      (setq useless-list (cons unread useless-list))
    (prog1 (car useless-list)
           (setq useless-list (cdr useless-list)))))
     ⇒ useless-stream

Now we read using the stream thus constructed:

(read 'useless-stream)
     ⇒ XY
useless-list
     ⇒ (40 41)

Note that the open and close parentheses remains in the list. The Lisp reader encountered the open parenthesis, decided that it ended the input, and unread it. Another attempt to read from the stream at this point would read ‘()’ and return nil.

17.3 Input Functions

This section describes the Lisp functions and variables that pertain to reading. In the functions below, stream stands for an input stream (see the previous section). If

stream is nil or omitted, it defaults to the value of standard-input.

An end-of-file error is signaled if reading encounters an unterminated list, vector, or string.

Function read &optional stream This function reads one textual Lisp expression from stream, returning it as a Lisp object. This is the basic Lisp input function.

Function read-from-string string &optional start end This function reads the first textual Lisp expression from the text in string. It returns a cons cell whose car is that expression, and whose cdr is an integer giving the position of the next remaining character in the string (i.e., the first one not read). If start is supplied, then reading begins at index start in the string (where the first character is at index 0). If end is also supplied, then reading stops just before that index, as if the rest of the string were not there. For example:

(read-from-string "(setq x 55) (setq y 5)")
     ⇒ ((setq x 55) . 11)
(read-from-string "\"A short string\"")
     ⇒ ("A short string" . 16)

;; Read starting at the first character.

(read-from-string "(list 112)" 0)
     ⇒ ((list 112) . 10)

;; Read starting at the second character.

(read-from-string "(list 112)" 1)
     ⇒ (list . 5)

;; Read starting at the seventh character, ;; and stopping at the ninth.

(read-from-string "(list 112)" 6 8)
     ⇒ (11 . 8)

Variable standard-input

This variable holds the default input stream—the stream that read uses when the stream argument is nil.

17.4 Output Streams

An output stream specifies what to do with the characters produced by printing. Most print functions accept an output stream as an optional argument. Here are the possible types of output stream:

buffer The output characters are inserted into buffer at point. Point advances as characters are inserted.

marker The output characters are inserted into the buffer that marker points into, at the marker position. The marker position advances as characters are inserted. The value of point in the buffer has no effect on printing when the stream is a marker. function The output characters are passed to function, which is responsible for storing

them away. It is called with a single character as argument, as many times as there are characters to be output, and is free to do anything at all with the characters it receives.

t The output characters are displayed in the echo area.

nil nil specified as an output stream means to the value of standard-output instead; that value is the default output stream, and must be a non-nil output stream.

symbol A symbol as output stream is equivalent to the symbol’s function definition (if any).

Many of the valid output streams are also valid as input streams. The difference between input and output streams is therefore mostly one of how you use a Lisp object, not a distinction of types of object.

Here is an example of a buffer used as an output stream. Point is initially located as shown immediately before the ‘h’ in ‘the’. At the end, point is located directly before that same ‘h’.

---------- Buffer: foo ----------
This is t⋆he contents of foo.
---------- Buffer: foo ----------
(print "This is the output" (get-buffer "foo"))
     ⇒ "This is the output"
---------- Buffer: foo ----------
This is t
"This is the output"
⋆he contents of foo.
---------- Buffer: foo ----------

Now we show a use of a marker as an output stream. Initially, the marker is in buffer foo, between the ‘t’ and the ‘h’ in the word ‘the’. At the end, the marker has advanced over the inserted text so that it remains positioned before the same ‘h’. Note that the location of point, shown in the usual fashion, has no effect.

---------- Buffer: foo ----------
"This is the ⋆output"
---------- Buffer: foo ----------
m
     ⇒ #<marker at 11 in foo>
(print "More output for foo." m)
     ⇒ "More output for foo."
---------- Buffer: foo ----------
"This is t
"More output for foo."
he ⋆output"
---------- Buffer: foo ----------

     m
          ⇒ #<marker at 35 in foo>

The following example shows output to the echo area:

     (print "Echo Area output" t)
          ⇒ "Echo Area output"
     ---------- Echo Area ----------
     "Echo Area output"
     ---------- Echo Area ----------

Finally, we show the use of a function as an output stream. The function eat-output takes each character that it is given and conses it onto the front of the list last-output (see Section 5.5 [Building Lists], page 82). At the end, the list contains all the characters output, but in reverse order.

     (setq last-output nil)
          ⇒ nil
     (defun eat-output (c)
       (setq last-output (cons c last-output)))
          ⇒ eat-output
     (print "This is the output" 'eat-output)
          ⇒ "This is the output"
     last-output
          ⇒ (?\n ?\" ?t ?u ?p ?t ?u ?o ?\  ?e ?h ?t
                     ?\  ?s ?i ?\  ?s ?i ?h ?T ?\" ?\n)

Now we can put the output in the proper order by reversing the list:

     (concat (nreverse last-output))
          ⇒ "
     \"This is the output\"
     "

Calling concat converts the list to a string so you can see its contents more clearly.

17.5 Output Functions

This section describes the Lisp functions for printing Lisp objects.

Some of the XEmacs printing functions add quoting characters to the output when necessary so that it can be read properly. The quoting characters used are ‘"’ and ‘\’; they distinguish strings from symbols, and prevent punctuation characters in strings and symbols from being taken as delimiters when reading. See Section 2.1 [Printed Representation], page 15, for full details. You specify quoting or no quoting by the choice of printing function.

If the text is to be read back into Lisp, then it is best to print with quoting characters to avoid ambiguity. Likewise, if the purpose is to describe a Lisp object clearly for a Lisp programmer. However, if the purpose of the output is to look nice for humans, then it is better to print without quoting.

Printing a self-referent Lisp object requires an infinite amount of text. In certain cases, trying to produce this text leads to a stack overflow. XEmacs detects such recursion and prints ‘#level’ instead of recursively printing an object already being printed. For example, here ‘#0’ indicates a recursive reference to the object at level 0 of the current print operation:

(setq foo (list nil))
     ⇒ (nil)
(setcar foo foo)
     ⇒ (#0)

In the functions below, stream stands for an output stream. (See the previous section for a description of output streams.) If stream is nil or omitted, it defaults to the value of

standard-output.

Function print object &optional stream The print function is a convenient way of printing. It outputs the printed representation of object to stream, printing in addition one newline before object and another after it. Quoting characters are used. print returns object. For example:

     (progn (print 'The\ cat\ in)
            (print "the hat")
            (print " came back"))

⊣

          ⊣ The\ cat\ in

⊣

          ⊣ "the hat"

⊣

          ⊣ " came back"

⊣

          ⇒ " came back"

Function prin1 object &optional stream This function outputs the printed representation of object to stream. It does not print newlines to separate output as print does, but it does use quoting characters just like print. It returns object.

     (progn (prin1 'The\ cat\ in)
            (prin1 "the hat")
            (prin1 " came back"))
          ⊣ The\ cat\ in"the hat"" came back"
          ⇒ " came back"

Function princ object &optional stream This function outputs the printed representation of object to stream. It returns object. This function is intended to produce output that is readable by people, not by read, so it doesn’t insert quoting characters and doesn’t put double-quotes around the contents of strings. It does not add any spacing between calls.

     (progn
       (princ 'The\ cat)
       (princ " in the \"hat\""))
          ⊣ The cat in the "hat"
          ⇒ " in the \"hat\""

Function terpri &optional stream This function outputs a newline to stream. The name stands for “terminate print”.

Function write-char character &optional stream This function outputs character to stream. It returns character.

Function prin1-to-string object &optional noescape This function returns a string containing the text that prin1 would have printed for the same argument.

     (prin1-to-string 'foo)
          ⇒ "foo"
     (prin1-to-string (mark-marker))
          ⇒ "#<marker at 2773 in strings.texi>"

If noescape is non-nil, that inhibits use of quoting characters in the output. (This argument is supported in Emacs versions 19 and later.)

     (prin1-to-string "foo")
          ⇒ "\"foo\""
     (prin1-to-string "foo" t)
          ⇒ "foo"

See format, in Section 4.7 [String Conversion], page 65, for other ways to obtain the printed representation of a Lisp object as a string.

17.6 Variables Affecting Output

Variable standard-output

The value of this variable is the default output stream—the stream that print functions use when the stream argument is nil.

Variable print-escape-newlines

If this variable is non-nil, then newline characters in strings are printed as ‘\n’ and formfeeds are printed as ‘\f’. Normally these characters are printed as actual newlines and formfeeds. This variable affects the print functions prin1 and print, as well as everything that uses them. It does not affect princ. Here is an example using prin1:

     (prin1 "a\nb")
          ⊣ "a

          ⊣ b"
          ⇒ "a

     b"
     (let ((print-escape-newlines t))
       (prin1 "a\nb"))
          ⊣ "a\nb"
          ⇒ "a
     b"

In the second expression, the local binding of print-escape-newlines is in effect during the call to prin1, but not during the printing of the result.

Variable print-readably

If non-nil, then all objects will be printed in a readable form. If an object has no readable representation, then an error is signalled. When print-readably is true, compiled-function objects will be written in ‘#[...]’ form instead of in ‘#<compiled-function [...]>’ form, and two-element lists of the form ‘(quote object)’ will be written as the equivalent ‘'object’. Do not set this variable; bind it instead.

Variable print-length

The value of this variable is the maximum number of elements of a list that will be printed. If a list being printed has more than this many elements, it is abbreviated with an ellipsis. If the value is nil (the default), then there is no limit.

     (setq print-length 2)
          ⇒ 2
     (print '(1 2 3 4 5))
          ⊣ (1 2 ...)

          ⇒ (1 2 ...)

Variable print-level

The value of this variable is the maximum depth of nesting of parentheses and brackets when printed. Any list or vector at a depth exceeding this limit is abbreviated with an ellipsis. A value of nil (which is the default) means no limit. This variable exists in version 19 and later versions.

Variable print-string-length

The value of this variable is the maximum number of characters of a string that will be printed. If a string being printed has more than this many characters, it is abbreviated with an ellipsis.

Variable print-gensym

If non-nil, then uninterned symbols will be printed specially. Uninterned symbols are those which are not present in obarray, that is, those which were made with make-symbol or by calling intern with a second argument. When print-gensym is true, such symbols will be preceded by ‘#:’, which causes the reader to create a new symbol instead of interning and returning an existing one. Beware: The ‘#:’ syntax creates a new symbol each time it is seen, so if you print an object which contains two pointers to the same uninterned symbol, read will not duplicate that structure. Also, since XEmacs has no real notion of packages, there is no way for the printer to distinguish between symbols interned in no obarray, and symbols interned in an alternate obarray.

Variable float-output-format

This variable holds the format descriptor string that Lisp uses to print floats. This is a ‘%’-spec like those accepted by printf in C, but with some restrictions.

It must start with the two characters ‘%.’. After that comes an integer precision specification, and then a letter which controls the format. The letters allowed are ‘e’, ‘f’ and ‘g’.

• Use ‘e’ for exponential notation ‘dig.digitseexpt’.

• Use ‘f’ for decimal point notation ‘DIGITS.DIGITS’.

• Use ‘g’ to choose the shorter of those two formats for the number at hand.

The precision in any of these cases is the number of digits following the decimal point. With ‘f’, a precision of 0 means to omit the decimal point. 0 is not allowed with ‘f’ or ‘g’. A value of nil means to use ‘%.16g’. Regardless of the value of float-output-format, a floating point number will never be printed in such a way that it is ambiguous with an integer; that is, a floating-point number will always be printed with a decimal point and/or an exponent, even if the digits following the decimal point are all zero. This is to preserve read-equivalence.

18 Minibuffers

A minibuffer is a special buffer that XEmacs commands use to read arguments more complicated than the single numeric prefix argument. These arguments include file names, buffer names, and command names (as in M-x). The minibuffer is displayed on the bottom line of the frame, in the same place as the echo area, but only while it is in use for reading an argument.

18.1 Introduction to Minibuffers

In most ways, a minibuffer is a normal XEmacs buffer. Most operations within a buffer, such as editing commands, work normally in a minibuffer. However, many operations for managing buffers do not apply to minibuffers. The name of a minibuffer always has the form ‘ *Minibuf-number’, and it cannot be changed. Minibuffers are displayed only in special windows used only for minibuffers; these windows always appear at the bottom of a frame. (Sometime frames have no minibuffer window, and sometimes a special kind of frame contains nothing but a minibuffer window; see Section 32.7 [Minibuffers and Frames], page 478.)

The minibuffer’s window is normally a single line. You can resize it temporarily with the window sizing commands; it reverts to its normal size when the minibuffer is exited. You can resize it permanently by using the window sizing commands in the frame’s other window, when the minibuffer is not active. If the frame contains just a minibuffer, you can change the minibuffer’s size by changing the frame’s size.

If a command uses a minibuffer while there is an active minibuffer, this is called a recursive minibuffer. The first minibuffer is named ‘ *Minibuf-0*’. Recursive minibuffers are named by incrementing the number at the end of the name. (The names begin with a space so that they won’t show up in normal buffer lists.) Of several recursive minibuffers, the innermost (or most recently entered) is the active minibuffer. We usually call this “the” minibuffer. You can permit or forbid recursive minibuffers by setting the variable

enable-recursive-minibuffers.

Like other buffers, a minibuffer may use any of several local keymaps (see Chapter 20 [Keymaps], page 315); these contain various exit commands and in some cases completion commands (see Section 18.5 [Completion], page 266).

• minibuffer-local-map is for ordinary input (no completion).

• minibuffer-local-ns-map is similar, except that ^{⟨}\underline{\overline{SPC}}^{⟩} exits just like ^{⟨}\underline{\overline{RET}}^{⟩}. This is used mainly for Mocklisp compatibility.

  • minibuffer-local-completion-map is for permissive completion.

• minibuffer-local-must-match-map is for strict completion and for cautious completion.

18.2 Reading Text Strings with the Minibuffer

Most often, the minibuffer is used to read text as a string. It can also be used to read a Lisp object in textual form. The most basic primitive for minibuffer input is read-from-minibuffer; it can do either one.

In most cases, you should not call minibuffer input functions in the middle of a Lisp function. Instead, do all minibuffer input as part of reading the arguments for a command, in the interactive spec. See Section 19.2 [Defining Commands], page 282.

Function read-from-minibuffer prompt-string &optional initial-contents

keymap read hist This function is the most general way to get input through the minibuffer. By default, it accepts arbitrary text and returns it as a string; however, if read is non-nil, then it uses read to convert the text into a Lisp object (see Section 17.3 [Input Functions], page 253). The first thing this function does is to activate a minibuffer and display it with prompt-string as the prompt. This value must be a string. Then, if initial-contents is a string, read-from-minibuffer inserts it into the minibuffer, leaving point at the end. The minibuffer appears with this text as its contents. The value of initial-contents may also be a cons cell of the form (string . position). This means to insert string in the minibuffer but put point position characters from the beginning, rather than at the end. If keymap is non-nil, that keymap is the local keymap to use in the minibuffer. If keymap is omitted or nil, the value of minibuffer-local-map is used as the keymap. Specifying a keymap is the most important way to customize the minibuffer for various applications such as completion. The argument hist specifies which history list variable to use for saving the input and for history commands used in the minibuffer. It defaults to minibufferhistory. See Section 18.4 [Minibuffer History], page 264. When the user types a command to exit the minibuffer, read-from-minibuffer uses the text in the minibuffer to produce its return value. Normally it simply makes a string containing that text. However, if read is non-nil, read-from-minibuffer reads the text and returns the resulting Lisp object, unevaluated. (See Section 17.3 [Input Functions], page 253, for information about reading.)

Function read-string prompt &optional initial This function reads a string from the minibuffer and returns it. The arguments prompt and initial are used as in read-from-minibuffer. The keymap used

is minibuffer-local-map.

This is a simplified interface to the read-from-minibuffer function:

     (read-string prompt initial)

≡

     (read-from-minibuffer prompt initial nil nil nil)

Variable minibuffer-local-map

This is the default local keymap for reading from the minibuffer. By default, it makes the following bindings:

⟨\underline{\overline{LFD}}^{⟩}  exit-minibuffer
⟨\underline{\overline{RET}}^{⟩}  exit-minibuffer

C-g       abort-recursive-edit
M-n       next-history-element
M-p       previous-history-element
M-r       next-matching-history-element
M-s       previous-matching-history-element

Function read-no-blanks-input prompt &optional initial This function reads a string from the minibuffer, but does not allow whitespace characters as part of the input: instead, those characters terminate the input. The arguments prompt and initial are used as in read-from-minibuffer. This is a simplified interface to the read-from-minibuffer function, and passes the value of the minibuffer-local-ns-map keymap as the keymap argument for that function. Since the keymap minibuffer-local-ns-map does not rebind C-q, it is possible to put a space into the string, by quoting it.

     (read-no-blanks-input prompt initial)

≡

     (read-from-minibuffer prompt initial minibuffer-local-ns-map)

Variable minibuffer-local-ns-map

This built-in variable is the keymap used as the minibuffer local keymap in the function read-no-blanks-input. By default, it makes the following bindings,

in addition to those of minibuffer-local-map:
⟨\underline{\overline{SPC}}^{⟩}  exit-minibuffer
⟨\underline{\overline{TAB}}^{⟩}  exit-minibuffer
?         self-insert-and-exit

18.3 Reading Lisp Objects with the Minibuffer

This section describes functions for reading Lisp objects with the minibuffer.

Function read-minibuffer prompt &optional initial This function reads a Lisp object in the minibuffer and returns it, without evaluating it. The arguments prompt and initial are used as in read-from-

minibuffer.

This is a simplified interface to the read-from-minibuffer function:

     (read-minibuffer prompt initial)

≡

     (read-from-minibuffer prompt initial nil t)

Here is an example in which we supply the string "(testing)" as initial input:

     (read-minibuffer
      "Enter an expression: " (format "%s" '(testing)))

;; Here is how the minibuffer is displayed:

     ---------- Buffer: Minibuffer ----------
     Enter an expression: (testing)⋆
     ---------- Buffer: Minibuffer ----------

The user can type ^{⟨}\underline{\overline{RET}}^{⟩} immediately to use the initial input as a default, or can edit the input.

Function eval-minibuffer prompt &optional initial This function reads a Lisp expression in the minibuffer, evaluates it, then returns the result. The arguments prompt and initial are used as in read-from-

minibuffer.

This function simply evaluates the result of a call to read-minibuffer:

     (eval-minibuffer prompt initial)

≡

     (eval (read-minibuffer prompt initial))

Function edit-and-eval-command prompt form

This function reads a Lisp expression in the minibuffer, and then evaluates it. The difference between this command and eval-minibuffer is that here the initial form is not optional and it is treated as a Lisp object to be converted to printed representation rather than as a string of text. It is printed with prin1, so if it is a string, double-quote characters (‘"’) appear in the initial text. See Section 17.5 [Output Functions], page 256. The first thing edit-and-eval-command does is to activate the minibuffer with prompt as the prompt. Then it inserts the printed representation of form in the minibuffer, and lets the user edit. When the user exits the minibuffer, the edited text is read with read and then evaluated. The resulting value becomes

the value of edit-and-eval-command.

In the following example, we offer the user an expression with initial text which is a valid form already:

     (edit-and-eval-command "Please edit: " '(forward-word 1))

;; After evaluation of the preceding expression, ;; the following appears in the minibuffer:

     ---------- Buffer: Minibuffer ----------
     Please edit: (forward-word 1)⋆
     ---------- Buffer: Minibuffer ----------

Typing ^{⟨}\underline{\overline{RET}}^{⟩} right away would exit the minibuffer and evaluate the expression, thus moving point forward one word. edit-and-eval-command returns t in this example.

18.4 Minibuffer History

A minibuffer history list records previous minibuffer inputs so the user can reuse them conveniently. A history list is actually a symbol, not a list; it is a variable whose value is a list of strings (previous inputs), most recent first.

There are many separate history lists, used for different kinds of inputs. It’s the Lisp programmer’s job to specify the right history list for each use of the minibuffer.

   The basic minibuffer input functions read-from-minibuffer and completing-read

both accept an optional argument named hist which is how you specify the history list. Here are the possible values:

variable Use variable (a symbol) as the history list.

(variable . startpos)

Use variable (a symbol) as the history list, and assume that the initial history position is startpos (an integer, counting from zero which specifies the most recent element of the history). If you specify startpos, then you should also specify that element of the history as the initial minibuffer contents, for consistency.

If you don’t specify hist, then the default history list minibuffer-history is used. For other standard history lists, see below. You can also create your own history list variable; just initialize it to nil before the first use.

   Both read-from-minibuffer and completing-read add new elements to the history

list automatically, and provide commands to allow the user to reuse items on the list. The only thing your program needs to do to use a history list is to initialize it and to pass its name to the input functions when you wish. But it is safe to modify the list by hand when the minibuffer input functions are not using it.

Variable minibuffer-history

The default history list for minibuffer history input.

Variable query-replace-history

A history list for arguments to query-replace (and similar arguments to other commands).

Variable file-name-history

A history list for file name arguments.

Variable regexp-history

A history list for regular expression arguments.

Variable extended-command-history

A history list for arguments that are names of extended commands.

Variable shell-command-history

A history list for arguments that are shell commands.

Variable read-expression-history

A history list for arguments that are Lisp expressions to evaluate.

Variable Info-minibuffer-history

A history list for Info mode’s minibuffer.

Variable Manual-page-minibuffer-history

A history list for manual-entry.

There are many other minibuffer history lists, defined by various libraries. An M-x apropos search for ‘history’ should prove fruitful in discovering them.

18.5 Completion

Completion is a feature that fills in the rest of a name starting from an abbreviation for it. Completion works by comparing the user’s input against a list of valid names and determining how much of the name is determined uniquely by what the user has typed. For example, when you type C-x b (switch-to-buffer) and then type the first few letters of the name of the buffer to which you wish to switch, and then type ^{⟨}\underline{\overline{TAB}}^{⟩} (minibuffercomplete), Emacs extends the name as far as it can.

Standard XEmacs commands offer completion for names of symbols, files, buffers, and processes; with the functions in this section, you can implement completion for other kinds of names.

The try-completion function is the basic primitive for completion: it returns the longest determined completion of a given initial string, with a given set of strings to match against.

The function completing-read provides a higher-level interface for completion. A call to completing-read specifies how to determine the list of valid names. The function then activates the minibuffer with a local keymap that binds a few keys to commands useful for completion. Other functions provide convenient simple interfaces for reading certain kinds of names with completion.

18.5.1 Basic Completion Functions

The two functions try-completion and all-completions have nothing in themselves to do with minibuffers. We describe them in this chapter so as to keep them near the higher-level completion features that do use the minibuffer.

Function try-completion string collection &optional predicate This function returns the longest common substring of all possible completions of string in collection. The value of collection must be an alist, an obarray, or a function that implements a virtual set of strings (see below).

Completion compares string against each of the permissible completions specified by collection; if the beginning of the permissible completion equals string, it matches. If no permissible completions match, try-completion returns nil. If only one permissible completion matches, and the match is exact, then trycompletion returns t. Otherwise, the value is the longest initial sequence common to all the permissible completions that match.

If collection is an alist (see Section 5.8 [Association Lists], page 92), the cars of the alist elements form the set of permissible completions.

If collection is an obarray (see Section 7.3 [Creating Symbols], page 113), the names of all symbols in the obarray form the set of permissible completions. The global variable obarray holds an obarray containing the names of all interned Lisp symbols.

Note that the only valid way to make a new obarray is to create it empty and then add symbols to it one by one using intern. Also, you cannot intern a given symbol in more than one obarray.

If the argument predicate is non-nil, then it must be a function of one argument. It is used to test each possible match, and the match is accepted only if predicate returns non-nil. The argument given to predicate is either a cons cell from the alist (the car of which is a string) or else it is a symbol (not a symbol name) from the obarray.

You can also use a symbol that is a function as collection. Then the function is solely responsible for performing completion; try-completion returns whatever this function returns. The function is called with three arguments: string, predicate and nil. (The reason for the third argument is so that the same function can be used in all-completions and do the appropriate thing in either case.) See Section 18.5.6 [Programmed Completion], page 274.

In the first of the following examples, the string ‘foo’ is matched by three of the alist cars. All of the matches begin with the characters ‘fooba’, so that is the result. In the second example, there is only one possible match, and it is exact, so the value is t.

     (try-completion
      "foo"
      '(("foobar1" 1) ("barfoo" 2) ("foobaz" 3) ("foobar2" 4)))
          ⇒ "fooba"
     (try-completion "foo" '(("barfoo" 2) ("foo" 3)))
          ⇒ t

In the following example, numerous symbols begin with the characters ‘forw’, and all of them begin with the word ‘forward’. In most of the symbols, this is followed with a ‘-’, but not in all, so no more than ‘forward’ can be completed.

     (try-completion "forw" obarray)
          ⇒ "forward"

Finally, in the following example, only two of the three possible matches pass the predicate test (the string ‘foobaz’ is too short). Both of those begin with the string ‘foobar’.

     (defun test (s)
       (> (length (car s)) 6))
          ⇒ test
     (try-completion
      "foo"
      '(("foobar1" 1) ("barfoo" 2) ("foobaz" 3) ("foobar2" 4))
      'test)
          ⇒ "foobar"

Function all-completions string collection &optional predicate nospace This function returns a list of all possible completions of string. The parameters to this function are the same as to try-completion.

If collection is a function, it is called with three arguments: string, predicate and t; then all-completions returns whatever the function returns. See Section 18.5.6 [Programmed Completion], page 274. If nospace is non-nil, completions that start with a space are ignored unless string also starts with a space. Here is an example, using the function test shown in the example for try-

completion:
     (defun test (s)
       (> (length (car s)) 6))
          ⇒ test
     (all-completions
      "foo"
      '(("foobar1" 1) ("barfoo" 2) ("foobaz" 3) ("foobar2" 4))
      'test)
          ⇒ ("foobar1" "foobar2")

Variable completion-ignore-case

If the value of this variable is non-nil, XEmacs does not consider case significant in completion.

18.5.2 Completion and the Minibuffer

This section describes the basic interface for reading from the minibuffer with completion.

Function completing-read prompt collection &optional predicate

require-match initial hist This function reads a string in the minibuffer, assisting the user by providing completion. It activates the minibuffer with prompt prompt, which must be a string. If initial is non-nil, completing-read inserts it into the minibuffer as part of the input. Then it allows the user to edit the input, providing several commands to attempt completion. The actual completion is done by passing collection and predicate to the function try-completion. This happens in certain commands bound in the local keymaps used for completion. If require-match is t, the usual minibuffer exit commands won’t exit unless the input completes to an element of collection. If require-match is neither nil nor t, then the exit commands won’t exit unless the input typed is itself an element of collection. If require-match is nil, the exit commands work regardless of the input in the minibuffer. The user can exit with null input by typing ^{⟨}\underline{\overline{RET}}^{⟩} with an empty minibuffer. Then completing-read returns nil. This is how the user requests whatever default the command uses for the value being read. The user can return using ⟨\underline{\overline{RET}}^{⟩} in this way regardless of the value of require-match. The function completing-read works by calling read-minibuffer. It uses

minibuffer-local-completion-map as the keymap if require-match is nil,

and uses minibuffer-local-must-match-map if require-match is non-nil. See

Section 18.5.3 [Completion Commands], page 269. The argument hist specifies which history list variable to use for saving the input and for minibuffer history commands. It defaults to minibuffer-history. See Section 18.4 [Minibuffer History], page 264. Completion ignores case when comparing the input against the possible matches, if the built-in variable completion-ignore-case is non-nil. See Section 18.5.1 [Basic Completion], page 266. Here’s an example of using completing-read:

     (completing-read
      "Complete a foo: "
      '(("foobar1" 1) ("barfoo" 2) ("foobaz" 3) ("foobar2" 4))
      nil t "fo")

;; After evaluation of the preceding expression, ;; the following appears in the minibuffer:

     ---------- Buffer: Minibuffer ----------
     Complete a foo: fo⋆
     ---------- Buffer: Minibuffer ----------

If the user then types ^{⟨}\underline{\overline{DEL}}^{⟩} ^{⟨}\underline{\overline{DEL}}^{⟩} b ^{⟨}\underline{\overline{RET}}^{⟩}, completing-read returns barfoo. The completing-read function binds three variables to pass information to the commands that actually do completion. These variables are minibuffer-

completion-table, minibuffer-completion-predicate and minibuffer-

completion-confirm. For more information about them, see Section 18.5.3 [Completion Commands], page 269.

18.5.3 Minibuffer Commands That Do Completion

This section describes the keymaps, commands and user options used in the minibuffer to do completion.

Variable minibuffer-local-completion-map

completing-read uses this value as the local keymap when an exact match of one of the completions is not required. By default, this keymap makes the following bindings:

?         minibuffer-completion-help
⟨\underline{\overline{SPC}}^{⟩}  minibuffer-complete-word
⟨\underline{\overline{TAB}}^{⟩}  minibuffer-complete

with other characters bound as in minibuffer-local-map (see Section 18.2 [Text from Minibuffer], page 261).

Variable minibuffer-local-must-match-map

completing-read uses this value as the local keymap when an exact match of one of the completions is required. Therefore, no keys are bound to exitminibuffer, the command that exits the minibuffer unconditionally. By default, this keymap makes the following bindings:

?         minibuffer-completion-help
⟨\underline{\overline{SPC}}^{⟩}  minibuffer-complete-word
⟨\underline{\overline{TAB}}^{⟩}  minibuffer-complete
⟨\underline{\overline{LFD}}^{⟩}  minibuffer-complete-and-exit
⟨\underline{\overline{RET}}^{⟩}  minibuffer-complete-and-exit

with other characters bound as in minibuffer-local-map.

Variable minibuffer-completion-table

The value of this variable is the alist or obarray used for completion in the minibuffer. This is the global variable that contains what completing-read passes to try-completion. It is used by minibuffer completion commands

such as minibuffer-complete-word.

Variable minibuffer-completion-predicate

This variable’s value is the predicate that completing-read passes to trycompletion. The variable is also used by the other minibuffer completion functions.

Command minibuffer-complete-word

This function completes the minibuffer contents by at most a single word. Even if the minibuffer contents have only one completion, minibuffer-complete-word does not add any characters beyond the first character that is not a word constituent. See Chapter 38 [Syntax Tables], page 569.

Command minibuffer-complete

This function completes the minibuffer contents as far as possible.

Command minibuffer-complete-and-exit

This function completes the minibuffer contents, and exits if confirmation is not required, i.e., if minibuffer-completion-confirm is non-nil. If confirmation is required, it is given by repeating this command immediately—the command is programmed to work without confirmation when run twice in succession.

Variable minibuffer-completion-confirm

When the value of this variable is non-nil, XEmacs asks for confirmation of a completion before exiting the minibuffer. The function minibuffer-complete-and-exit checks the value of this variable before it exits.

Command minibuffer-completion-help

This function creates a list of the possible completions of the current minibuffer contents. It works by calling all-completions using the value of the variable minibuffer-completion-table as the collection argument, and the value of minibuffer-completion-predicate as the predicate argument. The list of completions is displayed as text in a buffer named ‘*Completions*’.

Function display-completion-list completions

This function displays completions to the stream in standard-output, usually a buffer. (See Chapter 17 [Read and Print], page 251, for more information about streams.) The argument completions is normally a list of completions just returned by all-completions, but it does not have to be. Each element may be a symbol or a string, either of which is simply printed, or a list of two strings, which is printed as if the strings were concatenated. This function is called by minibuffer-completion-help. The most common way to use it is together with with-output-to-temp-buffer, like this:

     (with-output-to-temp-buffer "*Completions*"
       (display-completion-list
         (all-completions (buffer-string) my-alist)))

User Option completion-auto-help

If this variable is non-nil, the completion commands automatically display a list of possible completions whenever nothing can be completed because the next character is not uniquely determined.

18.5.4 High-Level Completion Functions

This section describes the higher-level convenient functions for reading certain sorts of names with completion.

In most cases, you should not call these functions in the middle of a Lisp function. When possible, do all minibuffer input as part of reading the arguments for a command, in the interactive spec. See Section 19.2 [Defining Commands], page 282.

Function read-buffer prompt &optional default existing This function reads the name of a buffer and returns it as a string. The argument default is the default name to use, the value to return if the user exits with an empty minibuffer. If non-nil, it should be a string or a buffer. It is mentioned in the prompt, but is not inserted in the minibuffer as initial input. If existing is non-nil, then the name specified must be that of an existing buffer. The usual commands to exit the minibuffer do not exit if the text is not valid, and ^{⟨}\underline{\overline{RET}}^{⟩} does completion to attempt to find a valid name. (However, default is not checked for validity; it is returned, whatever it is, if the user exits with the minibuffer empty.) In the following example, the user enters ‘minibuffer.t’, and then types ^{⟨}\underline{\overline{RET}}^{⟩}. The argument existing is t, and the only buffer name starting with the given input is ‘minibuffer.texi’, so that name is the value.

     (read-buffer "Buffer name? " "foo" t)

;; After evaluation of the preceding expression, ;; the following prompt appears, ;; with an empty minibuffer:

     ---------- Buffer: Minibuffer ----------
     Buffer name? (default foo) ⋆
     ---------- Buffer: Minibuffer ----------

;; The user types minibuffer.t ^{⟨}\underline{\overline{RET}}^{⟩}.

     ⇒ "minibuffer.texi"

Function read-command prompt

This function reads the name of a command and returns it as a Lisp symbol. The argument prompt is used as in read-from-minibuffer. Recall that a command is anything for which commandp returns t, and a command name is a symbol for which commandp returns t. See Section 19.3 [Interactive Call], page 286.

(read-command "Command name? ")

;; After evaluation of the preceding expression, ;; the following prompt appears with an empty minibuffer:

---------- Buffer: Minibuffer ----------
Command name?
---------- Buffer: Minibuffer ----------

If the user types forward-c ^{⟨}\underline{\overline{RET}}^{⟩}, then this function returns forward-char. The read-command function is a simplified interface to the function completingread. It uses the variable obarray so as to complete in the set of extant Lisp symbols, and it uses the commandp predicate so as to accept only command names:

(read-command prompt)

≡

(intern (completing-read prompt obarray
                         'commandp t nil))

Function read-variable prompt

This function reads the name of a user variable and returns it as a symbol.

(read-variable "Variable name? ")

;; After evaluation of the preceding expression, ;; the following prompt appears, ;; with an empty minibuffer:

---------- Buffer: Minibuffer ----------
Variable name? ⋆
---------- Buffer: Minibuffer ----------

If the user then types fill-p ^{⟨}\underline{\overline{RET}}^{⟩}, read-variable returns fill-prefix. This function is similar to read-command, but uses the predicate user-

variable-p instead of commandp:
(read-variable prompt)

≡

(intern
 (completing-read prompt obarray
                  'user-variable-p t nil))

18.5.5 Reading File Names

Here is another high-level completion function, designed for reading a file name. It provides special features including automatic insertion of the default directory.

Function read-file-name prompt &optional directory default existing initial This function reads a file name in the minibuffer, prompting with prompt and providing completion. If default is non-nil, then the function returns default if the user just types ^{⟨}\underline{\overline{RET}}^{⟩}. default is not checked for validity; it is returned, whatever it is, if the user exits with the minibuffer empty. If existing is non-nil, then the user must specify the name of an existing file; ⟨\underline{\overline{RET}}^{⟩} performs completion to make the name valid if possible, and then refuses to exit if it is not valid. If the value of existing is neither nil nor t, then ^{⟨}\underline{\overline{RET}}^{⟩} also requires confirmation after completion. If existing is nil, then the name of a nonexistent file is acceptable. The argument directory specifies the directory to use for completion of relative file names. If insert-default-directory is non-nil, directory is also inserted in the minibuffer as initial input. It defaults to the current buffer’s value of

default-directory.

If you specify initial, that is an initial file name to insert in the buffer (after with directory, if that is inserted). In this case, point goes at the beginning of initial. The default for initial is nil—don’t insert any file name. To see what initial does, try the command C-x C-v. Here is an example:

     (read-file-name "The file is ")

;; After evaluation of the preceding expression, ;; the following appears in the minibuffer:

     ---------- Buffer: Minibuffer ----------
     The file is /gp/gnu/elisp/⋆
     ---------- Buffer: Minibuffer ----------

Typing manual ^{⟨}\underline{\overline{TAB}}^{⟩} results in the following:

     ---------- Buffer: Minibuffer ----------
     The file is /gp/gnu/elisp/manual.texi⋆
     ---------- Buffer: Minibuffer ----------

If the user types ^{⟨}\underline{\overline{RET}}^{⟩}, read-file-name returns the file name as the string

"/gp/gnu/elisp/manual.texi".

User Option insert-default-directory

This variable is used by read-file-name. Its value controls whether readfile-name starts by placing the name of the default directory in the minibuffer, plus the initial file name if any. If the value of this variable is nil, then readfile-name does not place any initial input in the minibuffer (unless you specify initial input with the initial argument). In that case, the default directory is still used for completion of relative file names, but is not displayed. For example:

;; Here the minibuffer starts out with the default directory.

          (let ((insert-default-directory t))
            (read-file-name "The file is "))
          ---------- Buffer: Minibuffer ----------
          The file is ~lewis/manual/⋆
          ---------- Buffer: Minibuffer ----------

;; Here the minibuffer is empty and only the prompt ;; appears on its line.

          (let ((insert-default-directory nil))
            (read-file-name "The file is "))
          ---------- Buffer: Minibuffer ----------
          The file is ⋆
          ---------- Buffer: Minibuffer ----------

18.5.6 Programmed Completion

Sometimes it is not possible to create an alist or an obarray containing all the intended possible completions. In such a case, you can supply your own function to compute the completion of a given string. This is called programmed completion.

To use this feature, pass a symbol with a function definition as the collection argument to completing-read. The function completing-read arranges to pass your completion function along to try-completion and all-completions, which will then let your function do all the work.

The completion function should accept three arguments:

• The string to be completed.

• The predicate function to filter possible matches, or nil if none. Your function should call the predicate for each possible match, and ignore the possible match if the predicate returns nil.

• A flag specifying the type of operation.

There are three flag values for three operations:

• nil specifies try-completion. The completion function should return the completion of the specified string, or t if the string is an exact match already, or nil if the string matches no possibility.

• t specifies all-completions. The completion function should return a list of all possible completions of the specified string.

• lambda specifies a test for an exact match. The completion function should return t if the specified string is an exact match for some possibility; nil otherwise.

It would be consistent and clean for completion functions to allow lambda expressions (lists that are functions) as well as function symbols as collection, but this is impossible. Lists as completion tables are already assigned another meaning—as alists. It would be unreliable to fail to handle an alist normally because it is also a possible function. So you must arrange for any function you wish to use for completion to be encapsulated in a symbol.

Emacs uses programmed completion when completing file names. See Section 28.8.6 [File Name Completion], page 411.

18.6 Yes-or-No Queries

This section describes functions used to ask the user a yes-or-no question. The function y-or-n-p can be answered with a single character; it is useful for questions where an inadvertent wrong answer will not have serious consequences. yes-or-no-p is suitable for more momentous questions, since it requires three or four characters to answer. Variations of these functions can be used to ask a yes-or-no question using a dialog box, or optionally using one.

If either of these functions is called in a command that was invoked using the mouse, then it uses a dialog box or pop-up menu to ask the question. Otherwise, it uses keyboard input.

Strictly speaking, yes-or-no-p uses the minibuffer and y-or-n-p does not; but it seems best to describe them together.

Function y-or-n-p prompt

This function asks the user a question, expecting input in the echo area. It returns t if the user types y, nil if the user types n. This function also accepts ⟨\underline{\overline{SPC}}^{⟩} to mean yes and ^{⟨}\underline{\overline{DEL}}^{⟩} to mean no. It accepts C-] to mean “quit”, like C-g, because the question might look like a minibuffer and for that reason the user might try to use C-] to get out. The answer is a single character, with no ⟨\underline{\overline{RET}}^{⟩} needed to terminate it. Upper and lower case are equivalent. “Asking the question” means printing prompt in the echo area, followed by the string ‘(y or n) ’. If the input is not one of the expected answers (y, n, ⟨\underline{\overline{SPC}}^{⟩}, ^{⟨}\underline{\overline{DEL}}^{⟩}, or something that quits), the function responds ‘Please answer y or n.’, and repeats the request. This function does not actually use the minibuffer, since it does not allow editing of the answer. It actually uses the echo area (see Section 45.3 [The Echo Area], page 648), which uses the same screen space as the minibuffer. The cursor moves to the echo area while the question is being asked. The answers and their meanings, even ‘y’ and ‘n’, are not hardwired. The keymap query-replace-map specifies them. See Section 37.5 [Search and Replace], page 560. In the following example, the user first types q, which is invalid. At the next prompt the user types y.

     (y-or-n-p "Do you need a lift? ")

;; After evaluation of the preceding expression, ;; the following prompt appears in the echo area:

     ---------- Echo area ----------
     Do you need a lift? (y or n)
     ---------- Echo area ----------

;; If the user then types q, the following appears:

---------- Echo area ----------
Please answer y or n.  Do you need a lift? (y or n)
---------- Echo area ----------

;; When the user types a valid answer, ;; it is displayed after the question:

---------- Echo area ----------
Do you need a lift? (y or n) y
---------- Echo area ----------

We show successive lines of echo area messages, but only one actually appears on the screen at a time.

Function yes-or-no-p prompt

This function asks the user a question, expecting input in the minibuffer. It returns t if the user enters ‘yes’, nil if the user types ‘no’. The user must type ⟨\underline{\overline{RET}}^{⟩} to finalize the response. Upper and lower case are equivalent. yes-or-no-p starts by displaying prompt in the echo area, followed by ‘(yes or no) ’. The user must type one of the expected responses; otherwise, the function responds ‘Please answer yes or no.’, waits about two seconds and repeats the request. yes-or-no-p requires more work from the user than y-or-n-p and is appropriate for more crucial decisions. Here is an example:

(yes-or-no-p "Do you really want to remove everything? ")

;; After evaluation of the preceding expression, ;; the following prompt appears, ;; with an empty minibuffer:

---------- Buffer: minibuffer ----------
Do you really want to remove everything? (yes or no)
---------- Buffer: minibuffer ----------

If the user first types y ^{⟨}\underline{\overline{RET}}^{⟩}, which is invalid because this function demands the entire word ‘yes’, it responds by displaying these prompts, with a brief pause between them:

---------- Buffer: minibuffer ----------
Please answer yes or no.
Do you really want to remove everything? (yes or no)
---------- Buffer: minibuffer ----------

Function yes-or-no-p-dialog-box prompt

This function asks the user a “y or n” question with a popup dialog box. It returns t if the answer is “yes”. prompt is the string to display to ask the question.

The following functions ask a question either in the minibuffer or a dialog box, depending on whether the last user event (which presumably invoked this command) was a keyboard or mouse event. When XEmacs is running on a window system, the functions y-or-n-p and yes-or-no-p are replaced with the following functions, so that menu items bring up dialog boxes instead of minibuffer questions.

Function y-or-n-p-maybe-dialog-box prompt

This function asks user a “y or n” question, using either a dialog box or the minibuffer, as appropriate.

Function yes-or-no-p-maybe-dialog-box prompt

This function asks user a “yes or no” question, using either a dialog box or the minibuffer, as appropriate.

18.7 Asking Multiple Y-or-N Questions

When you have a series of similar questions to ask, such as “Do you want to save this buffer” for each buffer in turn, you should use map-y-or-n-p to ask the collection of questions, rather than asking each question individually. This gives the user certain convenient facilities such as the ability to answer the whole series at once.

Function map-y-or-n-p prompter actor list &optional help action-alist This function, new in Emacs 19, asks the user a series of questions, reading a single-character answer in the echo area for each one. The value of list specifies the objects to ask questions about. It should be either a list of objects or a generator function. If it is a function, it should expect no arguments, and should return either the next object to ask about, or nil meaning stop asking questions. The argument prompter specifies how to ask each question. If prompter is a string, the question text is computed like this:

(format prompter object) where object is the next object to ask about (as obtained from list). If not a string, prompter should be a function of one argument (the next object to ask about) and should return the question text. If the value is a string, that is the question to ask the user. The function can also return t meaning do act on this object (and don’t ask the user), or nil meaning ignore this object (and don’t ask the user). The argument actor says how to act on the answers that the user gives. It should be a function of one argument, and it is called with each object that the user says yes for. Its argument is always an object obtained from list. If the argument help is given, it should be a list of this form:

(singular plural action) where singular is a string containing a singular noun that describes the objects conceptually being acted on, plural is the corresponding plural noun, and action is a transitive verb describing what actor does. If you don’t specify help, the default is ("object" "objects" "act on").

Each time a question is asked, the user may enter y, Y, or ^{⟨}\underline{\overline{SPC}}^{⟩} to act on that object; n, N, or^{⟨}\underline{\overline{DEL}}^{⟩} to skip that object; ! to act on all following objects; ^{⟨}\underline{\overline{ESC}}^{⟩} or q to exit (skip all following objects); . (period) to act on the current object and then exit; or C-h to get help. These are the same answers that query-replace accepts. The keymap query-replace-map defines their meaning for map-y-or-n-p as well as for query-replace; see Section 37.5 [Search and Replace], page 560. You can use action-alist to specify additional possible answers and what they mean. It is an alist of elements of the form (char function help), each of which defines one additional answer. In this element, char is a character (the answer); function is a function of one argument (an object from list); help is a string. When the user responds with char, map-y-or-n-p calls function. If it returns non-nil, the object is considered “acted upon”, and map-y-or-n-p advances to the next object in list. If it returns nil, the prompt is repeated for the same object. If map-y-or-n-p is called in a command that was invoked using the mouse— more precisely, if last-nonmenu-event (see Section 19.4 [Command Loop Info], page 288) is either nil or a list—then it uses a dialog box or pop-up menu to ask the question. In this case, it does not use keyboard input or the echo area. You can force use of the mouse or use of keyboard input by binding last-nonmenu-event to a suitable value around the call. The return value of map-y-or-n-p is the number of objects acted on.

18.8 Minibuffer Miscellany

This section describes some basic functions and variables related to minibuffers.

Command exit-minibuffer

This command exits the active minibuffer. It is normally bound to keys in minibuffer local keymaps.

Command self-insert-and-exit

This command exits the active minibuffer after inserting the last character typed on the keyboard (found in last-command-char; see Section 19.4 [Command Loop Info], page 288).

Command previous-history-element n

This command replaces the minibuffer contents with the value of the nth previous (older) history element.

Command next-history-element n

This command replaces the minibuffer contents with the value of the nth more recent history element.

Command previous-matching-history-element pattern

This command replaces the minibuffer contents with the value of the previous (older) history element that matches pattern (a regular expression).

Command next-matching-history-element pattern

This command replaces the minibuffer contents with the value of the next (newer) history element that matches pattern (a regular expression).

Function minibuffer-prompt

This function returns the prompt string of the currently active minibuffer. If no minibuffer is active, it returns nil.

Function minibuffer-prompt-width

This function returns the display width of the prompt string of the currently active minibuffer. If no minibuffer is active, it returns 0.

Variable minibuffer-setup-hook

This is a normal hook that is run whenever the minibuffer is entered. See Section 26.4 [Hooks], page 378.

Variable minibuffer-exit-hook

This is a normal hook that is run whenever the minibuffer is exited. See Section 26.4 [Hooks], page 378.

Variable minibuffer-help-form

The current value of this variable is used to rebind help-form locally inside the minibuffer (see Section 27.5 [Help Functions], page 387).

Function active-minibuffer-window

This function returns the currently active minibuffer window, or nil if none is currently active.

Function minibuffer-window &optional frame

This function returns the minibuffer window used for frame frame. If frame is nil, that stands for the current frame. Note that the minibuffer window used by a frame need not be part of that frame—a frame that has no minibuffer of its own necessarily uses some other frame’s minibuffer window.

Function window-minibuffer-p window

This function returns non-nil if window is a minibuffer window.

It is not correct to determine whether a given window is a minibuffer by comparing it with the result of (minibuffer-window), because there can be more than one minibuffer window if there is more than one frame.

Function minibuffer-window-active-p window

This function returns non-nil if window, assumed to be a minibuffer window, is currently active.

Variable minibuffer-scroll-window

If the value of this variable is non-nil, it should be a window object. When the function scroll-other-window is called in the minibuffer, it scrolls this window.

Finally, some functions and variables deal with recursive minibuffers (see Section 19.10 [Recursive Editing], page 310):

Function minibuffer-depth

This function returns the current depth of activations of the minibuffer, a nonnegative integer. If no minibuffers are active, it returns zero.

User Option enable-recursive-minibuffers

If this variable is non-nil, you can invoke commands (such as find-file) that use minibuffers even while in the minibuffer window. Such invocation produces a recursive editing level for a new minibuffer. The outer-level minibuffer is invisible while you are editing the inner one. This variable only affects invoking the minibuffer while the minibuffer window is selected. If you switch windows while in the minibuffer, you can always invoke minibuffer commands while some other window is selected.

In FSF Emacs 19, if a command name has a property enable-recursive-minibuffers that is non-nil, then the command can use the minibuffer to read arguments even if it is invoked from the minibuffer. The minibuffer command next-matching-history-element (normally M-s in the minibuffer) uses this feature.

This is not implemented in XEmacs because it is a kludge. If you want to explicitly set the value of enable-recursive-minibuffers in this fashion, just use an evaluated interactive spec and bind enable-recursive-minibuffers while reading from the minibuffer.

See the definition of next-matching-history-element in ‘lisp/prim/minibuf.el’.

19 Command Loop

When you run XEmacs, it enters the editor command loop almost immediately. This loop reads events, executes their definitions, and displays the results. In this chapter, we describe how these things are done, and the subroutines that allow Lisp programs to do them.

19.1 Command Loop Overview

The command loop in XEmacs is a standard event loop, reading events one at a time with next-event and handling them with dispatch-event. An event is typically a single user action, such as a keypress, mouse movement, or menu selection; but they can also be notifications from the window system, informing XEmacs that (for example) part of its window was just uncovered and needs to be redrawn. See Section 19.5 [Events], page 290. Pending events are held in a first-in, first-out list called the event queue: events are read from the head of the list, and newly arriving events are added to the tail. In this way, events are always processed in the order in which they arrive.

dispatch-event does most of the work of handling user actions. The first thing it must do is put the events together into a key sequence, which is a sequence of events that translates into a command. It does this by consulting the active keymaps, which specify what the valid key sequences are and how to translate them into commands. See Section 20.8 [Key Lookup], page 324, for information on how this is done. The result of the translation should be a keyboard macro or an interactively callable function. If the key is M-x, then it reads the name of another command, which it then calls. This is done by the command execute-extended-command (see Section 19.3 [Interactive Call], page 286).

To execute a command requires first reading the arguments for it. This is done by calling command-execute (see Section 19.3 [Interactive Call], page 286). For commands written in Lisp, the interactive specification says how to read the arguments. This may use the prefix argument (see Section 19.9 [Prefix Command Arguments], page 308) or may read with prompting in the minibuffer (see Chapter 18 [Minibuffers], page 261). For example, the command find-file has an interactive specification which says to read a file name using the minibuffer. The command’s function body does not use the minibuffer; if you call this command from Lisp code as a function, you must supply the file name string as an ordinary Lisp function argument.

If the command is a string or vector (i.e., a keyboard macro) then execute-kbd-macro is used to execute it. You can call this function yourself (see Section 19.13 [Keyboard Macros], page 313).

To terminate the execution of a running command, type C-g. This character causes quitting (see Section 19.8 [Quitting], page 306).

Variable pre-command-hook

The editor command loop runs this normal hook before each command. At that time, this-command contains the command that is about to run, and lastcommand describes the previous command. See Section 26.4 [Hooks], page 378.

Variable post-command-hook

The editor command loop runs this normal hook after each command. (In FSF Emacs, it is also run when the command loop is entered, or reentered after an error or quit.) At that time, this-command describes the command that just ran, and last-command describes the command before that. See Section 26.4 [Hooks], page 378.

Quitting is suppressed while running pre-command-hook and post-command-hook. If an error happens while executing one of these hooks, it terminates execution of the hook, but that is all it does.

19.2 Defining Commands

A Lisp function becomes a command when its body contains, at top level, a form that calls the special form interactive. This form does nothing when actually executed, but its presence serves as a flag to indicate that interactive calling is permitted. Its argument controls the reading of arguments for an interactive call.

19.2.1 Using interactive

This section describes how to write the interactive form that makes a Lisp function an interactively-callable command.

Special Form interactive arg-descriptor This special form declares that the function in which it appears is a command, and that it may therefore be called interactively (via M-x or by entering a key sequence bound to it). The argument arg-descriptor declares how to compute the arguments to the command when the command is called interactively. A command may be called from Lisp programs like any other function, but then the caller supplies the arguments and arg-descriptor has no effect. The interactive form has its effect because the command loop (actually, its subroutine call-interactively) scans through the function definition looking for it, before calling the function. Once the function is called, all its body forms including the interactive form are executed, but at this time interactive simply returns nil without even evaluating its argument.

There are three possibilities for the argument arg-descriptor:

• It may be omitted or nil; then the command is called with no arguments. This leads quickly to an error if the command requires one or more arguments.

• It may be a Lisp expression that is not a string; then it should be a form that is evaluated to get a list of arguments to pass to the command. If this expression reads keyboard input (this includes using the minibuffer), keep in mind that the integer value of point or the mark before reading input may be incorrect after reading input. This is because the current buffer may be receiving subprocess output; if subprocess output arrives while the command is waiting for input, it could relocate point and the mark. Here’s an example of what not to do:

     (interactive
      (list (region-beginning) (region-end)
            (read-string "Foo: " nil 'my-history)))

Here’s how to avoid the problem, by examining point and the mark only after reading the keyboard input:

     (interactive
      (let ((string (read-string "Foo: " nil 'my-history)))
        (list (region-beginning) (region-end) string)))

• It may be a string; then its contents should consist of a code character followed by a prompt (which some code characters use and some ignore). The prompt ends either with the end of the string or with a newline. Here is a simple example:

     (interactive "bFrobnicate buffer: ")

The code letter ‘b’ says to read the name of an existing buffer, with completion. The buffer name is the sole argument passed to the command. The rest of the string is a prompt. If there is a newline character in the string, it terminates the prompt. If the string does not end there, then the rest of the string should contain another code character and prompt, specifying another argument. You can specify any number of arguments in this way. The prompt string can use ‘%’ to include previous argument values (starting with the first argument) in the prompt. This is done using format (see Section 4.10 [Formatting Strings], page 67). For example, here is how you could read the name of an existing buffer followed by a new name to give to that buffer:

     (interactive "bBuffer to rename: \nsRename buffer %s to: ")

If the first character in the string is ‘*’, then an error is signaled if the buffer is read-only. If the first character in the string is ‘@’, and if the key sequence used to invoke the command includes any mouse events, then the window associated with the first of those events is selected before the command is run. If the first character in the string is ‘_’, then this command will not cause the region to be deactivated when it completes; that is, zmacs-region-stays will be set to t when the command exits successfully. You can use ‘*’, ‘@’, and ‘_’ together; the order does not matter. Actual reading of arguments is controlled by the rest of the prompt string (starting with the first character that is not ‘*’, ‘@’, or ‘_’).

Function function-interactive function

This function retrieves the interactive specification of function, which may be any funcallable object. The specification will be returned as the list of the symbol interactive and the specs. If function is not interactive, nil will be returned.

19.2.2 Code Characters for interactive

The code character descriptions below contain a number of key words, defined here as follows:

Completion Provide completion. ^{⟨}\underline{\overline{TAB}}^{⟩}, ^{⟨}\underline{\overline{SPC}}^{⟩}, and ^{⟨}\underline{\overline{RET}}^{⟩} perform name completion because the argument is read using completing-read (see Section 18.5 [Completion], page 266). ? displays a list of possible completions.

Existing Require the name of an existing object. An invalid name is not accepted; the commands to exit the minibuffer do not exit if the current input is not valid.

Default A default value of some sort is used if the user enters no text in the minibuffer. The default depends on the code character.

No I/O This code letter computes an argument without reading any input. Therefore, it does not use a prompt string, and any prompt string you supply is ignored. Even though the code letter doesn’t use a prompt string, you must follow it with a newline if it is not the last code character in the string.

Prompt A prompt immediately follows the code character. The prompt ends either with the end of the string or with a newline.

Special This code character is meaningful only at the beginning of the interactive string, and it does not look for a prompt or a newline. It is a single, isolated character.

Here are the code character descriptions for use with interactive:

‘*’ Signal an error if the current buffer is read-only. Special.

‘@’ Select the window mentioned in the first mouse event in the key sequence that invoked this command. Special.

‘_’ Do not cause the region to be deactivated when this command completes. Special.

‘a’ A function name (i.e., a symbol satisfying fboundp). Existing, Completion, Prompt.

‘b’ The name of an existing buffer. By default, uses the name of the current buffer (see Chapter 30 [Buffers], page 431). Existing, Completion, Default, Prompt.

‘B’ A buffer name. The buffer need not exist. By default, uses the name of a recently used buffer other than the current buffer. Completion, Default, Prompt.

‘c’ A character. The cursor does not move into the echo area. Prompt.

‘C’ A command name (i.e., a symbol satisfying commandp). Existing, Completion, Prompt.

‘d’ The position of point, as an integer (see Section 34.1 [Point], page 487). No I/O.

‘D’ A directory name. The default is the current default directory of the current buffer, default-directory (see Section 50.3 [System Environment], page 697). Existing, Completion, Default, Prompt.

‘e’ The last mouse-button or misc-user event in the key sequence that invoked the command. No I/O.

You can use ‘e’ more than once in a single command’s interactive specification. If the key sequence that invoked the command has n mouse-button or misc-user events, the nth ‘e’ provides the nth such event.

‘f’ A file name of an existing file (see Section 28.8 [File Names], page 406). The default directory is default-directory. Existing, Completion, Default, Prompt.

‘F’ A file name. The file need not exist. Completion, Default, Prompt.

‘k’ A key sequence (see Section 20.1 [Keymap Terminology], page 315). This keeps reading events until a command (or undefined command) is found in the current key maps. The key sequence argument is represented as a vector of events. The cursor does not move into the echo area. Prompt. This kind of input is used by commands such as describe-key and global-

set-key.

‘K’ A key sequence, whose definition you intend to change. This works like ‘k’, except that it suppresses, for the last input event in the key sequence, the conversions that are normally used (when necessary) to convert an undefined key into a defined one.

‘m’ The position of the mark, as an integer. No I/O.

‘n’ A number read with the minibuffer. If the input is not a number, the user is asked to try again. The prefix argument, if any, is not used. Prompt.

‘N’ The raw prefix argument. If the prefix argument is nil, then read a number as with n. Requires a number. See Section 19.9 [Prefix Command Arguments], page 308. Prompt.

‘p’ The numeric prefix argument. (Note that this ‘p’ is lower case.) No I/O.

‘P’ The raw prefix argument. (Note that this ‘P’ is upper case.) No I/O.

‘r’ Point and the mark, as two numeric arguments, smallest first. This is the only code letter that specifies two successive arguments rather than one. No I/O.

‘s’ Arbitrary text, read in the minibuffer and returned as a string (see Section 18.2 [Text from Minibuffer], page 261). Terminate the input with either ^{⟨}\underline{\overline{LFD}}^{⟩} or ⟨\underline{\overline{RET}}^{⟩}. (C-q may be used to include either of these characters in the input.) Prompt.

‘S’ An interned symbol whose name is read in the minibuffer. Any whitespace character terminates the input. (Use C-q to include whitespace in the string.) Other characters that normally terminate a symbol (e.g., parentheses and brackets) do not do so here. Prompt.

‘v’ A variable declared to be a user option (i.e., satisfying the predicate uservariable-p). See Section 18.5.4 [High-Level Completion], page 271. Existing, Completion, Prompt.

‘x’ A Lisp object, specified with its read syntax, terminated with a ^{⟨}\underline{\overline{LFD}}^{⟩} or ⟨\underline{\overline{RET}}^{⟩}. The object is not evaluated. See Section 18.3 [Object from Minibuffer], page 263. Prompt.

‘X’ A Lisp form is read as with x, but then evaluated so that its value becomes the argument for the command. Prompt.

19.2.3 Examples of Using interactive

Here are some examples of interactive:

(defun foo1 ()              ; foo1 takes no arguments,

(interactive) ; just moves forward two words.

(forward-word 2))
 ⇒ foo1
(defun foo2 (n)             ; foo2 takes one argument,

(interactive "p") ; which is the numeric prefix.

(forward-word (* 2 n)))
 ⇒ foo2
(defun foo3 (n)             ; foo3 takes one argument,

(interactive "nCount:") ; which is read with the Minibuffer.

(forward-word (* 2 n)))
 ⇒ foo3
(defun three-b (b1 b2 b3)
"Select three existing buffers.
Put them into three windows, selecting the last one."
(interactive "bBuffer1:\nbBuffer2:\nbBuffer3:")
(delete-other-windows)
(split-window (selected-window) 8)
(switch-to-buffer b1)
(other-window 1)
(split-window (selected-window) 8)
(switch-to-buffer b2)
(other-window 1)
(switch-to-buffer b3))
 ⇒ three-b
(three-b "*scratch*" "declarations.texi" "*mail*")
 ⇒ nil

19.3 Interactive Call

After the command loop has translated a key sequence into a definition, it invokes that definition using the function command-execute. If the definition is a function that is a command, command-execute calls call-interactively, which reads the arguments and calls the command. You can also call these functions yourself.

Function commandp object

Returns t if object is suitable for calling interactively; that is, if object is a command. Otherwise, returns nil. The interactively callable objects include strings and vectors (treated as keyboard macros), lambda expressions that contain a top-level call to interactive, compiled-function objects made from such lambda expressions, autoload objects that are declared as interactive (non-nil fourth argument to autoload), and some of the primitive functions. A symbol is commandp if its function definition is commandp. Keys and keymaps are not commands. Rather, they are used to look up commands (see Chapter 20 [Keymaps], page 315). See documentation in Section 27.2 [Accessing Documentation], page 382, for a realistic example of using commandp.

Function call-interactively command &optional record-flag This function calls the interactively callable function command, reading arguments according to its interactive calling specifications. An error is signaled if command is not a function or if it cannot be called interactively (i.e., is not a command). Note that keyboard macros (strings and vectors) are not accepted, even though they are considered commands, because they are not functions. If record-flag is the symbol lambda, the interactive calling arguments for command are read and returned as a list, but the function is not called on them. If record-flag is t, then this command and its arguments are unconditionally added to the list command-history. Otherwise, the command is added only if it uses the minibuffer to read an argument. See Section 19.12 [Command History], page 312.

Function command-execute command &optional record-flag This function executes command as an editing command. The argument command must satisfy the commandp predicate; i.e., it must be an interactively callable function or a keyboard macro. A string or vector as command is executed with execute-kbd-macro. A function is passed to call-interactively, along with the optional record-flag. A symbol is handled by using its function definition in its place. A symbol with an autoload definition counts as a command if it was declared to stand for an interactively callable function. Such a definition is handled by loading the specified library and then rechecking the definition of the symbol.

Command execute-extended-command prefix-argument

This function reads a command name from the minibuffer using completingread (see Section 18.5 [Completion], page 266). Then it uses command-execute to call the specified command. Whatever that command returns becomes the

value of execute-extended-command.

If the command asks for a prefix argument, it receives the value prefix-argument. If execute-extended-command is called interactively, the current raw prefix argument is used for prefix-argument, and thus passed on to whatever command is run. execute-extended-command is the normal definition of M-x, so it uses the string ‘M-x ’ as a prompt. (It would be better to take the prompt from the events used to invoke execute-extended-command, but that is painful to implement.) A description of the value of the prefix argument, if any, also becomes part of the prompt.

     (execute-extended-command 1)
     ---------- Buffer: Minibuffer ----------
     1 M-x forward-word RET
     ---------- Buffer: Minibuffer ----------
          ⇒ t

Function interactive-p

This function returns t if the containing function (the one that called interactivep) was called interactively, with the function call-interactively. (It makes no difference whether call-interactively was called from Lisp or directly from the editor command loop.) If the containing function was called by Lisp evaluation (or with apply or funcall), then it was not called interactively. The most common use of interactive-p is for deciding whether to print an informative message. As a special exception, interactive-p returns nil whenever a keyboard macro is being run. This is to suppress the informative messages and speed execution of the macro. For example:

     (defun foo ()
       (interactive)
       (and (interactive-p)
            (message "foo")))
          ⇒ foo
     (defun bar ()
       (interactive)
       (setq foobar (list (foo) (interactive-p))))
          ⇒ bar
     ;; Type M-x foo.
          ⊣ foo
     ;; Type M-x bar.

;; This does not print anything.

     foobar ⇒ (nil t)

19.4 Information from the Command Loop

The editor command loop sets several Lisp variables to keep status records for itself and for commands that are run.

Variable last-command

This variable records the name of the previous command executed by the command loop (the one before the current command). Normally the value is a symbol with a function definition, but this is not guaranteed. The value is copied from this-command when a command returns to the command loop, except when the command specifies a prefix argument for the following command.

Variable this-command

This variable records the name of the command now being executed by the editor command loop. Like last-command, it is normally a symbol with a function definition.

The command loop sets this variable just before running a command, and copies its value into last-command when the command finishes (unless the command specifies a prefix argument for the following command).

Some commands set this variable during their execution, as a flag for whatever command runs next. In particular, the functions for killing text set thiscommand to kill-region so that any kill commands immediately following will know to append the killed text to the previous kill.

If you do not want a particular command to be recognized as the previous command in the case where it got an error, you must code that command to prevent this. One way is to set this-command to t at the beginning of the command, and set this-command back to its proper value at the end, like this:

(defun foo (args...)
  (interactive ...)
  (let ((old-this-command this-command))
    (setq this-command t)

…do the work…

    (setq this-command old-this-command)))

Function this-command-keys

This function returns a vector containing the key and mouse events that invoked the present command, plus any previous commands that generated the prefix argument for this command. (Note: this is not the same as in FSF Emacs, which can return a string.) See Section 19.5 [Events], page 290.

This function copies the vector and the events; it is safe to keep and modify them.

     (this-command-keys)

;; Now use C-u C-x C-e to evaluate that.

          ⇒ [#<keypress-event control-U> #<keypress-event control-X> #<keypress-event control-E>]

Variable last-command-event

This variable is set to the last input event that was read by the command loop as part of a command. The principal use of this variable is in self-insert-command, which uses it to decide which character to insert.

This variable is off limits: you may not set its value or modify the event that is its value, as it is destructively modified by read-key-sequence. If you want to keep a pointer to this value, you must use copy-event.

Note that this variable is an alias for last-command-char in FSF Emacs.

     last-command-event
     ;; Now type C-u C-x C-e.
          ⇒ #<keypress-event control-E>

Variable last-command-char

If the value of last-command-event is a keyboard event, then this is the nearest character equivalent to it (or nil if there is no character equivalent). lastcommand-char is the character that self-insert-command will insert in the buffer. Remember that there is not a one-to-one mapping between keyboard events and XEmacs characters: many keyboard events have no corresponding character, and when the Mule feature is available, most characters can not be input on standard keyboards, except possibly with help from an input method. So writing code that examines this variable to determine what key has been typed is bad practice, unless you are certain that it will be one of a small set of characters.

This variable exists for compatibility with Emacs version 18.

     last-command-char

;; Now use C-u C-x C-e to evaluate that.

          ⇒ ?\^E

Variable current-mouse-event

This variable holds the mouse-button event which invoked this command, or

nil. This is what (interactive "e") returns.

Variable echo-keystrokes

This variable determines how much time should elapse before command characters echo. Its value must be an integer, which specifies the number of seconds to wait before echoing. If the user types a prefix key (say C-x) and then delays this many seconds before continuing, the key C-x is echoed in the echo area. Any subsequent characters in the same command will be echoed as well.

If the value is zero, then command input is not echoed.

19.5 Events

The XEmacs command loop reads a sequence of events that represent keyboard or mouse activity. Unlike in Emacs 18 and in FSF Emacs, events are a primitive Lisp type that must be manipulated using their own accessor and settor primitives. This section describes the representation and meaning of input events in detail.

A key sequence that starts with a mouse event is read using the keymaps of the buffer in the window that the mouse was in, not the current buffer. This does not imply that clicking in a window selects that window or its buffer—that is entirely under the control of the command binding of the key sequence.

For information about how exactly the XEmacs command loop works, See Section 19.6 [Reading Input], page 301.

Function eventp object

This function returns non-nil if event is an input event.

19.5.1 Event Types

Events represent keyboard or mouse activity or status changes of various sorts, such as process input being available or a timeout being triggered. The different event types are as follows:

key-press event A key was pressed. Note that modifier keys such as “control”, “shift”, and “alt” do not generate events; instead, they are tracked internally by XEmacs, and non-modifier key presses generate events that specify both the key pressed and the modifiers that were held down at the time.

button-press event button-release event

A button was pressed or released. Along with the button that was pressed or released, button events specify the modifier keys that were held down at the time and the position of the pointer at the time.

motion event The pointer was moved. Along with the position of the pointer, these events also specify the modifier keys that were held down at the time.

misc-user event A menu item was selected, the scrollbar was used, or a drag or a drop occurred.

process event Input is available on a process.

timeout event A timeout has triggered.

magic event Some window-system-specific action (such as a frame being resized or a portion of a frame needing to be redrawn) has occurred. The contents of this event are not accessible at the E-Lisp level, but dispatch-event knows what to do with an event of this type.

eval event This is a special kind of event specifying that a particular function needs to be called when this event is dispatched. An event of this type is sometimes placed in the event queue when a magic event is processed. This kind of event should generally just be passed off to dispatch-event. See Section 19.6.3 [Dispatching an Event], page 303.

19.5.2 Contents of the Different Types of Events

Every event, no matter what type it is, contains a timestamp (which is typically an offset in milliseconds from when the X server was started) indicating when the event occurred. In addition, many events contain a channel, which specifies which frame the event occurred on, and/or a value indicating which modifier keys (shift, control, etc.) were held down at the time of the event.

The contents of each event are as follows: key-press event

channel

timestamp

key Which key was pressed. This is an integer (in the printing ASCII range: >32 and <127) or a symbol such as left or right. Note that many physical keys are actually treated as two separate keys, depending on whether the shift key is pressed; for example, the “a” key is treated as either “a” or “A” depending on the state of the shift key, and the “1” key is similarly treated as either “1” or “!” on most keyboards. In such cases, the shift key does not show up in the modifier list. For other keys, such as backspace, the shift key shows up as a regular modifier.

modifiers Which modifier keys were pressed. As mentioned above, the shift key is not treated as a modifier for many keys and will not show up in this list in such cases.

button-press event button-release event

channel

timestamp

button What button went down or up. Buttons are numbered starting at 1.

modifiers Which modifier keys were pressed. The special business mentioned above for the shift key does not apply to mouse events.

x y The position of the pointer (in pixels) at the time of the event.

pointer-motion event channel

timestamp

x y The position of the pointer (in pixels) after it moved.

modifiers Which modifier keys were pressed. The special business mentioned above for the shift key does not apply to mouse events.

misc-user event timestamp

function The E-Lisp function to call for this event. This is normally either

          eval or call-interactively.

object The object to pass to the function. This is normally the callback that was specified in the menu description. button What button went down or up. Buttons are numbered starting at

1.

modifiers Which modifier keys were pressed. The special business mentioned

above for the shift key does not apply to mouse events.

x y The position of the pointer (in pixels) at the time of the event.

process event timestamp process The Emacs “process” object in question.

timeout event timestamp function The E-Lisp function to call for this timeout. It is called with one

argument, the event.

object Some Lisp object associated with this timeout, to make it easier to tell them apart. The function and object for this event were specified when the timeout was set.

magic event timestamp (The rest of the information in this event is not user-accessible.)

eval event timestamp function An E-Lisp function to call when this event is dispatched. object The object to pass to the function. The function and object are set

when the event is created.

Function event-type event

Return the type of event. This will be a symbol; one of

key-press

A key was pressed.

button-press

A mouse button was pressed.

button-release

A mouse button was released.

motion The mouse moved.

misc-user

Some other user action happened; typically, this is a menu selection, scrollbar action, or drag and drop action. process Input is available from a subprocess.

timeout A timeout has expired.

eval This causes a specified action to occur when dispatched.

magic Some window-system-specific event has occurred.

19.5.3 Event Predicates

The following predicates return whether an object is an event of a particular type.

Function key-press-event-p object

This is true if object is a key-press event.

Function button-event-p object object

This is true if object is a mouse button-press or button-release event.

Function button-press-event-p object

This is true if object is a mouse button-press event.

Function button-release-event-p object

This is true if object is a mouse button-release event.

Function motion-event-p object

This is true if object is a mouse motion event.

Function mouse-event-p object

This is true if object is a mouse button-press, button-release or motion event.

Function eval-event-p object

This is true if object is an eval event.

Function misc-user-event-p object

This is true if object is a misc-user event.

Function process-event-p object

This is true if object is a process event.

Function timeout-event-p object

This is true if object is a timeout event.

Function event-live-p object

This is true if object is any event that has not been deallocated.

19.5.4 Accessing the Position of a Mouse Event

Unlike other events, mouse events (i.e. motion, button-press, button-release, and drag or drop type misc-user events) occur in a particular location on the screen. Many primitives are provided for determining exactly where the event occurred and what is under that location.

19.5.4.1 Frame-Level Event Position Info

The following functions return frame-level information about where a mouse event occurred.

Function event-frame event

This function returns the “channel” or frame that the given mouse motion, button press, button release, or misc-user event occurred in. This will be nil for non-mouse events.

Function event-x-pixel event

This function returns the X position in pixels of the given mouse event. The value returned is relative to the frame the event occurred in. This will signal an error if the event is not a mouse event.

Function event-y-pixel event

This function returns the Y position in pixels of the given mouse event. The value returned is relative to the frame the event occurred in. This will signal an error if the event is not a mouse event.

19.5.4.2 Window-Level Event Position Info

The following functions return window-level information about where a mouse event occurred.

Function event-window event

Given a mouse motion, button press, button release, or misc-user event, compute and return the window on which that event occurred. This may be nil if the event occurred in the border or over a toolbar. The modeline is considered to be within the window it describes.

Function event-buffer event

Given a mouse motion, button press, button release, or misc-user event, compute and return the buffer of the window on which that event occurred. This may be nil if the event occurred in the border or over a toolbar. The modeline is considered to be within the window it describes. This is equivalent to calling event-window and then calling window-buffer on the result if it is a window.

Function event-window-x-pixel event

This function returns the X position in pixels of the given mouse event. The value returned is relative to the window the event occurred in. This will signal an error if the event is not a mouse-motion, button-press, button-release, or misc-user event.

Function event-window-y-pixel event

This function returns the Y position in pixels of the given mouse event. The value returned is relative to the window the event occurred in. This will signal an error if the event is not a mouse-motion, button-press, button-release, or misc-user event.

19.5.4.3 Event Text Position Info

The following functions return information about the text (including the modeline) that a mouse event occurred over or near.

Function event-over-text-area-p event

Given a mouse-motion, button-press, button-release, or misc-user event, this function returns t if the event is over the text area of a window. Otherwise, nil is returned. The modeline is not considered to be part of the text area.

Function event-over-modeline-p event

Given a mouse-motion, button-press, button-release, or misc-user event, this function returns t if the event is over the modeline of a window. Otherwise, nil is returned.

Function event-x event

This function returns the X position of the given mouse-motion, button-press, button-release, or misc-user event in characters. This is relative to the window the event occurred over.

Function event-y event

This function returns the Y position of the given mouse-motion, button-press, button-release, or misc-user event in characters. This is relative to the window the event occurred over.

Function event-point event

This function returns the character position of the given mouse-motion, buttonpress, button-release, or misc-user event. If the event did not occur over a window, or did not occur over text, then this returns nil. Otherwise, it returns an index into the buffer visible in the event’s window.

Function event-closest-point event

This function returns the character position of the given mouse-motion, buttonpress, button-release, or misc-user event. If the event did not occur over a window or over text, it returns the closest point to the location of the event. If the Y pixel position overlaps a window and the X pixel position is to the left of that window, the closest point is the beginning of the line containing the Y position. If the Y pixel position overlaps a window and the X pixel position is to the right of that window, the closest point is the end of the line containing the Y position. If the Y pixel position is above a window, 0 is returned. If it is below a window, the value of (window-end) is returned.

19.5.4.4 Event Glyph Position Info

The following functions return information about the glyph (if any) that a mouse event occurred over.

Function event-over-glyph-p event

Given a mouse-motion, button-press, button-release, or misc-user event, this function returns t if the event is over a glyph. Otherwise, nil is returned.

Function event-glyph-extent event

If the given mouse-motion, button-press, button-release, or misc-user event happened on top of a glyph, this returns its extent; else nil is returned.

Function event-glyph-x-pixel event

Given a mouse-motion, button-press, button-release, or misc-user event over a glyph, this function returns the X position of the pointer relative to the upper left of the glyph. If the event is not over a glyph, it returns nil.

Function event-glyph-y-pixel event

Given a mouse-motion, button-press, button-release, or misc-user event over a glyph, this function returns the Y position of the pointer relative to the upper left of the glyph. If the event is not over a glyph, it returns nil.

19.5.4.5 Event Toolbar Position Info

Function event-over-toolbar-p event

Given a mouse-motion, button-press, button-release, or misc-user event, this function returns t if the event is over a toolbar. Otherwise, nil is returned.

Function event-toolbar-button event

If the given mouse-motion, button-press, button-release, or misc-user event happened on top of a toolbar button, this function returns the button. Otherwise, nil is returned.

19.5.4.6 Other Event Position Info

Function event-over-border-p event

Given a mouse-motion, button-press, button-release, or misc-user event, this function returns t if the event is over an internal toolbar. Otherwise, nil is returned.

19.5.5 Accessing the Other Contents of Events

The following functions allow access to the contents of events other than the position info described in the previous section.

Function event-timestamp event

This function returns the timestamp of the given event object.

Function event-device event

This function returns the device that the given event occurred on.

Function event-key event

This function returns the Keysym of the given key-press event. This will be the ASCII code of a printing character, or a symbol.

Function event-button event

This function returns the button-number of the given button-press or buttonrelease event.

Function event-modifiers event

This function returns a list of symbols, the names of the modifier keys which were down when the given mouse or keyboard event was produced.

Function event-modifier-bits event

This function returns a number representing the modifier keys which were down when the given mouse or keyboard event was produced.

Function event-function event

This function returns the callback function of the given timeout, misc-user, or eval event.

Function event-object event

This function returns the callback function argument of the given timeout, misc-user, or eval event.

Function event-process event

This function returns the process of the given process event.

19.5.6 Working With Events

XEmacs provides primitives for creating, copying, and destroying event objects. Many functions that return events take an event object as an argument and fill in the fields of this event; or they make accept either an event object or nil, creating the event object first in the latter case.

Function make-event &optional type plist This function creates a new event structure. If no arguments are specified, the created event will be empty. To specify the event type, use the type argument.

The allowed types are empty, key-press, button-press, button-release,
motion, or misc-user.

plist is a property list, the properties being compatible to those returned by event-properties. For events other than empty, it is mandatory to specify certain properties. For empty events, plist must be nil. The list is canonicalized, which means that if a property keyword is present more than once, only the first instance is taken into account. Specifying an unknown or illegal property signals an error. The following properties are allowed:

channel The event channel. This is a frame or a console. For mouse events

          (of type button-press, button-release and motion), this must

be a frame. For key-press events, it must be a console. If channel is unspecified by plist, it will be set to the selected frame or selected console, as appropriate. key The event key. This is either a symbol or a character. It is allowed

(and required) only for key-press events.

button The event button. This an integer, either 1, 2 or 3. It is allowed only for button-press and button-release events.

modifiers

The event modifiers. This is a list of modifier symbols. It is allowed for key-press, button-press, button-release and motion events.

x The event X coordinate. This is an integer. It is relative to the channel’s root window, and is allowed for button-press, buttonrelease and motion events.

y The event Y coordinate. This is an integer. It is relative to the channel’s root window, and is allowed for button-press, buttonrelease and motion events. This means that, for instance, to access the toolbar, the y property will have to be negative.

timestamp

The event timestamp, a non-negative integer. Allowed for all types of events.

WARNING: the event object returned by this function may be a reused one;

see the function deallocate-event.

The events created by make-event can be used as non-interactive arguments to the functions with an (interactive "e") specification.

Here are some basic examples of usage:

;; Create an empty event.

(make-event)
     ⇒ #<empty-event>

;; Try creating a key-press event.

(make-event 'key-press)
      error Undefined key for keypress event

;; Creating a key-press event, try 2

(make-event 'key-press '(key home))
     ⇒ #<keypress-event home>

;; Create a key-press event of dubious fame.

(make-event 'key-press '(key escape modifiers (meta alt control shift)))
     ⇒ #<keypress-event control-meta-alt-shift-escape>

;; Create a M-button1 event at coordinates defined by variables ;; x and y.

(make-event 'button-press `(button 1 modifiers (meta) x ,x y ,y))
     ⇒ #<buttondown-event meta-button1>

;; Create a similar button-release event.

(make-event 'button-release `(button 1 modifiers (meta) x ,x y ,x))
     ⇒ #<buttonup-event meta-button1up>

;; Create a mouse-motion event.

     (make-event 'motion '(x 20 y 30))
          ⇒ #<motion-event 20, 30>
     (event-properties (make-event 'motion '(x 20 y 30)))
          ⇒ (channel #<x-frame "emacs" 0x8e2> x 20 y 30
              modifiers nil timestamp 0)

In conjunction with event-properties, you can use make-event to create modified copies of existing events. For instance, the following code will return an equal copy of event:

     (make-event (event-type event)
                 (event-properties event))

Note, however, that you cannot use make-event as the generic replacement for copy-event, because it does not allow creating all of the event types. To create a modified copy of an event, you can use the canonicalization feature of plist. The following example creates a copy of event, but with modifiers reset to nil.

     (make-event (event-type event)
                 (append '(modifiers nil)
                         (event-properties event)))

Function copy-event event1 &optional event2 This function makes a copy of the given event object. If a second argument is given, the first event is copied into the second and the second is returned. If the second argument is not supplied (or is nil) then a new event will be made.

Function deallocate-event event

This function allows the given event structure to be reused. You MUST NOT use this event object after calling this function with it. You will lose. It is not necessary to call this function, as event objects are garbage-collected like all other objects; however, it may be more efficient to explicitly deallocate events when you are sure that that is safe.

19.5.7 Converting Events

XEmacs provides some auxiliary functions for converting between events and other ways of representing keys. These are useful when working with ASCII strings and with keymaps.

Function character-to-event ch &optional event device This function converts a numeric ASCII value to an event structure, replete with modifier bits. ch is the character to convert, and event is the event object to fill in. This function contains knowledge about what the codes “mean” – for example, the number 9 is converted to the character ^{⟨}\underline{\overline{Tab}}^{⟩}, not the distinct character ^{⟨}\underline{\overline{Control-I}}^{⟩}. Note that ch does not have to be a numeric value, but can be a symbol such as

clear or a list such as (control backspace).

If event is not nil, it is modified; otherwise, a new event object is created. In both cases, the event is returned. Optional third arg device is the device to store in the event; this also affects whether the high bit is interpreted as a meta key. A value of nil means use the selected device but always treat the high bit as meta.

Beware that character-to-event and event-to-character are not strictly

inverse functions, since events contain much more information than the ASCII character set can encode.

Function event-to-character event &optional allow-extra-modifiers

allow-meta allow-non-ascii This function returns the closest ASCII approximation to event. If the event isn’t a keypress, this returns nil. If allow-extra-modifiers is non-nil, then this is lenient in its translation; it will ignore modifier keys other than^{⟨}\underline{\overline{control}}^{⟩} and^{⟨}\underline{\overline{meta}}^{⟩}, and will ignore the^{⟨}\underline{\overline{shift}}^{⟩} modifier on those characters which have no shifted ASCII equivalent (^{⟨}\underline{\overline{Control-Shift-A}}^{⟩} for example, will be mapped to the same ASCII code as ^{⟨}\underline{\overline{Control-A}}^{⟩}). If allow-meta is non-nil, then the ^{⟨}\underline{\overline{Meta}}^{⟩} modifier will be represented by turning on the high bit of the byte returned; otherwise, nil will be returned for events containing the ^{⟨}\underline{\overline{Meta}}^{⟩} modifier. If allow-non-ascii is non-nil, then characters which are present in the prevailing character set (see Chapter 20 [Keymaps], page 315) will be returned as their code in that character set, instead of the return value being restricted to ASCII. Note that specifying both allow-meta and allow-non-ascii is ambiguous, as both use the high bit; ^{⟨}\underline{\overline{M-x}}^{⟩} and ^{⟨}\underline{\overline{oslash}}^{⟩} will be indistinguishable.

Function events-to-keys events &optional no-mice Given a vector of event objects, this function returns a vector of key descriptors, or a string (if they all fit in the ASCII range). Optional arg no-mice means that button events are not allowed.

19.6 Reading Input

The editor command loop reads keyboard input using the function next-event and constructs key sequences out of the events using dispatch-event. Lisp programs can also use the function read-key-sequence, which reads input a key sequence at a time. See also momentary-string-display in Section 45.8 [Temporary Displays], page 656, and sitfor in Section 19.7 [Waiting], page 305. See Section 50.8 [Terminal Input], page 706, for functions and variables for controlling terminal input modes and debugging terminal input.

For higher-level input facilities, see Chapter 18 [Minibuffers], page 261.

19.6.1 Key Sequence Input

Lisp programs can read input a key sequence at a time by calling read-key-sequence; for example, describe-key uses it to read the key to describe.

Function read-key-sequence prompt

This function reads a sequence of keystrokes or mouse clicks and returns it as a vector of events. It keeps reading events until it has accumulated a full key sequence; that is, enough to specify a non-prefix command using the currently active keymaps. The vector and the event objects it contains are freshly created, and will not be side-effected by subsequent calls to this function. The function read-key-sequence suppresses quitting: C-g typed while reading with this function works like any other character, and does not set quit-flag. See Section 19.8 [Quitting], page 306. The argument prompt is either a string to be displayed in the echo area as a prompt, or nil, meaning not to display a prompt. If the user selects a menu item while we are prompting for a key sequence, the returned value will be a vector of a single menu-selection event (a misc-user event). An error will be signalled if you pass this value to lookup-key or a related function. In the example below, the prompt ‘?’ is displayed in the echo area, and the user

types C-x C-f.
     (read-key-sequence "?")
     ---------- Echo Area ----------
     ?C-x C-f
     ---------- Echo Area ----------
          ⇒ [#<keypress-event control-X> #<keypress-event control-
     F>]

If an input character is an upper-case letter and has no key binding, but its lower-case equivalent has one, then read-key-sequence converts the character to lower case. Note that lookup-key does not perform case conversion in this way.

19.6.2 Reading One Event

The lowest level functions for command input are those which read a single event. These functions often make a distinction between command events, which are user actions (keystrokes and mouse actions), and other events, which serve as communication between XEmacs and the window system.

Function next-event &optional event prompt This function reads and returns the next available event from the window system or terminal driver, waiting if necessary until an event is available. Pass this object to dispatch-event to handle it. If an event object is supplied, it is filled in and returned; otherwise a new event object will be created. Events can come directly from the user, from a keyboard macro, or from

unread-command-events.

In most cases, the function next-command-event is more appropriate.

Function next-command-event &optional event

This function returns the next available “user” event from the window system or terminal driver. Pass this object to dispatch-event to handle it. If an event object is supplied, it is filled in and returned, otherwise a new event object will be created.

The event returned will be a keyboard, mouse press, or mouse release event. If there are non-command events available (mouse motion, sub-process output, etc) then these will be executed (with dispatch-event) and discarded. This function is provided as a convenience; it is equivalent to the Lisp code

     (while (progn
      (next-event event)
              (not (or (key-press-event-p event)
                       (button-press-event-p event)
                       (button-release-event-p event)
                       (menu-event-p event))))
        (dispatch-event event))

Here is what happens if you call next-command-event and then press the rightarrow function key:

     (next-command-event)
          ⇒ #<keypress-event right>

Function read-char

This function reads and returns a character of command input. If a mouse click is detected, an error is signalled. The character typed is returned as an ASCII value. This function is retained for compatibility with Emacs 18, and is most likely the wrong thing for you to be using: consider using next-command-event instead.

Function enqueue-eval-event function object

This function adds an eval event to the back of the queue. The eval event will be the next event read after all pending events.

19.6.3 Dispatching an Event

Function dispatch-event event

Given an event object returned by next-event, this function executes it. This is the basic function that makes XEmacs respond to user input; it also deals with notifications from the window system (such as Expose events).

19.6.4 Quoted Character Input

You can use the function read-quoted-char to ask the user to specify a character, and allow the user to specify a control or meta character conveniently, either literally or as an octal character code. The command quoted-insert uses this function.

Function read-quoted-char &optional prompt

This function is like read-char, except that if the first character read is an octal digit (0-7), it reads up to two more octal digits (but stopping if a nonoctal digit is found) and returns the character represented by those digits in octal.

Quitting is suppressed when the first character is read, so that the user can enter a C-g. See Section 19.8 [Quitting], page 306.

If prompt is supplied, it specifies a string for prompting the user. The prompt string is always displayed in the echo area, followed by a single ‘-’.

In the following example, the user types in the octal number 177 (which is 127 in decimal).

     (read-quoted-char "What character")
     ---------- Echo Area ----------
     What character-177
     ---------- Echo Area ----------

          ⇒ 127

19.6.5 Miscellaneous Event Input Features

This section describes how to “peek ahead” at events without using them up, how to check for pending input, and how to discard pending input.

See also the variables last-command-event and last-command-char (Section 19.4

[Command Loop Info], page 288).

Variable unread-command-events

This variable holds a list of events waiting to be read as command input. The events are used in the order they appear in the list, and removed one by one as they are used.

The variable is needed because in some cases a function reads a event and then decides not to use it. Storing the event in this variable causes it to be processed normally, by the command loop or by the functions to read command input.

For example, the function that implements numeric prefix arguments reads any number of digits. When it finds a non-digit event, it must unread the event so that it can be read normally by the command loop. Likewise, incremental search uses this feature to unread events with no special meaning in a search, because these events should exit the search and then execute normally.

Variable unread-command-event

This variable holds a single event to be read as command input.

This variable is mostly obsolete now that you can use unread-command-events instead; it exists only to support programs written for versions of XEmacs prior to 19.12.

Function input-pending-p

This function determines whether any command input is currently available to be read. It returns immediately, with value t if there is available input, nil otherwise. On rare occasions it may return t when no input is available.

Variable last-input-event

This variable is set to the last keyboard or mouse button event received. This variable is off limits: you may not set its value or modify the event that is its value, as it is destructively modified by read-key-sequence. If you want to keep a pointer to this value, you must use copy-event. Note that this variable is an alias for last-input-char in FSF Emacs. In the example below, a character is read (the character 1). It becomes the value of last-input-event, while C-e (from the C-x C-e command used to evaluate this expression) remains the value of last-command-event.

     (progn (print (next-command-event))
            (print last-command-event)
            last-input-event)
          ⊣ #<keypress-event 1>
          ⊣ #<keypress-event control-E>
          ⇒ #<keypress-event 1>

Variable last-input-char

If the value of last-input-event is a keyboard event, then this is the nearest ASCII equivalent to it. Remember that there is not a 1:1 mapping between keyboard events and ASCII characters: the set of keyboard events is much larger, so writing code that examines this variable to determine what key has been typed is bad practice, unless you are certain that it will be one of a small set of characters. This function exists for compatibility with Emacs version 18.

Function discard-input

This function discards the contents of the terminal input buffer and cancels any keyboard macro that might be in the process of definition. It returns nil. In the following example, the user may type a number of characters right after starting the evaluation of the form. After the sleep-for finishes sleeping, discard-input discards any characters typed during the sleep.

     (progn (sleep-for 2)
            (discard-input))
          ⇒ nil

19.7 Waiting for Elapsed Time or Input

The wait functions are designed to wait for a certain amount of time to pass or until there is input. For example, you may wish to pause in the middle of a computation to allow the user time to view the display. sit-for pauses and updates the screen, and returns immediately if input comes in, while sleep-for pauses without updating the screen.

Note that in FSF Emacs, the commands sit-for and sleep-for take two arguments to specify the time (one integer and one float value), instead of a single argument that can be either an integer or a float.

Function sit-for seconds &optional nodisp This function performs redisplay (provided there is no pending input from the user), then waits seconds seconds, or until input is available. The result is t if sit-for waited the full time with no input arriving (see input-pending-p in Section 19.6.5 [Peeking and Discarding], page 304). Otherwise, the value is

nil.

The argument seconds need not be an integer. If it is a floating point number, sit-for waits for a fractional number of seconds. Redisplay is normally preempted if input arrives, and does not happen at all if input is available before it starts. (You can force screen updating in such a case by using force-redisplay. See Section 45.1 [Refresh Screen], page 647.) If there is no input pending, you can force an update with no delay by using

(sit-for 0).

If nodisp is non-nil, then sit-for does not redisplay, but it still returns as soon as input is available (or when the timeout elapses). The usual purpose of sit-for is to give the user time to read text that you display.

Function sleep-for seconds

This function simply pauses for seconds seconds without updating the display. This function pays no attention to available input. It returns nil. The argument seconds need not be an integer. If it is a floating point number, sleep-for waits for a fractional number of seconds. Use sleep-for when you wish to guarantee a delay.

See Section 50.5 [Time of Day], page 702, for functions to get the current time.

19.8 Quitting

Typing C-g while a Lisp function is running causes XEmacs to quit whatever it is doing. This means that control returns to the innermost active command loop.

Typing C-g while the command loop is waiting for keyboard input does not cause a quit; it acts as an ordinary input character. In the simplest case, you cannot tell the difference, because C-g normally runs the command keyboard-quit, whose effect is to quit. However, when C-g follows a prefix key, the result is an undefined key. The effect is to cancel the prefix key as well as any prefix argument.

In the minibuffer, C-g has a different definition: it aborts out of the minibuffer. This means, in effect, that it exits the minibuffer and then quits. (Simply quitting would return to the command loop within the minibuffer.) The reason why C-g does not quit directly when the command reader is reading input is so that its meaning can be redefined in the minibuffer in this way. C-g following a prefix key is not redefined in the minibuffer, and it has its normal effect of canceling the prefix key and prefix argument. This too would not be possible if C-g always quit directly.

When C-g does directly quit, it does so by setting the variable quit-flag to t. XEmacs checks this variable at appropriate times and quits if it is not nil. Setting quit-flag non-nil in any way thus causes a quit.

At the level of C code, quitting cannot happen just anywhere; only at the special places that check quit-flag. The reason for this is that quitting at other places might leave an inconsistency in XEmacs’s internal state. Because quitting is delayed until a safe place, quitting cannot make XEmacs crash.

Certain functions such as read-key-sequence or read-quoted-char prevent quitting entirely even though they wait for input. Instead of quitting, C-g serves as the requested input. In the case of read-key-sequence, this serves to bring about the special behavior of C-g in the command loop. In the case of read-quoted-char, this is so that C-q can be used to quote a C-g.

You can prevent quitting for a portion of a Lisp function by binding the variable inhibit-quit to a non-nil value. Then, although C-g still sets quit-flag to t as usual, the usual result of this—a quit—is prevented. Eventually, inhibit-quit will become nil again, such as when its binding is unwound at the end of a let form. At that time, if quit-flag is still non-nil, the requested quit happens immediately. This behavior is ideal when you wish to make sure that quitting does not happen within a “critical section” of the program.

In some functions (such as read-quoted-char), C-g is handled in a special way that does not involve quitting. This is done by reading the input with inhibit-quit bound to t, and setting quit-flag to nil before inhibit-quit becomes nil again. This excerpt from the definition of read-quoted-char shows how this is done; it also shows that normal quitting is permitted after the first character of input.

     (defun read-quoted-char (&optional prompt)

"...documentation..."

       (let ((count 0) (code 0) char)
         (while (< count 3)
           (let ((inhibit-quit (zerop count))
                 (help-form nil))
             (and prompt (message "%s-" prompt))
             (setq char (read-char))
             (if inhibit-quit (setq quit-flag nil)))
           ...)
         (logand 255 code)))

Variable quit-flag

If this variable is non-nil, then XEmacs quits immediately, unless inhibitquit is non-nil. Typing C-g ordinarily sets quit-flag non-nil, regardless of

     inhibit-quit.

Variable inhibit-quit

This variable determines whether XEmacs should quit when quit-flag is set to a value other than nil. If inhibit-quit is non-nil, then quit-flag has no special effect.

Command keyboard-quit

This function signals the quit condition with (signal 'quit nil). This is the same thing that quitting does. (See signal in Section 9.5.3 [Errors], page 136.)

You can specify a character other than C-g to use for quitting. See the function setinput-mode in Section 50.8 [Terminal Input], page 706.

19.9 Prefix Command Arguments

Most XEmacs commands can use a prefix argument, a number specified before the command itself. (Don’t confuse prefix arguments with prefix keys.) The prefix argument is at all times represented by a value, which may be nil, meaning there is currently no prefix argument. Each command may use the prefix argument or ignore it.

There are two representations of the prefix argument: raw and numeric. The editor command loop uses the raw representation internally, and so do the Lisp variables that store the information, but commands can request either representation.

Here are the possible values of a raw prefix argument:

• nil, meaning there is no prefix argument. Its numeric value is 1, but numerous commands make a distinction between nil and the integer 1.

• An integer, which stands for itself.

• A list of one element, which is an integer. This form of prefix argument results from one or a succession of C-u’s with no digits. The numeric value is the integer in the list, but some commands make a distinction between such a list and an integer alone.

• The symbol -. This indicates that M-- or C-u - was typed, without following digits. The equivalent numeric value is −1, but some commands make a distinction between the integer −1 and the symbol -.

We illustrate these possibilities by calling the following function with various prefixes:

(defun display-prefix (arg)
  "Display the value of the raw prefix arg."
  (interactive "P")
  (message "%s" arg))

Here are the results of calling display-prefix with various raw prefix arguments:

        M-x display-prefix ⊣ nil
C-u     M-x display-prefix ⊣ (4)
C-u C-u M-x display-prefix ⊣ (16)
C-u 3   M-x display-prefix ⊣ 3

M-3     M-x display-prefix ⊣ 3      ; (Same as C-u 3.)
C-3     M-x display-prefix ⊣ 3      ; (Same as C-u 3.)
C-u -   M-x display-prefix ⊣ -
M--     M-x display-prefix ⊣ -      ; (Same as C-u -.)
C--     M-x display-prefix ⊣ -      ; (Same as C-u -.)
C-u - 7 M-x display-prefix ⊣ -7
M-- 7   M-x display-prefix ⊣ -7     ; (Same as C-u -7.)
C-- 7   M-x display-prefix ⊣ -7     ; (Same as C-u -7.)

XEmacs uses two variables to store the prefix argument: prefix-arg and currentprefix-arg. Commands such as universal-argument that set up prefix arguments for other commands store them in prefix-arg. In contrast, current-prefix-arg conveys the prefix argument to the current command, so setting it has no effect on the prefix arguments for future commands.

Normally, commands specify which representation to use for the prefix argument, either numeric or raw, in the interactive declaration. (See Section 19.2.1 [Using Interactive], page 282.) Alternatively, functions may look at the value of the prefix argument directly in the variable current-prefix-arg, but this is less clean.

Function prefix-numeric-value arg

This function returns the numeric meaning of a valid raw prefix argument value, arg. The argument may be a symbol, a number, or a list. If it is nil, the value 1 is returned; if it is -, the value −1 is returned; if it is a number, that number is returned; if it is a list, the car of that list (which should be a number) is returned.

Variable current-prefix-arg

This variable holds the raw prefix argument for the current command. Commands may examine it directly, but the usual way to access it is with (interactive

"P").

Variable prefix-arg

The value of this variable is the raw prefix argument for the next editing command. Commands that specify prefix arguments for the following command work by setting this variable.

Do not call the functions universal-argument, digit-argument, or negative-argument

unless you intend to let the user enter the prefix argument for the next command.

Command universal-argument

This command reads input and specifies a prefix argument for the following command. Don’t call this command yourself unless you know what you are doing.

Command digit-argument arg

This command adds to the prefix argument for the following command. The argument arg is the raw prefix argument as it was before this command; it is used to compute the updated prefix argument. Don’t call this command yourself unless you know what you are doing.

Command negative-argument arg

This command adds to the numeric argument for the next command. The argument arg is the raw prefix argument as it was before this command; its value is negated to form the new prefix argument. Don’t call this command yourself unless you know what you are doing.

19.10 Recursive Editing

The XEmacs command loop is entered automatically when XEmacs starts up. This top-level invocation of the command loop never exits; it keeps running as long as XEmacs does. Lisp programs can also invoke the command loop. Since this makes more than one activation of the command loop, we call it recursive editing. A recursive editing level has the effect of suspending whatever command invoked it and permitting the user to do arbitrary editing before resuming that command.

The commands available during recursive editing are the same ones available in the top-level editing loop and defined in the keymaps. Only a few special commands exit the recursive editing level; the others return to the recursive editing level when they finish. (The special commands for exiting are always available, but they do nothing when recursive editing is not in progress.)

All command loops, including recursive ones, set up all-purpose error handlers so that an error in a command run from the command loop will not exit the loop.

Minibuffer input is a special kind of recursive editing. It has a few special wrinkles, such as enabling display of the minibuffer and the minibuffer window, but fewer than you might suppose. Certain keys behave differently in the minibuffer, but that is only because of the minibuffer’s local map; if you switch windows, you get the usual XEmacs commands.

To invoke a recursive editing level, call the function recursive-edit. This function contains the command loop; it also contains a call to catch with tag exit, which makes it possible to exit the recursive editing level by throwing to exit (see Section 9.5.1 [Catch and Throw], page 134). If you throw a value other than t, then recursive-edit returns normally to the function that called it. The command C-M-c (exit-recursive-edit) does this. Throwing a t value causes recursive-edit to quit, so that control returns to the command loop one level up. This is called aborting, and is done by C-] (abort-recursive-

edit).

Most applications should not use recursive editing, except as part of using the minibuffer. Usually it is more convenient for the user if you change the major mode of the current buffer temporarily to a special major mode, which should have a command to go back to the previous mode. (The e command in Rmail uses this technique.) Or, if you wish to give the user different text to edit “recursively”, create and select a new buffer in a special mode. In this mode, define a command to complete the processing and go back to the previous buffer. (The m command in Rmail does this.)

Recursive edits are useful in debugging. You can insert a call to debug into a function definition as a sort of breakpoint, so that you can look around when the function gets there. debug invokes a recursive edit but also provides the other features of the debugger.

Recursive editing levels are also used when you type C-r in query-replace or use C-x

q (kbd-macro-query).

Function recursive-edit

This function invokes the editor command loop. It is called automatically by the initialization of XEmacs, to let the user begin editing. When called from a Lisp program, it enters a recursive editing level. In the following example, the function simple-rec first advances point one word, then enters a recursive edit, printing out a message in the echo area. The user can then do any editing desired, and then type C-M-c to exit and continue

executing simple-rec.
     (defun simple-rec ()
       (forward-word 1)
       (message "Recursive edit in progress")
       (recursive-edit)
       (forward-word 1))
          ⇒ simple-rec
     (simple-rec)
          ⇒ nil

Command exit-recursive-edit

This function exits from the innermost recursive edit (including minibuffer input). Its definition is effectively (throw 'exit nil).

Command abort-recursive-edit

This function aborts the command that requested the innermost recursive edit (including minibuffer input), by signaling quit after exiting the recursive edit. Its definition is effectively (throw 'exit t). See Section 19.8 [Quitting], page 306.

Command top-level

This function exits all recursive editing levels; it does not return a value, as it jumps completely out of any computation directly back to the main command loop.

Function recursion-depth

This function returns the current depth of recursive edits. When no recursive edit is active, it returns 0.

19.11 Disabling Commands

Disabling a command marks the command as requiring user confirmation before it can be executed. Disabling is used for commands which might be confusing to beginning users, to prevent them from using the commands by accident.

The low-level mechanism for disabling a command is to put a non-nildisabled property on the Lisp symbol for the command. These properties are normally set up by the user’s ‘.emacs’ file with Lisp expressions such as this:

(put 'upcase-region 'disabled t)

For a few commands, these properties are present by default and may be removed by the

‘.emacs’ file.

If the value of the disabled property is a string, the message saying the command is disabled includes that string. For example:

(put 'delete-region 'disabled
     "Text deleted this way cannot be yanked back!\n")

See section “Disabling” in The XEmacs User’s Manual, for the details on what happens when a disabled command is invoked interactively. Disabling a command has no effect on calling it as a function from Lisp programs.

Command enable-command command

Allow command to be executed without special confirmation from now on, and (if the user confirms) alter the user’s ‘.emacs’ file so that this will apply to future sessions.

Command disable-command command

Require special confirmation to execute command from now on, and (if the user confirms) alter the user’s ‘.emacs’ file so that this will apply to future sessions.

Variable disabled-command-hook

This normal hook is run instead of a disabled command, when the user invokes the disabled command interactively. The hook functions can use thiscommand-keys to determine what the user typed to run the command, and thus find the command itself. See Section 26.4 [Hooks], page 378. By default, disabled-command-hook contains a function that asks the user whether to proceed.

19.12 Command History

The command loop keeps a history of the complex commands that have been executed, to make it convenient to repeat these commands. A complex command is one for which the interactive argument reading uses the minibuffer. This includes any M-x command, any M-: command, and any command whose interactive specification reads an argument from the minibuffer. Explicit use of the minibuffer during the execution of the command itself does not cause the command to be considered complex.

Variable command-history

This variable’s value is a list of recent complex commands, each represented as a form to evaluate. It continues to accumulate all complex commands for the duration of the editing session, but all but the first (most recent) thirty elements are deleted when a garbage collection takes place (see Section B.3 [Garbage Collection], page 768).

     command-history
     ⇒ ((switch-to-buffer "chistory.texi")
         (describe-key "^X^[")
         (visit-tags-table "~/emacs/src/")
         (find-tag "repeat-complex-command"))

This history list is actually a special case of minibuffer history (see Section 18.4 [Minibuffer History], page 264), with one special twist: the elements are expressions rather than strings.

There are a number of commands devoted to the editing and recall of previous commands.

The commands repeat-complex-command, and list-command-history are described in

the user manual (see section “Repetition” in The XEmacs User’s Manual). Within the minibuffer, the history commands used are the same ones available in any minibuffer.

19.13 Keyboard Macros

A keyboard macro is a canned sequence of input events that can be considered a command and made the definition of a key. The Lisp representation of a keyboard macro is a string or vector containing the events. Don’t confuse keyboard macros with Lisp macros (see Chapter 12 [Macros], page 177).

Function execute-kbd-macro macro &optional count This function executes macro as a sequence of events. If macro is a string or vector, then the events in it are executed exactly as if they had been input by the user. The sequence is not expected to be a single key sequence; normally a keyboard macro definition consists of several key sequences concatenated. If macro is a symbol, then its function definition is used in place of macro. If that is another symbol, this process repeats. Eventually the result should be a string or vector. If the result is not a symbol, string, or vector, an error is signaled. The argument count is a repeat count; macro is executed that many times. If count is omitted or nil, macro is executed once. If it is 0, macro is executed over and over until it encounters an error or a failing search.

Variable executing-macro

This variable contains the string or vector that defines the keyboard macro that is currently executing. It is nil if no macro is currently executing. A command can test this variable to behave differently when run from an executing macro. Do not set this variable yourself.

Variable defining-kbd-macro

This variable indicates whether a keyboard macro is being defined. A command can test this variable to behave differently while a macro is being defined. The commands start-kbd-macro and end-kbd-macro set this variable—do not set it yourself.

Variable last-kbd-macro

This variable is the definition of the most recently defined keyboard macro. Its value is a string or vector, or nil.

314 XEmacs Lisp Reference Manual

The commands are described in the user’s manual (see section “Keyboard Macros” in The XEmacs User’s Manual).

20 Keymaps

The bindings between input events and commands are recorded in data structures called keymaps. Each binding in a keymap associates (or binds) an individual event type either with another keymap or with a command. When an event is bound to a keymap, that keymap is used to look up the next input event; this continues until a command is found. The whole process is called key lookup.

20.1 Keymap Terminology

A keymap is a table mapping event types to definitions (which can be any Lisp objects, though only certain types are meaningful for execution by the command loop). Given an event (or an event type) and a keymap, XEmacs can get the event’s definition. Events mapped in keymaps include keypresses, button presses, and button releases (see Section 19.5 [Events], page 290).

A sequence of input events that form a unit is called a key sequence, or key for short. A sequence of one event is always a key sequence, and so are some multi-event sequences.

A keymap determines a binding or definition for any key sequence. If the key sequence is a single event, its binding is the definition of the event in the keymap. The binding of a key sequence of more than one event is found by an iterative process: the binding of the first event is found, and must be a keymap; then the second event’s binding is found in that keymap, and so on until all the events in the key sequence are used up.

If the binding of a key sequence is a keymap, we call the key sequence a prefix key. Otherwise, we call it a complete key (because no more events can be added to it). If the binding is nil, we call the key undefined. Examples of prefix keys are C-c, C-x, and C-x 4. Examples of defined complete keys are X, ^{⟨}\underline{\overline{RET}}^{⟩}, and C-x 4 C-f. Examples of undefined complete keys are C-x C-g, and C-c 3. See Section 20.6 [Prefix Keys], page 319, for more details.

The rule for finding the binding of a key sequence assumes that the intermediate bindings (found for the events before the last) are all keymaps; if this is not so, the sequence of events does not form a unit—it is not really a key sequence. In other words, removing one or more events from the end of any valid key must always yield a prefix key. For example, C-f C-n is not a key; C-f is not a prefix key, so a longer sequence starting with C-f cannot be a key.

Note that the set of possible multi-event key sequences depends on the bindings for prefix keys; therefore, it can be different for different keymaps, and can change when bindings are changed. However, a one-event sequence is always a key sequence, because it does not depend on any prefix keys for its well-formedness.

At any time, several primary keymaps are active—that is, in use for finding key bindings. These are the global map, which is shared by all buffers; the local keymap, which is usually associated with a specific major mode; and zero or more minor mode keymaps, which belong to currently enabled minor modes. (Not all minor modes have keymaps.) The local keymap bindings shadow (i.e., take precedence over) the corresponding global bindings. The minor mode keymaps shadow both local and global keymaps. See Section 20.7 [Active Keymaps], page 320, for details.

20.2 Format of Keymaps

A keymap is a primitive type that associates events with their bindings. Note that this is different from Emacs 18 and FSF Emacs, where keymaps are lists.

Function keymapp object

This function returns t if object is a keymap, nil otherwise.

20.3 Creating Keymaps

Here we describe the functions for creating keymaps.

Function make-keymap &optional name This function constructs and returns a new keymap object. All entries in it are nil, meaning “command undefined”. Optional argument name specifies a name to assign to the keymap, as in setkeymap-name. This name is only a debugging convenience; it is not used except when printing the keymap.

Function make-sparse-keymap &optional name

This function constructs and returns a new keymap object. All entries in it are nil, meaning “command undefined”. The only difference between this function and make-keymap is that this function returns a “smaller” keymap (one that is expected to contain fewer entries). As keymaps dynamically resize, the distinction is not great. Optional argument name specifies a name to assign to the keymap, as in setkeymap-name. This name is only a debugging convenience; it is not used except when printing the keymap.

Function set-keymap-name keymap new-name

This function assigns a “name” to a keymap. The name is only a debugging convenience; it is not used except when printing the keymap.

Function keymap-name keymap

This function returns the “name” of a keymap, as assigned using set-keymap-

name.

Function copy-keymap keymap

This function returns a copy of keymap. Any keymaps that appear directly as bindings in keymap are also copied recursively, and so on to any number of levels. However, recursive copying does not take place when the definition of a character is a symbol whose function definition is a keymap; the same symbol appears in the new copy.

     (setq map (copy-keymap (current-local-map)))
     ⇒ #<keymap 3 entries 0x21f80>
     (eq map (current-local-map))
         ⇒ nil

20.4 Inheritance and Keymaps

A keymap can inherit the bindings of other keymaps. The other keymaps are called the keymap’s parents, and are set with set-keymap-parents. When searching for a binding for a key sequence in a particular keymap, that keymap itself will first be searched; then, if no binding was found in the map and it has parents, the first parent keymap will be searched; then that keymap’s parent will be searched, and so on, until either a binding for the key sequence is found, or a keymap without a parent is encountered. At this point, the search will continue with the next parent of the most recently encountered keymap that has another parent, etc. Essentially, a depth-first search of all the ancestors of the keymap is conducted.

(current-global-map) is the default parent of all keymaps.

Function set-keymap-parents keymap parents

This function sets the parent keymaps of keymap to the list parents.

If you change the bindings in one of the keymaps in parents using definekey or other key-binding functions, these changes are visible in keymap unless shadowed by bindings in that map or in earlier-searched ancestors. The converse is not true: if you use define-key to change keymap, that affects the bindings in that map, but has no effect on any of the keymaps in parents.

Function keymap-parents keymap

This function returns the list of parent keymaps of keymap, or nil if keymap has no parents.

As an alternative to specifying a parent, you can also specify a default binding that is used whenever a key is not otherwise bound in the keymap. This is useful for terminal emulators, for example, which may want to trap all keystrokes and pass them on in some modified format. Note that if you specify a default binding for a keymap, neither the keymap’s parents nor the current global map are searched for key bindings.

Function set-keymap-default-binding keymap command

This function sets the default binding of keymap to command, or nil if no default is desired.

Function keymap-default-binding keymap

This function returns the default binding of keymap, or nil if it has none.

20.5 Key Sequences

Contrary to popular belief, the world is not ASCII. When running under a window manager, XEmacs can tell the difference between, for example, the keystrokes control-h, control-shift-h, and backspace. You can, in fact, bind different commands to each of these.

A key sequence is a set of keystrokes. A keystroke is a keysym and some set of modifiers (such as ^{⟨}\underline{\overline{CONTROL}}^{⟩} and ^{⟨}\underline{\overline{META}}^{⟩}). A keysym is what is printed on the keys on your keyboard.

A keysym may be represented by a symbol, or (if and only if it is equivalent to an ASCII character in the range 32 - 255) by a character or its equivalent ASCII code. The A key may be represented by the symbol A, the character ?A, or by the number 65. The break key may be represented only by the symbol break.

A keystroke may be represented by a list: the last element of the list is the key (a symbol, character, or number, as above) and the preceding elements are the symbolic names of modifier keys (^{⟨}\underline{\overline{CONTROL}}^{⟩}, ^{⟨}\underline{\overline{META}}^{⟩}, ^{⟨}\underline{\overline{SUPER}}^{⟩}, ^{⟨}\underline{\overline{HYPER}}^{⟩}, ^{⟨}\underline{\overline{ALT}}^{⟩}, and ^{⟨}\underline{\overline{SHIFT}}^{⟩}). Thus, the sequence

control-b is represented by the forms (control b), (control ?b), and (control 98). A

keystroke may also be represented by an event object, as returned by the next-command-

event and read-key-sequence functions.

Note that in this context, the keystroke control-b is not represented by the number 2 (the ASCII code for ‘^B’) or the character ?\^B. See below.

The ^{⟨}\underline{\overline{SHIFT}}^{⟩} modifier is somewhat of a special case. You should not (and cannot) use (meta shift a) to mean (meta A), since for characters that have ASCII equivalents, the state of the shift key is implicit in the keysym (‘a’ vs. ‘A’). You also cannot say (shift =) to mean +, as that sort of thing varies from keyboard to keyboard. The ^{⟨}\underline{\overline{SHIFT}}^{⟩} modifier is for use only with characters that do not have a second keysym on the same key, such as

backspace and tab.

A key sequence is a vector of keystrokes. As a degenerate case, elements of this vector may also be keysyms if they have no modifiers. That is, the A keystroke is represented by all of these forms:

     A ?A 65 (A) (?A) (65)
     [A] [?A] [65] [(A)] [(?A)] [(65)]

the control-a keystroke is represented by these forms:

     (control A) (control ?A) (control 65)
     [(control A)] [(control ?A)] [(control 65)]

the key sequence control-c control-a is represented by these forms:

     [(control c) (control a)] [(control ?c) (control ?a)]
     [(control 99) (control 65)] etc.

Mouse button clicks work just like keypresses: (control button1) means pressing the left mouse button while holding down the control key. [(control c) (shift button3)] means control-c, hold ^{⟨}\underline{\overline{SHIFT}}^{⟩}, click right.

Commands may be bound to the mouse-button up-stroke rather than the down-stroke as well. button1 means the down-stroke, and button1up means the up-stroke. Different commands may be bound to the up and down strokes, though that is probably not what you want, so be careful.

For backward compatibility, a key sequence may also be represented by a string. In this case, it represents the key sequence(s) that would produce that sequence of ASCII characters in a purely ASCII world. For example, a string containing the ASCII backspace character, "\^H", would represent two key sequences: (control h) and backspace. Binding a command to this will actually bind both of those key sequences. Likewise for the following pairs:

     control h backspace
     control i    tab
     control m    return
     control j    linefeed

     control [    escape
     control @ control space

After binding a command to two key sequences with a form like

     (define-key global-map "\^X\^I" 'command-1)

it is possible to redefine only one of those sequences like so:

     (define-key global-map [(control x) (control i)] 'command-2)
     (define-key global-map [(control x) tab] 'command-3)

Of course, all of this applies only when running under a window system. If you’re talking to XEmacs through a TTY connection, you don’t get any of these features.

Function event-matches-key-specifier-p event key-specifier

This function returns non-nil if event matches key-specifier, which can be any valid form representing a key sequence. This can be useful, e.g., to determine if the user pressed help-char or quit-char.

20.6 Prefix Keys

A prefix key has an associated keymap that defines what to do with key sequences that start with the prefix key. For example, C-x is a prefix key, and it uses a keymap that is also stored in the variable ctl-x-map. Here is a list of the standard prefix keys of XEmacs and their keymaps:

• help-map is used for events that follow C-h.

• mode-specific-map is for events that follow C-c. This map is not actually mode specific; its name was chosen to be informative for the user in C-h b (display-bindings), where it describes the main use of the C-c prefix key.

• ctl-x-map is the map used for events that follow C-x. This map is also the function

    definition of Control-X-prefix.

• ctl-x-4-map is used for events that follow C-x 4.

• ctl-x-5-map is used for events that follow C-x 5.

• The prefix keys C-x n, C-x r and C-x a use keymaps that have no special name.

• esc-map is an evil hack that is present for compatibility purposes with Emacs 18. Defining a key in esc-map is equivalent to defining the same key in global-map but with the ^{⟨}\underline{\overline{META}}^{⟩} prefix added. You should not use this in your code. (This map is also the function definition of ESC-prefix.)

The binding of a prefix key is the keymap to use for looking up the events that follow the prefix key. (It may instead be a symbol whose function definition is a keymap. The effect is the same, but the symbol serves as a name for the prefix key.) Thus, the binding of C-x is the symbol Control-X-prefix, whose function definition is the keymap for C-x commands. (The same keymap is also the value of ctl-x-map.)

Prefix key definitions can appear in any active keymap. The definitions of C-c, C-x, C-h and ^{⟨}\underline{\overline{ESC}}^{⟩} as prefix keys appear in the global map, so these prefix keys are always available. Major and minor modes can redefine a key as a prefix by putting a prefix key definition for it in the local map or the minor mode’s map. See Section 20.7 [Active Keymaps], page 320.

If a key is defined as a prefix in more than one active map, then its various definitions are in effect merged: the commands defined in the minor mode keymaps come first, followed by those in the local map’s prefix definition, and then by those from the global map.

In the following example, we make C-p a prefix key in the local keymap, in such a way that C-p is identical to C-x. Then the binding for C-p C-f is the function find-file, just like C-x C-f. The key sequence C-p 6 is not found in any active keymap.

     (use-local-map (make-sparse-keymap))
         ⇒ nil
     (local-set-key "\C-p" ctl-x-map)
         ⇒ nil
     (key-binding "\C-p\C-f")
         ⇒ find-file
     (key-binding "\C-p6")
         ⇒ nil

Function define-prefix-command symbol &optional mapvar This function defines symbol as a prefix command: it creates a keymap and stores it as symbol’s function definition. Storing the symbol as the binding of a key makes the key a prefix key that has a name. If optional argument mapvar is not specified, it also sets symbol as a variable, to have the keymap as its value. (If mapvar is given and is not t, its value is stored as the value of symbol.) The function returns symbol. In Emacs version 18, only the function definition of symbol was set, not the value as a variable.

20.7 Active Keymaps

XEmacs normally contains many keymaps; at any given time, just a few of them are active in that they participate in the interpretation of user input. These are the global keymap, the current buffer’s local keymap, and the keymaps of any enabled minor modes.

The global keymap holds the bindings of keys that are defined regardless of the current buffer, such as C-f. The variable global-map holds this keymap, which is always active.

Each buffer may have another keymap, its local keymap, which may contain new or overriding definitions for keys. The current buffer’s local keymap is always active except

when overriding-local-map or overriding-terminal-local-map overrides it. Extents

and text properties can specify an alternative local map for certain parts of the buffer; see Section 40.10 [Extents and Events], page 597.

Each minor mode may have a keymap; if it does, the keymap is active when the minor mode is enabled.

   The variable overriding-local-map and overriding-terminal-local-map, if non-

nil, specify other local keymaps that override the buffer’s local map and all the minor mode keymaps.

All the active keymaps are used together to determine what command to execute when a key is entered. XEmacs searches these maps one by one, in order of decreasing precedence, until it finds a binding in one of the maps.

More specifically:

For key-presses, the order of keymaps searched is:

• the keymap property of any extent(s) or text properties at point;

• any applicable minor-mode maps;

• the current local map of the current buffer;

• the current global map.

For mouse-clicks, the order of keymaps searched is:

• the current local map of the mouse-grabbed-buffer if any;

• the keymap property of any extent(s) at the position of the click (this includes modeline extents);

• the modeline-map of the buffer corresponding to the modeline under the mouse (if the click happened over a modeline);

• the value of toolbar-map in the current buffer (if the click happened over a toolbar);

• the current local map of the buffer under the mouse (does not apply to toolbar clicks);

• any applicable minor-mode maps;

• the current global map.

   Note that if overriding-local-map or overriding-terminal-local-map is non-nil,

only those two maps and the current global map are searched.

The procedure for searching a single keymap is called key lookup; see Section 20.8 [Key Lookup], page 324.

Since every buffer that uses the same major mode normally uses the same local keymap, you can think of the keymap as local to the mode. A change to the local keymap of a buffer (using local-set-key, for example) is seen also in the other buffers that share that keymap.

The local keymaps that are used for Lisp mode, C mode, and several other major modes exist even if they have not yet been used. These local maps are the values of the variables lisp-mode-map, c-mode-map, and so on. For most other modes, which are less frequently used, the local keymap is constructed only when the mode is used for the first time in a session.

The minibuffer has local keymaps, too; they contain various completion and exit commands. See Section 18.1 [Intro to Minibuffers], page 261.

See Appendix E [Standard Keymaps], page 781, for a list of standard keymaps.

Function current-keymaps &optional event-or-keys This function returns a list of the current keymaps that will be searched for bindings. This lists keymaps such as the current local map and the minor-mode maps, but does not list the parents of those keymaps. event-or-keys controls which keymaps will be listed. If event-or-keys is a mouse event (or a vector whose last element is a mouse event), the keymaps for that mouse event will be listed. Otherwise, the keymaps for key presses will be listed.

Variable global-map

This variable contains the default global keymap that maps XEmacs keyboard input to commands. The global keymap is normally this keymap. The default global keymap is a full keymap that binds self-insert-command to all of the printing characters. It is normal practice to change the bindings in the global map, but you should not assign this variable any value other than the keymap it starts out with.

Function current-global-map

This function returns the current global keymap. This is the same as the value of global-map unless you change one or the other.

     (current-global-map)
     ⇒ #<keymap global-map 639 entries 0x221>

Function current-local-map

This function returns the current buffer’s local keymap, or nil if it has none. In the following example, the keymap for the ‘*scratch*’ buffer (using Lisp Interaction mode) has a number of entries, including one prefix key, C-x.

     (current-local-map)
     ⇒ #<keymap lisp-interaction-mode-map 5 entries 0x558>
     (describe-bindings-internal (current-local-map))

⇒ ; Inserted into the buffer:

     backspace backward-delete-char-untabify
     linefeed eval-print-last-sexp
     delete delete-char
     C-j eval-print-last-sexp
     C-x << Prefix Command >>
     M-tab lisp-complete-symbol
     M-; lisp-indent-for-comment
     M-C-i lisp-complete-symbol
     M-C-q indent-sexp
     M-C-x eval-defun
     Alt-backspace backward-kill-sexp
     Alt-delete kill-sexp
     C-x x edebug-defun

Function current-minor-mode-maps

This function returns a list of the keymaps of currently enabled minor modes.

Function use-global-map keymap

This function makes keymap the new current global keymap. It returns nil. It is very unusual to change the global keymap.

Function use-local-map keymap &optional buffer This function makes keymap the new local keymap of buffer. buffer defaults to the current buffer. If keymap is nil, then the buffer has no local keymap. use-local-map returns nil. Most major mode commands use this function.

Variable minor-mode-map-alist

This variable is an alist describing keymaps that may or may not be active according to the values of certain variables. Its elements look like this:

(variable . keymap) The keymap keymap is active whenever variable has a non-nil value. Typically variable is the variable that enables or disables a minor mode. See Section 26.2.2 [Keymaps and Minor Modes], page 372. Note that elements of minor-mode-map-alist do not have the same structure as elements of minor-mode-alist. The map must be the cdr of the element; a list with the map as the second element will not do. What’s more, the keymap itself must appear in the cdr. It does not work to store a variable in the cdr and make the map the value of that variable. When more than one minor mode keymap is active, their order of priority is the order of minor-mode-map-alist. But you should design minor modes so that they don’t interfere with each other. If you do this properly, the order will not matter. See also minor-mode-key-binding, above. See Section 26.2.2 [Keymaps and Minor Modes], page 372, for more information about minor modes.

Variable modeline-map

This variable holds the keymap consulted for mouse-clicks on the modeline of a window. This variable may be buffer-local; its value will be looked up in the buffer of the window whose modeline was clicked upon.

Variable toolbar-map

This variable holds the keymap consulted for mouse-clicks over a toolbar.

Variable mouse-grabbed-buffer

If non-nil, a buffer which should be consulted first for all mouse activity. When a mouse-click is processed, it will first be looked up in the local-map of this buffer, and then through the normal mechanism if there is no binding for that click. This buffer’s value of mode-motion-hook will be consulted instead of the mode-motion-hook of the buffer of the window under the mouse. You should bind this, not set it.

Variable overriding-local-map

If non-nil, this variable holds a keymap to use instead of the buffer’s local keymap and instead of all the minor mode keymaps. This keymap, if any, overrides all other maps that would have been active, except for the current global map.

Variable overriding-terminal-local-map

If non-nil, this variable holds a keymap to use instead of the buffer’s local keymap and instead of all the minor mode keymaps, but for the selected console only. (In other words, this variable is always console-local; putting a keymap here only applies to keystrokes coming from the selected console. See Chapter 33 [Consoles and Devices], page 483.) This keymap, if any, overrides all other maps that would have been active, except for the current global map.

20.8 Key Lookup

Key lookup is the process of finding the binding of a key sequence from a given keymap. Actual execution of the binding is not part of key lookup.

Key lookup uses just the event type of each event in the key sequence; the rest of the event is ignored. In fact, a key sequence used for key lookup may designate mouse events with just their types (symbols) instead of with entire mouse events (lists). See Section 19.5 [Events], page 290. Such a pseudo-key-sequence is insufficient for command-execute, but it is sufficient for looking up or rebinding a key.

When the key sequence consists of multiple events, key lookup processes the events sequentially: the binding of the first event is found, and must be a keymap; then the second event’s binding is found in that keymap, and so on until all the events in the key sequence are used up. (The binding thus found for the last event may or may not be a keymap.) Thus, the process of key lookup is defined in terms of a simpler process for looking up a single event in a keymap. How that is done depends on the type of object associated with the event in that keymap.

Let’s use the term keymap entry to describe the value found by looking up an event type in a keymap. (This doesn’t include the item string and other extra elements in menu key bindings because lookup-key and other key lookup functions don’t include them in the returned value.) While any Lisp object may be stored in a keymap as a keymap entry, not all make sense for key lookup. Here is a list of the meaningful kinds of keymap entries:

nil nil means that the events used so far in the lookup form an undefined key. When a keymap fails to mention an event type at all, and has no default binding, that is equivalent to a binding of nil for that event type.

keymap The events used so far in the lookup form a prefix key. The next event of the key sequence is looked up in keymap.

command The events used so far in the lookup form a complete key, and command is its binding. See Section 11.1 [What Is a Function], page 163.

array The array (either a string or a vector) is a keyboard macro. The events used so far in the lookup form a complete key, and the array is its binding. See Section 19.13 [Keyboard Macros], page 313, for more information. (Note that you cannot use a shortened form of a key sequence here, such as (control y); you must use the full form [(control y)]. See Section 20.5 [Key Sequences], page 317.)

list The meaning of a list depends on the types of the elements of the list.

• If the car of list is lambda, then the list is a lambda expression. This is presumed to be a command, and is treated as such (see above).

• If the car of list is a keymap and the cdr is an event type, then this is an

indirect entry:

(othermap . othertype) When key lookup encounters an indirect entry, it looks up instead the binding of othertype in othermap and uses that.

This feature permits you to define one key as an alias for another key. For example, an entry whose car is the keymap called esc-map and whose cdr is 32 (the code for^{⟨}\underline{\overline{SPC}}^{⟩}) means, “Use the global binding of Meta-^{⟨}\underline{\overline{SPC}}^{⟩}, whatever that may be.”

symbol The function definition of symbol is used in place of symbol. If that too is a symbol, then this process is repeated, any number of times. Ultimately this should lead to an object that is a keymap, a command or a keyboard macro. A list is allowed if it is a keymap or a command, but indirect entries are not understood when found via symbols. Note that keymaps and keyboard macros (strings and vectors) are not valid functions, so a symbol with a keymap, string, or vector as its function definition is invalid as a function. It is, however, valid as a key binding. If the definition is a keyboard macro, then the symbol is also valid as an argument to commandexecute (see Section 19.3 [Interactive Call], page 286). The symbol undefined is worth special mention: it means to treat the key as undefined. Strictly speaking, the key is defined, and its binding is the command undefined; but that command does the same thing that is done automatically for an undefined key: it rings the bell (by calling ding) but does not signal an error. undefined is used in local keymaps to override a global key binding and make the key “undefined” locally. A local binding of nil would fail to do this because it would not override the global binding.

anything else If any other type of object is found, the events used so far in the lookup form a complete key, and the object is its binding, but the binding is not executable as a command.

In short, a keymap entry may be a keymap, a command, a keyboard macro, a symbol that leads to one of them, or an indirection or nil.

20.9 Functions for Key Lookup

Here are the functions and variables pertaining to key lookup.

Function lookup-key keymap key &optional accept-defaults This function returns the definition of key in keymap. If the string or vector key is not a valid key sequence according to the prefix keys specified in keymap (which means it is “too long” and has extra events at the end), then the value is a number, the number of events at the front of key that compose a complete key. If accept-defaults is non-nil, then lookup-key considers default bindings as well as bindings for the specific events in key. Otherwise, lookup-key reports only bindings for the specific sequence key, ignoring default bindings except when you explicitly ask about them. All the other functions described in this chapter that look up keys use lookup-

key.

     (lookup-key (current-global-map) "\C-x\C-f")
         ⇒ find-file
     (lookup-key (current-global-map) "\C-x\C-f12345")
         ⇒ 2

If key begins with the character whose value is contained in meta-prefix-char, that character is implicitly removed and the ^{⟨}\underline{\overline{META}}^{⟩} modifier added to the key. Thus, the first example below is handled by conversion into the second example.

     (lookup-key (current-global-map) "\ef")
         ⇒ forward-word
     (lookup-key (current-global-map) "\M-f")
         ⇒ forward-word

Unlike read-key-sequence, this function does not modify the specified events in ways that discard information (see Section 19.6.1 [Key Sequence Input], page 301). In particular, it does not convert letters to lower case.

Command undefined

Used in keymaps to undefine keys. If a key sequence is defined to this, invoking this key sequence causes a “key undefined” error, just as if the key sequence had no binding.

Function key-binding key &optional accept-defaults This function returns the binding for key in the current keymaps, trying all the active keymaps. The result is nil if key is undefined in the keymaps. The argument accept-defaults controls checking for default bindings, as in

lookup-key (above).
     (key-binding "\C-x\C-f")
         ⇒ find-file
     (key-binding '(control home))
         ⇒ beginning-of-buffer
     (key-binding [escape escape escape])
         ⇒ keyboard-escape-quit

Function local-key-binding key &optional accept-defaults This function returns the binding for key in the current local keymap, or nil if it is undefined there. The argument accept-defaults controls checking for default bindings, as in

lookup-key (above).

Function global-key-binding key &optional accept-defaults This function returns the binding for command key in the current global keymap, or nil if it is undefined there. The argument accept-defaults controls checking for default bindings, as in

lookup-key (above).

Function minor-mode-key-binding key &optional accept-defaults This function returns a list of all the active minor mode bindings of key. More precisely, it returns an alist of pairs (modename . binding), where modename is the variable that enables the minor mode, and binding is key’s binding in that mode. If key has no minor-mode bindings, the value is nil. If the first binding is not a prefix command, all subsequent bindings from other minor modes are omitted, since they would be completely shadowed. Similarly, the list omits non-prefix bindings that follow prefix bindings. The argument accept-defaults controls checking for default bindings, as in

lookup-key (above).

Variable meta-prefix-char

This variable is the meta-prefix character code. It is used when translating a two-character sequence to a meta character so it can be looked up in a keymap. For useful results, the value should be a prefix event (see Section 20.6 [Prefix Keys], page 319). The default value is ?\^[ (integer 27), which is the ASCII character usually produced by the ^{⟨}\underline{\overline{ESC}}^{⟩} key. As long as the value of meta-prefix-char remains ?\^[, key lookup translates ⟨\underline{\overline{ESC}}^{⟩} b into M-b, which is normally defined as the backward-word command. However, if you set meta-prefix-char to ?\^X (i.e. the keystroke C-x) or its equivalent ASCII code 24, then XEmacs will translate C-x b (whose standard binding is the switch-to-buffer command) into M-b.

     meta-prefix-char                    ; The default value.

⇒ ?\^[ ; Under XEmacs 20. ⇒ 27 ; Under XEmacs 19.

     (key-binding "\eb")
          ⇒ backward-word

?\C-x ; The print representation

; of a character.

⇒ ?\^X ; Under XEmacs 20. ⇒ 24 ; Under XEmacs 19.

     (setq meta-prefix-char 24)
          ⇒ 24
     (key-binding "\C-xb")
          ⇒ backward-word            ; Now, typing C-x b is

; like typing M-b.

     (setq meta-prefix-char ?\e)          ; Avoid confusion!

; Restore the default value!

⇒ ?\^[ ; Under XEmacs 20. ⇒ 27 ; Under XEmacs 19.

20.10 Changing Key Bindings

The way to rebind a key is to change its entry in a keymap. If you change a binding in the global keymap, the change is effective in all buffers (though it has no direct effect in buffers that shadow the global binding with a local one). If you change the current buffer’s local map, that usually affects all buffers using the same major mode. The globalset-key and local-set-key functions are convenient interfaces for these operations (see Section 20.11 [Key Binding Commands], page 330). You can also use define-key, a more general function; then you must specify explicitly the map to change.

The way to specify the key sequence that you want to rebind is described above (see Section 20.5 [Key Sequences], page 317).

For the functions below, an error is signaled if keymap is not a keymap or if key is not a string or vector representing a key sequence. You can use event types (symbols) as shorthand for events that are lists.

Function define-key keymap key binding This function sets the binding for key in keymap. (If key is more than one event long, the change is actually made in another keymap reached from keymap.) The argument binding can be any Lisp object, but only certain types are meaningful. (For a list of meaningful types, see Section 20.8 [Key Lookup], page 324.) The value returned by define-key is binding.

Every prefix of key must be a prefix key (i.e., bound to a keymap) or undefined; otherwise an error is signaled.

If some prefix of key is undefined, then define-key defines it as a prefix key so that the rest of key may be defined as specified.

Here is an example that creates a sparse keymap and makes a number of bindings in it:

(setq map (make-sparse-keymap))
    ⇒ #<keymap 0 entries 0xbee>
(define-key map "\C-f" 'forward-char)
    ⇒ forward-char
map
    ⇒ #<keymap 1 entry 0xbee>
(describe-bindings-internal map)

⇒ ; (Inserted in buffer)

C-f             forward-char

;; Build sparse submap for C-x and bind f in that.

(define-key map "\C-xf" 'forward-word)
    ⇒ forward-word
map
    ⇒ #<keymap 2 entries 0xbee>
(describe-bindings-internal map)

⇒ ; (Inserted in buffer)

C-f             forward-char
C-x             << Prefix Command >>
C-x f           forward-word
;; Bind C-p to the ctl-x-map.
(define-key map "\C-p" ctl-x-map)
;; ctl-x-map
⇒ #<keymap Control-X-prefix 77 entries 0x3bf>

;; Bind C-f to foo in the ctl-x-map.
(define-key map "\C-p\C-f" 'foo)
⇒ foo
map
    ⇒ #<keymap 3 entries 0xbee>
(describe-bindings-internal map)

⇒ ; (Inserted in buffer)

C-f             forward-char
C-p             << Prefix command Control-X-prefix >>
C-x             << Prefix Command >>
C-p tab         indent-rigidly
C-p $           set-selective-display
C-p '           expand-abbrev
C-p (           start-kbd-macro
C-p )           end-kbd-macro
   ...
C-p C-x         exchange-point-and-mark
C-p C-z         suspend-or-iconify-emacs
C-p M-escape    repeat-complex-command
C-p M-C-[       repeat-complex-command
C-x f           forward-word
C-p 4 .         find-tag-other-window
   ...
C-p 4 C-o       display-buffer
C-p 5 0         delete-frame
   ...
C-p 5 C-f       find-file-other-frame
   ...
C-p a i g       inverse-add-global-abbrev
C-p a i l       inverse-add-mode-abbrev

Note that storing a new binding for C-p C-f actually works by changing an entry in ctlx-map, and this has the effect of changing the bindings of both C-p C-f and C-x C-f in the default global map.

Function substitute-key-definition olddef newdef keymap &optional

oldmap This function replaces olddef with newdef for any keys in keymap that were bound to olddef. In other words, olddef is replaced with newdef wherever it appears. The function returns nil.

For example, this redefines C-x C-f, if you do it in an XEmacs with standard bindings:

     (substitute-key-definition
      'find-file 'find-file-read-only (current-global-map))

If oldmap is non-nil, then its bindings determine which keys to rebind. The rebindings still happen in newmap, not in oldmap. Thus, you can change one map under the control of the bindings in another. For example,

     (substitute-key-definition
       'delete-backward-char 'my-funny-delete
       my-map global-map)

puts the special deletion command in my-map for whichever keys are globally bound to the standard deletion command.

Function suppress-keymap keymap &optional nodigits This function changes the contents of the full keymap keymap by making all the printing characters undefined. More precisely, it binds them to the command undefined. This makes ordinary insertion of text impossible. suppress-

keymap returns nil.

If nodigits is nil, then suppress-keymap defines digits to run digit-argument, and - to run negative-argument. Otherwise it makes them undefined like the rest of the printing characters. The suppress-keymap function does not make it impossible to modify a buffer, as it does not suppress commands such as yank and quoted-insert. To prevent any modification of a buffer, make it read-only (see Section 30.7 [Read Only Buffers], page 438). Since this function modifies keymap, you would normally use it on a newly created keymap. Operating on an existing keymap that is used for some other purpose is likely to cause trouble; for example, suppressing global-map would make it impossible to use most of XEmacs. Most often, suppress-keymap is used to initialize local keymaps of modes such as Rmail and Dired where insertion of text is not desirable and the buffer is read-only. Here is an example taken from the file ‘emacs/lisp/dired.el’, showing how the local keymap for Dired mode is set up:

       ...
       (setq dired-mode-map (make-keymap))
       (suppress-keymap dired-mode-map)
       (define-key dired-mode-map "r" 'dired-rename-file)
       (define-key dired-mode-map "\C-d" 'dired-flag-file-deleted)
       (define-key dired-mode-map "d" 'dired-flag-file-deleted)
       (define-key dired-mode-map "v" 'dired-view-file)
       (define-key dired-mode-map "e" 'dired-find-file)
       (define-key dired-mode-map "f" 'dired-find-file)
       ...

20.11 Commands for Binding Keys

This section describes some convenient interactive interfaces for changing key bindings. They work by calling define-key.

People often use global-set-key in their ‘.emacs’ file for simple customization. For example,

(global-set-key "\C-x\C-\\" 'next-line)

or

(global-set-key [(control ?x) (control ?\\)] 'next-line)

or

(global-set-key [?\C-x ?\C-\\] 'next-line)

redefines C-x C-\ to move down a line.

(global-set-key [(meta button1)] 'mouse-set-point)

redefines the first (leftmost) mouse button, typed with the Meta key, to set point where you click.

Command global-set-key key definition

This function sets the binding of key in the current global map to definition.

     (global-set-key key definition)

≡

     (define-key (current-global-map) key definition)

Command global-unset-key key

This function removes the binding of key from the current global map. One use of this function is in preparation for defining a longer key that uses key as a prefix—which would not be allowed if key has a non-prefix binding. For example:

     (global-unset-key "\C-l")
         ⇒ nil
     (global-set-key "\C-l\C-l" 'redraw-display)
         ⇒ nil

This function is implemented simply using define-key:

     (global-unset-key key)

≡

     (define-key (current-global-map) key nil)

Command local-set-key key definition

This function sets the binding of key in the current local keymap to definition.

     (local-set-key key definition)

≡

     (define-key (current-local-map) key definition)

Command local-unset-key key

This function removes the binding of key from the current local map.

     (local-unset-key key)

≡

     (define-key (current-local-map) key nil)

20.12 Scanning Keymaps

This section describes functions used to scan all the current keymaps, or all keys within a keymap, for the sake of printing help information.

Function accessible-keymaps keymap &optional prefix This function returns a list of all the keymaps that can be accessed (via prefix keys) from keymap. The value is an association list with elements of the form (key . map), where key is a prefix key whose definition in keymap is map.

The elements of the alist are ordered so that the key increases in length. The first element is always ([] . keymap), because the specified keymap is accessible from itself with a prefix of no events.

If prefix is given, it should be a prefix key sequence; then accessible-keymaps includes only the submaps whose prefixes start with prefix. These elements look just as they do in the value of (accessible-keymaps); the only difference is that some elements are omitted.

In the example below, the returned alist indicates that the key C-x, which is displayed as ‘[(control x)]’, is a prefix key whose definition is the keymap

#<keymap ((control x) #<keymap emacs-lisp-mode-map 8 entries 0x546>)

1 entry 0x8a2>. (The strange notation for the keymap’s name indicates that this is an internal submap of emacs-lisp-mode-map. This is because lisp-

interaction-mode-map has set up emacs-lisp-mode-map as its parent, and

lisp-interaction-mode-map defines no key sequences beginning with C-x.)

     (current-local-map)
     ⇒ #<keymap lisp-interaction-mode-map 5 entries 0x558>
     (accessible-keymaps (current-local-map))
     ⇒(([] . #<keymap lisp-interaction-mode-map 5 entries 0x558>)
         ([(control x)] .
          #<keymap ((control x) #<keymap emacs-lisp-mode-map 8 entries 0x546>)
                   1 entry 0x8a2>))

The following example shows the results of calling accessible-keymaps on a large, complex keymap. Notice how some keymaps were given explicit names using set-keymap-name; those submaps without explicit names are given descriptive names indicating their relationship to their enclosing keymap.

     (accessible-keymaps (current-global-map))
     ⇒ (([] . #<keymap global-map 639 entries 0x221>)
        ([(control c)] . #<keymap mode-specific-command-prefix 1 entry 0x3cb>)
        ([(control h)] . #<keymap help-map 33 entries 0x4ec>)
        ([(control x)] . #<keymap Control-X-prefix 77 entries 0x3bf>)
        ([(meta escape)] .
           #<keymap ((meta escape) #<keymap global-map 639 entries 0x221>)
                    3 entries 0x3e0>)
        ([(meta control \[)] .
           #<keymap ((meta escape) #<keymap global-map 639 entries 0x221>)
                    3 entries 0x3e0>)
        ([f1] . #<keymap help-map 33 entries 0x4ec>)
        ([(control x) \4] . #<keymap ctl-x-4-prefix 9 entries 0x3c5>)
        ([(control x) \5] . #<keymap ctl-x-5-prefix 8 entries 0x3c8>)
        ([(control x) \6] . #<keymap 13 entries 0x4d2>)
        ([(control x) a] .
           #<keymap (a #<keymap Control-X-prefix 77 entries 0x3bf>)
                    8 entries 0x3ef>)
        ([(control x) n] . #<keymap narrowing-prefix 3 entries 0x3dd>)
        ([(control x) r] . #<keymap rectangle-prefix 18 entries 0x3e9>)
        ([(control x) v] . #<keymap vc-prefix-map 13 entries 0x60e>)
        ([(control x) a i] .
          #<keymap (i #<keymap (a #<keymap Control-X-prefix 77 entries 0x3bf>)
                               8 entries 0x3ef>)
                   2 entries 0x3f5>))

Function map-keymap function keymap &optional sort-first This function applies function to each element of KEYMAP. function will be called with two arguments: a key-description list, and the binding. The order in which the elements of the keymap are passed to the function is unspecified. If the function inserts new elements into the keymap, it may or may not be called with them later. No element of the keymap will ever be passed to the function more than once. The function will not be called on elements of this keymap’s parents (see Section 20.4 [Inheritance and Keymaps], page 317) or upon keymaps which are contained within this keymap (multi-character definitions). It will be called on ⟨\underline{\overline{META}}^{⟩} characters since they are not really two-character sequences. If the optional third argument sort-first is non-nil, then the elements of the keymap will be passed to the mapper function in a canonical order. Otherwise, they will be passed in hash (that is, random) order, which is faster.

Function keymap-fullness keymap

This function returns the number of bindings in the keymap.

Function where-is-internal definition &optional keymaps firstonly

noindirect event-or-keys This function returns a list of key sequences (of any length) that are bound to definition in a set of keymaps.

The argument definition can be any object; it is compared with all keymap entries using eq. KEYMAPS can be either a keymap (meaning search in that keymap and the current global keymap) or a list of keymaps (meaning search in exactly those keymaps and no others). If KEYMAPS is nil, search in the currently applicable maps for EVENT-OR-KEYS. If keymap is a keymap, then the maps searched are keymap and the global keymap. If keymap is a list of keymaps, then the maps searched are exactly those keymaps, and no others. If keymap is nil, then the maps used are the current active keymaps for event-or-keys (this is equivalent to specifying (current-keymaps event-or-keys) as the argument to keymaps). If firstonly is non-nil, then the value is a single vector representing the first key sequence found, rather than a list of all possible key sequences. If noindirect is non-nil, where-is-internal doesn’t follow indirect keymap bindings. This makes it possible to search for an indirect definition itself. This function is used by where-is (see section “Help” in The XEmacs Reference Manual).

     (where-is-internal 'describe-function)
         ⇒ ([(control h) d] [(control h) f] [f1 d] [f1 f])

Function describe-bindings-internal map &optional all shadow prefix

mouse-only-p This function inserts (into the current buffer) a list of all defined keys and their definitions in map. Optional second argument all says whether to include even “uninteresting” definitions, i.e. symbols with a non-nilsuppress-keymap property. Third argument shadow is a list of keymaps whose bindings shadow those of map; if a binding is present in any shadowing map, it is not printed. Fourth argument prefix, if non-nil, should be a key sequence; only bindings which start with that key sequence will be printed. Fifth argument mouse-only-p says to only print bindings for mouse clicks.

describe-bindings-internal is used to implement the help command describe-
bindings.

Command describe-bindings prefix mouse-only-p

This function creates a listing of all defined keys and their definitions. It writes the listing in a buffer named ‘*Help*’ and displays it in a window. If prefix is non-nil, it should be a prefix key; then the listing includes only keys that start with prefix. When several characters with consecutive ASCII codes have the same definition, they are shown together, as ‘firstchar..lastchar’. In this instance, you need to know the ASCII codes to understand which characters this means. For example, in the default global map, the characters ‘^{⟨}\underline{\overline{SPC}}^{⟩} .. ~’ are described by a single line. ^{⟨}\underline{\overline{SPC}}^{⟩} is ASCII 32, ~ is ASCII 126, and the characters between them include all the normal printing characters, (e.g., letters, digits, punctuation, etc.); all these characters are bound to self-insert-command.

If the second argument (prefix arg, interactively) is non-nil then only the mouse bindings are displayed.

20.13 Other Keymap Functions

Function set-keymap-prompt keymap new-prompt

This function sets the “prompt” of keymap to string new-prompt, or nil if no prompt is desired. The prompt is shown in the echo-area when reading a key-sequence to be looked-up in this keymap.

Function keymap-prompt keymap &optional use-inherited This function returns the “prompt” of the given keymap. If use-inherited is non-nil, any parent keymaps will also be searched for a prompt.

336 XEmacs Lisp Reference Manual

21 Menus

21.1 Format of Menus

A menu is described using a menu description, which is a list of menu items, keywordvalue pairs, strings, and submenus. The menu description specifies which items are present in the menu, what function each item invokes, and whether the item is selectable or not. Pop-up menus are directly described with a menu description, while menubars are described slightly differently (see below).

The first element of a menu must be a string, which is the name of the menu. This is the string that will be displayed in the parent menu or menubar, if any. This string is not displayed in the menu itself, except in the case of the top level pop-up menu, where there is no parent. In this case, the string will be displayed at the top of the menu if

popup-menu-titles is non-nil.

Immediately following the first element there may optionally be up to four keyword-value pairs, as follows:

:included form

This can be used to control the visibility of a menu. The form is evaluated and the menu will be omitted if the result is nil.

:config symbol
This is an efficient shorthand for :included (memq symbol menubar-configuration).
See the variable menubar-configuration.

:filter function A menu filter is used to sensitize or incrementally create a submenu only when it is selected by the user and not every time the menubar is activated. The filter function is passed the list of menu items in the submenu and must return a list of menu items to be used for the menu. It is called only when the menu is about to be displayed, so other menus may already be displayed. Vile and terrible things will happen if a menu filter function changes the current buffer, window, or frame. It also should not raise, lower, or iconify any frames. Basically, the filter function should have no side-effects.

:accelerator key

A menu accelerator is a keystroke which can be pressed while the menu is visible which will immediately activate the item. key must be a char or the symbol name of a key. See Section 21.7 [Menu Accelerators], page 345.

The rest of the menu consists of elements as follows:

• A menu item, which is a vector in the following form:

[ name callback :keyword value :keyword value ... ] name is a string, the name of the menu item; it is the string to display on the menu. It is filtered through the resource database, so it is possible for resources to override what string is actually displayed. callback is a form that will be invoked when the menu item is selected. If the callback of a menu item is a symbol, then it must name a command. It will be invoked with call-interactively. If it is a list, then it is evaluated with eval. The valid keywords and their meanings are described below. Note that for compatibility purposes, the form

[ name callback active-p ] is also accepted and is equivalent to

[ name callback :active active-p ] and the form

[ name callback active-p suffix] is accepted and is equivalent to

[ name callback :active active-p :suffix suffix] However, these older forms are deprecated and should generally not be used.

• If an element of a menu is a string, then that string will be presented in the menu as unselectable text.

• If an element of a menu is a string consisting solely of hyphens, then that item will be presented as a solid horizontal line.

• If an element of a menu is a string beginning with ‘--:’, then a particular sort of horizontal line will be displayed, as follows:

‘"--:singleLine"’

A solid horizontal line. This is equivalent to a string consisting solely of hyphens.

‘"--:doubleLine"’

A solid double horizontal line.

‘"--:singleDashedLine"’

A dashed horizontal line.

‘"--:doubleDashedLine"’

A dashed double horizontal line.

‘"--:noLine"’

No line (but a small space is left).

‘"--:shadowEtchedIn"’

A solid horizontal line with a 3-d recessed appearance.

‘"--:shadowEtchedOut"’

A solid horizontal line with a 3-d pushed-out appearance.

‘"--:shadowDoubleEtchedIn"’

A solid double horizontal line with a 3-d recessed appearance.

‘"--:shadowDoubleEtchedOut"’

A solid double horizontal line with a 3-d pushed-out appearance.

‘"--:shadowEtchedInDash"’

A dashed horizontal line with a 3-d recessed appearance.

‘"--:shadowEtchedOutDash"’

A dashed horizontal line with a 3-d pushed-out appearance.

‘"--:shadowDoubleEtchedInDash"’

A dashed double horizontal line with a 3-d recessed appearance.

‘"--:shadowDoubleEtchedOutDash"’

A dashed double horizontal line with a 3-d pushed-out appearance.

• If an element of a menu is a list, it is treated as a submenu. The name of that submenu (the first element in the list) will be used as the name of the item representing this menu on the parent.

The possible keywords are as follows:

:active form form will be evaluated when the menu that this item is a part of is about to be displayed, and the item will be selectable only if the result is non-nil. If the item is unselectable, it will usually be displayed grayed-out to indicate this.

:suffix form form will be evaluated when the menu that this item is a part of is about to be displayed, and the resulting string is appended to the displayed name. This provides a convenient way of adding the name of a command’s “argument” to the menu, like ‘Kill Buffer NAME’.

:keys string Normally, the keyboard equivalents of commands in menus are displayed when the “callback” is a symbol. This can be used to specify keys for more complex menu items. It is passed through substitute-command-keys first.

:style style Specifies what kind of object this menu item is. style be one of the symbols

nil A normal menu item.

toggle A toggle button.

radio A radio button.

button A menubar button.

The only difference between toggle and radio buttons is how they are displayed. But for consistency, a toggle button should be used when there is one option whose value can be turned on or off, and radio buttons should be used when there is a set of mutually exclusive options. When using a group of radio buttons, you should arrange for no more than one to be marked as selected at a time.

:selected form Meaningful only when style is toggle, radio or button. This specifies whether the button will be in the selected or unselected state. form is evaluated, as for

:active.

:included form This can be used to control the visibility of a menu item. The form is evaluated and the menu item is only displayed if the result is non-nil. Note that this is different from :active: If :active evaluates to nil, the item will be displayed grayed out, while if :included evaluates to nil, the item will be omitted entirely.

:config symbol

This is an efficient shorthand for :included (memq symbol menubar-configuration).
See the variable menubar-configuration.

:accelerator key A menu accelerator is a keystroke which can be pressed while the menu is visible which will immediately activate the item. key must be a char or the symbol name of a key. See Section 21.7 [Menu Accelerators], page 345.

Variable menubar-configuration

This variable holds a list of symbols, against which the value of the :config tag for each menubar item will be compared. If a menubar item has a :config tag, then it is omitted from the menubar if that tag is not a member of the

menubar-configuration list.

For example:

("File"
:filter file-menu-filter ; file-menu-filter is a function that takes
; one argument (a list of menu items) and
; returns a list of menu items
[ "Save As..."    write-file]
[ "Revert Buffer" revert-buffer :active (buffer-modified-p) ]
[ "Read Only"     toggle-read-only :style toggle :selected buffer-read-only ]
)

21.2 Format of the Menubar

A menubar is a list of menus, menu items, and strings. The format is similar to that of a menu, except:

• The first item need not be a string, and is not treated specially.

• A string consisting solely of hyphens is not treated specially.

• If an element of a menubar is nil, then it is used to represent the division between the set of menubar items which are flush-left and those which are flush-right. (Note: this isn’t completely implemented yet.)

21.3 Menubar

Variable current-menubar

This variable holds the description of the current menubar. This may be bufferlocal. When the menubar is changed, the function set-menubar-dirty-flag has to be called in order for the menubar to be updated on the screen.

Constant default-menubar

This variable holds the menubar description of the menubar that is visible at startup. This is the value that current-menubar has at startup.

Function set-menubar-dirty-flag

This function tells XEmacs that the menubar widget has to be updated. Changes to the menubar will generally not be visible until this function is called.

The following convenience functions are provided for setting the menubar. They are equivalent to doing the appropriate action to change current-menubar, and then calling set-menubar-dirty-flag. Note that these functions copy their argument using copy-

sequence.

Function set-menubar menubar

This function sets the default menubar to be menubar (see Section 21.1 [Menu Format], page 337). This is the menubar that will be visible in buffers that have not defined their own, buffer-local menubar.

Function set-buffer-menubar menubar

This function sets the buffer-local menubar to be menubar. This does not change the menubar in any buffers other than the current one.

Miscellaneous:

Variable menubar-show-keybindings

If true, the menubar will display keyboard equivalents. If false, only the command names will be displayed.

Variable activate-menubar-hook

Function or functions called before a menubar menu is pulled down. These functions are called with no arguments, and should interrogate and modify the value of current-menubar as desired. The functions on this hook are invoked after the mouse goes down, but before the menu is mapped, and may be used to activate, deactivate, add, or delete items from the menus. However, using a filter (with the :filter keyword in a menu description) is generally a more efficient way of accomplishing the same thing, because the filter is invoked only when the actual menu goes down. With a complex menu, there can be a quite noticeable and sometimes aggravating delay if all menu modification is implemented using the activate-menubar-hook. See above. These functions may return the symbol t to assert that they have made no changes to the menubar. If any other value is returned, the menubar is recomputed. If t is returned but the menubar has been changed, then the changes may not show up right away. Returning nil when the menubar has not changed is not so bad; more computation will be done, but redisplay of the menubar will still be performed optimally.

Variable menu-no-selection-hook

Function or functions to call when a menu or dialog box is dismissed without a selection having been made.

21.4 Modifying Menus

The following functions are provided to modify the menubar of one of its submenus. Note that these functions modify the menu in-place, rather than copying it and making a new menu.

Some of these functions take a menu path, which is a list of strings identifying the menu to be modified. For example, ("File") names the top-level “File” menu. ("File" "Foo") names a hypothetical submenu of “File”.

Others take a menu item path, which is similar to a menu path but also specifies a particular item to be modified. For example, ("File" "Save") means the menu item called “Save” under the top-level “File” menu. ("Menu" "Foo" "Item") means the menu item called “Item” under the “Foo” submenu of “Menu”.

Function add-submenu menu-path submenu &optional before This function adds a menu to the menubar or one of its submenus. If the named menu exists already, it is changed. menu-path identifies the menu under which the new menu should be inserted. If menu-path is nil, then the menu will be added to the menubar itself. submenu is the new menu to add (see Section 21.1 [Menu Format], page 337). before, if provided, is the name of a menu before which this menu should be added, if this menu is not on its parent already. If the menu is already present, it will not be moved.

Function add-menu-button menu-path menu-leaf &optional before This function adds a menu item to some menu, creating the menu first if necessary. If the named item exists already, it is changed. menu-path identifies the menu under which the new menu item should be inserted. menu-leaf is a menubar leaf node (see Section 21.1 [Menu Format], page 337). before, if provided, is the name of a menu before which this item should be added, if this item is not on the menu already. If the item is already present, it will not be moved.

Function delete-menu-item menu-item-path

This function removes the menu item specified by menu-item-path from the menu hierarchy.

Function enable-menu-item menu-item-path

This function makes the menu item specified by menu-item-path be selectable.

Function disable-menu-item menu-item-path

This function makes the menu item specified by menu-item-path be unselectable.

Function relabel-menu-item menu-item-path new-name This function changes the string of the menu item specified by menu-item-path. new-name is the string that the menu item will be printed as from now on.

The following function can be used to search for a particular item in a menubar specification, given a path to the item.

Function find-menu-item menubar menu-item-path &optional parent This function searches menubar for the item given by menu-item-path starting from parent (nil means start at the top of menubar). This function returns (item . parent), where parent is the immediate parent of the item found (a menu description), and item is either a vector, list, or string, depending on the nature of the menu item.

This function signals an error if the item is not found.

The following deprecated functions are also documented, so that existing code can be understood. You should not use these functions in new code.

Function add-menu menu-path menu-name menu-items &optional before This function adds a menu to the menubar or one of its submenus. If the named menu exists already, it is changed. This is obsolete; use add-submenu instead.

menu-path identifies the menu under which the new menu should be inserted. If menu-path is nil, then the menu will be added to the menubar itself.

menu-name is the string naming the menu to be added; menu-items is a list of menu items, strings, and submenus. These two arguments are the same as the first and following elements of a menu description (see Section 21.1 [Menu Format], page 337).

before, if provided, is the name of a menu before which this menu should be added, if this menu is not on its parent already. If the menu is already present, it will not be moved.

Function add-menu-item menu-path item-name function enabled-p

&optional before This function adds a menu item to some menu, creating the menu first if necessary. If the named item exists already, it is changed. This is obsolete; use

add-menu-button instead.

menu-path identifies the menu under which the new menu item should be inserted. item-name, function, and enabled-p are the first, second, and third elements of a menu item vector (see Section 21.1 [Menu Format], page 337).

before, if provided, is the name of a menu item before which this item should be added, if this item is not on the menu already. If the item is already present, it will not be moved.

21.5 Menu Filters

The following filter functions are provided for use in default-menubar. You may want to use them in your own menubar description.

Function file-menu-filter menu-items

This function changes the arguments and sensitivity of these File menu items:

‘Delete Buffer’

Has the name of the current buffer appended to it.

‘Print Buffer’

Has the name of the current buffer appended to it.

‘Pretty-Print Buffer’

Has the name of the current buffer appended to it.

‘Save Buffer’

Has the name of the current buffer appended to it, and is sensitive only when the current buffer is modified.

‘Revert Buffer’

Has the name of the current buffer appended to it, and is sensitive only when the current buffer has a file.

‘Delete Frame’

Sensitive only when there is more than one visible frame.

Function edit-menu-filter menu-items

This function changes the arguments and sensitivity of these Edit menu items:

‘Cut’ Sensitive only when XEmacs owns the primary X Selection (if zmacs-regions is t, this is equivalent to saying that there is a region selected).

‘Copy’ Sensitive only when XEmacs owns the primary X Selection.

‘Clear’ Sensitive only when XEmacs owns the primary X Selection.

‘Paste’ Sensitive only when there is an owner for the X Clipboard Selection.

‘Undo’ Sensitive only when there is undo information. While in the midst of an undo, this is changed to ‘Undo More’.

Function buffers-menu-filter menu-items

This function sets up the Buffers menu. See Section 21.8 [Buffers Menu], page 347, for more information.

21.6 Pop-Up Menus

Function popup-menu menu-desc

This function pops up a menu specified by menu-desc, which is a menu description (see Section 21.1 [Menu Format], page 337). The menu is displayed at the current mouse position.

Function popup-menu-up-p

This function returns t if a pop-up menu is up, nil otherwise.

Variable popup-menu-titles

If true (the default), pop-up menus will have title bars at the top.

Some machinery is provided that attempts to provide a higher-level mechanism onto pop-up menus. This only works if you do not redefine the binding for button3.

Command popup-mode-menu

This function pops up a menu of global and mode-specific commands. The

menu is computed by combining global-popup-menu and mode-popup-menu.

This is the default binding for button3. You should generally not change this binding.

Variable global-popup-menu

This holds the global popup menu. This is present in all modes. (This is nil by default.)

Variable mode-popup-menu

The mode-specific popup menu. Automatically buffer local. This is appended to the default items in global-popup-menu.

Constant default-popup-menu

This holds the default value of mode-popup-menu.

Variable activate-popup-menu-hook

Function or functions run before a mode-specific popup menu is made visible. These functions are called with no arguments, and should interrogate and modify the value of global-popup-menu or mode-popup-menu as desired. Note: this hook is only run if you use popup-mode-menu for activating the global and mode-specific commands; if you have your own binding for button3, this hook won’t be run.

The following convenience functions are provided for displaying pop-up menus.

Function popup-buffer-menu event

This function pops up a copy of the ‘Buffers’ menu (from the menubar) where the mouse is clicked.

Function popup-menubar-menu event

This function pops up a copy of menu that also appears in the menubar.

21.7 Menu Accelerators

Menu accelerators are keyboard shortcuts for accessing the menubar. Accelerator keys can be specified for menus as well as for menu items. An accelerator key for a menu is used to activate that menu when it appears as a submenu of another menu. An accelerator key for a menu item is used to activate that item.

21.7.1 Creating Menu Accelerators

Menu accelerators are specified as part of the menubar format using the :accelerator tag to specify a key or by placing "% " in the menu or menu item name prior to the letter which is to be used as the accelerator key. The advantage of the second method is that the menu rendering code then knows to draw an underline under that character, which is the canonical way of indicating an accelerator key to a user.

For example, the command

     (add-submenu nil '("%_Test"
        ["One" (insert "1") :accelerator ?1 :active t]
        ["%_Two" (insert "2")]
        ["%_3" (insert "3")]))

will add a new menu to the top level menubar. The new menu can be reached by pressing "t" while the top level menubar is active. When the menu is active, pressing "1" will activate the first item and insert the character "1" into the buffer. Pressing "2" will activate the second item and insert the character "2" into the buffer. Pressing "3" will activate the third item and insert the character "3" into the buffer.

It is possible to activate the top level menubar itself using accelerator keys. See Section 21.7.3 [Menu Accelerator Functions], page 346.

21.7.2 Keyboard Menu Traversal

In addition to immediately activating a menu or menu item, the keyboard can be used to traverse the menus without activating items. The keyboard arrow keys, the return key and the escape key are defined to traverse the menus in a way that should be familiar to users of any of a certain family of popular PC operating systems.

This behavior can be changed by modifying the bindings in menu-accelerator-map. At this point, the online help is your best bet for more information about how to modify the menu traversal keys.

21.7.3 Menu Accelerator Functions

Function accelerate-menu

Make the menubar immediately active and place the cursor on the left most entry in the top level menu. Menu items can be selected as usual.

Variable menu-accelerator-enabled

Whether menu accelerator keys can cause the menubar to become active.

If menu-force or menu-fallback, then menu accelerator keys can be used to activate the top level menu. Once the menubar becomes active, the accelerator keys can be used regardless of the value of this variable.

menu-force is used to indicate that the menu accelerator key takes precedence over bindings in the current keymap(s). menu-fallback means that bindings in the current keymap take precedence over menu accelerator keys. Thus a top level menu with an accelerator of "T" would be activated on a keypress of Metat if menu-accelerator-enabled is menu-force. However, if menu-accelerator-enabled is menu-fallback, then Meta-t will not activate the menubar and will instead run the function transpose-words, to which it is normally bound. The default value is nil. See also menu-accelerator-modifiers and menu-accelerator-prefix.

Variable menu-accelerator-map

Keymap consulted to determine the commands to run in response to keypresses occurring while the menubar is active. See Section 21.7.2 [Keyboard Menu Traversal], page 346.

Variable menu-accelerator-modifiers

A list of modifier keys which must be pressed in addition to a valid menu accelerator in order for the top level menu to be activated in response to a keystroke. The default value of (meta) mirrors the usage of the alt key as a menu accelerator in popular PC operating systems. The modifier keys in menu-accelerator-modifiers must match exactly the modifiers present in the keypress. The only exception is that the shift modifier is accepted in conjunction with alphabetic keys even if it is not a menu accelerator modifier. See also menu-accelerator-enabled and menu-accelerator-prefix.

Variable menu-accelerator-prefix

Prefix key(s) that must be typed before menu accelerators will be activated. Must be a valid key descriptor. The default value is nil.

(setq menu-accelerator-prefix ?\C-x)
(setq menu-accelerator-modifiers '(meta control))
(setq menu-accelerator-enabled 'menu-force)
(add-submenu nil '("%_Test"
   ["One" (insert "1") :accelerator ?1 :active t]
   ["%_Two" (insert "2")]
   ["%_3" (insert "3")]))

will add the menu "Test" to the top level menubar. Pressing C-x followed by C-M-T will activate the menubar and display the "Test" menu. Pressing C-M-T by itself will not activate the menubar. Neither will pressing C-x followed by anything else.

21.8 Buffers Menu

The following options control how the ‘Buffers’ menu is displayed. This is a list of all (or a subset of) the buffers currently in existence, and is updated dynamically.

User Option buffers-menu-max-size

This user option holds the maximum number of entries which may appear on the ‘Buffers’ menu. If this is 10, then only the ten most-recently-selected buffers will be shown. If this is nil, then all buffers will be shown. Setting this to a large number or nil will slow down menu responsiveness.

Function format-buffers-menu-line buffer

This function returns a string to represent buffer in the ‘Buffers’ menu. nil means the buffer shouldn’t be listed. You can redefine this.

User Option complex-buffers-menu-p

If true, the ‘Buffers’ menu will contain several commands, as submenus of each buffer line. If this is false, then there will be only one command: select that buffer.

User Option buffers-menu-switch-to-buffer-function

This user option holds the function to call to select a buffer from the ‘Buffers’

menu. switch-to-buffer is a good choice, as is pop-to-buffer.

22 Dialog Boxes

22.1 Dialog Box Format

A dialog box description is a list.

• The first element of the list is a string to display in the dialog box.

• The rest of the elements are descriptions of the dialog box’s buttons. Each one is a vector of three elements: − The first element is the text of the button. − The second element is the callback. − The third element is t or nil, whether this button is selectable.

If the callback of a button is a symbol, then it must name a command. It will be invoked with call-interactively. If it is a list, then it is evaluated with eval.

One (and only one) of the buttons may be nil. This marker means that all following buttons should be flushright instead of flushleft.

The syntax, more precisely:

form :=  <something to pass to `eval'>
command :=  <a symbol or string, to pass to `call-interactively'>
callback  :=  command | form
active-p :=  <t, nil, or a form to evaluate to decide whether this
 button should be selectable>
name :=  <string>
partition :=  'nil'
button :=  '['  name callback active-p ']'
dialog :=  '(' name [ button ]+ [ partition [ button ]+ ] ')'

22.2 Dialog Box Functions

Function popup-dialog-box dbox-desc

This function pops up a dialog box. dbox-desc describes how the dialog box will appear (see Section 22.1 [Dialog Box Format], page 349).

See Section 18.6 [Yes-or-No Queries], page 275, for functions to ask a yes/no question using a dialog box.

350 XEmacs Lisp Reference Manual

23 Toolbar

23.1 Toolbar Intro

A toolbar is a bar of icons displayed along one edge of a frame. You can view a toolbar as a series of menu shortcuts – the most common menu options can be accessed with a single click rather than a series of clicks and/or drags to select the option from a menu. Consistent with this, a help string (called the help-echo) describing what an icon in the toolbar (called a toolbar button) does, is displayed in the minibuffer when the mouse is over the button.

In XEmacs, a toolbar can be displayed along any of the four edges of the frame, and two or more different edges can be displaying toolbars simultaneously. The contents, thickness, and visibility of the toolbars can be controlled separately, and the values can be per-buffer, per-frame, etc., using specifiers (see Chapter 41 [Specifiers], page 599).

Normally, there is one toolbar displayed in a frame. Usually, this is the standard toolbar, but certain modes will override this and substitute their own toolbar. In some cases (e.g. the VM package), a package will supply its own toolbar along a different edge from the standard toolbar, so that both can be visible at once. This standard toolbar is usually positioned along the top of the frame, but this can be changed using set-default-toolbar-position.

Note that, for each of the toolbar properties (contents, thickness, and visibility), there is a separate specifier for each of the four toolbar positions (top, bottom, left, and right), and an additional specifier for the “default” toolbar, i.e. the toolbar whose position is controlled

by set-default-toolbar-position. The way this works is that set-default-toolbar-

position arranges things so that the appropriate position-specific specifiers for the default position inherit from the corresponding default specifiers. That way, if the position-specific specifier does not give a value (which it usually doesn’t), then the value from the default specifier applies. If you want to control the default toolbar, you just change the default specifiers, and everything works. A package such as VM that wants to put its own toolbar in a different location from the default just sets the position-specific specifiers, and if the user sets the default toolbar to the same position, it will just not be visible.

23.2 Toolbar Descriptor Format

The contents of a toolbar are specified using a toolbar descriptor. The format of a toolbar descriptor is a list of toolbar button descriptors. Each toolbar button descriptor is a vector in one of the following formats:

• [glyph-list function enabled-p help]

• [:style 2d-or-3d]

• [:style 2d-or-3d :size width-or-height]

• [:size width-or-height :style 2d-or-3d]

Optionally, one of the toolbar button descriptors may be nil instead of a vector; this signifies the division between the toolbar buttons that are to be displayed flush-left, and the buttons to be displayed flush-right.

The first vector format above specifies a normal toolbar button; the others specify blank areas in the toolbar.

For the first vector format:

• glyph-list should be a list of one to six glyphs (as created by make-glyph) or a symbol whose value is such a list. The first glyph, which must be provided, is the glyph used to display the toolbar button when it is in the “up” (not pressed) state. The optional second glyph is for displaying the button when it is in the “down” (pressed) state. The optional third glyph is for when the button is disabled. The last three glyphs are for displaying the button in the “up”, “down”, and “disabled” states, respectively, but are used when the user has called for captioned toolbar buttons (using toolbar-buttons-captioned-p). The function toolbar-make-button-list is useful in creating these glyph lists.

• Even if you do not provide separate down-state and disabled-state glyphs, the user will still get visual feedback to indicate which state the button is in. Buttons in the up-state are displayed with a shadowed border that gives a raised appearance to the button. Buttons in the down-state are displayed with shadows that give a recessed appearance. Buttons in the disabled state are displayed with no shadows, giving a 2-d effect.

• If some of the toolbar glyphs are not provided, they inherit as follows:

          UP:                up
          DOWN:              down -> up
          DISABLED:          disabled -> up
          CAP-UP:            cap-up -> up
          CAP-DOWN:          cap-down -> cap-up -> down -> up
          CAP-DISABLED:      cap-disabled -> cap-up -> disabled -> up

• The second element function is a function to be called when the toolbar button is activated (i.e. when the mouse is released over the toolbar button, if the press occurred in the toolbar). It can be any form accepted by call-interactively, since this is how it is invoked.

• The third element enabled-p specifies whether the toolbar button is enabled (disabled buttons do nothing when they are activated, and are displayed differently; see above). It should be either a boolean or a form that evaluates to a boolean.

• The fourth element help, if non-nil, should be a string. This string is displayed in the echo area when the mouse passes over the toolbar button.

For the other vector formats (specifying blank areas of the toolbar):

• 2d-or-3d should be one of the symbols 2d or 3d, indicating whether the area is displayed with shadows (giving it a raised, 3-d appearance) or without shadows (giving it a flat appearance).

• width-or-height specifies the length, in pixels, of the blank area. If omitted, it defaults to a device-specific value (8 pixels for X devices).

Function toolbar-make-button-list up &optional down disabled cap-up

cap-down cap-disabled This function calls make-glyph on each arg and returns a list of the results. This is useful for setting the first argument of a toolbar button descriptor (typically, the result of this function is assigned to a symbol, which is specified as the first argument of the toolbar button descriptor).

Function check-toolbar-button-syntax button &optional noerror

Verify the syntax of entry button in a toolbar description list. If you want to verify the syntax of a toolbar description list as a whole, use check-valid-instantiator with a specifier type of toolbar.

23.3 Specifying the Toolbar

In order to specify the contents of a toolbar, set one of the specifier variables default-

toolbar, top-toolbar, bottom-toolbar, left-toolbar, or right-toolbar. These are

specifiers, which means you set them with set-specifier and query them with specifierspecs or specifier-instance. You will get an error if you try to set them using setq. The valid instantiators for these specifiers are toolbar descriptors, as described above. See Chapter 41 [Specifiers], page 599, for more information.

Most of the time, you will set default-toolbar, which allows the user to choose where the toolbar should go.

Specifier default-toolbar

The position of this toolbar is specified in the function default-toolbar-position. If the corresponding position-specific toolbar (e.g. top-toolbar if default-toolbar-position is top) does not specify a toolbar in a particular domain, then the value of default-toolbar in that domain, of any, will be used instead.

Note that the toolbar at any particular position will not be displayed unless its thickness (width or height, depending on orientation) is non-zero and its visibility status is true. The

thickness is controlled by the specifiers top-toolbar-height, bottom-toolbar-height,

left-toolbar-width, and right-toolbar-width, and the visibility status is controlled

by the specifiers top-toolbar-visible-p, bottom-toolbar-visible-p, left-toolbar-

visible-p, and right-toolbar-visible-p (see Section 23.4 [Other Toolbar Variables], page 354).

Function set-default-toolbar-position position

This function sets the position that the default-toolbar will be displayed at. Valid positions are the symbols top, bottom, left and right. What this actually does is set the fallback specifier for the position-specific specifier corresponding to the given position to default-toolbar, and set the fallbacks for the other position-specific specifiers to nil. It also does the same thing for the position-specific thickness and visibility specifiers, which inherit from one

of default-toolbar-height or default-toolbar-width, and from default-

toolbar-visible-p, respectively (see Section 23.4 [Other Toolbar Variables], page 354).

Function default-toolbar-position

This function returns the position that the default-toolbar will be displayed at.

You can also explicitly set a toolbar at a particular position. When redisplay determines what to display at a particular position in a particular domain (i.e. window), it first consults the position-specific toolbar. If that does not yield a toolbar descriptor, the defaulttoolbar is consulted if default-toolbar-position indicates this position.

Specifier top-toolbar

Specifier for the toolbar at the top of the frame.

Specifier bottom-toolbar

Specifier for the toolbar at the bottom of the frame.

Specifier left-toolbar

Specifier for the toolbar at the left edge of the frame.

Specifier right-toolbar

Specifier for the toolbar at the right edge of the frame.

Function toolbar-specifier-p object

This function returns non-nil if object is a toolbar specifier. Toolbar specifiers are the actual objects contained in the toolbar variables described above, and their valid instantiators are toolbar descriptors (see Section 23.2 [Toolbar Descriptor Format], page 351).

23.4 Other Toolbar Variables

The variables to control the toolbar thickness, visibility status, and captioned status are all specifiers. See Chapter 41 [Specifiers], page 599.

Specifier default-toolbar-height

This specifies the height of the default toolbar, if it’s oriented horizontally. The position of the default toolbar is specified by the function set-default-toolbar-position. If the corresponding position-specific toolbar thickness

specifier (e.g. top-toolbar-height if default-toolbar-position is top)

does not specify a thickness in a particular domain (a window or a frame),

then the value of default-toolbar-height or default-toolbar-width (de-

pending on the toolbar orientation) in that domain, if any, will be used instead.

Specifier default-toolbar-width

This specifies the width of the default toolbar, if it’s oriented vertically. This

behaves like default-toolbar-height.
Note that default-toolbar-height is only used when default-toolbar-position
is top or bottom, and default-toolbar-width is only used when default-toolbar-
position is left or right.

Specifier top-toolbar-height

This specifies the height of the top toolbar.

Specifier bottom-toolbar-height

This specifies the height of the bottom toolbar.

Specifier left-toolbar-width

This specifies the width of the left toolbar.

Specifier right-toolbar-width

This specifies the width of the right toolbar.

Note that all of the position-specific toolbar thickness specifiers have a fallback value of zero when they do not correspond to the default toolbar. Therefore, you will have to set a non-zero thickness value if you want a position-specific toolbar to be displayed.

Specifier default-toolbar-visible-p

This specifies whether the default toolbar is visible. The position of the default toolbar is specified by the function set-default-toolbar-position. If the corresponding position-specific toolbar visibility specifier (e.g. top-toolbar-

     visible-p if default-toolbar-position is top) does not specify a visible-p

value in a particular domain (a window or a frame), then the value of defaulttoolbar-visible-p in that domain, if any, will be used instead.

Specifier top-toolbar-visible-p

This specifies whether the top toolbar is visible.

Specifier bottom-toolbar-visible-p

This specifies whether the bottom toolbar is visible.

Specifier left-toolbar-visible-p

This specifies whether the left toolbar is visible.

Specifier right-toolbar-visible-p

This specifies whether the right toolbar is visible.

default-toolbar-visible-p and all of the position-specific toolbar visibility specifiers have a fallback value of true.

Internally, toolbar thickness and visibility specifiers are instantiated in both window and frame domains, for different purposes. The value in the domain of a frame’s selected window specifies the actual toolbar thickness or visibility that you will see in that frame. The value in the domain of a frame itself specifies the toolbar thickness or visibility that is used in frame geometry calculations.

Thus, for example, if you set the frame width to 80 characters and the left toolbar width for that frame to 68 pixels, then the frame will be sized to fit 80 characters plus a 68-pixel left toolbar. If you then set the left toolbar width to 0 for a particular buffer (or if that buffer does not specify a left toolbar or has a nil value specified for left-toolbar-visible-p), you will find that, when that buffer is displayed in the selected window, the window will have a width of 86 or 87 characters – the frame is sized for a 68-pixel left toolbar but the selected window specifies that the left toolbar is not visible, so it is expanded to take up the slack.

Specifier toolbar-buttons-captioned-p

Whether toolbar buttons are captioned. This affects which glyphs from a toolbar button descriptor are chosen. See Section 23.2 [Toolbar Descriptor Format], page 351.

356 XEmacs Lisp Reference Manual

You can also reset the toolbar to what it was when XEmacs started up.

Constant initial-toolbar-spec

The toolbar descriptor used to initialize default-toolbar at startup.

Chapter 24: scrollbars 357

24 scrollbars

Not yet documented.

358 XEmacs Lisp Reference Manual

25 Drag and Drop

WARNING: the Drag’n’Drop API is still under development and the interface may change! The current implementation is considered experimental.

Drag’n’drop is a way to transfer information between multiple applications. To do this several GUIs define their own protocols. Examples are OffiX, CDE, Motif, KDE, MSWindows, GNOME, and many more. To catch all these protocols, XEmacs provides a generic API.

One prime idea behind the API is to use a data interface that is transparent for all systems. The author thinks that this is best archived by using URL and MIME data, cause any internet enabled system must support these for email already. XEmacs also already provides powerful interfaces to support these types of data (tm and w3).

25.1 Supported Protocols

The current release of XEmacs only support a small set of Drag’n’drop protocols. Some of these only support limited options available in the API.

25.1.1 OffiX DND

WARNING: If you compile in OffiX, you may not be able to use multiple X displays successfully. If the two servers are from different vendors, the results may be unpredictable.

The OffiX Drag’n’Drop protocol is part of a X API/Widget library created by Cesar Crusius. It is based on X-Atoms and ClientMessage events, and works with any X platform supporting them.

OffiX is supported if ’offix is member of the variable dragdrop-protocols, or the feature ’offix is defined.

Unfortunately it uses it’s own data types. Examples are: File, Files, Exe, Link, URL, MIME. The API tries to choose the right type for the data that is dragged from XEmacs (well, not yet...).

XEmacs supports both MIME and URL drags and drops using this API. No application interaction is possible while dragging is in progress.

For information about the OffiX project have a look at http://leb.net/~offix/

25.1.2 CDE dt

CDE stands for Common Desktop Environment. It is based on the Motif widget library. It’s drag’n’drop protocol is also an abstraction of the Motif protocol (so it might be possible, that XEmacs will also support the Motif protocol soon).

CDE has three different types: file, buffer, and text. XEmacs only uses file and buffer drags. The API will disallow full URL drags, only file method URLs are passed through.

Buffer drags are always converted to plain text.

25.1.3 MSWindows OLE

Only allows file drags and drops.

25.1.4 Loose ends

The following protocols will be supported soon: Xdnd, Motif, Xde (if I get some specs), KDE OffiX (if KDE can find XEmacs windows).

In particular Xdnd will be one of the protocols that can benefit from the XEmacs API, cause it also uses MIME types to encode dragged data.

25.2 Drop Interface

For each activated low-level protocol, a internal routine will catch incoming drops and convert them to a dragdrop-drop type misc-user-event.

This misc-user-event has its function argument set to dragdrop-drop-dispatch and the object contains the data of the drop (converted to URL/MIME specific data). This function will search the variable experimental-dragdrop-drop-functions for a function that can handle the dropped data.

To modify the drop behavior, the user can modify the variable experimental-dragdrop-drop-functions. Each element of this list specifies a possible handler for dropped data. The first one that can handle the data will return t and exit. Another possibility is to set a extent-property with the same name. Extents are checked prior to the variable.

The customization group drag-n-drop shows all variables of user interest.

25.3 Drag Interface

This describes the drag API (not implemented yet).

26 Major and Minor Modes

A mode is a set of definitions that customize XEmacs and can be turned on and off while you edit. There are two varieties of modes: major modes, which are mutually exclusive and used for editing particular kinds of text, and minor modes, which provide features that users can enable individually.

This chapter describes how to write both major and minor modes, how to indicate them in the modeline, and how they run hooks supplied by the user. For related topics such as keymaps and syntax tables, see Chapter 20 [Keymaps], page 315, and Chapter 38 [Syntax Tables], page 569.

26.1 Major Modes

Major modes specialize XEmacs for editing particular kinds of text. Each buffer has only one major mode at a time.

The least specialized major mode is called Fundamental mode. This mode has no modespecific definitions or variable settings, so each XEmacs command behaves in its default manner, and each option is in its default state. All other major modes redefine various keys and options. For example, Lisp Interaction mode provides special key bindings for ^{⟨}\underline{\overline{LFD}}^{⟩}

(eval-print-last-sexp), ^{⟨}\underline{\overline{TAB}}^{⟩} (lisp-indent-line), and other keys.

When you need to write several editing commands to help you perform a specialized editing task, creating a new major mode is usually a good idea. In practice, writing a major mode is easy (in contrast to writing a minor mode, which is often difficult).

If the new mode is similar to an old one, it is often unwise to modify the old one to serve two purposes, since it may become harder to use and maintain. Instead, copy and rename an existing major mode definition and alter the copy—or define a derived mode (see Section 26.1.5 [Derived Modes], page 370). For example, Rmail Edit mode, which is in ‘emacs/lisp/rmailedit.el’, is a major mode that is very similar to Text mode except that it provides three additional commands. Its definition is distinct from that of Text mode, but was derived from it.

Rmail Edit mode is an example of a case where one piece of text is put temporarily into a different major mode so it can be edited in a different way (with ordinary XEmacs commands rather than Rmail). In such cases, the temporary major mode usually has a command to switch back to the buffer’s usual mode (Rmail mode, in this case). You might be tempted to present the temporary redefinitions inside a recursive edit and restore the usual ones when the user exits; but this is a bad idea because it constrains the user’s options when it is done in more than one buffer: recursive edits must be exited most-recently-entered first. Using alternative major modes avoids this limitation. See Section 19.10 [Recursive Editing], page 310.

The standard XEmacs Lisp library directory contains the code for several major

modes, in files including ‘text-mode.el’, ‘texinfo.el’, ‘lisp-mode.el’, ‘c-mode.el’,

and ‘rmail.el’. You can look at these libraries to see how modes are written. Text mode is perhaps the simplest major mode aside from Fundamental mode. Rmail mode is a complicated and specialized mode.

26.1.1 Major Mode Conventions

The code for existing major modes follows various coding conventions, including conventions for local keymap and syntax table initialization, global names, and hooks. Please follow these conventions when you define a new major mode:

• Define a command whose name ends in ‘-mode’, with no arguments, that switches to the new mode in the current buffer. This command should set up the keymap, syntax table, and local variables in an existing buffer without changing the buffer’s text.

• Write a documentation string for this command that describes the special commands available in this mode. C-h m (describe-mode) in your mode will display this string. The documentation string may include the special documentation substrings, ‘\[command]’, ‘\{keymap}’, and ‘\<keymap>’, that enable the documentation to adapt automatically to the user’s own key bindings. See Section 27.3 [Keys in Documentation], page 384.

• The major mode command should start by calling kill-all-local-variables. This is what gets rid of the local variables of the major mode previously in effect.

• The major mode command should set the variable major-mode to the major mode command symbol. This is how describe-mode discovers which documentation to print.

• The major mode command should set the variable mode-name to the “pretty” name of the mode, as a string. This appears in the mode line.

• Since all global names are in the same name space, all the global variables, constants, and functions that are part of the mode should have names that start with the major mode name (or with an abbreviation of it if the name is long). See Section A.1 [Style Tips], page 757.

• The major mode should usually have its own keymap, which is used as the local keymap in all buffers in that mode. The major mode function should call use-local-map to install this local map. See Section 20.7 [Active Keymaps], page 320, for more information. This keymap should be kept in a global variable named modename-mode-map. Normally the library that defines the mode sets this variable.

• The mode may have its own syntax table or may share one with other related modes. If it has its own syntax table, it should store this in a variable named modename-mode-syntax-table. See Chapter 38 [Syntax Tables], page 569.

• The mode may have its own abbrev table or may share one with other related modes. If it has its own abbrev table, it should store this in a variable named modenamemode-abbrev-table. See Section 39.2 [Abbrev Tables], page 579.

• Use defvar to set mode-related variables, so that they are not reinitialized if they already have a value. (Such reinitialization could discard customizations made by the user.)

• To make a buffer-local binding for an Emacs customization variable, use make-local-variable in the major mode command, not make-variable-buffer-local. The latter function would make the variable local to every buffer in which it is subsequently set, which would affect buffers that do not use this mode. It is undesirable for a mode to have such global effects. See Section 10.9 [Buffer-Local Variables], page 157.

It’s ok to use make-variable-buffer-local, if you wish, for a variable used only within a single Lisp package.

• Each major mode should have a mode hook named modename-mode-hook. The major mode command should run that hook, with run-hooks, as the very last thing it does. See Section 26.4 [Hooks], page 378.

• The major mode command may also run the hooks of some more basic modes. For

example, indented-text-mode runs text-mode-hook as well as indented-text-mode-

hook. It may run these other hooks immediately before the mode’s own hook (that is, after everything else), or it may run them earlier.

• If something special should be done if the user switches a buffer from this mode to any other major mode, the mode can set a local value for change-major-mode-hook.

• If this mode is appropriate only for specially-prepared text, then the major mode command symbol should have a property named mode-class with value special, put on as follows:

     (put 'funny-mode 'mode-class 'special)

This tells XEmacs that new buffers created while the current buffer has Funny mode should not inherit Funny mode. Modes such as Dired, Rmail, and Buffer List use this feature.

• If you want to make the new mode the default for files with certain recognizable names, add an element to auto-mode-alist to select the mode for those file names. If you define the mode command to autoload, you should add this element in the same file that calls autoload. Otherwise, it is sufficient to add the element in the file that contains the mode definition. See Section 26.1.3 [Auto Major Mode], page 367.

• In the documentation, you should provide a sample autoload form and an example of how to add to auto-mode-alist, that users can include in their ‘.emacs’ files.

• The top-level forms in the file defining the mode should be written so that they may be evaluated more than once without adverse consequences. Even if you never load the file more than once, someone else will.

Variable change-major-mode-hook

This normal hook is run by kill-all-local-variables before it does anything else. This gives major modes a way to arrange for something special to be done if the user switches to a different major mode. For best results, make this variable buffer-local, so that it will disappear after doing its job and will not interfere with the subsequent major mode. See Section 26.4 [Hooks], page 378.

26.1.2 Major Mode Examples

Text mode is perhaps the simplest mode besides Fundamental mode. Here are excerpts from ‘text-mode.el’ that illustrate many of the conventions listed above:

;; Create mode-specific tables.

 (defvar text-mode-syntax-table nil
   "Syntax table used while in text mode.")

(if text-mode-syntax-table

() ; Do not change the table if it is already set up.

(setq text-mode-syntax-table (make-syntax-table))
(modify-syntax-entry ?\" ".   " text-mode-syntax-table)
(modify-syntax-entry ?\\ ".   " text-mode-syntax-table)
(modify-syntax-entry ?' "w   " text-mode-syntax-table))
(defvar text-mode-abbrev-table nil
"Abbrev table used while in text mode.")
(define-abbrev-table 'text-mode-abbrev-table ())

(defvar text-mode-map nil) ; Create a mode-specific keymap.

(if text-mode-map

() ; Do not change the keymap if it is already set up.

(setq text-mode-map (make-sparse-keymap))
(define-key text-mode-map "\t" 'tab-to-tab-stop)
(define-key text-mode-map "\es" 'center-line)
(define-key text-mode-map "\eS" 'center-paragraph))

Here is the complete major mode function definition for Text mode:

(defun text-mode ()
"Major mode for editing text intended for humans to read.
Special commands: \\{text-mode-map}
Turning on text-mode runs the hook `text-mode-hook'."
(interactive)
(kill-all-local-variables)
(use-local-map text-mode-map)     ; This provides the local keymap.

(setq mode-name "Text") ; This name goes into the modeline.

(setq major-mode 'text-mode)      ; This is how describe-mode

; finds the doc string to print.

(setq local-abbrev-table text-mode-abbrev-table)
(set-syntax-table text-mode-syntax-table)
(run-hooks 'text-mode-hook))      ; Finally, this permits the user to

; customize the mode with a hook.

The three Lisp modes (Lisp mode, Emacs Lisp mode, and Lisp Interaction mode) have more features than Text mode and the code is correspondingly more complicated. Here are excerpts from ‘lisp-mode.el’ that illustrate how these modes are written.

;; Create mode-specific table variables.

(defvar lisp-mode-syntax-table nil "")
(defvar emacs-lisp-mode-syntax-table nil "")
(defvar lisp-mode-abbrev-table nil "")
(if (not emacs-lisp-mode-syntax-table) ; Do not change the table

; if it is already set.

  (let ((i 0))
    (setq emacs-lisp-mode-syntax-table (make-syntax-table))

;; Set syntax of chars up to 0 to class of chars that are ;; part of symbol names but not words. ;; (The number 0 is 48 in the ASCII character set.)

(while (< i ?0)
  (modify-syntax-entry i "_" emacs-lisp-mode-syntax-table)
  (setq i (1+ i)))
...

;; Set the syntax for other characters.

(modify-syntax-entry ?  "    " emacs-lisp-mode-syntax-table)
(modify-syntax-entry ?\t "    " emacs-lisp-mode-syntax-table)
...
(modify-syntax-entry ?\( "()  " emacs-lisp-mode-syntax-table)
(modify-syntax-entry ?\) ")(  " emacs-lisp-mode-syntax-table)
...))

;; Create an abbrev table for lisp-mode.

(define-abbrev-table 'lisp-mode-abbrev-table ())

Much code is shared among the three Lisp modes. The following function sets various variables; it is called by each of the major Lisp mode functions:

(defun lisp-mode-variables (lisp-syntax)

;; The lisp-syntax argument is nil in Emacs Lisp mode, ;; and t in the other two Lisp modes.

(cond (lisp-syntax
   (if (not lisp-mode-syntax-table)

;; The Emacs Lisp mode syntax table always exists, but ;; the Lisp Mode syntax table is created the first time a ;; mode that needs it is called. This is to save space.

       (progn (setq lisp-mode-syntax-table
                 (copy-syntax-table emacs-lisp-mode-syntax-table))

;; Change some entries for Lisp mode.

              (modify-syntax-entry ?\| "\"   "
                                   lisp-mode-syntax-table)
              (modify-syntax-entry ?\[ "_"
                                   lisp-mode-syntax-table)
              (modify-syntax-entry ?\] "_"
                                   lisp-mode-syntax-table)))
    (set-syntax-table lisp-mode-syntax-table)))
(setq local-abbrev-table lisp-mode-abbrev-table)
...)

Functions such as forward-paragraph use the value of the paragraph-start variable. Since Lisp code is different from ordinary text, the paragraph-start variable needs to be set specially to handle Lisp. Also, comments are indented in a special fashion in Lisp and the Lisp modes need their own mode-specific comment-indent-function. The code to set these variables is the rest of lisp-mode-variables.

(make-local-variable 'paragraph-start)

;; Having ‘^’ is not clean, but page-delimiter ;; has them too, and removing those is a pain.

(setq paragraph-start (concat "^$\\|" page-delimiter))
...
(make-local-variable 'comment-indent-function)
(setq comment-indent-function 'lisp-comment-indent))

Each of the different Lisp modes has a slightly different keymap. For example, Lisp mode binds C-c C-l to run-lisp, but the other Lisp modes do not. However, all Lisp modes have some commands in common. The following function adds these common commands to a given keymap.

(defun lisp-mode-commands (map)
(define-key map "\e\C-q" 'indent-sexp)
(define-key map "\177" 'backward-delete-char-untabify)
(define-key map "\t" 'lisp-indent-line))

Here is an example of using lisp-mode-commands to initialize a keymap, as part of the code for Emacs Lisp mode. First we declare a variable with defvar to hold the modespecific keymap. When this defvar executes, it sets the variable to nil if it was void. Then we set up the keymap if the variable is nil.

This code avoids changing the keymap or the variable if it is already set up. This lets the user customize the keymap.

(defvar emacs-lisp-mode-map () "")
(if emacs-lisp-mode-map
  ()
(setq emacs-lisp-mode-map (make-sparse-keymap))
(define-key emacs-lisp-mode-map "\e\C-x" 'eval-defun)
(lisp-mode-commands emacs-lisp-mode-map))

Finally, here is the complete major mode function definition for Emacs Lisp mode.

(defun emacs-lisp-mode ()
"Major mode for editing Lisp code to run in XEmacs.
Commands:
Delete converts tabs to spaces as it moves back.
Blank lines separate paragraphs.  Semicolons start comments.
\\{emacs-lisp-mode-map}
Entry to this mode runs the hook `emacs-lisp-mode-hook'."
(interactive)
(kill-all-local-variables)
(use-local-map emacs-lisp-mode-map)    ; This provides the local keymap.
(set-syntax-table emacs-lisp-mode-syntax-table)
(setq major-mode 'emacs-lisp-mode)     ; This is how describe-mode

; finds out what to describe.

(setq mode-name "Emacs-Lisp")          ; This goes into the modeline.

(lisp-mode-variables nil) ; This defines various variables.

(run-hooks 'emacs-lisp-mode-hook))     ; This permits the user to use a

; hook to customize the mode.

26.1.3 How XEmacs Chooses a Major Mode

Based on information in the file name or in the file itself, XEmacs automatically selects a major mode for the new buffer when a file is visited.

Command fundamental-mode

Fundamental mode is a major mode that is not specialized for anything in particular. Other major modes are defined in effect by comparison with this one—their definitions say what to change, starting from Fundamental mode. The fundamental-mode function does not run any hooks; you’re not supposed to customize it. (If you want Emacs to behave differently in Fundamental mode, change the global state of Emacs.)

Command normal-mode &optional find-file This function establishes the proper major mode and local variable bindings for the current buffer. First it calls set-auto-mode, then it runs hack-local-variables to parse, and bind or evaluate as appropriate, any local variables. If the find-file argument to normal-mode is non-nil, normal-mode assumes that the find-file function is calling it. In this case, it may process a local variables list at the end of the file and in the ‘-*-’ line. The variable enablelocal-variables controls whether to do so. If you run normal-mode interactively, the argument find-file is normally nil. In this case, normal-mode unconditionally processes any local variables list. See section “Local Variables in Files” in The XEmacs Reference Manual, for the syntax of the local variables section of a file. normal-mode uses condition-case around the call to the major mode function, so errors are caught and reported as a ‘File mode specification error’, followed by the original error message.

User Option enable-local-variables

This variable controls processing of local variables lists in files being visited. A value of t means process the local variables lists unconditionally; nil means ignore them; anything else means ask the user what to do for each file. The default value is t.

Variable ignored-local-variables

This variable holds a list of variables that should not be set by a local variables list. Any value specified for one of these variables is ignored.

In addition to this list, any variable whose name has a non-nil risky-local-variable property is also ignored.

User Option enable-local-eval

This variable controls processing of ‘Eval:’ in local variables lists in files being visited. A value of t means process them unconditionally; nil means ignore them; anything else means ask the user what to do for each file. The default value is maybe.

Function set-auto-mode

This function selects the major mode that is appropriate for the current buffer. It may base its decision on the value of the ‘-*-’ line, on the visited file name (using auto-mode-alist), or on the value of a local variable. However, this function does not look for the ‘mode:’ local variable near the end of a file; the hack-local-variables function does that. See section “How Major Modes are Chosen” in The XEmacs Reference Manual.

User Option default-major-mode

This variable holds the default major mode for new buffers. The standard value

is fundamental-mode.

If the value of default-major-mode is nil, XEmacs uses the (previously) current buffer’s major mode for the major mode of a new buffer. However, if the major mode symbol has a mode-class property with value special, then it is not used for new buffers; Fundamental mode is used instead. The modes that have this property are those such as Dired and Rmail that are useful only with text that has been specially prepared.

Function set-buffer-major-mode buffer

This function sets the major mode of buffer to the value of default-major-mode. If that variable is nil, it uses the current buffer’s major mode (if that is suitable). The low-level primitives for creating buffers do not use this function, but

medium-level commands such as switch-to-buffer and find-file-noselect

use it whenever they create buffers.

Variable initial-major-mode

The value of this variable determines the major mode of the initial ‘*scratch*’ buffer. The value should be a symbol that is a major mode command name.

The default value is lisp-interaction-mode.

Variable auto-mode-alist

This variable contains an association list of file name patterns (regular expressions; see Section 37.2 [Regular Expressions], page 550) and corresponding major mode functions. Usually, the file name patterns test for suffixes, such as ‘.el’ and ‘.c’, but this need not be the case. An ordinary element of the alist looks like (regexp . mode-function). For example,

     (("^/tmp/fol/" . text-mode)
      ("\\.texinfo\\'" . texinfo-mode)
      ("\\.texi\\'" . texinfo-mode)
      ("\\.el\\'" . emacs-lisp-mode)
      ("\\.c\\'" . c-mode)
      ("\\.h\\'" . c-mode)
      ...)

When you visit a file whose expanded file name (see Section 28.8.4 [File Name Expansion], page 409) matches a regexp, set-auto-mode calls the corresponding mode-function. This feature enables XEmacs to select the proper major mode for most files. If an element of auto-mode-alist has the form (regexp function t), then after calling function, XEmacs searches auto-mode-alist again for a match against the portion of the file name that did not match before. This match-again feature is useful for uncompression packages: an entry of the form ("\\.gz\\'" . function) can uncompress the file and then put the uncompressed file in the proper mode according to the name sans ‘.gz’. Here is an example of how to prepend several pattern pairs to auto-mode-alist. (You might use this sort of expression in your ‘.emacs’ file.)

     (setq auto-mode-alist
       (append

;; File name starts with a dot.

        '(("/\\.[^/]*\\'" . fundamental-mode)

;; File name has no dot.

          ("[^\\./]*\\'" . fundamental-mode)

;; File name ends in ‘.C’.

          ("\\.C\\'" . c++-mode))
        auto-mode-alist))

Variable interpreter-mode-alist

This variable specifies major modes to use for scripts that specify a command interpreter in an ‘#!’ line. Its value is a list of elements of the form (interpreter . mode); for example, ("perl" . perl-mode) is one element present by default. The element says to use mode mode if the file specifies interpreter. This variable is applicable only when the auto-mode-alist does not indicate which major mode to use.

Function hack-local-variables &optional force

This function parses, and binds or evaluates as appropriate, any local variables for the current buffer.

The handling of enable-local-variables documented for normal-mode ac-

tually takes place here. The argument force usually comes from the argument find-file given to normal-mode.

26.1.4 Getting Help about a Major Mode

The describe-mode function is used to provide information about major modes. It is normally called with C-h m. The describe-mode function uses the value of major-mode, which is why every major mode function needs to set the major-mode variable.

Command describe-mode

This function displays the documentation of the current major mode. The describe-mode function calls the documentation function using the value of major-mode as an argument. Thus, it displays the documentation string of the major mode function. (See Section 27.2 [Accessing Documentation], page 382.)

Variable major-mode

This variable holds the symbol for the current buffer’s major mode. This symbol should have a function definition that is the command to switch to that major mode. The describe-mode function uses the documentation string of the function as the documentation of the major mode.

26.1.5 Defining Derived Modes

It’s often useful to define a new major mode in terms of an existing one. An easy way

to do this is to use define-derived-mode.

Macro define-derived-mode variant parent name docstring body… This construct defines variant as a major mode command, using name as the string form of the mode name.

The new command variant is defined to call the function parent, then override certain aspects of that parent mode:

• The new mode has its own keymap, named variant-map. define-derived-mode initializes this map to inherit from parent-map, if it is not already set.

• The new mode has its own syntax table, kept in the variable variantsyntax-table. define-derived-mode initializes this variable by copying parent-syntax-table, if it is not already set.

• The new mode has its own abbrev table, kept in the variable variantabbrev-table. define-derived-mode initializes this variable by copying parent-abbrev-table, if it is not already set.

• The new mode has its own mode hook, variant-hook, which it runs in standard fashion as the very last thing that it does. (The new mode also runs the mode hook of parent as part of calling parent.)

In addition, you can specify how to override other aspects of parent with body. The command variant evaluates the forms in body after setting up all its usual overrides, just before running variant-hook.

The argument docstring specifies the documentation string for the new mode. If you omit docstring, define-derived-mode generates a documentation string.

Here is a hypothetical example:

     (define-derived-mode hypertext-mode
       text-mode "Hypertext"
       "Major mode for hypertext.
     \\{hypertext-mode-map}"
       (setq case-fold-search nil))
     (define-key hypertext-mode-map
       [down-mouse-3] 'do-hyper-link)

26.2 Minor Modes

A minor mode provides features that users may enable or disable independently of the choice of major mode. Minor modes can be enabled individually or in combination. Minor modes would be better named “Generally available, optional feature modes” except that such a name is unwieldy.

A minor mode is not usually a modification of single major mode. For example, Auto Fill mode may be used in any major mode that permits text insertion. To be general, a minor mode must be effectively independent of the things major modes do.

A minor mode is often much more difficult to implement than a major mode. One reason is that you should be able to activate and deactivate minor modes in any order. A minor mode should be able to have its desired effect regardless of the major mode and regardless of the other minor modes in effect.

Often the biggest problem in implementing a minor mode is finding a way to insert the necessary hook into the rest of XEmacs. Minor mode keymaps make this easier than it used to be.

26.2.1 Conventions for Writing Minor Modes

There are conventions for writing minor modes just as there are for major modes. Several of the major mode conventions apply to minor modes as well: those regarding the name of the mode initialization function, the names of global symbols, and the use of keymaps and other tables.

In addition, there are several conventions that are specific to minor modes.

• Make a variable whose name ends in ‘-mode’ to represent the minor mode. Its value should enable or disable the mode (nil to disable; anything else to enable.) We call this the mode variable. This variable is used in conjunction with the minor-mode-alist to display the minor mode name in the modeline. It can also enable or disable a minor mode keymap. Individual commands or hooks can also check the variable’s value. If you want the minor mode to be enabled separately in each buffer, make the variable buffer-local.

• Define a command whose name is the same as the mode variable. Its job is to enable and disable the mode by setting the variable. The command should accept one optional argument. If the argument is nil, it should toggle the mode (turn it on if it is off, and off if it is on). Otherwise, it should turn the mode on if the argument is a positive integer, a symbol other than nil or -, or a list whose car is such an integer or symbol; it should turn the mode off otherwise. Here is an example taken from the definition of transient-mark-mode. It shows the use of transient-mark-mode as a variable that enables or disables the mode’s behavior, and also shows the proper way to toggle, enable or disable the minor mode based on the raw prefix argument value.

     (setq transient-mark-mode
           (if (null arg) (not transient-mark-mode)
             (> (prefix-numeric-value arg) 0)))

• Add an element to minor-mode-alist for each minor mode (see Section 26.3.2 [Mod-

eline Variables], page 374). This element should be a list of the following form:

(mode-variable string) Here mode-variable is the variable that controls enabling of the minor mode, and string is a short string, starting with a space, to represent the mode in the modeline. These strings must be short so that there is room for several of them at once. When you add an element to minor-mode-alist, use assq to check for an existing element, to avoid duplication. For example:

         (or (assq 'leif-mode minor-mode-alist)
             (setq minor-mode-alist
                   (cons '(leif-mode " Leif") minor-mode-alist)))

26.2.2 Keymaps and Minor Modes

Each minor mode can have its own keymap, which is active when the mode is enabled. To set up a keymap for a minor mode, add an element to the alist minor-mode-map-alist. See Section 20.7 [Active Keymaps], page 320.

One use of minor mode keymaps is to modify the behavior of certain self-inserting characters so that they do something else as well as self-insert. In general, this is the only way to do that, since the facilities for customizing self-insert-command are limited to special cases (designed for abbrevs and Auto Fill mode). (Do not try substituting your own definition of self-insert-command for the standard one. The editor command loop handles this function specially.)

26.3 Modeline Format

Each Emacs window (aside from minibuffer windows) includes a modeline, which displays status information about the buffer displayed in the window. The modeline contains information about the buffer, such as its name, associated file, depth of recursive editing, and the major and minor modes.

This section describes how the contents of the modeline are controlled. It is in the chapter on modes because much of the information displayed in the modeline relates to the enabled major and minor modes.

modeline-format is a buffer-local variable that holds a template used to display the modeline of the current buffer. All windows for the same buffer use the same modelineformat and their modelines appear the same (except for scrolling percentages and line numbers).

The modeline of a window is normally updated whenever a different buffer is shown in the window, or when the buffer’s modified-status changes from nil to t or vice-versa. If you modify any of the variables referenced by modeline-format (see Section 26.3.2 [Modeline Variables], page 374), you may want to force an update of the modeline so as to display the new information.

Function redraw-modeline &optional all

Force redisplay of the current buffer’s modeline. If all is non-nil, then force redisplay of all modelines.

The modeline is usually displayed in inverse video. This is controlled using the modeline face. See Section 42.1 [Faces], page 615.

26.3.1 The Data Structure of the Modeline

The modeline contents are controlled by a data structure of lists, strings, symbols, and numbers kept in the buffer-local variable mode-line-format. The data structure is called a modeline construct, and it is built in recursive fashion out of simpler modeline constructs. The same data structure is used for constructing frame titles (see Section 32.3 [Frame Titles], page 476).

Variable modeline-format

The value of this variable is a modeline construct with overall responsibility for the modeline format. The value of this variable controls which other variables are used to form the modeline text, and where they appear.

A modeline construct may be as simple as a fixed string of text, but it usually specifies how to use other variables to construct the text. Many of these variables are themselves defined to have modeline constructs as their values.

The default value of modeline-format incorporates the values of variables such as modename and minor-mode-alist. Because of this, very few modes need to alter modelineformat. For most purposes, it is sufficient to alter the variables referenced by modeline-

format.

A modeline construct may be a list, a symbol, or a string. If the value is a list, each element may be a list, a symbol, or a string.

string A string as a modeline construct is displayed verbatim in the mode line except for %-constructs. Decimal digits after the ‘%’ specify the field width for space filling on the right (i.e., the data is left justified). See Section 26.3.3 [%-Constructs], page 376.

symbol A symbol as a modeline construct stands for its value. The value of symbol is used as a modeline construct, in place of symbol. However, the symbols t and nil are ignored; so is any symbol whose value is void. There is one exception: if the value of symbol is a string, it is displayed verbatim: the %-constructs are not recognized.

(string rest...) or (list rest...)

A list whose first element is a string or list means to process all the elements recursively and concatenate the results. This is the most common form of mode line construct.

(symbol then else)

A list whose first element is a symbol is a conditional. Its meaning depends on the value of symbol. If the value is non-nil, the second element, then, is processed recursively as a modeline element. But if the value of symbol is nil, the third element, else, is processed recursively. You may omit else; then the mode line element displays nothing if the value of symbol is nil.

(width rest...)

A list whose first element is an integer specifies truncation or padding of the results of rest. The remaining elements rest are processed recursively as modeline constructs and concatenated together. Then the result is space filled (if width is positive) or truncated (to −width columns, if width is negative) on the right. For example, the usual way to show what percentage of a buffer is above the top of the window is to use a list like this: (-3 "%p").

If you do alter modeline-format itself, the new value should use the same variables that appear in the default value (see Section 26.3.2 [Modeline Variables], page 374), rather than duplicating their contents or displaying the information in another fashion. This way, customizations made by the user or by Lisp programs (such as display-time and major modes) via changes to those variables remain effective.

Here is an example of a modeline-format that might be useful for shell-mode, since it contains the hostname and default directory.

(setq modeline-format
(list ""
'modeline-modified
"%b--"

(getenv "HOST") ; One element is not constant.

":"
'default-directory
"   "
'global-mode-string
"   %[("
'mode-name
'modeline-process
'minor-mode-alist
"%n"
")%]----"
'(line-number-mode "L%l--")
'(-3 . "%p")
"-%-"))

26.3.2 Variables Used in the Modeline

This section describes variables incorporated by the standard value of modeline-format into the text of the mode line. There is nothing inherently special about these variables; any other variables could have the same effects on the modeline if modeline-format were changed to use them.

Variable modeline-modified

This variable holds the value of the modeline construct that displays whether the current buffer is modified. The default value of modeline-modified is ("--%1*%1+-"). This means that the modeline displays ‘--**-’ if the buffer is modified, ‘-----’ if the buffer is not modified, ‘--%%-’ if the buffer is read only, and ‘--%*--’ if the buffer is read only and modified. Changing this variable does not force an update of the modeline.

Variable modeline-buffer-identification

This variable identifies the buffer being displayed in the window. Its default value is ("%F: %17b"), which means that it usually displays ‘Emacs:’ followed by seventeen characters of the buffer name. (In a terminal frame, it displays the frame name instead of ‘Emacs’; this has the effect of showing the frame number.) You may want to change this in modes such as Rmail that do not behave like a “normal” XEmacs.

Variable global-mode-string

This variable holds a modeline spec that appears in the mode line by default, just after the buffer name. The command display-time sets global-mode-string to refer to the variable display-time-string, which holds a string containing the time and load information. The ‘%M’ construct substitutes the value of global-mode-string, but this is obsolete, since the variable is included directly in the modeline.

Variable mode-name

This buffer-local variable holds the “pretty” name of the current buffer’s major mode. Each major mode should set this variable so that the mode name will appear in the modeline.

Variable minor-mode-alist

This variable holds an association list whose elements specify how the modeline should indicate that a minor mode is active. Each element of the minor-mode-alist should be a two-element list:

(minor-mode-variable modeline-string) More generally, modeline-string can be any mode line spec. It appears in the mode line when the value of minor-mode-variable is non-nil, and not otherwise. These strings should begin with spaces so that they don’t run together. Conventionally, the minor-mode-variable for a specific mode is set to a non-nil value when that minor mode is activated. The default value of minor-mode-alist is:

     minor-mode-alist
     ⇒ ((vc-mode vc-mode)
         (abbrev-mode " Abbrev")
         (overwrite-mode overwrite-mode)
         (auto-fill-function " Fill")
         (defining-kbd-macro " Def")
         (isearch-mode isearch-mode))

minor-mode-alist is not buffer-local. The variables mentioned in the alist should be buffer-local if the minor mode can be enabled separately in each buffer.

Variable modeline-process

This buffer-local variable contains the modeline information on process status in modes used for communicating with subprocesses. It is displayed immediately following the major mode name, with no intervening space. For example, its value in the ‘*shell*’ buffer is (": %s"), which allows the shell to display its status along with the major mode as: ‘(Shell: run)’. Normally this variable

is nil.

Variable default-modeline-format

This variable holds the default modeline-format for buffers that do not over-

ride it. This is the same as (default-value 'modeline-format).
The default value of default-modeline-format is:
(""
modeline-modified
modeline-buffer-identification
"   "
global-mode-string
"   %[("
mode-name
modeline-process
minor-mode-alist
"%n"
")%]----"
(line-number-mode "L%l--")
(-3 . "%p")
"-%-")

Variable vc-mode

The variable vc-mode, local in each buffer, records whether the buffer’s visited file is maintained with version control, and, if so, which kind. Its value is nil for no version control, or a string that appears in the mode line.

26.3.3 %-Constructs in the ModeLine

The following table lists the recognized %-constructs and what they mean. In any construct except ‘%%’, you can add a decimal integer after the ‘%’ to specify how many characters to display.

%b The current buffer name, obtained with the buffer-name function. See Section 30.3 [Buffer Names], page 433.

%f The visited file name, obtained with the buffer-file-name function. See Section 30.4 [Buffer File Name], page 434.

%F The name of the selected frame.

%c The current column number of point.

%l The current line number of point.

%* ‘%’ if the buffer is read only (see buffer-read-only); ‘*’ if the buffer is modified (see buffer-modified-p); ‘-’ otherwise. See Section 30.5 [Buffer Modification], page 436.

%+ ‘*’ if the buffer is modified (see buffer-modified-p); ‘%’ if the buffer is read only (see buffer-read-only); ‘-’ otherwise. This differs from ‘%*’ only for a modified read-only buffer. See Section 30.5 [Buffer Modification], page 436.

%& ‘*’ if the buffer is modified, and ‘-’ otherwise.

%s The status of the subprocess belonging to the current buffer, obtained with process-status. See Section 49.6 [Process Information], page 679.

%l the current line number.

%S the name of the selected frame; this is only meaningful under the X Window System. See Section 32.2.5 [Frame Name], page 475.

%t Whether the visited file is a text file or a binary file. (This is a meaningful distinction only on certain operating systems.)

%p The percentage of the buffer text above the top of window, or ‘Top’, ‘Bottom’ or ‘All’.

%P The percentage of the buffer text that is above the bottom of the window (which includes the text visible in the window, as well as the text above the top), plus ‘Top’ if the top of the buffer is visible on screen; or ‘Bottom’ or ‘All’.

%n ‘Narrow’ when narrowing is in effect; nothing otherwise (see narrow-to-region in Section 34.4 [Narrowing], page 496).

%[ An indication of the depth of recursive editing levels (not counting minibuffer levels): one ‘[’ for each editing level. See Section 19.10 [Recursive Editing], page 310.

%] One ‘]’ for each recursive editing level (not counting minibuffer levels).

%% The character ‘%’—this is how to include a literal ‘%’ in a string in which %- constructs are allowed.

%- Dashes sufficient to fill the remainder of the modeline.

The following two %-constructs are still supported, but they are obsolete, since you can get the same results with the variables mode-name and global-mode-string.

%m  The value of mode-name.

%M The value of global-mode-string. Currently, only display-time modifies the

          value of global-mode-string.

26.4 Hooks

A hook is a variable where you can store a function or functions to be called on a particular occasion by an existing program. XEmacs provides hooks for the sake of customization. Most often, hooks are set up in the ‘.emacs’ file, but Lisp programs can set them also. See Appendix F [Standard Hooks], page 785, for a list of standard hook variables.

Most of the hooks in XEmacs are normal hooks. These variables contain lists of functions to be called with no arguments. The reason most hooks are normal hooks is so that you can use them in a uniform way. You can usually tell when a hook is a normal hook, because its name ends in ‘-hook’.

The recommended way to add a hook function to a normal hook is by calling add-hook (see below). The hook functions may be any of the valid kinds of functions that funcall accepts (see Section 11.1 [What Is a Function], page 163). Most normal hook variables are initially void; add-hook knows how to deal with this.

As for abnormal hooks, those whose names end in ‘-function’ have a value that is a single function. Those whose names end in ‘-hooks’ have a value that is a list of functions. Any hook that is abnormal is abnormal because a normal hook won’t do the job; either the functions are called with arguments, or their values are meaningful. The name shows you that the hook is abnormal and that you should look at its documentation string to see how to use it properly.

Major mode functions are supposed to run a hook called the mode hook as the last step of initialization. This makes it easy for a user to customize the behavior of the mode, by overriding the local variable assignments already made by the mode. But hooks are used in other contexts too. For example, the hook suspend-hook runs just before XEmacs suspends itself (see Section 50.2.2 [Suspending XEmacs], page 696).

Here’s an expression that uses a mode hook to turn on Auto Fill mode when in Lisp Interaction mode:

     (add-hook 'lisp-interaction-mode-hook 'turn-on-auto-fill)

The next example shows how to use a hook to customize the way XEmacs formats C code. (People often have strong personal preferences for one format or another.) Here the hook function is an anonymous lambda expression.

     (add-hook 'c-mode-hook
       (function (lambda ()
                   (setq c-indent-level 4
                         c-argdecl-indent 0
                         c-label-offset -4
                         c-continued-statement-indent 0
                         c-brace-offset 0
                         comment-column 40))))
     (setq c++-mode-hook c-mode-hook)

The final example shows how the appearance of the modeline can be modified for a particular class of buffers only.

(add-hook 'text-mode-hook
  (function (lambda ()
              (setq modeline-format
                    '(modeline-modified
                      "Emacs: %14b"
                      "  "
                      default-directory
                      " "
                      global-mode-string
                      "%[("
                      mode-name
                      minor-mode-alist
                      "%n"
                      modeline-process
                      ") %]---"
                      (-3 . "%p")
                      "-%-")))))

At the appropriate time, XEmacs uses the run-hooks function to run particular hooks. This function calls the hook functions you have added with add-hooks.

Function run-hooks &rest hookvar This function takes one or more hook variable names as arguments, and runs each hook in turn. Each hookvar argument should be a symbol that is a hook variable. These arguments are processed in the order specified. If a hook variable has a non-nil value, that value may be a function or a list of functions. If the value is a function (either a lambda expression or a symbol with a function definition), it is called. If it is a list, the elements are called, in order. The hook functions are called with no arguments. For example, here’s how emacs-lisp-mode runs its mode hook:

     (run-hooks 'emacs-lisp-mode-hook)

Function add-hook hook function &optional append local This function is the handy way to add function function to hook variable hook. The argument function may be any valid Lisp function with the proper number of arguments. For example,

     (add-hook 'text-mode-hook 'my-text-hook-function)
adds my-text-hook-function to the hook called text-mode-hook.

You can use add-hook for abnormal hooks as well as for normal hooks. It is best to design your hook functions so that the order in which they are executed does not matter. Any dependence on the order is “asking for trouble.” However, the order is predictable: normally, function goes at the front of the hook list, so it will be executed first (barring another add-hook call). If the optional argument append is non-nil, the new hook function goes at the end of the hook list and will be executed last. If local is non-nil, that says to make the new hook function local to the current buffer. Before you can do this, you must make the hook itself buffer-local by

calling make-local-hook (not make-local-variable). If the hook itself is not

buffer-local, then the value of local makes no difference—the hook function is always global.

Function remove-hook hook function &optional local This function removes function from the hook variable hook. If local is non-nil, that says to remove function from the local hook list instead of from the global hook list. If the hook itself is not buffer-local, then the value of local makes no difference.

Function make-local-hook hook

This function makes the hook variable hook local to the current buffer. When a hook variable is local, it can have local and global hook functions, and runhooks runs all of them. This function works by making t an element of the buffer-local value. That serves as a flag to use the hook functions in the default value of the hook variable as well as those in the local value. Since run-hooks understands this flag, make-local-hook works with all normal hooks. It works for only some non-normal hooks—those whose callers have been updated to understand this meaning of t. Do not use make-local-variable directly for hook variables; it is not sufficient.

27 Documentation

XEmacs Lisp has convenient on-line help facilities, most of which derive their information from the documentation strings associated with functions and variables. This chapter describes how to write good documentation strings for your Lisp programs, as well as how to write programs to access documentation.

Note that the documentation strings for XEmacs are not the same thing as the XEmacs manual. Manuals have their own source files, written in the Texinfo language; documentation strings are specified in the definitions of the functions and variables they apply to. A collection of documentation strings is not sufficient as a manual because a good manual is not organized in that fashion; it is organized in terms of topics of discussion.

27.1 Documentation Basics

A documentation string is written using the Lisp syntax for strings, with double-quote characters surrounding the text of the string. This is because it really is a Lisp string object. The string serves as documentation when it is written in the proper place in the definition of a function or variable. In a function definition, the documentation string follows the argument list. In a variable definition, the documentation string follows the initial value of the variable.

When you write a documentation string, make the first line a complete sentence (or two complete sentences) since some commands, such as apropos, show only the first line of a multi-line documentation string. Also, you should not indent the second line of a documentation string, if you have one, because that looks odd when you use C-h f

(describe-function) or C-h v (describe-variable). See Section A.3 [Documentation

Tips], page 760.

Documentation strings may contain several special substrings, which stand for key bindings to be looked up in the current keymaps when the documentation is displayed. This allows documentation strings to refer to the keys for related commands and be accurate even when a user rearranges the key bindings. (See Section 27.2 [Accessing Documentation], page 382.)

Within the Lisp world, a documentation string is accessible through the function or variable that it describes:

• The documentation for a function is stored in the function definition itself (see Section 11.2 [Lambda Expressions], page 164). The function documentation knows how to extract it.

• The documentation for a variable is stored in the variable’s property list under the

    property name variable-documentation. The function documentation-property

knows how to extract it.

To save space, the documentation for preloaded functions and variables (including primitive functions and autoloaded functions) is stored in the internal doc file ‘DOC’. The documentation for functions and variables loaded during the XEmacs session from byte-compiled files is stored in those very same byte-compiled files (see Section 15.3 [Docs and Compilation], page 208).

XEmacs does not keep documentation strings in memory unless necessary. Instead, XEmacs maintains, for preloaded symbols, an integer offset into the internal doc file, and for symbols loaded from byte-compiled files, a list containing the filename of the bytecompiled file and an integer offset, in place of the documentation string. The functions documentation and documentation-property use that information to read the documentation from the appropriate file; this is transparent to the user.

For information on the uses of documentation strings, see section “Help” in The XEmacs Reference Manual.

The ‘emacs/lib-src’ directory contains two utilities that you can use to print nicelooking hardcopy for the file ‘emacs/etc/DOC-version’. These are ‘sorted-doc.c’ and

‘digest-doc.c’.

27.2 Access to Documentation Strings

Function documentation-property symbol property &optional verbatim This function returns the documentation string that is recorded in symbol’s property list under property property. It retrieves the text from a file if necessary, and runs substitute-command-keys to substitute actual key bindings. (This substitution is not done if verbatim is non-nil; the verbatim argument exists only as of Emacs 19.)

     (documentation-property 'command-line-processed
        'variable-documentation)
          ⇒ "t once command line has been processed"
     (symbol-plist 'command-line-processed)
          ⇒ (variable-documentation 188902)

Function documentation function &optional verbatim This function returns the documentation string of function. It reads the text from a file if necessary. Then (unless verbatim is non-nil) it calls substitutecommand-keys, to return a value containing the actual (current) key bindings.

The function documentation signals a void-function error if function has no function definition. However, it is ok if the function definition has no documentation string. In that case, documentation returns nil.

Here is an example of using the two functions, documentation and documentationproperty, to display the documentation strings for several symbols in a ‘*Help*’ buffer.

(defun describe-symbols (pattern)
  "Describe the XEmacs Lisp symbols matching PATTERN.
All symbols that have PATTERN in their name are described
in the `*Help*' buffer."
  (interactive "sDescribe symbols matching: ")
  (let ((describe-func
         (function
          (lambda (s)

;; Print description of symbol.

            (if (fboundp s)             ; It is a function.
                (princ
                 (format "%s\t%s\n%s\n\n" s
                   (if (commandp s)
                       (let ((keys (where-is-internal s)))
                         (if keys
                             (concat
                              "Keys: "
                              (mapconcat 'key-description
                                         keys " "))
                           "Keys: none"))
                     "Function")
                   (or (documentation s)
                       "not documented"))))

(if (boundp s) ; It is a variable.

                (princ
                 (format "%s\t%s\n%s\n\n" s
                   (if (user-variable-p s)
                       "Option " "Variable")
                   (or (documentation-property
                         s 'variable-documentation)
                       "not documented")))))))
        sym-list)

;; Build a list of symbols that match pattern.

    (mapatoms (function
               (lambda (sym)
                 (if (string-match pattern (symbol-name sym))
                     (setq sym-list (cons sym sym-list))))))

;; Display the data.

    (with-output-to-temp-buffer "*Help*"
      (mapcar describe-func (sort sym-list 'string<))
      (print-help-return-message))))

The describe-symbols function works like apropos, but provides more information.

(describe-symbols "goal")
---------- Buffer: *Help* ----------
goal-column     Option
*Semipermanent goal column for vertical motion, as set by C-x C-n, or nil.
set-goal-column Command: C-x C-n
Set the current horizontal position as a goal for C-n and C-p.
Those commands will move to this position in the line moved to
rather than trying to keep the same horizontal position.
With a non-nil argument, clears out the goal column
so that C-n and C-p resume vertical motion.
The goal column is stored in the variable `goal-column'.

temporary-goal-column   Variable
Current goal column for vertical motion.
It is the column where point was
at the start of current run of vertical motion commands.
When the `track-eol' feature is doing its job, the value is 9999.
---------- Buffer: *Help* ----------

Function Snarf-documentation filename

This function is used only during XEmacs initialization, just before the runnable XEmacs is dumped. It finds the file offsets of the documentation strings stored in the file filename, and records them in the in-core function definitions and variable property lists in place of the actual strings. See Section B.1 [Building XEmacs], page 765. XEmacs finds the file filename in the ‘lib-src’ directory. When the dumped XEmacs is later executed, the same file is found in the directory doc-directory. The usual value for filename is ‘DOC’, but this can be changed by modifying the

variable internal-doc-file-name.

Variable internal-doc-file-name

This variable holds the name of the file containing documentation strings of built-in symbols, usually ‘DOC’. The full pathname of the internal doc file is

‘(concat doc-directory internal-doc-file-name)’.

Variable doc-directory

This variable holds the name of the directory which contains the internal doc file that contains documentation strings for built-in and preloaded functions and variables. In most cases, this is the same as exec-directory. They may be different when you run XEmacs from the directory where you built it, without actually installing it. See exec-directory in Section 27.5 [Help Functions], page 387. In older Emacs versions, exec-directory was used for this.

Variable data-directory

This variable holds the name of the directory in which XEmacs finds certain system independent documentation and text files that come with XEmacs. In older Emacs versions, exec-directory was used for this.

27.3 Substituting Key Bindings in Documentation

When documentation strings refer to key sequences, they should use the current, actual key bindings. They can do so using certain special text sequences described below. Accessing documentation strings in the usual way substitutes current key binding information for these special sequences. This works by calling substitute-command-keys. You can also call that function yourself.

Here is a list of the special sequences and what they mean:

\[command] stands for a key sequence that will invoke command, or ‘M-x command’ if command has no key bindings.

\{mapvar} stands for a summary of the value of mapvar, which should be a keymap. The

     summary is made by describe-bindings.

\<mapvar> stands for no text itself. It is used for a side effect: it specifies mapvar as the keymap for any following ‘\[command]’ sequences in this documentation string.

\= quotes the following character and is discarded; this ‘\=\=’ puts ‘\=’ into the output, and ‘\=\[’ puts ‘\[’ into the output.

Please note: Each ‘\’ must be doubled when written in a string in XEmacs Lisp.

Function substitute-command-keys string

This function scans string for the above special sequences and replaces them by what they stand for, returning the result as a string. This permits display of documentation that refers accurately to the user’s own customized key bindings.

Here are examples of the special sequences:

(substitute-command-keys
   "To abort recursive edit, type: \\[abort-recursive-edit]")
⇒ "To abort recursive edit, type: C-]"
(substitute-command-keys
   "The keys that are defined for the minibuffer here are:
  \\{minibuffer-local-must-match-map}")
⇒ "The keys that are defined for the minibuffer here are:
?               minibuffer-completion-help
SPC             minibuffer-complete-word
TAB             minibuffer-complete
LFD             minibuffer-complete-and-exit
RET             minibuffer-complete-and-exit
C-g             abort-recursive-edit
"
(substitute-command-keys
   "To abort a recursive edit from the minibuffer, type\
\\<minibuffer-local-must-match-map>\\[abort-recursive-edit].")
⇒ "To abort a recursive edit from the minibuffer, type C-g."

(substitute-command-keys
  "Substrings of the form \\=\\{MAPVAR} are replaced by summaries
\(made by describe-bindings) of the value of MAPVAR, taken as a keymap.
Substrings of the form \\=\\<MAPVAR> specify to use the value of MAPVAR
as the keymap for future \\=\\[COMMAND] substrings.
\\=\\= quotes the following character and is discarded;
thus, \\=\\=\\=\\= puts \\=\\= into the output,
and \\=\\=\\=\\[ puts \\=\\[ into the output.")
⇒ "Substrings of the form \{MAPVAR} are replaced by summaries
(made by describe-bindings) of the value of MAPVAR, taken as a keymap.
Substrings of the form \<MAPVAR> specify to use the value of MAPVAR
as the keymap for future \[COMMAND] substrings.
\= quotes the following character and is discarded;
thus, \=\= puts \= into the output,
and \=\[ puts \[ into the output."

27.4 Describing Characters for Help Messages

These functions convert events, key sequences or characters to textual descriptions. These descriptions are useful for including arbitrary text characters or key sequences in messages, because they convert non-printing and whitespace characters to sequences of printing characters. The description of a non-whitespace printing character is the character itself.

Function key-description sequence

This function returns a string containing the XEmacs standard notation for the input events in sequence. The argument sequence may be a string, vector or list. See Section 19.5 [Events], page 290, for more information about valid events. See also the examples for single-key-description, below.

Function single-key-description key

This function returns a string describing key in the standard XEmacs notation for keyboard input. A normal printing character appears as itself, but a control character turns into a string starting with ‘C-’, a meta character turns into a string starting with ‘M-’, and space, linefeed, etc. appear as ‘SPC’, ‘LFD’, etc. A symbol appears as the name of the symbol. An event that is a list appears as the name of the symbol in the car of the list.

     (single-key-description ?\C-x)
          ⇒ "C-x"
     (key-description "\C-x \M-y \n \t \r \f123")
          ⇒ "C-x SPC M-y SPC LFD SPC TAB SPC RET SPC C-l 1 2 3"
     (single-key-description 'kp_next)
          ⇒ "kp_next"
     (single-key-description '(shift button1))
          ⇒ "Sh-button1"

Function text-char-description character

This function returns a string describing character in the standard XEmacs notation for characters that appear in text—like single-key-description, except that control characters are represented with a leading caret (which is how control characters in XEmacs buffers are usually displayed).

     (text-char-description ?\C-c)
          ⇒ "^C"
     (text-char-description ?\M-m)
          ⇒ "M-m"
     (text-char-description ?\C-\M-m)
          ⇒ "M-^M"

27.5 Help Functions

XEmacs provides a variety of on-line help functions, all accessible to the user as subcommands of the prefix C-h, or on some keyboards, help. For more information about them, see section “Help” in The XEmacs Reference Manual. Here we describe some program-level interfaces to the same information.

Command apropos regexp &optional do-all predicate This function finds all symbols whose names contain a match for the regular expression regexp, and returns a list of them (see Section 37.2 [Regular Expressions], page 550). It also displays the symbols in a buffer named ‘*Help*’, each with a one-line description. If do-all is non-nil, then apropos also shows key bindings for the functions that are found. If predicate is non-nil, it should be a function to be called on each symbol that has matched regexp. Only symbols for which predicate returns a non-nil value are listed or displayed. In the first of the following examples, apropos finds all the symbols with names containing ‘exec’. In the second example, it finds and returns only those symbols that are also commands. (We don’t show the output that results in the ‘*Help*’ buffer.)

     (apropos "exec")
          ⇒ (Buffer-menu-execute command-execute exec-directory
         exec-path execute-extended-command execute-kbd-macro
         executing-kbd-macro executing-macro)
     (apropos "exec" nil 'commandp)
          ⇒ (Buffer-menu-execute execute-extended-command)

apropos is used by various user-level commands, such as C-h a (hyper-

apropos), a graphical front-end to apropos; and C-h A (command-apropos),

which does an apropos over only those functions which are user commands. command-apropos calls apropos, specifying a predicate to restrict the output to symbols that are commands. The call to apropos looks like this:

     (apropos string t 'commandp)

Variable help-map

The value of this variable is a local keymap for characters following the Help key, C-h.

Prefix Command help-command

This symbol is not a function; its function definition is actually the keymap known as help-map. It is defined in ‘help.el’ as follows:

     (define-key global-map "\C-h" 'help-command)
     (fset 'help-command help-map)

Function print-help-return-message &optional function

This function builds a string that explains how to restore the previous state of the windows after a help command. After building the message, it applies function to it if function is non-nil. Otherwise it calls message to display it in the echo area. This function expects to be called inside a with-output-to-temp-buffer special form, and expects standard-output to have the value bound by that special form. For an example of its use, see the long example in Section 27.2 [Accessing Documentation], page 382.

Variable help-char

The value of this variable is the help character—the character that XEmacs recognizes as meaning Help. By default, it is the character ‘?\^H’ (ASCII 8), which is C-h. When XEmacs reads this character, if help-form is non-nil Lisp expression, it evaluates that expression, and displays the result in a window if it is a string. help-char can be a character or a key description such as help or (meta h). Usually the value of help-form’s value is nil. Then the help character has no special meaning at the level of command input, and it becomes part of a key sequence in the normal way. The standard key binding of C-h is a prefix key for several general-purpose help features. The help character is special after prefix keys, too. If it has no binding as a subcommand of the prefix key, it runs describe-prefix-bindings, which displays a list of all the subcommands of the prefix key.

Variable help-form

If this variable is non-nil, its value is a form to evaluate whenever the character help-char is read. If evaluating the form produces a string, that string is displayed. A command that calls next-command-event or next-event probably should bind help-form to a non-nil expression while it does input. (The exception is when C-h is meaningful input.) Evaluating this expression should result in a string that explains what the input is for and how to enter it properly. Entry to the minibuffer binds this variable to the value of minibuffer-help-form (see Section 18.8 [Minibuffer Misc], page 278).

Variable prefix-help-command

This variable holds a function to print help for a prefix character. The function is called when the user types a prefix key followed by the help character, and the help character has no binding after that prefix. The variable’s default value

is describe-prefix-bindings.

Function describe-prefix-bindings

This function calls describe-bindings to display a list of all the subcommands of the prefix key of the most recent key sequence. The prefix described consists of all but the last event of that key sequence. (The last event is, presumably, the help character.)

The following two functions are found in the library ‘helper’. They are for modes that want to provide help without relinquishing control, such as the “electric” modes. You must load that library with (require 'helper) in order to use them. Their names begin with ‘Helper’ to distinguish them from the ordinary help functions.

Command Helper-describe-bindings

This command pops up a window displaying a help buffer containing a listing of all of the key bindings from both the local and global keymaps. It works by

calling describe-bindings.

Command Helper-help

This command provides help for the current mode. It prompts the user in the minibuffer with the message ‘Help (Type ? for further options)’, and then provides assistance in finding out what the key bindings are, and what the mode is intended for. It returns nil. This can be customized by changing the map Helper-help-map.

27.6 Obsoleteness

As you add functionality to a package, you may at times want to replace an older function with a new one. To preserve compatibility with existing code, the older function needs to still exist; but users of that function should be told to use the newer one instead. XEmacs Lisp lets you mark a function or variable as obsolete, and indicate what should be used instead.

Function make-obsolete function new

This function indicates that function is an obsolete function, and the function new should be used instead. The byte compiler will issue a warning to this effect when it encounters a usage of the older function, and the help system will also note this in the function’s documentation. new can also be a string (if there is not a single function with the same functionality any more), and should be a descriptive statement, such as "use foo or bar instead" or "this function is unnecessary".

Function make-obsolete-variable variable new

This is like make-obsolete but is for variables instead of functions.

Function define-obsolete-function-alias oldfun newfun

This function combines make-obsolete and define-function, declaring oldfun to be an obsolete variant of newfun and defining oldfun as an alias for newfun.

Function define-obsolete-variable-alias oldvar newvar

     This is like define-obsolete-function-alias but for variables.

Note that you should not normally put obsoleteness information explicitly in a function or variable’s doc string. The obsoleteness information that you specify using the above functions will be displayed whenever the doc string is displayed, and by adding it explicitly the result is redundancy.

Also, if an obsolete function is substantially the same as a newer one but is not actually an alias, you should consider omitting the doc string entirely (use a null string ‘""’ as the doc string). That way, the user is told about the obsoleteness and is forced to look at the documentation of the new function, making it more likely that he will use the new function.

Function function-obsoleteness-doc function

If function is obsolete, this function returns a string describing this. This is the message that is printed out during byte compilation or in the function’s documentation. If function is not obsolete, nil is returned.

Function variable-obsoleteness-doc variable

This is like function-obsoleteness-doc but for variables.

The obsoleteness information is stored internally by putting a property byte-obsolete-info (for functions) or byte-obsolete-variable (for variables) on the symbol that specifies the obsolete function or variable. For more information, see the implementation of make-

obsolete and make-obsolete-variable in ‘lisp/bytecomp/bytecomp-runtime.el’.

28 Files

In XEmacs, you can find, create, view, save, and otherwise work with files and file directories. This chapter describes most of the file-related functions of XEmacs Lisp, but a few others are described in Chapter 30 [Buffers], page 431, and those related to backups and auto-saving are described in Chapter 29 [Backups and Auto-Saving], page 421.

Many of the file functions take one or more arguments that are file names. A file name is actually a string. Most of these functions expand file name arguments using expandfile-name, so that ‘~’ is handled correctly, as are relative file names (including ‘../’). These functions don’t recognize environment variable substitutions such as ‘$HOME’. See Section 28.8.4 [File Name Expansion], page 409.

28.1 Visiting Files

Visiting a file means reading a file into a buffer. Once this is done, we say that the buffer is visiting that file, and call the file “the visited file” of the buffer.

A file and a buffer are two different things. A file is information recorded permanently in the computer (unless you delete it). A buffer, on the other hand, is information inside of XEmacs that will vanish at the end of the editing session (or when you kill the buffer). Usually, a buffer contains information that you have copied from a file; then we say the buffer is visiting that file. The copy in the buffer is what you modify with editing commands. Such changes to the buffer do not change the file; therefore, to make the changes permanent, you must save the buffer, which means copying the altered buffer contents back into the file.

In spite of the distinction between files and buffers, people often refer to a file when they mean a buffer and vice-versa. Indeed, we say, “I am editing a file,” rather than, “I am editing a buffer that I will soon save as a file of the same name.” Humans do not usually need to make the distinction explicit. When dealing with a computer program, however, it is good to keep the distinction in mind.

28.1.1 Functions for Visiting Files

This section describes the functions normally used to visit files. For historical reasons, these functions have names starting with ‘find-’ rather than ‘visit-’. See Section 30.4 [Buffer File Name], page 434, for functions and variables that access the visited file name of a buffer or that find an existing buffer by its visited file name.

In a Lisp program, if you want to look at the contents of a file but not alter it, the fastest way is to use insert-file-contents in a temporary buffer. Visiting the file is not necessary and takes longer. See Section 28.3 [Reading from Files], page 396.

Command find-file filename

This command selects a buffer visiting the file filename, using an existing buffer if there is one, and otherwise creating a new buffer and reading the file into it. It also returns that buffer.

The body of the find-file function is very simple and looks like this:

     (switch-to-buffer (find-file-noselect filename))

(See switch-to-buffer in Section 31.7 [Displaying Buffers], page 453.) When find-file is called interactively, it prompts for filename in the minibuffer.

Function find-file-noselect filename &optional nowarn This function is the guts of all the file-visiting functions. It finds or creates a buffer visiting the file filename, and returns it. It uses an existing buffer if there is one, and otherwise creates a new buffer and reads the file into it. You may make the buffer current or display it in a window if you wish, but this function does not do so. When find-file-noselect uses an existing buffer, it first verifies that the file has not changed since it was last visited or saved in that buffer. If the file has changed, then this function asks the user whether to reread the changed file. If the user says ‘yes’, any changes previously made in the buffer are lost. If find-file-noselect needs to create a buffer, and there is no file named filename, it displays the message ‘New file’ in the echo area, and leaves the buffer empty. If no-warn is non-nil, various warnings that XEmacs normally gives (e.g. if another buffer is already visiting filename but filename has been removed from disk since that buffer was created) are suppressed.

The find-file-noselect function calls after-find-file after reading the

file (see Section 28.1.2 [Subroutines of Visiting], page 393). That function sets the buffer major mode, parses local variables, warns the user if there exists an auto-save file more recent than the file just visited, and finishes by running the

functions in find-file-hooks.

The find-file-noselect function returns the buffer that is visiting the file filename.

     (find-file-noselect "/etc/fstab")
          ⇒ #<buffer fstab>

Command find-file-other-window filename

This command selects a buffer visiting the file filename, but does so in a window other than the selected window. It may use another existing window or split a window; see Section 31.7 [Displaying Buffers], page 453. When this command is called interactively, it prompts for filename.

Command find-file-read-only filename

This command selects a buffer visiting the file filename, like find-file, but it marks the buffer as read-only. See Section 30.7 [Read Only Buffers], page 438, for related functions and variables. When this command is called interactively, it prompts for filename.

Command view-file filename

This command visits filename in View mode, and displays it in a recursive edit, returning to the previous buffer when done. View mode is a mode that allows you to skim rapidly through the file but does not let you modify it. Entering View mode runs the normal hook view-mode-hook. See Section 26.4 [Hooks], page 378. When view-file is called interactively, it prompts for filename.

Variable find-file-hooks

The value of this variable is a list of functions to be called after a file is visited. The file’s local-variables specification (if any) will have been processed before the hooks are run. The buffer visiting the file is current when the hook functions are run. This variable works just like a normal hook, but we think that renaming it would not be advisable.

Variable find-file-not-found-hooks

The value of this variable is a list of functions to be called when find-file or

find-file-noselect is passed a nonexistent file name. find-file-noselect

calls these functions as soon as it detects a nonexistent file. It calls them in the order of the list, until one of them returns non-nil. buffer-file-name is already set up. This is not a normal hook because the values of the functions are used and they may not all be called.

28.1.2 Subroutines of Visiting

The find-file-noselect function uses the create-file-buffer and after-find-

file functions as subroutines. Sometimes it is useful to call them directly.

Function create-file-buffer filename

This function creates a suitably named buffer for visiting filename, and returns it. It uses filename (sans directory) as the name if that name is free; otherwise, it appends a string such as ‘<2>’ to get an unused name. See also Section 30.9 [Creating Buffers], page 440. Please note: create-file-buffer does not associate the new buffer with a file and does not select the buffer. It also does not use the default major mode.

     (create-file-buffer "foo")
          ⇒ #<buffer foo>
     (create-file-buffer "foo")
          ⇒ #<buffer foo<2>>
     (create-file-buffer "foo")
          ⇒ #<buffer foo<3>>
This function is used by find-file-noselect. It uses generate-new-buffer

(see Section 30.9 [Creating Buffers], page 440).

Function after-find-file &optional error warn noauto This function sets the buffer major mode, and parses local variables (see Section 26.1.3 [Auto Major Mode], page 367). It is called by find-file-noselect and by the default revert function (see Section 29.3 [Reverting], page 428).

If reading the file got an error because the file does not exist, but its directory does exist, the caller should pass a non-nil value for error. In that case, afterfind-file issues a warning: ‘(New File)’. For more serious errors, the caller should usually not call after-find-file. If warn is non-nil, then this function issues a warning if an auto-save file exists and is more recent than the visited file. If noauto is non-nil, then this function does not turn on auto-save mode; otherwise, it does. The last thing after-find-file does is call all the functions in find-file-

hooks.

28.2 Saving Buffers

When you edit a file in XEmacs, you are actually working on a buffer that is visiting that file—that is, the contents of the file are copied into the buffer and the copy is what you edit. Changes to the buffer do not change the file until you save the buffer, which means copying the contents of the buffer into the file.

Command save-buffer &optional backup-option This function saves the contents of the current buffer in its visited file if the buffer has been modified since it was last visited or saved. Otherwise it does nothing. save-buffer is responsible for making backup files. Normally, backup-option is nil, and save-buffer makes a backup file only if this is the first save since visiting the file. Other values for backup-option request the making of backup files in other circumstances:

• With an argument of 4 or 64, reflecting 1 or 3 C-u’s, the save-buffer function marks this version of the file to be backed up when the buffer is next saved.

• With an argument of 16 or 64, reflecting 2 or 3 C-u’s, the save-buffer function unconditionally backs up the previous version of the file before saving it.

Command save-some-buffers &optional save-silently-p exiting This command saves some modified file-visiting buffers. Normally it asks the user about each buffer. But if save-silently-p is non-nil, it saves all the filevisiting buffers without querying the user. The optional exiting argument, if non-nil, requests this function to offer also to save certain other buffers that are not visiting files. These are buffers that have a non-nil local value of buffer-offer-save. (A user who says yes to saving one of these is asked to specify a file name to use.) The save-buffers-kill-emacs function passes a non-nil value for this argument.

Variable buffer-offer-save

When this variable is non-nil in a buffer, XEmacs offers to save the buffer on exit even if the buffer is not visiting a file. The variable is automatically local in all buffers. Normally, Mail mode (used for editing outgoing mail) sets this to t.

Command write-file filename

This function writes the current buffer into file filename, makes the buffer visit that file, and marks it not modified. Then it renames the buffer based on filename, appending a string like ‘<2>’ if necessary to make a unique buffer name. It does most of this work by calling set-visited-file-name and save-

buffer.

Variable write-file-hooks

The value of this variable is a list of functions to be called before writing out a buffer to its visited file. If one of them returns non-nil, the file is considered already written and the rest of the functions are not called, nor is the usual code for writing the file executed. If a function in write-file-hooks returns non-nil, it is responsible for making a backup file (if that is appropriate). To do so, execute the following code:

     (or buffer-backed-up (backup-buffer))

You might wish to save the file modes value returned by backup-buffer and use that to set the mode bits of the file that you write. This is what save-buffer normally does. Even though this is not a normal hook, you can use add-hook and remove-hook to manipulate the list. See Section 26.4 [Hooks], page 378.

Variable local-write-file-hooks

This works just like write-file-hooks, but it is intended to be made local to particular buffers. It’s not a good idea to make write-file-hooks local to a buffer—use this variable instead. The variable is marked as a permanent local, so that changing the major mode does not alter a buffer-local value. This is convenient for packages that read “file” contents in special ways, and set up hooks to save the data in a corresponding way.

Variable write-contents-hooks

This works just like write-file-hooks, but it is intended for hooks that pertain to the contents of the file, as opposed to hooks that pertain to where the file came from. Such hooks are usually set up by major modes, as buffer-local bindings for this variable. Switching to a new major mode always resets this variable.

Variable after-save-hook

This normal hook runs after a buffer has been saved in its visited file.

Variable file-precious-flag

If this variable is non-nil, then save-buffer protects against I/O errors while saving by writing the new file to a temporary name instead of the name it is supposed to have, and then renaming it to the intended name after it is clear there are no errors. This procedure prevents problems such as a lack of disk space from resulting in an invalid file. As a side effect, backups are necessarily made by copying. See Section 29.1.2 [Rename or Copy], page 422. Yet, at the same time, saving a precious file always breaks all hard links between the file you save and other file names. Some modes set this variable non-nil locally in particular buffers.

User Option require-final-newline

This variable determines whether files may be written out that do not end with a newline. If the value of the variable is t, then save-buffer silently adds a newline at the end of the file whenever the buffer being saved does not already end in one. If the value of the variable is non-nil, but not t, then save-buffer asks the user whether to add a newline each time the case arises. If the value of the variable is nil, then save-buffer doesn’t add newlines at all. nil is the default value, but a few major modes set it to t in particular buffers.

28.3 Reading from Files

You can copy a file from the disk and insert it into a buffer using the insert-file-contents function. Don’t use the user-level command insert-file in a Lisp program, as that sets the mark.

Function insert-file-contents filename &optional visit beg end replace This function inserts the contents of file filename into the current buffer after point. It returns a list of the absolute file name and the length of the data inserted. An error is signaled if filename is not the name of a file that can be read. The function insert-file-contents checks the file contents against the defined file formats, and converts the file contents if appropriate. See Section 28.13 [Format Conversion], page 417. It also calls the functions in the list afterinsert-file-functions; see Section 36.18.5 [Saving Properties], page 543. If visit is non-nil, this function additionally marks the buffer as unmodified and sets up various fields in the buffer so that it is visiting the file filename: these include the buffer’s visited file name and its last save file modtime. This feature is used by find-file-noselect and you probably should not use it yourself. If beg and end are non-nil, they should be integers specifying the portion of the file to insert. In this case, visit must be nil. For example,

     (insert-file-contents filename nil 0 500)

inserts the first 500 characters of a file. If the argument replace is non-nil, it means to replace the contents of the buffer (actually, just the accessible portion) with the contents of the file. This is better than simply deleting the buffer contents and inserting the whole file, because (1) it preserves some marker positions and (2) it puts less data in the undo list.

If you want to pass a file name to another process so that another program can read the file, use the function file-local-copy; see Section 28.11 [Magic File Names], page 414.

28.4 Writing to Files

You can write the contents of a buffer, or part of a buffer, directly to a file on disk using the append-to-file and write-region functions. Don’t use these functions to write to files that are being visited; that could cause confusion in the mechanisms for visiting.

Command append-to-file start end filename This function appends the contents of the region delimited by start and end in the current buffer to the end of file filename. If that file does not exist, it is created. If that file exists it is overwritten. This function returns nil.

An error is signaled if filename specifies a nonwritable file, or a nonexistent file in a directory where files cannot be created.

Command write-region start end filename &optional append visit This function writes the region delimited by start and end in the current buffer into the file specified by filename.

If start is a string, then write-region writes or appends that string, rather than text from the buffer.

If append is non-nil, then the specified text is appended to the existing file contents (if any).

If visit is t, then XEmacs establishes an association between the buffer and the file: the buffer is then visiting that file. It also sets the last file modification time for the current buffer to filename’s modtime, and marks the buffer as not modified. This feature is used by save-buffer, but you probably should not use it yourself.

If visit is a string, it specifies the file name to visit. This way, you can write the data to one file (filename) while recording the buffer as visiting another file (visit). The argument visit is used in the echo area message and also for file locking; visit is stored in buffer-file-name. This feature is used to implement file-precious-flag; don’t use it yourself unless you really know what you’re doing.

The function write-region converts the data which it writes to the appropriate file formats specified by buffer-file-format. See Section 28.13 [Format Conversion], page 417. It also calls the functions in the list write-region-annotate-functions; see Section 36.18.5 [Saving Properties], page 543.

Normally, write-region displays a message ‘Wrote file filename’ in the echo area. If visit is neither t nor nil nor a string, then this message is inhibited. This feature is useful for programs that use files for internal purposes, files that the user does not need to know about.

28.5 File Locks

When two users edit the same file at the same time, they are likely to interfere with each other. XEmacs tries to prevent this situation from arising by recording a file lock when a file is being modified. XEmacs can then detect the first attempt to modify a buffer visiting a file that is locked by another XEmacs process, and ask the user what to do.

File locks do not work properly when multiple machines can share file systems, such as with NFS. Perhaps a better file locking system will be implemented in the future. When file locks do not work, it is possible for two users to make changes simultaneously, but XEmacs can still warn the user who saves second. Also, the detection of modification of a buffer visiting a file changed on disk catches some cases of simultaneous editing; see Section 30.6 [Modification Time], page 437.

Function file-locked-p &optional filename This function returns nil if the file filename is not locked by this XEmacs process. It returns t if it is locked by this XEmacs, and it returns the name of the user who has locked it if it is locked by someone else.

     (file-locked-p "foo")
          ⇒ nil

Function lock-buffer &optional filename This function locks the file filename, if the current buffer is modified. The argument filename defaults to the current buffer’s visited file. Nothing is done if the current buffer is not visiting a file, or is not modified.

Function unlock-buffer

This function unlocks the file being visited in the current buffer, if the buffer is modified. If the buffer is not modified, then the file should not be locked, so this function does nothing. It also does nothing if the current buffer is not visiting a file.

Function ask-user-about-lock file other-user

This function is called when the user tries to modify file, but it is locked by another user named other-user. The value it returns determines what happens next:

• A value of t says to grab the lock on the file. Then this user may edit the file and other-user loses the lock.

• A value of nil says to ignore the lock and let this user edit the file anyway.

• This function may instead signal a file-locked error, in which case the change that the user was about to make does not take place. The error message for this error looks like this:

error File is locked: file other-user where file is the name of the file and other-user is the name of the user who has locked the file.

The default definition of this function asks the user to choose what to do. If you wish, you can replace the ask-user-about-lock function with your own version that decides in another way. The code for its usual definition is in

‘userlock.el’.

28.6 Information about Files

The functions described in this section all operate on strings that designate file names. All the functions have names that begin with the word ‘file’. These functions all return information about actual files or directories, so their arguments must all exist as actual files or directories unless otherwise noted.

28.6.1 Testing Accessibility

These functions test for permission to access a file in specific ways.

Function file-exists-p filename

This function returns t if a file named filename appears to exist. This does not mean you can necessarily read the file, only that you can find out its attributes. (On Unix, this is true if the file exists and you have execute permission on the containing directories, regardless of the protection of the file itself.) If the file does not exist, or if fascist access control policies prevent you from finding the attributes of the file, this function returns nil.

Function file-readable-p filename

This function returns t if a file named filename exists and you can read it. It returns nil otherwise.

     (file-readable-p "files.texi")
          ⇒ t
     (file-exists-p "/usr/spool/mqueue")
          ⇒ t
     (file-readable-p "/usr/spool/mqueue")
          ⇒ nil

Function file-executable-p filename

This function returns t if a file named filename exists and you can execute it. It returns nil otherwise. If the file is a directory, execute permission means you can check the existence and attributes of files inside the directory, and open those files if their modes permit.

Function file-writable-p filename

This function returns t if the file filename can be written or created by you, and nil otherwise. A file is writable if the file exists and you can write it. It is creatable if it does not exist, but the specified directory does exist and you can write in that directory. In the third example below, ‘foo’ is not writable because the parent directory does not exist, even though the user could create such a directory.

     (file-writable-p "~/foo")
          ⇒ t

     (file-writable-p "/foo")
          ⇒ nil
     (file-writable-p "~/no-such-dir/foo")
          ⇒ nil

Function file-accessible-directory-p dirname

This function returns t if you have permission to open existing files in the directory whose name as a file is dirname; otherwise (or if there is no such directory), it returns nil. The value of dirname may be either a directory name or the file name of a directory. Example: after the following,

     (file-accessible-directory-p "/foo")
          ⇒ nil

we can deduce that any attempt to read a file in ‘/foo/’ will give an error.

Function file-ownership-preserved-p filename

This function returns t if deleting the file filename and then creating it anew would keep the file’s owner unchanged.

Function file-newer-than-file-p filename1 filename2

This function returns t if the file filename1 is newer than file filename2. If filename1 does not exist, it returns nil. If filename2 does not exist, it returns

t.

In the following example, assume that the file ‘aug-19’ was written on the 19th, ‘aug-20’ was written on the 20th, and the file ‘no-file’ doesn’t exist at all.

     (file-newer-than-file-p "aug-19" "aug-20")
          ⇒ nil
     (file-newer-than-file-p "aug-20" "aug-19")
          ⇒ t
     (file-newer-than-file-p "aug-19" "no-file")
          ⇒ t
     (file-newer-than-file-p "no-file" "aug-19")
          ⇒ nil

You can use file-attributes to get a file’s last modification time as a list of two numbers. See Section 28.6.4 [File Attributes], page 401.

28.6.2 Distinguishing Kinds of Files

This section describes how to distinguish various kinds of files, such as directories, symbolic links, and ordinary files.

Function file-symlink-p filename

If the file filename is a symbolic link, the file-symlink-p function returns the file name to which it is linked. This may be the name of a text file, a directory, or even another symbolic link, or it may be a nonexistent file name. If the file filename is not a symbolic link (or there is no such file), file-

symlink-p returns nil.

(file-symlink-p "foo")
     ⇒ nil
(file-symlink-p "sym-link")
     ⇒ "foo"
(file-symlink-p "sym-link2")
     ⇒ "sym-link"
(file-symlink-p "/bin")
     ⇒ "/pub/bin"

Function file-directory-p filename

This function returns t if filename is the name of an existing directory, nil otherwise.

(file-directory-p "~rms")
     ⇒ t
(file-directory-p "~rms/lewis/files.texi")
     ⇒ nil
(file-directory-p "~rms/lewis/no-such-file")
     ⇒ nil
(file-directory-p "$HOME")
     ⇒ nil
(file-directory-p
 (substitute-in-file-name "$HOME"))
     ⇒ t

Function file-regular-p filename

This function returns t if the file filename exists and is a regular file (not a directory, symbolic link, named pipe, terminal, or other I/O device).

28.6.3 Truenames

The truename of a file is the name that you get by following symbolic links until none remain, then expanding to get rid of ‘.’ and ‘..’ as components. Strictly speaking, a file need not have a unique truename; the number of distinct truenames a file has is equal to the number of hard links to the file. However, truenames are useful because they eliminate symbolic links as a cause of name variation.

Function file-truename filename &optional default The function file-truename returns the true name of the file filename. This is the name that you get by following symbolic links until none remain. If the filename is relative, default is the directory to start with. If default is nil or missing, the current buffer’s value of default-directory is used.

See Section 30.4 [Buffer File Name], page 434, for related information.

28.6.4 Other Information about Files

This section describes the functions for getting detailed information about a file, other than its contents. This information includes the mode bits that control access permission, the owner and group numbers, the number of names, the inode number, the size, and the times of access and modification.

Function file-modes filename

This function returns the mode bits of filename, as an integer. The mode bits are also called the file permissions, and they specify access control in the usual Unix fashion. If the low-order bit is 1, then the file is executable by all users, if the second-lowest-order bit is 1, then the file is writable by all users, etc. The highest value returnable is 4095 (7777 octal), meaning that everyone has read, write, and execute permission, that the suid bit is set for both others and group, and that the sticky bit is set.

     (file-modes "~/junk/diffs")

⇒ 492 ; Decimal integer.

     (format "%o" 492)

⇒ "754" ; Convert to octal.

     (set-file-modes "~/junk/diffs" 438)
          ⇒ nil
     (format "%o" 438)

⇒ "666" ; Convert to octal.

     % ls -l diffs
       -rw-rw-rw-  1 lewis 0 3063 Oct 30 16:00 diffs

Function file-nlinks filename

This functions returns the number of names (i.e., hard links) that file filename has. If the file does not exist, then this function returns nil. Note that symbolic links have no effect on this function, because they are not considered to be names of the files they link to.

     % ls -l foo*
     -rw-rw-rw-  2 rms       4 Aug 19 01:27 foo
     -rw-rw-rw-  2 rms       4 Aug 19 01:27 foo1
     (file-nlinks "foo")
          ⇒ 2
     (file-nlinks "doesnt-exist")
          ⇒ nil

Function file-attributes filename

This function returns a list of attributes of file filename. If the specified file cannot be opened, it returns nil. The elements of the list, in order, are:

0. t for a directory, a string for a symbolic link (the name linked to), or nil for a text file.

1. The number of names the file has. Alternate names, also known as hard links, can be created by using the add-name-to-file function (see Section 28.7 [Changing File Attributes], page 404).

2. The file’s uid.

3. The file’s gid.

4. The time of last access, as a list of two integers. The first integer has the high-order 16 bits of time, the second has the low 16 bits. (This is similar to the value of current-time; see Section 50.5 [Time of Day], page 702.)

5. The time of last modification as a list of two integers (as above).

6. The time of last status change as a list of two integers (as above).

7. The size of the file in bytes.

8. The file’s modes, as a string of ten letters or dashes, as in ‘ls -l’.

9. t if the file’s gid would change if file were deleted and recreated; nil otherwise.

10. The file’s inode number.

11. The file system number of the file system that the file is in. This element and the file’s inode number together give enough information to distinguish any two files on the system—no two files can have the same values for both of these numbers.

For example, here are the file attributes for ‘files.texi’:

(file-attributes "files.texi")
⇒ (nil
1
2235
75
(8489 20284)
(8489 20284)
(8489 20285)
14906
"-rw-rw-rw-"
nil
129500
-32252)

and here is how the result is interpreted:

nil is neither a directory nor a symbolic link.

1 has only one name (the name ‘files.texi’ in the current default directory).

2235 is owned by the user with uid 2235.

75 is in the group with gid 75.

(8489 20284)

was last accessed on Aug 19 00:09. Use format-time-string to ! convert this number into a time string. See Section 50.6 [Time Conversion], page 703.

(8489 20284)

was last modified on Aug 19 00:09.

(8489 20285)

last had its inode changed on Aug 19 00:09.

14906 is 14906 characters long.

"-rw-rw-rw-"

has a mode of read and write access for the owner, group, and world.

nil would retain the same gid if it were recreated.

129500 has an inode number of 129500.

-32252 is on file system number -32252.

28.7 Changing File Names and Attributes

The functions in this section rename, copy, delete, link, and set the modes of files. In the functions that have an argument newname, if a file by the name of newname

already exists, the actions taken depend on the value of the argument ok-if-already-exists: • Signal a file-already-exists error if ok-if-already-exists is nil.

• Request confirmation if ok-if-already-exists is a number.

• Replace the old file without confirmation if ok-if-already-exists is any other value.

Command add-name-to-file oldname newname &optional

ok-if-already-exists This function gives the file named oldname the additional name newname. This means that newname becomes a new “hard link” to oldname. In the first part of the following example, we list two files, ‘foo’ and ‘foo3’.

     % ls -l fo*
     -rw-rw-rw-  1 rms       29 Aug 18 20:32 foo
     -rw-rw-rw-  1 rms       24 Aug 18 20:31 foo3
Then we evaluate the form (add-name-to-file "~/lewis/foo" "~/lewis/foo2").

Again we list the files. This shows two names, ‘foo’ and ‘foo2’.

     (add-name-to-file "~/lewis/foo1" "~/lewis/foo2")
          ⇒ nil
     % ls -l fo*
     -rw-rw-rw-  2 rms       29 Aug 18 20:32 foo
     -rw-rw-rw-  2 rms       29 Aug 18 20:32 foo2
     -rw-rw-rw-  1 rms       24 Aug 18 20:31 foo3

Finally, we evaluate the following:

     (add-name-to-file "~/lewis/foo" "~/lewis/foo3" t)

and list the files again. Now there are three names for one file: ‘foo’, ‘foo2’, and ‘foo3’. The old contents of ‘foo3’ are lost.

     (add-name-to-file "~/lewis/foo1" "~/lewis/foo3")
          ⇒ nil

     % ls -l fo*
     -rw-rw-rw-  3 rms       29 Aug 18 20:32 foo
     -rw-rw-rw-  3 rms       29 Aug 18 20:32 foo2
     -rw-rw-rw-  3 rms       29 Aug 18 20:32 foo3

This function is meaningless on VMS, where multiple names for one file are not allowed. See also file-nlinks in Section 28.6.4 [File Attributes], page 401.

Command rename-file filename newname &optional ok-if-already-exists This command renames the file filename as newname. If filename has additional names aside from filename, it continues to have those names. In fact, adding the name newname with add-name-to-file and then deleting filename has the same effect as renaming, aside from momentary intermediate states. In an interactive call, this function prompts for filename and newname in the minibuffer; also, it requests confirmation if newname already exists.

Command copy-file oldname newname &optional ok-if-exists time This command copies the file oldname to newname. An error is signaled if oldname does not exist. If time is non-nil, then this functions gives the new file the same last-modified time that the old one has. (This works on only some operating systems.) In an interactive call, this function prompts for filename and newname in the minibuffer; also, it requests confirmation if newname already exists.

Command delete-file filename

This command deletes the file filename, like the shell command ‘rm filename’. If the file has multiple names, it continues to exist under the other names. A suitable kind of file-error error is signaled if the file does not exist, or is not deletable. (On Unix, a file is deletable if its directory is writable.) See also delete-directory in Section 28.10 [Create/Delete Dirs], page 413.

Command make-symbolic-link filename newname &optional ok-if-exists This command makes a symbolic link to filename, named newname. This is like the shell command ‘ln -s filename newname’. In an interactive call, this function prompts for filename and newname in the minibuffer; also, it requests confirmation if newname already exists.

Function define-logical-name varname string

This function defines the logical name name to have the value string. It is available only on VMS.

Function set-file-modes filename mode

This function sets mode bits of filename to mode (which must be an integer). Only the low 12 bits of mode are used.

Function set-default-file-modes mode

This function sets the default file protection for new files created by XEmacs and its subprocesses. Every file created with XEmacs initially has this protection. On Unix, the default protection is the bitwise complement of the “umask” value.

The argument mode must be an integer. Only the low 9 bits of mode are used.

Saving a modified version of an existing file does not count as creating the file; it does not change the file’s mode, and does not use the default file protection.

Function default-file-modes

This function returns the current default protection value.

On MS-DOS, there is no such thing as an “executable” file mode bit. So Emacs considers a file executable if its name ends in ‘.com’, ‘.bat’ or ‘.exe’. This is reflected in the values

returned by file-modes and file-attributes.

28.8 File Names

Files are generally referred to by their names, in XEmacs as elsewhere. File names in XEmacs are represented as strings. The functions that operate on a file all expect a file name argument.

In addition to operating on files themselves, XEmacs Lisp programs often need to operate on the names; i.e., to take them apart and to use part of a name to construct related file names. This section describes how to manipulate file names.

The functions in this section do not actually access files, so they can operate on file names that do not refer to an existing file or directory.

On VMS, all these functions understand both VMS file-name syntax and Unix syntax. This is so that all the standard Lisp libraries can specify file names in Unix syntax and work properly on VMS without change. On MS-DOS, these functions understand MS-DOS file-name syntax as well as Unix syntax.

28.8.1 File Name Components

The operating system groups files into directories. To specify a file, you must specify the directory and the file’s name within that directory. Therefore, XEmacs considers a file name as having two main parts: the directory name part, and the nondirectory part (or file name within the directory). Either part may be empty. Concatenating these two parts reproduces the original file name.

On Unix, the directory part is everything up to and including the last slash; the nondirectory part is the rest. The rules in VMS syntax are complicated.

For some purposes, the nondirectory part is further subdivided into the name proper and the version number. On Unix, only backup files have version numbers in their names; on VMS, every file has a version number, but most of the time the file name actually used in XEmacs omits the version number. Version numbers are found mostly in directory lists.

Function file-name-directory filename

This function returns the directory part of filename (or nil if filename does not include a directory part). On Unix, the function returns a string ending in a slash. On VMS, it returns a string ending in one of the three characters ‘:’, ‘]’, or ‘>’.

(file-name-directory "lewis/foo")  ; Unix example
     ⇒ "lewis/"
(file-name-directory "foo")        ; Unix example
     ⇒ nil
(file-name-directory "[X]FOO.TMP") ; VMS example
     ⇒ "[X]"

Function file-name-nondirectory filename

This function returns the nondirectory part of filename.

(file-name-nondirectory "lewis/foo")
     ⇒ "foo"
(file-name-nondirectory "foo")
     ⇒ "foo"

;; The following example is accurate only on VMS.

(file-name-nondirectory "[X]FOO.TMP")
     ⇒ "FOO.TMP"

Function file-name-sans-versions filename &optional keep-backup-version This function returns filename without any file version numbers, backup version numbers, or trailing tildes.

If keep-backup-version is non-nil, we do not remove backup version numbers, only true file version numbers.

(file-name-sans-versions "~rms/foo.~1~")
     ⇒ "~rms/foo"
(file-name-sans-versions "~rms/foo~")
     ⇒ "~rms/foo"
(file-name-sans-versions "~rms/foo")
     ⇒ "~rms/foo"

;; The following example applies to VMS only.

(file-name-sans-versions "foo;23")
     ⇒ "foo"

Function file-name-sans-extension filename

This function returns filename minus its “extension,” if any. The extension, in a file name, is the part that starts with the last ‘.’ in the last name component. For example,

(file-name-sans-extension "foo.lose.c")
     ⇒ "foo.lose"
(file-name-sans-extension "big.hack/foo")
     ⇒ "big.hack/foo"

28.8.2 Directory Names

A directory name is the name of a directory. A directory is a kind of file, and it has a file name, which is related to the directory name but not identical to it. (This is not quite the same as the usual Unix terminology.) These two different names for the same entity are related by a syntactic transformation. On Unix, this is simple: a directory name ends in a slash, whereas the directory’s name as a file lacks that slash. On VMS, the relationship is more complicated.

The difference between a directory name and its name as a file is subtle but crucial. When an XEmacs variable or function argument is described as being a directory name, a file name of a directory is not acceptable.

The following two functions convert between directory names and file names. They do nothing special with environment variable substitutions such as ‘$HOME’, and the constructs ‘~’, and ‘..’.

Function file-name-as-directory filename

This function returns a string representing filename in a form that the operating system will interpret as the name of a directory. In Unix, this means appending a slash to the string. On VMS, the function converts a string of the form

     ‘[X]Y.DIR.1’ to the form ‘[X.Y]’.
          (file-name-as-directory "~rms/lewis")
               ⇒ "~rms/lewis/"

Function directory-file-name dirname

This function returns a string representing dirname in a form that the operating system will interpret as the name of a file. On Unix, this means removing a final slash from the string. On VMS, the function converts a string of the form

     ‘[X.Y]’ to ‘[X]Y.DIR.1’.
          (directory-file-name "~lewis/")
               ⇒ "~lewis"

Directory name abbreviations are useful for directories that are normally accessed through symbolic links. Sometimes the users recognize primarily the link’s name as “the name” of the directory, and find it annoying to see the directory’s “real” name. If you define the link name as an abbreviation for the “real” name, XEmacs shows users the abbreviation instead.

If you wish to convert a directory name to its abbreviation, use this function:

Function abbreviate-file-name dirname &optional hack-homedir This function applies abbreviations from directory-abbrev-alist to its argument, and substitutes ‘~’ for the user’s home directory. If hack-homedir is non-nil, then this also substitutes ‘~’ for the user’s home directory.

Variable directory-abbrev-alist

The variable directory-abbrev-alist contains an alist of abbreviations to use for file directories. Each element has the form (from . to), and says to replace from with to when it appears in a directory name. The from string is actually a regular expression; it should always start with ‘^’. The function abbreviate-file-name performs these substitutions. You can set this variable in ‘site-init.el’ to describe the abbreviations appropriate for your site. Here’s an example, from a system on which file system ‘/home/fsf’ and so on are normally accessed through symbolic links named ‘/fsf’ and so on.

     (("^/home/fsf" . "/fsf")
      ("^/home/gp" . "/gp")
      ("^/home/gd" . "/gd"))

28.8.3 Absolute and Relative File Names

All the directories in the file system form a tree starting at the root directory. A file name can specify all the directory names starting from the root of the tree; then it is called an absolute file name. Or it can specify the position of the file in the tree relative to a default directory; then it is called a relative file name. On Unix, an absolute file name starts with a slash or a tilde (‘~’), and a relative one does not. The rules on VMS are complicated.

Function file-name-absolute-p filename

This function returns t if file filename is an absolute file name, nil otherwise. On VMS, this function understands both Unix syntax and VMS syntax.

     (file-name-absolute-p "~rms/foo")
          ⇒ t
     (file-name-absolute-p "rms/foo")
          ⇒ nil
     (file-name-absolute-p "/user/rms/foo")
          ⇒ t

28.8.4 Functions that Expand Filenames

Expansion of a file name means converting a relative file name to an absolute one. Since this is done relative to a default directory, you must specify the default directory name as well as the file name to be expanded. Expansion also simplifies file names by eliminating redundancies such as ‘./’ and ‘name/../’.

Function expand-file-name filename &optional directory This function converts filename to an absolute file name. If directory is supplied, it is the directory to start with if filename is relative. (The value of directory should itself be an absolute directory name; it may start with ‘~’.) Otherwise, the current buffer’s value of default-directory is used. For example:

     (expand-file-name "foo")
          ⇒ "/xcssun/users/rms/lewis/foo"
     (expand-file-name "../foo")
          ⇒ "/xcssun/users/rms/foo"

     (expand-file-name "foo" "/usr/spool/")
          ⇒ "/usr/spool/foo"
     (expand-file-name "$HOME/foo")
          ⇒ "/xcssun/users/rms/lewis/$HOME/foo"

Filenames containing ‘.’ or ‘..’ are simplified to their canonical form:

     (expand-file-name "bar/../foo")
          ⇒ "/xcssun/users/rms/lewis/foo"

‘~/’ at the beginning is expanded into the user’s home directory. A ‘/’ or ‘~’ following a ‘/’. Note that expand-file-name does not expand environment variables; only

substitute-in-file-name does that.

Function file-relative-name filename &optional directory This function does the inverse of expansion—it tries to return a relative name that is equivalent to filename when interpreted relative to directory. If directory is nil or omitted, the value of default-directory is used.

     (file-relative-name "/foo/bar" "/foo/")
          ⇒ "bar")
     (file-relative-name "/foo/bar" "/hack/")
          ⇒ "../foo/bar")

Variable default-directory

The value of this buffer-local variable is the default directory for the current buffer. It should be an absolute directory name; it may start with ‘~’. This variable is local in every buffer. expand-file-name uses the default directory when its second argument is nil. On Unix systems, the value is always a string ending with a slash.

     default-directory
          ⇒ "/user/lewis/manual/"

Function substitute-in-file-name filename

This function replaces environment variable references in filename with the environment variable values. Following standard Unix shell syntax, ‘$’ is the prefix to substitute an environment variable value. The environment variable name is the series of alphanumeric characters (including underscores) that follow the ‘$’. If the character following the ‘$’ is a ‘{’, then the variable name is everything up to the matching ‘}’. Here we assume that the environment variable HOME, which holds the user’s home directory name, has value ‘/xcssun/users/rms’.

     (substitute-in-file-name "$HOME/foo")
          ⇒ "/xcssun/users/rms/foo"

After substitution, a ‘/’ or ‘~’ following a ‘/’ is taken to be the start of an absolute file name that overrides what precedes it, so everything before that ‘/’ or ‘~’ is deleted. For example:

     (substitute-in-file-name "bar/~/foo")
          ⇒ "~/foo"
     (substitute-in-file-name "/usr/local/$HOME/foo")
          ⇒ "/xcssun/users/rms/foo"

On VMS, ‘$’ substitution is not done, so this function does nothing on VMS except discard superfluous initial components as shown above.

28.8.5 Generating Unique File Names

Some programs need to write temporary files. Here is the usual way to construct a name for such a file:

(make-temp-name (expand-file-name name-of-application (temp-directory)))

Here we use (temp-directory) to specify a directory for temporary files—under Unix, it will normally evaluate to ‘"/tmp/"’. The job of make-temp-name is to prevent two different users or two different processes from trying to use the same name.

Function temp-directory

This function returns the name of the directory to use for temporary files. Under Unix, this will be the value of TMPDIR, defaulting to ‘/tmp’. On Windows, this will be obtained from the TEMP or TMP environment variables, defaulting to ‘/’. Note that the temp-directory function does not exist under FSF Emacs.

Function make-temp-name prefix

This function generates a temporary file name starting with prefix. The Emacs process number forms part of the result, so there is no danger of generating a name being used by another process.

     (make-temp-name "/tmp/foo")
          ⇒ "/tmp/fooGaAQjC"

In addition, this function makes an attempt to choose a name that does not specify an existing file. To make this work, prefix should be an absolute file name. To avoid confusion, each Lisp application should preferably use a unique prefix

to make-temp-name.

28.8.6 File Name Completion

This section describes low-level subroutines for completing a file name. For other completion functions, see Section 18.5 [Completion], page 266.

Function file-name-all-completions partial-filename directory

This function returns a list of all possible completions for a file whose name starts with partial-filename in directory directory. The order of the completions is the order of the files in the directory, which is unpredictable and conveys no useful information. The argument partial-filename must be a file name containing no directory part and no slash. The current buffer’s default directory is prepended to directory, if directory is not absolute.

In the following example, suppose that the current default directory, ‘~rms/lewis’, has five files whose names begin with ‘f’: ‘foo’, ‘file~’, ‘file.c’, ‘file.c.~1~’,

and ‘file.c.~2~’.
     (file-name-all-completions "f" "")
          ⇒ ("foo" "file~" "file.c.~2~"
                     "file.c.~1~" "file.c")
     (file-name-all-completions "fo" "")
          ⇒ ("foo")

Function file-name-completion filename directory

This function completes the file name filename in directory directory. It returns the longest prefix common to all file names in directory directory that start with filename.

If only one match exists and filename matches it exactly, the function returns t. The function returns nil if directory directory contains no name starting with filename.

In the following example, suppose that the current default directory has five files whose names begin with ‘f’: ‘foo’, ‘file~’, ‘file.c’, ‘file.c.~1~’, and

‘file.c.~2~’.
     (file-name-completion "fi" "")
          ⇒ "file"
     (file-name-completion "file.c.~1" "")
          ⇒ "file.c.~1~"
     (file-name-completion "file.c.~1~" "")
          ⇒ t
     (file-name-completion "file.c.~3" "")
          ⇒ nil

User Option completion-ignored-extensions

file-name-completion usually ignores file names that end in any string in this list. It does not ignore them when all the possible completions end in one of these suffixes or when a buffer showing all possible completions is displayed.

A typical value might look like this:

     completion-ignored-extensions
          ⇒ (".o" ".elc" "~" ".dvi")

28.9 Contents of Directories

A directory is a kind of file that contains other files entered under various names. Directories are a feature of the file system.

XEmacs can list the names of the files in a directory as a Lisp list, or display the names in a buffer using the ls shell command. In the latter case, it can optionally display information about each file, depending on the value of switches passed to the ls command.

Function directory-files directory &optional full-name match-regexp nosort

files-only This function returns a list of the names of the files in the directory directory. By default, the list is in alphabetical order. If full-name is non-nil, the function returns the files’ absolute file names. Otherwise, it returns just the names relative to the specified directory. If match-regexp is non-nil, this function returns only those file names that contain that regular expression—the other file names are discarded from the list. If nosort is non-nil, directory-files does not sort the list, so you get the file names in no particular order. Use this if you want the utmost possible speed and don’t care what order the files are processed in. If the order of processing is visible to the user, then the user will probably be happier if you do sort the names. If files-only is the symbol t, then only the “files” in the directory will be returned; subdirectories will be excluded. If files-only is not nil and not t, then only the subdirectories will be returned. Otherwise, if files-only is nil (the default) then both files and subdirectories will be returned.

     (directory-files "~lewis")
          ⇒ ("#foo#" "#foo.el#" "." ".."
              "dired-mods.el" "files.texi"
              "files.texi.~1~")

An error is signaled if directory is not the name of a directory that can be read.

Function insert-directory file switches &optional wildcard full-directory-p This function inserts (in the current buffer) a directory listing for directory file, formatted with ls according to switches. It leaves point after the inserted text. The argument file may be either a directory name or a file specification including wildcard characters. If wildcard is non-nil, that means treat file as a file specification with wildcards. If full-directory-p is non-nil, that means file is a directory and switches do not contain ‘-d’, so that the listing should show the full contents of the directory. (The ‘-d’ option to ls says to describe a directory itself rather than its contents.) This function works by running a directory listing program whose name is in the variable insert-directory-program. If wildcard is non-nil, it also runs the shell specified by shell-file-name, to expand the wildcards.

Variable insert-directory-program

This variable’s value is the program to run to generate a directory listing for

the function insert-directory.

28.10 Creating and Deleting Directories

Most XEmacs Lisp file-manipulation functions get errors when used on files that are directories. For example, you cannot delete a directory with delete-file. These special functions exist to create and delete directories.

Command make-directory dirname &optional parents This function creates a directory named dirname. Interactively, the default choice of directory to create is the current default directory for file names. That is useful when you have visited a file in a nonexistent directory. Non-interactively, optional argument parents says whether to create parent directories if they don’t exist. (Interactively, this always happens.)

Command delete-directory dirname

This function deletes the directory named dirname. The function delete-file does not work for files that are directories; you must use delete-directory in that case.

28.11 Making Certain File Names “Magic”

You can implement special handling for certain file names. This is called making those names magic. You must supply a regular expression to define the class of names (all those that match the regular expression), plus a handler that implements all the primitive XEmacs file operations for file names that do match.

The variable file-name-handler-alist holds a list of handlers, together with regular expressions that determine when to apply each handler. Each element has this form:

(regexp . handler) All the XEmacs primitives for file access and file name transformation check the given file name against file-name-handler-alist. If the file name matches regexp, the primitives handle that file by calling handler.

The first argument given to handler is the name of the primitive; the remaining arguments are the arguments that were passed to that operation. (The first of these arguments is typically the file name itself.) For example, if you do this:

     (file-exists-p filename)

and filename has handler handler, then handler is called like this:

     (funcall handler 'file-exists-p filename)

Here are the operations that a magic file name handler gets to handle:

add-name-to-file, copy-file, delete-directory, delete-file,
diff-latest-backup-file, directory-file-name, directory-files, dired-compress-
file, dired-uncache, expand-file-name,
file-accessible-directory-p, file-attributes, file-directory-p, file-executable-
p, file-exists-p, file-local-copy, file-modes, file-name-all-completions, file-
name-as-directory, file-name-completion, file-name-directory, file-name-nondirectory,
file-name-sans-versions, file-newer-than-file-p, file-readable-p, file-regular-
p, file-symlink-p, file-truename, file-writable-p, get-file-buffer, insert-
directory, insert-file-contents, load, make-directory, make-symbolic-link, rename-
file, set-file-modes, set-visited-file-modtime, unhandled-file-name-directory,
verify-visited-file-modtime, write-region.

Handlers for insert-file-contents typically need to clear the buffer’s modified flag, with (set-buffer-modified-p nil), if the visit argument is non-nil. This also has the effect of unlocking the buffer if it is locked.

The handler function must handle all of the above operations, and possibly others to be added in the future. It need not implement all these operations itself—when it has nothing special to do for a certain operation, it can reinvoke the primitive, to handle the operation “in the usual way”. It should always reinvoke the primitive for an operation it does not recognize. Here’s one way to do this:

(defun my-file-handler (operation &rest args)

;; First check for the specific operations

;; that we have special handling for.

  (cond ((eq operation 'insert-file-contents) ...)
        ((eq operation 'write-region) ...)
        ...

;; Handle any operation we don’t know about.

        (t (let ((inhibit-file-name-handlers
                 (cons 'my-file-handler
                       (and (eq inhibit-file-name-operation operation)
                            inhibit-file-name-handlers)))
                (inhibit-file-name-operation operation))
             (apply operation args)))))

When a handler function decides to call the ordinary Emacs primitive for the operation at hand, it needs to prevent the primitive from calling the same handler once again, thus leading to an infinite recursion. The example above shows how to do this, with the vari-

ables inhibit-file-name-handlers and inhibit-file-name-operation. Be careful to

use them exactly as shown above; the details are crucial for proper behavior in the case of multiple handlers, and for operations that have two file names that may each have handlers.

Variable inhibit-file-name-handlers

This variable holds a list of handlers whose use is presently inhibited for a certain operation.

Variable inhibit-file-name-operation

The operation for which certain handlers are presently inhibited.

Function find-file-name-handler file operation

This function returns the handler function for file name file, or nil if there is none. The argument operation should be the operation to be performed on the file—the value you will pass to the handler as its first argument when you call it. The operation is needed for comparison with inhibit-file-name-operation.

Function file-local-copy filename

This function copies file filename to an ordinary non-magic file, if it isn’t one already.

If filename specifies a “magic” file name, which programs outside Emacs cannot directly read or write, this copies the contents to an ordinary file and returns that file’s name.

If filename is an ordinary file name, not magic, then this function does nothing and returns nil.

Function unhandled-file-name-directory filename

This function returns the name of a directory that is not magic. It uses the directory part of filename if that is not magic. Otherwise, it asks the handler what to do. This is useful for running a subprocess; every subprocess must have a non-magic directory to serve as its current directory, and this function is a good way to come up with one.

28.12 Partial Files

28.12.1 Intro to Partial Files

A partial file is a section of a buffer (called the master buffer) that is placed in its own buffer and treated as its own file. Changes made to the partial file are not reflected in the master buffer until the partial file is “saved” using the standard buffer save commands. Partial files can be “reverted” (from the master buffer) just like normal files. When a file part is active on a master buffer, that section of the master buffer is marked as read-only. Two file parts on the same master buffer are not allowed to overlap. Partial file buffers are indicated by the words ‘File Part’ in the modeline.

The master buffer knows about all the partial files that are active on it, and thus killing or reverting the master buffer will be handled properly. When the master buffer is saved, if there are any unsaved partial files active on it then the user will be given the opportunity to first save these files.

When a partial file buffer is first modified, the master buffer is automatically marked as modified so that saving the master buffer will work correctly.

28.12.2 Creating a Partial File

Function make-file-part &optional start end name buffer Make a file part on buffer buffer out of the region. Call it name. This command creates a new buffer containing the contents of the region and marks the buffer as referring to the specified buffer, called the master buffer. When the file-part buffer is saved, its changes are integrated back into the master buffer. When the master buffer is deleted, all file parts are deleted with it. When called from a function, expects four arguments, start, end, name, and buffer, all of which are optional and default to the beginning of buffer, the end of buffer, a name generated from buffer name, and the current buffer, respectively.

28.12.3 Detached Partial Files

Every partial file has an extent in the master buffer associated with it (called the master extent), marking where in the master buffer the partial file begins and ends. If the text in master buffer that is contained by the extent is deleted, then the extent becomes “detached”, meaning that it no longer refers to a specific region of the master buffer. This can happen either when the text is deleted directly or when the master buffer is reverted. Neither of these should happen in normal usage because the master buffer should generally not be edited directly.

Before doing any operation that references a partial file’s master extent, XEmacs checks to make sure that the extent is not detached. If this is the case, XEmacs warns the user of this and the master extent is deleted out of the master buffer, disconnecting the file part. The file part’s filename is cleared and thus must be explicitly specified if the detached file part is to be saved.

28.13 File Format Conversion

The variable format-alist defines a list of file formats, which describe textual representations used in files for the data (text, text-properties, and possibly other information) in an Emacs buffer. Emacs performs format conversion if appropriate when reading and writing files.

Variable format-alist

This list contains one format definition for each defined file format.

Each format definition is a list of this form:

(name doc-string regexp from-fn to-fn modify mode-fn) Here is what the elements in a format definition mean:

name The name of this format.

doc-string A documentation string for the format.

regexp A regular expression which is used to recognize files represented in this format.

from-fn A function to call to decode data in this format (to convert file data into the usual Emacs data representation). The from-fn is called with two args, begin and end, which specify the part of the buffer it should convert. It should convert the text by editing it in place. Since this can change the length of the text, from-fn should return the modified end position. One responsibility of from-fn is to make sure that the beginning of the file no longer matches regexp. Otherwise it is likely to get called again.

to-fn A function to call to encode data in this format (to convert the usual Emacs data representation into this format). The to-fn is called with two args, begin and end, which specify the part of the buffer it should convert. There are two ways it can do the conversion:

• By editing the buffer in place. In this case, to-fn should return the endposition of the range of text, as modified.

• By returning a list of annotations. This is a list of elements of the form (position . string), where position is an integer specifying the relative position in the text to be written, and string is the annotation to add there. The list must be sorted in order of position when to-fn returns it.

When write-region actually writes the text from the buffer to the file, it intermixes the specified annotations at the corresponding positions. All this takes place without modifying the buffer.

modify A flag, t if the encoding function modifies the buffer, and nil if it works by returning a list of annotations.

mode A mode function to call after visiting a file converted from this format.

The function insert-file-contents automatically recognizes file formats when it reads the specified file. It checks the text of the beginning of the file against the regular expressions of the format definitions, and if it finds a match, it calls the decoding function for that format. Then it checks all the known formats over again. It keeps checking them until none of them is applicable.

Visiting a file, with find-file-noselect or the commands that use it, performs conversion likewise (because it calls insert-file-contents); it also calls the mode function for each format that it decodes. It stores a list of the format names in the buffer-local variable

buffer-file-format.

Variable buffer-file-format

This variable states the format of the visited file. More precisely, this is a list of the file format names that were decoded in the course of visiting the current buffer’s file. It is always local in all buffers.

When write-region writes data into a file, it first calls the encoding functions for the formats listed in buffer-file-format, in the order of appearance in the list.

Function format-write-file file format

This command writes the current buffer contents into the file file in format format, and makes that format the default for future saves of the buffer. The argument format is a list of format names.

Function format-find-file file format

This command finds the file file, converting it according to format format. It also makes format the default if the buffer is saved later.

The argument format is a list of format names. If format is nil, no conversion takes place. Interactively, typing just ^{⟨}\underline{\overline{RET}}^{⟩} for format specifies nil.

Function format-insert-file file format &optional beg end This command inserts the contents of file file, converting it according to format format. If beg and end are non-nil, they specify which part of the file to read, as in insert-file-contents (see Section 28.3 [Reading from Files], page 396).

The return value is like what insert-file-contents returns: a list of the absolute file name and the length of the data inserted (after conversion).

The argument format is a list of format names. If format is nil, no conversion takes place. Interactively, typing just ^{⟨}\underline{\overline{RET}}^{⟩} for format specifies nil.

Function format-find-file file format

This command finds the file file, converting it according to format format. It also makes format the default if the buffer is saved later. The argument format is a list of format names. If format is nil, no conversion takes place. Interactively, typing just ^{⟨}\underline{\overline{RET}}^{⟩} for format specifies nil.

Function format-insert-file file format &optional beg end This command inserts the contents of file file, converting it according to format format. If beg and end are non-nil, they specify which part of the file to read, as in insert-file-contents (see Section 28.3 [Reading from Files], page 396). The return value is like what insert-file-contents returns: a list of the absolute file name and the length of the data inserted (after conversion). The argument format is a list of format names. If format is nil, no conversion takes place. Interactively, typing just ^{⟨}\underline{\overline{RET}}^{⟩} for format specifies nil.

Variable auto-save-file-format

This variable specifies the format to use for auto-saving. Its value is a list of format names, just like the value of buffer-file-format; but it is used instead of buffer-file-format for writing auto-save files. This variable is always local in all buffers.

28.14 Files and MS-DOS

Emacs on MS-DOS makes a distinction between text files and binary files. This is necessary because ordinary text files on MS-DOS use a two character sequence between lines: carriage-return and linefeed (crlf). Emacs expects just a newline character (a linefeed) between lines. When Emacs reads or writes a text file on MS-DOS, it needs to convert the line separators. This means it needs to know which files are text files and which are binary. It makes this decision when visiting a file, and records the decision in the variable buffer-file-type for use when the file is saved.

See Section 49.3 [MS-DOS Subprocesses], page 677, for a related feature for subprocesses.

Variable buffer-file-type

This variable, automatically local in each buffer, records the file type of the buffer’s visited file. The value is nil for text, t for binary.

Function find-buffer-file-type filename

This function determines whether file filename is a text file or a binary file. It returns nil for text, t for binary.

User Option file-name-buffer-file-type-alist

This variable holds an alist for distinguishing text files from binary files. Each element has the form (regexp . type), where regexp is matched against the file name, and type may be is nil for text, t for binary, or a function to call to compute which. If it is a function, then it is called with a single argument (the file name) and should return t or nil.

User Option default-buffer-file-type

This variable specifies the default file type for files whose names don’t indicate anything in particular. Its value should be nil for text, or t for binary.

Command find-file-text filename

Like find-file, but treat the file as text regardless of its name.

Command find-file-binary filename

Like find-file, but treat the file as binary regardless of its name.

29 Backups and Auto-Saving

Backup files and auto-save files are two methods by which XEmacs tries to protect the user from the consequences of crashes or of the user’s own errors. Auto-saving preserves the text from earlier in the current editing session; backup files preserve file contents prior to the current session.

29.1 Backup Files

A backup file is a copy of the old contents of a file you are editing. XEmacs makes a backup file the first time you save a buffer into its visited file. Normally, this means that the backup file contains the contents of the file as it was before the current editing session. The contents of the backup file normally remain unchanged once it exists.

Backups are usually made by renaming the visited file to a new name. Optionally, you can specify that backup files should be made by copying the visited file. This choice makes a difference for files with multiple names; it also can affect whether the edited file remains owned by the original owner or becomes owned by the user editing it.

By default, XEmacs makes a single backup file for each file edited. You can alternatively request numbered backups; then each new backup file gets a new name. You can delete old numbered backups when you don’t want them any more, or XEmacs can delete them automatically.

29.1.1 Making Backup Files

Function backup-buffer

This function makes a backup of the file visited by the current buffer, if appropriate. It is called by save-buffer before saving the buffer the first time.

Variable buffer-backed-up

This buffer-local variable indicates whether this buffer’s file has been backed up on account of this buffer. If it is non-nil, then the backup file has been written. Otherwise, the file should be backed up when it is next saved (if backups are enabled). This is a permanent local; kill-local-variables does not alter it.

User Option make-backup-files

This variable determines whether or not to make backup files. If it is nonnil, then XEmacs creates a backup of each file when it is saved for the first time—provided that backup-inhibited is nil (see below). The following example shows how to change the make-backup-files variable only in the ‘RMAIL’ buffer and not elsewhere. Setting it nil stops XEmacs from making backups of the ‘RMAIL’ file, which may save disk space. (You would put this code in your ‘.emacs’ file.)

          (add-hook 'rmail-mode-hook
                    (function (lambda ()
                                (make-local-variable
                                 'make-backup-files)
                                (setq make-backup-files nil))))

Variable backup-enable-predicate

This variable’s value is a function to be called on certain occasions to decide whether a file should have backup files. The function receives one argument, a file name to consider. If the function returns nil, backups are disabled for that file. Otherwise, the other variables in this section say whether and how to make backups. The default value is this:

     (lambda (name)
       (or (< (length name) 5)
           (not (string-equal "/tmp/"
                              (substring name 0 5)))))

Variable backup-inhibited

If this variable is non-nil, backups are inhibited. It records the result of testing backup-enable-predicate on the visited file name. It can also coherently be used by other mechanisms that inhibit backups based on which file is visited. For example, VC sets this variable non-nil to prevent making backups for files managed with a version control system. This is a permanent local, so that changing the major mode does not lose its value. Major modes should not set this variable—they should set make-backup-files instead.

29.1.2 Backup by Renaming or by Copying?

There are two ways that XEmacs can make a backup file:

• XEmacs can rename the original file so that it becomes a backup file, and then write the buffer being saved into a new file. After this procedure, any other names (i.e., hard links) of the original file now refer to the backup file. The new file is owned by the user doing the editing, and its group is the default for new files written by the user in that directory.

• XEmacs can copy the original file into a backup file, and then overwrite the original file with new contents. After this procedure, any other names (i.e., hard links) of the original file still refer to the current version of the file. The file’s owner and group will be unchanged.

The first method, renaming, is the default. The variable backup-by-copying, if non-nil, says to use the second method, which is

to copy the original file and overwrite it with the new buffer contents. The variable fileprecious-flag, if non-nil, also has this effect (as a sideline of its main significance). See Section 28.2 [Saving Buffers], page 394.

Variable backup-by-copying

If this variable is non-nil, XEmacs always makes backup files by copying.

The following two variables, when non-nil, cause the second method to be used in certain special cases. They have no effect on the treatment of files that don’t fall into the special cases.

Variable backup-by-copying-when-linked

If this variable is non-nil, XEmacs makes backups by copying for files with multiple names (hard links). This variable is significant only if backup-by-copying is nil, since copying is always used when that variable is non-nil.

Variable backup-by-copying-when-mismatch

If this variable is non-nil, XEmacs makes backups by copying in cases where renaming would change either the owner or the group of the file. The value has no effect when renaming would not alter the owner or group of the file; that is, for files which are owned by the user and whose group matches the default for a new file created there by the user. This variable is significant only if backup-by-copying is nil, since copying is always used when that variable is non-nil.

29.1.3 Making and Deleting Numbered Backup Files

If a file’s name is ‘foo’, the names of its numbered backup versions are ‘foo.~v~’, for various integers v, like this: ‘foo.~1~’, ‘foo.~2~’, ‘foo.~3~’, …, ‘foo.~259~’, and so on.

User Option version-control

This variable controls whether to make a single non-numbered backup file or multiple numbered backups.

nil Make numbered backups if the visited file already has numbered backups; otherwise, do not.

never Do not make numbered backups.

anything else Make numbered backups.

The use of numbered backups ultimately leads to a large number of backup versions, which must then be deleted. XEmacs can do this automatically or it can ask the user whether to delete them.

User Option kept-new-versions

The value of this variable is the number of newest versions to keep when a new numbered backup is made. The newly made backup is included in the count. The default value is 2.

User Option kept-old-versions

The value of this variable is the number of oldest versions to keep when a new numbered backup is made. The default value is 2.

If there are backups numbered 1, 2, 3, 5, and 7, and both of these variables have the value 2, then the backups numbered 1 and 2 are kept as old versions and those numbered 5 and 7 are kept as new versions; backup version 3 is excess. The function find-backup-file-name (see Section 29.1.4 [Backup Names], page 424) is responsible for determining which backup versions to delete, but does not delete them itself.

User Option trim-versions-without-asking

If this variable is non-nil, then saving a file deletes excess backup versions silently. Otherwise, it asks the user whether to delete them.

User Option dired-kept-versions

This variable specifies how many of the newest backup versions to keep in the Dired command . (dired-clean-directory). That’s the same thing keptnew-versions specifies when you make a new backup file. The default value is 2.

29.1.4 Naming Backup Files

The functions in this section are documented mainly because you can customize the naming conventions for backup files by redefining them. If you change one, you probably need to change the rest.

Function backup-file-name-p filename

This function returns a non-nil value if filename is a possible name for a backup file. A file with the name filename need not exist; the function just checks the name.

     (backup-file-name-p "foo")
          ⇒ nil
     (backup-file-name-p "foo~")
          ⇒ 3

The standard definition of this function is as follows:

     (defun backup-file-name-p (file)
       "Return non-nil if FILE is a backup file \
     name (numeric or not)..."
       (string-match "~$" file))

Thus, the function returns a non-nil value if the file name ends with a ‘~’. (We use a backslash to split the documentation string’s first line into two lines in the text, but produce just one line in the string itself.) This simple expression is placed in a separate function to make it easy to redefine for customization.

Function make-backup-file-name filename

This function returns a string that is the name to use for a non-numbered backup file for file filename. On Unix, this is just filename with a tilde appended. The standard definition of this function is as follows:

     (defun make-backup-file-name (file)
       "Create the non-numeric backup file name for FILE.
...
     "
       (concat file "~"))

You can change the backup-file naming convention by redefining this function. The following example redefines make-backup-file-name to prepend a ‘.’ in addition to appending a tilde:

     (defun make-backup-file-name (filename)
       (concat "." filename "~"))
     (make-backup-file-name "backups.texi")
          ⇒ ".backups.texi~"

Function find-backup-file-name filename

This function computes the file name for a new backup file for filename. It may also propose certain existing backup files for deletion. find-backup-file-name returns a list whose car is the name for the new backup file and whose cdr is a list of backup files whose deletion is proposed.

Two variables, kept-old-versions and kept-new-versions, determine which

backup versions should be kept. This function keeps those versions by excluding them from the cdr of the value. See Section 29.1.3 [Numbered Backups], page 423. In this example, the value says that ‘~rms/foo.~5~’ is the name to use for the new backup file, and ‘~rms/foo.~3~’ is an “excess” version that the caller should consider deleting now.

     (find-backup-file-name "~rms/foo")
          ⇒ ("~rms/foo.~5~" "~rms/foo.~3~")

Function file-newest-backup filename

This function returns the name of the most recent backup file for filename, or nil if that file has no backup files. Some file comparison commands use this function so that they can automatically compare a file with its most recent backup.

29.2 Auto-Saving

XEmacs periodically saves all files that you are visiting; this is called auto-saving. Autosaving prevents you from losing more than a limited amount of work if the system crashes. By default, auto-saves happen every 300 keystrokes, or after around 30 seconds of idle time. See section “Auto-Saving: Protection Against Disasters” in The XEmacs Reference Manual, for information on auto-save for users. Here we describe the functions used to implement auto-saving and the variables that control them.

Variable buffer-auto-save-file-name

This buffer-local variable is the name of the file used for auto-saving the current buffer. It is nil if the buffer should not be auto-saved.

     buffer-auto-save-file-name
     => "/xcssun/users/rms/lewis/#files.texi#"

Command auto-save-mode arg

When used interactively without an argument, this command is a toggle switch: it turns on auto-saving of the current buffer if it is off, and vice-versa. With an argument arg, the command turns auto-saving on if the value of arg is t, a nonempty list, or a positive integer. Otherwise, it turns auto-saving off.

Function auto-save-file-name-p filename

This function returns a non-nil value if filename is a string that could be the name of an auto-save file. It works based on knowledge of the naming convention for auto-save files: a name that begins and ends with hash marks (‘#’) is a possible auto-save file name. The argument filename should not contain a directory part.

     (make-auto-save-file-name)
          ⇒ "/xcssun/users/rms/lewis/#files.texi#"
     (auto-save-file-name-p "#files.texi#")
          ⇒ 0
     (auto-save-file-name-p "files.texi")
          ⇒ nil

The standard definition of this function is as follows:

     (defun auto-save-file-name-p (filename)
       "Return non-nil if FILENAME can be yielded by..."
       (string-match "^#.*#$" filename))

This function exists so that you can customize it if you wish to change the naming convention for auto-save files. If you redefine it, be sure to redefine the

function make-auto-save-file-name correspondingly.

Function make-auto-save-file-name

This function returns the file name to use for auto-saving the current buffer. This is just the file name with hash marks (‘#’) appended and prepended to it. This function does not look at the variable auto-save-visited-file-name (described below); you should check that before calling this function.

     (make-auto-save-file-name)
          ⇒ "/xcssun/users/rms/lewis/#backup.texi#"

The standard definition of this function is as follows:

     (defun make-auto-save-file-name ()
       "Return file name to use for auto-saves \
     of current buffer.
...
     "
       (if buffer-file-name
           (concat
            (file-name-directory buffer-file-name)
            "#"
            (file-name-nondirectory buffer-file-name)
            "#")
         (expand-file-name
          (concat "#%" (buffer-name) "#"))))

This exists as a separate function so that you can redefine it to customize the naming convention for auto-save files. Be sure to change auto-save-file-name-p in a corresponding way.

Variable auto-save-visited-file-name

If this variable is non-nil, XEmacs auto-saves buffers in the files they are visiting. That is, the auto-save is done in the same file that you are editing.

Normally, this variable is nil, so auto-save files have distinct names that are

created by make-auto-save-file-name.

When you change the value of this variable, the value does not take effect until the next time auto-save mode is reenabled in any given buffer. If auto-save mode is already enabled, auto-saves continue to go in the same file name until

auto-save-mode is called again.

Function recent-auto-save-p

This function returns t if the current buffer has been auto-saved since the last time it was read in or saved.

Function set-buffer-auto-saved

This function marks the current buffer as auto-saved. The buffer will not be auto-saved again until the buffer text is changed again. The function returns

nil.

User Option auto-save-interval

The value of this variable is the number of characters that XEmacs reads from the keyboard between auto-saves. Each time this many more characters are read, auto-saving is done for all buffers in which it is enabled.

User Option auto-save-timeout

The value of this variable is the number of seconds of idle time that should cause auto-saving. Each time the user pauses for this long, XEmacs auto-saves any buffers that need it. (Actually, the specified timeout is multiplied by a factor depending on the size of the current buffer.)

Variable auto-save-hook

This normal hook is run whenever an auto-save is about to happen.

User Option auto-save-default

If this variable is non-nil, buffers that are visiting files have auto-saving enabled by default. Otherwise, they do not.

Command do-auto-save &optional no-message current-only This function auto-saves all buffers that need to be auto-saved. It saves all buffers for which auto-saving is enabled and that have been changed since the previous auto-save.

Normally, if any buffers are auto-saved, a message that says ‘Auto-saving...’ is displayed in the echo area while auto-saving is going on. However, if nomessage is non-nil, the message is inhibited.

If current-only is non-nil, only the current buffer is auto-saved.

Function delete-auto-save-file-if-necessary

This function deletes the current buffer’s auto-save file if delete-auto-save-files is non-nil. It is called every time a buffer is saved.

Variable delete-auto-save-files

This variable is used by the function delete-auto-save-file-if-necessary.

If it is non-nil, Emacs deletes auto-save files when a true save is done (in the visited file). This saves disk space and unclutters your directory.

Function rename-auto-save-file

This function adjusts the current buffer’s auto-save file name if the visited file name has changed. It also renames an existing auto-save file. If the visited file name has not changed, this function does nothing.

Variable buffer-saved-size

The value of this buffer-local variable is the length of the current buffer as of the last time it was read in, saved, or auto-saved. This is used to detect a substantial decrease in size, and turn off auto-saving in response. If it is -1, that means auto-saving is temporarily shut off in this buffer due to a substantial deletion. Explicitly saving the buffer stores a positive value in this variable, thus reenabling auto-saving. Turning auto-save mode off or on also alters this variable.

Variable auto-save-list-file-name

This variable (if non-nil) specifies a file for recording the names of all the autosave files. Each time XEmacs does auto-saving, it writes two lines into this file for each buffer that has auto-saving enabled. The first line gives the name of the visited file (it’s empty if the buffer has none), and the second gives the name of the auto-save file. If XEmacs exits normally, it deletes this file. If XEmacs crashes, you can look in the file to find all the auto-save files that might contain work that was otherwise lost. The recover-session command uses these files. The default name for this file is in your home directory and starts with ‘.saves-’. It also contains the XEmacs process id and the host name.

29.3 Reverting

If you have made extensive changes to a file and then change your mind about them, you can get rid of them by reading in the previous version of the file with the revert-buffer command. See section “Reverting a Buffer” in The XEmacs Reference Manual.

Command revert-buffer &optional check-auto-save noconfirm This command replaces the buffer text with the text of the visited file on disk. This action undoes all changes since the file was visited or saved. If the argument check-auto-save is non-nil, and the latest auto-save file is more recent than the visited file, revert-buffer asks the user whether to use that instead. Otherwise, it always uses the text of the visited file itself. Interactively, check-auto-save is set if there is a numeric prefix argument. Normally, revert-buffer asks for confirmation before it changes the buffer; but if the argument noconfirm is non-nil, revert-buffer does not ask for confirmation.

Reverting tries to preserve marker positions in the buffer by using the replacement feature of insert-file-contents. If the buffer contents and the file contents are identical before the revert operation, reverting preserves all the markers. If they are not identical, reverting does change the buffer; then it preserves the markers in the unchanged text (if any) at the beginning and end of the buffer. Preserving any additional markers would be problematical.

You can customize how revert-buffer does its work by setting these variables— typically, as buffer-local variables.

Variable revert-buffer-function

The value of this variable is the function to use to revert this buffer. If non-nil, it is called as a function with no arguments to do the work of reverting. If the value is nil, reverting works the usual way. Modes such as Dired mode, in which the text being edited does not consist of a file’s contents but can be regenerated in some other fashion, give this variable a buffer-local value that is a function to regenerate the contents.

Variable revert-buffer-insert-file-contents-function

The value of this variable, if non-nil, is the function to use to insert the updated contents when reverting this buffer. The function receives two arguments: first the file name to use; second, t if the user has asked to read the auto-save file.

Variable before-revert-hook

This normal hook is run by revert-buffer before actually inserting the mod-

ified contents—but only if revert-buffer-function is nil.

Font Lock mode uses this hook to record that the buffer contents are no longer fontified.

Variable after-revert-hook

This normal hook is run by revert-buffer after actually inserting the modified

contents—but only if revert-buffer-function is nil.

Font Lock mode uses this hook to recompute the fonts for the updated buffer contents.

430 XEmacs Lisp Reference Manual

30 Buffers

A buffer is a Lisp object containing text to be edited. Buffers are used to hold the contents of files that are being visited; there may also be buffers that are not visiting files. While several buffers may exist at one time, exactly one buffer is designated the current buffer at any time. Most editing commands act on the contents of the current buffer. Each buffer, including the current buffer, may or may not be displayed in any windows.

30.1 Buffer Basics

Buffers in Emacs editing are objects that have distinct names and hold text that can be edited. Buffers appear to Lisp programs as a special data type. You can think of the contents of a buffer as an extendable string; insertions and deletions may occur in any part of the buffer. See Chapter 36 [Text], page 511.

A Lisp buffer object contains numerous pieces of information. Some of this information is directly accessible to the programmer through variables, while other information is accessible only through special-purpose functions. For example, the visited file name is directly accessible through a variable, while the value of point is accessible only through a primitive function.

Buffer-specific information that is directly accessible is stored in buffer-local variable bindings, which are variable values that are effective only in a particular buffer. This feature allows each buffer to override the values of certain variables. Most major modes override variables such as fill-column or comment-column in this way. For more information about buffer-local variables and functions related to them, see Section 10.9 [Buffer-Local Variables], page 157.

For functions and variables related to visiting files in buffers, see Section 28.1 [Visiting Files], page 391 and Section 28.2 [Saving Buffers], page 394. For functions and variables related to the display of buffers in windows, see Section 31.6 [Buffers and Windows], page 453.

Function bufferp object

This function returns t if object is a buffer, nil otherwise.

30.2 The Current Buffer

There are, in general, many buffers in an Emacs session. At any time, one of them is designated as the current buffer. This is the buffer in which most editing takes place, because most of the primitives for examining or changing text in a buffer operate implicitly on the current buffer (see Chapter 36 [Text], page 511). Normally the buffer that is displayed on the screen in the selected window is the current buffer, but this is not always so: a Lisp program can designate any buffer as current temporarily in order to operate on its contents, without changing what is displayed on the screen.

The way to designate a current buffer in a Lisp program is by calling set-buffer. The specified buffer remains current until a new one is designated.

When an editing command returns to the editor command loop, the command loop designates the buffer displayed in the selected window as current, to prevent confusion: the buffer that the cursor is in when Emacs reads a command is the buffer that the command will apply to. (See Chapter 19 [Command Loop], page 281.) Therefore, set-buffer is not the way to switch visibly to a different buffer so that the user can edit it. For this, you must use the functions described in Section 31.7 [Displaying Buffers], page 453.

However, Lisp functions that change to a different current buffer should not depend on the command loop to set it back afterwards. Editing commands written in XEmacs Lisp can be called from other programs as well as from the command loop. It is convenient for the caller if the subroutine does not change which buffer is current (unless, of course, that is the subroutine’s purpose). Therefore, you should normally use set-buffer within a saveexcursion that will restore the current buffer when your function is done (see Section 34.3 [Excursions], page 495). Here is an example, the code for the command append-to-buffer (with the documentation string abridged):

     (defun append-to-buffer (buffer start end)
       "Append to specified buffer the text of the region.
...
     "
       (interactive "BAppend to buffer: \nr")
       (let ((oldbuf (current-buffer)))
         (save-excursion
           (set-buffer (get-buffer-create buffer))
           (insert-buffer-substring oldbuf start end))))

This function binds a local variable to the current buffer, and then save-excursion records the values of point, the mark, and the original buffer. Next, set-buffer makes another buffer current. Finally, insert-buffer-substring copies the string from the original current buffer to the new current buffer.

If the buffer appended to happens to be displayed in some window, the next redisplay will show how its text has changed. Otherwise, you will not see the change immediately on the screen. The buffer becomes current temporarily during the execution of the command, but this does not cause it to be displayed.

If you make local bindings (with let or function arguments) for a variable that may also have buffer-local bindings, make sure that the same buffer is current at the beginning and at the end of the local binding’s scope. Otherwise you might bind it in one buffer and unbind it in another! There are two ways to do this. In simple cases, you may see that nothing ever changes the current buffer within the scope of the binding. Otherwise, use save-excursion to make sure that the buffer current at the beginning is current again whenever the variable is unbound.

It is not reliable to change the current buffer back with set-buffer, because that won’t do the job if a quit happens while the wrong buffer is current. Here is what not to do:

     (let (buffer-read-only
           (obuf (current-buffer)))
       (set-buffer ...)
       ...
       (set-buffer obuf))

Using save-excursion, as shown below, handles quitting, errors, and throw, as well as ordinary evaluation.

(let (buffer-read-only)
  (save-excursion
    (set-buffer ...)
    ...))

Function current-buffer

This function returns the current buffer.

     (current-buffer)
          ⇒ #<buffer buffers.texi>

Function set-buffer buffer-or-name This function makes buffer-or-name the current buffer. It does not display the buffer in the currently selected window or in any other window, so the user cannot necessarily see the buffer. But Lisp programs can in any case work on it.

This function returns the buffer identified by buffer-or-name. An error is signaled if buffer-or-name does not identify an existing buffer.

30.3 Buffer Names

Each buffer has a unique name, which is a string. Many of the functions that work on buffers accept either a buffer or a buffer name as an argument. Any argument called buffer-or-name is of this sort, and an error is signaled if it is neither a string nor a buffer. Any argument called buffer must be an actual buffer object, not a name.

Buffers that are ephemeral and generally uninteresting to the user have names starting with a space, so that the list-buffers and buffer-menu commands don’t mention them. A name starting with space also initially disables recording undo information; see Section 36.9 [Undo], page 523.

Function buffer-name &optional buffer This function returns the name of buffer as a string. If buffer is not supplied, it defaults to the current buffer.

If buffer-name returns nil, it means that buffer has been killed. See Section 30.10 [Killing Buffers], page 441.

     (buffer-name)
          ⇒ "buffers.texi"
     (setq foo (get-buffer "temp"))
          ⇒ #<buffer temp>
     (kill-buffer foo)
          ⇒ nil
     (buffer-name foo)
          ⇒ nil
     foo
          ⇒ #<killed buffer>

Command rename-buffer newname &optional unique This function renames the current buffer to newname. An error is signaled if newname is not a string, or if there is already a buffer with that name. The function returns nil. Ordinarily, rename-buffer signals an error if newname is already in use. However, if unique is non-nil, it modifies newname to make a name that is not in use. Interactively, you can make unique non-nil with a numeric prefix argument. One application of this command is to rename the ‘*shell*’ buffer to some other name, thus making it possible to create a second shell buffer under the

name ‘*shell*’.

Function get-buffer buffer-or-name This function returns the buffer specified by buffer-or-name. If buffer-or-name is a string and there is no buffer with that name, the value is nil. If bufferor-name is a buffer, it is returned as given. (That is not very useful, so the argument is usually a name.) For example:

     (setq b (get-buffer "lewis"))
          ⇒ #<buffer lewis>
     (get-buffer b)
          ⇒ #<buffer lewis>
     (get-buffer "Frazzle-nots")
          ⇒ nil

See also the function get-buffer-create in Section 30.9 [Creating Buffers], page 440.

Function generate-new-buffer-name starting-name &optional ignore This function returns a name that would be unique for a new buffer—but does not create the buffer. It starts with starting-name, and produces a name not currently in use for any buffer by appending a number inside of ‘<...>’. If ignore is given, it specifies a name that is okay to use (if it is in the sequence to be tried), even if a buffer with that name exists. See the related function generate-new-buffer in Section 30.9 [Creating Buffers], page 440.

30.4 Buffer File Name

The buffer file name is the name of the file that is visited in that buffer. When a buffer is not visiting a file, its buffer file name is nil. Most of the time, the buffer name is the same as the nondirectory part of the buffer file name, but the buffer file name and the buffer name are distinct and can be set independently. See Section 28.1 [Visiting Files], page 391.

Function buffer-file-name &optional buffer

This function returns the absolute file name of the file that buffer is visiting. If buffer is not visiting any file, buffer-file-name returns nil. If buffer is not supplied, it defaults to the current buffer.

     (buffer-file-name (other-buffer))
          ⇒ "/usr/user/lewis/manual/files.texi"

Variable buffer-file-name

This buffer-local variable contains the name of the file being visited in the current buffer, or nil if it is not visiting a file. It is a permanent local, unaffected

by kill-local-variables.
     buffer-file-name
          ⇒ "/usr/user/lewis/manual/buffers.texi"

It is risky to change this variable’s value without doing various other things. See the definition of set-visited-file-name in ‘files.el’; some of the things done there, such as changing the buffer name, are not strictly necessary, but others are essential to avoid confusing XEmacs.

Variable buffer-file-truename

This buffer-local variable holds the truename of the file visited in the current buffer, or nil if no file is visited. It is a permanent local, unaffected by killlocal-variables. See Section 28.6.3 [Truenames], page 401.

Variable buffer-file-number

This buffer-local variable holds the file number and directory device number of the file visited in the current buffer, or nil if no file or a nonexistent file is visited. It is a permanent local, unaffected by kill-local-variables. See Section 28.6.3 [Truenames], page 401. The value is normally a list of the form (filenum devnum). This pair of numbers uniquely identifies the file among all files accessible on the system. See the function file-attributes, in Section 28.6.4 [File Attributes], page 401, for more information about them.

Function get-file-buffer filename

This function returns the buffer visiting file filename. If there is no such buffer, it returns nil. The argument filename, which must be a string, is expanded (see Section 28.8.4 [File Name Expansion], page 409), then compared against the visited file names of all live buffers.

     (get-file-buffer "buffers.texi")
         ⇒ #<buffer buffers.texi>

In unusual circumstances, there can be more than one buffer visiting the same file name. In such cases, this function returns the first such buffer in the buffer list.

Command set-visited-file-name filename

If filename is a non-empty string, this function changes the name of the file visited in current buffer to filename. (If the buffer had no visited file, this gives it one.) The next time the buffer is saved it will go in the newly-specified file. This command marks the buffer as modified, since it does not (as far as XEmacs knows) match the contents of filename, even if it matched the former visited file.

If filename is nil or the empty string, that stands for “no visited file”. In this case, set-visited-file-name marks the buffer as having no visited file.

When the function set-visited-file-name is called interactively, it prompts for filename in the minibuffer.

See also clear-visited-file-modtime and verify-visited-file-modtime

in Section 30.5 [Buffer Modification], page 436.

Variable list-buffers-directory

This buffer-local variable records a string to display in a buffer listing in place of the visited file name, for buffers that don’t have a visited file name. Dired buffers use this variable.

30.5 Buffer Modification

XEmacs keeps a flag called the modified flag for each buffer, to record whether you have changed the text of the buffer. This flag is set to t whenever you alter the contents of the buffer, and cleared to nil when you save it. Thus, the flag shows whether there are unsaved changes. The flag value is normally shown in the modeline (see Section 26.3.2 [Modeline Variables], page 374), and controls saving (see Section 28.2 [Saving Buffers], page 394) and auto-saving (see Section 29.2 [Auto-Saving], page 425).

Some Lisp programs set the flag explicitly. For example, the function set-visited-file-name sets the flag to t, because the text does not match the newly-visited file, even if it is unchanged from the file formerly visited.

The functions that modify the contents of buffers are described in Chapter 36 [Text], page 511.

Function buffer-modified-p &optional buffer

This function returns t if the buffer buffer has been modified since it was last read in from a file or saved, or nil otherwise. If buffer is not supplied, the current buffer is tested.

Function set-buffer-modified-p flag

This function marks the current buffer as modified if flag is non-nil, or as unmodified if the flag is nil.

Another effect of calling this function is to cause unconditional redisplay of the modeline for the current buffer. In fact, the function redraw-modeline works by doing this:

     (set-buffer-modified-p (buffer-modified-p))

Command not-modified &optional arg This command marks the current buffer as unmodified, and not needing to be saved. (If arg is non-nil, the buffer is instead marked as modified.) Don’t use this function in programs, since it prints a message in the echo area; use

set-buffer-modified-p (above) instead.

Function buffer-modified-tick &optional buffer

This function returns buffer‘s modification-count. This is a counter that increments every time the buffer is modified. If buffer is nil (or omitted), the current buffer is used.

30.6 Comparison of Modification Time

Suppose that you visit a file and make changes in its buffer, and meanwhile the file itself is changed on disk. At this point, saving the buffer would overwrite the changes in the file. Occasionally this may be what you want, but usually it would lose valuable information. XEmacs therefore checks the file’s modification time using the functions described below before saving the file.

Function verify-visited-file-modtime buffer

This function compares what buffer has recorded for the modification time of its visited file against the actual modification time of the file as recorded by the operating system. The two should be the same unless some other process has written the file since XEmacs visited or saved it. The function returns t if the last actual modification time and XEmacs’s recorded modification time are the same, nil otherwise.

Function clear-visited-file-modtime

This function clears out the record of the last modification time of the file being visited by the current buffer. As a result, the next attempt to save this buffer will not complain of a discrepancy in file modification times. This function is called in set-visited-file-name and other exceptional places where the usual test to avoid overwriting a changed file should not be done.

Function visited-file-modtime

This function returns the buffer’s recorded last file modification time, as a list of the form (high . low). (This is the same format that file-attributes uses to return time values; see Section 28.6.4 [File Attributes], page 401.)

Function set-visited-file-modtime &optional time

This function updates the buffer’s record of the last modification time of the visited file, to the value specified by time if time is not nil, and otherwise to the last modification time of the visited file. If time is not nil, it should have the form (high . low) or (high low), in either case containing two integers, each of which holds 16 bits of the time. This function is useful if the buffer was not read from the file normally, or if the file itself has been changed for some known benign reason.

Function ask-user-about-supersession-threat filename

This function is used to ask a user how to proceed after an attempt to modify an obsolete buffer visiting file filename. An obsolete buffer is an unmodified buffer for which the associated file on disk is newer than the last save-time of the buffer. This means some other program has probably altered the file.

Depending on the user’s answer, the function may return normally, in which case the modification of the buffer proceeds, or it may signal a file-supersession error with data (filename), in which case the proposed buffer modification is not allowed. This function is called automatically by XEmacs on the proper occasions. It exists so you can customize XEmacs by redefining it. See the file ‘userlock.el’ for the standard definition. See also the file locking mechanism in Section 28.5 [File Locks], page 398.

30.7 Read-Only Buffers

If a buffer is read-only, then you cannot change its contents, although you may change your view of the contents by scrolling and narrowing.

Read-only buffers are used in two kinds of situations:

• A buffer visiting a write-protected file is normally read-only.

Here, the purpose is to show the user that editing the buffer with the aim of saving it in the file may be futile or undesirable. The user who wants to change the buffer text despite this can do so after clearing the read-only flag with C-x C-q.

• Modes such as Dired and Rmail make buffers read-only when altering the contents with the usual editing commands is probably a mistake. The special commands of these modes bind buffer-read-only to nil (with let) or bind inhibit-read-only to t around the places where they change the text.

Variable buffer-read-only

This buffer-local variable specifies whether the buffer is read-only. The buffer is read-only if this variable is non-nil.

Variable inhibit-read-only

If this variable is non-nil, then read-only buffers and read-only characters may be modified. Read-only characters in a buffer are those that have non-nilread-only properties (either text properties or extent properties). See Section 40.6 [Extent Properties], page 591, for more information about text properties and extent properties. If inhibit-read-only is t, all read-only character properties have no effect. If inhibit-read-only is a list, then read-only character properties have no effect if they are members of the list (comparison is done with eq).

Command toggle-read-only

This command changes whether the current buffer is read-only. It is intended for interactive use; don’t use it in programs. At any given point in a program, you should know whether you want the read-only flag on or off; so you can set buffer-read-only explicitly to the proper value, t or nil.

Function barf-if-buffer-read-only

This function signals a buffer-read-only error if the current buffer is readonly. See Section 19.3 [Interactive Call], page 286, for another way to signal an error if the current buffer is read-only.

30.8 The Buffer List

The buffer list is a list of all live buffers. Creating a buffer adds it to this list, and killing a buffer deletes it. The order of the buffers in the list is based primarily on how recently each buffer has been displayed in the selected window. Buffers move to the front of the list when they are selected and to the end when they are buried. Several functions, notably other-buffer, use this ordering. A buffer list displayed for the user also follows this order.

Every frame has its own order for the buffer list. Switching to a new buffer inside of a particular frame changes the buffer list order for that frame, but does not affect the buffer list order of any other frames. In addition, there is a global, non-frame buffer list order that is independent of the buffer list orders for any particular frame.

Note that the different buffer lists all contain the same elements. It is only the order of those elements that is different.

Function buffer-list &optional frame This function returns a list of all buffers, including those whose names begin with a space. The elements are actual buffers, not their names. The order of the list is specific to frame, which defaults to the current frame. If frame is t, the global, non-frame ordering is returned instead.

     (buffer-list)
          ⇒ (#<buffer buffers.texi>
              #<buffer  *Minibuf-1*> #<buffer buffer.c>
              #<buffer *Help*> #<buffer TAGS>)

;; Note that the name of the minibuffer

;; begins with a space!

     (mapcar (function buffer-name) (buffer-list))
         ⇒ ("buffers.texi" " *Minibuf-1*"
             "buffer.c" "*Help*" "TAGS")

Buffers appear earlier in the list if they were current more recently. This list is a copy of a list used inside XEmacs; modifying it has no effect on the buffers.

Function other-buffer &optional buffer-or-name frame visible-ok This function returns the first buffer in the buffer list other than buffer-or-name, in frame’s ordering for the buffer list. (frame defaults to the current frame. If frame is t, then the global, non-frame ordering is used.) Usually this is the buffer most recently shown in the selected window, aside from buffer-or-name. Buffers are moved to the front of the list when they are selected and to the end when they are buried. Buffers whose names start with a space are not considered. If buffer-or-name is not supplied (or if it is not a buffer), then other-buffer returns the first buffer on the buffer list that is not visible in any window in a visible frame. If the selected frame has a non-nil buffer-predicate property, then otherbuffer uses that predicate to decide which buffers to consider. It calls the predicate once for each buffer, and if the value is nil, that buffer is ignored. See Section 32.2.3 [X Frame Properties], page 473. If visible-ok is nil, other-buffer avoids returning a buffer visible in any window on any visible frame, except as a last resort. If visible-ok is non-nil, then it does not matter whether a buffer is displayed somewhere or not. If no suitable buffer exists, the buffer ‘*scratch*’ is returned (and created, if necessary). Note that in FSF Emacs 19, there is no frame argument, and visible-ok is the second argument instead of the third. FSF Emacs 19.

Command list-buffers &optional files-only This function displays a listing of the names of existing buffers. It clears the buffer ‘*Buffer List*’, then inserts the listing into that buffer and displays it in a window. list-buffers is intended for interactive use, and is described fully in The XEmacs Reference Manual. It returns nil.

Command bury-buffer &optional buffer-or-name This function puts buffer-or-name at the end of the buffer list without changing the order of any of the other buffers on the list. This buffer therefore becomes the least desirable candidate for other-buffer to return. If buffer-or-name is nil or omitted, this means to bury the current buffer. In addition, if the buffer is displayed in the selected window, this switches to some other buffer (obtained using other-buffer) in the selected window. But if the buffer is displayed in some other window, it remains displayed there. If you wish to replace a buffer in all the windows that display it, use replacebuffer-in-windows. See Section 31.6 [Buffers and Windows], page 453.

30.9 Creating Buffers

This section describes the two primitives for creating buffers. get-buffer-create creates a buffer if it finds no existing buffer with the specified name; generate-new-buffer always creates a new buffer and gives it a unique name.

Other functions you can use to create buffers include with-output-to-temp-buffer (see Section 45.8 [Temporary Displays], page 656) and create-file-buffer (see Section 28.1 [Visiting Files], page 391). Starting a subprocess can also create a buffer (see Chapter 49 [Processes], page 673).

Function get-buffer-create name

This function returns a buffer named name. It returns an existing buffer with that name, if one exists; otherwise, it creates a new buffer. The buffer does not become the current buffer—this function does not change which buffer is current. An error is signaled if name is not a string.

     (get-buffer-create "foo")
          ⇒ #<buffer foo>

The major mode for the new buffer is set to Fundamental mode. The variable default-major-mode is handled at a higher level. See Section 26.1.3 [Auto Major Mode], page 367.

Function generate-new-buffer name

This function returns a newly created, empty buffer, but does not make it current. If there is no buffer named name, then that is the name of the new buffer. If that name is in use, this function adds suffixes of the form ‘<n>’ to name, where n is an integer. It tries successive integers starting with 2 until it finds an available name.

An error is signaled if name is not a string.

     (generate-new-buffer "bar")
          ⇒ #<buffer bar>
     (generate-new-buffer "bar")
          ⇒ #<buffer bar<2>>
     (generate-new-buffer "bar")
          ⇒ #<buffer bar<3>>

The major mode for the new buffer is set to Fundamental mode. The variable default-major-mode is handled at a higher level. See Section 26.1.3 [Auto Major Mode], page 367.

See the related function generate-new-buffer-name in Section 30.3 [Buffer Names], page 433.

30.10 Killing Buffers

Killing a buffer makes its name unknown to XEmacs and makes its text space available for other use.

The buffer object for the buffer that has been killed remains in existence as long as anything refers to it, but it is specially marked so that you cannot make it current or display it. Killed buffers retain their identity, however; two distinct buffers, when killed, remain distinct according to eq.

If you kill a buffer that is current or displayed in a window, XEmacs automatically selects or displays some other buffer instead. This means that killing a buffer can in general change the current buffer. Therefore, when you kill a buffer, you should also take the precautions associated with changing the current buffer (unless you happen to know that the buffer being killed isn’t current). See Section 30.2 [Current Buffer], page 431.

If you kill a buffer that is the base buffer of one or more indirect buffers, the indirect buffers are automatically killed as well.

The buffer-name of a killed buffer is nil. To test whether a buffer has been killed, you can either use this feature or the function buffer-live-p.

Function buffer-live-p buffer

This function returns nil if buffer is deleted, and t otherwise.

Command kill-buffer buffer-or-name This function kills the buffer buffer-or-name, freeing all its memory for use as space for other buffers. (Emacs version 18 and older was unable to return the memory to the operating system.) It returns nil. Any processes that have this buffer as the process-buffer are sent the SIGHUP signal, which normally causes them to terminate. (The basic meaning of SIGHUP is that a dialup line has been disconnected.) See Section 49.5 [Deleting Processes], page 678. If the buffer is visiting a file and contains unsaved changes, kill-buffer asks the user to confirm before the buffer is killed. It does this even if not called interactively. To prevent the request for confirmation, clear the modified flag before calling kill-buffer. See Section 30.5 [Buffer Modification], page 436. Killing a buffer that is already dead has no effect.

     (kill-buffer "foo.unchanged")
          ⇒ nil
     (kill-buffer "foo.changed")
     ---------- Buffer: Minibuffer ----------
     Buffer foo.changed modified; kill anyway? (yes or no) yes
     ---------- Buffer: Minibuffer ----------

          ⇒ nil

Variable kill-buffer-query-functions

After confirming unsaved changes, kill-buffer calls the functions in the list kill-buffer-query-functions, in order of appearance, with no arguments. The buffer being killed is the current buffer when they are called. The idea is that these functions ask for confirmation from the user for various nonstandard reasons. If any of them returns nil, kill-buffer spares the buffer’s life.

Variable kill-buffer-hook

This is a normal hook run by kill-buffer after asking all the questions it is going to ask, just before actually killing the buffer. The buffer to be killed is current when the hook functions run. See Section 26.4 [Hooks], page 378.

Variable buffer-offer-save

This variable, if non-nil in a particular buffer, tells save-buffers-kill-emacs and save-some-buffers to offer to save that buffer, just as they offer to save file-visiting buffers. The variable buffer-offer-save automatically becomes buffer-local when set for any reason. See Section 10.9 [Buffer-Local Variables], page 157.

30.11 Indirect Buffers

An indirect buffer shares the text of some other buffer, which is called the base buffer of the indirect buffer. In some ways it is the analogue, for buffers, of a symbolic link among files. The base buffer may not itself be an indirect buffer.

The text of the indirect buffer is always identical to the text of its base buffer; changes made by editing either one are visible immediately in the other. This includes the text properties as well as the characters themselves.

But in all other respects, the indirect buffer and its base buffer are completely separate. They have different names, different values of point, different narrowing, different markers and overlays (though inserting or deleting text in either buffer relocates the markers and overlays for both), different major modes, and different local variables.

An indirect buffer cannot visit a file, but its base buffer can. If you try to save the indirect buffer, that actually works by saving the base buffer.

Killing an indirect buffer has no effect on its base buffer. Killing the base buffer effectively kills the indirect buffer in that it cannot ever again be the current buffer.

Command make-indirect-buffer base-buffer name

This creates an indirect buffer named name whose base buffer is base-buffer.

The argument base-buffer may be a buffer or a string.

If base-buffer is an indirect buffer, its base buffer is used as the base for the new buffer.

Function buffer-base-buffer buffer

This function returns the base buffer of buffer. If buffer is not indirect, the value is nil. Otherwise, the value is another buffer, which is never an indirect buffer.

444 XEmacs Lisp Reference Manual

31 Windows

This chapter describes most of the functions and variables related to Emacs windows. See Chapter 45 [Display], page 647, for information on how text is displayed in windows.

31.1 Basic Concepts of Emacs Windows

A window in XEmacs is the physical area of the screen in which a buffer is displayed. The term is also used to refer to a Lisp object that represents that screen area in XEmacs Lisp. It should be clear from the context which is meant.

XEmacs groups windows into frames. A frame represents an area of screen available for XEmacs to use. Each frame always contains at least one window, but you can subdivide it vertically or horizontally into multiple nonoverlapping Emacs windows.

In each frame, at any time, one and only one window is designated as selected within the frame. The frame’s cursor appears in that window. At ant time, one frame is the selected frame; and the window selected within that frame is the selected window. The selected window’s buffer is usually the current buffer (except when set-buffer has been used). See Section 30.2 [Current Buffer], page 431.

For practical purposes, a window exists only while it is displayed in a frame. Once removed from the frame, the window is effectively deleted and should not be used, even though there may still be references to it from other Lisp objects. Restoring a saved window configuration is the only way for a window no longer on the screen to come back to life. (See Section 31.3 [Deleting Windows], page 449.)

Each window has the following attributes:

• containing frame

• window height

• window width

• window edges with respect to the frame or screen

• the buffer it displays

• position within the buffer at the upper left of the window

• amount of horizontal scrolling, in columns

• point

• the mark

• how recently the window was selected

Users create multiple windows so they can look at several buffers at once. Lisp libraries use multiple windows for a variety of reasons, but most often to display related information. In Rmail, for example, you can move through a summary buffer in one window while the other window shows messages one at a time as they are reached.

The meaning of “window” in XEmacs is similar to what it means in the context of general-purpose window systems such as X, but not identical. The X Window System places X windows on the screen; XEmacs uses one or more X windows as frames, and subdivides them into Emacs windows. When you use XEmacs on a character-only terminal, XEmacs treats the whole terminal screen as one frame.

Most window systems support arbitrarily located overlapping windows. In contrast, Emacs windows are tiled; they never overlap, and together they fill the whole screen or frame. Because of the way in which XEmacs creates new windows and resizes them, you can’t create every conceivable tiling of windows on an Emacs frame. See Section 31.2 [Splitting Windows], page 446, and Section 31.13 [Size of Window], page 464.

See Chapter 45 [Display], page 647, for information on how the contents of the window’s buffer are displayed in the window.

Function windowp object

This function returns t if object is a window.

31.2 Splitting Windows

The functions described here are the primitives used to split a window into two windows. Two higher level functions sometimes split a window, but not always: pop-to-buffer and display-buffer (see Section 31.7 [Displaying Buffers], page 453).

The functions described here do not accept a buffer as an argument. The two “halves” of the split window initially display the same buffer previously visible in the window that was split.

Function one-window-p &optional no-mini all-frames This function returns non-nil if there is only one window. The argument nomini, if non-nil, means don’t count the minibuffer even if it is active; otherwise, the minibuffer window is included, if active, in the total number of windows which is compared against one. The argument all-frame controls which set of windows are counted.

• If it is nil or omitted, then count only the selected frame, plus the minibuffer it uses (which may be on another frame).

• If it is t, then windows on all frames that currently exist (including invisible and iconified frames) are counted.

• If it is the symbol visible, then windows on all visible frames are counted.

• If it is the number 0, then windows on all visible and iconified frames are counted.

• If it is any other value, then precisely the windows in window’s frame are counted, excluding the minibuffer in use if it lies in some other frame.

Command split-window &optional window size horizontal This function splits window into two windows. The original window window remains the selected window, but occupies only part of its former screen area. The rest is occupied by a newly created window which is returned as the value of this function. If horizontal is non-nil, then window splits into two side by side windows. The original window window keeps the leftmost size columns, and gives the rest of the columns to the new window. Otherwise, it splits into windows one above the other, and window keeps the upper size lines and gives the rest of the lines to the new window. The original window is therefore the left-hand or upper of the two, and the new window is the right-hand or lower.

If window is omitted or nil, then the selected window is split. If size is omitted or nil, then window is divided evenly into two parts. (If there is an odd line, it is allocated to the new window.) When split-window is called interactively, all its arguments are nil.

The following example starts with one window on a frame that is 50 lines high by 80 columns wide; then the window is split.

     (setq w (selected-window))
          ⇒ #<window 8 on windows.texi>
     (window-edges)          ; Edges in order:

⇒ (0 0 80 50) ; left–top–right–bottom

;; Returns window created

     (setq w2 (split-window w 15))
          ⇒ #<window 28 on windows.texi>
     (window-edges w2)

⇒ (0 15 80 50) ; Bottom window;

; top is line 15

     (window-edges w)

⇒ (0 0 80 15) ; Top window

The frame looks like this:

              __________
             |          |  line 0
             |    w     |
             |__________|
             |          |  line 15
             |    w2    |
             |__________|
                           line 50
      column 0   column 80

Next, the top window is split horizontally:

     (setq w3 (split-window w 35 t))
          ⇒ #<window 32 on windows.texi>
     (window-edges w3)

⇒ (35 0 80 15) ; Left edge at column 35

     (window-edges w)

⇒ (0 0 35 15) ; Right edge at column 35

     (window-edges w2)

⇒ (0 15 80 50) ; Bottom window unchanged Now, the screen looks like this:

          column 35
              __________
             |   |      |  line 0
             | w |  w3  |
             |___|______|
             |          |  line 15
             |    w2    |
             |__________|
                           line 50
      column 0   column 80

Normally, Emacs indicates the border between two side-by-side windows with a scroll bar (see Section 32.2.3 [X Frame Properties], page 473) or ‘|’ characters. The display table can specify alternative border characters; see Section 45.11 [Display Tables], page 659.

Command split-window-vertically &optional size

This function splits the selected window into two windows, one above the other, leaving the selected window with size lines. This function is simply an interface to split-windows. Here is the complete function definition for it:

     (defun split-window-vertically (&optional arg)
       "Split current window into two windows, one above the other."
       (interactive "P")
       (split-window nil (and arg (prefix-numeric-value arg))))

Command split-window-horizontally &optional size

This function splits the selected window into two windows side-by-side, leaving the selected window with size columns. This function is simply an interface to split-windows. Here is the complete definition for split-window-horizontally (except for part of the documentation string):

     (defun split-window-horizontally (&optional arg)
       "Split selected window into two windows, side by side..."
       (interactive "P")
       (split-window nil (and arg (prefix-numeric-value arg)) t))

Function one-window-p &optional no-mini all-frames This function returns non-nil if there is only one window. The argument nomini, if non-nil, means don’t count the minibuffer even if it is active; otherwise, the minibuffer window is included, if active, in the total number of windows, which is compared against one. The argument all-frames specifies which frames to consider. Here are the possible values and their meanings:

nil Count the windows in the selected frame, plus the minibuffer used by that frame even if it lies in some other frame. t Count all windows in all existing frames.

visible Count all windows in all visible frames.

0 Count all windows in all visible or iconified frames.

anything else Count precisely the windows in the selected frame, and no others.

31.3 Deleting Windows

A window remains visible on its frame unless you delete it by calling certain functions that delete windows. A deleted window cannot appear on the screen, but continues to exist as a Lisp object until there are no references to it. There is no way to cancel the deletion of a window aside from restoring a saved window configuration (see Section 31.16 [Window Configurations], page 468). Restoring a window configuration also deletes any windows that aren’t part of that configuration.

When you delete a window, the space it took up is given to one adjacent sibling. (In Emacs version 18, the space was divided evenly among all the siblings.)

Function window-live-p window

This function returns nil if window is deleted, and t otherwise. Warning: Erroneous information or fatal errors may result from using a deleted window as if it were live.

Command delete-window &optional window This function removes window from the display. If window is omitted, then the selected window is deleted. An error is signaled if there is only one window

when delete-window is called.

This function returns nil. When delete-window is called interactively, window defaults to the selected window.

Command delete-other-windows &optional window

This function makes window the only window on its frame, by deleting the other windows in that frame. If window is omitted or nil, then the selected window is used by default. The result is nil.

Command delete-windows-on buffer &optional frame This function deletes all windows showing buffer. If there are no windows showing buffer, it does nothing. delete-windows-on operates frame by frame. If a frame has several windows showing different buffers, then those showing buffer are removed, and the others expand to fill the space. If all windows in some frame are showing buffer (including the case where there is only one window), then the frame reverts to having a single window showing another buffer chosen with other-buffer. See Section 30.8 [The Buffer List], page 439. The argument frame controls which frames to operate on:

• If it is nil, operate on the selected frame.

• If it is t, operate on all frames.

• If it is visible, operate on all visible frames.

• 0 If it is 0, operate on all visible or iconified frames.

• If it is a frame, operate on that frame.

This function always returns nil.

31.4 Selecting Windows

When a window is selected, the buffer in the window becomes the current buffer, and the cursor will appear in it.

Function selected-window &optional device

This function returns the selected window. This is the window in which the cursor appears and to which many commands apply. Each separate device can have its own selected window, which is remembered as focus changes from device to device. Optional argument device specifies which device to return the selected window for, and defaults to the selected device.

Function select-window window &optional norecord This function makes window the selected window. The cursor then appears in window (on redisplay). The buffer being displayed in window is immediately designated the current buffer. If optional argument norecord is non-nil then the global and per-frame buffer orderings are not modified, as by the function record-buffer. The return value is window.

     (setq w (next-window))
     (select-window w)
          ⇒ #<window 65 on windows.texi>

Macro save-selected-window forms…

This macro records the selected window, executes forms in sequence, then restores the earlier selected window. It does not save or restore anything about the sizes, arrangement or contents of windows; therefore, if the forms change them, the changes are permanent.

The following functions choose one of the windows on the screen, offering various criteria for the choice.

Function get-lru-window &optional frame

This function returns the window least recently “used” (that is, selected). The selected window is always the most recently used window. The selected window can be the least recently used window if it is the only window. A newly created window becomes the least recently used window until it is selected. A minibuffer window is never a candidate. The argument frame controls which windows are considered.

• If it is nil, consider windows on the selected frame.

• If it is t, consider windows on all frames.

• If it is visible, consider windows on all visible frames.

• If it is 0, consider windows on all visible or iconified frames.

• If it is a frame, consider windows on that frame.

Function get-largest-window &optional frame

This function returns the window with the largest area (height times width). If there are no side-by-side windows, then this is the window with the most lines. A minibuffer window is never a candidate. If there are two windows of the same size, then the function returns the window that is first in the cyclic ordering of windows (see following section), starting from the selected window. The argument frame controls which set of windows are considered. See get-

lru-window, above.

31.5 Cyclic Ordering of Windows

When you use the command C-x o (other-window) to select the next window, it moves through all the windows on the screen in a specific cyclic order. For any given configuration of windows, this order never varies. It is called the cyclic ordering of windows.

This ordering generally goes from top to bottom, and from left to right. But it may go down first or go right first, depending on the order in which the windows were split.

If the first split was vertical (into windows one above each other), and then the subwindows were split horizontally, then the ordering is left to right in the top of the frame, and then left to right in the next lower part of the frame, and so on. If the first split was horizontal, the ordering is top to bottom in the left part, and so on. In general, within each set of siblings at any level in the window tree, the order is left to right, or top to bottom.

Function next-window &optional window minibuf all-frames This function returns the window following window in the cyclic ordering of windows. This is the window that C-x o would select if typed when window is selected. If window is the only window visible, then this function returns window. If omitted, window defaults to the selected window. The value of the argument minibuf determines whether the minibuffer is included in the window order. Normally, when minibuf is nil, the minibuffer is included if it is currently active; this is the behavior of C-x o. (The minibuffer window is active while the minibuffer is in use. See Chapter 18 [Minibuffers], page 261.) If minibuf is t, then the cyclic ordering includes the minibuffer window even if it is not active. If minibuf is neither t nor nil, then the minibuffer window is not included even if it is active. The argument all-frames specifies which frames to consider. Here are the possible values and their meanings: nil Consider all the windows in window’s frame, plus the minibuffer

used by that frame even if it lies in some other frame.

t Consider all windows in all existing frames.

visible Consider all windows in all visible frames. (To get useful results, you must ensure window is in a visible frame.)

0 Consider all windows in all visible or iconified frames.

anything else Consider precisely the windows in window’s frame, and no others.

This example assumes there are two windows, both displaying the buffer

‘windows.texi’:
     (selected-window)
          ⇒ #<window 56 on windows.texi>
     (next-window (selected-window))
          ⇒ #<window 52 on windows.texi>
     (next-window (next-window (selected-window)))
          ⇒ #<window 56 on windows.texi>

Function previous-window &optional window minibuf all-frames This function returns the window preceding window in the cyclic ordering of windows. The other arguments specify which windows to include in the cycle,

as in next-window.

Command other-window count &optional frame This function selects the countth following window in the cyclic order. If count is negative, then it selects the −countth preceding window. It returns nil. In an interactive call, count is the numeric prefix argument. The argument frame controls which set of windows are considered.

• If it is nil or omitted, then windows on the selected frame are considered.

• If it is a frame, then windows on that frame are considered.

• If it is t, then windows on all frames that currently exist (including invisible and iconified frames) are considered.

• If it is the symbol visible, then windows on all visible frames are considered.

• If it is the number 0, then windows on all visible and iconified frames are considered.

• If it is any other value, then the behavior is undefined.

Function walk-windows proc &optional minibuf all-frames This function cycles through all windows, calling proc once for each window with the window as its sole argument. The optional arguments minibuf and all-frames specify the set of windows to include in the scan. See next-window, above, for details.

31.6 Buffers and Windows

This section describes low-level functions to examine windows or to display buffers in windows in a precisely controlled fashion. See the following section for related functions that find a window to use and specify a buffer for it. The functions described there are easier to use than these, but they employ heuristics in choosing or creating a window; use these functions when you need complete control.

Function set-window-buffer window buffer-or-name This function makes window display buffer-or-name as its contents. It returns

     nil.
          (set-window-buffer (selected-window) "foo")
               ⇒ nil

Function window-buffer &optional window This function returns the buffer that window is displaying. If window is omitted, this function returns the buffer for the selected window.

          (window-buffer)
               ⇒ #<buffer windows.texi>

Function get-buffer-window buffer-or-name &optional frame This function returns a window currently displaying buffer-or-name, or nil if there is none. If there are several such windows, then the function returns the first one in the cyclic ordering of windows, starting from the selected window. See Section 31.5 [Cyclic Window Ordering], page 451. The argument all-frames controls which windows to consider. • If it is nil, consider windows on the selected frame.

• If it is t, consider windows on all frames.

• If it is visible, consider windows on all visible frames.

• If it is 0, consider windows on all visible or iconified frames.

• If it is a frame, consider windows on that frame.

31.7 Displaying Buffers in Windows

In this section we describe convenient functions that choose a window automatically and use it to display a specified buffer. These functions can also split an existing window in certain circumstances. We also describe variables that parameterize the heuristics used for choosing a window. See the preceding section for low-level functions that give you more precise control.

Do not use the functions in this section in order to make a buffer current so that a Lisp program can access or modify it; they are too drastic for that purpose, since they change the display of buffers in windows, which is gratuitous and will surprise the user. Instead, use set-buffer (see Section 30.2 [Current Buffer], page 431) and save-excursion (see Section 34.3 [Excursions], page 495), which designate buffers as current for programmed access without affecting the display of buffers in windows.

Command switch-to-buffer buffer-or-name &optional norecord This function makes buffer-or-name the current buffer, and also displays the buffer in the selected window. This means that a human can see the buffer and subsequent keyboard commands will apply to it. Contrast this with setbuffer, which makes buffer-or-name the current buffer but does not display it in the selected window. See Section 30.2 [Current Buffer], page 431. If buffer-or-name does not identify an existing buffer, then a new buffer by that name is created. The major mode for the new buffer is set according to the variable default-major-mode. See Section 26.1.3 [Auto Major Mode], page 367. Normally the specified buffer is put at the front of the buffer list. This affects the operation of other-buffer. However, if norecord is non-nil, this is not done. See Section 30.8 [The Buffer List], page 439. The switch-to-buffer function is often used interactively, as the binding of C-x b. It is also used frequently in programs. It always returns nil.

Command switch-to-buffer-other-window buffer-or-name

This function makes buffer-or-name the current buffer and displays it in a window not currently selected. It then selects that window. The handling of the buffer is the same as in switch-to-buffer. The currently selected window is absolutely never used to do the job. If it is the only window, then it is split to make a distinct window for this purpose. If the selected window is already displaying the buffer, then it continues to do so, but another window is nonetheless found to display it in as well.

Function pop-to-buffer buffer-or-name &optional other-window on-frame This function makes buffer-or-name the current buffer and switches to it in some window, preferably not the window previously selected. The “popped-to” window becomes the selected window within its frame. If the variable pop-up-frames is non-nil, pop-to-buffer looks for a window in any visible frame already displaying the buffer; if there is one, it returns that window and makes it be selected within its frame. If there is none, it creates a new frame and displays the buffer in it. If pop-up-frames is nil, then pop-to-buffer operates entirely within the selected frame. (If the selected frame has just a minibuffer, pop-to-buffer operates within the most recently selected frame that was not just a minibuffer.) If the variable pop-up-windows is non-nil, windows may be split to create a new window that is different from the original window. For details, see Section 31.8 [Choosing Window], page 455. If other-window is non-nil, pop-to-buffer finds or creates another window even if buffer-or-name is already visible in the selected window. Thus bufferor-name could end up displayed in two windows. On the other hand, if bufferor-name is already displayed in the selected window and other-window is nil, then the selected window is considered sufficient display for buffer-or-name, so that nothing needs to be done.

All the variables that affect display-buffer affect pop-to-buffer as well. See Section 31.8 [Choosing Window], page 455. If buffer-or-name is a string that does not name an existing buffer, a buffer by that name is created. The major mode for the new buffer is set according to the variable default-major-mode. See Section 26.1.3 [Auto Major Mode], page 367. If on-frame is non-nil, it is the frame to pop to this buffer on. An example use of this function is found at the end of Section 49.9.2 [Filter Functions], page 684.

Command replace-buffer-in-windows buffer

This function replaces buffer with some other buffer in all windows displaying it. The other buffer used is chosen with other-buffer. In the usual applications of this function, you don’t care which other buffer is used; you just want to make sure that buffer is no longer displayed. This function returns nil.

31.8 Choosing a Window for Display

This section describes the basic facility that chooses a window to display a buffer in— display-buffer. All the higher-level functions and commands use this subroutine. Here we describe how to use display-buffer and how to customize it.

Command display-buffer buffer-or-name &optional not-this-window This command makes buffer-or-name appear in some window, like pop-to-buffer, but it does not select that window and does not make the buffer current. The identity of the selected window is unaltered by this function. If not-this-window is non-nil, it means to display the specified buffer in a window other than the selected one, even if it is already on display in the selected window. This can cause the buffer to appear in two windows at once. Otherwise, if buffer-or-name is already being displayed in any window, that is good enough, so this function does nothing. display-buffer returns the window chosen to display buffer-or-name. Precisely how display-buffer finds or creates a window depends on the variables described below.

A window can be marked as “dedicated” to a particular buffer. Then XEmacs will not automatically change which buffer appears in the window, such as display-buffer might normally do.

Function window-dedicated-p window

This function returns window’s dedicated object, usually t or nil.

Function set-window-buffer-dedicated window buffer

This function makes window display buffer and be dedicated to that buffer. Then XEmacs will not automatically change which buffer appears in window. If buffer is nil, this function makes window not be dedicated (but doesn’t change which buffer appears in it currently).

User Option pop-up-windows

This variable controls whether display-buffer makes new windows. If it is non-nil and there is only one window, then that window is split. If it is nil, then display-buffer does not split the single window, but uses it whole.

User Option split-height-threshold

This variable determines when display-buffer may split a window, if there are multiple windows. display-buffer always splits the largest window if it has at least this many lines. If the largest window is not this tall, it is split only if it is the sole window and pop-up-windows is non-nil.

User Option pop-up-frames

This variable controls whether display-buffer makes new frames. If it is non-nil, display-buffer looks for an existing window already displaying the desired buffer, on any visible frame. If it finds one, it returns that window. Otherwise it makes a new frame. The variables pop-up-windows and split-

height-threshold do not matter if pop-up-frames is non-nil.

If pop-up-frames is nil, then display-buffer either splits a window or reuses one. See Chapter 32 [Frames], page 471, for more information.

Variable pop-up-frame-function

This variable specifies how to make a new frame if pop-up-frames is non-nil. Its value should be a function of no arguments. When display-buffer makes a new frame, it does so by calling that function, which should return a frame. The default value of the variable is a function that creates a frame using properties

from pop-up-frame-plist.

Variable pop-up-frame-plist

This variable holds a plist specifying frame properties used when displaybuffer makes a new frame. See Section 32.2 [Frame Properties], page 471, for more information about frame properties.

Variable special-display-buffer-names

A list of buffer names for buffers that should be displayed specially. If the buffer’s name is in this list, display-buffer handles the buffer specially. By default, special display means to give the buffer a dedicated frame. If an element is a list, instead of a string, then the car of the list is the buffer name, and the rest of the list says how to create the frame. There are two possibilities for the rest of the list. It can be a plist, specifying frame properties, or it can contain a function and arguments to give to it. (The function’s first argument is always the buffer to be displayed; the arguments from the list come after that.)

Variable special-display-regexps

A list of regular expressions that specify buffers that should be displayed specially. If the buffer’s name matches any of the regular expressions in this list, display-buffer handles the buffer specially.

By default, special display means to give the buffer a dedicated frame. If an element is a list, instead of a string, then the car of the list is the regular expression, and the rest of the list says how to create the frame. See above,

under special-display-buffer-names.

Variable special-display-function

This variable holds the function to call to display a buffer specially. It receives the buffer as an argument, and should return the window in which it is displayed. The default value of this variable is special-display-popup-frame.

Function special-display-popup-frame buffer

This function makes buffer visible in a frame of its own. If buffer is already displayed in a window in some frame, it makes the frame visible and raises it, to use that window. Otherwise, it creates a frame that will be dedicated to buffer. This function uses an existing window displaying buffer whether or not it is in a frame of its own; but if you set up the above variables in your init file, before buffer was created, then presumably the window was previously made by this function.

User Option special-display-frame-plist

This variable holds frame properties for special-display-popup-frame to use when it creates a frame.

Variable same-window-buffer-names

A list of buffer names for buffers that should be displayed in the selected window. If the buffer’s name is in this list, display-buffer handles the buffer by switching to it in the selected window.

Variable same-window-regexps

A list of regular expressions that specify buffers that should be displayed in the selected window. If the buffer’s name matches any of the regular expressions in this list, display-buffer handles the buffer by switching to it in the selected window.

Variable display-buffer-function

This variable is the most flexible way to customize the behavior of displaybuffer. If it is non-nil, it should be a function that display-buffer calls to do the work. The function should accept two arguments, the same two arguments that display-buffer received. It should choose or create a window, display the specified buffer, and then return the window. This hook takes precedence over all the other options and hooks described above.

A window can be marked as “dedicated” to its buffer. Then display-buffer does not try to use that window.

Function window-dedicated-p window

This function returns t if window is marked as dedicated; otherwise nil.

Function set-window-dedicated-p window flag

This function marks window as dedicated if flag is non-nil, and nondedicated otherwise.

31.9 Windows and Point

Each window has its own value of point, independent of the value of point in other windows displaying the same buffer. This makes it useful to have multiple windows showing one buffer.

• The window point is established when a window is first created; it is initialized from the buffer’s point, or from the window point of another window opened on the buffer if such a window exists.

• Selecting a window sets the value of point in its buffer to the window’s value of point. Conversely, deselecting a window sets the window’s value of point from that of the buffer. Thus, when you switch between windows that display a given buffer, the point value for the selected window is in effect in the buffer, while the point values for the other windows are stored in those windows.

• As long as the selected window displays the current buffer, the window’s point and the buffer’s point always move together; they remain equal.

• See Chapter 34 [Positions], page 487, for more details on buffer positions.

As far as the user is concerned, point is where the cursor is, and when the user switches to another buffer, the cursor jumps to the position of point in that buffer.

Function window-point window

This function returns the current position of point in window. For a nonselected window, this is the value point would have (in that window’s buffer) if that window were selected. When window is the selected window and its buffer is also the current buffer, the value returned is the same as point in that buffer. Strictly speaking, it would be more correct to return the “top-level” value of point, outside of any save-excursion forms. But that value is hard to find.

Function set-window-point window position

This function positions point in window at position position in window’s buffer.

31.10 The Window Start Position

Each window contains a marker used to keep track of a buffer position that specifies where in the buffer display should start. This position is called the display-start position of the window (or just the start). The character after this position is the one that appears at the upper left corner of the window. It is usually, but not inevitably, at the beginning of a text line.

Function window-start &optional window This function returns the display-start position of window window. If window is nil, the selected window is used. For example,

     (window-start)
          ⇒ 7058

When you create a window, or display a different buffer in it, the display-start position is set to a display-start position recently used for the same buffer, or 1 if the buffer doesn’t have any. For a realistic example, see the description of count-lines in Section 34.2.4 [Text Lines], page 490.

Function window-end &optional window This function returns the position of the end of the display in window window. If window is nil, the selected window is used. Simply changing the buffer text or moving point does not update the value that window-end returns. The value is updated only when Emacs redisplays and redisplay actually finishes. If the last redisplay of window was preempted, and did not finish, Emacs does not know the position of the end of display in that window. In that case, this function returns a value that is not correct. In a future version, window-end will return nil in that case.

Function set-window-start window position &optional noforce This function sets the display-start position of window to position in window’s buffer. It returns position. The display routines insist that the position of point be visible when a buffer is displayed. Normally, they change the display-start position (that is, scroll the window) whenever necessary to make point visible. However, if you specify the start position with this function using nil for noforce, it means you want display to start at position even if that would put the location of point off the screen. If this does place point off screen, the display routines move point to the left margin on the middle line in the window. For example, if point is 1 and you set the start of the window to 2, then point would be “above” the top of the window. The display routines will automatically move point if it is still 1 when redisplay occurs. Here is an example:

;; Here is what ‘foo’ looks like before executing

     ;; the set-window-start expression.
     ---------- Buffer: foo ----------
     ⋆This is the contents of buffer foo.
     2
     3
     4
     5
     6
     ---------- Buffer: foo ----------
     (set-window-start
      (selected-window)
      (1+ (window-start)))
     ⇒ 2

;; Here is what ‘foo’ looks like after executing

          ;; the set-window-start expression.
          ---------- Buffer: foo ----------
          his is the contents of buffer foo.
          2
          3
          ⋆4
          5
          6
          ---------- Buffer: foo ----------

If noforce is non-nil, and position would place point off screen at the next redisplay, then redisplay computes a new window-start position that works well with point, and thus position is not used.

Function pos-visible-in-window-p &optional position window

This function returns t if position is within the range of text currently visible on the screen in window. It returns nil if position is scrolled vertically out of view. The argument position defaults to the current position of point; window, to the selected window. Here is an example:

          (or (pos-visible-in-window-p
               (point) (selected-window))
              (recenter 0))

The pos-visible-in-window-p function considers only vertical scrolling. If position is out of view only because window has been scrolled horizontally, pos-visible-in-window-p returns t. See Section 31.12 [Horizontal Scrolling], page 462.

31.11 Vertical Scrolling

Vertical scrolling means moving the text up or down in a window. It works by changing the value of the window’s display-start location. It may also change the value of windowpoint to keep it on the screen.

In the commands scroll-up and scroll-down, the directions “up” and “down” refer to the motion of the text in the buffer at which you are looking through the window. Imagine that the text is written on a long roll of paper and that the scrolling commands move the paper up and down. Thus, if you are looking at text in the middle of a buffer and repeatedly call scroll-down, you will eventually see the beginning of the buffer.

Some people have urged that the opposite convention be used: they imagine that the window moves over text that remains in place. Then “down” commands would take you to the end of the buffer. This view is more consistent with the actual relationship between windows and the text in the buffer, but it is less like what the user sees. The position of a window on the terminal does not move, and short scrolling commands clearly move the text up or down on the screen. We have chosen names that fit the user’s point of view.

The scrolling functions (aside from scroll-other-window) have unpredictable results if the current buffer is different from the buffer that is displayed in the selected window. See Section 30.2 [Current Buffer], page 431.

Command scroll-up &optional count This function scrolls the text in the selected window upward count lines. If count is negative, scrolling is actually downward. If count is nil (or omitted), then the length of scroll is next-screen-context-lines lines less than the usable height of the window (not counting its modeline).

scroll-up returns nil.

Command scroll-down &optional count This function scrolls the text in the selected window downward count lines. If count is negative, scrolling is actually upward. If count is omitted or nil, then the length of the scroll is next-screen-context-lines lines less than the usable height of the window (not counting its mode line).

scroll-down returns nil.

Command scroll-other-window &optional count

This function scrolls the text in another window upward count lines. Negative values of count, or nil, are handled as in scroll-up. You can specify a buffer to scroll with the variable other-window-scroll-buffer. When the selected window is the minibuffer, the next window is normally the one at the top left corner. You can specify a different window to scroll with the variable minibuffer-scroll-window. This variable has no effect when any other window is selected. See Section 18.8 [Minibuffer Misc], page 278. When the minibuffer is active, it is the next window if the selected window is the one at the bottom right corner. In this case, scroll-other-window attempts to scroll the minibuffer. If the minibuffer contains just one line, it has nowhere to scroll to, so the line reappears after the echo area momentarily displays the message “Beginning of buffer”.

Variable other-window-scroll-buffer

If this variable is non-nil, it tells scroll-other-window which buffer to scroll.

User Option scroll-step

This variable controls how scrolling is done automatically when point moves off the screen. If the value is zero, then redisplay scrolls the text to center point vertically in the window. If the value is a positive integer n, then redisplay brings point back on screen by scrolling n lines in either direction, if possible; otherwise, it centers point. The default value is zero.

User Option scroll-conservatively

This variable controls how many lines Emacs tries to scroll before recentering. If you set it to a small number, then when you move point a short distance off the screen, XEmacs will scroll the screen just far enough to bring point back on screen, provided that does not exceed scroll-conservatively lines. This variable overrides the redisplay preemption.

User Option next-screen-context-lines

The value of this variable is the number of lines of continuity to retain when scrolling by full screens. For example, scroll-up with an argument of nil scrolls so that this many lines at the bottom of the window appear instead at the top. The default value is 2.

Command recenter &optional count This function scrolls the selected window to put the text where point is located at a specified vertical position within the window. If count is a nonnegative number, it puts the line containing point count lines down from the top of the window. If count is a negative number, then it counts upward from the bottom of the window, so that −1 stands for the last usable line in the window. If count is a non-nil list, then it stands for the line in the middle of the window. If count is nil, recenter puts the line containing point in the middle of the window, then clears and redisplays the entire selected frame. When recenter is called interactively, count is the raw prefix argument. Thus, typing C-u as the prefix sets the count to a non-nil list, while typing C-u 4 sets count to 4, which positions the current line four lines from the top. With an argument of zero, recenter positions the current line at the top of the window. This action is so handy that some people make a separate key binding to do this. For example,

     (defun line-to-top-of-window ()
       "Scroll current line to top of window.
     Replaces three keystroke sequence C-u 0 C-l."
       (interactive)
       (recenter 0))
     (global-set-key [kp-multiply] 'line-to-top-of-window)

31.12 Horizontal Scrolling

Because we read English first from top to bottom and second from left to right, horizontal scrolling is not like vertical scrolling. Vertical scrolling involves selection of a contiguous portion of text to display. Horizontal scrolling causes part of each line to go off screen. The amount of horizontal scrolling is therefore specified as a number of columns rather than as a position in the buffer. It has nothing to do with the display-start position returned by

window-start.

Usually, no horizontal scrolling is in effect; then the leftmost column is at the left edge of the window. In this state, scrolling to the right is meaningless, since there is no data to the left of the screen to be revealed by it; so this is not allowed. Scrolling to the left is allowed; it scrolls the first columns of text off the edge of the window and can reveal additional columns on the right that were truncated before. Once a window has a nonzero amount of leftward horizontal scrolling, you can scroll it back to the right, but only so far as to reduce the net horizontal scroll to zero. There is no limit to how far left you can scroll, but eventually all the text will disappear off the left edge.

Command scroll-left count

This function scrolls the selected window count columns to the left (or to the right if count is negative). The return value is the total amount of leftward horizontal scrolling in effect after the change—just like the value returned by

window-hscroll (below).

Command scroll-right count

This function scrolls the selected window count columns to the right (or to the left if count is negative). The return value is the total amount of leftward horizontal scrolling in effect after the change—just like the value returned by

window-hscroll (below).

Once you scroll a window as far right as it can go, back to its normal position where the total leftward scrolling is zero, attempts to scroll any farther right have no effect.

Function window-hscroll &optional window This function returns the total leftward horizontal scrolling of window—the number of columns by which the text in window is scrolled left past the left margin. The value is never negative. It is zero when no horizontal scrolling has been done in window (which is usually the case). If window is nil, the selected window is used.

     (window-hscroll)
          ⇒ 0
     (scroll-left 5)
          ⇒ 5
     (window-hscroll)
          ⇒ 5

Function set-window-hscroll window columns

This function sets the number of columns from the left margin that window is scrolled to the value of columns. The argument columns should be zero or positive; if not, it is taken as zero. The value returned is columns.

     (set-window-hscroll (selected-window) 10)
          ⇒ 10

Here is how you can determine whether a given position position is off the screen due to horizontal scrolling:

(defun hscroll-on-screen (window position)
  (save-excursion
    (goto-char position)
    (and
     (>= (- (current-column) (window-hscroll window)) 0)
     (< (- (current-column) (window-hscroll window))
        (window-width window)))))

31.13 The Size of a Window

An Emacs window is rectangular, and its size information consists of the height (in lines or pixels) and the width (in character positions or pixels). The modeline is included in the height. The pixel width and height values include scrollbars and margins, while the line/character-position values do not.

Note that the height in lines, and the width in characters, are determined by dividing the corresponding pixel value by the height or width of the default font in that window (if this is a variable-width font, the average width is used). The resulting values may or may not represent the actual number of lines in the window, or the actual number of character positions in any particular line, esp. if there are pixmaps or various different fonts in the window.

The following functions return size information about a window:

Function window-height &optional window This function returns the number of lines in window, including its modeline but not including the horizontal scrollbar, if any (this is different from windowpixel-height). If window is nil, the function uses the selected window.

          (window-height)
               ⇒ 40
          (split-window-vertically)
               ⇒ #<window on "windows.texi" 0x679b>
          (window-height)
               ⇒ 20

Function window-width &optional window This function returns the number of columns in window, not including any left margin, right margin, or vertical scrollbar (this is different from window-pixel-width). If window is nil, the function uses the selected window.

          (window-width)
               ⇒ 80
          (window-height)
               ⇒ 40
          (split-window-horizontally)
               ⇒ #<window on "windows.texi" 0x7538>
          (window-width)
               ⇒ 39

Note that after splitting the window into two side-by-side windows, the width of each window is less the half the width of the original window because a vertical scrollbar appeared between the windows, occupying two columns worth of space. Also, the height shrunk by one because horizontal scrollbars appeared that weren’t there before. (Horizontal scrollbars appear only when lines are truncated, not when they wrap. This is usually the case for horizontally split windows but not for full-frame windows. You can change this using the

variables truncate-lines and truncate-partial-width-windows.)

Function window-pixel-height &optional window

This function returns the height of window in pixels, including its modeline and horizontal scrollbar, if any. If window is nil, the function uses the selected window.

     (window-pixel-height)
          ⇒ 600
     (split-window-vertically)
          ⇒ #<window on "windows.texi" 0x68a6>
     (window-pixel-height)
          ⇒ 300

Function window-pixel-width &optional window

This function returns the width of window in pixels, including any left margin, right margin, or vertical scrollbar that may be displayed alongside it. If window is nil, the function uses the selected window.

     (window-pixel-width)
          ⇒ 735
     (window-pixel-height)
          ⇒ 600
     (split-window-horizontally)
          ⇒ #<window on "windows.texi" 0x7538>
     (window-pixel-width)
          ⇒ 367
     (window-pixel-height)
          ⇒ 600

Function window-text-area-pixel-height &optional window

This function returns the height in pixels of the text displaying portion of window, which defaults to the selected window. Unlike window-pixel-height, the space occupied by the modeline and horizontal scrollbar, if any, is not counted.

Function window-text-area-pixel-width &optional window

This function returns the width in pixels of the text displaying portion of window, which defaults to the selected window. Unlike window-pixel-width, the space occupied by the vertical scrollbar and divider, if any, is not counted.

Function window-displayed-text-pixel-height &optional window

noclipped This function returns the height in pixels of the text displayed in window, which defaults to the selected window. Unlike window-text-area-pixel-height, any blank space below the end of the buffer is not included. If optional argument noclipped is non-nil, any space occupied by clipped lines will not be included.

31.14 The Position of a Window

XEmacs provides functions to determine the absolute location of windows within a frame, and the relative location of a window in comparison to other windows in the same frame.

Function window-pixel-edges &optional window

This function returns a list of the pixel edge coordinates of window. If window is nil, the selected window is used. The order of the list is (left top right bottom), all elements relative to 0, 0 at the top left corner of the frame. The element right of the value is one more than the rightmost pixel used by window (including any left margin, right margin, or vertical scrollbar displayed alongside it), and bottom is one more than the bottommost pixel used by window (including any modeline or horizontal scrollbar displayed above or below it). The frame area does not include any frame menubars or toolbars that may be displayed; thus, for example, if there is only one window on the frame, the values for left and top will always be 0. If window is at the upper left corner of its frame, right and bottom are the

same as the values returned by (window-pixel-width) and (window-pixel-

height) respectively, and top and bottom are zero.

There is no longer a function window-edges because it does not make sense in a world with variable-width and variable-height lines, as are allowed in XEmacs.

Function window-highest-p window

This function returns non-nil if window is along the top of its frame.

Function window-lowest-p window

This function returns non-nil if window is along the bottom of its frame.

Function window-text-area-pixel-edges &optional window

This function allows one to determine the location of the text-displaying portion of window, which defaults to the selected window, with respect to the top left corner of the window. It returns a list of integer pixel positions (left top right bottom), all relative to (0,0) at the top left corner of the window.

31.15 Changing the Size of a Window

The window size functions fall into two classes: high-level commands that change the size of windows and low-level functions that access window size. XEmacs does not permit overlapping windows or gaps between windows, so resizing one window affects other windows.

Command enlarge-window size &optional horizontal window This function makes the selected window size lines taller, stealing lines from neighboring windows. It takes the lines from one window at a time until that window is used up, then takes from another. If a window from which lines are stolen shrinks below window-min-height lines, that window disappears. If horizontal is non-nil, this function makes window wider by size columns, stealing columns instead of lines. If a window from which columns are stolen shrinks below window-min-width columns, that window disappears. If the requested size would exceed that of the window’s frame, then the function makes the window occupy the entire height (or width) of the frame.

If size is negative, this function shrinks the window by −size lines or columns. If that makes the window smaller than the minimum size (window-min-height

and window-min-width), enlarge-window deletes the window.

If window is non-nil, it specifies a window to change instead of the selected window.

enlarge-window returns nil.

Command enlarge-window-horizontally columns

This function makes the selected window columns wider. It could be defined as follows:

     (defun enlarge-window-horizontally (columns)
       (enlarge-window columns t))

Command enlarge-window-pixels count &optional side window This function makes the selected window count pixels larger. When called from Lisp, optional second argument side non-nil means to grow sideways count pixels, and optional third argument window specifies the window to change instead of the selected window.

Command shrink-window size &optional horizontal window This function is like enlarge-window but negates the argument size, making the selected window smaller by giving lines (or columns) to the other windows.

If the window shrinks below window-min-height or window-min-width, then

it disappears. If size is negative, the window is enlarged by −size lines or columns. If window is non-nil, it specifies a window to change instead of the selected window.

Command shrink-window-horizontally columns

This function makes the selected window columns narrower. It could be defined as follows:

     (defun shrink-window-horizontally (columns)
       (shrink-window columns t))

Command shrink-window-pixels count &optional side window This function makes the selected window count pixels smaller. When called from Lisp, optional second argument side non-nil means to shrink sideways count pixels, and optional third argument window specifies the window to change instead of the selected window.

The following two variables constrain the window-size-changing functions to a minimum height and width.

User Option window-min-height

The value of this variable determines how short a window may become before it is automatically deleted. Making a window smaller than window-min-height automatically deletes it, and no window may be created shorter than this. The absolute minimum height is two (allowing one line for the mode line, and one line for the buffer display). Actions that change window sizes reset this variable to two if it is less than two. The default value is 4.

User Option window-min-width

The value of this variable determines how narrow a window may become before it automatically deleted. Making a window smaller than window-min-width automatically deletes it, and no window may be created narrower than this. The absolute minimum width is one; any value below that is ignored. The default value is 10.

Variable window-size-change-functions

This variable holds a list of functions to be called if the size of any window changes for any reason. The functions are called just once per redisplay, and just once for each frame on which size changes have occurred. Each function receives the frame as its sole argument. There is no direct way to find out which windows changed size, or precisely how; however, if your sizechange function keeps track, after each change, of the windows that interest you, you can figure out what has changed by comparing the old size data with the new. Creating or deleting windows counts as a size change, and therefore causes these functions to be called. Changing the frame size also counts, because it changes the sizes of the existing windows. It is not a good idea to use save-window-excursion in these functions, because that always counts as a size change, and it would cause these functions to be called over and over. In most cases, save-selected-window is what you need here.

31.16 Window Configurations

A window configuration records the entire layout of a frame—all windows, their sizes, which buffers they contain, what part of each buffer is displayed, and the values of point and the mark. You can bring back an entire previous layout by restoring a window configuration previously saved.

If you want to record all frames instead of just one, use a frame configuration instead of a window configuration. See Section 32.11 [Frame Configurations], page 481.

Function current-window-configuration

This function returns a new object representing XEmacs’s current window configuration, namely the number of windows, their sizes and current buffers, which window is the selected window, and for each window the displayed buffer, the display-start position, and the positions of point and the mark. An exception is made for point in the current buffer, whose value is not saved.

Function set-window-configuration configuration

This function restores the configuration of XEmacs’s windows and buffers to the state specified by configuration. The argument configuration must be a value that was previously returned by current-window-configuration.

This function always counts as a window size change and triggers execution of the window-size-change-functions. (It doesn’t know how to tell whether the new configuration actually differs from the old one.) Here is a way of using this function to get the same effect as save-window-

excursion:
     (let ((config (current-window-configuration)))
       (unwind-protect
           (progn (split-window-vertically nil)
                  ...)
         (set-window-configuration config)))

Special Form save-window-excursion forms…

This special form records the window configuration, executes forms in sequence, then restores the earlier window configuration. The window configuration includes the value of point and the portion of the buffer that is visible. It also includes the choice of selected window. However, it does not include the value of point in the current buffer; use save-excursion if you wish to preserve that. Don’t use this construct when save-selected-window is all you need. Exit from save-window-excursion always triggers execution of the windowsize-change-functions. (It doesn’t know how to tell whether the restored configuration actually differs from the one in effect at the end of the forms.) The return value is the value of the final form in forms. For example:

     (split-window)
          ⇒ #<window 25 on control.texi>
     (setq w (selected-window))
          ⇒ #<window 19 on control.texi>
     (save-window-excursion
       (delete-other-windows w)
       (switch-to-buffer "foo")
       'do-something)
          ⇒ do-something

;; The frame is now split again.

Function window-configuration-p object

This function returns t if object is a window configuration.

Primitives to look inside of window configurations would make sense, but none are implemented. It is not clear they are useful enough to be worth implementing.

470 XEmacs Lisp Reference Manual

32 Frames

A frame is a rectangle on the screen that contains one or more XEmacs windows. A frame initially contains a single main window (plus perhaps a minibuffer window), which you can subdivide vertically or horizontally into smaller windows.

When XEmacs runs on a text-only terminal, it starts with one TTY frame. If you create additional ones, XEmacs displays one and only one at any given time—on the terminal screen, of course.

When XEmacs communicates directly with an X server, it does not have a TTY frame; instead, it starts with a single X window frame. It can display multiple X window frames at the same time, each in its own X window.

Function framep object

This predicate returns t if object is a frame, and nil otherwise.

See Chapter 45 [Display], page 647, for related information.

32.1 Creating Frames

To create a new frame, call the function make-frame.

Function make-frame &optional props device This function creates a new frame on device, if device permits creation of frames. (An X server does; an ordinary terminal does not (yet).) device defaults to the selected device if omitted. See Chapter 33 [Consoles and Devices], page 483. The argument props is a property list (a list of alternating keyword-value specifications) of properties for the new frame. (An alist is accepted for backward compatibility but should not be passed in.) Any properties not mentioned in props default according to the value of the variable default-frame-plist. For X devices, properties not specified in default-frame-plist default in turn from default-x-frame-plist and, if not specified there, from the X resources. For TTY devices, default-tty-frame-plist is consulted as well as default-

frame-plist.

The set of possible properties depends in principle on what kind of window system XEmacs uses to display its frames. See Section 32.2.3 [X Frame Properties], page 473, for documentation of individual properties you can specify when creating an X window frame.

32.2 Frame Properties

A frame has many properties that control its appearance and behavior. Just what properties a frame has depends on which display mechanism it uses.

Frame properties exist for the sake of window systems. A terminal frame has few properties, mostly for compatibility’s sake; only the height, width and buffer-predicate properties really do something.

32.2.1 Access to Frame Properties

These functions let you read and change the properties of a frame.

Function frame-properties &optional frame

This function returns a plist listing all the properties of frame and their values.

Function frame-property frame property &optional default This function returns frame’s value for the property property.

Function set-frame-properties frame plist

This function alters the properties of frame frame based on the elements of property list plist. If you don’t mention a property in plist, its value doesn’t change.

Function set-frame-property frame prop val

This function sets the property prop of frame frame to the value val.

32.2.2 Initial Frame Properties

You can specify the properties for the initial startup frame by setting initial-frame-plist in your ‘.emacs’ file.

Variable initial-frame-plist

This variable’s value is a plist of alternating property-value pairs used when creating the initial X window frame. XEmacs creates the initial frame before it reads your ‘~/.emacs’ file. After reading that file, XEmacs checks initial-frame-plist, and applies the property settings in the altered value to the already created initial frame. If these settings affect the frame geometry and appearance, you’ll see the frame appear with the wrong ones and then change to the specified ones. If that bothers you, you can specify the same geometry and appearance with X resources; those do take affect before the frame is created. See section “X Resources” in The XEmacs User’s Manual. X resource settings typically apply to all frames. If you want to specify some X resources solely for the sake of the initial frame, and you don’t want them to apply to subsequent frames, here’s how to achieve this: specify properties in default-frame-plist to override the X resources for subsequent frames; then, to prevent these from affecting the initial frame, specify the same properties in initial-frame-plist with values that match the X resources.

If these properties specify a separate minibuffer-only frame via a minibuffer property of nil, and you have not yet created one, XEmacs creates one for you.

Variable minibuffer-frame-plist

This variable’s value is a plist of properties used when creating an initial minibuffer-only frame—if such a frame is needed, according to the properties for the main initial frame.

Variable default-frame-plist

This is a plist specifying default values of frame properties for subsequent XEmacs frames (not the initial ones).

See also special-display-frame-plist, in Section 31.8 [Choosing Window], page 455. If you use options that specify window appearance when you invoke XEmacs, they

take effect by adding elements to default-frame-plist. One exception is ‘-geometry’, which adds the specified position to initial-frame-plist instead. See section “Command Arguments” in The XEmacs User’s Manual.

32.2.3 X Window Frame Properties

Just what properties a frame has depends on what display mechanism it uses. Here is a table of the properties of an X window frame; of these, name, height, width, and buffer-predicate provide meaningful information in non-X frames.

name The name of the frame. Most window managers display the frame’s name in the frame’s border, at the top of the frame. If you don’t specify a name, and you have more than one frame, XEmacs sets the frame name based on the buffer displayed in the frame’s selected window. If you specify the frame name explicitly when you create the frame, the name is also used (instead of the name of the XEmacs executable) when looking up X resources for the frame.

display The display on which to open this frame. It should be a string of the form "host:dpy.screen", just like the DISPLAY environment variable.

left The screen position of the left edge, in pixels, with respect to the left edge of the screen. The value may be a positive number pos, or a list of the form (+ pos) which permits specifying a negative pos value. A negative number −pos, or a list of the form (- pos), actually specifies the position of the right edge of the window with respect to the right edge of the screen. A positive value of pos counts toward the left. If the property is a negative integer −pos then pos is positive!

top The screen position of the top edge, in pixels, with respect to the top edge of the screen. The value may be a positive number pos, or a list of the form (+ pos) which permits specifying a negative pos value. A negative number −pos, or a list of the form (- pos), actually specifies the position of the bottom edge of the window with respect to the bottom edge of the screen. A positive value of pos counts toward the top. If the property is a negative integer −pos then pos is positive!

icon-left

The screen position of the left edge of the frame’s icon, in pixels, counting from the left edge of the screen. This takes effect if and when the frame is iconified.

icon-top The screen position of the top edge of the frame’s icon, in pixels, counting from the top edge of the screen. This takes effect if and when the frame is iconified.

user-position

Non-nil if the screen position of the frame was explicitly requested by the user (for example, with the ‘-geometry’ option). Nothing automatically makes this property non-nil; it is up to Lisp programs that call make-frame to specify this property as well as specifying the left and top properties.

height The height of the frame contents, in characters. (To get the height in pixels, call frame-pixel-height; see Section 32.2.4 [Size and Position], page 475.)

width The width of the frame contents, in characters. (To get the height in pixels, call frame-pixel-width; see Section 32.2.4 [Size and Position], page 475.)

window-id

The number of the X window for the frame.

minibuffer

Whether this frame has its own minibuffer. The value t means yes, nil means no, only means this frame is just a minibuffer. If the value is a minibuffer window (in some other frame), the new frame uses that minibuffer. (Minibufferonly and minibuffer-less frames are not yet implemented in XEmacs.)

buffer-predicate

The buffer-predicate function for this frame. The function other-buffer uses this predicate (from the selected frame) to decide which buffers it should consider, if the predicate is not nil. It calls the predicate with one arg, a buffer, once for each buffer; if the predicate returns a non-nil value, it considers that buffer.

scroll-bar-width

The width of the vertical scroll bar, in pixels.

cursor-color

The color for the cursor that shows point.

border-color

The color for the border of the frame.

border-width

The width in pixels of the window border.

internal-border-width

The distance in pixels between text and border.

unsplittable

If non-nil, this frame’s window is never split automatically.

inter-line-space

The space in pixels between adjacent lines of text. (Not currently implemented.)

modeline Whether the frame has a modeline.

32.2.4 Frame Size And Position

You can read or change the size and position of a frame using the frame properties left, top, height, and width. Whatever geometry properties you don’t specify are chosen by the window manager in its usual fashion.

Here are some special features for working with sizes and positions:

Function set-frame-position frame left top

This function sets the position of the top left corner of frame to left and top. These arguments are measured in pixels, and count from the top left corner of the screen. Negative property values count up or rightward from the top left corner of the screen.

Function frame-height &optional frame Function frame-width &optional frame

These functions return the height and width of frame, measured in lines and columns. If you don’t supply frame, they use the selected frame.

Function frame-pixel-height &optional frame Function frame-pixel-width &optional frame

These functions return the height and width of frame, measured in pixels. If you don’t supply frame, they use the selected frame.

Function set-frame-size frame cols rows &optional pretend This function sets the size of frame, measured in characters; cols and rows specify the new width and height. (If pretend is non-nil, it means that redisplay should act as if the frame’s size is cols by rows, but the actual size of the frame should not be changed. You should not normally use this option.)

You can also use the functions set-frame-height and set-frame-width to set the height and width individually. The frame is the first argument and the size (in rows or columns) is the second. (There is an optional third argument, pretend, which has the same purpose as the corresponding argument in set-frame-size.)

32.2.5 The Name of a Frame (As Opposed to Its Title)

Under X, every frame has a name, which is not the same as the title of the frame. A frame’s name is used to look up its resources and does not normally change over the lifetime of a frame. It is perfectly allowable, and quite common, for multiple frames to have the same name.

Function frame-name &optional frame This function returns the name of frame, which defaults to the selected frame if not specified. The name of a frame can also be obtained from the frame’s properties. See Section 32.2 [Frame Properties], page 471.

Variable default-frame-name

This variable holds the default name to assign to newly-created frames. This can be overridden by arguments to make-frame. This must be a string.

32.3 Frame Titles

Every frame has a title; most window managers display the frame title at the top of the frame. You can specify an explicit title with the name frame property. But normally you don’t specify this explicitly, and XEmacs computes the title automatically.

XEmacs computes the frame title based on a template stored in the variable frame-

title-format.

Variable frame-title-format

This variable specifies how to compute a title for a frame when you have not explicitly specified one.

The variable’s value is actually a modeline construct, just like modelineformat. See Section 26.3.1 [Modeline Data], page 373.

Variable frame-icon-title-format

This variable specifies how to compute the title for an iconified frame, when you have not explicitly specified the frame title. This title appears in the icon itself.

Function x-set-frame-icon-pixmap frame pixmap &optional mask This function sets the icon of the given frame to the given image instance, which should be an image instance object (as returned by make-image-instance), a glyph object (as returned by make-glyph), or nil. If a glyph object is given, the glyph will be instantiated on the frame to produce an image instance object.

If the given image instance has a mask, that will be used as the icon mask; however, not all window managers support this.

The window manager is also not required to support color pixmaps, only bitmaps (one plane deep).

If the image instance does not have a mask, then the optional third argument may be the image instance to use as the mask (it must be one plane deep). See Chapter 43 [Glyphs], page 625.

32.4 Deleting Frames

Frames remain potentially visible until you explicitly delete them. A deleted frame cannot appear on the screen, but continues to exist as a Lisp object until there are no references to it.

Command delete-frame &optional frame This function deletes the frame frame. By default, frame is the selected frame.

Function frame-live-p frame

The function frame-live-p returns non-nil if the frame frame has not been deleted.

32.5 Finding All Frames

Function frame-list

The function frame-list returns a list of all the frames that have not been deleted. It is analogous to buffer-list for buffers. The list that you get is newly created, so modifying the list doesn’t have any effect on the internals of XEmacs.

Function device-frame-list &optional device

This function returns a list of all frames on device. If device is nil, the selected device will be used.

Function visible-frame-list &optional device

This function returns a list of just the currently visible frames. If device is specified only frames on that device will be returned. See Section 32.9 [Visibility of Frames], page 480. (TTY frames always count as “visible”, even though only the selected one is actually displayed.)

Function next-frame &optional frame minibuf The function next-frame lets you cycle conveniently through all the frames from an arbitrary starting point. It returns the “next” frame after frame in the cycle. If frame is omitted or nil, it defaults to the selected frame.

The second argument, minibuf, says which frames to consider:

nil Exclude minibuffer-only frames.

visible Consider all visible frames.

0 Consider all visible or iconified frames.

a window Consider only the frames using that particular window as their minibuffer.

the symbol visible Include all visible frames.

0 Include all visible and iconified frames.

anything else Consider all frames.

Function previous-frame &optional frame minibuf Like next-frame, but cycles through all frames in the opposite direction.

See also next-window and previous-window, in Section 31.5 [Cyclic Window Ordering], page 451.

32.6 Frames and Windows

Each window is part of one and only one frame; you can get the frame with window-

frame.

Function frame-root-window &optional frame

This returns the root window of frame frame. frame defaults to the selected frame if not specified.

Function window-frame &optional window This function returns the frame that window is on. window defaults to the selected window if omitted.

All the non-minibuffer windows in a frame are arranged in a cyclic order. The order runs from the frame’s top window, which is at the upper left corner, down and to the right, until it reaches the window at the lower right corner (always the minibuffer window, if the frame has one), and then it moves back to the top.

Function frame-top-window frame

This returns the topmost, leftmost window of frame frame.

At any time, exactly one window on any frame is selected within the frame. The significance of this designation is that selecting the frame also selects this window. You can get the frame’s current selected window with frame-selected-window.

Function frame-selected-window &optional frame

This function returns the window on frame that is selected within frame. frame defaults to the selected frame if not specified.

Conversely, selecting a window for XEmacs with select-window also makes that window selected within its frame. See Section 31.4 [Selecting Windows], page 450.

Another function that (usually) returns one of the windows in a frame is minibufferwindow. See Section 18.8 [Minibuffer Misc], page 278.

32.7 Minibuffers and Frames

Normally, each frame has its own minibuffer window at the bottom, which is used whenever that frame is selected. If the frame has a minibuffer, you can get it with minibufferwindow (see Section 18.8 [Minibuffer Misc], page 278).

However, you can also create a frame with no minibuffer. Such a frame must use the minibuffer window of some other frame. When you create the frame, you can specify explicitly the minibuffer window to use (in some other frame). If you don’t, then the minibuffer is found in the frame which is the value of the variable default-minibuffer-frame. Its value should be a frame which does have a minibuffer.

Variable default-minibuffer-frame

This variable specifies the frame to use for the minibuffer window, by default.

32.8 Input Focus

At any time, one frame in XEmacs is the selected frame. The selected window always resides on the selected frame. As the focus moves from device to device, the selected frame on each device is remembered and restored when the focus moves back to that device.

Function selected-frame &optional device This function returns the selected frame on device. If device is not specified, the selected device will be used. If no frames exist on the device, nil is returned.

The X server normally directs keyboard input to the X window that the mouse is in. Some window managers use mouse clicks or keyboard events to shift the focus to various X windows, overriding the normal behavior of the server.

Lisp programs can switch frames “temporarily” by calling the function select-frame. This does not override the window manager; rather, it escapes from the window manager’s control until that control is somehow reasserted.

When using a text-only terminal, there is no window manager; therefore, select-frame is the only way to switch frames, and the effect lasts until overridden by a subsequent call to select-frame. Only the selected terminal frame is actually displayed on the terminal. Each terminal screen except for the initial one has a number, and the number of the selected frame appears in the mode line after the word ‘XEmacs’ (see Section 26.3.2 [Modeline Variables], page 374).

Function select-frame frame

This function selects frame frame, temporarily disregarding the focus of the X server if any. The selection of frame lasts until the next time the user does something to select a different frame, or until the next time this function is called. Note that select-frame does not actually cause the window-system focus to

be set to this frame, or the select-frame-hook or deselect-frame-hook to

be run, until the next time that XEmacs is waiting for an event. Also note that when the variable focus-follows-mouse is non-nil, the frame selection is temporary and is reverted when the current command terminates, much like the buffer selected by set-buffer. In order to effect a permanent focus change use focus-frame.

Function focus-frame frame

This function selects frame and gives it the window system focus. The operation

of focus-frame is not affected by the value of focus-follows-mouse.

Macro save-selected-frame forms…

This macro records the selected frame, executes forms in sequence, then restores the earlier selected frame. The value returned is the value of the last form.

Macro with-selected-frame frame forms…

This macro records the selected frame, then selects frame and executes forms in sequence. After the last form is finished, the earlier selected frame is restored. The value returned is the value of the last form.

32.9 Visibility of Frames

An X window frame may be visible, invisible, or iconified. If it is visible, you can see its contents. If it is iconified, the frame’s contents do not appear on the screen, but an icon does. If the frame is invisible, it doesn’t show on the screen, not even as an icon.

Visibility is meaningless for TTY frames, since only the selected one is actually displayed in any case.

Command make-frame-visible &optional frame

This function makes frame frame visible. If you omit frame, it makes the selected frame visible.

Command make-frame-invisible &optional frame

This function makes frame frame invisible.

Command iconify-frame &optional frame This function iconifies frame frame.

Command deiconify-frame &optional frame

This function de-iconifies frame frame. Under X, this is equivalent to make-

frame-visible.

Function frame-visible-p frame

This returns whether frame is currently “visible” (actually in use for display). A frame that is not visible is not updated, and, if it works through a window system, may not show at all.

Function frame-iconified-p frame

This returns whether frame is iconified. Not all window managers use icons; some merely unmap the window, so this function is not the inverse of framevisible-p. It is possible for a frame to not be visible and not be iconified either. However, if the frame is iconified, it will not be visible. (Under FSF Emacs, the functionality of this function is obtained through frame-visible-p.)

Function frame-totally-visible-p frame

This returns whether frame is not obscured by any other X windows. On TTY frames, this is the same as frame-visible-p.

32.10 Raising and Lowering Frames

The X Window System uses a desktop metaphor. Part of this metaphor is the idea that windows are stacked in a notional third dimension perpendicular to the screen surface, and thus ordered from “highest” to “lowest”. Where two windows overlap, the one higher up covers the one underneath. Even a window at the bottom of the stack can be seen if no other window overlaps it.

A window’s place in this ordering is not fixed; in fact, users tend to change the order frequently. Raising a window means moving it “up”, to the top of the stack. Lowering a window means moving it to the bottom of the stack. This motion is in the notional third dimension only, and does not change the position of the window on the screen.

You can raise and lower XEmacs’s X windows with these functions:

Command raise-frame &optional frame This function raises frame frame.

Command lower-frame &optional frame This function lowers frame frame.

You can also specify auto-raise (raising automatically when a frame is selected) or autolower (lowering automatically when it is deselected). Under X, most ICCCM-compliant window managers will have an option to do this for you, but the following variables are provided in case you’re using a broken WM. (Under FSF Emacs, the same functionality is provided through the auto-raise and auto-lower frame properties.)

Variable auto-raise-frame

This variable’s value is t if frames will be raised to the top when selected.

Variable auto-lower-frame

This variable’s value is t if frames will be lowered to the bottom when no longer selected.

Auto-raising and auto-lowering is implemented through functions attached to selectframe-hook and deselect-frame-hook (see Section 32.12 [Frame Hooks], page 481). Under normal circumstances, you should not call these functions directly.

Function default-select-frame-hook

This hook function implements the auto-raise-frame variable; it is for use as

the value of select-frame-hook.

Function default-deselect-frame-hook

This hook function implements the auto-lower-frame variable; it is for use as

the value of deselect-frame-hook.

32.11 Frame Configurations

A frame configuration records the current arrangement of frames, all their properties, and the window configuration of each one.

Function current-frame-configuration

This function returns a frame configuration list that describes the current arrangement of frames and their contents.

Function set-frame-configuration configuration

This function restores the state of frames described in configuration.

32.12 Hooks for Customizing Frame Behavior

XEmacs provides many hooks that are called at various times during a frame’s lifetime. See Section 26.4 [Hooks], page 378.

Variable create-frame-hook

This hook is called each time a frame is created. The functions are called with one argument, the newly-created frame.

Variable delete-frame-hook

This hook is called each time a frame is deleted. The functions are called with one argument, the about-to-be-deleted frame.

Variable select-frame-hook

This is a normal hook that is run just after a frame is selected. The function default-select-frame-hook, which implements auto-raising (see Section 32.10 [Raising and Lowering], page 480), is normally attached to this hook. Note that calling select-frame does not necessarily set the focus: The actual window-system focus will not be changed until the next time that XEmacs is waiting for an event, and even then, the window manager may refuse the focus-change request.

Variable deselect-frame-hook

This is a normal hook that is run just before a frame is deselected (and another frame is selected). The function default-deselect-frame-hook, which implements auto-lowering (see Section 32.10 [Raising and Lowering], page 480), is normally attached to this hook.

Variable map-frame-hook

This hook is called each time a frame is mapped (i.e. made visible). The functions are called with one argument, the newly mapped frame.

Variable unmap-frame-hook

This hook is called each time a frame is unmapped (i.e. made invisible or iconified). The functions are called with one argument, the newly unmapped frame.

33 Consoles and Devices

A console is an object representing a single input connection to XEmacs, such as an X display or a TTY connection. It is possible for XEmacs to have frames on multiple consoles at once (even on heterogeneous types – you can simultaneously have a frame on an X display and a TTY connection). Normally, there is only one console in existence.

A device is an object representing a single output device, such as a particular screen on an X display. (Usually there is exactly one device per X console connection, but there may be more than one if you have a multi-headed X display. For TTY connections, there is always exactly one device per console.)

Each device has one or more frames in which text can be displayed. For X displays and the like, a frame corresponds to the normal window-system concept of a window. Frames can overlap, be displayed at various locations within the display, be resized, etc. For TTY, only one frame can be displayed at a time, and it occupies the entire TTY display area.

However, you can still define multiple frames and switch between them. Their contents are entirely separate from each other. These sorts of frames resemble the “virtual console” capability provided under Linux or the multiple screens provided by the multiplexing program ‘screen’ under Unix.

When you start up XEmacs, an initial console and device are created to receive input and display frames on. This will either be an X display or a TTY connection, depending on what mode you started XEmacs in (this is determined by the ‘DISPLAY’ environment variable, the ‘-nw’, ‘-t’ and ‘-display’ command-line options, etc.).

You can connect to other X displays and TTY connections by creating new console objects, and to other X screens on an existing display by creating new device objects, as described below. Many functions (for example the frame-creation functions) take an optional device argument specifying which device the function pertains to. If the argument is omitted, it defaults to the selected device (see below).

Function consolep object

This returns non-nil if object is a console.

Function devicep object

This returns non-nil if object is a device.

33.1 Basic Console Functions

Function console-list

This function returns a list of all existing consoles.

Function console-device-list &optional console

This function returns a list of all devices on console. If console is nil, the selected console will be used.

33.2 Basic Device Functions

Function device-list

This function returns a list of all existing devices.

Function device-or-frame-p object

This function returns non-nil if object is a device or frame. This function is useful because devices and frames are similar in many respects and many functions can operate on either one.

Function device-frame-list device

This function returns a list of all frames on device.

Function frame-device frame

This function returns the device that frame is on.

33.3 Console Types and Device Classes

Every device is of a particular type, which describes how the connection to that device is made and how the device operates, and a particular class, which describes other characteristics of the device (currently, the color capabilities of the device).

The currently-defined device types are

x A connection to an X display (such as ‘willow:0’).

tty A connection to a tty (such as ‘/dev/ttyp3’).

stream A stdio connection. This describes a device for which input and output is only possible in a stream-like fashion, such as when XEmacs in running in batch mode. The very first device created by XEmacs is a terminal device and is used to print out messages of various sorts (for example, the help message when you use the ‘-help’ command-line option).

The currently-defined device classes are

color A color device.

grayscale

A grayscale device (a device that can display multiple shades of gray, but no color).

mono A device that can only display two colors (e.g. black and white).

Function device-type device

This function returns the type of device. This is a symbol whose name is one of the device types mentioned above.

Function device-or-frame-type device-or-frame

This function returns the type of device-or-frame.

Function device-class device

This function returns the class (color behavior) of device. This is a symbol whose name is one of the device classes mentioned above.

Function valid-device-type-p device-type

This function returns whether device-type (which should be a symbol) species a valid device type.

Function valid-device-class-p device-class

This function returns whether device-class (which should be a symbol) species a valid device class.

Variable terminal-device

This variable holds the initial terminal device object, which represents XEmacs’s stdout.

33.4 Connecting to a Console or Device

Function make-device &optional type device-data This function creates a new device.

The following two functions create devices of specific types and are written in terms of

make-device.

Function make-tty-device &optional tty terminal-type This function creates a new tty device on tty. This also creates the tty’s first frame. tty should be a string giving the name of a tty device file (e.g. ‘/dev/ttyp3’ under SunOS et al.), as returned by the ‘tty’ command issued from the Unix shell. A value of nil means use the stdin and stdout as passed to XEmacs from the shell. If terminal-type is non-nil, it should be a string specifying the type of the terminal attached to the specified tty. If it is nil, the terminal type will be inferred from the ‘TERM’ environment variable.

Function make-x-device &optional display argv-list This function creates a new device connected to display. Optional argument argv-list is a list of strings describing command line options.

Function delete-device device

This function deletes device, permanently eliminating it from use. This disconnects XEmacs’s connection to the device.

Variable create-device-hook

This variable, if non-nil, should contain a list of functions, which are called when a device is created.

Variable delete-device-hook

This variable, if non-nil, should contain a list of functions, which are called when a device is deleted.

Function console-live-p object

This function returns non-nil if object is a console that has not been deleted.

Function device-live-p object

This function returns non-nil if object is a device that has not been deleted.

Function device-x-display device

This function returns the X display which device is connected to, if device is an X device.

33.5 The Selected Console and Device

Function select-console console

This function selects the console console. Subsequent editing commands apply to its selected device, selected frame, and selected window. The selection of console lasts until the next time the user does something to select a different console, or until the next time this function is called.

Function selected-console

This function returns the console which is currently active.

Function select-device device

This function selects the device device.

Function selected-device &optional console This function returns the device which is currently active. If optional console is non-nil, this function returns the device that would be currently active if console were the selected console.

33.6 Console and Device I/O

Function console-disable-input console

This function disables input on console console.

Function console-enable-input console

This function enables input on console console.

Each device has a baud rate value associated with it. On most systems, changing this value will affect the amount of padding and other strategic decisions made during redisplay.

Function device-baud-rate &optional device

This function returns the output baud rate of device.

Function set-device-baud-rate device rate

This function sets the output baud rate of device to rate.

34 Positions

A position is the index of a character in the text of a buffer. More precisely, a position identifies the place between two characters (or before the first character, or after the last character), so we can speak of the character before or after a given position. However, we often speak of the character “at” a position, meaning the character after that position.

Positions are usually represented as integers starting from 1, but can also be represented as markers—special objects that relocate automatically when text is inserted or deleted so they stay with the surrounding characters. See Chapter 35 [Markers], page 499.

34.1 Point

Point is a special buffer position used by many editing commands, including the selfinserting typed characters and text insertion functions. Other commands move point through the text to allow editing and insertion at different places.

Like other positions, point designates a place between two characters (or before the first character, or after the last character), rather than a particular character. Usually terminals display the cursor over the character that immediately follows point; point is actually before the character on which the cursor sits.

The value of point is a number between 1 and the buffer size plus 1. If narrowing is in effect (see Section 34.4 [Narrowing], page 496), then point is constrained to fall within the accessible portion of the buffer (possibly at one end of it).

Each buffer has its own value of point, which is independent of the value of point in other buffers. Each window also has a value of point, which is independent of the value of point in other windows on the same buffer. This is why point can have different values in various windows that display the same buffer. When a buffer appears in only one window, the buffer’s point and the window’s point normally have the same value, so the distinction is rarely important. See Section 31.9 [Window Point], page 458, for more details.

Function point &optional buffer This function returns the value of point in buffer, as an integer. buffer defaults to the current buffer if omitted.

          (point) ⇒ 175

Function point-min &optional buffer This function returns the minimum accessible value of point in buffer. This is normally 1, but if narrowing is in effect, it is the position of the start of the region that you narrowed to. (See Section 34.4 [Narrowing], page 496.) buffer defaults to the current buffer if omitted.

Function point-max &optional buffer This function returns the maximum accessible value of point in buffer. This is (1+ (buffer-size buffer)), unless narrowing is in effect, in which case it is the position of the end of the region that you narrowed to. (see Section 34.4 [Narrowing], page 496). buffer defaults to the current buffer if omitted.

Function buffer-end flag &optional buffer This function returns (point-min buffer) if flag is less than 1, (point-max buffer) otherwise. The argument flag must be a number. buffer defaults to the current buffer if omitted.

Function buffer-size &optional buffer This function returns the total number of characters in buffer. In the absence of any narrowing (see Section 34.4 [Narrowing], page 496), point-max returns a value one larger than this. buffer defaults to the current buffer if omitted.

     (buffer-size)
          ⇒ 35
     (point-max) ⇒ 36

Variable buffer-saved-size

The value of this buffer-local variable is the former length of the current buffer, as of the last time it was read in, saved or auto-saved.

34.2 Motion

Motion functions change the value of point, either relative to the current value of point, relative to the beginning or end of the buffer, or relative to the edges of the selected window. See Section 34.1 [Point], page 487.

34.2.1 Motion by Characters

These functions move point based on a count of characters. goto-char is the fundamental primitive; the other functions use that.

Command goto-char position &optional buffer This function sets point in buffer to the value position. If position is less than 1, it moves point to the beginning of the buffer. If position is greater than the length of the buffer, it moves point to the end. buffer defaults to the current buffer if omitted. If narrowing is in effect, position still counts from the beginning of the buffer, but point cannot go outside the accessible portion. If position is out of range, goto-char moves point to the beginning or the end of the accessible portion. When this function is called interactively, position is the numeric prefix argument, if provided; otherwise it is read from the minibuffer. goto-char returns position.

Command forward-char &optional count buffer This function moves point count characters forward, towards the end of the buffer (or backward, towards the beginning of the buffer, if count is negative). If the function attempts to move point past the beginning or end of the buffer (or the limits of the accessible portion, when narrowing is in effect), an error

is signaled with error code beginning-of-buffer or end-of-buffer. buffer

defaults to the current buffer if omitted. In an interactive call, count is the numeric prefix argument.

Command backward-char &optional count buffer This function moves point count characters backward, towards the beginning of the buffer (or forward, towards the end of the buffer, if count is negative). If the function attempts to move point past the beginning or end of the buffer (or the limits of the accessible portion, when narrowing is in effect), an error

is signaled with error code beginning-of-buffer or end-of-buffer. buffer

defaults to the current buffer if omitted. In an interactive call, count is the numeric prefix argument.

34.2.2 Motion by Words

These functions for parsing words use the syntax table to decide whether a given character is part of a word. See Chapter 38 [Syntax Tables], page 569.

Command forward-word count &optional buffer This function moves point forward count words (or backward if count is negative). Normally it returns t. If this motion encounters the beginning or end of the buffer, or the limits of the accessible portion when narrowing is in effect, point stops there and the value is nil. buffer defaults to the current buffer if omitted. In an interactive call, count is set to the numeric prefix argument.

Command backward-word count &optional buffer This function is just like forward-word, except that it moves backward until encountering the front of a word, rather than forward. buffer defaults to the current buffer if omitted. In an interactive call, count is set to the numeric prefix argument. This function is rarely used in programs, as it is more efficient to call forwardword with a negative argument.

Variable words-include-escapes

This variable affects the behavior of forward-word and everything that uses it. If it is non-nil, then characters in the “escape” and “character quote” syntax classes count as part of words. Otherwise, they do not.

34.2.3 Motion to an End of the Buffer

To move point to the beginning of the buffer, write:

(goto-char (point-min))

Likewise, to move to the end of the buffer, use:

(goto-char (point-max))

Here are two commands that users use to do these things. They are documented here to warn you not to use them in Lisp programs, because they set the mark and display messages in the echo area.

Command beginning-of-buffer &optional n

This function moves point to the beginning of the buffer (or the limits of the accessible portion, when narrowing is in effect), setting the mark at the previous position. If n is non-nil, then it puts point n tenths of the way from the beginning of the buffer.

In an interactive call, n is the numeric prefix argument, if provided; otherwise n defaults to nil.

Don’t use this function in Lisp programs!

Command end-of-buffer &optional n

This function moves point to the end of the buffer (or the limits of the accessible portion, when narrowing is in effect), setting the mark at the previous position. If n is non-nil, then it puts point n tenths of the way from the end of the buffer.

In an interactive call, n is the numeric prefix argument, if provided; otherwise n defaults to nil.

Don’t use this function in Lisp programs!

34.2.4 Motion by Text Lines

Text lines are portions of the buffer delimited by newline characters, which are regarded as part of the previous line. The first text line begins at the beginning of the buffer, and the last text line ends at the end of the buffer whether or not the last character is a newline. The division of the buffer into text lines is not affected by the width of the window, by line continuation in display, or by how tabs and control characters are displayed.

Command goto-line line

This function moves point to the front of the lineth line, counting from line 1 at beginning of the buffer. If line is less than 1, it moves point to the beginning of the buffer. If line is greater than the number of lines in the buffer, it moves point to the end of the buffer—that is, the end of the last line of the buffer. This is the only case in which goto-line does not necessarily move to the beginning of a line.

If narrowing is in effect, then line still counts from the beginning of the buffer, but point cannot go outside the accessible portion. So goto-line moves point to the beginning or end of the accessible portion, if the line number specifies an inaccessible position.

The return value of goto-line is the difference between line and the line number of the line to which point actually was able to move (in the full buffer, before taking account of narrowing). Thus, the value is positive if the scan encounters the real end of the buffer. The value is zero if scan encounters the end of the accessible portion but not the real end of the buffer.

In an interactive call, line is the numeric prefix argument if one has been provided. Otherwise line is read in the minibuffer.

Command beginning-of-line &optional count buffer This function moves point to the beginning of the current line. With an argument count not nil or 1, it moves forward count−1 lines and then to the beginning of the line. buffer defaults to the current buffer if omitted.

If this function reaches the end of the buffer (or of the accessible portion, if narrowing is in effect), it positions point there. No error is signaled.

Command end-of-line &optional count buffer This function moves point to the end of the current line. With an argument count not nil or 1, it moves forward count−1 lines and then to the end of the line. buffer defaults to the current buffer if omitted.

If this function reaches the end of the buffer (or of the accessible portion, if narrowing is in effect), it positions point there. No error is signaled.

Command forward-line &optional count buffer This function moves point forward count lines, to the beginning of the line. If count is negative, it moves point −count lines backward, to the beginning of a line. If count is zero, it moves point to the beginning of the current line. buffer defaults to the current buffer if omitted.

If forward-line encounters the beginning or end of the buffer (or of the accessible portion) before finding that many lines, it sets point there. No error is signaled.

forward-line returns the difference between count and the number of lines actually moved. If you attempt to move down five lines from the beginning of a buffer that has only three lines, point stops at the end of the last line, and the value will be 2.

In an interactive call, count is the numeric prefix argument.

Function count-lines start end

This function returns the number of lines between the positions start and end in the current buffer. If start and end are equal, then it returns 0. Otherwise it returns at least 1, even if start and end are on the same line. This is because the text between them, considered in isolation, must contain at least one line unless it is empty.

Here is an example of using count-lines:

     (defun current-line ()
       "Return the vertical position of point..."
       (+ (count-lines (window-start) (point))
          (if (= (current-column) 0) 1 0)
          -1))

Also see the functions bolp and eolp in Section 36.1 [Near Point], page 511. These functions do not move point, but test whether it is already at the beginning or end of a line.

34.2.5 Motion by Screen Lines

The line functions in the previous section count text lines, delimited only by newline characters. By contrast, these functions count screen lines, which are defined by the way the text appears on the screen. A text line is a single screen line if it is short enough to fit the width of the selected window, but otherwise it may occupy several screen lines.

In some cases, text lines are truncated on the screen rather than continued onto additional screen lines. In these cases, vertical-motion moves point much like forward-line. See Section 45.2 [Truncation], page 648.

Because the width of a given string depends on the flags that control the appearance of certain characters, vertical-motion behaves differently, for a given piece of text, depending on the buffer it is in, and even on the selected window (because the width, the truncation flag, and display table may vary between windows). See Section 45.10 [Usual Display], page 658.

These functions scan text to determine where screen lines break, and thus take time proportional to the distance scanned. If you intend to use them heavily, Emacs provides caches which may improve the performance of your code. See Section 34.2.4 [Text Lines], page 490.

Function vertical-motion count &optional window pixels This function moves point to the start of the frame line count frame lines down from the frame line containing point. If count is negative, it moves up instead. The optional second argument window may be used to specify a window other than the selected window in which to perform the motion. Normally, vertical-motion returns the number of lines moved. The value may be less in absolute value than count if the beginning or end of the buffer was reached. If the optional third argument, pixels is non-nil, the vertical pixel height of the motion which took place is returned instead of the actual number of lines moved. A motion of zero lines returns the height of the current line. Note that vertical-motion sets window’s buffer’s point, not window’s point. (This differs from FSF Emacs, which buggily always sets current buffer’s point, regardless of window.)

Function vertical-motion-pixels count &optional window how This function moves point to the start of the frame line pixels vertical pixels down from the frame line containing point, or up if pixels is negative. The optional second argument window is the window to move in, and defaults to the selected window. The optional third argument how specifies the stopping condition. A negative integer indicates that the motion should be no more than pixels. A positive value indicates that the motion should be at least pixels. Any other value indicates that the motion should be as close as possible to pixels.

Command move-to-window-line count &optional window This function moves point with respect to the text currently displayed in window, which defaults to the selected window. It moves point to the beginning of the screen line count screen lines from the top of the window. If count is negative, that specifies a position −count lines from the bottom (or the last line of the buffer, if the buffer ends above the specified screen position). If count is nil, then point moves to the beginning of the line in the middle of the window. If the absolute value of count is greater than the size of the window, then point moves to the place that would appear on that screen line if the window were tall enough. This will probably cause the next redisplay to scroll to bring that location onto the screen. In an interactive call, count is the numeric prefix argument. The value returned is the window line number point has moved to, with the top line in the window numbered 0.

34.2.6 Moving over Balanced Expressions

Here are several functions concerned with balanced-parenthesis expressions (also called sexps in connection with moving across them in XEmacs). The syntax table controls how these functions interpret various characters; see Chapter 38 [Syntax Tables], page 569. See Section 38.5 [Parsing Expressions], page 575, for lower-level primitives for scanning sexps or parts of sexps. For user-level commands, see section “Lists and Sexps” in XEmacs Reference Manual.

Command forward-list &optional arg

This function moves forward across arg balanced groups of parentheses. (Other syntactic entities such as words or paired string quotes are ignored.) arg defaults to 1 if omitted. If arg is negative, move backward across that many groups of parentheses.

Command backward-list &optional arg

This function moves backward across arg balanced groups of parentheses. (Other syntactic entities such as words or paired string quotes are ignored.) arg defaults to 1 if omitted. If arg is negative, move forward across that many groups of parentheses.

Command up-list arg

This function moves forward out of arg levels of parentheses. A negative argument means move backward but still to a less deep spot.

Command down-list arg

This function moves forward into arg levels of parentheses. A negative argument means move backward but still go deeper in parentheses (−arg levels).

Command forward-sexp &optional arg

This function moves forward across arg balanced expressions. Balanced expressions include both those delimited by parentheses and other kinds, such as words and string constants. arg defaults to 1 if omitted. If arg is negative, move backward across that many balanced expressions. For example,

     ---------- Buffer: foo ----------
     (concat⋆ "foo " (car x) y z)
     ---------- Buffer: foo ----------

     (forward-sexp 3)
          ⇒ nil
     ---------- Buffer: foo ----------
     (concat "foo " (car x) y⋆ z)
     ---------- Buffer: foo ----------

Command backward-sexp &optional arg

This function moves backward across arg balanced expressions. arg defaults to 1 if omitted. If arg is negative, move forward across that many balanced expressions.

Command beginning-of-defun &optional arg

This function moves back to the argth beginning of a defun. If arg is negative, this actually moves forward, but it still moves to the beginning of a defun, not to the end of one. arg defaults to 1 if omitted.

Command end-of-defun &optional arg

This function moves forward to the argth end of a defun. If arg is negative, this actually moves backward, but it still moves to the end of a defun, not to the beginning of one. arg defaults to 1 if omitted.

User Option defun-prompt-regexp

If non-nil, this variable holds a regular expression that specifies what text can appear before the open-parenthesis that starts a defun. That is to say, a defun begins on a line that starts with a match for this regular expression, followed by a character with open-parenthesis syntax.

34.2.7 Skipping Characters

The following two functions move point over a specified set of characters. For example, they are often used to skip whitespace. For related functions, see Section 38.4 [Motion and Syntax], page 575.

Function skip-chars-forward character-set &optional limit buffer This function moves point in buffer forward, skipping over a given set of characters. It examines the character following point, then advances point if the character matches character-set. This continues until it reaches a character that does not match. The function returns nil. buffer defaults to the current buffer if omitted. The argument character-set is like the inside of a ‘[...]’ in a regular expression except that ‘]’ is never special and ‘\’ quotes ‘^’, ‘-’ or ‘\’. Thus, "a-zA-Z" skips over all letters, stopping before the first non-letter, and "^a-zA-Z" skips nonletters stopping before the first letter. See Section 37.2 [Regular Expressions], page 550. If limit is supplied (it must be a number or a marker), it specifies the maximum position in the buffer that point can be skipped to. Point will stop at or before limit.

In the following example, point is initially located directly before the ‘T’. After the form is evaluated, point is located at the end of that line (between the ‘t’ of ‘hat’ and the newline). The function skips all letters and spaces, but not newlines.

     ---------- Buffer: foo ----------
     I read "⋆The cat in the hat
     comes back" twice.
     ---------- Buffer: foo ----------
     (skip-chars-forward "a-zA-Z ")
          ⇒ nil
     ---------- Buffer: foo ----------
     I read "The cat in the hat⋆
     comes back" twice.
     ---------- Buffer: foo ----------

Function skip-chars-backward character-set &optional limit buffer This function moves point backward, skipping characters that match characterset, until limit. It just like skip-chars-forward except for the direction of motion.

34.3 Excursions

It is often useful to move point “temporarily” within a localized portion of the program, or to switch buffers temporarily. This is called an excursion, and it is done with the saveexcursion special form. This construct saves the current buffer and its values of point and the mark so they can be restored after the completion of the excursion.

The forms for saving and restoring the configuration of windows are described elsewhere (see Section 31.16 [Window Configurations], page 468 and see Section 32.11 [Frame Configurations], page 481).

Special Form save-excursion forms…

The save-excursion special form saves the identity of the current buffer and the values of point and the mark in it, evaluates forms, and finally restores the buffer and its saved values of point and the mark. All three saved values are restored even in case of an abnormal exit via throw or error (see Section 9.5 [Nonlocal Exits], page 134). The save-excursion special form is the standard way to switch buffers or move point within one part of a program and avoid affecting the rest of the program. It is used more than 500 times in the Lisp sources of XEmacs. save-excursion does not save the values of point and the mark for other buffers, so changes in other buffers remain in effect after save-excursion exits. Likewise, save-excursion does not restore window-buffer correspondences altered by functions such as switch-to-buffer. One way to restore these correspondences, and the selected window, is to use save-window-excursion inside save-excursion (see Section 31.16 [Window Configurations], page 468).

The value returned by save-excursion is the result of the last of forms, or nil if no forms are given.

     (save-excursion

forms) ≡

     (let ((old-buf (current-buffer))
           (old-pnt (point-marker))
           (old-mark (copy-marker (mark-marker))))
       (unwind-protect
           (progn forms)
         (set-buffer old-buf)
         (goto-char old-pnt)
         (set-marker (mark-marker) old-mark)))

Special Form save-current-buffer forms…

This special form is similar to save-excursion but it only saves and restores the current buffer. Beginning with XEmacs 20.3, save-current-buffer is a primitive.

Special Form with-current-buffer buffer forms…

This special form evaluates forms with buffer as the current buffer. It returns the value of the last form.

Special Form with-temp-file file forms…

This special form creates a new buffer, evaluates forms there, and writes the buffer to file. It returns the value of the last form evaluated.

Special Form save-selected-window forms…

This special form is similar to save-excursion but it saves and restores the selected window and nothing else.

34.4 Narrowing

Narrowing means limiting the text addressable by XEmacs editing commands to a limited range of characters in a buffer. The text that remains addressable is called the accessible portion of the buffer.

Narrowing is specified with two buffer positions which become the beginning and end of the accessible portion. For most editing commands and most Emacs primitives, these positions replace the values of the beginning and end of the buffer. While narrowing is in effect, no text outside the accessible portion is displayed, and point cannot move outside the accessible portion.

Values such as positions or line numbers, which usually count from the beginning of the buffer, do so despite narrowing, but the functions which use them refuse to operate on text that is inaccessible.

The commands for saving buffers are unaffected by narrowing; they save the entire buffer regardless of any narrowing.

Command narrow-to-region start end &optional buffer This function sets the accessible portion of buffer to start at start and end at end. Both arguments should be character positions. buffer defaults to the current buffer if omitted. In an interactive call, start and end are set to the bounds of the current region (point and the mark, with the smallest first).

Command narrow-to-page &optional move-count This function sets the accessible portion of the current buffer to include just the current page. An optional first argument move-count non-nil means to move forward or backward by move-count pages and then narrow. The variable page-delimiter specifies where pages start and end (see Section 37.8 [Standard Regexps], page 566). In an interactive call, move-count is set to the numeric prefix argument.

Command widen &optional buffer This function cancels any narrowing in buffer, so that the entire contents are accessible. This is called widening. It is equivalent to the following expression:

     (narrow-to-region 1 (1+ (buffer-size)))

buffer defaults to the current buffer if omitted.

Special Form save-restriction body…

This special form saves the current bounds of the accessible portion, evaluates the body forms, and finally restores the saved bounds, thus restoring the same state of narrowing (or absence thereof) formerly in effect. The state of narrowing is restored even in the event of an abnormal exit via throw or error (see Section 9.5 [Nonlocal Exits], page 134). Therefore, this construct is a clean way to narrow a buffer temporarily. The value returned by save-restriction is that returned by the last form in body, or nil if no body forms were given. Caution: it is easy to make a mistake when using the save-restriction construct. Read the entire description here before you try it. If body changes the current buffer, save-restriction still restores the restrictions on the original buffer (the buffer whose restrictions it saved from), but it does not restore the identity of the current buffer. save-restriction does not restore point and the mark; use save-excursion for that. If you use both save-restriction and save-excursion together, save-excursion should come first (on the outside). Otherwise, the old point value would be restored with temporary narrowing still in effect. If the old point value were outside the limits of the temporary narrowing, this would fail to restore it accurately. The save-restriction special form records the values of the beginning and end of the accessible portion as distances from the beginning and end of the buffer. In other words, it records the amount of inaccessible text before and after the accessible portion.

This method yields correct results if body does further narrowing. However, save-restriction can become confused if the body widens and then make changes outside the range of the saved narrowing. When this is what you want to do, save-restriction is not the right tool for the job. Here is what you must use instead:

(let ((beg (point-min-marker))
      (end (point-max-marker)))
  (unwind-protect
      (progn body)
    (save-excursion
      (set-buffer (marker-buffer beg))
      (narrow-to-region beg end))))

Here is a simple example of correct use of save-restriction:

---------- Buffer: foo ----------
This is the contents of foo
This is the contents of foo
This is the contents of foo⋆
---------- Buffer: foo ----------
(save-excursion
  (save-restriction
    (goto-char 1)
    (forward-line 2)
    (narrow-to-region 1 (point))
    (goto-char (point-min))
    (replace-string "foo" "bar")))
---------- Buffer: foo ----------
This is the contents of bar
This is the contents of bar
This is the contents of foo⋆
---------- Buffer: foo ----------

35 Markers

A marker is a Lisp object used to specify a position in a buffer relative to the surrounding text. A marker changes its offset from the beginning of the buffer automatically whenever text is inserted or deleted, so that it stays with the two characters on either side of it.

35.1 Overview of Markers

A marker specifies a buffer and a position in that buffer. The marker can be used to represent a position in the functions that require one, just as an integer could be used. See Chapter 34 [Positions], page 487, for a complete description of positions.

A marker has two attributes: the marker position, and the marker buffer. The marker position is an integer that is equivalent (at a given time) to the marker as a position in that buffer. But the marker’s position value can change often during the life of the marker. Insertion and deletion of text in the buffer relocate the marker. The idea is that a marker positioned between two characters remains between those two characters despite insertion and deletion elsewhere in the buffer. Relocation changes the integer equivalent of the marker.

Deleting text around a marker’s position leaves the marker between the characters immediately before and after the deleted text. Inserting text at the position of a marker normally leaves the marker in front of the new text—unless it is inserted with insertbefore-markers (see Section 36.4 [Insertion], page 513).

Insertion and deletion in a buffer must check all the markers and relocate them if necessary. This slows processing in a buffer with a large number of markers. For this reason, it is a good idea to make a marker point nowhere if you are sure you don’t need it any more. Unreferenced markers are garbage collected eventually, but until then will continue to use time if they do point somewhere.

Because it is common to perform arithmetic operations on a marker position, most of the arithmetic operations (including + and -) accept markers as arguments. In such cases, the marker stands for its current position.

Note that you can use extents to achieve the same functionality, and more, as markers. (Markers were defined before extents, which is why they both continue to exist.) A zerolength extent with the detachable property removed is almost identical to a marker. (See Section 40.3 [Extent Endpoints], page 586, for more information on zero-length extents.)

In particular:

• In order to get marker-like behavior in a zero-length extent, the detachable property must be removed (otherwise, the extent will disappear when text near it is deleted) and exactly one endpoint must be closed (if both endpoints are closed, the extent will expand to contain text inserted where it is located).

• If a zero-length extent has the end-open property but not the start-open property (this is the default), text inserted at the extent’s location causes the extent to move forward, just like a marker.

• If a zero-length extent has the start-open property but not the end-open property, text inserted at the extent’s location causes the extent to remain before the text, like what happens to markers when insert-before-markers is used.

• Markers end up after or before inserted text depending on whether insert or insertbefore-markers was called. These functions do not affect zero-length extents differently; instead, the presence or absence of the start-open and end-open extent properties determines this, as just described.

• Markers are automatically removed from a buffer when they are no longer in use. Extents remain around until explicitly removed from a buffer.

• Many functions are provided for listing the extents in a buffer or in a region of a buffer. No such functions exist for markers.

Here are examples of creating markers, setting markers, and moving point to markers:

;; Make a new marker that initially does not point anywhere:

(setq m1 (make-marker))
     ⇒ #<marker in no buffer>

;; Set m1 to point between the 99th and 100th characters ;; in the current buffer:

(set-marker m1 100)
     ⇒ #<marker at 100 in markers.texi>

;; Now insert one character at the beginning of the buffer:

(goto-char (point-min))
     ⇒ 1
(insert "Q")
     ⇒ nil

;; m1 is updated appropriately.

m1
     ⇒ #<marker at 101 in markers.texi>

;; Two markers that point to the same position ;; are not eq, but they are equal.

(setq m2 (copy-marker m1))
     ⇒ #<marker at 101 in markers.texi>
(eq m1 m2)
     ⇒ nil
(equal m1 m2)
     ⇒ t

;; When you are finished using a marker, make it point nowhere.

(set-marker m1 nil)
     ⇒ #<marker in no buffer>

35.2 Predicates on Markers

You can test an object to see whether it is a marker, or whether it is either an integer or a marker or either an integer, a character, or a marker. The latter tests are useful in connection with the arithmetic functions that work with any of markers, integers, or characters.

Function markerp object

This function returns t if object is a marker, nil otherwise. Note that integers are not markers, even though many functions will accept either a marker or an integer.

Function integer-or-marker-p object

This function returns t if object is an integer or a marker, nil otherwise.

Function integer-char-or-marker-p object

This function returns t if object is an integer, a character, or a marker, nil otherwise.

Function number-or-marker-p object

This function returns t if object is a number (either kind) or a marker, nil otherwise.

Function number-char-or-marker-p object

This function returns t if object is a number (either kind), a character, or a marker, nil otherwise.

35.3 Functions That Create Markers

When you create a new marker, you can make it point nowhere, or point to the present position of point, or to the beginning or end of the accessible portion of the buffer, or to the same place as another given marker.

Function make-marker

This functions returns a newly created marker that does not point anywhere.

     (make-marker)
          ⇒ #<marker in no buffer>

Function point-marker &optional dont-copy-p buffer This function returns a marker that points to the present position of point in buffer, which defaults to the current buffer. See Section 34.1 [Point], page 487. For an example, see copy-marker, below.

Internally, a marker corresponding to point is always maintained. Normally the marker returned by point-marker is a copy; you may modify it with reckless abandon. However, if optional argument dont-copy-p is non-nil, then the real point-marker is returned; modifying the position of this marker will move point. It is illegal to change the buffer of it, or make it point nowhere.

Function point-min-marker &optional buffer

This function returns a new marker that points to the beginning of the accessible portion of buffer, which defaults to the current buffer. This will be the beginning of the buffer unless narrowing is in effect. See Section 34.4 [Narrowing], page 496.

Function point-max-marker &optional buffer

This function returns a new marker that points to the end of the accessible portion of buffer, which defaults to the current buffer. This will be the end of the buffer unless narrowing is in effect. See Section 34.4 [Narrowing], page 496.

Here are examples of this function and point-min-marker, shown in a buffer containing a version of the source file for the text of this chapter.

     (point-min-marker)
          ⇒ #<marker at 1 in markers.texi>
     (point-max-marker)
          ⇒ #<marker at 15573 in markers.texi>
     (narrow-to-region 100 200)
          ⇒ nil
     (point-min-marker)
          ⇒ #<marker at 100 in markers.texi>
     (point-max-marker)
          ⇒ #<marker at 200 in markers.texi>

Function copy-marker marker-or-integer If passed a marker as its argument, copy-marker returns a new marker that points to the same place and the same buffer as does marker-or-integer. If passed an integer as its argument, copy-marker returns a new marker that points to position marker-or-integer in the current buffer.

If passed an integer argument less than 1, copy-marker returns a new marker that points to the beginning of the current buffer. If passed an integer argument greater than the length of the buffer, copy-marker returns a new marker that points to the end of the buffer.

An error is signaled if marker is neither a marker nor an integer.

     (setq p (point-marker))
          ⇒ #<marker at 2139 in markers.texi>
     (setq q (copy-marker p))
          ⇒ #<marker at 2139 in markers.texi>
     (eq p q) ⇒ nil
     (equal p q) ⇒ t
     (point) ⇒ 2139
     (set-marker p 3000)
          ⇒ #<marker at 3000 in markers.texi>
     (point) ⇒ 2139
     (setq p (point-marker t))
          ⇒ #<marker at 2139 in markers.texi>

(set-marker p 3000)
     ⇒ #<marker at 3000 in markers.texi>
(point) ⇒ 3000
(copy-marker 0)
     ⇒ #<marker at 1 in markers.texi>
(copy-marker 20000)
     ⇒ #<marker at 7572 in markers.texi>

35.4 Information from Markers

This section describes the functions for accessing the components of a marker object.

Function marker-position marker

This function returns the position that marker points to, or nil if it points nowhere.

Function marker-buffer marker

This function returns the buffer that marker points into, or nil if it points nowhere.

(setq m (make-marker))
     ⇒ #<marker in no buffer>
(marker-position m)
     ⇒ nil
(marker-buffer m)
     ⇒ nil
(set-marker m 3770 (current-buffer))
     ⇒ #<marker at 3770 in markers.texi>
(marker-buffer m)
     ⇒ #<buffer markers.texi>
(marker-position m)
     ⇒ 3770

Two distinct markers are considered equal (even though not eq) to each other if they have the same position and buffer, or if they both point nowhere.

35.5 Changing Marker Positions

This section describes how to change the position of an existing marker. When you do this, be sure you know whether the marker is used outside of your program, and, if so, what effects will result from moving it—otherwise, confusing things may happen in other parts of Emacs.

Function set-marker marker position &optional buffer This function moves marker to position in buffer. If buffer is not provided, it defaults to the current buffer.

If position is less than 1, set-marker moves marker to the beginning of the buffer. If position is greater than the size of the buffer, set-marker moves marker to the end of the buffer. If position is nil or a marker that points nowhere, then marker is set to point nowhere. The value returned is marker.

          (setq m (point-marker))
               ⇒ #<marker at 4714 in markers.texi>
          (set-marker m 55)
               ⇒ #<marker at 55 in markers.texi>
          (setq b (get-buffer "foo"))
               ⇒ #<buffer foo>
          (set-marker m 0 b)
               ⇒ #<marker at 1 in foo>

Function move-marker marker position &optional buffer This is another name for set-marker.

35.6 The Mark

One special marker in each buffer is designated the mark. It records a position for the user for the sake of commands such as C-w and C-x ^{⟨}\underline{\overline{TAB}}^{⟩}. Lisp programs should set the mark only to values that have a potential use to the user, and never for their own internal purposes. For example, the replace-regexp command sets the mark to the value of point before doing any replacements, because this enables the user to move back there conveniently after the replace is finished.

Once the mark “exists” in a buffer, it normally never ceases to exist. However, it may become inactive, and usually does so after each command (other than simple motion commands and some commands that explicitly activate the mark). When the mark is active, the region between point and the mark is called the active region and is highlighted specially.

Many commands are designed so that when called interactively they operate on the text between point and the mark. Such commands work only when an active region exists, i.e. when the mark is active. (The reason for this is to prevent you from accidentally deleting or changing large chunks of your text.) If you are writing such a command, don’t examine the mark directly; instead, use interactive with the ‘r’ specification. This provides the values of point and the mark as arguments to the command in an interactive call, but permits other Lisp programs to specify arguments explicitly, and automatically signals an error if the command is called interactively when no active region exists. See Section 19.2.2 [Interactive Codes], page 283.

Each buffer has its own value of the mark that is independent of the value of the mark in other buffers. (When a buffer is created, the mark exists but does not point anywhere. We consider this state as “the absence of a mark in that buffer.”) However, only one active region can exist at a time. Activating the mark in one buffer automatically deactivates an active mark in any other buffer. Note that the user can explicitly activate a mark at any time by using the command activate-region (normally bound to M-C-z) or by using the command exchange-point-and-mark (normally bound to C-x C-x), which has the side effect of activating the mark.

Some people do not like active regions, so they disable this behavior by setting the variable zmacs-regions to nil. This makes the mark always active (except when a buffer is just created and the mark points nowhere), and turns off the highlighting of the region between point and the mark. Commands that explicitly retrieve the value of the mark should make sure that they behave correctly and consistently irrespective of the setting of zmacs-regions; some primitives are provided to ensure this behavior.

In addition to the mark, each buffer has a mark ring which is a list of markers containing previous values of the mark. When editing commands change the mark, they should normally save the old value of the mark on the mark ring. The variable mark-ring-max specifies the maximum number of entries in the mark ring; once the list becomes this long, adding a new element deletes the last element.

Function mark &optional force buffer This function returns buffer’s mark position as an integer. buffer defaults to the current buffer if omitted. If the mark is inactive, mark normally returns nil. However, if force is non-nil, then mark returns the mark position anyway—or nil, if the mark is not yet set for the buffer. (Remember that if zmacs-regions is nil, the mark is always active as long as it exists, and the force argument will have no effect.) If you are using this in an editing command, you are most likely making a mistake; see the documentation of set-mark below.

Function mark-marker inactive-p buffer This function returns buffer’s mark. buffer defaults to the current buffer if omitted. This is the very marker that records the mark location inside XEmacs, not a copy. Therefore, changing this marker’s position will directly affect the position of the mark. Don’t do it unless that is the effect you want. If the mark is inactive, mark-marker normally returns nil. However, if force is non-nil, then mark-marker returns the mark anyway.

     (setq m (mark-marker))
          ⇒ #<marker at 3420 in markers.texi>
     (set-marker m 100)
          ⇒ #<marker at 100 in markers.texi>
     (mark-marker)
          ⇒ #<marker at 100 in markers.texi>

Like any marker, this marker can be set to point at any buffer you like. We don’t recommend that you make it point at any buffer other than the one of which it is the mark. If you do, it will yield perfectly consistent, but rather odd, results.

Function set-mark position &optional buffer This function sets buffer’s mark to position, and activates the mark. buffer defaults to the current buffer if omitted. The old value of the mark is not pushed onto the mark ring.

Please note: Use this function only if you want the user to see that the mark has moved, and you want the previous mark position to be lost. Normally, when a new mark is set, the old one should go on the mark-ring. For this reason, most applications should use push-mark and pop-mark, not set-mark. Novice XEmacs Lisp programmers often try to use the mark for the wrong purposes. The mark saves a location for the user’s convenience. An editing command should not alter the mark unless altering the mark is part of the user-level functionality of the command. (And, in that case, this effect should be documented.) To remember a location for internal use in the Lisp program, store it in a Lisp variable. For example:

     (let ((beg (point)))
       (forward-line 1)
       (delete-region beg (point))).

Command exchange-point-and-mark &optional dont-activate-region This function exchanges the positions of point and the mark. It is intended for interactive use. The mark is also activated unless dont-activate-region is non-nil.

Function push-mark &optional position nomsg activate buffer This function sets buffer’s mark to position, and pushes a copy of the previous mark onto mark-ring. buffer defaults to the current buffer if omitted. If position is nil, then the value of point is used. push-mark returns nil. If the last global mark pushed was not in buffer, also push position on the global mark ring (see below). The function push-mark normally does not activate the mark. To do that, specify t for the argument activate. A ‘Mark set’ message is displayed unless nomsg is non-nil.

Function pop-mark

This function pops off the top element of mark-ring and makes that mark become the buffer’s actual mark. This does not move point in the buffer, and it does nothing if mark-ring is empty. It deactivates the mark. The return value is not meaningful.

Variable mark-ring

The value of this buffer-local variable is the list of saved former marks of the current buffer, most recent first.

     mark-ring
     ⇒ (#<marker at 11050 in markers.texi>
         #<marker at 10832 in markers.texi>
         ...)

User Option mark-ring-max

The value of this variable is the maximum size of mark-ring. If more marks than this are pushed onto the mark-ring, push-mark discards an old mark when it adds a new one.

In additional to a per-buffer mark ring, there is a global mark ring. Marks are pushed onto the global mark ring the first time you set a mark after switching buffers.

Variable global-mark-ring

The value of this variable is the list of saved former global marks, most recent first.

User Option mark-ring-max

The value of this variable is the maximum size of global-mark-ring. If more marks than this are pushed onto the global-mark-ring, push-mark discards an old mark when it adds a new one.

Command pop-global-mark

This function pops a mark off the global mark ring and jumps to that location.

35.7 The Region

The text between point and the mark is known as the region. Various functions operate on text delimited by point and the mark, but only those functions specifically related to the region itself are described here.

When zmacs-regions is non-nil (this is the default), the concept of an active region exists. The region is active when the corresponding mark is active. Note that only one active region at a time can exist – i.e. only one buffer’s region is active at a time. See Section 35.6 [The Mark], page 504, for more information about active regions.

User Option zmacs-regions

If non-nil (the default), active regions are used. See Section 35.6 [The Mark], page 504, for a detailed explanation of what this means.

A number of functions are provided for explicitly determining the bounds of the region and whether it is active. Few programs need to use these functions, however. A command designed to operate on a region should normally use interactive with the ‘r’ specification to find the beginning and end of the region. This lets other Lisp programs specify the bounds explicitly as arguments and automatically respects the user’s setting for zmacs-regions. (See Section 19.2.2 [Interactive Codes], page 283.)

Function region-beginning &optional buffer

This function returns the position of the beginning of buffer’s region (as an integer). This is the position of either point or the mark, whichever is smaller. buffer defaults to the current buffer if omitted. If the mark does not point anywhere, an error is signaled. Note that this function ignores whether the region is active.

Function region-end &optional buffer This function returns the position of the end of buffer’s region (as an integer). This is the position of either point or the mark, whichever is larger. buffer defaults to the current buffer if omitted. If the mark does not point anywhere, an error is signaled. Note that this function ignores whether the region is active.

Function region-exists-p

This function is non-nil if the region exists. If active regions are in use (i.e. zmacs-regions is true), this means that the region is active. Otherwise, this means that the user has pushed a mark in this buffer at some point in the past. If this function returns nil, a function that uses the ‘r’ interactive specification will cause an error when called interactively.

Function region-active-p

If zmacs-regions is true, this is equivalent to region-exists-p. Otherwise, this function always returns false. This function is used by commands such as

fill-paragraph-or-region and capitalize-region-or-word, which operate

either on the active region or on something else (e.g. the word or paragraph at point).

Variable zmacs-region-stays

If a command sets this variable to true, the currently active region will remain activated when the command finishes. (Normally the region is deactivated when each command terminates.) If zmacs-regions is false, however, this has no effect. Under normal circumstances, you do not need to set this; use the interactive specification ‘_’ instead, if you want the region to remain active.

Function zmacs-activate-region

This function activates the region in the current buffer (this is equivalent to activating the current buffer’s mark). This will normally also highlight the text in the active region and set zmacs-region-stays to t. (If zmacs-regions is false, however, this function has no effect.)

Function zmacs-deactivate-region

This function deactivates the region in the current buffer (this is equivalent to deactivating the current buffer’s mark). This will normally also unhighlight the text in the active region and set zmacs-region-stays to nil. (If zmacs-regions is false, however, this function has no effect.)

Function zmacs-update-region

This function updates the active region, if it’s currently active. (If there is no active region, this function does nothing.) This has the effect of updating the highlighting on the text in the region; but you should never need to call this except under rather strange circumstances. The command loop automatically calls it when appropriate. Calling this function will call the hook

zmacs-update-region-hook, if the region is active.

Variable zmacs-activate-region-hook

This normal hook is called when a region becomes active. (Usually this happens as a result of a command that activates the region, such as set-mark-

command, activate-region, or exchange-point-and-mark.) Note that call-

ing ‘zmacs-activate-region’ will call this hook, even if the region is already active. If zmacs-regions is false, however, this hook will never get called under any circumstances.

Variable zmacs-deactivate-region-hook

This normal hook is called when an active region becomes inactive. (Calling ‘zmacs-deactivate-region’ when the region is inactive will not cause this hook to be called.) If zmacs-regions is false, this hook will never get called.

Variable zmacs-update-region-hook

This normal hook is called when an active region is "updated" by zmacsupdate-region. This normally gets called at the end of each command that sets zmacs-region-stays to t, indicating that the region should remain activated. The motion commands do this.

510 XEmacs Lisp Reference Manual

36 Text

This chapter describes the functions that deal with the text in a buffer. Most examine, insert, or delete text in the current buffer, often in the vicinity of point. Many are interactive. All the functions that change the text provide for undoing the changes (see Section 36.9 [Undo], page 523).

Many text-related functions operate on a region of text defined by two buffer positions passed in arguments named start and end. These arguments should be either markers (see Chapter 35 [Markers], page 499) or numeric character positions (see Chapter 34 [Positions], page 487). The order of these arguments does not matter; it is all right for start to be the end of the region and end the beginning. For example, (delete-region 1 10) and (delete-region 10 1) are equivalent. An args-out-of-range error is signaled if either start or end is outside the accessible portion of the buffer. In an interactive call, point and the mark are used for these arguments.

Throughout this chapter, “text” refers to the characters in the buffer, together with their properties (when relevant).

36.1 Examining Text Near Point

Many functions are provided to look at the characters around point. Several simple functions are described here. See also looking-at in Section 37.3 [Regexp Search], page 557.

Many of these functions take an optional buffer argument. In all such cases, the current buffer will be used if this argument is omitted. (In FSF Emacs, and earlier versions of XEmacs, these functions usually did not have these optional buffer arguments and always operated on the current buffer.)

Function char-after position &optional buffer This function returns the character in the buffer at (i.e., immediately after) position position. If position is out of range for this purpose, either before the beginning of the buffer, or at or beyond the end, then the value is nil. If optional argument buffer is nil, the current buffer is assumed. In the following example, assume that the first character in the buffer is ‘@’:

          (char-to-string (char-after 1))
               ⇒ "@"

Function following-char &optional buffer

This function returns the character following point in the buffer. This is similar to (char-after (point)). However, if point is at the end of the buffer, then the result of following-char is 0. If optional argument buffer is nil, the current buffer is assumed. Remember that point is always between characters, and the terminal cursor normally appears over the character following point. Therefore, the character returned by following-char is the character the cursor is over. In this example, point is between the ‘a’ and the ‘c’.

     ---------- Buffer: foo ----------
     Gentlemen may cry ``Pea⋆ce! Peace!,''
     but there is no peace.
     ---------- Buffer: foo ----------
     (char-to-string (preceding-char))
          ⇒ "a"
     (char-to-string (following-char))
          ⇒ "c"

Function preceding-char &optional buffer

This function returns the character preceding point in the buffer. See above, under following-char, for an example. If point is at the beginning of the buffer, preceding-char returns 0. If optional argument buffer is nil, the current buffer is assumed.

Function bobp &optional buffer This function returns t if point is at the beginning of the buffer. If narrowing is in effect, this means the beginning of the accessible portion of the text. If optional argument buffer is nil, the current buffer is assumed. See also pointmin in Section 34.1 [Point], page 487.

Function eobp &optional buffer This function returns t if point is at the end of the buffer. If narrowing is in effect, this means the end of accessible portion of the text. If optional argument buffer is nil, the current buffer is assumed. See also point-max in See Section 34.1 [Point], page 487.

Function bolp &optional buffer This function returns t if point is at the beginning of a line. If optional argument buffer is nil, the current buffer is assumed. See Section 34.2.4 [Text Lines], page 490. The beginning of the buffer (or its accessible portion) always counts as the beginning of a line.

Function eolp &optional buffer This function returns t if point is at the end of a line. The end of the buffer is always considered the end of a line. If optional argument buffer is nil, the current buffer is assumed. The end of the buffer (or of its accessible portion) is always considered the end of a line.

36.2 Examining Buffer Contents

This section describes two functions that allow a Lisp program to convert any portion of the text in the buffer into a string.

Function buffer-substring start end &optional buffer Function buffer-string start end &optional buffer

These functions are equivalent and return a string containing a copy of the text of the region defined by positions start and end in the buffer. If the arguments are not positions in the accessible portion of the buffer, buffer-substring signals an args-out-of-range error. If optional argument buffer is nil, the current buffer is assumed. If the region delineated by start and end contains duplicable extents, they will be remembered in the string. See Section 40.9 [Duplicable Extents], page 596. It is not necessary for start to be less than end; the arguments can be given in either order. But most often the smaller argument is written first.

     ---------- Buffer: foo ----------
     This is the contents of buffer foo
     ---------- Buffer: foo ----------
     (buffer-substring 1 10)
     ⇒ "This is t"
     (buffer-substring (point-max) 10)
     ⇒ "he contents of buffer foo
     "

36.3 Comparing Text

This function lets you compare portions of the text in a buffer, without copying them into strings first.

Function compare-buffer-substrings buffer1 start1 end1 buffer2 start2

end2 This function lets you compare two substrings of the same buffer or two different buffers. The first three arguments specify one substring, giving a buffer and two positions within the buffer. The last three arguments specify the other substring in the same way. You can use nil for buffer1, buffer2, or both to stand for the current buffer. The value is negative if the first substring is less, positive if the first is greater, and zero if they are equal. The absolute value of the result is one plus the index of the first differing characters within the substrings. This function ignores case when comparing characters if case-fold-search is non-nil. It always ignores text properties. Suppose the current buffer contains the text ‘foobarbar haha!rara!’; then in this example the two substrings are ‘rbar ’ and ‘rara!’. The value is 2 because the first substring is greater at the second character.

     (compare-buffer-substring nil 6 11 nil 16 21)
          ⇒ 2

36.4 Inserting Text

Insertion means adding new text to a buffer. The inserted text goes at point—between the character before point and the character after point.

Insertion relocates markers that point at positions after the insertion point, so that they stay with the surrounding text (see Chapter 35 [Markers], page 499). When a marker points at the place of insertion, insertion normally doesn’t relocate the marker, so that it points to the beginning of the inserted text; however, certain special functions such as insert-before-markers relocate such markers to point after the inserted text.

Some insertion functions leave point before the inserted text, while other functions leave it after. We call the former insertion after point and the latter insertion before point.

If a string with non-nil extent data is inserted, the remembered extents will also be inserted. See Section 40.9 [Duplicable Extents], page 596.

Insertion functions signal an error if the current buffer is read-only. These functions copy text characters from strings and buffers along with their properties.

The inserted characters have exactly the same properties as the characters they were copied from. By contrast, characters specified as separate arguments, not part of a string or buffer, inherit their text properties from the neighboring text.

Function insert &rest args This function inserts the strings and/or characters args into the current buffer, at point, moving point forward. In other words, it inserts the text before point. An error is signaled unless all args are either strings or characters. The value

is nil.

Function insert-before-markers &rest args

This function inserts the strings and/or characters args into the current buffer, at point, moving point forward. An error is signaled unless all args are either strings or characters. The value is nil. This function is unlike the other insertion functions in that it relocates markers initially pointing at the insertion point, to point after the inserted text.

Function insert-string string &optional buffer This function inserts string into buffer before point. buffer defaults to the current buffer if omitted. This function is chiefly useful if you want to insert a string in a buffer other than the current one (otherwise you could just use

insert).

Function insert-char character count &optional buffer This function inserts count instances of character into buffer before point. count must be a number, and character must be a character. The value is nil. If optional argument buffer is nil, the current buffer is assumed. (In FSF Emacs, the third argument is called inherit and refers to text properties.)

Function insert-buffer-substring from-buffer-or-name &optional start

end This function inserts a portion of buffer from-buffer-or-name (which must already exist) into the current buffer before point. The text inserted is the region from start and end. (These arguments default to the beginning and end of the accessible portion of that buffer.) This function returns nil.

In this example, the form is executed with buffer ‘bar’ as the current buffer. We assume that buffer ‘bar’ is initially empty.

     ---------- Buffer: foo ----------
     We hold these truths to be self-evident, that all
     ---------- Buffer: foo ----------
     (insert-buffer-substring "foo" 1 20)
          ⇒ nil
     ---------- Buffer: bar ----------
     We hold these truth⋆
     ---------- Buffer: bar ----------

36.5 User-Level Insertion Commands

This section describes higher-level commands for inserting text, commands intended primarily for the user but useful also in Lisp programs.

Command insert-buffer from-buffer-or-name This command inserts the entire contents of from-buffer-or-name (which must exist) into the current buffer after point. It leaves the mark after the inserted text. The value is nil.

Command self-insert-command count

This command inserts the last character typed; it does so count times, before point, and returns nil. Most printing characters are bound to this command. In routine use, self-insert-command is the most frequently called function in XEmacs, but programs rarely use it except to install it on a keymap. In an interactive call, count is the numeric prefix argument. This command calls auto-fill-function whenever that is non-nil and the character inserted is a space or a newline (see Section 36.13 [Auto Filling], page 529). This command performs abbrev expansion if Abbrev mode is enabled and the inserted character does not have word-constituent syntax. (See Chapter 39 [Abbrevs], page 579, and Section 38.2.1 [Syntax Class Table], page 570.) This is also responsible for calling blink-paren-function when the inserted character has close parenthesis syntax (see Section 45.9 [Blinking], page 658).

Command newline &optional number-of-newlines This command inserts newlines into the current buffer before point. If numberof-newlines is supplied, that many newline characters are inserted. This function calls auto-fill-function if the current column number is greater than the value of fill-column and number-of-newlines is nil. Typically what auto-fill-function does is insert a newline; thus, the overall result in this case is to insert two newlines at different places: one at point, and another earlier in the line. newline does not auto-fill if number-of-newlines is non-nil.

This command indents to the left margin if that is not zero. See Section 36.12 [Margins], page 528. The value returned is nil. In an interactive call, count is the numeric prefix argument.

Command split-line

This command splits the current line, moving the portion of the line after point down vertically so that it is on the next line directly below where it was before. Whitespace is inserted as needed at the beginning of the lower line, using the indent-to function. split-line returns the position of point. Programs hardly ever use this function.

Variable overwrite-mode

This variable controls whether overwrite mode is in effect: a non-nil value enables the mode. It is automatically made buffer-local when set in any fashion.

36.6 Deleting Text

Deletion means removing part of the text in a buffer, without saving it in the kill ring (see Section 36.8 [The Kill Ring], page 519). Deleted text can’t be yanked, but can be reinserted using the undo mechanism (see Section 36.9 [Undo], page 523). Some deletion functions do save text in the kill ring in some special cases.

All of the deletion functions operate on the current buffer, and all return a value of nil.

Function erase-buffer &optional buffer This function deletes the entire text of buffer, leaving it empty. If the buffer is read-only, it signals a buffer-read-only error. Otherwise, it deletes the text without asking for any confirmation. It returns nil. buffer defaults to the current buffer if omitted. Normally, deleting a large amount of text from a buffer inhibits further autosaving of that buffer “because it has shrunk”. However, erase-buffer does not do this, the idea being that the future text is not really related to the former text, and its size should not be compared with that of the former text.

Command delete-region start end &optional buffer This command deletes the text in buffer in the region defined by start and end. The value is nil. If optional argument buffer is nil, the current buffer is assumed.

Command delete-char count &optional killp This command deletes count characters directly after point, or before point if count is negative. If killp is non-nil, then it saves the deleted characters in the kill ring. In an interactive call, count is the numeric prefix argument, and killp is the unprocessed prefix argument. Therefore, if a prefix argument is supplied, the text is saved in the kill ring. If no prefix argument is supplied, then one character is deleted, but not saved in the kill ring. The value returned is always nil.

Command delete-backward-char count &optional killp

This command deletes count characters directly before point, or after point if count is negative. If killp is non-nil, then it saves the deleted characters in the kill ring.

In an interactive call, count is the numeric prefix argument, and killp is the unprocessed prefix argument. Therefore, if a prefix argument is supplied, the text is saved in the kill ring. If no prefix argument is supplied, then one character is deleted, but not saved in the kill ring.

The value returned is always nil.

Command backward-delete-char-untabify count &optional killp

This command deletes count characters backward, changing tabs into spaces. When the next character to be deleted is a tab, it is first replaced with the proper number of spaces to preserve alignment and then one of those spaces is deleted instead of the tab. If killp is non-nil, then the command saves the deleted characters in the kill ring.

Conversion of tabs to spaces happens only if count is positive. If it is negative, exactly −count characters after point are deleted.

In an interactive call, count is the numeric prefix argument, and killp is the unprocessed prefix argument. Therefore, if a prefix argument is supplied, the text is saved in the kill ring. If no prefix argument is supplied, then one character is deleted, but not saved in the kill ring.

The value returned is always nil.

36.7 User-Level Deletion Commands

This section describes higher-level commands for deleting text, commands intended primarily for the user but useful also in Lisp programs.

Command delete-horizontal-space

This function deletes all spaces and tabs around point. It returns nil.

In the following examples, we call delete-horizontal-space four times, once on each line, with point between the second and third characters on the line each time.

     ---------- Buffer: foo ----------
     I ⋆thought
     I ⋆  thought
     We⋆ thought
     Yo⋆u thought
     ---------- Buffer: foo ----------

     (delete-horizontal-space)   ; Four times.
          ⇒ nil
     ---------- Buffer: foo ----------
     Ithought
     Ithought
     Wethought
     You thought
     ---------- Buffer: foo ----------

Command delete-indentation &optional join-following-p This function joins the line point is on to the previous line, deleting any whitespace at the join and in some cases replacing it with one space. If join-following-p is non-nil, delete-indentation joins this line to the following line instead. The value is nil. If there is a fill prefix, and the second of the lines being joined starts with the prefix, then delete-indentation deletes the fill prefix before joining the lines. See Section 36.12 [Margins], page 528. In the example below, point is located on the line starting ‘events’, and it makes no difference if there are trailing spaces in the preceding line.

     ---------- Buffer: foo ----------
     When in the course of human
     ⋆  events, it becomes necessary
     ---------- Buffer: foo ----------
     (delete-indentation)
          ⇒ nil
     ---------- Buffer: foo ----------
     When in the course of human⋆ events, it becomes necessary
     ---------- Buffer: foo ----------

After the lines are joined, the function fixup-whitespace is responsible for deciding whether to leave a space at the junction.

Function fixup-whitespace

This function replaces all the white space surrounding point with either one space or no space, according to the context. It returns nil. At the beginning or end of a line, the appropriate amount of space is none. Before a character with close parenthesis syntax, or after a character with open parenthesis or expression-prefix syntax, no space is also appropriate. Otherwise, one space is appropriate. See Section 38.2.1 [Syntax Class Table], page 570. In the example below, fixup-whitespace is called the first time with point before the word ‘spaces’ in the first line. For the second invocation, point is directly after the ‘(’.

     ---------- Buffer: foo ----------
     This has too many  ⋆spaces
     This has too many spaces at the start of (⋆ this list)
     ---------- Buffer: foo ----------

          (fixup-whitespace)
               ⇒ nil
          (fixup-whitespace)
               ⇒ nil
          ---------- Buffer: foo ----------
          This has too many spaces
          This has too many spaces at the start of (this list)
          ---------- Buffer: foo ----------

Command just-one-space

This command replaces any spaces and tabs around point with a single space.

It returns nil.

Command delete-blank-lines

This function deletes blank lines surrounding point. If point is on a blank line with one or more blank lines before or after it, then all but one of them are deleted. If point is on an isolated blank line, then it is deleted. If point is on a nonblank line, the command deletes all blank lines following it.

A blank line is defined as a line containing only tabs and spaces.

     delete-blank-lines returns nil.

36.8 The Kill Ring

Kill functions delete text like the deletion functions, but save it so that the user can reinsert it by yanking. Most of these functions have ‘kill-’ in their name. By contrast, the functions whose names start with ‘delete-’ normally do not save text for yanking (though they can still be undone); these are “deletion” functions.

Most of the kill commands are primarily for interactive use, and are not described here. What we do describe are the functions provided for use in writing such commands. You can use these functions to write commands for killing text. When you need to delete text for internal purposes within a Lisp function, you should normally use deletion functions, so as not to disturb the kill ring contents. See Section 36.6 [Deletion], page 516.

Killed text is saved for later yanking in the kill ring. This is a list that holds a number of recent kills, not just the last text kill. We call this a “ring” because yanking treats it as having elements in a cyclic order. The list is kept in the variable kill-ring, and can be operated on with the usual functions for lists; there are also specialized functions, described in this section, that treat it as a ring.

Some people think this use of the word “kill” is unfortunate, since it refers to operations that specifically do not destroy the entities “killed”. This is in sharp contrast to ordinary life, in which death is permanent and “killed” entities do not come back to life. Therefore, other metaphors have been proposed. For example, the term “cut ring” makes sense to people who, in pre-computer days, used scissors and paste to cut up and rearrange manuscripts. However, it would be difficult to change the terminology now.

36.8.1 Kill Ring Concepts

The kill ring records killed text as strings in a list, most recent first. A short kill ring, for example, might look like this:

("some text" "a different piece of text" "even older text")

When the list reaches kill-ring-max entries in length, adding a new entry automatically deletes the last entry.

When kill commands are interwoven with other commands, each kill command makes a new entry in the kill ring. Multiple kill commands in succession build up a single entry in the kill ring, which would be yanked as a unit; the second and subsequent consecutive kill commands add text to the entry made by the first one.

For yanking, one entry in the kill ring is designated the “front” of the ring. Some yank commands “rotate” the ring by designating a different element as the “front.” But this virtual rotation doesn’t change the list itself—the most recent entry always comes first in the list.

36.8.2 Functions for Killing

kill-region is the usual subroutine for killing text. Any command that calls this function is a “kill command” (and should probably have ‘kill’ in its name). kill-region puts the newly killed text in a new element at the beginning of the kill ring or adds it to the most recent element. It uses the last-command variable to determine whether the previous command was a kill command, and if so appends the killed text to the most recent entry.

Command kill-region start end

This function kills the text in the region defined by start and end. The text is deleted but saved in the kill ring, along with its text properties. The value is always nil.

In an interactive call, start and end are point and the mark.

If the buffer is read-only, kill-region modifies the kill ring just the same, then signals an error without modifying the buffer. This is convenient because it lets the user use all the kill commands to copy text into the kill ring from a read-only buffer.

Command copy-region-as-kill start end

This command saves the region defined by start and end on the kill ring (including text properties), but does not delete the text from the buffer. It returns nil. It also indicates the extent of the text copied by moving the cursor momentarily, or by displaying a message in the echo area.

The command does not set this-command to kill-region, so a subsequent kill command does not append to the same kill ring entry.

Don’t call copy-region-as-kill in Lisp programs unless you aim to support Emacs 18. For Emacs 19, it is better to use kill-new or kill-append instead. See Section 36.8.4 [Low-Level Kill Ring], page 521.

36.8.3 Functions for Yanking

Yanking means reinserting an entry of previously killed text from the kill ring. The text properties are copied too.

Command yank &optional arg This command inserts before point the text in the first entry in the kill ring. It positions the mark at the beginning of that text, and point at the end. If arg is a list (which occurs interactively when the user types C-u with no digits), then yank inserts the text as described above, but puts point before the yanked text and puts the mark after it. If arg is a number, then yank inserts the argth most recently killed text—the argth element of the kill ring list. yank does not alter the contents of the kill ring or rotate it. It returns nil.

Command yank-pop arg

This command replaces the just-yanked entry from the kill ring with a different entry from the kill ring. This is allowed only immediately after a yank or another yank-pop. At such a time, the region contains text that was just inserted by yanking. yank-pop deletes that text and inserts in its place a different piece of killed text. It does not add the deleted text to the kill ring, since it is already in the kill ring somewhere. If arg is nil, then the replacement text is the previous element of the kill ring. If arg is numeric, the replacement is the argth previous kill. If arg is negative, a more recent kill is the replacement. The sequence of kills in the kill ring wraps around, so that after the oldest one comes the newest one, and before the newest one goes the oldest. The value is always nil.

36.8.4 Low-Level Kill Ring

These functions and variables provide access to the kill ring at a lower level, but still convenient for use in Lisp programs. They take care of interaction with X Window selections. They do not exist in Emacs version 18.

Function current-kill n &optional do-not-move The function current-kill rotates the yanking pointer which designates the “front” of the kill ring by n places (from newer kills to older ones), and returns the text at that place in the ring. If the optional second argument do-not-move is non-nil, then current-kill doesn’t alter the yanking pointer; it just returns the nth kill, counting from the current yanking pointer. If n is zero, indicating a request for the latest kill, current-kill calls the value of interprogram-paste-function (documented below) before consulting the kill ring.

Function kill-new string

This function puts the text string into the kill ring as a new entry at the front of the ring. It discards the oldest entry if appropriate. It also invokes the value

of interprogram-cut-function (see below).

Function kill-append string before-p This function appends the text string to the first entry in the kill ring. Normally string goes at the end of the entry, but if before-p is non-nil, it goes at the beginning. This function also invokes the value of interprogram-cut-function (see below).

Variable interprogram-paste-function

This variable provides a way of transferring killed text from other programs, when you are using a window system. Its value should be nil or a function of no arguments.

If the value is a function, current-kill calls it to get the “most recent kill”. If the function returns a non-nil value, then that value is used as the “most recent kill”. If it returns nil, then the first element of kill-ring is used.

The normal use of this hook is to get the X server’s primary selection as the most recent kill, even if the selection belongs to another X client. See Section 51.1 [X Selections], page 713.

Variable interprogram-cut-function

This variable provides a way of communicating killed text to other programs, when you are using a window system. Its value should be nil or a function of one argument.

If the value is a function, kill-new and kill-append call it with the new first element of the kill ring as an argument.

The normal use of this hook is to set the X server’s primary selection to the newly killed text.

36.8.5 Internals of the Kill Ring

The variable kill-ring holds the kill ring contents, in the form of a list of strings. The most recent kill is always at the front of the list.

The kill-ring-yank-pointer variable points to a link in the kill ring list, whose car is the text to yank next. We say it identifies the “front” of the ring. Moving kill-ring-yank-pointer to a different link is called rotating the kill ring. We call the kill ring a “ring” because the functions that move the yank pointer wrap around from the end of the list to the beginning, or vice-versa. Rotation of the kill ring is virtual; it does not change the value of kill-ring.

Both kill-ring and kill-ring-yank-pointer are Lisp variables whose values are normally lists. The word “pointer” in the name of the kill-ring-yank-pointer indicates that the variable’s purpose is to identify one element of the list for use by the next yank command.

The value of kill-ring-yank-pointer is always eq to one of the links in the kill ring list. The element it identifies is the car of that link. Kill commands, which change the kill ring, also set this variable to the value of kill-ring. The effect is to rotate the ring so that the newly killed text is at the front.

Here is a diagram that shows the variable kill-ring-yank-pointer pointing to the sec-

ond entry in the kill ring ("some text" "a different piece of text" "yet older text").
     kill-ring       kill-ring-yank-pointer
       |               |
       |     ______--->  ____________
        --> |___|___|------> |___|___|--> |___|___|--> nil
              |                |            |
              |                |            |
              |                |             -->"yet older text"
              |                |
              |                 --> "a different piece of text"
              |
               --> "some text"

This state of affairs might occur after C-y (yank) immediately followed by M-y (yank-pop).

Variable kill-ring

This variable holds the list of killed text sequences, most recently killed first.

Variable kill-ring-yank-pointer

This variable’s value indicates which element of the kill ring is at the “front” of the ring for yanking. More precisely, the value is a tail of the value of killring, and its car is the kill string that C-y should yank.

User Option kill-ring-max

The value of this variable is the maximum length to which the kill ring can grow, before elements are thrown away at the end. The default value for kill-

     ring-max is 30.

36.9 Undo

Most buffers have an undo list, which records all changes made to the buffer’s text so that they can be undone. (The buffers that don’t have one are usually special-purpose buffers for which XEmacs assumes that undoing is not useful.) All the primitives that modify the text in the buffer automatically add elements to the front of the undo list, which is in the

variable buffer-undo-list.

Variable buffer-undo-list

This variable’s value is the undo list of the current buffer. A value of t disables the recording of undo information.

Here are the kinds of elements an undo list can have:

integer This kind of element records a previous value of point. Ordinary cursor motion does not get any sort of undo record, but deletion commands use these entries to record where point was before the command.

(beg . end)

This kind of element indicates how to delete text that was inserted. Upon insertion, the text occupied the range beg–end in the buffer.

(text . position)

This kind of element indicates how to reinsert text that was deleted. The deleted text itself is the string text. The place to reinsert it is (abs position).

(t high . low)

This kind of element indicates that an unmodified buffer became modified. The elements high and low are two integers, each recording 16 bits of the visited file’s modification time as of when it was previously visited or saved. primitiveundo uses those values to determine whether to mark the buffer as unmodified once again; it does so only if the file’s modification time matches those numbers.

(nil property value beg . end)

This kind of element records a change in a text property. Here’s how you might undo the change:

          (put-text-property beg end property value)

position This element indicates where point was at an earlier time. Undoing this element sets point to position. Deletion normally creates an element of this kind as well as a reinsertion element.

nil This element is a boundary. The elements between two boundaries are called a change group; normally, each change group corresponds to one keyboard command, and undo commands normally undo an entire group as a unit.

Function undo-boundary

This function places a boundary element in the undo list. The undo command stops at such a boundary, and successive undo commands undo to earlier and earlier boundaries. This function returns nil. The editor command loop automatically creates an undo boundary before each key sequence is executed. Thus, each undo normally undoes the effects of one command. Self-inserting input characters are an exception. The command loop makes a boundary for the first such character; the next 19 consecutive self-inserting input characters do not make boundaries, and then the 20th does, and so on as long as self-inserting characters continue. All buffer modifications add a boundary whenever the previous undoable change was made in some other buffer. This way, a command that modifies several buffers makes a boundary in each buffer it changes. Calling this function explicitly is useful for splitting the effects of a command into more than one unit. For example, query-replace calls undo-boundary after each replacement, so that the user can undo individual replacements one by one.

Function primitive-undo count list

This is the basic function for undoing elements of an undo list. It undoes the first count elements of list, returning the rest of list. You could write this function in Lisp, but it is convenient to have it in C. primitive-undo adds elements to the buffer’s undo list when it changes the buffer. Undo commands avoid confusion by saving the undo list value at the beginning of a sequence of undo operations. Then the undo operations use and update the saved value. The new elements added by undoing are not part of this saved value, so they don’t interfere with continuing to undo.

36.10 Maintaining Undo Lists

This section describes how to enable and disable undo information for a given buffer. It also explains how the undo list is truncated automatically so it doesn’t get too big.

Recording of undo information in a newly created buffer is normally enabled to start with; but if the buffer name starts with a space, the undo recording is initially disabled. You can explicitly enable or disable undo recording with the following two functions, or by

setting buffer-undo-list yourself.

Command buffer-enable-undo &optional buffer-or-name This command enables recording undo information for buffer buffer-or-name, so that subsequent changes can be undone. If no argument is supplied, then the current buffer is used. This function does nothing if undo recording is already enabled in the buffer. It returns nil. In an interactive call, buffer-or-name is the current buffer. You cannot specify any other buffer.

Function buffer-disable-undo &optional buffer Function buffer-flush-undo &optional buffer

This function discards the undo list of buffer, and disables further recording of undo information. As a result, it is no longer possible to undo either previous changes or any subsequent changes. If the undo list of buffer is already disabled, this function has no effect. This function returns nil. It cannot be called interactively. The name buffer-flush-undo is not considered obsolete, but the preferred name buffer-disable-undo is new as of Emacs versions 19.

As editing continues, undo lists get longer and longer. To prevent them from using up all available memory space, garbage collection trims them back to size limits you can set. (For this purpose, the “size” of an undo list measures the cons cells that make up the list, plus the strings of deleted text.) Two variables control the range of acceptable sizes:

undo-limit and undo-strong-limit.

Variable undo-limit

This is the soft limit for the acceptable size of an undo list. The change group at which this size is exceeded is the last one kept.

Variable undo-strong-limit

This is the upper limit for the acceptable size of an undo list. The change group at which this size is exceeded is discarded itself (along with all older change groups). There is one exception: the very latest change group is never discarded no matter how big it is.

36.11 Filling

Filling means adjusting the lengths of lines (by moving the line breaks) so that they are nearly (but no greater than) a specified maximum width. Additionally, lines can be justified, which means inserting spaces to make the left and/or right margins line up precisely. The width is controlled by the variable fill-column. For ease of reading, lines should be no longer than 70 or so columns.

You can use Auto Fill mode (see Section 36.13 [Auto Filling], page 529) to fill text automatically as you insert it, but changes to existing text may leave it improperly filled. Then you must fill the text explicitly.

Most of the commands in this section return values that are not meaningful. All the functions that do filling take note of the current left margin, current right margin, and current justification style (see Section 36.12 [Margins], page 528). If the current justification style is none, the filling functions don’t actually do anything.

Several of the filling functions have an argument justify. If it is non-nil, that requests some kind of justification. It can be left, right, full, or center, to request a specific style of justification. If it is t, that means to use the current justification style for this part

of the text (see current-justification, below).

When you call the filling functions interactively, using a prefix argument implies the value full for justify.

Command fill-paragraph justify

This command fills the paragraph at or after point. If justify is non-nil, each line is justified as well. It uses the ordinary paragraph motion commands to find paragraph boundaries. See section “Paragraphs” in The XEmacs User’s Manual.

Command fill-region start end &optional justify This command fills each of the paragraphs in the region from start to end. It justifies as well if justify is non-nil. The variable paragraph-separate controls how to distinguish paragraphs. See Section 37.8 [Standard Regexps], page 566.

Command fill-individual-paragraphs start end &optional justify

mail-flag This command fills each paragraph in the region according to its individual fill prefix. Thus, if the lines of a paragraph were indented with spaces, the filled paragraph will remain indented in the same fashion. The first two arguments, start and end, are the beginning and end of the region to be filled. The third and fourth arguments, justify and mail-flag, are optional. If justify is non-nil, the paragraphs are justified as well as filled. If mail-flag is non-nil, it means the function is operating on a mail message and therefore should not fill the header lines. Ordinarily, fill-individual-paragraphs regards each change in indentation

as starting a new paragraph. If fill-individual-varying-indent is non-nil,

then only separator lines separate paragraphs. That mode can handle indented paragraphs with additional indentation on the first line.

User Option fill-individual-varying-indent

This variable alters the action of fill-individual-paragraphs as described above.

Command fill-region-as-paragraph start end &optional justify This command considers a region of text as a paragraph and fills it. If the region was made up of many paragraphs, the blank lines between paragraphs are removed. This function justifies as well as filling when justify is non-nil. In an interactive call, any prefix argument requests justification. In Adaptive Fill mode, which is enabled by default, fill-region-as-paragraph on an indented paragraph when there is no fill prefix uses the indentation of the second line of the paragraph as the fill prefix.

Command justify-current-line how eop nosqueeze

This command inserts spaces between the words of the current line so that the line ends exactly at fill-column. It returns nil. The argument how, if non-nil specifies explicitly the style of justification. It can be left, right, full, center, or none. If it is t, that means to do follow specified justification style (see current-justification, below). nil means to do full justification. If eop is non-nil, that means do left-justification when current-justification specifies full justification. This is used for the last line of a paragraph; even if the paragraph as a whole is fully justified, the last line should not be. If nosqueeze is non-nil, that means do not change interior whitespace.

User Option default-justification

This variable’s value specifies the style of justification to use for text that doesn’t specify a style with a text property. The possible values are left, right, full, center, or none. The default value is left.

Function current-justification

This function returns the proper justification style to use for filling the text around point.

Variable fill-paragraph-function

This variable provides a way for major modes to override the filling of paragraphs. If the value is non-nil, fill-paragraph calls this function to do the work. If the function returns a non-nil value, fill-paragraph assumes the job is done, and immediately returns that value. The usual use of this feature is to fill comments in programming language modes. If the function needs to fill a paragraph in the usual way, it can do so as follows:

     (let ((fill-paragraph-function nil))
       (fill-paragraph arg))

Variable use-hard-newlines

If this variable is non-nil, the filling functions do not delete newlines that have the hard text property. These “hard newlines” act as paragraph separators.

36.12 Margins for Filling

User Option fill-prefix

This variable specifies a string of text that appears at the beginning of normal text lines and should be disregarded when filling them. Any line that fails to start with the fill prefix is considered the start of a paragraph; so is any line that starts with the fill prefix followed by additional whitespace. Lines that start with the fill prefix but no additional whitespace are ordinary text lines that can be filled together. The resulting filled lines also start with the fill prefix.

The fill prefix follows the left margin whitespace, if any.

User Option fill-column

This buffer-local variable specifies the maximum width of filled lines. Its value should be an integer, which is a number of columns. All the filling, justification and centering commands are affected by this variable, including Auto Fill mode (see Section 36.13 [Auto Filling], page 529).

As a practical matter, if you are writing text for other people to read, you should set fill-column to no more than 70. Otherwise the line will be too long for people to read comfortably, and this can make the text seem clumsy.

Variable default-fill-column

The value of this variable is the default value for fill-column in buffers that do not override it. This is the same as (default-value 'fill-column).

The default value for default-fill-column is 70.

Command set-left-margin from to margin

This sets the left-margin property on the text from from to to to the value margin. If Auto Fill mode is enabled, this command also refills the region to fit the new margin.

Command set-right-margin from to margin

This sets the right-margin property on the text from from to to to the value margin. If Auto Fill mode is enabled, this command also refills the region to fit the new margin.

Function current-left-margin

This function returns the proper left margin value to use for filling the text around point. The value is the sum of the left-margin property of the character at the start of the current line (or zero if none), and the value of the variable

left-margin.

Function current-fill-column

This function returns the proper fill column value to use for filling the text around point. The value is the value of the fill-column variable, minus the value of the right-margin property of the character after point.

Command move-to-left-margin &optional n force

This function moves point to the left margin of the current line. The column moved to is determined by calling the function current-left-margin. If the argument n is non-nil, move-to-left-margin moves forward n−1 lines first. If force is non-nil, that says to fix the line’s indentation if that doesn’t match the left margin value.

Function delete-to-left-margin from to

This function removes left margin indentation from the text between from and to. The amount of indentation to delete is determined by calling currentleft-margin. In no case does this function delete non-whitespace.

Function indent-to-left-margin

This is the default indent-line-function, used in Fundamental mode, Text mode, etc. Its effect is to adjust the indentation at the beginning of the current line to the value specified by the variable left-margin. This may involve either inserting or deleting whitespace.

Variable left-margin

This variable specifies the base left margin column. In Fundamental mode, ⟨\underline{\overline{LFD}}^{⟩} indents to this column. This variable automatically becomes buffer-local when set in any fashion.

36.13 Auto Filling

Auto Fill mode is a minor mode that fills lines automatically as text is inserted. This section describes the hook used by Auto Fill mode. For a description of functions that you can call explicitly to fill and justify existing text, see Section 36.11 [Filling], page 526.

Auto Fill mode also enables the functions that change the margins and justification style to refill portions of the text. See Section 36.12 [Margins], page 528.

Variable auto-fill-function

The value of this variable should be a function (of no arguments) to be called after self-inserting a space or a newline. It may be nil, in which case nothing special is done in that case. The value of auto-fill-function is do-auto-fill when Auto-Fill mode is enabled. That is a function whose sole purpose is to implement the usual strategy for breaking a line.

In older Emacs versions, this variable was named auto-fill-hook, but since it is not called with the standard convention for hooks, it was renamed to auto-fill-function in version 19.

36.14 Sorting Text

The sorting functions described in this section all rearrange text in a buffer. This is in contrast to the function sort, which rearranges the order of the elements of a list (see Section 5.6.3 [Rearrangement], page 87). The values returned by these functions are not meaningful.

Function sort-subr reverse nextrecfun endrecfun &optional startkeyfun

endkeyfun This function is the general text-sorting routine that divides a buffer into records and sorts them. Most of the commands in this section use this function.

To understand how sort-subr works, consider the whole accessible portion of the buffer as being divided into disjoint pieces called sort records. The records may or may not be contiguous; they may not overlap. A portion of each sort record (perhaps all of it) is designated as the sort key. Sorting rearranges the records in order by their sort keys.

Usually, the records are rearranged in order of ascending sort key. If the first argument to the sort-subr function, reverse, is non-nil, the sort records are rearranged in order of descending sort key.

The next four arguments to sort-subr are functions that are called to move point across a sort record. They are called many times from within sort-subr.

1. nextrecfun is called with point at the end of a record. This function moves point to the start of the next record. The first record is assumed to start at the position of point when sort-subr is called. Therefore, you should usually move point to the beginning of the buffer before calling sort-subr.

This function can indicate there are no more sort records by leaving point at the end of the buffer.

2. endrecfun is called with point within a record. It moves point to the end of the record.

3. startkeyfun is called to move point from the start of a record to the start of the sort key. This argument is optional; if it is omitted, the whole record is the sort key. If supplied, the function should either return a non-nil value to be used as the sort key, or return nil to indicate that the sort key is in the buffer starting at point. In the latter case, endkeyfun is called to find the end of the sort key.

4. endkeyfun is called to move point from the start of the sort key to the end of the sort key. This argument is optional. If startkeyfun returns nil and this argument is omitted (or nil), then the sort key extends to the end of the record. There is no need for endkeyfun if startkeyfun returns a non-nil value.

As an example of sort-subr, here is the complete function definition for sort-

lines:

;; Note that the first two lines of doc string ;; are effectively one line when viewed by a user.

 (defun sort-lines (reverse beg end)
   "Sort lines in region alphabetically.
 Called from a program, there are three arguments:

     REVERSE (non-nil means reverse order),
     and BEG and END (the region to sort)."
       (interactive "P\nr")
       (save-restriction
         (narrow-to-region beg end)
         (goto-char (point-min))
         (sort-subr reverse
                    'forward-line
                    'end-of-line)))

Here forward-line moves point to the start of the next record, and endof-line moves point to the end of record. We do not pass the arguments startkeyfun and endkeyfun, because the entire record is used as the sort key. The sort-paragraphs function is very much the same, except that its sortsubr call looks like this:

     (sort-subr reverse
                (function
                 (lambda ()
                   (skip-chars-forward "\n \t\f")))
                'forward-paragraph)

Command sort-regexp-fields reverse record-regexp key-regexp start end This command sorts the region between start and end alphabetically as specified by record-regexp and key-regexp. If reverse is a negative integer, then sorting is in reverse order. Alphabetical sorting means that two sort keys are compared by comparing the first characters of each, the second characters of each, and so on. If a mismatch is found, it means that the sort keys are unequal; the sort key whose character is less at the point of first mismatch is the lesser sort key. The individual characters are compared according to their numerical values. Since Emacs uses the ASCII character set, the ordering in that set determines alphabetical order. The value of the record-regexp argument specifies how to divide the buffer into sort records. At the end of each record, a search is done for this regular expression, and the text that matches it is the next record. For example, the regular expression ‘^.+$’, which matches lines with at least one character besides a newline, would make each such line into a sort record. See Section 37.2 [Regular Expressions], page 550, for a description of the syntax and meaning of regular expressions. The value of the key-regexp argument specifies what part of each record is the sort key. The key-regexp could match the whole record, or only a part. In the latter case, the rest of the record has no effect on the sorted order of records, but it is carried along when the record moves to its new position. The key-regexp argument can refer to the text matched by a subexpression of record-regexp, or it can be a regular expression on its own. If key-regexp is:

‘\digit’ then the text matched by the digitth ‘\(...\)’ parenthesis grouping in record-regexp is the sort key.

‘\&’ then the whole record is the sort key.

a regular expression then sort-regexp-fields searches for a match for the regular expression within the record. If such a match is found, it is the sort key. If there is no match for key-regexp within a record then that record is ignored, which means its position in the buffer is not changed. (The other records may move around it.)

For example, if you plan to sort all the lines in the region by the first word on each line starting with the letter ‘f’, you should set record-regexp to ‘^.*$’ and set key-regexp to ‘\<f\w*\>’. The resulting expression looks like this:

     (sort-regexp-fields nil "^.*$" "\\<f\\w*\\>"
                         (region-beginning)
                         (region-end))

If you call sort-regexp-fields interactively, it prompts for record-regexp and key-regexp in the minibuffer.

Command sort-lines reverse start end This command alphabetically sorts lines in the region between start and end. If reverse is non-nil, the sort is in reverse order.

Command sort-paragraphs reverse start end

This command alphabetically sorts paragraphs in the region between start and end. If reverse is non-nil, the sort is in reverse order.

Command sort-pages reverse start end This command alphabetically sorts pages in the region between start and end. If reverse is non-nil, the sort is in reverse order.

Command sort-fields field start end This command sorts lines in the region between start and end, comparing them alphabetically by the fieldth field of each line. Fields are separated by whitespace and numbered starting from 1. If field is negative, sorting is by the −fieldth field from the end of the line. This command is useful for sorting tables.

Command sort-numeric-fields field start end

This command sorts lines in the region between start and end, comparing them numerically by the fieldth field of each line. The specified field must contain a number in each line of the region. Fields are separated by whitespace and numbered starting from 1. If field is negative, sorting is by the −fieldth field from the end of the line. This command is useful for sorting tables.

Command sort-columns reverse &optional beg end This command sorts the lines in the region between beg and end, comparing them alphabetically by a certain range of columns. The column positions of beg and end bound the range of columns to sort on. If reverse is non-nil, the sort is in reverse order.

One unusual thing about this command is that the entire line containing position beg, and the entire line containing position end, are included in the region sorted. Note that sort-columns uses the sort utility program, and so cannot work properly on text containing tab characters. Use M-x untabify to convert tabs to spaces before sorting.

36.15 Counting Columns

The column functions convert between a character position (counting characters from the beginning of the buffer) and a column position (counting screen characters from the beginning of a line).

A character counts according to the number of columns it occupies on the screen. This means control characters count as occupying 2 or 4 columns, depending upon the value of ctl-arrow, and tabs count as occupying a number of columns that depends on the value of tab-width and on the column where the tab begins. See Section 45.10 [Usual Display], page 658.

Column number computations ignore the width of the window and the amount of horizontal scrolling. Consequently, a column value can be arbitrarily high. The first (or leftmost) column is numbered 0.

Function current-column

This function returns the horizontal position of point, measured in columns, counting from 0 at the left margin. The column position is the sum of the widths of all the displayed representations of the characters between the start of the current line and point. For an example of using current-column, see the description of count-lines in Section 34.2.4 [Text Lines], page 490.

Function move-to-column column &optional force This function moves point to column in the current line. The calculation of column takes into account the widths of the displayed representations of the characters between the start of the line and point. If column column is beyond the end of the line, point moves to the end of the line. If column is negative, point moves to the beginning of the line. If it is impossible to move to column column because that is in the middle of a multicolumn character such as a tab, point moves to the end of that character. However, if force is non-nil, and column is in the middle of a tab, then moveto-column converts the tab into spaces so that it can move precisely to column column. Other multicolumn characters can cause anomalies despite force, since there is no way to split them. The argument force also has an effect if the line isn’t long enough to reach column column; in that case, it says to add whitespace at the end of the line to reach that column. If column is not an integer, an error is signaled. The return value is the column number actually moved to.

36.16 Indentation

The indentation functions are used to examine, move to, and change whitespace that is at the beginning of a line. Some of the functions can also change whitespace elsewhere on a line. Columns and indentation count from zero at the left margin.

36.16.1 Indentation Primitives

This section describes the primitive functions used to count and insert indentation. The functions in the following sections use these primitives.

Function current-indentation

This function returns the indentation of the current line, which is the horizontal position of the first nonblank character. If the contents are entirely blank, then this is the horizontal position of the end of the line.

Command indent-to column &optional minimum This function indents from point with tabs and spaces until column is reached. If minimum is specified and non-nil, then at least that many spaces are inserted even if this requires going beyond column. Otherwise the function does nothing if point is already beyond column. The value is the column at which the inserted indentation ends.

User Option indent-tabs-mode

If this variable is non-nil, indentation functions can insert tabs as well as spaces. Otherwise, they insert only spaces. Setting this variable automatically makes it local to the current buffer.

36.16.2 Indentation Controlled by Major Mode

An important function of each major mode is to customize the ^{⟨}\underline{\overline{TAB}}^{⟩} key to indent properly for the language being edited. This section describes the mechanism of the ^{⟨}\underline{\overline{TAB}}^{⟩} key and how to control it. The functions in this section return unpredictable values.

Variable indent-line-function

This variable’s value is the function to be used by^{⟨}\underline{\overline{TAB}}^{⟩} (and various commands) to indent the current line. The command indent-according-to-mode does no more than call this function.

In Lisp mode, the value is the symbol lisp-indent-line; in C mode, cindent-line; in Fortran mode, fortran-indent-line. In Fundamental mode, Text mode, and many other modes with no standard for indentation, the value is indent-to-left-margin (which is the default value).

Command indent-according-to-mode

This command calls the function in indent-line-function to indent the current line in a way appropriate for the current major mode.

Command indent-for-tab-command

This command calls the function in indent-line-function to indent the current line; except that if that function is indent-to-left-margin, it calls insert-tab instead. (That is a trivial command that inserts a tab character.)

Command newline-and-indent

This function inserts a newline, then indents the new line (the one following the newline just inserted) according to the major mode.

It does indentation by calling the current indent-line-function. In programming language modes, this is the same thing^{⟨}\underline{\overline{TAB}}^{⟩} does, but in some text modes, where ^{⟨}\underline{\overline{TAB}}^{⟩} inserts a tab, newline-and-indent indents to the column specified

by left-margin.

Command reindent-then-newline-and-indent

This command reindents the current line, inserts a newline at point, and then reindents the new line (the one following the newline just inserted).

This command does indentation on both lines according to the current major mode, by calling the current value of indent-line-function. In programming language modes, this is the same thing ^{⟨}\underline{\overline{TAB}}^{⟩} does, but in some text modes, where ^{⟨}\underline{\overline{TAB}}^{⟩} inserts a tab, reindent-then-newline-and-indent indents to the column specified by left-margin.

36.16.3 Indenting an Entire Region

This section describes commands that indent all the lines in the region. They return unpredictable values.

Command indent-region start end to-column This command indents each nonblank line starting between start (inclusive) and end (exclusive). If to-column is nil, indent-region indents each nonblank line by calling the current mode’s indentation function, the value of indent-line-

function.

If to-column is non-nil, it should be an integer specifying the number of columns of indentation; then this function gives each line exactly that much indentation, by either adding or deleting whitespace.

If there is a fill prefix, indent-region indents each line by making it start with the fill prefix.

Variable indent-region-function

The value of this variable is a function that can be used by indent-region as a short cut. You should design the function so that it will produce the same results as indenting the lines of the region one by one, but presumably faster.

If the value is nil, there is no short cut, and indent-region actually works line by line.

A short-cut function is useful in modes such as C mode and Lisp mode, where the indent-line-function must scan from the beginning of the function definition: applying it to each line would be quadratic in time. The short cut can update the scan information as it moves through the lines indenting them; this takes linear time. In a mode where indenting a line individually is fast, there is no need for a short cut. indent-region with a non-nil argument to-column has a different meaning and does not use this variable.

Command indent-rigidly start end count

This command indents all lines starting between start (inclusive) and end (exclusive) sideways by count columns. This “preserves the shape” of the affected region, moving it as a rigid unit. Consequently, this command is useful not only for indenting regions of unindented text, but also for indenting regions of formatted code. For example, if count is 3, this command adds 3 columns of indentation to each of the lines beginning in the region specified.

In Mail mode, C-c C-y (mail-yank-original) uses indent-rigidly to indent

the text copied from the message being replied to.

Function indent-code-rigidly start end columns &optional

nochange-regexp This is like indent-rigidly, except that it doesn’t alter lines that start within strings or comments. In addition, it doesn’t alter a line if nochange-regexp matches at the beginning of the line (if nochange-regexp is non-nil).

36.16.4 Indentation Relative to Previous Lines

This section describes two commands that indent the current line based on the contents of previous lines.

Command indent-relative &optional unindented-ok This command inserts whitespace at point, extending to the same column as the next indent point of the previous nonblank line. An indent point is a nonwhitespace character following whitespace. The next indent point is the first one at a column greater than the current column of point. For example, if point is underneath and to the left of the first non-blank character of a line of text, it moves to that column by inserting whitespace. If the previous nonblank line has no next indent point (i.e., none at a great enough column position), indent-relative either does nothing (if unindentedok is non-nil) or calls tab-to-tab-stop. Thus, if point is underneath and to the right of the last column of a short line of text, this command ordinarily moves point to the next tab stop by inserting whitespace. The return value of indent-relative is unpredictable. In the following example, point is at the beginning of the second line:

                 This line is indented twelve spaces.
     ⋆The quick brown fox jumped.

Evaluation of the expression (indent-relative nil) produces the following:

                 This line is indented twelve spaces.
                 ⋆The quick brown fox jumped.

In this example, point is between the ‘m’ and ‘p’ of ‘jumped’:

                 This line is indented twelve spaces.
     The quick brown fox jum⋆ped.

Evaluation of the expression (indent-relative nil) produces the following:

                 This line is indented twelve spaces.
     The quick brown fox jum ⋆ped.

Command indent-relative-maybe

This command indents the current line like the previous nonblank line. It calls indent-relative with t as the unindented-ok argument. The return value is unpredictable. If the previous nonblank line has no indent points beyond the current column, this command does nothing.

36.16.5 Adjustable “Tab Stops”

This section explains the mechanism for user-specified “tab stops” and the mechanisms that use and set them. The name “tab stops” is used because the feature is similar to that of the tab stops on a typewriter. The feature works by inserting an appropriate number of spaces and tab characters to reach the next tab stop column; it does not affect the display of tab characters in the buffer (see Section 45.10 [Usual Display], page 658). Note that the ⟨\underline{\overline{TAB}}^{⟩} character as input uses this tab stop feature only in a few major modes, such as Text mode.

Command tab-to-tab-stop

This command inserts spaces or tabs up to the next tab stop column defined by tab-stop-list. It searches the list for an element greater than the current column number, and uses that element as the column to indent to. It does nothing if no such element is found.

User Option tab-stop-list

This variable is the list of tab stop columns used by tab-to-tab-stops. The elements should be integers in increasing order. The tab stop columns need not be evenly spaced. Use M-x edit-tab-stops to edit the location of tab stops interactively.

36.16.6 Indentation-Based Motion Commands

These commands, primarily for interactive use, act based on the indentation in the text.

Command back-to-indentation

This command moves point to the first non-whitespace character in the current line (which is the line in which point is located). It returns nil.

Command backward-to-indentation arg

This command moves point backward arg lines and then to the first nonblank character on that line. It returns nil.

Command forward-to-indentation arg

This command moves point forward arg lines and then to the first nonblank character on that line. It returns nil.

36.17 Case Changes

The case change commands described here work on text in the current buffer. See Section 4.11 [Character Case], page 70, for case conversion commands that work on strings and characters. See Section 4.12 [Case Tables], page 71, for how to customize which characters are upper or lower case and how to convert them.

Command capitalize-region start end

This function capitalizes all words in the region defined by start and end. To capitalize means to convert each word’s first character to upper case and convert the rest of each word to lower case. The function returns nil.

If one end of the region is in the middle of a word, the part of the word within the region is treated as an entire word.

When capitalize-region is called interactively, start and end are point and the mark, with the smallest first.

     ---------- Buffer: foo ----------
     This is the contents of the 5th foo.
     ---------- Buffer: foo ----------
     (capitalize-region 1 44)
     ⇒ nil
     ---------- Buffer: foo ----------
     This Is The Contents Of The 5th Foo.
     ---------- Buffer: foo ----------

Command downcase-region start end

This function converts all of the letters in the region defined by start and end to lower case. The function returns nil.

When downcase-region is called interactively, start and end are point and the mark, with the smallest first.

Command upcase-region start end

This function converts all of the letters in the region defined by start and end to upper case. The function returns nil.

When upcase-region is called interactively, start and end are point and the mark, with the smallest first.

Command capitalize-word count

This function capitalizes count words after point, moving point over as it does. To capitalize means to convert each word’s first character to upper case and convert the rest of each word to lower case. If count is negative, the function capitalizes the −count previous words but does not move point. The value is

nil.

If point is in the middle of a word, the part of the word before point is ignored when moving forward. The rest is treated as an entire word.

When capitalize-word is called interactively, count is set to the numeric prefix argument.

Command downcase-word count

This function converts the count words after point to all lower case, moving point over as it does. If count is negative, it converts the −count previous words but does not move point. The value is nil.

When downcase-word is called interactively, count is set to the numeric prefix argument.

Command upcase-word count

This function converts the count words after point to all upper case, moving point over as it does. If count is negative, it converts the −count previous words but does not move point. The value is nil.

When upcase-word is called interactively, count is set to the numeric prefix argument.

36.18 Text Properties

Text properties are an alternative interface to extents (see Chapter 40 [Extents], page 585), and are built on top of them. They are useful when you want to view textual properties as being attached to the characters themselves rather than to intervals of characters. The text property interface is compatible with FSF Emacs.

Each character position in a buffer or a string can have a text property list, much like the property list of a symbol (see Section 5.9 [Property Lists], page 96). The properties belong to a particular character at a particular place, such as, the letter ‘T’ at the beginning of this sentence or the first ‘o’ in ‘foo’—if the same character occurs in two different places, the two occurrences generally have different properties.

Each property has a name and a value. Both of these can be any Lisp object, but the name is normally a symbol. The usual way to access the property list is to specify a name and ask what value corresponds to it.

Note that FSF Emacs also looks at the category property to find defaults for text properties. We consider this too bogus to implement.

Copying text between strings and buffers preserves the properties along with the characters; this includes such diverse functions as substring, insert, and buffer-substring.

36.18.1 Examining Text Properties

The simplest way to examine text properties is to ask for the value of a particular property of a particular character. For that, use get-text-property. Use text-properties-at to get the entire property list of a character. See Section 36.18.3 [Property Search], page 542, for functions to examine the properties of a number of characters at once.

These functions handle both strings and buffers. (Keep in mind that positions in a string start from 0, whereas positions in a buffer start from 1.)

Function get-text-property pos prop &optional object This function returns the value of the prop property of the character after position pos in object (a buffer or string). The argument object is optional and defaults to the current buffer.

Function get-char-property pos prop &optional object This function is like get-text-property, except that it checks all extents, not just text-property extents.

Function text-properties-at position &optional object This function returns the entire property list of the character at position in the string or buffer object. If object is nil, it defaults to the current buffer.

Variable default-text-properties

This variable holds a property list giving default values for text properties. Whenever a character does not specify a value for a property, the value stored in this list is used instead. Here is an example:

          (setq default-text-properties '(foo 69))

;; Make sure character 1 has no properties of its own.

          (set-text-properties 1 2 nil)

;; What we get, when we ask, is the default value.

          (get-text-property 1 'foo)
               ⇒ 69

36.18.2 Changing Text Properties

The primitives for changing properties apply to a specified range of text. The function set-text-properties (see end of section) sets the entire property list of the text in that range; more often, it is useful to add, change, or delete just certain properties specified by name.

Since text properties are considered part of the buffer’s contents, and can affect how the buffer looks on the screen, any change in the text properties is considered a buffer modification. Buffer text property changes are undoable (see Section 36.9 [Undo], page 523).

Function put-text-property start end prop value &optional object This function sets the prop property to value for the text between start and end in the string or buffer object. If object is nil, it defaults to the current buffer.

Function add-text-properties start end props &optional object This function modifies the text properties for the text between start and end in the string or buffer object. If object is nil, it defaults to the current buffer.

The argument props specifies which properties to change. It should have the form of a property list (see Section 5.9 [Property Lists], page 96): a list whose elements include the property names followed alternately by the corresponding values.

The return value is t if the function actually changed some property’s value; nil otherwise (if props is nil or its values agree with those in the text).

For example, here is how to set the comment and face properties of a range of text:

     (add-text-properties start end
                          '(comment t face highlight))

Function remove-text-properties start end props &optional object This function deletes specified text properties from the text between start and end in the string or buffer object. If object is nil, it defaults to the current buffer.

The argument props specifies which properties to delete. It should have the form of a property list (see Section 5.9 [Property Lists], page 96): a list whose elements are property names alternating with corresponding values. But only the names matter—the values that accompany them are ignored. For example, here’s how to remove the face property.

     (remove-text-properties start end '(face nil))

The return value is t if the function actually changed some property’s value; nil otherwise (if props is nil or if no character in the specified text had any of those properties).

Function set-text-properties start end props &optional object This function completely replaces the text property list for the text between start and end in the string or buffer object. If object is nil, it defaults to the current buffer.

The argument props is the new property list. It should be a list whose elements are property names alternating with corresponding values.

After set-text-properties returns, all the characters in the specified range have identical properties.

If props is nil, the effect is to get rid of all properties from the specified range of text. Here’s an example:

     (set-text-properties start end nil)

See also the function buffer-substring-without-properties (see Section 36.2 [Buffer Contents], page 512) which copies text from the buffer but does not copy its properties.

36.18.3 Property Search Functions

In typical use of text properties, most of the time several or many consecutive characters have the same value for a property. Rather than writing your programs to examine characters one by one, it is much faster to process chunks of text that have the same property value.

Here are functions you can use to do this. They use eq for comparing property values. In all cases, object defaults to the current buffer.

For high performance, it’s very important to use the limit argument to these functions, especially the ones that search for a single property—otherwise, they may spend a long time scanning to the end of the buffer, if the property you are interested in does not change.

Remember that a position is always between two characters; the position returned by these functions is between two characters with different properties.

Function next-property-change pos &optional object limit The function scans the text forward from position pos in the string or buffer object till it finds a change in some text property, then returns the position of the change. In other words, it returns the position of the first character beyond pos whose properties are not identical to those of the character just after pos.

If limit is non-nil, then the scan ends at position limit. If there is no property change before that point, next-property-change returns limit.

The value is nil if the properties remain unchanged all the way to the end of object and limit is nil. If the value is non-nil, it is a position greater than or equal to pos. The value equals pos only when limit equals pos.

Here is an example of how to scan the buffer by chunks of text within which all properties are constant:

     (while (not (eobp))
       (let ((plist (text-properties-at (point)))
             (next-change
              (or (next-property-change (point) (current-buffer))
                  (point-max))))

Process text from point to next-change…

         (goto-char next-change)))

Function next-single-property-change pos prop &optional object limit

The function scans the text forward from position pos in the string or buffer object till it finds a change in the prop property, then returns the position of the change. In other words, it returns the position of the first character beyond pos whose prop property differs from that of the character just after pos.

If limit is non-nil, then the scan ends at position limit. If there is no property change before that point, next-single-property-change returns limit.

The value is nil if the property remains unchanged all the way to the end of object and limit is nil. If the value is non-nil, it is a position greater than or equal to pos; it equals pos only if limit equals pos.

Function previous-property-change pos &optional object limit

This is like next-property-change, but scans back from pos instead of forward. If the value is non-nil, it is a position less than or equal to pos; it equals pos only if limit equals pos.

Function previous-single-property-change pos prop &optional object

limit This is like next-single-property-change, but scans back from pos instead of forward. If the value is non-nil, it is a position less than or equal to pos; it equals pos only if limit equals pos.

Function text-property-any start end prop value &optional object This function returns non-nil if at least one character between start and end has a property prop whose value is value. More precisely, it returns the position of the first such character. Otherwise, it returns nil. The optional fifth argument, object, specifies the string or buffer to scan. Positions are relative to object. The default for object is the current buffer.

Function text-property-not-all start end prop value &optional object This function returns non-nil if at least one character between start and end has a property prop whose value differs from value. More precisely, it returns the position of the first such character. Otherwise, it returns nil. The optional fifth argument, object, specifies the string or buffer to scan. Positions are relative to object. The default for object is the current buffer.

36.18.4 Properties with Special Meanings

The predefined properties are the same as those for extents. See Section 40.6 [Extent Properties], page 591.

36.18.5 Saving Text Properties in Files

You can save text properties in files, and restore text properties when inserting the files, using these two hooks:

Variable write-region-annotate-functions

This variable’s value is a list of functions for write-region to run to encode text properties in some fashion as annotations to the text being written in the file. See Section 28.4 [Writing to Files], page 397. Each function in the list is called with two arguments: the start and end of the region to be written. These functions should not alter the contents of the buffer. Instead, they should return lists indicating annotations to write in the file in addition to the text in the buffer. Each function should return a list of elements of the form (position . string), where position is an integer specifying the relative position in the text to be written, and string is the annotation to add there.

Each list returned by one of these functions must be already sorted in increasing order by position. If there is more than one function, write-region merges the lists destructively into one sorted list.

When write-region actually writes the text from the buffer to the file, it intermixes the specified annotations at the corresponding positions. All this takes place without modifying the buffer.

Variable after-insert-file-functions

This variable holds a list of functions for insert-file-contents to call after inserting a file’s contents. These functions should scan the inserted text for annotations, and convert them to the text properties they stand for.

Each function receives one argument, the length of the inserted text; point indicates the start of that text. The function should scan that text for annotations, delete them, and create the text properties that the annotations specify. The function should return the updated length of the inserted text, as it stands after those changes. The value returned by one function becomes the argument to the next function.

These functions should always return with point at the beginning of the inserted text.

The intended use of after-insert-file-functions is for converting some sort of textual annotations into actual text properties. But other uses may be possible.

We invite users to write Lisp programs to store and retrieve text properties in files, using these hooks, and thus to experiment with various data formats and find good ones. Eventually we hope users will produce good, general extensions we can install in Emacs.

We suggest not trying to handle arbitrary Lisp objects as property names or property values—because a program that general is probably difficult to write, and slow. Instead, choose a set of possible data types that are reasonably flexible, and not too hard to encode.

See Section 28.13 [Format Conversion], page 417, for a related feature.

36.19 Substituting for a Character Code

The following functions replace characters within a specified region based on their character codes.

Function subst-char-in-region start end old-char new-char &optional

noundo This function replaces all occurrences of the character old-char with the character new-char in the region of the current buffer defined by start and end.

If noundo is non-nil, then subst-char-in-region does not record the change for undo and does not mark the buffer as modified. This feature is used for controlling selective display (see Section 45.6 [Selective Display], page 654).

subst-char-in-region does not move point and returns nil.

     ---------- Buffer: foo ----------
     This is the contents of the buffer before.
     ---------- Buffer: foo ----------
     (subst-char-in-region 1 20 ?i ?X)
          ⇒ nil
     ---------- Buffer: foo ----------
     ThXs Xs the contents of the buffer before.
     ---------- Buffer: foo ----------

Function translate-region start end table

This function applies a translation table to the characters in the buffer between positions start and end. The translation table table is a string; (aref table ochar) gives the translated character corresponding to ochar. If the length of table is less than 256, any characters with codes larger than the length of table are not altered by the translation. The return value of translate-region is the number of characters that were actually changed by the translation. This does not count characters that were mapped into themselves in the translation table.

36.20 Registers

A register is a sort of variable used in XEmacs editing that can hold a marker, a string, a rectangle, a window configuration (of one frame), or a frame configuration (of all frames). Each register is named by a single character. All characters, including control and meta characters (but with the exception of C-g), can be used to name registers. Thus, there are 255 possible registers. A register is designated in Emacs Lisp by a character that is its name.

The functions in this section return unpredictable values unless otherwise stated.

Variable register-alist

This variable is an alist of elements of the form (name . contents). Normally, there is one element for each XEmacs register that has been used. The object name is a character (an integer) identifying the register. The object contents is a string, marker, or list representing the register contents. A string represents text stored in the register. A marker represents a position. A list represents a rectangle; its elements are strings, one per line of the rectangle.

Function get-register reg

This function returns the contents of the register reg, or nil if it has no contents.

Function set-register reg value

This function sets the contents of register reg to value. A register can be set to any value, but the other register functions expect only certain data types. The return value is value.

Command view-register reg

This command displays what is contained in register reg.

Command insert-register reg &optional beforep This command inserts contents of register reg into the current buffer. Normally, this command puts point before the inserted text, and the mark after it. However, if the optional second argument beforep is non-nil, it puts the mark before and point after. You can pass a non-nil second argument beforep to this function interactively by supplying any prefix argument. If the register contains a rectangle, then the rectangle is inserted with its upper left corner at point. This means that text is inserted in the current line and underneath it on successive lines. If the register contains something other than saved text (a string) or a rectangle (a list), currently useless things happen. This may be changed in the future.

36.21 Transposition of Text

This subroutine is used by the transposition commands.

Function transpose-regions start1 end1 start2 end2 &optional

leave-markers This function exchanges two nonoverlapping portions of the buffer. Arguments start1 and end1 specify the bounds of one portion and arguments start2 and end2 specify the bounds of the other portion. Normally, transpose-regions relocates markers with the transposed text; a marker previously positioned within one of the two transposed portions moves along with that portion, thus remaining between the same two characters in their new position. However, if leave-markers is non-nil, transpose-regions does not do this—it leaves all markers unrelocated.

36.22 Change Hooks

These hook variables let you arrange to take notice of all changes in all buffers (or in a particular buffer, if you make them buffer-local).

The functions you use in these hooks should save and restore the match data if they do anything that uses regular expressions; otherwise, they will interfere in bizarre ways with the editing operations that call them.

Buffer changes made while executing the following hooks don’t themselves cause any change hooks to be invoked.

Variable before-change-functions

This variable holds a list of a functions to call before any buffer modification. Each function gets two arguments, the beginning and end of the region that is about to change, represented as integers. The buffer that is about to change is always the current buffer.

Variable after-change-functions

This variable holds a list of a functions to call after any buffer modification. Each function receives three arguments: the beginning and end of the region just changed, and the length of the text that existed before the change. (To get the current length, subtract the region beginning from the region end.) All three arguments are integers. The buffer that’s about to change is always the current buffer.

Variable before-change-function

This obsolete variable holds one function to call before any buffer modification (or nil for no function). It is called just like the functions in before-change-

functions.

Variable after-change-function

This obsolete variable holds one function to call after any buffer modification (or nil for no function). It is called just like the functions in after-change-

functions.

Variable first-change-hook

This variable is a normal hook that is run whenever a buffer is changed that was previously in the unmodified state.

548 XEmacs Lisp Reference Manual

37 Searching and Matching

XEmacs provides two ways to search through a buffer for specified text: exact string searches and regular expression searches. After a regular expression search, you can examine the match data to determine which text matched the whole regular expression or various portions of it.

The ‘skip-chars...’ functions also perform a kind of searching. See Section 34.2.7 [Skipping Characters], page 494.

37.1 Searching for Strings

These are the primitive functions for searching through the text in a buffer. They are meant for use in programs, but you may call them interactively. If you do so, they prompt for the search string; limit and noerror are set to nil, and repeat is set to 1.

Command search-forward string &optional limit noerror repeat This function searches forward from point for an exact match for string. If successful, it sets point to the end of the occurrence found, and returns the new value of point. If no match is found, the value and side effects depend on noerror (see below). In the following example, point is initially at the beginning of the line. Then (search-forward "fox") moves point after the last letter of ‘fox’:

     ---------- Buffer: foo ----------
     ⋆The quick brown fox jumped over the lazy dog.
     ---------- Buffer: foo ----------
     (search-forward "fox")
          ⇒ 20
     ---------- Buffer: foo ----------
     The quick brown fox⋆ jumped over the lazy dog.
     ---------- Buffer: foo ----------

The argument limit specifies the upper bound to the search. (It must be a position in the current buffer.) No match extending after that position is accepted. If limit is omitted or nil, it defaults to the end of the accessible portion of the buffer. What happens when the search fails depends on the value of noerror. If noerror is nil, a search-failed error is signaled. If noerror is t, search-forward returns nil and does nothing. If noerror is neither nil nor t, then searchforward moves point to the upper bound and returns nil. (It would be more consistent now to return the new position of point in that case, but some programs may depend on a value of nil.) If repeat is supplied (it must be a positive number), then the search is repeated that many times (each time starting at the end of the previous time’s match). If these successive searches succeed, the function succeeds, moving point and returning its new value. Otherwise the search fails.

Command search-backward string &optional limit noerror repeat This function searches backward from point for string. It is just like searchforward except that it searches backwards and leaves point at the beginning of the match.

Command word-search-forward string &optional limit noerror repeat This function searches forward from point for a “word” match for string. If it finds a match, it sets point to the end of the match found, and returns the new value of point. Word matching regards string as a sequence of words, disregarding punctuation that separates them. It searches the buffer for the same sequence of words. Each word must be distinct in the buffer (searching for the word ‘ball’ does not match the word ‘balls’), but the details of punctuation and spacing are ignored (searching for ‘ball boy’ does match ‘ball. Boy!’). In this example, point is initially at the beginning of the buffer; the search leaves it between the ‘y’ and the ‘!’.

     ---------- Buffer: foo ----------
     ⋆He said "Please!  Find
     the ball boy!"
     ---------- Buffer: foo ----------
     (word-search-forward "Please find the ball, boy.")
          ⇒ 35
     ---------- Buffer: foo ----------
     He said "Please!  Find
     the ball boy⋆!"
     ---------- Buffer: foo ----------

If limit is non-nil (it must be a position in the current buffer), then it is the upper bound to the search. The match found must not extend after that position. If noerror is nil, then word-search-forward signals an error if the search fails. If noerror is t, then it returns nil instead of signaling an error. If noerror is neither nil nor t, it moves point to limit (or the end of the buffer) and returns

nil.

If repeat is non-nil, then the search is repeated that many times. Point is positioned at the end of the last match.

Command word-search-backward string &optional limit noerror repeat This function searches backward from point for a word match to string. This function is just like word-search-forward except that it searches backward and normally leaves point at the beginning of the match.

37.2 Regular Expressions

A regular expression (regexp, for short) is a pattern that denotes a (possibly infinite) set of strings. Searching for matches for a regexp is a very powerful operation. This section explains how to write regexps; the following section says how to search using them.

To gain a thorough understanding of regular expressions and how to use them to best advantage, we recommend that you study Mastering Regular Expressions, by Jeffrey E.F. Friedl, O’Reilly and Associates, 1997. (It’s known as the "Hip Owls" book, because of the picture on its cover.) You might also read the manuals to ⟨undefined⟩ [(gawk)Top], page ⟨undefined⟩, ⟨undefined⟩ [(ed)Top], page ⟨undefined⟩, sed, grep, ⟨undefined⟩ [(perl)Top], page ⟨undefined⟩, ⟨undefined⟩ [(regex)Top], page ⟨undefined⟩, ⟨undefined⟩ [(rx)Top], page ⟨undefined⟩, pcre, and ⟨undefined⟩ [(flex)Top], page ⟨undefined⟩. All of these programs and libraries make effective use of regular expressions.

The XEmacs regular expression syntax most closely resembles that of ed, or grep, the GNU versions of which all utilize the GNU regex library. XEmacs’ version of regex has recently been extended with some Perl–like capabilities, which are described in the next section.

37.2.1 Syntax of Regular Expressions

Regular expressions have a syntax in which a few characters are special constructs and the rest are ordinary. An ordinary character is a simple regular expression that matches that character and nothing else. The special characters are ‘.’, ‘*’, ‘+’, ‘?’, ‘[’, ‘]’, ‘^’, ‘$’, and ‘\’; no new special characters will be defined in the future. Any other character appearing in a regular expression is ordinary, unless a ‘\’ precedes it.

For example, ‘f’ is not a special character, so it is ordinary, and therefore ‘f’ is a regular expression that matches the string ‘f’ and no other string. (It does not match the string ‘ff’.) Likewise, ‘o’ is a regular expression that matches only ‘o’.

Any two regular expressions a and b can be concatenated. The result is a regular expression that matches a string if a matches some amount of the beginning of that string and b matches the rest of the string.

As a simple example, we can concatenate the regular expressions ‘f’ and ‘o’ to get the regular expression ‘fo’, which matches only the string ‘fo’. Still trivial. To do something more powerful, you need to use one of the special characters. Here is a list of them:

. (Period) is a special character that matches any single character except a newline. Using concatenation, we can make regular expressions like ‘a.b’, which matches any three-character string that begins with ‘a’ and ends with ‘b’.

* is not a construct by itself; it is a quantifying suffix operator that means to repeat the preceding regular expression as many times as possible. In ‘fo*’, the ‘*’ applies to the ‘o’, so ‘fo*’ matches one ‘f’ followed by any number of ‘o’s. The case of zero ‘o’s is allowed: ‘fo*’ does match ‘f’. ‘*’ always applies to the smallest possible preceding expression. Thus, ‘fo*’ has a repeating ‘o’, not a repeating ‘fo’. The matcher processes a ‘*’ construct by matching, immediately, as many repetitions as can be found; it is "greedy". Then it continues with the rest of the pattern. If that fails, backtracking occurs, discarding some of the matches of the ‘*’-modified construct in case that makes it possible to match the rest of the pattern. For example, in matching ‘ca*ar’ against the string ‘caaar’, the ‘a*’ first tries to match all three ‘a’s; but the rest of the pattern is ‘ar’ and there is only ‘r’ left to match, so this try fails. The next alternative is for ‘a*’ to match only two ‘a’s. With this choice, the rest of the regexp matches successfully. Nested repetition operators can be extremely slow if they specify backtracking loops. For example, it could take hours for the regular expression ‘\(x+y*\)*a’

to match the sequence ‘xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxz’. The slow-

ness is because Emacs must try each imaginable way of grouping the 35 ‘x’’s before concluding that none of them can work. To make sure your regular expressions run fast, check nested repetitions carefully.

+ is a quantifying suffix operator similar to ‘*’ except that the preceding expression must match at least once. It is also "greedy". So, for example, ‘ca+r’ matches the strings ‘car’ and ‘caaaar’ but not the string ‘cr’, whereas ‘ca*r’ matches all three strings.

? is a quantifying suffix operator similar to ‘*’, except that the preceding expression can match either once or not at all. For example, ‘ca?r’ matches ‘car’ or ‘cr’, but does not match anything else.

*? works just like ‘*’, except that rather than matching the longest match, it matches the shortest match. ‘*?’ is known as a non-greedy quantifier, a regexp construct borrowed from Perl. This construct is very useful for when you want to match the text inside a pair of delimiters. For instance, ‘/\*.*?\*/’ will match C comments in a string. This could not be so elegantly achieved without the use of a non-greedy quantifier. This construct has not been available prior to XEmacs 20.4. It is not available in FSF Emacs.

+? is the ‘+’ analog to ‘*?’.

\{n,m\} serves as an interval quantifier, analogous to ‘*’ or ‘+’, but specifies that the expression must match at least n times, but no more than m times. This syntax is supported by most Unix regexp utilities, and has been introduced to XEmacs for the version 20.3.

[ ... ] ‘[’ begins a character set, which is terminated by a ‘]’. In the simplest case, the characters between the two brackets form the set. Thus, ‘[ad]’ matches either one ‘a’ or one ‘d’, and ‘[ad]*’ matches any string composed of just ‘a’s and ‘d’s (including the empty string), from which it follows that ‘c[ad]*r’ matches ‘cr’, ‘car’, ‘cdr’, ‘caddaar’, etc. The usual regular expression special characters are not special inside a character set. A completely different set of special characters exists inside character sets: ‘]’, ‘-’ and ‘^’. ‘-’ is used for ranges of characters. To write a range, write two characters with a ‘-’ between them. Thus, ‘[a-z]’ matches any lower case letter. Ranges may be intermixed freely with individual characters, as in ‘[a-z$%.]’, which matches any lower case letter or ‘$’, ‘%’, or a period. To include a ‘]’ in a character set, make it the first character. For example, ‘[]a]’ matches ‘]’ or ‘a’. To include a ‘-’, write ‘-’ as the first character in the set, or put it immediately after a range. (You can replace one individual character c with the range ‘c-c’ to make a place to put the ‘-’.) There is no way to write a set containing just ‘-’ and ‘]’. To include ‘^’ in a set, put it anywhere but at the beginning of the set.

[^... ] ‘[^’ begins a complement character set, which matches any character except the ones specified. Thus, ‘[^a-z0-9A-Z]’ matches all characters except letters and digits. ‘^’ is not special in a character set unless it is the first character. The character following the ‘^’ is treated as if it were first (thus, ‘-’ and ‘]’ are not special there). Note that a complement character set can match a newline, unless newline is mentioned as one of the characters not to match.

^is a special character that matches the empty string, but only at the beginning of a line in the text being matched. Otherwise it fails to match anything. Thus, ‘^foo’ matches a ‘foo’ that occurs at the beginning of a line. When matching a string instead of a buffer, ‘^’ matches at the beginning of the string or after a newline character ‘\n’.

$ is similar to ‘^’ but matches only at the end of a line. Thus, ‘x+$’ matches a string of one ‘x’ or more at the end of a line. When matching a string instead of a buffer, ‘$’ matches at the end of the string or before a newline character ‘\n’.

\ has two functions: it quotes the special characters (including ‘\’), and it introduces additional special constructs. Because ‘\’ quotes special characters, ‘\$’ is a regular expression that matches only ‘$’, and ‘\[’ is a regular expression that matches only ‘[’, and so on. Note that ‘\’ also has special meaning in the read syntax of Lisp strings (see Section 2.4.8 [String Type], page 25), and must be quoted with ‘\’. For example, the regular expression that matches the ‘\’ character is ‘\\’. To write a Lisp string that contains the characters ‘\\’, Lisp syntax requires you to quote each ‘\’ with another ‘\’. Therefore, the read syntax for a regular expression matching ‘\’ is "\\\\".

Please note: For historical compatibility, special characters are treated as ordinary ones if they are in contexts where their special meanings make no sense. For example, ‘*foo’ treats ‘*’ as ordinary since there is no preceding expression on which the ‘*’ can act. It is poor practice to depend on this behavior; quote the special character anyway, regardless of where it appears.

For the most part, ‘\’ followed by any character matches only that character. However, there are several exceptions: characters that, when preceded by ‘\’, are special constructs. Such characters are always ordinary when encountered on their own. Here is a table of ‘\’ constructs:

\| specifies an alternative. Two regular expressions a and b with ‘\|’ in between form an expression that matches anything that either a or b matches.

Thus, ‘foo\|bar’ matches either ‘foo’ or ‘bar’ but no other string. ‘\|’ applies to the largest possible surrounding expressions. Only a surrounding ‘\( ... \)’ grouping can limit the grouping power of ‘\|’. Full backtracking capability exists to handle multiple uses of ‘\|’.

\( ... \) is a grouping construct that serves three purposes:

1. To enclose a set of ‘\|’ alternatives for other operations. Thus, ‘\(foo\|bar\)x’ matches either ‘foox’ or ‘barx’.

2. To enclose an expression for a suffix operator such as ‘*’ to act on. Thus, ‘ba\(na\)*’ matches ‘bananana’, etc., with any (zero or more) number of ‘na’ strings.

3. To record a matched substring for future reference.

This last application is not a consequence of the idea of a parenthetical grouping; it is a separate feature that happens to be assigned as a second meaning to the same ‘\( ... \)’ construct because there is no conflict in practice between the two meanings. Here is an explanation of this feature:

\digit matches the same text that matched the digitth occurrence of a ‘\( ... \)’ construct. In other words, after the end of a ‘\( ... \)’ construct. the matcher remembers the beginning and end of the text matched by that construct. Then, later on in the regular expression, you can use ‘\’ followed by digit to match that same text, whatever it may have been. The strings matching the first nine ‘\( ... \)’ constructs appearing in a regular expression are assigned numbers 1 through 9 in the order that the open parentheses appear in the regular expression. So you can use ‘\1’ through ‘\9’ to refer to the text matched by the corresponding ‘\( ... \)’ constructs. For example, ‘\(.*\)\1’ matches any newline-free string that is composed of two identical halves. The ‘\(.*\)’ matches the first half, which may be anything, but the ‘\1’ that follows must match the same exact text.

\(?: ... \)

is called a shy grouping operator, and it is used just like ‘\( ... \)’, except that it does not cause the match substring to be recorded for future reference. This is useful when you need to use a lot of nested grouping ‘\( ... \)’ constructs to express complex alternation, but only want to memoize, or capture, one or two of the subexpression matches. Since ‘\(?: ... \)’ doesn’t capture a sub-match, it also doesn’t need to be counted when you count ‘\( ... \)’ groups to figure the ‘match-string’ index. That turns out to be a very convenient characteristic. This situation occurs where parts of a regular expression have been automaticly generated by a program that builds them from lists of strings, and the static code following the matching operation must access a specific match number. Here’s an example that shows this. We will assume that (require 'regexp-opt) has been executed already, to ensure that ‘regexp-opt.el’, which is part of the xemacs-devel package, is loaded. In a real program, lets pretend that varnames would be a list of strings holding the names of some variables extracted somehow from the text of a program source you are editing and running this function on. For the purposes of this illustration, we can just bind it in the let* expression.

     (let* ((varnames '("k" "n" "i" "j" "varname"))
            (keys-regexp (regexp-opt
                          (mapcar #'symbol-name
                                  '(if then else elif
                                    case in of do while
                                    with for next unless
                                    cond begin end))))
           (varname-regexp (regexp-opt varnames))
           (contrived-regexp (concat "\\(" keys-regexp "\\)"
                                     "\\s-(\\s-\\("
                                     varname-regexp
                                     "\\)\\s-)"))
           (keyname "")
           (varname ""))

;; In the body of this particular defun, we:

       (re-search-forward contrived-regexp nil t)

;; … and it finds a match. Now we want to extract the ;; text that it matched on, and save it into keyname

       ;; and varname.
       (setq keyname (match-string 1)
             varname (match-string 2))

;; … and then do something with those values.

       (list keyname varname))

;; Here’s something for it to match, so you can try it with

     ;; C-x C-e
     ;; while ( j ) do ...

Here you should see that if the regular expression returned by regexp-opt did not use ‘\(?: ... \)’ for grouping, and instead used ‘\( ... \)’, it would be necessary to count the number of opening parentheses in the keys-regexp and to use that figure to calculate which match number is matched by the varnameregexp. It is much more convenient to be able to just ask for the second match string. The shy grouping operator has been borrowed from Perl, and has not been available prior to XEmacs 20.3, nor is it available in FSF Emacs.

\w matches any word-constituent character. The editor syntax table determines which characters these are. See Chapter 38 [Syntax Tables], page 569.

\W matches any character that is not a word constituent.

\scode matches any character whose syntax is code. Here code is a character that represents a syntax code: thus, ‘w’ for word constituent, ‘-’ for whitespace, ‘(’ for open parenthesis, etc. See Chapter 38 [Syntax Tables], page 569, for a list of syntax codes and the characters that stand for them.

\Scode matches any character whose syntax is not code.

The following regular expression constructs match the empty string—that is, they don’t use up any characters—but whether they match depends on the context.

\` matches the empty string, but only at the beginning of the buffer or string being matched against.

\' matches the empty string, but only at the end of the buffer or string being matched against.

\= matches the empty string, but only at point. (This construct is not defined when matching against a string.)

\b matches the empty string, but only at the beginning or end of a word. Thus, ‘\bfoo\b’ matches any occurrence of ‘foo’ as a separate word. ‘\bballs?\b’ matches ‘ball’ or ‘balls’ as a separate word.

\B matches the empty string, but not at the beginning or end of a word.

\< matches the empty string, but only at the beginning of a word.

\> matches the empty string, but only at the end of a word.

Not every string is a valid regular expression. For example, a string with unbalanced square brackets is invalid (with a few exceptions, such as ‘[]]’), and so is a string that ends with a single ‘\’. If an invalid regular expression is passed to any of the search functions, an invalid-regexp error is signaled.

Function regexp-quote string

This function returns a regular expression string that matches exactly string and nothing else. This allows you to request an exact string match when calling a function that wants a regular expression.

          (regexp-quote "^The cat$")
               ⇒ "\\^The cat\\$"

One use of regexp-quote is to combine an exact string match with context described as a regular expression. For example, this searches for the string that is the value of string, surrounded by whitespace:

          (re-search-forward
           (concat "\\s-" (regexp-quote string) "\\s-"))

37.2.2 Complex Regexp Example

Here is a complicated regexp, used by XEmacs to recognize the end of a sentence together with any whitespace that follows. It is the value of the variable sentence-end.

First, we show the regexp as a string in Lisp syntax to distinguish spaces from tab characters. The string constant begins and ends with a double-quote. ‘\"’ stands for a double-quote as part of the string, ‘\\’ for a backslash as part of the string, ‘\t’ for a tab and ‘\n’ for a newline.

     "[.?!][]\"')}]*\\($\\| $\\|\t\\|  \\)[ \t\n]*"

In contrast, if you evaluate the variable sentence-end, you will see the following:

sentence-end

⇒

"[.?!][]\"')}]*\\($\\| $\\|  \\|  \\)[
]*"

In this output, tab and newline appear as themselves.

This regular expression contains four parts in succession and can be deciphered as follows:

[.?!] The first part of the pattern is a character set that matches any one of three characters: period, question mark, and exclamation mark. The match must begin with one of these three characters.

[]\"')}]*

The second part of the pattern matches any closing braces and quotation marks, zero or more of them, that may follow the period, question mark or exclamation mark. The \" is Lisp syntax for a double-quote in a string. The ‘*’ at the end indicates that the immediately preceding regular expression (a character set, in this case) may be repeated zero or more times.

\\($\\| $\\|\t\\|  \\)

The third part of the pattern matches the whitespace that follows the end of a sentence: the end of a line, or a tab, or two spaces. The double backslashes mark the parentheses and vertical bars as regular expression syntax; the parentheses delimit a group and the vertical bars separate alternatives. The dollar sign is used to match the end of a line.

[ \t\n]* Finally, the last part of the pattern matches any additional whitespace beyond the minimum needed to end a sentence.

37.3 Regular Expression Searching

In XEmacs, you can search for the next match for a regexp either incrementally or not. Incremental search commands are described in the The XEmacs Reference Manual. See section “Regular Expression Search” in The XEmacs Reference Manual. Here we describe only the search functions useful in programs. The principal one is re-search-forward.

Command re-search-forward regexp &optional limit noerror repeat This function searches forward in the current buffer for a string of text that is matched by the regular expression regexp. The function skips over any amount of text that is not matched by regexp, and leaves point at the end of the first match found. It returns the new value of point. If limit is non-nil (it must be a position in the current buffer), then it is the upper bound to the search. No match extending after that position is accepted. What happens when the search fails depends on the value of noerror. If noerror

is nil, a search-failed error is signaled. If noerror is t, re-search-forward

does nothing and returns nil. If noerror is neither nil nor t, then re-search-forward moves point to limit (or the end of the buffer) and returns nil. If repeat is supplied (it must be a positive number), then the search is repeated that many times (each time starting at the end of the previous time’s match).

If these successive searches succeed, the function succeeds, moving point and returning its new value. Otherwise the search fails. In the following example, point is initially before the ‘T’. Evaluating the search call moves point to the end of that line (between the ‘t’ of ‘hat’ and the newline).

     ---------- Buffer: foo ----------
     I read "⋆The cat in the hat
     comes back" twice.
     ---------- Buffer: foo ----------
     (re-search-forward "[a-z]+" nil t 5)
          ⇒ 27
     ---------- Buffer: foo ----------
     I read "The cat in the hat⋆
     comes back" twice.
     ---------- Buffer: foo ----------

Command re-search-backward regexp &optional limit noerror repeat This function searches backward in the current buffer for a string of text that is matched by the regular expression regexp, leaving point at the beginning of the first text found. This function is analogous to re-search-forward, but they are not simple mirror images. re-search-forward finds the match whose beginning is as close as possible to the starting point. If re-search-backward were a perfect mirror image, it would find the match whose end is as close as possible. However, in fact it finds the match whose beginning is as close as possible. The reason is that matching a regular expression at a given spot always works from beginning to end, and starts at a specified beginning position. A true mirror-image of re-search-forward would require a special feature for matching regexps from end to beginning. It’s not worth the trouble of implementing that.

Function string-match regexp string &optional start This function returns the index of the start of the first match for the regular expression regexp in string, or nil if there is no match. If start is non-nil, the search starts at that index in string. For example,

     (string-match
      "quick" "The quick brown fox jumped quickly.")
          ⇒ 4
     (string-match
      "quick" "The quick brown fox jumped quickly." 8)
          ⇒ 27

The index of the first character of the string is 0, the index of the second character is 1, and so on. After this function returns, the index of the first character beyond the match is available as (match-end 0). See Section 37.6 [Match Data], page 561.

     (string-match
      "quick" "The quick brown fox jumped quickly." 8)
          ⇒ 27
     (match-end 0)
          ⇒ 32

Function split-string string &optional pattern This function splits string to substrings delimited by pattern, and returns a list of substrings. If pattern is omitted, it defaults to ‘[ \f\t\n\r\v]+’, which means that it splits string by white–space.

     (split-string "foo bar")
          ⇒ ("foo" "bar")
     (split-string "something")
          ⇒ ("something")
     (split-string "a:b:c" ":")
          ⇒ ("a" "b" "c")
     (split-string ":a::b:c" ":")
          ⇒ ("" "a" "" "b" "c")

Function split-path path

This function splits a search path into a list of strings. The path components are separated with the characters specified with path-separator. Under Unix, path-separator will normally be ‘:’, while under Windows, it will be ‘;’.

Function looking-at regexp

This function determines whether the text in the current buffer directly following point matches the regular expression regexp. “Directly following” means precisely that: the search is “anchored” and it can succeed only starting with the first character following point. The result is t if so, nil otherwise. This function does not move point, but it updates the match data, which you can access using match-beginning and match-end. See Section 37.6 [Match Data], page 561. In this example, point is located directly before the ‘T’. If it were anywhere else, the result would be nil.

     ---------- Buffer: foo ----------
     I read "⋆The cat in the hat
     comes back" twice.
     ---------- Buffer: foo ----------
     (looking-at "The cat in the hat$")
          ⇒ t

37.4 POSIX Regular Expression Searching

The usual regular expression functions do backtracking when necessary to handle the ‘\|’ and repetition constructs, but they continue this only until they find some match. Then they succeed and report the first match found.

This section describes alternative search functions which perform the full backtracking specified by the POSIX standard for regular expression matching. They continue backtracking until they have tried all possibilities and found all matches, so they can report the longest match, as required by POSIX. This is much slower, so use these functions only when you really need the longest match.

In Emacs versions prior to 19.29, these functions did not exist, and the functions described above implemented full POSIX backtracking.

Function posix-search-forward regexp &optional limit noerror repeat This is like re-search-forward except that it performs the full backtracking specified by the POSIX standard for regular expression matching.

Function posix-search-backward regexp &optional limit noerror repeat This is like re-search-backward except that it performs the full backtracking specified by the POSIX standard for regular expression matching.

Function posix-looking-at regexp

This is like looking-at except that it performs the full backtracking specified by the POSIX standard for regular expression matching.

Function posix-string-match regexp string &optional start This is like string-match except that it performs the full backtracking specified by the POSIX standard for regular expression matching.

37.5 Search and Replace

Function perform-replace from-string replacements query-flag regexp-flag

delimited-flag &optional repeat-count map This function is the guts of query-replace and related commands. It searches for occurrences of from-string and replaces some or all of them. If query-flag is nil, it replaces all occurrences; otherwise, it asks the user what to do about each one.

If regexp-flag is non-nil, then from-string is considered a regular expression; otherwise, it must match literally. If delimited-flag is non-nil, then only replacements surrounded by word boundaries are considered.

The argument replacements specifies what to replace occurrences with. If it is a string, that string is used. It can also be a list of strings, to be used in cyclic order.

If repeat-count is non-nil, it should be an integer. Then it specifies how many times to use each of the strings in the replacements list before advancing cyclicly to the next one.

Normally, the keymap query-replace-map defines the possible user responses for queries. The argument map, if non-nil, is a keymap to use instead of

query-replace-map.

Variable query-replace-map

This variable holds a special keymap that defines the valid user responses for

     query-replace and related functions, as well as y-or-n-p and map-y-or-n-p.

It is unusual in two ways:

• The “key bindings” are not commands, just symbols that are meaningful to the functions that use this map.

• Prefix keys are not supported; each key binding must be for a single event key sequence. This is because the functions don’t use read key sequence to get the input; instead, they read a single event and look it up “by hand.”

Here are the meaningful “bindings” for query-replace-map. Several of them are meaningful only for query-replace and friends.

act Do take the action being considered—in other words, “yes.”

skip Do not take action for this question—in other words, “no.”

exit Answer this question “no,” and give up on the entire series of questions, assuming that the answers will be “no.”

act-and-exit

Answer this question “yes,” and give up on the entire series of questions, assuming that subsequent answers will be “no.”

act-and-show

Answer this question “yes,” but show the results—don’t advance yet to the next question.

automatic

Answer this question and all subsequent questions in the series with “yes,” without further user interaction.

backup Move back to the previous place that a question was asked about.

edit Enter a recursive edit to deal with this question—instead of any other action that would normally be taken.

delete-and-edit

Delete the text being considered, then enter a recursive edit to replace it.

recenter Redisplay and center the window, then ask the same question again.

quit Perform a quit right away. Only y-or-n-p and related functions use this answer.

help Display some help, then ask again.

37.6 The Match Data

XEmacs keeps track of the positions of the start and end of segments of text found during a regular expression search. This means, for example, that you can search for a complex pattern, such as a date in an Rmail message, and then extract parts of the match under control of the pattern.

Because the match data normally describe the most recent search only, you must be careful not to do another search inadvertently between the search you wish to refer back to and the use of the match data. If you can’t avoid another intervening search, you must save and restore the match data around it, to prevent it from being overwritten.

37.6.1 Simple Match Data Access

This section explains how to use the match data to find out what was matched by the last search or match operation.

You can ask about the entire matching text, or about a particular parenthetical subexpression of a regular expression. The count argument in the functions below specifies which. If count is zero, you are asking about the entire match. If count is positive, it specifies which subexpression you want.

Recall that the subexpressions of a regular expression are those expressions grouped with escaped parentheses, ‘\(...\)’. The countth subexpression is found by counting occurrences of ‘\(’ from the beginning of the whole regular expression. The first subexpression is numbered 1, the second 2, and so on. Only regular expressions can have subexpressions— after a simple string search, the only information available is about the entire match.

Function match-string count &optional in-string This function returns, as a string, the text matched in the last search or match operation. It returns the entire text if count is zero, or just the portion corresponding to the countth parenthetical subexpression, if count is positive. If count is out of range, or if that subexpression didn’t match anything, the value

is nil.

If the last such operation was done against a string with string-match, then you should pass the same string as the argument in-string. Otherwise, after a buffer search or match, you should omit in-string or pass nil for it; but you should make sure that the current buffer when you call match-string is the one in which you did the searching or matching.

Function match-beginning count

This function returns the position of the start of text matched by the last regular expression searched for, or a subexpression of it. If count is zero, then the value is the position of the start of the entire match. Otherwise, count specifies a subexpression in the regular expression, and the value of the function is the starting position of the match for that subexpression. The value is nil for a subexpression inside a ‘\|’ alternative that wasn’t used in the match.

Function match-end count

This function is like match-beginning except that it returns the position of the end of the match, rather than the position of the beginning.

Here is an example of using the match data, with a comment showing the positions within the text:

(string-match "\\(qu\\)\\(ick\\)"
              "The quick fox jumped quickly.")
              ;0123456789
     ⇒ 4
(match-string 0 "The quick fox jumped quickly.")
     ⇒ "quick"
(match-string 1 "The quick fox jumped quickly.")
     ⇒ "qu"
(match-string 2 "The quick fox jumped quickly.")
     ⇒ "ick"

(match-beginning 1) ; The beginning of the match ⇒ 4 ; with ‘qu’ is at index 4.

(match-beginning 2) ; The beginning of the match ⇒ 6 ; with ‘ick’ is at index 6.

(match-end 1) ; The end of the match ⇒ 6 ; with ‘qu’ is at index 6.

(match-end 2) ; The end of the match

⇒ 9 ; with ‘ick’ is at index 9. Here is another example. Point is initially located at the beginning of the line. Searching

moves point to between the space and the word ‘in’. The beginning of the entire match is at the 9th character of the buffer (‘T’), and the beginning of the match for the first subexpression is at the 13th character (‘c’).

(list
  (re-search-forward "The \\(cat \\)")
  (match-beginning 0)
  (match-beginning 1))
    ⇒ (9 9 13)
---------- Buffer: foo ----------
I read "The cat ⋆in the hat comes back" twice.
        ^^
        9  13
---------- Buffer: foo ----------

(In this case, the index returned is a buffer position; the first character of the buffer counts as 1.)

37.6.2 Replacing the Text That Matched

This function replaces the text matched by the last search with replacement.

Function replace-match replacement &optional fixedcase literal string This function replaces the text in the buffer (or in string) that was matched by the last search. It replaces that text with replacement. If you did the last search in a buffer, you should specify nil for string. Then replace-match does the replacement by editing the buffer; it leaves point at the end of the replacement text, and returns t.

If you did the search in a string, pass the same string as string. Then replacematch does the replacement by constructing and returning a new string.

If fixedcase is non-nil, then the case of the replacement text is not changed; otherwise, the replacement text is converted to a different case depending upon the capitalization of the text to be replaced. If the original text is all upper case, the replacement text is converted to upper case. If the first word of the original text is capitalized, then the first word of the replacement text is capitalized. If the original text contains just one word, and that word is a capital letter, replace-match considers this a capitalized first word rather than all upper case.

If case-replace is nil, then case conversion is not done, regardless of the value of fixed-case. See Section 37.7 [Searching and Case], page 565.

If literal is non-nil, then replacement is inserted exactly as it is, the only alterations being case changes as needed. If it is nil (the default), then the character ‘\’ is treated specially. If a ‘\’ appears in replacement, then it must be part of one of the following sequences:

‘\&’ ‘\&’ stands for the entire text being replaced.

‘\n’ ‘\n’, where n is a digit, stands for the text that matched the nth subexpression in the original regexp. Subexpressions are those expressions grouped inside ‘\(...\)’.

‘\\’ ‘\\’ stands for a single ‘\’ in the replacement text.

37.6.3 Accessing the Entire Match Data

The functions match-data and set-match-data read or write the entire match data, all at once.

Function match-data

This function returns a newly constructed list containing all the information on what text the last search matched. Element zero is the position of the beginning of the match for the whole expression; element one is the position of the end of the match for the expression. The next two elements are the positions of the beginning and end of the match for the first subexpression, and so on. In general, element number 2n corresponds to (match-beginning n); and element number 2n + 1 corresponds to (match-end n).

All the elements are markers or nil if matching was done on a buffer, and all are integers or nil if matching was done on a string with string-match. (In Emacs 18 and earlier versions, markers were used even for matching on a string, except in the case of the integer 0.)

As always, there must be no possibility of intervening searches between the call to a search function and the call to match-data that is intended to access the match data for that search.

     (match-data)
          ⇒ (#<marker at 9 in foo>
               #<marker at 17 in foo>
               #<marker at 13 in foo>
               #<marker at 17 in foo>)

Function set-match-data match-list

This function sets the match data from the elements of match-list, which should be a list that was the value of a previous call to match-data. If match-list refers to a buffer that doesn’t exist, you don’t get an error; that sets the match data in a meaningless but harmless way.

store-match-data is an alias for set-match-data.

37.6.4 Saving and Restoring the Match Data

When you call a function that may do a search, you may need to save and restore the match data around that call, if you want to preserve the match data from an earlier search for later use. Here is an example that shows the problem that arises if you fail to save the match data:

(re-search-forward "The \\(cat \\)")
     ⇒ 48

(foo) ; Perhaps foo does

; more searching.

(match-end 0)

⇒ 61 ; Unexpected result—not 48! You can save and restore the match data with save-match-data:

Macro save-match-data body…

This special form executes body, saving and restoring the match data around it.

You can use set-match-data together with match-data to imitate the effect of the special form save-match-data. This is useful for writing code that can run in Emacs 18. Here is how:

(let ((data (match-data)))
  (unwind-protect

... ; May change the original match data.

    (set-match-data data)))

Emacs automatically saves and restores the match data when it runs process filter functions (see Section 49.9.2 [Filter Functions], page 684) and process sentinels (see Section 49.10 [Sentinels], page 687).

37.7 Searching and Case

By default, searches in Emacs ignore the case of the text they are searching through; if you specify searching for ‘FOO’, then ‘Foo’ or ‘foo’ is also considered a match. Regexps, and in particular character sets, are included: thus, ‘[aB]’ would match ‘a’ or ‘A’ or ‘b’ or ‘B’.

If you do not want this feature, set the variable case-fold-search to nil. Then all letters must match exactly, including case. This is a buffer-local variable; altering the variable affects only the current buffer. (See Section 10.9.1 [Intro to Buffer-Local], page 157.) Alternatively, you may change the value of default-case-fold-search, which is the default value of case-fold-search for buffers that do not override it.

Note that the user-level incremental search feature handles case distinctions differently. When given a lower case letter, it looks for a match of either case, but when given an upper case letter, it looks for an upper case letter only. But this has nothing to do with the searching functions Lisp functions use.

User Option case-replace

This variable determines whether the replacement functions should preserve case. If the variable is nil, that means to use the replacement text verbatim. A non-nil value means to convert the case of the replacement text according to the text being replaced. The function replace-match is where this variable actually has its effect. See Section 37.6.2 [Replacing Match], page 563.

User Option case-fold-search

This buffer-local variable determines whether searches should ignore case. If the variable is nil they do not ignore case; otherwise they do ignore case.

Variable default-case-fold-search

The value of this variable is the default value for case-fold-search in buffers that do not override it. This is the same as (default-value 'case-fold-

search).

37.8 Standard Regular Expressions Used in Editing

This section describes some variables that hold regular expressions used for certain purposes in editing:

Variable page-delimiter

This is the regexp describing line-beginnings that separate pages. The default value is "^\014" (i.e., "^^L" or "^\C-l"); this matches a line that starts with a formfeed character.

The following two regular expressions should not assume the match always starts at the beginning of a line; they should not use ‘^’ to anchor the match. Most often, the paragraph commands do check for a match only at the beginning of a line, which means that ‘^’ would be superfluous. When there is a nonzero left margin, they accept matches that start after the left margin. In that case, a ‘^’ would be incorrect. However, a ‘^’ is harmless in modes where a left margin is never used.

Variable paragraph-separate

This is the regular expression for recognizing the beginning of a line that separates paragraphs. (If you change this, you may have to change paragraphstart also.) The default value is "[ \t\f]*$", which matches a line that consists entirely of spaces, tabs, and form feeds (after its left margin).

Variable paragraph-start

This is the regular expression for recognizing the beginning of a line that starts or separates paragraphs. The default value is "[ \t\n\f]", which matches a line starting with a space, tab, newline, or form feed (after its left margin).

Variable sentence-end

This is the regular expression describing the end of a sentence. (All paragraph boundaries also end sentences, regardless.) The default value is:

     "[.?!][]\"')}]*\\($\\| $\\|\t\\| \\)[ \t\n]*"

This means a period, question mark or exclamation mark, followed optionally by a closing parenthetical character, followed by tabs, spaces or new lines. For a detailed explanation of this regular expression, see Section 37.2.2 [Regexp Example], page 556.

568 XEmacs Lisp Reference Manual

38 Syntax Tables

A syntax table specifies the syntactic textual function of each character. This information is used by the parsing commands, the complex movement commands, and others to determine where words, symbols, and other syntactic constructs begin and end. The current syntax table controls the meaning of the word motion functions (see Section 34.2.2 [Word Motion], page 489) and the list motion functions (see Section 34.2.6 [List Motion], page 493) as well as the functions in this chapter.

38.1 Syntax Table Concepts

Under XEmacs 20, a syntax table is a particular subtype of the primitive char table type (see Section 4.13 [Char Tables], page 73), and each element of the char table is an integer that encodes the syntax of the character in question, or a cons of such an integer and a matching character (for characters with parenthesis syntax).

Under XEmacs 19, a syntax table is a vector of 256 elements; it contains one entry for each of the 256 possible characters in an 8-bit byte. Each element is an integer that encodes the syntax of the character in question. (The matching character, if any, is embedded in the bits of this integer.)

Syntax tables are used only for moving across text, not for the Emacs Lisp reader. XEmacs Lisp uses built-in syntactic rules when reading Lisp expressions, and these rules cannot be changed.

Each buffer has its own major mode, and each major mode has its own idea of the syntactic class of various characters. For example, in Lisp mode, the character ‘;’ begins a comment, but in C mode, it terminates a statement. To support these variations, XEmacs makes the choice of syntax table local to each buffer. Typically, each major mode has its own syntax table and installs that table in each buffer that uses that mode. Changing this table alters the syntax in all those buffers as well as in any buffers subsequently put in that mode. Occasionally several similar modes share one syntax table. See Section 26.1.2 [Example Major Modes], page 363, for an example of how to set up a syntax table.

A syntax table can inherit the data for some characters from the standard syntax table, while specifying other characters itself. The “inherit” syntax class means “inherit this character’s syntax from the standard syntax table.” Most major modes’ syntax tables inherit the syntax of character codes 0 through 31 and 128 through 255. This is useful with character sets such as ISO Latin-1 that have additional alphabetic characters in the range 128 to 255. Just changing the standard syntax for these characters affects all major modes.

Function syntax-table-p object

This function returns t if object is a vector of length 256 elements. This means that the vector may be a syntax table. However, according to this test, any vector of length 256 is considered to be a syntax table, no matter what its contents.

38.2 Syntax Descriptors

This section describes the syntax classes and flags that denote the syntax of a character, and how they are represented as a syntax descriptor, which is a Lisp string that you pass to modify-syntax-entry to specify the desired syntax.

XEmacs defines a number of syntax classes. Each syntax table puts each character into one class. There is no necessary relationship between the class of a character in one syntax table and its class in any other table.

Each class is designated by a mnemonic character, which serves as the name of the class when you need to specify a class. Usually the designator character is one that is frequently in that class; however, its meaning as a designator is unvarying and independent of what syntax that character currently has.

A syntax descriptor is a Lisp string that specifies a syntax class, a matching character (used only for the parenthesis classes) and flags. The first character is the designator for a syntax class. The second character is the character to match; if it is unused, put a space there. Then come the characters for any desired flags. If no matching character or flags are needed, one character is sufficient.

For example, the descriptor for the character ‘*’ in C mode is ‘. 23’ (i.e., punctuation, matching character slot unused, second character of a comment-starter, first character of an comment-ender), and the entry for ‘/’ is ‘. 14’ (i.e., punctuation, matching character slot unused, first character of a comment-starter, second character of a comment-ender).

38.2.1 Table of Syntax Classes

Here is a table of syntax classes, the characters that stand for them, their meanings, and examples of their use.

whitespace character Syntax class

Whitespace characters (designated with ‘ ’ or ‘-’) separate symbols and words from each other. Typically, whitespace characters have no other syntactic significance, and multiple whitespace characters are syntactically equivalent to a single one. Space, tab, newline and formfeed are almost always classified as whitespace.

word constituent Syntax class

Word constituents (designated with ‘w’) are parts of normal English words and are typically used in variable and command names in programs. All upper- and lower-case letters, and the digits, are typically word constituents.

symbol constituent Syntax class

Symbol constituents (designated with ‘_’) are the extra characters that are used in variable and command names along with word constituents. For example, the symbol constituents class is used in Lisp mode to indicate that certain characters may be part of symbol names even though they are not part of English words. These characters are ‘$&*+-_<>’. In standard C, the only non-word-constituent character that is valid in symbols is underscore (‘_’).

punctuation character Syntax class

Punctuation characters (‘.’) are those characters that are used as punctuation in English, or are used in some way in a programming language to separate symbols from one another. Most programming language modes, including Emacs Lisp mode, have no characters in this class since the few characters that are not symbol or word constituents all have other uses.

open parenthesis character Syntax class close parenthesis character Syntax class

Open and close parenthesis characters are characters used in dissimilar pairs to surround sentences or expressions. Such a grouping is begun with an open parenthesis character and terminated with a close. Each open parenthesis character matches a particular close parenthesis character, and vice versa. Normally, XEmacs indicates momentarily the matching open parenthesis when you insert a close parenthesis. See Section 45.9 [Blinking], page 658.

The class of open parentheses is designated with ‘(’, and that of close parentheses with ‘)’.

In English text, and in C code, the parenthesis pairs are ‘()’, ‘[]’, and ‘{}’. In XEmacs Lisp, the delimiters for lists and vectors (‘()’ and ‘[]’) are classified as parenthesis characters.

string quote Syntax class

String quote characters (designated with ‘"’) are used in many languages, including Lisp and C, to delimit string constants. The same string quote character appears at the beginning and the end of a string. Such quoted strings do not nest.

The parsing facilities of XEmacs consider a string as a single token. The usual syntactic meanings of the characters in the string are suppressed.

The Lisp modes have two string quote characters: double-quote (‘"’) and vertical bar (‘|’). ‘|’ is not used in XEmacs Lisp, but it is used in Common Lisp. C also has two string quote characters: double-quote for strings, and single-quote (‘'’) for character constants.

English text has no string quote characters because English is not a programming language. Although quotation marks are used in English, we do not want them to turn off the usual syntactic properties of other characters in the quotation.

Syntax class escape

An escape character (designated with ‘\’) starts an escape sequence such as is used in C string and character constants. The character ‘\’ belongs to this class in both C and Lisp. (In C, it is used thus only inside strings, but it turns out to cause no trouble to treat it this way throughout C code.)

Characters in this class count as part of words if words-include-escapes is non-nil. See Section 34.2.2 [Word Motion], page 489.

character quote Syntax class

A character quote character (designated with ‘/’) quotes the following character so that it loses its normal syntactic meaning. This differs from an escape character in that only the character immediately following is ever affected. Characters in this class count as part of words if words-include-escapes is non-nil. See Section 34.2.2 [Word Motion], page 489. This class is used for backslash in TEX mode.

paired delimiter Syntax class

Paired delimiter characters (designated with ‘$’) are like string quote characters except that the syntactic properties of the characters between the delimiters are not suppressed. Only TEX mode uses a paired delimiter presently—the ‘$’ that both enters and leaves math mode.

expression prefix Syntax class

An expression prefix operator (designated with ‘'’) is used for syntactic operators that are part of an expression if they appear next to one. These characters in Lisp include the apostrophe, ‘'’ (used for quoting), the comma, ‘,’ (used in macros), and ‘#’ (used in the read syntax for certain data types).

comment starter Syntax class comment ender Syntax class

The comment starter and comment ender characters are used in various languages to delimit comments. These classes are designated with ‘<’ and ‘>’, respectively. English text has no comment characters. In Lisp, the semicolon (‘;’) starts a comment and a newline or formfeed ends one.

Syntax class inherit

This syntax class does not specify a syntax. It says to look in the standard syntax table to find the syntax of this character. The designator for this syntax code is ‘@’.

38.2.2 Syntax Flags

In addition to the classes, entries for characters in a syntax table can include flags. There are six possible flags, represented by the characters ‘1’, ‘2’, ‘3’, ‘4’, ‘b’ and ‘p’.

All the flags except ‘p’ are used to describe multi-character comment delimiters. The digit flags indicate that a character can also be part of a comment sequence, in addition to the syntactic properties associated with its character class. The flags are independent of the class and each other for the sake of characters such as ‘*’ in C mode, which is a punctuation character, and the second character of a start-of-comment sequence (‘/*’), and the first character of an end-of-comment sequence (‘*/’).

The flags for a character c are:

• ‘1’ means c is the start of a two-character comment-start sequence.

• ‘2’ means c is the second character of such a sequence.

• ‘3’ means c is the start of a two-character comment-end sequence.

• ‘4’ means c is the second character of such a sequence.

• ‘b’ means that c as a comment delimiter belongs to the alternative “b” comment style. Emacs supports two comment styles simultaneously in any one syntax table. This is for the sake of C++. Each style of comment syntax has its own comment-start sequence and its own comment-end sequence. Each comment must stick to one style or the other; thus, if it starts with the comment-start sequence of style “b”, it must also end with the comment-end sequence of style “b”. The two comment-start sequences must begin with the same character; only the second character may differ. Mark the second character of the “b”-style comment-start sequence with the ‘b’ flag. A comment-end sequence (one or two characters) applies to the “b” style if its first character has the ‘b’ flag set; otherwise, it applies to the “a” style. The appropriate comment syntax settings for C++ are as follows:

‘/’  ‘124b’

‘*’ ‘23’

newline ‘>b’

This defines four comment-delimiting sequences:

‘/*’ This is a comment-start sequence for “a” style because the second character, ‘*’, does not have the ‘b’ flag.

‘//’ This is a comment-start sequence for “b” style because the second character, ‘/’, does have the ‘b’ flag.

‘*/’ This is a comment-end sequence for “a” style because the first character, ‘*’, does not have the ‘b’ flag

newline This is a comment-end sequence for “b” style, because the newline charac-

ter has the ‘b’ flag.

• ‘p’ identifies an additional “prefix character” for Lisp syntax. These characters are treated as whitespace when they appear between expressions. When they appear within an expression, they are handled according to their usual syntax codes. The function backward-prefix-chars moves back over these characters, as well as over characters whose primary syntax class is prefix (‘'’). See Section 38.4 [Motion and Syntax], page 575.

38.3 Syntax Table Functions

In this section we describe functions for creating, accessing and altering syntax tables.

Function make-syntax-table &optional table

This function creates a new syntax table. Character codes 0 through 31 and 128 through 255 are set up to inherit from the standard syntax table. The other character codes are set up by copying what the standard syntax table says about them. Most major mode syntax tables are created in this way.

Function copy-syntax-table &optional table

This function constructs a copy of table and returns it. If table is not supplied (or is nil), it returns a copy of the current syntax table. Otherwise, an error is signaled if table is not a syntax table.

Command modify-syntax-entry char syntax-descriptor &optional table This function sets the syntax entry for char according to syntax-descriptor. The syntax is changed only for table, which defaults to the current buffer’s syntax table, and not in any other syntax table. The argument syntax-descriptor specifies the desired syntax; this is a string beginning with a class designator character, and optionally containing a matching character and flags as well. See Section 38.2 [Syntax Descriptors], page 570. This function always returns nil. The old syntax information in the table for this character is discarded. An error is signaled if the first character of the syntax descriptor is not one of the twelve syntax class designator characters. An error is also signaled if char is not a character. Examples:

;; Put the space character in class whitespace.

     (modify-syntax-entry ?\  " ")
          ⇒ nil

;; Make ‘$’ an open parenthesis character,

;; with ‘^’ as its matching close.

     (modify-syntax-entry ?$ "(^")
          ⇒ nil

;; Make ‘^’ a close parenthesis character,

;; with ‘$’ as its matching open.

     (modify-syntax-entry ?^")$")
          ⇒ nil

;; Make ‘/’ a punctuation character,

;; the first character of a start-comment sequence, ;; and the second character of an end-comment sequence. ;; This is used in C mode.

     (modify-syntax-entry ?/ ". 14")
          ⇒ nil

Function char-syntax character

This function returns the syntax class of character, represented by its mnemonic designator character. This only returns the class, not any matching parenthesis or flags. An error is signaled if char is not a character. The following examples apply to C mode. The first example shows that the syntax class of space is whitespace (represented by a space). The second example shows that the syntax of ‘/’ is punctuation. This does not show the fact that it is also part of comment-start and -end sequences. The third example shows that open parenthesis is in the class of open parentheses. This does not show the fact that it has a matching character, ‘)’.

     (char-to-string (char-syntax ?\ ))
          ⇒ " "
     (char-to-string (char-syntax ?/))
          ⇒ "."
     (char-to-string (char-syntax ?\())
          ⇒ "("

Function set-syntax-table table &optional buffer This function makes table the syntax table for buffer, which defaults to the current buffer if omitted. It returns table.

Function syntax-table &optional buffer This function returns the syntax table for buffer, which defaults to the current buffer if omitted.

38.4 Motion and Syntax

This section describes functions for moving across characters in certain syntax classes. None of these functions exists in Emacs version 18 or earlier.

Function skip-syntax-forward syntaxes &optional limit buffer This function moves point forward across characters having syntax classes mentioned in syntaxes. It stops when it encounters the end of the buffer, or position limit (if specified), or a character it is not supposed to skip. Optional argument buffer defaults to the current buffer if omitted.

Function skip-syntax-backward syntaxes &optional limit buffer This function moves point backward across characters whose syntax classes are mentioned in syntaxes. It stops when it encounters the beginning of the buffer, or position limit (if specified), or a character it is not supposed to skip. Optional argument buffer defaults to the current buffer if omitted.

Function backward-prefix-chars &optional buffer

This function moves point backward over any number of characters with expression prefix syntax. This includes both characters in the expression prefix syntax class, and characters with the ‘p’ flag. Optional argument buffer defaults to the current buffer if omitted.

38.5 Parsing Balanced Expressions

Here are several functions for parsing and scanning balanced expressions, also known as sexps, in which parentheses match in pairs. The syntax table controls the interpretation of characters, so these functions can be used for Lisp expressions when in Lisp mode and for C expressions when in C mode. See Section 34.2.6 [List Motion], page 493, for convenient higher-level functions for moving over balanced expressions.

Function parse-partial-sexp start limit &optional target-depth stop-before

state stop-comment buffer This function parses a sexp in the current buffer starting at start, not scanning past limit. It stops at position limit or when certain criteria described below are met, and sets point to the location where parsing stops. It returns a value describing the status of the parse at the point where it stops. If state is nil, start is assumed to be at the top level of parenthesis structure, such as the beginning of a function definition. Alternatively, you might wish to resume parsing in the middle of the structure. To do this, you must provide a state argument that describes the initial status of parsing. If the third argument target-depth is non-nil, parsing stops if the depth in parentheses becomes equal to target-depth. The depth starts at 0, or at whatever is given in state. If the fourth argument stop-before is non-nil, parsing stops when it comes to any character that starts a sexp. If stop-comment is non-nil, parsing stops when it comes to the start of a comment. The fifth argument state is an eight-element list of the same form as the value of this function, described below. The return value of one call may be used to initialize the state of the parse on another call to parse-partial-sexp. The result is a list of eight elements describing the final state of the parse:

0. The depth in parentheses, counting from 0.

1. The character position of the start of the innermost parenthetical grouping containing the stopping point; nil if none.

2. The character position of the start of the last complete subexpression terminated; nil if none.

3. Non-nil if inside a string. More precisely, this is the character that will terminate the string.

4. t if inside a comment (of either style).

5. t if point is just after a quote character.

6. The minimum parenthesis depth encountered during this scan.

7. t if inside a comment of style “b”.

Elements 0, 3, 4, 5 and 7 are significant in the argument state. This function is most often used to compute indentation for languages that have nested parentheses.

Function scan-lists from count depth &optional buffer noerror This function scans forward count balanced parenthetical groupings from character number from. It returns the character position where the scan stops. If depth is nonzero, parenthesis depth counting begins from that value. The only candidates for stopping are places where the depth in parentheses becomes zero; scan-lists counts count such places and then stops. Thus, a positive value for depth means go out depth levels of parenthesis. Scanning ignores comments if parse-sexp-ignore-comments is non-nil.

If the scan reaches the beginning or end of the buffer (or its accessible portion), and the depth is not zero, an error is signaled. If the depth is zero but the count is not used up, nil is returned.

If optional arg buffer is non-nil, scanning occurs in that buffer instead of in the current buffer.

If optional arg noerror is non-nil, scan-lists will return nil instead of signalling an error.

Function scan-sexps from count &optional buffer noerror This function scans forward count sexps from character position from. It returns the character position where the scan stops.

Scanning ignores comments if parse-sexp-ignore-comments is non-nil.

If the scan reaches the beginning or end of (the accessible part of) the buffer in the middle of a parenthetical grouping, an error is signaled. If it reaches the beginning or end between groupings but before count is used up, nil is returned.

If optional arg buffer is non-nil, scanning occurs in that buffer instead of in the current buffer.

If optional arg noerror is non-nil, scan-sexps will return nil instead of signalling an error.

Variable parse-sexp-ignore-comments

If the value is non-nil, then comments are treated as whitespace by the functions in this section and by forward-sexp.

In older Emacs versions, this feature worked only when the comment terminator is something like ‘*/’, and appears only to end a comment. In languages where newlines terminate comments, it was necessary make this variable nil, since not every newline is the end of a comment. This limitation no longer exists.

You can use forward-comment to move forward or backward over one comment or several comments.

Function forward-comment count &optional buffer This function moves point forward across count comments (backward, if count is negative). If it finds anything other than a comment or whitespace, it stops, leaving point at the place where it stopped. It also stops after satisfying count.

Optional argument buffer defaults to the current buffer.

To move forward over all comments and whitespace following point, use (forwardcomment (buffer-size)). (buffer-size) is a good argument to use, because the number of comments in the buffer cannot exceed that many.

38.6 Some Standard Syntax Tables

Most of the major modes in XEmacs have their own syntax tables. Here are several of them:

Function standard-syntax-table

This function returns the standard syntax table, which is the syntax table used in Fundamental mode.

Variable text-mode-syntax-table

The value of this variable is the syntax table used in Text mode.

Variable c-mode-syntax-table

The value of this variable is the syntax table for C-mode buffers.

Variable emacs-lisp-mode-syntax-table

The value of this variable is the syntax table used in Emacs Lisp mode by editing commands. (It has no effect on the Lisp read function.)

38.7 Syntax Table Internals

Each element of a syntax table is an integer that encodes the syntax of one character: the syntax class, possible matching character, and flags. Lisp programs don’t usually work with the elements directly; the Lisp-level syntax table functions usually work with syntax descriptors (see Section 38.2 [Syntax Descriptors], page 570).

The low 8 bits of each element of a syntax table indicate the syntax class. Integer Class 0 whitespace 1 punctuation 2 word 3 symbol 4 open parenthesis 5 close parenthesis 6 expression prefix 7 string quote 8 paired delimiter 9 escape 10 character quote 11 comment-start 12 comment-end 13 inherit

The next 8 bits are the matching opposite parenthesis (if the character has parenthesis syntax); otherwise, they are not meaningful. The next 6 bits are the flags.

39 Abbrevs And Abbrev Expansion

An abbreviation or abbrev is a string of characters that may be expanded to a longer string. The user can insert the abbrev string and find it replaced automatically with the expansion of the abbrev. This saves typing.

The set of abbrevs currently in effect is recorded in an abbrev table. Each buffer has a local abbrev table, but normally all buffers in the same major mode share one abbrev table. There is also a global abbrev table. Normally both are used.

An abbrev table is represented as an obarray containing a symbol for each abbreviation. The symbol’s name is the abbreviation; its value is the expansion; its function definition is the hook function to do the expansion (see Section 39.3 [Defining Abbrevs], page 580); its property list cell contains the use count, the number of times the abbreviation has been expanded. Because these symbols are not interned in the usual obarray, they will never appear as the result of reading a Lisp expression; in fact, normally they are never used except by the code that handles abbrevs. Therefore, it is safe to use them in an extremely nonstandard way. See Section 7.3 [Creating Symbols], page 113.

For the user-level commands for abbrevs, see section “Abbrev Mode” in The XEmacs Reference Manual.

39.1 Setting Up Abbrev Mode

Abbrev mode is a minor mode controlled by the value of the variable abbrev-mode.

Variable abbrev-mode

A non-nil value of this variable turns on the automatic expansion of abbrevs when their abbreviations are inserted into a buffer. If the value is nil, abbrevs may be defined, but they are not expanded automatically.

This variable automatically becomes local when set in any fashion.

Variable default-abbrev-mode

This is the value of abbrev-mode for buffers that do not override it. This is the

     same as (default-value 'abbrev-mode).

39.2 Abbrev Tables

This section describes how to create and manipulate abbrev tables.

Function make-abbrev-table

This function creates and returns a new, empty abbrev table—an obarray containing no symbols. It is a vector filled with zeros.

Function clear-abbrev-table table

This function undefines all the abbrevs in abbrev table table, leaving it empty.

The function returns nil.

Function define-abbrev-table tabname definitions

This function defines tabname (a symbol) as an abbrev table name, i.e., as a variable whose value is an abbrev table. It defines abbrevs in the table according to definitions, a list of elements of the form (abbrevname expansion hook usecount). The value is always nil.

Variable abbrev-table-name-list

This is a list of symbols whose values are abbrev tables. define-abbrev-table adds the new abbrev table name to this list.

Function insert-abbrev-table-description name &optional human

This function inserts before point a description of the abbrev table named name. The argument name is a symbol whose value is an abbrev table. The value is always nil. If human is non-nil, the description is human-oriented. Otherwise the description is a Lisp expression—a call to define-abbrev-table that would define name exactly as it is currently defined.

39.3 Defining Abbrevs

These functions define an abbrev in a specified abbrev table. define-abbrev is the low-level basic function, while add-abbrev is used by commands that ask for information from the user.

Function add-abbrev table type arg This function adds an abbreviation to abbrev table table based on information from the user. The argument type is a string describing in English the kind of abbrev this will be (typically, "global" or "mode-specific"); this is used in prompting the user. The argument arg is the number of words in the expansion. The return value is the symbol that internally represents the new abbrev, or nil if the user declines to confirm redefining an existing abbrev.

Function define-abbrev table name expansion hook This function defines an abbrev in table named name, to expand to expansion, and call hook. The return value is an uninterned symbol that represents the abbrev inside XEmacs; its name is name. The argument name should be a string. The argument expansion should be a string, or nil to undefine the abbrev. The argument hook is a function or nil. If hook is non-nil, then it is called with no arguments after the abbrev is replaced with expansion; point is located at the end of expansion when hook is called. The use count of the abbrev is initialized to zero.

User Option only-global-abbrevs

If this variable is non-nil, it means that the user plans to use global abbrevs only. This tells the commands that define mode-specific abbrevs to define global ones instead. This variable does not alter the behavior of the functions in this section; it is examined by their callers.

39.4 Saving Abbrevs in Files

A file of saved abbrev definitions is actually a file of Lisp code. The abbrevs are saved in the form of a Lisp program to define the same abbrev tables with the same contents. Therefore, you can load the file with load (see Section 14.1 [How Programs Do Loading], page 195). However, the function quietly-read-abbrev-file is provided as a more convenient interface.

User-level facilities such as save-some-buffers can save abbrevs in a file automatically, under the control of variables described here.

User Option abbrev-file-name

This is the default file name for reading and saving abbrevs.

Function quietly-read-abbrev-file filename

This function reads abbrev definitions from a file named filename, previously written with write-abbrev-file. If filename is nil, the file specified in abbrev-file-name is used. save-abbrevs is set to t so that changes will be saved.

This function does not display any messages. It returns nil.

User Option save-abbrevs

A non-nil value for save-abbrev means that XEmacs should save abbrevs when files are saved. abbrev-file-name specifies the file to save the abbrevs in.

Variable abbrevs-changed

This variable is set non-nil by defining or altering any abbrevs. This serves as a flag for various XEmacs commands to offer to save your abbrevs.

Command write-abbrev-file filename

Save all abbrev definitions, in all abbrev tables, in the file filename, in the form of a Lisp program that when loaded will define the same abbrevs. This function returns nil.

39.5 Looking Up and Expanding Abbreviations

Abbrevs are usually expanded by commands for interactive use, including self-insert-command. This section describes the subroutines used in writing such functions, as well as the variables they use for communication.

Function abbrev-symbol abbrev &optional table This function returns the symbol representing the abbrev named abbrev. The value returned is nil if that abbrev is not defined. The optional second argument table is the abbrev table to look it up in. If table is nil, this function tries first the current buffer’s local abbrev table, and second the global abbrev table.

Function abbrev-expansion abbrev &optional table This function returns the string that abbrev would expand into (as defined by the abbrev tables used for the current buffer). The optional argument table specifies the abbrev table to use, as in abbrev-symbol.

Command expand-abbrev

This command expands the abbrev before point, if any. If point does not follow an abbrev, this command does nothing. The command returns t if it did expansion, nil otherwise.

Command abbrev-prefix-mark &optional arg

Mark current point as the beginning of an abbrev. The next call to expandabbrev will use the text from here to point (where it is then) as the abbrev to expand, rather than using the previous word as usual.

User Option abbrev-all-caps

When this is set non-nil, an abbrev entered entirely in upper case is expanded using all upper case. Otherwise, an abbrev entered entirely in upper case is expanded by capitalizing each word of the expansion.

Variable abbrev-start-location

This is the buffer position for expand-abbrev to use as the start of the next abbrev to be expanded. (nil means use the word before point instead.) abbrevstart-location is set to nil each time expand-abbrev is called. This variable

is also set by abbrev-prefix-mark.

Variable abbrev-start-location-buffer

The value of this variable is the buffer for which abbrev-start-location has been set. Trying to expand an abbrev in any other buffer clears abbrev-start-

location. This variable is set by abbrev-prefix-mark.

Variable last-abbrev

This is the abbrev-symbol of the last abbrev expanded. This information is left by expand-abbrev for the sake of the unexpand-abbrev command.

Variable last-abbrev-location

This is the location of the last abbrev expanded. This contains information left

by expand-abbrev for the sake of the unexpand-abbrev command.

Variable last-abbrev-text

This is the exact expansion text of the last abbrev expanded, after case conversion (if any). Its value is nil if the abbrev has already been unexpanded. This contains information left by expand-abbrev for the sake of the unexpandabbrev command.

Variable pre-abbrev-expand-hook

This is a normal hook whose functions are executed, in sequence, just before any expansion of an abbrev. See Section 26.4 [Hooks], page 378. Since it is a normal hook, the hook functions receive no arguments. However, they can find the abbrev to be expanded by looking in the buffer before point.

The following sample code shows a simple use of pre-abbrev-expand-hook. If the user terminates an abbrev with a punctuation character, the hook function asks for confirmation. Thus, this hook allows the user to decide whether to expand the abbrev, and aborts expansion if it is not confirmed.

(add-hook 'pre-abbrev-expand-hook 'query-if-not-space)

;; This is the function invoked by pre-abbrev-expand-hook.

;; If the user terminated the abbrev with a space, the function does ;; nothing (that is, it returns so that the abbrev can expand). If the ;; user entered some other character, this function asks whether ;; expansion should continue.

;; If the user answers the prompt with y, the function returns ;; nil (because of the not function), but that is ;; acceptable; the return value has no effect on expansion.

(defun query-if-not-space ()
  (if (/= ?\  (preceding-char))
      (if (not (y-or-n-p "Do you want to expand this abbrev? "))
          (error "Not expanding this abbrev"))))

39.6 Standard Abbrev Tables

Here we list the variables that hold the abbrev tables for the preloaded major modes of XEmacs.

Variable global-abbrev-table

This is the abbrev table for mode-independent abbrevs. The abbrevs defined in it apply to all buffers. Each buffer may also have a local abbrev table, whose abbrev definitions take precedence over those in the global table.

Variable local-abbrev-table

The value of this buffer-local variable is the (mode-specific) abbreviation table of the current buffer.

Variable fundamental-mode-abbrev-table

This is the local abbrev table used in Fundamental mode; in other words, it is the local abbrev table in all buffers in Fundamental mode.

Variable text-mode-abbrev-table

This is the local abbrev table used in Text mode.

Variable c-mode-abbrev-table

This is the local abbrev table used in C mode.

Variable lisp-mode-abbrev-table

This is the local abbrev table used in Lisp mode and Emacs Lisp mode.

584 XEmacs Lisp Reference Manual

40 Extents

An extent is a region of text (a start position and an end position) that is displayed in a particular face and can have certain other properties such as being read-only. Extents can overlap each other. XEmacs efficiently handles buffers with large numbers of extents in them.

Function extentp object

This returns t if object is an extent.

40.1 Introduction to Extents

An extent is a region of text within a buffer or string that has certain properties associated with it. The properties of an extent primarily affect the way the text contained in the extent is displayed. Extents can freely overlap each other in a buffer or string. Extents are invisible to functions that merely examine the text of a buffer or string.

Please note: An alternative way to add properties to a buffer or string is to use text properties. See Section 36.18 [Text Properties], page 539.

An extent is logically a Lisp object consisting of a start position, an end position, a buffer or string to which these positions refer, and a property list. As text is inserted into a buffer, the start and end positions of the extent are automatically adjusted as necessary to keep the extent referring to the same text in the buffer. If text is inserted at the boundary of an extent, the extent’s start-open and end-open properties control whether the text is included as part of the extent. If the text bounded by an extent is deleted, the extent becomes detached; its start and end positions are no longer meaningful, but it maintains all its other properties and can later be reinserted into a buffer. (None of these considerations apply to strings, because text cannot be inserted into or deleted from a string.)

Each extent has a face or list of faces associated with it, which controls the way in which the text bounded by the extent is displayed. If an extent’s face is nil or its properties are partially undefined, the corresponding properties from the default face for the frame is used. If two or more extents overlap, or if a list of more than one face is specified for a particular extent, the corresponding faces are merged to determine the text’s displayed properties. Every extent has a priority that determines which face takes precedence if the faces conflict. (If two extents have the same priority, the one that comes later in the display order takes precedence. See Section 40.3 [Extent Endpoints], page 586.) Higher-numbered priority values correspond to a higher priority, and priority values can be negative. Every extent is created with a priority of 0, but this can be changed with set-extent-priority. Within a single extent with a list of faces, faces earlier in the list have a higher priority than faces later in the list.

Extents can be set to respond specially to key and mouse events within the extent. An extent’s keymap property controls the effect of key and mouse strokes within the extent’s text, and the mouse-face property controls whether the extent is highlighted when the mouse moves over it. See Section 40.10 [Extents and Events], page 597.

An extent can optionally have a begin-glyph or end-glyph associated with it. A beginglyph or end-glyph is a pixmap or string that will be displayed either at the start or end of an extent or in the margin of the line that the start or end of the extent lies in, depending on the extent’s layout policy. Begin-glyphs and end-glyphs are used to implement annotations, and you should use the annotation API functions in preference to the lower-level extent functions. For more information, See Chapter 44 [Annotations], page 641.

If an extent has its detachable property set, it will become detached (i.e. no longer in the buffer) when all its text its deleted. Otherwise, it will simply shrink down to zerolength and sit it the same place in the buffer. By default, the detachable property is set on newly-created extents. See Section 40.7 [Detached Extents], page 595.

If an extent has its duplicable property set, it will be remembered when a string is created from text bounded by the extent. When the string is re-inserted into a buffer, the extent will also be re-inserted. This mechanism is used in the kill, yank, and undo commands. See Section 40.9 [Duplicable Extents], page 596.

40.2 Creating and Modifying Extents

Function make-extent from to &optional object This function makes an extent for the range [from, to) in object (a buffer or string). object defaults to the current buffer. Insertions at point to will be outside of the extent; insertions at from will be inside the extent, causing the extent to grow (see Section 40.3 [Extent Endpoints], page 586). This is the same way that markers behave. The extent is initially detached if both from and to are nil, and in this case object defaults to nil, meaning the extent is in no buffer or string (see Section 40.7 [Detached Extents], page 595).

Function delete-extent extent

This function removes extent from its buffer and destroys it. This does not modify the buffer’s text, only its display properties. The extent cannot be used thereafter. To remove an extent in such a way that it can be re-inserted later, use detach-extent. See Section 40.7 [Detached Extents], page 595.

Function extent-object extent

This function returns the buffer or string that extent is in. If the return value is nil, this means that the extent is detached; however, a detached extent will not necessarily return a value of nil.

Function extent-live-p extent

This function returns nil if extent is deleted, and t otherwise.

40.3 Extent Endpoints

Every extent has a start position and an end position, and logically affects the characters between those positions. Normally the start and end positions must both be valid positions in the extent’s buffer or string. However, both endpoints can be nil, meaning the extent is detached. See Section 40.7 [Detached Extents], page 595.

Whether the extent overlaps its endpoints is governed by its start-open and end-open properties. Insertion of a character at a closed endpoint will expand the extent to include that character; insertion at an open endpoint will not. Similarly, functions such as extentat that scan over all extents overlapping a particular position will include extents with a closed endpoint at that position, but not extents with an open endpoint.

Note that the start-closed and end-closed properties are equivalent to start-open and end-open with the opposite sense.

Both endpoints can be equal, in which case the extent includes no characters but still exists in the buffer or string. Zero-length extents are used to represent annotations (see Chapter 44 [Annotations], page 641) and can be used as a more powerful form of a marker. Deletion of all the characters in an extent may or may not result in a zero-length extent; this depends on the detachable property (see Section 40.7 [Detached Extents], page 595). Insertion at the position of a zero-length extent expands the extent if both endpoints are closed; goes before the extent if it has the start-open property; and goes after the extent if it has the end-open property. Zero-length extents with both the start-open and end-open properties are treated as if their starting point were closed. Deletion of a character on a side of a zero-length extent whose corresponding endpoint is closed causes the extent to be detached if its detachable property is set; if the corresponding endpoint is open, the extent remains in the buffer, moving as necessary.

Extents are ordered within a buffer or string by increasing start position, and then by decreasing end position (this is called the display order).

Function extent-start-position extent

This function returns the start position of extent.

Function extent-end-position extent

This function returns the end position of extent.

Function extent-length extent

This function returns the length of extent in characters. If the extent is detached, this returns 0. If the extent is not detached, this is equivalent to

          (- (extent-end-position extent) (extent-start-position extent))

Function set-extent-endpoints extent start end &optional buffer-or-string This function sets the start and end position of extent to start and end. If both are nil, this is equivalent to detach-extent. buffer-or-string specifies the new buffer or string that the extent should be in, and defaults to extent’s buffer or string. (If nil, and extent is in no buffer and no string, it defaults to the current buffer.) See documentation on detach-extent for a discussion of undo recording.

40.4 Finding Extents

The following functions provide a simple way of determining the extents in a buffer or string. A number of more sophisticated primitives for mapping over the extents in a range of a buffer or string are also provided (see Section 40.5 [Mapping Over Extents], page 589). When reading through this section, keep in mind the way that extents are ordered (see Section 40.3 [Extent Endpoints], page 586).

Function extent-list &optional buffer-or-string from to flags This function returns a list of the extents in buffer-or-string. buffer-or-string defaults to the current buffer if omitted. from and to can be used to limit the range over which extents are returned; if omitted, all extents in the buffer or string are returned. More specifically, if a range is specified using from and to, only extents that overlap the range (i.e. begin or end inside of the range) are included in the list. from and to default to the beginning and end of buffer-or-string, respectively. flags controls how end cases are treated. For a discussion of this, and exactly what “overlap” means, see map-extents.

Functions that create extents must be prepared for the possibility that there are other extents in the same area, created by other functions. To deal with this, functions typically mark their own extents by setting a particular property on them. The following function makes it easier to locate those extents.

Function extent-at pos &optional object property before at-flag This function finds the “smallest” extent (i.e., the last one in the display order) at (i.e., overlapping) pos in object (a buffer or string) having property set. object defaults to the current buffer. property defaults to nil, meaning that any extent will do. Returns nil if there is no matching extent at pos. If the fourth argument before is not nil, it must be an extent; any returned extent will precede that extent. This feature allows extent-at to be used by a loop over extents. at-flag controls how end cases are handled (i.e. what “at” really means), and should be one of:

nil

after An extent is at pos if it covers the character after pos. This is consistent with the way that text properties work.

before An extent is at pos if it covers the character before pos.

at An extent is at pos if it overlaps or abuts pos. This includes all zero-length extents at pos.

Note that in all cases, the start-openness and end-openness of the extents considered is ignored. If you want to pay attention to those properties, you should use map-extents, which gives you more control.

The following low-level functions are provided for explicitly traversing the extents in a buffer according to the display order. These functions are mostly intended for debugging – in normal operation, you should probably use mapcar-extents or map-extents, or loop using the before argument to extent-at, rather than creating a loop using next-extent.

Function next-extent extent

Given an extent extent, this function returns the next extent in the buffer or string’s display order. If extent is a buffer or string, this returns the first extent in the buffer or string.

Function previous-extent extent

Given an extent extent, this function returns the previous extent in the buffer or string’s display order. If extent is a buffer or string, this returns the last extent in the buffer or string.

40.5 Mapping Over Extents

The most basic and general function for mapping over extents is called map-extents. You should read through the definition of this function to familiarize yourself with the concepts and optional arguments involved. However, in practice you may find it more convenient to use the function mapcar-extents or to create a loop using the before argument to extent-at (see Section 40.4 [Finding Extents], page 587).

Function map-extents function &optional object from to maparg flags

property value This function maps function over the extents which overlap a region in object. object is normally a buffer or string but could be an extent (see below). The region is normally bounded by [from, to) (i.e. the beginning of the region is closed and the end of the region is open), but this can be changed with the flags argument (see below for a complete discussion).

function is called with the arguments (extent, maparg). The arguments object, from, to, maparg, and flags are all optional and default to the current buffer, the beginning of object, the end of object, nil, and nil, respectively. map-extents returns the first non-nil result produced by function, and no more calls to function are made after it returns non-nil.

If object is an extent, from and to default to the extent’s endpoints, and the mapping omits that extent and its predecessors. This feature supports restarting a loop based on map-extents. Note: object must be attached to a buffer or string, and the mapping is done over that buffer or string.

An extent overlaps the region if there is any point in the extent that is also in the region. (For the purpose of overlap, zero-length extents and regions are treated as closed on both ends regardless of their endpoints’ specified open/closedness.) Note that the endpoints of an extent or region are considered to be in that extent or region if and only if the corresponding end is closed. For example, the extent [5,7] overlaps the region [2,5] because 5 is in both the extent and the region. However, (5,7] does not overlap [2,5] because 5 is not in the extent, and neither [5,7] nor (5,7] overlaps the region [2,5) because 5 is not in the region.

The optional flags can be a symbol or a list of one or more symbols, modifying the behavior of map-extents. Allowed symbols are:

end-closed

The region’s end is closed.

start-open

The region’s start is open.

all-extents-closed

Treat all extents as closed on both ends for the purpose of determining whether they overlap the region, irrespective of their actual open- or closedness.

all-extents-open

Treat all extents as open on both ends.

all-extents-closed-open

Treat all extents as start-closed, end-open.

all-extents-open-closed

Treat all extents as start-open, end-closed.

start-in-region

In addition to the above conditions for extent overlap, the extent’s start position must lie within the specified region. Note that, for this condition, open start positions are treated as if 0.5 was added to the endpoint’s value, and open end positions are treated as if 0.5 was subtracted from the endpoint’s value.

end-in-region

The extent’s end position must lie within the region.

start-and-end-in-region

Both the extent’s start and end positions must lie within the region.

start-or-end-in-region

Either the extent’s start or end position must lie within the region.

negate-in-region

The condition specified by a *-in-region flag must not hold for the extent to be considered.

At most one of all-extents-closed, all-extents-open, all-extents-
closed-open, and all-extents-open-closed may be specified.
At most one of start-in-region, end-in-region, start-and-end-in-region,
and start-or-end-in-region may be specified.

If optional arg property is non-nil, only extents with that property set on them will be visited. If optional arg value is non-nil, only extents whose value for that property is eq to value will be visited.

If you want to map over extents and accumulate a list of results, the following function may be more convenient than map-extents.

Function mapcar-extents function &optional predicate buffer-or-string

from to flags property value This function applies function to all extents which overlap a region in bufferor-string. The region is delimited by from and to. function is called with one argument, the extent. A list of the values returned by function is returned. An optional predicate may be used to further limit the extents over which function is mapped. The optional arguments flags, property, and value may also be used to control the extents passed to predicate or function, and have the same meaning as in map-extents.

Function map-extent-children function &optional object from to maparg

flags property value This function is similar to map-extents, but differs in that:

• It only visits extents which start in the given region.

• After visiting an extent e, it skips all other extents which start inside e but end before e’s end.

Thus, this function may be used to walk a tree of extents in a buffer:

     (defun walk-extents (buffer &optional ignore)
       (map-extent-children 'walk-extents buffer))

Function extent-in-region-p extent &optional from to flags This function returns t if map-extents would visit extent if called with the given arguments.

40.6 Properties of Extents

Each extent has a property list associating property names with values. Some property names have predefined meanings, and can usually only assume particular values. Assigning other values to such a property either cause the value to be converted into a legal value (e.g., assigning anything but nil to a Boolean property will cause the value of t to be assigned to the property) or will cause an error. Property names without predefined meanings can be assigned any value. An undefined property is equivalent to a property with a value of nil, or with a particular default value in the case of properties with predefined meanings. Note that, when an extent is created, the end-open and detachable properties are set on it.

If an extent has a parent, all of its properties actually derive from that parent (or from the root ancestor if the parent in turn has a parent), and setting a property of the extent actually sets that property on the parent. See Section 40.8 [Extent Parents], page 596.

Function extent-property extent property

This function returns the value of property in extent. If property is undefined, nil is returned.

Function extent-properties extent

This function returns a list of all of extent’s properties that do not have the value of nil (or the default value, for properties with predefined meanings).

Function set-extent-property extent property value

This function sets property to value in extent. (If property has a predefined meaning, only certain values are allowed, and some values may be converted to others before being stored.)

Function set-extent-properties extent plist

Change some properties of extent. plist is a property list. This is useful to change many extent properties at once.

The following table lists the properties with predefined meanings, along with their allowable values.

detached (Boolean) Whether the extent is detached. Setting this is the same as calling detach-extent. See Section 40.7 [Detached Extents], page 595.

destroyed

(Boolean) Whether the extent has been deleted. Setting this is the same as

calling delete-extent.

priority (integer) The extent’s redisplay priority. Defaults to 0. See Section 40.1 [Intro to Extents], page 585. This property can also be set with set-extent-priority and accessed with extent-priority.

start-open

(Boolean) Whether the start position of the extent is open, meaning that characters inserted at that position go outside of the extent. See Section 40.3 [Extent Endpoints], page 586.

start-closed

(Boolean) Same as start-open but with the opposite sense. Setting this property clears start-open and vice-versa.

end-open (Boolean) Whether the end position of the extent is open, meaning that characters inserted at that position go outside of the extent. This is t by default. See Section 40.3 [Extent Endpoints], page 586.

end-closed

(Boolean) Same as end-open but with the opposite sense. Setting this property clears end-open and vice-versa.

read-only

(Boolean) Whether text within this extent will be unmodifiable.

face (face, face name, list of faces or face names, or nil) The face in which to display the extent’s text. This property can also be set with set-extent-face and accessed with extent-face. Note that if a list of faces is specified, the faces are merged together, with faces earlier in the list having priority over faces later in the list.

mouse-face

(face, face name, list of faces or face names, or nil) The face used to display the extent when the mouse moves over it. This property can also be set with

set-extent-mouse-face and accessed with extent-mouse-face. Note that if

a list of faces is specified, the faces are merged together, with faces earlier in the list having priority over faces later in the list. See Section 40.10 [Extents and Events], page 597. pointer (pointer glyph) The glyph used as the pointer when the mouse moves over the

extent. This takes precedence over the text-pointer-glyph and nontextpointer-glyph variables. If for any reason this glyph is an invalid pointer, the standard glyphs will be used as fallbacks. See Section 43.4 [Mouse Pointer], page 638.

detachable

(Boolean) Whether this extent becomes detached when all of the text it covers is deleted. This is t by default. See Section 40.7 [Detached Extents], page 595.

duplicable

(Boolean) Whether this extent should be copied into strings, so that kill, yank, and undo commands will restore or copy it. See Section 40.9 [Duplicable Extents], page 596.

unique (Boolean) Meaningful only in conjunction with duplicable. When this is set, there may be only one instance of this extent attached at a time. See Section 40.9 [Duplicable Extents], page 596.

invisible

(Boolean) If t, text under this extent will not be displayed – it will look as if the text is not there at all.

keymap (keymap or nil) This keymap is consulted for mouse clicks on this extent or keypresses made while point is within the extent. See Section 40.10 [Extents and Events], page 597.

copy-function

This is a hook that is run when a duplicable extent is about to be copied from a buffer to a string (or the kill ring). See Section 40.9 [Duplicable Extents], page 596.

paste-function

This is a hook that is run when a duplicable extent is about to be copied from a string (or the kill ring) into a buffer. See Section 40.9 [Duplicable Extents], page 596.

begin-glyph

(glyph or nil) This extent’s begin glyph. See Chapter 44 [Annotations], page 641.

end-glyph

(glyph or nil) This extent’s end glyph. See Chapter 44 [Annotations], page 641.

begin-glyph-layout
(text, whitespace, inside-margin, or outside-margin) The layout policy

for this extent’s begin glyph. Defaults to text. See Chapter 44 [Annotations], page 641.

end-glyph-layout
(text, whitespace, inside-margin, or outside-margin) The layout policy

for this extent’s end glyph. Defaults to text. See Chapter 44 [Annotations], page 641.

initial-redisplay-function

(any funcallable object) The function to be called the first time (a part of) the extent is redisplayed. It will be called with the extent as its argument. This is used by lazy-shot to implement lazy font-locking. The functionality is still experimental, and may change without further notice.

The following convenience functions are provided for accessing particular properties of an extent.

Function extent-face extent

This function returns the face property of extent. This might also return a list of face names. Do not modify this list directly! Instead, use set-extent-face. Note that you can use eq to compare lists of faces as returned by extent-face. In other words, if you set the face of two different extents to two lists that are equal but not eq, then the return value of extent-face on the two extents will return the identical list.

Function extent-mouse-face extent

This function returns the mouse-face property of extent. This might also return a list of face names. Do not modify this list directly! Instead, use

set-extent-mouse-face.

Note that you can use eq to compare lists of faces as returned by extent-

mouse-face, just like for extent-face.

Function extent-priority extent

This function returns the priority property of extent.

Function extent-keymap extent

This function returns the keymap property of extent.

Function extent-begin-glyph-layout extent

This function returns the begin-glyph-layout property of extent, i.e. the layout policy associated with the extent’s begin glyph.

Function extent-end-glyph-layout extent

This function returns the end-glyph-layout property of extent, i.e. the layout policy associated with the extent’s end glyph.

Function extent-begin-glyph extent

This function returns the begin-glyph property of extent, i.e. the glyph object displayed at the beginning of extent. If there is none, nil is returned.

Function extent-end-glyph extent

This function returns the end-glyph property of extent, i.e. the glyph object displayed at the end of extent. If there is none, nil is returned.

The following convenience functions are provided for setting particular properties of an extent.

Function set-extent-priority extent pri

This function sets the priority property of extent to pri.

Function set-extent-face extent face

This function sets the face property of extent to face.

Function set-extent-mouse-face extent face

This function sets the mouse-face property of extent to face.

Function set-extent-keymap extent keymap

This function sets the keymap property of extent to keymap. keymap must be either a keymap object, or nil.

Function set-extent-begin-glyph-layout extent layout

This function sets the begin-glyph-layout property of extent to layout.

Function set-extent-end-glyph-layout extent layout

This function sets the end-glyph-layout property of extent to layout.

Function set-extent-begin-glyph extent begin-glyph &optional layout This function sets the begin-glyph and glyph-layout properties of extent to begin-glyph and layout, respectively. (layout defaults to text if not specified.)

Function set-extent-end-glyph extent end-glyph &optional layout This function sets the end-glyph and glyph-layout properties of extent to end-glyph and layout, respectively. (layout defaults to text if not specified.)

Function set-extent-initial-redisplay-function extent function

This function sets the initial-redisplay-function property of the extent to function.

40.7 Detached Extents

A detached extent is an extent that is not attached to a buffer or string but can be re-inserted. Detached extents have a start position and end position of nil. Extents can be explicitly detached using detach-extent. An extent is also detached when all of its characters are all killed by a deletion, if its detachable property is set; if this property is not set, the extent becomes a zero-length extent. (Zero-length extents with the detachable property set behave specially. See Section 40.3 [Extent Endpoints], page 586.)

Function detach-extent extent

This function detaches extent from its buffer or string. If extent has the duplicable property, its detachment is tracked by the undo mechanism. See Section 40.9 [Duplicable Extents], page 596.

Function extent-detached-p extent

This function returns nil if extent is detached, and t otherwise.

Function copy-extent extent &optional object This function makes a copy of extent. It is initially detached. Optional argument object defaults to extent’s object (normally a buffer or string, but could be nil).

Function insert-extent extent &optional start end no-hooks object This function inserts extent from start to end in object (a buffer or string). If extent is detached from a different buffer or string, or in most cases when extent is already attached, the extent will first be copied as if with copy-extent. This function operates the same as if insert were called on a string whose extent data calls for extent to be inserted, except that if no-hooks is non-nil, extent’s paste-function will not be invoked. See Section 40.9 [Duplicable Extents], page 596.

40.8 Extent Parents

An extent can have a parent extent set for it. If this is the case, the extent derives all its properties from that extent and has no properties of its own. The only “properties” that the extent keeps are the buffer or string it refers to and the start and end points. (More correctly, the extent’s own properties are shadowed. If you later change the extent to have no parent, its own properties will become visible again.)

It is possible for an extent’s parent to itself have a parent, and so on. Through this, a whole tree of extents can be created, all deriving their properties from one root extent. Note, however, that you cannot create an inheritance loop – this is explicitly disallowed.

Parent extents are used to implement the extents over the modeline.

Function set-extent-parent extent parent

This function sets the parent of extent to parent. If parent is nil, the extent is set to have no parent.

Function extent-parent extent

This function return the parents (if any) of extent, or nil.

Function extent-children extent

This function returns a list of the children (if any) of extent. The children of an extent are all those extents whose parent is that extent. This function does not recursively trace children of children.

Function extent-descendants extent

This function returns a list of all descendants of extent, including extent. This recursively applies extent-children to any children of extent, until no more children can be found.

40.9 Duplicable Extents

If an extent has the duplicable property, it will be copied into strings, so that kill, yank, and undo commands will restore or copy it.

Specifically:

• When a string is created using buffer-substring or buffer-string, any duplicable extents in the region corresponding to the string will be copied into the string (see Section 36.2 [Buffer Contents], page 512). When the string in inserted into a buffer using

insert, insert-before-markers, insert-buffer or insert-buffer-substring, the

extents in the string will be copied back into the buffer (see Section 36.4 [Insertion], page 513). The extents in a string can, of course, be retrieved explicitly using the standard extent primitives over the string.

• Similarly, when text is copied or cut into the kill ring, any duplicable extents will be remembered and reinserted later when the text is pasted back into a buffer.

• When concat is called on strings, the extents in the strings are copied into the resulting string.

• When substring is called on a string, the relevant extents are copied into the resulting string.

• When a duplicable extent is detached by detach-extent or string deletion, or inserted by insert-extent or string insertion, the action is recorded by the undo mechanism so that it can be undone later. Note that if an extent gets detached and then a later undo causes the extent to get reinserted, the new extent will not be ‘eq’ to the original extent.

• Extent motion, face changes, and attachment via make-extent are not recorded by the undo mechanism. This means that extent changes which are to be undo-able must be performed by character editing, or by insertion and detachment of duplicable extents.

• A duplicable extent’s copy-function property, if non-nil, should be a function, and will be run when a duplicable extent is about to be copied from a buffer to a string (or the kill ring). It is called with three arguments: the extent and the buffer positions within it which are being copied. If this function returns nil, then the extent will not be copied; otherwise it will.

• A duplicable extent’s paste-function property, if non-nil, should be a function, and will be run when a duplicable extent is about to be copied from a string (or the kill ring) into a buffer. It is called with three arguments: the original extent and the buffer positions which the copied extent will occupy. (This hook is run after the corresponding text has already been inserted into the buffer.) Note that the extent argument may be detached when this function is run. If this function returns nil, no extent will be inserted. Otherwise, there will be an extent covering the range in question. Note: if the extent to be copied is already attached to the buffer and overlaps the new range, the extent will simply be extended and the paste-function will not be called.

40.10 Interaction of Extents with Keyboard and Mouse Events

If an extent has the mouse-face property set, it will be highlighted when the mouse passes over it. Highlighting is accomplished by merging the extent’s face with the face or faces specified by the mouse-face property. The effect is as if a pseudo-extent with the mouse-face face were inserted after the extent in the display order (see Section 40.3 [Extent Endpoints], page 586, display order).

Variable mouse-highlight-priority

This variable holds the priority to use when merging in the highlighting pseudoextent. The default is 1000. This is purposely set very high so that the highlighting pseudo-extent shows up even if there are other extents with various priorities at the same location.

You can also explicitly cause an extent to be highlighted. Only one extent at a time can be highlighted in this fashion, and any other highlighted extent will be de-highlighted.

Function highlight-extent extent &optional highlight-p This function highlights (if highlight-p is non-nil) or de-highlights (if highlightp is nil) extent, if extent has the mouse-face property. (Nothing happens if extent does not have the mouse-face property.)

Function force-highlight-extent extent &optional highlight-p This function is similar to highlight-extent but highlights or de-highlights the extent regardless of whether it has the mouse-face property.

If an extent has a keymap property, this keymap will be consulted for mouse clicks on the extent and keypresses made while point is within the extent. The behavior of mouse clicks and keystrokes not defined in the keymap is as normal for the buffer.

40.11 Atomic Extents

If the Lisp file ‘atomic-extents’ is loaded, then the atomic extent facility is available. An atomic extent is an extent for which point cannot be positioned anywhere within it. This ensures that when selecting text, either all or none of the extent is selected.

To make an extent atomic, set its atomic property.

41 Specifiers

A specifier is an object used to keep track of a property whose value may vary depending on the particular situation (e.g. particular buffer displayed in a particular window) that it is used in. The value of many built-in properties, such as the font, foreground, background, and such properties of a face and variables such as modeline-shadow-thickness and toptoolbar-height, is actually a specifier object. The specifier object, in turn, is “instanced” in a particular situation to yield the real value of the property in that situation.

Function specifierp object

This function returns non-nil if object is a specifier.

41.1 Introduction to Specifiers

Sometimes you may want the value of a property to vary depending on the context the property is used in. A simple example of this in XEmacs is buffer-local variables. For example, the variable modeline-format, which controls the format of the modeline, can have different values depending on the particular buffer being edited. The variable has a default value which most modes will use, but a specialized package such as Calendar might change the variable so as to tailor the modeline to its own purposes.

Other properties (such as those that can be changed by the modify-frame-parameters function, for example the color of the text cursor) can have frame-local values, although it might also make sense for them to have buffer-local values. In other cases, you might want the property to vary depending on the particular window within the frame that applies (e.g. the top or bottom window in a split frame), the device type that that frame appears on (X or tty), etc. Perhaps you can envision some more complicated scenario where you want a particular value in a specified buffer, another value in all other buffers displayed on a particular frame, another value in all other buffers displayed in all other frames on any mono (two-color, e.g. black and white only) displays, and a default value in all other circumstances.

A specifier is a generalization of this, allowing a great deal of flexibility in controlling exactly what value a property has in which circumstances. It is most commonly used for display properties, such as an image or the foreground color of a face. As a simple example, you can specify that the foreground of the default face be

• blue for a particular buffer

• green for all other buffers

As a more complicated example, you could specify that the foreground of the default face be

• forest green for all buffers displayed in a particular Emacs window, or green if the X server doesn’t recognize the color ‘forest green’

• blue for all buffers displayed in a particular frame

• red for all other buffers displayed on a color device

• white for all other buffers

41.2 In-Depth Overview of a Specifier

A specifier object encapsulates a set of specifications, each of which says what its value should be if a particular condition applies. For example, one specification might be “The value should be darkseagreen2 on X devices” another might be “The value should be blue in the *Help* buffer”. In specifier terminology, these conditions are called locales and the values are called instantiators. Given a specifier, a logical question is “What is its value in a particular situation?” This involves looking through the specifications to see which ones apply to this particular situation, and perhaps preferring one over another if more than one applies. In specifier terminology, a “particular situation” is called a domain, and determining its value in a particular domain is called instancing. Most of the time, a domain is identified by a particular window. For example, if the redisplay engine is drawing text in the default face in a particular window, it retrieves the specifier for the foreground color of the default face and instances it in the domain given by that window; in other words, it asks the specifier, “What is your value in this window?”.

More specifically, a specifier contains a set of specifications, each of which associates a locale (a window object, a buffer object, a frame object, a device object, or the symbol global) with an inst-list, which is a list of one or more inst-pairs. (For each possible locale, there can be at most one specification containing that locale.) Each inst-pair is a cons of a tag set (an unordered list of zero or more symbols, or tags) and an instantiator (the allowed form of this varies depending on the type of specifier). In a given specification, there may be more than one inst-pair with the same tag set; this is unlike for locales.

The tag set is used to restrict the sorts of devices over which the instantiator is valid and to uniquely identify instantiators added by a particular application, so that different applications can work on the same specifier and not interfere with each other. Each tag can have a predicate associated with it, which is a function of one argument (a device) that specifies whether the tag matches that particular device. (If a tag does not have a predicate, it matches all devices.) All tags in a tag set must match a device for the associated inst-pair to be instantiable over that device. (A null tag set is perfectly valid.)

The valid device types (normally x, tty, and stream) and device classes (normally color, grayscale, and mono) can always be used as tags, and match devices of the associated type or class (see Chapter 33 [Consoles and Devices], page 483). User-defined tags may be defined, with an optional predicate specified. An application can create its own tag, use it to mark all its instantiators, and be fairly confident that it will not interfere with other applications that modify the same specifier – Functions that add a specification to a specifier usually only overwrite existing inst-pairs with the same tag set as was given, and a particular tag or tag set can be specified when removing instantiators.

When a specifier is instanced in a domain, both the locale and the tag set can be viewed as specifying necessary conditions that must apply in that domain for an instantiator to be considered as a possible result of the instancing. More specific locales always override more general locales (thus, there is no particular ordering of the specifications in a specifier); however, the tag sets are simply considered in the order that the inst-pairs occur in the specification’s inst-list.

Note also that the actual object that results from the instancing (called an instance object) may not be the same as the instantiator from which it was derived. For some specifier types (such as integer specifiers and boolean specifiers), the instantiator will be returned directly as the instance object. For other types, however, this is not the case. For example, for font specifiers, the instantiator is a font-description string and the instance object is a font-instance object, which describes how the font is displayed on a particular device. A font-instance object encapsulates such things as the actual font name used to display the font on that device (a font-description string under X is usually a wildcard specification that may resolve to different font names, with possibly different foundries, widths, etc., on different devices), the extra properties of that font on that device, etc. Furthermore, this conversion (called instantiation) might fail – a font or color might not exist on a particular device, for example.

41.3 How a Specifier Is Instanced

Instancing of a specifier in a particular window domain proceeds as follows:

• First, XEmacs searches for a specification whose locale is the same as the window. If that fails, the search is repeated, looking for a locale that is the same as the window’s buffer. If that fails, the search is repeated using the window’s frame, then using the device that frame is on. Finally, the specification whose locale is the symbol global (if there is such a specification) is considered.

• The inst-pairs contained in the specification that was found are considered in their order in the inst-list, looking for one whose tag set matches the device that is derived from the window domain. (The tag set is an unordered list of zero or more tag symbols. For all tags that have predicates associated with them, the predicate must match the device.)

• If a matching tag set is found, the corresponding instantiator is passed to the specifier’s instantiation method, which is specific to the type of the specifier. If it succeeds, the resulting instance object is returned as the result of the instancing and the instancing is done. Otherwise, the operation continues, looking for another matching inst-pair in the current specification.

• When there are no more inst-pairs to be considered in the current specification, the search starts over, looking for another specification as in the first step above.

• If all specifications are exhausted and no instance object can be derived, the instancing fails. (Actually, this is not completely true. Some specifier objects for built-in properties have a fallback value, which is either an inst-list or another specifier object, that is consulted if the instancing is about to fail. If it is an inst-list, the searching proceeds using the inst-pairs in that list. If it is a specifier, the entire instancing starts over using that specifier instead of the given one. Fallback values are set by the C code and cannot be modified, except perhaps indirectly, using any Lisp functions. The purpose of them is to supply some values to make sure that instancing of built-in properties can’t fail and to implement some basic specifier inheritance, such as the fact that faces inherit their properties from the default face.)

It is also possible to instance a specifier over a frame domain or device domain instead of over a window domain. The C code, for example, instances the top-toolbar-height variable over a frame domain in order to determine the height of a frame’s top toolbar.

Instancing over a frame or device is similar to instancing over a window except that specifications for locales that cannot be derived from the domain are ignored. Specifically, instancing over a frame looks first for frame locales, then device locales, then the global locale. Instancing over a device domain looks only for device locales and the global locale.

41.4 Specifier Types

There are various different types of specifiers. The type of a specifier controls what sorts of instantiators are valid, how an instantiator is instantiated, etc. Here is a list of built-in specifier types:

boolean The valid instantiators are the symbols t and nil. Instance objects are the same as instantiators so no special instantiation function is needed.

integer The valid instantiators are integers. Instance objects are the same as instantiators so no special instantiation function is needed. modeline-shadow-thickness is an example of an integer specifier (negative thicknesses indicate that the shadow is drawn recessed instead of raised).

natnum The valid instantiators are natnums (non-negative integers). Instance objects are the same as instantiators so no special instantiation function is needed. Natnum specifiers are used for dimension variables such as top-toolbar-height.

generic All Lisp objects are valid instantiators. Instance objects are the same as instantiators so no special instantiation function is needed.

font The valid instantiators are strings describing fonts or vectors indicating inheritance from the font of some face. Instance objects are font-instance objects, which are specific to a particular device. The instantiation method for font specifiers can fail, unlike for integer, natnum, boolean, and generic specifiers.

color The valid instantiators are strings describing colors or vectors indicating inheritance from the foreground or background of some face. Instance objects are color-instance objects, which are specific to a particular device. The instantiation method for color specifiers can fail, as for font specifiers.

image Images are perhaps the most complicated type of built-in specifier. The valid instantiators are strings (a filename, inline data for a pixmap, or text to be displayed in a text glyph) or vectors describing inline data of various sorts or indicating inheritance from the background-pixmap property of some face. Instance objects are either strings (for text images), image-instance objects (for pixmap images), or subwindow objects (for subwindow images). The instantiation method for image specifiers can fail, as for font and color specifiers.

face-boolean

The valid instantiators are the symbols t and nil and vectors indicating inheritance from a boolean property of some face. Specifiers of this sort are used for all of the built-in boolean properties of faces. Instance objects are either the symbol t or the symbol nil. toolbar The valid instantiators are toolbar descriptors, which are lists of toolbar-button

descriptors (each of which is a vector of two or four elements). See Chapter 23 [Toolbar], page 351, for more information.

Color and font instance objects can also be used in turn as instantiators for a new color or font instance object. Since these instance objects are device-specific, the instantiator can be used directly as the new instance object, but only if they are of the same device. If the devices differ, the base color or font of the instantiating object is effectively used instead as the instantiator.

See Chapter 42 [Faces and Window-System Objects], page 615, for more information on fonts, colors, and face-boolean specifiers. See Chapter 43 [Glyphs], page 625, for more information about image specifiers. See Chapter 23 [Toolbar], page 351, for more information on toolbar specifiers.

Function specifier-type specifier

This function returns the type of specifier. The returned value will be a symbol: one of integer, boolean, etc., as listed in the above table.

Functions are also provided to query whether an object is a particular kind of specifier:

Function boolean-specifier-p object

This function returns non-nil if object is a boolean specifier.

Function integer-specifier-p object

This function returns non-nil if object is an integer specifier.

Function natnum-specifier-p object

This function returns non-nil if object is a natnum specifier.

Function generic-specifier-p object

This function returns non-nil if object is a generic specifier.

Function face-boolean-specifier-p object

This function returns non-nil if object is a face-boolean specifier.

Function toolbar-specifier-p object

This function returns non-nil if object is a toolbar specifier.

Function font-specifier-p object

This function returns non-nil if object is a font specifier.

Function color-specifier-p object

This function returns non-nil if object is a color specifier.

Function image-specifier-p object

This function returns non-nil if object is an image specifier.

41.5 Adding specifications to a Specifier

Function add-spec-to-specifier specifier instantiator &optional locale

tag-set how-to-add This function adds a specification to specifier. The specification maps from locale (which should be a window, buffer, frame, device, or the symbol global, and defaults to global) to instantiator, whose allowed values depend on the type of the specifier. Optional argument tag-set limits the instantiator to apply only to the specified tag set, which should be a list of tags all of which must match the device being instantiated over (tags are a device type, a device class, or tags defined with define-specifier-tag). Specifying a single symbol for tag-set is equivalent to specifying a one-element list containing that symbol. Optional argument how-to-add specifies what to do if there are already specifications in the specifier. It should be one of

prepend Put at the beginning of the current list of instantiators for locale.

append Add to the end of the current list of instantiators for locale.

remove-tag-set-prepend

This is the default. Remove any existing instantiators whose tag set is the same as tag-set; then put the new instantiator at the beginning of the current list.

remove-tag-set-append

Remove any existing instantiators whose tag set is the same as tag-set; then put the new instantiator at the end of the current list.

remove-locale

Remove all previous instantiators for this locale before adding the new spec.

remove-locale-type

Remove all specifications for all locales of the same type as locale (this includes locale itself) before adding the new spec.

remove-all

Remove all specifications from the specifier before adding the new spec.

remove-tag-set-prepend is the default.

You can retrieve the specifications for a particular locale or locale type with

the function specifier-spec-list or specifier-specs.

Function add-spec-list-to-specifier specifier spec-list &optional

how-to-add This function adds a spec-list (a list of specifications) to specifier. The format of a spec-list is

((locale (tag-set . instantiator) ...) ...) where • locale := a window, a buffer, a frame, a device, or global

• tag-set := an unordered list of zero or more tags, each of which is a symbol • tag := a device class (see Chapter 33 [Consoles and Devices], page 483), a

device type, or a tag defined with define-specifier-tag

• instantiator := format determined by the type of specifier

The pair (tag-set . instantiator) is called an inst-pair. A list of inst-pairs is called an inst-list. The pair (locale . inst-list) is called a specification. A spec-list, then, can be viewed as a list of specifications. how-to-add specifies how to combine the new specifications with the existing ones, and has the same semantics as for add-spec-to-specifier. In many circumstances, the higher-level function set-specifier is more convenient and should be used instead.

Macro let-specifier specifier-list &rest body This special form temporarily adds specifications to specifiers, evaluates forms in body and restores the specifiers to their previous states. The specifiers and their temporary specifications are listed in specifier-list. The format of specifier-list is

((specifier value &optional locale tag-set how-to-add) ...) specifier is the specifier to be temporarily modified. value is the instantiator to be temporarily added to specifier in locale. locale, tag-set and how-to-add have the same meaning as in add-spec-to-specifier. This special form is implemented as a macro; the code resulting from macro expansion will add specifications to specifiers using add-spec-to-specifier. After forms in body are evaluated, the temporary specifications are removed and old specifier spec-lists are restored. locale, tag-set and how-to-add may be omitted, and default to nil. The value of the last form in body is returned. NOTE: If you want the specifier’s instance to change in all circumstances, use (selected-window) as the locale. If locale is nil or omitted, it defaults to

global.

The following example removes the 3D modeline effect in the currently selected window for the duration of a second:

     (let-specifier ((modeline-shadow-thickness 0 (selected-window)))
       (sit-for 1))

Function set-specifier specifier value &optional how-to-add This function adds some specifications to specifier. value can be a single instantiator or tagged instantiator (added as a global specification), a list of tagged and/or untagged instantiators (added as a global specification), a cons of a locale and instantiator or locale and instantiator list, a list of such conses, or nearly any other reasonable form. More specifically, value can be anything

accepted by canonicalize-spec-list.

how-to-add is the same as in add-spec-to-specifier.

Note that set-specifier is exactly complementary to specifier-specs except in the case where specifier has no specs at all in it but nil is a valid instantiator (in that case, specifier-specs will return nil (meaning no specs) and set-specifier will interpret the nil as meaning “I’m adding a global instantiator and its value is nil”), or in strange cases where there is an ambiguity between a spec-list and an inst-list, etc. (The built-in specifier types are designed in such a way as to avoid any such ambiguities.) If you want to work with spec-lists, you should probably not use these functions, but should use the lower-level functions specifier-spec-list and add-spec-list-to-specifier. These functions always work with fully-qualified speclists; thus, there is no ambiguity.

Function canonicalize-inst-pair inst-pair specifier-type &optional noerror This function canonicalizes the given inst-pair. specifier-type specifies the type of specifier that this spec-list will be used for. Canonicalizing means converting to the full form for an inst-pair, i.e. (tag-set . instantiator). A single, untagged instantiator is given a tag set of nil (the empty set), and a single tag is converted into a tag set consisting only of that tag. If noerror is non-nil, signal an error if the inst-pair is invalid; otherwise return

t.

Function canonicalize-inst-list inst-list specifier-type &optional noerror This function canonicalizes the given inst-list (a list of inst-pairs). specifier-type specifies the type of specifier that this inst-list will be used for. Canonicalizing means converting to the full form for an inst-list, i.e. ((tag-set . instantiator) ...). This function accepts a single inst-pair or any abbreviation thereof or a list of (possibly abbreviated) inst-pairs. (See canonicalize-inst-

pair.)

If noerror is non-nil, signal an error if the inst-list is invalid; otherwise return

t.

Function canonicalize-spec spec specifier-type &optional noerror This function canonicalizes the given spec (a specification). specifier-type specifies the type of specifier that this spec-list will be used for. Canonicalizing means converting to the full form for a spec, i.e. (locale (tagset . instantiator) ...). This function accepts a possibly abbreviated inst-list or a cons of a locale and a possibly abbreviated inst-list. (See canonicalize-

inst-list.)

If noerror is nil, signal an error if the specification is invalid; otherwise return

t.

Function canonicalize-spec-list spec-list specifier-type &optional noerror This function canonicalizes the given spec-list (a list of specifications). specifier-type specifies the type of specifier that this spec-list will be used for.

Canonicalizing means converting to the full form for a spec-list, i.e. ((locale (tag-set . instantiator) ...) ...). This function accepts a possibly abbreviated specification or a list of such things. (See canonicalize-spec.) This is the function used to convert spec-lists accepted by set-specifier and such

into a form suitable for add-spec-list-to-specifier.

This function tries extremely hard to resolve any ambiguities, and the built-in specifier types (font, image, toolbar, etc.) are designed so that there won’t be any ambiguities. If noerror is nil, signal an error if the spec-list is invalid; otherwise return t.

41.6 Retrieving the Specifications from a Specifier

Function specifier-spec-list specifier &optional locale tag-set exact-p This function returns the spec-list of specifications for specifier in locale. If locale is a particular locale (a window, buffer, frame, device, or the symbol global), a spec-list consisting of the specification for that locale will be returned. If locale is a locale type (i.e. a symbol window, buffer, frame, or device), a spec-list of the specifications for all locales of that type will be returned. If locale is nil or the symbol all, a spec-list of all specifications in specifier will be returned. locale can also be a list of locales, locale types, and/or all; the result is as if specifier-spec-list were called on each element of the list and the results concatenated together. Only instantiators where tag-set (a list of zero or more tags) is a subset of (or possibly equal to) the instantiator’s tag set are returned. (The default value of nil is a subset of all tag sets, so in this case no instantiators will be screened out.) If exact-p is non-nil, however, tag-set must be equal to an instantiator’s tag set for the instantiator to be returned.

Function specifier-specs specifier &optional locale tag-set exact-p This function returns the specification(s) for specifier in locale. If locale is a single locale or is a list of one element containing a single locale, then a “short form” of the instantiators for that locale will be returned. Otherwise, this function is identical to specifier-spec-list. The “short form” is designed for readability and not for ease of use in Lisp programs, and is as follows:

1. If there is only one instantiator, then an inst-pair (i.e. cons of tag and instantiator) will be returned; otherwise a list of inst-pairs will be returned.

2. For each inst-pair returned, if the instantiator’s tag is any, the tag will be removed and the instantiator itself will be returned instead of the inst-pair.

3. If there is only one instantiator, its value is nil, and its tag is any, a one-element list containing nil will be returned rather than just nil, to distinguish this case from there being no instantiators at all.

Function specifier-fallback specifier

This function returns the fallback value for specifier. Fallback values are provided by the C code for certain built-in specifiers to make sure that instancing won’t fail even if all specs are removed from the specifier, or to implement simple inheritance behavior (e.g. this method is used to ensure that faces other than default inherit their attributes from default). By design, you cannot change the fallback value, and specifiers created with make-specifier will never have a fallback (although a similar, Lisp-accessible capability may be provided in the future to allow for inheritance). The fallback value will be an inst-list that is instanced like any other inst-list, a specifier of the same type as specifier (results in inheritance), or nil for no fallback. When you instance a specifier, you can explicitly request that the fallback not be consulted. (The C code does this, for example, when merging faces.) See

specifier-instance.

41.7 Working With Specifier Tags

A specifier tag set is an entity that is attached to an instantiator and can be used to restrict the scope of that instantiator to a particular device class or device type and/or to mark instantiators added by a particular package so that they can be later removed.

A specifier tag set consists of a list of zero of more specifier tags, each of which is a symbol that is recognized by XEmacs as a tag. (The valid device types and device classes are always tags, as are any tags defined by define-specifier-tag.) It is called a “tag set” (as opposed to a list) because the order of the tags or the number of times a particular tag occurs does not matter.

Each tag has a predicate associated with it, which specifies whether that tag applies to a particular device. The tags which are device types and classes match devices of that type or class. User-defined tags can have any predicate, or none (meaning that all devices match). When attempting to instance a specifier, a particular instantiator is only considered if the device of the domain being instanced over matches all tags in the tag set attached to that instantiator.

Most of the time, a tag set is not specified, and the instantiator gets a null tag set, which matches all devices.

Function valid-specifier-tag-p tag

This function returns non-nil if tag is a valid specifier tag.

Function valid-specifier-tag-set-p tag-set

This function returns non-nil if tag-set is a valid specifier tag set.

Function canonicalize-tag-set tag-set

This function canonicalizes the given tag set. Two canonicalized tag sets can be compared with equal to see if they represent the same tag set. (Specifically, canonicalizing involves sorting by symbol name and removing duplicates.)

Function device-matches-specifier-tag-set-p device tag-set

This function returns non-nil if device matches specifier tag set tag-set. This means that device matches each tag in the tag set.

Function define-specifier-tag tag &optional predicate This function defines a new specifier tag. If predicate is specified, it should be a function of one argument (a device) that specifies whether the tag matches that particular device. If predicate is omitted, the tag matches all devices. You can redefine an existing user-defined specifier tag. However, you cannot redefine the built-in specifier tags (the device types and classes) or the symbols

nil, t, all, or global.

Function device-matching-specifier-tag-list &optional device

This function returns a list of all specifier tags matching device. device defaults to the selected device if omitted.

Function specifier-tag-list

This function returns a list of all currently-defined specifier tags. This includes the built-in ones (the device types and classes).

Function specifier-tag-predicate tag

This function returns the predicate for the given specifier tag.

41.8 Functions for Instancing a Specifier

Function specifier-instance specifier &optional domain default no-fallback This function instantiates specifier (return its value) in domain. If no instance can be generated for this domain, return default. domain should be a window, frame, or device. Other values that are legal as a locale (e.g. a buffer) are not valid as a domain because they do not provide enough information to identify a particular device (see valid-specifier-domain-p). domain defaults to the selected window if omitted. Instantiating a specifier in a particular domain means determining the specifier’s “value” in that domain. This is accomplished by searching through the specifications in the specifier that correspond to all locales that can be derived from the given domain, from specific to general. In most cases, the domain is an Emacs window. In that case specifications are searched for as follows:

1. A specification whose locale is the window itself;

2. A specification whose locale is the window’s buffer;

3. A specification whose locale is the window’s frame;

4. A specification whose locale is the window’s frame’s device;

5. A specification whose locale is the symbol global.

If all of those fail, then the C-code-provided fallback value for this specifier is consulted (see specifier-fallback). If it is an inst-list, then this function attempts to instantiate that list just as when a specification is located in the first five steps above. If the fallback is a specifier, specifier-instance is called recursively on this specifier and the return value used. Note, however, that if the optional argument no-fallback is non-nil, the fallback value will not be consulted. Note that there may be more than one specification matching a particular locale; all such specifications are considered before looking for any specifications for more general locales. Any particular specification that is found may be rejected because it is tagged to a particular device class (e.g. color) or device type (e.g.

x) or both and the device for the given domain does not match this, or because the specification is not valid for the device of the given domain (e.g. the font or color name does not exist for this particular X server). The returned value is dependent on the type of specifier. For example, for a font specifier (as returned by the face-font function), the returned value will be a font-instance object. For images, the returned value will be a string, pixmap, or subwindow.

Function specifier-instance-from-inst-list specifier domain inst-list

&optional default This function attempts to convert a particular inst-list into an instance. This attempts to instantiate inst-list in the given domain, as if inst-list existed in a specification in specifier. If the instantiation fails, default is returned. In most circumstances, you should not use this function; use specifier-instance instead.

41.9 Example of Specifier Usage

Now let us present an example to clarify the theoretical discussions we have been through. In this example, we will use the general specifier functions for clarity. Keep in mind that many types of specifiers, and some other types of objects that are associated with specifiers (e.g. faces), provide convenience functions making it easier to work with objects of that type.

Let us consider the background color of the default face. A specifier is used to specify how that color will appear in different domains. First, let’s retrieve the specifier:

(setq sp (face-property 'default 'background))
    ⇒ #<color-specifier 0x3da>
(specifier-specs sp)
    ⇒ ((#<buffer "device.c"> (nil . "forest green"))
                 (#<window on "Makefile" 0x8a2b> (nil . "hot pink"))
                 (#<x-frame "emacs" 0x4ac> (nil . "puke orange")
   (nil . "moccasin"))
                 (#<x-frame "VM" 0x4ac> (nil . "magenta"))
 (global ((tty) . "cyan") (nil . "white"))
                )

Then, say we want to determine what the background color of the default face is for the window currently displaying the buffer ‘*scratch*’. We call

(get-buffer-window "*scratch*")
    ⇒ #<window on "*scratch*" 0x4ad>
(window-frame (get-buffer-window "*scratch*"))
    ⇒ #<x-frame "emacs" 0x4ac>
(specifier-instance sp (get-buffer-window "*scratch*"))
    ⇒ #<color-instance moccasin 47=(FFFF,E4E4,B5B5) 0x6309>

Note that we passed a window to specifier-instance, not a buffer. We cannot pass a buffer because a buffer by itself does not provide enough information. The buffer might not be displayed anywhere at all, or could be displayed in many different frames on different devices.

The result is arrived at like this:

1. First, we look for a specification matching the buffer displayed in the window, i.e. ‘*scratch’. There are none, so we proceed.

2. Then, we look for a specification matching the window itself. Again, there are none.

3. Then, we look for a specification matching the window’s frame. The specification (#<xframe "emacs" 0x4ac> . "puke orange") is found. We call the instantiation method for colors, passing it the locale we were searching over (i.e. the window, in this case) and the instantiator (‘"puke orange"’). However, the particular device which this window is on (let’s say it’s an X connection) doesn’t recognize the color ‘"puke orange"’, so the specification is rejected.

4. So we continue looking for a specification matching the window’s frame. We find ‘(#<x-frame "emacs" 0x4ac> . "moccasin")’. Again, we call the instantiation method for colors. This time, the X server our window is on recognizes the color ‘moccasin’, and so the instantiation method succeeds and returns a color instance.

41.10 Creating New Specifier Objects

Function make-specifier type

This function creates a new specifier. A specifier is an object that can be used to keep track of a property whose value can be per-buffer, per-window, per-frame, or per-device, and can further be restricted to a particular device-type or device-class. Specifiers are used, for example, for the various built-in properties of a face; this allows a face to have different values in different frames, buffers, etc. For more information, see ‘specifier-instance’, ‘specifier-specs’, and ‘add-spec-to-specifier’; or, for a detailed description of specifiers, including how they are instantiated over a particular domain (i.e. how their value in that domain is determined), see the chapter on specifiers in the XEmacs Lisp Reference Manual. type specifies the particular type of specifier, and should be one of the symbols

generic, integer, natnum, boolean, color, font, image, face-boolean, or
toolbar.

For more information on particular types of specifiers, see the functions

generic-specifier-p, integer-specifier-p, natnum-specifier-p, boolean-
specifier-p, color-specifier-p, font-specifier-p, image-specifier-p,
face-boolean-specifier-p, and toolbar-specifier-p.

Function make-specifier-and-init type spec-list &optional

dont-canonicalize This function creates and initialize a new specifier. This is a front-end onto make-specifier that allows you to create a specifier and add specs to it at the same time. type specifies the specifier type. spec-list supplies the specification(s) to be added to the specifier. Normally, almost any reasonable abbreviation of the full spec-list form is accepted, and is converted to the full form; however, if optional argument dont-canonicalize is non-nil, this conversion is not performed, and the spec-list must already be in full form.

See canonicalize-spec-list.

41.11 Functions for Checking the Validity of Specifier Components

Function valid-specifier-domain-p domain

This function returns non-nil if domain is a valid specifier domain. A domain is used to instance a specifier (i.e. determine the specifier’s value in that domain). Valid domains are a window, frame, or device. (nil is not valid.)

Function valid-specifier-locale-p locale

This function returns non-nil if locale is a valid specifier locale. Valid locales are a device, a frame, a window, a buffer, and global. (nil is not valid.)

Function valid-specifier-locale-type-p locale-type

Given a specifier locale-type, this function returns non-nil if it is valid. Valid locale types are the symbols global, device, frame, window, and buffer. (Note, however, that in functions that accept either a locale or a locale type, global is considered an individual locale.)

Function valid-specifier-type-p specifier-type

Given a specifier-type, this function returns non-nil if it is valid. Valid types are

generic, integer, boolean, color, font, image, face-boolean, and toolbar.

Function valid-specifier-tag-p tag

This function returns non-nil if tag is a valid specifier tag.

Function valid-instantiator-p instantiator specifier-type This function returns non-nil if instantiator is valid for specifier-type.

Function valid-inst-list-p inst-list type

This function returns non-nil if inst-list is valid for specifier type type.

Function valid-spec-list-p spec-list type

This function returns non-nil if spec-list is valid for specifier type type.

Function check-valid-instantiator instantiator specifier-type This function signals an error if instantiator is invalid for specifier-type.

Function check-valid-inst-list inst-list type

This function signals an error if inst-list is invalid for specifier type type.

Function check-valid-spec-list spec-list type

This function signals an error if spec-list is invalid for specifier type type.

41.12 Other Functions for Working with Specifications in a Specifier

Function copy-specifier specifier &optional dest locale tag-set exact-p

how-to-add This function copies specifier to dest, or creates a new one if dest is nil. If dest is nil or omitted, a new specifier will be created and the specifications copied into it. Otherwise, the specifications will be copied into the existing specifier in dest. If locale is nil or the symbol all, all specifications will be copied. If locale is a particular locale, the specification for that particular locale will be copied. If locale is a locale type, the specifications for all locales of that type will be copied. locale can also be a list of locales, locale types, and/or all; this is equivalent to calling copy-specifier for each of the elements of the list. See specifier-spec-list for more information about locale. Only instantiators where tag-set (a list of zero or more tags) is a subset of (or possibly equal to) the instantiator’s tag set are copied. (The default value of nil is a subset of all tag sets, so in this case no instantiators will be screened out.) If exact-p is non-nil, however, tag-set must be equal to an instantiator’s tag set for the instantiator to be copied. Optional argument how-to-add specifies what to do with existing specifications in dest. If nil, then whichever locales or locale types are copied will first be completely erased in dest. Otherwise, it is the same as in add-spec-to-specifier.

Function remove-specifier specifier &optional locale tag-set exact-p This function removes specification(s) for specifier. If locale is a particular locale (a buffer, window, frame, device, or the symbol global), the specification for that locale will be removed. If instead, locale is a locale type (i.e. a symbol buffer, window, frame, or device), the specifications for all locales of that type will be removed. If locale is nil or the symbol all, all specifications will be removed. locale can also be a list of locales, locale types, and/or all; this is equivalent to calling remove-specifier for each of the elements in the list. Only instantiators where tag-set (a list of zero or more tags) is a subset of (or possibly equal to) the instantiator’s tag set are removed. (The default value of nil is a subset of all tag sets, so in this case no instantiators will be screened out.) If exact-p is non-nil, however, tag-set must be equal to an instantiator’s tag set for the instantiator to be removed.

Function map-specifier specifier func &optional locale maparg This function applies func to the specification(s) for locale in specifier. If locale is a locale, func will be called for that locale. If locale is a locale type, func will be mapped over all locales of that type. If locale is nil or the symbol all, func will be mapped over all locales in specifier. func is called with four arguments: the specifier, the locale being mapped over, the inst-list for that locale, and the optional maparg. If any invocation of func returns non-nil, the mapping will stop and the returned value becomes the value returned from map-specifier. Otherwise, map-specifier returns nil.

Function specifier-locale-type-from-locale locale

Given a specifier locale, this function returns its type.

42 Faces and Window-System Objects

42.1 Faces

A face is a named collection of graphical properties: font, foreground color, background color, background pixmap, optional underlining, and (on TTY devices) whether the text is to be highlighted, dimmed, blinking, or displayed in reverse video. Faces control the display of text on the screen. Every face has a name, which is a symbol such as default

or modeline.

Each built-in property of a face is controlled using a specifier, which allows it to have separate values in particular buffers, frames, windows, and devices and to further vary according to device type (X or TTY) and device class (color, mono, or grayscale). See Chapter 41 [Specifiers], page 599, for more information.

The face named default is used for ordinary text. The face named modeline is used for displaying the modeline. The face named highlight is used for highlighted extents (see Chapter 40 [Extents], page 585). The faces named left-margin and right-margin are used for the left and right margin areas, respectively (see Chapter 44 [Annotations], page 641). The face named zmacs-region is used for the highlighted region between point and mark.

42.1.1 Merging Faces for Display

Here are all the ways to specify which face to use for display of text:

• With defaults. Each frame has a default face, which is used for all text that doesn’t somehow specify another face. The face named default applies to the text area, while the faces left-margin and right-margin apply to the left and right margin areas.

• With text properties. A character may have a face property; if so, it’s displayed with that face. (Text properties are actually implemented in terms of extents.) See Section 36.18 [Text Properties], page 539.

• With extents. An extent may have a face property, which applies to all the text covered by the extent; in addition, if the highlight property is set, the highlight property applies when the mouse moves over the extent or if the extent is explicitly highlighted. See Chapter 40 [Extents], page 585.

• With annotations. Annotations that are inserted into a buffer can specify their own face. (Annotations are actually implemented in terms of extents.) See Chapter 44 [Annotations], page 641.

If these various sources together specify more than one face for a particular character, XEmacs merges the properties of the various faces specified. Extents, text properties, and annotations all use the same underlying representation (as extents). When multiple extents cover one character, an extent with higher priority overrides those with lower priority. See Chapter 40 [Extents], page 585. If no extent covers a particular character, the default face is used.

If a background pixmap is specified, it determines what will be displayed in the background of text characters. If the background pixmap is actually a pixmap, with its colors specified, those colors are used; if it is a bitmap, the face’s foreground and background colors are used to color it.

42.1.2 Basic Functions for Working with Faces

The properties a face can specify include the font, the foreground color, the background color, the background pixmap, the underlining, the display table, and (for TTY devices) whether the text is to be highlighted, dimmed, blinking, or displayed in reverse video. The face can also leave these unspecified, causing them to assume the value of the corresponding property of the default face.

Here are the basic primitives for working with faces.

Function make-face name &optional doc-string temporary This function defines and returns a new face named name, initially with all properties unspecified. It does nothing if there is already a face named name. Optional argument doc-string specifies an explanatory string used for descriptive purposes. If optional argument temporary is non-nil, the face will automatically disappear when there are no more references to it anywhere in text or Lisp code (otherwise, the face will continue to exist indefinitely even if it is not used).

Function face-list &optional temporary This function returns a list of the names of all defined faces. If temporary is nil, only the permanent faces are included. If it is t, only the temporary faces are included. If it is any other non-nil value both permanent and temporary are included.

Function facep object This function returns whether the given object is a face.

Function copy-face old-face new-name &optional locale how-to-add This function defines a new face named new-name which is a copy of the existing face named old-face. If there is already a face named new-name, then it alters the face to have the same properties as old-face. locale and how-to-add let you copy just parts of the old face rather than the whole face, and are as in copy-specifier (see Chapter 41 [Specifiers], page 599).

42.1.3 Face Properties

You can examine and modify the properties of an existing face with the following functions.

The following symbols have predefined meanings:

foreground

The foreground color of the face.

background

The background color of the face. font The font used to display text covered by this face.

display-table

The display table of the face.

background-pixmap

The pixmap displayed in the background of the face. Only used by faces on X devices.

underline

Underline all text covered by this face.

highlight

Highlight all text covered by this face. Only used by faces on TTY devices.

dim Dim all text covered by this face. Only used by faces on TTY devices.

blinking Blink all text covered by this face. Only used by faces on TTY devices.

reverse Reverse the foreground and background colors. Only used by faces on TTY devices.

doc-string

Description of what the face’s normal use is. NOTE: This is not a specifier, unlike all the other built-in properties, and cannot contain locale-specific values.

Function set-face-property face property value &optional locale tag

how-to-add This function changes a property of a face. For built-in properties, the actual value of the property is a specifier and you cannot change this; but you can change the specifications within the specifier, and that is what this function will do. For user-defined properties, you can use this function to either change the actual value of the property or, if this value is a specifier, change the specifications within it. If property is a built-in property, the specifications to be added to this property can be supplied in many different ways:

If value is a simple instantiator (e.g. a string naming a font or color) or a list of instantiators, then the instantiator(s) will be added as a specification of the property for the given locale (which defaults to global if omitted). If value is a list of specifications (each of which is a cons of a locale and a list of instantiators), then locale must be nil (it does not make sense to explicitly specify a locale in this case), and specifications will be added as given. If value is a specifier (as would be returned by face-property if no locale argument is given), then some or all of the specifications in the specifier will be added to the property. In this case, the function is really equivalent to copy-specifier and locale has the same semantics (if it is a particular locale, the specification for the locale will be copied; if a locale type, specifications for all locales of that type will be copied; if nil or all, then all specifications will be copied). how-to-add should be either nil or one of the symbols prepend, append,

remove-tag-set-prepend, remove-tag-set-append, remove-locale, remove-
locale-type, or remove-all. See copy-specifier and add-spec-to-specifier

for a description of what each of these means. Most of the time, you do not need to worry about this argument; the default behavior usually is fine. In general, it is OK to pass an instance object (e.g. as returned by faceproperty-instance) as an instantiator in place of an actual instantiator. In such a case, the instantiator used to create that instance object will be used (for example, if you set a font-instance object as the value of the font property, then the font name used to create that object will be used instead). If some cases, however, doing this conversion does not make sense, and this will be noted in the documentation for particular types of instance objects. If property is not a built-in property, then this function will simply set its value if locale is nil. However, if locale is given, then this function will attempt to add value as the instantiator for the given locale, using add-spec-to-specifier. If the value of the property is not a specifier, it will automatically be converted into a generic specifier.

Function face-property face property &optional locale This function returns face’s value of the given property. If locale is omitted, the face’s actual value for property will be returned. For built-in properties, this will be a specifier object of a type appropriate to the property (e.g. a font or color specifier). For other properties, this could be anything. If locale is supplied, then instead of returning the actual value, the specification(s) for the given locale or locale type will be returned. This will only work if the actual value of property is a specifier (this will always be the case for built-in properties, but not or not may apply to user-defined properties). If the actual value of property is not a specifier, this value will simply be returned regardless of locale. The return value will be a list of instantiators (e.g. strings specifying a font or color name), or a list of specifications, each of which is a cons of a locale and a list of instantiators. Specifically, if locale is a particular locale (a buffer, window, frame, device, or global), a list of instantiators for that locale will be returned. Otherwise, if locale is a locale type (one of the symbols buffer, window, frame, or device), the specifications for all locales of that type will be returned. Finally, if locale is all, the specifications for all locales of all types will be returned. The specifications in a specifier determine what the value of property will be in a particular domain or set of circumstances, which is typically a particular Emacs window along with the buffer it contains and the frame and device it lies within. The value is derived from the instantiator associated with the most specific locale (in the order buffer, window, frame, device, and global) that matches the domain in question. In other words, given a domain (i.e. an Emacs window, usually), the specifier for property will first be searched for a specification whose locale is the buffer contained within that window; then for a specification whose locale is the window itself; then for a specification whose locale is the frame that the window is contained within; etc. The first instantiator that is valid for the domain (usually this means that the instantiator is recognized by the device [i.e. the X server or TTY device] that the domain is on). The function face-property-instance actually does all this, and is used to determine how to display the face.

Function face-property-instance face property &optional domain default

no-fallback This function returns the instance of face’s property in the specified domain.

Under most circumstances, domain will be a particular window, and the returned instance describes how the specified property actually is displayed for that window and the particular buffer in it. Note that this may not be the same as how the property appears when the buffer is displayed in a different window or frame, or how the property appears in the same window if you switch to another buffer in that window; and in those cases, the returned instance would be different.

The returned instance will typically be a color-instance, font-instance, or pixmap-instance object, and you can query it using the appropriate objectspecific functions. For example, you could use color-instance-rgb-components to find out the RGB (red, green, and blue) components of how the background property of the highlight face is displayed in a particular window. The results might be different from the results you would get for another window (perhaps the user specified a different color for the frame that window is on; or perhaps the same color was specified but the window is on a different X server, and that X server has different RGB values for the color from this one).

domain defaults to the selected window if omitted.

domain can be a frame or device, instead of a window. The value returned for a such a domain is used in special circumstances when a more specific domain does not apply; for example, a frame value might be used for coloring a toolbar, which is conceptually attached to a frame rather than a particular window. The value is also useful in determining what the value would be for a particular window within the frame or device, if it is not overridden by a more specific specification.

If property does not name a built-in property, its value will simply be returned unless it is a specifier object, in which case it will be instanced using specifier-

instance.

Optional arguments default and no-fallback are the same as in specifierinstance. See Chapter 41 [Specifiers], page 599.

42.1.4 Face Convenience Functions

Function set-face-foreground face color &optional locale tag how-to-add Function set-face-background face color &optional locale tag how-to-add

These functions set the foreground (respectively, background) color of face face to color. The argument color should be a string (the name of a color) or a color object as returned by make-color (see Section 42.3 [Colors], page 623).

Function set-face-background-pixmap face pixmap &optional locale tag

how-to-add This function sets the background pixmap of face face to pixmap. The argument pixmap should be a string (the name of a bitmap or pixmap file; the directories listed in the variable x-bitmap-file-path will be searched) or a glyph object as returned by make-glyph (see Chapter 43 [Glyphs], page 625). The argument may also be a list of the form (width height data) where width and height are the size in pixels, and data is a string, containing the raw bits of the bitmap.

Function set-face-font face font &optional locale tag how-to-add This function sets the font of face face. The argument font should be a string or a font object as returned by make-font (see Section 42.2 [Fonts], page 621).

Function set-face-underline-p face underline-p &optional locale tag

how-to-add This function sets the underline property of face face.

Function face-foreground face &optional locale Function face-background face &optional locale

These functions return the foreground (respectively, background) color specifier of face face. See Section 42.3 [Colors], page 623.

Function face-background-pixmap face &optional locale

This function return the background-pixmap glyph object of face face.

Function face-font face &optional locale This function returns the font specifier of face face. (Note: This is not the same as the function face-font in FSF Emacs.) See Section 42.2 [Fonts], page 621.

Function face-font-name face &optional domain This function returns the name of the font of face face, or nil if it is unspecified. This is basically equivalent to (font-name (face-font face) domain) except that it does not cause an error if face’s font is nil. (This function is named face-font in FSF Emacs.)

Function face-underline-p face &optional locale This function returns the underline property of face face.

Function face-foreground-instance face &optional domain Function face-background-instance face &optional domain

These functions return the foreground (respectively, background) color specifier of face face. See Section 42.3 [Colors], page 623.

Function face-background-pixmap-instance face &optional domain

This function return the background-pixmap glyph object of face face.

Function face-font-instance face &optional domain This function returns the font specifier of face face. See Section 42.2 [Fonts], page 621.

42.1.5 Other Face Display Functions

Function invert-face face &optional locale Swap the foreground and background colors of face face. If the face doesn’t specify both foreground and background, then its foreground and background are set to the default background and foreground.

Function face-equal face1 face2 &optional domain This returns t if the faces face1 and face2 will display in the same way. domain

is as in face-property-instance.

Function face-differs-from-default-p face &optional domain

This returns t if the face face displays differently from the default face. domain

is as in face-property-instance.

42.2 Fonts

This section describes how to work with font specifier and font instance objects, which encapsulate fonts in the window system.

42.2.1 Font Specifiers

Function font-specifier-p object

This predicate returns t if object is a font specifier, and nil otherwise.

42.2.2 Font Instances

Function font-instance-p object

This predicate returns t if object is a font instance, and nil otherwise.

Function make-font-instance name &optional device noerror This function creates a new font-instance object of the specified name. device specifies the device this object applies to and defaults to the selected device. An error is signalled if the font is unknown or cannot be allocated; however, if noerror is non-nil, nil is simply returned in this case. The returned object is a normal, first-class lisp object. The way you “deallocate” the font is the way you deallocate any other lisp object: you drop all pointers to it and allow it to be garbage collected. When these objects are GCed, the underlying X data is deallocated as well.

42.2.3 Font Instance Names

Function list-fonts pattern &optional device This function returns a list of font names matching the given pattern. device specifies which device to search for names, and defaults to the currently selected device.

Function font-instance-name font-instance

This function returns the name used to allocate font-instance.

Function font-instance-truename font-instance

This function returns the canonical name of the given font instance. Font names are patterns which may match any number of fonts, of which the first found is used. This returns an unambiguous name for that font (but not necessarily its only unambiguous name).

42.2.4 Font Instance Size

Function x-font-size font

This function returns the nominal size of the given font. This is done by parsing its name, so it’s likely to lose. X fonts can be specified (by the user) in either pixels or 10ths of points, and this returns the first one it finds, so you have to decide which units the returned value is measured in yourself ...

Function x-find-larger-font font &optional device This function loads a new, slightly larger version of the given font (or font name). Returns the font if it succeeds, nil otherwise. If scalable fonts are available, this returns a font which is 1 point larger. Otherwise, it returns the next larger version of this font that is defined.

Function x-find-smaller-font font &optional device This function loads a new, slightly smaller version of the given font (or font name). Returns the font if it succeeds, nil otherwise. If scalable fonts are available, this returns a font which is 1 point smaller. Otherwise, it returns the next smaller version of this font that is defined.

42.2.5 Font Instance Characteristics

Function font-instance-properties font

This function returns the properties (an alist or nil) of font-instance.

Function x-make-font-bold font &optional device Given an X font specification, this attempts to make a “bold” font. If it fails, it returns nil.

Function x-make-font-unbold font &optional device Given an X font specification, this attempts to make a non-bold font. If it fails, it returns nil.

Function x-make-font-italic font &optional device Given an X font specification, this attempts to make an “italic” font. If it fails, it returns nil.

Function x-make-font-unitalic font &optional device

Given an X font specification, this attempts to make a non-italic font. If it fails, it returns nil.

Function x-make-font-bold-italic font &optional device

Given an X font specification, this attempts to make a “bold-italic” font. If it fails, it returns nil.

42.2.6 Font Convenience Functions

Function font-name font &optional domain This function returns the name of the font in the specified domain, if any. font should be a font specifier object and domain is normally a window and defaults to the selected window if omitted. This is equivalent to using specifier-

instance and applying font-instance-name to the result.

Function font-truename font &optional domain This function returns the truename of the font in the specified domain, if any. font should be a font specifier object and domain is normally a window and defaults to the selected window if omitted. This is equivalent to using specifier-

instance and applying font-instance-truename to the result.

Function font-properties font &optional domain This function returns the properties of the font in the specified domain, if any. font should be a font specifier object and domain is normally a window and defaults to the selected window if omitted. This is equivalent to using

specifier-instance and applying font-instance-properties to the result.

42.3 Colors

42.3.1 Color Specifiers

Function color-specifier-p object

This function returns non-nil if object is a color specifier.

42.3.2 Color Instances

A color-instance object is an object describing the way a color specifier is instanced in a particular domain. Functions such as face-background-instance return a color-instance object. For example,

(face-background-instance 'default (next-window))
    ⇒ #<color-instance moccasin 47=(FFFF,E4E4,B5B5) 0x678d>

The color-instance object returned describes the way the background color of the default face is displayed in the next window after the selected one.

Function color-instance-p object

This function returns non-nil if object is a color-instance.

42.3.3 Color Instance Properties

Function color-instance-name color-instance

This function returns the name used to allocate color-instance.

Function color-instance-rgb-components color-instance

This function returns a three element list containing the red, green, and blue color components of color-instance.

     (color-instance-rgb-components
       (face-background-instance 'default (next-window)))
         ⇒ (65535 58596 46517)

42.3.4 Color Convenience Functions

Function color-name color &optional domain This function returns the name of the color in the specified domain, if any. color should be a color specifier object and domain is normally a window and defaults to the selected window if omitted. This is equivalent to using specifier-

instance and applying color-instance-name to the result.

Function color-rgb-components color &optional domain

This function returns the RGB components of the color in the specified domain, if any. color should be a color specifier object and domain is normally a window and defaults to the selected window if omitted. This is equivalent to using

specifier-instance and applying color-instance-rgb-components to the

result.

     (color-rgb-components (face-background 'default (next-window)))
         ⇒ (65535 58596 46517)

43 Glyphs

A glyph is an object that is used for pixmaps and images of all sorts, as well as for things that “act” like pixmaps, such as non-textual strings (annotations) displayed in a buffer or in the margins. It is used in begin-glyphs and end-glyphs attached to extents, marginal and textual annotations, overlay arrows (overlay-arrow-* variables), toolbar buttons, mouse pointers, frame icons, truncation and continuation markers, and the like. (Basically, any place there is an image or something that acts like an image, there will be a glyph object representing it.)

The actual image that is displayed (as opposed to its position or clipping) is defined by an image specifier object contained within the glyph. The separation between an image specifier object and a glyph object is made because the glyph includes other properties than just the actual image: e.g. the face it is displayed in (for text images), the alignment of the image (when it is in a buffer), etc.

Function glyphp object

This function returns t if object is a glyph.

43.1 Glyph Functions

43.1.1 Creating Glyphs

Function make-glyph &optional spec-list type This function creates a new glyph object of type type. spec-list is used to initialize the glyph’s image. It is typically an image instantiator (a string or a vector; Section 43.2.1 [Image Specifiers], page 630), but can also be a list of such instantiators (each one in turn is tried until an image is successfully produced), a cons of a locale (frame, buffer, etc.) and an instantiator, a list of such conses, or any other form accepted by canonicalize-spec-list. See Chapter 41 [Specifiers], page 599, for more information about specifiers. type specifies the type of the glyph, which specifies in which contexts the glyph can be used, and controls the allowable image types into which the glyph’s image can be instantiated. type should be one of buffer (used for glyphs in an extent, the modeline, the toolbar, or elsewhere in a buffer), pointer (used for the mouse-pointer), or icon (used for a frame’s icon), and defaults to buffer. See Section 43.3 [Glyph Types], page 638.

Function make-glyph-internal &optional type

This function creates a new, uninitialized glyph of type type.

Function make-pointer-glyph &optional spec-list

This function is equivalent to calling make-glyph with a type of pointer.

Function make-icon-glyph &optional spec-list This function is equivalent to calling make-glyph with a type of icon.

43.1.2 Glyph Properties

Each glyph has a list of properties, which control all of the aspects of the glyph’s appearance. The following symbols have predefined meanings:

image The image used to display the glyph.

baseline Percent above baseline that glyph is to be displayed. Only for glyphs displayed inside of a buffer.

contrib-p

Whether the glyph contributes to the height of the line it’s on. Only for glyphs displayed inside of a buffer.

face Face of this glyph (not a specifier).

Function set-glyph-property glyph property value &optional locale

tag-set how-to-add This function changes a property of a glyph. For built-in properties, the actual value of the property is a specifier and you cannot change this; but you can change the specifications within the specifier, and that is what this function will do. For user-defined properties, you can use this function to either change the actual value of the property or, if this value is a specifier, change the specifications within it. If property is a built-in property, the specifications to be added to this property can be supplied in many different ways:

• If value is a simple instantiator (e.g. a string naming a pixmap filename) or a list of instantiators, then the instantiator(s) will be added as a specification of the property for the given locale (which defaults to global if omitted).

• If value is a list of specifications (each of which is a cons of a locale and a list of instantiators), then locale must be nil (it does not make sense to explicitly specify a locale in this case), and specifications will be added as given.

• If value is a specifier (as would be returned by glyph-property if no locale argument is given), then some or all of the specifications in the specifier will be added to the property. In this case, the function is really equivalent to copy-specifier and locale has the same semantics (if it is a particular locale, the specification for the locale will be copied; if a locale type, specifications for all locales of that type will be copied; if nil or all, then all specifications will be copied).

how-to-add should be either nil or one of the symbols prepend, append,

remove-tag-set-prepend, remove-tag-set-append, remove-locale, remove-
locale-type, or remove-all. See copy-specifier and add-spec-to-specifier

for a description of what each of these means. Most of the time, you do not need to worry about this argument; the default behavior usually is fine. In general, it is OK to pass an instance object (e.g. as returned by glyphproperty-instance) as an instantiator in place of an actual instantiator. In such a case, the instantiator used to create that instance object will be used (for example, if you set a font-instance object as the value of the font property, then the font name used to create that object will be used instead). If some cases, however, doing this conversion does not make sense, and this will be noted in the documentation for particular types of instance objects.

If property is not a built-in property, then this function will simply set its value if locale is nil. However, if locale is given, then this function will attempt to add value as the instantiator for the given locale, using add-spec-to-specifier. If the value of the property is not a specifier, it will automatically be converted into a generic specifier.

Function glyph-property glyph property &optional locale This function returns glyph’s value of the given property.

If locale is omitted, the glyph’s actual value for property will be returned. For built-in properties, this will be a specifier object of a type appropriate to the property (e.g. a font or color specifier). For other properties, this could be anything.

If locale is supplied, then instead of returning the actual value, the specification(s) for the given locale or locale type will be returned. This will only work if the actual value of property is a specifier (this will always be the case for built-in properties, but may or may not apply to user-defined properties). If the actual value of property is not a specifier, this value will simply be returned regardless of locale.

The return value will be a list of instantiators (e.g. vectors specifying pixmap data), or a list of specifications, each of which is a cons of a locale and a list of instantiators. Specifically, if locale is a particular locale (a buffer, window, frame, device, or global), a list of instantiators for that locale will be returned. Otherwise, if locale is a locale type (one of the symbols buffer, window, frame, or device), the specifications for all locales of that type will be returned. Finally, if locale is all, the specifications for all locales of all types will be returned.

The specifications in a specifier determine what the value of property will be in a particular domain or set of circumstances, which is typically a particular Emacs window along with the buffer it contains and the frame and device it lies within. The value is derived from the instantiator associated with the most specific locale (in the order buffer, window, frame, device, and global) that matches the domain in question. In other words, given a domain (i.e. an Emacs window, usually), the specifier for property will first be searched for a specification whose locale is the buffer contained within that window; then for a specification whose locale is the window itself; then for a specification whose locale is the frame that the window is contained within; etc. The first instantiator that is valid for the domain (usually this means that the instantiator is recognized by the device [i.e. the X server or TTY device] that the domain is on). The function glyph-property-instance actually does all this, and is used to determine how to display the glyph.

Function glyph-property-instance glyph property &optional domain

default no-fallback This function returns the instance of glyph’s property in the specified domain. Under most circumstances, domain will be a particular window, and the returned instance describes how the specified property actually is displayed for that window and the particular buffer in it. Note that this may not be the same as how the property appears when the buffer is displayed in a different window or frame, or how the property appears in the same window if you switch to another buffer in that window; and in those cases, the returned instance would be different. The returned instance is an image-instance object, and you can query it using the appropriate image instance functions. For example, you could use imageinstance-depth to find out the depth (number of color planes) of a pixmap displayed in a particular window. The results might be different from the results you would get for another window (perhaps the user specified a different image for the frame that window is on; or perhaps the same image was specified but the window is on a different X server, and that X server has different color capabilities from this one). domain defaults to the selected window if omitted. domain can be a frame or device, instead of a window. The value returned for such a domain is used in special circumstances when a more specific domain does not apply; for example, a frame value might be used for coloring a toolbar, which is conceptually attached to a frame rather than a particular window. The value is also useful in determining what the value would be for a particular window within the frame or device, if it is not overridden by a more specific specification. If property does not name a built-in property, its value will simply be returned unless it is a specifier object, in which case it will be instanced using specifier-

instance.

Optional arguments default and no-fallback are the same as in specifierinstance. See Chapter 41 [Specifiers], page 599.

Function remove-glyph-property glyph property &optional locale tag-set

exact-p This function removes a property from a glyph. For built-in properties, this is analogous to remove-specifier. See Chapter 41 [Specifiers], page 599, for the meaning of the locale, tag-set, and exact-p arguments.

43.1.3 Glyph Convenience Functions

The following functions are provided for working with specific properties of a glyph. Note that these are exactly like calling the general functions described above and passing in the appropriate value for property.

Remember that if you want to determine the “value” of a specific glyph property, you probably want to use the *-instance functions. For example, to determine whether a glyph contributes to its line height, use glyph-contrib-p-instance, not glyph-contrib-p. (The latter will return a boolean specifier or a list of specifications, and you probably aren’t concerned with these.)

Function glyph-image glyph &optional locale This function is equivalent to calling glyph-property with a property of image. The return value will be an image specifier if locale is nil or omitted; otherwise, it will be a specification or list of specifications.

Function set-glyph-image glyph spec &optional locale tag-set how-to-add This function is equivalent to calling set-glyph-property with a property of

image.

Function glyph-image-instance glyph &optional domain default

no-fallback This function returns the instance of glyph’s image in the given domain, and is equivalent to calling glyph-property-instance with a property of image. The return value will be an image instance. Normally domain will be a window or nil (meaning the selected window), and an instance object describing how the image appears in that particular window and buffer will be returned.

Function glyph-contrib-p glyph &optional locale This function is equivalent to calling glyph-property with a property of contrib-p. The return value will be a boolean specifier if locale is nil or omitted; otherwise, it will be a specification or list of specifications.

Function set-glyph-contrib-p glyph spec &optional locale tag-set

how-to-add This function is equivalent to calling set-glyph-property with a property of

contrib-p.

Function glyph-contrib-p-instance glyph &optional domain default

no-fallback This function returns whether the glyph contributes to its line height in the given domain, and is equivalent to calling glyph-property-instance with a property of contrib-p. The return value will be either nil or t. (Normally domain will be a window or nil, meaning the selected window.)

Function glyph-baseline glyph &optional locale This function is equivalent to calling glyph-property with a property of baseline. The return value will be a specifier if locale is nil or omitted; otherwise, it will be a specification or list of specifications.

Function set-glyph-baseline glyph spec &optional locale tag-set

how-to-add This function is equivalent to calling set-glyph-property with a property of

baseline.

Function glyph-baseline-instance glyph &optional domain default

no-fallback This function returns the instance of glyph’s baseline value in the given domain, and is equivalent to calling glyph-property-instance with a property of baseline. The return value will be an integer or nil. Normally domain will be a window or nil (meaning the selected window), and an instance object describing the baseline value appears in that particular window and buffer will be returned.

Function glyph-face glyph

This function returns the face of glyph. (Remember, this is not a specifier, but a simple property.)

Function set-glyph-face glyph face

This function changes the face of glyph to face.

43.1.4 Glyph Dimensions

Function glyph-width glyph &optional window This function returns the width of glyph on window. This may not be exact as it does not take into account all of the context that redisplay will.

Function glyph-ascent glyph &optional window This function returns the ascent value of glyph on window. This may not be exact as it does not take into account all of the context that redisplay will.

Function glyph-descent glyph &optional window This function returns the descent value of glyph on window. This may not be exact as it does not take into account all of the context that redisplay will.

Function glyph-height glyph &optional window This function returns the height of glyph on window. (This is equivalent to the sum of the ascent and descent values.) This may not be exact as it does not take into account all of the context that redisplay will.

43.2 Images

43.2.1 Image Specifiers

An image specifier is used to describe the actual image of a glyph. It works like other specifiers (see Chapter 41 [Specifiers], page 599), in that it contains a number of specifications describing how the image should appear in a variety of circumstances. These specifications are called image instantiators. When XEmacs wants to display the image, it instantiates the image into an image instance. Image instances are their own primitive object type (similar to font instances and color instances), describing how the image appears in a particular domain. (On the other hand, image instantiators, which are just descriptions of how the image should appear, are represented using strings or vectors.)

Function image-specifier-p object

This function returns non-nil if object is an image specifier. Usually, an image specifier results from calling glyph-image on a glyph.

Function make-image-specifier spec-list

This function creates a new image specifier object and initializes it according to spec-list. It is unlikely that you will ever want to do this, but this function is provided for completeness and for experimentation purposes. See Chapter 41 [Specifiers], page 599.

Image instantiators come in many formats: xbm, xpm, gif, jpeg, etc. This describes the format of the data describing the image. The resulting image instances also come in many types – mono-pixmap, color-pixmap, text, pointer, etc. This refers to the behavior of the image and the sorts of places it can appear. (For example, a color-pixmap image has fixed colors specified for it, while a mono-pixmap image comes in two unspecified shades “foreground” and “background” that are determined from the face of the glyph or surrounding text; a text image appears as a string of text and has an unspecified foreground, background, and font; a pointer image behaves like a mono-pixmap image but can only be used as a mouse pointer [mono-pixmap images cannot be used as mouse pointers]; etc.) It is important to keep the distinction between image instantiator format and image instance type in mind. Typically, a given image instantiator format can result in many different image instance types (for example, xpm can be instanced as color-pixmap, mono-pixmap, or pointer; whereas cursor-font can be instanced only as pointer), and a particular image instance type can be generated by many different image instantiator formats (e.g. color-pixmap can be generated by xpm, gif, jpeg, etc.).

See Section 43.2.3 [Image Instances], page 635, for a more detailed discussion of image instance types.

An image instantiator should be a string or a vector of the form

[format :keyword value ...] i.e. a format symbol followed by zero or more alternating keyword-value pairs. The

format field should be a symbol, one of

nothing (Don’t display anything; no keywords are valid for this. Can only be instanced

          as nothing.)

string (Display this image as a text string. Can only be instanced as text, although support for instancing as mono-pixmap should be added.)

formatted-string

(Display this image as a text string with replaceable fields, similar to a modeline format string; not currently implemented.)

xbm (An X bitmap; only if X support was compiled into this XEmacs. Can be

          instanced as mono-pixmap, color-pixmap, or pointer.)

xpm (An XPM pixmap; only if XPM support was compiled into this XEmacs. Can be instanced as color-pixmap, mono-pixmap, or pointer. XPM is an add-on library for X that was designed to rectify the shortcomings of the XBM format. Most implementations of X include the XPM library as a standard part. If your vendor does not, it is highly recommended that you download it and install it. You can get it from the standard XEmacs FTP site, among other places.)

xface (An X-Face bitmap, used to encode people’s faces in e-mail messages; only if X-Face support was compiled into this XEmacs. Can be instanced as mono-

pixmap, color-pixmap, or pointer.)

gif (A GIF87 or GIF89 image; only if GIF support was compiled into this XEmacs. Can be instanced as color-pixmap. Note that XEmacs includes GIF decoding functions as a standard part of it, so if you have X support, you will normally have GIF support, unless you explicitly disable it at configure time.)

jpeg (A JPEG-format image; only if JPEG support was compiled into this XEmacs. Can be instanced as color-pixmap. If you have the JPEG libraries present on your system when XEmacs is built, XEmacs will automatically detect this and use them, unless you explicitly disable it at configure time.)

png (A PNG/GIF24 image; only if PNG support was compiled into this XEmacs. Can be instanced as color-pixmap.)

tiff (A TIFF-format image; only if TIFF support was compiled into this XEmacs. Not currently implemented.)

cursor-font

(One of the standard cursor-font names, such as ‘watch’ or ‘right_ptr’ under X. Under X, this is, more specifically, any of the standard cursor names from appendix B of the Xlib manual [also known as the file ‘<X11/cursorfont.h>’] minus the ‘XC_’ prefix. On other window systems, the valid names will be specific to the type of window system. Can only be instanced as pointer.)

font (A glyph from a font; i.e. the name of a font, and glyph index into it of the

form ‘font fontname index [[mask-font] mask-index]’. Only if X support

was compiled into this XEmacs. Currently can only be instanced as pointer, although this should probably be fixed.)

subwindow

(An embedded X window; not currently implemented.)

autodetect

(XEmacs tries to guess what format the data is in. If X support exists, the data string will be checked to see if it names a filename. If so, and this filename contains XBM or XPM data, the appropriate sort of pixmap or pointer will be created. [This includes picking up any specified hotspot or associated mask file.] Otherwise, if pointer is one of the allowable image-instance types and the string names a valid cursor-font name, the image will be created as a pointer. Otherwise, the image will be displayed as text. If no X support exists, the image will always be displayed as text.)

The valid keywords are:

:data (Inline data. For most formats above, this should be a string. For XBM images, this should be a list of three elements: width, height, and a string of bit data. This keyword is not valid for instantiator format nothing.)

:file (Data is contained in a file. The value is the name of this file. If both :data and :file are specified, the image is created from what is specified in :data and the string in :file becomes the value of the image-instance-file-name function when applied to the resulting image-instance. This keyword is not

valid for instantiator formats nothing, string, formatted-string, cursor-
font, font, and autodetect.)
:foreground
:background

(For xbm, xface, cursor-font, and font. These keywords allow you to explicitly specify foreground and background colors. The argument should be anything acceptable to make-color-instance. This will cause what would be a mono-pixmap to instead be colorized as a two-color color-pixmap, and specifies the foreground and/or background colors for a pointer instead of black and white.)

:mask-data

(For xbm and xface. This specifies a mask to be used with the bitmap. The format is a list of width, height, and bits, like for :data.)

:mask-file

(For xbm and xface. This specifies a file containing the mask data. If neither a mask file nor inline mask data is given for an XBM image, and the XBM image comes from a file, XEmacs will look for a mask file with the same name as the image file but with ‘Mask’ or ‘msk’ appended. For example, if you specify the XBM file ‘left_ptr’ [usually located in ‘/usr/include/X11/bitmaps’], the associated mask file ‘left_ptrmsk’ will automatically be picked up.)

:hotspot-x
:hotspot-y

(For xbm and xface. These keywords specify a hotspot if the image is instantiated as a pointer. Note that if the XBM image file specifies a hotspot, it will automatically be picked up if no explicit hotspot is given.)

:color-symbols

(Only for xpm. This specifies an alist that maps strings that specify symbolic color names to the actual color to be used for that symbolic color (in the form of a string or a color-specifier object). If this is not specified, the contents of xpm-color-symbols are used to generate the alist.)

If instead of a vector, the instantiator is a string, it will be converted into a vector by looking it up according to the specs in the console-type-image-conversion-list for the console type of the domain (usually a window; sometimes a frame or device) over which the image is being instantiated.

If the instantiator specifies data from a file, the data will be read in at the time that the instantiator is added to the image specifier (which may be well before the image is actually displayed), and the instantiator will be converted into one of the inline-data forms, with the filename retained using a :file keyword. This implies that the file must exist when the instantiator is added to the image, but does not need to exist at any other time (e.g. it may safely be a temporary file).

Function valid-image-instantiator-format-p format

This function returns non-nil if format is a valid image instantiator format. Note that the return value for many formats listed above depends on whether XEmacs was compiled with support for that format.

Function image-instantiator-format-list

This function return a list of valid image-instantiator formats.

Variable xpm-color-symbols

This variable holds definitions of logical color-names used when reading XPM files. Elements of this list should be of the form (color-name form-to-evaluate). The color-name should be a string, which is the name of the color to define; the form-to-evaluate should evaluate to a color specifier object, or a string to be passed to make-color-instance (see Section 42.3 [Colors], page 623). If a loaded XPM file references a symbolic color called color-name, it will display as the computed color instead.

The default value of this variable defines the logical color names ‘"foreground"’ and ‘"background"’ to be the colors of the default face.

Variable x-bitmap-file-path

A list of the directories in which X bitmap files may be found. If nil, this is initialized from the ‘"*bitmapFilePath"’ resource. This is used by the make-image-instance function (however, note that if the environment variable ‘XBMLANGPATH’ is set, it is consulted first).

43.2.2 Image Instantiator Conversion

Function set-console-type-image-conversion-list console-type list

This function sets the image-conversion-list for consoles of the given consoletype. The image-conversion-list specifies how image instantiators that are strings should be interpreted. Each element of the list should be a list of two elements (a regular expression string and a vector) or a list of three elements (the preceding two plus an integer index into the vector). The string is converted to the vector associated with the first matching regular expression. If a vector index is specified, the string itself is substituted into that position in the vector.

Note: The conversion above is applied when the image instantiator is added to an image specifier, not when the specifier is actually instantiated. Therefore, changing the image-conversion-list only affects newly-added instantiators. Existing instantiators in glyphs and image specifiers will not be affected.

Function console-type-image-conversion-list console-type

This function returns the image-conversion-list for consoles of the given consoletype.

43.2.3 Image Instances

Image-instance objects encapsulate the way a particular image (pixmap, etc.) is displayed on a particular device.

In most circumstances, you do not need to directly create image instances; use a glyph instead. However, it may occasionally be useful to explicitly create image instances, if you want more control over the instantiation process.

Function image-instance-p object

This function returns non-nil if object is an image instance.

43.2.3.1 Image Instance Types

Image instances come in a number of different types. The type of an image instance specifies the nature of the image: Whether it is a text string, a mono pixmap, a color pixmap, etc.

The valid image instance types are

nothing Nothing is displayed.

text Displayed as text. The foreground and background colors and the font of the text are specified independent of the pixmap. Typically these attributes will come from the face of the surrounding text, unless a face is specified for the glyph in which the image appears.

mono-pixmap

Displayed as a mono pixmap (a pixmap with only two colors where the foreground and background can be specified independent of the pixmap; typically the pixmap assumes the foreground and background colors of the text around it, unless a face is specified for the glyph in which the image appears).

color-pixmap

Displayed as a color pixmap.

pointer Used as the mouse pointer for a window.

subwindow

A child window that is treated as an image. This allows (e.g.) another program to be responsible for drawing into the window. Not currently implemented.

Function valid-image-instance-type-p type

This function returns non-nil if type is a valid image instance type.

Function image-instance-type-list

This function returns a list of the valid image instance types.

Function image-instance-type image-instance

This function returns the type of the given image instance. The return value will

     be one of nothing, text, mono-pixmap, color-pixmap, pointer, or subwindow.

Function text-image-instance-p object

This function returns non-nil if object is an image instance of type text.

Function mono-pixmap-image-instance-p object

This function returns non-nil if object is an image instance of type mono-

pixmap.

Function color-pixmap-image-instance-p object

This function returns non-nil if object is an image instance of type color-

pixmap.

Function pointer-image-instance-p object

This function returns non-nil if object is an image instance of type pointer.

Function subwindow-image-instance-p object

This function returns non-nil if object is an image instance of type subwindow.

Function nothing-image-instance-p object

This function returns non-nil if object is an image instance of type nothing.

43.2.3.2 Image Instance Functions

Function make-image-instance data &optional device dest-types no-error This function creates a new image-instance object. data is an image instantiator, which describes the image (see Section 43.2.1 [Image Specifiers], page 630). dest-types should be a list of allowed image instance types that can be generated. The dest-types list is unordered. If multiple destination types are possible for a given instantiator, the “most natural” type for the instantiator’s format is chosen. (For XBM, the most natural types are mono-pixmap, followed by color-pixmap, followed by pointer. For the other normal image formats, the most natural types are color-pixmap, followed by mono-pixmap, followed by pointer. For the string and formatted-string formats, the most natural types are text, followed by mono-pixmap (not currently implemented), followed by color-pixmap (not currently implemented). The other formats can only be instantiated as one type. (If you want to control more specifically the order of the types into which an image is instantiated, just call make-image-instance repeatedly until it succeeds, passing less and less preferred destination types each time. If dest-types is omitted, all possible types are allowed. no-error controls what happens when the image cannot be generated. If nil, an error message is generated. If t, no messages are generated and this function returns nil. If anything else, a warning message is generated and this function returns nil.

Function colorize-image-instance image-instance foreground background This function makes the image instance be displayed in the given colors. Image instances come in two varieties: bitmaps, which are 1 bit deep which are rendered in the prevailing foreground and background colors; and pixmaps, which are of arbitrary depth (including 1) and which have the colors explicitly specified. This function converts a bitmap to a pixmap. If the image instance was a pixmap already, nothing is done (and nil is returned). Otherwise t is returned.

Function image-instance-name image-instance

This function returns the name of the given image instance.

Function image-instance-string image-instance

This function returns the string of the given image instance. This will only be non-nil for text image instances.

Function image-instance-file-name image-instance

This function returns the file name from which image-instance was read, if known.

Function image-instance-mask-file-name image-instance

This function returns the file name from which image-instance’s mask was read, if known.

Function image-instance-depth image-instance

This function returns the depth of the image instance. This is 0 for a mono pixmap, or a positive integer for a color pixmap.

Function image-instance-height image-instance

This function returns the height of the image instance, in pixels.

Function image-instance-width image-instance

This function returns the width of the image instance, in pixels.

Function image-instance-hotspot-x image-instance

This function returns the X coordinate of the image instance’s hotspot, if known. This is a point relative to the origin of the pixmap. When an image is used as a mouse pointer, the hotspot is the point on the image that sits over the location that the pointer points to. This is, for example, the tip of the arrow or the center of the crosshairs. This will always be nil for a non-pointer image instance.

Function image-instance-hotspot-y image-instance

This function returns the Y coordinate of the image instance’s hotspot, if known.

Function image-instance-foreground image-instance

This function returns the foreground color of image-instance, if applicable. This will be a color instance or nil. (It will only be non-nil for colorized mono pixmaps and for pointers.)

Function image-instance-background image-instance

This function returns the background color of image-instance, if applicable. This will be a color instance or nil. (It will only be non-nil for colorized mono pixmaps and for pointers.)

43.3 Glyph Types

Each glyph has a particular type, which controls how the glyph’s image is generated. Each glyph type has a corresponding list of allowable image instance types that can be generated. When you call glyph-image-instance to retrieve the image instance of a glyph, XEmacs does the equivalent of calling make-image-instance and passing in dest-types the list of allowable image instance types for the glyph’s type.

• buffer glyphs can be used as the begin-glyph or end-glyph of an extent, in the modeline, and in the toolbar. Their image can be instantiated as nothing, mono-pixmap,

color-pixmap, text, and subwindow.

• pointer glyphs can be used to specify the mouse pointer. Their image can be instantiated as pointer.

• icon glyphs can be used to specify the icon used when a frame is iconified. Their image

can be instantiated as mono-pixmap and color-pixmap.

Function glyph-type glyph

This function returns the type of the given glyph. The return value will be a

symbol, one of buffer, pointer, or icon.

Function valid-glyph-type-p glyph-type

Given a glyph-type, this function returns non-nil if it is valid.

Function glyph-type-list

This function returns a list of valid glyph types.

Function buffer-glyph-p object

This function returns non-nil if object is a glyph of type buffer.

Function icon-glyph-p object

This function returns non-nil if object is a glyph of type icon.

Function pointer-glyph-p object

This function returns non-nil if object is a glyph of type pointer.

43.4 Mouse Pointer

The shape of the mouse pointer when over a particular section of a frame is controlled using various glyph variables. Since the image of a glyph is a specifier, it can be controlled on a per-buffer, per-frame, per-window, or per-device basis.

You should use set-glyph-image to set the following variables, not setq.

Glyph text-pointer-glyph

This variable specifies the shape of the mouse pointer when over text.

Glyph nontext-pointer-glyph

This variable specifies the shape of the mouse pointer when over a buffer, but not over text. If unspecified in a particular domain, text-pointer-glyph is used.

Glyph modeline-pointer-glyph

This variable specifies the shape of the mouse pointer when over the modeline. If unspecified in a particular domain, nontext-pointer-glyph is used.

Glyph selection-pointer-glyph

This variable specifies the shape of the mouse pointer when over a selectable text region. If unspecified in a particular domain, text-pointer-glyph is used.

Glyph gc-pointer-glyph

This variable specifies the shape of the mouse pointer when a garbage collection is in progress. If the selected window is on a window system and this glyph specifies a value (i.e. a pointer image instance) in the domain of the selected window, the pointer will be changed as specified during garbage collection. Otherwise, a message will be printed in the echo area, as controlled by gc-

message.

Glyph busy-pointer-glyph

This variable specifies the shape of the mouse pointer when XEmacs is busy. If unspecified in a particular domain, the pointer is not changed when XEmacs is busy.

Glyph menubar-pointer-glyph

This variable specifies the shape of the mouse pointer when over the menubar. If unspecified in a particular domain, the window-system-provided default pointer is used.

Glyph scrollbar-pointer-glyph

This variable specifies the shape of the mouse pointer when over a scrollbar. If unspecified in a particular domain, the window-system-provided default pointer is used.

Glyph toolbar-pointer-glyph

This variable specifies the shape of the mouse pointer when over a toolbar. If unspecified in a particular domain, nontext-pointer-glyph is used.

Internally, these variables are implemented in default-mouse-motion-handler, and thus only take effect when the mouse moves. That function calls set-frame-pointer, which sets the current mouse pointer for a frame.

Function set-frame-pointer frame image-instance This function sets the mouse pointer of frame to the given pointer image instance. You should not call this function directly. (If you do, the pointer will change again the next time the mouse moves.)

43.5 Redisplay Glyphs

Glyph truncation-glyph

This variable specifies what is displayed at the end of truncated lines.

Glyph continuation-glyph

This variable specifies what is displayed at the end of wrapped lines.

Glyph octal-escape-glyph

This variable specifies what to prefix character codes displayed in octal with.

Glyph hscroll-glyph

This variable specifies what to display at the beginning of horizontally scrolled lines.

Glyph invisible-text-glyph

This variable specifies what to use to indicate the presence of invisible text. This is the glyph that is displayed when an ellipsis is called for, according

to selective-display-ellipses or buffer-invisibility-spec). Normally

this is three dots (“...”).

Glyph control-arrow-glyph

This variable specifies what to use as an arrow for control characters.

43.6 Subwindows

Subwindows are not currently implemented.

Function subwindowp object

This function returns non-nil if object is a subwindow.

44 Annotations

An annotation is a pixmap or string that is not part of a buffer’s text but is displayed next to a particular location in a buffer. Annotations can be displayed intermixed with text, in any whitespace at the beginning or end of a line, or in a special area at the left or right side of the frame called a margin, whose size is controllable. Annotations are implemented using extents (see Chapter 40 [Extents], page 585); but you can work with annotations without knowing how extents work.

44.1 Annotation Basics

Marginal annotations are notes associated with a particular location in a buffer. They may be displayed in a margin created on the left-hand or right-hand side of the frame, in any whitespace at the beginning or end of a line, or inside of the text itself. Every annotation may have an associated action to be performed when the annotation is selected. The term annotation is used to refer to an individual note. The term margin is generically used to refer to the whitespace before the first character on a line or after the last character on a line.

Each annotation has the following characteristics:

glyph This is a glyph object and is used as the displayed representation of the annotation.

down-glyph If given, this glyph is used as the displayed representation of the annotation when the mouse is pressed down over the annotation.

face The face with which to display the glyph.

side Which side of the text (left or right) the annotation is displayed at.

action If non-nil, this field must contain a function capable of being the first argument to funcall. This function is normally evaluated with a single argument, the value of the data field, each time the annotation is selected. However, if the with-event parameter to make-annotation is non-nil, the function is called with two arguments. The first argument is the same as before, and the second argument is the event (a button-up event, usually) that activated the annotation.

data Not used internally. This field can contain any E-Lisp object. It is passed as the first argument to action described above.

menu A menu displayed when the right mouse button is pressed over the annotation.

The margin is divided into outside and inside. The outside margin is space on the left or right side of the frame which normal text cannot be displayed in. The inside margin is that space between the leftmost or rightmost point at which text can be displayed and where the first or last character actually is.

There are four different layout types which affect the exact location an annotation appears.

outside-margin

The annotation is placed in the outside margin area. as close as possible to the edge of the frame. If the outside margin is not wide enough for an annotation to fit, it is not displayed.

inside-margin

The annotation is placed in the inside margin area, as close as possible to the edge of the frame. If the inside margin is not wide enough for the annotation to fit, it will be displayed using any available outside margin space if and only if

the specifier use-left-overflow or use-right-overflow (depending on which

side the annotation appears in) is non-nil.

whitespace

The annotation is placed in the inside margin area, as close as possible to the first or last non-whitespace character on a line. If the inside margin is not wide enough for the annotation to fit, it will be displayed if and only if

the specifier use-left-overflow or use-right-overflow (depending on which

side the annotation appears in) is non-nil.

text The annotation is placed at the position it is inserted. It will create enough space for itself inside of the text area. It does not take up a place in the logical buffer, only in the display of the buffer.

The current layout policy is that all whitespace annotations are displayed first. Next, all inside-margin annotations are displayed using any remaining space. Finally as many outside-margin annotations are displayed as possible. The text annotations will always display as they create their own space to display in.

44.2 Annotation Primitives

Function make-annotation glyph &optional position layout buffer

with-event d-glyph rightp This function creates a marginal annotation at position pos in buffer. The annotation is displayed using glyph, which should be a glyph object or a string, and is positioned using layout policy layout. If pos is nil, point is used. If layout is nil, whitespace is used. If buffer is nil, the current buffer is used. If with-event is non-nil, then when an annotation is activated, the triggering event is passed as the second arg to the annotation function. If d-glyph is nonnil then it is used as the glyph that will be displayed when button1 is down. If rightp is non-nil then the glyph will be displayed on the right side of the buffer instead of the left. The newly created annotation is returned.

Function delete-annotation annotation

This function removes annotation from its buffer. This does not modify the buffer text.

Function annotationp annotation

This function returns t if annotation is an annotation, nil otherwise.

44.3 Annotation Properties

Function annotation-glyph annotation

This function returns the glyph object used to display annotation.

Function set-annotation-glyph annotation glyph &optional layout side This function sets the glyph of annotation to glyph, which should be a glyph object. If layout is non-nil, set the layout policy of annotation to layout. If side is left or right, change the side of the buffer at which the annotation is displayed to the given side. The new value of annotation-glyph is returned.

Function annotation-down-glyph annotation

This function returns the glyph used to display annotation when the left mouse button is depressed on the annotation.

Function set-annotation-down-glyph annotation glyph

This function returns the glyph used to display annotation when the left mouse button is depressed on the annotation to glyph, which should be a glyph object.

Function annotation-face annotation

This function returns the face associated with annotation.

Function set-annotation-face annotation face

This function sets the face associated with annotation to face.

Function annotation-layout annotation

This function returns the layout policy of annotation.

Function set-annotation-layout annotation layout

This function sets the layout policy of annotation to layout.

Function annotation-side annotation

This function returns the side of the buffer that annotation is displayed on. Return value is a symbol, either left or right.

Function annotation-data annotation

This function returns the data associated with annotation.

Function set-annotation-data annotation data

This function sets the data field of annotation to data. data is returned.

Function annotation-action annotation

This function returns the action associated with annotation.

Function set-annotation-action annotation action

This function sets the action field of annotation to action. action is returned..

Function annotation-menu annotation

This function returns the menu associated with annotation.

Function set-annotation-menu annotation menu

This function sets the menu associated with annotation to menu. This menu will be displayed when the right mouse button is pressed over the annotation.

Function annotation-visible annotation

This function returns t if there is enough available space to display annotation, nil otherwise.

Function annotation-width annotation

This function returns the width of annotation in pixels.

Function hide-annotation annotation

This function removes annotation’s glyph, making it invisible.

Function reveal-annotation annotation

This function restores annotation’s glyph, making it visible.

44.4 Locating Annotations

Function annotations-in-region start end buffer

This function returns a list of all annotations in buffer which are between start and end inclusively.

Function annotations-at &optional position buffer This function returns a list of all annotations at position in buffer. If position is nil point is used. If buffer is nil the current buffer is used.

Function annotation-list &optional buffer

This function returns a list of all annotations in buffer. If buffer is nil, the current buffer is used.

Function all-annotations

This function returns a list of all annotations in all buffers in existence.

44.5 Margin Primitives

The margin widths are controllable on a buffer-local, window-local, frame-local, devicelocal, or device-type-local basis through the use of specifiers. See Chapter 41 [Specifiers], page 599.

Specifier left-margin-width

This is a specifier variable controlling the width of the left outside margin, in characters. Use set-specifier to change its value.

Specifier right-margin-width

This is a specifier variable controlling the width of the right outside margin, in characters. Use set-specifier to change its value.

Specifier use-left-overflow

If non-nil, use the left outside margin as extra whitespace when displaying whitespace and inside-margin annotations. Defaults to nil. This is a specifier variable; use set-specifier to change its value.

Specifier use-right-overflow

If non-nil, use the right outside margin as extra whitespace when displaying whitespace and inside-margin annotations. Defaults to nil. This is a specifier variable; use set-specifier to change its value.

Function window-left-margin-pixel-width &optional window

This function returns the width in pixels of the left outside margin of window. If window is nil, the selected window is assumed.

Function window-right-margin-pixel-width &optional window

This function returns the width in pixels of the right outside margin of window. If window is nil, the selected window is assumed.

The margin colors are controlled by the faces left-margin and right-margin. These

can be set using the X resources Emacs.left-margin.background and Emacs.left-

margin.foreground; likewise for the right margin.

44.6 Annotation Hooks

The following three hooks are provided for use with the marginal annotations:

before-delete-annotation-hook

This hook is called immediately before an annotation is destroyed. It is passed a single argument, the annotation being destroyed.

after-delete-annotation-hook

This normal hook is called immediately after an annotation is destroyed.

make-annotation-hook

This hook is called immediately after an annotation is created. It is passed a single argument, the newly created annotation.

646 XEmacs Lisp Reference Manual

45 Emacs Display

This chapter describes a number of other features related to the display that XEmacs presents to the user.

45.1 Refreshing the Screen

The function redraw-frame redisplays the entire contents of a given frame. See Chapter 32 [Frames], page 471.

Function redraw-frame frame

This function clears and redisplays frame frame.

Even more powerful is redraw-display:

Command redraw-display &optional device This function redraws all frames on device marked as having their image garbled. device defaults to the selected device. If device is t, all devices will have their frames checked.

Processing user input takes absolute priority over redisplay. If you call these functions when input is available, they do nothing immediately, but a full redisplay does happen eventually—after all the input has been processed.

Normally, suspending and resuming XEmacs also refreshes the screen. Some terminal emulators record separate contents for display-oriented programs such as XEmacs and for ordinary sequential display. If you are using such a terminal, you might want to inhibit the redisplay on resumption. See Section 50.2.2 [Suspending XEmacs], page 696.

Variable no-redraw-on-reenter

This variable controls whether XEmacs redraws the entire screen after it has been suspended and resumed. Non-nil means yes, nil means no.

The above functions do not actually cause the display to be updated; rather, they clear out the internal display records that XEmacs maintains, so that the next time the display is updated it will be redrawn from scratch. Normally this occurs the next time that nextevent or sit-for is called; however, a display update will not occur if there is input pending. See Chapter 19 [Command Loop], page 281.

Function force-cursor-redisplay

This function causes an immediate update of the cursor on the selected frame.

(This function does not exist in FSF Emacs.)

45.2 Truncation

When a line of text extends beyond the right edge of a window, the line can either be truncated or continued on the next line. When a line is truncated, this is normally shown with a ‘\’ in the rightmost column of the window on X displays, and with a ‘$’ on TTY devices. When a line is continued or “wrapped” onto the next line, this is shown with a curved arrow in the rightmost column of the window (or with a ‘\’ on TTY devices). The additional screen lines used to display a long text line are called continuation lines.

Normally, whenever line truncation is in effect for a particular window, a horizontal scrollbar is displayed in that window if the device supports scrollbars. See Chapter 24 [Scrollbars], page 357.

Note that continuation is different from filling; continuation happens on the screen only, not in the buffer contents, and it breaks a line precisely at the right margin, not at a word boundary. See Section 36.11 [Filling], page 526.

User Option truncate-lines

This buffer-local variable controls how XEmacs displays lines that extend beyond the right edge of the window. If it is non-nil, then XEmacs does not display continuation lines; rather each line of text occupies exactly one screen line, and a backslash appears at the edge of any line that extends to or beyond the edge of the window. The default is nil. If the variable truncate-partial-width-windows is non-nil, then truncation is always used for side-by-side windows (within one frame) regardless of the

     value of truncate-lines.

User Option default-truncate-lines

This variable is the default value for truncate-lines, for buffers that do not have local values for it.

User Option truncate-partial-width-windows

This variable controls display of lines that extend beyond the right edge of the window, in side-by-side windows (see Section 31.2 [Splitting Windows], page 446). If it is non-nil, these lines are truncated; otherwise, truncatelines says what to do with them.

The backslash and curved arrow used to indicate truncated or continued lines are only defaults, and can be changed. These images are actually glyphs (see Chapter 43 [Glyphs], page 625). XEmacs provides a great deal of flexibility in how glyphs can be controlled. (This differs from FSF Emacs, which uses display tables to control these images.)

For details, Section 43.5 [Redisplay Glyphs], page 639.

45.3 The Echo Area

The echo area is used for displaying messages made with the message primitive, and for echoing keystrokes. It is not the same as the minibuffer, despite the fact that the minibuffer appears (when active) in the same place on the screen as the echo area. The XEmacs Reference Manual specifies the rules for resolving conflicts between the echo area and the minibuffer for use of that screen space (see section “The Minibuffer” in The XEmacs Reference Manual). Error messages appear in the echo area; see Section 9.5.3 [Errors], page 136.

You can write output in the echo area by using the Lisp printing functions with t as the stream (see Section 17.5 [Output Functions], page 256), or as follows:

Function message string &rest arguments This function displays a one-line message in the echo area. The argument string is similar to a C language printf control string. See format in Section 4.7 [String Conversion], page 65, for the details on the conversion specifications. message returns the constructed string. In batch mode, message prints the message text on the standard error stream, followed by a newline. If string is nil, message clears the echo area. If the minibuffer is active, this brings the minibuffer contents back onto the screen immediately.

     (message "Minibuffer depth is %d."
              (minibuffer-depth))
      ⊣ Minibuffer depth is 0.
     ⇒ "Minibuffer depth is 0."
     ---------- Echo Area ----------
     Minibuffer depth is 0.
     ---------- Echo Area ----------

In addition to only displaying a message, XEmacs allows you to label your messages, giving you fine-grained control of their display. Message label is a symbol denoting the message type. Some standard labels are:

• message—default label used by the message function;

• error—default label used for reporting errors;

• progress—progress indicators like ‘Converting... 45%’ (not logged by default);

• prompt—prompt-like messages like ‘Isearch: foo’ (not logged by default);

• command—helper command messages like ‘Mark set’ (not logged by default);

• no-log—messages that should never be logged

Several messages may be stacked in the echo area at once. Lisp programs may access these messages, or remove them as appropriate, via the message stack.

Function display-message label message &optional frame stdout-p This function displays message (a string) labeled as label, as described above. The frame argument specifies the frame to whose minibuffer the message should be printed. This is currently unimplemented. The stdout-p argument is used internally.

     (display-message 'command "Mark set")

Function lmessage label string &rest arguments This function displays a message string with label label. It is similar to message in that it accepts a printf-like strings and any number of arguments.

;; Display a command message.

     (lmessage 'command "Comment column set to %d" comment-column)

;; Display a progress message.

     (lmessage 'progress "Fontifying %s... (%d)" buffer percentage)

;; Display a message that should not be logged.

     (lmessage 'no-log "Done")

Function clear-message &optional label frame stdout-p no-restore This function remove any message with the given label from the message-stack, erasing it from the echo area if it’s currently displayed there. If a message remains at the head of the message-stack and no-restore is nil, it will be displayed. The string which remains in the echo area will be returned, or nil if the message-stack is now empty. If label is nil, the entire message-stack is cleared.

;; Show a message, wait for 2 seconds, and restore old minibuffer ;; contents.

     (message "A message")
      ⊣ A message
     ⇒ "A Message"
     (lmessage 'my-label "Newsflash!  Newsflash!")
      ⊣ Newsflash!  Newsflash!
     ⇒ "Newsflash!  Newsflash!"
     (sit-for 2)
     (clear-message 'my-label)
      ⊣ A message
     ⇒ "A message"

Unless you need the return value or you need to specify a label, you should just

use (message nil).

Function current-message &optional frame

This function returns the current message in the echo area, or nil. The frame argument is currently unused.

Some of the messages displayed in the echo area are also recorded in the ‘ *Message-Log*’ buffer. Exactly which messages will be recorded can be tuned using the following variables.

User Option log-message-max-size

This variable specifies the maximum size of the ‘ *Message-log*’ buffer.

Variable log-message-ignore-labels

This variable specifies the labels whose messages will not be logged. It should be a list of symbols.

Variable log-message-ignore-regexps

This variable specifies the regular expressions matching messages that will not be logged. It should be a list of regular expressions. Normally, packages that generate messages that might need to be ignored should label them with progress, prompt, or no-log, so they can be filtered by log-

message-ignore-labels.

Variable echo-keystrokes

This variable determines how much time should elapse before command characters echo. Its value must be a number, which specifies the number of seconds to wait before echoing. If the user types a prefix key (such as C-x) and then delays this many seconds before continuing, the prefix key is echoed in the echo area. Any subsequent characters in the same command will be echoed as well.

If the value is zero, then command input is not echoed.

Variable cursor-in-echo-area

This variable controls where the cursor appears when a message is displayed in the echo area. If it is non-nil, then the cursor appears at the end of the message. Otherwise, the cursor appears at point—not in the echo area at all.

The value is normally nil; Lisp programs bind it to t for brief periods of time.

45.4 Warnings

XEmacs contains a facility for unified display of various warnings. Unlike errors, warnings are displayed in the situations when XEmacs encounters a problem that is recoverable, but which should be fixed for safe future operation.

For example, warnings are printed by the startup code when it encounters problems with X keysyms, when there is an error in ‘.emacs’, and in other problematic situations. Unlike messages, warnings are displayed in a separate buffer, and include an explanatory message that may span across several lines. Here is an example of how a warning is displayed:

(1) (initialization/error) An error has occurred while loading ~/.emacs:
Symbol's value as variable is void: bogus-variable
To ensure normal operation, you should investigate the cause of the error
in your initialization file and remove it.  Use the `-debug-init' option
to XEmacs to view a complete error backtrace.

Each warning has a class and a priority level. The class is a symbol describing what sort

of warning this is, such as initialization, resource or key-mapping.

The warning priority level specifies how important the warning is. The recognized warning levels, in increased order of priority, are: debug, info, notice, warning, error,

critical, alert and emergency.

Function display-warning class message &optional level This function displays a warning message message (a string). class should be a warning class symbol, as described above, or a list of such symbols. level describes the warning priority level. If unspecified, it default to warning.

     (display-warning 'resource
       "Bad resource specification encountered:
     something like
         Emacs*foo: bar
     You should replace the * with a . in order to get proper behavior when
     you use the specifier and/or `set-face-*' functions.")
     ---------- Warning buffer ----------
     (1) (resource/warning) Bad resource specification encountered:
     something like
         Emacs*foo: bar
     You should replace the * with a . in order to get proper behavior when
     you use the specifier and/or `set-face-*' functions.
     ---------- Warning buffer ----------

Function lwarn class level message &rest args This function displays a formatted labeled warning message. As above, class should be the warning class symbol, or a list of such symbols, and level should specify the warning priority level (warning by default). Unlike in display-warning, message may be a formatted message, which will be, together with the rest of the arguments, passed to format.

     (lwarn 'message-log 'warning
       "Error caught in `remove-message-hook': %s"
       (error-message-string e))

Variable log-warning-minimum-level

This variable specifies the minimum level of warnings that should be generated. Warnings with level lower than defined by this variable are completely ignored, as if they never happened.

Variable display-warning-minimum-level

This variable specifies the minimum level of warnings that should be displayed. Unlike log-warning-minimum-level, setting this function does not suppress warnings entirely—they are still generated in the ‘*Warnings*’ buffer, only they are not displayed by default.

Variable log-warning-suppressed-classes

This variable specifies a list of classes that should not be logged or displayed. If any of the class symbols associated with a warning is the same as any of the symbols listed here, the warning will be completely ignored, as it they never happened.

Variable display-warning-suppressed-classes

This variable specifies a list of classes that should not be logged or displayed. If any of the class symbols associated with a warning is the same as any of the symbols listed here, the warning will not be displayed. The warning will still logged in the *Warnings* buffer (unless also contained in ‘log-warning-suppressed-classes’), but the buffer will not be automatically popped up.

45.5 Invisible Text

You can make characters invisible, so that they do not appear on the screen, with the invisible property. This can be either a text property or a property of an overlay.

In the simplest case, any non-nil invisible property makes a character invisible. This is the default case—if you don’t alter the default value of buffer-invisibility-spec, this is how the invisibility property works. This feature is much like selective display (see Section 45.6 [Selective Display], page 654), but more general and cleaner.

More generally, you can use the variable buffer-invisibility-spec to control which values of the invisible property make text invisible. This permits you to classify the text into different subsets in advance, by giving them different invisible values, and subsequently make various subsets visible or invisible by changing the value of buffer-

invisibility-spec.

Controlling visibility with buffer-invisibility-spec is especially useful in a program to display the list of entries in a data base. It permits the implementation of convenient filtering commands to view just a part of the entries in the data base. Setting this variable is very fast, much faster than scanning all the text in the buffer looking for properties to change.

Variable buffer-invisibility-spec

This variable specifies which kinds of invisible properties actually make a character invisible.

t A character is invisible if its invisible property is non-nil. This is the default.

a list Each element of the list makes certain characters invisible. Ultimately, a character is invisible if any of the elements of this list applies to it. The list can have two kinds of elements:

atom A character is invisible if its invisible property value is atom or if it is a list with atom as a member.

               (atom . t)

A character is invisible if its invisible property value is atom or if it is a list with atom as a member. Moreover, if this character is at the end of a line and is followed by a visible newline, it displays an ellipsis.

Ordinarily, commands that operate on text or move point do not care whether the text is invisible. However, the user-level line motion commands explicitly ignore invisible newlines.

45.6 Selective Display

Selective display is a pair of features that hide certain lines on the screen.

The first variant, explicit selective display, is designed for use in a Lisp program. The program controls which lines are hidden by altering the text. Outline mode has traditionally used this variant. It has been partially replaced by the invisible text feature (see Section 45.5 [Invisible Text], page 653); there is a new version of Outline mode which uses that instead.

In the second variant, the choice of lines to hide is made automatically based on indentation. This variant is designed to be a user-level feature.

The way you control explicit selective display is by replacing a newline (control-j) with a carriage return (control-m). The text that was formerly a line following that newline is now invisible. Strictly speaking, it is temporarily no longer a line at all, since only newlines can separate lines; it is now part of the previous line.

Selective display does not directly affect editing commands. For example, C-f (forwardchar) moves point unhesitatingly into invisible text. However, the replacement of newline characters with carriage return characters affects some editing commands. For example, next-line skips invisible lines, since it searches only for newlines. Modes that use selective display can also define commands that take account of the newlines, or that make parts of the text visible or invisible.

When you write a selectively displayed buffer into a file, all the control-m’s are output as newlines. This means that when you next read in the file, it looks OK, with nothing invisible. The selective display effect is seen only within XEmacs.

Variable selective-display

This buffer-local variable enables selective display. This means that lines, or portions of lines, may be made invisible.

• If the value of selective-display is t, then any portion of a line that follows a control-m is not displayed.

• If the value of selective-display is a positive integer, then lines that start with more than that many columns of indentation are not displayed.

When some portion of a buffer is invisible, the vertical movement commands operate as if that portion did not exist, allowing a single next-line command to skip any number of invisible lines. However, character movement commands (such as forward-char) do not skip the invisible portion, and it is possible (if tricky) to insert or delete text in an invisible portion.

In the examples below, we show the display appearance of the buffer foo, which changes with the value of selective-display. The contents of the buffer do not change.

(setq selective-display nil)
     ⇒ nil
---------- Buffer: foo ----------
1 on this column
 2on this column
  3n this column
  3n this column
 2on this column
1 on this column
---------- Buffer: foo ----------
(setq selective-display 2)
     ⇒ 2
---------- Buffer: foo ----------
1 on this column
 2on this column
 2on this column
1 on this column
---------- Buffer: foo ----------

Variable selective-display-ellipses

If this buffer-local variable is non-nil, then XEmacs displays ‘...’ at the end of a line that is followed by invisible text. This example is a continuation of the previous one.

(setq selective-display-ellipses t)
     ⇒ t
---------- Buffer: foo ----------
1 on this column
 2on this column ...
 2on this column
1 on this column
---------- Buffer: foo ----------

You can use a display table to substitute other text for the ellipsis (‘...’). See Section 45.11 [Display Tables], page 659.

45.7 The Overlay Arrow

The overlay arrow is useful for directing the user’s attention to a particular line in a buffer. For example, in the modes used for interface to debuggers, the overlay arrow indicates the line of code about to be executed.

Variable overlay-arrow-string

This variable holds the string to display to call attention to a particular line, or nil if the arrow feature is not in use. Despite its name, the value of this variable can be either a string or a glyph (see Chapter 43 [Glyphs], page 625).

Variable overlay-arrow-position

This variable holds a marker that indicates where to display the overlay arrow. It should point at the beginning of a line. The arrow text appears at the beginning of that line, overlaying any text that would otherwise appear. Since the arrow is usually short, and the line usually begins with indentation, normally nothing significant is overwritten. The overlay string is displayed only in the buffer that this marker points into. Thus, only one buffer can have an overlay arrow at any given time.

You can do the same job by creating an extent with a begin-glyph property. See Section 40.6 [Extent Properties], page 591.

45.8 Temporary Displays

Temporary displays are used by commands to put output into a buffer and then present it to the user for perusal rather than for editing. Many of the help commands use this feature.

Special Form with-output-to-temp-buffer buffer-name forms…

This function executes forms while arranging to insert any output they print into the buffer named buffer-name. The buffer is then shown in some window for viewing, displayed but not selected. The string buffer-name specifies the temporary buffer, which need not already exist. The argument must be a string, not a buffer. The buffer is erased initially (with no questions asked), and it is marked as unmodified after with-output-

to-temp-buffer exits.
with-output-to-temp-buffer binds standard-output to the temporary

buffer, then it evaluates the forms in forms. Output using the Lisp output functions within forms goes by default to that buffer (but screen display and messages in the echo area, although they are “output” in the general sense of the word, are not affected). See Section 17.5 [Output Functions], page 256. The value of the last form in forms is returned.

     ---------- Buffer: foo ----------
      This is the contents of foo.
     ---------- Buffer: foo ----------
     (with-output-to-temp-buffer "foo"
         (print 20)
         (print standard-output))
     ⇒ #<buffer foo>
     ---------- Buffer: foo ----------
     20
     #<buffer foo>
     ---------- Buffer: foo ----------

Variable temp-buffer-show-function

If this variable is non-nil, with-output-to-temp-buffer calls it as a function to do the job of displaying a help buffer. The function gets one argument, which is the buffer it should display.

In Emacs versions 18 and earlier, this variable was called temp-buffer-show-

hook.

Function momentary-string-display string position &optional char

message This function momentarily displays string in the current buffer at position. It has no effect on the undo list or on the buffer’s modification status.

The momentary display remains until the next input event. If the next input event is char, momentary-string-display ignores it and returns. Otherwise, that event remains buffered for subsequent use as input. Thus, typing char will simply remove the string from the display, while typing (say) C-f will remove the string from the display and later (presumably) move point forward. The argument char is a space by default.

The return value of momentary-string-display is not meaningful.

You can do the same job in a more general way by creating an extent with a begin-glyph property. See Section 40.6 [Extent Properties], page 591.

If message is non-nil, it is displayed in the echo area while string is displayed in the buffer. If it is nil, a default message says to type char to continue.

In this example, point is initially located at the beginning of the second line:

     ---------- Buffer: foo ----------
     This is the contents of foo.
     ⋆Second line.
     ---------- Buffer: foo ----------
     (momentary-string-display
       "**** Important Message! ****"
       (point) ?\r
       "Type RET when done reading")
     ⇒ t
     ---------- Buffer: foo ----------
     This is the contents of foo.
     **** Important Message! ****Second line.
     ---------- Buffer: foo ----------
     ---------- Echo Area ----------
     Type RET when done reading
     ---------- Echo Area ----------

This function works by actually changing the text in the buffer. As a result, if you later undo in this buffer, you will see the message come and go.

45.9 Blinking Parentheses

This section describes the mechanism by which XEmacs shows a matching open parenthesis when the user inserts a close parenthesis.

Variable blink-paren-function

The value of this variable should be a function (of no arguments) to be called whenever a character with close parenthesis syntax is inserted. The value of blink-paren-function may be nil, in which case nothing is done.

Please note: This variable was named blink-paren-hook in older Emacs versions, but since it is not called with the standard convention for hooks, it was renamed to blink-paren-function in version 19.

Variable blink-matching-paren

If this variable is nil, then blink-matching-open does nothing.

Variable blink-matching-paren-distance

This variable specifies the maximum distance to scan for a matching parenthesis before giving up.

Variable blink-matching-paren-delay

This variable specifies the number of seconds for the cursor to remain at the matching parenthesis. A fraction of a second often gives good results, but the default is 1, which works on all systems.

Function blink-matching-open

This function is the default value of blink-paren-function. It assumes that point follows a character with close parenthesis syntax and moves the cursor momentarily to the matching opening character. If that character is not already on the screen, it displays the character’s context in the echo area. To avoid long delays, this function does not search farther than blink-matching-paren-distance characters. Here is an example of calling this function explicitly.

     (defun interactive-blink-matching-open ()
       "Indicate momentarily the start of sexp before point."
       (interactive)
       (let ((blink-matching-paren-distance
              (buffer-size))
             (blink-matching-paren t))
         (blink-matching-open)))

45.10 Usual Display Conventions

The usual display conventions define how to display each character code. You can override these conventions by setting up a display table (see Section 45.11 [Display Tables], page 659). Here are the usual display conventions:

• Character codes 32 through 126 map to glyph codes 32 through 126. Normally this means they display as themselves.

• Character code 9 is a horizontal tab. It displays as whitespace up to a position deter-

    mined by tab-width.

• Character code 10 is a newline.

• All other codes in the range 0 through 31, and code 127, display in one of two ways according to the value of ctl-arrow. If it is non-nil, these codes map to sequences of two glyphs, where the first glyph is the ASCII code for ‘^’. (A display table can specify a glyph to use instead of ‘^’.) Otherwise, these codes map just like the codes in the range 128 to 255.

• Character codes 128 through 255 map to sequences of four glyphs, where the first glyph is the ASCII code for ‘\’, and the others are digit characters representing the code in octal. (A display table can specify a glyph to use instead of ‘\’.)

The usual display conventions apply even when there is a display table, for any character whose entry in the active display table is nil. Thus, when you set up a display table, you need only specify the characters for which you want unusual behavior.

These variables affect the way certain characters are displayed on the screen. Since they change the number of columns the characters occupy, they also affect the indentation functions.

User Option ctl-arrow

This buffer-local variable controls how control characters are displayed. If it is non-nil, they are displayed as a caret followed by the character: ‘^A’. If it is nil, they are displayed as a backslash followed by three octal digits: ‘\001’.

Variable default-ctl-arrow

The value of this variable is the default value for ctl-arrow in buffers that do not override it. See Section 10.9.3 [Default Value], page 159.

User Option tab-width

The value of this variable is the spacing between tab stops used for displaying tab characters in Emacs buffers. The default is 8. Note that this feature is completely independent from the user-settable tab stops used by the command tab-to-tab-stop. See Section 36.16.5 [Indent Tabs], page 537.

45.11 Display Tables

You can use the display table feature to control how all 256 possible character codes display on the screen. This is useful for displaying European languages that have letters not in the ASCII character set.

The display table maps each character code into a sequence of runes, each rune being an image that takes up one character position on the screen. You can also define how to display each rune on your terminal, using the rune table.

45.11.1 Display Table Format

A display table is an array of 256 elements. (In FSF Emacs, a display table is 262 elements. The six extra elements specify the truncation and continuation glyphs, etc. This method is very kludgey, and in XEmacs the variables truncation-glyph, continuationglyph, etc. are used. See Section 45.2 [Truncation], page 648.)

Function make-display-table

This creates and returns a display table. The table initially has nil in all elements.

The 256 elements correspond to character codes; the nth element says how to display the character code n. The value should be nil, a string, a glyph, or a vector of strings and glyphs (see Section 45.11.3 [Character Descriptors], page 660). If an element is nil, it says to display that character according to the usual display conventions (see Section 45.10 [Usual Display], page 658).

If you use the display table to change the display of newline characters, the whole buffer will be displayed as one long “line.”

For example, here is how to construct a display table that mimics the effect of setting ctl-arrow to a non-nil value:

     (setq disptab (make-display-table))
     (let ((i 0))
       (while (< i 32)
         (or (= i ?\t) (= i ?\n)
             (aset disptab i (concat "^" (char-to-string (+ i 64)))))
         (setq i (1+ i)))
       (aset disptab 127 "^?"))

45.11.2 Active Display Table

The active display table is controlled by the variable current-display-table. This is a specifier, which means that you can specify separate values for it in individual buffers, windows, frames, and devices, as well as a global value. It also means that you cannot set this variable using setq; use set-specifier instead. See Chapter 41 [Specifiers], page 599.

(FSF Emacs uses window-display-table, buffer-display-table, standard-display-

table, etc. to control the display table. However, specifiers are a cleaner and more powerful way of doing the same thing. FSF Emacs also uses a different format for the contents of a display table, using additional indirection to a “glyph table” and such. Note that “glyph” has a different meaning in XEmacs.)

Individual faces can also specify an overriding display table; this is set using set-face-display-table. See Section 42.1 [Faces], page 615.

If no display table can be determined for a particular window, then XEmacs uses the usual display conventions. See Section 45.10 [Usual Display], page 658.

45.11.3 Character Descriptors

Each element of the display-table vector describes how to display a particular character and is called a character descriptor. A character descriptor can be: a string Display this particular string wherever the character is to be displayed.

a glyph Display this particular glyph wherever the character is to be displayed.

a vector The vector may contain strings and/or glyphs. Display the elements of the vector one after another wherever the character is to be displayed.

nil Display according to the standard interpretation (see Section 45.10 [Usual Display], page 658).

45.12 Beeping

You can make XEmacs ring a bell, play a sound, or blink the screen to attract the user’s attention. Be conservative about how often you do this; frequent bells can become irritating. Also be careful not to use beeping alone when signaling an error is appropriate. (See Section 9.5.3 [Errors], page 136.)

Function ding &optional dont-terminate sound device This function beeps, or flashes the screen (see visible-bell below). It also terminates any keyboard macro currently executing unless dont-terminate is non-nil. If sound is specified, it should be a symbol specifying which sound to make. This sound will be played if visible-bell is nil. (This only works if sound support was compiled into the executable and you are running on the console of a Sun SparcStation, SGI, HP9000s700, or Linux PC. Otherwise you just get a beep.) The optional third argument specifies what device to make the sound on, and defaults to the selected device.

Function beep &optional dont-terminate sound device This is a synonym for ding.

User Option visible-bell

This variable determines whether XEmacs should flash the screen to represent a bell. Non-nil means yes, nil means no. On TTY devices, this is effective only if the Termcap entry for the terminal type has the visible bell flag (‘vb’) set.

Variable sound-alist

This variable holds an alist associating names with sounds. When beep or ding is called with one of the name symbols, the associated sound will be generated instead of the standard beep. Each element of sound-alist is a list describing a sound. The first element of the list is the name of the sound being defined. Subsequent elements of the list are alternating keyword/value pairs:

sound A string of raw sound data, or the name of another sound to play. The symbol t here means use the default X beep.

volume An integer from 0-100, defaulting to bell-volume.

pitch If using the default X beep, the pitch (Hz) to generate. duration If using the default X beep, the duration (milliseconds).

For compatibility, elements of ‘sound-alist’ may also be:

  • ( sound-name . <sound> )
  • ( sound-name <volume> <sound> )

You should probably add things to this list by calling the function load-sound-

file.

Caveats: − You can only play audio data if running on the console screen of a Sun

SparcStation, SGI, or HP9000s700. − The pitch, duration, and volume options are available everywhere, but

many X servers ignore the ‘pitch’ option.

The following beep-types are used by XEmacs itself:

auto-save-error

when an auto-save does not succeed

command-error

when the XEmacs command loop catches an error

undefined-key

when you type a key that is undefined

undefined-click

when you use an undefined mouse-click combination

no-completion

during completing-read

y-or-n-p when you type something other than ’y’ or ’n’

yes-or-no-p

when you type something other than ’yes’ or ’no’

default used when nothing else is appropriate.

Other lisp packages may use other beep types, but these are the ones that the C kernel of XEmacs uses.

User Option bell-volume

This variable specifies the default volume for sounds, from 0 to 100.

Command load-default-sounds

This function loads and installs some sound files as beep-types.

Command load-sound-file filename sound-name &optional volume This function reads in an audio file and adds it to sound-alist. The sound file must be in the Sun/NeXT U-LAW format. sound-name should be a symbol, specifying the name of the sound. If volume is specified, the sound will be played at that volume; otherwise, the value of bell-volume will be used.

Function play-sound sound &optional volume device This function plays sound sound, which should be a symbol mentioned in sound-alist. If volume is specified, it overrides the value (if any) specified in sound-alist. device specifies the device to play the sound on, and defaults to the selected device.

Command play-sound-file file &optional volume device This function plays the named sound file at volume volume, which defaults to bell-volume. device specifies the device to play the sound on, and defaults to the selected device.

664 XEmacs Lisp Reference Manual

46 Hash Tables

Function hashtablep object

This function returns non-nil if object is a hash table.

46.1 Introduction to Hash Tables

A hash table is a data structure that provides mappings from arbitrary Lisp objects (called keys) to other arbitrary Lisp objects (called values). There are many ways other than hash tables of implementing the same sort of mapping, e.g. association lists (see Section 5.8 [Association Lists], page 92) and property lists (see Section 5.9 [Property Lists], page 96), but hash tables provide much faster lookup.

When you create a hash table, you specify a size, which indicates the expected number of elements that the table will hold. You are not bound by this size, however; hash tables automatically resize themselves if the number of elements becomes too large.

(Internally, hash tables are hashed using a modification of the linear probing hash table method. This method hashes each key to a particular spot in the hash table, and then scans forward sequentially until a blank entry is found. To look up a key, hash to the appropriate spot, then search forward for the key until either a key is found or a blank entry stops the search. The modification actually used is called double hashing and involves moving forward by a fixed increment, whose value is computed from the original hash value, rather than always moving forward by one. This eliminates problems with clustering that can arise from the simple linear probing method. For more information, see Algorithms (second edition) by Robert Sedgewick, pp. 236-241.)

Function make-hashtable size &optional test-fun This function makes a hash table of initial size size. Comparison between keys is normally done with eql; i.e. two keys must be the same object to be considered equivalent. However, you can explicitly specify the comparison function using test-fun, which must be one of eq, eql, or equal. Note that currently, eq and eql are the same. This will change when bignums are implemented.

Function copy-hashtable old-table

This function makes a new hash table which contains the same keys and values as the given table. The keys and values will not themselves be copied.

Function hashtable-fullness table

This function returns number of entries in table.

46.2 Working With Hash Tables

Function puthash key val table This function hashes key to val in table.

Function gethash key table &optional default This function finds the hash value for key in table. If there is no corresponding value, default is returned (defaults to nil).

Function remhash key table This function removes the hash value for key in table.

Function clrhash table

This function flushes table. Afterwards, the hash table will contain no entries.

Function maphash function table This function maps function over entries in table, calling it with two args, each key and value in the table.

46.3 Weak Hash Tables

A weak hash table is a special variety of hash table whose elements do not count as GC referents. For any key-value pair in such a hash table, if either the key or value (or in some cases, if one particular one of the two) has no references to it outside of weak hash tables (and similar structures such as weak lists), the pair will be removed from the table, and the key and value collected. A non-weak hash table (or any other pointer) would prevent the objects from being collected.

Weak hash tables are useful for keeping track of information in a non-obtrusive way, for example to implement caching. If the cache contains objects such as buffers, markers, image instances, etc. that will eventually disappear and get garbage-collected, using a weak hash table ensures that these objects are collected normally rather than remaining around forever, long past their actual period of use. (Otherwise, you’d have to explicitly map over the hash table every so often and remove unnecessary elements.)

There are three types of weak hash tables:

fully weak hash tables In these hash tables, a pair disappears if either the key or the value is unreferenced outside of the table.

key-weak hash tables In these hash tables, a pair disappears if the key is unreferenced outside of the table, regardless of how the value is referenced.

value-weak hash tables In these hash tables, a pair disappears if the value is unreferenced outside of the table, regardless of how the key is referenced.

Also see Section 5.10 [Weak Lists], page 98.

Function make-weak-hashtable size &optional test-fun This function makes a fully weak hash table of initial size size. test-fun is as in

     make-hashtable.

Function make-key-weak-hashtable size &optional test-fun

This function makes a key-weak hash table of initial size size. test-fun is as in

     make-hashtable.

Function make-value-weak-hashtable size &optional test-fun

This function makes a value-weak hash table of initial size size. test-fun is as

in make-hashtable.

668 XEmacs Lisp Reference Manual

47 Range Tables

A range table is a table that efficiently associated values with ranges of integers. Note that range tables have a read syntax, like this:

#s(range-table data ((-3 2) foo (5 20) bar))

This maps integers in the range (-3, 2) to foo and integers in the range (5, 20) to bar.

Function range-table-p object

Return non-nil if object is a range table.

47.1 Introduction to Range Tables

Function make-range-table

Make a new, empty range table.

Function copy-range-table old-table

Make a new range table which contains the same values for the same ranges as the given table. The values will not themselves be copied.

47.2 Working With Range Tables

Function get-range-table pos table &optional default This function finds value for position pos in table. If there is no corresponding value, return default (defaults to nil).

Function put-range-table start end val table This function sets the value for range (start, end) to be val in table.

Function remove-range-table start end table

This function removes the value for range (start, end) in table.

Function clear-range-table table

This function flushes table.

Function map-range-table function table

This function maps function over entries in table, calling it with three args, the beginning and end of the range and the corresponding value.

670 XEmacs Lisp Reference Manual

48 Databases

Function databasep object

This function returns non-nil if object is a database.

48.1 Connecting to a Database

Function open-database file &optional type subtype access mode This function opens database file, using database method type and subtype, with access rights access and permissions mode. access can be any combination of r w and +, for read, write, and creation flags.

type can have the value 'dbm or 'berkeley_db to select the type of database file to use. (Note: XEmacs may not support both of these types.)

For a type of 'dbm, there are no subtypes, so subtype should by nil.

For a type of 'berkeley_db, the following subtypes are available: 'hash, 'btree, and 'recno. See the manpages for the Berkeley DB functions to more information about these types.

Function close-database obj

This function closes database obj.

Function database-live-p obj

This function returns t iff obj is an active database, else nil.

48.2 Working With a Database

Function get-database key dbase &optional default This function finds the value for key in database. If there is no corresponding value, default is returned (nil if default is omitted).

Function map-database function dbase This function maps function over entries in database, calling it with two args, each key and value in the database.

Function put-database key val dbase &optional replace This function stores key and val in database. If optional fourth arg replace is non-nil, replace any existing entry in the database.

Function remove-database key dbase

This function removes key from database.

48.3 Other Database Functions

Function database-file-name obj

This function returns the filename associated with the database obj.

Function database-last-error &optional obj

This function returns the last error associated with database obj.

Function database-subtype obj

This function returns the subtype of database obj, if any.

Function database-type obj

This function returns the type of database obj.

49 Processes

In the terminology of operating systems, a process is a space in which a program can execute. XEmacs runs in a process. XEmacs Lisp programs can invoke other programs in processes of their own. These are called subprocesses or child processes of the XEmacs process, which is their parent process.

A subprocess of XEmacs may be synchronous or asynchronous, depending on how it is created. When you create a synchronous subprocess, the Lisp program waits for the subprocess to terminate before continuing execution. When you create an asynchronous subprocess, it can run in parallel with the Lisp program. This kind of subprocess is represented within XEmacs by a Lisp object which is also called a “process”. Lisp programs can use this object to communicate with the subprocess or to control it. For example, you can send signals, obtain status information, receive output from the process, or send input to it.

Function processp object

This function returns t if object is a process, nil otherwise.

49.1 Functions that Create Subprocesses

There are three functions that create a new subprocess in which to run a program. One of them, start-process, creates an asynchronous process and returns a process object (see Section 49.4 [Asynchronous Processes], page 677). The other two, call-process and call-process-region, create a synchronous process and do not return a process object (see Section 49.2 [Synchronous Processes], page 674).

Synchronous and asynchronous processes are explained in following sections. Since the three functions are all called in a similar fashion, their common arguments are described here.

In all cases, the function’s program argument specifies the program to be run. An error is signaled if the file is not found or cannot be executed. If the file name is relative, the variable exec-path contains a list of directories to search. Emacs initializes exec-path when it starts up, based on the value of the environment variable PATH. The standard file name constructs, ‘~’, ‘.’, and ‘..’, are interpreted as usual in exec-path, but environment variable substitutions (‘$HOME’, etc.) are not recognized; use substitute-in-file-name to perform them (see Section 28.8.4 [File Name Expansion], page 409).

Each of the subprocess-creating functions has a buffer-or-name argument which specifies where the standard output from the program will go. If buffer-or-name is nil, that says to discard the output unless a filter function handles it. (See Section 49.9.2 [Filter Functions], page 684, and Chapter 17 [Read and Print], page 251.) Normally, you should avoid having multiple processes send output to the same buffer because their output would be intermixed randomly.

All three of the subprocess-creating functions have a &rest argument, args. The args must all be strings, and they are supplied to program as separate command line arguments. Wildcard characters and other shell constructs are not allowed in these strings, since they are passed directly to the specified program.

Please note: The argument program contains only the name of the program; it may not contain any command-line arguments. You must use args to provide those.

The subprocess gets its current directory from the value of default-directory (see Section 28.8.4 [File Name Expansion], page 409).

The subprocess inherits its environment from XEmacs; but you can specify overrides for it with process-environment. See Section 50.3 [System Environment], page 697.

Variable exec-directory

The value of this variable is the name of a directory (a string) that contains programs that come with XEmacs, that are intended for XEmacs to invoke. The program wakeup is an example of such a program; the display-time command uses it to get a reminder once per minute.

User Option exec-path

The value of this variable is a list of directories to search for programs to run in subprocesses. Each element is either the name of a directory (i.e., a string), or nil, which stands for the default directory (which is the value of default-

directory).
The value of exec-path is used by call-process and start-process when

the program argument is not an absolute file name.

49.2 Creating a Synchronous Process

After a synchronous process is created, XEmacs waits for the process to terminate before continuing. Starting Dired is an example of this: it runs ls in a synchronous process, then modifies the output slightly. Because the process is synchronous, the entire directory listing arrives in the buffer before XEmacs tries to do anything with it.

While Emacs waits for the synchronous subprocess to terminate, the user can quit by typing C-g. The first C-g tries to kill the subprocess with a SIGINT signal; but it waits until the subprocess actually terminates before quitting. If during that time the user types another C-g, that kills the subprocess instantly with SIGKILL and quits immediately. See Section 19.8 [Quitting], page 306.

The synchronous subprocess functions returned nil in version 18. In version 19, they return an indication of how the process terminated.

Function call-process program &optional infile destination display &rest

args This function calls program in a separate process and waits for it to finish.

The standard input for the process comes from file infile if infile is not nil and from ‘/dev/null’ otherwise. The argument destination says where to put the process output. Here are the possibilities:

a buffer Insert the output in that buffer, before point. This includes both the standard output stream and the standard error stream of the process. a string Find or create a buffer with that name, then insert the output in

that buffer, before point.

t Insert the output in the current buffer, before point.

nil Discard the output.

0 Discard the output, and return immediately without waiting for the subprocess to finish.

In this case, the process is not truly synchronous, since it can run in parallel with Emacs; but you can think of it as synchronous in that Emacs is essentially finished with the subprocess as soon as this function returns.

(real-destination error-destination)

Keep the standard output stream separate from the standard error stream; deal with the ordinary output as specified by realdestination, and dispose of the error output according to errordestination. The value nil means discard it, t means mix it with the ordinary output, and a string specifies a file name to redirect error output into.

You can’t directly specify a buffer to put the error output in; that is too difficult to implement. But you can achieve this result by sending the error output to a temporary file and then inserting the file into a buffer.

If display is non-nil, then call-process redisplays the buffer as output is inserted. Otherwise the function does no redisplay, and the results become visible on the screen only when XEmacs redisplays that buffer in the normal course of events.

The remaining arguments, args, are strings that specify command line arguments for the program.

The value returned by call-process (unless you told it not to wait) indicates the reason for process termination. A number gives the exit status of the subprocess; 0 means success, and any other value means failure. If the process terminated with a signal, call-process returns a string describing the signal.

In the examples below, the buffer ‘foo’ is current.

     (call-process "pwd" nil t)
          ⇒ nil
     ---------- Buffer: foo ----------
     /usr/user/lewis/manual
     ---------- Buffer: foo ----------

     (call-process "grep" nil "bar" nil "lewis" "/etc/passwd")
          ⇒ nil
     ---------- Buffer: bar ----------
     lewis:5LTsHm66CSWKg:398:21:Bil Lewis:/user/lewis:/bin/csh
     ---------- Buffer: bar ----------

The insert-directory function contains a good example of the use of call-

process:
     (call-process insert-directory-program nil t nil switches
                   (if full-directory-p
                       (concat (file-name-as-directory file) ".")
                     file))

Function call-process-region start end program &optional delete

destination display &rest args This function sends the text between start to end as standard input to a process running program. It deletes the text sent if delete is non-nil; this is useful when buffer is t, to insert the output in the current buffer.

The arguments destination and display control what to do with the output from the subprocess, and whether to update the display as it comes in. For details, see the description of call-process, above. If destination is the integer 0, callprocess-region discards the output and returns nil immediately, without waiting for the subprocess to finish.

The remaining arguments, args, are strings that specify command line arguments for the program.

The return value of call-process-region is just like that of call-process: nil if you told it to return without waiting; otherwise, a number or string which indicates how the subprocess terminated.

In the following example, we use call-process-region to run the cat utility, with standard input being the first five characters in buffer ‘foo’ (the word ‘input’). cat copies its standard input into its standard output. Since the argument destination is t, this output is inserted in the current buffer.

     ---------- Buffer: foo ----------
     input⋆
     ---------- Buffer: foo ----------
     (call-process-region 1 6 "cat" nil t)
          ⇒ nil
     ---------- Buffer: foo ----------
     inputinput⋆
     ---------- Buffer: foo ----------
The shell-command-on-region command uses call-process-region like

this:

     (call-process-region
      start end
      shell-file-name      ; Name of program.

nil ; Do not delete region.

      buffer               ; Send output to buffer.

nil ; No redisplay during output. "-c" command) ; Arguments for the shell.

49.3 MS-DOS Subprocesses

On MS-DOS, you must indicate whether the data going to and from a synchronous subprocess are text or binary. Text data requires translation between the end-of-line convention used within Emacs (a single newline character) and the convention used outside Emacs (the two-character sequence, crlf).

The variable binary-process-input applies to input sent to the subprocess, and binary-process-output applies to output received from it. A non-nil value means the data is non-text; nil means the data is text, and calls for conversion.

Variable binary-process-input

If this variable is nil, convert newlines to crlf sequences in the input to a synchronous subprocess.

Variable binary-process-output

If this variable is nil, convert crlf sequences to newlines in the output from a synchronous subprocess.

See Section 28.14 [Files and MS-DOS], page 419, for related information.

49.4 Creating an Asynchronous Process

After an asynchronous process is created, Emacs and the Lisp program both continue running immediately. The process may thereafter run in parallel with Emacs, and the two may communicate with each other using the functions described in following sections. Here we describe how to create an asynchronous process with start-process.

Function start-process name buffer-or-name program &rest args This function creates a new asynchronous subprocess and starts the program program running in it. It returns a process object that stands for the new subprocess in Lisp. The argument name specifies the name for the process object; if a process with this name already exists, then name is modified (by adding ‘<1>’, etc.) to be unique. The buffer buffer-or-name is the buffer to associate with the process. The remaining arguments, args, are strings that specify command line arguments for the program. In the example below, the first process is started and runs (rather, sleeps) for 100 seconds. Meanwhile, the second process is started, and given the name ‘my-process<1>’ for the sake of uniqueness. It inserts the directory listing at the end of the buffer ‘foo’, before the first process finishes. Then it finishes, and a message to that effect is inserted in the buffer. Much later, the first process finishes, and another message is inserted in the buffer for it.

     (start-process "my-process" "foo" "sleep" "100")
          ⇒ #<process my-process>
     (start-process "my-process" "foo" "ls" "-l" "/user/lewis/bin")
          ⇒ #<process my-process<1>>
     ---------- Buffer: foo ----------
     total 2
     lrwxrwxrwx  1 lewis     14 Jul 22 10:12 gnuemacs --> /emacs
     -rwxrwxrwx  1 lewis     19 Jul 30 21:02 lemon
     Process my-process<1> finished
     Process my-process finished
     ---------- Buffer: foo ----------

Function start-process-shell-command name buffer-or-name command

&rest command-args This function is like start-process except that it uses a shell to execute the specified command. The argument command is a shell command name, and command-args are the arguments for the shell command.

Variable process-connection-type

This variable controls the type of device used to communicate with asynchronous subprocesses. If it is non-nil, then ptys are used, when available. Otherwise, pipes are used. ptys are usually preferable for processes visible to the user, as in Shell mode, because they allow job control (C-c, C-z, etc.) to work between the process and its children whereas pipes do not. For subprocesses used for internal purposes by programs, it is often better to use a pipe, because they are more efficient. In addition, the total number of ptys is limited on many systems and it is good not to waste them.

The value process-connection-type is used when start-process is called.

So you can specify how to communicate with one subprocess by binding the variable around the call to start-process.

     (let ((process-connection-type nil))  ; Use a pipe.
       (start-process ...))

To determine whether a given subprocess actually got a pipe or a pty, use the function process-tty-name (see Section 49.6 [Process Information], page 679).

49.5 Deleting Processes

Deleting a process disconnects XEmacs immediately from the subprocess, and removes it from the list of active processes. It sends a signal to the subprocess to make the subprocess terminate, but this is not guaranteed to happen immediately. The process object itself continues to exist as long as other Lisp objects point to it.

You can delete a process explicitly at any time. Processes are deleted automatically after they terminate, but not necessarily right away. If you delete a terminated process explicitly before it is deleted automatically, no harm results.

Variable delete-exited-processes

This variable controls automatic deletion of processes that have terminated (due to calling exit or to a signal). If it is nil, then they continue to exist until the user runs list-processes. Otherwise, they are deleted immediately after they exit.

Function delete-process name

This function deletes the process associated with name, killing it with a SIGHUP signal. The argument name may be a process, the name of a process, a buffer, or the name of a buffer.

     (delete-process "*shell*")
          ⇒ nil

Function process-kill-without-query process &optional require-query-p This function declares that XEmacs need not query the user if process is still running when XEmacs is exited. The process will be deleted silently. If requirequery-p is non-nil, then XEmacs will query the user (this is the default). The return value is t if a query was formerly required, and nil otherwise.

     (process-kill-without-query (get-process "shell"))
          ⇒ t

49.6 Process Information

Several functions return information about processes. list-processes is provided for interactive use.

Command list-processes

This command displays a listing of all living processes. In addition, it finally deletes any process whose status was ‘Exited’ or ‘Signaled’. It returns nil.

Function process-list

This function returns a list of all processes that have not been deleted.

     (process-list)
          ⇒ (#<process display-time> #<process shell>)

Function get-process name

This function returns the process named name, or nil if there is none. An error is signaled if name is not a string.

     (get-process "shell")
          ⇒ #<process shell>

Function process-command process

This function returns the command that was executed to start process. This is a list of strings, the first string being the program executed and the rest of the strings being the arguments that were given to the program.

     (process-command (get-process "shell"))
          ⇒ ("/bin/csh" "-i")

Function process-id process

This function returns the pid of process. This is an integer that distinguishes the process process from all other processes running on the same computer at the current time. The pid of a process is chosen by the operating system kernel when the process is started and remains constant as long as the process exists.

Function process-name process

This function returns the name of process.

Function process-status process-name

This function returns the status of process-name as a symbol. The argument process-name must be a process, a buffer, a process name (string) or a buffer name (string). The possible values for an actual subprocess are:

run for a process that is running.

stop for a process that is stopped but continuable.

exit for a process that has exited.

signal for a process that has received a fatal signal.

open for a network connection that is open.

closed for a network connection that is closed. Once a connection is closed, you cannot reopen it, though you might be able to open a new connection to the same place.

nil if process-name is not the name of an existing process.

     (process-status "shell")
          ⇒ run
     (process-status (get-buffer "*shell*"))
          ⇒ run
     x
          ⇒ #<process xx<1>>
     (process-status x)
          ⇒ exit

For a network connection, process-status returns one of the symbols open or closed. The latter means that the other side closed the connection, or XEmacs

did delete-process.

In earlier Emacs versions (prior to version 19), the status of a network connection was run if open, and exit if closed.

Function process-kill-without-query-p process

This function returns whether process will be killed without querying the user, if it is running when XEmacs is exited. The default value is nil.

Function process-exit-status process

This function returns the exit status of process or the signal number that killed it. (Use the result of process-status to determine which of those it is.) If process has not yet terminated, the value is 0.

Function process-tty-name process

This function returns the terminal name that process is using for its communication with Emacs—or nil if it is using pipes instead of a terminal (see processconnection-type in Section 49.4 [Asynchronous Processes], page 677).

49.7 Sending Input to Processes

Asynchronous subprocesses receive input when it is sent to them by XEmacs, which is done with the functions in this section. You must specify the process to send input to, and the input data to send. The data appears on the “standard input” of the subprocess.

Some operating systems have limited space for buffered input in a pty. On these systems, Emacs sends an eof periodically amidst the other characters, to force them through. For most programs, these eofs do no harm.

Function process-send-string process-name string

This function sends process-name the contents of string as standard input. The argument process-name must be a process or the name of a process. If it is nil, the current buffer’s process is used. The function returns nil.

     (process-send-string "shell<1>" "ls\n")
          ⇒ nil
     ---------- Buffer: *shell* ----------
     ...
     introduction.texi               syntax-tables.texi~
     introduction.texi~              text.texi
     introduction.txt                text.texi~
     ...
     ---------- Buffer: *shell* ----------

Command process-send-region process-name start end This function sends the text in the region defined by start and end as standard input to process-name, which is a process or a process name. (If it is nil, the current buffer’s process is used.) An error is signaled unless both start and end are integers or markers that indicate positions in the current buffer. (It is unimportant which number is larger.)

Function process-send-eof &optional process-name This function makes process-name see an end-of-file in its input. The eof comes after any text already sent to it.

If process-name is not supplied, or if it is nil, then this function sends the eof to the current buffer’s process. An error is signaled if the current buffer has no process.

The function returns process-name.

          (process-send-eof "shell")
               ⇒ "shell"

49.8 Sending Signals to Processes

Sending a signal to a subprocess is a way of interrupting its activities. There are several different signals, each with its own meaning. The set of signals and their names is defined by the operating system. For example, the signal SIGINT means that the user has typed C-c, or that some analogous thing has happened.

Each signal has a standard effect on the subprocess. Most signals kill the subprocess, but some stop or resume execution instead. Most signals can optionally be handled by programs; if the program handles the signal, then we can say nothing in general about its effects.

The set of signals and their names is defined by the operating system; XEmacs has facilities for sending only a few of the signals that are defined. XEmacs can send signals only to its own subprocesses.

You can send signals explicitly by calling the functions in this section. XEmacs also sends signals automatically at certain times: killing a buffer sends a SIGHUP signal to all its associated processes; killing XEmacs sends a SIGHUP signal to all remaining processes. (SIGHUP is a signal that usually indicates that the user hung up the phone.)

Each of the signal-sending functions takes two optional arguments: process-name and current-group.

The argument process-name must be either a process, the name of one, or nil. If it is nil, the process defaults to the process associated with the current buffer. An error is signaled if process-name does not identify a process.

The argument current-group is a flag that makes a difference when you are running a jobcontrol shell as an XEmacs subprocess. If it is non-nil, then the signal is sent to the current process-group of the terminal that XEmacs uses to communicate with the subprocess. If the process is a job-control shell, this means the shell’s current subjob. If it is nil, the signal is sent to the process group of the immediate subprocess of XEmacs. If the subprocess is a job-control shell, this is the shell itself.

The flag current-group has no effect when a pipe is used to communicate with the subprocess, because the operating system does not support the distinction in the case of pipes. For the same reason, job-control shells won’t work when a pipe is used. See processconnection-type in Section 49.4 [Asynchronous Processes], page 677.

Function interrupt-process &optional process-name current-group This function interrupts the process process-name by sending the signal SIGINT. Outside of XEmacs, typing the “interrupt character” (normally C-c on some systems, and DEL on others) sends this signal. When the argument currentgroup is non-nil, you can think of this function as “typing C-c” on the terminal by which XEmacs talks to the subprocess.

Function kill-process &optional process-name current-group This function kills the process process-name by sending the signal SIGKILL. This signal kills the subprocess immediately, and cannot be handled by the subprocess.

Function quit-process &optional process-name current-group This function sends the signal SIGQUIT to the process process-name. This signal is the one sent by the “quit character” (usually C-b or C-\) when you are not inside XEmacs.

Function stop-process &optional process-name current-group This function stops the process process-name by sending the signal SIGTSTP. Use continue-process to resume its execution. On systems with job control, the “stop character” (usually C-z) sends this signal (outside of XEmacs). When current-group is non-nil, you can think of this function as “typing C-z” on the terminal XEmacs uses to communicate with the subprocess.

Function continue-process &optional process-name current-group This function resumes execution of the process process by sending it the signal SIGCONT. This presumes that process-name was stopped previously.

Function signal-process pid signal

This function sends a signal to process pid, which need not be a child of XEmacs. The argument signal specifies which signal to send; it should be an integer.

49.9 Receiving Output from Processes

There are two ways to receive the output that a subprocess writes to its standard output stream. The output can be inserted in a buffer, which is called the associated buffer of the process, or a function called the filter function can be called to act on the output. If the process has no buffer and no filter function, its output is discarded.

49.9.1 Process Buffers

A process can (and usually does) have an associated buffer, which is an ordinary Emacs buffer that is used for two purposes: storing the output from the process, and deciding when to kill the process. You can also use the buffer to identify a process to operate on, since in normal practice only one process is associated with any given buffer. Many applications of processes also use the buffer for editing input to be sent to the process, but this is not built into XEmacs Lisp.

Unless the process has a filter function (see Section 49.9.2 [Filter Functions], page 684), its output is inserted in the associated buffer. The position to insert the output is determined by the process-mark, which is then updated to point to the end of the text just inserted. Usually, but not always, the process-mark is at the end of the buffer.

Function process-buffer process

This function returns the associated buffer of the process process.

     (process-buffer (get-process "shell"))
          ⇒ #<buffer *shell*>

Function process-mark process

This function returns the process marker for process, which is the marker that says where to insert output from the process. If process does not have a buffer, process-mark returns a marker that points nowhere. Insertion of process output in a buffer uses this marker to decide where to insert, and updates it to point after the inserted text. That is why successive batches of output are inserted consecutively. Filter functions normally should use this marker in the same fashion as is done by direct insertion of output in the buffer. A good example of a filter function that uses process-mark is found at the end of the following section. When the user is expected to enter input in the process buffer for transmission to the process, the process marker is useful for distinguishing the new input from previous output.

Function set-process-buffer process buffer

This function sets the buffer associated with process to buffer. If buffer is nil, the process becomes associated with no buffer.

Function get-buffer-process buffer-or-name

This function returns the process associated with buffer-or-name. If there are several processes associated with it, then one is chosen. (Presently, the one chosen is the one most recently created.) It is usually a bad idea to have more than one process associated with the same buffer.

     (get-buffer-process "*shell*")
          ⇒ #<process shell>

Killing the process’s buffer deletes the process, which kills the subprocess with a SIGHUP signal (see Section 49.8 [Signals to Processes], page 682).

49.9.2 Process Filter Functions

A process filter function is a function that receives the standard output from the associated process. If a process has a filter, then all output from that process is passed to the filter. The process buffer is used directly for output from the process only when there is no filter.

A filter function must accept two arguments: the associated process and a string, which is the output. The function is then free to do whatever it chooses with the output.

A filter function runs only while XEmacs is waiting (e.g., for terminal input, or for time to elapse, or for process output). This avoids the timing errors that could result from running filters at random places in the middle of other Lisp programs. You may explicitly cause Emacs to wait, so that filter functions will run, by calling sit-for or sleepfor (see Section 19.7 [Waiting], page 305), or accept-process-output (see Section 49.9.3 [Accepting Output], page 686). Emacs is also waiting when the command loop is reading input.

Quitting is normally inhibited within a filter function—otherwise, the effect of typing C-g at command level or to quit a user command would be unpredictable. If you want to permit quitting inside a filter function, bind inhibit-quit to nil. See Section 19.8 [Quitting], page 306.

If an error happens during execution of a filter function, it is caught automatically, so that it doesn’t stop the execution of whatever program was running when the filter function was started. However, if debug-on-error is non-nil, the error-catching is turned off. This makes it possible to use the Lisp debugger to debug the filter function. See Section 16.1 [Debugger], page 217.

Many filter functions sometimes or always insert the text in the process’s buffer, mimicking the actions of XEmacs when there is no filter. Such filter functions need to use set-buffer in order to be sure to insert in that buffer. To avoid setting the current buffer semipermanently, these filter functions must use unwind-protect to make sure to restore the previous current buffer. They should also update the process marker, and in some cases update the value of point. Here is how to do these things:

     (defun ordinary-insertion-filter (proc string)
       (let ((old-buffer (current-buffer)))
         (unwind-protect
             (let (moving)
               (set-buffer (process-buffer proc))
               (setq moving (= (point) (process-mark proc)))
               (save-excursion

;; Insert the text, moving the process-marker.

                 (goto-char (process-mark proc))
                 (insert string)
                 (set-marker (process-mark proc) (point)))
               (if moving (goto-char (process-mark proc))))
           (set-buffer old-buffer))))

The reason to use an explicit unwind-protect rather than letting save-excursion restore the current buffer is so as to preserve the change in point made by goto-char.

To make the filter force the process buffer to be visible whenever new text arrives, insert the following line just before the unwind-protect:

     (display-buffer (process-buffer proc))

To force point to move to the end of the new output no matter where it was previously, eliminate the variable moving and call goto-char unconditionally.

In earlier Emacs versions, every filter function that did regexp searching or matching had to explicitly save and restore the match data. Now Emacs does this automatically; filter functions never need to do it explicitly. See Section 37.6 [Match Data], page 561.

A filter function that writes the output into the buffer of the process should check whether the buffer is still alive. If it tries to insert into a dead buffer, it will get an error.

If the buffer is dead, (buffer-name (process-buffer process)) returns nil.

The output to the function may come in chunks of any size. A program that produces the same output twice in a row may send it as one batch of 200 characters one time, and five batches of 40 characters the next.

Function set-process-filter process filter

This function gives process the filter function filter. If filter is nil, then the process will have no filter. If filter is t, then no output from the process will be accepted until the filter is changed. (Output received during this time is not discarded, but is queued, and will be processed as soon as the filter is changed.)

Function process-filter process

This function returns the filter function of process, or nil if it has none. t means that output processing has been stopped.

Here is an example of use of a filter function:

(defun keep-output (process output)
   (setq kept (cons output kept)))
     ⇒ keep-output
(setq kept nil)
     ⇒ nil
(set-process-filter (get-process "shell") 'keep-output)
     ⇒ keep-output
(process-send-string "shell" "ls ~/other\n")
     ⇒ nil
kept
     ⇒ ("lewis@slug[8] % "
"FINAL-W87-SHORT.MSS    backup.otl              kolstad.mss~
address.txt             backup.psf              kolstad.psf
backup.bib~             david.mss               resume-Dec-86.mss~
backup.err              david.psf               resume-Dec.psf
backup.mss              dland                   syllabus.mss
"
"#backups.mss#          backup.mss~             kolstad.mss
")

49.9.3 Accepting Output from Processes

Output from asynchronous subprocesses normally arrives only while XEmacs is waiting for some sort of external event, such as elapsed time or terminal input. Occasionally it is useful in a Lisp program to explicitly permit output to arrive at a specific point, or even to wait until output arrives from a process.

Function accept-process-output &optional process seconds millisec This function allows XEmacs to read pending output from processes. The output is inserted in the associated buffers or given to their filter functions. If process is non-nil then this function does not return until some output has been received from process. The arguments seconds and millisec let you specify timeout periods. The former specifies a period measured in seconds and the latter specifies one measured in milliseconds. The two time periods thus specified are added together, and accept-process-output returns after that much time whether or not there has been any subprocess output. Note that seconds is allowed to be a floatingpoint number; thus, there is no need to ever use millisec. (It is retained for compatibility purposes.) The function accept-process-output returns non-nil if it did get some output, or nil if the timeout expired before output arrived.

49.10 Sentinels: Detecting Process Status Changes

A process sentinel is a function that is called whenever the associated process changes status for any reason, including signals (whether sent by XEmacs or caused by the process’s own actions) that terminate, stop, or continue the process. The process sentinel is also called if the process exits. The sentinel receives two arguments: the process for which the event occurred, and a string describing the type of event.

The string describing the event looks like one of the following:

  • "finished\n".
  • "exited abnormally with code exitcode\n".

• "name-of-signal\n".

  • "name-of-signal (core dumped)\n".

A sentinel runs only while XEmacs is waiting (e.g., for terminal input, or for time to elapse, or for process output). This avoids the timing errors that could result from running them at random places in the middle of other Lisp programs. A program can wait, so that sentinels will run, by calling sit-for or sleep-for (see Section 19.7 [Waiting], page 305), or accept-process-output (see Section 49.9.3 [Accepting Output], page 686). Emacs is also waiting when the command loop is reading input.

Quitting is normally inhibited within a sentinel—otherwise, the effect of typing C-g at command level or to quit a user command would be unpredictable. If you want to permit quitting inside a sentinel, bind inhibit-quit to nil. See Section 19.8 [Quitting], page 306.

A sentinel that writes the output into the buffer of the process should check whether the buffer is still alive. If it tries to insert into a dead buffer, it will get an error. If the buffer

is dead, (buffer-name (process-buffer process)) returns nil.

If an error happens during execution of a sentinel, it is caught automatically, so that it doesn’t stop the execution of whatever programs was running when the sentinel was started. However, if debug-on-error is non-nil, the error-catching is turned off. This makes it possible to use the Lisp debugger to debug the sentinel. See Section 16.1 [Debugger], page 217.

In earlier Emacs versions, every sentinel that did regexp searching or matching had to explicitly save and restore the match data. Now Emacs does this automatically; sentinels never need to do it explicitly. See Section 37.6 [Match Data], page 561.

Function set-process-sentinel process sentinel

This function associates sentinel with process. If sentinel is nil, then the process will have no sentinel. The default behavior when there is no sentinel is to insert a message in the process’s buffer when the process status changes.

     (defun msg-me (process event)
        (princ
          (format "Process: %s had the event `%s'" process event)))
     (set-process-sentinel (get-process "shell") 'msg-me)
          ⇒ msg-me
     (kill-process (get-process "shell"))
          ⊣ Process: #<process shell> had the event `killed'
          ⇒ #<process shell>

Function process-sentinel process

This function returns the sentinel of process, or nil if it has none.

Function waiting-for-user-input-p

While a sentinel or filter function is running, this function returns non-nil if XEmacs was waiting for keyboard input from the user at the time the sentinel or filter function was called, nil if it was not.

49.11 Process Window Size

Function set-process-window-size process height width

This function tells process that its logical window size is height by width characters. This is principally useful with pty’s.

49.12 Transaction Queues

You can use a transaction queue for more convenient communication with subprocesses using transactions. First use tq-create to create a transaction queue communicating with a specified process. Then you can call tq-enqueue to send a transaction.

Function tq-create process

This function creates and returns a transaction queue communicating with process. The argument process should be a subprocess capable of sending and receiving streams of bytes. It may be a child process, or it may be a TCP connection to a server, possibly on another machine.

Function tq-enqueue queue question regexp closure fn This function sends a transaction to queue queue. Specifying the queue has the effect of specifying the subprocess to talk to.

The argument question is the outgoing message that starts the transaction. The argument fn is the function to call when the corresponding answer comes back; it is called with two arguments: closure, and the answer received.

The argument regexp is a regular expression that should match the entire answer, but nothing less; that’s how tq-enqueue determines where the answer ends. The return value of tq-enqueue itself is not meaningful.

Function tq-close queue

Shut down transaction queue queue, waiting for all pending transactions to complete, and then terminate the connection or child process.

Transaction queues are implemented by means of a filter function. See Section 49.9.2 [Filter Functions], page 684.

49.13 Network Connections

XEmacs Lisp programs can open TCP network connections to other processes on the same machine or other machines. A network connection is handled by Lisp much like a subprocess, and is represented by a process object. However, the process you are communicating with is not a child of the XEmacs process, so you can’t kill it or send it signals. All you can do is send and receive data. delete-process closes the connection, but does not kill the process at the other end; that process must decide what to do about closure of the connection.

You can distinguish process objects representing network connections from those representing subprocesses with the process-status function. It always returns either open or closed for a network connection, and it never returns either of those values for a real subprocess. See Section 49.6 [Process Information], page 679.

Function open-network-stream name buffer-or-name host service This function opens a TCP connection for a service to a host. It returns a process object to represent the connection. The name argument specifies the name for the process object. It is modified as necessary to make it unique. The buffer-or-name argument is the buffer to associate with the connection. Output from the connection is inserted in the buffer, unless you specify a filter function to handle the output. If buffer-or-name is nil, it means that the connection is not associated with any buffer. The arguments host and service specify where to connect to; host is the host name or IP address (a string), and service is the name of a defined network service (a string) or a port number (an integer).

690 XEmacs Lisp Reference Manual

50 Operating System Interface

This chapter is about starting and getting out of Emacs, access to values in the operating system environment, and terminal input, output, and flow control.

See Section B.1 [Building XEmacs], page 765, for related information. See also Chapter 45 [Display], page 647, for additional operating system status information pertaining to the terminal and the screen.

50.1 Starting Up XEmacs

This section describes what XEmacs does when it is started, and how you can customize these actions.

50.1.1 Summary: Sequence of Actions at Start Up

The order of operations performed (in ‘startup.el’) by XEmacs when it is started up is as follows:

1. It loads the initialization library for the window system, if you are using a window system. This library’s name is ‘term/windowsystem-win.el’.

2. It processes the initial options. (Some of them are handled even earlier than this.)

3. It initializes the X window frame and faces, if appropriate.

4. It runs the normal hook before-init-hook.

5. It loads the library ‘site-start’, unless the option ‘-no-site-file’ was specified. The library’s file name is usually ‘site-start.el’.

6. It loads the file ‘~/.emacs’ unless ‘-q’ was specified on the command line. (This is not done in ‘-batch’ mode.) The ‘-u’ option can specify the user name whose home directory should be used instead of ‘~’.

7. It loads the library ‘default’ unless inhibit-default-init is non-nil. (This is not done in ‘-batch’ mode or if ‘-q’ was specified on the command line.) The library’s file name is usually ‘default.el’.

8. It runs the normal hook after-init-hook.

9. It sets the major mode according to initial-major-mode, provided the buffer ‘*scratch*’ is still current and still in Fundamental mode.

10. It loads the terminal-specific Lisp file, if any, except when in batch mode or using a window system.

11. It displays the initial echo area message, unless you have suppressed that with inhibit-

startup-echo-area-message.

12. It processes the action arguments from the command line.

13. It runs term-setup-hook.

14. It calls frame-notice-user-settings, which modifies the parameters of the selected frame according to whatever the init files specify.

15. It runs window-setup-hook. See Section 50.1.3 [Terminal-Specific], page 693.

16. It displays copyleft, nonwarranty, and basic use information, provided there were no remaining command line arguments (a few steps above) and the value of inhibit-

startup-message is nil.

User Option inhibit-startup-message

This variable inhibits the initial startup messages (the nonwarranty, etc.). If it is non-nil, then the messages are not printed. This variable exists so you can set it in your personal init file, once you are familiar with the contents of the startup message. Do not set this variable in the init file of a new user, or in a way that affects more than one user, because that would prevent new users from receiving the information they are supposed to see.

User Option inhibit-startup-echo-area-message

This variable controls the display of the startup echo area message. You can suppress the startup echo area message by adding text with this form to your

‘.emacs’ file:
     (setq inhibit-startup-echo-area-message

"your-login-name")

Simply setting inhibit-startup-echo-area-message to your login name is

not sufficient to inhibit the message; Emacs explicitly checks whether ‘.emacs’ contains an expression as shown above. Your login name must appear in the expression as a Lisp string constant. This way, you can easily inhibit the message for yourself if you wish, but thoughtless copying of your ‘.emacs’ file will not inhibit the message for someone else.

50.1.2 The Init File: ‘.emacs’

When you start XEmacs, it normally attempts to load the file ‘.emacs’ from your home directory. This file, if it exists, must contain Lisp code. It is called your init file. The command line switches ‘-q’ and ‘-u’ affect the use of the init file; ‘-q’ says not to load an init file, and ‘-u’ says to load a specified user’s init file instead of yours. See section “Entering XEmacs” in The XEmacs User’s Manual.

A site may have a default init file, which is the library named ‘default.el’. XEmacs finds the ‘default.el’ file through the standard search path for libraries (see Section 14.1 [How Programs Do Loading], page 195). The XEmacs distribution does not come with this file; sites may provide one for local customizations. If the default init file exists, it is loaded whenever you start Emacs, except in batch mode or if ‘-q’ is specified. But your own personal init file, if any, is loaded first; if it sets inhibit-default-init to a non-nil value, then XEmacs does not subsequently load the ‘default.el’ file.

Another file for site-customization is ‘site-start.el’. Emacs loads this before the user’s init file. You can inhibit the loading of this file with the option ‘-no-site-file’.

Variable site-run-file

This variable specifies the site-customization file to load before the user’s init file. Its normal value is "site-start".

If there is a great deal of code in your ‘.emacs’ file, you should move it into another file named ‘something.el’, byte-compile it (see Chapter 15 [Byte Compilation], page 205), and make your ‘.emacs’ file load the other file using load (see Chapter 14 [Loading], page 195).

See section “Init File Examples” in The XEmacs User’s Manual, for examples of how to make various commonly desired customizations in your ‘.emacs’ file.

User Option inhibit-default-init

This variable prevents XEmacs from loading the default initialization library file for your session of XEmacs. If its value is non-nil, then the default library is not loaded. The default value is nil.

Variable before-init-hook Variable after-init-hook

These two normal hooks are run just before, and just after, loading of the user’s

     init file, ‘default.el’, and/or ‘site-start.el’.

50.1.3 Terminal-Specific Initialization

Each terminal type can have its own Lisp library that XEmacs loads when run on that type of terminal. For a terminal type named termtype, the library is called ‘term/termtype’. XEmacs finds the file by searching the load-path directories as it does for other files, and trying the ‘.elc’ and ‘.el’ suffixes. Normally, terminal-specific Lisp library is located in ‘emacs/lisp/term’, a subdirectory of the ‘emacs/lisp’ directory in which most XEmacs Lisp libraries are kept.

The library’s name is constructed by concatenating the value of the variable termfile-prefix and the terminal type. Normally, term-file-prefix has the value "term/"; changing this is not recommended.

The usual function of a terminal-specific library is to enable special keys to send sequences that XEmacs can recognize. It may also need to set or add to function-key-map if the Termcap entry does not specify all the terminal’s function keys. See Section 50.8 [Terminal Input], page 706.

When the name of the terminal type contains a hyphen, only the part of the name before the first hyphen is significant in choosing the library name. Thus, terminal types ‘aaa-48’ and ‘aaa-30-rv’ both use the ‘term/aaa’ library. If necessary, the library can evaluate (getenv "TERM") to find the full name of the terminal type.

Your ‘.emacs’ file can prevent the loading of the terminal-specific library by setting the variable term-file-prefix to nil. This feature is useful when experimenting with your own peculiar customizations.

You can also arrange to override some of the actions of the terminal-specific library by setting the variable term-setup-hook. This is a normal hook which XEmacs runs using run-hooks at the end of XEmacs initialization, after loading both your ‘.emacs’ file and any terminal-specific libraries. You can use this variable to define initializations for terminals that do not have their own libraries. See Section 26.4 [Hooks], page 378.

Variable term-file-prefix

If the term-file-prefix variable is non-nil, XEmacs loads a terminal-specific initialization file as follows:

     (load (concat term-file-prefix (getenv "TERM")))

You may set the term-file-prefix variable to nil in your ‘.emacs’ file if you do not wish to load the terminal-initialization file. To do this, put the following

in your ‘.emacs’ file: (setq term-file-prefix nil).

Variable term-setup-hook

This variable is a normal hook that XEmacs runs after loading your ‘.emacs’ file, the default initialization file (if any) and the terminal-specific Lisp file.

You can use term-setup-hook to override the definitions made by a terminalspecific file.

Variable window-setup-hook

This variable is a normal hook which XEmacs runs after loading your ‘.emacs’ file and the default initialization file (if any), after loading terminal-specific Lisp code, and after running the hook term-setup-hook.

50.1.4 Command Line Arguments

You can use command line arguments to request various actions when you start XEmacs. Since you do not need to start XEmacs more than once per day, and will often leave your XEmacs session running longer than that, command line arguments are hardly ever used. As a practical matter, it is best to avoid making the habit of using them, since this habit would encourage you to kill and restart XEmacs unnecessarily often. These options exist for two reasons: to be compatible with other editors (for invocation by other programs) and to enable shell scripts to run specific Lisp programs.

This section describes how Emacs processes command line arguments, and how you can customize them.

Function command-line

This function parses the command line that XEmacs was called with, processes it, loads the user’s ‘.emacs’ file and displays the startup messages.

Variable command-line-processed

The value of this variable is t once the command line has been processed.

If you redump XEmacs by calling dump-emacs, you may wish to set this variable to nil first in order to cause the new dumped XEmacs to process its new command line arguments.

Variable command-switch-alist

The value of this variable is an alist of user-defined command-line options and associated handler functions. This variable exists so you can add elements to it.

A command line option is an argument on the command line of the form: -option

The elements of the command-switch-alist look like this:

(option . handler-function) The handler-function is called to handle option and receives the option name as its sole argument. In some cases, the option is followed in the command line by an argument. In these cases, the handler-function can find all the remaining command-line arguments in the variable command-line-args-left. (The entire list of commandline arguments is in command-line-args.) The command line arguments are parsed by the command-line-1 function in the ‘startup.el’ file. See also section “Command Line Switches and Arguments” in The XEmacs User’s Manual.

Variable command-line-args

The value of this variable is the list of command line arguments passed to XEmacs.

Variable command-line-functions

This variable’s value is a list of functions for handling an unrecognized command-line argument. Each time the next argument to be processed has no special meaning, the functions in this list are called, in order of appearance, until one of them returns a non-nil value. These functions are called with no arguments. They can access the commandline argument under consideration through the variable argi. The remaining arguments (not including the current one) are in the variable command-line-

args-left.

When a function recognizes and processes the argument in argi, it should return a non-nil value to say it has dealt with that argument. If it has also dealt with some of the following arguments, it can indicate that by deleting

them from command-line-args-left.

If all of these functions return nil, then the argument is used as a file name to visit.

50.2 Getting out of XEmacs

There are two ways to get out of XEmacs: you can kill the XEmacs job, which exits permanently, or you can suspend it, which permits you to reenter the XEmacs process later. As a practical matter, you seldom kill XEmacs—only when you are about to log out. Suspending is much more common.

50.2.1 Killing XEmacs

Killing XEmacs means ending the execution of the XEmacs process. The parent process normally resumes control. The low-level primitive for killing XEmacs is kill-emacs.

Function kill-emacs &optional exit-data This function exits the XEmacs process and kills it.

If exit-data is an integer, then it is used as the exit status of the XEmacs process. (This is useful primarily in batch operation; see Section 50.11 [Batch Mode], page 711.) If exit-data is a string, its contents are stuffed into the terminal input buffer so that the shell (or whatever program next reads input) can read them.

All the information in the XEmacs process, aside from files that have been saved, is lost when the XEmacs is killed. Because killing XEmacs inadvertently can lose a lot of work, XEmacs queries for confirmation before actually terminating if you have buffers that need saving or subprocesses that are running. This is done in the function save-buffers-kill-

emacs.

Variable kill-emacs-query-functions

After asking the standard questions, save-buffers-kill-emacs calls the functions in the list kill-buffer-query-functions, in order of appearance, with no arguments. These functions can ask for additional confirmation from the user. If any of them returns non-nil, XEmacs is not killed.

Variable kill-emacs-hook

This variable is a normal hook; once save-buffers-kill-emacs is finished with all file saving and confirmation, it runs the functions in this hook.

50.2.2 Suspending XEmacs

Suspending XEmacs means stopping XEmacs temporarily and returning control to its superior process, which is usually the shell. This allows you to resume editing later in the same XEmacs process, with the same buffers, the same kill ring, the same undo history, and so on. To resume XEmacs, use the appropriate command in the parent shell—most likely

fg.

Some operating systems do not support suspension of jobs; on these systems, “suspension” actually creates a new shell temporarily as a subprocess of XEmacs. Then you would exit the shell to return to XEmacs.

Suspension is not useful with window systems such as X, because the XEmacs job may not have a parent that can resume it again, and in any case you can give input to some other job such as a shell merely by moving to a different window. Therefore, suspending is not allowed when XEmacs is an X client.

Function suspend-emacs string

This function stops XEmacs and returns control to the superior process. If and when the superior process resumes XEmacs, suspend-emacs returns nil to its caller in Lisp. If string is non-nil, its characters are sent to be read as terminal input by XEmacs’s superior shell. The characters in string are not echoed by the superior shell; only the results appear. Before suspending, suspend-emacs runs the normal hook suspend-hook. In Emacs version 18, suspend-hook was not a normal hook; its value was a single function, and if its value was non-nil, then suspend-emacs returned immediately without actually suspending anything.

After the user resumes XEmacs, suspend-emacs runs the normal hook suspendresume-hook. See Section 26.4 [Hooks], page 378.

The next redisplay after resumption will redraw the entire screen, unless the variable no-redraw-on-reenter is non-nil (see Section 45.1 [Refresh Screen], page 647).

In the following example, note that ‘pwd’ is not echoed after XEmacs is suspended. But it is read and executed by the shell.

(suspend-emacs)
     ⇒ nil
(add-hook 'suspend-hook
          (function (lambda ()
                      (or (y-or-n-p
                            "Really suspend? ")
                          (error "Suspend cancelled")))))
     ⇒ (lambda nil
          (or (y-or-n-p "Really suspend? ")
              (error "Suspend cancelled")))
(add-hook 'suspend-resume-hook
          (function (lambda () (message "Resumed!"))))
     ⇒ (lambda nil (message "Resumed!"))
(suspend-emacs "pwd")
     ⇒ nil
---------- Buffer: Minibuffer ----------
Really suspend? y
---------- Buffer: Minibuffer ----------
---------- Parent Shell ----------
lewis@slug[23] % /user/lewis/manual
lewis@slug[24] % fg
---------- Echo Area ----------
Resumed!

Variable suspend-hook

This variable is a normal hook run before suspending.

Variable suspend-resume-hook

This variable is a normal hook run after suspending.

50.3 Operating System Environment

XEmacs provides access to variables in the operating system environment through various functions. These variables include the name of the system, the user’s uid, and so on.

Variable system-type

The value of this variable is a symbol indicating the type of operating system XEmacs is operating on. Here is a table of the possible values:

aix-v3  AIX.
berkeley-unix

Berkeley BSD.

dgux Data General DGUX operating system.

gnu A GNU system using the GNU HURD and Mach.

hpux Hewlett-Packard HPUX operating system.

irix Silicon Graphics Irix system.

linux A GNU system using the Linux kernel.

ms-dos Microsoft MS-DOS “operating system.”

next-mach

NeXT Mach-based system.

rtu Masscomp RTU, UCB universe.

unisoft-unix

UniSoft UniPlus.

usg-unix-v

AT&T System V.

vax-vms VAX VMS.
windows-nt

Microsoft windows NT.

xenix SCO Xenix 386.

We do not wish to add new symbols to make finer distinctions unless it is absolutely necessary! In fact, we hope to eliminate some of these alternatives in the future. We recommend using system-configuration to distinguish between different operating systems.

Variable system-configuration

This variable holds the three-part configuration name for the hardware/software configuration of your system, as a string. The convenient way to test parts of this string is with string-match.

Function system-name

This function returns the name of the machine you are running on.

     (system-name)
          ⇒ "prep.ai.mit.edu"

The symbol system-name is a variable as well as a function. In fact, the function returns whatever value the variable system-name currently holds. Thus, you can set the variable system-name in case Emacs is confused about the name of your system. The variable is also useful for constructing frame titles (see Section 32.3 [Frame Titles], page 476).

Variable mail-host-address

If this variable is non-nil, it is used instead of system-name for purposes of generating email addresses. For example, it is used when constructing the default value of user-mail-address. See Section 50.4 [User Identification], page 700. (Since this is done when XEmacs starts up, the value actually used is the one saved when XEmacs was dumped. See Section B.1 [Building XEmacs], page 765.)

Function getenv var

This function returns the value of the environment variable var, as a string. Within XEmacs, the environment variable values are kept in the Lisp variable

process-environment.
     (getenv "USER")
          ⇒ "lewis"
     lewis@slug[10] % printenv
     PATH=.:/user/lewis/bin:/usr/bin:/usr/local/bin
     USER=lewis
     TERM=ibmapa16
     SHELL=/bin/csh
     HOME=/user/lewis

Command setenv variable value This command sets the value of the environment variable named variable to value. Both arguments should be strings. This function works by modifying process-environment; binding that variable with let is also reasonable practice.

Variable process-environment

This variable is a list of strings, each describing one environment variable. The functions getenv and setenv work by means of this variable.

     process-environment
     ⇒ ("l=/usr/stanford/lib/gnuemacs/lisp"
         "PATH=.:/user/lewis/bin:/usr/class:/nfsusr/local/bin"
         "USER=lewis"
         "TERM=ibmapa16"
         "SHELL=/bin/csh"
         "HOME=/user/lewis")

Variable path-separator

This variable holds a string which says which character separates directories in a search path (as found in an environment variable). Its value is ":" for Unix and GNU systems, and ";" for MS-DOS and Windows NT.

Variable invocation-name

This variable holds the program name under which Emacs was invoked. The value is a string, and does not include a directory name.

Variable invocation-directory

This variable holds the directory from which the Emacs executable was invoked, or perhaps nil if that directory cannot be determined.

Variable installation-directory

If non-nil, this is a directory within which to look for the ‘lib-src’ and ‘etc’ subdirectories. This is non-nil when Emacs can’t find those directories in their standard installed locations, but can find them in a directory related somehow to the one containing the Emacs executable.

Function load-average &optional use-floats This function returns a list of the current 1-minute, 5-minute and 15-minute load averages. The values are integers that are 100 times the system load averages. (The load averages indicate the number of processes trying to run.) When use-floats is non-nil, floats will be returned instead of integers. These floats are not multiplied by 100.

     (load-average)
          ⇒ (169 158 164)
     (load-average t)
          ⇒ (1.69921875 1.58984375 1.640625)
     lewis@rocky[5] % uptime
       8:06pm  up 16 day(s), 21:57,  40 users,
      load average: 1.68, 1.59, 1.64

If the 5-minute or 15-minute load averages are not available, return a shortened list, containing only those averages which are available. On some systems, this function may require special privileges to run, or it may be unimplemented for the particular system type. In that case, the function will signal an error.

Function emacs-pid

This function returns the process id of the Emacs process.

Function setprv privilege-name &optional setp getprv This function sets or resets a VMS privilege. (It does not exist on Unix.) The first arg is the privilege name, as a string. The second argument, setp, is t or nil, indicating whether the privilege is to be turned on or off. Its default is nil. The function returns t if successful, nil otherwise. If the third argument, getprv, is non-nil, setprv does not change the privilege, but returns t or nil indicating whether the privilege is currently enabled.

50.4 User Identification

Variable user-mail-address

This holds the nominal email address of the user who is using Emacs. When Emacs starts up, it computes a default value that is usually right, but users often set this themselves when the default value is not right.

Function user-login-name &optional uid

If you don’t specify uid, this function returns the name under which the user is logged in. If the environment variable LOGNAME is set, that value is used. Otherwise, if the environment variable USER is set, that value is used. Otherwise, the value is based on the effective uid, not the real uid.

If you specify uid, the value is the user name that corresponds to uid (which should be an integer).

     (user-login-name)
          ⇒ "lewis"

Function user-real-login-name

This function returns the user name corresponding to Emacs’s real uid. This ignores the effective uid and ignores the environment variables LOGNAME and

USER.

Variable user-full-name

This variable holds the name of the user running this Emacs. It is initialized at startup time from the value of NAME environment variable. You can change the value of this variable to alter the result of the user-full-name function.

Function user-full-name &optional user

This function returns the full name of user. If user is nil, it defaults to the user running this Emacs. In that case, the value of user-full-name variable, if non-nil, will be used.

If user is specified explicitly, user-full-name variable is ignored.

     (user-full-name)
          ⇒ "Hrvoje Niksic"
     (setq user-full-name "Hrvoje \"Niksa\" Niksic")
     (user-full-name)
          ⇒ "Hrvoje \"Niksa\" Niksic"
     (user-full-name "hniksic")
          ⇒ "Hrvoje Niksic"
The symbols user-login-name, user-real-login-name and user-full-name are vari-

ables as well as functions. The functions return the same values that the variables hold. These variables allow you to “fake out” Emacs by telling the functions what to return. The variables are also useful for constructing frame titles (see Section 32.3 [Frame Titles], page 476).

Function user-real-uid

This function returns the real uid of the user.

     (user-real-uid)
          ⇒ 19

Function user-uid

This function returns the effective uid of the user.

Function user-home-directory

This function returns the “HOME” directory of the user, and is intended to replace occurrences of “(getenv "HOME")”. Under Unix systems, the following is done:

1. Return the value of “(getenv "HOME")”, if set.

2. Return “/”, as a fallback, but issue a warning. (Future versions of XEmacs will also attempt to lookup the HOME directory via getpwent(), but this has not yet been implemented.)

Under MS Windows, this is done:

1. Return the value of “(getenv "HOME")”, if set.

2. If the environment variables HOMEDRIVE and HOMEDIR are both set, return the concatenation (the following description uses MS Windows environment variable substitution syntax): %HOMEDRIVE%%HOMEDIR%.

3. Return “C:\”, as a fallback, but issue a warning.

50.5 Time of Day

This section explains how to determine the current time and the time zone.

Function current-time-string &optional time-value

This function returns the current time and date as a humanly-readable string. The format of the string is unvarying; the number of characters used for each part is always the same, so you can reliably use substring to extract pieces of it. It is wise to count the characters from the beginning of the string rather than from the end, as additional information may be added at the end. The argument time-value, if given, specifies a time to format instead of the current time. The argument should be a list whose first two elements are integers. Thus, you can use times obtained from current-time (see below) and from file-attributes (see Section 28.6.4 [File Attributes], page 401).

     (current-time-string)
          ⇒ "Wed Oct 14 22:21:05 1987"

Function current-time

This function returns the system’s time value as a list of three integers: (high low microsec). The integers high and low combine to give the number of seconds since 0:00 January 1, 1970, which is high ∗ 2^{16} + low. The third element, microsec, gives the microseconds since the start of the current second (or 0 for systems that return time only on the resolution of a second). The first two elements can be compared with file time values such as you get with the function file-attributes. See Section 28.6.4 [File Attributes], page 401.

Function current-time-zone &optional time-value This function returns a list describing the time zone that the user is in. The value has the form (offset name). Here offset is an integer giving the number of seconds ahead of UTC (east of Greenwich). A negative value means west of Greenwich. The second element, name is a string giving the name of the time zone. Both elements change when daylight savings time begins or ends; if the user has specified a time zone that does not use a seasonal time adjustment, then the value is constant through time. If the operating system doesn’t supply all the information necessary to compute the value, both elements of the list are nil. The argument time-value, if given, specifies a time to analyze instead of the current time. The argument should be a cons cell containing two integers, or a list whose first two elements are integers. Thus, you can use times obtained from current-time (see above) and from file-attributes (see Section 28.6.4 [File Attributes], page 401).

50.6 Time Conversion

These functions convert time values (lists of two or three integers) to strings or to calendrical information. There is also a function to convert calendrical information to a time value. You can get time values from the functions current-time (see Section 50.5 [Time of Day], page 702) and file-attributes (see Section 28.6.4 [File Attributes], page 401).

Function format-time-string format-string &optional time This function converts time to a string according to format-string. If time is omitted, it defaults to the current time. The argument format-string may contain ‘%’-sequences which say to substitute parts of the time. Here is a table of what the ‘%’-sequences mean:

‘%a’ This stands for the abbreviated name of the day of week.

‘%A’ This stands for the full name of the day of week.

‘%b’ This stands for the abbreviated name of the month.

‘%B’ This stands for the full name of the month.

‘%c’ This is a synonym for ‘%x %X’.

‘%C’ This has a locale-specific meaning. In the default locale (named C), it is equivalent to ‘%A, %B %e, %Y’.

‘%d’ This stands for the day of month, zero-padded.

‘%D’ This is a synonym for ‘%m/%d/%y’.

‘%e’ This stands for the day of month, blank-padded.

‘%h’ This is a synonym for ‘%b’.

‘%H’ This stands for the hour (00-23).

‘%I’ This stands for the hour (00-12).

‘%j’ This stands for the day of the year (001-366).

‘%k’ This stands for the hour (0-23), blank padded.

‘%l’ This stands for the hour (1-12), blank padded.

‘%m’ This stands for the month (01-12). ‘%M’ This stands for the minute (00-59). ‘%n’ This stands for a newline. ‘%p’ This stands for ‘AM’ or ‘PM’, as appropriate. ‘%r’ This is a synonym for ‘%I:%M:%S %p’. ‘%R’ This is a synonym for ‘%H:%M’. ‘%S’ This stands for the seconds (00-60). ‘%t’ This stands for a tab character. ‘%T’ This is a synonym for ‘%H:%M:%S’. ‘%U’ This stands for the week of the year (01-52), assuming that weeks

start on Sunday.

‘%w’ This stands for the numeric day of week (0-6). Sunday is day 0. ‘%W’ This stands for the week of the year (01-52), assuming that weeks

start on Monday.

‘%x’ This has a locale-specific meaning. In the default locale (named

C), it is equivalent to ‘%D’.

‘%X’ This has a locale-specific meaning. In the default locale (named

C), it is equivalent to ‘%T’.

‘%y’ This stands for the year without century (00-99). ‘%Y’ This stands for the year with century. ‘%Z’ This stands for the time zone abbreviation.

Function decode-time time

This function converts a time value into calendrical information. The return value is a list of nine elements, as follows:

(seconds minutes hour day month year dow dst zone) Here is what the elements mean: sec The number of seconds past the minute, as an integer between 0

and 59.

minute The number of minutes past the hour, as an integer between 0 and

59.

hour The hour of the day, as an integer between 0 and 23. day The day of the month, as an integer between 1 and 31. month The month of the year, as an integer between 1 and 12. year The year, an integer typically greater than 1900. dow The day of week, as an integer between 0 and 6, where 0 stands for

Sunday. dst t if daylight savings time is effect, otherwise nil.

zone An integer indicating the time zone, as the number of seconds east of Greenwich.

Note that Common Lisp has different meanings for dow and zone.

Function encode-time seconds minutes hour day month year &optional

zone This function is the inverse of decode-time. It converts seven items of calendrical data into a time value. For the meanings of the arguments, see the table

above under decode-time.

Year numbers less than 100 are treated just like other year numbers. If you want them to stand for years above 1900, you must alter them yourself before

you call encode-time.

The optional argument zone defaults to the current time zone and its daylight savings time rules. If specified, it can be either a list (as you would get from current-time-zone) or an integer (as you would get from decode-time). The specified zone is used without any further alteration for daylight savings time.

50.7 Timers for Delayed Execution

You can set up a timer to call a function at a specified future time.

Function add-timeout secs function object &optional resignal This function adds a timeout, to be signaled after the timeout period has elapsed. secs is a number of seconds, expressed as an integer or a float. function will be called after that many seconds have elapsed, with one argument, the given object. If the optional resignal argument is provided, then after this timeout expires, ‘add-timeout’ will automatically be called again with resignal as the first argument. This function returns an object which is the id of this particular timeout. You can pass that object to disable-timeout to turn off the timeout before it has been signalled. The number of seconds may be expressed as a floating-point number, in which case some fractional part of a second will be used. Caveat: the usable timeout granularity will vary from system to system. Adding a timeout causes a timeout event to be returned by next-event, and the function will be invoked by dispatch-event, so if XEmacs is in a tight loop, the function will not be invoked until the next call to sit-for or until the return to top-level (the same is true of process filters). WARNING: if you are thinking of calling add-timeout from inside of a callback function as a way of resignalling a timeout, think again. There is a race condition. That’s why the resignal argument exists. (NOTE: In FSF Emacs, this function is called run-at-time and has different semantics.)

Function disable-timeout id

Cancel the requested action for id, which should be a value previously returned by add-timeout. This cancels the effect of that call to add-timeout; the arrival of the specified time will not cause anything special to happen. (NOTE: In FSF Emacs, this function is called cancel-timer.)

50.8 Terminal Input

This section describes functions and variables for recording or manipulating terminal input. See Chapter 45 [Display], page 647, for related functions.

50.8.1 Input Modes

Function set-input-mode interrupt flow meta quit-char This function sets the mode for reading keyboard input. If interrupt is non-null, then XEmacs uses input interrupts. If it is nil, then it uses cbreak mode. When XEmacs communicates directly with X, it ignores this argument and uses interrupts if that is the way it knows how to communicate. If flow is non-nil, then XEmacs uses xon/xoff (C-q, C-s) flow control for output to the terminal. This has no effect except in cbreak mode. See Section 50.10 [Flow Control], page 710. The default setting is system dependent. Some systems always use cbreak mode regardless of what is specified. The argument meta controls support for input character codes above 127. If meta is t, XEmacs converts characters with the 8th bit set into Meta characters. If meta is nil, XEmacs disregards the 8th bit; this is necessary when the terminal uses it as a parity bit. If meta is neither t nor nil, XEmacs uses all 8 bits of input unchanged. This is good for terminals using European 8-bit character sets. If quit-char is non-nil, it specifies the character to use for quitting. Normally this character is C-g. See Section 19.8 [Quitting], page 306.

The current-input-mode function returns the input mode settings XEmacs is currently using.

Function current-input-mode

This function returns current mode for reading keyboard input. It returns a list, corresponding to the arguments of set-input-mode, of the form (interrupt flow meta quit) in which:

interrupt is non-nil when XEmacs is using interrupt-driven input. If nil, Emacs is using cbreak mode.

flow is non-nil if XEmacs uses xon/xoff (C-q, C-s) flow control for output to the terminal. This value has no effect unless interrupt is non-nil. meta is t if XEmacs treats the eighth bit of input characters as the

meta bit; nil means XEmacs clears the eighth bit of every input character; any other value means XEmacs uses all eight bits as the basic character code.

quit is the character XEmacs currently uses for quitting, usually C-g.

50.8.2 Translating Input Events

This section describes features for translating input events into other input events before they become part of key sequences.

Variable function-key-map

This variable holds a keymap that describes the character sequences sent by function keys on an ordinary character terminal. This keymap uses the same data structure as other keymaps, but is used differently: it specifies translations to make while reading events. If function-key-map “binds” a key sequence k to a vector v, then when k appears as a subsequence anywhere in a key sequence, it is replaced with the events in v. For example, VT100 terminals send ^{⟨}\underline{\overline{ESC}}^{⟩} O P when the keypad PF1 key is pressed. Therefore, we want XEmacs to translate that sequence of events into the single event pf1. We accomplish this by “binding” ^{⟨}\underline{\overline{ESC}}^{⟩} O P to [pf1] in function-key-map, when using a VT100. Thus, typing C-c ^{⟨}\underline{\overline{PF1}}^{⟩} sends the character sequence C-c ^{⟨}\underline{\overline{ESC}}^{⟩} O P; later the function read-key-sequence translates this back into C-c ^{⟨}\underline{\overline{PF1}}^{⟩}, which it returns as the vector [?\C-c pf1]. Entries in function-key-map are ignored if they conflict with bindings made in the minor mode, local, or global keymaps. The intent is that the character sequences that function keys send should not have command bindings in their own right. The value of function-key-map is usually set up automatically according to the terminal’s Terminfo or Termcap entry, but sometimes those need help from terminal-specific Lisp files. XEmacs comes with terminal-specific files for many common terminals; their main purpose is to make entries in function-key-map beyond those that can be deduced from Termcap and Terminfo. See Section 50.1.3 [Terminal-Specific], page 693. Emacs versions 18 and earlier used totally different means of detecting the character sequences that represent function keys.

Variable key-translation-map

This variable is another keymap used just like function-key-map to translate input events into other events. It differs from function-key-map in two ways:

  • key-translation-map goes to work after function-key-map is finished;

it receives the results of translation by function-key-map.

• key-translation-map overrides actual key bindings.

The intent of key-translation-map is for users to map one character set to another, including ordinary characters normally bound to self-insert-command.

You can use function-key-map or key-translation-map for more than simple aliases, by using a function, instead of a key sequence, as the “translation” of a key. Then this function is called to compute the translation of that key.

The key translation function receives one argument, which is the prompt that was specified in read-key-sequence—or nil if the key sequence is being read by the editor command loop. In most cases you can ignore the prompt value.

If the function reads input itself, it can have the effect of altering the event that follows. For example, here’s how to define C-c h to turn the character that follows into a Hyper character:

(defun hyperify (prompt)
  (let ((e (read-event)))
    (vector (if (numberp e)
                (logior (lsh 1 20) e)
              (if (memq 'hyper (event-modifiers e))
                  e
                (add-event-modifier "H-" e))))))
(defun add-event-modifier (string e)
  (let ((symbol (if (symbolp e) e (car e))))
    (setq symbol (intern (concat string
                                 (symbol-name symbol))))
    (if (symbolp e)
        symbol
      (cons symbol (cdr e)))))
(define-key function-key-map "\C-ch" 'hyperify)

The ‘iso-transl’ library uses this feature to provide a way of inputting non-ASCII Latin-1 characters.

50.8.3 Recording Input

Function recent-keys &optional number This function returns a vector containing recent input events from the keyboard or mouse. By default, 100 events are recorded, which is how many recent-keys returns.

All input events are included, whether or not they were used as parts of key sequences. Thus, you always get the last 100 inputs, not counting keyboard macros. (Events from keyboard macros are excluded because they are less interesting for debugging; it should be enough to see the events that invoked the macros.)

If number is specified, not more than number events will be returned. You may change the number of stored events using set-recent-keys-ring-size.

Function recent-keys-ring-size

This function returns the number of recent events stored internally. This is also the maximum number of events recent-keys can return. By default, 100 events are stored.

Function set-recent-keys-ring-size size

This function changes the number of events stored by XEmacs and returned by

recent-keys.

For example, (set-recent-keys-ring-size 250) will make XEmacs remember last 250 events and will make recent-keys return last 250 events by default.

Command open-dribble-file filename

This function opens a dribble file named filename. When a dribble file is open, each input event from the keyboard or mouse (but not those from keyboard macros) is written in that file. A non-character event is expressed using its printed representation surrounded by ‘<...>’. You close the dribble file by calling this function with an argument of nil. This function is normally used to record the input necessary to trigger an XEmacs bug, for the sake of a bug report.

     (open-dribble-file "~/dribble")
          ⇒ nil

See also the open-termscript function (see Section 50.9 [Terminal Output], page 709).

50.9 Terminal Output

The terminal output functions send output to the terminal or keep track of output sent to the terminal. The function device-baud-rate tells you what XEmacs thinks is the output speed of the terminal.

Function device-baud-rate &optional device

This function’s value is the output speed of the terminal associated with device, as far as XEmacs knows. device defaults to the selected device (usually the only device) if omitted. Changing this value does not change the speed of actual data transmission, but the value is used for calculations such as padding. This value has no effect for window-system devices. (This is different in FSF Emacs, where the baud rate also affects decisions about whether to scroll part of the screen or repaint, even when using a window system.) The value is measured in bits per second.

XEmacs attempts to automatically initialize the baud rate by querying the terminal. If you are running across a network, however, and different parts of the network work are at different baud rates, the value returned by XEmacs may be different from the value used by your local terminal. Some network protocols communicate the local terminal speed to the remote machine, so that XEmacs and other programs can get the proper value, but others do not. If XEmacs has the wrong value, it makes decisions that are less than optimal. To

fix the problem, use set-device-baud-rate.

Function set-device-baud-rate &optional device

This function sets the output speed of device. See device-baud-rate. device defaults to the selected device (usually the only device) if omitted.

Function send-string-to-terminal char-or-string &optional stdout-p

device This function sends char-or-string to the terminal without alteration. Control characters in char-or-string have terminal-dependent effects. If device is nil, this function writes to XEmacs’s stderr, or to stdout if stdout-p is non-nil. Otherwise, device should be a tty or stream device, and the function writes to the device’s normal or error output, according to stdout-p. One use of this function is to define function keys on terminals that have downloadable function key definitions. For example, this is how on certain terminals to define function key 4 to move forward four characters (by transmitting the characters C-u C-f to the computer):

     (send-string-to-terminal "\eF4\^U\^F")
          ⇒ nil

Command open-termscript filename

This function is used to open a termscript file that will record all the characters sent by XEmacs to the terminal. (If there are multiple tty or stream devices, all characters sent to all such devices are recorded.) The function returns nil. Termscript files are useful for investigating problems where XEmacs garbles the screen, problems that are due to incorrect Termcap entries or to undesirable settings of terminal options more often than to actual XEmacs bugs. Once you are certain which characters were actually output, you can determine reliably whether they correspond to the Termcap specifications in use. A nil value for filename stops recording terminal output. See also open-dribble-file in Section 50.8 [Terminal Input], page 706.

     (open-termscript "../junk/termscript")
          ⇒ nil

50.10 Flow Control

This section attempts to answer the question “Why does XEmacs choose to use flowcontrol characters in its command character set?” For a second view on this issue, read the comments on flow control in the ‘emacs/INSTALL’ file from the distribution; for help with Termcap entries and DEC terminal concentrators, see ‘emacs/etc/TERMS’.

At one time, most terminals did not need flow control, and none used C-s and C-q for flow control. Therefore, the choice of C-s and C-q as command characters was uncontroversial. XEmacs, for economy of keystrokes and portability, used nearly all the ASCII control characters, with mnemonic meanings when possible; thus, C-s for search and C-q for quote.

Later, some terminals were introduced which required these characters for flow control. They were not very good terminals for full-screen editing, so XEmacs maintainers did not pay attention. In later years, flow control with C-s and C-q became widespread among terminals, but by this time it was usually an option. And the majority of users, who can turn flow control off, were unwilling to switch to less mnemonic key bindings for the sake of flow control.

So which usage is “right”, XEmacs’s or that of some terminal and concentrator manufacturers? This question has no simple answer.

One reason why we are reluctant to cater to the problems caused by C-s and C-q is that they are gratuitous. There are other techniques (albeit less common in practice) for flow control that preserve transparency of the character stream. Note also that their use for flow control is not an official standard. Interestingly, on the model 33 teletype with a paper tape punch (which is very old), C-s and C-q were sent by the computer to turn the punch on and off!

As X servers and other window systems replace character-only terminals, this problem is gradually being cured. For the mean time, XEmacs provides a convenient way of enabling flow control if you want it: call the function enable-flow-control.

Function enable-flow-control

This function enables use of C-s and C-q for output flow control, and provides the characters C-\ and C-^as aliases for them using keyboard-translate-table (see Section 50.8.2 [Translating Input], page 707).

You can use the function enable-flow-control-on in your ‘.emacs’ file to enable flow control automatically on certain terminal types.

Function enable-flow-control-on &rest termtypes

This function enables flow control, and the aliases C-\ and C-^, if the terminal type is one of termtypes. For example:

          (enable-flow-control-on "vt200" "vt300" "vt101" "vt131")

Here is how enable-flow-control does its job:

1. It sets cbreak mode for terminal input, and tells the operating system to handle flow

    control, with (set-input-mode nil t).
 2. It sets up keyboard-translate-table to translate C-\ and C-^into C-s and C-q.

Except at its very lowest level, XEmacs never knows that the characters typed were anything but C-s and C-q, so you can in effect type them as C-\ and C-^even when they are input for other commands. See Section 50.8.2 [Translating Input], page 707.

If the terminal is the source of the flow control characters, then once you enable kernel flow control handling, you probably can make do with less padding than normal for that terminal. You can reduce the amount of padding by customizing the Termcap entry. You can also reduce it by setting baud-rate to a smaller value so that XEmacs uses a smaller speed when calculating the padding needed. See Section 50.9 [Terminal Output], page 709.

50.11 Batch Mode

The command line option ‘-batch’ causes XEmacs to run noninteractively. In this mode, XEmacs does not read commands from the terminal, it does not alter the terminal modes, and it does not expect to be outputting to an erasable screen. The idea is that you specify Lisp programs to run; when they are finished, XEmacs should exit. The way to specify the programs to run is with ‘-l file’, which loads the library named file, and ‘-f function’, which calls function with no arguments.

Any Lisp program output that would normally go to the echo area, either using message or using prin1, etc., with t as the stream, goes instead to XEmacs’s standard error descriptor when in batch mode. Thus, XEmacs behaves much like a noninteractive application program. (The echo area output that XEmacs itself normally generates, such as command echoing, is suppressed entirely.)

Function noninteractive

This function returns non-nil when XEmacs is running in batch mode.

Variable noninteractive

This variable is non-nil when XEmacs is running in batch mode. Setting this variable to nil, however, will not change whether XEmacs is running in batch mode, and will not change the return value of the noninteractive function.

51 Functions Specific to the X Window System

XEmacs provides the concept of devices, which generalizes connections to an X server, a TTY device, etc. Most information about an X server that XEmacs is connected to can be determined through general console and device functions. See Chapter 33 [Consoles and Devices], page 483. However, there are some features of the X Window System that do not generalize well, and they are covered specially here.

51.1 X Selections

The X server records a set of selections which permit transfer of data between application programs. The various selections are distinguished by selection types, represented in XEmacs by symbols. X clients including XEmacs can read or set the selection for any given type.

Function x-own-selection data &optional type This function sets a “selection” in the X server. It takes two arguments: a value, data, and the selection type type to assign it to. data may be a string, a cons of two markers, or an extent. In the latter cases, the selection is considered to be the text between the markers, or between the extent’s endpoints. Each possible type has its own selection value, which changes independently. The usual values of type are PRIMARY and SECONDARY; these are symbols with upper-case names, in accord with X Windows conventions. The default is

PRIMARY.

(In FSF Emacs, this function is called x-set-selection and takes different arguments.)

Function x-get-selection

This function accesses selections set up by XEmacs or by other X clients. It returns the value of the current primary selection.

Function x-disown-selection &optional secondary-p Assuming we own the selection, this function disowns it. If secondary-p is non-nil, the secondary selection instead of the primary selection is discarded.

The X server also has a set of numbered cut buffers which can store text or other data being moved between applications. Cut buffers are considered obsolete, but XEmacs supports them for the sake of X clients that still use them.

Function x-get-cutbuffer &optional n

This function returns the contents of cut buffer number n. (This function is called x-get-cut-buffer in FSF Emacs.)

Function x-store-cutbuffer string

This function stores string into the first cut buffer (cut buffer 0), moving the other values down through the series of cut buffers, kill-ring-style. (This function is called x-set-cut-buffer in FSF Emacs.)

51.2 X Server

This section describes how to access and change the overall status of the X server XEmacs is using.

51.2.1 Resources

Function default-x-device

This function return the default X device for resourcing. This is the first-created X device that still exists.

Function x-get-resource name class type &optional locale device noerror

This function retrieves a resource value from the X resource manager. • The first arg is the name of the resource to retrieve, such as ‘"font"’.

• The second arg is the class of the resource to retrieve, like ‘"Font"’.

• The third arg should be one of the symbols string, integer, natnum, or boolean, specifying the type of object that the database is searched for.

• The fourth arg is the locale to search for the resources on, and can currently be a a buffer, a frame, a device, or the symbol global. If omitted, it defaults to global.

• The fifth arg is the device to search for the resources on. (The resource database for a particular device is constructed by combining non-device-specific resources such any command-line resources specified and any appdefaults files found [or the fallback resources supplied by XEmacs, if no appdefaults file is found] with device-specific resources such as those supplied using ‘xrdb’.) If omitted, it defaults to the device of locale, if a device can be derived (i.e. if locale is a frame or device), and otherwise defaults to

the value of default-x-device.

• The sixth arg noerror, if non-nil, means do not signal an error if a bogus resource specification was retrieved (e.g. if a non-integer was given when an integer was requested). In this case, a warning is issued instead.

The resource names passed to this function are looked up relative to the locale. If you want to search for a subresource, you just need to specify the resource levels in name and class. For example, name could be ‘"modeline.attributeFont"’,

and class ‘"Face.AttributeFont"’.

Specifically,

1. If locale is a buffer, a call

         (x-get-resource "foreground" "Foreground" 'string some-

buffer) is an interface to a C call something like

         XrmGetResource (db, "xemacs.buffer.buffer-name.foreground",
     "Emacs.EmacsLocaleType.EmacsBuffer.Foreground",
     "String");

2. If locale is a frame, a call

             (x-get-resource "foreground" "Foreground" 'string some-

frame) is an interface to a C call something like

             XrmGetResource (db, "xemacs.frame.frame-name.foreground",
         "Emacs.EmacsLocaleType.EmacsFrame.Foreground",
         "String");

3. If locale is a device, a call

             (x-get-resource "foreground" "Foreground" 'string some-

device) is an interface to a C call something like

             XrmGetResource (db, "xemacs.device.device-name.foreground",
         "Emacs.EmacsLocaleType.EmacsDevice.Foreground",
         "String");

4. If locale is the symbol global, a call

             (x-get-resource "foreground" "Foreground" 'string 'global)

is an interface to a C call something like

             XrmGetResource (db, "xemacs.foreground",
         "Emacs.Foreground",
         "String");

Note that for global, no prefix is added other than that of the application itself; thus, you can use this locale to retrieve arbitrary application resources, if you really want to.

The returned value of this function is nil if the queried resource is not found. If type is string, a string is returned, and if it is integer, an integer is returned. If type is boolean, then the returned value is the list (t) for true, (nil) for false, and is nil to mean “unspecified”.

Function x-put-resource resource-line &optional device This function adds a resource to the resource database for device. resource-line specifies the resource to add and should be a standard resource specification.

Variable x-emacs-application-class

This variable holds The X application class of the XEmacs process. This controls, among other things, the name of the “app-defaults” file that XEmacs will use. For changes to this variable to take effect, they must be made before the connection to the X server is initialized, that is, this variable may only be changed before XEmacs is dumped, or by setting it in the file

‘lisp/term/x-win.el’.

By default, this variable is nil at startup. When the connection to the X server is first initialized, the X resource database will be consulted and the value will be set according to whether any resources are found for the application class “XEmacs”.

51.2.2 Data about the X Server

This section describes functions and a variable that you can use to get information about the capabilities and origin of the X server corresponding to a particular device. The device argument is generally optional and defaults to the selected device.

Function x-server-version &optional device

This function returns the list of version numbers of the X server device is on. The returned value is a list of three integers: the major and minor version numbers of the X protocol in use, and the vendor-specific release number.

Function x-server-vendor &optional device

This function returns the vendor supporting the X server device is on.

Function x-display-visual-class &optional device

This function returns the visual class of the display device is on. The value is

one of the symbols static-gray, gray-scale, static-color, pseudo-color,

true-color, and direct-color. (Note that this is different from previous versions of XEmacs, which returned StaticGray, GrayScale, etc.)

51.2.3 Restricting Access to the Server by Other Apps

Function x-grab-keyboard &optional device

This function grabs the keyboard on the given device (defaulting to the selected one). So long as the keyboard is grabbed, all keyboard events will be delivered to XEmacs – it is not possible for other X clients to eavesdrop on them. Ungrab

the keyboard with x-ungrab-keyboard (use an unwind-protect). Returns t

if the grab was successful; nil otherwise.

Function x-ungrab-keyboard &optional device

This function releases a keyboard grab made with x-grab-keyboard.

Function x-grab-pointer &optional device cursor ignore-keyboard This function grabs the pointer and restricts it to its current window. If optional device argument is nil, the selected device will be used. If optional cursor argument is non-nil, change the pointer shape to that until x-ungrab-pointer is called (it should be an object returned by the make-cursor function). If the second optional argument ignore-keyboard is non-nil, ignore all keyboard events during the grab. Returns t if the grab is successful, nil otherwise.

Function x-ungrab-pointer &optional device

This function releases a pointer grab made with x-grab-pointer. If optional first arg device is nil the selected device is used. If it is t the pointer will be released on all X devices.

51.3 Miscellaneous X Functions and Variables

Variable x-bitmap-file-path

This variable holds a list of the directories in which X bitmap files may be found. If nil, this is initialized from the ‘"*bitmapFilePath"’ resource. This is used by the make-image-instance function (however, note that if the environment variable ‘XBMLANGPATH’ is set, it is consulted first).

Variable x-library-search-path

This variable holds the search path used by read-color to find ‘rgb.txt’.

Function x-valid-keysym-name-p keysym

This function returns true if keysym names a keysym that the X library knows about. Valid keysyms are listed in the files ‘/usr/include/X11/keysymdef.h’ and in ‘/usr/lib/X11/XKeysymDB’, or whatever the equivalents are on your system.

Function x-window-id &optional frame This function returns the ID of the X11 window. This gives us a chance to manipulate the Emacs window from within a different program. Since the ID is an unsigned long, we return it as a string.

Variable x-allow-sendevents

If non-nil, synthetic events are allowed. nil means they are ignored. Beware: allowing XEmacs to process SendEvents opens a big security hole.

Function x-debug-mode arg &optional device With a true arg, make the connection to the X server synchronous. With false, make it asynchronous. Synchronous connections are much slower, but are useful for debugging. (If you get X errors, make the connection synchronous, and use a debugger to set a breakpoint on x_error_handler. Your backtrace of the C stack will now be useful. In asynchronous mode, the stack above x_ error_handler isn’t helpful because of buffering.) If device is not specified, the selected device is assumed. Calling this function is the same as calling the C function XSynchronize, or starting the program with the ‘-sync’ command line argument.

Variable x-debug-events

If non-zero, debug information about events that XEmacs sees is displayed. Information is displayed on stderr. Currently defined values are:

• 1 == non-verbose output

• 2 == verbose output 718 XEmacs Lisp Reference Manual

52 ToolTalk Support

52.1 XEmacs ToolTalk API Summary

The XEmacs Lisp interface to ToolTalk is similar, at least in spirit, to the standard C ToolTalk API. Only the message and pattern parts of the API are supported at present; more of the API could be added if needed. The Lisp interface departs from the C API in a few ways:

• ToolTalk is initialized automatically at XEmacs startup-time. Messages can only be sent other ToolTalk applications connected to the same X11 server that XEmacs is running on.

• There are fewer entry points; polymorphic functions with keyword arguments are used instead.

• The callback interface is simpler and marginally less functional. A single callback may be associated with a message or a pattern; the callback is specified with a Lisp symbol (the symbol should have a function binding).

• The session attribute for messages and patterns is always initialized to the default session.

• Anywhere a ToolTalk enum constant, e.g. ‘TT_SESSION’, is valid, one can substitute the corresponding symbol, e.g. 'TT_SESSION. This simplifies building lists that represent messages and patterns.

52.2 Sending Messages

52.2.1 Example of Sending Messages

Here’s a simple example that sends a query to another application and then displays its reply. Both the query and the reply are stored in the first argument of the message.

 (defun tooltalk-random-query-handler (msg)
   (let ((state (get-tooltalk-message-attribute msg 'state)))
     (cond
       ((eq state 'TT_HANDLED)
        (message (get-tooltalk-message-attribute msg arg_val 0)))
       ((memq state '(TT_FAILED TT_REJECTED))
        (message "Random query turns up nothing")))))
 (defvar random-query-message
   '(   class TT_REQUEST
        scope TT_SESSION
      address TT_PROCEDURE
           op "random-query"
         args '((TT_INOUT "?" "string"))
     callback tooltalk-random-query-handler))

(let ((m (make-tooltalk-message random-query-message)))
  (send-tooltalk-message m))

52.2.2 Elisp Interface for Sending Messages

Function make-tooltalk-message attributes

Create a ToolTalk message and initialize its attributes. The value of attributes must be a list of alternating keyword/values, where keywords are symbols that name valid message attributes. For example:

       (make-tooltalk-message
         '(class TT_NOTICE
           scope TT_SESSION
           address TT_PROCEDURE
           op "do-something"
           args ("arg1" 12345 (TT_INOUT "arg3" "string"))))

Values must always be strings, integers, or symbols that represent ToolTalk constants. Attribute names are the same as those supported by set-tooltalk-

message-attribute, plus args.

The value of args should be a list of message arguments where each message argument has the following form:

     ‘(mode [value [type]])’ or just ‘value’

Where mode is one of TT_IN, TT_OUT, or TT_INOUT and type is a string. If type isn’t specified then int is used if value is a number; otherwise string is used. If type is string then value is converted to a string (if it isn’t a string already) with prin1-to-string. If only a value is specified then mode defaults to TT_ IN. If mode is TT_OUT then value and type don’t need to be specified. You can find out more about the semantics and uses of ToolTalk message arguments in chapter 4 of the ToolTalk Programmer’s Guide.

Function send-tooltalk-message msg

Send the message on its way. Once the message has been sent it’s almost always a good idea to get rid of it with destroy-tooltalk-message.

Function return-tooltalk-message msg &optional mode

Send a reply to this message. The second argument can be reply, reject or fail; the default is reply. Before sending a reply, all message arguments whose mode is TT_INOUT or TT_OUT should have been filled in – see set-tooltalk-

message-attribute.

Function get-tooltalk-message-attribute msg attribute &optional argn

Returns the indicated ToolTalk message attribute. Attributes are identified by symbols with the same name (underscores and all) as the suffix of the ToolTalk ‘tt_message_<attribute>’ function that extracts the value. String attribute values are copied and enumerated type values (except disposition) are converted to symbols; e.g. ‘TT_HANDLER’ is 'TT_HANDLER, ‘uid’ and ‘gid’ are represented by fixnums (small integers), ‘opnum’ is converted to a string, and ‘disposition’ is converted to a fixnum. We convert ‘opnum’ (a C int) to a string (e.g. 123 ⇒ "123") because there’s no guarantee that opnums will fit within the range of XEmacs Lisp integers.

[TBD] Use the plist attribute instead of C API user attribute for user-defined message data. To retrieve the value of a message property, specify the indicator for argn. For example, to get the value of a property called rflag, use

        (get-tooltalk-message-attribute msg 'plist 'rflag)

To get the value of a message argument use one of the arg_val (strings), arg_ ival (integers), or arg_bval (strings with embedded nulls), attributes. For example, to get the integer value of the third argument:

        (get-tooltalk-message-attribute msg 'arg_ival 2)

As you can see, argument numbers are zero-based. The type of each arguments can be retrieved with the arg_type attribute; however ToolTalk doesn’t define any semantics for the string value of arg_type. Conventionally string is used for strings and int for 32 bit integers. Note that XEmacs Lisp stores the lengths of strings explicitly (unlike C) so treating the value returned by arg_bval like a string is fine.

Function set-tooltalk-message-attribute value msg attribute &optional

argn Initialize one ToolTalk message attribute. Attribute names and values are the same as for get-tooltalk-message-attribute. A property list is provided for user data (instead of the user

message attribute); see get-tooltalk-message-attribute.

Callbacks are handled slightly differently than in the C ToolTalk API. The value of callback should be the name of a function of one argument. It will be called each time the state of the message changes. This is usually used to notice when the message’s state has changed to TT_HANDLED (or TT_FAILED), so that reply argument values can be used. If one of the argument attributes is specified as arg_val, arg_ival, or arg_ bval, then argn must be the number of an already created argument. Arguments can be added to a message with add-tooltalk-message-arg.

Function add-tooltalk-message-arg msg mode type &optional value Append one new argument to the message. mode must be one of TT_IN, TT_ INOUT, or TT_OUT, type must be a string, and value can be a string or an integer. ToolTalk doesn’t define any semantics for type, so only the participants in the protocol you’re using need to agree what types mean (if anything). Conventionally string is used for strings and int for 32 bit integers. Arguments can initialized by providing a value or with set-tooltalk-message-attribute; the latter is necessary if you want to initialize the argument with a string that can contain embedded nulls (use arg_bval).

Function create-tooltalk-message

Create a new ToolTalk message. The message’s session attribute is initialized to the default session. Other attributes can be initialized with set-tooltalk-

message-attribute. make-tooltalk-message is the preferred way to create

and initialize a message.

Function destroy-tooltalk-message msg

Apply ‘tt_message_destroy’ to the message. It’s not necessary to destroy messages after they’ve been processed by a message or pattern callback, the Lisp/ToolTalk callback machinery does this for you.

52.3 Receiving Messages

52.3.1 Example of Receiving Messages

Here’s a simple example of a handler for a message that tells XEmacs to display a string in the mini-buffer area. The message operation is called ‘emacs-display-string’. Its first (0th) argument is the string to display.

(defun tooltalk-display-string-handler (msg)
  (message (get-tooltalk-message-attribute msg 'arg_val 0)))
(defvar display-string-pattern
  '(category TT_HANDLE
       scope TT_SESSION
          op "emacs-display-string"
    callback tooltalk-display-string-handler))
(let ((p (make-tooltalk-pattern display-string-pattern)))
  (register-tooltalk-pattern p))

52.3.2 Elisp Interface for Receiving Messages

Function make-tooltalk-pattern attributes

Create a ToolTalk pattern and initialize its attributes. The value of attributes must be a list of alternating keyword/values, where keywords are symbols that name valid pattern attributes or lists of valid attributes. For example:

       (make-tooltalk-pattern
         '(category TT_OBSERVE
              scope TT_SESSION
                 op ("operation1" "operation2")
               args ("arg1" 12345 (TT_INOUT "arg3" "string"))))

Attribute names are the same as those supported by add-tooltalk-pattern-

attribute, plus 'args.

Values must always be strings, integers, or symbols that represent ToolTalk constants or lists of same. When a list of values is provided all of the list elements are added to the attribute. In the example above, messages whose ‘op’ attribute is ‘"operation1"’ or ‘"operation2"’ would match the pattern. The value of args should be a list of pattern arguments where each pattern argument has the following form:

     ‘(mode [value [type]])’ or just ‘value’

Where mode is one of TT_IN, TT_OUT, or TT_INOUT and type is a string. If type isn’t specified then int is used if value is a number; otherwise string is used. If type is string then value is converted to a string (if it isn’t a string already) with prin1-to-string. If only a value is specified then mode defaults to TT_ IN. If mode is TT_OUT then value and type don’t need to be specified. You can find out more about the semantics and uses of ToolTalk pattern arguments in chapter 3 of the ToolTalk Programmer’s Guide.

Function register-tooltalk-pattern pat

XEmacs will begin receiving messages that match this pattern.

Function unregister-tooltalk-pattern pat

XEmacs will stop receiving messages that match this pattern.

Function add-tooltalk-pattern-attribute value pat indicator

Add one value to the indicated pattern attribute. The names of attributes are the same as the ToolTalk accessors used to set them less the ‘tooltalk_pattern_’ prefix and the ‘_add’ suffix. For example, the name of the attribute for the

‘tt_pattern_disposition_add’ attribute is disposition. The category at-

tribute is handled specially, since a pattern can only be a member of one

category (TT_OBSERVE or TT_HANDLE).

Callbacks are handled slightly differently than in the C ToolTalk API. The value of callback should be the name of a function of one argument. It will be called each time the pattern matches an incoming message.

Function add-tooltalk-pattern-arg pat mode type value

Add one fully-specified argument to a ToolTalk pattern. mode must be one of TT_IN, TT_INOUT, or TT_OUT. type must be a string. value can be an integer, string or nil. If value is an integer then an integer argument (‘tt_pattern_iarg_add’) is added; otherwise a string argument is added. At present there’s no way to add a binary data argument.

Function create-tooltalk-pattern

Create a new ToolTalk pattern and initialize its session attribute to be the default session.

Function destroy-tooltalk-pattern pat

Apply ‘tt_pattern_destroy’ to the pattern. This effectively unregisters the pattern.

Function describe-tooltalk-message msg &optional stream

Print the message’s attributes and arguments to stream. This is often useful for debugging.

724 XEmacs Lisp Reference Manual

53 LDAP Support

XEmacs can be linked with a LDAP client library to provide Elisp primitives to access directory servers using the Lightweight Directory Access Protocol.

53.1 Building XEmacs with LDAP support

LDAP support must be added to XEmacs at build time since it requires linking to an external LDAP client library. As of 21.0, XEmacs has been successfully built and tested with

  • University of Michigan’s LDAP 3.3 (http://www.umich.edu/~dirsvcs/ldap/)
  • LDAP SDK 1.0 from Netscape Corp. (http://developer.netscape.com/)

Other libraries conforming to RFC 1823 will probably work also but may require some minor tweaking at C level.

The standard XEmacs configure script autodetects an installed LDAP library provided the library itself and the corresponding header files can be found in the library and include paths. A successful detection will be signalled in the final output of the configure script.

53.2 XEmacs LDAP API

XEmacs LDAP API consists of two layers: a low-level layer which tries to stay as close as possible to the C API (where practical) and a higher-level layer which provides more convenient primitives to effectively use LDAP.

As of XEmacs 21.0, only interfaces to basic LDAP search functions are provided, broader support is planned in future versions.

53.2.1 LDAP Variables

Variable ldap-default-host

The default LDAP server

Variable ldap-default-port

Default TCP port for LDAP connections. Initialized from the LDAP library.

Default value is 389.

Variable ldap-default-base

Default base for LDAP searches. This is a string using the syntax of RFC 1779. For instance, "oME, c" limits the search to the Acme organization in the United States.

Variable ldap-host-parameters-alist

An alist of per host options for LDAP transactions. The list elements look like (HOST PROP1 VAL1 PROP2 VAL2 ...) host is the name of an LDAP server. propn and valn are property/value pairs describing parameters for the server. Valid properties: binddn The distinguished name of the user to bind as. This may look like

‘c, ome, cnnny Bugs’, see RFC 1779 for details.

passwd The password to use for authentication.

auth The authentication method to use, possible values depend on the LDAP library XEmacs was compiled with, they may include

          simple, krbv41 and krbv42.

base The base for the search. This may look like ‘c, ome’, see RFC 1779 for syntax details.

scope One of the symbols base, onelevel or subtree indicating the scope of the search limited to a base object, to a single level or to the whole subtree.

deref The dereference policy is one of the symbols never, always, search or find and defines how aliases are dereferenced.

never Aliases are never dereferenced

always Aliases are always dereferenced

search Aliases are dereferenced when searching

find Aliases are dereferenced when locating the base object for the search

timelimit

The timeout limit for the connection in seconds.

sizelimit

The maximum number of matches to return for searches performed on this connection.

53.2.2 The High-Level LDAP API

As of this writing the high-level Lisp LDAP API only provides for LDAP searches. Further support is planned in the future.

The ldap-search function provides the most convenient interface to perform LDAP searches. It opens a connection to a host, performs the query and cleanly closes the connection thus insulating the user from all the details of the low-level interface such as LDAP Lisp objects see Section 53.2.3 [The Low-Level LDAP API], page 726

Function ldap-search filter &optional host attributes attrsonly Perform an LDAP search. filter is the search filter see Section 53.3 [Syntax of Search Filters], page 728 host is the LDAP host on which to perform the search attributes is the specific attributes to retrieve, nil means retrieve all attrsonly if non-nil retrieves the attributes only without their associated values. Additional search parameters can be specified through ldap-host-parameters-

alist.

53.2.3 The Low-Level LDAP API 53.2.3.1 The LDAP Lisp Object

An internal built-in ldap lisp object represents a LDAP connection.

Function ldapp object This function returns non-nil if object is a ldap object.

Function ldap-host ldap

Return the server host of the connection represented by ldap

Function ldap-live-p ldap

Return non-nil if ldap is an active LDAP connection

53.2.3.2 Opening and Closing a LDAP Connection

Function ldap-open host &optional plist Open a LDAP connection to host. plist is a property list containing additional parameters for the connection. Valid keys in that list are:

port The TCP port to use for the connection if different from ldapdefault-port or the library builtin value

auth The authentication method to use, possible values depend on the LDAP library XEmacs was compiled with, they may include

          simple, krbv41 and krbv42.

binddn The distinguished name of the user to bind as. This may look like ‘c, ome, cnnny Bugs’, see RFC 1779 for details.

passwd The password to use for authentication.

deref The dereference policy is one of the symbols never, always, search or find and defines how aliases are dereferenced.

never Aliases are never dereferenced

always Aliases are always dereferenced

search Aliases are dereferenced when searching

find Aliases are dereferenced when locating the base object for the search

The default is never.

timelimit

The timeout limit for the connection in seconds.

sizelimit

The maximum number of matches to return for searches performed on this connection.

Function ldap-close ldap

Close the connection represented by ldap

53.2.3.3 Searching on a LDAP Server (Low-level)

ldap-search-internal is the low-level primitive to perform a search on a LDAP server. It works directly on an open LDAP connection thus requiring a preliminary call to ldapopen. Multiple searches can be made on the same connection, then the session must be closed with ldap-close.

Function ldap-search-internal ldap filter base scope attrs attrsonly Perform a search on an open connection ldap created with ldap-open. filter is a filter string for the search see Section 53.3 [Syntax of Search Filters], page 728 base is the distinguished name at which to start the search. scope is one of the symbols base, onelevel or subtree indicating the scope of the search limited to a base object, to a single level or to the whole subtree. The default is subtree. attrs is a list of strings indicating which attributes to retrieve for each matching entry. If nil all available attributes are returned. If attrsonly is non-nil then only the attributes are retrieved, not their associated values The function returns a list of matching entries. Each entry being itself an alist of attribute/values.

53.3 Syntax of Search Filters

LDAP search functions use RFC1558 syntax to describe the search filter. In that syntax simple filters have the form:

     (<attr> <filtertype> <value>)

<attr> is an attribute name such as cn for Common Name, o for Organization, etc... <value> is the corresponding value. This is generally an exact string but may also

contain * characters as wildcards

filtertype is one = ~=, <=, >= which respectively describe equality, approximate equality, inferiority and superiority.

Thus (cn=John Smith) matches all records having a canonical name equal to John Smith.

A special case is the presence filter (<attr>=* which matches records containing a particular attribute. For instance (mail=*) matches all records containing a mail attribute.

Simple filters can be connected together with the logical operators &, | and ! which stand for the usual and, or and not operators.

   (&(objectClass=Person)(mail=*)(|(sn=Smith)(givenname=John))) matches records

of class Person containing a mail attribute and corresponding to people whose last name is Smith or whose first name is John.

54 Internationalization

54.1 I18N Levels 1 and 2

XEmacs is now compliant with I18N levels 1 and 2. Specifically, this means that it is 8-bit clean and correctly handles time and date functions. XEmacs will correctly display the entire ISO-Latin 1 character set.

The compose key may now be used to create any character in the ISO-Latin 1 character set not directly available via the keyboard.. In order for the compose key to work it is necessary to load the file ‘x-compose.el’. At any time while composing a character, C-h will display all valid completions and the character which would be produced.

54.2 I18N Level 3

54.2.1 Level 3 Basics

XEmacs now provides alpha-level functionality for I18N Level 3. This means that everything necessary for full messaging is available, but not every file has been converted.

The two message files which have been created are ‘src/emacs.po’ and ‘lisp/packages/mh-e.po’. Both files need to be converted using msgfmt, and the resulting ‘.mo’ files placed in some locale’s LC_MESSAGES directory. The test “translations” in these files are the original messages prefixed by TRNSLT_.

The domain for a variable is stored on the variable’s property list under the property name variable-domain. The function documentation-property uses this information when translating a variable’s documentation.

54.2.2 Level 3 Primitives

Function gettext string

This function looks up string in the default message domain and returns its translation. If I18N3 was not enabled when XEmacs was compiled, it just returns string.

Function dgettext domain string This function looks up string in the specified message domain and returns its translation. If I18N3 was not enabled when XEmacs was compiled, it just returns string.

Function bind-text-domain domain pathname

This function associates a pathname with a message domain. Here’s how the path to message file is constructed under SunOS 5.x:

          {pathname}/{LANG}/LC_MESSAGES/{domain}.mo

If I18N3 was not enabled when XEmacs was compiled, this function does nothing.

Special Form domain string

This function specifies the text domain used for translating documentation strings and interactive prompts of a function. For example, write:

          (defun foo (arg) "Doc string" (domain "emacs-foo") ...)

to specify emacs-foo as the text domain of the function foo. The “call” to domain is actually a declaration rather than a function; when actually called, domain just returns nil.

Function domain-of function

This function returns the text domain of function; it returns nil if it is the default domain. If I18N3 was not enabled when XEmacs was compiled, it always returns nil.

54.2.3 Dynamic Messaging

The format function has been extended to permit you to change the order of parameter insertion. For example, the conversion format %1$s inserts parameter one as a string, while %2$s inserts parameter two. This is useful when creating translations which require you to change the word order.

54.2.4 Domain Specification

The default message domain of XEmacs is ‘emacs’. For add-on packages, it is best to use a different domain. For example, let us say we want to convert the “gorilla” package to use the domain ‘emacs-gorilla’. To translate the message “What gorilla?”, use dgettext as follows:

     (dgettext "emacs-gorilla" "What gorilla?")

A function (or macro) which has a documentation string or an interactive prompt needs to be associated with the domain in order for the documentation or prompt to be translated. This is done with the domain special form as follows:

     (defun scratch (location)
       "Scratch the specified location."
       (domain "emacs-gorilla")
       (interactive "sScratch: ")
       ... )

It is most efficient to specify the domain in the first line of the function body, before the

interactive form.

For variables and constants which have documentation strings, specify the domain after the documentation.

Special Form defvar symbol [value [doc-string [domain]]]

Example:

          (defvar weight 250 "Weight of gorilla, in pounds." "emacs-gorilla")

Special Form defconst symbol [value [doc-string [domain]]]

Example:

          (defconst limbs 4 "Number of limbs" "emacs-gorilla")

Autoloaded functions which are specified in ‘loaddefs.el’ do not need to have a domain specification, because their documentation strings are extracted into the main message base. However, for autoloaded functions which are specified in a separate package, use following syntax:

Function autoload symbol filename &optional docstring interactive macro

domain Example:

          (autoload 'explore "jungle" "Explore the jungle." nil nil "emacs-gorilla")

54.2.5 Documentation String Extraction

The utility ‘etc/make-po’ scans the file DOC to extract documentation strings and creates a message file doc.po. This file may then be inserted within emacs.po.

Currently, make-po is hard-coded to read from DOC and write to doc.po. In order to extract documentation strings from an add-on package, first run make-docfile on the package to produce the DOC file. Then run make-po -p with the -p argument to indicate that we are extracting documentation for an add-on package.

(The -p argument is a kludge to make up for a subtle difference between pre-loaded documentation and add-on documentation: For add-on packages, the final carriage returns in the strings produced by make-docfile must be ignored.)

54.3 I18N Level 4

The Asian-language support in XEmacs is called “MULE”. See Chapter 55 [MULE], page 733.

732 XEmacs Lisp Reference Manual

55 MULE

MULE is the name originally given to the version of GNU Emacs extended for multilingual (and in particular Asian-language) support. “MULE” is short for “MUlti-Lingual Emacs”. It was originally called Nemacs (“Nihon Emacs” where “Nihon” is the Japanese word for “Japan”), when it only provided support for Japanese. XEmacs refers to its multi-lingual support as MULE support since it is based on MULE.

55.1 Internationalization Terminology

In internationalization terminology, a string of text is divided up into characters, which are the printable units that make up the text. A single character is (for example) a capital ‘A’, the number ‘2’, a Katakana character, a Kanji ideograph (an ideograph is a “picture” character, such as is used in Japanese Kanji, Chinese Hanzi, and Korean Hangul; typically there are thousands of such ideographs in each language), etc. The basic property of a character is its shape. Note that the same character may be drawn by two different people (or in two different fonts) in slightly different ways, although the basic shape will be the same.

In some cases, the differences will be significant enough that it is actually possible to identify two or more distinct shapes that both represent the same character. For example, the lowercase letters ‘a’ and ‘g’ each have two distinct possible shapes – the ‘a’ can optionally have a curved tail projecting off the top, and the ‘g’ can be formed either of two loops, or of one loop and a tail hanging off the bottom. Such distinct possible shapes of a character are called glyphs. The important characteristic of two glyphs making up the same character is that the choice between one or the other is purely stylistic and has no linguistic effect on a word (this is the reason why a capital ‘A’ and lowercase ‘a’ are different characters rather than different glyphs – e.g. ‘Aspen’ is a city while ‘aspen’ is a kind of tree).

Note that character and glyph are used differently here than elsewhere in XEmacs. A character set is simply a set of related characters. ASCII, for example, is a set of 94

characters (or 128, if you count non-printing characters). Other character sets are ISO8859-1 (ASCII plus various accented characters and other international symbols), JISX0201 (ASCII, more or less, plus half-width Katakana), JISX0208 (Japanese Kanji), JISX0212 (a second set of less-used Japanese Kanji), GB2312 (Mainland Chinese Hanzi), etc.

Every character set has one or more orderings, which can be viewed as a way of assigning a number (or set of numbers) to each character in the set. For most character sets, there is a standard ordering, and in fact all of the character sets mentioned above define a particular ordering. ASCII, for example, places letters in their “natural” order, puts uppercase letters before lowercase letters, numbers before letters, etc. Note that for many of the Asian character sets, there is no natural ordering of the characters. The actual orderings are based on one or more salient characteristic, of which there are many to choose from – e.g. number of strokes, common radicals, phonetic ordering, etc.

The set of numbers assigned to any particular character are called the character’s position codes. The number of position codes required to index a particular character in a character set is called the dimension of the character set. ASCII, being a relatively small character set, is of dimension one, and each character in the set is indexed using a single position code, in the range 0 through 127 (if non-printing characters are included) or 33 through 126 (if only the printing characters are considered). JISX0208, i.e. Japanese Kanji, has thousands of characters, and is of dimension two – every character is indexed by two position codes, each in the range 33 through 126. (Note that the choice of the range here is somewhat arbitrary. Although a character set such as JISX0208 defines an ordering of all its characters, it does not define the actual mapping between numbers and characters. You could just as easily index the characters in JISX0208 using numbers in the range 0 through 93, 1 through 94, 2 through 95, etc. The reason for the actual range chosen is so that the position codes match up with the actual values used in the common encodings.)

An encoding is a way of numerically representing characters from one or more character sets into a stream of like-sized numerical values called words; typically these are 8-bit, 16-bit, or 32-bit quantities. If an encoding encompasses only one character set, then the position codes for the characters in that character set could be used directly. (This is the case with ASCII, and as a result, most people do not understand the difference between a character set and an encoding.) This is not possible, however, if more than one character set is to be used in the encoding. For example, printed Japanese text typically requires characters from multiple character sets – ASCII, JISX0208, and JISX0212, to be specific. Each of these is indexed using one or more position codes in the range 33 through 126, so the position codes could not be used directly or there would be no way to tell which character was meant. Different Japanese encodings handle this differently – JIS uses special escape characters to denote different character sets; EUC sets the high bit of the position codes for JISX0208 and JISX0212, and puts a special extra byte before each JISX0212 character; etc. (JIS, EUC, and most of the other encodings you will encounter are 7-bit or 8-bit encodings. There is one common 16-bit encoding, which is Unicode; this strives to represent all the world’s characters in a single large character set. 32-bit encodings are generally used internally in programs to simplify the code that manipulates them; however, they are not much used externally because they are not very space-efficient.)

Encodings are classified as either modal or non-modal. In a modal encoding, there are multiple states that the encoding can be in, and the interpretation of the values in the stream depends on the current global state of the encoding. Special values in the encoding, called escape sequences, are used to change the global state. JIS, for example, is a modal encoding. The bytes ‘ESC $ B’ indicate that, from then on, bytes are to be interpreted as position codes for JISX0208, rather than as ASCII. This effect is cancelled using the bytes ‘ESC ( B’, which mean “switch from whatever the current state is to ASCII”. To switch to JISX0212, the escape sequence ‘ESC $ ( D’. (Note that here, as is common, the escape sequences do in fact begin with ‘ESC’. This is not necessarily the case, however.)

A non-modal encoding has no global state that extends past the character currently being interpreted. EUC, for example, is a non-modal encoding. Characters in JISX0208 are encoded by setting the high bit of the position codes, and characters in JISX0212 are encoded by doing the same but also prefixing the character with the byte 0x8F.

The advantage of a modal encoding is that it is generally more space-efficient, and is easily extendable because there are essentially an arbitrary number of escape sequences that can be created. The disadvantage, however, is that it is much more difficult to work with if it is not being processed in a sequential manner. In the non-modal EUC encoding, for example, the byte 0x41 always refers to the letter ‘A’; whereas in JIS, it could either be the letter ‘A’, or one of the two position codes in a JISX0208 character, or one of the two position codes in a JISX0212 character. Determining exactly which one is meant could be difficult and time-consuming if the previous bytes in the string have not already been processed.

Non-modal encodings are further divided into fixed-width and variable-width formats. A fixed-width encoding always uses the same number of words per character, whereas a variable-width encoding does not. EUC is a good example of a variable-width encoding: one to three bytes are used per character, depending on the character set. 16-bit and 32-bit encodings are nearly always fixed-width, and this is in fact one of the main reasons for using an encoding with a larger word size. The advantages of fixed-width encodings should be obvious. The advantages of variable-width encodings are that they are generally more space-efficient and allow for compatibility with existing 8-bit encodings such as ASCII.

Note that the bytes in an 8-bit encoding are often referred to as octets rather than simply as bytes. This terminology dates back to the days before 8-bit bytes were universal, when some computers had 9-bit bytes, others had 10-bit bytes, etc.

55.2 Charsets

A charset in MULE is an object that encapsulates a particular character set as well as an ordering of those characters. Charsets are permanent objects and are named using symbols, like faces.

Function charsetp object

This function returns non-nil if object is a charset.

55.2.1 Charset Properties

Charsets have the following properties:

name A symbol naming the charset. Every charset must have a different name; this allows a charset to be referred to using its name rather than the actual charset object.

doc-string

A documentation string describing the charset.

registry A regular expression matching the font registry field for this character set. For example, both the ascii and latin-iso8859-1 charsets use the registry "ISO8859-1". This field is used to choose an appropriate font when the user gives a general font specification such as ‘-*-courier-medium-r-*-140-*’, i.e. a 14-point upright medium-weight Courier font.

dimension

Number of position codes used to index a character in the character set. XEmacs/MULE can only handle character sets of dimension 1 or 2. This property defaults to 1.

chars Number of characters in each dimension. In XEmacs/MULE, the only allowed values are 94 or 96. (There are a couple of pre-defined character sets, such as ASCII, that do not follow this, but you cannot define new ones like this.) Defaults to 94. Note that if the dimension is 2, the character set thus described is 94x94 or 96x96.

columns Number of columns used to display a character in this charset. Only used in TTY mode. (Under X, the actual width of a character can be derived from the font used to display the characters.) If unspecified, defaults to the dimension. (This is almost always the correct value, because character sets with dimension 2 are usually ideograph character sets, which need two columns to display the intricate ideographs.)

direction

A symbol, either l2r (left-to-right) or r2l (right-to-left). Defaults to l2r. This specifies the direction that the text should be displayed in, and will be left-to-right for most charsets but right-to-left for Hebrew and Arabic. (Right-to-left display is not currently implemented.)

final Final byte of the standard ISO 2022 escape sequence designating this charset. Must be supplied. Each combination of (dimension, chars) defines a separate namespace for final bytes, and each charset within a particular namespace must have a different final byte. Note that ISO 2022 restricts the final byte to the range 0x30 - 0x7E if dimension == 1, and 0x30 - 0x5F if dimension == 2. Note also that final bytes in the range 0x30 - 0x3F are reserved for user-defined (not official) character sets. For more information on ISO 2022, see Section 55.6 [Coding Systems], page 743.

graphic 0 (use left half of font on output) or 1 (use right half of font on output). Defaults to 0. This specifies how to convert the position codes that index a character in a character set into an index into the font used to display the character set. With graphic set to 0, position codes 33 through 126 map to font indices 33 through 126; with it set to 1, position codes 33 through 126 map to font indices 161 through 254 (i.e. the same number but with the high bit set). For example, for a font whose registry is ISO8859-1, the left half of the font (octets 0x20 - 0x7F) is the ascii charset, while the right half (octets 0xA0 - 0xFF) is the

latin-iso8859-1 charset.
ccl-program

A compiled CCL program used to convert a character in this charset into an index into the font. This is in addition to the graphic property. If a CCL program is defined, the position codes of a character will first be processed according to graphic and then passed through the CCL program, with the resulting values used to index the font.

This is used, for example, in the Big5 character set (used in Taiwan). This character set is not ISO-2022-compliant, and its size (94x157) does not fit within the maximum 96x96 size of ISO-2022-compliant character sets. As a result, XEmacs/MULE splits it (in a rather complex fashion, so as to group the most commonly used characters together) into two charset objects (big5-1 and big5-2), each of size 94x94, and each charset object uses a CCL program to convert the modified position codes back into standard Big5 indices to retrieve a character from a Big5 font.

Most of the above properties can only be changed when the charset is created. See Section 55.2.3 [Charset Property Functions], page 737.

55.2.2 Basic Charset Functions

Function find-charset charset-or-name This function retrieves the charset of the given name. If charset-or-name is a charset object, it is simply returned. Otherwise, charset-or-name should be a symbol. If there is no such charset, nil is returned. Otherwise the associated charset object is returned.

Function get-charset name

This function retrieves the charset of the given name. Same as find-charset except an error is signalled if there is no such charset instead of returning nil.

Function charset-list

This function returns a list of the names of all defined charsets.

Function make-charset name doc-string props This function defines a new character set. This function is for use with Mule support. name is a symbol, the name by which the character set is normally referred. doc-string is a string describing the character set. props is a property list, describing the specific nature of the character set. The recognized proper-

ties are registry, dimension, columns, chars, final, graphic, direction,

and ccl-program, as previously described.

Function make-reverse-direction-charset charset new-name

This function makes a charset equivalent to charset but which goes in the opposite direction. new-name is the name of the new charset. The new charset is returned.

Function charset-from-attributes dimension chars final &optional

direction This function returns a charset with the given dimension, chars, final, and direction. If direction is omitted, both directions will be checked (left-to-right will be returned if character sets exist for both directions).

Function charset-reverse-direction-charset charset

This function returns the charset (if any) with the same dimension, number of characters, and final byte as charset, but which is displayed in the opposite direction.

55.2.3 Charset Property Functions

All of these functions accept either a charset name or charset object.

Function charset-property charset prop

This function returns property prop of charset. See Section 55.2.1 [Charset Properties], page 735.

Convenience functions are also provided for retrieving individual properties of a charset.

Function charset-name charset

This function returns the name of charset. This will be a symbol.

Function charset-doc-string charset

This function returns the doc string of charset.

Function charset-registry charset

This function returns the registry of charset.

Function charset-dimension charset

This function returns the dimension of charset.

Function charset-chars charset

This function returns the number of characters per dimension of charset.

Function charset-columns charset

This function returns the number of display columns per character (in TTY mode) of charset.

Function charset-direction charset

This function returns the display direction of charset – either l2r or r2l.

Function charset-final charset

This function returns the final byte of the ISO 2022 escape sequence designating charset.

Function charset-graphic charset

This function returns either 0 or 1, depending on whether the position codes of characters in charset map to the left or right half of their font, respectively.

Function charset-ccl-program charset

This function returns the CCL program, if any, for converting position codes of characters in charset into font indices.

The only property of a charset that can currently be set after the charset has been created is the CCL program.

Function set-charset-ccl-program charset ccl-program

This function sets the ccl-program property of charset to ccl-program.

55.2.4 Predefined Charsets

The following charsets are predefined in the C code.

Name                    Type  Fi Gr Dir Registry
--------------------------------------------------------------
ascii                    94    B  0  l2r ISO8859-1
control-1                94       0  l2r ---
latin-iso8859-1          94    A  1  l2r ISO8859-1
latin-iso8859-2          96    B  1  l2r ISO8859-2
latin-iso8859-3          96    C  1  l2r ISO8859-3
latin-iso8859-4          96    D  1  l2r ISO8859-4
cyrillic-iso8859-5       96    L  1  l2r ISO8859-5
arabic-iso8859-6         96    G  1  r2l ISO8859-6
greek-iso8859-7          96    F  1  l2r ISO8859-7
hebrew-iso8859-8         96    H  1  r2l ISO8859-8
latin-iso8859-9          96    M  1  l2r ISO8859-9
thai-tis620              96    T  1  l2r TIS620
katakana-jisx0201        94    I  1  l2r JISX0201.1976
latin-jisx0201           94    J  0  l2r JISX0201.1976
japanese-jisx0208-1978   94x94 @  0  l2r JISX0208.1978
japanese-jisx0208        94x94 B  0  l2r JISX0208.19(83|90)
japanese-jisx0212        94x94 D  0  l2r JISX0212
chinese-gb2312           94x94 A  0  l2r GB2312
chinese-cns11643-1       94x94 G  0  l2r CNS11643.1
chinese-cns11643-2       94x94 H  0  l2r CNS11643.2
chinese-big5-1           94x94 0  0  l2r Big5
chinese-big5-2           94x94 1  0  l2r Big5
korean-ksc5601           94x94 C  0  l2r KSC5601
composite                96x96    0  l2r ---

The following charsets are predefined in the Lisp code.

Name                     Type  Fi Gr Dir Registry
--------------------------------------------------------------
arabic-digit             94    2  0  l2r MuleArabic-0
arabic-1-column          94    3  0  r2l MuleArabic-1
arabic-2-column          94    4  0  r2l MuleArabic-2
sisheng                  94    0  0  l2r sisheng_cwnn\|OMRON_UDC_ZH
chinese-cns11643-3       94x94 I  0  l2r CNS11643.1
chinese-cns11643-4       94x94 J  0  l2r CNS11643.1
chinese-cns11643-5       94x94 K  0  l2r CNS11643.1
chinese-cns11643-6       94x94 L  0  l2r CNS11643.1
chinese-cns11643-7       94x94 M  0  l2r CNS11643.1
ethiopic                 94x94 2  0  l2r Ethio
ascii-r2l                94    B  0  r2l ISO8859-1
ipa                      96    0  1  l2r MuleIPA
vietnamese-lower         96    1  1  l2r VISCII1.1
vietnamese-upper         96    2  1  l2r VISCII1.1

For all of the above charsets, the dimension and number of columns are the same.

Note that ASCII, Control-1, and Composite are handled specially. This is why some of the fields are blank; and some of the filled-in fields (e.g. the type) are not really accurate.

55.3 MULE Characters

Function make-char charset arg1 &optional arg2 This function makes a multi-byte character from charset and octets arg1 and arg2.

Function char-charset ch

This function returns the character set of char ch.

Function char-octet ch &optional n This function returns the octet (i.e. position code) numbered n (should be 0 or

1) of char ch. n defaults to 0 if omitted.

Function find-charset-region start end &optional buffer This function returns a list of the charsets in the region between start and end. buffer defaults to the current buffer if omitted.

Function find-charset-string string

This function returns a list of the charsets in string.

55.4 Composite Characters

Composite characters are not yet completely implemented.

Function make-composite-char string

This function converts a string into a single composite character. The character is the result of overstriking all the characters in the string.

Function composite-char-string ch

This function returns a string of the characters comprising a composite character.

Function compose-region start end &optional buffer This function composes the characters in the region from start to end in buffer into one composite character. The composite character replaces the composed characters. buffer defaults to the current buffer if omitted.

Function decompose-region start end &optional buffer This function decomposes any composite characters in the region from start to end in buffer. This converts each composite character into one or more characters, the individual characters out of which the composite character was formed. Non-composite characters are left as-is. buffer defaults to the current buffer if omitted.

55.5 ISO 2022

This section briefly describes the ISO 2022 encoding standard. For more thorough understanding, please refer to the original document of ISO 2022.

Character sets (charsets) are classified into the following four categories, according to the number of characters of charset: 94-charset, 96-charset, 94x94-charset, and 96x96-charset.

94-charset ASCII(B), left(J) and right(I) half of JISX0201, ...

96-charset Latin-1(A), Latin-2(B), Latin-3(C), ...

94x94-charset GB2312(A), JISX0208(B), KSC5601(C), ...

96x96-charset none for the moment

The character in parentheses after the name of each charset is the final character F, which can be regarded as the identifier of the charset. ECMA allocates F to each charset. F is in the range of 0x30..0x7F, but 0x30..0x3F are only for private use.

Note: ECMA = European Computer Manufacturers Association

There are four registers of charsets, called G0 thru G3. You can designate (or assign) any charset to one of these registers.

The code space contained within one octet (of size 256) is divided into 4 areas: C0, GL, C1, and GR. GL and GR are the areas into which a register of charset can be invoked into.

     C0: 0x00 - 0x1F
     GL: 0x20 - 0x7F
     C1: 0x80 - 0x9F
     GR: 0xA0 - 0xFF

Usually, in the initial state, G0 is invoked into GL, and G1 is invoked into GR.

ISO 2022 distinguishes 7-bit environments and 8-bit environments. In 7-bit environments, only C0 and GL are used.

Charset designation is done by escape sequences of the form:

     ESC [I] I F

where I is an intermediate character in the range 0x20 - 0x2F, and F is the final character identifying this charset.

The meaning of intermediate characters are:

     $ [0x24]: indicate charset of dimension 2 (94x94 or 96x96).
     ( [0x28]: designate to G0 a 94-charset whose final byte is F.
     ) [0x29]: designate to G1 a 94-charset whose final byte is F.
     * [0x2A]: designate to G2 a 94-charset whose final byte is F.
     + [0x2B]: designate to G3 a 94-charset whose final byte is F.
     - [0x2D]: designate to G1 a 96-charset whose final byte is F.
     . [0x2E]: designate to G2 a 96-charset whose final byte is F.
     / [0x2F]: designate to G3 a 96-charset whose final byte is F.

The following rule is not allowed in ISO 2022 but can be used in Mule.

, [0x2C]: designate to G0 a 96-charset whose final byte is F.

Here are examples of designations:

ESC ( B :              designate to G0 ASCII
ESC - A :              designate to G1 Latin-1
ESC $ ( A or ESC $ A : designate to G0 GB2312
ESC $ ( B or ESC $ B : designate to G0 JISX0208
ESC $ ) C :            designate to G1 KSC5601

To use a charset designated to G2 or G3, and to use a charset designated to G1 in a 7-bit environment, you must explicitly invoke G1, G2, or G3 into GL. There are two types of invocation, Locking Shift (forever) and Single Shift (one character only).

Locking Shift is done as follows:

LS0 or SI (0x0F): invoke G0 into GL
LS1 or SO (0x0E): invoke G1 into GL
LS2:  invoke G2 into GL
LS3:  invoke G3 into GL
LS1R: invoke G1 into GR
LS2R: invoke G2 into GR
LS3R: invoke G3 into GR

Single Shift is done as follows:

SS2 or ESC N: invoke G2 into GL
SS3 or ESC O: invoke G3 into GL

(#### Ben says: I think the above is slightly incorrect. It appears that SS2 invokes G2 into GR and SS3 invokes G3 into GR, whereas ESC N and ESC O behave as indicated. The above definitions will not parse EUC-encoded text correctly, and it looks like the code in mule-coding.c has similar problems.)

You may realize that there are a lot of ISO-2022-compliant ways of encoding multilingual text. Now, in the world, there exist many coding systems such as X11’s Compound Text, Japanese JUNET code, and so-called EUC (Extended UNIX Code); all of these are variants of ISO 2022.

In Mule, we characterize ISO 2022 by the following attributes:

1. Initial designation to G0 thru G3.

2. Allow designation of short form for Japanese and Chinese.

3. Should we designate ASCII to G0 before control characters?

4. Should we designate ASCII to G0 at the end of line?

5. 7-bit environment or 8-bit environment.

6. Use Locking Shift or not.

7. Use ASCII or JIS0201-1976-Roman.

8. Use JISX0208-1983 or JISX0208-1976.

(The last two are only for Japanese.)

By specifying these attributes, you can create any variant of ISO 2022.

Here are several examples:

junet -- Coding system used in JUNET.
1. G0 <- ASCII, G1..3 <- never used
2. Yes.
3. Yes.
4. Yes.
5. 7-bit environment
6. No.
7. Use ASCII
8. Use JISX0208-1983
ctext -- Compound Text
1. G0 <- ASCII, G1 <- Latin-1, G2,3 <- never used
2. No.
3. No.
4. Yes.
5. 8-bit environment
6. No.
7. Use ASCII
8. Use JISX0208-1983
euc-china -- Chinese EUC.  Although many people call this
as "GB encoding", the name may cause misunderstanding.
1. G0 <- ASCII, G1 <- GB2312, G2,3 <- never used
2. No.
3. Yes.
4. Yes.
5. 8-bit environment
6. No.
7. Use ASCII
8. Use JISX0208-1983
korean-mail -- Coding system used in Korean network.
1. G0 <- ASCII, G1 <- KSC5601, G2,3 <- never used
2. No.
3. Yes.
4. Yes.
5. 7-bit environment
6. Yes.
7. No.
8. No.

Mule creates all these coding systems by default.

55.6 Coding Systems

A coding system is an object that defines how text containing multiple character sets is encoded into a stream of (typically 8-bit) bytes. The coding system is used to decode the stream into a series of characters (which may be from multiple charsets) when the text is read from a file or process, and is used to encode the text back into the same format when it is written out to a file or process.

For example, many ISO-2022-compliant coding systems (such as Compound Text, which is used for inter-client data under the X Window System) use escape sequences to switch between different charsets – Japanese Kanji, for example, is invoked with ‘ESC $ ( B’; ASCII is invoked with ‘ESC ( B’; and Cyrillic is invoked with ‘ESC - L’. See make-coding-system for more information.

Coding systems are normally identified using a symbol, and the symbol is accepted in place of the actual coding system object whenever a coding system is called for. (This is similar to how faces and charsets work.)

Function coding-system-p object

This function returns non-nil if object is a coding system.

55.6.1 Coding System Types

nil
autodetect

Automatic conversion. XEmacs attempts to detect the coding system used in the file.

no-conversion

No conversion. Use this for binary files and such. On output, graphic characters that are not in ASCII or Latin-1 will be replaced by a ‘?’. (For a no-conversion-encoded buffer, these characters will only be present if you explicitly insert them.)

shift-jis

Shift-JIS (a Japanese encoding commonly used in PC operating systems).

iso2022 Any ISO-2022-compliant encoding. Among other things, this includes JIS (the Japanese encoding commonly used for e-mail), national variants of EUC (the standard Unix encoding for Japanese and other languages), and Compound Text (an encoding used in X11). You can specify more specific information about the conversion with the flags argument.

big5 Big5 (the encoding commonly used for Taiwanese).

ccl The conversion is performed using a user-written pseudo-code program. CCL (Code Conversion Language) is the name of this pseudo-code.

internal Write out or read in the raw contents of the memory representing the buffer’s text. This is primarily useful for debugging purposes, and is only enabled when XEmacs has been compiled with DEBUG_XEMACS set (the ‘--debug’ configure option). Warning: Reading in a file using internal conversion can result in an internal inconsistency in the memory representing a buffer’s text, which will produce unpredictable results and may cause XEmacs to crash. Under normal circumstances you should never use internal conversion.

55.6.2 EOL Conversion

nil Automatically detect the end-of-line type (LF, CRLF, or CR). Also generate subsidiary coding systems named name-unix, name-dos, and name-mac, that are identical to this coding system but have an EOL-TYPE value of lf, crlf, and cr, respectively.

lf The end of a line is marked externally using ASCII LF. Since this is also the way that XEmacs represents an end-of-line internally, specifying this option results in no end-of-line conversion. This is the standard format for Unix text files.

crlf The end of a line is marked externally using ASCII CRLF. This is the standard format for MS-DOS text files.

cr The end of a line is marked externally using ASCII CR. This is the standard format for Macintosh text files.

t Automatically detect the end-of-line type but do not generate subsidiary coding systems. (This value is converted to nil when stored internally, and coding-

system-property will return nil.)

55.6.3 Coding System Properties

mnemonic String to be displayed in the modeline when this coding system is active.

eol-type End-of-line conversion to be used. It should be one of the types listed in Section 55.6.2 [EOL Conversion], page 744.

post-read-conversion

Function called after a file has been read in, to perform the decoding. Called with two arguments, beg and end, denoting a region of the current buffer to be decoded.

pre-write-conversion

Function called before a file is written out, to perform the encoding. Called with two arguments, beg and end, denoting a region of the current buffer to be encoded.

The following additional properties are recognized if type is iso2022:

charset-g0
charset-g1
charset-g2
charset-g3

The character set initially designated to the G0 - G3 registers. The value should be one of

• A charset object (designate that character set)

• nil (do not ever use this register)

• t (no character set is initially designated to the register, but may be later on; this automatically sets the corresponding force-g*-on-output property)

force-g0-on-output
force-g1-on-output
force-g2-on-output
force-g3-on-output

If non-nil, send an explicit designation sequence on output before using the specified register.

short If non-nil, use the short forms ‘ESC $ @’, ‘ESC $ A’, and ‘ESC $ B’ on output in place of the full designation sequences ‘ESC $ ( @’, ‘ESC $ ( A’, and ‘ESC $ ( B’.

no-ascii-eol

If non-nil, don’t designate ASCII to G0 at each end of line on output. Setting this to non-nil also suppresses other state-resetting that normally happens at the end of a line.

no-ascii-cntl

If non-nil, don’t designate ASCII to G0 before control chars on output.

seven If non-nil, use 7-bit environment on output. Otherwise, use 8-bit environment.

lock-shift

If non-nil, use locking-shift (SO/SI) instead of single-shift or designation by escape sequence.

no-iso6429

If non-nil, don’t use ISO6429’s direction specification.

escape-quoted

If non-nil, literal control characters that are the same as the beginning of a recognized ISO 2022 or ISO 6429 escape sequence (in particular, ESC (0x1B), SO (0x0E), SI (0x0F), SS2 (0x8E), SS3 (0x8F), and CSI (0x9B)) are “quoted” with an escape character so that they can be properly distinguished from an escape sequence. (Note that doing this results in a non-portable encoding.) This encoding flag is used for byte-compiled files. Note that ESC is a good choice for a quoting character because there are no escape sequences whose second byte is a character from the Control-0 or Control-1 character sets; this is explicitly disallowed by the ISO 2022 standard.

input-charset-conversion

A list of conversion specifications, specifying conversion of characters in one charset to another when decoding is performed. Each specification is a list of two elements: the source charset, and the destination charset.

output-charset-conversion

A list of conversion specifications, specifying conversion of characters in one charset to another when encoding is performed. The form of each specification

is the same as for input-charset-conversion.

The following additional properties are recognized (and required) if type is ccl:

decode CCL program used for decoding (converting to internal format).

encode CCL program used for encoding (converting to external format).

55.6.4 Basic Coding System Functions

Function find-coding-system coding-system-or-name This function retrieves the coding system of the given name. If coding-system-or-name is a coding-system object, it is simply returned. Otherwise, coding-system-or-name should be a symbol. If there is no such coding system, nil is returned. Otherwise the associated coding system object is returned.

Function get-coding-system name

This function retrieves the coding system of the given name. Same as findcoding-system except an error is signalled if there is no such coding system instead of returning nil.

Function coding-system-list

This function returns a list of the names of all defined coding systems.

Function coding-system-name coding-system

This function returns the name of the given coding system.

Function make-coding-system name type &optional doc-string props This function registers symbol name as a coding system. type describes the conversion method used and should be one of the types listed in Section 55.6.1 [Coding System Types], page 744. doc-string is a string describing the coding system. props is a property list, describing the specific nature of the character set. Recognized properties are as in Section 55.6.3 [Coding System Properties], page 745.

Function copy-coding-system old-coding-system new-name This function copies old-coding-system to new-name. If new-name does not name an existing coding system, a new one will be created.

Function subsidiary-coding-system coding-system eol-type

This function returns the subsidiary coding system of coding-system with eol type eol-type.

55.6.5 Coding System Property Functions

Function coding-system-doc-string coding-system

This function returns the doc string for coding-system.

Function coding-system-type coding-system

This function returns the type of coding-system.

Function coding-system-property coding-system prop

This function returns the prop property of coding-system.

55.6.6 Encoding and Decoding Text

Function decode-coding-region start end coding-system &optional buffer This function decodes the text between start and end which is encoded in coding-system. This is useful if you’ve read in encoded text from a file without decoding it (e.g. you read in a JIS-formatted file but used the binary or no-conversion coding system, so that it shows up as ‘^[$B!<!+^[(B’). The length of the encoded text is returned. buffer defaults to the current buffer if unspecified.

Function encode-coding-region start end coding-system &optional buffer This function encodes the text between start and end using coding-system. This will, for example, convert Japanese characters into stuff such as ‘^[$B!<!+^[(B’ if you use the JIS encoding. The length of the encoded text is returned. buffer defaults to the current buffer if unspecified.

55.6.7 Detection of Textual Encoding

Function coding-category-list

This function returns a list of all recognized coding categories.

Function set-coding-priority-list list

This function changes the priority order of the coding categories. list should be a list of coding categories, in descending order of priority. Unspecified coding categories will be lower in priority than all specified ones, in the same relative order they were in previously.

Function coding-priority-list

This function returns a list of coding categories in descending order of priority.

Function set-coding-category-system coding-category coding-system This function changes the coding system associated with a coding category.

Function coding-category-system coding-category

This function returns the coding system associated with a coding category.

Function detect-coding-region start end &optional buffer This function detects coding system of the text in the region between start and end. Returned value is a list of possible coding systems ordered by priority. If only ASCII characters are found, it returns autodetect or one of its subsidiary coding systems according to a detected end-of-line type. Optional arg buffer defaults to the current buffer.

55.6.8 Big5 and Shift-JIS Functions

These are special functions for working with the non-standard Shift-JIS and Big5 encodings.

Function decode-shift-jis-char code

This function decodes a JISX0208 character of Shift-JIS coding-system. code is the character code in Shift-JIS as a cons of type bytes. The corresponding character is returned.

Function encode-shift-jis-char ch

This function encodes a JISX0208 character ch to SHIFT-JIS coding-system. The corresponding character code in SHIFT-JIS is returned as a cons of two bytes.

Function decode-big5-char code

This function decodes a Big5 character code of BIG5 coding-system. code is the character code in BIG5. The corresponding character is returned.

Function encode-big5-char ch

This function encodes the Big5 character char to BIG5 coding-system. The corresponding character code in Big5 is returned.

55.7 CCL

CCL (Code Conversion Language) is a simple structured programming language designed for character coding conversions. A CCL program is compiled to CCL code (represented by a vector of integers) and executed by the CCL interpreter embedded in Emacs. The CCL interpreter implements a virtual machine with 8 registers called r0, ..., r7, a number of control structures, and some I/O operators. Take care when using registers r0 (used in implicit set statements) and especially r7 (used internally by several statements and operations, especially for multiple return values and I/O operations).

CCL is used for code conversion during process I/O and file I/O for non-ISO2022 coding systems. (It is the only way for a user to specify a code conversion function.) It is also used for calculating the code point of an X11 font from a character code. However, since CCL is designed as a powerful programming language, it can be used for more generic calculation where efficiency is demanded. A combination of three or more arithmetic operations can be calculated faster by CCL than by Emacs Lisp.

   Warning: The code in ‘src/mule-ccl.c’ and ‘$packages/lisp/mule-base/mule-ccl.el’

is the definitive description of CCL’s semantics. The previous version of this section contained several typos and obsolete names left from earlier versions of MULE, and many may remain. (I am not an experienced CCL programmer; the few who know CCL well find writing English painful.)

A CCL program transforms an input data stream into an output data stream. The input stream, held in a buffer of constant bytes, is left unchanged. The buffer may be filled by an external input operation, taken from an Emacs buffer, or taken from a Lisp string. The output buffer is a dynamic array of bytes, which can be written by an external output operation, inserted into an Emacs buffer, or returned as a Lisp string.

A CCL program is a (Lisp) list containing two or three members. The first member is the buffer magnification, which indicates the required minimum size of the output buffer as a multiple of the input buffer. It is followed by the main block which executes while there is input remaining, and an optional EOF block which is executed when the input is exhausted. Both the main block and the EOF block are CCL blocks.

A CCL block is either a CCL statement or list of CCL statements. A CCL statement is either a set statement (either an integer or an assignment, which is a list of a register to receive the assignment, an assignment operator, and an expression) or a control statement (a list starting with a keyword, whose allowable syntax depends on the keyword).

55.7.1 CCL Syntax

The full syntax of a CCL program in BNF notation: CCL PROGRAM :=

(BUFFER MAGNIFICATION

CCL MAIN BLOCK [ CCL EOF BLOCK ])

BUFFER MAGNIFICATION := integer CCL MAIN BLOCK := CCL BLOCK CCL EOF BLOCK := CCL BLOCK

CCL BLOCK :=

STATEMENT | (STATEMENT [STATEMENT ...]) STATEMENT :=

SET | IF | BRANCH | LOOP | REPEAT | BREAK | READ | WRITE | CALL | END

SET := (REG = EXPRESSION) | (REG ASSIGNMENT OPERATOR EXPRESSION) | integer

EXPRESSION := ARG | (EXPRESSION OPERATOR ARG)

IF := (if EXPRESSION CCL BLOCK [CCL BLOCK]) BRANCH := (branch EXPRESSION CCL BLOCK [CCL BLOCK ...]) LOOP := (loop STATEMENT [STATEMENT ...]) BREAK := (break) REPEAT :=

(repeat)

| (write-repeat [REG | integer | string])

| (write-read-repeat REG [integer | ARRAY]) READ :=

(read REG ...)

| (read-if (REG OPERATOR ARG) CCL BLOCK CCL BLOCK)

| (read-branch REG CCL BLOCK [CCL BLOCK ...]) WRITE :=

(write REG ...)

| (write EXPRESSION)

| (write integer) | (write string) | (write REG ARRAY)

| string CALL := (call ccl-program-name) END := (end)

REG := r0 | r1 | r2 | r3 | r4 | r5 | r6 | r7 ARG := REG | integer OPERATOR :=

     + | - | * | / | % | & | ’|’ | ^| << | >> | <8 | >8 | //

| < | > | == | <= | >= | != | de-sjis | en-sjis ASSIGNMENT OPERATOR :=

+= | -= | *= | /= | %= | &= | ’|=’ | ^= | <<= | >>= ARRAY := ’[’ integer ... ’]’

55.7.2 CCL Statements

The Emacs Code Conversion Language provides the following statement types: set, if, branch, loop, repeat, break, read, write, call, and end.

Set statement:

The set statement has three variants with the syntaxes ‘(reg = expression)’, ‘(reg assignment operator expression)’, and ‘integer’. The assignment operator variation of the set statement works the same way as the corresponding C expression statement does. The assignment operators are +=, -=, *=, /=, %=, &=, |=, ^=, <<=, and >>=, and they have the same meanings as in C. A "naked integer" integer is equivalent to a set statement of the form (r0 = integer).

I/O statements:

The read statement takes one or more registers as arguments. It reads one byte (a C char) from the input into each register in turn.

The write takes several forms. In the form ‘(write reg ...)’ it takes one or more registers as arguments and writes each in turn to the output. The integer in a register (interpreted as an Emchar) is encoded to multibyte form (ie, Bufbytes) and written to the current output buffer. If it is less than 256, it is written as is. The forms ‘(write expression)’ and ‘(write integer)’ are treated analogously. The form ‘(write string)’ writes the constant string to the output. A "naked string" ‘string’ is equivalent to the statement ‘(write string)’. The form ‘(write reg array)’ writes the regth element of the array to the output.

Conditional statements:

The if statement takes an expression, a CCL block, and an optional second CCL block as arguments. If the expression evaluates to non-zero, the first CCL block is executed. Otherwise, if there is a second CCL block, it is executed.

The read-if variant of the if statement takes an expression, a CCL block, and an optional second CCL block as arguments. The expression must have the form (reg operator operand) (where operand is a register or an integer). The read-if statement first reads from the input into the first register operand in the expression, then conditionally executes a CCL block just as the if statement does.

The branch statement takes an expression and one or more CCL blocks as arguments. The CCL blocks are treated as a zero-indexed array, and the branch statement uses the expression as the index of the CCL block to execute. Null CCL blocks may be used as no-ops, continuing execution with the statement following the branch statement in the containing CCL block. Out-of-range values for the EXPRESSION are also treated as no-ops.

The read-branch variant of the branch statement takes an register, a CCL block, and an optional second CCL block as arguments. The read-branch statement first reads from the input into the register, then conditionally executes a CCL block just as the branch statement does.

Loop control statements:

The loop statement creates a block with an implied jump from the end of the block back to its head. The loop is exited on a break statement, and continued without executing the tail by a repeat statement.

The break statement, written ‘(break)’, terminates the current loop and continues with the next statement in the current block.

The repeat statement has three variants, repeat, write-repeat, and write-read-repeat. Each continues the current loop from its head, possibly after performing I/O. repeat takes no arguments and does no I/O before jumping. write-repeat takes a single argument (a register, an integer, or a string), writes it to the output, then jumps. writeread-repeat takes one or two arguments. The first must be a register. The second may be an integer or an array; if absent, it is implicitly set to the first (register) argument. write-read-repeat writes its second argument to the output, then reads from the input into the register, and finally jumps. See the write and read statements for the semantics of the I/O operations for each type of argument.

Other control statements:

The call statement, written ‘(call ccl-program-name)’, executes a CCL program as a subroutine. It does not return a value to the caller, but can modify the register status.

The end statement, written ‘(end)’, terminates the CCL program successfully, and returns to caller (which may be a CCL program). It does not alter the status of the registers.

55.7.3 CCL Expressions

CCL, unlike Lisp, uses infix expressions. The simplest CCL expressions consist of a single operand, either a register (one of r0, ..., r0) or an integer. Complex expressions are lists of the form ( expression operator operand ). Unlike C, assignments are not expressions.

In the following table, X is the target resister for a set. In subexpressions, this is implicitly r7. This means that >8, //, de-sjis, and en-sjis cannot be used freely in subexpressions, since they return parts of their values in r7. Y may be an expression, register, or integer, while Z must be a register or an integer.

Name  Operator Code C-like Description
CCL PLUS  +  0x00 X = Y + Z
CCL MINUS  -  0x01 X = Y - Z
CCL MUL  *  0x02 X = Y * Z
CCL DIV  /  0x03 X = Y / Z
CCL MOD  %  0x04 X = Y % Z
CCL AND  &  0x05 X = Y & Z
CCL OR  |  0x06 X = Y | Z
CCL XOR  ^0x07 X = Y ^Z
CCL LSH  <<  0x08 X = Y << Z
CCL RSH  >>  0x09 X = Y >> Z
CCL LSH8  <8  0x0A X = (Y << 8) | Z
CCL RSH8  >8  0x0B X = Y >> 8, r[7] = Y & 0xFF
CCL DIVMOD  //  0x0C X = Y / Z, r[7] = Y % Z
CCL LS  <  0x10 X = (X < Y)
CCL GT  >  0x11 X = (X > Y)
CCL EQ  ==  0x12 X = (X == Y)
CCL LE  <=  0x13 X = (X <= Y)
CCL GE  >=  0x14 X = (X >= Y)
CCL NE  !=  0x15 X = (X != Y)
CCL ENCODE SJIS en-sjis 0x16 X = HIGHER BYTE (SJIS (Y, Z))

r[7] = LOWER BYTE (SJIS (Y, Z)

CCL DECODE SJIS de-sjis 0x17 X = HIGHER BYTE (DE-SJIS (Y, Z))

r[7] = LOWER BYTE (DE-SJIS (Y, Z))

The CCL operators are as in C, with the addition of CCL LSH8, CCL RSH8, CCL DIVMOD, CCL ENCODE SJIS, and CCL DECODE SJIS. The CCL ENCODE SJIS and CCL DECODE SJIS treat their first and second bytes as the high and low bytes of a two-byte character code. (SJIS stands for Shift JIS, an encoding of Japanese characters used by Microsoft. CCL ENCODE SJIS is a complicated transformation of the Japanese standard JIS encoding to Shift JIS. CCL DECODE SJIS is its inverse.) It is somewhat odd to represent the SJIS operations in infix form.

55.7.4 Calling CCL

CCL programs are called automatically during Emacs buffer I/O when the external representation has a coding system type of shift-jis, big5, or ccl. The program is specified by the coding system (see Section 55.6 [Coding Systems], page 743). You can also call CCL programs from other CCL programs, and from Lisp using these functions:

Function ccl-execute ccl-program status Execute ccl-program with registers initialized by status. ccl-program is a vector of compiled CCL code created by ccl-compile. It is an error for the program to try to execute a CCL I/O command. status must be a vector of nine values, specifying the initial value for the R0, R1 .. R7 registers and for the instruction counter IC. A nil value for a register initializer causes the register to be set to

0. A nil value for the IC initializer causes execution to start at the beginning of the program. When the program is done, status is modified (by side-effect) to contain the ending values for the corresponding registers and IC.

Function ccl-execute-on-string ccl-program status str &optional continue Execute ccl-program with initial status on string. ccl-program is a vector of compiled CCL code created by ccl-compile. status must be a vector of nine values, specifying the initial value for the R0, R1 .. R7 registers and for the instruction counter IC. A nil value for a register initializer causes the register to be set to 0. A nil value for the IC initializer causes execution to start at the beginning of the program. An optional fourth argument continue, if nonnil, causes the IC to remain on the unsatisfied read operation if the program terminates due to exhaustion of the input buffer. Otherwise the IC is set to the end of the program. When the program is done, status is modified (by side-effect) to contain the ending values for the corresponding registers and IC. Returns the resulting string.

To call a CCL program from another CCL program, it must first be registered:

Function register-ccl-program name ccl-program

Register name for CCL program program in ccl-program-table. program should be the compiled form of a CCL program, or nil. Return index number of the registered CCL program.

Information about the processor time used by the CCL interpreter can be obtained using these functions:

Function ccl-elapsed-time

Returns the elapsed processor time of the CCL interpreter as cons of user and system time, as floating point numbers measured in seconds. If only one overall value can be determined, the return value will be a cons of that value and 0.

Function ccl-reset-elapsed-time

Resets the CCL interpreter’s internal elapsed time registers.

55.7.5 CCL Examples

This section is not yet written.

55.8 Category Tables

A category table is a type of char table used for keeping track of categories. Categories are used for classifying characters for use in regexps – you can refer to a category rather than having to use a complicated [] expression (and category lookups are significantly faster).

There are 95 different categories available, one for each printable character (including space) in the ASCII charset. Each category is designated by one such character, called a category designator. They are specified in a regexp using the syntax ‘\cX’, where X is a category designator. (This is not yet implemented.)

A category table specifies, for each character, the categories that the character is in. Note that a character can be in more than one category. More specifically, a category table maps from a character to either the value nil (meaning the character is in no categories) or a 95-element bit vector, specifying for each of the 95 categories whether the character is in that category.

Special Lisp functions are provided that abstract this, so you do not have to directly manipulate bit vectors.

Function category-table-p obj

This function returns t if arg is a category table.

Function category-table &optional buffer

This function returns the current category table. This is the one specified by the current buffer, or by buffer if it is non-nil.

Function standard-category-table

This function returns the standard category table. This is the one used for new buffers.

Function copy-category-table &optional table

This function constructs a new category table and return it. It is a copy of the table, which defaults to the standard category table.

Function set-category-table table &optional buffer This function selects a new category table for buffer. One argument, a category table. buffer defaults to the current buffer if omitted.

Function category-designator-p obj

This function returns t if arg is a category designator (a char in the range ‘' '’ to ‘'~'’).

Function category-table-value-p obj

This function returns t if arg is a category table value. Valid values are nil or a bit vector of size 95.

756 XEmacs Lisp Reference Manual

Appendix A Tips and Standards

This chapter describes no additional features of XEmacs Lisp. Instead it gives advice on making effective use of the features described in the previous chapters.

A.1 Writing Clean Lisp Programs

Here are some tips for avoiding common errors in writing Lisp code intended for widespread use:

• Since all global variables share the same name space, and all functions share another name space, you should choose a short word to distinguish your program from other Lisp programs. Then take care to begin the names of all global variables, constants, and functions with the chosen prefix. This helps avoid name conflicts. This recommendation applies even to names for traditional Lisp primitives that are not primitives in XEmacs Lisp—even to cadr. Believe it or not, there is more than one plausible way to define cadr. Play it safe; append your name prefix to produce a

name like foo-cadr or mylib-cadr instead.

If you write a function that you think ought to be added to Emacs under a certain name, such as twiddle-files, don’t call it by that name in your program. Call it mylib-

twiddle-files in your program, and send mail to ‘bug-gnu-emacs@prep.ai.mit.edu’

suggesting we add it to Emacs. If and when we do, we can change the name easily enough. If one prefix is insufficient, your package may use two or three alternative common prefixes, so long as they make sense. Separate the prefix from the rest of the symbol name with a hyphen, ‘-’. This will be consistent with XEmacs itself and with most Emacs Lisp programs.

• It is often useful to put a call to provide in each separate library program, at least if there is more than one entry point to the program.

• If a file requires certain other library programs to be loaded beforehand, then the comments at the beginning of the file should say so. Also, use require to make sure they are loaded.

• If one file foo uses a macro defined in another file bar, foo should contain this expression

before the first use of the macro:

     (eval-when-compile (require 'bar))

(And bar should contain (provide 'bar), to make the require work.) This will cause bar to be loaded when you byte-compile foo. Otherwise, you risk compiling foo without the necessary macro loaded, and that would produce compiled code that won’t work right. See Section 12.3 [Compiling Macros], page 178. Using eval-when-compile avoids loading bar when the compiled version of foo is used.

• If you define a major mode, make sure to run a hook variable using run-hooks, just as the existing major modes do. See Section 26.4 [Hooks], page 378.

• If the purpose of a function is to tell you whether a certain condition is true or false, give the function a name that ends in ‘p’. If the name is one word, add just ‘p’; if the name is multiple words, add ‘-p’. Examples are framep and frame-live-p.

• If a user option variable records a true-or-false condition, give it a name that ends in

‘-flag’.

• Please do not define C-c letter as a key in your major modes. These sequences are reserved for users; they are the only sequences reserved for users, so we cannot do without them.

Instead, define sequences consisting of C-c followed by a non-letter. These sequences are reserved for major modes.

Changing all the major modes in Emacs 18 so they would follow this convention was a lot of work. Abandoning this convention would make that work go to waste, and inconvenience users.

• Sequences consisting of C-c followed by {, }, <, >, : or ; are also reserved for major modes.

• Sequences consisting of C-c followed by any other punctuation character are allocated for minor modes. Using them in a major mode is not absolutely prohibited, but if you do that, the major mode binding may be shadowed from time to time by minor modes.

• You should not bind C-h following any prefix character (including C-c). If you don’t bind C-h, it is automatically available as a help character for listing the subcommands of the prefix character.

• You should not bind a key sequence ending in ^{⟨}\underline{\overline{ESC}}^{⟩} except following another ^{⟨}\underline{\overline{ESC}}^{⟩}. (That is, it is ok to bind a sequence ending in ^{⟨}\underline{\overline{ESC}}^{⟩} ^{⟨}\underline{\overline{ESC}}^{⟩}.)

The reason for this rule is that a non-prefix binding for ^{⟨}\underline{\overline{ESC}}^{⟩} in any context prevents recognition of escape sequences as function keys in that context.

• Applications should not bind mouse events based on button 1 with the shift key held down. These events include S-mouse-1, M-S-mouse-1, C-S-mouse-1, and so on. They are reserved for users.

• Modes should redefine mouse-2 as a command to follow some sort of reference in the text of a buffer, if users usually would not want to alter the text in that buffer by hand. Modes such as Dired, Info, Compilation, and Occur redefine it in this way.

• When a package provides a modification of ordinary Emacs behavior, it is good to include a command to enable and disable the feature, Provide a command named whatever-mode which turns the feature on or off, and make it autoload (see Section 14.2 [Autoload], page 198). Design the package so that simply loading it has no visible effect—that should not enable the feature. Users will request the feature by invoking the command.

• It is a bad idea to define aliases for the Emacs primitives. Use the standard names instead.

• Redefining an Emacs primitive is an even worse idea. It may do the right thing for a particular program, but there is no telling what other programs might break as a result.

• If a file does replace any of the functions or library programs of standard XEmacs, prominent comments at the beginning of the file should say which functions are replaced, and how the behavior of the replacements differs from that of the originals.

• Please keep the names of your XEmacs Lisp source files to 13 characters or less. This way, if the files are compiled, the compiled files’ names will be 14 characters or less, which is short enough to fit on all kinds of Unix systems.

• Don’t use next-line or previous-line in programs; nearly always, forward-line is more convenient as well as more predictable and robust. See Section 34.2.4 [Text Lines], page 490.

• Don’t call functions that set the mark, unless setting the mark is one of the intended features of your program. The mark is a user-level feature, so it is incorrect to change the mark except to supply a value for the user’s benefit. See Section 35.6 [The Mark], page 504. In particular, don’t use these functions:

  • beginning-of-buffer, end-of-buffer
  • replace-string, replace-regexp

If you just want to move point, or replace a certain string, without any of the other features intended for interactive users, you can replace these functions with one or two lines of simple Lisp code.

• Use lists rather than vectors, except when there is a particular reason to use a vector. Lisp has more facilities for manipulating lists than for vectors, and working with lists is usually more convenient. Vectors are advantageous for tables that are substantial in size and are accessed in random order (not searched front to back), provided there is no need to insert or delete elements (only lists allow that).

• The recommended way to print a message in the echo area is with the message function, not princ. See Section 45.3 [The Echo Area], page 648.

• When you encounter an error condition, call the function error (or signal). The function error does not return. See Section 9.5.3.1 [Signaling Errors], page 137. Do not use message, throw, sleep-for, or beep to report errors.

• An error message should start with a capital letter but should not end with a period.

• Try to avoid using recursive edits. Instead, do what the Rmail e command does: use a new local keymap that contains one command defined to switch back to the old local keymap. Or do what the edit-options command does: switch to another buffer and let the user switch back at will. See Section 19.10 [Recursive Editing], page 310.

• In some other systems there is a convention of choosing variable names that begin and end with ‘*’. We don’t use that convention in Emacs Lisp, so please don’t use it in your programs. (Emacs uses such names only for program-generated buffers.) The users will find Emacs more coherent if all libraries use the same conventions.

• Indent each function with C-M-q (indent-sexp) using the default indentation parameters.

• Don’t make a habit of putting close-parentheses on lines by themselves; Lisp programmers find this disconcerting. Once in a while, when there is a sequence of many consecutive close-parentheses, it may make sense to split them in one or two significant places.

• Please put a copyright notice on the file if you give copies to anyone. Use the same lines that appear at the top of the Lisp files in XEmacs itself. If you have not signed papers to assign the copyright to the Foundation, then place your name in the copyright notice in place of the Foundation’s name.

A.2 Tips for Making Compiled Code Fast

Here are ways of improving the execution speed of byte-compiled Lisp programs.

• Use the ‘profile’ library to profile your program. See the file ‘profile.el’ for instructions.

• Use iteration rather than recursion whenever possible. Function calls are slow in XEmacs Lisp even when a compiled function is calling another compiled function.

• Using the primitive list-searching functions memq, member, assq, or assoc is even faster than explicit iteration. It may be worth rearranging a data structure so that one of these primitive search functions can be used.

• Certain built-in functions are handled specially in byte-compiled code, avoiding the need for an ordinary function call. It is a good idea to use these functions rather than alternatives. To see whether a function is handled specially by the compiler, examine its byte-compile property. If the property is non-nil, then the function is handled specially. For example, the following input will show you that aref is compiled specially (see Section 6.3 [Array Functions], page 104) while elt is not (see Section 6.1 [Sequence Functions], page 101):

     (get 'aref 'byte-compile)
          ⇒ byte-compile-two-args
     (get 'elt 'byte-compile)
          ⇒ nil

• If calling a small function accounts for a substantial part of your program’s running time, make the function inline. This eliminates the function call overhead. Since making a function inline reduces the flexibility of changing the program, don’t do it unless it gives a noticeable speedup in something slow enough that users care about the speed. See Section 11.9 [Inline Functions], page 175.

A.3 Tips for Documentation Strings

Here are some tips for the writing of documentation strings.

• Every command, function, or variable intended for users to know about should have a documentation string.

• An internal variable or subroutine of a Lisp program might as well have a documentation string. In earlier Emacs versions, you could save space by using a comment instead of a documentation string, but that is no longer the case.

• The first line of the documentation string should consist of one or two complete sentences that stand on their own as a summary. M-x apropos displays just the first line, and if it doesn’t stand on its own, the result looks bad. In particular, start the first line with a capital letter and end with a period. The documentation string can have additional lines that expand on the details of how to use the function or variable. The additional lines should be made up of complete sentences also, but they may be filled if that looks good.

• For consistency, phrase the verb in the first sentence of a documentation string as an infinitive with “to” omitted. For instance, use “Return the cons of A and B.” in preference to “Returns the cons of A and B.” Usually it looks good to do likewise for the rest of the first paragraph. Subsequent paragraphs usually look better if they have proper subjects.

• Write documentation strings in the active voice, not the passive, and in the present tense, not the future. For instance, use “Return a list containing A and B.” instead of “A list containing A and B will be returned.”

• Avoid using the word “cause” (or its equivalents) unnecessarily. Instead of, “Cause Emacs to display text in boldface,” write just “Display text in boldface.”

• Do not start or end a documentation string with whitespace.

• Format the documentation string so that it fits in an Emacs window on an 80-column screen. It is a good idea for most lines to be no wider than 60 characters. The first line can be wider if necessary to fit the information that ought to be there. However, rather than simply filling the entire documentation string, you can make it much more readable by choosing line breaks with care. Use blank lines between topics if the documentation string is long.

• Do not indent subsequent lines of a documentation string so that the text is lined up in the source code with the text of the first line. This looks nice in the source code, but looks bizarre when users view the documentation. Remember that the indentation before the starting double-quote is not part of the string!

• A variable’s documentation string should start with ‘*’ if the variable is one that users would often want to set interactively. If the value is a long list, or a function, or if the variable would be set only in init files, then don’t start the documentation string with ‘*’. See Section 10.5 [Defining Variables], page 149.

• The documentation string for a variable that is a yes-or-no flag should start with words such as “Non-nil means…”, to make it clear that all non-nil values are equivalent and indicate explicitly what nil and non-nil mean.

• When a function’s documentation string mentions the value of an argument of the function, use the argument name in capital letters as if it were a name for that value. Thus, the documentation string of the function / refers to its second argument as ‘DIVISOR’, because the actual argument name is divisor. Also use all caps for meta-syntactic variables, such as when you show the decomposition of a list or vector into subunits, some of which may vary.

• When a documentation string refers to a Lisp symbol, write it as it would be printed (which usually means in lower case), with single-quotes around it. For example: ‘`lambda'’. There are two exceptions: write t and nil without single-quotes.

• Don’t write key sequences directly in documentation strings. Instead, use the ‘\\[...]’ construct to stand for them. For example, instead of writing ‘C-f’, write ‘\\[forward-char]’.

When Emacs displays the documentation string, it substitutes whatever key is currently bound to forward-char. (This is normally ‘C-f’, but it may be some other character if the user has moved key bindings.) See Section 27.3 [Keys in Documentation], page 384.

• In documentation strings for a major mode, you will want to refer to the key bindings of that mode’s local map, rather than global ones. Therefore, use the construct ‘\\<...>’ once in the documentation string to specify which key map to use. Do this before the first use of ‘\\[...]’. The text inside the ‘\\<...>’ should be the name of the variable containing the local keymap for the major mode. It is not practical to use ‘\\[...]’ very many times, because display of the documentation string will become slow. So use this to describe the most important commands in your major mode, and then use ‘\\{...}’ to display the rest of the mode’s keymap.

A.4 Tips on Writing Comments

We recommend these conventions for where to put comments and how to indent them:

‘;’ Comments that start with a single semicolon, ‘;’, should all be aligned to the same column on the right of the source code. Such comments usually explain how the code on the same line does its job. In Lisp mode and related modes, the M-; (indent-for-comment) command automatically inserts such a ‘;’ in the right place, or aligns such a comment if it is already present. This and following examples are taken from the Emacs sources.

     (setq base-version-list                 ; there was a base
           (assoc (substring fn 0 start-vn)  ; version to which
                  file-version-assoc-list))  ; this looks like
                                             ; a subversion

‘;;’ Comments that start with two semicolons, ‘;;’, should be aligned to the same level of indentation as the code. Such comments usually describe the purpose of the following lines or the state of the program at that point. For example:

     (prog1 (setq auto-fill-function
                  ...
                  ...
       ;; update modeline
       (redraw-modeline)))

Every function that has no documentation string (because it is use only internally within the package it belongs to), should have instead a two-semicolon comment right before the function, explaining what the function does and how to call it properly. Explain precisely what each argument means and how the function interprets its possible values.

‘;;;’ Comments that start with three semicolons, ‘;;;’, should start at the left margin. Such comments are used outside function definitions to make general statements explaining the design principles of the program. For example:

     ;;; This Lisp code is run in XEmacs
     ;;; when it is to operate as a server
     ;;; for other processes.

Another use for triple-semicolon comments is for commenting out lines within a function. We use triple-semicolons for this precisely so that they remain at the left margin.

               (defun foo (a)
               ;;; This is no longer necessary.
               ;;;  (force-mode-line-update)
                 (message "Finished with %s" a))

‘;;;;’ Comments that start with four semicolons, ‘;;;;’, should be aligned to the left

margin and are used for headings of major sections of a program. For example:

               ;;;; The kill ring

The indentation commands of the Lisp modes in XEmacs, such as M-; (indent-for-comment) and^{⟨}\underline{\overline{TAB}}^{⟩} (lisp-indent-line) automatically indent comments according to these conventions, depending on the number of semicolons. See section “Manipulating Comments” in The XEmacs Reference Manual.

A.5 Conventional Headers for XEmacs Libraries

XEmacs has conventions for using special comments in Lisp libraries to divide them into sections and give information such as who wrote them. This section explains these conventions. First, an example:

     ;;; lisp-mnt.el --- minor mode for Emacs Lisp maintainers
     ;; Copyright (C) 1992 Free Software Foundation, Inc.
     ;; Author: Eric S. Raymond <esr@snark.thyrsus.com>
     ;; Maintainer: Eric S. Raymond <esr@snark.thyrsus.com>
     ;; Created: 14 Jul 1992
     ;; Version: 1.2
     ;; Keywords: docs
     ;; This file is part of XEmacs.

copying permissions... The very first line should have this format:

;;; filename --- description The description should be complete in one line.

After the copyright notice come several header comment lines, each beginning with ‘;; header-name:’. Here is a table of the conventional possibilities for header-name:

‘Author’ This line states the name and net address of at least the principal author of the library. If there are multiple authors, you can list them on continuation lines led by ;; and a tab character, like this:

               ;; Author: Ashwin Ram <Ram-Ashwin@cs.yale.edu>
               ;;      Dave Sill <de5@ornl.gov>
               ;;      Dave Brennan <brennan@hal.com>
               ;;      Eric Raymond <esr@snark.thyrsus.com>

‘Maintainer’

This line should contain a single name/address as in the Author line, or an address only, or the string ‘FSF’. If there is no maintainer line, the person(s) in the Author field are presumed to be the maintainers. The example above is mildly bogus because the maintainer line is redundant. The idea behind the ‘Author’ and ‘Maintainer’ lines is to make possible a Lisp function to “send mail to the maintainer” without having to mine the name out by hand. Be sure to surround the network address with ‘<...>’ if you include the person’s full name as well as the network address.

‘Created’ This optional line gives the original creation date of the file. For historical interest only.

‘Version’ If you wish to record version numbers for the individual Lisp program, put them in this line.

‘Adapted-By’

In this header line, place the name of the person who adapted the library for installation (to make it fit the style conventions, for example).

‘Keywords’

This line lists keywords for the finder-by-keyword help command. This field is important; it’s how people will find your package when they’re looking for things by topic area. To separate the keywords, you can use spaces, commas, or both.

Just about every Lisp library ought to have the ‘Author’ and ‘Keywords’ header comment lines. Use the others if they are appropriate. You can also put in header lines with other header names—they have no standard meanings, so they can’t do any harm.

We use additional stylized comments to subdivide the contents of the library file. Here is a table of them:

‘;;; Commentary:’

This begins introductory comments that explain how the library works. It should come right after the copying permissions.

‘;;; Change log:’

This begins change log information stored in the library file (if you store the change history there). For most of the Lisp files distributed with XEmacs, the change history is kept in the file ‘ChangeLog’ and not in the source file at all; these files do not have a ‘;;; Change log:’ line.

‘;;; Code:’

This begins the actual code of the program.

‘;;; filename ends here’

This is the footer line; it appears at the very end of the file. Its purpose is to enable people to detect truncated versions of the file from the lack of a footer line.

Appendix B Building XEmacs; Allocation of Objects

This chapter describes how the runnable XEmacs executable is dumped with the preloaded Lisp libraries in it and how storage is allocated.

There is an entire separate document, the XEmacs Internals Manual, devoted to the internals of XEmacs from the perspective of the C programmer. It contains much more detailed information about the build process, the allocation and garbage-collection process, and other aspects related to the internals of XEmacs.

B.1 Building XEmacs

This section explains the steps involved in building the XEmacs executable. You don’t have to know this material to build and install XEmacs, since the makefiles do all these things automatically. This information is pertinent to XEmacs maintenance.

The XEmacs Internals Manual contains more information about this.

Compilation of the C source files in the ‘src’ directory produces an executable file called ‘temacs’, also called a bare impure XEmacs. It contains the XEmacs Lisp interpreter and I/O routines, but not the editing commands.

Before XEmacs is actually usable, a number of Lisp files need to be loaded. These define all the editing commands, plus most of the startup code and many very basic Lisp primitives. This is accomplished by loading the file ‘loadup.el’, which in turn loads all of the other standardly-loaded Lisp files.

It takes a substantial time to load the standard Lisp files. Luckily, you don’t have to do this each time you run XEmacs; ‘temacs’ can dump out an executable program called ‘xemacs’ that has these files preloaded. ‘xemacs’ starts more quickly because it does not need to load the files. This is the XEmacs executable that is normally installed.

To create ‘xemacs’, use the command ‘temacs -batch -l loadup dump’. The purpose of ‘-batch’ here is to tell ‘temacs’ to run in non-interactive, command-line mode. (‘temacs’ can only run in this fashion. Part of the code required to initialize frames and faces is in Lisp, and must be loaded before XEmacs is able to create any frames.) The argument ‘dump’ tells ‘loadup.el’ to dump a new executable named ‘xemacs’.

The dumping process is highly system-specific, and some operating systems don’t support dumping. On those systems, you must start XEmacs with the ‘temacs -batch -l loadup run-temacs’ command each time you use it. This takes a substantial time, but since you need to start Emacs once a day at most—or once a week if you never log out—the extra time is not too severe a problem. (In older versions of Emacs, you started Emacs

from ‘temacs’ using ‘temacs -l loadup’.)

You are free to start XEmacs directly from ‘temacs’ if you want, even if there is already a dumped ‘xemacs’. Normally you wouldn’t want to do that; but the Makefiles do this when you rebuild XEmacs using ‘make all-elc’, which builds XEmacs and simultaneously compiles any out-of-date Lisp files. (You need ‘xemacs’ in order to compile Lisp files. However, you also need the compiled Lisp files in order to dump out ‘xemacs’. If both of these are missing or corrupted, you are out of luck unless you’re able to bootstrap ‘xemacs’ from ‘temacs’. Note that ‘make all-elc’ actually loads the alternative loadup file ‘loadup-el.el’, which works like ‘loadup.el’ but disables the pure-copying process and forces XEmacs to ignore any compiled Lisp files even if they exist.)

You can specify additional files to preload by writing a library named ‘site-load.el’ that loads them. You may need to increase the value of PURESIZE, in ‘src/puresize.h’, to make room for the additional files. You should not modify this file directly, however; instead, use the ‘--puresize’ configuration option. (If you run out of pure space while dumping ‘xemacs’, you will be told how much pure space you actually will need.) However, the advantage of preloading additional files decreases as machines get faster. On modern machines, it is often not advisable, especially if the Lisp code is on a file system local to the machine running XEmacs.

You can specify other Lisp expressions to execute just before dumping by putting them in a library named ‘site-init.el’. However, if they might alter the behavior that users expect from an ordinary unmodified XEmacs, it is better to put them in ‘default.el’, so that users can override them if they wish. See Section 50.1.1 [Start-up Summary], page 691.

Before ‘loadup.el’ dumps the new executable, it finds the documentation strings for primitive and preloaded functions (and variables) in the file where they are stored, by calling Snarf-documentation (see Section 27.2 [Accessing Documentation], page 382). These strings were moved out of the ‘xemacs’ executable to make it smaller. See Section 27.1 [Documentation Basics], page 381.

Function dump-emacs to-file from-file This function dumps the current state of XEmacs into an executable file to-file. It takes symbols from from-file (this is normally the executable file ‘temacs’).

If you use this function in an XEmacs that was already dumped, you must set command-line-processed to nil first for good results. See Section 50.1.4 [Command Line Arguments], page 694.

Function run-emacs-from-temacs &rest args

This is the function that implements the ‘run-temacs’ command-line argument. It is called from ‘loadup.el’ as appropriate. You should most emphatically not call this yourself; it will reinitialize your XEmacs process and you’ll be sorry.

Command emacs-version

This function returns a string describing the version of XEmacs that is running.

It is useful to include this string in bug reports.

          (emacs-version)
            ⇒ "XEmacs 20.1 [Lucid] (i586-unknown-linux2.0.29)
                           of Mon Apr  7 1997 on altair.xemacs.org"

Called interactively, the function prints the same information in the echo area.

Variable emacs-build-time

The value of this variable is the time at which XEmacs was built at the local site.

          emacs-build-time "Mon Apr  7 20:28:52 1997"

⇒

Variable emacs-version

The value of this variable is the version of Emacs being run. It is a string, e.g.

"20.1 XEmacs Lucid".

The following two variables did not exist before FSF GNU Emacs version 19.23 and XEmacs version 19.10, which reduces their usefulness at present, but we hope they will be convenient in the future.

Variable emacs-major-version

The major version number of Emacs, as an integer. For XEmacs version 20.1, the value is 20.

Variable emacs-minor-version

The minor version number of Emacs, as an integer. For XEmacs version 20.1, the value is 1.

B.2 Pure Storage

XEmacs Lisp uses two kinds of storage for user-created Lisp objects: normal storage and pure storage. Normal storage is where all the new data created during an XEmacs session is kept; see the following section for information on normal storage. Pure storage is used for certain data in the preloaded standard Lisp files—data that should never change during actual use of XEmacs.

Pure storage is allocated only while ‘temacs’ is loading the standard preloaded Lisp libraries. In the file ‘xemacs’, it is marked as read-only (on operating systems that permit this), so that the memory space can be shared by all the XEmacs jobs running on the machine at once. Pure storage is not expandable; a fixed amount is allocated when XEmacs is compiled, and if that is not sufficient for the preloaded libraries, ‘temacs’ aborts with an error message. If that happens, you must increase the compilation parameter PURESIZE using the ‘--puresize’ option to ‘configure’. This normally won’t happen unless you try to preload additional libraries or add features to the standard ones.

Function purecopy object

This function makes a copy of object in pure storage and returns it. It copies strings by simply making a new string with the same characters in pure storage. It recursively copies the contents of vectors and cons cells. It does not make copies of other objects such as symbols, but just returns them unchanged. It signals an error if asked to copy markers.

This function is a no-op except while XEmacs is being built and dumped; it is usually called only in the file ‘xemacs/lisp/prim/loaddefs.el’, but a few packages call it just in case you decide to preload them.

Variable pure-bytes-used

The value of this variable is the number of bytes of pure storage allocated so far. Typically, in a dumped XEmacs, this number is very close to the total amount of pure storage available—if it were not, we would preallocate less.

Variable purify-flag

This variable determines whether defun should make a copy of the function definition in pure storage. If it is non-nil, then the function definition is copied into pure storage.

This flag is t while loading all of the basic functions for building XEmacs initially (allowing those functions to be sharable and non-collectible). Dumping XEmacs as an executable always writes nil in this variable, regardless of the value it actually has before and after dumping.

You should not change this flag in a running XEmacs.

B.3 Garbage Collection

When a program creates a list or the user defines a new function (such as by loading a library), that data is placed in normal storage. If normal storage runs low, then XEmacs asks the operating system to allocate more memory in blocks of 2k bytes. Each block is used for one type of Lisp object, so symbols, cons cells, markers, etc., are segregated in distinct blocks in memory. (Vectors, long strings, buffers and certain other editing types, which are fairly large, are allocated in individual blocks, one per object, while small strings are packed into blocks of 8k bytes. [More correctly, a string is allocated in two sections: a fixed size chunk containing the length, list of extents, etc.; and a chunk containing the actual characters in the string. It is this latter chunk that is either allocated individually or packed into 8k blocks. The fixed size chunk is packed into 2k blocks, as for conses, markers, etc.])

It is quite common to use some storage for a while, then release it by (for example) killing a buffer or deleting the last pointer to an object. XEmacs provides a garbage collector to reclaim this abandoned storage. (This name is traditional, but “garbage recycler” might be a more intuitive metaphor for this facility.)

The garbage collector operates by finding and marking all Lisp objects that are still accessible to Lisp programs. To begin with, it assumes all the symbols, their values and associated function definitions, and any data presently on the stack, are accessible. Any objects that can be reached indirectly through other accessible objects are also accessible.

When marking is finished, all objects still unmarked are garbage. No matter what the Lisp program or the user does, it is impossible to refer to them, since there is no longer a way to reach them. Their space might as well be reused, since no one will miss them. The second (“sweep”) phase of the garbage collector arranges to reuse them.

The sweep phase puts unused cons cells onto a free list for future allocation; likewise for symbols, markers, extents, events, floats, compiled-function objects, and the fixed-size portion of strings. It compacts the accessible small string-chars chunks so they occupy fewer 8k blocks; then it frees the other 8k blocks. Vectors, buffers, windows, and other large objects are individually allocated and freed using malloc and free.

Common Lisp note: unlike other Lisps, XEmacs Lisp does not call the garbage collector when the free list is empty. Instead, it simply requests the operating system to allocate more storage, and processing continues until gc-cons-threshold bytes have been used.

This means that you can make sure that the garbage collector will not run during a certain portion of a Lisp program by calling the garbage collector explicitly just before it (provided that portion of the program does not use so much space as to force a second garbage collection).

Command garbage-collect

This command runs a garbage collection, and returns information on the amount of space in use. (Garbage collection can also occur spontaneously if you use more than gc-cons-threshold bytes of Lisp data since the previous garbage collection.)

garbage-collect returns a list containing the following information:

((used-conses . free-conses) (used-syms . free-syms) (used-markers . free-markers) used-string-chars used-vector-slots (plist))

⇒ ((73362 . 8325) (13718 . 164)
(5089 . 5098) 949121 118677
(conses-used 73362 conses-free 8329 cons-storage 658168
symbols-used 13718 symbols-free 164 symbol-storage 335216
bit-vectors-used 0 bit-vectors-total-length 0
bit-vector-storage 0 vectors-used 7882
vectors-total-length 118677 vector-storage 537764
compiled-functions-used 1336 compiled-functions-free 37
compiled-function-storage 44440 short-strings-used 28829
long-strings-used 2 strings-free 7722
short-strings-total-length 916657 short-string-storage 1179648
long-strings-total-length 32464 string-header-storage 441504
floats-used 3 floats-free 43 float-storage 2044 markers-used 5089
markers-free 5098 marker-storage 245280 events-used 103
events-free 835 event-storage 110656 extents-used 10519
extents-free 2718 extent-storage 372736
extent-auxiliarys-used 111 extent-auxiliarys-freed 3
extent-auxiliary-storage 4440 window-configurations-used 39
window-configurations-on-free-list 5
window-configurations-freed 10 window-configuration-storage 9492
popup-datas-used 3 popup-data-storage 72 toolbar-buttons-used 62
toolbar-button-storage 4960 toolbar-datas-used 12
toolbar-data-storage 240 symbol-value-buffer-locals-used 182
symbol-value-buffer-local-storage 5824
symbol-value-lisp-magics-used 22
symbol-value-lisp-magic-storage 1496
symbol-value-varaliases-used 43
symbol-value-varalias-storage 1032 opaque-lists-used 2
opaque-list-storage 48 color-instances-used 12
color-instance-storage 288 font-instances-used 5
font-instance-storage 180 opaques-used 11 opaque-storage 312
range-tables-used 1 range-table-storage 16 faces-used 34
face-storage 2584 glyphs-used 124 glyph-storage 4464
specifiers-used 775 specifier-storage 43869 weak-lists-used 786
weak-list-storage 18864 char-tables-used 40
char-table-storage 41920 buffers-used 25 buffer-storage 7000
extent-infos-used 457 extent-infos-freed 73
extent-info-storage 9140 keymaps-used 275 keymap-storage 12100
consoles-used 4 console-storage 384 command-builders-used 2
command-builder-storage 120 devices-used 2 device-storage 344
frames-used 3 frame-storage 624 image-instances-used 47
image-instance-storage 3008 windows-used 27 windows-freed 2
window-storage 9180 lcrecord-lists-used 15
lcrecord-list-storage 360 hashtables-used 631
hashtable-storage 25240 streams-used 1 streams-on-free-list 3
streams-freed 12 stream-storage 91))

Here is a table explaining each element: used-conses

The number of cons cells in use.

free-conses The number of cons cells for which space has been obtained from the operating system, but that are not currently being used.

used-syms The number of symbols in use.

free-syms The number of symbols for which space has been obtained from the operating system, but that are not currently being used.

used-markers The number of markers in use.

free-markers The number of markers for which space has been obtained from the operating system, but that are not currently being used.

used-string-chars The total size of all strings, in characters.

used-vector-slots The total number of elements of existing vectors.

plist A list of alternating keyword/value pairs providing more detailed information. (As you can see above, quite a lot of information is provided.)

User Option gc-cons-threshold

The value of this variable is the number of bytes of storage that must be allocated for Lisp objects after one garbage collection in order to trigger another garbage collection. A cons cell counts as eight bytes, a string as one byte per character plus a few bytes of overhead, and so on; space allocated to the contents of buffers does not count. Note that the subsequent garbage collection does not happen immediately when the threshold is exhausted, but only the next time the Lisp evaluator is called. The initial threshold value is 500,000. If you specify a larger value, garbage collection will happen less often. This reduces the amount of time spent garbage collecting, but increases total memory use. You may want to do this when running a program that creates lots of Lisp data. You can make collections more frequent by specifying a smaller value, down to 10,000. A value less than 10,000 will remain in effect only until the subsequent garbage collection, at which time garbage-collect will set the threshold back to 10,000. (This does not apply if XEmacs was configured with ‘--debug’. Therefore, be careful when setting gc-cons-threshold in that case!)

Function memory-limit

This function returns the address of the last byte XEmacs has allocated, divided by 1024. We divide the value by 1024 to make sure it fits in a Lisp integer. You can use this to get a general idea of how your actions affect the memory usage.

Variable pre-gc-hook

This is a normal hook to be run just before each garbage collection. Interrupts, garbage collection, and errors are inhibited while this hook runs, so be extremely careful in what you add here. In particular, avoid consing, and do not interact with the user.

Variable post-gc-hook

This is a normal hook to be run just after each garbage collection. Interrupts, garbage collection, and errors are inhibited while this hook runs, so be extremely careful in what you add here. In particular, avoid consing, and do not interact with the user.

Variable gc-message

This is a string to print to indicate that a garbage collection is in progress. This is printed in the echo area. If the selected frame is on a window system and gcpointer-glyph specifies a value (i.e. a pointer image instance) in the domain of the selected frame, the mouse cursor will change instead of this message being printed.

Glyph gc-pointer-glyph

This holds the pointer glyph used to indicate that a garbage collection is in progress. If the selected window is on a window system and this glyph specifies a value (i.e. a pointer image instance) in the domain of the selected window, the cursor will be changed as specified during garbage collection. Otherwise, a message will be printed in the echo area, as controlled by gc-message. See Chapter 43 [Glyphs], page 625.

If XEmacs was configured with ‘--debug’, you can set the following two variables to get direct information about all the allocation that is happening in a segment of Lisp code.

Variable debug-allocation

If non-zero, print out information to stderr about all objects allocated.

Variable debug-allocation-backtrace

Length (in stack frames) of short backtrace printed out by debug-allocation.

Appendix C Standard Errors

Here is the complete list of the error symbols in standard Emacs, grouped by concept. The list includes each symbol’s message (on the error-message property of the symbol) and a cross reference to a description of how the error can occur.

Each error symbol has an error-conditions property that is a list of symbols. Normally this list includes the error symbol itself and the symbol error. Occasionally it includes additional symbols, which are intermediate classifications, narrower than error but broader than a single error symbol. For example, all the errors in accessing files have the condition

file-error.

As a special exception, the error symbol quit does not have the condition error, because quitting is not considered an error.

See Section 9.5.3 [Errors], page 136, for an explanation of how errors are generated and handled.

symbol string; reference.

error     "error"

See Section 9.5.3 [Errors], page 136.

quit      "Quit"

See Section 19.8 [Quitting], page 306.

args-out-of-range
          "Args out of range"

See Chapter 6 [Sequences Arrays Vectors], page 101.

arith-error
          "Arithmetic error"

See / and % in Chapter 3 [Numbers], page 45.

beginning-of-buffer
          "Beginning of buffer"

See Section 34.2 [Motion], page 488.

buffer-read-only
          "Buffer is read-only"

See Section 30.7 [Read Only Buffers], page 438.

cyclic-function-indirection
          "Symbol's chain of function indirections contains a loop"

See Section 8.2.4 [Function Indirection], page 123.

domain-error
          "Arithmetic domain error"

end-of-buffer
          "End of buffer"

See Section 34.2 [Motion], page 488.

end-of-file
"End of file during parsing"

This is not a file-error. See Section 17.3 [Input Functions], page 253.

file-error

This error and its subcategories do not have error-strings, because the error message is constructed from the data items alone when the error condition

file-error is present.

See Chapter 28 [Files], page 391.

file-locked
This is a file-error.

See Section 28.5 [File Locks], page 398.

file-already-exists
This is a file-error.

See Section 28.4 [Writing to Files], page 397.

file-supersession
This is a file-error.

See Section 30.6 [Modification Time], page 437.

invalid-function
"Invalid function"

See Section 8.2.3 [Classifying Lists], page 123.

invalid-read-syntax
"Invalid read syntax"

See Section 17.3 [Input Functions], page 253.

invalid-regexp
"Invalid regexp"

See Section 37.2 [Regular Expressions], page 550.

mark-inactive
"The mark is not active now"
no-catch  "No catch for tag"

See Section 9.5.1 [Catch and Throw], page 134.

overflow-error
"Arithmetic overflow error"
protected-field
"Attempt to modify a protected field"
range-error
"Arithmetic range error"

search-failed
"Search failed"

See Chapter 37 [Searching and Matching], page 549.

setting-constant
"Attempt to set a constant symbol"

See Section 10.2 [Variables that Never Change], page 145.

singularity-error
"Arithmetic singularity error"
tooltalk-error
"ToolTalk error"

See Chapter 52 [ToolTalk Support], page 719.

undefined-keystroke-sequence
"Undefined keystroke sequence"
void-function
"Symbol's function definition is void"

See Section 11.8 [Function Cells], page 173.

void-variable
"Symbol's value as variable is void"

See Section 10.6 [Accessing Variables], page 151.

wrong-number-of-arguments
"Wrong number of arguments"

See Section 8.2.3 [Classifying Lists], page 123.

wrong-type-argument
"Wrong type argument"

See Section 2.7 [Type Predicates], page 35.

These error types, which are all classified as special cases of arith-error, can occur on certain systems for invalid use of mathematical functions.

domain-error
"Arithmetic domain error"

See Section 3.9 [Math Functions], page 56.

overflow-error
"Arithmetic overflow error"

See Section 3.9 [Math Functions], page 56.

range-error
"Arithmetic range error"

See Section 3.9 [Math Functions], page 56.

singularity-error
"Arithmetic singularity error"

See Section 3.9 [Math Functions], page 56.

underflow-error
          "Arithmetic underflow error"

See Section 3.9 [Math Functions], page 56.

Appendix D Buffer-Local Variables

The table below lists the general-purpose Emacs variables that are automatically local (when set) in each buffer. Many Lisp packages define such variables for their internal use; we don’t list them here.

abbrev-mode

see Chapter 39 [Abbrevs], page 579

auto-fill-function

see Section 36.13 [Auto Filling], page 529

buffer-auto-save-file-name

see Section 29.2 [Auto-Saving], page 425

buffer-backed-up

see Section 29.1 [Backup Files], page 421

buffer-display-table

see Section 45.11 [Display Tables], page 659

buffer-file-format

see Section 28.13 [Format Conversion], page 417

buffer-file-name

see Section 30.4 [Buffer File Name], page 434

buffer-file-number

see Section 30.4 [Buffer File Name], page 434

buffer-file-truename

see Section 30.4 [Buffer File Name], page 434

buffer-file-type

see Section 28.14 [Files and MS-DOS], page 419

buffer-invisibility-spec

see Section 45.5 [Invisible Text], page 653

buffer-offer-save

see Section 28.2 [Saving Buffers], page 394

buffer-read-only

see Section 30.7 [Read Only Buffers], page 438

buffer-saved-size

see Section 34.1 [Point], page 487

buffer-undo-list

see Section 36.9 [Undo], page 523

cache-long-line-scans

see Section 34.2.4 [Text Lines], page 490

case-fold-search

see Section 37.7 [Searching and Case], page 565

ctl-arrow

see Section 45.10 [Usual Display], page 658

comment-column

see section “Comments” in The XEmacs User’s Manual

default-directory

see Section 50.3 [System Environment], page 697

defun-prompt-regexp

see Section 34.2.6 [List Motion], page 493

fill-column

see Section 36.13 [Auto Filling], page 529

goal-column

see section “Moving Point” in The XEmacs User’s Manual

left-margin

see Section 36.16 [Indentation], page 534

local-abbrev-table

see Chapter 39 [Abbrevs], page 579

local-write-file-hooks

see Section 28.2 [Saving Buffers], page 394

major-mode

see Section 26.1.4 [Mode Help], page 369

mark-active

see Section 35.6 [The Mark], page 504

mark-ring

see Section 35.6 [The Mark], page 504

minor-modes

see Section 26.2 [Minor Modes], page 371

modeline-format

see Section 26.3.1 [Modeline Data], page 373

modeline-buffer-identification

see Section 26.3.2 [Modeline Variables], page 374

modeline-format

see Section 26.3.1 [Modeline Data], page 373

modeline-modified

see Section 26.3.2 [Modeline Variables], page 374

modeline-process

see Section 26.3.2 [Modeline Variables], page 374

mode-name

see Section 26.3.2 [Modeline Variables], page 374

overwrite-mode

see Section 36.4 [Insertion], page 513

paragraph-separate

see Section 37.8 [Standard Regexps], page 566

paragraph-start

see Section 37.8 [Standard Regexps], page 566

point-before-scroll

Used for communication between mouse commands and scroll-bar commands.

require-final-newline

see Section 36.4 [Insertion], page 513

selective-display

see Section 45.6 [Selective Display], page 654

selective-display-ellipses

see Section 45.6 [Selective Display], page 654

tab-width

see Section 45.10 [Usual Display], page 658

truncate-lines

see Section 45.2 [Truncation], page 648

vc-mode see Section 26.3.2 [Modeline Variables], page 374 780 XEmacs Lisp Reference Manual

Appendix E Standard Keymaps

The following symbols are used as the names for various keymaps. Some of these exist when XEmacs is first started, others are loaded only when their respective mode is used. This is not an exhaustive list.

Almost all of these maps are used as local maps. Indeed, of the modes that presently exist, only Vip mode and Terminal mode ever change the global keymap.

bookmark-map

A keymap containing bindings to bookmark functions.

Buffer-menu-mode-map

A keymap used by Buffer Menu mode.

c++-mode-map

A keymap used by C++ mode.

c-mode-map

A keymap used by C mode. A sparse keymap used by C mode.

command-history-map

A keymap used by Command History mode.

ctl-x-4-map

A keymap for subcommands of the prefix C-x 4.

ctl-x-5-map

A keymap for subcommands of the prefix C-x 5.

ctl-x-map

A keymap for C-x commands.

debugger-mode-map

A keymap used by Debugger mode.

dired-mode-map

A keymap for dired-mode buffers.

edit-abbrevs-map

A keymap used in edit-abbrevs.

edit-tab-stops-map
          A keymap used in edit-tab-stops.
electric-buffer-menu-mode-map

A keymap used by Electric Buffer Menu mode.

electric-history-map

A keymap used by Electric Command History mode.

emacs-lisp-mode-map

A keymap used by Emacs Lisp mode.

help-map A keymap for characters following the Help key.

Helper-help-map

A keymap used by the help utility package. It has the same keymap in its value cell and in its function cell.

Info-edit-map

A keymap used by the e command of Info.

Info-mode-map

A keymap containing Info commands.

isearch-mode-map

A keymap that defines the characters you can type within incremental search.

itimer-edit-map

A keymap used when in Itimer Edit mode.

lisp-interaction-mode-map

A keymap used by Lisp mode.

lisp-mode-map

A keymap used by Lisp mode. A keymap for minibuffer input with completion.

minibuffer-local-isearch-map

A keymap for editing isearch strings in the minibuffer.

minibuffer-local-map

Default keymap to use when reading from the minibuffer.

minibuffer-local-must-match-map

A keymap for minibuffer input with completion, for exact match.

mode-specific-map

The keymap for characters following C-c. Note, this is in the global map. This map is not actually mode specific: its name was chosen to be informative for the user in C-h b (display-bindings), where it describes the main use of the C-c prefix key.

modeline-map

The keymap consulted for mouse-clicks on the modeline of a window.

objc-mode-map

A keymap used in Objective C mode as a local map.

occur-mode-map

A local keymap used by Occur mode.

overriding-local-map

A keymap that overrides all other local keymaps.

query-replace-map

A local keymap used for responses in query-replace and related commands; also for y-or-n-p and map-y-or-n-p. The functions that use this map do not support prefix keys; they look up one event at a time.

read-expression-map

The minibuffer keymap used for reading Lisp expressions.

read-shell-command-map

The minibuffer keymap used by shell-command and related commands.

shared-lisp-mode-map

A keymap for commands shared by all sorts of Lisp modes.

text-mode-map

A keymap used by Text mode.

toolbar-map

The keymap consulted for mouse-clicks over a toolbar.

view-mode-map

A keymap used by View mode.

784 XEmacs Lisp Reference Manual

Appendix F Standard Hooks

The following is a list of hook variables that let you provide functions to be called from within Emacs on suitable occasions.

Most of these variables have names ending with ‘-hook’. They are normal hooks, run by means of run-hooks. The value of such a hook is a list of functions. The recommended way to put a new function on such a hook is to call add-hook. See Section 26.4 [Hooks], page 378, for more information about using hooks.

The variables whose names end in ‘-function’ have single functions as their values. Usually there is a specific reason why the variable is not a normal hook, such as the need to pass arguments to the function. (In older Emacs versions, some of these variables had names ending in ‘-hook’ even though they were not normal hooks.)

The variables whose names end in ‘-hooks’ or ‘-functions’ have lists of functions as their values, but these functions are called in a special way (they are passed arguments, or else their values are used).

activate-menubar-hook
activate-popup-menu-hook
ad-definition-hooks
adaptive-fill-function
add-log-current-defun-function
after-change-functions
after-delete-annotation-hook
after-init-hook
after-insert-file-functions
after-revert-hook
after-save-hook
after-set-visited-file-name-hooks
after-write-file-hooks
auto-fill-function
auto-save-hook
before-change-functions
before-delete-annotation-hook
before-init-hook
before-revert-hook
blink-paren-function
buffers-menu-switch-to-buffer-function
c++-mode-hook
c-delete-function
c-mode-common-hook
c-mode-hook

c-special-indent-hook
calendar-load-hook
change-major-mode-hook
command-history-hook
comment-indent-function
compilation-buffer-name-function
compilation-exit-message-function
compilation-finish-function
compilation-parse-errors-function
compilation-mode-hook
create-console-hook
create-device-hook
create-frame-hook
dabbrev-friend-buffer-function
dabbrev-select-buffers-function
delete-console-hook
delete-device-hook
delete-frame-hook
deselect-frame-hook
diary-display-hook
diary-hook
dired-after-readin-hook
dired-before-readin-hook
dired-load-hook
dired-mode-hook
disabled-command-hook
display-buffer-function
ediff-after-setup-control-frame-hook
ediff-after-setup-windows-hook
ediff-before-setup-control-frame-hook
ediff-before-setup-windows-hook
ediff-brief-help-message-function
ediff-cleanup-hook
ediff-control-frame-position-function
ediff-display-help-hook
ediff-focus-on-regexp-matches-function
ediff-forward-word-function
ediff-hide-regexp-matches-function

ediff-keymap-setup-hook
ediff-load-hook
ediff-long-help-message-function
ediff-make-wide-display-function
ediff-merge-split-window-function
ediff-meta-action-function
ediff-meta-redraw-function
ediff-mode-hook
ediff-prepare-buffer-hook
ediff-quit-hook
ediff-registry-setup-hook
ediff-select-hook
ediff-session-action-function
ediff-session-group-setup-hook
ediff-setup-diff-regions-function
ediff-show-registry-hook
ediff-show-session-group-hook
ediff-skip-diff-region-function
ediff-split-window-function
ediff-startup-hook
ediff-suspend-hook
ediff-toggle-read-only-function
ediff-unselect-hook
ediff-window-setup-function
edit-picture-hook
electric-buffer-menu-mode-hook
electric-command-history-hook
electric-help-mode-hook
emacs-lisp-mode-hook
fill-paragraph-function
find-file-hooks
find-file-not-found-hooks
first-change-hook
font-lock-after-fontify-buffer-hook
font-lock-beginning-of-syntax-function
font-lock-mode-hook
fume-found-function-hook
fume-list-mode-hook

fume-rescan-buffer-hook
fume-sort-function
gnus-startup-hook
hack-local-variables-hook
highlight-headers-follow-url-function
hyper-apropos-mode-hook
indent-line-function
indent-mim-hook
indent-region-function
initial-calendar-window-hook
isearch-mode-end-hook
isearch-mode-hook
java-mode-hook
kill-buffer-hook
kill-buffer-query-functions
kill-emacs-hook
kill-emacs-query-functions
kill-hooks
LaTeX-mode-hook
latex-mode-hook
ledit-mode-hook
lisp-indent-function
lisp-interaction-mode-hook
lisp-mode-hook
list-diary-entries-hook
load-read-function
log-message-filter-function
m2-mode-hook
mail-citation-hook
mail-mode-hook
mail-setup-hook
make-annotation-hook
makefile-mode-hook
map-frame-hook
mark-diary-entries-hook
medit-mode-hook
menu-no-selection-hook
mh-compose-letter-hook

mh-folder-mode-hook
mh-letter-mode-hook
mim-mode-hook
minibuffer-exit-hook
minibuffer-setup-hook
mode-motion-hook
mouse-enter-frame-hook
mouse-leave-frame-hook
mouse-track-cleanup-hook
mouse-track-click-hook
mouse-track-down-hook
mouse-track-drag-hook
mouse-track-drag-up-hook
mouse-track-up-hook
mouse-yank-function
news-mode-hook
news-reply-mode-hook
news-setup-hook
nongregorian-diary-listing-hook
nongregorian-diary-marking-hook
nroff-mode-hook
objc-mode-hook
outline-mode-hook
perl-mode-hook
plain-TeX-mode-hook
post-command-hook
post-gc-hook
pre-abbrev-expand-hook
pre-command-hook
pre-display-buffer-function
pre-gc-hook
pre-idle-hook
print-diary-entries-hook
prolog-mode-hook
protect-innocence-hook
remove-message-hook
revert-buffer-function
revert-buffer-insert-contents-function

rmail-edit-mode-hook
rmail-mode-hook
rmail-retry-setup-hook
rmail-summary-mode-hook
scheme-indent-hook
scheme-mode-hook
scribe-mode-hook
select-frame-hook
send-mail-function
shell-mode-hook
shell-set-directory-error-hook
special-display-function
suspend-hook
suspend-resume-hook
temp-buffer-show-function
term-setup-hook
terminal-mode-hook
terminal-mode-break-hook
TeX-mode-hook
tex-mode-hook
text-mode-hook
today-visible-calendar-hook
today-invisible-calendar-hook
tooltalk-message-handler-hook
tooltalk-pattern-handler-hook
tooltalk-unprocessed-message-hook
unmap-frame-hook
vc-checkin-hook
vc-checkout-writable-buffer-hook
vc-log-after-operation-hook
vc-make-buffer-writable-hook
view-hook
vm-arrived-message-hook
vm-arrived-messages-hook
vm-chop-full-name-function
vm-display-buffer-hook
vm-edit-message-hook
vm-forward-message-hook

vm-iconify-frame-hook
vm-inhibit-write-file-hook
vm-key-functions
vm-mail-hook
vm-mail-mode-hook
vm-menu-setup-hook
vm-mode-hook
vm-quit-hook
vm-rename-current-buffer-function
vm-reply-hook
vm-resend-bounced-message-hook
vm-resend-message-hook
vm-retrieved-spooled-mail-hook
vm-select-message-hook
vm-select-new-message-hook
vm-select-unread-message-hook
vm-send-digest-hook
vm-summary-mode-hook
vm-summary-pointer-update-hook
vm-summary-redo-hook
vm-summary-update-hook
vm-undisplay-buffer-hook
vm-visit-folder-hook
window-setup-hook
write-contents-hooks
write-file-data-hooks
write-file-hooks
write-region-annotate-functions
x-lost-selection-hooks
x-sent-selection-hooks
zmacs-activate-region-hook
zmacs-deactivate-region-hook
zmacs-update-region-hook

792 XEmacs Lisp Reference Manual

Index 793

Index

All variables, functions, keys, programs, files, and concepts are in this one index. All names and concepts are permuted, so they appear several times, one for each per-

mutation of the parts of the name. For example, function-name would appear as functionname and name, function-. Key entries are not permuted, however.

#

#$ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 #@count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

$

$ in display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 $ in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555

%

% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 % in format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

&

& in replacement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 &define (Edebug). . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 &not (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 &optional . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 &optional (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . 244 &or (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 &rest. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 &rest (Edebug). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244

’

’ for quoting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

(

( in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 (?: in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 (cf. no-redraw-on-reenter), resume . . . . . . . . . . . . 649 (cf. no-redraw-on-reenter), suspend . . . . . . . . . . . 649 (documentation) file, DOC . . . . . . . . . . . . . . . . . . . 383 (Edebug), &define . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 (Edebug), &not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 (Edebug), &optional. . . . . . . . . . . . . . . . . . . . . . . . . . 244 (Edebug), &or. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 (Edebug), &rest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 (Edebug), ‘ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 (Edebug), anonymous lambda expressions. . . . . 231 (Edebug), backquote . . . . . . . . . . . . . . . . . . . . . . . . . 247 (Edebug), cl.el. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 (Edebug), Common Lisp. . . . . . . . . . . . . . . . . . . . . . 231 (Edebug), current buffer point and mark . . . . . . 242 (Edebug), dotted lists . . . . . . . . . . . . . . . . . . . . . . . . 246 (Edebug), eval-current-buffer . . . . . . . . . . . . . . . . . 231 (Edebug), eval-defun . . . . . . . . . . . . . . . . . . . . . . . . . 231 (Edebug), eval-expression . . . . . . . . . . . . . . . . . . . . . 232 (Edebug), eval-region . . . . . . . . . . . . . . . . . . . . . . . . . 231 (Edebug), evaluation list. . . . . . . . . . . . . . . . . . . . . . 238 (Edebug), interactive commands . . . . . . . . . . . . . . 231 (Edebug), keyboard macros . . . . . . . . . . . . . . . . . . . 233 (Edebug), lambda-list . . . . . . . . . . . . . . . . . . . . . . . . 245 (Edebug), lexical binding . . . . . . . . . . . . . . . . . . . . . 237 (Edebug), printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 (Edebug), reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 (Edebug), save-excursion . . . . . . . . . . . . . . . . . . . . . 242 (Edebug), special forms. . . . . . . . . . . . . . . . . . . . . . . 231 (Edebug), syntax error . . . . . . . . . . . . . . . . . . . . . . . 246 (Edebug), window configuration . . . . . . . . . . . . . . 242 (for printing), stream . . . . . . . . . . . . . . . . . . . . . . . . . 256 (for reading), stream . . . . . . . . . . . . . . . . . . . . . . . . . 253 (for X windows), selection . . . . . . . . . . . . . . . . . . . . 715 (in a buffer), restriction. . . . . . . . . . . . . . . . . . . . . . . 498 (in a specifier), domain . . . . . . . . . . . . . . . . . . . . . . . 602 (in a specifier), fallback . . . . . . . . . . . . . . . . . . . . . . . 603 (in a specifier), inst-list . . . . . . . . . . . . . . . . . . . . . . . 602 (in a specifier), inst-pair . . . . . . . . . . . . . . . . . . . . . . 602 (in a specifier), instance . . . . . . . . . . . . . . . . . . . . . . 602 (in a specifier), instancing . . . . . . . . . . . . . . . . . . . . 602 (in a specifier), instantiator . . . . . . . . . . . . . . . . . . . 602 (in a specifier), locale . . . . . . . . . . . . . . . . . . . . . . . . . 602 (in a specifier), specification . . . . . . . . . . . . . . . . . . 602 (in a specifier), tag . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 (in a specifier), tag set . . . . . . . . . . . . . . . . . . . . . . . . 602 (in buffer), position. . . . . . . . . . . . . . . . . . . . . . . . . . . 489 (in file name), version number . . . . . . . . . . . . . . . . 408 (in obarray), bucket . . . . . . . . . . . . . . . . . . . . . . . . . . 115 (in regexp), character set . . . . . . . . . . . . . . . . . . . . . 554 (input), ISO Latin-1 characters . . . . . . . . . . . . . . . 710 (input), Latin-1 character set . . . . . . . . . . . . . . . . . 710 (list substitution), ‘. . . . . . . . . . . . . . . . . . . . . . . . . . . 181 (list substitution), backquote . . . . . . . . . . . . . . . . . 181 (mouse), cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 (mouse), pointer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 (of a buffer), accessible portion . . . . . . . . . . . . . . . 498 (of buffer), modification flag . . . . . . . . . . . . . . . . . . 438 (of file name), directory part. . . . . . . . . . . . . . . . . . 408 (of file name), nondirectory part . . . . . . . . . . . . . . 408 (of file), truename . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 (of list), element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 (symbol), property list cell. . . . . . . . . . . . . . . . . . . . 113 (with Backquote), , . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 (with Backquote), ,@ . . . . . . . . . . . . . . . . . . . . . . . . . 181 (with backquote), splicing . . . . . . . . . . . . . . . . . . . . 181

)

) in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

* in interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285

* in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 *? in regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 *scratch* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370

,

, (with Backquote) . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 ,@ (with Backquote). . . . . . . . . . . . . . . . . . . . . . . . . . 181

.

. in lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 . in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 .emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 .emacs customization . . . . . . . . . . . . . . . . . . . . . . . . . 365

/

/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 /= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

;

; in comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

?

? in character constant . . . . . . . . . . . . . . . . . . . . . . . . 23 ? in minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 ? in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554

@

@ in interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285

[

[ in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554

]

] in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554

‘

‘ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 ‘ (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 ‘ (list substitution) . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 ‘, edebug- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247

in interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285

|

| in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555

+

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 + in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 +? in regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554

=

= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

>

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 >= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

^

^in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555

\

\ in character constant . . . . . . . . . . . . . . . . . . . . . . . . 23 \ in display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 \ in printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 \ in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 \ in replacement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 \ in strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 \ in symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 \’ in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \‘ in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \{n,m\} in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 \= in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \> in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \< in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \b. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \b in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \B in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \n. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \n in print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 \n in replacement . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 \r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \s in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 \S in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 \t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \w in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 \W in regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 <= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

1

1 character set (input), Latin- . . . . . . . . . . . . . . . . 710 1 characters (input), ISO Latin- . . . . . . . . . . . . . . 710 1, ISO Latin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 1- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 1+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

4

4, C-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 4-map, ctl-x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783

5

5, C-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 5-map, ctl-x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783

A

abbrev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 abbrev table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 abbrev tables in modes . . . . . . . . . . . . . . . . . . . . . . . 364 abbrev, add- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 abbrev, define- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 abbrev, expand- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbrev, last-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbrev-alist, directory- . . . . . . . . . . . . . . . . . . . . . . . 410 abbrev-all-caps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbrev-expand-hook, pre-. . . . . . . . . . . . . . . . . . . . . 584 abbrev-expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbrev-file, quietly-read- . . . . . . . . . . . . . . . . . . . . . . 583 abbrev-file, write- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 abbrev-file-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 abbrev-location, last- . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbrev-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 abbrev-mode, default- . . . . . . . . . . . . . . . . . . . . . . . . 581 abbrev-prefix-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbrev-start-location . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbrev-start-location-buffer . . . . . . . . . . . . . . . . . . . 584 abbrev-symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 abbrev-table, c-mode- . . . . . . . . . . . . . . . . . . . . . . . . 585 abbrev-table, clear- . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 abbrev-table, define- . . . . . . . . . . . . . . . . . . . . . . . . . . 582 abbrev-table, fundamental-mode- . . . . . . . . . . . . . 585 abbrev-table, global-. . . . . . . . . . . . . . . . . . . . . . . . . . 585 abbrev-table, lisp-mode- . . . . . . . . . . . . . . . . . . . . . . 585 abbrev-table, local- . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 abbrev-table, make- . . . . . . . . . . . . . . . . . . . . . . . . . . 581 abbrev-table, text-mode- . . . . . . . . . . . . . . . . . . . . . 585 abbrev-table-description, insert- . . . . . . . . . . . . . . 582 abbrev-table-name-list . . . . . . . . . . . . . . . . . . . . . . . . 582 abbrev-text, last- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 abbreviate-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . 410 abbreviation, directory name . . . . . . . . . . . . . . . . . 410 abbrevs, only-global- . . . . . . . . . . . . . . . . . . . . . . . . . 582 abbrevs, save- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 abbrevs-changed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 abbrevs-map, edit- . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 abort-recursive-edit . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 aborting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 about-lock, ask-user- . . . . . . . . . . . . . . . . . . . . . . . . . 400 about-supersession-threat, ask-user-. . . . . . . . . . . 439 abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 absolute file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 absolute-p, file-name- . . . . . . . . . . . . . . . . . . . . . . . . . 411 accelerate-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 accelerator-enabled, menu-. . . . . . . . . . . . . . . . . . . . 348 accelerator-map, menu- . . . . . . . . . . . . . . . . . . . . . . . 349 accelerator-modifiers, menu- . . . . . . . . . . . . . . . . . . 349 accelerator-prefix, menu-. . . . . . . . . . . . . . . . . . . . . . 349 accelerators, keyboard menu . . . . . . . . . . . . . . . . . . 347 accelerators, menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 accept-process-output . . . . . . . . . . . . . . . . . . . . . . . . 688 access, environment variable . . . . . . . . . . . . . . . . . . 701 accessibility of a file . . . . . . . . . . . . . . . . . . . . . . . . . . 401 accessibility, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 accessible portion (of a buffer) . . . . . . . . . . . . . . . . 498 accessible-directory-p, file- . . . . . . . . . . . . . . . . . . . . 402 accessible-keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 according-to-mode, indent- . . . . . . . . . . . . . . . . . . . 536 acos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 acosh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 action, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 action, set-annotation-. . . . . . . . . . . . . . . . . . . . . . . . 645 activate-menubar-hook . . . . . . . . . . . . . . . . . . . . . . . 343 activate-popup-menu-hook . . . . . . . . . . . . . . . . . . . . 347 activate-region, zmacs- . . . . . . . . . . . . . . . . . . . . . . . 510 activate-region-hook, zmacs- . . . . . . . . . . . . . . . . . . 510 active display table . . . . . . . . . . . . . . . . . . . . . . . . . . . 662 active keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 active-minibuffer-window . . . . . . . . . . . . . . . . . . . . . 281 active-p, minibuffer-window- . . . . . . . . . . . . . . . . . . 282 active-p, region- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 add-abbrev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 add-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 add-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 add-menu-button . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 add-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 add-name-to-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 add-spec-list-to-specifier . . . . . . . . . . . . . . . . . . . . . . 606 add-spec-to-specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 606 add-submenu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 add-text-properties . . . . . . . . . . . . . . . . . . . . . . . . . . . 543 add-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 add-to-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 add-tooltalk-message-arg . . . . . . . . . . . . . . . . . . . . . 723 add-tooltalk-pattern-arg . . . . . . . . . . . . . . . . . . . . . . 725 add-tooltalk-pattern-attribute . . . . . . . . . . . . . . . . 725 address field of register . . . . . . . . . . . . . . . . . . . . . . . . 24 address, mail-host- . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 address, user-mail- . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 after, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 after, edebug-print-trace- . . . . . . . . . . . . . . . . 240, 250 after-change-function . . . . . . . . . . . . . . . . . . . . . . . . . 549 after-change-functions . . . . . . . . . . . . . . . . . . . . . . . . 549 after-find-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 after-init-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 after-insert-file-functions . . . . . . . . . . . . . . . . . . . . . . 546 after-load-alist. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 after-revert-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 after-save-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 age, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 alias, define-obsolete-function- . . . . . . . . . . . . . . . . 391 alias, define-obsolete-variable-. . . . . . . . . . . . . . . . . 392 alias, variable- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 aliases, for variables . . . . . . . . . . . . . . . . . . . . . . . . . . 163 aliases, variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 alist, after-load- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 alist, auto-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 alist, command-switch- . . . . . . . . . . . . . . . . . . . . . . . 696 alist, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 alist, destructive-plist-to- . . . . . . . . . . . . . . . . . . . . . 100 alist, directory-abbrev- . . . . . . . . . . . . . . . . . . . . . . . 410 alist, file-name-buffer-file-type-. . . . . . . . . . . . . . . . 421 alist, format- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 alist, interpreter-mode- . . . . . . . . . . . . . . . . . . . . . . . 371 alist, ldap-host-parameters- . . . . . . . . . . . . . . . . . . . 727 alist, minor-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 alist, minor-mode-map- . . . . . . . . . . . . . . . . . . . . . . . 325 alist, plist-to-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 alist, register- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 alist, sound- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 alist-to-plist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 alist-to-plist, destructive- . . . . . . . . . . . . . . . . . . . . . 100 alists, copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 all, text-property-not- . . . . . . . . . . . . . . . . . . . . . . . . 545 all-annotations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 all-caps, abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 all-completions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 all-completions, file-name- . . . . . . . . . . . . . . . . . . . . 413 all-defs, edebug- . . . . . . . . . . . . . . . . . . . . . . . . . 231, 249 all-forms, edebug-. . . . . . . . . . . . . . . . . . . . . . . . 231, 250 all-local-variables, kill- . . . . . . . . . . . . . . . . . . . . . . . . 161 allocate more storage, CL note— . . . . . . . . . . . . . 770 allocation, debug- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 allocation, memory . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 allocation-backtrace, debug- . . . . . . . . . . . . . . . . . . 774 allow-sendevents, x- . . . . . . . . . . . . . . . . . . . . . . . . . . 719 already-exists, file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 alternative, regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 analysis, performance . . . . . . . . . . . . . . . . . . . . . . . . . 240 and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 and, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 and, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 and-compile, eval- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 and-eval-command, edit-. . . . . . . . . . . . . . . . . . . . . . 266 and-exit, minibuffer-complete- . . . . . . . . . . . . . . . . 272 and-exit, self-insert- . . . . . . . . . . . . . . . . . . . . . . . . . . 280 and-indent, newline- . . . . . . . . . . . . . . . . . . . . . . . . . . 537 and-indent, reindent-then-newline- . . . . . . . . . . . . 537 and-init, make-specifier- . . . . . . . . . . . . . . . . . . . . . . 614 and-mark, exchange-point-. . . . . . . . . . . . . . . . . . . . 508 annotate-functions, write-region- . . . . . . . . . . . . . . 545 annotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 annotation hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 annotation, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 annotation, hide- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 annotation, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 annotation, reveal- . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 annotation-action. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-action, set-. . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-data, set- . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-down-glyph . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-down-glyph, set- . . . . . . . . . . . . . . . . . . 645 annotation-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-face, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-glyph, set- . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-layout, set-. . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 annotation-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-menu, set- . . . . . . . . . . . . . . . . . . . . . . . . 646 annotation-side . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 annotation-visible . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 annotation-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 annotationp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 annotations, all- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 annotations-at. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 annotations-in-region . . . . . . . . . . . . . . . . . . . . . . . . . 646 anonymous function . . . . . . . . . . . . . . . . . . . . . . . . . . 174 anonymous lambda expressions (Edebug) . . . . . 231 another buffer, changing to . . . . . . . . . . . . . . . . . . . 433 any, text-property- . . . . . . . . . . . . . . . . . . . . . . . . . . . 545 API, Drag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 API, Drop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 apostrophe for quoting . . . . . . . . . . . . . . . . . . . . . . . 129 apostrophe, quoting using . . . . . . . . . . . . . . . . . . . . 129 append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 append, kill-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 append-to-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 application-class, x-emacs- . . . . . . . . . . . . . . . . . . . . 717 apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 apply, and debugging . . . . . . . . . . . . . . . . . . . . . . . . . 227 apropos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 area, cursor-in-echo- . . . . . . . . . . . . . . . . . . . . . . . . . . 653 area, echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 area-message, inhibit-startup-echo-. . . . . . . . . . . . 694 area-p, event-over-text- . . . . . . . . . . . . . . . . . . . . . . . 298 area-pixel-edges, window-text- . . . . . . . . . . . . . . . . 468 area-pixel-height, window-text- . . . . . . . . . . . . . . . 467 area-pixel-width, window-text- . . . . . . . . . . . . . . . . 467 aref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 arg, add-tooltalk-message- . . . . . . . . . . . . . . . . . . . . 723 arg, add-tooltalk-pattern-. . . . . . . . . . . . . . . . . . . . . 725 arg, CL note—default optional. . . . . . . . . . . . . . . . 168 arg, current-prefix- . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 arg, prefix- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 arglist, compiled-function- . . . . . . . . . . . . . . . . . . . . 213 args, command-line- . . . . . . . . . . . . . . . . . . . . . . . . . . 697 argument binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 argument descriptors . . . . . . . . . . . . . . . . . . . . . . . . . 284 argument evaluation form . . . . . . . . . . . . . . . . . . . . 284 argument evaluation, macro . . . . . . . . . . . . . . . . . . 183 argument prompt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 argument string, default . . . . . . . . . . . . . . . . . . . . . . 286 argument unreading, prefix . . . . . . . . . . . . . . . . . . . 306 argument usage, numeric prefix . . . . . . . . . . . . . . . 287 argument usage, raw prefix . . . . . . . . . . . . . . . . . . . 287 argument, digit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 argument, evaluated expression . . . . . . . . . . . . . . . 288 argument, execute with prefix . . . . . . . . . . . . . . . . 289 argument, marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 argument, negative- . . . . . . . . . . . . . . . . . . . . . . . . . . 312 argument, numeric prefix . . . . . . . . . . . . . . . . . . . . . 310 argument, position . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 argument, prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 argument, raw prefix . . . . . . . . . . . . . . . . . . . . . . . . . 310 argument, region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 argument, universal- . . . . . . . . . . . . . . . . . . . . . . . . . . 312 argument, wrong-type-. . . . . . . . . . . . . . . . . . . . . . . . . 37 arguments, binding . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 arguments, command line. . . . . . . . . . . . . . . . . . . . . 696 arguments, complex . . . . . . . . . . . . . . . . . . . . . . . . . . 263 arguments, optional . . . . . . . . . . . . . . . . . . . . . . . . . . 168 arguments, program . . . . . . . . . . . . . . . . . . . . . . . . . . 675 arguments, reading . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 arguments, reading interactive . . . . . . . . . . . . . . . . 286 arguments, repositioning format . . . . . . . . . . . . . . . 71 arguments, rest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 arguments, wrong-number-of- . . . . . . . . . . . . . . . . . 167 arith-error example . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 arith-error in division . . . . . . . . . . . . . . . . . . . . . . . . . . 53 arithmetic shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 array elements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 arrayp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 arrays, character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 arrow, ctl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 arrow, default-ctl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 arrow, overlay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 657 arrow-glyph, control- . . . . . . . . . . . . . . . . . . . . . . . . . 642 arrow-position, overlay- . . . . . . . . . . . . . . . . . . . . . . . 658 arrow-string, overlay- . . . . . . . . . . . . . . . . . . . . . . . . . 657 ascent, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 ASCII character codes . . . . . . . . . . . . . . . . . . . . . . . . . 20 aset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 ash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 asin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 asinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 ask-user-about-lock . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 ask-user-about-supersession-threat . . . . . . . . . . . . 439 asking the user questions . . . . . . . . . . . . . . . . . . . . . 277 asking, trim-versions-without- . . . . . . . . . . . . . . . . 426 assoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 association list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 association lists, property lists vs . . . . . . . . . . . . . 118 assq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 asynchronous subprocess. . . . . . . . . . . . . . . . . . . . . . 679 at input, peeking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 at, annotations- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 at, extent-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 at, looking-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 at, posix-looking- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 at, text-properties- . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 atan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 atanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 atom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 80 atomic extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 atoms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 attribute, add-tooltalk-pattern- . . . . . . . . . . . . . . . 725 attribute, get-tooltalk-message- . . . . . . . . . . . . . . . 722 attribute, set-tooltalk-message- . . . . . . . . . . . . . . . 723 attributes of text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 attributes, charset-from- . . . . . . . . . . . . . . . . . . . . . . 739 attributes, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 attributes, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 Auto Fill mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 Auto Fill mode, newline and . . . . . . . . . . . . . . . . . . 517 auto-fill-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 auto-help, completion- . . . . . . . . . . . . . . . . . . . . . . . . 273 auto-lower-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 auto-mode, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 auto-mode-alist. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 auto-raise-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 auto-save, do- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 auto-save-default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 auto-save-file, rename- . . . . . . . . . . . . . . . . . . . . . . . . 430 auto-save-file-format . . . . . . . . . . . . . . . . . . . . . . . . . . 421 auto-save-file-if-necessary, delete- . . . . . . . . . . . . . 429 auto-save-file-name, buffer- . . . . . . . . . . . . . . . . . . . 427 auto-save-file-name, make- . . . . . . . . . . . . . . . . . . . . 428 auto-save-file-name-p . . . . . . . . . . . . . . . . . . . . . . . . . 428 auto-save-files, delete- . . . . . . . . . . . . . . . . . . . . . . . . 430 auto-save-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 auto-save-interval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 auto-save-list-file-name . . . . . . . . . . . . . . . . . . . . . . . 430 auto-save-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 auto-save-p, recent-. . . . . . . . . . . . . . . . . . . . . . . . . . . 429 auto-save-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 auto-save-visited-file-name . . . . . . . . . . . . . . . . . . . . 428 auto-saved, set-buffer- . . . . . . . . . . . . . . . . . . . . . . . . 429 auto-saving. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 autoload. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 733 autoload errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 autoload, function cell in . . . . . . . . . . . . . . . . . . . . . 201 autoloads, update-directory- . . . . . . . . . . . . . . . . . . 201 autoloads, update-file- . . . . . . . . . . . . . . . . . . . . . . . . 201 automatic, filling,. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 automatically buffer-local . . . . . . . . . . . . . . . . . . . . . 159 available fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 available, closures not . . . . . . . . . . . . . . . . . . . . . . . . 157 available, fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 average, load- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 avoidance, undo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546

B

back-to-indentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 539 backed-up, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 background pixmap . . . . . . . . . . . . . . . . . . . . . . . . . . 617 background, face-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 background, image-instance- . . . . . . . . . . . . . . . . . . 639 background, set-face- . . . . . . . . . . . . . . . . . . . . . . . . . 622 background-instance, face- . . . . . . . . . . . . . . . . . . . . 622 background-pixmap, face-. . . . . . . . . . . . . . . . . . . . . 622 background-pixmap, set-face- . . . . . . . . . . . . . . . . . 622 background-pixmap-instance, face- . . . . . . . . . . . . 623 backquote (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . 247 backquote (list substitution) . . . . . . . . . . . . . . . . . . 181 Backquote), , (with . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Backquote), ,@ (with . . . . . . . . . . . . . . . . . . . . . . . . . 181 backquote), splicing (with . . . . . . . . . . . . . . . . . . . . 181 backslash in character constant . . . . . . . . . . . . . . . . 23 backslash in strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 backslash in symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . 23 backspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 backtrace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 backtrace, debug-allocation- . . . . . . . . . . . . . . . . . . 774 backtrace-debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 backtrace-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 backtracking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 backtracking, preventing . . . . . . . . . . . . . . . . . . . . . . 245 backup file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 backup files, how to make them . . . . . . . . . . . . . . . 424 backup, file-newest-. . . . . . . . . . . . . . . . . . . . . . . . . . . 427 backup-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 backup-by-copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . 424 backup-by-copying-when-linked . . . . . . . . . . . . . . . 425 backup-by-copying-when-mismatch . . . . . . . . . . . 425 backup-enable-predicate . . . . . . . . . . . . . . . . . . . . . . 424 backup-file-name, find- . . . . . . . . . . . . . . . . . . . . . . . 427 backup-file-name, make- . . . . . . . . . . . . . . . . . . . . . . 426 backup-file-name-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 backup-files, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 backup-inhibited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 424 backward, posix-search-. . . . . . . . . . . . . . . . . . . . . . . 562 backward, re-search- . . . . . . . . . . . . . . . . . . . . . . . . . . 560 backward, search- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 backward, skip-chars- . . . . . . . . . . . . . . . . . . . . . . . . . 497 backward, skip-syntax- . . . . . . . . . . . . . . . . . . . . . . . 577 backward, word-search- . . . . . . . . . . . . . . . . . . . . . . . 552 backward-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 backward-char, delete- . . . . . . . . . . . . . . . . . . . . . . . . 519 backward-delete-char-untabify . . . . . . . . . . . . . . . . 519 backward-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 backward-prefix-chars . . . . . . . . . . . . . . . . . . . . . . . . 577 backward-sexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 backward-to-indentation . . . . . . . . . . . . . . . . . . . . . . 540 backward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 balancing parentheses . . . . . . . . . . . . . . . . . . . . . . . . 660 barf-if-buffer-read-only. . . . . . . . . . . . . . . . . . . . . . . . 440 base buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 base, ldap-default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 base-buffer, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 baseline, glyph-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 baseline, set-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 baseline-instance, glyph- . . . . . . . . . . . . . . . . . . . . . . 632 batch mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714 batch-byte-compile . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 batch-byte-recompile-directory . . . . . . . . . . . . . . . . 209 baud-rate, device- . . . . . . . . . . . . . . . . . . . . . . . 488, 711 baud-rate, set-device- . . . . . . . . . . . . . . . . . . . . 488, 712 beep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 beeping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 before point, insertion . . . . . . . . . . . . . . . . . . . . . . . . 516 before, edebug-print-trace- . . . . . . . . . . . . . . . 240, 250 before-change-function. . . . . . . . . . . . . . . . . . . . . . . . 549 before-change-functions . . . . . . . . . . . . . . . . . . . . . . . 548 before-init-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 before-markers, insert- . . . . . . . . . . . . . . . . . . . . . . . . 516 before-revert-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 begin-glyph, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . 596 begin-glyph, set-extent-. . . . . . . . . . . . . . . . . . . . . . . 597 begin-glyph-layout, extent- . . . . . . . . . . . . . . . . . . . 596 begin-glyph-layout, set-extent- . . . . . . . . . . . . . . . . 597 beginning of line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 beginning of line in regexp . . . . . . . . . . . . . . . . . . . . 555 beginning, match- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 beginning, region- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 beginning-of-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 beginning-of-defun. . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 beginning-of-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 bell character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 bell, visible- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 bell-volume. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 664 big5-char, decode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 big5-char, encode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 binary files and text files. . . . . . . . . . . . . . . . . . . . . . 421 binary files, text files and . . . . . . . . . . . . . . . . . . . . . 421 binary, find-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 binary-process-input . . . . . . . . . . . . . . . . . . . . . . . . . . 679 binary-process-output . . . . . . . . . . . . . . . . . . . . . . . . 679 bind-text-domain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 731 binding (Edebug), lexical . . . . . . . . . . . . . . . . . . . . . 237 binding arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 binding local variables . . . . . . . . . . . . . . . . . . . . . . . . 148 binding of a key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 binding, argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 binding, current . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 binding, deep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 binding, global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 binding, global-key- . . . . . . . . . . . . . . . . . . . . . . . . . . 328 binding, key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 binding, key- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 binding, keymap-default- . . . . . . . . . . . . . . . . . . . . . 319 binding, local. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 binding, local-key- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 binding, minor-mode-key-. . . . . . . . . . . . . . . . . . . . . 328 binding, set-keymap-default- . . . . . . . . . . . . . . . . . . 319 binding, shallow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 bindings, changing key. . . . . . . . . . . . . . . . . . . . . . . . 329 bindings, describe-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 bindings, describe-prefix- . . . . . . . . . . . . . . . . . . . . . 391 bindings, Helper-describe- . . . . . . . . . . . . . . . . . . . . 391 bindings, inheriting a keymap’s . . . . . . . . . . . . . . . 319 bindings, replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 bindings-internal, describe- . . . . . . . . . . . . . . . . . . . 336 bit vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 bit vector length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 bit vectors, copying. . . . . . . . . . . . . . . . . . . . . . . . . . . 111 bit-vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 bit-vector, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 bit-vector-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 bitmap-file-path, x- . . . . . . . . . . . . . . . . . . . . . . 636, 719 bitp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 bits, event-modifier- . . . . . . . . . . . . . . . . . . . . . . . . . . 300 bitwise and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 bitwise exclusive or . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 bitwise not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 bitwise or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 blank-lines, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 blanks-input, read-no- . . . . . . . . . . . . . . . . . . . . . . . . 265 blink-matching-open . . . . . . . . . . . . . . . . . . . . . . . . . . 660 blink-matching-paren . . . . . . . . . . . . . . . . . . . . . . . . . 660 blink-matching-paren-delay . . . . . . . . . . . . . . . . . . . 660 blink-matching-paren-distance . . . . . . . . . . . . . . . . 660 blink-paren-function . . . . . . . . . . . . . . . . . . . . . . . . . . 660 blink-paren-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 blinking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 bobp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 body of function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 bold. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 bold, x-make-font- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 bold-italic, x-make-font- . . . . . . . . . . . . . . . . . . . . . . 625 bolp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 bookmark-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 boolean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 boolean-specifier-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 boolean-specifier-p, face-. . . . . . . . . . . . . . . . . . . . . . 605 bootstrapping XEmacs from temacs . . . . . . . . . . 767 border-p, event-over- . . . . . . . . . . . . . . . . . . . . . . . . . 299 bottom-toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 bottom-toolbar-height . . . . . . . . . . . . . . . . . . . . . . . . 356 bottom-toolbar-visible-p . . . . . . . . . . . . . . . . . . . . . . 357 boundary, undo- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 boundp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 boundp, default-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 box diagrams, for lists . . . . . . . . . . . . . . . . . . . . . . . . . 25 box representation for lists. . . . . . . . . . . . . . . . . . . . . 79 box, cons cell as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 box, dialog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 box, popup-dialog- . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 box, y-or-n-p-maybe-dialog- . . . . . . . . . . . . . . . . . . 279 box, yes-or-no-p-dialog- . . . . . . . . . . . . . . . . . . . . . . . 279 box, yes-or-no-p-maybe-dialog-. . . . . . . . . . . . . . . . 279 boxed, for lists, diagrams,. . . . . . . . . . . . . . . . . . . . . . 25 boxes, lists represented as. . . . . . . . . . . . . . . . . . . . . . 79 break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 break condition, global . . . . . . . . . . . . . . . . . . . . . . . 235 break-condition, edebug-global- . . . . . . . . . . 235, 251 break-condition, edebug-set-global- . . . . . . . . . . . 235 breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 breakpoints, embedded . . . . . . . . . . . . . . . . . . . . . . . 235 bucket (in obarray) . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 buffer (Edebug), eval-current-. . . . . . . . . . . . . . . . . 231 buffer contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 buffer contents, evaluation of . . . . . . . . . . . . . . . . . 122 buffer excursion, current . . . . . . . . . . . . . . . . . . . . . . 497 buffer file name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 buffer input stream . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 buffer list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 buffer mark, current . . . . . . . . . . . . . . . . . . . . . . . . . . 507 buffer marker, end of . . . . . . . . . . . . . . . . . . . . . . . . . 504 buffer modification . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 buffer names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 buffer output stream . . . . . . . . . . . . . . . . . . . . . . . . . 256 buffer point and mark (Edebug), current . . . . . . 242 buffer position, current . . . . . . . . . . . . . . . . . . . . . . . 489 buffer text notation. . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 buffer text, comparing . . . . . . . . . . . . . . . . . . . . . . . . 515 buffer), accessible portion (of a . . . . . . . . . . . . . . . 498 buffer), modification flag (of . . . . . . . . . . . . . . . . . . 438 buffer), position (in. . . . . . . . . . . . . . . . . . . . . . . . . . . 489 buffer), restriction (in a. . . . . . . . . . . . . . . . . . . . . . . 498 buffer, abbrev-start-location- . . . . . . . . . . . . . . . . . 584 buffer, backup- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 buffer, base. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 buffer, beginning-of- . . . . . . . . . . . . . . . . . . . . . . . . . . 492 buffer, buffer-base- . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 buffer, bury-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 buffer, changing to another . . . . . . . . . . . . . . . . . . . 433 buffer, create-file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 buffer, current . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 buffer, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 buffer, cut . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 buffer, display- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 457 buffer, displaying a . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 buffer, end-of- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 buffer, erase- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 buffer, eval- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 buffer, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 buffer, file name of . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 buffer, generate-new- . . . . . . . . . . . . . . . . . . . . . . . . . 443 buffer, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 buffer, get-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 buffer, insert-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 buffer, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 buffer, lock- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 buffer, make-indirect- . . . . . . . . . . . . . . . . . . . . . . . . . 445 buffer, marker- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 buffer, mouse-grabbed- . . . . . . . . . . . . . . . . . . . . . . . 325 buffer, obsolete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 buffer, other- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 buffer, other-window-scroll- . . . . . . . . . . . . . . . . . . . 463 buffer, pop-to-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 456 buffer, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 buffer, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 buffer, read-only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 buffer, rename- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 buffer, revert- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 430 buffer, save- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 buffer, save-current- . . . . . . . . . . . . . . . . . . . . . . . . . . 498 buffer, selecting a. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 buffer, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 buffer, set-process- . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 buffer, set-window- . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 buffer, switch-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 456 buffer, switching to a . . . . . . . . . . . . . . . . . . . . . . . . . 455 buffer, unlock-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 buffer, window-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 buffer, with-current- . . . . . . . . . . . . . . . . . . . . . . . . . . 498 buffer, with-output-to-temp- . . . . . . . . . . . . . . . . . . 658 buffer-auto-save-file-name. . . . . . . . . . . . . . . . . . . . . 427 buffer-auto-saved, set- . . . . . . . . . . . . . . . . . . . . . . . . 429 buffer-backed-up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 buffer-base-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 buffer-case-table, describe- . . . . . . . . . . . . . . . . . . . . . 75 buffer-create, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 buffer-dedicated, set-window- . . . . . . . . . . . . . . . . . 457 buffer-disable-undo . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 buffer-enable-undo. . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 buffer-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 buffer-file-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 buffer-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . 436, 437 buffer-file-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 buffer-file-truename. . . . . . . . . . . . . . . . . . . . . . . . . . . 437 buffer-file-type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 buffer-file-type, default-. . . . . . . . . . . . . . . . . . . . . . . 422 buffer-file-type, find-. . . . . . . . . . . . . . . . . . . . . . . . . . 421 buffer-file-type-alist, file-name-. . . . . . . . . . . . . . . . 421 buffer-flush-undo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 buffer-function, buffers-menu-switch-to- . . . . . . . 350 buffer-function, display- . . . . . . . . . . . . . . . . . . . . . . 459 buffer-function, revert-. . . . . . . . . . . . . . . . . . . . . . . . 431 buffer-glyph-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 buffer-hook, kill-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 buffer-identification, modeline- . . . . . . . . . . . . . . . . 377 buffer-in-windows, replace-. . . . . . . . . . . . . . . . . . . . 457 buffer-insert-file-contents-function, revert- . . . . . 431 buffer-invisibility-spec . . . . . . . . . . . . . . . . . . . . . . . . 655 buffer-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 buffer-live-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 buffer-local variables. . . . . . . . . . . . . . . . . . . . . . . . . . 159 buffer-local variables in modes . . . . . . . . . . . . . . . . 364 buffer-local, automatically . . . . . . . . . . . . . . . . . . . . 159 buffer-local, make-variable- . . . . . . . . . . . . . . . . . . . 160 buffer-local, variables, . . . . . . . . . . . . . . . . . . . . . . . . 159 buffer-local-variables. . . . . . . . . . . . . . . . . . . . . . . . . . 160 buffer-major-mode, set- . . . . . . . . . . . . . . . . . . . . . . . 370 buffer-menu, popup- . . . . . . . . . . . . . . . . . . . . . . . . . . 347 Buffer-menu-mode-map . . . . . . . . . . . . . . . . . . . . . . . 783 buffer-menu-mode-map, electric- . . . . . . . . . . . . . . 783 buffer-menubar, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 343 buffer-modified-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 buffer-modified-p, set- . . . . . . . . . . . . . . . . . . . . . . . . 438 buffer-modified-tick. . . . . . . . . . . . . . . . . . . . . . . . . . . 439 buffer-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 buffer-name, generate-new- . . . . . . . . . . . . . . . . . . . 436 buffer-names, same-window- . . . . . . . . . . . . . . . . . . 459 buffer-names, special-display- . . . . . . . . . . . . . . . . . 458 buffer-offer-save. . . . . . . . . . . . . . . . . . . . . . . . . . 396, 444 buffer-other-window, switch-to- . . . . . . . . . . . . . . . 456 buffer-points, edebug-save-displayed- . . . . . 242, 250 buffer-process, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 buffer-query-functions, kill- . . . . . . . . . . . . . . . . . . . 444 buffer-read-only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 buffer-read-only, barf-if- . . . . . . . . . . . . . . . . . . . . . . 440 buffer-saved-size . . . . . . . . . . . . . . . . . . . . . . . . . 430, 490 buffer-show-function, temp- . . . . . . . . . . . . . . . . . . . 659 buffer-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 buffer-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 buffer-substring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 buffer-substring, insert- . . . . . . . . . . . . . . . . . . . . . . . 516 buffer-substrings, compare- . . . . . . . . . . . . . . . . . . . 515 buffer-undo-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 buffer-window, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 bufferp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 buffers in interactive, read-only . . . . . . . . . . . . . . . 285 buffers menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349 buffers, controlled in windows. . . . . . . . . . . . . . . . . 455 buffers, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 buffers, indirect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 buffers, killing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 buffers, list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 buffers, save-some-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 buffers-directory, list- . . . . . . . . . . . . . . . . . . . . . . . . . 438 buffers-menu-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 buffers-menu-line, format- . . . . . . . . . . . . . . . . . . . . 350 buffers-menu-max-size . . . . . . . . . . . . . . . . . . . . . . . . 349 buffers-menu-p, complex- . . . . . . . . . . . . . . . . . . . . . 350 buffers-menu-switch-to-buffer-function . . . . . . . . 350 build-time, emacs- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 building lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 building XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 767 built-in function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 bury-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 busy-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 button type, toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 button, add-menu- . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 button, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 button, event-toolbar- . . . . . . . . . . . . . . . . . . . . . . . . 299 button-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 button-list, toolbar-make- . . . . . . . . . . . . . . . . . . . . 354 button-press-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . 296 button-release-event-p . . . . . . . . . . . . . . . . . . . . . . . . 296 button-syntax, check-toolbar- . . . . . . . . . . . . . . . . . 355 buttons-captioned-p, toolbar- . . . . . . . . . . . . . . . . . 357 bvconcat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 byte-code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207, 210 byte-code function. . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 byte-code interpreter . . . . . . . . . . . . . . . . . . . . . . . . . 210 byte-code, disassembled. . . . . . . . . . . . . . . . . . . . . . . 214 byte-code, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 byte-compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 byte-compile, batch- . . . . . . . . . . . . . . . . . . . . . . . . . . 209 byte-compile-dynamic . . . . . . . . . . . . . . . . . . . . . . . . 211 byte-compile-dynamic-docstrings. . . . . . . . . . . . . . 210 byte-compile-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 byte-compiling macros . . . . . . . . . . . . . . . . . . . . . . . . 180 byte-compiling require . . . . . . . . . . . . . . . . . . . . . . . . 203 byte-recompile-directory . . . . . . . . . . . . . . . . . . . . . . 209 byte-recompile-directory, batch-. . . . . . . . . . . . . . . 209 byte-recompile-directory-ignore-errors-p . . . . . . . 210 bytecode, fetch- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 bytes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 bytes-used, pure- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769

C

c, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-c. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 C-h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-M-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 c-mode-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . . . . 585 c-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 c-mode-syntax-table . . . . . . . . . . . . . . . . . . . . . . . . . . 580 C-q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 C-s. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 C-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-x 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-x 5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-x a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-x n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 C-x r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 c++-mode-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 caaaar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caaadr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caadar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caadr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cadaar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cadadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cadar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caddar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cadddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 caddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 call debugging, function . . . . . . . . . . . . . . . . . . . . . . 220 call evaluation, macro . . . . . . . . . . . . . . . . . . . . . . . . 126 call stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 call, debug-on-next- . . . . . . . . . . . . . . . . . . . . . . . . . . 227 call, function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 call, interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 call, macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 call-interactively. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 call-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 call-process-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . 678 calling a function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 cancel-debug-on-entry . . . . . . . . . . . . . . . . . . . . . . . . 221 canonicalize-inst-list . . . . . . . . . . . . . . . . . . . . . . . . . . 608 canonicalize-inst-pair . . . . . . . . . . . . . . . . . . . . . . . . . 608 canonicalize-lax-plist . . . . . . . . . . . . . . . . . . . . . . . . . 100 canonicalize-plist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 canonicalize-spec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 608 canonicalize-spec-list . . . . . . . . . . . . . . . . . . . . . . . . . 608 canonicalize-tag-set . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 capitalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 capitalize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 capitalize-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 capitalize-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 caps, abbrev-all- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 captioned-p, toolbar-buttons- . . . . . . . . . . . . . . . . . 357 car . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 car-safe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 case changes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 case in replacements . . . . . . . . . . . . . . . . . . . . . . . . . . 565 case key sequence, upper . . . . . . . . . . . . . . . . . . . . . 304 case of letters, CL note— . . . . . . . . . . . . . . . . . . . . . . 23 case, character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 case, completion-ignore- . . . . . . . . . . . . . . . . . . . . . . 270 case, condition-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 case, lower. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 case, searching and . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 case, upper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 case-fold-search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 case-fold-search, default- . . . . . . . . . . . . . . . . . . . . . . 568 case-replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 case-syntax, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 case-syntax-delims, set- . . . . . . . . . . . . . . . . . . . . . . . . 75 case-syntax-pair, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 75 case-table, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 case-table, describe-buffer- . . . . . . . . . . . . . . . . . . . . . 75 case-table, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 case-table, set-standard- . . . . . . . . . . . . . . . . . . . . . . . 75 case-table, standard-. . . . . . . . . . . . . . . . . . . . . . . . . . . 75 case-table-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 catch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 catch, no- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 category-designator-p . . . . . . . . . . . . . . . . . . . . . . . . . 757 category-list, coding- . . . . . . . . . . . . . . . . . . . . . . . . . 750 category-system, coding-. . . . . . . . . . . . . . . . . . . . . . 750 category-system, set-coding- . . . . . . . . . . . . . . . . . . 750 category-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 category-table, copy- . . . . . . . . . . . . . . . . . . . . . . . . . 757 category-table, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 category-table, standard- . . . . . . . . . . . . . . . . . . . . . 757 category-table-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 category-table-value-p . . . . . . . . . . . . . . . . . . . . . . . . 757 cbreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 ccl-elapsed-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 756 ccl-execute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 755 ccl-execute-on-string . . . . . . . . . . . . . . . . . . . . . . . . . . 756 ccl-program, charset- . . . . . . . . . . . . . . . . . . . . . . . . . 740 ccl-program, register- . . . . . . . . . . . . . . . . . . . . . . . . . 756 ccl-program, set-charset-. . . . . . . . . . . . . . . . . . . . . . 740 ccl-reset-elapsed-time . . . . . . . . . . . . . . . . . . . . . . . . . 756 cdaaar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdaadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdaar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdadar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdaddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdadr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cddaar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cddadr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cddar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdddar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cddddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cdddr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 cddr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 CDE dt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 cdr-safe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 cell (symbol), property list. . . . . . . . . . . . . . . . . . . . 113 cell as box, cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 cell in autoload, function . . . . . . . . . . . . . . . . . . . . . 201 cell, function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 cell, print name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 cell, value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 cell, void function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 cells, cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 cells, lists and cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 centering point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 change hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 change, next-property-. . . . . . . . . . . . . . . . . . . . . . . . 544 change, next-single-property- . . . . . . . . . . . . . . . . . 544 change, previous-property- . . . . . . . . . . . . . . . . . . . . 545 change, previous-single-property- . . . . . . . . . . . . . 545 change-function, after- . . . . . . . . . . . . . . . . . . . . . . . . 549 change-function, before- . . . . . . . . . . . . . . . . . . . . . . 549 change-functions, after- . . . . . . . . . . . . . . . . . . . . . . . 549 change-functions, before- . . . . . . . . . . . . . . . . . . . . . 548 change-functions, window-size- . . . . . . . . . . . . . . . . 470 change-hook, first- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 549 change-major-mode-hook . . . . . . . . . . . . . . . . . . . . . 365 changed, abbrevs- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 changes, case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 changes, hooks for text . . . . . . . . . . . . . . . . . . . . . . . 548 changing key bindings . . . . . . . . . . . . . . . . . . . . . . . . 329 changing to another buffer . . . . . . . . . . . . . . . . . . . . 433 changing window size . . . . . . . . . . . . . . . . . . . . . . . . . 468 changing, window size,. . . . . . . . . . . . . . . . . . . . . . . . 468 char quitting, read-quoted- . . . . . . . . . . . . . . . . . . . 309 char table type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 char, backward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 char, decode-big5- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 char, decode-shift-jis- . . . . . . . . . . . . . . . . . . . . . . . . . 751 char, delete previous. . . . . . . . . . . . . . . . . . . . . . . . . . 519 char, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 char, delete-backward- . . . . . . . . . . . . . . . . . . . . . . . . 519 char, encode-big5- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 char, encode-shift-jis- . . . . . . . . . . . . . . . . . . . . . . . . . 751 char, following- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 char, forward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 char, goto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 char, help- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 char, insert- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 char, int- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 char, last-command- . . . . . . . . . . . . . . . . . . . . . . . . . . 292 char, last-input- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 char, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 char, make-composite- . . . . . . . . . . . . . . . . . . . . . . . . 742 char, meta-prefix- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 char, preceding- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 char, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 char, read-quoted- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 char, string-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 char, write-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 char-after . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 char-charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 char-description, text- . . . . . . . . . . . . . . . . . . . . . . . . 388 char-equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 char-in-region, subst- . . . . . . . . . . . . . . . . . . . . . . . . . 546 char-int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 char-int confoundance disease . . . . . . . . . . . . . . . . . . 20 char-int-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 char-int-p, char-or- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 char-octet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 char-or-char-int-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 char-or-marker-p, integer-. . . . . . . . . . . . . . . . . . . . . 503 char-or-marker-p, number- . . . . . . . . . . . . . . . . . . . . 503 char-or-string-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 char-p, integer-or- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 char-property, get-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 char-string, composite- . . . . . . . . . . . . . . . . . . . . . . . 742 char-syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 576 char-table, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 char-table, get-range- . . . . . . . . . . . . . . . . . . . . . . . . . . 77 char-table, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 char-table, map- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 char-table, put- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 char-table, reset- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 char-table-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 char-table-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 char-table-type-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 char-table-type-p, valid- . . . . . . . . . . . . . . . . . . . . . . . 76 char-table-value, check-valid-. . . . . . . . . . . . . . . . . . . 77 char-table-value-p, valid-. . . . . . . . . . . . . . . . . . . . . . . 77 char-to-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 char-untabify, backward-delete- . . . . . . . . . . . . . . . 519 char= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 character arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 character case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 character code, octal. . . . . . . . . . . . . . . . . . . . . . . . . . . 23 character codes, ASCII . . . . . . . . . . . . . . . . . . . . . . . . 20 character constant, ? in. . . . . . . . . . . . . . . . . . . . . . . . 23 character constant, \ in . . . . . . . . . . . . . . . . . . . . . . . . 23 character constant, backslash in. . . . . . . . . . . . . . . . 23 character constant, question mark in . . . . . . . . . . . 23 character descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . 662 character input, octal. . . . . . . . . . . . . . . . . . . . . . . . . 306 character input, quoted . . . . . . . . . . . . . . . . . . . . . . . 306 character insertion. . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 character printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 character printing, control . . . . . . . . . . . . . . . . . . . . 388 character printing, meta . . . . . . . . . . . . . . . . . . . . . . 388 character quote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 character set (in regexp) . . . . . . . . . . . . . . . . . . . . . . 554 character set (input), Latin-1 . . . . . . . . . . . . . . . . . 710 character to string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 character, bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 character, close parenthesis . . . . . . . . . . . . . . . . . . . 573 character, event-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 character, open parenthesis . . . . . . . . . . . . . . . . . . . 573 character, punctuation. . . . . . . . . . . . . . . . . . . . . . . . 573 character, quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 character, string to . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 character, whitespace . . . . . . . . . . . . . . . . . . . . . . . . . 572 character-to-event . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 characteristics of font instances . . . . . . . . . . . . . . . 624 characteristics, font instance . . . . . . . . . . . . . . . . . . 624 characterp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 characters (input), ISO Latin-1 . . . . . . . . . . . . . . . 710 characters for interactive codes . . . . . . . . . . . . . . . 285 characters in display, control. . . . . . . . . . . . . . . . . . 661 characters in printing, escape . . . . . . . . . . . . . . . . . 258 characters in printing, quoting . . . . . . . . . . . . . . . . 258 characters, control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 characters, escape . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 characters, flow control . . . . . . . . . . . . . . . . . . . . . . . 712 characters, printed representation for . . . . . . . . . . 21 characters, read syntax for . . . . . . . . . . . . . . . . . . . . . 21 characters, reading, control . . . . . . . . . . . . . . . . . . . 306 characters, reading, nonprinting. . . . . . . . . . . . . . . 306 characters, replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 characters, skipping . . . . . . . . . . . . . . . . . . . . . . . . . . 496 characters, syntax for . . . . . . . . . . . . . . . . . . . . . . . . . . 21 chars, backward-prefix- . . . . . . . . . . . . . . . . . . . . . . . 577 chars, charset- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 chars-backward, skip- . . . . . . . . . . . . . . . . . . . . . . . . . 497 chars-forward, skip-. . . . . . . . . . . . . . . . . . . . . . . . . . . 496 charset type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 charset, char-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 charset, charset-reverse-direction- . . . . . . . . . . . . . 739 charset, find- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 charset, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 charset, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 charset, make-reverse-direction- . . . . . . . . . . . . . . . 739 charset-ccl-program. . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-ccl-program, set-. . . . . . . . . . . . . . . . . . . . . . 740 charset-chars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-dimension . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-direction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-doc-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-final . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-from-attributes . . . . . . . . . . . . . . . . . . . . . . . 739 charset-graphic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 charset-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-region, find- . . . . . . . . . . . . . . . . . . . . . . . . . . 742 charset-registry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 charset-reverse-direction-charset . . . . . . . . . . . . . . 739 charset-string, find- . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 charsetp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 737 check-toolbar-button-syntax . . . . . . . . . . . . . . . . . . 355 check-valid-char-table-value . . . . . . . . . . . . . . . . . . . . 77 check-valid-inst-list . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 check-valid-instantiator . . . . . . . . . . . . . . . . . . . . . . . 614 check-valid-plist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 check-valid-spec-list . . . . . . . . . . . . . . . . . . . . . . . . . . 615 checking, type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 child process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 children, extent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 children, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 children, map-extent- . . . . . . . . . . . . . . . . . . . . . . . . . 593 children, of extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 circle, edebug-print- . . . . . . . . . . . . . . . . . . . . . 239, 251 circular structures, printing . . . . . . . . . . . . . . . . . . . 239 CL note—allocate more storage. . . . . . . . . . . . . . . 770 CL note—case of letters . . . . . . . . . . . . . . . . . . . . . . . 23 CL note—default optional arg . . . . . . . . . . . . . . . . 168 CL note—integers vrs eq. . . . . . . . . . . . . . . . . . . . . . . 49 CL note—lack union, set. . . . . . . . . . . . . . . . . . . . . . . 92 CL note—no continuable errors . . . . . . . . . . . . . . . 140 CL note—only throw in Emacs . . . . . . . . . . . . . . . 137 CL note—rplaca vrs setcar. . . . . . . . . . . . . . . . . . . . . 87 CL note—set local. . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 CL note—special forms compared. . . . . . . . . . . . . 128 CL note—special variables . . . . . . . . . . . . . . . . . . . . 156 CL note—symbol in obarrays . . . . . . . . . . . . . . . . . 116 cl-read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 cl-specs.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 cl.el (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 class property, mode- . . . . . . . . . . . . . . . . . . . . . . . . . 365 class, device- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 class, x-display-visual- . . . . . . . . . . . . . . . . . . . . . . . . 718 class, x-emacs-application- . . . . . . . . . . . . . . . . . . . . 717 class-p, valid-device- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 classes, display-warning-suppressed-. . . . . . . . . . . 654 classes, log-warning-suppressed-. . . . . . . . . . . . . . . 654 classes, syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 cleanup forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 cleanup, error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 clear-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 clear-hashing, locate-file- . . . . . . . . . . . . . . . . . . . . . 199 clear-message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 652 clear-range-table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 clear-visited-file-modtime . . . . . . . . . . . . . . . . . . . . . 439 close parenthesis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 close parenthesis character . . . . . . . . . . . . . . . . . . . . 573 close, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 close, tq- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 close-database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 closest-point, event- . . . . . . . . . . . . . . . . . . . . . . . . . . 298 closures not available . . . . . . . . . . . . . . . . . . . . . . . . . 157 clrhash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 code description, interactive . . . . . . . . . . . . . . . . . . 285 code function, byte- . . . . . . . . . . . . . . . . . . . . . . . . . . 212 code interpreter, byte- . . . . . . . . . . . . . . . . . . . . . . . . 210 code, byte- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207, 210 code, disassembled byte- . . . . . . . . . . . . . . . . . . . . . . 214 code, make-byte- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 code, octal character. . . . . . . . . . . . . . . . . . . . . . . . . . . 23 code-rigidly, indent- . . . . . . . . . . . . . . . . . . . . . . . . . . 538 codes, ASCII character . . . . . . . . . . . . . . . . . . . . . . . . 20 codes, characters for interactive . . . . . . . . . . . . . . . 285 codes, description for interactive . . . . . . . . . . . . . . 285 codes, interactive, description of . . . . . . . . . . . . . . 285 coding standards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 759 coding style, standards of . . . . . . . . . . . . . . . . . . . . . 759 coding system type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 coding-category-list. . . . . . . . . . . . . . . . . . . . . . . . . . . 750 coding-category-system . . . . . . . . . . . . . . . . . . . . . . . 750 coding-category-system, set- . . . . . . . . . . . . . . . . . . 750 coding-priority-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . 750 coding-priority-list, set- . . . . . . . . . . . . . . . . . . . . . . . 750 coding-region, decode- . . . . . . . . . . . . . . . . . . . . . . . . 750 coding-region, detect-. . . . . . . . . . . . . . . . . . . . . . . . . 750 coding-region, encode- . . . . . . . . . . . . . . . . . . . . . . . . 750 coding-system, copy- . . . . . . . . . . . . . . . . . . . . . . . . . 749 coding-system, find- . . . . . . . . . . . . . . . . . . . . . . . . . . 749 coding-system, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 coding-system, make- . . . . . . . . . . . . . . . . . . . . . . . . . 749 coding-system, subsidiary- . . . . . . . . . . . . . . . . . . . . 749 coding-system-doc-string . . . . . . . . . . . . . . . . . . . . . 749 coding-system-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 coding-system-name . . . . . . . . . . . . . . . . . . . . . . . . . . 749 coding-system-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 746 coding-system-property . . . . . . . . . . . . . . . . . . . . . . . 749 coding-system-type . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 collect, garbage- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 771 collection, marker garbage . . . . . . . . . . . . . . . . . . . . 501 collector, garbage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 color instance type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 color instances. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 color-instance-name . . . . . . . . . . . . . . . . . . . . . . . . . . 626 color-instance-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626 color-instance-rgb-components . . . . . . . . . . . . . . . . 626 color-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626 color-pixmap-image-instance-p . . . . . . . . . . . . . . . . 638 color-rgb-components . . . . . . . . . . . . . . . . . . . . . . . . . 626 color-specifier-p . . . . . . . . . . . . . . . . . . . . . . . . . . 605, 625 color-symbols, xpm- . . . . . . . . . . . . . . . . . . . . . . . . . . 636 colorize-image-instance . . . . . . . . . . . . . . . . . . . . . . . 638 colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 column, current-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 column, current-fill- . . . . . . . . . . . . . . . . . . . . . . . . . . 530 column, default-fill-. . . . . . . . . . . . . . . . . . . . . . . . . . . 530 column, fill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 column, move-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 columns, charset- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 columns, counting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 columns, display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 columns, sort- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 command descriptions . . . . . . . . . . . . . . . . . . . . . . . . . 12 command history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 command history, record. . . . . . . . . . . . . . . . . . . . . . 289 command in keymap . . . . . . . . . . . . . . . . . . . . . . . . . 326 command key input, waiting for . . . . . . . . . . . . . . 307 command line arguments . . . . . . . . . . . . . . . . . . . . . 696 command line options . . . . . . . . . . . . . . . . . . . . . . . . 696 command line, options on. . . . . . . . . . . . . . . . . . . . . 696 command line, switches on. . . . . . . . . . . . . . . . . . . . 696 command list, debugger . . . . . . . . . . . . . . . . . . . . . . 223 command loop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 command loop, editor . . . . . . . . . . . . . . . . . . . . . . . . 283 command loop, recursive . . . . . . . . . . . . . . . . . . . . . 312 command name, read . . . . . . . . . . . . . . . . . . . . . . . . . 289 command override, self-insert- . . . . . . . . . . . . . . . . 332 command repetition, kill. . . . . . . . . . . . . . . . . . . . . . 291 command, complex . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 command, current . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 command, define-prefix- . . . . . . . . . . . . . . . . . . . . . . 322 command, disable-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 command, disabled . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 command, edit-and-eval-. . . . . . . . . . . . . . . . . . . . . . 266 command, enable- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 command, execute-extended- . . . . . . . . . . . . . . . . . 289 command, help- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 command, indent-for-tab-. . . . . . . . . . . . . . . . . . . . . 537 command, keystroke . . . . . . . . . . . . . . . . . . . . . . . . . . 166 command, last- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 command, minor modes, self-insert- . . . . . . . . . . . 374 command, prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 command, prefix-help- . . . . . . . . . . . . . . . . . . . . . . . . 390 command, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 command, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 command, self-insert- . . . . . . . . . . . . . . . . . . . . . . . . . 517 command, start-process-shell-. . . . . . . . . . . . . . . . . 680 command, this-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 command-char, last- . . . . . . . . . . . . . . . . . . . . . . . . . . 292 command-debug-status . . . . . . . . . . . . . . . . . . . . . . . 227 command-event, last- . . . . . . . . . . . . . . . . . . . . . . . . . 291 command-event, next- . . . . . . . . . . . . . . . . . . . . . . . . 305 command-event, unread-. . . . . . . . . . . . . . . . . . . . . . 307 command-events, unread- . . . . . . . . . . . . . . . . . . . . . 306 command-execute . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 command-history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 command-history, extended- . . . . . . . . . . . . . . . . . . 267 command-history, shell-. . . . . . . . . . . . . . . . . . . . . . . 267 command-history-map . . . . . . . . . . . . . . . . . . . . . . . . 783 command-hook, disabled- . . . . . . . . . . . . . . . . . . . . . 314 command-hook, post-. . . . . . . . . . . . . . . . . . . . . . . . . 284 command-hook, pre-. . . . . . . . . . . . . . . . . . . . . . . . . . 283 command-keys, substitute-. . . . . . . . . . . . . . . . . . . . 387 command-keys, this- . . . . . . . . . . . . . . . . . . . . . . . . . . 291 command-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 command-line-args . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 command-line-functions . . . . . . . . . . . . . . . . . . . . . . 697 command-line-processed . . . . . . . . . . . . . . . . . . . . . . 696 command-map, read-shell- . . . . . . . . . . . . . . . . . . . . 785 command-switch-alist. . . . . . . . . . . . . . . . . . . . . . . . . 696 commandp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 commandp example . . . . . . . . . . . . . . . . . . . . . . . . . . 274 commands (Edebug), interactive . . . . . . . . . . . . . . 231 commands, defining . . . . . . . . . . . . . . . . . . . . . . . . . . 284 commands, history of . . . . . . . . . . . . . . . . . . . . . . . . . 314 comment ender . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 comment starter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 comment syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 comment, ; in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 comment, forward- . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 comment, inside . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 comments, header . . . . . . . . . . . . . . . . . . . . . . . . . . . . 765 comments, library header . . . . . . . . . . . . . . . . . . . . . 765 comments, parse-sexp-ignore- . . . . . . . . . . . . . . . . . 579 comments, skipping. . . . . . . . . . . . . . . . . . . . . . . . . . . 579 Common Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Common Lisp (Edebug) . . . . . . . . . . . . . . . . . . . . . . 231 compare-buffer-substrings. . . . . . . . . . . . . . . . . . . . . 515 compared, CL note—special forms . . . . . . . . . . . . 128 comparing buffer text. . . . . . . . . . . . . . . . . . . . . . . . . 515 comparison of modification time . . . . . . . . . . . . . . 439 comparison of, modification time, . . . . . . . . . . . . . 439 comparison, lexical . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 compilation functions . . . . . . . . . . . . . . . . . . . . . . . . . 208 compilation, library . . . . . . . . . . . . . . . . . . . . . . . . . . 209 compilation, macro . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 compile, batch-byte- . . . . . . . . . . . . . . . . . . . . . . . . . . 209 compile, byte- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 compile, eval-and- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 compile, eval-when-. . . . . . . . . . . . . . . . . . . . . . . . . . . 212 compile-defun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 compile-dynamic, byte- . . . . . . . . . . . . . . . . . . . . . . . 211 compile-dynamic-docstrings, byte- . . . . . . . . . . . . 210 compile-file, byte- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 compiled function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 compiled-function-arglist . . . . . . . . . . . . . . . . . . . . . 213 compiled-function-constants . . . . . . . . . . . . . . . . . . 213 compiled-function-doc-string . . . . . . . . . . . . . . . . . . 213 compiled-function-domain . . . . . . . . . . . . . . . . . . . . 214 compiled-function-instructions . . . . . . . . . . . . . . . . 213 compiled-function-interactive . . . . . . . . . . . . . . . . . 213 compiled-function-p . . . . . . . . . . . . . . . . . . . . . . . . . . 166 compiled-function-stack-size . . . . . . . . . . . . . . . . . . 213 compiling macros, byte- . . . . . . . . . . . . . . . . . . . . . . 180 compiling require, byte- . . . . . . . . . . . . . . . . . . . . . . 203 complement, two’s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 complete key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 complete subexpression, previous . . . . . . . . . . . . . 578 complete, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . . . 272 complete-and-exit, minibuffer- . . . . . . . . . . . . . . . . 272 complete-word, minibuffer- . . . . . . . . . . . . . . . . . . . 272 completing-read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 completion subroutines, file name . . . . . . . . . . . . . 413 completion, file name . . . . . . . . . . . . . . . . . . . . . . . . . 413 completion, file-name- . . . . . . . . . . . . . . . . . . . . . . . . 414 completion, interactive . . . . . . . . . . . . . . . . . . . . . . . 286 completion, obarray in. . . . . . . . . . . . . . . . . . . . . . . . 268 completion, programmed . . . . . . . . . . . . . . . . . . . . . 276 completion, try- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 completion-auto-help . . . . . . . . . . . . . . . . . . . . . . . . . 273 completion-confirm, minibuffer- . . . . . . . . . . . . . . . 272 completion-help, minibuffer- . . . . . . . . . . . . . . . . . . 272 completion-ignore-case. . . . . . . . . . . . . . . . . . . . . . . . 270 completion-ignored-extensions . . . . . . . . . . . . . . . . 414 completion-list, display- . . . . . . . . . . . . . . . . . . . . . . 273 completion-map, minibuffer-local- . . . . . . . . 271, 784 completion-predicate, minibuffer- . . . . . . . . . . . . . 272 completion-table, minibuffer- . . . . . . . . . . . . . . . . . 272 completions, all- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 completions, file-name-all- . . . . . . . . . . . . . . . . . . . . 413 complex arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 complex command. . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 complex-buffers-menu-p . . . . . . . . . . . . . . . . . . . . . . 350 components, color-instance-rgb-. . . . . . . . . . . . . . . 626 components, color-rgb- . . . . . . . . . . . . . . . . . . . . . . . 626 components, symbol . . . . . . . . . . . . . . . . . . . . . . . . . . 113 compose-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 composite-char, make- . . . . . . . . . . . . . . . . . . . . . . . . 742 composite-char-string . . . . . . . . . . . . . . . . . . . . . . . . . 742 concat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 concatenating lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 concatenating strings . . . . . . . . . . . . . . . . . . . . . . . . . . 63 cond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 condition name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 condition, edebug-global-break- . . . . . . . . . . 235, 251 condition, edebug-set-global-break- . . . . . . . . . . . 235 condition, global break . . . . . . . . . . . . . . . . . . . . . . . 235 condition-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 conditional evaluation . . . . . . . . . . . . . . . . . . . . . . . . 132 conditions, error- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 configuration (Edebug), window . . . . . . . . . . . . . . 242 configuration, current-frame-. . . . . . . . . . . . . . . . . . 483 configuration, current-window-. . . . . . . . . . . . . . . . 470 configuration, frame . . . . . . . . . . . . . . . . . . . . . . . . . . 483 configuration, menubar- . . . . . . . . . . . . . . . . . . . . . . 342 configuration, set-frame- . . . . . . . . . . . . . . . . . . . . . . 483 configuration, set-window- . . . . . . . . . . . . . . . . . . . . 470 configuration, system- . . . . . . . . . . . . . . . . . . . . . . . . 700 configuration-p, window- . . . . . . . . . . . . . . . . . . . . . 471 configurations, window . . . . . . . . . . . . . . . . . . . . . . . 470 confirm, minibuffer-completion- . . . . . . . . . . . . . . . 272 confoundance disease, char-int . . . . . . . . . . . . . . . . . 20 connection, network . . . . . . . . . . . . . . . . . . . . . . . . . . 691 connection-type, process- . . . . . . . . . . . . . . . . . . . . . 680 cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 cons cell as box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 cons cells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 cons cells, lists and . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 cons-threshold, gc-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 773 conservatively, scroll- . . . . . . . . . . . . . . . . . . . . . . . . . 463 consing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 console, select- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 console, selected- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 console-device-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 console-disable-input . . . . . . . . . . . . . . . . . . . . . . . . . 488 console-enable-input . . . . . . . . . . . . . . . . . . . . . . . . . . 488 console-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 console-live-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 console-type-image-conversion-list. . . . . . . . . . . . . 636 console-type-image-conversion-list, set- . . . . . . . . 636 consolep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 consoles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 consp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 constant, ? in character. . . . . . . . . . . . . . . . . . . . . . . . 23 constant, \ in character . . . . . . . . . . . . . . . . . . . . . . . . 23 constant, backslash in character. . . . . . . . . . . . . . . . 23 constant, question mark in character . . . . . . . . . . . 23 constant, setting- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 constants, compiled-function- . . . . . . . . . . . . . . . . . 213 constituent, symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . 572 constituent, word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 construct, modeline. . . . . . . . . . . . . . . . . . . . . . . . . . . 375 containing parentheses, innermost . . . . . . . . . . . . 578 contents, buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 contents, evaluation of buffer . . . . . . . . . . . . . . . . . 122 contents, insert-file-. . . . . . . . . . . . . . . . . . . . . . . . . . . 398 contents-function, revert-buffer-insert-file- . . . . . 431 contents-hooks, write- . . . . . . . . . . . . . . . . . . . . . . . . 397 context-lines, next-screen- . . . . . . . . . . . . . . . . . . . . 464 continuable errors, CL note—no . . . . . . . . . . . . . . 140 continuation lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 continuation-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 continue-kbd-macro, edebug- . . . . . . . . . . . . . . . . . 251 continue-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 contrib-p, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 contrib-p, set-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . 631 contrib-p-instance, glyph-. . . . . . . . . . . . . . . . . . . . . 631 control character printing . . . . . . . . . . . . . . . . . . . . . 388 control characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 control characters in display . . . . . . . . . . . . . . . . . . 661 control characters, flow . . . . . . . . . . . . . . . . . . . . . . . 712 control characters, reading . . . . . . . . . . . . . . . . . . . . 306 control structures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 control structures, special forms for . . . . . . . . . . . 131 control, enable-flow- . . . . . . . . . . . . . . . . . . . . . . . . . . 713 control, version- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 control-arrow-glyph. . . . . . . . . . . . . . . . . . . . . . . . . . . 642 control-on, enable-flow- . . . . . . . . . . . . . . . . . . . . . . . 713 Control-X-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 controlled in windows, buffers, . . . . . . . . . . . . . . . . 455 controlling precisely, windows, . . . . . . . . . . . . . . . . 455 conventions for writing minor modes . . . . . . . . . . 373 conventions, documentation. . . . . . . . . . . . . . . . . . . 383 conventions, minor mode . . . . . . . . . . . . . . . . . . . . . 373 conversion of image instantiators. . . . . . . . . . . . . . 636 conversion of strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 conversion, file format . . . . . . . . . . . . . . . . . . . . . . . . 419 conversion, image instantiator . . . . . . . . . . . . . . . . 636 conversion, rounding without . . . . . . . . . . . . . . . . . . 54 conversion-list, console-type-image- . . . . . . . . . . . 636 conversion-list, set-console-type-image- . . . . . . . . 636 conversions, rounding in . . . . . . . . . . . . . . . . . . . . . . . 51 copy, file-local- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 417 copy-alist. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 copy-category-table . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 copy-coding-system . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 copy-event . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 copy-extent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 copy-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 copy-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 copy-hashtable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 copy-keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 copy-marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 504 copy-range-table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 copy-region-as-kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 copy-sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 copy-specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 copy-syntax-table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 576 copying alists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 copying bit vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 copying files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 copying lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 copying sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 copying strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 copying vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 copying, backup-by- . . . . . . . . . . . . . . . . . . . . . . . . . . 424 copying-when-linked, backup-by-. . . . . . . . . . . . . . 425 copying-when-mismatch, backup-by- . . . . . . . . . . 425 cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 cosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 count, edebug-display-freq- . . . . . . . . . . . . . . . . . . . 240 count-lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 count-loop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 counting columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 counts, frequency . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 coverage testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 coverage, edebug-test- . . . . . . . . . . . . . . . . . . . . . . . . 250 create, get-buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 create, tq- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 create-device-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 create-file-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 create-frame-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 create-tooltalk-message . . . . . . . . . . . . . . . . . . . . . . . 723 create-tooltalk-pattern. . . . . . . . . . . . . . . . . . . . . . . . 725 creating buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 creating keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 creating, buffers, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 ctl-arrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 ctl-arrow, default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 ctl-x-4-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 ctl-x-5-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 ctl-x-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 cube-root. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 current binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 current buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 current buffer excursion . . . . . . . . . . . . . . . . . . . . . . 497 current buffer mark. . . . . . . . . . . . . . . . . . . . . . . . . . . 507 current buffer point and mark (Edebug) . . . . . . 242 current buffer position . . . . . . . . . . . . . . . . . . . . . . . . 489 current command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 current stack frame . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 current-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 current-buffer (Edebug), eval-. . . . . . . . . . . . . . . . . 231 current-buffer, save- . . . . . . . . . . . . . . . . . . . . . . . . . . 498 current-buffer, with- . . . . . . . . . . . . . . . . . . . . . . . . . . 498 current-case-table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 current-column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 current-fill-column . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 current-frame-configuration . . . . . . . . . . . . . . . . . . . 483 current-global-map . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 current-indentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 current-input-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 current-justification. . . . . . . . . . . . . . . . . . . . . . . . . . . 529 current-keymaps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 current-kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 current-left-margin . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 current-line, justify- . . . . . . . . . . . . . . . . . . . . . . . . . . 529 current-local-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 current-menubar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 current-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 652 current-minor-mode-maps . . . . . . . . . . . . . . . . . . . . 324 current-mouse-event . . . . . . . . . . . . . . . . . . . . . . . . . . 292 current-prefix-arg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 current-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 current-time-string . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 current-time-zone. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 current-window-configuration . . . . . . . . . . . . . . . . . 470 cursor (mouse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 cursor, mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 cursor-in-echo-area . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 cursor-redisplay, force- . . . . . . . . . . . . . . . . . . . . . . . . 649 cust-print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 customization, .emacs . . . . . . . . . . . . . . . . . . . . . . . . 365 cut buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 cut-function, interprogram- . . . . . . . . . . . . . . . . . . . 524 cutbuffer, x-get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 cutbuffer, x-store- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 cyclic ordering of windows . . . . . . . . . . . . . . . . . . . . 453 cyclic, ordering of windows, . . . . . . . . . . . . . . . . . . . 453 cyclic, window ordering, . . . . . . . . . . . . . . . . . . . . . . 453

D

data type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 data, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 data, match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 data, match- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 data, save-match-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 data, set-annotation- . . . . . . . . . . . . . . . . . . . . . . . . . 645 data, set-match- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 data, store-match- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 data-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 database, close-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database, map- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database, open- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database, put- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database, remove- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database-file-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . 674 database-last-error. . . . . . . . . . . . . . . . . . . . . . . . . . . . 674 database-live-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 database-subtype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 674 database-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 674 databasep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 deactivate-region, zmacs- . . . . . . . . . . . . . . . . . . . . . 510 deactivate-region-hook, zmacs- . . . . . . . . . . . . . . . . 511 deallocate-event . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 debug, backtrace-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 debug, error in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 debug, lambda in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 debug-allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 debug-allocation-backtrace. . . . . . . . . . . . . . . . . . . . 774 debug-events, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 debug-ignored-errors. . . . . . . . . . . . . . . . . . . . . . . . . . 220 debug-mode, x-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 debug-on-entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 debug-on-entry, cancel- . . . . . . . . . . . . . . . . . . . . . . . 221 debug-on-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 debug-on-error use. . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 debug-on-next-call . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 debug-on-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 debug-on-signal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 debug-status, command- . . . . . . . . . . . . . . . . . . . . . . 227 debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219, 226 debugger command list . . . . . . . . . . . . . . . . . . . . . . . 223 debugger, Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 debugger-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . 783 debugging errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 debugging specific functions . . . . . . . . . . . . . . . . . . 220 debugging, apply, and . . . . . . . . . . . . . . . . . . . . . . . . 227 debugging, error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 debugging, eval, and . . . . . . . . . . . . . . . . . . . . . . . . . . 227 debugging, funcall, and . . . . . . . . . . . . . . . . . . . . . . . 227 debugging, function call . . . . . . . . . . . . . . . . . . . . . . 220 decimal, integer to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 decode-big5-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 decode-coding-region . . . . . . . . . . . . . . . . . . . . . . . . . 750 decode-shift-jis-char . . . . . . . . . . . . . . . . . . . . . . . . . . 751 decode-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 706 decoding file formats . . . . . . . . . . . . . . . . . . . . . . . . . 419 decompose-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 decrement field of register. . . . . . . . . . . . . . . . . . . . . . 24 dedicated window. . . . . . . . . . . . . . . . . . . . . . . . 457, 459 dedicated, set-window-buffer- . . . . . . . . . . . . . . . . . 457 dedicated-p, set-window- . . . . . . . . . . . . . . . . . . . . . 460 dedicated-p, window- . . . . . . . . . . . . . . . . . . . . 457, 459 deep binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 def-edebug-spec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 defalias. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 default argument string. . . . . . . . . . . . . . . . . . . . . . . 286 default init file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 default optional arg, CL note— . . . . . . . . . . . . . . . 168 default value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 default, auto-save- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 default, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 default, setq- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 default-abbrev-mode. . . . . . . . . . . . . . . . . . . . . . . . . . 581 default-base, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 default-binding, keymap- . . . . . . . . . . . . . . . . . . . . . 319 default-binding, set-keymap- . . . . . . . . . . . . . . . . . . 319 default-boundp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 default-buffer-file-type . . . . . . . . . . . . . . . . . . . . . . . . 422 default-case-fold-search . . . . . . . . . . . . . . . . . . . . . . . 568 default-ctl-arrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 default-deselect-frame-hook . . . . . . . . . . . . . . . . . . . 483 default-directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 default-directory, insert- . . . . . . . . . . . . . . . . . . . . . . 275 default-file-modes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 default-file-modes, set-. . . . . . . . . . . . . . . . . . . . . . . . 408 default-fill-column . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 default-frame-name . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 default-frame-plist . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475 default-host, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 default-init, inhibit- . . . . . . . . . . . . . . . . . . . . . . . . . . 695 default-justification . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 default-major-mode . . . . . . . . . . . . . . . . . . . . . . . . . . 370 default-menubar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 343 default-minibuffer-frame . . . . . . . . . . . . . . . . . . . . . . 480 default-modeline-format . . . . . . . . . . . . . . . . . . . . . . 378 default-p, face-differs-from- . . . . . . . . . . . . . . . . . . . 623 default-popup-menu . . . . . . . . . . . . . . . . . . . . . . . . . . 347 default-port, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 default-select-frame-hook . . . . . . . . . . . . . . . . . . . . . 483 default-sounds, load- . . . . . . . . . . . . . . . . . . . . . . . . . 664 default-text-properties . . . . . . . . . . . . . . . . . . . . . . . . 542 default-toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 default-toolbar-height . . . . . . . . . . . . . . . . . . . . . . . . 356 default-toolbar-position . . . . . . . . . . . . . . . . . . . . . . . 355 default-toolbar-position, set- . . . . . . . . . . . . . . . . . . 355 default-toolbar-visible-p . . . . . . . . . . . . . . . . . . . . . . 357 default-toolbar-width . . . . . . . . . . . . . . . . . . . . . . . . . 356 default-truncate-lines . . . . . . . . . . . . . . . . . . . . . . . . . 650 default-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 default-x-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 default.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 defconst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153, 733 defcustom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 defgroup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 define-abbrev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 define-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 define-derived-mode . . . . . . . . . . . . . . . . . . . . . . . . . . 372 define-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 define-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 define-logical-name . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 define-obsolete-function-alias. . . . . . . . . . . . . . . . . . 391 define-obsolete-variable-alias . . . . . . . . . . . . . . . . . . 392 define-prefix-command. . . . . . . . . . . . . . . . . . . . . . . . 322 define-specifier-tag. . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 defined error, user- . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 defining a function . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 defining commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 defining, commands, . . . . . . . . . . . . . . . . . . . . . . . . . . 284 defining-kbd-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 definition of a symbol. . . . . . . . . . . . . . . . . . . . . . . . . 114 definition, format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 definition, function . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 definition, substitute-key- . . . . . . . . . . . . . . . . . . . . . 331 definition, variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 defmacro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 defs, edebug-all- . . . . . . . . . . . . . . . . . . . . . . . . . 231, 249 defsubst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 defun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 defun (Edebug), eval-. . . . . . . . . . . . . . . . . . . . . . . . . 231 defun, beginning-of- . . . . . . . . . . . . . . . . . . . . . . . . . . 496 defun, compile- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 defun, end-of- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 defun-prompt-regexp . . . . . . . . . . . . . . . . . . . . . . . . . 496 defvar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151, 733 defvaralias. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 deiconify-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 delay, blink-matching-paren- . . . . . . . . . . . . . . . . . . 660 delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 delete previous char . . . . . . . . . . . . . . . . . . . . . . . . . . 519 delete-annotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 delete-auto-save-file-if-necessary. . . . . . . . . . . . . . . 429 delete-auto-save-files . . . . . . . . . . . . . . . . . . . . . . . . . . 430 delete-backward-char . . . . . . . . . . . . . . . . . . . . . . . . . 519 delete-blank-lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 delete-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 delete-char-untabify, backward- . . . . . . . . . . . . . . . 519 delete-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 delete-device-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 delete-directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 delete-exited-processes . . . . . . . . . . . . . . . . . . . . . . . . 681 delete-extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 delete-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 delete-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 delete-frame-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 delete-horizontal-space. . . . . . . . . . . . . . . . . . . . . . . . 519 delete-indentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 520 delete-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 delete-other-windows . . . . . . . . . . . . . . . . . . . . . . . . . 451 delete-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 delete-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 delete-to-left-margin . . . . . . . . . . . . . . . . . . . . . . . . . . 531 delete-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 delete-windows-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 deleting files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 deleting processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 680 deleting whitespace . . . . . . . . . . . . . . . . . . . . . . . . . . . 519 deleting windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 deletion of elements. . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 deletion of frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 deletion vs killing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 deletion, tab. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519 delimiter, page-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 delimiter, paired. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 delims, set-case-syntax- . . . . . . . . . . . . . . . . . . . . . . . . 75 delq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 demibold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 depth, image-instance-. . . . . . . . . . . . . . . . . . . . . . . . 639 depth, max-lisp-eval- . . . . . . . . . . . . . . . . . . . . . . . . . 123 depth, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 depth, parenthesis . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 depth, recursion- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 derived-mode, define- . . . . . . . . . . . . . . . . . . . . . . . . . 372 descendants, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . 598 descent, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 describe-bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 describe-bindings, Helper- . . . . . . . . . . . . . . . . . . . . 391 describe-bindings-internal. . . . . . . . . . . . . . . . . . . . . 336 describe-buffer-case-table . . . . . . . . . . . . . . . . . . . . . . 75 describe-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 describe-prefix-bindings. . . . . . . . . . . . . . . . . . . . . . . 391 describe-tooltalk-message . . . . . . . . . . . . . . . . . . . . . 725 description for interactive codes. . . . . . . . . . . . . . . 285 description format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 description of, codes, interactive,. . . . . . . . . . . . . . 285 description, insert-abbrev-table- . . . . . . . . . . . . . . 582 description, interactive code . . . . . . . . . . . . . . . . . . 285 description, key- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 description, single-key- . . . . . . . . . . . . . . . . . . . . . . . 388 description, text-char- . . . . . . . . . . . . . . . . . . . . . . . . 388 descriptions, command. . . . . . . . . . . . . . . . . . . . . . . . . 12 descriptions, function . . . . . . . . . . . . . . . . . . . . . . . . . . 12 descriptions, macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 descriptions, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 descriptions, special form . . . . . . . . . . . . . . . . . . . . . . 12 descriptions, variable . . . . . . . . . . . . . . . . . . . . . . . . . . 14 descriptor, character . . . . . . . . . . . . . . . . . . . . . . . . . . 662 descriptor, syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 descriptors, argument. . . . . . . . . . . . . . . . . . . . . . . . . 284 deselect-frame-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 484 deselect-frame-hook, default-. . . . . . . . . . . . . . . . . . 483 designator-p, category- . . . . . . . . . . . . . . . . . . . . . . . 757 destroy-tooltalk-message . . . . . . . . . . . . . . . . . . . . . . 724 destroy-tooltalk-pattern . . . . . . . . . . . . . . . . . . . . . . 725 destructive-alist-to-plist. . . . . . . . . . . . . . . . . . . . . . . 100 destructive-plist-to-alist. . . . . . . . . . . . . . . . . . . . . . . 100 detach-extent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 detached extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 detached-p, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 detect-coding-region . . . . . . . . . . . . . . . . . . . . . . . . . . 750 device, default-x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 device, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device, event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 device, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 device, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device, make-tty- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device, make-x-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device, select- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 device, selected- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 device, terminal- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device-baud-rate . . . . . . . . . . . . . . . . . . . . . . . . . 488, 711 device-baud-rate, set- . . . . . . . . . . . . . . . . . . . . 488, 712 device-class. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 device-class-p, valid- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device-frame-list . . . . . . . . . . . . . . . . . . . . . . . . . 479, 486 device-hook, create- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device-hook, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 device-list, console- . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 device-live-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device-matches-specifier-tag-set-p . . . . . . . . . . . . . 611 device-matching-specifier-tag-list . . . . . . . . . . . . . . 611 device-or-frame-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 device-or-frame-type . . . . . . . . . . . . . . . . . . . . . . . . . . 486 device-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 device-type-p, valid- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 device-x-display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 devicep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 devices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 dgettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 731 diagrams, boxed, for lists . . . . . . . . . . . . . . . . . . . . . . 25 diagrams, for lists, box . . . . . . . . . . . . . . . . . . . . . . . . 25 dialog box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 dialog-box, popup- . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 dialog-box, y-or-n-p-maybe- . . . . . . . . . . . . . . . . . . 279 dialog-box, yes-or-no-p- . . . . . . . . . . . . . . . . . . . . . . . 279 dialog-box, yes-or-no-p-maybe-. . . . . . . . . . . . . . . . 279 differs-from-default-p, face- . . . . . . . . . . . . . . . . . . . 623 digit-argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 dimension, charset- . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 ding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 direction, charset- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 direction-charset, charset-reverse- . . . . . . . . . . . . . 739 direction-charset, make-reverse- . . . . . . . . . . . . . . . 739 directories, program . . . . . . . . . . . . . . . . . . . . . . . . . . 676 directory name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 directory name abbreviation . . . . . . . . . . . . . . . . . . 410 directory part (of file name) . . . . . . . . . . . . . . . . . . 408 directory, batch-byte-recompile-. . . . . . . . . . . . . . . 209 directory, byte-recompile- . . . . . . . . . . . . . . . . . . . . . 209 directory, data- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 directory, default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 directory, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 directory, doc- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 directory, exec- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 directory, file name of . . . . . . . . . . . . . . . . . . . . . . . . 410 directory, file names in. . . . . . . . . . . . . . . . . . . . . . . . 414 directory, file-name- . . . . . . . . . . . . . . . . . . . . . . . . . . 409 directory, file-name-as-. . . . . . . . . . . . . . . . . . . . . . . . 410 directory, insert-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 directory, insert-default- . . . . . . . . . . . . . . . . . . . . . . 275 directory, installation- . . . . . . . . . . . . . . . . . . . . . . . . 702 directory, invocation- . . . . . . . . . . . . . . . . . . . . . . . . . 702 directory, list-buffers- . . . . . . . . . . . . . . . . . . . . . . . . . 438 directory, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 directory, temp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 directory, unhandled-file-name- . . . . . . . . . . . . . . . 418 directory, user-home- . . . . . . . . . . . . . . . . . . . . . . . . . 704 directory-abbrev-alist . . . . . . . . . . . . . . . . . . . . . . . . . 410 directory-autoloads, update- . . . . . . . . . . . . . . . . . . 201 directory-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 directory-files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 directory-ignore-errors-p, byte-recompile- . . . . . 210 directory-oriented functions . . . . . . . . . . . . . . . . . . . 414 directory-p, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 directory-p, file-accessible- . . . . . . . . . . . . . . . . . . . . 402 directory-program, insert- . . . . . . . . . . . . . . . . . . . . 415 dired-kept-versions . . . . . . . . . . . . . . . . . . . . . . . . . . . 426 dired-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 dirty-flag, set-menubar- . . . . . . . . . . . . . . . . . . . . . . . 343 disable undo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 disable-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 disable-input, console- . . . . . . . . . . . . . . . . . . . . . . . . 488 disable-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 disable-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 disable-undo, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . 527 disabled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 disabled command. . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 disabled-command-hook . . . . . . . . . . . . . . . . . . . . . . 314 disassemble. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 disassembled byte-code . . . . . . . . . . . . . . . . . . . . . . . 214 discard input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 discard-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 disease, char-int confoundance . . . . . . . . . . . . . . . . . 20 disown-selection, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 dispatch-event . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 dispatching an event. . . . . . . . . . . . . . . . . . . . . . . . . . 305 display columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 display lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 display order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 display table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 display table, active . . . . . . . . . . . . . . . . . . . . . . . . . . 662 display update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 display, $ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 display, \ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 display, control characters in. . . . . . . . . . . . . . . . . . 661 display, device-x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 display, error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 display, momentary-string-. . . . . . . . . . . . . . . . . . . . 659 display, redraw- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 display, refresh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 display, selective . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 display, selective- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 display, update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 display-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 457 display-buffer-function . . . . . . . . . . . . . . . . . . . . . . . . 459 display-buffer-names, special- . . . . . . . . . . . . . . . . . 458 display-completion-list . . . . . . . . . . . . . . . . . . . . . . . . 273 display-ellipses, selective- . . . . . . . . . . . . . . . . . . . . . 657 display-frame-plist, special- . . . . . . . . . . . . . . . . . . . 459 display-freq-count, edebug- . . . . . . . . . . . . . . . . . . . 240 display-function, special-. . . . . . . . . . . . . . . . . . . . . . 459 display-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 display-popup-frame, special- . . . . . . . . . . . . . . . . . 459 display-regexps, special- . . . . . . . . . . . . . . . . . . . . . . 458 display-table, make- . . . . . . . . . . . . . . . . . . . . . . . . . . 662 display-visual-class, x- . . . . . . . . . . . . . . . . . . . . . . . . 718 display-warning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 display-warning-minimum-level . . . . . . . . . . . . . . . 654 display-warning-suppressed-classes . . . . . . . . . . . . 654 displayed-buffer-points, edebug-save- . . . . . 242, 250 displayed-text-pixel-height, window- . . . . . . . . . . 467 displaying a buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 distance, blink-matching-paren- . . . . . . . . . . . . . . . 660 division, arith-error in . . . . . . . . . . . . . . . . . . . . . . . . . 53 DND, OffiX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 do-auto-save. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 DOC (documentation) file . . . . . . . . . . . . . . . . . . . . 383 doc string, string, writing a . . . . . . . . . . . . . . . . . . . 383 doc, function-obsoleteness-. . . . . . . . . . . . . . . . . . . . 392 doc, variable-obsoleteness- . . . . . . . . . . . . . . . . . . . . 392 doc-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 doc-file-name, internal- . . . . . . . . . . . . . . . . . . . . . . . 386 doc-string, charset- . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 doc-string, coding-system- . . . . . . . . . . . . . . . . . . . . 749 doc-string, compiled-function- . . . . . . . . . . . . . . . . 213 docstrings, byte-compile-dynamic- . . . . . . . . . . . . 210 documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 documentation conventions . . . . . . . . . . . . . . . . . . . 383 documentation for major mode . . . . . . . . . . . . . . . 371 documentation notation . . . . . . . . . . . . . . . . . . . . . . . 11 documentation of function . . . . . . . . . . . . . . . . . . . . 169 documentation string, writing a. . . . . . . . . . . . . . . 383 documentation strings . . . . . . . . . . . . . . . . . . . . . . . . 383 documentation strings, keys in . . . . . . . . . . . . . . . . 386 documentation, dynamic loading of . . . . . . . . . . . 210 documentation, keys in . . . . . . . . . . . . . . . . . . . . . . . 386 documentation, Snarf- . . . . . . . . . . . . . . . . . . . . . . . . 386 documentation, substituting keys in. . . . . . . . . . . 386 documentation, variable- . . . . . . . . . . . . . . . . . . . . . 383 documentation-property . . . . . . . . . . . . . . . . . . . . . . 384 domain. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 732 domain (in a specifier) . . . . . . . . . . . . . . . . . . . . . . . . 602 domain, bind-text- . . . . . . . . . . . . . . . . . . . . . . . . . . . 731 domain, compiled-function- . . . . . . . . . . . . . . . . . . . 214 domain, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 domain-of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 732 domain-p, valid-specifier- . . . . . . . . . . . . . . . . . . . . . 614 DOS and file modes, MS- . . . . . . . . . . . . . . . . . . . . . 408 DOS file types, MS- . . . . . . . . . . . . . . . . . . . . . . . . . . 421 DOS, file modes and MS- . . . . . . . . . . . . . . . . . . . . . 408 DOS, file types on MS- . . . . . . . . . . . . . . . . . . . . . . . 421 dotted lists (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . 246 dotted pair notation . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 double-quote in strings . . . . . . . . . . . . . . . . . . . . . . . . 27 down, scroll- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 down-glyph, annotation- . . . . . . . . . . . . . . . . . . . . . . 645 down-glyph, set-annotation- . . . . . . . . . . . . . . . . . . 645 down-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 downcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 downcase-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 downcase-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 downcasing in lookup-key. . . . . . . . . . . . . . . . . . . . . 304 drag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 drag and drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 Drag API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 dribble file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 dribble-file, open-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 Drop API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 drop, drag and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 dt, CDE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 dump-emacs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 duplicable extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 duplicable, extent, . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 dynamic loading of documentation . . . . . . . . . . . . 210 dynamic loading of functions . . . . . . . . . . . . . . . . . 211 dynamic scoping. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 dynamic, byte-compile- . . . . . . . . . . . . . . . . . . . . . . . 211 dynamic-docstrings, byte-compile- . . . . . . . . . . . . 210

E

echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 echo-area, cursor-in- . . . . . . . . . . . . . . . . . . . . . . . . . . 653 echo-area-message, inhibit-startup-. . . . . . . . . . . . 694 echo-keystrokes . . . . . . . . . . . . . . . . . . . . . . . . . . 292, 653 edebug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 Edebug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 Edebug execution modes. . . . . . . . . . . . . . . . . . . . . . 232 Edebug mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 Edebug specification list . . . . . . . . . . . . . . . . . . . . . . 243 edebug-‘. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 edebug-all-defs. . . . . . . . . . . . . . . . . . . . . . . . . . . 231, 249 edebug-all-forms . . . . . . . . . . . . . . . . . . . . . . . . . 231, 250 edebug-continue-kbd-macro . . . . . . . . . . . . . . . . . . . 251 edebug-display-freq-count . . . . . . . . . . . . . . . . . . . . . 240 edebug-eval-top-level-form . . . . . . . . . . . . . . . . . . . . 231 edebug-global-break-condition. . . . . . . . . . . . 235, 251 edebug-initial-mode . . . . . . . . . . . . . . . . . . . . . . . . . . 250 edebug-on-error. . . . . . . . . . . . . . . . . . . . . . . . . . 236, 251 edebug-on-quit . . . . . . . . . . . . . . . . . . . . . . . . . . 236, 251 edebug-print-circle . . . . . . . . . . . . . . . . . . . . . . . 239, 251 edebug-print-length . . . . . . . . . . . . . . . . . . . . . . 239, 251 edebug-print-level. . . . . . . . . . . . . . . . . . . . . . . . 239, 251 edebug-print-trace-after . . . . . . . . . . . . . . . . . . 240, 250 edebug-print-trace-before . . . . . . . . . . . . . . . . 240, 250 edebug-save-displayed-buffer-points . . . . . . 242, 250 edebug-save-windows . . . . . . . . . . . . . . . . . . . . 242, 250 edebug-set-global-break-condition . . . . . . . . . . . . . 235 edebug-setup-hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 edebug-spec, def- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 edebug-test-coverage. . . . . . . . . . . . . . . . . . . . . . . . . . 250 edebug-trace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240, 250 edebug-tracing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 edebug-unwrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 edebug-unwrap-results . . . . . . . . . . . . . . . . . . . 248, 251 edges, window-pixel-. . . . . . . . . . . . . . . . . . . . . . . . . . 468 edges, window-text-area-pixel- . . . . . . . . . . . . . . . . 468 edit, abort-recursive- . . . . . . . . . . . . . . . . . . . . . . . . . 313 edit, exit-recursive- . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 edit, recursive- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 edit-abbrevs-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 edit-and-eval-command . . . . . . . . . . . . . . . . . . . . . . . 266 edit-map, Info- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 edit-map, itimer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 edit-menu-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 edit-tab-stops-map . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 editing level, recursive . . . . . . . . . . . . . . . . . . . . . . . . 312 editing types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 editing, exit recursive . . . . . . . . . . . . . . . . . . . . . . . . . 312 editing, regexps used standardly in . . . . . . . . . . . 568 editing, standard regexps used in . . . . . . . . . . . . . 568 editor command loop . . . . . . . . . . . . . . . . . . . . . . . . . 283 effect, side. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 eighth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 elapsed-time, ccl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 756 elapsed-time, ccl-reset-. . . . . . . . . . . . . . . . . . . . . . . . 756 elc-files, load-ignore- . . . . . . . . . . . . . . . . . . . . . . . . . . 200 electric-buffer-menu-mode-map . . . . . . . . . . . . . . . 783 electric-future-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 electric-history-map . . . . . . . . . . . . . . . . . . . . . . . . . . 783 element (of list) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 element, next-history- . . . . . . . . . . . . . . . . . . . . . . . . 281 element, next-matching-history-. . . . . . . . . . . . . . . 281 element, previous-history- . . . . . . . . . . . . . . . . . . . . 281 element, previous-matching-history-. . . . . . . . . . . 281 elements of sequences. . . . . . . . . . . . . . . . . . . . . . . . . 105 elements, array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 elements, deletion of . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 elements, list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 ellipses, selective-display- . . . . . . . . . . . . . . . . . . . . . 657 elt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 Emacs, CL note—only throw in. . . . . . . . . . . . . . . 137 emacs, dump- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 emacs, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 emacs, suspend- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 emacs-application-class, x- . . . . . . . . . . . . . . . . . . . . 717 emacs-build-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 emacs-from-temacs, run- . . . . . . . . . . . . . . . . . . . . . . 768 emacs-hook, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 emacs-lisp-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . 783 emacs-lisp-mode-syntax-table . . . . . . . . . . . . . . . . . 580 emacs-major-version . . . . . . . . . . . . . . . . . . . . . . . . . . 769 emacs-minor-version . . . . . . . . . . . . . . . . . . . . . . . . . . 769 emacs-pid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 emacs-query-functions, kill- . . . . . . . . . . . . . . . . . . . 698 emacs-version. . . . . . . . . . . . . . . . . . . . . . . . . . . . 768, 769 EMACSLOADPATH environment variable …. 198 embedded breakpoints . . . . . . . . . . . . . . . . . . . . . . . . 235 empty list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 enable-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 enable-flow-control . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 enable-flow-control-on . . . . . . . . . . . . . . . . . . . . . . . . 713 enable-input, console-. . . . . . . . . . . . . . . . . . . . . . . . . 488 enable-local-eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 enable-local-variables . . . . . . . . . . . . . . . . . . . . . . . . . 369 enable-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 enable-predicate, backup- . . . . . . . . . . . . . . . . . . . . . 424 enable-recursive-minibuffers. . . . . . . . . . . . . . . . . . . 282 enable-undo, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . 527 enabled, menu-accelerator-. . . . . . . . . . . . . . . . . . . . 348 encode-big5-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 encode-coding-region . . . . . . . . . . . . . . . . . . . . . . . . . 750 encode-shift-jis-char . . . . . . . . . . . . . . . . . . . . . . . . . . 751 encode-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 encoding file formats . . . . . . . . . . . . . . . . . . . . . . . . . 419 end of buffer marker . . . . . . . . . . . . . . . . . . . . . . . . . . 504 end position, extent . . . . . . . . . . . . . . . . . . . . . . . . . . 588 end, buffer-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 end, match- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 end, region- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 end, sentence- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 end, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 end-glyph, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 end-glyph, set-extent- . . . . . . . . . . . . . . . . . . . . . . . . 597 end-glyph-layout, extent- . . . . . . . . . . . . . . . . . . . . . 596 end-glyph-layout, set-extent-. . . . . . . . . . . . . . . . . . 597 end-of-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 end-of-defun. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 end-of-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 end-of-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 end-position, extent-. . . . . . . . . . . . . . . . . . . . . . . . . . 589 ender, comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 endpoint, extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 endpoints, set-extent-. . . . . . . . . . . . . . . . . . . . . . . . . 589 enlarge-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 enlarge-window-horizontally . . . . . . . . . . . . . . . . . . 469 enlarge-window-pixels . . . . . . . . . . . . . . . . . . . . . . . . 469 enqueue, tq- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 enqueue-eval-event . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 entry, cancel-debug-on- . . . . . . . . . . . . . . . . . . . . . . . 221 entry, debug-on- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 entry, keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 entry, modify-syntax- . . . . . . . . . . . . . . . . . . . . . . . . . 576 environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 environment variable access. . . . . . . . . . . . . . . . . . . 701 environment variable, EMACSLOADPATH … 198 environment variable, HOME . . . . . . . . . . . . . . . . . 675 environment variable, PATH . . . . . . . . . . . . . . . . . . 675 environment variable, TERM . . . . . . . . . . . . . . . . . 696 environment variables, subprocesses. . . . . . . . . . . 676 environment, operating system. . . . . . . . . . . . . . . . 699 environment, process-. . . . . . . . . . . . . . . . . . . . . . . . . 701 eobp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 eof, process-send- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 eolp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 eq, CL note—integers vrs . . . . . . . . . . . . . . . . . . . . . . 49 eq, lax-plists- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 eq, old- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 eq, plists- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 120 equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 equal, char-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 equal, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 equal, lax-plists- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 equal, plists- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 120 equal, string- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 equality, number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 equality, string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 equality, symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 erase-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 error (Edebug), syntax . . . . . . . . . . . . . . . . . . . . . . . 246 error cleanup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 error debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 error display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 error example, arith- . . . . . . . . . . . . . . . . . . . . . . . . . 142 error handler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 error in debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 error in division, arith- . . . . . . . . . . . . . . . . . . . . . . . . 53 error message notation. . . . . . . . . . . . . . . . . . . . . . . . . 12 error name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 error symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 error use, debug-on- . . . . . . . . . . . . . . . . . . . . . . . . . . 140 error with require, load . . . . . . . . . . . . . . . . . . . . . . . 203 error, database-last- . . . . . . . . . . . . . . . . . . . . . . . . . . 674 error, debug-on- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 error, edebug-on- . . . . . . . . . . . . . . . . . . . . . . . . 236, 251 error, evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 error, file mode specification . . . . . . . . . . . . . . . . . . 369 error, file open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 error, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 error, invalid prefix key . . . . . . . . . . . . . . . . . . . . . . . 330 error, key sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 error, Lisp nesting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 error, peculiar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 error, user-defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 error, variable limit . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 error-conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 errors, autoload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 errors, CL note—no continuable . . . . . . . . . . . . . . 140 errors, debug-ignored- . . . . . . . . . . . . . . . . . . . . . . . . 220 errors, debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 errors, handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 errors, load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 errors, signaling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 errors-p, byte-recompile-directory-ignore- . . . . . 210 ESC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 esc-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 ESC-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 escape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22, 573 escape characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 escape characters in printing. . . . . . . . . . . . . . . . . . 258 escape sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 escape-glyph, octal- . . . . . . . . . . . . . . . . . . . . . . . . . . 642 escape-newlines, print-. . . . . . . . . . . . . . . . . . . . . . . . 260 escapes, words-include- . . . . . . . . . . . . . . . . . . . . . . . 491 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 eval, and debugging . . . . . . . . . . . . . . . . . . . . . . . . . . 227 eval, enable-local- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 eval-and-compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 eval-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 eval-command, edit-and-. . . . . . . . . . . . . . . . . . . . . . 266 eval-current-buffer (Edebug) . . . . . . . . . . . . . . . . . . 231 eval-defun (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . 231 eval-depth, max-lisp- . . . . . . . . . . . . . . . . . . . . . . . . . 123 eval-event, enqueue- . . . . . . . . . . . . . . . . . . . . . . . . . . 305 eval-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 eval-expression (Edebug) . . . . . . . . . . . . . . . . . . . . . 232 eval-minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 eval-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 eval-region (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . 231 eval-top-level-form, edebug-. . . . . . . . . . . . . . . . . . . 231 eval-when-compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 evaluated expression argument . . . . . . . . . . . . . . . . 288 evaluating form, self- . . . . . . . . . . . . . . . . . . . . . . . . . 124 evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 evaluation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 evaluation form, argument . . . . . . . . . . . . . . . . . . . . 284 evaluation list (Edebug) . . . . . . . . . . . . . . . . . . . . . . 238 evaluation notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 evaluation of buffer contents . . . . . . . . . . . . . . . . . . 122 evaluation, conditional. . . . . . . . . . . . . . . . . . . . . . . . 132 evaluation, function form . . . . . . . . . . . . . . . . . . . . . 126 evaluation, list form . . . . . . . . . . . . . . . . . . . . . . . . . . 125 evaluation, literal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 evaluation, macro argument . . . . . . . . . . . . . . . . . . 183 evaluation, macro call . . . . . . . . . . . . . . . . . . . . . . . . 126 evaluation, recursive . . . . . . . . . . . . . . . . . . . . . . . . . . 121 evaluation, special form. . . . . . . . . . . . . . . . . . . . . . . 127 evaluation, suspend. . . . . . . . . . . . . . . . . . . . . . . . . . . 313 evaluation, symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 evaluation, vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 event printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 event standard notation, XEmacs . . . . . . . . . . . . . 388 event, character-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 event, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 event, current-mouse-. . . . . . . . . . . . . . . . . . . . . . . . . 292 event, deallocate- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 event, dispatch- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 event, dispatching an . . . . . . . . . . . . . . . . . . . . . . . . . 305 event, enqueue-eval- . . . . . . . . . . . . . . . . . . . . . . . . . . 305 event, last-command- . . . . . . . . . . . . . . . . . . . . . . . . . 291 event, last-input- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 event, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 event, next- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 event, next-command- . . . . . . . . . . . . . . . . . . . . . . . . 305 event, unread-command-. . . . . . . . . . . . . . . . . . . . . . 307 event-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 event-button . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 event-closest-point . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 event-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 event-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 event-glyph-extent. . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-glyph-x-pixel . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-glyph-y-pixel . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-live-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-matches-key-specifier-p . . . . . . . . . . . . . . . . . 321 event-modifier-bits. . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 event-modifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 event-object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 event-over-border-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-over-glyph-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 event-over-modeline-p . . . . . . . . . . . . . . . . . . . . . . . . 298 event-over-text-area-p . . . . . . . . . . . . . . . . . . . . . . . . 298 event-over-toolbar-p . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-p, button- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, button-press- . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, button-release- . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, eval- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, key-press-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, misc-user- . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, motion- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, mouse- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, process-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-p, timeout- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 event-point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 event-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 event-timestamp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-to-character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 event-toolbar-button. . . . . . . . . . . . . . . . . . . . . . . . . . 299 event-type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 event-window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 event-window-x-pixel . . . . . . . . . . . . . . . . . . . . . . . . . 297 event-window-y-pixel . . . . . . . . . . . . . . . . . . . . . . . . . 297 event-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 event-x-pixel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 event-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 event-y-pixel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 eventp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 events, input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 events, stopping on . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 events, translating input . . . . . . . . . . . . . . . . . . . . . . 709 events, unread-command- . . . . . . . . . . . . . . . . . . . . . 306 events, x-debug- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 events-to-keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 examining windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 example, arith-error . . . . . . . . . . . . . . . . . . . . . . . . . . 142 example, commandp . . . . . . . . . . . . . . . . . . . . . . . . . . 274 example, print. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 example, syntax table . . . . . . . . . . . . . . . . . . . . . . . . 366 example, throw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 example, user-variable-p . . . . . . . . . . . . . . . . . . . . . . 274 examples of using interactive . . . . . . . . . . . . . . . . . 288 examples of using, interactive, . . . . . . . . . . . . . . . . 288 exchange-point-and-mark . . . . . . . . . . . . . . . . . . . . . 508 exclusive or, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 exclusive or, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 excursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 excursion (Edebug), save-. . . . . . . . . . . . . . . . . . . . . 242 excursion, current buffer . . . . . . . . . . . . . . . . . . . . . . 497 excursion, mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 excursion, point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 excursion, save-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 excursion, save-window- . . . . . . . . . . . . . . . . . . . . . . 471 excursions, window . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 exec-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 exec-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 executable-p, file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 execute program. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 execute with prefix argument . . . . . . . . . . . . . . . . . 289 execute, ccl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 755 execute, command- . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 execute-extended-command . . . . . . . . . . . . . . . . . . . 289 execute-kbd-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 execute-on-string, ccl- . . . . . . . . . . . . . . . . . . . . . . . . 756 executing-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 execution modes, Edebug . . . . . . . . . . . . . . . . . . . . . 232 execution speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 execution, keyboard macro . . . . . . . . . . . . . . . . . . . 289 exists, file-already-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 exists-p, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 exists-p, region- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 exit recursive editing . . . . . . . . . . . . . . . . . . . . . . . . . 312 exit, minibuffer-complete-and- . . . . . . . . . . . . . . . . 272 exit, self-insert-and- . . . . . . . . . . . . . . . . . . . . . . . . . . 280 exit-hook, minibuffer-. . . . . . . . . . . . . . . . . . . . . . . . . 281 exit-minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 exit-recursive-edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 exit-status, process- . . . . . . . . . . . . . . . . . . . . . . . . . . 683 exited-processes, delete- . . . . . . . . . . . . . . . . . . . . . . 681 exiting XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 exits, nonlocal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 exp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 expand-abbrev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 expand-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 expand-hook, pre-abbrev-. . . . . . . . . . . . . . . . . . . . . 584 expansion of file names . . . . . . . . . . . . . . . . . . . . . . . 411 expansion of macros . . . . . . . . . . . . . . . . . . . . . . . . . . 179 expansion, abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 expansion, macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 explicit, filling, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 expression (Edebug), eval- . . . . . . . . . . . . . . . . . . . . 232 expression argument, evaluated . . . . . . . . . . . . . . . 288 expression in hook, lambda . . . . . . . . . . . . . . . . . . . 380 expression motion, Lisp. . . . . . . . . . . . . . . . . . . . . . . 495 expression prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 expression searching, regular . . . . . . . . . . . . . . . . . . 559 expression, lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 expression, regular. . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 expression, value of . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 expression-history, read- . . . . . . . . . . . . . . . . . . . . . . 267 expression-map, read- . . . . . . . . . . . . . . . . . . . . . . . . 785 expressions (Edebug), anonymous lambda. . . . . 231 expt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 extended-command, execute- . . . . . . . . . . . . . . . . . 289 extended-command-history . . . . . . . . . . . . . . . . . . . 267 extension, file-name-sans- . . . . . . . . . . . . . . . . . . . . . 409 extensions, completion-ignored- . . . . . . . . . . . . . . . 414 extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156, 587 extent children . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent end position . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent endpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent priority. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 extent property. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 extent replica. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent start position . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent, atomic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 extent, children, of . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent, detach- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 extent, detached . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 extent, duplicable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent, event-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 extent, force-highlight-. . . . . . . . . . . . . . . . . . . . . . . . 600 extent, highlight- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 extent, insert- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent, next- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 extent, parent, of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent, previous- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591 extent, priority of an . . . . . . . . . . . . . . . . . . . . . . . . . 587 extent, property of an . . . . . . . . . . . . . . . . . . . . . . . . 593 extent, unique . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent, zero-length. . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent-at . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 extent-begin-glyph. . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 extent-begin-glyph, set-. . . . . . . . . . . . . . . . . . . . . . . 597 extent-begin-glyph-layout . . . . . . . . . . . . . . . . . . . . . 596 extent-begin-glyph-layout, set- . . . . . . . . . . . . . . . . 597 extent-children . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent-children, map- . . . . . . . . . . . . . . . . . . . . . . . . . 593 extent-descendants . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent-detached-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 extent-end-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 extent-end-glyph, set- . . . . . . . . . . . . . . . . . . . . . . . . 597 extent-end-glyph-layout. . . . . . . . . . . . . . . . . . . . . . . 596 extent-end-glyph-layout, set-. . . . . . . . . . . . . . . . . . 597 extent-end-position . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 extent-endpoints, set-. . . . . . . . . . . . . . . . . . . . . . . . . 589 extent-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 extent-face, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 extent-in-region-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 extent-initial-redisplay-function, set- . . . . . . . . . . 597 extent-keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 extent-keymap, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . 597 extent-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 extent-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 extent-live-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent-mouse-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 extent-mouse-face, set- . . . . . . . . . . . . . . . . . . . . . . . 597 extent-object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extent-parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent-parent, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 extent-priority. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 extent-priority, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 extent-properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 extent-properties, set- . . . . . . . . . . . . . . . . . . . . . . . . 594 extent-property. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 extent-property, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 593 extent-start-position . . . . . . . . . . . . . . . . . . . . . . . . . . 589 extentp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 extents, locating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 extents, map-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591 extents, mapcar- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 592 extents, mapping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591 extents, markers vs. . . . . . . . . . . . . . . . . . . . . . . . . . . 501 extents, order of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 extents, unique . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598

F

face type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 face, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 face, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 face, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 face, extent-mouse- . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 face, glyph-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 face, invert- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 face, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 face, set-annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . 645 face, set-extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 face, set-extent-mouse- . . . . . . . . . . . . . . . . . . . . . . . 597 face, set-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 face-background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 face-background, set- . . . . . . . . . . . . . . . . . . . . . . . . . 622 face-background-instance . . . . . . . . . . . . . . . . . . . . . 622 face-background-pixmap . . . . . . . . . . . . . . . . . . . . . . 622 face-background-pixmap, set- . . . . . . . . . . . . . . . . . 622 face-background-pixmap-instance . . . . . . . . . . . . . 623 face-boolean-specifier-p . . . . . . . . . . . . . . . . . . . . . . . 605 face-differs-from-default-p. . . . . . . . . . . . . . . . . . . . . 623 face-equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 face-font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 face-font, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 face-font-instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 face-font-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 face-foreground . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 face-foreground, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 621 face-foreground-instance . . . . . . . . . . . . . . . . . . . . . . 622 face-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 face-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 face-property, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 face-property-instance . . . . . . . . . . . . . . . . . . . . . . . . 621 face-underline-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 face-underline-p, set- . . . . . . . . . . . . . . . . . . . . . . . . . 622 facep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 faces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 617 failed, search- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 fallback (in a specifier) . . . . . . . . . . . . . . . . . . . . . . . 603 fallback, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . . . 603 fallback, specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 fboundp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 fceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 feature, unload- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 featurep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203, 205 features, providing. . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 features, requiring . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 fetch-bytecode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 ffloor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 field of register, address. . . . . . . . . . . . . . . . . . . . . . . . 24 field of register, decrement . . . . . . . . . . . . . . . . . . . . . 24 field width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 fields, sort- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 fields, sort-numeric- . . . . . . . . . . . . . . . . . . . . . . . . . . 534 fields, sort-regexp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 fifth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 file accessibility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 file age . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 file attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 file format conversion . . . . . . . . . . . . . . . . . . . . . . . . . 419 file formats, decoding . . . . . . . . . . . . . . . . . . . . . . . . . 419 file formats, encoding . . . . . . . . . . . . . . . . . . . . . . . . . 419 file hard link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 file locks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 file message, new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 file mode specification error. . . . . . . . . . . . . . . . . . . 369 file mode, visited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 file modes and MS-DOS . . . . . . . . . . . . . . . . . . . . . . 408 file modes, MS-DOS and. . . . . . . . . . . . . . . . . . . . . . 408 file modification time . . . . . . . . . . . . . . . . . . . . . . . . . 402 file name completion subroutines. . . . . . . . . . . . . . 413 file name of buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 file name of directory . . . . . . . . . . . . . . . . . . . . . . . . . 410 file name), directory part (of. . . . . . . . . . . . . . . . . . 408 file name), nondirectory part (of . . . . . . . . . . . . . . 408 file name), version number (in . . . . . . . . . . . . . . . . 408 file name, absolute. . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 file name, buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 file name, completion, . . . . . . . . . . . . . . . . . . . . . . . . 413 file name, relative . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 file names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 file names in directory . . . . . . . . . . . . . . . . . . . . . . . . 414 file names, expansion of. . . . . . . . . . . . . . . . . . . . . . . 411 file names, magic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 file open error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 file symbolic links . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 file types on MS-DOS . . . . . . . . . . . . . . . . . . . . . . . . 421 file types, MS-DOS . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 file with multiple names . . . . . . . . . . . . . . . . . . . . . . 406 file), truename (of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 file, accessibility of a. . . . . . . . . . . . . . . . . . . . . . . . . . 401 file, add-name-to-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 file, after-find- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 file, append-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 file, backup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 file, byte-compile- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 file, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 file, default init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 file, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 file, DOC (documentation) . . . . . . . . . . . . . . . . . . . 383 file, dribble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 file, end-of- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 file, find- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 file, format-find- . . . . . . . . . . . . . . . . . . . . . . . . . 420, 421 file, format-insert- . . . . . . . . . . . . . . . . . . . . . . . 420, 421 file, format-write-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 file, init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 file, load-sound- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 664 file, locate- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 file, open-dribble-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 file, play-sound- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 file, quietly-read-abbrev- . . . . . . . . . . . . . . . . . . . . . . 583 file, rename- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 file, rename-auto-save- . . . . . . . . . . . . . . . . . . . . . . . . 430 file, site-run- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 file, termscript. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 file, view- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 file, visited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 file, with-temp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 file, write- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 file, write-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 file-accessible-directory-p . . . . . . . . . . . . . . . . . . . . . 402 file-already-exists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 file-attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 file-autoloads, update- . . . . . . . . . . . . . . . . . . . . . . . . 201 file-binary, find- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 file-buffer, create-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 file-buffer, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 file-clear-hashing, locate- . . . . . . . . . . . . . . . . . . . . . 199 file-contents, insert-. . . . . . . . . . . . . . . . . . . . . . . . . . . 398 file-contents-function, revert-buffer-insert- . . . . . 431 file-directory-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 file-error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 file-executable-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 file-exists-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 file-format, auto-save- . . . . . . . . . . . . . . . . . . . . . . . . 421 file-format, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 file-functions, after-insert- . . . . . . . . . . . . . . . . . . . . 546 file-hooks, find- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 file-hooks, local-write- . . . . . . . . . . . . . . . . . . . . . . . . 397 file-hooks, write-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 file-if-necessary, delete-auto-save- . . . . . . . . . . . . . 429 file-local-copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 417 file-locked . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 file-locked-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 file-menu-filter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 file-modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 file-modes, default- . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 file-modes, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 file-modes, set-default-. . . . . . . . . . . . . . . . . . . . . . . . 408 file-modtime, clear-visited-. . . . . . . . . . . . . . . . . . . . 439 file-modtime, set-visited-. . . . . . . . . . . . . . . . . . . . . . 439 file-modtime, verify-visited- . . . . . . . . . . . . . . . . . . . 439 file-modtime, visited- . . . . . . . . . . . . . . . . . . . . . . . . . 439 file-name, abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 file-name, abbreviate-. . . . . . . . . . . . . . . . . . . . . . . . . 410 file-name, auto-save-list- . . . . . . . . . . . . . . . . . . . . . . 430 file-name, auto-save-visited-. . . . . . . . . . . . . . . . . . . 428 file-name, buffer- . . . . . . . . . . . . . . . . . . . . . . . . 436, 437 file-name, buffer-auto-save- . . . . . . . . . . . . . . . . . . . 427 file-name, database- . . . . . . . . . . . . . . . . . . . . . . . . . . 674 file-name, directory- . . . . . . . . . . . . . . . . . . . . . . . . . . 410 file-name, expand- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 file-name, find-backup- . . . . . . . . . . . . . . . . . . . . . . . 427 file-name, image-instance- . . . . . . . . . . . . . . . . . . . . 639 file-name, image-instance-mask-. . . . . . . . . . . . . . . 639 file-name, internal-doc- . . . . . . . . . . . . . . . . . . . . . . . 386 file-name, make-auto-save- . . . . . . . . . . . . . . . . . . . . 428 file-name, make-backup- . . . . . . . . . . . . . . . . . . . . . . 426 file-name, read-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 file-name, set-visited- . . . . . . . . . . . . . . . . . . . . . . . . . 437 file-name, substitute-in-. . . . . . . . . . . . . . . . . . . . . . . 412 file-name-absolute-p . . . . . . . . . . . . . . . . . . . . . . . . . . 411 file-name-all-completions. . . . . . . . . . . . . . . . . . . . . . 413 file-name-as-directory . . . . . . . . . . . . . . . . . . . . . . . . . 410 file-name-buffer-file-type-alist . . . . . . . . . . . . . . . . . 421 file-name-completion. . . . . . . . . . . . . . . . . . . . . . . . . . 414 file-name-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . 409 file-name-directory, unhandled- . . . . . . . . . . . . . . . 418 file-name-handler, find- . . . . . . . . . . . . . . . . . . . . . . . 417 file-name-handlers, inhibit- . . . . . . . . . . . . . . . . . . . 417 file-name-history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 file-name-nondirectory . . . . . . . . . . . . . . . . . . . . . . . . 409 file-name-operation, inhibit- . . . . . . . . . . . . . . . . . . 417 file-name-p, auto-save- . . . . . . . . . . . . . . . . . . . . . . . . 428 file-name-p, backup- . . . . . . . . . . . . . . . . . . . . . . . . . . 426 file-name-sans-extension . . . . . . . . . . . . . . . . . . . . . . 409 file-name-sans-versions. . . . . . . . . . . . . . . . . . . . . . . . 409 file-newer-than-file-p . . . . . . . . . . . . . . . . . . . . . . . . . . 402 file-newest-backup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 file-nlinks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 file-noselect, find-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 file-not-found-hooks, find- . . . . . . . . . . . . . . . . . . . . 395 file-number, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 file-other-window, find- . . . . . . . . . . . . . . . . . . . . . . . 394 file-ownership-preserved-p . . . . . . . . . . . . . . . . . . . . 402 file-p, file-newer-than- . . . . . . . . . . . . . . . . . . . . . . . . 402 file-part, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 file-path, x-bitmap- . . . . . . . . . . . . . . . . . . . . . . 636, 719 file-precious-flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 file-prefix, term- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 file-read-only, find- . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 file-readable-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 file-regular-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 file-relative-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 file-supersession . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 file-symlink-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 file-text, find- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 file-truename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 file-truename, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . 437 file-type, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 file-type, default-buffer-. . . . . . . . . . . . . . . . . . . . . . . 422 file-type, find-buffer-. . . . . . . . . . . . . . . . . . . . . . . . . . 421 file-type-alist, file-name-buffer-. . . . . . . . . . . . . . . . 421 file-writable-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 files and binary files, text . . . . . . . . . . . . . . . . . . . . . 421 files and text files, binary . . . . . . . . . . . . . . . . . . . . . 421 files, binary files and text . . . . . . . . . . . . . . . . . . . . . 421 files, copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 files, delete-auto-save- . . . . . . . . . . . . . . . . . . . . . . . . 430 files, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 files, directory- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 files, finding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 files, how to make them, backup . . . . . . . . . . . . . . 424 files, linking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 files, load-ignore-elc- . . . . . . . . . . . . . . . . . . . . . . . . . . 200 files, make-backup- . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 files, partial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 files, renaming. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 files, setting modes of. . . . . . . . . . . . . . . . . . . . . . . . . 406 files, text files and binary . . . . . . . . . . . . . . . . . . . . . 421 files, text properties in. . . . . . . . . . . . . . . . . . . . . . . . 545 files, visiting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 Fill mode, Auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 Fill mode, newline and Auto . . . . . . . . . . . . . . . . . . 517 fill-column. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 fill-column, current- . . . . . . . . . . . . . . . . . . . . . . . . . . 530 fill-column, default-. . . . . . . . . . . . . . . . . . . . . . . . . . . 530 fill-function, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 fill-individual-paragraphs . . . . . . . . . . . . . . . . . . . . . 528 fill-individual-varying-indent . . . . . . . . . . . . . . . . . . 529 fill-paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 fill-paragraph-function . . . . . . . . . . . . . . . . . . . . . . . . 529 fill-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 fill-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 fill-region-as-paragraph . . . . . . . . . . . . . . . . . . . . . . . 529 fillarray . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 filling a paragraph . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 filling, automatic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 filling, explicit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 filter function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 filter, buffers-menu-. . . . . . . . . . . . . . . . . . . . . . . . . . . 346 filter, edit-menu- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 filter, file-menu- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 filter, process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 filter, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 filter, set-process-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 filters, menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 final, charset-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 final-newline, require-. . . . . . . . . . . . . . . . . . . . . . . . . 398 find-backup-file-name . . . . . . . . . . . . . . . . . . . . . . . . . 427 find-buffer-file-type . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 find-charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 find-charset-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 find-charset-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 find-coding-system. . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 find-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 find-file, after- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 find-file, format- . . . . . . . . . . . . . . . . . . . . . . . . . 420, 421 find-file-binary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 find-file-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 find-file-name-handler . . . . . . . . . . . . . . . . . . . . . . . . 417 find-file-noselect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 find-file-not-found-hooks . . . . . . . . . . . . . . . . . . . . . . 395 find-file-other-window . . . . . . . . . . . . . . . . . . . . . . . . 394 find-file-read-only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 find-file-text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 find-larger-font, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 find-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 find-smaller-font, x- . . . . . . . . . . . . . . . . . . . . . . . . . . 624 finding files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 finding windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 first . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 first-change-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 549 fixup-whitespace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 520 flag (of buffer), modification . . . . . . . . . . . . . . . . . . 438 flag, file-precious-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 flag, purify- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 flag, quit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 flag, set-menubar-dirty- . . . . . . . . . . . . . . . . . . . . . . . 343 flags, syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 float. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 float-output-format . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 floating point, IEEE . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 floating-point numbers, printing . . . . . . . . . . . . . . 261 floatp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 floor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 flow control characters . . . . . . . . . . . . . . . . . . . . . . . . 712 flow-control, enable- . . . . . . . . . . . . . . . . . . . . . . . . . . 713 flow-control-on, enable- . . . . . . . . . . . . . . . . . . . . . . . 713 flush input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 flush-undo, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 fmakunbound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 focus, input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 focus-frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 fold-search, case- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 fold-search, default-case- . . . . . . . . . . . . . . . . . . . . . . 568 following-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 font instance characteristics. . . . . . . . . . . . . . . . . . . 624 font instance name . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 font instance size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 font instance type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 font instances, characteristics of . . . . . . . . . . . . . . 624 font, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 font, set-face-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 font, x-find-larger- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 font, x-find-smaller- . . . . . . . . . . . . . . . . . . . . . . . . . . 624 font-bold, x-make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 font-bold-italic, x-make- . . . . . . . . . . . . . . . . . . . . . . 625 font-instance, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 font-instance, make- . . . . . . . . . . . . . . . . . . . . . . . . . . 623 font-instance-name . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 font-instance-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 font-instance-properties . . . . . . . . . . . . . . . . . . . . . . . 624 font-instance-truename . . . . . . . . . . . . . . . . . . . . . . . 624 font-italic, x-make- . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 font-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 font-name, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 font-properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 font-size, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 font-specifier-p. . . . . . . . . . . . . . . . . . . . . . . . . . . 605, 623 font-truename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 font-unbold, x-make- . . . . . . . . . . . . . . . . . . . . . . . . . 624 font-unitalic, x-make- . . . . . . . . . . . . . . . . . . . . . . . . . 625 fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 623 fonts available . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 fonts, available . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 fonts, list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 foo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 for, sit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 for, sleep- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 for-tab-command, indent-. . . . . . . . . . . . . . . . . . . . . 537 for-user-input-p, waiting- . . . . . . . . . . . . . . . . . . . . . 690 force-cursor-redisplay . . . . . . . . . . . . . . . . . . . . . . . . . 649 force-highlight-extent . . . . . . . . . . . . . . . . . . . . . . . . . 600 forcing redisplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 foreground, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 foreground, image-instance- . . . . . . . . . . . . . . . . . . . 639 foreground, set-face- . . . . . . . . . . . . . . . . . . . . . . . . . . 621 foreground-instance, face-. . . . . . . . . . . . . . . . . . . . . 622 form descriptions, special . . . . . . . . . . . . . . . . . . . . . . 12 form evaluation, function . . . . . . . . . . . . . . . . . . . . . 126 form evaluation, list . . . . . . . . . . . . . . . . . . . . . . . . . . 125 form evaluation, special. . . . . . . . . . . . . . . . . . . . . . . 127 form, argument evaluation . . . . . . . . . . . . . . . . . . . . 284 form, edebug-eval-top-level-. . . . . . . . . . . . . . . . . . . 231 form, help- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 form, minibuffer-help- . . . . . . . . . . . . . . . . . . . . . . . . 281 form, self-evaluating . . . . . . . . . . . . . . . . . . . . . . . . . . 124 form, top-level. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 format arguments, repositioning . . . . . . . . . . . . . . . 71 format conversion, file . . . . . . . . . . . . . . . . . . . . . . . . 419 format definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 format of keymaps. . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 format of menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 format of the menubar. . . . . . . . . . . . . . . . . . . . . . . . 342 format precision . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 format specification. . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 format, % in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 format, auto-save-file- . . . . . . . . . . . . . . . . . . . . . . . . 421 format, buffer-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 format, default-modeline- . . . . . . . . . . . . . . . . . . . . . 378 format, description . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 format, float-output- . . . . . . . . . . . . . . . . . . . . . . . . . 261 format, frame-icon-title- . . . . . . . . . . . . . . . . . . . . . . 478 format, frame-title- . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 format, keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 format, menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 format, menubar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 format, modeline- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 375 format, Shell mode modeline- . . . . . . . . . . . . . . . . . 376 format-alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419 format-buffers-menu-line . . . . . . . . . . . . . . . . . . . . . . 350 format-find-file . . . . . . . . . . . . . . . . . . . . . . . . . . 420, 421 format-insert-file . . . . . . . . . . . . . . . . . . . . . . . . . 420, 421 format-list, image-instantiator-. . . . . . . . . . . . . . . . 636 format-p, valid-image-instantiator- . . . . . . . . . . . . 636 format-time-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 705 format-write-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 formats, decoding file . . . . . . . . . . . . . . . . . . . . . . . . . 419 formats, encoding file . . . . . . . . . . . . . . . . . . . . . . . . . 419 formatted numbers, precision of. . . . . . . . . . . . . . . . 72 formatting strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 formatting them, strings, . . . . . . . . . . . . . . . . . . . . . . 69 formatting, multilingual string . . . . . . . . . . . . . . . . . 71 formfeed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 forms (Edebug), special. . . . . . . . . . . . . . . . . . . . . . . 231 forms compared, CL note—special . . . . . . . . . . . . 128 forms for control structures, special . . . . . . . . . . . 131 forms, cleanup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 forms, edebug-all-. . . . . . . . . . . . . . . . . . . . . . . . 231, 250 forms, protected . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 forms, special. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 forward, posix-search- . . . . . . . . . . . . . . . . . . . . . . . . 562 forward, re-search-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 forward, search- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 forward, skip-chars-. . . . . . . . . . . . . . . . . . . . . . . . . . . 496 forward, skip-syntax- . . . . . . . . . . . . . . . . . . . . . . . . . 577 forward, word-search-. . . . . . . . . . . . . . . . . . . . . . . . . 552 forward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 forward-comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 forward-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 forward-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 forward-sexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 forward-to-indentation . . . . . . . . . . . . . . . . . . . . . . . . 540 forward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 found-hooks, find-file-not- . . . . . . . . . . . . . . . . . . . . 395 fourth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 frame configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 frame hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 frame name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame of terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 frame position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame visibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, auto-lower- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 frame, auto-raise-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 frame, backtrace- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 frame, current stack . . . . . . . . . . . . . . . . . . . . . . . . . . 222 frame, default-minibuffer-. . . . . . . . . . . . . . . . . . . . . 480 frame, deiconify- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 frame, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 frame, focus- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 frame, iconified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, iconify- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, invisible . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, lower- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 frame, lowering a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 frame, next- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 frame, position of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame, previous- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 frame, raise- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 frame, raising a. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, redraw-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 frame, save-selected- . . . . . . . . . . . . . . . . . . . . . . . . . . 481 frame, select- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 frame, selected . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 frame, selected-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 frame, size of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame, special-display-popup- . . . . . . . . . . . . . . . . . 459 frame, terminal . . . . . . . . . . . . . . . . . . . . . . . . . . 448, 473 frame, visible . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 frame, with-selected- . . . . . . . . . . . . . . . . . . . . . . . . . 481 frame, X window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 frame-configuration, current-. . . . . . . . . . . . . . . . . . 483 frame-configuration, set- . . . . . . . . . . . . . . . . . . . . . . 483 frame-device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 frame-function, pop-up- . . . . . . . . . . . . . . . . . . . . . . 458 frame-height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame-hook, create- . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 frame-hook, default-deselect-. . . . . . . . . . . . . . . . . . 483 frame-hook, default-select- . . . . . . . . . . . . . . . . . . . . 483 frame-hook, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 frame-hook, deselect- . . . . . . . . . . . . . . . . . . . . . . . . . 484 frame-hook, map- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 frame-hook, select- . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 frame-hook, unmap- . . . . . . . . . . . . . . . . . . . . . . . . . . 484 frame-icon-pixmap, x-set- . . . . . . . . . . . . . . . . . . . . . 478 frame-icon-title-format. . . . . . . . . . . . . . . . . . . . . . . . 478 frame-iconified-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame-invisible, make- . . . . . . . . . . . . . . . . . . . . . . . . 482 frame-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 frame-list, device-. . . . . . . . . . . . . . . . . . . . . . . . 479, 486 frame-list, visible- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 frame-live-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 frame-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame-name, default- . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame-p, device-or- . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 frame-pixel-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame-pixel-width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame-plist, default- . . . . . . . . . . . . . . . . . . . . . . . . . . 475 frame-plist, initial- . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 frame-plist, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . 474 frame-plist, pop-up- . . . . . . . . . . . . . . . . . . . . . . . . . . 458 frame-plist, special-display- . . . . . . . . . . . . . . . . . . . 459 frame-pointer, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 frame-position, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame-properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 frame-properties, set- . . . . . . . . . . . . . . . . . . . . . . . . . 474 frame-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 frame-property, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 474 frame-root-window . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 frame-selected-window . . . . . . . . . . . . . . . . . . . . . . . . 480 frame-size, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 frame-title-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 frame-top-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 frame-totally-visible-p . . . . . . . . . . . . . . . . . . . . . . . . 482 frame-type, device-or- . . . . . . . . . . . . . . . . . . . . . . . . 486 frame-visible, make- . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame-visible-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 frame-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 framep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 frames, deletion of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 frames, pop-up- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 free list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 freq-count, edebug-display- . . . . . . . . . . . . . . . . . . . 240 frequency counts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 fround. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 fset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 ftp-login. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ftruncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 full-name, user-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 fullness, hashtable- . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 fullness, keymap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 funcall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 funcall, and debugging. . . . . . . . . . . . . . . . . . . . . . . . 227 function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165, 175 function call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 function call debugging . . . . . . . . . . . . . . . . . . . . . . . 220 function cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 function cell in autoload . . . . . . . . . . . . . . . . . . . . . . 201 function cell, void . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 function definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 function descriptions. . . . . . . . . . . . . . . . . . . . . . . . . . . 12 function form evaluation . . . . . . . . . . . . . . . . . . . . . . 126 function indirection, symbol . . . . . . . . . . . . . . . . . . 125 function input stream . . . . . . . . . . . . . . . . . . . . . . . . 254 function invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 function name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 function output stream . . . . . . . . . . . . . . . . . . . . . . . 257 function quoting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 function, after-change- . . . . . . . . . . . . . . . . . . . . . . . . 549 function, anonymous . . . . . . . . . . . . . . . . . . . . . . . . . 174 function, auto-fill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 function, before-change- . . . . . . . . . . . . . . . . . . . . . . 549 function, blink-paren-. . . . . . . . . . . . . . . . . . . . . . . . . 660 function, body of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 function, buffers-menu-switch-to-buffer- . . . . . . . 350 function, built-in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 function, byte-code . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 function, calling a . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 function, compiled . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 function, define- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 function, defining a . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 function, display-buffer- . . . . . . . . . . . . . . . . . . . . . . 459 function, documentation of . . . . . . . . . . . . . . . . . . . 169 function, event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 function, fill-paragraph- . . . . . . . . . . . . . . . . . . . . . . 529 function, filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 function, indent-line- . . . . . . . . . . . . . . . . . . . . . . . . . 536 function, indent-region- . . . . . . . . . . . . . . . . . . . . . . . 537 function, indirect- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 function, interactive . . . . . . . . . . . . . . . . . . . . . . . . . . 284 function, interprogram-cut- . . . . . . . . . . . . . . . . . . . 524 function, interprogram-paste- . . . . . . . . . . . . . . . . . 524 function, invalid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 function, invalid- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 function, key translation . . . . . . . . . . . . . . . . . . . . . . 710 function, load-read-. . . . . . . . . . . . . . . . . . . . . . . . . . . 199 function, named . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 function, pop-up-frame- . . . . . . . . . . . . . . . . . . . . . . 458 function, revert-buffer-. . . . . . . . . . . . . . . . . . . . . . . . 431 function, revert-buffer-insert-file-contents- . . . . . 431 function, set-extent-initial-redisplay- . . . . . . . . . . 597 function, special-display-. . . . . . . . . . . . . . . . . . . . . . 459 function, symbol- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 function, temp-buffer-show- . . . . . . . . . . . . . . . . . . . 659 function, void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 function, void-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 function-alias, define-obsolete- . . . . . . . . . . . . . . . . 391 function-arglist, compiled- . . . . . . . . . . . . . . . . . . . . 213 function-constants, compiled- . . . . . . . . . . . . . . . . . 213 function-doc-string, compiled- . . . . . . . . . . . . . . . . 213 function-domain, compiled- . . . . . . . . . . . . . . . . . . . 214 function-instructions, compiled- . . . . . . . . . . . . . . . 213 function-interactive . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 function-interactive, compiled- . . . . . . . . . . . . . . . . 213 function-key-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 function-obsoleteness-doc . . . . . . . . . . . . . . . . . . . . . 392 function-p, compiled- . . . . . . . . . . . . . . . . . . . . . . . . . 166 function-stack-size, compiled- . . . . . . . . . . . . . . . . . 213 functionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 functions in modes . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 functions, after-change- . . . . . . . . . . . . . . . . . . . . . . . 549 functions, after-insert-file- . . . . . . . . . . . . . . . . . . . . 546 functions, before-change- . . . . . . . . . . . . . . . . . . . . . 548 functions, command-line- . . . . . . . . . . . . . . . . . . . . . 697 functions, compilation . . . . . . . . . . . . . . . . . . . . . . . . 208 functions, debugging specific . . . . . . . . . . . . . . . . . . 220 functions, directory-oriented . . . . . . . . . . . . . . . . . . 414 functions, dynamic loading of . . . . . . . . . . . . . . . . . 211 functions, inline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 functions, kill-buffer-query- . . . . . . . . . . . . . . . . . . . 444 functions, kill-emacs-query- . . . . . . . . . . . . . . . . . . . 698 functions, making them interactive. . . . . . . . . . . . 284 functions, mapping . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 functions, mathematical . . . . . . . . . . . . . . . . . . . . . . . 58 functions, transcendental . . . . . . . . . . . . . . . . . . . . . . 58 functions, window-size-change- . . . . . . . . . . . . . . . . 470 functions, write-region-annotate- . . . . . . . . . . . . . . 545 Fundamental mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 fundamental-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 fundamental-mode-abbrev-table. . . . . . . . . . . . . . . 585 future-map, electric- . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

G

g, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 garbage collection, marker . . . . . . . . . . . . . . . . . . . . 501 garbage collector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 garbage-collect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 771 gc-cons-threshold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 773 gc-hook, post- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 gc-hook, pre- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 gc-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 gc-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . . . 641, 774 generate-new-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . 443 generate-new-buffer-name. . . . . . . . . . . . . . . . . . . . . 436 generic-specifier-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 gensym, print-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 get, lax-plist- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 get, plist-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 get-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 get-buffer-create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 get-buffer-process. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 get-buffer-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 get-char-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 get-char-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 get-charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 get-coding-system . . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 get-cutbuffer, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 get-database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 get-file-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 get-largest-window . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 get-lru-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 get-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 get-range-char-table . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 get-range-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 get-register . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 get-resource, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 get-selection, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 get-text-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 get-tooltalk-message-attribute . . . . . . . . . . . . . . . . 722 getenv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 getf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 gethash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 731 global binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 global break condition . . . . . . . . . . . . . . . . . . . . . . . . 235 global keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 global mark ring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 global variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 global-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 global-abbrevs, only- . . . . . . . . . . . . . . . . . . . . . . . . . 582 global-break-condition, edebug- . . . . . . . . . . 235, 251 global-break-condition, edebug-set- . . . . . . . . . . . 235 global-key-binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 global-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 global-map, current- . . . . . . . . . . . . . . . . . . . . . . . . . . 324 global-map, use- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 global-mark, pop- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 global-mark-ring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 global-mode-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 global-popup-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 global-set-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 global-unset-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 glyph type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 glyph, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 glyph, annotation-down- . . . . . . . . . . . . . . . . . . . . . . 645 glyph, busy-pointer- . . . . . . . . . . . . . . . . . . . . . . . . . . 641 glyph, continuation- . . . . . . . . . . . . . . . . . . . . . . . . . . 642 glyph, control-arrow- . . . . . . . . . . . . . . . . . . . . . . . . . 642 glyph, extent-begin- . . . . . . . . . . . . . . . . . . . . . . . . . . 596 glyph, extent-end- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 glyph, gc-pointer- . . . . . . . . . . . . . . . . . . . . . . . . 641, 774 glyph, hscroll- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 glyph, invisible-text-. . . . . . . . . . . . . . . . . . . . . . . . . . 642 glyph, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 glyph, make-icon- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 glyph, make-pointer- . . . . . . . . . . . . . . . . . . . . . . . . . 627 glyph, menubar-pointer- . . . . . . . . . . . . . . . . . . . . . . 641 glyph, modeline-pointer- . . . . . . . . . . . . . . . . . . . . . . 641 glyph, nontext-pointer- . . . . . . . . . . . . . . . . . . . . . . . 640 glyph, octal-escape- . . . . . . . . . . . . . . . . . . . . . . . . . . 642 glyph, scrollbar-pointer- . . . . . . . . . . . . . . . . . . . . . . 641 glyph, selection-pointer- . . . . . . . . . . . . . . . . . . . . . . 641 glyph, set-annotation- . . . . . . . . . . . . . . . . . . . . . . . . 645 glyph, set-annotation-down- . . . . . . . . . . . . . . . . . . 645 glyph, set-extent-begin-. . . . . . . . . . . . . . . . . . . . . . . 597 glyph, set-extent-end- . . . . . . . . . . . . . . . . . . . . . . . . 597 glyph, text-pointer-. . . . . . . . . . . . . . . . . . . . . . . . . . . 640 glyph, toolbar-pointer-. . . . . . . . . . . . . . . . . . . . . . . . 641 glyph, truncation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 glyph-ascent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 glyph-baseline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 glyph-baseline, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 glyph-baseline-instance . . . . . . . . . . . . . . . . . . . . . . . 632 glyph-contrib-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 glyph-contrib-p, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 631 glyph-contrib-p-instance . . . . . . . . . . . . . . . . . . . . . . 631 glyph-descent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 glyph-extent, event- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 glyph-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 glyph-face, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 glyph-height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 glyph-image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 glyph-image, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 glyph-image-instance . . . . . . . . . . . . . . . . . . . . . . . . . 631 glyph-internal, make- . . . . . . . . . . . . . . . . . . . . . . . . . 627 glyph-layout, extent-begin- . . . . . . . . . . . . . . . . . . . 596 glyph-layout, extent-end- . . . . . . . . . . . . . . . . . . . . . 596 glyph-layout, set-extent-begin- . . . . . . . . . . . . . . . . 597 glyph-layout, set-extent-end-. . . . . . . . . . . . . . . . . . 597 glyph-p, buffer-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 glyph-p, event-over- . . . . . . . . . . . . . . . . . . . . . . . . . . 298 glyph-p, icon- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 glyph-p, pointer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 glyph-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 glyph-property, remove- . . . . . . . . . . . . . . . . . . . . . . 630 glyph-property, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . 628 glyph-property-instance. . . . . . . . . . . . . . . . . . . . . . . 629 glyph-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 glyph-type-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 glyph-type-p, valid-. . . . . . . . . . . . . . . . . . . . . . . . . . . 640 glyph-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 glyph-x-pixel, event- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 glyph-y-pixel, event- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 glyphp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 glyphs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 goto-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 goto-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 grab-keyboard, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 grab-pointer, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 grabbed-buffer, mouse- . . . . . . . . . . . . . . . . . . . . . . . 325 graphic, charset-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 grouping, regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556

H

h, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 hack-local-variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 handler, error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 handler, find-file-name- . . . . . . . . . . . . . . . . . . . . . . . 417 handlers, inhibit-file-name- . . . . . . . . . . . . . . . . . . . 417 handling errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 hard link, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 hard-newlines, use- . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 hash notation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 hash table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 hash table type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 hash table, weak. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 hashing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 hashing, locate-file-clear- . . . . . . . . . . . . . . . . . . . . . 199 hashing, symbol name . . . . . . . . . . . . . . . . . . . . . . . . 115 hashtable, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 hashtable, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 hashtable, make-key-weak-. . . . . . . . . . . . . . . . . . . . 668 hashtable, make-value-weak- . . . . . . . . . . . . . . . . . . 669 hashtable, make-weak- . . . . . . . . . . . . . . . . . . . . . . . . 668 hashtable-fullness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 hashtablep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 header comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 765 header comments, library . . . . . . . . . . . . . . . . . . . . . 765 height, bottom-toolbar- . . . . . . . . . . . . . . . . . . . . . . . 356 height, default-toolbar- . . . . . . . . . . . . . . . . . . . . . . . 356 height, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 height, frame-pixel- . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 height, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 height, image-instance- . . . . . . . . . . . . . . . . . . . . . . . 639 height, top-toolbar-. . . . . . . . . . . . . . . . . . . . . . . . . . . 356 height, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 height, window-displayed-text-pixel- . . . . . . . . . . 467 height, window-min- . . . . . . . . . . . . . . . . . . . . . . . . . . 469 height, window-pixel- . . . . . . . . . . . . . . . . . . . . . . . . . 466 height, window-text-area-pixel- . . . . . . . . . . . . . . . 467 height-threshold, split-. . . . . . . . . . . . . . . . . . . . . . . . 458 help for major mode . . . . . . . . . . . . . . . . . . . . . . . . . . 371 help, completion-auto- . . . . . . . . . . . . . . . . . . . . . . . . 273 help, Helper- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 help, minibuffer-completion- . . . . . . . . . . . . . . . . . . 272 help, mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 help-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 help-command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 help-command, prefix- . . . . . . . . . . . . . . . . . . . . . . . . 390 help-form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 help-form, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . . 281 help-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389, 783 help-map, Helper- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 help-return-message, print- . . . . . . . . . . . . . . . . . . . 390 Helper-describe-bindings . . . . . . . . . . . . . . . . . . . . . . 391 Helper-help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 Helper-help-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 hexadecimal, integer to . . . . . . . . . . . . . . . . . . . . . . . . 70 hide-annotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 highest-p, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 highlight-extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 600 highlight-extent, force-. . . . . . . . . . . . . . . . . . . . . . . . 600 highlight-priority, mouse- . . . . . . . . . . . . . . . . . . . . . 600 history list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 history of commands . . . . . . . . . . . . . . . . . . . . . . . . . 314 history, command . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 history, command-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 history, extended-command- . . . . . . . . . . . . . . . . . . 267 history, file-name- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 history, Info-minibuffer- . . . . . . . . . . . . . . . . . . . . . . 268 history, Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 history, load- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 history, Manual-page-minibuffer- . . . . . . . . . . . . . . 268 history, minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 history, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 history, query-replace- . . . . . . . . . . . . . . . . . . . . . . . . 267 history, read-expression- . . . . . . . . . . . . . . . . . . . . . . 267 history, record command. . . . . . . . . . . . . . . . . . . . . . 289 history, regexp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 history, shell-command-. . . . . . . . . . . . . . . . . . . . . . . 267 history-element, next- . . . . . . . . . . . . . . . . . . . . . . . . 281 history-element, next-matching-. . . . . . . . . . . . . . . 281 history-element, previous- . . . . . . . . . . . . . . . . . . . . 281 history-element, previous-matching-. . . . . . . . . . . 281 history-map, command- . . . . . . . . . . . . . . . . . . . . . . 783 history-map, electric- . . . . . . . . . . . . . . . . . . . . . . . . . 783 HOME environment variable. . . . . . . . . . . . . . . . . . 675 home-directory, user- . . . . . . . . . . . . . . . . . . . . . . . . . 704 hook, activate-menubar- . . . . . . . . . . . . . . . . . . . . . . 343 hook, activate-popup-menu- . . . . . . . . . . . . . . . . . . 347 hook, add- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 hook, after-init- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 hook, after-revert- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 hook, after-save-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 hook, auto-save- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 hook, before-init- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 hook, before-revert-. . . . . . . . . . . . . . . . . . . . . . . . . . . 431 hook, blink-paren- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 hook, change-major-mode- . . . . . . . . . . . . . . . . . . . . 365 hook, create-device- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 hook, create-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 hook, default-deselect-frame-. . . . . . . . . . . . . . . . . . 483 hook, default-select-frame- . . . . . . . . . . . . . . . . . . . . 483 hook, delete-device- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 hook, delete-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 hook, deselect-frame- . . . . . . . . . . . . . . . . . . . . . . . . . 484 hook, disabled-command- . . . . . . . . . . . . . . . . . . . . . 314 hook, edebug-setup- . . . . . . . . . . . . . . . . . . . . . . . . . . 249 hook, first-change- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 549 hook, kill-buffer-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 hook, kill-emacs- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 hook, lambda expression in . . . . . . . . . . . . . . . . . . . 380 hook, major mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 hook, make-local-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 hook, map-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 hook, menu-no-selection- . . . . . . . . . . . . . . . . . . . . . 344 hook, minibuffer-exit-. . . . . . . . . . . . . . . . . . . . . . . . . 281 hook, minibuffer-setup- . . . . . . . . . . . . . . . . . . . . . . . 281 hook, mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 hook, post-command-. . . . . . . . . . . . . . . . . . . . . . . . . 284 hook, post-gc- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 hook, pre-abbrev-expand-. . . . . . . . . . . . . . . . . . . . . 584 hook, pre-command-. . . . . . . . . . . . . . . . . . . . . . . . . . 283 hook, pre-gc- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 hook, remove- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 hook, select-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 hook, suspend- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 699 hook, suspend-resume-. . . . . . . . . . . . . . . . . . . . . . . . 699 hook, term-setup- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 hook, unmap-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . 484 hook, window-setup-. . . . . . . . . . . . . . . . . . . . . . . . . . 696 hook, zmacs-activate-region- . . . . . . . . . . . . . . . . . . 510 hook, zmacs-deactivate-region- . . . . . . . . . . . . . . . . 511 hook, zmacs-update-region- . . . . . . . . . . . . . . . . . . . 511 hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 hooks for loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 hooks for text changes . . . . . . . . . . . . . . . . . . . . . . . . 548 hooks, annotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 hooks, change . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 hooks, find-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 hooks, find-file-not-found- . . . . . . . . . . . . . . . . . . . . 395 hooks, frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 hooks, loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 hooks, local-write-file- . . . . . . . . . . . . . . . . . . . . . . . . 397 hooks, run- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 hooks, write-contents- . . . . . . . . . . . . . . . . . . . . . . . . 397 hooks, write-file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 horizontal position. . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 horizontal scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 horizontal-space, delete- . . . . . . . . . . . . . . . . . . . . . . 519 horizontally, enlarge-window- . . . . . . . . . . . . . . . . . 469 horizontally, shrink-window- . . . . . . . . . . . . . . . . . . 469 horizontally, split-window- . . . . . . . . . . . . . . . . . . . . 450 host, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 host, ldap-default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 host-address, mail- . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 host-parameters-alist, ldap- . . . . . . . . . . . . . . . . . . . 727 hotspot-x, image-instance- . . . . . . . . . . . . . . . . . . . . 639 hotspot-y, image-instance- . . . . . . . . . . . . . . . . . . . . 639 hscroll, set-window- . . . . . . . . . . . . . . . . . . . . . . . . . . 465 hscroll, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 hscroll-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642

I

icon-glyph, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 icon-glyph-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 icon-pixmap, x-set-frame- . . . . . . . . . . . . . . . . . . . . . 478 icon-title-format, frame- . . . . . . . . . . . . . . . . . . . . . . 478 iconified frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 iconified-p, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 iconify-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 id, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 id, x-window-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 identification, modeline-buffer- . . . . . . . . . . . . . . . . 377 identity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 IEEE floating point. . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 if. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 if-buffer-read-only, barf- . . . . . . . . . . . . . . . . . . . . . . 440 if-necessary, delete-auto-save-file- . . . . . . . . . . . . . 429 ignore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 ignore-case, completion- . . . . . . . . . . . . . . . . . . . . . . 270 ignore-comments, parse-sexp- . . . . . . . . . . . . . . . . . 579 ignore-elc-files, load- . . . . . . . . . . . . . . . . . . . . . . . . . . 200 ignore-errors-p, byte-recompile-directory- . . . . . 210 ignore-labels, log-message- . . . . . . . . . . . . . . . . . . . . 652 ignore-regexps, log-message- . . . . . . . . . . . . . . . . . . 652 ignored-errors, debug- . . . . . . . . . . . . . . . . . . . . . . . . 220 ignored-extensions, completion- . . . . . . . . . . . . . . . 414 ignored-local-variables . . . . . . . . . . . . . . . . . . . . . . . . 369 image instance type . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 image instance types . . . . . . . . . . . . . . . . . . . . . . . . . 637 image instances. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 image instantiator conversion . . . . . . . . . . . . . . . . . 636 image instantiators, conversion of . . . . . . . . . . . . . 636 image specifiers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 image, glyph-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 image, set-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 image-conversion-list, console-type- . . . . . . . . . . . 636 image-conversion-list, set-console-type- . . . . . . . . 636 image-instance, colorize- . . . . . . . . . . . . . . . . . . . . . . 638 image-instance, glyph- . . . . . . . . . . . . . . . . . . . . . . . . 631 image-instance, make- . . . . . . . . . . . . . . . . . . . . . . . . 638 image-instance-background . . . . . . . . . . . . . . . . . . . 639 image-instance-depth . . . . . . . . . . . . . . . . . . . . . . . . . 639 image-instance-file-name . . . . . . . . . . . . . . . . . . . . . . 639 image-instance-foreground . . . . . . . . . . . . . . . . . . . . 639 image-instance-height. . . . . . . . . . . . . . . . . . . . . . . . . 639 image-instance-hotspot-x . . . . . . . . . . . . . . . . . . . . . 639 image-instance-hotspot-y . . . . . . . . . . . . . . . . . . . . . 639 image-instance-mask-file-name . . . . . . . . . . . . . . . . 639 image-instance-name . . . . . . . . . . . . . . . . . . . . . . . . . 639 image-instance-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 image-instance-p, color-pixmap- . . . . . . . . . . . . . . 638 image-instance-p, mono-pixmap- . . . . . . . . . . . . . . 638 image-instance-p, nothing- . . . . . . . . . . . . . . . . . . . . 638 image-instance-p, pointer- . . . . . . . . . . . . . . . . . . . . 638 image-instance-p, subwindow-. . . . . . . . . . . . . . . . . 638 image-instance-p, text- . . . . . . . . . . . . . . . . . . . . . . . 637 image-instance-string . . . . . . . . . . . . . . . . . . . . . . . . . 639 image-instance-type . . . . . . . . . . . . . . . . . . . . . . . . . . 637 image-instance-type-list. . . . . . . . . . . . . . . . . . . . . . . 637 image-instance-type-p, valid-. . . . . . . . . . . . . . . . . . 637 image-instance-width . . . . . . . . . . . . . . . . . . . . . . . . . 639 image-instantiator-format-list . . . . . . . . . . . . . . . . . 636 image-instantiator-format-p, valid- . . . . . . . . . . . . 636 image-specifier, make- . . . . . . . . . . . . . . . . . . . . . . . . 633 image-specifier-p . . . . . . . . . . . . . . . . . . . . . . . . . 605, 633 implicit progn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 inc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 include-escapes, words- . . . . . . . . . . . . . . . . . . . . . . . 491 inclusive or, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 indent, fill-individual-varying-. . . . . . . . . . . . . . . . . 529 indent, newline-and- . . . . . . . . . . . . . . . . . . . . . . . . . . 537 indent, reindent-then-newline-and- . . . . . . . . . . . . 537 indent-according-to-mode . . . . . . . . . . . . . . . . . . . . . 536 indent-code-rigidly . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 indent-for-tab-command . . . . . . . . . . . . . . . . . . . . . . 537 indent-line-function. . . . . . . . . . . . . . . . . . . . . . . . . . . 536 indent-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 indent-region-function . . . . . . . . . . . . . . . . . . . . . . . . 537 indent-relative . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 indent-relative-maybe. . . . . . . . . . . . . . . . . . . . . . . . . 539 indent-rigidly. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 indent-tabs-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 indent-to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 indent-to-left-margin . . . . . . . . . . . . . . . . . . . . . . . . . 531 indentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 indentation, back-to- . . . . . . . . . . . . . . . . . . . . . . . . . 539 indentation, backward-to-. . . . . . . . . . . . . . . . . . . . . 540 indentation, current-. . . . . . . . . . . . . . . . . . . . . . . . . . 536 indentation, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . 520 indentation, forward-to- . . . . . . . . . . . . . . . . . . . . . . 540 indentation, tabs stops for . . . . . . . . . . . . . . . . . . . . 539 indenting with parentheses. . . . . . . . . . . . . . . . . . . . 578 indirect buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 indirect specifications . . . . . . . . . . . . . . . . . . . . . . . . . 246 indirect variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 indirect, variables, . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 indirect-buffer, make- . . . . . . . . . . . . . . . . . . . . . . . . . 445 indirect-function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 indirect-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 indirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 indirection, symbol function . . . . . . . . . . . . . . . . . . 125 individual-paragraphs, fill- . . . . . . . . . . . . . . . . . . . . 528 individual-varying-indent, fill-. . . . . . . . . . . . . . . . . 529 infinite loop, quitting from. . . . . . . . . . . . . . . . . . . . 220 infinite loop, stopping an . . . . . . . . . . . . . . . . . . . . . 220 infinite loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 infinite recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 infinite, loops, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 infinity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 infinity, negative . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 infinity, positive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Info-edit-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 Info-minibuffer-history. . . . . . . . . . . . . . . . . . . . . . . . 268 Info-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 information, saving window . . . . . . . . . . . . . . . . . . . 470 inherit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 inheritance, keymap . . . . . . . . . . . . . . . . . . . . . . . . . . 319 inheriting a keymap’s bindings . . . . . . . . . . . . . . . . 319 inhibit-default-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 inhibit-file-name-handlers . . . . . . . . . . . . . . . . . . . . . 417 inhibit-file-name-operation . . . . . . . . . . . . . . . . . . . . 417 inhibit-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 inhibit-read-only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 inhibit-startup-echo-area-message . . . . . . . . . . . . . 694 inhibit-startup-message . . . . . . . . . . . . . . . . . . . . . . . 694 inhibited, backup- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 424 init file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 init file, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 init, inhibit-default- . . . . . . . . . . . . . . . . . . . . . . . . . . 695 init, make-specifier-and- . . . . . . . . . . . . . . . . . . . . . . 614 init-hook, after- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 init-hook, before- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 init.el, site- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 initial-frame-plist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 initial-major-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 initial-mode, edebug- . . . . . . . . . . . . . . . . . . . . . . . . . 250 initial-redisplay-function, set-extent- . . . . . . . . . . 597 initial-toolbar-spec . . . . . . . . . . . . . . . . . . . . . . . . . . . 358 initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 initialization, terminal-specific . . . . . . . . . . . . . . . . 695 inline functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 innermost containing parentheses . . . . . . . . . . . . . 578 input events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 input events, translating . . . . . . . . . . . . . . . . . . . . . . 709 input focus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 input modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 input modes, terminal . . . . . . . . . . . . . . . . . . . . . . . . 708 input stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 input stream, buffer . . . . . . . . . . . . . . . . . . . . . . . . . . 253 input stream, function . . . . . . . . . . . . . . . . . . . . . . . . 254 input stream, marker . . . . . . . . . . . . . . . . . . . . . . . . . 253 input stream, nil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 input stream, string . . . . . . . . . . . . . . . . . . . . . . . . . . 254 input stream, t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 input, binary-process- . . . . . . . . . . . . . . . . . . . . . . . . 679 input, console-disable- . . . . . . . . . . . . . . . . . . . . . . . . 488 input, console-enable-. . . . . . . . . . . . . . . . . . . . . . . . . 488 input, discard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 input, discard-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 input, flush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 input, key sequence. . . . . . . . . . . . . . . . . . . . . . . . . . . 304 input, minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 input, next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 input, octal character. . . . . . . . . . . . . . . . . . . . . . . . . 306 input, peeking at . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 input, process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 input, quoted character . . . . . . . . . . . . . . . . . . . . . . . 306 input, read-no-blanks- . . . . . . . . . . . . . . . . . . . . . . . . 265 input, standard- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 input, terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 input, waiting for command key . . . . . . . . . . . . . . 307 input-char, last- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 input-event, last- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 input-mode, current- . . . . . . . . . . . . . . . . . . . . . . . . . 708 input-mode, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 input-p, waiting-for-user- . . . . . . . . . . . . . . . . . . . . . 690 input-pending-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 insert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 insert suppression, quoted-. . . . . . . . . . . . . . . . . . . . 332 insert-abbrev-table-description . . . . . . . . . . . . . . . . 582 insert-and-exit, self- . . . . . . . . . . . . . . . . . . . . . . . . . . 280 insert-before-markers . . . . . . . . . . . . . . . . . . . . . . . . . 516 insert-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 insert-buffer-substring . . . . . . . . . . . . . . . . . . . . . . . . 516 insert-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 insert-command override, self- . . . . . . . . . . . . . . . . 332 insert-command, minor modes, self- . . . . . . . . . . . 374 insert-command, self- . . . . . . . . . . . . . . . . . . . . . . . . . 517 insert-default-directory . . . . . . . . . . . . . . . . . . . . . . . 275 insert-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 415 insert-directory-program . . . . . . . . . . . . . . . . . . . . . . 415 insert-extent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 insert-file, format- . . . . . . . . . . . . . . . . . . . . . . . 420, 421 insert-file-contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 insert-file-contents-function, revert-buffer- . . . . . 431 insert-file-functions, after- . . . . . . . . . . . . . . . . . . . . 546 insert-register . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 insert-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 inserting killed text. . . . . . . . . . . . . . . . . . . . . . . . . . . 523 insertion before point . . . . . . . . . . . . . . . . . . . . . . . . . 516 insertion of text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 insertion, before point,. . . . . . . . . . . . . . . . . . . . . . . . 516 insertion, character . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 insertion, self- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 insertion, text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 inside comment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 inside margin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 inside string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 inst-list (in a specifier). . . . . . . . . . . . . . . . . . . . . . . . 602 inst-list, canonicalize-. . . . . . . . . . . . . . . . . . . . . . . . . 608 inst-list, check-valid-. . . . . . . . . . . . . . . . . . . . . . . . . . 615 inst-list, specifier,. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 inst-list, specifier-instance-from- . . . . . . . . . . . . . . 612 inst-list-p, valid- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 inst-pair (in a specifier) . . . . . . . . . . . . . . . . . . . . . . . 602 inst-pair, canonicalize- . . . . . . . . . . . . . . . . . . . . . . . . 608 inst-pair, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 installation-directory . . . . . . . . . . . . . . . . . . . . . . . . . 702 instance (in a specifier) . . . . . . . . . . . . . . . . . . . . . . . 602 instance characteristics, font . . . . . . . . . . . . . . . . . . 624 instance name, font. . . . . . . . . . . . . . . . . . . . . . . . . . . 624 instance size, font . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 instance type, color . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 instance type, font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 instance type, image . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 instance types, image . . . . . . . . . . . . . . . . . . . . . . . . . 637 instance, colorize-image- . . . . . . . . . . . . . . . . . . . . . . 638 instance, face-background- . . . . . . . . . . . . . . . . . . . . 622 instance, face-background-pixmap- . . . . . . . . . . . . 623 instance, face-font- . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 instance, face-foreground-. . . . . . . . . . . . . . . . . . . . . 622 instance, face-property- . . . . . . . . . . . . . . . . . . . . . . . 621 instance, glyph-baseline- . . . . . . . . . . . . . . . . . . . . . . 632 instance, glyph-contrib-p-. . . . . . . . . . . . . . . . . . . . . 631 instance, glyph-image- . . . . . . . . . . . . . . . . . . . . . . . . 631 instance, glyph-property- . . . . . . . . . . . . . . . . . . . . . 629 instance, make-font- . . . . . . . . . . . . . . . . . . . . . . . . . . 623 instance, make-image- . . . . . . . . . . . . . . . . . . . . . . . . 638 instance, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 instance, specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 instance-background, image- . . . . . . . . . . . . . . . . . . 639 instance-depth, image-. . . . . . . . . . . . . . . . . . . . . . . . 639 instance-file-name, image- . . . . . . . . . . . . . . . . . . . . 639 instance-foreground, image- . . . . . . . . . . . . . . . . . . . 639 instance-from-inst-list, specifier- . . . . . . . . . . . . . . 612 instance-height, image- . . . . . . . . . . . . . . . . . . . . . . . 639 instance-hotspot-x, image- . . . . . . . . . . . . . . . . . . . . 639 instance-hotspot-y, image- . . . . . . . . . . . . . . . . . . . . 639 instance-mask-file-name, image-. . . . . . . . . . . . . . . 639 instance-name, color- . . . . . . . . . . . . . . . . . . . . . . . . . 626 instance-name, font- . . . . . . . . . . . . . . . . . . . . . . . . . . 624 instance-name, image- . . . . . . . . . . . . . . . . . . . . . . . . 639 instance-p, color- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626 instance-p, color-pixmap-image- . . . . . . . . . . . . . . 638 instance-p, font- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 instance-p, image- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 instance-p, mono-pixmap-image- . . . . . . . . . . . . . . 638 instance-p, nothing-image- . . . . . . . . . . . . . . . . . . . . 638 instance-p, pointer-image- . . . . . . . . . . . . . . . . . . . . 638 instance-p, subwindow-image-. . . . . . . . . . . . . . . . . 638 instance-p, text-image- . . . . . . . . . . . . . . . . . . . . . . . 637 instance-properties, font- . . . . . . . . . . . . . . . . . . . . . 624 instance-rgb-components, color-. . . . . . . . . . . . . . . 626 instance-string, image-. . . . . . . . . . . . . . . . . . . . . . . . 639 instance-truename, font- . . . . . . . . . . . . . . . . . . . . . . 624 instance-type, image- . . . . . . . . . . . . . . . . . . . . . . . . . 637 instance-type-list, image- . . . . . . . . . . . . . . . . . . . . . 637 instance-type-p, valid-image-. . . . . . . . . . . . . . . . . . 637 instance-width, image-. . . . . . . . . . . . . . . . . . . . . . . . 639 instances, characteristics of font . . . . . . . . . . . . . . 624 instances, color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 instances, image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 instancing (in a specifier) . . . . . . . . . . . . . . . . . . . . . 602 instancing, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . 602 instantiator (in a specifier). . . . . . . . . . . . . . . . . . . . 602 instantiator conversion, image . . . . . . . . . . . . . . . . 636 instantiator, check-valid-. . . . . . . . . . . . . . . . . . . . . . 614 instantiator, specifier,. . . . . . . . . . . . . . . . . . . . . . . . . 602 instantiator-format-list, image-. . . . . . . . . . . . . . . . 636 instantiator-format-p, valid-image- . . . . . . . . . . . . 636 instantiator-p, valid-. . . . . . . . . . . . . . . . . . . . . . . . . . 614 instantiators, conversion of image . . . . . . . . . . . . . 636 instructions, compiled-function- . . . . . . . . . . . . . . . 213 int confoundance disease, char- . . . . . . . . . . . . . . . . 20 int, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 int, string-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 int-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 int-p, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 int-p, char-or-char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 int-to-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 integer to decimal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 integer to hexadecimal . . . . . . . . . . . . . . . . . . . . . . . . . 70 integer to octal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 integer to string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 integer-char-or-marker-p . . . . . . . . . . . . . . . . . . . . . . 503 integer-or-char-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 integer-or-marker-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 integer-specifier-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 integerp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 integers vrs eq, CL note— . . . . . . . . . . . . . . . . . . . . . 49 interaction-mode-map, lisp- . . . . . . . . . . . . . . . . . . . 784 interactive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 interactive arguments, reading . . . . . . . . . . . . . . . . 286 interactive call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 interactive code description . . . . . . . . . . . . . . . . . . . 285 interactive codes, characters for . . . . . . . . . . . . . . . 285 interactive codes, description for . . . . . . . . . . . . . . 285 interactive commands (Edebug). . . . . . . . . . . . . . . 231 interactive completion . . . . . . . . . . . . . . . . . . . . . . . . 286 interactive function . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 interactive, * in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 interactive, @ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 interactive, in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 interactive, compiled-function- . . . . . . . . . . . . . . . . 213 interactive, description of, codes,. . . . . . . . . . . . . . 285 interactive, examples of using . . . . . . . . . . . . . . . . . 288 interactive, function- . . . . . . . . . . . . . . . . . . . . . . . . . 285 interactive, functions, making them . . . . . . . . . . . 284 interactive, read-only buffers in . . . . . . . . . . . . . . . 285 interactive-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 interactively, call- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 intern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 intern-soft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 internal, describe-bindings- . . . . . . . . . . . . . . . . . . . 336 internal, ldap-search- . . . . . . . . . . . . . . . . . . . . . . . . . 730 internal, make-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . 627 internal, where-is- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 internal-doc-file-name. . . . . . . . . . . . . . . . . . . . . . . . . 386 internals, syntax table . . . . . . . . . . . . . . . . . . . . . . . . 580 interning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 interpreter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 interpreter, byte-code. . . . . . . . . . . . . . . . . . . . . . . . . 210 interpreter-mode-alist. . . . . . . . . . . . . . . . . . . . . . . . . 371 interprogram-cut-function . . . . . . . . . . . . . . . . . . . . 524 interprogram-paste-function . . . . . . . . . . . . . . . . . . 524 interrupt-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 interval, auto-save- . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 invalid function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 invalid prefix key error. . . . . . . . . . . . . . . . . . . . . . . . 330 invalid-function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 invalid-read-syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 invalid-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 invert-face. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 invisibility-spec, buffer- . . . . . . . . . . . . . . . . . . . . . . . 655 invisible frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 invisible text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 invisible, make-frame- . . . . . . . . . . . . . . . . . . . . . . . . 482 invisible-text-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 invocation, function . . . . . . . . . . . . . . . . . . . . . . . . . . 171 invocation-directory . . . . . . . . . . . . . . . . . . . . . . . . . . 702 invocation-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 is-internal, where- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 isearch-map, minibuffer-local- . . . . . . . . . . . . . . . . . 784 isearch-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 ISO Latin 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 ISO Latin-1 characters (input) . . . . . . . . . . . . . . . . 710 iso-syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 iso-transl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 italic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 italic, x-make-font- . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 italic, x-make-font-bold- . . . . . . . . . . . . . . . . . . . . . . 625 item, add-menu- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 item, delete-menu-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 item, disable-menu-. . . . . . . . . . . . . . . . . . . . . . . . . . . 345 item, enable-menu- . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 item, find-menu-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 item, relabel-menu- . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 iteration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 itimer-edit-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784

J

jis-char, decode-shift- . . . . . . . . . . . . . . . . . . . . . . . . . 751 jis-char, encode-shift- . . . . . . . . . . . . . . . . . . . . . . . . . 751 joining lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 just-one-space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 justification, current- . . . . . . . . . . . . . . . . . . . . . . . . . 529 justification, default- . . . . . . . . . . . . . . . . . . . . . . . . . 529 justify-current-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . 529

K

kbd-macro, defining-. . . . . . . . . . . . . . . . . . . . . . . . . . 316 kbd-macro, edebug-continue- . . . . . . . . . . . . . . . . . 251 kbd-macro, execute- . . . . . . . . . . . . . . . . . . . . . . . . . . 315 kbd-macro, last- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 kept-new-versions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 kept-old-versions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 kept-versions, dired- . . . . . . . . . . . . . . . . . . . . . . . . . . 426 key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 key binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 key bindings, changing. . . . . . . . . . . . . . . . . . . . . . . . 329 key error, invalid prefix . . . . . . . . . . . . . . . . . . . . . . . 330 key input, waiting for command . . . . . . . . . . . . . . 307 key lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 key sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 key sequence error . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 key sequence input . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 key sequence, upper case . . . . . . . . . . . . . . . . . . . . . 304 key sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 key translation function . . . . . . . . . . . . . . . . . . . . . . 710 key, binding of a. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 key, complete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 key, define- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 key, downcasing in lookup- . . . . . . . . . . . . . . . . . . . 304 key, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 key, global-set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 key, global-unset- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 key, local-set-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 key, local-unset- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 key, lookup- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327 key, prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 key, preventing prefix . . . . . . . . . . . . . . . . . . . . . . . . . 327 key, undefined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 key-binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 key-binding, global- . . . . . . . . . . . . . . . . . . . . . . . . . . 328 key-binding, local- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 key-binding, minor-mode-. . . . . . . . . . . . . . . . . . . . . 328 key-definition, substitute- . . . . . . . . . . . . . . . . . . . . . 331 key-description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 key-description, single- . . . . . . . . . . . . . . . . . . . . . . . 388 key-map, function- . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 key-press-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 key-sequence, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 key-specifier-p, event-matches- . . . . . . . . . . . . . . . . 321 key-translation-map . . . . . . . . . . . . . . . . . . . . . . . . . . 710 key-weak-hashtable, make-. . . . . . . . . . . . . . . . . . . . 668 keybindings, menubar-show- . . . . . . . . . . . . . . . . . . 343 keyboard macro execution . . . . . . . . . . . . . . . . . . . . 289 keyboard macro termination . . . . . . . . . . . . . . . . . . 663 keyboard macro, terminate . . . . . . . . . . . . . . . . . . . 307 keyboard macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 keyboard macros (Edebug) . . . . . . . . . . . . . . . . . . . 233 keyboard menu accelerators. . . . . . . . . . . . . . . . . . . 347 keyboard, x-grab- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 keyboard, x-ungrab- . . . . . . . . . . . . . . . . . . . . . . . . . . 718 keyboard-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 keymap entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keymap in keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap inheritance . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 keymap parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 keymap’s bindings, inheriting a . . . . . . . . . . . . . . . 319 keymap, active . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 keymap, command in . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keymap, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 keymap, global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 keymap, keymap in . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap, lambda in . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap, list in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap, local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 keymap, major mode . . . . . . . . . . . . . . . . . . . . . . . . . 323 keymap, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keymap, make-sparse- . . . . . . . . . . . . . . . . . . . . . . . . 318 keymap, map- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 keymap, nil in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap, parent of a . . . . . . . . . . . . . . . . . . . . . . . . . . 319 keymap, set-extent-. . . . . . . . . . . . . . . . . . . . . . . . . . . 597 keymap, string in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 keymap, suppress- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 keymap, symbol in . . . . . . . . . . . . . . . . . . . . . . . . . . . 327 keymap, undefined in . . . . . . . . . . . . . . . . . . . . . . . . . 327 keymap-default-binding. . . . . . . . . . . . . . . . . . . . . . . 319 keymap-default-binding, set- . . . . . . . . . . . . . . . . . . 319 keymap-fullness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 keymap-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keymap-name, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keymap-parents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 keymap-parents, set-. . . . . . . . . . . . . . . . . . . . . . . . . . 319 keymap-prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 keymap-prompt, set- . . . . . . . . . . . . . . . . . . . . . . . . . 337 keymapp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keymaps in modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 keymaps, accessible- . . . . . . . . . . . . . . . . . . . . . . . . . . 334 keymaps, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keymaps, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 keymaps, format of . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 keys in documentation strings . . . . . . . . . . . . . . . . 386 keys in documentation, substituting. . . . . . . . . . . 386 keys in, documentation, . . . . . . . . . . . . . . . . . . . . . . 386 keys, events-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 keys, recent-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 keys, substitute-command-. . . . . . . . . . . . . . . . . . . . 387 keys, this-command- . . . . . . . . . . . . . . . . . . . . . . . . . . 291 keys, unbinding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 keys-ring-size, recent- . . . . . . . . . . . . . . . . . . . . . . . . . 711 keys-ring-size, set-recent- . . . . . . . . . . . . . . . . . . . . . 711 keystroke . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 keystroke command . . . . . . . . . . . . . . . . . . . . . . . . . . 166 keystrokes, echo-. . . . . . . . . . . . . . . . . . . . . . . . . 292, 653 keysym-name-p, x-valid- . . . . . . . . . . . . . . . . . . . . . . 719 keywordp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 keywordp, lambda-list- . . . . . . . . . . . . . . . . . . . . . . . 246 kill command repetition . . . . . . . . . . . . . . . . . . . . . . 291 kill ring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 kill, copy-region-as- . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 kill, current-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 kill-all-local-variables . . . . . . . . . . . . . . . . . . . . . . . . . 161 kill-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 kill-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 kill-buffer-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 444 kill-buffer-query-functions . . . . . . . . . . . . . . . . . . . . 444 kill-emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 kill-emacs-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 kill-emacs-query-functions . . . . . . . . . . . . . . . . . . . . 698 kill-local-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 kill-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 kill-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 kill-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 kill-ring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 kill-ring-max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 kill-ring-yank-pointer . . . . . . . . . . . . . . . . . . . . . . . . . 525 kill-without-query, process- . . . . . . . . . . . . . . . . . . . 681 kill-without-query-p, process- . . . . . . . . . . . . . . . . . 683 killed text, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . 523 killing buffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 killing XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 killing, buffers, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 killing, deletion vs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518

L

labels, log-message-ignore- . . . . . . . . . . . . . . . . . . . . 652 lack union, set, CL note— . . . . . . . . . . . . . . . . . . . . . 92 lambda expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 lambda expression in hook . . . . . . . . . . . . . . . . . . . . 380 lambda expressions (Edebug), anonymous. . . . . 231 lambda in debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 lambda in keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 lambda list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 lambda-list (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . 245 lambda-list-keywordp . . . . . . . . . . . . . . . . . . . . . . . . . 246 larger-font, x-find- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 largest-window, get- . . . . . . . . . . . . . . . . . . . . . . . . . . 453 last-abbrev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 last-abbrev-location . . . . . . . . . . . . . . . . . . . . . . . . . . 584 last-abbrev-text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 last-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 last-command-char . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 last-command-event . . . . . . . . . . . . . . . . . . . . . . . . . . 291 last-error, database- . . . . . . . . . . . . . . . . . . . . . . . . . . 674 last-input-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 last-input-event . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 last-kbd-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 Latin 1, ISO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Latin-1 character set (input). . . . . . . . . . . . . . . . . . 710 Latin-1 characters (input), ISO . . . . . . . . . . . . . . . 710 lax-plist, canonicalize- . . . . . . . . . . . . . . . . . . . . . . . . 100 lax-plist-get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 lax-plist-member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 lax-plist-put . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 lax-plist-remprop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 lax-plists-eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 lax-plists-equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 layout policy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 layout types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 layout, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 layout, extent-begin-glyph- . . . . . . . . . . . . . . . . . . . 596 layout, extent-end-glyph- . . . . . . . . . . . . . . . . . . . . . 596 layout, screen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 layout, set-annotation-. . . . . . . . . . . . . . . . . . . . . . . . 645 layout, set-extent-begin-glyph- . . . . . . . . . . . . . . . . 597 layout, set-extent-end-glyph-. . . . . . . . . . . . . . . . . . 597 lazy loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 LDAP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 ldap-close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 ldap-default-base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 ldap-default-host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 ldap-default-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 ldap-host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 ldap-host-parameters-alist . . . . . . . . . . . . . . . . . . . . 727 ldap-live-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 ldap-open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 ldap-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728 ldap-search-internal. . . . . . . . . . . . . . . . . . . . . . . . . . . 730 ldapp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 left, scroll- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 left-margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 left-margin, current- . . . . . . . . . . . . . . . . . . . . . . . . . . 530 left-margin, delete-to- . . . . . . . . . . . . . . . . . . . . . . . . 531 left-margin, indent-to- . . . . . . . . . . . . . . . . . . . . . . . . 531 left-margin, move-to- . . . . . . . . . . . . . . . . . . . . . . . . . 531 left-margin, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 left-margin-pixel-width, window- . . . . . . . . . . . . . . 647 left-margin-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 left-overflow, use- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 left-toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 left-toolbar-visible-p . . . . . . . . . . . . . . . . . . . . . . . . . . 357 left-toolbar-width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 length extent, zero- . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 length, bit vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 length, edebug-print- . . . . . . . . . . . . . . . . . . . . 239, 251 length, extent-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 length, list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 length, maximum when printing, string . . . . . . . 261 length, print- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 length, print-string- . . . . . . . . . . . . . . . . . . . . . . . . . . 261 length, sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 length, string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 length, vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 lessp, string-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 let* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 let-specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 letters, CL note—case of. . . . . . . . . . . . . . . . . . . . . . . 23 level form, top- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 level, display-warning-minimum- . . . . . . . . . . . . . . 654 level, edebug-print- . . . . . . . . . . . . . . . . . . . . . . 239, 251 level, log-warning-minimum- . . . . . . . . . . . . . . . . . . 654 level, print-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 level, recursive editing . . . . . . . . . . . . . . . . . . . . . . . . 312 level, top- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 level-form, edebug-eval-top-. . . . . . . . . . . . . . . . . . . 231 lexical binding (Edebug) . . . . . . . . . . . . . . . . . . . . . . 237 lexical comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 library compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 library header comments. . . . . . . . . . . . . . . . . . . . . . 765 library, Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 library-search-path, x- . . . . . . . . . . . . . . . . . . . . . . . . 719 limit error, variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 limit, memory- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 773 limit, undo- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 limit, undo-strong- . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 limits, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 line arguments, command. . . . . . . . . . . . . . . . . . . . . 696 line in regexp, beginning of . . . . . . . . . . . . . . . . . . . 555 line options, command. . . . . . . . . . . . . . . . . . . . . . . . 696 line wrapping. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 line, beginning of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 line, beginning-of- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 line, command- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 line, end-of- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 line, format-buffers-menu- . . . . . . . . . . . . . . . . . . . . 350 line, forward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 line, goto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 line, justify-current- . . . . . . . . . . . . . . . . . . . . . . . . . . 529 line, move-to-window- . . . . . . . . . . . . . . . . . . . . . . . . 494 line, options on command. . . . . . . . . . . . . . . . . . . . . 696 line, split- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 line, switches on command. . . . . . . . . . . . . . . . . . . . 696 line, window top. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 line-args, command- . . . . . . . . . . . . . . . . . . . . . . . . . . 697 line-function, indent- . . . . . . . . . . . . . . . . . . . . . . . . . 536 line-functions, command- . . . . . . . . . . . . . . . . . . . . . 697 line-processed, command- . . . . . . . . . . . . . . . . . . . . . 696 lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 492 lines in region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 lines, continuation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 lines, count- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 lines, default-truncate-. . . . . . . . . . . . . . . . . . . . . . . . 650 lines, delete-blank- . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 lines, display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 lines, next-screen-context- . . . . . . . . . . . . . . . . . . . . 464 lines, sort-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 lines, truncate- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 link, file hard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 link, make-symbolic-. . . . . . . . . . . . . . . . . . . . . . . . . . 407 linked, backup-by-copying-when-. . . . . . . . . . . . . . 425 linking files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 links, file symbolic . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 Lisp (Edebug), Common. . . . . . . . . . . . . . . . . . . . . . 231 Lisp debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 Lisp expression motion . . . . . . . . . . . . . . . . . . . . . . . 495 Lisp history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Lisp library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 Lisp nesting error . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 Lisp object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Lisp printer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Lisp reader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 Lisp, Common. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 lisp-eval-depth, max- . . . . . . . . . . . . . . . . . . . . . . . . . 123 lisp-interaction-mode-map . . . . . . . . . . . . . . . . . . . . 784 lisp-mode-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . 585 lisp-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 lisp-mode-map, emacs- . . . . . . . . . . . . . . . . . . . . . . . 783 lisp-mode-map, shared- . . . . . . . . . . . . . . . . . . . . . . . 785 lisp-mode-syntax-table, emacs-. . . . . . . . . . . . . . . . 580 lisp-mode.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 366 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 84 list (Edebug), evaluation. . . . . . . . . . . . . . . . . . . . . . 238 list (Edebug), lambda-. . . . . . . . . . . . . . . . . . . . . . . . 245 list (in a specifier), inst- . . . . . . . . . . . . . . . . . . . . . . 602 list cell (symbol), property. . . . . . . . . . . . . . . . . . . . 113 list elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 list form evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 list in keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 list length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 list motion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 list structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 list type, weak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 list), element (of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 list, abbrev-table-name- . . . . . . . . . . . . . . . . . . . . . . 582 list, add-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 list, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 list, association . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 list, backward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 list, buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 list, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 list, buffer-undo- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 list, canonicalize-inst-. . . . . . . . . . . . . . . . . . . . . . . . . 608 list, canonicalize-spec- . . . . . . . . . . . . . . . . . . . . . . . . 608 list, char-table-type- . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 list, charset- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 list, check-valid-inst-. . . . . . . . . . . . . . . . . . . . . . . . . . 615 list, check-valid-spec- . . . . . . . . . . . . . . . . . . . . . . . . . 615 list, coding-category- . . . . . . . . . . . . . . . . . . . . . . . . . 750 list, coding-priority- . . . . . . . . . . . . . . . . . . . . . . . . . . 750 list, coding-system- . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 list, console- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 list, console-device- . . . . . . . . . . . . . . . . . . . . . . . . . . . 485 list, console-type-image-conversion- . . . . . . . . . . . 636 list, debugger command . . . . . . . . . . . . . . . . . . . . . . 223 list, device- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 list, device-frame-. . . . . . . . . . . . . . . . . . . . . . . . 479, 486 list, device-matching-specifier-tag- . . . . . . . . . . . . 611 list, display-completion- . . . . . . . . . . . . . . . . . . . . . . 273 list, down-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 list, Edebug specification . . . . . . . . . . . . . . . . . . . . . 243 list, empty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 list, extent-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 list, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 list, forward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 list, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 list, free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 list, glyph-type- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 list, history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 list, image-instance-type- . . . . . . . . . . . . . . . . . . . . . 637 list, image-instantiator-format-. . . . . . . . . . . . . . . . 636 list, lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 list, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 list, make-weak- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 list, membership in a . . . . . . . . . . . . . . . . . . . . . . . . . . 92 list, process-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 list, property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 list, reversing a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 list, set-coding-priority- . . . . . . . . . . . . . . . . . . . . . . . 750 list, set-console-type-image-conversion- . . . . . . . . 636 list, set-weak-list-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 list, specifier, inst- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 list, specifier-instance-from-inst- . . . . . . . . . . . . . . 612 list, specifier-spec- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 list, specifier-tag- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 list, symbol, property . . . . . . . . . . . . . . . . . . . . . . . . . 118 list, tab-stop-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 539 list, toolbar-make-button- . . . . . . . . . . . . . . . . . . . . 354 list, up- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 list, visible-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 list, weak. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 list, weak-list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 list-buffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442 list-buffers-directory . . . . . . . . . . . . . . . . . . . . . . . . . . 438 list-file-name, auto-save- . . . . . . . . . . . . . . . . . . . . . . 430 list-fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 list-keywordp, lambda- . . . . . . . . . . . . . . . . . . . . . . . 246 list-list, set-weak-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 list-list, weak- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 list-p, valid-inst- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 list-p, valid-spec- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 list-p, weak- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 list-processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 list-to-specifier, add-spec- . . . . . . . . . . . . . . . . . . . . . 606 list-type, weak- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 listp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 lists (Edebug), dotted . . . . . . . . . . . . . . . . . . . . . . . . 246 lists and cons cells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 lists as sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 lists represented as boxes . . . . . . . . . . . . . . . . . . . . . . 79 lists vs association lists, property . . . . . . . . . . . . . 118 lists, . in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 lists, box diagrams, for . . . . . . . . . . . . . . . . . . . . . . . . 25 lists, box representation for . . . . . . . . . . . . . . . . . . . . 79 lists, building . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 lists, concatenating . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 lists, copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 lists, diagrams, boxed, for. . . . . . . . . . . . . . . . . . . . . . 25 lists, joining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 lists, modification of . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 lists, nil and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 lists, nil in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 lists, property lists vs association . . . . . . . . . . . . . 118 lists, rearrangement of . . . . . . . . . . . . . . . . . . . . . . . . . 89 lists, scan-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 lists, sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 literal evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 live-p, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 live-p, console- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 live-p, database- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 live-p, device- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 live-p, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 live-p, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 live-p, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 live-p, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 live-p, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 lmessage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 ln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 load error with require. . . . . . . . . . . . . . . . . . . . . . . . 203 load errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 load-alist, after- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 load-average . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 load-default-sounds . . . . . . . . . . . . . . . . . . . . . . . . . . . 664 load-history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 load-ignore-elc-files . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 load-in-progress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 load-path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 load-read-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 load-sound-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 664 load-warn-when-source-newer . . . . . . . . . . . . . . . . . 200 load-warn-when-source-only. . . . . . . . . . . . . . . . . . . 200 load.el, site- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 loading hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 loading of documentation, dynamic . . . . . . . . . . . 210 loading of functions, dynamic . . . . . . . . . . . . . . . . . 211 loading, hooks for . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 loading, lazy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 loading, mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 loading, repeated . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 loadup.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 767 local binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 local keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 local variable, permanent . . . . . . . . . . . . . . . . . . . . . 161 local variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 local variables in modes, buffer-. . . . . . . . . . . . . . . 364 local variables, binding . . . . . . . . . . . . . . . . . . . . . . . 148 local variables, buffer- . . . . . . . . . . . . . . . . . . . . . . . . 159 local, automatically buffer- . . . . . . . . . . . . . . . . . . . 159 local, CL note—set . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 local, make-variable-buffer- . . . . . . . . . . . . . . . . . . . 160 local, variables, buffer-. . . . . . . . . . . . . . . . . . . . . . . . 159 local-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 local-completion-map, minibuffer- . . . . . . . . 271, 784 local-copy, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 417 local-eval, enable- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 local-hook, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 local-isearch-map, minibuffer- . . . . . . . . . . . . . . . . . 784 local-key-binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 local-map, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 local-map, minibuffer- . . . . . . . . . . . . . . . . . . . 264, 784 local-map, overriding-. . . . . . . . . . . . . . . . . . . . 325, 784 local-map, overriding-terminal- . . . . . . . . . . . . . . . 325 local-map, use- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 local-must-match-map, minibuffer- . . . . . . . 272, 784 local-ns-map, minibuffer- . . . . . . . . . . . . . . . . . . . . . 265 local-set-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 local-unset-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 local-variable, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 local-variable, make-. . . . . . . . . . . . . . . . . . . . . . . . . . 159 local-variable-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 local-variables, buffer- . . . . . . . . . . . . . . . . . . . . . . . . 160 local-variables, enable-. . . . . . . . . . . . . . . . . . . . . . . . 369 local-variables, hack- . . . . . . . . . . . . . . . . . . . . . . . . . 371 local-variables, ignored-. . . . . . . . . . . . . . . . . . . . . . . 369 local-variables, kill-all- . . . . . . . . . . . . . . . . . . . . . . . . 161 local-write-file-hooks. . . . . . . . . . . . . . . . . . . . . . . . . . 397 locale (in a specifier) . . . . . . . . . . . . . . . . . . . . . . . . . 602 locale, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 locale, specifier-locale-type-from-. . . . . . . . . . . . . . 616 locale-p, valid-specifier- . . . . . . . . . . . . . . . . . . . . . . . 614 locale-type-from-locale, specifier-. . . . . . . . . . . . . . 616 locale-type-p, valid-specifier- . . . . . . . . . . . . . . . . . . 614 locate-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 locate-file-clear-hashing . . . . . . . . . . . . . . . . . . . . . . . 199 locating, extents, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 location, abbrev-start- . . . . . . . . . . . . . . . . . . . . . . . . 584 location, last-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . 584 location-buffer, abbrev-start- . . . . . . . . . . . . . . . . . 584 lock, ask-user-about- . . . . . . . . . . . . . . . . . . . . . . . . . 400 lock-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 locked, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 locked-p, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 locks, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 log-message-ignore-labels . . . . . . . . . . . . . . . . . . . . . 652 log-message-ignore-regexps. . . . . . . . . . . . . . . . . . . . 652 log-message-max-size . . . . . . . . . . . . . . . . . . . . . . . . . 652 log-warning-minimum-level . . . . . . . . . . . . . . . . . . . 654 log-warning-suppressed-classes . . . . . . . . . . . . . . . . 654 log10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 logand. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 logb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 logical and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 logical exclusive or. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 logical inclusive or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 logical not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 logical shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 logical-name, define- . . . . . . . . . . . . . . . . . . . . . . . . . . 407 login, ftp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 login-name, user- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 login-name, user-real-. . . . . . . . . . . . . . . . . . . . . . . . . 703 logior. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 lognot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 logxor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 looking-at . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 looking-at, posix- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 lookup, key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 lookup-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327 lookup-key, downcasing in . . . . . . . . . . . . . . . . . . . . 304 loop, command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 loop, count- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 loop, editor command . . . . . . . . . . . . . . . . . . . . . . . . 283 loop, quitting from infinite. . . . . . . . . . . . . . . . . . . . 220 loop, recursive command . . . . . . . . . . . . . . . . . . . . . 312 loop, recursive, command . . . . . . . . . . . . . . . . . . . . . 312 loop, stopping an infinite . . . . . . . . . . . . . . . . . . . . . 220 loops, infinite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 lower case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 lower-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 lower-frame, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 lowering a frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 lowest-p, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 lru-window, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 lsh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 lwarn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 654

M

M-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 M-x, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 Maclisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 macro argument evaluation . . . . . . . . . . . . . . . . . . . 183 macro call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 macro call evaluation . . . . . . . . . . . . . . . . . . . . . . . . . 126 macro compilation. . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 macro descriptions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 macro execution, keyboard . . . . . . . . . . . . . . . . . . . 289 macro expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 macro termination, keyboard . . . . . . . . . . . . . . . . . 663 macro, defining-kbd-. . . . . . . . . . . . . . . . . . . . . . . . . . 316 macro, edebug-continue-kbd- . . . . . . . . . . . . . . . . . 251 macro, execute-kbd- . . . . . . . . . . . . . . . . . . . . . . . . . . 315 macro, executing- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 macro, last-kbd- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 macro, terminate keyboard . . . . . . . . . . . . . . . . . . . 307 macroexpand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 macros (Edebug), keyboard . . . . . . . . . . . . . . . . . . . 233 macros, byte-compiling . . . . . . . . . . . . . . . . . . . . . . . 180 macros, expansion of . . . . . . . . . . . . . . . . . . . . . . . . . 179 macros, keyboard. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 magic file names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 mail-address, user- . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 mail-host-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 major mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 major mode hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 major mode keymap . . . . . . . . . . . . . . . . . . . . . . . . . . 323 major mode, documentation for. . . . . . . . . . . . . . . 371 major mode, help for . . . . . . . . . . . . . . . . . . . . . . . . . 371 major-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 major-mode, default- . . . . . . . . . . . . . . . . . . . . . . . . . 370 major-mode, initial- . . . . . . . . . . . . . . . . . . . . . . . . . . 370 major-mode, set-buffer- . . . . . . . . . . . . . . . . . . . . . . . 370 major-mode-hook, change- . . . . . . . . . . . . . . . . . . . . 365 major-version, emacs- . . . . . . . . . . . . . . . . . . . . . . . . 769 make them, backup files, how to . . . . . . . . . . . . . . 424 make-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 make-annotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 make-auto-save-file-name . . . . . . . . . . . . . . . . . . . . . 428 make-backup-file-name . . . . . . . . . . . . . . . . . . . . . . . 426 make-backup-files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 make-bit-vector. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 make-button-list, toolbar- . . . . . . . . . . . . . . . . . . . . 354 make-byte-code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 make-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 make-char-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 make-charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 739 make-coding-system . . . . . . . . . . . . . . . . . . . . . . . . . . 749 make-composite-char . . . . . . . . . . . . . . . . . . . . . . . . . 742 make-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 make-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 make-display-table. . . . . . . . . . . . . . . . . . . . . . . . . . . . 662 make-event . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 make-extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 make-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 618 make-file-part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 make-font-bold, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 make-font-bold-italic, x- . . . . . . . . . . . . . . . . . . . . . . 625 make-font-instance . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 make-font-italic, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 make-font-unbold, x- . . . . . . . . . . . . . . . . . . . . . . . . . 624 make-font-unitalic, x- . . . . . . . . . . . . . . . . . . . . . . . . . 625 make-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 make-frame-invisible . . . . . . . . . . . . . . . . . . . . . . . . . . 482 make-frame-visible. . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 make-glyph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 make-glyph-internal . . . . . . . . . . . . . . . . . . . . . . . . . . 627 make-hashtable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 make-icon-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 make-image-instance. . . . . . . . . . . . . . . . . . . . . . . . . . 638 make-image-specifier. . . . . . . . . . . . . . . . . . . . . . . . . . 633 make-indirect-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . 445 make-key-weak-hashtable . . . . . . . . . . . . . . . . . . . . . 668 make-keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 make-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 make-local-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 make-local-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 make-marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 make-obsolete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 make-obsolete-variable . . . . . . . . . . . . . . . . . . . . . . . . 391 make-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . 627 make-range-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 make-reverse-direction-charset . . . . . . . . . . . . . . . . 739 make-sparse-keymap . . . . . . . . . . . . . . . . . . . . . . . . . . 318 make-specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613 make-specifier-and-init. . . . . . . . . . . . . . . . . . . . . . . . 614 make-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 make-symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 make-symbolic-link . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 make-syntax-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 make-temp-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 make-tooltalk-message . . . . . . . . . . . . . . . . . . . . . . . . 722 make-tooltalk-pattern . . . . . . . . . . . . . . . . . . . . . . . . 724 make-tty-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 make-value-weak-hashtable . . . . . . . . . . . . . . . . . . . 669 make-variable-buffer-local. . . . . . . . . . . . . . . . . . . . . 160 make-vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 make-weak-hashtable . . . . . . . . . . . . . . . . . . . . . . . . . 668 make-weak-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 make-x-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 making them interactive, functions, . . . . . . . . . . . 284 makunbound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Manual-page-minibuffer-history . . . . . . . . . . . . . . . 268 map, bookmark-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 map, Buffer-menu-mode- . . . . . . . . . . . . . . . . . . . . . 783 map, c-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 map, c++-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 map, command-history- . . . . . . . . . . . . . . . . . . . . . . 783 map, ctl-x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 map, ctl-x-4- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 map, ctl-x-5- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 map, current-global- . . . . . . . . . . . . . . . . . . . . . . . . . . 324 map, current-local- . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 map, debugger-mode-. . . . . . . . . . . . . . . . . . . . . . . . . 783 map, dired-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 map, edit-abbrevs- . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 map, edit-tab-stops- . . . . . . . . . . . . . . . . . . . . . . . . . . 783 map, electric-buffer-menu-mode- . . . . . . . . . . . . . . 783 map, electric-future- . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 map, electric-history- . . . . . . . . . . . . . . . . . . . . . . . . . 783 map, emacs-lisp-mode- . . . . . . . . . . . . . . . . . . . . . . . 783 map, esc-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 map, function-key- . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 map, global-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 map, help-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389, 783 map, Helper-help- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, Info-edit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, Info-mode-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, isearch-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, itimer-edit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, key-translation- . . . . . . . . . . . . . . . . . . . . . . . . . 710 map, lisp-interaction-mode- . . . . . . . . . . . . . . . . . . . 784 map, lisp-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, menu-accelerator- . . . . . . . . . . . . . . . . . . . . . . . 349 map, minibuffer-local- . . . . . . . . . . . . . . . . . . . 264, 784 map, minibuffer-local-completion- . . . . . . . . 271, 784 map, minibuffer-local-isearch- . . . . . . . . . . . . . . . . . 784 map, minibuffer-local-must-match- . . . . . . . 272, 784 map, minibuffer-local-ns- . . . . . . . . . . . . . . . . . . . . . 265 map, mode-specific-. . . . . . . . . . . . . . . . . . . . . . 321, 784 map, modeline- . . . . . . . . . . . . . . . . . . . . . . . . . . 325, 784 map, objc-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, occur-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 map, overriding-local-. . . . . . . . . . . . . . . . . . . . 325, 784 map, overriding-terminal-local- . . . . . . . . . . . . . . . 325 map, query-replace-. . . . . . . . . . . . . . . . . . . . . . 563, 784 map, read-expression- . . . . . . . . . . . . . . . . . . . . . . . . 785 map, read-shell-command- . . . . . . . . . . . . . . . . . . . . 785 map, shared-lisp-mode- . . . . . . . . . . . . . . . . . . . . . . . 785 map, text-mode-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 map, toolbar- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325, 785 map, use-global- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 map, use-local- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 map, view-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 map-alist, minor-mode- . . . . . . . . . . . . . . . . . . . . . . . 325 map-char-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 map-database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 map-extent-children . . . . . . . . . . . . . . . . . . . . . . . . . . 593 map-extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591 map-frame-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 map-keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 map-range-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 map-specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 map-y-or-n-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 mapatoms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 mapcar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 mapcar-extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 592 mapconcat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 maphash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 mapping functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 mapping, extents, . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591 maps, current-minor-mode- . . . . . . . . . . . . . . . . . . . 324 margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 margin width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 margin, current-left- . . . . . . . . . . . . . . . . . . . . . . . . . . 530 margin, delete-to-left- . . . . . . . . . . . . . . . . . . . . . . . . 531 margin, indent-to-left- . . . . . . . . . . . . . . . . . . . . . . . . 531 margin, inside . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 margin, left- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 margin, move-to-left- . . . . . . . . . . . . . . . . . . . . . . . . . 531 margin, outside. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 margin, set-left- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 margin, set-right- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 margin-pixel-width, window-left- . . . . . . . . . . . . . . 647 margin-pixel-width, window-right- . . . . . . . . . . . . 647 margin-width, left- . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 margin-width, right- . . . . . . . . . . . . . . . . . . . . . . . . . . 646 mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 mark (Edebug), current buffer point and . . . . . . 242 mark excursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 mark in character constant, question . . . . . . . . . . . 23 mark ring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 mark ring, global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 mark, abbrev-prefix- . . . . . . . . . . . . . . . . . . . . . . . . . . 584 mark, current buffer . . . . . . . . . . . . . . . . . . . . . . . . . . 507 mark, exchange-point-and-. . . . . . . . . . . . . . . . . . . . 508 mark, pop- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 mark, pop-global- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 mark, process-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 mark, push- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 mark, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 mark, the . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 mark-marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 mark-ring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 mark-ring, global- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 mark-ring-max. . . . . . . . . . . . . . . . . . . . . . . . . . . 508, 509 marker argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 marker garbage collection. . . . . . . . . . . . . . . . . . . . . 501 marker input stream. . . . . . . . . . . . . . . . . . . . . . . . . . 253 marker output stream . . . . . . . . . . . . . . . . . . . . . . . . 256 marker relocation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 marker, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 504 marker, end of buffer . . . . . . . . . . . . . . . . . . . . . . . . . 504 marker, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 marker, mark- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 marker, move- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 marker, point- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 marker, point-max- . . . . . . . . . . . . . . . . . . . . . . . . . . . 504 marker, point-min- . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 marker, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 marker-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 marker-p, integer-char-or-. . . . . . . . . . . . . . . . . . . . . 503 marker-p, integer-or- . . . . . . . . . . . . . . . . . . . . . . . . . 503 marker-p, number-char-or- . . . . . . . . . . . . . . . . . . . . 503 marker-p, number-or- . . . . . . . . . . . . . . . . . . . . . . . . . 503 marker-position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 markerp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 markers as numbers . . . . . . . . . . . . . . . . . . . . . . . . . . 501 markers vs. extents. . . . . . . . . . . . . . . . . . . . . . . . . . . 501 markers, insert-before- . . . . . . . . . . . . . . . . . . . . . . . . 516 mask-file-name, image-instance-. . . . . . . . . . . . . . . 639 match data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 563 match, posix-string- . . . . . . . . . . . . . . . . . . . . . . . . . . 562 match, replace- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 match, string- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 match-beginning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 match-data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 match-data, save-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 match-data, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 match-data, store- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 match-end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 match-map, minibuffer-local-must- . . . . . . . 272, 784 match-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 matches-key-specifier-p, event- . . . . . . . . . . . . . . . . 321 matches-specifier-tag-set-p, device-. . . . . . . . . . . . 611 matching, parenthesis. . . . . . . . . . . . . . . . . . . . . . . . . 660 matching-history-element, next-. . . . . . . . . . . . . . . 281 matching-history-element, previous-. . . . . . . . . . . 281 matching-open, blink- . . . . . . . . . . . . . . . . . . . . . . . . 660 matching-paren, blink-. . . . . . . . . . . . . . . . . . . . . . . . 660 matching-paren-delay, blink- . . . . . . . . . . . . . . . . . . 660 matching-paren-distance, blink- . . . . . . . . . . . . . . . 660 matching-specifier-tag-list, device- . . . . . . . . . . . . 611 mathematical functions . . . . . . . . . . . . . . . . . . . . . . . . 58 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 max, kill-ring- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 max, mark-ring- . . . . . . . . . . . . . . . . . . . . . . . . . 508, 509 max, point- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 max-lisp-eval-depth. . . . . . . . . . . . . . . . . . . . . . . . . . . 123 max-marker, point- . . . . . . . . . . . . . . . . . . . . . . . . . . . 504 max-size, buffers-menu- . . . . . . . . . . . . . . . . . . . . . . . 349 max-size, log-message- . . . . . . . . . . . . . . . . . . . . . . . . 652 max-specpdl-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 maximum when printing, string length, . . . . . . . 261 maybe, indent-relative- . . . . . . . . . . . . . . . . . . . . . . . 539 maybe-dialog-box, y-or-n-p- . . . . . . . . . . . . . . . . . . 279 maybe-dialog-box, yes-or-no-p-. . . . . . . . . . . . . . . . 279 member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 member, lax-plist- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 member, plist- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 membership in a list . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 memory allocation. . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 memory-limit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 773 memq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 menu accelerators . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 menu accelerators, keyboard . . . . . . . . . . . . . . . . . . 347 menu filters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 menu format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 menu, accelerate- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 menu, add- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 menu, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 menu, buffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349 menu, default-popup- . . . . . . . . . . . . . . . . . . . . . . . . . 347 menu, global-popup-. . . . . . . . . . . . . . . . . . . . . . . . . . 347 menu, mode-popup- . . . . . . . . . . . . . . . . . . . . . . . . . . 347 menu, pop-up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 menu, popup- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 menu, popup-buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . 347 menu, popup-menubar- . . . . . . . . . . . . . . . . . . . . . . . 347 menu, popup-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . 347 menu, set-annotation- . . . . . . . . . . . . . . . . . . . . . . . . 646 menu-accelerator-enabled . . . . . . . . . . . . . . . . . . . . . 348 menu-accelerator-map . . . . . . . . . . . . . . . . . . . . . . . . 349 menu-accelerator-modifiers. . . . . . . . . . . . . . . . . . . . 349 menu-accelerator-prefix . . . . . . . . . . . . . . . . . . . . . . . 349 menu-button, add- . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 menu-filter, buffers-. . . . . . . . . . . . . . . . . . . . . . . . . . . 346 menu-filter, edit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 menu-filter, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 menu-hook, activate-popup- . . . . . . . . . . . . . . . . . . 347 menu-item, add- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 menu-item, delete-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 menu-item, disable-. . . . . . . . . . . . . . . . . . . . . . . . . . . 345 menu-item, enable- . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 menu-item, find-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 menu-item, relabel- . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 menu-line, format-buffers- . . . . . . . . . . . . . . . . . . . . 350 menu-max-size, buffers- . . . . . . . . . . . . . . . . . . . . . . . 349 menu-mode-map, Buffer- . . . . . . . . . . . . . . . . . . . . . 783 menu-mode-map, electric-buffer- . . . . . . . . . . . . . . 783 menu-no-selection-hook . . . . . . . . . . . . . . . . . . . . . . . 344 menu-p, complex-buffers- . . . . . . . . . . . . . . . . . . . . . 350 menu-switch-to-buffer-function, buffers- . . . . . . . 350 menu-titles, popup-. . . . . . . . . . . . . . . . . . . . . . . . . . . 347 menu-up-p, popup- . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 menubar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 menubar format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 menubar, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 menubar, default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 343 menubar, format of the . . . . . . . . . . . . . . . . . . . . . . . 342 menubar, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 343 menubar, set-buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . 343 menubar-configuration . . . . . . . . . . . . . . . . . . . . . . . . 342 menubar-dirty-flag, set- . . . . . . . . . . . . . . . . . . . . . . . 343 menubar-hook, activate- . . . . . . . . . . . . . . . . . . . . . . 343 menubar-menu, popup- . . . . . . . . . . . . . . . . . . . . . . . 347 menubar-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . 641 menubar-show-keybindings . . . . . . . . . . . . . . . . . . . 343 menus, format of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 message notation, error . . . . . . . . . . . . . . . . . . . . . . . . 12 message, clear- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 652 message, create-tooltalk-. . . . . . . . . . . . . . . . . . . . . . 723 message, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 652 message, describe-tooltalk-. . . . . . . . . . . . . . . . . . . . 725 message, destroy-tooltalk- . . . . . . . . . . . . . . . . . . . . 724 message, display- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 message, gc- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 message, inhibit-startup-. . . . . . . . . . . . . . . . . . . . . . 694 message, inhibit-startup-echo-area-. . . . . . . . . . . . 694 message, make-tooltalk- . . . . . . . . . . . . . . . . . . . . . . 722 message, new file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 message, print-help-return- . . . . . . . . . . . . . . . . . . . 390 message, return-tooltalk- . . . . . . . . . . . . . . . . . . . . . 722 message, send-tooltalk- . . . . . . . . . . . . . . . . . . . . . . . 722 message, ToolTalk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 message-arg, add-tooltalk- . . . . . . . . . . . . . . . . . . . . 723 message-attribute, get-tooltalk- . . . . . . . . . . . . . . . 722 message-attribute, set-tooltalk- . . . . . . . . . . . . . . . 723 message-ignore-labels, log- . . . . . . . . . . . . . . . . . . . . 652 message-ignore-regexps, log- . . . . . . . . . . . . . . . . . . 652 message-max-size, log- . . . . . . . . . . . . . . . . . . . . . . . . 652 messages, receiving ToolTalk. . . . . . . . . . . . . . . . . . 724 messages, sending ToolTalk . . . . . . . . . . . . . . . . . . . 721 meta character printing. . . . . . . . . . . . . . . . . . . . . . . 388 meta-prefix-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 min, point- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 min-height, window- . . . . . . . . . . . . . . . . . . . . . . . . . . 469 min-marker, point- . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 min-width, window- . . . . . . . . . . . . . . . . . . . . . . . . . . 470 minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 minibuffer history . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 minibuffer input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 minibuffer window. . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 minibuffer, ? in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 minibuffer, eval- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 minibuffer, exit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 minibuffer, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 minibuffer, read-from- . . . . . . . . . . . . . . . . . . . . . . . . 264 minibuffer, SPC in . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 minibuffer, TAB in . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 minibuffer-complete . . . . . . . . . . . . . . . . . . . . . . . . . . 272 minibuffer-complete-and-exit. . . . . . . . . . . . . . . . . . 272 minibuffer-complete-word . . . . . . . . . . . . . . . . . . . . . 272 minibuffer-completion-confirm . . . . . . . . . . . . . . . . 272 minibuffer-completion-help . . . . . . . . . . . . . . . . . . . 272 minibuffer-completion-predicate. . . . . . . . . . . . . . . 272 minibuffer-completion-table . . . . . . . . . . . . . . . . . . . 272 minibuffer-depth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 minibuffer-exit-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 281 minibuffer-frame, default-. . . . . . . . . . . . . . . . . . . . . 480 minibuffer-frame-plist. . . . . . . . . . . . . . . . . . . . . . . . . 474 minibuffer-help-form. . . . . . . . . . . . . . . . . . . . . . . . . . 281 minibuffer-history . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 minibuffer-history, Info- . . . . . . . . . . . . . . . . . . . . . . 268 minibuffer-history, Manual-page- . . . . . . . . . . . . . . 268 minibuffer-local-completion-map . . . . . . . . . 271, 784 minibuffer-local-isearch-map . . . . . . . . . . . . . . . . . . 784 minibuffer-local-map . . . . . . . . . . . . . . . . . . . . . 264, 784 minibuffer-local-must-match-map . . . . . . . . 272, 784 minibuffer-local-ns-map. . . . . . . . . . . . . . . . . . . . . . . 265 minibuffer-p, window- . . . . . . . . . . . . . . . . . . . . . . . . 281 minibuffer-prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 minibuffer-prompt-width . . . . . . . . . . . . . . . . . . . . . 281 minibuffer-scroll-window . . . . . . . . . . . . . . . . . . . . . . 282 minibuffer-setup-hook . . . . . . . . . . . . . . . . . . . . . . . . 281 minibuffer-window. . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 minibuffer-window, active- . . . . . . . . . . . . . . . . . . . . 281 minibuffer-window-active-p . . . . . . . . . . . . . . . . . . . 282 minibuffers, enable-recursive- . . . . . . . . . . . . . . . . . 282 minimum window size . . . . . . . . . . . . . . . . . . . . . . . . 469 minimum-level, display-warning- . . . . . . . . . . . . . . 654 minimum-level, log-warning- . . . . . . . . . . . . . . . . . . 654 minor mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 minor mode conventions . . . . . . . . . . . . . . . . . . . . . . 373 minor modes, conventions for writing . . . . . . . . . 373 minor modes, self-insert-command, . . . . . . . . . . . 374 minor-mode-alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 minor-mode-key-binding . . . . . . . . . . . . . . . . . . . . . . 328 minor-mode-map-alist . . . . . . . . . . . . . . . . . . . . . . . . 325 minor-mode-maps, current- . . . . . . . . . . . . . . . . . . . 324 minor-version, emacs-. . . . . . . . . . . . . . . . . . . . . . . . . 769 misc-user-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 mismatch, backup-by-copying-when- . . . . . . . . . . 425 mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 mode conventions, minor . . . . . . . . . . . . . . . . . . . . . 373 mode help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 mode hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 mode hook, major . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 mode keymap, major . . . . . . . . . . . . . . . . . . . . . . . . . 323 mode loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 mode modeline-format, Shell. . . . . . . . . . . . . . . . . . 376 mode specification error, file . . . . . . . . . . . . . . . . . . 369 mode variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 mode, abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 mode, Auto Fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 mode, auto-save- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 mode, batch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714 mode, current-input- . . . . . . . . . . . . . . . . . . . . . . . . . 708 mode, default-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . 581 mode, default-major- . . . . . . . . . . . . . . . . . . . . . . . . . 370 mode, define-derived- . . . . . . . . . . . . . . . . . . . . . . . . . 372 mode, describe-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 mode, documentation for major. . . . . . . . . . . . . . . 371 mode, Edebug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 mode, edebug-initial- . . . . . . . . . . . . . . . . . . . . . . . . . 250 mode, Fundamental . . . . . . . . . . . . . . . . . . . . . . . . . . 363 mode, fundamental- . . . . . . . . . . . . . . . . . . . . . . . . . . 369 mode, help for major . . . . . . . . . . . . . . . . . . . . . . . . . 371 mode, indent-according-to- . . . . . . . . . . . . . . . . . . . 536 mode, indent-tabs- . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 mode, initial-major- . . . . . . . . . . . . . . . . . . . . . . . . . . 370 mode, major . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 mode, major-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 mode, minor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 mode, newline and Auto Fill . . . . . . . . . . . . . . . . . . 517 mode, normal-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 mode, Outline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 mode, overwrite- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 mode, set-auto-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 mode, set-buffer-major- . . . . . . . . . . . . . . . . . . . . . . . 370 mode, set-input- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 mode, vc- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 mode, visited file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 mode, x-debug-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 mode-abbrev-table, c- . . . . . . . . . . . . . . . . . . . . . . . . 585 mode-abbrev-table, fundamental- . . . . . . . . . . . . . 585 mode-abbrev-table, lisp- . . . . . . . . . . . . . . . . . . . . . . 585 mode-abbrev-table, text- . . . . . . . . . . . . . . . . . . . . . 585 mode-alist, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 mode-alist, interpreter- . . . . . . . . . . . . . . . . . . . . . . . 371 mode-alist, minor- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 mode-class property . . . . . . . . . . . . . . . . . . . . . . . . . . 365 mode-hook, change-major- . . . . . . . . . . . . . . . . . . . . 365 mode-key-binding, minor-. . . . . . . . . . . . . . . . . . . . . 328 mode-map, Buffer-menu- . . . . . . . . . . . . . . . . . . . . . 783 mode-map, c- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 mode-map, c++- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 mode-map, debugger-. . . . . . . . . . . . . . . . . . . . . . . . . 783 mode-map, dired- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 783 mode-map, electric-buffer-menu- . . . . . . . . . . . . . . 783 mode-map, emacs-lisp- . . . . . . . . . . . . . . . . . . . . . . . 783 mode-map, Info-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 mode-map, isearch- . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 mode-map, lisp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 mode-map, lisp-interaction- . . . . . . . . . . . . . . . . . . . 784 mode-map, objc- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 mode-map, occur- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 mode-map, shared-lisp- . . . . . . . . . . . . . . . . . . . . . . . 785 mode-map, text-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 mode-map, view- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 mode-map-alist, minor- . . . . . . . . . . . . . . . . . . . . . . . 325 mode-maps, current-minor- . . . . . . . . . . . . . . . . . . . 324 mode-menu, popup- . . . . . . . . . . . . . . . . . . . . . . . . . . 347 mode-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 mode-popup-menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 mode-specific-map . . . . . . . . . . . . . . . . . . . . . . . 321, 784 mode-string, global- . . . . . . . . . . . . . . . . . . . . . . . . . . 377 mode-syntax-table, c-. . . . . . . . . . . . . . . . . . . . . . . . . 580 mode-syntax-table, emacs-lisp-. . . . . . . . . . . . . . . . 580 mode-syntax-table, text- . . . . . . . . . . . . . . . . . . . . . . 580 mode.el, lisp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 366 modeline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 374 modeline construct . . . . . . . . . . . . . . . . . . . . . . . . . . . 375 modeline, percent symbol in . . . . . . . . . . . . . . . . . . 375 modeline, redraw- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 374 modeline-buffer-identification . . . . . . . . . . . . . . . . . 377 modeline-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 375 modeline-format, default- . . . . . . . . . . . . . . . . . . . . . 378 modeline-format, Shell mode. . . . . . . . . . . . . . . . . . 376 modeline-map . . . . . . . . . . . . . . . . . . . . . . . . . . . 325, 784 modeline-modified . . . . . . . . . . . . . . . . . . . . . . . . . . . . 376 modeline-p, event-over- . . . . . . . . . . . . . . . . . . . . . . . 298 modeline-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . 641 modeline-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 modes and MS-DOS, file. . . . . . . . . . . . . . . . . . . . . . 408 modes of files, setting. . . . . . . . . . . . . . . . . . . . . . . . . 406 modes, abbrev tables in. . . . . . . . . . . . . . . . . . . . . . . 364 modes, buffer-local variables in . . . . . . . . . . . . . . . 364 modes, conventions for writing minor . . . . . . . . . 373 modes, default-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 modes, Edebug execution . . . . . . . . . . . . . . . . . . . . . 232 modes, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 modes, functions in . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 modes, input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 modes, keymaps in . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 modes, MS-DOS and file. . . . . . . . . . . . . . . . . . . . . . 408 modes, self-insert-command, minor . . . . . . . . . . . 374 modes, set-default-file-. . . . . . . . . . . . . . . . . . . . . . . . 408 modes, set-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 modes, syntax tables in . . . . . . . . . . . . . . . . . . . . . . . 364 modes, terminal input . . . . . . . . . . . . . . . . . . . . . . . . 708 modification flag (of buffer) . . . . . . . . . . . . . . . . . . . 438 modification of lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 modification time, comparison of. . . . . . . . . . . . . . 439 modification time, file . . . . . . . . . . . . . . . . . . . . . . . . 402 modification, buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . 438 modified, modeline-. . . . . . . . . . . . . . . . . . . . . . . . . . . 376 modified, not- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 modified-p, buffer-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 modified-p, set-buffer- . . . . . . . . . . . . . . . . . . . . . . . . 438 modified-tick, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . 439 modified-tick, string- . . . . . . . . . . . . . . . . . . . . . . . . . . 69 modifier-bits, event- . . . . . . . . . . . . . . . . . . . . . . . . . . 300 modifiers, event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 modifiers, menu-accelerator- . . . . . . . . . . . . . . . . . . 349 modify-syntax-entry . . . . . . . . . . . . . . . . . . . . . . . . . . 576 modifying, strings, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 modtime, clear-visited-file-. . . . . . . . . . . . . . . . . . . . 439 modtime, set-visited-file-. . . . . . . . . . . . . . . . . . . . . . 439 modtime, verify-visited-file- . . . . . . . . . . . . . . . . . . . 439 modtime, visited-file- . . . . . . . . . . . . . . . . . . . . . . . . . 439 modulus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 momentary-string-display . . . . . . . . . . . . . . . . . . . . . 659 mono-pixmap-image-instance-p . . . . . . . . . . . . . . . 638 more storage, CL note—allocate . . . . . . . . . . . . . . 770 motion, Lisp expression. . . . . . . . . . . . . . . . . . . . . . . 495 motion, list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 motion, sexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 motion, vertical-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 motion-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 motion-pixels, vertical- . . . . . . . . . . . . . . . . . . . . . . . 494 mouse cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 mouse pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 mouse-event, current-. . . . . . . . . . . . . . . . . . . . . . . . . 292 mouse-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 mouse-face, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 mouse-face, set-extent- . . . . . . . . . . . . . . . . . . . . . . . 597 mouse-grabbed-buffer. . . . . . . . . . . . . . . . . . . . . . . . . 325 mouse-highlight-priority . . . . . . . . . . . . . . . . . . . . . . 600 move-marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 move-to-column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 move-to-left-margin. . . . . . . . . . . . . . . . . . . . . . . . . . . 531 move-to-window-line. . . . . . . . . . . . . . . . . . . . . . . . . . 494 MS-DOS and file modes . . . . . . . . . . . . . . . . . . . . . . 408 MS-DOS file types. . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 MS-DOS, file modes and. . . . . . . . . . . . . . . . . . . . . . 408 MS-DOS, file types on . . . . . . . . . . . . . . . . . . . . . . . . 421 MSWindows OLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 multilingual string formatting. . . . . . . . . . . . . . . . . . 71 multiple names, file with. . . . . . . . . . . . . . . . . . . . . . 406 multiple windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 must-match-map, minibuffer-local- . . . . . . . 272, 784

N

n, C-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 n-p, map-y-or-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 n-p, y-or-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 n-p-maybe-dialog-box, y-or- . . . . . . . . . . . . . . . . . . 279 name abbreviation, directory . . . . . . . . . . . . . . . . . 410 name cell, print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 name completion subroutines, file . . . . . . . . . . . . . 413 name hashing, symbol . . . . . . . . . . . . . . . . . . . . . . . . 115 name of buffer, file . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 name of directory, file . . . . . . . . . . . . . . . . . . . . . . . . 410 name), directory part (of file. . . . . . . . . . . . . . . . . . 408 name), nondirectory part (of file . . . . . . . . . . . . . . 408 name), version number (in file . . . . . . . . . . . . . . . . 408 name, abbrev-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 name, abbreviate-file-. . . . . . . . . . . . . . . . . . . . . . . . . 410 name, absolute file. . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 name, auto-save-list-file- . . . . . . . . . . . . . . . . . . . . . . 430 name, auto-save-visited-file-. . . . . . . . . . . . . . . . . . . 428 name, buffer file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 name, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 name, buffer-auto-save-file- . . . . . . . . . . . . . . . . . . . 427 name, buffer-file- . . . . . . . . . . . . . . . . . . . . . . . . 436, 437 name, charset-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 name, coding-system-. . . . . . . . . . . . . . . . . . . . . . . . . 749 name, color- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626 name, color-instance- . . . . . . . . . . . . . . . . . . . . . . . . . 626 name, completion, file . . . . . . . . . . . . . . . . . . . . . . . . 413 name, condition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 name, database-file- . . . . . . . . . . . . . . . . . . . . . . . . . . 674 name, default-frame- . . . . . . . . . . . . . . . . . . . . . . . . . 477 name, define-logical- . . . . . . . . . . . . . . . . . . . . . . . . . . 407 name, directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 410 name, directory-file- . . . . . . . . . . . . . . . . . . . . . . . . . . 410 name, error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 name, expand-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 name, face-font- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 name, file-relative-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 name, find-backup-file- . . . . . . . . . . . . . . . . . . . . . . . 427 name, font instance. . . . . . . . . . . . . . . . . . . . . . . . . . . 624 name, font-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 name, font-instance- . . . . . . . . . . . . . . . . . . . . . . . . . . 624 name, frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 name, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 name, function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 name, generate-new-buffer- . . . . . . . . . . . . . . . . . . . 436 name, image-instance- . . . . . . . . . . . . . . . . . . . . . . . . 639 name, image-instance-file- . . . . . . . . . . . . . . . . . . . . 639 name, image-instance-mask-file-. . . . . . . . . . . . . . . 639 name, internal-doc-file- . . . . . . . . . . . . . . . . . . . . . . . 386 name, invocation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 name, keymap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 name, make-auto-save-file- . . . . . . . . . . . . . . . . . . . . 428 name, make-backup-file- . . . . . . . . . . . . . . . . . . . . . . 426 name, make-temp-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 name, mode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 377 name, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 name, process-tty- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 name, read command . . . . . . . . . . . . . . . . . . . . . . . . . 289 name, read-file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 name, relative file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 name, set-keymap- . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 name, set-visited-file- . . . . . . . . . . . . . . . . . . . . . . . . . 437 name, substitute-in-file-. . . . . . . . . . . . . . . . . . . . . . . 412 name, symbol-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 name, system- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 name, user-full-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 name, user-login- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 name, user-real-login-. . . . . . . . . . . . . . . . . . . . . . . . . 703 name-absolute-p, file- . . . . . . . . . . . . . . . . . . . . . . . . . 411 name-all-completions, file- . . . . . . . . . . . . . . . . . . . . 413 name-as-directory, file-. . . . . . . . . . . . . . . . . . . . . . . . 410 name-buffer-file-type-alist, file-. . . . . . . . . . . . . . . . 421 name-completion, file- . . . . . . . . . . . . . . . . . . . . . . . . 414 name-directory, file- . . . . . . . . . . . . . . . . . . . . . . . . . . 409 name-directory, unhandled-file- . . . . . . . . . . . . . . . 418 name-handler, find-file- . . . . . . . . . . . . . . . . . . . . . . . 417 name-handlers, inhibit-file- . . . . . . . . . . . . . . . . . . . 417 name-history, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 name-list, abbrev-table- . . . . . . . . . . . . . . . . . . . . . . 582 name-nondirectory, file-. . . . . . . . . . . . . . . . . . . . . . . 409 name-operation, inhibit-file- . . . . . . . . . . . . . . . . . . 417 name-p, auto-save-file- . . . . . . . . . . . . . . . . . . . . . . . . 428 name-p, backup-file- . . . . . . . . . . . . . . . . . . . . . . . . . . 426 name-p, x-valid-keysym- . . . . . . . . . . . . . . . . . . . . . . 719 name-sans-extension, file- . . . . . . . . . . . . . . . . . . . . . 409 name-sans-versions, file- . . . . . . . . . . . . . . . . . . . . . . 409 name-to-file, add-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 named function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 names in directory, file. . . . . . . . . . . . . . . . . . . . . . . . 414 names, buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 names, expansion of file. . . . . . . . . . . . . . . . . . . . . . . 411 names, file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 names, file with multiple. . . . . . . . . . . . . . . . . . . . . . 406 names, magic file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416 names, same-window-buffer- . . . . . . . . . . . . . . . . . . 459 names, special-display-buffer- . . . . . . . . . . . . . . . . . 458 NaN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 narrow-to-page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 narrow-to-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 narrowing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 narrowing, point with . . . . . . . . . . . . . . . . . . . . . . . . 489 natnum-specifier-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 natnump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 natural numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 nconc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 necessary, delete-auto-save-file-if- . . . . . . . . . . . . . 429 negative infinity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 negative-argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 nesting error, Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 network connection . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 network-stream, open- . . . . . . . . . . . . . . . . . . . . . . . . 691 new file message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 new, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 524 new-buffer, generate- . . . . . . . . . . . . . . . . . . . . . . . . . 443 new-buffer-name, generate- . . . . . . . . . . . . . . . . . . . 436 new-versions, kept- . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 newer, load-warn-when-source-. . . . . . . . . . . . . . . . 200 newer-than-file-p, file- . . . . . . . . . . . . . . . . . . . . . . . . 402 newest-backup, file-. . . . . . . . . . . . . . . . . . . . . . . . . . . 427 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22, 517 newline and Auto Fill mode . . . . . . . . . . . . . . . . . . 517 newline in print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 newline in strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 newline, require-final-. . . . . . . . . . . . . . . . . . . . . . . . . 398 newline-and-indent . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 newline-and-indent, reindent-then- . . . . . . . . . . . . 537 newlines, print-escape-. . . . . . . . . . . . . . . . . . . . . . . . 260 newlines, use-hard- . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 next input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 next-call, debug-on- . . . . . . . . . . . . . . . . . . . . . . . . . . 227 next-command-event . . . . . . . . . . . . . . . . . . . . . . . . . 305 next-event. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 next-extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 590 next-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 next-history-element . . . . . . . . . . . . . . . . . . . . . . . . . . 281 next-matching-history-element . . . . . . . . . . . . . . . . 281 next-property-change . . . . . . . . . . . . . . . . . . . . . . . . . 544 next-screen-context-lines . . . . . . . . . . . . . . . . . . . . . . 464 next-single-property-change . . . . . . . . . . . . . . . . . . . 544 next-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 nil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 nil and lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 nil in keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 nil in lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 nil input stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 nil output stream. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 nil, uses of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ninth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 nlinks, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 nlistp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 no continuable errors, CL note— . . . . . . . . . . . . . 140 no questions, yes-or- . . . . . . . . . . . . . . . . . . . . . . . . . . 277 no-blanks-input, read- . . . . . . . . . . . . . . . . . . . . . . . . 265 no-catch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 no-p, yes-or-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 no-p-dialog-box, yes-or- . . . . . . . . . . . . . . . . . . . . . . . 279 no-p-maybe-dialog-box, yes-or-. . . . . . . . . . . . . . . . 279 no-redraw-on-reenter . . . . . . . . . . . . . . . . . . . . . . . . . 649 no-redraw-on-reenter), resume (cf. . . . . . . . . . . . . 649 no-redraw-on-reenter), suspend (cf. . . . . . . . . . . . 649 no-selection-hook, menu- . . . . . . . . . . . . . . . . . . . . . 344 nondirectory part (of file name). . . . . . . . . . . . . . . 408 nondirectory, file-name-. . . . . . . . . . . . . . . . . . . . . . . 409 noninteractive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714 noninteractive use . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714 nonlocal exits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 nonprinting characters, reading . . . . . . . . . . . . . . . 306 nontext-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . . 640 normal-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 369 noselect, find-file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 not. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 not available, closures . . . . . . . . . . . . . . . . . . . . . . . . 157 not, bitwise. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 not, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 not-all, text-property- . . . . . . . . . . . . . . . . . . . . . . . . 545 not-found-hooks, find-file- . . . . . . . . . . . . . . . . . . . . 395 not-modified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 notation, buffer text . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 notation, documentation . . . . . . . . . . . . . . . . . . . . . . . 11 notation, dotted pair . . . . . . . . . . . . . . . . . . . . . . . . . . 26 notation, error message . . . . . . . . . . . . . . . . . . . . . . . . 12 notation, evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 notation, hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 notation, printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 notation, XEmacs event standard . . . . . . . . . . . . . 388 note—allocate more storage, CL . . . . . . . . . . . . . . 770 note—case of letters, CL. . . . . . . . . . . . . . . . . . . . . . . 23 note—default optional arg, CL. . . . . . . . . . . . . . . . 168 note—integers vrs eq, CL . . . . . . . . . . . . . . . . . . . . . . 49 note—lack union, set, CL . . . . . . . . . . . . . . . . . . . . . . 92 note—no continuable errors, CL . . . . . . . . . . . . . . 140 note—only throw in Emacs, CL. . . . . . . . . . . . . . . 137 note—rplaca vrs setcar, CL . . . . . . . . . . . . . . . . . . . . 87 note—set local, CL . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 note—special forms compared, CL . . . . . . . . . . . . 128 note—special variables, CL . . . . . . . . . . . . . . . . . . . 156 note—symbol in obarrays, CL . . . . . . . . . . . . . . . . 116 nothing-image-instance-p . . . . . . . . . . . . . . . . . . . . . 638 nreverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 ns-map, minibuffer-local- . . . . . . . . . . . . . . . . . . . . . 265 nth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 nthcdr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 number (in file name), version . . . . . . . . . . . . . . . . 408 number equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 number, buffer-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . 437 number, string to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 number, string-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 number-char-or-marker-p . . . . . . . . . . . . . . . . . . . . . 503 number-of-arguments, wrong- . . . . . . . . . . . . . . . . . 167 number-or-marker-p . . . . . . . . . . . . . . . . . . . . . . . . . . 503 number-to-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 numberp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 numbers, markers as. . . . . . . . . . . . . . . . . . . . . . . . . . 501 numbers, natural . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 numbers, precision of formatted. . . . . . . . . . . . . . . . 72 numbers, printing floating-point . . . . . . . . . . . . . . 261 numbers, printing, floating-point . . . . . . . . . . . . . . 261 numbers, random . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 numeric prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 numeric prefix argument . . . . . . . . . . . . . . . . . . . . . . 310 numeric prefix argument usage. . . . . . . . . . . . . . . . 287 numeric-fields, sort- . . . . . . . . . . . . . . . . . . . . . . . . . . 534 numeric-value, prefix-. . . . . . . . . . . . . . . . . . . . . . . . . 311

O

obarray. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115, 117 obarray in completion . . . . . . . . . . . . . . . . . . . . . . . . 268 obarray), bucket (in . . . . . . . . . . . . . . . . . . . . . . . . . . 115 obarrays, CL note—symbol in . . . . . . . . . . . . . . . . 116 objc-mode-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 object to string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 object, event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 object, extent-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 object, Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 object, string to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 objects, window-system . . . . . . . . . . . . . . . . . . . . . . . 617 oblique . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 obsolete buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 obsolete, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 obsolete-function-alias, define- . . . . . . . . . . . . . . . . 391 obsolete-variable, make- . . . . . . . . . . . . . . . . . . . . . . 391 obsolete-variable-alias, define-. . . . . . . . . . . . . . . . . 392 obsoleteness-doc, function-. . . . . . . . . . . . . . . . . . . . 392 obsoleteness-doc, variable- . . . . . . . . . . . . . . . . . . . . 392 occur-mode-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 784 octal character code . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 octal character input . . . . . . . . . . . . . . . . . . . . . . . . . 306 octal, integer to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 octal-escape-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 octet, char-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 offer-save, buffer- . . . . . . . . . . . . . . . . . . . . . . . . 396, 444 OffiX DND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 old-eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 old-versions, kept- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 OLE, MSWindows. . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 one-space, just- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 one-window-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . 448, 450 only buffer, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 only buffers in interactive, read- . . . . . . . . . . . . . . 285 only throw in Emacs, CL note— . . . . . . . . . . . . . . 137 only, barf-if-buffer-read- . . . . . . . . . . . . . . . . . . . . . . 440 only, buffer, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 only, buffer-read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 only, find-file-read- . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 only, inhibit-read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 only, load-warn-when-source- . . . . . . . . . . . . . . . . . 200 only, toggle-read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 only-global-abbrevs . . . . . . . . . . . . . . . . . . . . . . . . . . . 582 open error, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 395 open parenthesis character. . . . . . . . . . . . . . . . . . . . 573 open, blink-matching- . . . . . . . . . . . . . . . . . . . . . . . . 660 open, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 open-database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 open-dribble-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 open-network-stream . . . . . . . . . . . . . . . . . . . . . . . . . 691 open-termscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 operating system environment . . . . . . . . . . . . . . . . 699 operation, inhibit-file-name- . . . . . . . . . . . . . . . . . . 417 option descriptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 option, user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 optional arg, CL note—default. . . . . . . . . . . . . . . . 168 optional arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 options on command line . . . . . . . . . . . . . . . . . . . . . 696 options, command line. . . . . . . . . . . . . . . . . . . . . . . . 696 or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 or, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 or, bitwise exclusive . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 or, logical exclusive . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 or, logical inclusive . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 or-char-int-p, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 or-char-p, integer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 or-frame-p, device- . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 or-frame-type, device- . . . . . . . . . . . . . . . . . . . . . . . . 486 or-marker-p, integer- . . . . . . . . . . . . . . . . . . . . . . . . . 503 or-marker-p, integer-char-. . . . . . . . . . . . . . . . . . . . . 503 or-marker-p, number- . . . . . . . . . . . . . . . . . . . . . . . . . 503 or-marker-p, number-char- . . . . . . . . . . . . . . . . . . . . 503 or-n-p, map-y-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 or-n-p, y-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 or-n-p-maybe-dialog-box, y- . . . . . . . . . . . . . . . . . . 279 or-no questions, yes- . . . . . . . . . . . . . . . . . . . . . . . . . . 277 or-no-p, yes-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 or-no-p-dialog-box, yes- . . . . . . . . . . . . . . . . . . . . . . . 279 or-no-p-maybe-dialog-box, yes-. . . . . . . . . . . . . . . . 279 or-string-p, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 order of extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 order, display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 order, extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 ordering of windows, cyclic . . . . . . . . . . . . . . . . . . . 453 ordering, cyclic, window . . . . . . . . . . . . . . . . . . . . . . 453 oriented functions, directory- . . . . . . . . . . . . . . . . . 414 other-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 other-window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 other-window, find-file- . . . . . . . . . . . . . . . . . . . . . . . 394 other-window, scroll- . . . . . . . . . . . . . . . . . . . . . . . . . 463 other-window, switch-to-buffer- . . . . . . . . . . . . . . . 456 other-window-scroll-buffer . . . . . . . . . . . . . . . . . . . . 463 other-windows, delete- . . . . . . . . . . . . . . . . . . . . . . . . 451 Outline mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 output from processes . . . . . . . . . . . . . . . . . . . . . . . . 685 output stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 output stream, buffer . . . . . . . . . . . . . . . . . . . . . . . . . 256 output stream, function . . . . . . . . . . . . . . . . . . . . . . 257 output stream, marker. . . . . . . . . . . . . . . . . . . . . . . . 256 output stream, nil . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 output stream, t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 output, accept-process- . . . . . . . . . . . . . . . . . . . . . . . 688 output, binary-process- . . . . . . . . . . . . . . . . . . . . . . . 679 output, process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 output, standard-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 output, terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 output-format, float- . . . . . . . . . . . . . . . . . . . . . . . . . 261 output-to-temp-buffer, with- . . . . . . . . . . . . . . . . . . 658 outside margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 over-border-p, event- . . . . . . . . . . . . . . . . . . . . . . . . . 299 over-glyph-p, event- . . . . . . . . . . . . . . . . . . . . . . . . . . 298 over-modeline-p, event- . . . . . . . . . . . . . . . . . . . . . . . 298 over-text-area-p, event- . . . . . . . . . . . . . . . . . . . . . . . 298 over-toolbar-p, event- . . . . . . . . . . . . . . . . . . . . . . . . . 299 overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 overflow, use-left- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 overflow, use-right- . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 overlay arrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 657 overlay-arrow-position . . . . . . . . . . . . . . . . . . . . . . . . 658 overlay-arrow-string . . . . . . . . . . . . . . . . . . . . . . . . . . 657 override, self-insert-command . . . . . . . . . . . . . . . . . 332 overriding-local-map . . . . . . . . . . . . . . . . . . . . . 325, 784 overriding-terminal-local-map . . . . . . . . . . . . . . . . . 325 overwrite-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 own-selection, x-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 ownership-preserved-p, file- . . . . . . . . . . . . . . . . . . . 402

P

p example, user-variable- . . . . . . . . . . . . . . . . . . . . . 274 p, auto-save-file-name- . . . . . . . . . . . . . . . . . . . . . . . . 428 p, backup-file-name- . . . . . . . . . . . . . . . . . . . . . . . . . . 426 p, bit-vector- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 p, boolean-specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . 605 p, bottom-toolbar-visible-. . . . . . . . . . . . . . . . . . . . . 357 p, buffer-glyph-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 p, buffer-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 443 p, buffer-modified-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 p, button-event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, button-press-event- . . . . . . . . . . . . . . . . . . . . . . . . 296 p, button-release-event- . . . . . . . . . . . . . . . . . . . . . . . 296 p, byte-recompile-directory-ignore-errors- . . . . . 210 p, case-table- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 p, category-designator- . . . . . . . . . . . . . . . . . . . . . . . 757 p, category-table-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 p, category-table-value- . . . . . . . . . . . . . . . . . . . . . . . 757 p, char-int- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 p, char-or-char-int- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 p, char-or-string- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 p, char-table- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 p, coding-system-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 746 p, color-instance- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626 p, color-pixmap-image-instance- . . . . . . . . . . . . . . 638 p, color-specifier- . . . . . . . . . . . . . . . . . . . . . . . . 605, 625 p, compiled-function- . . . . . . . . . . . . . . . . . . . . . . . . . 166 p, complex-buffers-menu- . . . . . . . . . . . . . . . . . . . . . 350 p, console-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 p, database-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 p, default-toolbar-visible- . . . . . . . . . . . . . . . . . . . . . 357 p, device-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 p, device-matches-specifier-tag-set-. . . . . . . . . . . . 611 p, device-or-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 p, eval-event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, event-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, event-matches-key-specifier- . . . . . . . . . . . . . . . . 321 p, event-over-border- . . . . . . . . . . . . . . . . . . . . . . . . . 299 p, event-over-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . 298 p, event-over-modeline- . . . . . . . . . . . . . . . . . . . . . . . 298 p, event-over-text-area- . . . . . . . . . . . . . . . . . . . . . . . 298 p, event-over-toolbar- . . . . . . . . . . . . . . . . . . . . . . . . . 299 p, extent-detached- . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 p, extent-in-region- . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 p, extent-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 p, face-boolean-specifier-. . . . . . . . . . . . . . . . . . . . . . 605 p, face-differs-from-default- . . . . . . . . . . . . . . . . . . . 623 p, face-underline- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 p, file-accessible-directory- . . . . . . . . . . . . . . . . . . . . 402 p, file-directory- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 p, file-executable-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 p, file-exists- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 p, file-locked- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 p, file-name-absolute- . . . . . . . . . . . . . . . . . . . . . . . . . 411 p, file-newer-than-file- . . . . . . . . . . . . . . . . . . . . . . . . 402 p, file-ownership-preserved- . . . . . . . . . . . . . . . . . . . 402 p, file-readable-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 p, file-regular- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 p, file-symlink- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 p, file-writable- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 p, font-instance- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 p, font-specifier- . . . . . . . . . . . . . . . . . . . . . . . . . 605, 623 p, frame-iconified- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 p, frame-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 p, frame-totally-visible- . . . . . . . . . . . . . . . . . . . . . . . 482 p, frame-visible- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 p, generic-specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 p, glyph-contrib- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 p, icon-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 p, image-instance- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 637 p, image-specifier- . . . . . . . . . . . . . . . . . . . . . . . 605, 633 p, input-pending- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 p, integer-char-or-marker-. . . . . . . . . . . . . . . . . . . . . 503 p, integer-or-char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 p, integer-or-marker- . . . . . . . . . . . . . . . . . . . . . . . . . 503 p, integer-specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 p, interactive- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 p, key-press-event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, ldap-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 p, left-toolbar-visible-. . . . . . . . . . . . . . . . . . . . . . . . . 357 p, local-variable-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 p, map-y-or-n-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 p, minibuffer-window-active- . . . . . . . . . . . . . . . . . . 282 p, misc-user-event- . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, mono-pixmap-image-instance- . . . . . . . . . . . . . . 638 p, motion-event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, mouse-event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, natnum-specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . 605 p, nothing-image-instance- . . . . . . . . . . . . . . . . . . . . 638 p, number-char-or-marker- . . . . . . . . . . . . . . . . . . . . 503 p, number-or-marker- . . . . . . . . . . . . . . . . . . . . . . . . . 503 p, one-window- . . . . . . . . . . . . . . . . . . . . . . . . . . 448, 450 p, pointer-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 p, pointer-image-instance- . . . . . . . . . . . . . . . . . . . . 638 p, popup-menu-up- . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 p, pos-visible-in-window- . . . . . . . . . . . . . . . . . . . . . 462 p, process-event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, process-kill-without-query- . . . . . . . . . . . . . . . . . 683 p, range-table-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 p, recent-auto-save-. . . . . . . . . . . . . . . . . . . . . . . . . . . 429 p, region-active- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 p, region-exists- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 p, right-toolbar-visible- . . . . . . . . . . . . . . . . . . . . . . . 357 p, set-buffer-modified- . . . . . . . . . . . . . . . . . . . . . . . . 438 p, set-face-underline- . . . . . . . . . . . . . . . . . . . . . . . . . 622 p, set-glyph-contrib- . . . . . . . . . . . . . . . . . . . . . . . . . . 631 p, set-window-dedicated- . . . . . . . . . . . . . . . . . . . . . 460 p, subwindow-image-instance-. . . . . . . . . . . . . . . . . 638 p, syntax-table- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 p, text-image-instance- . . . . . . . . . . . . . . . . . . . . . . . 637 p, timeout-event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 p, toolbar-buttons-captioned- . . . . . . . . . . . . . . . . . 357 p, toolbar-specifier- . . . . . . . . . . . . . . . . . . . . . . 356, 605 p, top-toolbar-visible-. . . . . . . . . . . . . . . . . . . . . . . . . 357 p, user-variable- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 p, valid-char-table-type- . . . . . . . . . . . . . . . . . . . . . . . 76 p, valid-char-table-value-. . . . . . . . . . . . . . . . . . . . . . . 77 p, valid-device-class- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 p, valid-device-type- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 p, valid-glyph-type-. . . . . . . . . . . . . . . . . . . . . . . . . . . 640 p, valid-image-instance-type-. . . . . . . . . . . . . . . . . . 637 p, valid-image-instantiator-format- . . . . . . . . . . . . 636 p, valid-inst-list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 p, valid-instantiator-. . . . . . . . . . . . . . . . . . . . . . . . . . 614 p, valid-plist- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 p, valid-spec-list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 p, valid-specifier-domain- . . . . . . . . . . . . . . . . . . . . . 614 p, valid-specifier-locale- . . . . . . . . . . . . . . . . . . . . . . . 614 p, valid-specifier-locale-type- . . . . . . . . . . . . . . . . . . 614 p, valid-specifier-tag-. . . . . . . . . . . . . . . . . . . . . 610, 614 p, valid-specifier-tag-set- . . . . . . . . . . . . . . . . . . . . . . 610 p, valid-specifier-type- . . . . . . . . . . . . . . . . . . . . . . . . 614 p, waiting-for-user-input- . . . . . . . . . . . . . . . . . . . . . 690 p, weak-list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 p, window-configuration- . . . . . . . . . . . . . . . . . . . . . 471 p, window-dedicated- . . . . . . . . . . . . . . . . . . . . 457, 459 p, window-highest- . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 p, window-live- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 p, window-lowest- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 p, window-minibuffer- . . . . . . . . . . . . . . . . . . . . . . . . 281 p, x-valid-keysym-name- . . . . . . . . . . . . . . . . . . . . . . 719 p, y-or-n-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 p, yes-or-no-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 p-dialog-box, yes-or-no- . . . . . . . . . . . . . . . . . . . . . . . 279 p-instance, glyph-contrib-. . . . . . . . . . . . . . . . . . . . . 631 p-maybe-dialog-box, y-or-n- . . . . . . . . . . . . . . . . . . 279 p-maybe-dialog-box, yes-or-no-. . . . . . . . . . . . . . . . 279 padding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 page, narrow-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 page-delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 page-minibuffer-history, Manual- . . . . . . . . . . . . . . 268 pages, sort-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 pair (in a specifier), inst- . . . . . . . . . . . . . . . . . . . . . 602 pair notation, dotted . . . . . . . . . . . . . . . . . . . . . . . . . . 26 pair, canonicalize-inst- . . . . . . . . . . . . . . . . . . . . . . . . 608 pair, set-case-syntax- . . . . . . . . . . . . . . . . . . . . . . . . . . 75 pair, specifier, inst- . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 paired delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 paragraph, fill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 paragraph, fill-region-as- . . . . . . . . . . . . . . . . . . . . . . 529 paragraph, filling a . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 paragraph-function, fill- . . . . . . . . . . . . . . . . . . . . . . 529 paragraph-separate . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 paragraph-start. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 paragraphs, fill-individual- . . . . . . . . . . . . . . . . . . . . 528 paragraphs, sort- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 parameters-alist, ldap-host- . . . . . . . . . . . . . . . . . . . 727 paren, blink-matching-. . . . . . . . . . . . . . . . . . . . . . . . 660 paren-delay, blink-matching- . . . . . . . . . . . . . . . . . . 660 paren-distance, blink-matching- . . . . . . . . . . . . . . . 660 paren-function, blink-. . . . . . . . . . . . . . . . . . . . . . . . . 660 paren-hook, blink- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 parent of a keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 parent process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 parent, extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 parent, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 parent, keymap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 parent, of extent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 parent, set-extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 parentheses, balancing . . . . . . . . . . . . . . . . . . . . . . . . 660 parentheses, indenting with . . . . . . . . . . . . . . . . . . . 578 parentheses, innermost containing . . . . . . . . . . . . 578 parenthesis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 parenthesis character, close . . . . . . . . . . . . . . . . . . . 573 parenthesis character, open . . . . . . . . . . . . . . . . . . . 573 parenthesis depth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 parenthesis matching . . . . . . . . . . . . . . . . . . . . . . . . . 660 parenthesis syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . 573 parenthesis, close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 660 parents, keymap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 parents, set-keymap-. . . . . . . . . . . . . . . . . . . . . . . . . . 319 parse state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 parse-partial-sexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 parse-sexp-ignore-comments . . . . . . . . . . . . . . . . . . 579 parsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 parsing, text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 part (of file name), directory. . . . . . . . . . . . . . . . . . 408 part (of file name), nondirectory . . . . . . . . . . . . . . 408 part, make-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 partial files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 418 partial-sexp, parse- . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 partial-width-windows, truncate-. . . . . . . . . . . . . . 650 paste-function, interprogram- . . . . . . . . . . . . . . . . . 524 PATH environment variable . . . . . . . . . . . . . . . . . . 675 path, exec- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 path, load- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 path, split- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 path, x-bitmap-file- . . . . . . . . . . . . . . . . . . . . . . 636, 719 path, x-library-search- . . . . . . . . . . . . . . . . . . . . . . . . 719 path-separator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 pattern, create-tooltalk- . . . . . . . . . . . . . . . . . . . . . . 725 pattern, destroy-tooltalk- . . . . . . . . . . . . . . . . . . . . . 725 pattern, make-tooltalk- . . . . . . . . . . . . . . . . . . . . . . . 724 pattern, register-tooltalk- . . . . . . . . . . . . . . . . . . . . . 725 pattern, ToolTalk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 724 pattern, unregister-tooltalk- . . . . . . . . . . . . . . . . . . 725 pattern-arg, add-tooltalk-. . . . . . . . . . . . . . . . . . . . . 725 pattern-attribute, add-tooltalk- . . . . . . . . . . . . . . . 725 pausing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 peculiar error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 peeking at input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 pending-p, input- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 percent symbol in modeline . . . . . . . . . . . . . . . . . . . 375 perform-replace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 performance analysis . . . . . . . . . . . . . . . . . . . . . . . . . 240 permanent local variable. . . . . . . . . . . . . . . . . . . . . . 161 permission . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 404 pid, emacs- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 pipes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 680 pixel, event-glyph-x- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 pixel, event-glyph-y- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 pixel, event-window-x- . . . . . . . . . . . . . . . . . . . . . . . . 297 pixel, event-window-y- . . . . . . . . . . . . . . . . . . . . . . . . 297 pixel, event-x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 pixel, event-y- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 pixel-edges, window-. . . . . . . . . . . . . . . . . . . . . . . . . . 468 pixel-edges, window-text-area- . . . . . . . . . . . . . . . . 468 pixel-height, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 pixel-height, window- . . . . . . . . . . . . . . . . . . . . . . . . . 466 pixel-height, window-displayed-text- . . . . . . . . . . 467 pixel-height, window-text-area- . . . . . . . . . . . . . . . 467 pixel-width, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 pixel-width, window- . . . . . . . . . . . . . . . . . . . . . . . . . 467 pixel-width, window-left-margin- . . . . . . . . . . . . . . 647 pixel-width, window-right-margin- . . . . . . . . . . . . 647 pixel-width, window-text-area- . . . . . . . . . . . . . . . . 467 pixels, enlarge-window- . . . . . . . . . . . . . . . . . . . . . . . 469 pixels, shrink-window- . . . . . . . . . . . . . . . . . . . . . . . . 469 pixels, vertical-motion- . . . . . . . . . . . . . . . . . . . . . . . 494 pixmap, background . . . . . . . . . . . . . . . . . . . . . . . . . . 617 pixmap, face-background-. . . . . . . . . . . . . . . . . . . . . 622 pixmap, set-face-background- . . . . . . . . . . . . . . . . . 622 pixmap, x-set-frame-icon- . . . . . . . . . . . . . . . . . . . . . 478 pixmap-image-instance-p, color- . . . . . . . . . . . . . . 638 pixmap-image-instance-p, mono- . . . . . . . . . . . . . . 638 pixmap-instance, face-background- . . . . . . . . . . . . 623 play-sound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 play-sound-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 plist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 plist, alist-to-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 plist, canonicalize- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 plist, canonicalize-lax- . . . . . . . . . . . . . . . . . . . . . . . . 100 plist, check-valid- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 plist, default-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . 475 plist, destructive-alist-to- . . . . . . . . . . . . . . . . . . . . . 100 plist, initial-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 plist, minibuffer-frame- . . . . . . . . . . . . . . . . . . . . . . . 474 plist, pop-up-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . 458 plist, special-display-frame- . . . . . . . . . . . . . . . . . . . 459 plist, symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 plist, symbol-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 plist-get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 plist-get, lax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 plist-member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 plist-member, lax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 plist-p, valid- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 plist-put. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 plist-put, lax-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 plist-remprop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 plist-remprop, lax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 plist-to-alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 plist-to-alist, destructive- . . . . . . . . . . . . . . . . . . . . . 100 plists-eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 120 plists-eq, lax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 plists-equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 120 plists-equal, lax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 point and mark (Edebug), current buffer . . . . . . 242 point excursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 point in window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 point numbers, printing floating-. . . . . . . . . . . . . . 261 point numbers, printing, floating- . . . . . . . . . . . . . 261 point with narrowing . . . . . . . . . . . . . . . . . . . . . . . . . 489 point, centering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 point, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 point, event-closest- . . . . . . . . . . . . . . . . . . . . . . . . . . 298 point, IEEE floating . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 point, insertion before . . . . . . . . . . . . . . . . . . . . . . . . 516 point, insertion, before. . . . . . . . . . . . . . . . . . . . . . . . 516 point, set-window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 point, window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 point, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 point-and-mark, exchange-. . . . . . . . . . . . . . . . . . . . 508 point-marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 point-max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 point-max-marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . 504 point-min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 point-min-marker. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 pointer (mouse) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 pointer, kill-ring-yank-. . . . . . . . . . . . . . . . . . . . . . . . 525 pointer, mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 pointer, set-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 pointer, x-grab- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 pointer, x-ungrab- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 pointer-glyph, busy- . . . . . . . . . . . . . . . . . . . . . . . . . . 641 pointer-glyph, gc- . . . . . . . . . . . . . . . . . . . . . . . . 641, 774 pointer-glyph, make- . . . . . . . . . . . . . . . . . . . . . . . . . 627 pointer-glyph, menubar- . . . . . . . . . . . . . . . . . . . . . . 641 pointer-glyph, modeline- . . . . . . . . . . . . . . . . . . . . . . 641 pointer-glyph, nontext- . . . . . . . . . . . . . . . . . . . . . . . 640 pointer-glyph, scrollbar- . . . . . . . . . . . . . . . . . . . . . . 641 pointer-glyph, selection- . . . . . . . . . . . . . . . . . . . . . . 641 pointer-glyph, text-. . . . . . . . . . . . . . . . . . . . . . . . . . . 640 pointer-glyph, toolbar-. . . . . . . . . . . . . . . . . . . . . . . . 641 pointer-glyph-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 pointer-image-instance-p. . . . . . . . . . . . . . . . . . . . . . 638 points, edebug-save-displayed-buffer- . . . . . 242, 250 points, stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 policy, layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 pop, yank- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 pop-global-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 pop-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 pop-to-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 456 pop-up menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 pop-up-frame-function . . . . . . . . . . . . . . . . . . . . . . . . 458 pop-up-frame-plist. . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 pop-up-frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 pop-up-windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 popup-buffer-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 popup-dialog-box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 popup-frame, special-display- . . . . . . . . . . . . . . . . . 459 popup-menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 popup-menu, default- . . . . . . . . . . . . . . . . . . . . . . . . . 347 popup-menu, global-. . . . . . . . . . . . . . . . . . . . . . . . . . 347 popup-menu, mode- . . . . . . . . . . . . . . . . . . . . . . . . . . 347 popup-menu-hook, activate- . . . . . . . . . . . . . . . . . . 347 popup-menu-titles . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 popup-menu-up-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 popup-menubar-menu . . . . . . . . . . . . . . . . . . . . . . . . 347 popup-mode-menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 port, ldap-default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 portion (of a buffer), accessible . . . . . . . . . . . . . . . 498 pos-visible-in-window-p . . . . . . . . . . . . . . . . . . . . . . . 462 position (in buffer) . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 position argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 position in window . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 position of frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 position of window . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 position, current buffer . . . . . . . . . . . . . . . . . . . . . . . 489 position, default-toolbar- . . . . . . . . . . . . . . . . . . . . . 355 position, extent end . . . . . . . . . . . . . . . . . . . . . . . . . . 588 position, extent start . . . . . . . . . . . . . . . . . . . . . . . . . 588 position, extent-end-. . . . . . . . . . . . . . . . . . . . . . . . . . 589 position, extent-start- . . . . . . . . . . . . . . . . . . . . . . . . 589 position, frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 position, horizontal . . . . . . . . . . . . . . . . . . . . . . . . . . . 535 position, marker- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 position, overlay-arrow- . . . . . . . . . . . . . . . . . . . . . . . 658 position, set-default-toolbar- . . . . . . . . . . . . . . . . . . 355 position, set-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 position, window. . . . . . . . . . . . . . . . . . . . . . . . . 460, 467 positive infinity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 posix-looking-at . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 posix-search-backward . . . . . . . . . . . . . . . . . . . . . . . . 562 posix-search-forward. . . . . . . . . . . . . . . . . . . . . . . . . . 562 posix-string-match. . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 post-command-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 284 post-gc-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 pre-abbrev-expand-hook . . . . . . . . . . . . . . . . . . . . . . 584 pre-command-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 pre-gc-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 774 preceding-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 precious-flag, file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 precisely, windows, controlling . . . . . . . . . . . . . . . . 455 precision of formatted numbers . . . . . . . . . . . . . . . . 72 precision, format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 predicate, backup-enable- . . . . . . . . . . . . . . . . . . . . . 424 predicate, minibuffer-completion- . . . . . . . . . . . . . 272 predicate, specifier-tag- . . . . . . . . . . . . . . . . . . . . . . . 611 predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 predicates, type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 prefix argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 prefix argument unreading . . . . . . . . . . . . . . . . . . . . 306 prefix argument usage, numeric . . . . . . . . . . . . . . . 287 prefix argument usage, raw . . . . . . . . . . . . . . . . . . . 287 prefix argument, execute with . . . . . . . . . . . . . . . . 289 prefix argument, numeric . . . . . . . . . . . . . . . . . . . . . 310 prefix argument, raw . . . . . . . . . . . . . . . . . . . . . . . . . 310 prefix command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 prefix key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 prefix key error, invalid . . . . . . . . . . . . . . . . . . . . . . . 330 prefix key, preventing . . . . . . . . . . . . . . . . . . . . . . . . . 327 prefix, Control-X- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 prefix, ESC- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 prefix, expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 prefix, fill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 prefix, menu-accelerator-. . . . . . . . . . . . . . . . . . . . . . 349 prefix, numeric . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 prefix, term-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 prefix-arg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 prefix-arg, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 prefix-bindings, describe- . . . . . . . . . . . . . . . . . . . . . 391 prefix-char, meta- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 prefix-chars, backward- . . . . . . . . . . . . . . . . . . . . . . . 577 prefix-command, define- . . . . . . . . . . . . . . . . . . . . . . 322 prefix-help-command . . . . . . . . . . . . . . . . . . . . . . . . . 390 prefix-mark, abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . 584 prefix-numeric-value . . . . . . . . . . . . . . . . . . . . . . . . . . 311 preserved-p, file-ownership- . . . . . . . . . . . . . . . . . . . 402 press-event-p, button- . . . . . . . . . . . . . . . . . . . . . . . . 296 press-event-p, key-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 preventing backtracking . . . . . . . . . . . . . . . . . . . . . . 245 preventing prefix key . . . . . . . . . . . . . . . . . . . . . . . . . 327 previous char, delete. . . . . . . . . . . . . . . . . . . . . . . . . . 519 previous complete subexpression . . . . . . . . . . . . . . 578 previous-extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591 previous-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 previous-history-element . . . . . . . . . . . . . . . . . . . . . . 281 previous-matching-history-element . . . . . . . . . . . . 281 previous-property-change . . . . . . . . . . . . . . . . . . . . . 545 previous-single-property-change . . . . . . . . . . . . . . . 545 previous-window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 primitive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 primitive type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 primitive types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 primitive-undo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 prin1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 prin1-to-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 princ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 print example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 print name cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 print, \n in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 print, cust- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 print, newline in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 print-circle, edebug- . . . . . . . . . . . . . . . . . . . . . 239, 251 print-escape-newlines . . . . . . . . . . . . . . . . . . . . . . . . . 260 print-gensym . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 print-help-return-message. . . . . . . . . . . . . . . . . . . . . 390 print-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 print-length, edebug- . . . . . . . . . . . . . . . . . . . . 239, 251 print-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 print-level, edebug- . . . . . . . . . . . . . . . . . . . . . . 239, 251 print-readably . . . . . . . . . . . . . . . . . . . . . . . . . . . 239, 261 print-string-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 print-trace-after, edebug- . . . . . . . . . . . . . . . . 240, 250 print-trace-before, edebug- . . . . . . . . . . . . . . . 240, 250 printed representation . . . . . . . . . . . . . . . . . . . . . . . . . 17 printed representation for characters . . . . . . . . . . . 21 printer, Lisp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 printing (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 printing circular structures . . . . . . . . . . . . . . . . . . . 239 printing floating-point numbers . . . . . . . . . . . . . . . 261 printing limits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 printing notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 printing readably . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 printing uninterned symbols . . . . . . . . . . . . . . . . . . 261 printing), stream (for . . . . . . . . . . . . . . . . . . . . . . . . . 256 printing, \ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 printing, character. . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 printing, control character . . . . . . . . . . . . . . . . . . . . 388 printing, escape characters in . . . . . . . . . . . . . . . . . 258 printing, event. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 printing, floating-point numbers, . . . . . . . . . . . . . . 261 printing, meta character . . . . . . . . . . . . . . . . . . . . . . 388 printing, quoting characters in . . . . . . . . . . . . . . . . 258 printing, string length, maximum when . . . . . . . 261 printing, uninterned symbols, . . . . . . . . . . . . . . . . . 261 priority of an extent . . . . . . . . . . . . . . . . . . . . . . . . . . 587 priority, extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 priority, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 priority, mouse-highlight- . . . . . . . . . . . . . . . . . . . . . 600 priority, set-extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 priority-list, coding- . . . . . . . . . . . . . . . . . . . . . . . . . . 750 priority-list, set-coding- . . . . . . . . . . . . . . . . . . . . . . . 750 process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 process filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 process input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 process output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 process sentinel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 689 process signals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 process window size . . . . . . . . . . . . . . . . . . . . . . . . . . 690 process, call- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 process, child . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 process, continue- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 process, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 process, event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 process, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 process, get-buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 process, interrupt- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 process, kill-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 process, modeline- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 process, parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 process, quit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 process, signal- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 process, start- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 679 process, stop-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 process-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 process-buffer, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 process-command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 process-connection-type. . . . . . . . . . . . . . . . . . . . . . . 680 process-environment . . . . . . . . . . . . . . . . . . . . . . . . . . 701 process-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 process-exit-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 process-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 process-filter, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 process-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 process-input, binary- . . . . . . . . . . . . . . . . . . . . . . . . 679 process-kill-without-query . . . . . . . . . . . . . . . . . . . . 681 process-kill-without-query-p . . . . . . . . . . . . . . . . . . 683 process-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 process-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 process-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 process-output, accept- . . . . . . . . . . . . . . . . . . . . . . . 688 process-output, binary- . . . . . . . . . . . . . . . . . . . . . . . 679 process-region, call-. . . . . . . . . . . . . . . . . . . . . . . . . . . 678 process-send-eof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 process-send-region . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 process-send-string . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 process-sentinel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 process-sentinel, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 690 process-shell-command, start-. . . . . . . . . . . . . . . . . 680 process-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 process-tty-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 process-window-size, set- . . . . . . . . . . . . . . . . . . . . . 690 processed, command-line- . . . . . . . . . . . . . . . . . . . . . 696 processes, delete-exited- . . . . . . . . . . . . . . . . . . . . . . 681 processes, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 680 processes, list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 processes, output from. . . . . . . . . . . . . . . . . . . . . . . . 685 processp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 profile.el. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 profiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 prog1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 prog2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 progn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 progn, implicit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 program arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 program directories . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 program, charset-ccl- . . . . . . . . . . . . . . . . . . . . . . . . . 740 program, execute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 program, insert-directory- . . . . . . . . . . . . . . . . . . . . 415 program, register-ccl- . . . . . . . . . . . . . . . . . . . . . . . . . 756 program, set-charset-ccl-. . . . . . . . . . . . . . . . . . . . . . 740 programmed completion . . . . . . . . . . . . . . . . . . . . . . 276 programming types . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 programs, timing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 progress, load-in- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 prompt, argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 prompt, keymap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 prompt, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . . . . 281 prompt, set-keymap- . . . . . . . . . . . . . . . . . . . . . . . . . 337 prompt-regexp, defun- . . . . . . . . . . . . . . . . . . . . . . . . 496 prompt-width, minibuffer- . . . . . . . . . . . . . . . . . . . . 281 properties in files, text. . . . . . . . . . . . . . . . . . . . . . . . 545 properties of strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 properties of text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 properties, add-text-. . . . . . . . . . . . . . . . . . . . . . . . . . 543 properties, default-text- . . . . . . . . . . . . . . . . . . . . . . 542 properties, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 properties, font- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 properties, font-instance- . . . . . . . . . . . . . . . . . . . . . 624 properties, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 properties, remove-text- . . . . . . . . . . . . . . . . . . . . . . 543 properties, saving text . . . . . . . . . . . . . . . . . . . . . . . . 545 properties, set-extent- . . . . . . . . . . . . . . . . . . . . . . . . 594 properties, set-frame- . . . . . . . . . . . . . . . . . . . . . . . . . 474 properties, set-text-. . . . . . . . . . . . . . . . . . . . . . . . . . . 543 properties, string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 properties, text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 properties-at, text- . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 property list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 property list cell (symbol) . . . . . . . . . . . . . . . . . . . . 113 property list, symbol . . . . . . . . . . . . . . . . . . . . . . . . . 118 property lists vs association lists . . . . . . . . . . . . . . 118 property of an extent . . . . . . . . . . . . . . . . . . . . . . . . . 593 property, charset-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 property, coding-system-. . . . . . . . . . . . . . . . . . . . . . 749 property, documentation- . . . . . . . . . . . . . . . . . . . . . 384 property, extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 property, extent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 property, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 620 property, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 property, get-char-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 property, get-text- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 property, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 629 property, mode-class . . . . . . . . . . . . . . . . . . . . . . . . . . 365 property, put-text- . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 property, remove-glyph- . . . . . . . . . . . . . . . . . . . . . . 630 property, set-extent- . . . . . . . . . . . . . . . . . . . . . . . . . . 593 property, set-face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 property, set-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . 474 property, set-glyph-. . . . . . . . . . . . . . . . . . . . . . . . . . . 628 property-any, text- . . . . . . . . . . . . . . . . . . . . . . . . . . . 545 property-change, next-. . . . . . . . . . . . . . . . . . . . . . . . 544 property-change, next-single- . . . . . . . . . . . . . . . . . 544 property-change, previous- . . . . . . . . . . . . . . . . . . . . 545 property-change, previous-single- . . . . . . . . . . . . . 545 property-instance, face- . . . . . . . . . . . . . . . . . . . . . . . 621 property-instance, glyph- . . . . . . . . . . . . . . . . . . . . . 629 property-not-all, text- . . . . . . . . . . . . . . . . . . . . . . . . 545 protect, unwind-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 protected forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 provide. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 providing features . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 ptys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 680 punctuation character . . . . . . . . . . . . . . . . . . . . . . . . 573 pure storage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 pure-bytes-used. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 purecopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 purify-flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 770 push-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 put. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 put, lax-plist-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 put, plist- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 put-char-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 put-database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 put-range-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 put-resource, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 717 put-text-property. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 putf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 puthash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667

Q

q, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 query, process-kill-without- . . . . . . . . . . . . . . . . . . . 681 query-functions, kill-buffer- . . . . . . . . . . . . . . . . . . . 444 query-functions, kill-emacs- . . . . . . . . . . . . . . . . . . . 698 query-p, process-kill-without- . . . . . . . . . . . . . . . . . 683 query-replace-history . . . . . . . . . . . . . . . . . . . . . . . . . 267 query-replace-map . . . . . . . . . . . . . . . . . . . . . . . 563, 784 querying the user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 question mark in character constant. . . . . . . . . . . . 23 questions, asking the user. . . . . . . . . . . . . . . . . . . . . 277 questions, yes-or-no. . . . . . . . . . . . . . . . . . . . . . . . . . . 277 queue, transaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 quietly-read-abbrev-file . . . . . . . . . . . . . . . . . . . . . . . 583 quit, debug-on- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 quit, edebug-on- . . . . . . . . . . . . . . . . . . . . . . . . . 236, 251 quit, inhibit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 quit, keyboard- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 quit-flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 quit-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 quitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 quitting from infinite loop . . . . . . . . . . . . . . . . . . . . 220 quitting, read-quoted-char . . . . . . . . . . . . . . . . . . . . 309 quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 quote character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 quote in strings, double- . . . . . . . . . . . . . . . . . . . . . . . 27 quote, character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 quote, regexp- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 quote, string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 573 quoted character input . . . . . . . . . . . . . . . . . . . . . . . 306 quoted-char quitting, read- . . . . . . . . . . . . . . . . . . . 309 quoted-char, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 quoted-insert suppression . . . . . . . . . . . . . . . . . . . . . 332 quoting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 quoting characters in printing. . . . . . . . . . . . . . . . . 258 quoting using apostrophe . . . . . . . . . . . . . . . . . . . . . 129 quoting, ’ for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 quoting, apostrophe for . . . . . . . . . . . . . . . . . . . . . . . 129 quoting, function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

R

r, C-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 raise-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 raise-frame, auto-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 raising a frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 random. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 random numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 range table type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Range Tables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-char-table, get- . . . . . . . . . . . . . . . . . . . . . . . . . . 77 range-table, clear- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-table, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-table, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-table, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-table, map-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-table, put- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-table, remove- . . . . . . . . . . . . . . . . . . . . . . . . . . 671 range-table-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 rassoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 rassq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 rate, device-baud- . . . . . . . . . . . . . . . . . . . . . . . 488, 711 rate, set-device-baud- . . . . . . . . . . . . . . . . . . . . 488, 712 raw prefix argument . . . . . . . . . . . . . . . . . . . . . . . . . . 310 raw prefix argument usage . . . . . . . . . . . . . . . . . . . . 287 re-search-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 re-search-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 read command name . . . . . . . . . . . . . . . . . . . . . . . . . 289 read syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 read syntax for characters . . . . . . . . . . . . . . . . . . . . . 21 read, cl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 read, completing- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 read-abbrev-file, quietly- . . . . . . . . . . . . . . . . . . . . . . 583 read-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 read-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 read-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 read-expression-history . . . . . . . . . . . . . . . . . . . . . . . 267 read-expression-map . . . . . . . . . . . . . . . . . . . . . . . . . . 785 read-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 read-from-minibuffer. . . . . . . . . . . . . . . . . . . . . . . . . . 264 read-from-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 read-function, load-. . . . . . . . . . . . . . . . . . . . . . . . . . . 199 read-key-sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 read-minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 read-no-blanks-input . . . . . . . . . . . . . . . . . . . . . . . . . 265 read-only buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 read-only buffers in interactive . . . . . . . . . . . . . . . . 285 read-only, barf-if-buffer- . . . . . . . . . . . . . . . . . . . . . . 440 read-only, buffer, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 read-only, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 read-only, find-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 read-only, inhibit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 read-only, toggle- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 read-quoted-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 read-quoted-char quitting . . . . . . . . . . . . . . . . . . . . . 309 read-shell-command-map . . . . . . . . . . . . . . . . . . . . . 785 read-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 read-syntax, invalid- . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 read-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 readable-p, file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 readably, print- . . . . . . . . . . . . . . . . . . . . . . . . . . 239, 261 readably, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 reader, Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 reading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 reading (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 reading interactive arguments. . . . . . . . . . . . . . . . . 286 reading symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 reading), stream (for . . . . . . . . . . . . . . . . . . . . . . . . . 253 reading, arguments, . . . . . . . . . . . . . . . . . . . . . . . . . . 263 reading, control characters, . . . . . . . . . . . . . . . . . . . 306 reading, nonprinting characters,. . . . . . . . . . . . . . . 306 real-login-name, user-. . . . . . . . . . . . . . . . . . . . . . . . . 703 real-uid, user- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 rearrangement of lists. . . . . . . . . . . . . . . . . . . . . . . . . . 89 rebinding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 receiving ToolTalk messages . . . . . . . . . . . . . . . . . . 724 recent-auto-save-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 recent-keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 recent-keys-ring-size . . . . . . . . . . . . . . . . . . . . . . . . . . 711 recent-keys-ring-size, set- . . . . . . . . . . . . . . . . . . . . . 711 recenter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 recompile-directory, batch-byte-. . . . . . . . . . . . . . . 209 recompile-directory, byte- . . . . . . . . . . . . . . . . . . . . . 209 recompile-directory-ignore-errors-p, byte- . . . . . 210 record command history . . . . . . . . . . . . . . . . . . . . . . 289 recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 recursion, infinite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 recursion-depth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 recursive command loop . . . . . . . . . . . . . . . . . . . . . . 312 recursive editing level. . . . . . . . . . . . . . . . . . . . . . . . . 312 recursive editing, exit . . . . . . . . . . . . . . . . . . . . . . . . . 312 recursive evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . 121 recursive, command loop, . . . . . . . . . . . . . . . . . . . . . 312 recursive-edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 recursive-edit, abort- . . . . . . . . . . . . . . . . . . . . . . . . . 313 recursive-edit, exit- . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 recursive-minibuffers, enable- . . . . . . . . . . . . . . . . . 282 redisplay, force-cursor- . . . . . . . . . . . . . . . . . . . . . . . . 649 redisplay, forcing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 redisplay, resize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 redisplay-function, set-extent-initial- . . . . . . . . . . 597 redo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 redraw-display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 redraw-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 redraw-modeline. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 374 redraw-on-reenter), resume (cf. no- . . . . . . . . . . . 649 redraw-on-reenter), suspend (cf. no- . . . . . . . . . . 649 redraw-on-reenter, no- . . . . . . . . . . . . . . . . . . . . . . . . 649 reenter), resume (cf. no-redraw-on- . . . . . . . . . . . 649 reenter), suspend (cf. no-redraw-on- . . . . . . . . . . 649 reenter, no-redraw-on- . . . . . . . . . . . . . . . . . . . . . . . . 649 refresh display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 regexp alternative . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 regexp grouping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 regexp searching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 regexp), character set (in . . . . . . . . . . . . . . . . . . . . . 554 regexp, $ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 regexp, ( in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 regexp, (?: in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 regexp, ) in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 556 regexp, * in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 regexp, *? in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 regexp, . in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 553 regexp, ? in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 regexp, [ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 regexp, ] in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 regexp, | in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 regexp, + in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 regexp, +? in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 regexp, ^in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 regexp, \ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 555 regexp, \’ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \‘ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \{n,m\} in . . . . . . . . . . . . . . . . . . . . . . . . . . . 554 regexp, \= in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \> in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \< in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \b in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \B in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \s in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 regexp, \S in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, \w in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 regexp, \W in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 557 regexp, beginning of line in . . . . . . . . . . . . . . . . . . . 555 regexp, defun-prompt- . . . . . . . . . . . . . . . . . . . . . . . . 496 regexp, invalid- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexp, searching for. . . . . . . . . . . . . . . . . . . . . . . . . . 559 regexp-fields, sort- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 regexp-history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 regexp-quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 558 regexps used in editing, standard . . . . . . . . . . . . . 568 regexps used standardly in editing . . . . . . . . . . . . 568 regexps, log-message-ignore- . . . . . . . . . . . . . . . . . . 652 regexps, same-window- . . . . . . . . . . . . . . . . . . . . . . . 459 regexps, special-display- . . . . . . . . . . . . . . . . . . . . . . 458 region (Edebug), eval- . . . . . . . . . . . . . . . . . . . . . . . . 231 region argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 region, annotations-in-. . . . . . . . . . . . . . . . . . . . . . . . 646 region, call-process-. . . . . . . . . . . . . . . . . . . . . . . . . . . 678 region, capitalize-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 region, compose- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 region, decode-coding- . . . . . . . . . . . . . . . . . . . . . . . . 750 region, decompose- . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 region, delete- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 region, detect-coding-. . . . . . . . . . . . . . . . . . . . . . . . . 750 region, downcase-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 region, encode-coding- . . . . . . . . . . . . . . . . . . . . . . . . 750 region, eval- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 region, fill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 region, find-charset- . . . . . . . . . . . . . . . . . . . . . . . . . . 742 region, indent-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 537 region, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 region, lines in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 493 region, narrow-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 region, process-send-. . . . . . . . . . . . . . . . . . . . . . . . . . 683 region, subst-char-in- . . . . . . . . . . . . . . . . . . . . . . . . . 546 region, the . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 region, translate- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 region, upcase- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 region, write- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 region, zmacs-activate- . . . . . . . . . . . . . . . . . . . . . . . 510 region, zmacs-deactivate- . . . . . . . . . . . . . . . . . . . . . 510 region, zmacs-update- . . . . . . . . . . . . . . . . . . . . . . . . 510 region-active-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 region-annotate-functions, write- . . . . . . . . . . . . . . 545 region-as-kill, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . 522 region-as-paragraph, fill- . . . . . . . . . . . . . . . . . . . . . . 529 region-beginning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 region-end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 region-exists-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 region-function, indent- . . . . . . . . . . . . . . . . . . . . . . . 537 region-hook, zmacs-activate- . . . . . . . . . . . . . . . . . . 510 region-hook, zmacs-deactivate- . . . . . . . . . . . . . . . . 511 region-hook, zmacs-update- . . . . . . . . . . . . . . . . . . . 511 region-p, extent-in- . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 region-stays, zmacs- . . . . . . . . . . . . . . . . . . . . . . . . . . 510 regions, transpose-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 regions, zmacs- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 register, address field of. . . . . . . . . . . . . . . . . . . . . . . . 24 register, decrement field of . . . . . . . . . . . . . . . . . . . . . 24 register, get-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 register, insert- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 register, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 register, view- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 register-alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 register-ccl-program . . . . . . . . . . . . . . . . . . . . . . . . . . 756 register-tooltalk-pattern . . . . . . . . . . . . . . . . . . . . . . 725 registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 registry, charset-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 regular expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 regular expression searching . . . . . . . . . . . . . . . . . . 559 regular-p, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 reindent-then-newline-and-indent . . . . . . . . . . . . . 537 relabel-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 relative file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 411 relative, indent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 relative-maybe, indent- . . . . . . . . . . . . . . . . . . . . . . . 539 relative-name, file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 release-event-p, button- . . . . . . . . . . . . . . . . . . . . . . . 296 relocation, marker . . . . . . . . . . . . . . . . . . . . . . . . . . . . 501 remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 remassoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 remassq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 remhash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 remove-database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 remove-glyph-property. . . . . . . . . . . . . . . . . . . . . . . . 630 remove-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 remove-range-table . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 remove-specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 remove-text-properties . . . . . . . . . . . . . . . . . . . . . . . . 543 remprop, lax-plist- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 remprop, plist- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 remrassoc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 remrassq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 rename-auto-save-file . . . . . . . . . . . . . . . . . . . . . . . . . 430 rename-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 rename-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 renaming files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406 repeated loading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 repetition, kill command. . . . . . . . . . . . . . . . . . . . . . 291 replace bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 331 replace characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 replace, case- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 replace, perform- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 replace-buffer-in-windows . . . . . . . . . . . . . . . . . . . . . 457 replace-history, query- . . . . . . . . . . . . . . . . . . . . . . . . 267 replace-map, query-. . . . . . . . . . . . . . . . . . . . . . 563, 784 replace-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 replacement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 562 replacement, & in . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 replacement, \ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 replacement, \n in . . . . . . . . . . . . . . . . . . . . . . . . . . . . 566 replacements, case in . . . . . . . . . . . . . . . . . . . . . . . . . 565 replica, extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 repositioning format arguments . . . . . . . . . . . . . . . . 71 representation for characters, printed . . . . . . . . . . 21 representation for lists, box . . . . . . . . . . . . . . . . . . . . 79 representation, printed. . . . . . . . . . . . . . . . . . . . . . . . . 17 represented as boxes, lists. . . . . . . . . . . . . . . . . . . . . . 79 require . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 require, byte-compiling . . . . . . . . . . . . . . . . . . . . . . . 203 require, load error with . . . . . . . . . . . . . . . . . . . . . . . 203 require-final-newline . . . . . . . . . . . . . . . . . . . . . . . . . . 398 requiring features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 reset-char-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 reset-elapsed-time, ccl-. . . . . . . . . . . . . . . . . . . . . . . . 756 resize redisplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 resizing, window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 resource type, X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 resource, x-get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 resource, x-put- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 717 rest arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 restriction (in a buffer) . . . . . . . . . . . . . . . . . . . . . . . 498 restriction, save- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 results, edebug-unwrap- . . . . . . . . . . . . . . . . . . 248, 251 resume (cf. no-redraw-on-reenter). . . . . . . . . . . . . 649 resume-hook, suspend-. . . . . . . . . . . . . . . . . . . . . . . . 699 return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 return-message, print-help- . . . . . . . . . . . . . . . . . . . 390 return-tooltalk-message . . . . . . . . . . . . . . . . . . . . . . . 722 reveal-annotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 reverse-direction-charset, charset- . . . . . . . . . . . . . 739 reverse-direction-charset, make- . . . . . . . . . . . . . . . 739 reversing a list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 revert-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 430 revert-buffer-function . . . . . . . . . . . . . . . . . . . . . . . . . 431 revert-buffer-insert-file-contents-function . . . . . . 431 revert-hook, after- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 revert-hook, before-. . . . . . . . . . . . . . . . . . . . . . . . . . . 431 rgb-components, color- . . . . . . . . . . . . . . . . . . . . . . . 626 rgb-components, color-instance-. . . . . . . . . . . . . . . 626 right, scroll- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 right-margin, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 right-margin-pixel-width, window- . . . . . . . . . . . . 647 right-margin-width . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 right-overflow, use- . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 right-toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 right-toolbar-visible-p . . . . . . . . . . . . . . . . . . . . . . . . 357 right-toolbar-width . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 rigidly, indent- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 rigidly, indent-code- . . . . . . . . . . . . . . . . . . . . . . . . . . 538 ring, global mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 ring, global-mark- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 ring, kill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 ring, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 ring, mark. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 506 ring, mark- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 508 ring-max, kill- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 ring-max, mark- . . . . . . . . . . . . . . . . . . . . . . . . . 508, 509 ring-size, recent-keys- . . . . . . . . . . . . . . . . . . . . . . . . . 711 ring-size, set-recent-keys- . . . . . . . . . . . . . . . . . . . . . 711 ring-yank-pointer, kill-. . . . . . . . . . . . . . . . . . . . . . . . 525 rm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 root, cube- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 root-window, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . 480 round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 rounding in conversions . . . . . . . . . . . . . . . . . . . . . . . . 51 rounding without conversion . . . . . . . . . . . . . . . . . . . 54 rplaca . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 rplaca vrs setcar, CL note— . . . . . . . . . . . . . . . . . . . 87 rplacd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 run time stack. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 run time stack, tag on . . . . . . . . . . . . . . . . . . . . . . . . 137 run-emacs-from-temacs . . . . . . . . . . . . . . . . . . . . . . . 768 run-file, site- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 run-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 runnable temacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 767

S

s, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 safe, car- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 safe, cdr- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 same-window-buffer-names. . . . . . . . . . . . . . . . . . . . 459 same-window-regexps . . . . . . . . . . . . . . . . . . . . . . . . . 459 sans-extension, file-name- . . . . . . . . . . . . . . . . . . . . . 409 sans-versions, file-name- . . . . . . . . . . . . . . . . . . . . . . 409 save, buffer-offer- . . . . . . . . . . . . . . . . . . . . . . . . 396, 444 save, do-auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 save-abbrevs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 save-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 save-current-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 save-default, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 save-displayed-buffer-points, edebug- . . . . . 242, 250 save-excursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 save-excursion (Edebug) . . . . . . . . . . . . . . . . . . . . . . 242 save-file, rename-auto- . . . . . . . . . . . . . . . . . . . . . . . . 430 save-file-format, auto- . . . . . . . . . . . . . . . . . . . . . . . . 421 save-file-if-necessary, delete-auto- . . . . . . . . . . . . . 429 save-file-name, buffer-auto- . . . . . . . . . . . . . . . . . . . 427 save-file-name, make-auto- . . . . . . . . . . . . . . . . . . . . 428 save-file-name-p, auto- . . . . . . . . . . . . . . . . . . . . . . . . 428 save-files, delete-auto- . . . . . . . . . . . . . . . . . . . . . . . . 430 save-hook, after-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 save-hook, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 save-interval, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 save-list-file-name, auto- . . . . . . . . . . . . . . . . . . . . . . 430 save-match-data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 save-mode, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 save-p, recent-auto-. . . . . . . . . . . . . . . . . . . . . . . . . . . 429 save-restriction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 save-selected-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 save-selected-window. . . . . . . . . . . . . . . . . . . . . 452, 498 save-some-buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 save-timeout, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 save-visited-file-name, auto-. . . . . . . . . . . . . . . . . . . 428 save-window-excursion . . . . . . . . . . . . . . . . . . . . . . . . 471 save-windows, edebug- . . . . . . . . . . . . . . . . . . . 242, 250 saved, set-buffer-auto- . . . . . . . . . . . . . . . . . . . . . . . . 429 saved-size, buffer- . . . . . . . . . . . . . . . . . . . . . . . . 430, 490 saving text properties. . . . . . . . . . . . . . . . . . . . . . . . . 545 saving window information . . . . . . . . . . . . . . . . . . . 470 saving, auto- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 427 scan-lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 scan-sexps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 scope. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 scoping, dynamic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 screen layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 screen-context-lines, next- . . . . . . . . . . . . . . . . . . . . 464 scroll-buffer, other-window- . . . . . . . . . . . . . . . . . . . 463 scroll-conservatively . . . . . . . . . . . . . . . . . . . . . . . . . . 463 scroll-down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 scroll-left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 scroll-other-window . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 scroll-right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 scroll-step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 scroll-up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 scroll-window, minibuffer- . . . . . . . . . . . . . . . . . . . . 282 scrollbar-pointer-glyph. . . . . . . . . . . . . . . . . . . . . . . . 641 scrollbars. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 scrolling vertically . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 scrolling, horizontal. . . . . . . . . . . . . . . . . . . . . . . . . . . 464 scrolling, vertical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 search, case-fold- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 568 search, default-case-fold- . . . . . . . . . . . . . . . . . . . . . . 568 search, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728 search, string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 search, word. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 search-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 search-backward, posix-. . . . . . . . . . . . . . . . . . . . . . . 562 search-backward, re- . . . . . . . . . . . . . . . . . . . . . . . . . . 560 search-backward, word- . . . . . . . . . . . . . . . . . . . . . . . 552 search-failed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 search-forward. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 search-forward, posix- . . . . . . . . . . . . . . . . . . . . . . . . 562 search-forward, re-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 search-forward, word-. . . . . . . . . . . . . . . . . . . . . . . . . 552 search-internal, ldap- . . . . . . . . . . . . . . . . . . . . . . . . . 730 search-path, x-library- . . . . . . . . . . . . . . . . . . . . . . . . 719 searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 searching and case . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 searching for regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 559 searching, regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 559 searching, regular expression . . . . . . . . . . . . . . . . . . 559 second. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 select-console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 select-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 select-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 select-frame-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 select-frame-hook, default- . . . . . . . . . . . . . . . . . . . . 483 select-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 selected frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 selected window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 selected-console. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 selected-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 selected-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 selected-frame, save- . . . . . . . . . . . . . . . . . . . . . . . . . . 481 selected-frame, with- . . . . . . . . . . . . . . . . . . . . . . . . . 481 selected-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 selected-window, frame- . . . . . . . . . . . . . . . . . . . . . . 480 selected-window, save- . . . . . . . . . . . . . . . . . . . 452, 498 selecting a buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 selecting windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 selection (for X windows) . . . . . . . . . . . . . . . . . . . . . 715 selection, x-disown- . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 selection, x-get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 selection, x-own-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 selection-hook, menu-no- . . . . . . . . . . . . . . . . . . . . . 344 selection-pointer-glyph. . . . . . . . . . . . . . . . . . . . . . . . 641 selective display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 selective-display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 selective-display-ellipses . . . . . . . . . . . . . . . . . . . . . . 657 self-evaluating form. . . . . . . . . . . . . . . . . . . . . . . . . . . 124 self-insert-and-exit. . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 self-insert-command . . . . . . . . . . . . . . . . . . . . . . . . . . 517 self-insert-command override. . . . . . . . . . . . . . . . . . 332 self-insert-command, minor modes . . . . . . . . . . . . 374 self-insertion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 517 send-eof, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 send-region, process-. . . . . . . . . . . . . . . . . . . . . . . . . . 683 send-string, process- . . . . . . . . . . . . . . . . . . . . . . . . . . 683 send-string-to-terminal . . . . . . . . . . . . . . . . . . . . . . . 712 send-tooltalk-message. . . . . . . . . . . . . . . . . . . . . . . . . 722 sendevents, x-allow- . . . . . . . . . . . . . . . . . . . . . . . . . . 719 sending signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 sending ToolTalk messages . . . . . . . . . . . . . . . . . . . 721 sentence-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 sentinel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 689 sentinel, process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 689 sentinel, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 sentinel, set-process- . . . . . . . . . . . . . . . . . . . . . . . . . . 690 separate, paragraph-. . . . . . . . . . . . . . . . . . . . . . . . . . 568 separator, path- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 sequence error, key . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 sequence input, key. . . . . . . . . . . . . . . . . . . . . . . . . . . 304 sequence length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 sequence, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 sequence, escape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 sequence, key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 sequence, read-key- . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 sequence, upper case key . . . . . . . . . . . . . . . . . . . . . 304 sequencep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 sequences, copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 sequences, elements of . . . . . . . . . . . . . . . . . . . . . . . . 105 sequences, key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 server-vendor, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 server-version, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 set (in a specifier), tag . . . . . . . . . . . . . . . . . . . . . . . . 602 set (in regexp), character . . . . . . . . . . . . . . . . . . . . . 554 set (input), Latin-1 character . . . . . . . . . . . . . . . . . 710 set local, CL note— . . . . . . . . . . . . . . . . . . . . . . . . . . 155 set, canonicalize-tag- . . . . . . . . . . . . . . . . . . . . . . . . . 610 set, CL note—lack union, . . . . . . . . . . . . . . . . . . . . . . 92 set, specifier, tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 set-annotation-action . . . . . . . . . . . . . . . . . . . . . . . . . 645 set-annotation-data. . . . . . . . . . . . . . . . . . . . . . . . . . . 645 set-annotation-down-glyph. . . . . . . . . . . . . . . . . . . . 645 set-annotation-face . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 set-annotation-glyph. . . . . . . . . . . . . . . . . . . . . . . . . . 645 set-annotation-layout . . . . . . . . . . . . . . . . . . . . . . . . . 645 set-annotation-menu . . . . . . . . . . . . . . . . . . . . . . . . . . 646 set-auto-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 set-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 435 set-buffer-auto-saved. . . . . . . . . . . . . . . . . . . . . . . . . . 429 set-buffer-major-mode . . . . . . . . . . . . . . . . . . . . . . . . 370 set-buffer-menubar . . . . . . . . . . . . . . . . . . . . . . . . . . . 343 set-buffer-modified-p. . . . . . . . . . . . . . . . . . . . . . . . . . 438 set-case-syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 set-case-syntax-delims . . . . . . . . . . . . . . . . . . . . . . . . . 75 set-case-syntax-pair. . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 set-case-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 set-category-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 set-charset-ccl-program . . . . . . . . . . . . . . . . . . . . . . . 740 set-coding-category-system . . . . . . . . . . . . . . . . . . . 750 set-coding-priority-list . . . . . . . . . . . . . . . . . . . . . . . . 750 set-console-type-image-conversion-list . . . . . . . . . 636 set-default. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 set-default-file-modes . . . . . . . . . . . . . . . . . . . . . . . . . 408 set-default-toolbar-position . . . . . . . . . . . . . . . . . . . 355 set-device-baud-rate . . . . . . . . . . . . . . . . . . . . . 488, 712 set-extent-begin-glyph . . . . . . . . . . . . . . . . . . . . . . . . 597 set-extent-begin-glyph-layout . . . . . . . . . . . . . . . . . 597 set-extent-end-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . 597 set-extent-end-glyph-layout . . . . . . . . . . . . . . . . . . . 597 set-extent-endpoints . . . . . . . . . . . . . . . . . . . . . . . . . . 589 set-extent-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 set-extent-initial-redisplay-function . . . . . . . . . . . 597 set-extent-keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 set-extent-mouse-face . . . . . . . . . . . . . . . . . . . . . . . . . 597 set-extent-parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 set-extent-priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 set-extent-properties . . . . . . . . . . . . . . . . . . . . . . . . . . 594 set-extent-property . . . . . . . . . . . . . . . . . . . . . . . . . . . 593 set-face-background . . . . . . . . . . . . . . . . . . . . . . . . . . 622 set-face-background-pixmap . . . . . . . . . . . . . . . . . . 622 set-face-font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 set-face-foreground . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 set-face-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 619 set-face-underline-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 set-file-modes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 407 set-frame-configuration . . . . . . . . . . . . . . . . . . . . . . . 483 set-frame-icon-pixmap, x- . . . . . . . . . . . . . . . . . . . . . 478 set-frame-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 set-frame-position . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 set-frame-properties . . . . . . . . . . . . . . . . . . . . . . . . . . 474 set-frame-property. . . . . . . . . . . . . . . . . . . . . . . . . . . . 474 set-frame-size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 set-global-break-condition, edebug- . . . . . . . . . . . 235 set-glyph-baseline . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 set-glyph-contrib-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 set-glyph-face . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 set-glyph-image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 631 set-glyph-property . . . . . . . . . . . . . . . . . . . . . . . . . . . . 628 set-input-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 set-key, global- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 set-key, local-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 set-keymap-default-binding . . . . . . . . . . . . . . . . . . . 319 set-keymap-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 set-keymap-parents . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 set-keymap-prompt . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 set-left-margin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 set-mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507 set-marker. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 505 set-match-data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 set-menubar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 343 set-menubar-dirty-flag . . . . . . . . . . . . . . . . . . . . . . . . 343 set-p, device-matches-specifier-tag-. . . . . . . . . . . . 611 set-p, valid-specifier-tag- . . . . . . . . . . . . . . . . . . . . . . 610 set-process-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 686 set-process-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 set-process-sentinel . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 set-process-window-size . . . . . . . . . . . . . . . . . . . . . . . 690 set-recent-keys-ring-size. . . . . . . . . . . . . . . . . . . . . . . 711 set-register . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 set-right-margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 530 set-specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 set-standard-case-table. . . . . . . . . . . . . . . . . . . . . . . . . 75 set-syntax-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577 set-text-properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . 543 set-tooltalk-message-attribute. . . . . . . . . . . . . . . . . 723 set-visited-file-modtime . . . . . . . . . . . . . . . . . . . . . . . 439 set-visited-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . 437 set-weak-list-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 set-window-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 set-window-buffer-dedicated . . . . . . . . . . . . . . . . . . 457 set-window-configuration . . . . . . . . . . . . . . . . . . . . . 470 set-window-dedicated-p . . . . . . . . . . . . . . . . . . . . . . . 460 set-window-hscroll . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 set-window-point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 set-window-start. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 setcar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 setcar, CL note—rplaca vrs . . . . . . . . . . . . . . . . . . . . 87 setcdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 setenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 701 setplist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 setprv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 setq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 setq-default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 sets, lists as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 setting modes of files . . . . . . . . . . . . . . . . . . . . . . . . . 406 setting-constant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 setup-hook, edebug- . . . . . . . . . . . . . . . . . . . . . . . . . . 249 setup-hook, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . 281 setup-hook, term- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 setup-hook, window-. . . . . . . . . . . . . . . . . . . . . . . . . . 696 seventh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 sexp motion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 sexp, backward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 sexp, forward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 sexp, parse-partial- . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 sexp-ignore-comments, parse- . . . . . . . . . . . . . . . . . 579 sexps, scan- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 shadowing of variables . . . . . . . . . . . . . . . . . . . . . . . . 148 shallow binding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 shared-lisp-mode-map . . . . . . . . . . . . . . . . . . . . . . . . 785 Shell mode modeline-format . . . . . . . . . . . . . . . . . . 376 shell-command, start-process-. . . . . . . . . . . . . . . . . 680 shell-command-history . . . . . . . . . . . . . . . . . . . . . . . . 267 shell-command-map, read- . . . . . . . . . . . . . . . . . . . . 785 shift, arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 shift, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 shift-jis-char, decode- . . . . . . . . . . . . . . . . . . . . . . . . . 751 shift-jis-char, encode- . . . . . . . . . . . . . . . . . . . . . . . . . 751 show-function, temp-buffer- . . . . . . . . . . . . . . . . . . . 659 show-keybindings, menubar- . . . . . . . . . . . . . . . . . . 343 shrink-window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 shrink-window-horizontally . . . . . . . . . . . . . . . . . . . 469 shrink-window-pixels . . . . . . . . . . . . . . . . . . . . . . . . . 469 side effect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 side, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645 signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 signal, debug-on- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 signal-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 signaling errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 signals, process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 signals, sending. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 684 sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 single-key-description . . . . . . . . . . . . . . . . . . . . . . . . . 388 single-property-change, next- . . . . . . . . . . . . . . . . . 544 single-property-change, previous- . . . . . . . . . . . . . 545 sinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 sit-for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 site-init.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 site-load.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 site-run-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 694 site-start.el . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 sixth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 size of frame. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 size of window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 size, buffer-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 490 size, buffer-saved- . . . . . . . . . . . . . . . . . . . . . . . . 430, 490 size, buffers-menu-max- . . . . . . . . . . . . . . . . . . . . . . . 349 size, changing window . . . . . . . . . . . . . . . . . . . . . . . . 468 size, changing, window. . . . . . . . . . . . . . . . . . . . . . . . 468 size, compiled-function-stack- . . . . . . . . . . . . . . . . . 213 size, font instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 size, frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 size, log-message-max- . . . . . . . . . . . . . . . . . . . . . . . . 652 size, max-specpdl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 size, minimum window. . . . . . . . . . . . . . . . . . . . . . . . 469 size, process window . . . . . . . . . . . . . . . . . . . . . . . . . . 690 size, recent-keys-ring- . . . . . . . . . . . . . . . . . . . . . . . . . 711 size, set-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 size, set-process-window- . . . . . . . . . . . . . . . . . . . . . 690 size, set-recent-keys-ring- . . . . . . . . . . . . . . . . . . . . . 711 size, window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 size, x-font- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 size-change-functions, window- . . . . . . . . . . . . . . . . 470 skip-chars-backward . . . . . . . . . . . . . . . . . . . . . . . . . . 497 skip-chars-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 skip-syntax-backward . . . . . . . . . . . . . . . . . . . . . . . . . 577 skip-syntax-forward. . . . . . . . . . . . . . . . . . . . . . . . . . . 577 skipping characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 496 skipping comments . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 sleep-for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 smaller-font, x-find- . . . . . . . . . . . . . . . . . . . . . . . . . . 624 Snarf-documentation . . . . . . . . . . . . . . . . . . . . . . . . . 386 soft, intern- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 some-buffers, save-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 sort, stable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 sort-columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 sort-fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 sort-lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 sort-numeric-fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 sort-pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 sort-paragraphs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534 sort-regexp-fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 sort-subr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532 sorting lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 sorting text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 sound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 sound, play- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 sound-alist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 sound-file, load- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 664 sound-file, play- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 sounds, load-default- . . . . . . . . . . . . . . . . . . . . . . . . . 664 source-newer, load-warn-when-. . . . . . . . . . . . . . . . 200 source-only, load-warn-when- . . . . . . . . . . . . . . . . . 200 space, delete-horizontal- . . . . . . . . . . . . . . . . . . . . . . 519 space, just-one- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 521 sparse-keymap, make- . . . . . . . . . . . . . . . . . . . . . . . . 318 SPC in minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 spec, buffer-invisibility- . . . . . . . . . . . . . . . . . . . . . . . 655 spec, canonicalize- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 608 spec, def-edebug- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 spec, initial-toolbar- . . . . . . . . . . . . . . . . . . . . . . . . . . 358 spec-list, canonicalize- . . . . . . . . . . . . . . . . . . . . . . . . 608 spec-list, check-valid- . . . . . . . . . . . . . . . . . . . . . . . . . 615 spec-list, specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 spec-list-p, valid- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 spec-list-to-specifier, add- . . . . . . . . . . . . . . . . . . . . . 606 spec-to-specifier, add-. . . . . . . . . . . . . . . . . . . . . . . . . 606 special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 special form descriptions . . . . . . . . . . . . . . . . . . . . . . . 12 special form evaluation . . . . . . . . . . . . . . . . . . . . . . . 127 special forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 special forms (Edebug) . . . . . . . . . . . . . . . . . . . . . . . 231 special forms compared, CL note— . . . . . . . . . . . 128 special forms for control structures. . . . . . . . . . . . 131 special variables, CL note— . . . . . . . . . . . . . . . . . . 156 special-display-buffer-names . . . . . . . . . . . . . . . . . . 458 special-display-frame-plist . . . . . . . . . . . . . . . . . . . . 459 special-display-function . . . . . . . . . . . . . . . . . . . . . . . 459 special-display-popup-frame . . . . . . . . . . . . . . . . . . 459 special-display-regexps. . . . . . . . . . . . . . . . . . . . . . . . 458 specific functions, debugging . . . . . . . . . . . . . . . . . . 220 specific initialization, terminal- . . . . . . . . . . . . . . . 695 specific-map, mode-. . . . . . . . . . . . . . . . . . . . . . 321, 784 specification (in a specifier) . . . . . . . . . . . . . . . . . . . 602 specification error, file mode . . . . . . . . . . . . . . . . . . 369 specification list, Edebug . . . . . . . . . . . . . . . . . . . . . 243 specification, format . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 specification, specifier, . . . . . . . . . . . . . . . . . . . . . . . . 602 specifications, indirect . . . . . . . . . . . . . . . . . . . . . . . . 246 specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 601 specifier type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 specifier), domain (in a . . . . . . . . . . . . . . . . . . . . . . . 602 specifier), fallback (in a . . . . . . . . . . . . . . . . . . . . . . . 603 specifier), inst-list (in a . . . . . . . . . . . . . . . . . . . . . . . 602 specifier), inst-pair (in a . . . . . . . . . . . . . . . . . . . . . . 602 specifier), instance (in a . . . . . . . . . . . . . . . . . . . . . . 602 specifier), instancing (in a . . . . . . . . . . . . . . . . . . . . 602 specifier), instantiator (in a . . . . . . . . . . . . . . . . . . . 602 specifier), locale (in a . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier), specification (in a . . . . . . . . . . . . . . . . . . 602 specifier), tag (in a . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier), tag set (in a . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, add-spec-list-to- . . . . . . . . . . . . . . . . . . . . . 606 specifier, add-spec-to-. . . . . . . . . . . . . . . . . . . . . . . . . 606 specifier, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 specifier, domain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, fallback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 603 specifier, inst-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, inst-pair . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, instance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, instancing. . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, instantiator . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, let- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 specifier, locale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613 specifier, make-image- . . . . . . . . . . . . . . . . . . . . . . . . 633 specifier, map-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 specifier, remove- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 specifier, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 607 specifier, specification. . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier, tag set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 specifier-and-init, make- . . . . . . . . . . . . . . . . . . . . . . 614 specifier-domain-p, valid- . . . . . . . . . . . . . . . . . . . . . 614 specifier-fallback. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 specifier-instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 specifier-instance-from-inst-list . . . . . . . . . . . . . . . . 612 specifier-locale-p, valid- . . . . . . . . . . . . . . . . . . . . . . . 614 specifier-locale-type-from-locale . . . . . . . . . . . . . . . 616 specifier-locale-type-p, valid- . . . . . . . . . . . . . . . . . . 614 specifier-p, boolean- . . . . . . . . . . . . . . . . . . . . . . . . . . 605 specifier-p, color- . . . . . . . . . . . . . . . . . . . . . . . . 605, 625 specifier-p, event-matches-key- . . . . . . . . . . . . . . . . 321 specifier-p, face-boolean-. . . . . . . . . . . . . . . . . . . . . . 605 specifier-p, font- . . . . . . . . . . . . . . . . . . . . . . . . . 605, 623 specifier-p, generic- . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 specifier-p, image- . . . . . . . . . . . . . . . . . . . . . . . 605, 633 specifier-p, integer- . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 specifier-p, natnum- . . . . . . . . . . . . . . . . . . . . . . . . . . 605 specifier-p, toolbar- . . . . . . . . . . . . . . . . . . . . . . 356, 605 specifier-spec-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 specifier-specs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 specifier-tag, define- . . . . . . . . . . . . . . . . . . . . . . . . . . 611 specifier-tag-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 specifier-tag-list, device-matching- . . . . . . . . . . . . 611 specifier-tag-p, valid-. . . . . . . . . . . . . . . . . . . . . 610, 614 specifier-tag-predicate . . . . . . . . . . . . . . . . . . . . . . . . 611 specifier-tag-set-p, device-matches-. . . . . . . . . . . . 611 specifier-tag-set-p, valid- . . . . . . . . . . . . . . . . . . . . . . 610 specifier-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 specifier-type-p, valid- . . . . . . . . . . . . . . . . . . . . . . . . 614 specifierp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 601 specifiers, image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 specpdl-size, max- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 specs, specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 609 specs.el, cl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 speed, execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 speedups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 splicing (with backquote) . . . . . . . . . . . . . . . . . . . . . 181 split-height-threshold . . . . . . . . . . . . . . . . . . . . . . . . . 458 split-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 split-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 split-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 split-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 split-window-horizontally . . . . . . . . . . . . . . . . . . . . . 450 split-window-vertically. . . . . . . . . . . . . . . . . . . . . . . . 450 splitting windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 splitting, window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 sqrt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 stable sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 stack frame, current . . . . . . . . . . . . . . . . . . . . . . . . . . 222 stack, call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 stack, run time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 stack, tag on run time . . . . . . . . . . . . . . . . . . . . . . . . 137 stack-size, compiled-function- . . . . . . . . . . . . . . . . . 213 standard notation, XEmacs event . . . . . . . . . . . . . 388 standard regexps used in editing . . . . . . . . . . . . . . 568 standard-case-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 standard-case-table, set- . . . . . . . . . . . . . . . . . . . . . . . 75 standard-category-table. . . . . . . . . . . . . . . . . . . . . . . 757 standard-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 standard-output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 standard-syntax-table . . . . . . . . . . . . . . . . . . . . . . . . 580 standardly in editing, regexps used . . . . . . . . . . . 568 standards of coding style . . . . . . . . . . . . . . . . . . . . . 759 standards, coding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 759 start position, extent . . . . . . . . . . . . . . . . . . . . . . . . . 588 start up of XEmacs. . . . . . . . . . . . . . . . . . . . . . . . . . . 693 start, paragraph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 start, set-window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 start, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 start-location, abbrev- . . . . . . . . . . . . . . . . . . . . . . . . 584 start-location-buffer, abbrev- . . . . . . . . . . . . . . . . . 584 start-position, extent- . . . . . . . . . . . . . . . . . . . . . . . . 589 start-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 679 start-process-shell-command . . . . . . . . . . . . . . . . . . 680 start.el, site-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 starter, comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 startup-echo-area-message, inhibit-. . . . . . . . . . . . 694 startup-message, inhibit-. . . . . . . . . . . . . . . . . . . . . . 694 startup.el. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 693 state, parse. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 status, command-debug- . . . . . . . . . . . . . . . . . . . . . . 227 status, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 682 status, process-exit- . . . . . . . . . . . . . . . . . . . . . . . . . . 683 stays, zmacs-region- . . . . . . . . . . . . . . . . . . . . . . . . . . 510 step, scroll-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 stop points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 stop, tab-to-tab-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 539 stop-list, tab-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 539 stop-process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 685 stopping an infinite loop . . . . . . . . . . . . . . . . . . . . . . 220 stopping on events. . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 stops for indentation, tabs . . . . . . . . . . . . . . . . . . . . 539 stops-map, edit-tab- . . . . . . . . . . . . . . . . . . . . . . . . . . 783 storage, CL note—allocate more . . . . . . . . . . . . . . 770 storage, pure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 store-cutbuffer, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 store-match-data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 567 stream (for printing). . . . . . . . . . . . . . . . . . . . . . . . . . 256 stream (for reading) . . . . . . . . . . . . . . . . . . . . . . . . . . 253 stream, buffer input . . . . . . . . . . . . . . . . . . . . . . . . . . 253 stream, buffer output . . . . . . . . . . . . . . . . . . . . . . . . . 256 stream, function input . . . . . . . . . . . . . . . . . . . . . . . . 254 stream, function output . . . . . . . . . . . . . . . . . . . . . . 257 stream, input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 stream, marker input . . . . . . . . . . . . . . . . . . . . . . . . . 253 stream, marker output. . . . . . . . . . . . . . . . . . . . . . . . 256 stream, nil input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 stream, nil output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 stream, open-network- . . . . . . . . . . . . . . . . . . . . . . . . 691 stream, output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 stream, string input . . . . . . . . . . . . . . . . . . . . . . . . . . 254 stream, t input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 stream, t output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 string equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 string formatting, multilingual . . . . . . . . . . . . . . . . . 71 string in keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 326 string input stream . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 string length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 string length, maximum when printing . . . . . . . . 261 string properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 string quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 573 string search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 551 string to character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 string to number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 string to object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 string, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 string, ccl-execute-on- . . . . . . . . . . . . . . . . . . . . . . . . 756 string, char-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 string, character to . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 string, charset-doc- . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 string, coding-system-doc- . . . . . . . . . . . . . . . . . . . . 749 string, compiled-function-doc- . . . . . . . . . . . . . . . . 213 string, composite-char- . . . . . . . . . . . . . . . . . . . . . . . 742 string, current-time- . . . . . . . . . . . . . . . . . . . . . . . . . . 704 string, default argument . . . . . . . . . . . . . . . . . . . . . . 286 string, find-charset- . . . . . . . . . . . . . . . . . . . . . . . . . . . 742 string, format-time- . . . . . . . . . . . . . . . . . . . . . . . . . . 705 string, global-mode- . . . . . . . . . . . . . . . . . . . . . . . . . . 377 string, image-instance-. . . . . . . . . . . . . . . . . . . . . . . . 639 string, insert-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 string, inside . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578 string, int-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 string, integer to. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 string, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 string, match- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 564 string, number-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 string, object to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 string, overlay-arrow- . . . . . . . . . . . . . . . . . . . . . . . . . 657 string, prin1-to- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 string, process-send- . . . . . . . . . . . . . . . . . . . . . . . . . . 683 string, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 string, read-from-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 string, split- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561 string, string, writing a doc . . . . . . . . . . . . . . . . . . . 383 string, writing a doc string . . . . . . . . . . . . . . . . . . . 383 string, writing a documentation. . . . . . . . . . . . . . . 383 string-display, momentary-. . . . . . . . . . . . . . . . . . . . 659 string-equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 string-length, print- . . . . . . . . . . . . . . . . . . . . . . . . . . 261 string-lessp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 string-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 string-match, posix- . . . . . . . . . . . . . . . . . . . . . . . . . . 562 string-modified-tick . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 string-p, char-or- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 string-to-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 string-to-int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 string-to-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 string-to-terminal, send- . . . . . . . . . . . . . . . . . . . . . . 712 string= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 string< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 stringp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 strings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 strings, \ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 strings, backslash in . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 strings, concatenating. . . . . . . . . . . . . . . . . . . . . . . . . . 63 strings, conversion of . . . . . . . . . . . . . . . . . . . . . . . . . . 67 strings, copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 strings, documentation . . . . . . . . . . . . . . . . . . . . . . . 383 strings, double-quote in . . . . . . . . . . . . . . . . . . . . . . . . 27 strings, formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 strings, formatting them . . . . . . . . . . . . . . . . . . . . . . . 69 strings, keys in documentation . . . . . . . . . . . . . . . . 386 strings, modifying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 strings, newline in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 strings, properties of . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 strong-limit, undo- . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 structure, list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 structures, control . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 structures, printing circular . . . . . . . . . . . . . . . . . . . 239 structures, special forms for control . . . . . . . . . . . 131 style, standards of coding . . . . . . . . . . . . . . . . . . . . . 759 subexpression, previous complete . . . . . . . . . . . . . 578 submenu, add- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 subprocess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 675 subprocess, asynchronous . . . . . . . . . . . . . . . . . . . . . 679 subprocess, synchronous . . . . . . . . . . . . . . . . . . . . . . 676 subprocesses, environment variables, . . . . . . . . . . 676 subr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 subr, sort- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 532 subroutines, file name completion . . . . . . . . . . . . . 413 subrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 subsidiary-coding-system . . . . . . . . . . . . . . . . . . . . . 749 subst-char-in-region . . . . . . . . . . . . . . . . . . . . . . . . . . 546 substitute-command-keys . . . . . . . . . . . . . . . . . . . . . 387 substitute-in-file-name . . . . . . . . . . . . . . . . . . . . . . . . 412 substitute-key-definition . . . . . . . . . . . . . . . . . . . . . . 331 substituting keys in documentation . . . . . . . . . . . 386 substitution), ‘ (list. . . . . . . . . . . . . . . . . . . . . . . . . . . 181 substitution), backquote (list . . . . . . . . . . . . . . . . . 181 substring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 substring, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 514 substring, insert-buffer- . . . . . . . . . . . . . . . . . . . . . . . 516 substrings, compare-buffer- . . . . . . . . . . . . . . . . . . . 515 subtype, database- . . . . . . . . . . . . . . . . . . . . . . . . . . . 674 subwindow type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 subwindow-image-instance-p . . . . . . . . . . . . . . . . . . 638 subwindowp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 supersession, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 supersession-threat, ask-user-about-. . . . . . . . . . . 439 suppress-keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 suppressed-classes, display-warning-. . . . . . . . . . . 654 suppressed-classes, log-warning-. . . . . . . . . . . . . . . 654 suppression, quoted-insert . . . . . . . . . . . . . . . . . . . . 332 suppression, yank. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 suspend (cf. no-redraw-on-reenter) . . . . . . . . . . . . 649 suspend evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 suspend-emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 698 suspend-hook. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 699 suspend-resume-hook . . . . . . . . . . . . . . . . . . . . . . . . . 699 suspending XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . 698 switch-alist, command- . . . . . . . . . . . . . . . . . . . . . . . 696 switch-to-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 456 switch-to-buffer-function, buffers-menu- . . . . . . . 350 switch-to-buffer-other-window . . . . . . . . . . . . . . . . 456 switches on command line . . . . . . . . . . . . . . . . . . . . 696 switching to a buffer. . . . . . . . . . . . . . . . . . . . . . . . . . 455 symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 symbol components. . . . . . . . . . . . . . . . . . . . . . . . . . . 113 symbol constituent . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 symbol equality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 symbol evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 symbol function indirection . . . . . . . . . . . . . . . . . . . 125 symbol in keymap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327 symbol in modeline, percent . . . . . . . . . . . . . . . . . . 375 symbol in obarrays, CL note—. . . . . . . . . . . . . . . . 116 symbol name hashing . . . . . . . . . . . . . . . . . . . . . . . . . 115 symbol, abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 symbol, definition of a . . . . . . . . . . . . . . . . . . . . . . . . 114 symbol, error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 symbol, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 symbol, plist, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 symbol, property list, . . . . . . . . . . . . . . . . . . . . . . . . . 118 symbol, uninterned . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 symbol-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 symbol-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 symbol-plist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 symbol-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 symbolic links, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 symbolic-link, make-. . . . . . . . . . . . . . . . . . . . . . . . . . 407 symbolp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 symbols, \ in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 symbols, backslash in . . . . . . . . . . . . . . . . . . . . . . . . . . 23 symbols, printing uninterned. . . . . . . . . . . . . . . . . . 261 symbols, printing, uninterned . . . . . . . . . . . . . . . . . 261 symbols, reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 symbols, xpm-color- . . . . . . . . . . . . . . . . . . . . . . . . . . 636 symlink-p, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 402 synchronous subprocess. . . . . . . . . . . . . . . . . . . . . . . 676 syntax classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 syntax descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 syntax error (Edebug) . . . . . . . . . . . . . . . . . . . . . . . . 246 syntax flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 syntax for characters . . . . . . . . . . . . . . . . . . . . . . . . . . 21 syntax for characters, read . . . . . . . . . . . . . . . . . . . . . 21 syntax table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 syntax table example . . . . . . . . . . . . . . . . . . . . . . . . . 366 syntax table internals. . . . . . . . . . . . . . . . . . . . . . . . . 580 syntax tables in modes . . . . . . . . . . . . . . . . . . . . . . . 364 syntax, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 576 syntax, check-toolbar-button- . . . . . . . . . . . . . . . . . 355 syntax, comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 574 syntax, invalid-read- . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 syntax, iso- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 syntax, parenthesis . . . . . . . . . . . . . . . . . . . . . . . . . . . 573 syntax, read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 syntax, set-case- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 syntax-backward, skip- . . . . . . . . . . . . . . . . . . . . . . . 577 syntax-delims, set-case- . . . . . . . . . . . . . . . . . . . . . . . . 75 syntax-entry, modify- . . . . . . . . . . . . . . . . . . . . . . . . . 576 syntax-forward, skip- . . . . . . . . . . . . . . . . . . . . . . . . . 577 syntax-pair, set-case- . . . . . . . . . . . . . . . . . . . . . . . . . . 75 syntax-table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577 syntax-table, c-mode-. . . . . . . . . . . . . . . . . . . . . . . . . 580 syntax-table, copy- . . . . . . . . . . . . . . . . . . . . . . . . . . . 576 syntax-table, emacs-lisp-mode-. . . . . . . . . . . . . . . . 580 syntax-table, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . 575 syntax-table, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577 syntax-table, standard- . . . . . . . . . . . . . . . . . . . . . . . 580 syntax-table, text-mode- . . . . . . . . . . . . . . . . . . . . . . 580 syntax-table-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 system environment, operating. . . . . . . . . . . . . . . . 699 system objects, window- . . . . . . . . . . . . . . . . . . . . . . 617 system type, coding . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 system types, window. . . . . . . . . . . . . . . . . . . . . . . . . . 36 system, coding-category-. . . . . . . . . . . . . . . . . . . . . . 750 system, copy-coding- . . . . . . . . . . . . . . . . . . . . . . . . . 749 system, find-coding- . . . . . . . . . . . . . . . . . . . . . . . . . . 749 system, get-coding- . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 system, make-coding- . . . . . . . . . . . . . . . . . . . . . . . . . 749 system, set-coding-category- . . . . . . . . . . . . . . . . . . 750 system, subsidiary-coding- . . . . . . . . . . . . . . . . . . . . 749 system-configuration. . . . . . . . . . . . . . . . . . . . . . . . . . 700 system-doc-string, coding- . . . . . . . . . . . . . . . . . . . . 749 system-list, coding- . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 system-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 system-name, coding-. . . . . . . . . . . . . . . . . . . . . . . . . 749 system-p, coding-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 746 system-property, coding-. . . . . . . . . . . . . . . . . . . . . . 749 system-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 system-type, coding-. . . . . . . . . . . . . . . . . . . . . . . . . . 749

T

t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 t and truth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 t input stream. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 t output stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 tab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 tab deletion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519 TAB in minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 tab, vertical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 tab-command, indent-for-. . . . . . . . . . . . . . . . . . . . . 537 tab-stop, tab-to-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 539 tab-stop-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 539 tab-stops-map, edit- . . . . . . . . . . . . . . . . . . . . . . . . . . 783 tab-to-tab-stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 539 tab-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 table example, syntax . . . . . . . . . . . . . . . . . . . . . . . . 366 table internals, syntax . . . . . . . . . . . . . . . . . . . . . . . . 580 table type, char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 table type, hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 table type, range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 table, abbrev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 table, active display . . . . . . . . . . . . . . . . . . . . . . . . . . 662 table, c-mode-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . 585 table, c-mode-syntax-. . . . . . . . . . . . . . . . . . . . . . . . . 580 table, category-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 table, clear-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 table, clear-range- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table, copy-category- . . . . . . . . . . . . . . . . . . . . . . . . . 757 table, copy-range- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table, copy-syntax- . . . . . . . . . . . . . . . . . . . . . . . . . . . 576 table, current-case- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 table, define-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . 582 table, describe-buffer-case- . . . . . . . . . . . . . . . . . . . . . 75 table, display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 table, emacs-lisp-mode-syntax-. . . . . . . . . . . . . . . . 580 table, fundamental-mode-abbrev- . . . . . . . . . . . . . 585 table, get-char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 table, get-range- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table, get-range-char- . . . . . . . . . . . . . . . . . . . . . . . . . . 77 table, global-abbrev-. . . . . . . . . . . . . . . . . . . . . . . . . . 585 table, hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 667 table, lisp-mode-abbrev- . . . . . . . . . . . . . . . . . . . . . . 585 table, local-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 table, make-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . 581 table, make-char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 table, make-display- . . . . . . . . . . . . . . . . . . . . . . . . . . 662 table, make-range-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table, make-syntax-. . . . . . . . . . . . . . . . . . . . . . . . . . . 575 table, map-char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 table, map-range-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table, minibuffer-completion- . . . . . . . . . . . . . . . . . 272 table, put-char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 table, put-range- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table, remove-range- . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table, reset-char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 table, set-case-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 table, set-category- . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 table, set-standard-case- . . . . . . . . . . . . . . . . . . . . . . . 75 table, set-syntax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577 table, standard-case-. . . . . . . . . . . . . . . . . . . . . . . . . . . 75 table, standard-category- . . . . . . . . . . . . . . . . . . . . . 757 table, standard-syntax- . . . . . . . . . . . . . . . . . . . . . . . 580 table, syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 table, syntax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 577 table, text-mode-abbrev- . . . . . . . . . . . . . . . . . . . . . 585 table, text-mode-syntax- . . . . . . . . . . . . . . . . . . . . . . 580 table, weak hash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 table, weak, hash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 table-description, insert-abbrev- . . . . . . . . . . . . . . 582 table-name-list, abbrev- . . . . . . . . . . . . . . . . . . . . . . 582 table-p, case- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 table-p, category-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 757 table-p, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 table-p, range-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 table-p, syntax- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 table-type, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 table-type-list, char- . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 table-type-p, valid-char- . . . . . . . . . . . . . . . . . . . . . . . 76 table-value, check-valid-char-. . . . . . . . . . . . . . . . . . . 77 table-value-p, category- . . . . . . . . . . . . . . . . . . . . . . . 757 table-value-p, valid-char-. . . . . . . . . . . . . . . . . . . . . . . 77 tables in modes, abbrev. . . . . . . . . . . . . . . . . . . . . . . 364 tables in modes, syntax . . . . . . . . . . . . . . . . . . . . . . . 364 Tables, Range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 tabs stops for indentation. . . . . . . . . . . . . . . . . . . . . 539 tabs-mode, indent- . . . . . . . . . . . . . . . . . . . . . . . . . . . 536 tag (in a specifier) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 tag on run time stack. . . . . . . . . . . . . . . . . . . . . . . . . 137 tag set (in a specifier) . . . . . . . . . . . . . . . . . . . . . . . . 602 tag set, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 tag, define-specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . 611 tag, specifier, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 tag-list, device-matching-specifier- . . . . . . . . . . . . 611 tag-list, specifier- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 611 tag-p, valid-specifier-. . . . . . . . . . . . . . . . . . . . . 610, 614 tag-predicate, specifier- . . . . . . . . . . . . . . . . . . . . . . . 611 tag-set, canonicalize- . . . . . . . . . . . . . . . . . . . . . . . . . 610 tag-set-p, device-matches-specifier-. . . . . . . . . . . . 611 tag-set-p, valid-specifier- . . . . . . . . . . . . . . . . . . . . . . 610 tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 tanh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 TCP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 temacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 767 temacs, bootstrapping XEmacs from . . . . . . . . . . 767 temacs, run-emacs-from- . . . . . . . . . . . . . . . . . . . . . . 768 temacs, runnable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 767 temp-buffer, with-output-to- . . . . . . . . . . . . . . . . . . 658 temp-buffer-show-function . . . . . . . . . . . . . . . . . . . . 659 temp-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 temp-file, with- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 temp-name, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 413 tenth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 TERM environment variable. . . . . . . . . . . . . . . . . . 696 term-file-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 term-setup-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 Termcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 terminal frame. . . . . . . . . . . . . . . . . . . . . . . . . . . 448, 473 terminal input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 terminal input modes . . . . . . . . . . . . . . . . . . . . . . . . . 708 terminal output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711 terminal, frame of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 terminal, send-string-to- . . . . . . . . . . . . . . . . . . . . . . 712 terminal-device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 terminal-local-map, overriding- . . . . . . . . . . . . . . . 325 terminal-specific initialization . . . . . . . . . . . . . . . . . 695 terminate keyboard macro . . . . . . . . . . . . . . . . . . . . 307 termination, keyboard macro . . . . . . . . . . . . . . . . . 663 termscript file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 termscript, open- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 712 terpri . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 test-coverage, edebug- . . . . . . . . . . . . . . . . . . . . . . . . 250 testing types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 testing, coverage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 text changes, hooks for . . . . . . . . . . . . . . . . . . . . . . . 548 text files and binary files. . . . . . . . . . . . . . . . . . . . . . 421 text files, binary files and . . . . . . . . . . . . . . . . . . . . . 421 text insertion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 text notation, buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 text parsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 571 text properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 text properties in files . . . . . . . . . . . . . . . . . . . . . . . . 545 text properties, saving . . . . . . . . . . . . . . . . . . . . . . . . 545 text, attributes of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 text, comparing buffer . . . . . . . . . . . . . . . . . . . . . . . . 515 text, find-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422 text, inserting killed . . . . . . . . . . . . . . . . . . . . . . . . . . 523 text, insertion of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 515 text, invisible. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 text, last-abbrev- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 584 text, properties of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 text, sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 531 text-area-p, event-over- . . . . . . . . . . . . . . . . . . . . . . . 298 text-area-pixel-edges, window- . . . . . . . . . . . . . . . . 468 text-area-pixel-height, window- . . . . . . . . . . . . . . . 467 text-area-pixel-width, window- . . . . . . . . . . . . . . . . 467 text-char-description. . . . . . . . . . . . . . . . . . . . . . . . . . 388 text-domain, bind- . . . . . . . . . . . . . . . . . . . . . . . . . . . 731 text-glyph, invisible-. . . . . . . . . . . . . . . . . . . . . . . . . . 642 text-image-instance-p . . . . . . . . . . . . . . . . . . . . . . . . . 637 text-mode-abbrev-table . . . . . . . . . . . . . . . . . . . . . . . 585 text-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 text-mode-syntax-table . . . . . . . . . . . . . . . . . . . . . . . 580 text-pixel-height, window-displayed- . . . . . . . . . . 467 text-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 text-properties, add-. . . . . . . . . . . . . . . . . . . . . . . . . . 543 text-properties, default- . . . . . . . . . . . . . . . . . . . . . . 542 text-properties, remove- . . . . . . . . . . . . . . . . . . . . . . 543 text-properties, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . 543 text-properties-at. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 text-property, get- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 text-property, put- . . . . . . . . . . . . . . . . . . . . . . . . . . . 542 text-property-any. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 545 text-property-not-all. . . . . . . . . . . . . . . . . . . . . . . . . . 545 than-file-p, file-newer- . . . . . . . . . . . . . . . . . . . . . . . . 402 then-newline-and-indent, reindent- . . . . . . . . . . . . 537 third . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 this-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 this-command-keys . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 threat, ask-user-about-supersession-. . . . . . . . . . . 439 threshold, gc-cons-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 773 threshold, split-height-. . . . . . . . . . . . . . . . . . . . . . . . 458 throw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 throw example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 throw in Emacs, CL note—only. . . . . . . . . . . . . . . 137 tick, buffer-modified- . . . . . . . . . . . . . . . . . . . . . . . . . 439 tick, string-modified- . . . . . . . . . . . . . . . . . . . . . . . . . . 69 tiled windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 time stack, run . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 time stack, tag on run . . . . . . . . . . . . . . . . . . . . . . . . 137 time, ccl-elapsed- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 756 time, ccl-reset-elapsed-. . . . . . . . . . . . . . . . . . . . . . . . 756 time, comparison of modification. . . . . . . . . . . . . . 439 time, comparison of, modification . . . . . . . . . . . . . 439 time, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 time, decode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 706 time, emacs-build- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 time, encode- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 time, file modification . . . . . . . . . . . . . . . . . . . . . . . . 402 time-string, current- . . . . . . . . . . . . . . . . . . . . . . . . . . 704 time-string, format- . . . . . . . . . . . . . . . . . . . . . . . . . . 705 time-zone, current- . . . . . . . . . . . . . . . . . . . . . . . . . . . 704 timeout, add-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 707 timeout, auto-save- . . . . . . . . . . . . . . . . . . . . . . . . . . . 429 timeout, disable- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708 timeout-event-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 timestamp, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 timing programs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 759 title-format, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 title-format, frame-icon- . . . . . . . . . . . . . . . . . . . . . . 478 titles, popup-menu-. . . . . . . . . . . . . . . . . . . . . . . . . . . 347 toggle-read-only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 353 toolbar button type. . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 toolbar, bottom-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar, default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 355 toolbar, left-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar, right- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar, top- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar-button, event- . . . . . . . . . . . . . . . . . . . . . . . . 299 toolbar-button-syntax, check- . . . . . . . . . . . . . . . . . 355 toolbar-buttons-captioned-p . . . . . . . . . . . . . . . . . . 357 toolbar-height, bottom- . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar-height, default- . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar-height, top-. . . . . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar-make-button-list . . . . . . . . . . . . . . . . . . . . . . 354 toolbar-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325, 785 toolbar-p, event-over- . . . . . . . . . . . . . . . . . . . . . . . . . 299 toolbar-pointer-glyph . . . . . . . . . . . . . . . . . . . . . . . . . 641 toolbar-position, default- . . . . . . . . . . . . . . . . . . . . . 355 toolbar-position, set-default- . . . . . . . . . . . . . . . . . . 355 toolbar-spec, initial- . . . . . . . . . . . . . . . . . . . . . . . . . . 358 toolbar-specifier-p . . . . . . . . . . . . . . . . . . . . . . . 356, 605 toolbar-visible-p, bottom-. . . . . . . . . . . . . . . . . . . . . 357 toolbar-visible-p, default- . . . . . . . . . . . . . . . . . . . . . 357 toolbar-visible-p, left-. . . . . . . . . . . . . . . . . . . . . . . . . 357 toolbar-visible-p, right- . . . . . . . . . . . . . . . . . . . . . . . 357 toolbar-visible-p, top-. . . . . . . . . . . . . . . . . . . . . . . . . 357 toolbar-width, default-. . . . . . . . . . . . . . . . . . . . . . . . 356 toolbar-width, left- . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 toolbar-width, right-. . . . . . . . . . . . . . . . . . . . . . . . . . 357 ToolTalk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 ToolTalk message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 721 ToolTalk messages, receiving. . . . . . . . . . . . . . . . . . 724 ToolTalk messages, sending . . . . . . . . . . . . . . . . . . . 721 ToolTalk pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 724 tooltalk-message, create-. . . . . . . . . . . . . . . . . . . . . . 723 tooltalk-message, describe-. . . . . . . . . . . . . . . . . . . . 725 tooltalk-message, destroy- . . . . . . . . . . . . . . . . . . . . 724 tooltalk-message, make- . . . . . . . . . . . . . . . . . . . . . . 722 tooltalk-message, return- . . . . . . . . . . . . . . . . . . . . . 722 tooltalk-message, send- . . . . . . . . . . . . . . . . . . . . . . . 722 tooltalk-message-arg, add- . . . . . . . . . . . . . . . . . . . . 723 tooltalk-message-attribute, get- . . . . . . . . . . . . . . . 722 tooltalk-message-attribute, set- . . . . . . . . . . . . . . . 723 tooltalk-pattern, create- . . . . . . . . . . . . . . . . . . . . . . 725 tooltalk-pattern, destroy- . . . . . . . . . . . . . . . . . . . . . 725 tooltalk-pattern, make- . . . . . . . . . . . . . . . . . . . . . . . 724 tooltalk-pattern, register- . . . . . . . . . . . . . . . . . . . . . 725 tooltalk-pattern, unregister- . . . . . . . . . . . . . . . . . . 725 tooltalk-pattern-arg, add-. . . . . . . . . . . . . . . . . . . . . 725 tooltalk-pattern-attribute, add- . . . . . . . . . . . . . . . 725 top line, window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 top-level. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 top-level form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 top-level-form, edebug-eval-. . . . . . . . . . . . . . . . . . . 231 top-toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 top-toolbar-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . 356 top-toolbar-visible-p . . . . . . . . . . . . . . . . . . . . . . . . . . 357 top-window, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 totally-visible-p, frame- . . . . . . . . . . . . . . . . . . . . . . . 482 tq-close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 tq-create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 tq-enqueue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 trace, edebug- . . . . . . . . . . . . . . . . . . . . . . . . . . . 240, 250 trace-after, edebug-print- . . . . . . . . . . . . . . . . 240, 250 trace-before, edebug-print- . . . . . . . . . . . . . . . 240, 250 tracing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 tracing, edebug- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 transaction queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . 690 transcendental functions . . . . . . . . . . . . . . . . . . . . . . . 58 transl, iso-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 translate-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 547 translating input events . . . . . . . . . . . . . . . . . . . . . . 709 translation function, key . . . . . . . . . . . . . . . . . . . . . . 710 translation-map, key- . . . . . . . . . . . . . . . . . . . . . . . . . 710 transpose-regions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 trim-versions-without-asking . . . . . . . . . . . . . . . . . . 426 true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 truename (of file) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 truename, buffer-file- . . . . . . . . . . . . . . . . . . . . . . . . . 437 truename, file-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 403 truename, font-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 truename, font-instance- . . . . . . . . . . . . . . . . . . . . . . 624 truncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 truncate-lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 truncate-lines, default-. . . . . . . . . . . . . . . . . . . . . . . . 650 truncate-partial-width-windows . . . . . . . . . . . . . . . 650 truncation-glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 truth value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 truth, t and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 try-completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 tty-device, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 tty-name, process- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 two’s complement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 type checking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, buffer-file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 type, char table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 type, char-table- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 type, charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 type, coding system . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 type, coding-system-. . . . . . . . . . . . . . . . . . . . . . . . . . 749 type, color instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 type, database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 type, database- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 674 type, default-buffer-file-. . . . . . . . . . . . . . . . . . . . . . . 422 type, device- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486 type, device-or-frame- . . . . . . . . . . . . . . . . . . . . . . . . 486 type, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 type, face. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 type, find-buffer-file-. . . . . . . . . . . . . . . . . . . . . . . . . . 421 type, font instance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 type, hash table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 type, image instance . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, image-instance- . . . . . . . . . . . . . . . . . . . . . . . . . 637 type, primitive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 type, process-connection- . . . . . . . . . . . . . . . . . . . . . 680 type, range table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 type, specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, specifier-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 605 type, subwindow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, system- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 type, toolbar button . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type, weak list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 type, weak-list- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 type, X resource . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 type-alist, file-name-buffer-file-. . . . . . . . . . . . . . . . 421 type-argument, wrong-. . . . . . . . . . . . . . . . . . . . . . . . . 37 type-from-locale, specifier-locale-. . . . . . . . . . . . . . 616 type-image-conversion-list, console- . . . . . . . . . . . 636 type-image-conversion-list, set-console- . . . . . . . . 636 type-list, char-table- . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 type-list, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 type-list, image-instance- . . . . . . . . . . . . . . . . . . . . . 637 type-of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 type-p, valid-char-table- . . . . . . . . . . . . . . . . . . . . . . . 76 type-p, valid-device- . . . . . . . . . . . . . . . . . . . . . . . . . . 487 type-p, valid-glyph-. . . . . . . . . . . . . . . . . . . . . . . . . . . 640 type-p, valid-image-instance-. . . . . . . . . . . . . . . . . . 637 type-p, valid-specifier- . . . . . . . . . . . . . . . . . . . . . . . . 614 type-p, valid-specifier-locale- . . . . . . . . . . . . . . . . . . 614 types on MS-DOS, file . . . . . . . . . . . . . . . . . . . . . . . . 421 types, editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 types, image instance . . . . . . . . . . . . . . . . . . . . . . . . . 637 types, layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 types, MS-DOS file . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 types, primitive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 types, programming . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 types, testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 types, window system. . . . . . . . . . . . . . . . . . . . . . . . . . 36

U

uid, user-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 uid, user-real- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 unbinding keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 unbold, x-make-font- . . . . . . . . . . . . . . . . . . . . . . . . . 624 undefined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 328 undefined in keymap . . . . . . . . . . . . . . . . . . . . . . . . . 327 undefined key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 underline-p, face- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 underline-p, set-face- . . . . . . . . . . . . . . . . . . . . . . . . . 622 undo avoidance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 undo, buffer-disable- . . . . . . . . . . . . . . . . . . . . . . . . . . 527 undo, buffer-enable- . . . . . . . . . . . . . . . . . . . . . . . . . . 527 undo, buffer-flush- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 undo, disable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 undo, primitive- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 undo-boundary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 undo-limit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 undo-list, buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 undo-strong-limit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 527 unexec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768 ungrab-keyboard, x- . . . . . . . . . . . . . . . . . . . . . . . . . . 718 ungrab-pointer, x- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 unhandled-file-name-directory . . . . . . . . . . . . . . . . 418 unintern. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 uninterned symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 uninterned symbols, printing. . . . . . . . . . . . . . . . . . 261 union, set, CL note—lack . . . . . . . . . . . . . . . . . . . . . . 92 unique extents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 unique, extent, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598 unitalic, x-make-font- . . . . . . . . . . . . . . . . . . . . . . . . . 625 universal-argument . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 unload-feature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 unloading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 unlock-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 400 unmap-frame-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 unread-command-event . . . . . . . . . . . . . . . . . . . . . . . 307 unread-command-events . . . . . . . . . . . . . . . . . . . . . . 306 unreading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 unreading, prefix argument . . . . . . . . . . . . . . . . . . . 306 unregister-tooltalk-pattern . . . . . . . . . . . . . . . . . . . . 725 unset-key, global- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 unset-key, local- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 untabify, backward-delete-char- . . . . . . . . . . . . . . . 519 unwind-protect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 unwinding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 unwrap, edebug-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 unwrap-results, edebug- . . . . . . . . . . . . . . . . . . 248, 251 up menu, pop-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346 up of XEmacs, start . . . . . . . . . . . . . . . . . . . . . . . . . . 693 up, buffer-backed- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423 up, scroll- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 463 up-frame-function, pop- . . . . . . . . . . . . . . . . . . . . . . 458 up-frame-plist, pop- . . . . . . . . . . . . . . . . . . . . . . . . . . 458 up-frames, pop- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 up-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 up-p, popup-menu- . . . . . . . . . . . . . . . . . . . . . . . . . . . 347 up-windows, pop-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 upcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 upcase-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 540 upcase-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 update display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 update, display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 649 update-directory-autoloads . . . . . . . . . . . . . . . . . . . 201 update-file-autoloads . . . . . . . . . . . . . . . . . . . . . . . . . 201 update-region, zmacs- . . . . . . . . . . . . . . . . . . . . . . . . 510 update-region-hook, zmacs- . . . . . . . . . . . . . . . . . . . 511 upper case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 upper case key sequence . . . . . . . . . . . . . . . . . . . . . . 304 usage, numeric prefix argument . . . . . . . . . . . . . . . 287 usage, raw prefix argument . . . . . . . . . . . . . . . . . . . 287 use, debug-on-error . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 use, noninteractive . . . . . . . . . . . . . . . . . . . . . . . . . . . 714 use-global-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 use-hard-newlines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 use-left-overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 use-local-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 use-right-overflow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 used in editing, standard regexps . . . . . . . . . . . . . 568 used standardly in editing, regexps . . . . . . . . . . . 568 used, pure-bytes- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 769 user option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 user questions, asking the. . . . . . . . . . . . . . . . . . . . . 277 user, querying the . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 user-about-lock, ask- . . . . . . . . . . . . . . . . . . . . . . . . . 400 user-about-supersession-threat, ask-. . . . . . . . . . . 439 user-defined error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 user-event-p, misc- . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 user-full-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 user-home-directory . . . . . . . . . . . . . . . . . . . . . . . . . . 704 user-input-p, waiting-for- . . . . . . . . . . . . . . . . . . . . . 690 user-login-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 user-mail-address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 user-real-login-name . . . . . . . . . . . . . . . . . . . . . . . . . . 703 user-real-uid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 user-uid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 user-variable-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 user-variable-p example. . . . . . . . . . . . . . . . . . . . . . . 274 uses of, nil, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 using apostrophe, quoting . . . . . . . . . . . . . . . . . . . . 129 using interactive, examples of . . . . . . . . . . . . . . . . . 288 using, interactive, examples of . . . . . . . . . . . . . . . . 288

V

valid-char-table-type-p . . . . . . . . . . . . . . . . . . . . . . . . . 76 valid-char-table-value, check-. . . . . . . . . . . . . . . . . . . 77 valid-char-table-value-p . . . . . . . . . . . . . . . . . . . . . . . . 77 valid-device-class-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 valid-device-type-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 valid-glyph-type-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640 valid-image-instance-type-p . . . . . . . . . . . . . . . . . . . 637 valid-image-instantiator-format-p . . . . . . . . . . . . . 636 valid-inst-list, check-. . . . . . . . . . . . . . . . . . . . . . . . . . 615 valid-inst-list-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 valid-instantiator, check-. . . . . . . . . . . . . . . . . . . . . . 614 valid-instantiator-p . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 valid-keysym-name-p, x- . . . . . . . . . . . . . . . . . . . . . . 719 valid-plist, check- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 valid-plist-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 valid-spec-list, check- . . . . . . . . . . . . . . . . . . . . . . . . . 615 valid-spec-list-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 614 valid-specifier-domain-p. . . . . . . . . . . . . . . . . . . . . . . 614 valid-specifier-locale-p . . . . . . . . . . . . . . . . . . . . . . . . 614 valid-specifier-locale-type-p . . . . . . . . . . . . . . . . . . . 614 valid-specifier-tag-p . . . . . . . . . . . . . . . . . . . . . . 610, 614 valid-specifier-tag-set-p . . . . . . . . . . . . . . . . . . . . . . . 610 valid-specifier-type-p . . . . . . . . . . . . . . . . . . . . . . . . . 614 value cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 value of expression . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 value, check-valid-char-table-. . . . . . . . . . . . . . . . . . . 77 value, default. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 value, default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 value, prefix-numeric-. . . . . . . . . . . . . . . . . . . . . . . . . 311 value, symbol- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 value, truth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 value-p, category-table- . . . . . . . . . . . . . . . . . . . . . . . 757 value-p, valid-char-table-. . . . . . . . . . . . . . . . . . . . . . . 77 value-weak-hashtable, make- . . . . . . . . . . . . . . . . . . 669 values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 variable access, environment . . . . . . . . . . . . . . . . . . 701 variable aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 variable definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 variable descriptions . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 variable limit error . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 variable, EMACSLOADPATH environment … 198 variable, global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 variable, HOME environment . . . . . . . . . . . . . . . . . 675 variable, indirect-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 variable, kill-local- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 variable, make-local-. . . . . . . . . . . . . . . . . . . . . . . . . . 159 variable, make-obsolete- . . . . . . . . . . . . . . . . . . . . . . 391 variable, mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 variable, PATH environment . . . . . . . . . . . . . . . . . . 675 variable, permanent local . . . . . . . . . . . . . . . . . . . . . 161 variable, read- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 variable, TERM environment . . . . . . . . . . . . . . . . . 696 variable, void . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 variable, void- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 variable-alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 variable-alias, define-obsolete-. . . . . . . . . . . . . . . . . 392 variable-buffer-local, make- . . . . . . . . . . . . . . . . . . . 160 variable-documentation . . . . . . . . . . . . . . . . . . . . . . . 383 variable-obsoleteness-doc . . . . . . . . . . . . . . . . . . . . . 392 variable-p example, user- . . . . . . . . . . . . . . . . . . . . . 274 variable-p, local-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 variable-p, user- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 variables in modes, buffer-local . . . . . . . . . . . . . . . 364 variables, aliases, for. . . . . . . . . . . . . . . . . . . . . . . . . . 163 variables, binding local . . . . . . . . . . . . . . . . . . . . . . . 148 variables, buffer-local . . . . . . . . . . . . . . . . . . . . . . . . . 159 variables, buffer-local- . . . . . . . . . . . . . . . . . . . . . . . . 160 variables, CL note—special . . . . . . . . . . . . . . . . . . . 156 variables, enable-local-. . . . . . . . . . . . . . . . . . . . . . . . 369 variables, hack-local- . . . . . . . . . . . . . . . . . . . . . . . . . 371 variables, ignored-local-. . . . . . . . . . . . . . . . . . . . . . . 369 variables, indirect . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 variables, kill-all-local- . . . . . . . . . . . . . . . . . . . . . . . . 161 variables, local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 variables, shadowing of . . . . . . . . . . . . . . . . . . . . . . . 148 variables, subprocesses, environment . . . . . . . . . . 676 varying-indent, fill-individual-. . . . . . . . . . . . . . . . . 529 vc-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 vconcat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108, 109 vector evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 vector length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 vector length, bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 vector, bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 vector, bit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 vector, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 vector, make-bit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 vector-p, bit- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 vectorp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 vectors, copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 vectors, copying bit. . . . . . . . . . . . . . . . . . . . . . . . . . . 111 vendor, x-server- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 verify-visited-file-modtime . . . . . . . . . . . . . . . . . . . . 439 version number (in file name) . . . . . . . . . . . . . . . . . 408 version, emacs- . . . . . . . . . . . . . . . . . . . . . . . . . . 768, 769 version, emacs-major- . . . . . . . . . . . . . . . . . . . . . . . . 769 version, emacs-minor-. . . . . . . . . . . . . . . . . . . . . . . . . 769 version, x-server- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 version-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 versions, dired-kept- . . . . . . . . . . . . . . . . . . . . . . . . . . 426 versions, file-name-sans- . . . . . . . . . . . . . . . . . . . . . . 409 versions, kept-new- . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 versions, kept-old- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 versions-without-asking, trim- . . . . . . . . . . . . . . . . 426 vertical scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 vertical tab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 vertical-motion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 494 vertical-motion-pixels . . . . . . . . . . . . . . . . . . . . . . . . . 494 vertically, scrolling. . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 vertically, split-window- . . . . . . . . . . . . . . . . . . . . . . 450 view-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 view-mode-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 785 view-register . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 548 visibility, frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 visible frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 visible, annotation- . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 visible, make-frame- . . . . . . . . . . . . . . . . . . . . . . . . . . 482 visible-bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 663 visible-frame-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 visible-in-window-p, pos- . . . . . . . . . . . . . . . . . . . . . 462 visible-p, bottom-toolbar-. . . . . . . . . . . . . . . . . . . . . 357 visible-p, default-toolbar- . . . . . . . . . . . . . . . . . . . . . 357 visible-p, frame- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 482 visible-p, frame-totally- . . . . . . . . . . . . . . . . . . . . . . . 482 visible-p, left-toolbar-. . . . . . . . . . . . . . . . . . . . . . . . . 357 visible-p, right-toolbar- . . . . . . . . . . . . . . . . . . . . . . . 357 visible-p, top-toolbar-. . . . . . . . . . . . . . . . . . . . . . . . . 357 visited file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 visited file mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 370 visited-file-modtime . . . . . . . . . . . . . . . . . . . . . . . . . . 439 visited-file-modtime, clear-. . . . . . . . . . . . . . . . . . . . 439 visited-file-modtime, set-. . . . . . . . . . . . . . . . . . . . . . 439 visited-file-modtime, verify- . . . . . . . . . . . . . . . . . . . 439 visited-file-name, auto-save-. . . . . . . . . . . . . . . . . . . 428 visited-file-name, set- . . . . . . . . . . . . . . . . . . . . . . . . . 437 visiting files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 visual-class, x-display- . . . . . . . . . . . . . . . . . . . . . . . . 718 void function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 void function cell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 void variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 void-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 void-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 volume, bell- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 664 vrs eq, CL note—integers . . . . . . . . . . . . . . . . . . . . . . 49 vrs setcar, CL note—rplaca . . . . . . . . . . . . . . . . . . . . 87 vs association lists, property lists . . . . . . . . . . . . . 118 vs killing, deletion . . . . . . . . . . . . . . . . . . . . . . . . . . . . 518 vs. extents, markers . . . . . . . . . . . . . . . . . . . . . . . . . . 501

W

waiting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 waiting for command key input . . . . . . . . . . . . . . . 307 waiting-for-user-input-p. . . . . . . . . . . . . . . . . . . . . . . 690 wakeup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 676 walk-windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 warn-when-source-newer, load-. . . . . . . . . . . . . . . . 200 warn-when-source-only, load- . . . . . . . . . . . . . . . . . 200 warning, display- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 warning-minimum-level, display- . . . . . . . . . . . . . . 654 warning-minimum-level, log- . . . . . . . . . . . . . . . . . . 654 warning-suppressed-classes, display-. . . . . . . . . . . 654 warning-suppressed-classes, log-. . . . . . . . . . . . . . . 654 weak hash table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 weak list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 weak list type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 weak, hash table, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 668 weak-hashtable, make- . . . . . . . . . . . . . . . . . . . . . . . . 668 weak-hashtable, make-key-. . . . . . . . . . . . . . . . . . . . 668 weak-hashtable, make-value- . . . . . . . . . . . . . . . . . . 669 weak-list, make- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 weak-list-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 weak-list-list, set-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 weak-list-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 weak-list-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 when printing, string length, maximum . . . . . . . 261 when-compile, eval-. . . . . . . . . . . . . . . . . . . . . . . . . . . 212 when-linked, backup-by-copying-. . . . . . . . . . . . . . 425 when-mismatch, backup-by-copying- . . . . . . . . . . 425 when-source-newer, load-warn-. . . . . . . . . . . . . . . . 200 when-source-only, load-warn- . . . . . . . . . . . . . . . . . 200 where-is-internal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 whitespace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 whitespace character . . . . . . . . . . . . . . . . . . . . . . . . . 572 whitespace, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . 519 whitespace, fixup- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 520 widen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 widening . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 width, annotation-. . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 width, default-toolbar-. . . . . . . . . . . . . . . . . . . . . . . . 356 width, field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 width, frame-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 width, frame-pixel- . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 width, glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 632 width, image-instance-. . . . . . . . . . . . . . . . . . . . . . . . 639 width, left-margin- . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 width, left-toolbar- . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 width, margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 646 width, minibuffer-prompt- . . . . . . . . . . . . . . . . . . . . 281 width, right-margin- . . . . . . . . . . . . . . . . . . . . . . . . . . 646 width, right-toolbar-. . . . . . . . . . . . . . . . . . . . . . . . . . 357 width, tab- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661 width, window- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 width, window-left-margin-pixel- . . . . . . . . . . . . . . 647 width, window-min- . . . . . . . . . . . . . . . . . . . . . . . . . . 470 width, window-pixel- . . . . . . . . . . . . . . . . . . . . . . . . . 467 width, window-right-margin-pixel- . . . . . . . . . . . . 647 width, window-text-area-pixel- . . . . . . . . . . . . . . . . 467 width-windows, truncate-partial-. . . . . . . . . . . . . . 650 window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 window configuration (Edebug) . . . . . . . . . . . . . . . 242 window configurations . . . . . . . . . . . . . . . . . . . . . . . . 470 window excursions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 497 window frame, X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 window information, saving . . . . . . . . . . . . . . . . . . . 470 window ordering, cyclic . . . . . . . . . . . . . . . . . . . . . . . 453 window point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 window position . . . . . . . . . . . . . . . . . . . . . . . . . 460, 467 window resizing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 window size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 window size, changing . . . . . . . . . . . . . . . . . . . . . . . . 468 window size, minimum. . . . . . . . . . . . . . . . . . . . . . . . 469 window size, process . . . . . . . . . . . . . . . . . . . . . . . . . . 690 window splitting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 window system types . . . . . . . . . . . . . . . . . . . . . . . . . . 36 window top line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 window, active-minibuffer- . . . . . . . . . . . . . . . . . . . . 281 window, dedicated . . . . . . . . . . . . . . . . . . . . . . . 457, 459 window, delete-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 window, enlarge- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 window, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 window, find-file-other- . . . . . . . . . . . . . . . . . . . . . . . 394 window, frame-root- . . . . . . . . . . . . . . . . . . . . . . . . . . 480 window, frame-selected- . . . . . . . . . . . . . . . . . . . . . . 480 window, frame-top- . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 window, get-buffer- . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 window, get-largest- . . . . . . . . . . . . . . . . . . . . . . . . . . 453 window, get-lru- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 window, minibuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 window, minibuffer- . . . . . . . . . . . . . . . . . . . . . . . . . . 281 window, minibuffer-scroll- . . . . . . . . . . . . . . . . . . . . 282 window, next- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 window, other- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 window, point in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 window, position in. . . . . . . . . . . . . . . . . . . . . . . . . . . 460 window, position of . . . . . . . . . . . . . . . . . . . . . . . . . . . 467 window, previous- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 window, save-selected- . . . . . . . . . . . . . . . . . . . 452, 498 window, scroll-other- . . . . . . . . . . . . . . . . . . . . . . . . . 463 window, select- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 window, selected . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 window, selected- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 window, shrink- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 window, size of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 window, split- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 window, switch-to-buffer-other- . . . . . . . . . . . . . . . 456 window-active-p, minibuffer- . . . . . . . . . . . . . . . . . . 282 window-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 window-buffer, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . 455 window-buffer-dedicated, set- . . . . . . . . . . . . . . . . . 457 window-buffer-names, same- . . . . . . . . . . . . . . . . . . 459 window-configuration, current-. . . . . . . . . . . . . . . . 470 window-configuration, set- . . . . . . . . . . . . . . . . . . . . 470 window-configuration-p . . . . . . . . . . . . . . . . . . . . . . . 471 window-dedicated-p. . . . . . . . . . . . . . . . . . . . . . 457, 459 window-dedicated-p, set- . . . . . . . . . . . . . . . . . . . . . 460 window-displayed-text-pixel-height . . . . . . . . . . . . 467 window-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 window-excursion, save- . . . . . . . . . . . . . . . . . . . . . . 471 window-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 window-height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 window-highest-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 window-horizontally, enlarge- . . . . . . . . . . . . . . . . . 469 window-horizontally, shrink- . . . . . . . . . . . . . . . . . . 469 window-horizontally, split- . . . . . . . . . . . . . . . . . . . . 450 window-hscroll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 window-hscroll, set- . . . . . . . . . . . . . . . . . . . . . . . . . . 465 window-id, x-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 window-left-margin-pixel-width . . . . . . . . . . . . . . . 647 window-line, move-to- . . . . . . . . . . . . . . . . . . . . . . . . 494 window-live-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 window-lowest-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 window-min-height . . . . . . . . . . . . . . . . . . . . . . . . . . . 469 window-min-width. . . . . . . . . . . . . . . . . . . . . . . . . . . . 470 window-minibuffer-p. . . . . . . . . . . . . . . . . . . . . . . . . . 281 window-p, one- . . . . . . . . . . . . . . . . . . . . . . . . . . 448, 450 window-p, pos-visible-in- . . . . . . . . . . . . . . . . . . . . . 462 window-pixel-edges . . . . . . . . . . . . . . . . . . . . . . . . . . . 468 window-pixel-height . . . . . . . . . . . . . . . . . . . . . . . . . . 466 window-pixel-width. . . . . . . . . . . . . . . . . . . . . . . . . . . 467 window-pixels, enlarge- . . . . . . . . . . . . . . . . . . . . . . . 469 window-pixels, shrink- . . . . . . . . . . . . . . . . . . . . . . . . 469 window-point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 window-point, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 window-regexps, same- . . . . . . . . . . . . . . . . . . . . . . . 459 window-right-margin-pixel-width. . . . . . . . . . . . . . 647 window-scroll-buffer, other- . . . . . . . . . . . . . . . . . . . 463 window-setup-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . 696 window-size, set-process- . . . . . . . . . . . . . . . . . . . . . 690 window-size-change-functions . . . . . . . . . . . . . . . . . 470 window-start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 window-start, set- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 461 window-system objects . . . . . . . . . . . . . . . . . . . . . . . 617 window-text-area-pixel-edges . . . . . . . . . . . . . . . . . 468 window-text-area-pixel-height. . . . . . . . . . . . . . . . . 467 window-text-area-pixel-width . . . . . . . . . . . . . . . . . 467 window-vertically, split- . . . . . . . . . . . . . . . . . . . . . . 450 window-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 466 window-x-pixel, event- . . . . . . . . . . . . . . . . . . . . . . . . 297 window-y-pixel, event- . . . . . . . . . . . . . . . . . . . . . . . . 297 windowp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 windows), selection (for X . . . . . . . . . . . . . . . . . . . . 715 windows, buffers, controlled in . . . . . . . . . . . . . . . . 455 windows, controlling precisely. . . . . . . . . . . . . . . . . 455 windows, cyclic ordering of . . . . . . . . . . . . . . . . . . . 453 windows, cyclic, ordering of . . . . . . . . . . . . . . . . . . . 453 windows, delete-other- . . . . . . . . . . . . . . . . . . . . . . . . 451 windows, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 451 windows, edebug-save- . . . . . . . . . . . . . . . . . . . 242, 250 windows, examining . . . . . . . . . . . . . . . . . . . . . . . . . . 455 windows, finding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 windows, multiple . . . . . . . . . . . . . . . . . . . . . . . . . . . . 447 windows, pop-up-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 windows, replace-buffer-in-. . . . . . . . . . . . . . . . . . . . 457 windows, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 452 windows, splitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 windows, tiled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 448 windows, truncate-partial-width-. . . . . . . . . . . . . . 650 windows, walk- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 454 Windows, X- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 windows-on, delete-. . . . . . . . . . . . . . . . . . . . . . . . . . . 451 with multiple names, file. . . . . . . . . . . . . . . . . . . . . . 406 with narrowing, point . . . . . . . . . . . . . . . . . . . . . . . . 489 with parentheses, indenting . . . . . . . . . . . . . . . . . . . 578 with prefix argument, execute . . . . . . . . . . . . . . . . 289 with require, load error . . . . . . . . . . . . . . . . . . . . . . . 203 with-current-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 with-output-to-temp-buffer . . . . . . . . . . . . . . . . . . . 658 with-selected-frame . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 with-temp-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 498 without conversion, rounding . . . . . . . . . . . . . . . . . . 54 without-asking, trim-versions- . . . . . . . . . . . . . . . . 426 without-query, process-kill- . . . . . . . . . . . . . . . . . . . 681 without-query-p, process-kill- . . . . . . . . . . . . . . . . . 683 word constituent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572 word search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 552 word, backward-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 word, capitalize- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 word, downcase- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 word, forward- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 491 word, minibuffer-complete- . . . . . . . . . . . . . . . . . . . 272 word, upcase- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 541 word-search-backward . . . . . . . . . . . . . . . . . . . . . . . . 552 word-search-forward . . . . . . . . . . . . . . . . . . . . . . . . . . 552 words-include-escapes. . . . . . . . . . . . . . . . . . . . . . . . . 491 wrapping, line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 650 writable-p, file- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 write-abbrev-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583 write-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 write-contents-hooks. . . . . . . . . . . . . . . . . . . . . . . . . . 397 write-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 write-file, format-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 420 write-file-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 write-file-hooks, local- . . . . . . . . . . . . . . . . . . . . . . . . 397 write-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399 write-region-annotate-functions . . . . . . . . . . . . . . . 545 writing a doc string, string, . . . . . . . . . . . . . . . . . . . 383 writing a documentation string . . . . . . . . . . . . . . . 383 writing minor modes, conventions for . . . . . . . . . 373 wrong-number-of-arguments . . . . . . . . . . . . . . . . . . 167 wrong-type-argument . . . . . . . . . . . . . . . . . . . . . . . . . . 37

X

X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 x 4, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 x 5, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 x a, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 x n, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 x r, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 X resource type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 X window frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 473 X windows), selection (for . . . . . . . . . . . . . . . . . . . . 715 x, C- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 x, C-M- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 x, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 x, image-instance-hotspot- . . . . . . . . . . . . . . . . . . . . 639 x, M-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 x-4-map, ctl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 x-5-map, ctl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 x-allow-sendevents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 x-bitmap-file-path . . . . . . . . . . . . . . . . . . . . . . . 636, 719 x-debug-events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 x-debug-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 x-device, default- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 x-device, make-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 x-disown-selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 x-display, device- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488 x-display-visual-class . . . . . . . . . . . . . . . . . . . . . . . . . 718 x-emacs-application-class . . . . . . . . . . . . . . . . . . . . . 717 x-find-larger-font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 x-find-smaller-font . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 x-font-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 x-get-cutbuffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 x-get-resource . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716 x-get-selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 x-grab-keyboard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 x-grab-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 x-library-search-path . . . . . . . . . . . . . . . . . . . . . . . . . 719 x-make-font-bold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 x-make-font-bold-italic. . . . . . . . . . . . . . . . . . . . . . . . 625 x-make-font-italic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 x-make-font-unbold . . . . . . . . . . . . . . . . . . . . . . . . . . . 624 x-make-font-unitalic . . . . . . . . . . . . . . . . . . . . . . . . . . 625 x-map, ctl- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321, 783 x-own-selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 x-pixel, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 x-pixel, event-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 x-pixel, event-window- . . . . . . . . . . . . . . . . . . . . . . . . 297 X-prefix, Control- . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 x-put-resource . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 717 x-server-vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 x-server-version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 x-set-frame-icon-pixmap . . . . . . . . . . . . . . . . . . . . . . 478 x-store-cutbuffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 x-ungrab-keyboard . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 x-ungrab-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 718 x-valid-keysym-name-p . . . . . . . . . . . . . . . . . . . . . . . 719 x-window-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 X-Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 715 XEmacs event standard notation . . . . . . . . . . . . . 388 XEmacs from temacs, bootstrapping . . . . . . . . . . 767 XEmacs, building . . . . . . . . . . . . . . . . . . . . . . . . . . . . 767 XEmacs, exiting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 XEmacs, killing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 697 XEmacs, start up of . . . . . . . . . . . . . . . . . . . . . . . . . . 693 XEmacs, suspending. . . . . . . . . . . . . . . . . . . . . . . . . . 698 xpm-color-symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . 636

Y

y, event-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 y, image-instance-hotspot- . . . . . . . . . . . . . . . . . . . . 639 y-or-n-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 y-or-n-p, map-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 y-or-n-p-maybe-dialog-box . . . . . . . . . . . . . . . . . . . . 279 y-pixel, event- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 y-pixel, event-glyph- . . . . . . . . . . . . . . . . . . . . . . . . . . 299 y-pixel, event-window- . . . . . . . . . . . . . . . . . . . . . . . . 297 yank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 yank suppression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 yank-pointer, kill-ring-. . . . . . . . . . . . . . . . . . . . . . . . 525 yank-pop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 yes-or-no questions . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 yes-or-no-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 yes-or-no-p-dialog-box . . . . . . . . . . . . . . . . . . . . . . . . 279 yes-or-no-p-maybe-dialog-box . . . . . . . . . . . . . . . . . 279

Z

zero-length extent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 588 zerop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 zmacs-activate-region . . . . . . . . . . . . . . . . . . . . . . . . . 510 zmacs-activate-region-hook . . . . . . . . . . . . . . . . . . . 510 zmacs-deactivate-region. . . . . . . . . . . . . . . . . . . . . . . 510 zmacs-deactivate-region-hook . . . . . . . . . . . . . . . . . 511 zmacs-region-stays. . . . . . . . . . . . . . . . . . . . . . . . . . . . 510 zmacs-regions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 509 zmacs-update-region . . . . . . . . . . . . . . . . . . . . . . . . . . 510 zmacs-update-region-hook . . . . . . . . . . . . . . . . . . . . 511 zone, current-time- . . . . . . . . . . . . . . . . . . . . . . . . . . . 704

Short Contents

GNU GENERAL PUBLIC LICENSE . . . . . . . . . . . . . . . . . . . . 1 1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2 Lisp Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3 Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4 Strings and Characters . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5 Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 6 Sequences, Arrays, and Vectors . . . . . . . . . . . . . . . . . . . . 101 7 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 8 Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 9 Control Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 10 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 11 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 12 Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 13 Writing Customization Definitions . . . . . . . . . . . . . . . . . . 185 14 Loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 15 Byte Compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 16 Debugging Lisp Programs. . . . . . . . . . . . . . . . . . . . . . . . 217 17 Reading and Printing Lisp Objects . . . . . . . . . . . . . . . . . 251 18 Minibuffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 19 Command Loop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 20 Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 21 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 22 Dialog Boxes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349 23 Toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 24 scrollbars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 25 Drag and Drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 26 Major and Minor Modes. . . . . . . . . . . . . . . . . . . . . . . . . 361 27 Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 28 Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 29 Backups and Auto-Saving. . . . . . . . . . . . . . . . . . . . . . . . 421 30 Buffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 31 Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445 32 Frames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 33 Consoles and Devices . . . . . . . . . . . . . . . . . . . . . . . . . . . 483

ii XEmacs Lisp Reference Manual

34 Positions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 35 Markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 36 Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 37 Searching and Matching . . . . . . . . . . . . . . . . . . . . . . . . . 549 38 Syntax Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 39 Abbrevs And Abbrev Expansion . . . . . . . . . . . . . . . . . . . 579 40 Extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 41 Specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 599 42 Faces and Window-System Objects . . . . . . . . . . . . . . . . . 615 43 Glyphs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 44 Annotations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 45 Emacs Display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 46 Hash Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 47 Range Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 669 48 Databases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 49 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673 50 Operating System Interface. . . . . . . . . . . . . . . . . . . . . . . 691 51 Functions Specific to the X Window System . . . . . . . . . . . 713 52 ToolTalk Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719 53 LDAP Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 725 54 Internationalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 55 MULE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 733 Appendix A Tips and Standards . . . . . . . . . . . . . . . . . . . . . 757 Appendix B Building XEmacs; Allocation of Objects . . . . . . . 765 Appendix C Standard Errors . . . . . . . . . . . . . . . . . . . . . . . . 773 Appendix D Buffer-Local Variables. . . . . . . . . . . . . . . . . . . . 777 Appendix E Standard Keymaps . . . . . . . . . . . . . . . . . . . . . . 781 Appendix F Standard Hooks . . . . . . . . . . . . . . . . . . . . . . . . 785 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 793

Table of Contents

GNU GENERAL PUBLIC LICENSE . . . . . . . . . . 1

Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION

AND MODIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 How to Apply These Terms to Your New Programs . . . . . . . . . . . . . 6

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

1.1 Caveats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.2 Lisp History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.3 Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

1.3.1 Some Terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.3.2 nil and t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.3.3 Evaluation Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.3.4 Printing Notation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.3.5 Error Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.3.6 Buffer Text Notation. . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.3.7 Format of Descriptions . . . . . . . . . . . . . . . . . . . . . . . . . 12

1.3.7.1 A Sample Function Description . . . . . . . . 12 1.3.7.2 A Sample Variable Description . . . . . . . . . 14

1.4 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2 Lisp Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.1 Printed Representation and Read Syntax. . . . . . . . . . . . . . . . . 15 2.2 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.3 Primitive Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.4 Programming Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

2.4.1 Integer Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.4.2 Floating Point Type . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.4.3 Character Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.4.4 Symbol Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.4.5 Sequence Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.4.6 Cons Cell and List Types . . . . . . . . . . . . . . . . . . . . . . 22

2.4.6.1 Dotted Pair Notation. . . . . . . . . . . . . . . . . . 24 2.4.6.2 Association List Type . . . . . . . . . . . . . . . . . 24

2.4.7 Array Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 2.4.8 String Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 2.4.9 Vector Type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.4.10 Bit Vector Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.4.11 Function Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2.4.12 Macro Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2.4.13 Primitive Function Type . . . . . . . . . . . . . . . . . . . . . . 27 2.4.14 Compiled-Function Type . . . . . . . . . . . . . . . . . . . . . . 28 2.4.15 Autoload Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.4.16 Char Table Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.4.17 Hash Table Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.4.18 Range Table Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.4.19 Weak List Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

2.5 Editing Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.5.1 Buffer Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.5.2 Marker Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 2.5.3 Extent Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 2.5.4 Window Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.5.5 Frame Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.5.6 Device Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.5.7 Console Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 2.5.8 Window Configuration Type. . . . . . . . . . . . . . . . . . . . 32 2.5.9 Event Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 2.5.10 Process Type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 2.5.11 Stream Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 2.5.12 Keymap Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 2.5.13 Syntax Table Type . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 2.5.14 Display Table Type . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.5.15 Database Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.5.16 Charset Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.5.17 Coding System Type. . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.5.18 ToolTalk Message Type . . . . . . . . . . . . . . . . . . . . . . . 34 2.5.19 ToolTalk Pattern Type. . . . . . . . . . . . . . . . . . . . . . . . 34

2.6 Window-System Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.6.1 Face Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.6.2 Glyph Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.6.3 Specifier Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.6.4 Font Instance Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.6.5 Color Instance Type . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.6.6 Image Instance Type. . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.6.7 Toolbar Button Type . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.6.8 Subwindow Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.6.9 X Resource Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

2.7 Type Predicates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.8 Equality Predicates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

3 Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

3.1 Integer Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 3.2 Floating Point Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.3 Type Predicates for Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.4 Comparison of Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 3.5 Numeric Conversions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 3.6 Arithmetic Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.7 Rounding Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.8 Bitwise Operations on Integers . . . . . . . . . . . . . . . . . . . . . . . . . . 53 3.9 Standard Mathematical Functions . . . . . . . . . . . . . . . . . . . . . . . 56

3.10 Random Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

4 Strings and Characters . . . . . . . . . . . . . . . . . . . . 59

4.1 String and Character Basics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 4.2 The Predicates for Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 4.3 Creating Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 4.4 The Predicates for Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 4.5 Character Codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 4.6 Comparison of Characters and Strings . . . . . . . . . . . . . . . . . . . 63 4.7 Conversion of Characters and Strings . . . . . . . . . . . . . . . . . . . . 65 4.8 Modifying Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 4.9 String Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.10 Formatting Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.11 Character Case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.12 The Case Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.13 The Char Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

4.13.1 Char Table Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.13.2 Working With Char Tables . . . . . . . . . . . . . . . . . . . . 74

5 Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

5.1 Lists and Cons Cells. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 5.2 Lists as Linked Pairs of Boxes . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 5.3 Predicates on Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.4 Accessing Elements of Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 5.5 Building Cons Cells and Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 5.6 Modifying Existing List Structure . . . . . . . . . . . . . . . . . . . . . . . 85

5.6.1 Altering List Elements with setcar . . . . . . . . . . . . . 85 5.6.2 Altering the CDR of a List . . . . . . . . . . . . . . . . . . . . . 86 5.6.3 Functions that Rearrange Lists . . . . . . . . . . . . . . . . . 87

5.7 Using Lists as Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 5.8 Association Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 5.9 Property Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

5.9.1 Working With Normal Plists . . . . . . . . . . . . . . . . . . . 96 5.9.2 Working With Lax Plists . . . . . . . . . . . . . . . . . . . . . . . 97 5.9.3 Converting Plists To/From Alists . . . . . . . . . . . . . . . 98

5.10 Weak Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

6 Sequences, Arrays, and Vectors . . . . . . . . . . . 101

6.1 Sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 6.2 Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 6.3 Functions that Operate on Arrays . . . . . . . . . . . . . . . . . . . . . . 104 6.4 Vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 6.5 Functions That Operate on Vectors . . . . . . . . . . . . . . . . . . . . . 106 6.6 Bit Vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 6.7 Functions That Operate on Bit Vectors . . . . . . . . . . . . . . . . . 108

7 Symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

7.1 Symbol Components. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.2 Defining Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 7.3 Creating and Interning Symbols . . . . . . . . . . . . . . . . . . . . . . . . 113 7.4 Symbol Properties. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

7.4.1 Property Lists and Association Lists . . . . . . . . . . . 116 7.4.2 Property List Functions for Symbols . . . . . . . . . . . 116 7.4.3 Property Lists Outside Symbols . . . . . . . . . . . . . . . 117

8 Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

8.1 Eval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 8.2 Kinds of Forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

8.2.1 Self-Evaluating Forms. . . . . . . . . . . . . . . . . . . . . . . . . 122 8.2.2 Symbol Forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 8.2.3 Classification of List Forms . . . . . . . . . . . . . . . . . . . . 123 8.2.4 Symbol Function Indirection. . . . . . . . . . . . . . . . . . . 123 8.2.5 Evaluation of Function Forms . . . . . . . . . . . . . . . . . 124 8.2.6 Lisp Macro Evaluation . . . . . . . . . . . . . . . . . . . . . . . . 124 8.2.7 Special Forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 8.2.8 Autoloading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

8.3 Quoting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

9 Control Structures. . . . . . . . . . . . . . . . . . . . . . . 129

9.1 Sequencing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 9.2 Conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 9.3 Constructs for Combining Conditions . . . . . . . . . . . . . . . . . . . 132 9.4 Iteration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 9.5 Nonlocal Exits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

9.5.1 Explicit Nonlocal Exits: catch and throw . . . . . . 134 9.5.2 Examples of catch and throw . . . . . . . . . . . . . . . . . 135 9.5.3 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

9.5.3.1 How to Signal an Error . . . . . . . . . . . . . . . 137 9.5.3.2 How XEmacs Processes Errors . . . . . . . . 138 9.5.3.3 Writing Code to Handle Errors. . . . . . . . 138 9.5.3.4 Error Symbols and Condition Names . . 140

9.5.4 Cleaning Up from Nonlocal Exits . . . . . . . . . . . . . . 142

10 Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145

10.1 Global Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 10.2 Variables That Never Change . . . . . . . . . . . . . . . . . . . . . . . . . 145 10.3 Local Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 10.4 When a Variable is “Void” . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 10.5 Defining Global Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 10.6 Accessing Variable Values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 10.7 How to Alter a Variable Value . . . . . . . . . . . . . . . . . . . . . . . . 152 10.8 Scoping Rules for Variable Bindings . . . . . . . . . . . . . . . . . . . 154

10.8.1 Scope. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 10.8.2 Extent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 10.8.3 Implementation of Dynamic Scoping . . . . . . . . . . 156 10.8.4 Proper Use of Dynamic Scoping . . . . . . . . . . . . . . 156

10.9 Buffer-Local Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 10.9.1 Introduction to Buffer-Local Variables. . . . . . . . . 157 10.9.2 Creating and Deleting Buffer-Local Bindings… 157 10.9.3 The Default Value of a Buffer-Local Variable . . 159

10.10 Variable Aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

11 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163

11.1 What Is a Function? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 11.2 Lambda Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

11.2.1 Components of a Lambda Expression . . . . . . . . . 164 11.2.2 A Simple Lambda-Expression Example . . . . . . . . 165 11.2.3 Advanced Features of Argument Lists . . . . . . . . . 165 11.2.4 Documentation Strings of Functions. . . . . . . . . . . 167

11.3 Naming a Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 11.4 Defining Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 11.5 Calling Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 11.6 Mapping Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 11.7 Anonymous Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 11.8 Accessing Function Cell Contents. . . . . . . . . . . . . . . . . . . . . . 173 11.9 Inline Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 11.10 Other Topics Related to Functions . . . . . . . . . . . . . . . . . . . 176

12 Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

12.1 A Simple Example of a Macro. . . . . . . . . . . . . . . . . . . . . . . . . 177 12.2 Expansion of a Macro Call . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 12.3 Macros and Byte Compilation. . . . . . . . . . . . . . . . . . . . . . . . . 178 12.4 Defining Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 12.5 Backquote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 12.6 Common Problems Using Macros. . . . . . . . . . . . . . . . . . . . . . 180

12.6.1 Evaluating Macro Arguments Repeatedly. . . . . . 180 12.6.2 Local Variables in Macro Expansions. . . . . . . . . . 182 12.6.3 Evaluating Macro Arguments in Expansion …. 182 12.6.4 How Many Times is the Macro Expanded? …. 183

13 Writing Customization Definitions . . . . . . . 185

13.1 Common Keywords for All Kinds of Items . . . . . . . . . . . . . 185 13.2 Defining Custom Groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 13.3 Defining Customization Variables. . . . . . . . . . . . . . . . . . . . . . 187 13.4 Customization Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

13.4.1 Simple Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 13.4.2 Composite Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 13.4.3 Splicing into Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 13.4.4 Type Keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192

14 Loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

14.1 How Programs Do Loading . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 14.2 Autoload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 14.3 Repeated Loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 14.4 Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 14.5 Unloading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 14.6 Hooks for Loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

15 Byte Compilation. . . . . . . . . . . . . . . . . . . . . . . 205

15.1 Performance of Byte-Compiled Code. . . . . . . . . . . . . . . . . . . 205 15.2 The Compilation Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 15.3 Documentation Strings and Compilation . . . . . . . . . . . . . . . 208 15.4 Dynamic Loading of Individual Functions . . . . . . . . . . . . . . 209 15.5 Evaluation During Compilation . . . . . . . . . . . . . . . . . . . . . . . 209 15.6 Compiled-Function Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 15.7 Disassembled Byte-Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212

16 Debugging Lisp Programs . . . . . . . . . . . . . . . 217

16.1 The Lisp Debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 16.1.1 Entering the Debugger on an Error . . . . . . . . . . . 217 16.1.2 Debugging Infinite Loops. . . . . . . . . . . . . . . . . . . . . 218 16.1.3 Entering the Debugger on a Function Call . . . . . 218 16.1.4 Explicit Entry to the Debugger . . . . . . . . . . . . . . . 220 16.1.5 Using the Debugger. . . . . . . . . . . . . . . . . . . . . . . . . . 220 16.1.6 Debugger Commands . . . . . . . . . . . . . . . . . . . . . . . . 221 16.1.7 Invoking the Debugger . . . . . . . . . . . . . . . . . . . . . . . 222 16.1.8 Internals of the Debugger . . . . . . . . . . . . . . . . . . . . 224

16.2 Debugging Invalid Lisp Syntax . . . . . . . . . . . . . . . . . . . . . . . . 226 16.2.1 Excess Open Parentheses. . . . . . . . . . . . . . . . . . . . . 226 16.2.2 Excess Close Parentheses. . . . . . . . . . . . . . . . . . . . . 226

16.3 Debugging Problems in Compilation . . . . . . . . . . . . . . . . . . . 227 16.4 Edebug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

16.4.1 Using Edebug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 16.4.2 Instrumenting for Edebug . . . . . . . . . . . . . . . . . . . . 229 16.4.3 Edebug Execution Modes . . . . . . . . . . . . . . . . . . . . 230 16.4.4 Jumping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 16.4.5 Miscellaneous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 16.4.6 Breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

16.4.6.1 Global Break Condition . . . . . . . . . . . . . 233 16.4.6.2 Embedded Breakpoints . . . . . . . . . . . . . . 233

16.4.7 Trapping Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 16.4.8 Edebug Views . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 16.4.9 Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 16.4.10 Evaluation List Buffer . . . . . . . . . . . . . . . . . . . . . . 235 16.4.11 Reading in Edebug . . . . . . . . . . . . . . . . . . . . . . . . . 237 16.4.12 Printing in Edebug . . . . . . . . . . . . . . . . . . . . . . . . . 237 16.4.13 Tracing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 16.4.14 Coverage Testing . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 16.4.15 The Outside Context . . . . . . . . . . . . . . . . . . . . . . . 239

16.4.15.1 Checking Whether to Stop. . . . . . . . . . 239 16.4.15.2 Edebug Display Update . . . . . . . . . . . . 239 16.4.15.3 Edebug Recursive Edit . . . . . . . . . . . . . 240

16.4.16 Instrumenting Macro Calls . . . . . . . . . . . . . . . . . . 240 16.4.16.1 Specification List . . . . . . . . . . . . . . . . . . 241 16.4.16.2 Backtracking . . . . . . . . . . . . . . . . . . . . . . 244 16.4.16.3 Debugging Backquote . . . . . . . . . . . . . . 245 16.4.16.4 Specification Examples . . . . . . . . . . . . . 246

16.4.17 Edebug Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247

17 Reading and Printing Lisp Objects . . . . . . 251

17.1 Introduction to Reading and Printing . . . . . . . . . . . . . . . . . . 251 17.2 Input Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 17.3 Input Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 17.4 Output Streams. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 17.5 Output Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 17.6 Variables Affecting Output. . . . . . . . . . . . . . . . . . . . . . . . . . . . 258

18 Minibuffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261

18.1 Introduction to Minibuffers . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 18.2 Reading Text Strings with the Minibuffer . . . . . . . . . . . . . . 261 18.3 Reading Lisp Objects with the Minibuffer . . . . . . . . . . . . . . 263 18.4 Minibuffer History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 18.5 Completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266

18.5.1 Basic Completion Functions . . . . . . . . . . . . . . . . . . 266 18.5.2 Completion and the Minibuffer . . . . . . . . . . . . . . . 268 18.5.3 Minibuffer Commands That Do Completion … 269 18.5.4 High-Level Completion Functions . . . . . . . . . . . . . 271 18.5.5 Reading File Names . . . . . . . . . . . . . . . . . . . . . . . . . 273 18.5.6 Programmed Completion. . . . . . . . . . . . . . . . . . . . . 274

18.6 Yes-or-No Queries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 18.7 Asking Multiple Y-or-N Questions . . . . . . . . . . . . . . . . . . . . . 277 18.8 Minibuffer Miscellany. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278

19 Command Loop . . . . . . . . . . . . . . . . . . . . . . . . 281

19.1 Command Loop Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 19.2 Defining Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282

19.2.1 Using interactive . . . . . . . . . . . . . . . . . . . . . . . . . . 282 19.2.2 Code Characters for interactive . . . . . . . . . . . . 283 19.2.3 Examples of Using interactive . . . . . . . . . . . . . . 286

19.3 Interactive Call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 19.4 Information from the Command Loop . . . . . . . . . . . . . . . . . 288 19.5 Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290

19.5.1 Event Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 19.5.2 Contents of the Different Types of Events . . . . . 291 19.5.3 Event Predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 19.5.4 Accessing the Position of a Mouse Event . . . . . . 294

19.5.4.1 Frame-Level Event Position Info . . . . . 295 19.5.4.2 Window-Level Event Position Info. … 295 19.5.4.3 Event Text Position Info. . . . . . . . . . . . . 296 19.5.4.4 Event Glyph Position Info . . . . . . . . . . . 296 19.5.4.5 Event Toolbar Position Info. . . . . . . . . . 297 19.5.4.6 Other Event Position Info . . . . . . . . . . . 297

19.5.5 Accessing the Other Contents of Events . . . . . . . 297 19.5.6 Working With Events . . . . . . . . . . . . . . . . . . . . . . . . 298 19.5.7 Converting Events . . . . . . . . . . . . . . . . . . . . . . . . . . . 300

19.6 Reading Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 19.6.1 Key Sequence Input . . . . . . . . . . . . . . . . . . . . . . . . . 301 19.6.2 Reading One Event . . . . . . . . . . . . . . . . . . . . . . . . . . 302 19.6.3 Dispatching an Event . . . . . . . . . . . . . . . . . . . . . . . . 303 19.6.4 Quoted Character Input . . . . . . . . . . . . . . . . . . . . . 303 19.6.5 Miscellaneous Event Input Features . . . . . . . . . . . 304

19.7 Waiting for Elapsed Time or Input . . . . . . . . . . . . . . . . . . . . 305 19.8 Quitting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 19.9 Prefix Command Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . 308 19.10 Recursive Editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 19.11 Disabling Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 19.12 Command History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 19.13 Keyboard Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313

20 Keymaps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315

20.1 Keymap Terminology. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 20.2 Format of Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 20.3 Creating Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 20.4 Inheritance and Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 20.5 Key Sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 20.6 Prefix Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 20.7 Active Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 20.8 Key Lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 20.9 Functions for Key Lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 20.10 Changing Key Bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327 20.11 Commands for Binding Keys . . . . . . . . . . . . . . . . . . . . . . . . . 330 20.12 Scanning Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 20.13 Other Keymap Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335

21 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337

21.1 Format of Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 337 21.2 Format of the Menubar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 21.3 Menubar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 21.4 Modifying Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 21.5 Menu Filters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 21.6 Pop-Up Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 344 21.7 Menu Accelerators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345

21.7.1 Creating Menu Accelerators . . . . . . . . . . . . . . . . . . 346 21.7.2 Keyboard Menu Traversal . . . . . . . . . . . . . . . . . . . . 346 21.7.3 Menu Accelerator Functions . . . . . . . . . . . . . . . . . . 346

21.8 Buffers Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 347

22 Dialog Boxes . . . . . . . . . . . . . . . . . . . . . . . . . . . 349

22.1 Dialog Box Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349 22.2 Dialog Box Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 349

23 Toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351

23.1 Toolbar Intro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 23.2 Toolbar Descriptor Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 23.3 Specifying the Toolbar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 353 23.4 Other Toolbar Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354

24 scrollbars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357

25 Drag and Drop . . . . . . . . . . . . . . . . . . . . . . . . . 359

25.1 Supported Protocols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 25.1.1 OffiX DND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 25.1.2 CDE dt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 25.1.3 MSWindows OLE . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 25.1.4 Loose ends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360

25.2 Drop Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360 25.3 Drag Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 360

26 Major and Minor Modes . . . . . . . . . . . . . . . . 361

26.1 Major Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 26.1.1 Major Mode Conventions . . . . . . . . . . . . . . . . . . . . 362 26.1.2 Major Mode Examples . . . . . . . . . . . . . . . . . . . . . . . 363 26.1.3 How XEmacs Chooses a Major Mode . . . . . . . . . 367 26.1.4 Getting Help about a Major Mode . . . . . . . . . . . . 369 26.1.5 Defining Derived Modes. . . . . . . . . . . . . . . . . . . . . . 370

26.2 Minor Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 26.2.1 Conventions for Writing Minor Modes. . . . . . . . . 371 26.2.2 Keymaps and Minor Modes . . . . . . . . . . . . . . . . . . 372

26.3 Modeline Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 372 26.3.1 The Data Structure of the Modeline . . . . . . . . . . 373 26.3.2 Variables Used in the Modeline . . . . . . . . . . . . . . . 374 26.3.3 %-Constructs in the ModeLine . . . . . . . . . . . . . . . . 376

26.4 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378

27 Documentation . . . . . . . . . . . . . . . . . . . . . . . . . 381

27.1 Documentation Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 27.2 Access to Documentation Strings . . . . . . . . . . . . . . . . . . . . . . 382 27.3 Substituting Key Bindings in Documentation. . . . . . . . . . . 384 27.4 Describing Characters for Help Messages . . . . . . . . . . . . . . . 386 27.5 Help Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 27.6 Obsoleteness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389

28 Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391

28.1 Visiting Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 391 28.1.1 Functions for Visiting Files . . . . . . . . . . . . . . . . . . . 391 28.1.2 Subroutines of Visiting. . . . . . . . . . . . . . . . . . . . . . . 393

28.2 Saving Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 394 28.3 Reading from Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 396 28.4 Writing to Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397 28.5 File Locks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 398 28.6 Information about Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 399

28.6.1 Testing Accessibility . . . . . . . . . . . . . . . . . . . . . . . . . 399 28.6.2 Distinguishing Kinds of Files . . . . . . . . . . . . . . . . . 400 28.6.3 Truenames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 401 28.6.4 Other Information about Files . . . . . . . . . . . . . . . . 401

28.7 Changing File Names and Attributes . . . . . . . . . . . . . . . . . . 404 28.8 File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406

28.8.1 File Name Components . . . . . . . . . . . . . . . . . . . . . . 406 28.8.2 Directory Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . 408 28.8.3 Absolute and Relative File Names . . . . . . . . . . . . 409 28.8.4 Functions that Expand Filenames. . . . . . . . . . . . . 409 28.8.5 Generating Unique File Names . . . . . . . . . . . . . . . 411 28.8.6 File Name Completion . . . . . . . . . . . . . . . . . . . . . . . 411

28.9 Contents of Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 412 28.10 Creating and Deleting Directories . . . . . . . . . . . . . . . . . . . . 413 28.11 Making Certain File Names “Magic” . . . . . . . . . . . . . . . . . 414 28.12 Partial Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 416

28.12.1 Intro to Partial Files. . . . . . . . . . . . . . . . . . . . . . . . 416 28.12.2 Creating a Partial File . . . . . . . . . . . . . . . . . . . . . . 416 28.12.3 Detached Partial Files . . . . . . . . . . . . . . . . . . . . . . 416

28.13 File Format Conversion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 417 28.14 Files and MS-DOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 419

29 Backups and Auto-Saving . . . . . . . . . . . . . . . 421

29.1 Backup Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421 29.1.1 Making Backup Files . . . . . . . . . . . . . . . . . . . . . . . . 421 29.1.2 Backup by Renaming or by Copying? . . . . . . . . . 422 29.1.3 Making and Deleting Numbered Backup Files. . 423 29.1.4 Naming Backup Files . . . . . . . . . . . . . . . . . . . . . . . . 424

29.2 Auto-Saving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 425 29.3 Reverting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 428

30 Buffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431

30.1 Buffer Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 30.2 The Current Buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 431 30.3 Buffer Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 433 30.4 Buffer File Name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 434 30.5 Buffer Modification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 436 30.6 Comparison of Modification Time . . . . . . . . . . . . . . . . . . . . . 437 30.7 Read-Only Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 438 30.8 The Buffer List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 439 30.9 Creating Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 440 30.10 Killing Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 441 30.11 Indirect Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 442

31 Windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 445

31.1 Basic Concepts of Emacs Windows . . . . . . . . . . . . . . . . . . . . 445 31.2 Splitting Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 446 31.3 Deleting Windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 449 31.4 Selecting Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 450 31.5 Cyclic Ordering of Windows . . . . . . . . . . . . . . . . . . . . . . . . . . 451 31.6 Buffers and Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 453 31.7 Displaying Buffers in Windows . . . . . . . . . . . . . . . . . . . . . . . . 453 31.8 Choosing a Window for Display . . . . . . . . . . . . . . . . . . . . . . . 455 31.9 Windows and Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 458 31.10 The Window Start Position . . . . . . . . . . . . . . . . . . . . . . . . . . 458 31.11 Vertical Scrolling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 460 31.12 Horizontal Scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 462 31.13 The Size of a Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 464 31.14 The Position of a Window . . . . . . . . . . . . . . . . . . . . . . . . . . . 465 31.15 Changing the Size of a Window . . . . . . . . . . . . . . . . . . . . . . 466 31.16 Window Configurations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 468

32 Frames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471

32.1 Creating Frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471 32.2 Frame Properties. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 471

32.2.1 Access to Frame Properties. . . . . . . . . . . . . . . . . . . 472 32.2.2 Initial Frame Properties. . . . . . . . . . . . . . . . . . . . . . 472 32.2.3 X Window Frame Properties . . . . . . . . . . . . . . . . . 473 32.2.4 Frame Size And Position . . . . . . . . . . . . . . . . . . . . . 475 32.2.5 The Name of a Frame (As Opposed to Its Title)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 475

32.3 Frame Titles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 32.4 Deleting Frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 476 32.5 Finding All Frames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 477 32.6 Frames and Windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 32.7 Minibuffers and Frames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 478 32.8 Input Focus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 479 32.9 Visibility of Frames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 480 32.10 Raising and Lowering Frames . . . . . . . . . . . . . . . . . . . . . . . . 480 32.11 Frame Configurations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 481 32.12 Hooks for Customizing Frame Behavior . . . . . . . . . . . . . . . 481

33 Consoles and Devices . . . . . . . . . . . . . . . . . . . 483

33.1 Basic Console Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 483 33.2 Basic Device Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 484 33.3 Console Types and Device Classes . . . . . . . . . . . . . . . . . . . . . 484 33.4 Connecting to a Console or Device . . . . . . . . . . . . . . . . . . . . 485 33.5 The Selected Console and Device . . . . . . . . . . . . . . . . . . . . . . 486 33.6 Console and Device I/O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 486

34 Positions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487

34.1 Point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 487 34.2 Motion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 488

34.2.1 Motion by Characters. . . . . . . . . . . . . . . . . . . . . . . . 488 34.2.2 Motion by Words . . . . . . . . . . . . . . . . . . . . . . . . . . . . 489 34.2.3 Motion to an End of the Buffer . . . . . . . . . . . . . . . 489 34.2.4 Motion by Text Lines . . . . . . . . . . . . . . . . . . . . . . . . 490 34.2.5 Motion by Screen Lines . . . . . . . . . . . . . . . . . . . . . . 492 34.2.6 Moving over Balanced Expressions . . . . . . . . . . . . 493 34.2.7 Skipping Characters . . . . . . . . . . . . . . . . . . . . . . . . . 494

34.3 Excursions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 495 34.4 Narrowing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 496

35 Markers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499

35.1 Overview of Markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 499 35.2 Predicates on Markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 500 35.3 Functions That Create Markers . . . . . . . . . . . . . . . . . . . . . . . 501 35.4 Information from Markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 35.5 Changing Marker Positions . . . . . . . . . . . . . . . . . . . . . . . . . . . 503 35.6 The Mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 504 35.7 The Region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 507

36 Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 511

36.1 Examining Text Near Point . . . . . . . . . . . . . . . . . . . . . . . . . . . 511 36.2 Examining Buffer Contents . . . . . . . . . . . . . . . . . . . . . . . . . . . 512 36.3 Comparing Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 36.4 Inserting Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 513 36.5 User-Level Insertion Commands . . . . . . . . . . . . . . . . . . . . . . . 515 36.6 Deleting Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 516 36.7 User-Level Deletion Commands . . . . . . . . . . . . . . . . . . . . . . . 517 36.8 The Kill Ring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 519

36.8.1 Kill Ring Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . 520 36.8.2 Functions for Killing . . . . . . . . . . . . . . . . . . . . . . . . . 520 36.8.3 Functions for Yanking . . . . . . . . . . . . . . . . . . . . . . . 521 36.8.4 Low-Level Kill Ring . . . . . . . . . . . . . . . . . . . . . . . . . 521 36.8.5 Internals of the Kill Ring. . . . . . . . . . . . . . . . . . . . . 522

36.9 Undo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 523 36.10 Maintaining Undo Lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 525 36.11 Filling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 526 36.12 Margins for Filling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 528 36.13 Auto Filling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 36.14 Sorting Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 529 36.15 Counting Columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 533 36.16 Indentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 534

36.16.1 Indentation Primitives . . . . . . . . . . . . . . . . . . . . . . 534 36.16.2 Indentation Controlled by Major Mode. . . . . . . 534 36.16.3 Indenting an Entire Region. . . . . . . . . . . . . . . . . . 535 36.16.4 Indentation Relative to Previous Lines . . . . . . . 536 36.16.5 Adjustable “Tab Stops” . . . . . . . . . . . . . . . . . . . . . 537 36.16.6 Indentation-Based Motion Commands. . . . . . . . 537

36.17 Case Changes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 538 36.18 Text Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 539

36.18.1 Examining Text Properties . . . . . . . . . . . . . . . . . . 540 36.18.2 Changing Text Properties . . . . . . . . . . . . . . . . . . . 540 36.18.3 Property Search Functions . . . . . . . . . . . . . . . . . . 542 36.18.4 Properties with Special Meanings . . . . . . . . . . . . 543 36.18.5 Saving Text Properties in Files . . . . . . . . . . . . . . 543

36.19 Substituting for a Character Code . . . . . . . . . . . . . . . . . . . . 544 36.20 Registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 545 36.21 Transposition of Text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546 36.22 Change Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 546

37 Searching and Matching . . . . . . . . . . . . . . . . 549

37.1 Searching for Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 549 37.2 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 550

37.2.1 Syntax of Regular Expressions . . . . . . . . . . . . . . . . 551 37.2.2 Complex Regexp Example. . . . . . . . . . . . . . . . . . . . 556

37.3 Regular Expression Searching . . . . . . . . . . . . . . . . . . . . . . . . . 557 37.4 POSIX Regular Expression Searching . . . . . . . . . . . . . . . . . . 559 37.5 Search and Replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 560 37.6 The Match Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 561

37.6.1 Simple Match Data Access . . . . . . . . . . . . . . . . . . . 562 37.6.2 Replacing the Text That Matched . . . . . . . . . . . . 563 37.6.3 Accessing the Entire Match Data . . . . . . . . . . . . . 564 37.6.4 Saving and Restoring the Match Data . . . . . . . . . 565

37.7 Searching and Case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 565 37.8 Standard Regular Expressions Used in Editing . . . . . . . . . 566

38 Syntax Tables . . . . . . . . . . . . . . . . . . . . . . . . . . 569

38.1 Syntax Table Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 569 38.2 Syntax Descriptors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 570

38.2.1 Table of Syntax Classes . . . . . . . . . . . . . . . . . . . . . . 570 38.2.2 Syntax Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 572

38.3 Syntax Table Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 573 38.4 Motion and Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 575 38.5 Parsing Balanced Expressions . . . . . . . . . . . . . . . . . . . . . . . . . 575 38.6 Some Standard Syntax Tables. . . . . . . . . . . . . . . . . . . . . . . . . 578 38.7 Syntax Table Internals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 578

39 Abbrevs And Abbrev Expansion. . . . . . . . . 579

39.1 Setting Up Abbrev Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 39.2 Abbrev Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 579 39.3 Defining Abbrevs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 580 39.4 Saving Abbrevs in Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 581 39.5 Looking Up and Expanding Abbreviations . . . . . . . . . . . . . 581 39.6 Standard Abbrev Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 583

40 Extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585

40.1 Introduction to Extents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 585 40.2 Creating and Modifying Extents. . . . . . . . . . . . . . . . . . . . . . . 586 40.3 Extent Endpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 586 40.4 Finding Extents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 587 40.5 Mapping Over Extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 589 40.6 Properties of Extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 591 40.7 Detached Extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 595 40.8 Extent Parents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 40.9 Duplicable Extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 596 40.10 Interaction of Extents with Keyboard and Mouse Events

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 597 40.11 Atomic Extents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 598

41 Specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 599

41.1 Introduction to Specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 599 41.2 In-Depth Overview of a Specifier . . . . . . . . . . . . . . . . . . . . . . 600 41.3 How a Specifier Is Instanced . . . . . . . . . . . . . . . . . . . . . . . . . . 601 41.4 Specifier Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 602 41.5 Adding specifications to a Specifier . . . . . . . . . . . . . . . . . . . . 604 41.6 Retrieving the Specifications from a Specifier . . . . . . . . . . . 607 41.7 Working With Specifier Tags . . . . . . . . . . . . . . . . . . . . . . . . . . 608 41.8 Functions for Instancing a Specifier . . . . . . . . . . . . . . . . . . . . 609 41.9 Example of Specifier Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . 610 41.10 Creating New Specifier Objects . . . . . . . . . . . . . . . . . . . . . . 611 41.11 Functions for Checking the Validity of Specifier Components

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 612 41.12 Other Functions for Working with Specifications in a

Specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 613

42 Faces and Window-System Objects . . . . . . 615

42.1 Faces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 615 42.1.1 Merging Faces for Display . . . . . . . . . . . . . . . . . . . . 615 42.1.2 Basic Functions for Working with Faces . . . . . . . 616 42.1.3 Face Properties. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 616 42.1.4 Face Convenience Functions . . . . . . . . . . . . . . . . . . 619 42.1.5 Other Face Display Functions . . . . . . . . . . . . . . . . 621

42.2 Fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 42.2.1 Font Specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 42.2.2 Font Instances. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 621 42.2.3 Font Instance Names . . . . . . . . . . . . . . . . . . . . . . . . 622 42.2.4 Font Instance Size . . . . . . . . . . . . . . . . . . . . . . . . . . . 622 42.2.5 Font Instance Characteristics . . . . . . . . . . . . . . . . . 622 42.2.6 Font Convenience Functions . . . . . . . . . . . . . . . . . . 623

42.3 Colors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 42.3.1 Color Specifiers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 42.3.2 Color Instances. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 623 42.3.3 Color Instance Properties . . . . . . . . . . . . . . . . . . . . 624 42.3.4 Color Convenience Functions . . . . . . . . . . . . . . . . . 624

43 Glyphs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625

43.1 Glyph Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 43.1.1 Creating Glyphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 625 43.1.2 Glyph Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . 626 43.1.3 Glyph Convenience Functions . . . . . . . . . . . . . . . . 628 43.1.4 Glyph Dimensions . . . . . . . . . . . . . . . . . . . . . . . . . . . 630

43.2 Images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 630 43.2.1 Image Specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 630 43.2.2 Image Instantiator Conversion . . . . . . . . . . . . . . . . 634 43.2.3 Image Instances . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 635

43.2.3.1 Image Instance Types . . . . . . . . . . . . . . . 635

43.2.3.2 Image Instance Functions . . . . . . . . . . . . 636

43.3 Glyph Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 638 43.4 Mouse Pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 638 43.5 Redisplay Glyphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 639 43.6 Subwindows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 640

44 Annotations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641

44.1 Annotation Basics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 641 44.2 Annotation Primitives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 642 44.3 Annotation Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 643 44.4 Locating Annotations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 44.5 Margin Primitives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 644 44.6 Annotation Hooks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 645

45 Emacs Display . . . . . . . . . . . . . . . . . . . . . . . . . 647

45.1 Refreshing the Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 647 45.2 Truncation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 45.3 The Echo Area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 648 45.4 Warnings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 651 45.5 Invisible Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 653 45.6 Selective Display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 654 45.7 The Overlay Arrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 655 45.8 Temporary Displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 656 45.9 Blinking Parentheses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 658 45.10 Usual Display Conventions. . . . . . . . . . . . . . . . . . . . . . . . . . . 658 45.11 Display Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 659

45.11.1 Display Table Format . . . . . . . . . . . . . . . . . . . . . . . 660 45.11.2 Active Display Table. . . . . . . . . . . . . . . . . . . . . . . . 660 45.11.3 Character Descriptors. . . . . . . . . . . . . . . . . . . . . . . 660

45.12 Beeping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 661

46 Hash Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 665

46.1 Introduction to Hash Tables . . . . . . . . . . . . . . . . . . . . . . . . . . 665 46.2 Working With Hash Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . 665 46.3 Weak Hash Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 666

47 Range Tables. . . . . . . . . . . . . . . . . . . . . . . . . . . 669

47.1 Introduction to Range Tables . . . . . . . . . . . . . . . . . . . . . . . . . 669 47.2 Working With Range Tables . . . . . . . . . . . . . . . . . . . . . . . . . . 669

48 Databases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671

48.1 Connecting to a Database . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 48.2 Working With a Database . . . . . . . . . . . . . . . . . . . . . . . . . . . . 671 48.3 Other Database Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 672

49 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 673

49.1 Functions that Create Subprocesses. . . . . . . . . . . . . . . . . . . . 673 49.2 Creating a Synchronous Process . . . . . . . . . . . . . . . . . . . . . . . 674 49.3 MS-DOS Subprocesses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 677 49.4 Creating an Asynchronous Process . . . . . . . . . . . . . . . . . . . . 677 49.5 Deleting Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 678 49.6 Process Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 679 49.7 Sending Input to Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . 681 49.8 Sending Signals to Processes . . . . . . . . . . . . . . . . . . . . . . . . . . 682 49.9 Receiving Output from Processes . . . . . . . . . . . . . . . . . . . . . . 683

49.9.1 Process Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 683 49.9.2 Process Filter Functions. . . . . . . . . . . . . . . . . . . . . . 684 49.9.3 Accepting Output from Processes . . . . . . . . . . . . . 686

49.10 Sentinels: Detecting Process Status Changes . . . . . . . . . . 687 49.11 Process Window Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 49.12 Transaction Queues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 688 49.13 Network Connections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 689

50 Operating System Interface . . . . . . . . . . . . . 691

50.1 Starting Up XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 691 50.1.1 Summary: Sequence of Actions at Start Up. … 691 50.1.2 The Init File: ‘.emacs’ . . . . . . . . . . . . . . . . . . . . . . . 692 50.1.3 Terminal-Specific Initialization . . . . . . . . . . . . . . . 693 50.1.4 Command Line Arguments . . . . . . . . . . . . . . . . . . . 694

50.2 Getting out of XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 50.2.1 Killing XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 695 50.2.2 Suspending XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . 696

50.3 Operating System Environment . . . . . . . . . . . . . . . . . . . . . . . 697 50.4 User Identification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 700 50.5 Time of Day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 702 50.6 Time Conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 703 50.7 Timers for Delayed Execution . . . . . . . . . . . . . . . . . . . . . . . . . 705 50.8 Terminal Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 706

50.8.1 Input Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 706 50.8.2 Translating Input Events . . . . . . . . . . . . . . . . . . . . . 707 50.8.3 Recording Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 708

50.9 Terminal Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 709 50.10 Flow Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 710 50.11 Batch Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 711

51 Functions Specific to the X Window System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713

51.1 X Selections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 713 51.2 X Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714

51.2.1 Resources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 714 51.2.2 Data about the X Server . . . . . . . . . . . . . . . . . . . . . 716 51.2.3 Restricting Access to the Server by Other Apps

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 716

51.3 Miscellaneous X Functions and Variables. . . . . . . . . . . . . . . 717

52 ToolTalk Support . . . . . . . . . . . . . . . . . . . . . . . 719

52.1 XEmacs ToolTalk API Summary . . . . . . . . . . . . . . . . . . . . . . 719 52.2 Sending Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 719

52.2.1 Example of Sending Messages . . . . . . . . . . . . . . . . 719 52.2.2 Elisp Interface for Sending Messages . . . . . . . . . . 720

52.3 Receiving Messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 722 52.3.1 Example of Receiving Messages . . . . . . . . . . . . . . . 722 52.3.2 Elisp Interface for Receiving Messages. . . . . . . . . 722

53 LDAP Support . . . . . . . . . . . . . . . . . . . . . . . . . 725

53.1 Building XEmacs with LDAP support . . . . . . . . . . . . . . . . . 725 53.2 XEmacs LDAP API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 725

53.2.1 LDAP Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 725 53.2.2 The High-Level LDAP API. . . . . . . . . . . . . . . . . . . 726 53.2.3 The Low-Level LDAP API . . . . . . . . . . . . . . . . . . . 726

53.2.3.1 The LDAP Lisp Object. . . . . . . . . . . . . . 727 53.2.3.2 Opening and Closing a LDAP Connection

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 727 53.2.3.3 Searching on a LDAP Server (Low-level)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728

53.3 Syntax of Search Filters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 728

54 Internationalization . . . . . . . . . . . . . . . . . . . . . 729

54.1 I18N Levels 1 and 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 54.2 I18N Level 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729

54.2.1 Level 3 Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 54.2.2 Level 3 Primitives . . . . . . . . . . . . . . . . . . . . . . . . . . . 729 54.2.3 Dynamic Messaging. . . . . . . . . . . . . . . . . . . . . . . . . . 730 54.2.4 Domain Specification . . . . . . . . . . . . . . . . . . . . . . . . 730 54.2.5 Documentation String Extraction . . . . . . . . . . . . . 731

54.3 I18N Level 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 731

55 MULE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 733

55.1 Internationalization Terminology . . . . . . . . . . . . . . . . . . . . . . 733 55.2 Charsets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 735

55.2.1 Charset Properties. . . . . . . . . . . . . . . . . . . . . . . . . . . 735 55.2.2 Basic Charset Functions . . . . . . . . . . . . . . . . . . . . . 737 55.2.3 Charset Property Functions . . . . . . . . . . . . . . . . . . 737 55.2.4 Predefined Charsets . . . . . . . . . . . . . . . . . . . . . . . . . 739

55.3 MULE Characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 55.4 Composite Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 740 55.5 ISO 2022 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 741 55.6 Coding Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 743

55.6.1 Coding System Types. . . . . . . . . . . . . . . . . . . . . . . . 744 55.6.2 EOL Conversion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 744 55.6.3 Coding System Properties . . . . . . . . . . . . . . . . . . . . 745 55.6.4 Basic Coding System Functions. . . . . . . . . . . . . . . 747 55.6.5 Coding System Property Functions . . . . . . . . . . . 747 55.6.6 Encoding and Decoding Text . . . . . . . . . . . . . . . . . 748 55.6.7 Detection of Textual Encoding. . . . . . . . . . . . . . . . 748 55.6.8 Big5 and Shift-JIS Functions . . . . . . . . . . . . . . . . . 748

55.7 CCL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 749 55.7.1 CCL Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 750 55.7.2 CCL Statements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 751 55.7.3 CCL Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 752 55.7.4 Calling CCL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 753 55.7.5 CCL Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 754

55.8 Category Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 754

Appendix A Tips and Standards . . . . . . . . . . . 757

A.1 Writing Clean Lisp Programs . . . . . . . . . . . . . . . . . . . . . . . . . . 757 A.2 Tips for Making Compiled Code Fast . . . . . . . . . . . . . . . . . . 760 A.3 Tips for Documentation Strings. . . . . . . . . . . . . . . . . . . . . . . . 760 A.4 Tips on Writing Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 762 A.5 Conventional Headers for XEmacs Libraries. . . . . . . . . . . . . 763

Appendix B Building XEmacs; Allocation of Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 765

B.1 Building XEmacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 765 B.2 Pure Storage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 767 B.3 Garbage Collection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 768

Appendix C Standard Errors . . . . . . . . . . . . . . . 773

Appendix D Buffer-Local Variables . . . . . . . . . 777

Appendix E Standard Keymaps . . . . . . . . . . . . 781

xxii XEmacs Lisp Reference Manual

Appendix F Standard Hooks . . . . . . . . . . . . . . . 785 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 793